A Complete Bibliography of Publications in Marine Ecology Progress Series: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

22 August 2024 Version 1.08

Title word cross-reference

3/4 [FIS04]. ^{+} [CACDC07]. ^{1} [COB^{+}03]. ^{13} [AW03a, AHC^{+}06, AHSPG08, AZKJAM06, BW01, BGS^{+}02b, CF02, DLLB03, Fre03, HH00a, HGMG^{+}04, HCCY02, KPM00, KP01a, KECJ06, KW05, KSK07, LHB09, MSF^{+}07, MH08a, MRHP01, MD07, MBM^{+}00, NEMF09, NHN^{+}06, PCL08, RSN02, RH03, RW08, Rol00, SAT^{+}03, SJE^{+}06, SR03, SKO^{+}08, TPSW00, TBI07, WPRS05]. ^{14} [MRHP01, MOOB06, MSF04, STF^{+}09, TKR06]. ^{1}5 [BMK07, Fre03, SJE^{+}06, TBI07, VS08, AHSPG08, CF02, CVF07, DLLB03, DPD^{+}08, EKBM05, GLS02, GV01, HH05a, HH00a, HGMG^{+}04, KPM00, KP01a, LHB09, LWK^{+}07, MSF^{+}07, MH08a, NEMF09, ORPND01, PCL08, RSN02, RH03, RW08, Rol00, SAT^{+}03, SMM^{+}04, STF^{+}09, WPRS05]. ^{1}8 [BW01, KSK07, RSY^{+}06]. ^{210} [ASLD09]. ^{33} [XYH^{+}09]. ^{34} [HCCY02]. ^{86} [MC05c, SZL^{+}05]. ^{87} [MC05c, SZL^{+}05]. _{2} [CGCMPA01, CGF04, CCB^{+}09a, DGBS08, DHT^{+}08, FPC05, FHL^{+}09, Fen08, FE07, FKSW09, GÁSPR03, GPM08, HLD^{+}07, IHK08, KFU^{+}08, KGB^{+}01, KS04c, Kur08, LRN^{+}09, MRHP01, MMC^{+}07, MD07, PZ07, RR05, RNDF^{+}04, RFG^{+}09, SFF^{+}07, SHL^{+}03, SFKT09, TDSM02, VVTH07, WDM^{+}09, ZER^{+}01]. _{3} [KYYM00]. _{4} [KCTG06, KFU^{+}08, KYYM00]. _{4}^{+} [HH05a]. a [BH04, CSD04, HHAH00, HHKH05, KKA^{+}06, MFS07, SMHL00, SdSC01, WKN^{+}03, YVB^{+}02, YMC05]. α [DBG^{+}09]. ≈ [MFE^{+}05]. β [DBG^{+}09]. : [CIT04, HYLT03, KYYM00, SCIT09, TSI02]. δ [AHC^{+}06, AHSPG08, BMK07, BW01, BGS^{+}02b, CF02, CVF07, DBG^{+}09, DLLB03, Fre03, HH00a, HGMG^{+}04, HCCY02, KPM00, KP01a, KSK07, LHB09, LWK^{+}07, MSF^{+}07, MH08a, NEMF09, RSN02, RH03, RW08, Rol00, RSY^{+}06, SAT^{+}03, SJE^{+}06, SR03, SKO^{+}08, TPSW00, WPRS05]. ϵ [DBG^{+}09]. f [CPSC01]. γ [DBG^{+}09]. μ [FLUF^{+}04]. P/B [NPN06]. × [TLLKC04].

-ATPase [CACDC07]. -diversity [DBG^{+}09]. -driven [DHT^{+}08, Kur08]. -fixing [FPC05]. -induced [MMC^{+}07]. -MGDG [ZER^{+}01]. -ratio [CPSC01]. -reducing [VFJ06].

/K [CACDC07].

0-group [TdLvdVR08, vdVBW00].

1-group [AP00a]. 101 [BWK^{+}08]. 106 [Ano01d]. 109Cd [ABY00]. 1254 [MFA03]. 137Cs [MLR02]. 13^{o} [THS^{+}02]. 155^{o} [SFKT09]. 160^{o} [CIHF01]. 16S [ISSK03]. 170^{o} [DGJ^{+}03]. 18S [PSPD07]. 191 [Ano00c, Ano00b]. 1990s [Car05, Tyn04]. 1997/1998 [GGM^{+}00, KJMP05]. 1997/98 [IG04]. 1998 [MLG^{+}03a, WG01].

2-D [Lei01]. 2-dimensional [EHWS07, NBDG05]. 2-ethylhexyl [SFHS00]. 2-heptanone [NCP^{+}08]. 2001 [Ano02h]. 2002 [Ano02g]. 2003 [Ano04d]. 2003/2004 [Ano04a]. 2006 [Ano07f, tAHSHA07]. 203 [Ano01f]. 20th [TLIL^{+}09]. 210Po [DGB^{+}02]. 211 [Ano01e]. 218 [Ano01d]. 221 [Ano02h]. 223 [Ano02h]. 224 [Ano02c, Ano02d]. 225 [Ano02g]. 236 [Ano02f]. 239 [Ano02e]. 23^{o} [MFE^{+}05]. 240 [Ano03c]. 241 [Ano03b]. 245 [Ano03e]. 247 [Ano03d]. 252 [Ano03f]. 260 [NH03a]. 263 [Ano04d, MBS07]. 267 [Ano06]. 271 [Ano04e]. 279 [Ano04c]. 29 [Ano02f]. 290 [Ano03c]. 291 [Ano05c]. 297 [Ano05d]. 2D [COSM03]. 2E [CBO04]. 2nd [NYK03]. 2nd-order [NYK03].

3-dimensional [DFW^{+}03]. 30-yr [WWHE04]. 313 [tAHSHA07]. 32.5^{o} [GUL^{+}07]. 324 [Ano07d]. 328 [Ano07f]. 329 [Ano07e]. 33^{o} [RGL01]. 341 [Ano07c]. 343 [Ano07h]. 349 [Ano10b, KH07c]. 35^{o} [WKL01]. 393 [Ano10a]. 3D [MWS^{+}04, RGD^{+}08].

4-keto-myxoxanthophyll-like [SGK04]. 4-Nonylphenol [MM05]. 450 [SRLA04]. 4E [CBO04]. 4E-decadienal [CBO04].

50^{o} [PSW^{+}08]. 53^{o} [KBP06].

77 [DCT^{+}04]. 77^{o} [BRS07].

8-oxodGuo [ABB^{+}00]. 80th [ZK09]. 85 [Ano07h].

9^{o} [THS^{+}02].

= [SJ06].

Îles [RATD07]. Îles-de-la-Madeleine [RATD07].

a-proteobacterium [WNWH02]. A. [AKL^{+}04, CAT^{+}06, CCM01, DLT^{+}06, GDR^{+}04a, MDR^{+}06, PPS00, RPS^{+}04, SICT03, TK04]. Ålesund [MAJ^{+}06]. aalge [Bar02b, ÖBFO01, RMT^{+}05, ZPP09]. abalone [BT00, CMS^{+}01, DB02, MBBP08, MARM06, MKK^{+}09, OKH^{+}07, PBFB08, PS06, RWA^{+}02, SJ09b, TSE^{+}08]. Abatus [TR05]. abilities [FBJ00a, NB06b, SP01c]. ability [BK04, MK02, RFB07, SMM^{+}09, SLB07]. Abiotic [BPBH09, PMH07, ABB02, BMV02, GS09, KvKB06, LSH00, MUT04, PZA09]. ablation [RCC01]. ablation-ICP-MS [RCC01]. abnormalities [GMS02]. Abore [FPKC05]. Above [KD00, CC00, LKR^{+}06, TBK^{+}03]. Above- [KD00, CC00]. Abra [GDR^{+}04a, HDW04, MDR^{+}06, MDBG07]. abrades [IC06]. Abrolhos [WAK^{+}09]. Absence [RBB^{+}08, BJN^{+}05]. absorbance [SA06]. absorbing [Zam03b, Zam07]. Absorption [STJ03, Enr05, OMT06, PCYC02, SMSJ04, TKTJ07]. Abudefduf [EM05]. Abundance [BPA^{+}08, CRV03, DNL08b, FF04, FFA03, Hob00, HOVG07, MS05, MMK^{+}08, NGMT01, WK02, XBW07, ABF07, Ark09, BB03a, BJC02, BM02b, BC01a, BL00b, BLT03, BBA^{+}05, BFLÁ01, CM04a, CCC09, CH07a, CF02, CMD^{+}06, CC03a, CM08, CH05, DE01, Das03, DMS^{+}05, DCAL^{+}02, DJ04, DWL09, DRTM09, DF06, DCC^{+}03, EJM01, EGJ07, FTDK07, FE02, FHVMDS07, FVS^{+}08, FGA^{+}04, FTO^{+}09, FHQ^{+}06, FT00, GHV02, GDG02, GDSB04, GMHA09, GL04b, GGG^{+}04, GEC08, GDL04, GUL^{+}07, GBW02, GUT04, GCM01, HLVU00, HN03, HAC06a, HCM01, HD06b, HB08b, Hyr01, HK03c, JRP01, Jef03, JAA^{+}01, JRS^{+}06, KHT^{+}02, KNTK05, KKTN05, KR04, Kim02, KSG03, KB05, LG00a, LH09, LEB07, LS02, LPM^{+}06, LVA07, LME05, MRN07, MBVBC09, MTH04, MHSS08, MS04b, MWS^{+}04, MC08b, MBS03, MBS07, MBG^{+}08]. abundance [NA04a, NW02, PBAD05, PMS^{+}08, PNK^{+}02, Pee02, PPB^{+}09a, PRM^{+}09b, PMR06, PBS^{+}00, RL09a, RH06, RBG01, RCP^{+}05, RAE07, RM09a, RFG^{+}09, RK09, RMAY04, RKJ01, SESP03, SHB03, SMH09, SMP^{+}09, SV01, SA06, STAS06, SGH^{+}06, SAC^{+}03, TGGA04, TSHT02, TT05b, TNSM00b, Vli05, WHM^{+}01, WA05, WPH01, WS06, Zam02, dSHCR07]. abundance-body [DJ04]. Abundance-distribution [FF04]. Abundance-occupancy [FFA03]. abundances [BVMBC07, HB07b, KHB00a, LOT^{+}00, Lat07, OPSM04]. abundant [BK06b, TWT08, WS06]. abuse [WGS^{+}02]. abyss [GSP^{+}02]. Abyssal [DDF01, AW03a, DPGN08, DPGN09, SW08c]. Acanthaster [GHdN07, LRM00, Pra01, San00]. Acanthina [LF02]. Acanthochromis [KHP09]. Acanthogobius [NW05a]. Acanthopagrus [EG03, HJW08]. Acanthopleura [NW06]. Acanthurus [RAC^{+}05, RCP^{+}05]. Acartia [BK05b, BGF^{+}06, BD02, BP03, CT05, CAT^{+}06, CW03b, CD03c, FGM^{+}00, Gor03, GFC^{+}02, HK03b, HSCH07, ISG04, JKS^{+}01, JTLG09, JFA02, KSKL^{+}03, LQRGT06, LWH06, MGZ^{+}03, OSA05, PK05b, PI07a, SSB09, SKK^{+}06, TT05a, TKTJ07, WTP08]. accelerated [WM04]. accessibility [MWS^{+}04, NBN^{+}08]. accident [PE09]. acclimation [BRBPL06, BWvdE^{+}06, LN05, OEDSJ02, RLBB04, RG09, SSGB^{+}08, SYPO09]. acclimatisation [MND^{+}02]. acclimatisation/adaptation [MND^{+}02]. Acclimatization [EG08]. accomplishments [EP00b]. account [HM01, SMP^{+}09]. Accounting [MTF^{+}08]. accumulated [NATR^{+}05]. Accumulation [HJ03, SSO^{+}01, TDP^{+}00, BW04, BTF^{+}02, CG04, DT08, GdLB^{+}09, JK06, JSH01, KS06, KSS^{+}05, KSKL^{+}03, MM06b, RSL09a, RG09, RGT01, RGT06, SFPGA05, VD00]. Accuracy [PSC^{+}04]. Accurate [HD06b]. aceratus [PSPZ09]. acerosa [GTE^{+}09]. Acharax [NSSI06]. achieve [SPD03]. acid [BPC^{+}02, BLS06, BRI^{+}07, BDGS07, BSIS07, CLP08, CPG^{+}09, CG04, ESDKM09, FBHB03, GSL^{+}06, Gor03, GEHZ07, GKWK02, GNAD^{+}03, GNHB^{+}05, GWBdN09, GFN^{+}09, GRM00, HBW^{+}05, HYR^{+}09, HPBT03, HPB^{+}04, IFL02, JBS08, JQ05, KKK^{+}05, KCV^{+}06, KFK^{+}07, KSIG08, KKIV01, KGS^{+}08, LCGN02, LYL^{+}07, LBPH06, LHK04, LHK05, MTH04, MGA^{+}01, MMC^{+}08b, MOPP05, NFN^{+}09, NWIT08, OTB^{+}00, PPB^{+}01, PB07, PJN01, PJN03, PGS^{+}00, RF08b, RGF^{+}03, RLDL02, RBGM06, RMC^{+}07, SMM^{+}04, SZ09, SKDP02, SPGS03, TBBI04, TGNT05, TBI^{+}09, WHP00, WP03, WISH09, WIB08]. acid-base [MOPP05, PB07]. acid-producing [LHK05]. acidification [AMB08, AC08, BF08a, BD07, BVHG08a, BVHG08b, DHT^{+}08, HOT08, Kur08, MMC^{+}07, MMA^{+}09, MAKMK09, MCN09, Pör08, WDM^{+}09]. acidified [IHK08, Kro05]. acids [BN09, BKLGN05, CCM01, CKO^{+}07, DK06, DPGN08, GNAD^{+}04, GKM02, GMR^{+}02, HRR^{+}01, HK03b, HMY^{+}08, HCK^{+}05, ITN^{+}06, ISK07, JQ04, KSI09, LBR04, LGM^{+}09, MCR03, MT00, MMBT03, ODA^{+}09, PTD05, PWAT01, PAB^{+}02, RTH06, SKFPS02, SFPGA05, TBI08, TMM05, VK02, WCBD01]. acorn [CMD^{+}06, GSK^{+}08, VBB06]. acoroides [LVRN02]. Acoustic [BAB^{+}01, CES^{+}09, JBS^{+}09, NMOO03, Tya09, WHCP^{+}04, BBA^{+}05, CHT06, CKTE01, HALL05, JdSM09, KTK^{+}08a, PBH^{+}03, PTSJ06, SAG08, SJL05, TTD^{+}06, TLC06, VCSL^{+}09]. acoustically [BBABL01]. Acoustics [SN09, BBDW09]. Acquisition [BC06a, ACT09, BWPB04, MAH^{+}08, RL03]. Acropora [AM06, BHH05, BMWH08, CSvO09, GDBF06, GDBF07, GNC07, HNYI07, Lir00, LLCHG01, MDW^{+}09, MBW07, NvWY05, NKS03, OYNO08, RR05, RFPB^{+}04, SHSF08, WM05]. across [Alt06, AR00, BBS05, BRG^{+}02, Bla09, CIFN00, CC07, CTD^{+}09, CEBR07, DWL09, FHM02, FWC02, FB07, GUL^{+}07, HPS07, HS08b, HTL^{+}02, Jos09, LBH^{+}04, LPHH^{+}09, MCS05, MBB^{+}04, NBB^{+}04, OPSM04, RS06, SH07b, SFT^{+}05, SCS^{+}07, TJS08, TCW^{+}08, TH06, UBRvO06, VLP^{+}04, VCA^{+}09, Vie09, VMH^{+}08, WHB^{+}06, WMCF08, WDB^{+}03, YQW^{+}01, ZSR03]. act [GSB^{+}06, HG08, HDQ04]. Actinia [TLP^{+}03]. Actiniaria [TLP^{+}03]. action [BBM^{+}02a, BOR08, ZMN07]. actions [PKM03]. Activated [CH00a, VWF^{+}01, PFP03, VH03]. Active [CNC00, GMVF08, HMH^{+}08, PC02a, SS02a, BBDW09, BSP03, CB00b, DBvOM06, TBVJ08]. actively [UÓ03]. activities [AG00, AWA^{+}05, CAA00, DAMJ08, EP00a, HW00, Kar07, PPI04, RL09b, SVC^{+}01, VDA^{+}03]. activity [AH07, AC02, AJ03, AFHA05, BS04c, CHSW03, CHT06, CPP^{+}04, CC03b, DDQ04, DR01, FB03a, FB03b, GBMHD08, GBG00a, GUT04, HHB^{+}00, HSCH07, HM05c, HMO^{+}04, IRL^{+}09, JAA^{+}01, KKB^{+}06, LKDO07, LS05, LYL^{+}07, LP01b, LSG05b, MDBG07, MPK03, MMBT03, MG09b, MLG03b, NW06, OHL^{+}06, PWM^{+}08, PCS^{+}07, QDDP06, RWPJ04, SP03a, SFHS00, SW08a, SN04, SAMN04, SML05, SWEBS00, TWYS03, TPQ^{+}03, TRS^{+}06, ULL^{+}05, WTF^{+}05, YYK^{+}08, YM09, YLv^{+}05, ZK01]. actual [BDC00a, Bar02c]. aculeata [DHR04a]. aculeatus [JHJ^{+}03, JHJS04]. acuminata [RH09, SPP^{+}09]. Acuña [Ano00c, Ano00b]. acuta [PEGG^{+}08]. acute [TGM07]. acutely [NSFW09]. acutorostrata [ABD^{+}03b, JTR05, MFG^{+}04]. acutus [STA^{+}05, TSW^{+}04]. ad [SNP03]. adaptation [FA08, HR02a, MND^{+}02, RCC^{+}04, SP02, UP07, vdVBW00]. adaptations [BAM^{+}04, DG05, RBG04, SP01c]. Adaptive [JS08a, Kaw01, SL05d, BGD^{+}03, Hid03]. added [NFV^{+}06, dBGLT08]. addition [CF02, GFF08, HJ07, JHTV02, LFSR06, MRP09, SN04, SCD^{+}09, SBC^{+}06, Sve02]. additions [FBKB03, GSW01a, MK03c]. Additive [Jos09, JSMB05, SMP01]. address [BWRP06, KBSWG09]. addressing [FGA^{+}04]. adductor [BG04c]. adducts [ABB^{+}00]. Adelaide [BB03a, SBFP08]. adeliae [CKF^{+}03, TSEH08]. Adélie [CKF^{+}03, CEO06, MHB^{+}09, TSEH08, WAN^{+}01]. adherens [CÁ04]. adipose [GNAD^{+}03, TBBI04]. adjacent [BIL^{+}02, CCML07, CHG05, DSF^{+}04, DGC^{+}05b, HB07a, LSG08, NvdV02, NvdV04a, PCSS01, PJJ^{+}05, RZ00, SWL^{+}06, YQC^{+}00]. adjoining [CRS^{+}07, GQR06, HCR01]. Adriatic [UB03, BTB^{+}09, CSPA09, CGH^{+}03, DSDS04, FGPS04, IRL^{+}09, MFR01, SKBP09, SPC^{+}00]. adsorption [EP01]. adspersus [NMJ02]. adult [BCB08a, EM05, FKSW09, FS07, Gil02a, GAB^{+}03, HANM^{+}05, HTM^{+}02, HM01, HMO^{+}04, LAK07, LTCP03, NFN^{+}09, NA04b, PBF07, RBHS06, TCRT^{+}07, TWMW06, TLC06, Ver05, AM04]. adult-juvenile [PBF07]. adults [BCB^{+}08b, CSTC01, EM04, Gil02a, Gil02b, TLLCT02]. Advances [GNPB07, KJRG09, MNT09]. advantage [MP05]. advantages [FMW07, GFM^{+}02, WM02b]. Advection [LF08, FBJ00b, JBH00, Kap06, KJMP05, NNLV06, PT04, SWD^{+}02, WM06]. advection-diffusion [SWD^{+}02]. advective [BP06, EWBH04, SWD^{+}02, VPSO01]. advise [WGCM^{+}03]. Aegean [RFLK03, SMA^{+}03, SHT^{+}04b]. aeglefinus [BW01, Cal05, LB05a, MGSJ08, PBCB03, RTL05]. Aequipecten [KMHS04]. Aequorea [MR07a, MR07b, MR07c, Pur03]. aerial [BVMBC07, CCB^{+}09a, FNB07, YG09]. aerobic [DER01, MCN09, RL09b, WBW^{+}06]. aestuarii [Mis04]. aff. [MC05b]. affect [BPB07, BvK07, BK06a, BBST07, DGdG^{+}09, FHVMDS07, HRR^{+}07, LHQ05a, LZW^{+}08, LÁMBF09, MVN^{+}06, MBHN07, MPK03, NL05, SKU^{+}08, TN04, VFM06, jMFVPL05]. affected [AGCF^{+}08, FHPW07, SJPC02a, TMFF00, WO05]. affecting [AS06b, CWCP^{+}06, CGM00, CDML06, DCD^{+}03, PBBC06, SFL^{+}09]. affects [BR09, Bul06, CRSL05, CLL^{+}09, CTB07, DQ05, DHT^{+}08, GG09, HHF^{+}07, HCB^{+}07, HLTT09, MSK00a, MMC^{+}07, MFA03, MAW06, MM08, MB05b, OLG^{+}09, RGC05, SBR09, SKK^{+}06, SLB^{+}08, WKH07, WCZ^{+}09, vdMPEM08]. affiliation [Bat09]. affinis [BEE06, EEA01, FVAB00, Kar07, KSKL^{+}03, LG08a, NWÓ03, SKK^{+}06, SKU^{+}08, WÓ02, WS01, vdBÓME01]. affinities [BBG^{+}04, SC03]. AFLP [MS03]. Africa [KHT^{+}02, Ano01a, EB04b, ESB^{+}04, HMK06, KPN03, Kor01, MP00, SPP^{+}09, XBW07, ZMTB07]. African [BA00, GB02b, NPM01, PBT^{+}09, RMC^{+}05, RF08b, RRDHG08, RBG^{+}07, Ryp08, SSB02, SAMN04, TMFB06]. africana [NPVW06]. after [AHDT^{+}09, Bai00, BVC^{+}02, BJN^{+}05, BB02a, Boo02, CCJT09, DWB04, FRJ09, GdLB^{+}09, GSC^{+}01, HR08, IG04, KHIA02, LS08a, MAVO^{+}06, NAB^{+}06, PKR07, PMGE01, PD01, San00, SB03a, SAC^{+}03, Str02, SAM^{+}06, TGGA04, VJ03, Yap09]. Ag [GFL02, LLY^{+}01, LW05, NW04, RSL09a, YLL^{+}04]. against [AE06, BP06, DDHQ06, FI09, GHdN07, LS09a, NDB04, NCP^{+}08, OP02, PCL^{+}09, PdNS06a, PLC^{+}02, PW07, PLG09, RWP07, RB08, VBAR07, WP00a, WP00b, Win09]. Agaricia [PV05, RPS^{+}04, ZE01]. agaricites [RPS^{+}04]. Age [CDAM06, GJB04, HYR^{+}09, LGB03a, LGB03b, LR06, LDR00, MHB01a, RSSA07, APG^{+}09, ASLD09, BRI^{+}07, BGSBC02, BM00, BK06a, CW03a, CLN00, CR07a, CRAP03, CPG^{+}09, Enb05, FHB^{+}02, FGSH^{+}05, GSM06, GBW02, HSJ08, HMY^{+}08, HDP08, JSWH03, KS08, KKL07, KdLCM09, LSP08, LTFH06, MB02b, MGW^{+}09, MAW01, MF07, OTBR00, PBS^{+}00, RT00a, RH08, She08, SMGCW00, SAP02, WDT^{+}08a, WIB08]. age- [WIB08]. Age-0 [LGB03a, LGB03b, LR06, RSSA07, CPG^{+}09, PBS^{+}00, RH08]. age-based [CRAP03]. age-class [BRI^{+}07]. Age-dependent [MHB01a]. age-selective [Enb05]. age-stage [HDP08]. age-structure [CW03a]. aged [KM03b]. ageing [PBH^{+}06, PTP01]. Agelas [ALPK00]. agent [HSCD01, Ryp08, WNWH02, AHDT^{+}09]. agents [Bel08b, TSF04]. ages [RGDI06]. agglutinated [GPB02]. aggregate [BGH06, FRJ09, HGKJ01, NB06a]. aggregated [BJLH04]. aggregates [GHP03, HWG06, KPT01, KT01, MGRP07]. aggregating [GTD^{+}08, JLW04, MVD^{+}08b]. aggregation [AP06, BPE^{+}06, CHW06, DK03, DNL08b, FK00, Gla07, GGMGR07, HG08, Hen09, KBSWG09, LGB^{+}04, LGS07, LWH08, MRS00, MY04, MDHL08, Nem05, PBS^{+}00, SML^{+}09, SPL00, SSS06, UDB^{+}09, WWHE04]. Aggregations [PKA06, PBS^{+}00, BUF02, BHD04, DUSJ^{+}09, GLG^{+}05, HFT^{+}09, HDDH06, IMW^{+}02, LS04, MC07, MWG02, PCCL05, PCB08, Zam03a]. aggregative [FE02, HDG^{+}08]. Aggression [TLP^{+}03, SKP^{+}07]. agonistic [MAH^{+}08]. agulhensis [RVMI^{+}03]. aided [RDG^{+}07, RGD^{+}08]. aids [Bye00, RB05b]. Aiguillon [HGRB04]. Ainigmaptilon [OLGG^{+}02]. air [CDW^{+}09, HDDH06, QWY07, SFF^{+}07]. air-breathing [CDW^{+}09]. airborne [APH^{+}04, JCMR^{+}03]. akashiwo [BGC06, JKK^{+}05, YNM^{+}07]. al [Ano01d, Ano01e, Ano01f, Ano02c, Ano02d, Ano02f, Ano02h, Ano03c, Ano03b, Ano03d, Ano03e, Ano03f, Ano04c, Ano04e, Ano05c, Ano05d, Ano06, Ano07c, Ano07e, Ano07f, Ano07h, Ano10b, CPB04b, GHP02, GNAD^{+}04, Ike08, KH07c, Pea03, TBBI04, vWNYS05]. al. [BBE03, BD11, BL07, CST08, LTMF01, MMPD09, RN06, She08]. alalunga [ACJ^{+}04]. alarms [CKTE01]. alascanus [RM09a]. Alaska [AP05a, JTN09, PBS^{+}00, BP02a, BRI^{+}07, DABC02, GRRH07, GSM^{+}02, HE08, HCBP05, IFL02, Kon00, KG07, MHSS08, NEK^{+}07, PS01, Pur03, Ran02, SO01b, SOMM06, SH00, WISH09, Wil09, WJB09, WTR02, WWS^{+}05]. Alaskan [PCH^{+}05]. albacore [ACJ^{+}04]. Albatross [Hyr01, PWC^{+}09, CWT00, HGB01, NLP^{+}01, PPS^{+}07, PSC05, HGB01, JRP01]. albatrosses [AFHA05, CPP^{+}04, GWWC00, HFA02, NRP^{+}02, PCW05]. Alboran [FC04, ME01]. Albula [HALL05]. Alcanivorax [AGCF^{+}08]. Alcian [EP01]. alcids [DHHY09]. alcohols [SKFPS02]. Aldehyde [FI09, CBO04]. Aldehyde-induced [FI09]. aldehydes [PLC^{+}02]. Alderia [BK06a, Kru01]. Aleutian [WCH09, HSS09, RHPB08]. aleuticus [AS04, AS06a]. Alexandrium [AKRA02, BGF^{+}06, CVF^{+}03, CCM01, CD03c, FLR05, FGM^{+}00, GFM^{+}02, ITE^{+}04, KS06, SPP^{+}09, SSO^{+}01, TCD01, TJ02, VGMC01]. alga [BCBVM05, BCBVM06, BMC^{+}09, BMS04, BS04a, BNP07, BBC08, CÁ04, CCL^{+}04, CB05b, CK09b, DEAI06, DdNS06, HDW02, JKK^{+}05, JM03a, Kam00, KLFH04, MS03, NDB04, NB06b, NCHP05, NP05, NCP^{+}08, PdNS06a, PB00a, PCT00, PCC01, PVGO01, PWdN^{+}00, RK08, SG06b, SRS05, SS05, TG03, TJA08, WdNS00, YM09]. alga- [PWdN^{+}00]. Algae [JB07a, Air00, BP02b, BCPB^{+}01, BC01a, BBM02b, Bul06, CD07a, CNPS05, CBMB03, CHG05, CHV^{+}09, EM04, FBKB03, FK04, GDS02, HP06, HK02a, HCV03a, HSW^{+}08, Her06, IOB04, IC06, JM03a, KSS^{+}08, KSS^{+}05, KGB^{+}01, LJ08, MCCV08, MACS07, MRP09, MBH^{+}00, MAJ^{+}06, MS01b, NWHH01, NHN^{+}06, OTB^{+}00, PdNS06b, RKGH01, SAHF03, STRD02, SGV^{+}02, SHA^{+}05, SJK^{+}06, SSSV00, TSJH05, VSL^{+}04, VAB09, WW06a, WWV08, vDSD02]. algae-epiphyte-seagrass [IOB04]. Algal [GLJ07, YBY^{+}09, AGB03, Arr05, BGJ^{+}07, BGC06, BAMJ09, BBC08, Bur02, CEF04, CSG01, DPM02, GWH03, GMC^{+}09, GP01, GC03, HJ03, HDQ04, HK03a, HSD03, IC02, JM03b, KCS04, KSIG08, KGRW06, KB04, KB05, LRHV07, LPK^{+}07, MSS^{+}03, MRB09, MG00, PFW^{+}02, PB00b, RHI^{+}01, RdNS02, SS09a, SASR03, SL07, SAC^{+}00, SZZ08, SCM09b, SLHT05, TLTB03, TSJH05, TH06, WG01, WGP^{+}00, Zhu06]. algicida [JKK^{+}05, JKS^{+}07]. Algicidal [SBH^{+}02]. algorithm [BT04]. algorithms [PBL^{+}09]. algosus [CC04]. alien [BOR08, CLC04, CD07b, SSB02]. align [BMdNP09]. Aliphatic [GGGVGG03]. Alkaline [AC02, IRL^{+}09, SN04, VDA^{+}03, FB03b, OHL^{+}06, SAMN04]. alkylpyrrole [KKBH03]. Allee [GL04a, GL04b, HJM07, SRC00]. allele [CHH^{+}05, PW07]. Allelochemical [DDHQ06, TJ02]. allelochemicals [SFG05]. Allelopathic [EP00a, FLG03, YNM^{+}07, LHK05]. Allelopathy [SH01, DDQ04, HP06, ULL^{+}05]. alleviation [CGP08]. allocation [BD00, DZ06, FBBH07, KS08, KHF^{+}00, LD00, MVHV06, MG00, PUB^{+}01, RHG04, SAJG06, TBAT02]. Allochthonous [CHV^{+}09, Hor06, SR09]. allometric [VFF09]. Allometry [PW08, ABBF01]. allopatric [LFP^{+}04, LQRGT06]. allorecognition [NL05]. allows [BGGR02, LPK^{+}03, TBAT02]. alloxanthin [AG04]. Allozyme [TA01a, CNT^{+}04, EGPT03, VBA06]. Allozymes [GPK^{+}04]. Almeria [VLP^{+}04]. Almeria-Oran [VLP^{+}04]. Along [ANCC07, APA07, AI04, ABD^{+}02, BAM^{+}04, BA01, BJ08, BFK^{+}05b, BAB04, BRS07, BFE^{+}03, BES^{+}08, CCR08, CL05, CLMR07, CLRM08, CH07d, CM08, CAA00, CV05, DCAL^{+}02, DGJ^{+}03, DCF^{+}05, DFT^{+}07, DGNvdV05, DESK03, EMB^{+}09, FZT^{+}05, FS07, GhGD^{+}02, GRPE08, GVF^{+}03, GFP^{+}05, GUL^{+}07, HKZ04, HM05c, HDCA06, ILL06, JMM^{+}09, JNMB08, Joh02, KTK08b, KNU02, KUN02, KCEB09, LFRG^{+}05, LNV^{+}05, LTNC07, MCCV08, MTL^{+}03, MVBGM04, MDBC03, MdL06, MLW03, MLG03b, NvdVVD06, PCSS01, PMM^{+}00, PCD04, PPL^{+}02, RWA^{+}02, RGMJM07, RHPB08, RBG01, SSGB^{+}08, SRGS01, SFKT09, SS03, SLM07, TRDV00, TSMC09, TNSM00a, UB01, Uth06, VHB^{+}07, VGMC01, WGV^{+}02, YI07, YMHV02, vWCAB07]. Along-shore [ANCC07]. Alongshore [Kap06, WKN^{+}03, PW07]. Alopias [NMOO03]. Alpha [DFV01]. alphabet [BEKT09]. alpina [LPM^{+}06]. also [VGW00]. alter [JFD09, PFT^{+}08, SCD^{+}09]. alterations [VMV^{+}09]. altered [BG04c, DH09a]. altering [JQ05, MHT07]. Alternate [MVK^{+}09, BGDS05, GBR01, RC08]. alternating [OHS^{+}07, TIM^{+}01]. Alternative [NNLF09, BBNK02, BBM02b, DJM^{+}05, FSB07, IMS05, MC05a, SBJ06, Wen00, WDR^{+}05]. alternatives [MD07]. alterniflora [CLH^{+}07, CS03c, HH05a, MJG01, PMH08, PAMJ03, WWAC00]. alters [FTP^{+}09, FSW09, FIS04, HJ07, RC05, TSSQ06, THJK06]. alveolata [DBCB05, DBCB06, DCRT07, RN06]. Alvinella [FCP^{+}07, PZMG05]. alvinellid [LJM03]. Alvinellidae [CTVP03, FCP^{+}07, PZMG05]. Alvinocarididae [CY06]. Alvinocaris [CY06]. always [HMJ^{+}09, MHM^{+}08, RFLC07, SBJ02]. Am [VF09]. Amakusa [NTN08].

Amazon [BBBSPH03, DL01]. Ambassis [NWT00]. Ambient [TJM00, WS09, EG05, FBBV09, Hil09, KNK^{+}00, LC03, LGVF00, PAW00a]. ambiguities [MSF04]. Ambiguous [KLFH04]. ambush [RCCK09]. ambush-foraging [RCCK09]. ameliorate [KHGA02]. ameliorates [KF01]. amelioration [MHT07]. America [BM07c, DC02, JBCL03, WGV^{+}02]. American [BH09a, BHH07, DCA^{+}09, GdLB^{+}09, JPC00, JSIT02, JSIT04, JSIT06, JSIT07, LGD04, MGC03, RBSP00, STAS06, SM01b, TB08, TDC^{+}07, Thi05, TML03, VGB^{+}05, WA00]. americana [EMPF07, KS04b, NH03a, NH03b, SR08b]. americanus [BWCW02, BHH07, GSM06, JPC00, KMAC^{+}02, MPM^{+}02, OA02, RBSP00, TC03a, Tay03, Tay05, WA00]. Amery [PHO^{+}07]. Amino [JQ05, CCM01, CKO^{+}07, GKM02, GRM00, GMR^{+}02, ITN^{+}06, JQ04, LBR04, LGM^{+}09, MGA^{+}01, RGF^{+}03, SMM^{+}04, SFPGA05, TMM05, VK02]. Ammodytes [EvdKB^{+}08, FHN^{+}09, HGG^{+}05, MCW^{+}07, WWHE04, vDvHT^{+}09]. ammonia [BCM^{+}09, NBB06]. Ammonium [Cam01, MFE^{+}05, SSH09, TC03b, AG02b, BP01, BHV^{+}02, BOM^{+}08, BL01, CIFN00, CT06b, CT09, DGW01, KGTH01, Kev01, ORPND01, SFPGA05, TRPNE06]. Amoco [PD01]. Amoebophrya [PCYC02]. among [AGCF^{+}08, BK04, BJC02, BG04b, BHH07, Chi09, CK09b, DTAF00, GGS09, GGFSP06, GC06b, GRM00, HH00a, IFL02, JKS^{+}01, KCS04, Kru01, LJM03, LSS^{+}09, MFCM05, MH09, MB07a, MACS07, MS01a, NMBH02, NMJ02, OPSB07, OGL09b, PMG05a, PMS^{+}06, PDS^{+}07, PWdN^{+}00, PCL08, PMV06, PS01, PCRR08, RSL09b, RZ00, SAHF03, ST06a, She06, SJ09a, SM09, SC07, SBF08, SSW08, SAG03, SM05, TRDV00, Tit01, UDB^{+}09, WFT^{+}08, ZE01]. Ampelisca [PD01]. Amphibalanus [MMA^{+}09]. amphibious [CCB^{+}09a]. Amphibolis [LV02, LMMT09, MCL07, SC01]. Amphidinium [JKS^{+}01]. amphidromous [HLFG^{+}07, ST02]. Amphipholis [BHFC08, HSD03, LFP^{+}04, SDR01]. Amphipod [BCL^{+}01, CRH01, DG05, DJM^{+}05, EEA01, FDDS09, ITN^{+}06, KSS^{+}05, Kum06, LH09, LP01a, MHCB08, MvdBÓ00, MR06c, NWÓ03, OAM02, PCT00, Pol00, PR01, SV07, Sot03, WÓ02, YSS03]. Amphipod-based [BCL^{+}01]. Amphipoda [JYS06, MS00a, PMM^{+}00, PD01, RJ03]. amphipods [AK02, BEE06, DH01, NBDG05, PWdN^{+}00, SCH^{+}07, WA05, WS01, WO05, YI03a, YI06, vdBÓME01]. Amphiprion [SYWM05]. amphitrite [DP09, GQ00, HMR00, HTZ^{+}07, MMA^{+}09, MWTB00, OTBR00, PA05, RBW03, THC^{+}05, TOBR07]. Amphiura [OKP06]. Ampithoe [PMM^{+}00, Sot03]. amplified [CRC02, MS03, SAG03]. amplitude [KGB^{+}01, MS00b]. Amsterdam [BDG^{+}04, GBG00a]. amurensis [HRDA05, Kim06, Pen00, RJH02]. Anaerobic [TBK^{+}03, PMT^{+}02, RL09b, TRPNE06, VNJ06, WBW^{+}06]. analis [DL02]. analoga [DTB07]. analogues [EOL^{+}05]. analyse [HMHB03b]. analysed [MBR^{+}05, NR00]. analyses [AKRA02, ACJ^{+}04, BDR06, BDC00b, CDK^{+}07, EDBvdL08, GVKF03, GFN^{+}09, HHF^{+}07, HTM^{+}02, HSS^{+}04, HBW^{+}05, Hir06, HH00a, HD07b, ISK07, JBS08, JPPW02, JWM02, KKL^{+}03, KKIV01, LCGN02, MS01b, PJN01, PJN03, SJ06, STM06, TBCB^{+}05]. Analysing [RFSC08]. Analysis [BSP^{+}08, CG05, LW04, AZKJAM06, AYSL09, AL03a, AGM06, BCL03, BCM^{+}07, BMW^{+}00, BJR04, BCC08, BLH02, CHLV07, CF02, CS02a, COB^{+}03, CRAP03, CES^{+}09, CRC02, CVSM09, CNT^{+}04, CWB^{+}06, DSS^{+}09, DM05b, DJM^{+}05, DHZ02, DCA^{+}09, ESS09, FHB^{+}02, Fre03, GVM^{+}05, GFB^{+}07, GY00, GMM03, GDR^{+}04a, GPM^{+}09, GCM01, HARC05, HM06a, HSC^{+}09, HWvdM03, HEN^{+}08, HBRW04, JAA^{+}01, JF07, KPM00, KCH^{+}06, KSHP00, KMB^{+}08, KMK01, LEF^{+}06, LNB^{+}03, LSP08, LHK01, LPBD00, LWB^{+}00, LBH01, MDBG07, MB05a, MHH^{+}05, MRQM06, MS05, MAWH03, NSEL00, NEK^{+}07, NEMF09, NLCS05, NWIT08, NBDG05, OW02, PKLD09, QFJ^{+}07, QMM^{+}08, RSR02, RL05a, RFSC08, RMAY04, ROBFR08, SFW01, SO08, SPI^{+}07, Sha00, Sha09b, SW02b, SW02a, SGCD01, SAG03, SMP01, SPGS03, SKO^{+}08, TK08, TNHL02, THHY05, TSEH08, TTN01, TA01a, TCWR08]. analysis [TBI^{+}09, UG04, UYNO01, VCA^{+}09, WR07, WJDA06, WWR09, dlMPN^{+}03, GCSSJV07]. Analytical [TBH^{+}09]. analyzed [KH00a]. analyzing [CG05]. anammox [TRPNE06]. Anchoa [JH04, KSHP00, Lap01a, Lap01b, MHHS06, Pee02, SBJ02]. anchoveta [CSH00, CCK^{+}09, KCG07]. anchovies [Kim06]. anchovy [CGH^{+}03, DTV08, HFMP03, JH04, KBB09, KSHP00, Lap01a, Lap01b, MHHS06, OSF08, ONKK03, OSP01, Pee02, SBJ02, TW04a, TW04b, TAM03, TOK^{+}04, TAM04, TAO07, TKO08, WHM^{+}01, WIDM09]. Ancient [RP02]. Andaman [MBB^{+}04, NBB^{+}04, SNH04]. Andrews [SAH09]. anemone [DKR09, GBI^{+}06, MSC04, TLP^{+}03, WG02c]. anemonefishes [GBI^{+}06]. anemones [HF03, SGV^{+}02]. angelfishes [EJM01]. angiosperm [AAHDS06, FJV^{+}02, SGB09, vDvTJD^{+}09, vTWMG08]. angiosperms [KHB00b]. angle [YA09]. anglica [HGK02, WPB08a]. angling [WMB00]. Angola [KT03]. Anguilla [ALOT01, AMT03b, AKL^{+}04, DLT^{+}06, DT06, EDB^{+}04, EDE05, EBLE07, GMA^{+}09, HYLT03, HT07, JT02, JT05, JSIT02, JSIT04, JSIT06, JSIT07, KYML02, KOY^{+}05, MPR^{+}06, OYT^{+}02, PMV06, RLFEF03, SAW^{+}01, STCJ01, SICT03, SAM^{+}01, SSM^{+}01, TTL06, TA01b, TSI02, Wes03, YOM^{+}09]. anguillid [KAM^{+}06, TCW^{+}07]. angulata [HFM^{+}04]. angustirostris [GAMB05]. Anilocra [ÖNCNG05]. animal [CG05, SFH^{+}00a, SM00a, SK02]. animal-community [SFH^{+}00a]. animal-sediment [SK02]. animals [BKP07, HZ04, TAS07, TCS^{+}05, Tya09]. Anisomysis [JR01]. annelid [SJS^{+}01]. Annelida [Iye02, MJG^{+}00, MG05b]. annelids [CC07]. Annella [PPS00]. Annual [AMR01a, BD06, CID^{+}06, HHG^{+}07, LBBH08, LMD^{+}02, LNH00, MHABGC04, NPM01, AS04, BM09, CW03a, CB01, FEMW07, HGS^{+}02, HRCB02, Hip09, HAD07, IMTW09, JNS04, JSIT04, KHH08, LG00a, LCL^{+}03, LCGN02, MRQ07, SBR09, SSRG05, SMSD07, SMG00, TM02b, WM04, WP03, WAB^{+}02]. annularis [FBM08, MW02a, MW02b, PB08a]. anomalies [COPB06, EBR^{+}02, LGC^{+}06]. Anomalocera [TCC04]. anomalous [MBG^{+}08]. anomalously [BWL06]. Anomura [Don04]. anophagefferens [BMS04, BM07b, CCL^{+}04, GLCL04, SLT^{+}01, TGSW06]. Anoplopoma [SO01a]. Anotrichium [JM03a]. ANOVA [dM01a]. anoxia [DP09, DEvdZ03, ISG04, TW00]. Anoxic [HTJ00, MHK^{+}04, CGCMPA01, DER01, dZCVC01]. Ant [PMPI03]. Antarctic [FHQ^{+}06, HVCT09, AD00, ASP^{+}03, AHC^{+}06, AVNN05a, AVNN05b, AIM^{+}05, AAJ07, ASP^{+}08, BA01, BWL06, BGK09, BND02, BKH^{+}09, BLDG01, Bow05, BCPB06, BSM02, BW00, BP01, BPT07, BWvdE^{+}06, CWC00, CBDH02, Chi09, Cla08, CDW^{+}09, CTB07, DL03b, FHM02, FMWGB08, FNDP08, GVKF03, GVF^{+}03, GRP^{+}06, GGS09, GKWK02, GTPC04, GDL^{+}01, GP03, HRQ^{+}02, HRR^{+}07, HWB^{+}09, HDDR08, HH00a, HL00, HHAH00, HHKH05, HKSW01, IBB^{+}07, JT08, KPKD06, Kas00, KTH^{+}04, KCK^{+}06, KTK^{+}02, KR01, KS00a, LCGN02, LGC^{+}08, LC07b, LC09a, LAG08, MVHV06, MRQM06, MAH^{+}08, MGC^{+}09, MABB03, MO05, MSD05, MGPAE01, NBDG05, OLGG^{+}02, OGA03, PTK^{+}09, PCC06, PWM^{+}08, PAM^{+}09, PTP01, PBS^{+}05, QR03, SAB^{+}05, SDPT06, SN02, STA^{+}05, Sma07, SRB04, SRS04, SPGS03, SZL^{+}05, TDT^{+}03, TCRT^{+}07, TGA02, TGGA04, TR05, Thu07, TYC00, VBV^{+}00]. Antarctic [WBL^{+}06, WAW^{+}09, WCM08, WAN^{+}01, vLvSGS05]. Antarctica [ASP^{+}03, AWP^{+}01, BB03a, CIHF01, CKF^{+}03, EBA^{+}03, FCP02b, HLPHA^{+}01, INH^{+}09, KKLL02, KKKK03, LBvdH03, LBBH08, MHB^{+}09, PHO^{+}07, RHHH09, STE04a, SHA^{+}05, SBFP08, SKE05, Sta08, TDT^{+}03, VAP^{+}07, WBS^{+}07, WAN^{+}01, WRW03, AMR00b, AMR00c, HRQ^{+}02, OGA03, SDPT06, SC01, SLKF03, SNGR00, TS05]. antarcticum [MVHV06, OLGG^{+}02]. antennular [MCM05]. antennule [FSW02]. anthesis [vTWMG08]. Anthophyta [RZEACGF00]. Anthoptilum [PCS09]. Anthozoa [FGO09, TLP^{+}03]. anthozoans [JLJ01, PLM03]. Anthropogenic [Hil09, PUL07, BSW^{+}09, BCPB^{+}01, CBP^{+}05, Fro00, GMTN00, JBG00, JDTB06, LWK^{+}07, MPB^{+}09, NMJ02, OGL^{+}09a, PBF08, SCS^{+}07, SCAF03, Tya09]. anthropogenically [GB03, GCU07]. Anti [CYC^{+}09, HLY05, WTP08, BIC^{+}03, BI03, FK08, MKKK08, PCL^{+}09, RST04]. Anti-grazing [WTP08]. anti-herbivore [PCL^{+}09]. anti-herbivory [MKKK08]. Anti-predator [CYC^{+}09, HLY05, FK08, RST04]. anti-predatory [BIC^{+}03, BI03]. Antibacterial [GA08b]. antibodies [PHH^{+}05]. anticipatory [PWC^{+}09]. antifoulants [HDQ04]. Antifouling [QDDP06, DBD^{+}05, DdNS06, GWBdN09, LYL^{+}07, NGdN^{+}07, RGEB00]. Antilarval [DDQ04]. antillarum [MRN07, MAG07b, TBSJ^{+}04, THBSJ05]. Antilles [NvdV02]. antimicrobial [DDQ04]. antioxidant [CNPS05, CSvO09, PEG^{+}00, RCC^{+}04]. antioxidative [Rij03]. antipatharians [RSNG09, TPBF09]. Antipredatory [EMPF07]. antiqua [SJ02]. Antofagasta [APM^{+}01]. any [RZLS09]. Apalachicola [MIL^{+}00, MIH^{+}00, MILH00, MIH01, PI07a]. apart [BRSM07]. aperiodic [Sha00]. apex [DFH05, FD02, VACT^{+}08]. Aphanius [Mal02]. Aphanizomenon [RAE07]. aphotic [HJ03]. Apical [SZ08a]. apiculata [ATTX00]. Aplacophora [KCB06a]. aplanic [HDW04]. Aplysia [GP01, RdNS02]. Aplysina [PFP03]. Apo [RAM03]. apocryphal [RJS04]. Apogon [FOH^{+}08]. apogonae [ÖNCNG05]. Apogonidae [GJ05, MB02a]. apoptotic [MM05]. apostome [OHS^{+}04]. apparent [CDO08, FCH^{+}09, GPSG06]. Appearance [KNT^{+}07]. appendage [PL00]. appendages [KPW03]. appendicularian [LSG05b, LUA99, LUIAH03, SVL^{+}04, TCG03, TGNT05]. appendicularians [FLUF^{+}04, LUAIH03, STI03, SLC05, VTS^{+}02]. apple [Swa04]. Applicability [GMP07, Osb07]. applicable [CW01]. Application [CCL^{+}08, EB05, Kam00, KS05, OW02, RDG^{+}07, ARP02, BDGS07, BMT07, CH07b, CCB06, DK06, FGS^{+}02, HRC^{+}07, HCD^{+}09, JSMB05, Joh04, JH04, LALB08, NGBG07, PKS07, RM09a]. Applications [Lóp07, DKA^{+}06, JDVG08, LNV^{+}03, VCSL^{+}09]. applied [ATT06b, CD05, EOK^{+}08, FCC^{+}08, PTS00]. Applying [NZ08]. appraisal [Rue01]. approach [APG^{+}09, Bea04, BAB^{+}06, CLABRO^{+}03, CES^{+}03, CM04b, CRCN07, EvBO^{+}08, FCC^{+}08, FEFT08, FGS^{+}02, FBT^{+}01, FBN^{+}07, GTD^{+}08, GÁSPR03, GN04, GDF03, GG09, GGGVGG03, GDL^{+}01, HRQ^{+}02, HHH02, HPBT03, HBTD04, HK04b, HBRW04, KHD07, KHIH08, Kra07, LFRG^{+}05, LPD09, LR08, LPM^{+}06, MRQM06, MS01a, MLS^{+}06, MDS^{+}07, NHS^{+}08, NRC07, OT08, PGCC^{+}06, PCS^{+}04a, PNCS^{+}08, PNCT^{+}08, PSRB00, PMP04, RMM04, RF08b, SDFM00, SM00a, SYVC00, SZ09, TSW^{+}04, TH06, WDB^{+}05, ZAH03]. approaches [BRF03, BS04b, SAG08, SRP^{+}06, TBH^{+}09, dlSPLV09]. Appropriate [NNG03]. April [MLG^{+}03a]. Aptenodytes [HHF^{+}07, MHB^{+}09, RBMB00]. Aqaba [AQHTL02, ARAMPW02, DH07b, KHGA02, KK02, LKW03, PDG^{+}02, RBRH02, BTKG^{+}08, GS02, SBBS^{+}02]. aquaculture [BPN^{+}09, CSD^{+}07, DAMJ08, DSJT^{+}06, HDC^{+}08, KEB07, LBB08, NLB08, PPSP07, WDRH07, ZHWCS08]. aquaedulcis [BCB^{+}08b]. aquarium [SBSPS04, ZM06]. aquarium-traded [ZM06]. Aquatic [NCC^{+}07, AG00, AB04, BRJ^{+}03, CH06c, CKO^{+}07, MMB03, SB05b, Via07, YG09, vDSD02]. aqueous [BFK06, Wan02a]. aquifer [TACM01]. Arabia [RSNG09]. Arabian [AHSPG08, HGAS^{+}01, NSJ02, PR05, SDW^{+}01]. aragonite [Tak02, TTM^{+}05]. Arbacia [CL05]. arborea [RC04b]. arboreum [TAS04]. ARC [EKN08]. Arcachon [DAM^{+}08]. Archaea [Kel04, WYB^{+}04]. Archaeomysis [TK04]. archaic [BDDL03]. Archipelago [AAPO00, BKSS07, CR07b, LPBD00, SHS08, dVCH^{+}06, FCS^{+}09, WJ06, Zam03a, AMR00c, HSS09, LGC^{+}08, LB05b, MW08, MSSP06, NM09, PR09, RHPB08, SC03, TBSJ^{+}04, TH06, WCZ^{+}09]. Archipelago-wide [SHS08]. architectural [BB02b]. architecture [ITN09, PBPLB06, YAF^{+}07]. archival [GNL08, JT02, LS06, LALB08, NS02, SAS^{+}06, SSR^{+}03, VCSL^{+}09]. archive [SHS^{+}05]. archiving [CG05]. Arctic [BSA^{+}08, FA08, JPMGS08, KVV^{+}00, QS05, AJ03, BFK05a, BFK06, BK04, BSKR^{+}06, BRS07, BSR09, BGS02a, BSIS07, CK05, CAA07, DFPG^{+}03, FHG^{+}03, FFML02, GGK^{+}07, GRPT^{+}00, GKWR02, GKB09, GKWK02, JGH05, LTNH00, MNTS08, MGF^{+}03, MRG08, MG00, MSJ^{+}09, NCMC09, NMS^{+}02, Pol00, PKA06, QASA08, RKGH01, SDS^{+}03, SKFPS02, SSJ^{+}02, SBR09, STH^{+}06, VMdBV06, VNJ06, VFJ06, VJKF07, WA05, WKPK05, ZZT05]. Arctica [HWvdM03, Bar02b, CM01, HANM^{+}05, SSJ^{+}02]. Arcto [Hel00]. Arctocephalus [BDG^{+}04, BKH^{+}09, BLDG01, GdLB^{+}09, GDL^{+}01, SRB04, TML03]. Area [CMD^{+}09, OT08, SHMN07, WWMF06, AP05b, BT09, BCL03, BSVM^{+}05, BHN09, CBAd00, Chi02, COSM03, DMH01, DNL08a, DSDS04, FF02, GSRH02, GSFN00, HPH^{+}01, HANM^{+}05, IG04, JLJ01, KMHS04, KAR04, KHvW02, KNF^{+}03, LGB03a, LAV04, NZ08, RLBB04, RGS^{+}09, RGCL^{+}05, SSB02, SRN02, SH06b, SWG^{+}09, YHSH02]. areal [MMA03a]. areas [BGQR01, BPG08, Ber00, CH07c, CSS00, CWDG05, CTFC08, DDQ^{+}00, DSF^{+}04, ESS09, FVB^{+}09, FBJ00b, FDM^{+}09, FBM07, FT00, GAAB^{+}08, GCSSJV07, HSWG^{+}06, HF09, HLH^{+}07, HCS07b, HFH09, IR02, JBH00, Kap09, KS04a, KH07b, KSWP04, KGF^{+}09, LHJD^{+}03, LH08, LBR^{+}09, LW04, MC01a, MJS07, MTR^{+}07, MG05a, MS00b, PH05, PvDMB08, PRL^{+}09, QMF05, RSZI01, RZ00, SMHL00, SH06c, TMAN08, TP00b, VDA^{+}03, FC00, HSTS^{+}09, MHP07, RRDHG08]. arenaceodentata [LLY^{+}01]. arenaria [ARB04, Hun04, MB08, SMT08, TE00]. Arenicola [CBA02, MGM05, ODA^{+}09, SMB^{+}03, SP02, TA03]. areolata [RZEACGF00]. argentatus [HW00]. argentimaculatus [XW02]. Argentina [VMH04, GSK^{+}08, GZY06, LFSR06, LME05, RTPS07]. Argentine [GZY06]. Argentinean [RMC^{+}07]. argenvillei [SSB02]. Argopecten [BCBQ01, HD07a, MB07b, SCM09b]. argus [BFLÁ01, BL09, CH05, EJKN03, NPLZ00, PE05]. Argyrosomus [TLFS06]. Ariake [KKA^{+}06, IT06b, YTK^{+}05]. arid [CM08, FH08b]. arietina [MJG^{+}00, RGA^{+}02]. arise [vWCAB07]. arm [MAH^{+}08]. arm-fencing [MAH^{+}08]. armament [FD08]. armata [PdNS06a]. armatus [HH07a]. armhook [ZKHP07]. armorhead [Hum00]. Aroclor [MFA03]. aromatic [CWR06, NAB^{+}06]. array [BHFC08, EB05]. arrays [OWBD07]. arrival [BOR08, TJM03]. arriving [MBN^{+}09]. arrow [JSWH03, SJL05]. art [CBEL06]. Artedidraconidae [ORD00]. artefacts [GRLML09, HJ05]. Artemia [RMC^{+}07, SW04, SM05]. arthropod [Hir06]. Article [SLL07]. articles [RBL^{+}04, Rii04b]. artificial [ABC^{+}05, ASES07, BKP07, BS08b, BCPB06, BFLÁ01, CUB08, FB00, Goo09, HAC06a, KOTL07, LÁMBF09, MHP03, OM04, PFZS^{+}06, PGPL03, RS05]. artificially [YOM^{+}09]. artisanal [FVB^{+}09, GAAB^{+}08, MRGC09]. Arutyunov [SLP^{+}09]. asbestinum [HBC09]. Ascension [GBGH02]. Aschers [KDAK01]. Ascidiacea [LLRDT05, dCLLT^{+}02]. ascidian [BSDR01, Car00, CLC04, MC07, MPK03, STE04a, SBL08]. ascidians [BDR06, LL03, SL05b, SL05c, TLLCT02, YAF^{+}07]. Ascophyllum [AVA^{+}09, DJM07, DJM08, PB00a, TP00a]. asexual [BRAS03, Gil02b, KAPTK02, McG02, Pur07, SBM02]. Asian [CML08, DTAF00, KUOI06, LVRN02, MCE01a, O’C07]. asinina [CMS^{+}01]. Asparagopsis [PdNS06a]. Aspects [HM06b, Kin01, AHW03, BA04, ED02, HVCT09, KT03, KSWP04, PPCU07]. aspera [BDW^{+}00, BDDG02, KS08]. aspergillosis [BMK07]. Aspergillus [Ryp08]. asperum [SK04]. assay [GMP07, GHD^{+}02, HJH^{+}02, NBDG05]. assays [OHL^{+}06, SN04, WP00a, WP00b]. assemblage [AHSPG08, BS05a, BB02a, CCJT09, CEF04, DFSS08, DGJ^{+}03, EH00, EÖKS^{+}02, FPKC05, GBdS^{+}05, GB01, GBNÖ08, HGA07, JKQ02, KF02, KFT08, KGRW06, KWL^{+}05, RP06, SF06, SMC^{+}05, SLC05, TDSM02, WP07]. Assemblages [MBB^{+}04, AB00, ABC^{+}05, AO01, AWA^{+}05, AVA^{+}09, ABU01, ARP09, AGDP^{+}06, AAF^{+}04, BF07, BSP^{+}08, BBB^{+}06, Bel07, BCO07, BMC^{+}09, BC06b, BdBK^{+}00, BH06, BES^{+}08, Bul05, CBM^{+}03, CHLV07, CDO08, Chi02, CS03a, CCU07, CDK^{+}07, CM05, CRZ04, CAA07, DP01, DPH02, DFJSJ^{+}05, DBN^{+}09, DVNvdV07, DGdG^{+}09, DM09, FBT^{+}01, FTBC05, FNDP08, GVKF03, GYLÖ06, Gla01, GC01, GS05, GP08, GGGVGG03, GTFB03, HDP02, HP09, HKH04, HKM00, HLBS07, HGG^{+}03, HK04a, HKK^{+}06, HLGCO06, HSCF06, Hir06, HK02c, HvW08, HW03, IC02, JMdG^{+}06, JMM^{+}07, Kam00, KKLL02, Kel02, KL03, KCS04, KRVB05, Kor01, Las03, LV02, LC09a, LJM^{+}00, Ley05, LÅCN01, LG08b, MBVBC09, Mag02, MGL^{+}07, MCCV08, MCC^{+}00, MG05a, MGD^{+}09, MSSP06, MM07b, MVK^{+}09, MRB09, MMA^{+}04, MVdMFF08, MG00, MGHA05]. assemblages [MJKH03, NMDS00, ÓUA05, PGR^{+}01, PPCU07, PDO01, PKRK07, PMNP^{+}01, PAR08, PR05, RFN02, RCDC06, RU02, RS05, SAB^{+}05, San00, SASR03, SPP^{+}09, SCZ^{+}06, SWL07, SA03b, SLM07, SRS04, Sta08, STRB06, SJ04, TCD01, TBCB^{+}05, TAFBC07, TDP^{+}00, TBSJ^{+}04, TH06, USG^{+}08, VLK04, VL00, VJ03, VBMBC08a, VBMBC08b, Vie09, VMH04, VMH^{+}08, WA03, WRW03, YYL^{+}00, YBGP06, Zhu06]. Assembly [Con05, SAF^{+}09]. assess [AMG^{+}07, BH08, BH02, MGTK00, MK02]. assessed [BBS05, DR01, HLBS07, SML05]. Assessing [AWA^{+}05, BRF03, BHV^{+}02, BDÁM04, CM04b, FPKC05, GP08, HMY^{+}08, HJ05, JRT^{+}06, JVJ09, LLH06, LHT06, PRB^{+}08, PCS^{+}04a, RCC01, RCWYK07, SH05, SAF^{+}09, VBMBC08a, Fox04, GTD^{+}08, GDBF07, MWD^{+}09, OSF08, PHH^{+}05, RM09b, SSL00b, SJ05]. Assessment [AG04, Eva08, KPA^{+}09, LVvdLH07, PEdlP^{+}02, Ano02i, AFPS08, BDGS07, BSW^{+}09, BCO07, DSDS04, DGCM09, FHDA^{+}07, FH08d, Gor03, HCB^{+}07, HSBD08, Hum00, JDVG08, KWP07, KFH04, Kin04, MAA^{+}07, NT06, RRS^{+}09, RW06a, RQF^{+}00, SdV06, SP08a, SMA^{+}03, SMG02, TBCB^{+}05]. Assigning [OP08]. assignment [MPR^{+}06]. Assimilation [MFJT08, PLM03, SLL02, WLT^{+}03, ABY00, CNW^{+}07b, GCL04, KG00b, LWH06, NATR^{+}05, RW01, RPS^{+}06b, RAAT^{+}07, TÁCTR06, XW01]. associated [BJH04, CWCP^{+}06, CCU07, CFMB00, CAS00, CHP^{+}05, CH00b, CM05, DDQ^{+}00, DP01, DOML^{+}07, EH00, ENC^{+}04, EB04b, ENAT07, FMCB05, FF02, GZG^{+}07, GLG^{+}05, GB05a, HS08a, HAMW09, IMW^{+}02, IHW^{+}01, JTVC08, Jen05, Joh03, JNFL04, Kel02, Kel04, KGB^{+}01, Kum06, LHQ03, LGQ^{+}00, LC09a, LG08b, MMC08a, MTH04, MRP09, MSUK01, MFE^{+}05, MJKH04, NDOS03, NK07, PGR^{+}01, PPCU07, PAK^{+}00, PCJ06, PQRAC08, PWdN^{+}00, PZA09, RYF01, Rit06, RBG^{+}07, RSAK02, RTH06, RFH^{+}06, SM08a, SBH^{+}02, SHJC08, SPRK02, TGY^{+}06, VHV^{+}04, WM04, WAB^{+}02, WTF^{+}05, WPC^{+}03, WYB^{+}04, WdNS00, YAF^{+}07, YFH^{+}02, ZLGC04, ZAZL08, dBS01]. Association [ADF^{+}09, EC07a, RFG04, KWHS^{+}06, LH02, ML07, NDV^{+}00b, PJP06, PBS^{+}00, SG05, SZ09, SH00, VK02, VBB06]. associational [Lev08]. Associations [BCC^{+}03, ACPS09, Bat09, BMdNP09, BIL^{+}02, CMGBJ05, CTD^{+}09, DE01, DMA02, FWC02, GA08a, HVWH06, JLJ01, MFCM05, ML06b, MMMB^{+}05, NA04b, Vli05]. assumptions [CA00, HCS07b]. Astaxanthin [SAHK06, AVS03]. Asterias [DCT^{+}04, GWH03, HRDA05, MJS07, MJS09, RJH02, VTY06, WBDR06]. Asterina [KAPTK02]. Asteroidea [HPBT03, HBTD04, KAPTK02, VTY06]. Astrangia [DC07a]. astreoides [BG04a, GEC08, RL05b]. asymbiotic [ACT09, RCR03]. asymmetrical [Fre07, TBCB^{+}05]. asymmetry [ZE01]. Asynchronous [MH08b, MH09]. At-sea [NCN^{+}09, SG05, BKH^{+}09, CRLP09, TND^{+}05]. Atherina [ADR^{+}05, CRC02]. Atherinidae [CRC02]. Atka [NS02]. Atlantic [BI02, BKSS07, CAS00, EL05, FR07, GMD08, HKH04, HB07a, HKK^{+}06, HH07a, KH08a, KRP^{+}06, LSE06, MGREA07, PMG^{+}06, SFF^{+}07, SOB^{+}06, TPQ^{+}03, TCC04, TCS^{+}09, TBH06, ZAHJ^{+}06, AW03a, AED^{+}07, AMS01, AH08, ABD^{+}03b, ACPS09, AL09, BCHS09, BAM^{+}04, Bar02b, BCC^{+}03, BCCM03, BM03, BRIP00, BIL01, BIL^{+}02, BW01, BAB04, BdBK^{+}00, BVI^{+}05, BIG^{+}06, BPK^{+}00, BPS^{+}00, BEKT09, BM07c, BCLS06, BGH06, CSB03, CW03a, CJFM09, CAD00, CRV03, CHH^{+}05, CIHL05, CS02a, CSMBH08, CCML07, CCCH01, CPG^{+}09, COSM03, CC05, DDF01, DBG^{+}09, DC02, DGC^{+}05a, DCM^{+}06, DÁSG^{+}01, EOK^{+}08, FHL^{+}09, FPWM03, FRvdMF02, FMZvW02, FFYS^{+}07, FKI^{+}09, GLB^{+}04, GBR^{+}04, GSMD00, GFB^{+}07, GAM02, GPB02, GLP^{+}08, GC07, GCVR01, GUT04, HPH^{+}01, HANM^{+}05, Har06b, HS08b, HMH^{+}08, HHD^{+}09, HK04a, HSHA06, tAHSHA07, HLY05, HD06a]. Atlantic [HRC^{+}07, JERS03, JVT01, KFL09, KJB^{+}06, KBP06, KvNVH00, KW00b, KdLCM09, LTFJ00, LTG^{+}01, LTMF01, LRN^{+}09, LNB^{+}03, LAV04, LAK07, LG02, LALB08, LSHL03, MPM^{+}03, MPSJ06, MG09a, MBG^{+}03, MRS00, MTL^{+}03, MB02b, MVBGM04, MCP^{+}06, MFA03, MST03, MSF^{+}09, MSSP06, MLS^{+}08, MFP04, MZD^{+}05, MGSJ08, MFP^{+}05, OJPM01, OÁSG^{+}08, PSW^{+}08, PMPI03, PTM08, PBCB03, PPB^{+}09a, PCRR08, QA01, RJS04, RR05, RR06, RMT^{+}05, RCF07, RGCC^{+}08, RSD^{+}08, RFG^{+}09, RHSB08, RF06, SLS^{+}00, SLM^{+}02, SRJ09a, SRJ09b, SAC^{+}00, SEH07a, SCGC07, SC06b, SHSV04, SMFP08, SWEBS00, SAC^{+}03, SBHE05, SOB^{+}06, SRJ07b, Tay05, TB08, TFSJ00, TPBF09, TWMW06, TR06, TSJH05, THBSJ05, ULL03, VDA^{+}03, VTY06, VAJ^{+}03, WHL^{+}08, WW07, WFC03, WGV^{+}02, WPIB09, WBJ^{+}07, WWWC02, XZUK05, YW04, vDHHM08, vdMM03, vHGH01]. atlanticum [PMNL07]. atmosphere [KFU^{+}08]. Atmospheric [OYT^{+}02, KS04c, RHHH09]. Atol [WGR07]. Atoll [SLGD00, AAPO00, FPF01, PCBS05, SSBZ07, RK09, SPCG09]. atolls [DFH05, DFSS08, MHP07]. atomus [SML05]. ATPase [CACDC07]. atrazine [JMHS03]. Atrina [GFP^{+}05, HTL^{+}02, MNP02]. Attached [DCO^{+}08, AY03, FK00, RCWYK07, VK05]. Attachment [HTZ^{+}07, HMR00, HS01, QRS00, SF07a, TL08, TJA08, ZMN07]. attack [BB02b, vHGH01]. attempt [Kin03]. attendance [SAB^{+}02, STAS06]. Attenuation [PLMS04, BN02, MHDS09, VMN^{+}03]. attract [DUSJ^{+}09, VNdG^{+}06]. attracted [JTB^{+}03]. Attraction [DSJBS^{+}02, HCB^{+}07, MCBT07, NPLZ00, SD01b]. attributes [DT06, Las03, Ley05, LAV04, PJ06a]. Auditory [EM05]. auklet [AS04, AS06a, HRCB02]. auklets [Dav00b, GRIR06, HBRJ06, WIDM09]. auks [KKW^{+}03]. auratus [DWB04, HJG03, NJPI03, PBH^{+}03, TSHT02]. Aurelia [MR07a, MR07b, MR07c, PBS^{+}00, Pur03, Pur07, PHS09]. Aureococcus [BMS04, BM07b, CCL^{+}04, GLCL04, SLT^{+}01, TGSW06]. aurita [MR07a, MR07b, MR07c, SZGS08, SSR^{+}09]. austral [JMM^{+}09, MGPAE01, PTK^{+}09, RHC^{+}08, WSA^{+}06]. Australia [AGDP^{+}06, BB09, BDR06, BML^{+}08, BB08a, Bra08, Cam01, CWBS03, CS03a, CMS^{+}01, DCAL^{+}02, EC07b, FWC02, FBJ00b, HGB01, HDRC^{+}08, JRP01, JBH00, JMHS03, Kev01, KS00a, KCSPT09, LHT^{+}02, Ley05, LPS08, MGL^{+}07, MB02a, McI03, MBP^{+}06, MJKH03, MK03c, MBG^{+}08, MS03, MBN^{+}09, NJPI03, NSEL00, OHT02, OKM06, PKM04, PCMB06, PPJ^{+}05, PK00, PCL08, RKLK07, SSHW^{+}07, SH05, SNP03, SR01, TK01, TC01a, TA01a, TWT08, UST^{+}04, UB01, Uth06, WHM^{+}01, WAK^{+}09, WKL01, WH03, YBGP06]. Australian [BSK06, CBE04, CRBE04, CRCV04, CVSM09, Eyr00, EF02, HJW08, HRDA05, JSWH03, KHH08, LAH07, PMG05a, SJ09a, SBH^{+}02, SdNH^{+}06, TLFS06, TBID09, VHH^{+}02]. australis [AML^{+}03, ABWK06, AKL^{+}04, EB04b, GdLB^{+}09, GMS02, HB03, JR01, LC07a, MJG01, Min02, OHT02, OKM06, PTSJ06, PAMJ03, SC01, STCJ01, SMG02, SPM03, TML03, TA01a, VGB^{+}05, WWAC00, WHM^{+}01]. Austrovenus [Mar04]. Author [Ano00a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b]. autogenic [DEAI06]. automated [PBL^{+}09, SJL05]. Automatic [CWB^{+}06, HD05, HD06b]. autonomy [GDSB04]. autotroph [DTD03, Tan02]. autotrophic [BDÁM04, LJK01a]. autumn [BNP^{+}09, GJG^{+}03, GGC^{+}05, GGDHRA00, GRG^{+}02, HWBZ05, LMKG01, MGF^{+}03, MFP04, WLS^{+}09a, WW07]. Auxis [MOP^{+}08]. Availability [TFL^{+}04, AQSP07, ACLF09, BPB07, BHH09, CLMR07, CLRM08, CGH^{+}03, DD07, DF06, FPB02, FEO07, Gla07, HLR07, HHJ04, HC05, ILL06, KHWH03, KPW07, KM01, KUN02, LF02, LLL^{+}09, MCL07, MDR^{+}06, Mar04, MRC^{+}08, MVB^{+}05, Mun07, NPTW05, RGR^{+}01, RWH00, RLDL02, SF06, STV^{+}09, SvNvdMvdV09, SO01a, SHM02, SIK^{+}02, Swa04, TN07, TDL05, VMRL04, VML05, VCBH06, WIDM09, WS06, WR00b]. available [HBP^{+}05, SW04, VMN^{+}03, WL03]. Aveiro [BMG00]. aversions [LH06]. avian [AL08]. avirostris [ASC06]. avoid [BWO04, Lus03, WH09]. avoidance [KKG06, Kro05, TK03b]. avoiding [AM01, Rit02]. AVS [LLY^{+}01]. ’Await’ [ANF^{+}00]. away [KNF^{+}03]. axes [NvdVVD06]. Axial [LLJ05]. azooxanthellate [FH08c]. Azores [FCS^{+}09, MSH08, MSSP06, MMK^{+}08, WCZ^{+}09]. aztecus [RMKR00, RRRM08].

B [MLW03, SML05, IP01, ABB^{+}00, LFSR06, MGL^{+}00, SMC^{+}05]. b-diversity [IP01]. B. [KT06]. Bacillariophyceae [BVH02, FDBT06, OPSM04, RDD^{+}02, SP00, TDvdW^{+}01]. bacillariophyte [YNM^{+}07]. Bacillus [NSFW09]. back [AE02, ADK^{+}06, BWW^{+}04, DGdG^{+}09]. back-reef [AE02, ADK^{+}06, DGdG^{+}09]. backed [SG05]. background [GLCL04, Joh04, OGL09b]. backtracking [CDJ^{+}07]. Bacteria [MMB^{+}02, BBV^{+}01, CBAd00, CWCP^{+}06, CZW^{+}08, CNC00, CPH^{+}02, GSW01a, GG09, HAMW09, JCK03, KECJ06, KCB06a, LCS^{+}01, MN00a, NWHH01, NBB^{+}04, NHN^{+}06, NCP^{+}08, PdNS06a, PPI04, PBW^{+}07, Rit06, RSAK02, SLGD00, SSGJ00, SPBM05, SBH^{+}02, TGY^{+}06, TBC^{+}00, VFJ06, VF09, WW04, ZLGC04, ZHS^{+}05]. Bacterial [GGG^{+}04, HK03c, LWB09, MLS^{+}08, SV01, TDC^{+}03, VGS^{+}01, AGCF^{+}08, Bat07, BWSH00, CPSC01, CAA00, GB06, HF06, HSD03, HH03b, HAD07, ISSK03, JQ05, Kir02b, Kir02c, LMCQ02, LYL^{+}07, MLW^{+}02, MHDS09, MGPAE01, MFP04, NFN03, NCHP05, NCP^{+}08, Pet02a, PPCJ^{+}06, SJPC02a, TUK09, TPQ^{+}03, TNHL02, VLP^{+}04, VTDS09, YFH^{+}02, YYL^{+}00]. bacterially [CB00b]. bacterially-active [CB00b]. Bacterioplankton [GS02, CSL^{+}05, HQDH06, HSCF06, HGKK06, KvNVH00, NLCS05, Pas02, PWB^{+}03, SAB^{+}05, SBvD05, VPS^{+}02, WLJ08]. bacterium [KHWH03, SDW^{+}01]. bacterivorous [MR04]. Bacterivory [SJK^{+}06, FH08a, GQ00]. Baffin [AG02b, CMG04, LSdM^{+}03, NGMT01, SSH^{+}04, VMN^{+}03]. bags [HTJ00]. Bahamas [BBC03, CSS00, KM01, LYL^{+}07, QDDP06]. bahamense [PBBC06, PBBC06]. Bahía [GGMGR07]. bahianus [RAC^{+}05, RCP^{+}05]. Baikal [GNHB^{+}05, WBS^{+}04]. Bait [HCB^{+}07, HLBS07, SAH^{+}05]. Baited [SB00, CDS07, FTDK07, HLBS07, JTB^{+}03, MGL^{+}07, WMB00]. Baja [GGDHRA00, IOSPT^{+}04, SVG07, BHN09, GGR06, HLGCO06, JPL04, LTPL05, LH09]. Balaena [LSMR05]. Balaenoptera [ABD^{+}03b, Ano01a, JTR05, MFG^{+}04, RBHS06, RBGM06]. balance [AMP^{+}02, AMR00a, AMR01a, BVN^{+}03, CAT^{+}06, CLN00, CDFRL03, HHIS06, MOPP05, TN07]. balanced [ULL^{+}05]. Balancing [ZMN07, GPW^{+}02]. balanoides [BBT02, BG07b, DBT02b, FKSW09, JÅC^{+}01, PSS06, SJH^{+}04, VBB06]. Balanophyllia [GAZ02]. Balanus [BJLH04, DP09, GSK^{+}08, GQ00, HMR00, HTZ^{+}07, Kad03, MWTB00, OTBR00, PA05, RBW03, THC^{+}05, TL08, TOBR07, WR00a]. Balearic [AO02, CMMN^{+}01, LIM^{+}06]. Baleen [MAC^{+}07]. Balistidae [BJC02]. ballast [DRG^{+}02, DBvOM06, HCP^{+}06, HTBT07, MBN^{+}09, SLEH01, TMD^{+}07, WKL^{+}03]. ballasting [WSL06]. balthica [ABB02, BDH01, BD03, BDJ09, BPCvdM06, BPvdM07, BdVD^{+}01, DB07a, GFL02, HW02, Hid03, LLSL06, ÓUA05, SMHC03, Tal01]. Baltic [HVW^{+}07, HLH04, LLSL06, RKJ01, SRGS01, SSSV00, SHS06, AVS03, BGK^{+}06, BGKN00, BEE06, CMB^{+}09, CH07b, DGBS08, DMV07, EDBvdL08, EÖM08, GEHZ07, HP06, HJ03, HCS^{+}07a, HBT^{+}08, HMKT02, HMHB03a, HMHB03b, HNMC09, JNS04, JLCA01, JCG^{+}02, Kaa04, KLRS00, KOTL07, KWV^{+}05, KM00, KMR01, KSKK^{+}07, Kre01, LSH^{+}07a, LG08a, MHP^{+}00, MvdBÓ00, MSH^{+}03a, MPDS07, MLW03, NNW04, NN06, NWÓ03, NB06a, ÖBFO01, ÖCOB06, PKS00, PBW^{+}07, RSK04, RMH07, Rol00, RE00b, RAE07, RHGB02, SAJW04, SHNV09, SGK04, SMHC03, SO08, SBC^{+}06, Tal01, TA07a, TKK05, VMCG09, VFV01, VV04, VCBH06, WVL01, WNS^{+}05, WR00b, vdBÓME01]. bamboo [ASLD09, DS09b, STF^{+}09, Thr09]. band [AS06b, MW02a, SKR08]. Banda [MMA03b]. Bank [DLB04, HKK^{+}06, SF05, CLO^{+}09, MSD05, SIK^{+}08, Bri03, BCLS06, GYD04, GJB04, HSTS^{+}09, HCV03b, JDCB09, KHvW02, KLW^{+}09, LB05a, LAV04, LHB^{+}06, Mag02, MJS07, MJS09, ML06b, ML07, MBS03, MBS07, MGSJ08, SBL^{+}00, SH06c, TCS^{+}09, TT02]. banks [LPK^{+}07, ZCR^{+}09, SC06a]. banksii [BMC^{+}09, Kev01]. Banyuls [MGA^{+}01]. Banyuls-sur-Mer [MGA^{+}01]. bar [KWHD08, MIH^{+}00, ZS03]. bar-bellied [KWHD08]. bar-built [MIH^{+}00]. bar-tailed [ZS03]. Barataria [JBA02, LB08a]. Barbados [RC04a, VKH08]. Barbara [ABWK06, NW02]. barbatus [LRGdS00, PEdlP^{+}02, SSL^{+}07]. barcoding [BAB^{+}06]. bare [O’D08]. Barents [DDL^{+}03, HMD^{+}07, JGH05, Bar02b, Bar07, DD09a, DD09b, Hel00, HSO04, SHFP^{+}03, TRH^{+}06, WHL00]. barium [HW09, MT05, WT06]. Barkley [MZJ06]. Barnacle [OTBR00, Sha09a, Sha09b, BCABC00, BJLH04, BBSO03, BNP07, CMD^{+}06, DW09, DBT02b, FKSW09, GSK^{+}08, GQ00, HMR00, Jar00, Jef03, JÅC^{+}01, Kad03, MMA^{+}09, Mil07a, MWTB00, Nod04, PA05, PSS06, QRS00, RW01, RBW03, RAA02, SM01a, TP07, THC^{+}05, TL08, TCW^{+}08, TLNH01, VBB06, WR00a, ZA07, Zab09]. barnacles [BWW^{+}04, BCS09, BG07b, CML08, DMM^{+}03, JC08, KT04, LNV^{+}05, MMV05, MM08, Ros01, TPGQ07, WTF^{+}05, YAF^{+}07, ZBNW09]. Barokinesis [Hue08a]. barracuda [FS08]. barramundi [MC05c]. barrens [BS00, DHR04b, LJ09]. Barrier [BM02a, Ber02a, Ber02b, BBA06b, BB08a, Boo02, CDS07, CWE^{+}08, CMD^{+}09, Dun02, EC07b, Gru03, JMHS03, KHP09, Kri02, LRM00, MB02a, NMDS00, PKM04, PCMB06, SRDG05, UBRvO06, Uth06, WNWH02, WHB09, HF09, LVvdLH07, PQRAC08, SLHT05]. barriers [SWM03]. Bartholomew [BBE03]. bartramii [CCC09, Par06, WKIK04]. base [JWM09, MOPP05, PB07]. based [Bea04, BMSH02, BCL^{+}01, BS04b, BS05b, BCC08, BL09, CN04, CPB04a, CRAP03, CDJ^{+}07, DGVK^{+}04, DC06, DFW^{+}03, DMA03, DH09b, FMWGB08, FM02, Fre03, FBD06, FWD^{+}01, GAMB05, GLV^{+}09, GDR^{+}04b, HDM04, HD07b, Hum00, ISSK03, JBS08, JPPW02, JWM02, JH04, KSR^{+}07, KTF^{+}01, KLW^{+}09, Kup03, LGC^{+}08, Lee00, LHB09, Lei01, LBR^{+}09, MMVB07, MCC^{+}00, MLS^{+}06, MS05, Mil07b, MSH^{+}03b, MTCU08, Nae06, NCN^{+}09, PCCL05, PG01, PBCB03, PD07, PRM^{+}09b, PRM^{+}09a, PTCH01, PKLD09, RDS^{+}09, RGEB00, RGDI06, RZS05, RG09, SAG08, SD09, SMM^{+}04, SRZZ01, SPM03, SB07, SHSF08, TBB^{+}04, WGCM^{+}03, dSHCR07]. Baseline [ZK01, CRT^{+}08]. baselines [Rui04]. Basin [GES^{+}06, HSHA06, tAHSHA07, HVW^{+}07, MB07b, BF08a, THJK06, FDD^{+}03, KT03, PPI04]. Basin-scale [MB07b]. basins [Joh07]. basis [BMW^{+}00, BDDG02, PMGE01]. basking [NWJ^{+}08, Sim00, SSR^{+}03]. BaSO [KCTG06]. bass [BM02d, Har00, IT06b, Kri02, LTCP03, LW04, NH03a, NH03b, SRZZ01, SMF^{+}02, SKO^{+}08, HDRC^{+}08, Kev01, KHH08]. bassana [Mon07]. bassanus [HPW^{+}00, HPH^{+}01]. bassensis [PGR^{+}01]. Bassia [PGR^{+}01]. basta [CWCP^{+}06]. bat [CD08]. batch [Pep06]. bathyal [CMMN^{+}01, GPB02, KJB^{+}06, PMNP^{+}01]. Bathygobius [Sha00]. Bathymetric [BF04, JTB^{+}03, HHKH05, ISY07, LG01, MRJ05, MC06, REM^{+}06]. Bathymetry [Roy02, dBTO^{+}09, KTF^{+}01, SCKC07]. Bathymodiolus [MC05b]. Bathypecten [BDDL03]. bathypelagic [ISYM06, Ike08]. Bathysiphon [GPB02]. Batillaria [KT06, KHK^{+}04]. battagliai [BBM^{+}02a]. Battle [Zab09]. Bay [AZKJAM06, AG02b, BP03, BMF06, Cam01, CRV03, CMG04, CDT^{+}03, CT01, DAM^{+}08, EB04a, FFN^{+}08, FTK^{+}04, GCHC00, HHIS06, HGAS^{+}01, Hen09, HGRB04, HE08, HAD07, HB08b, IOSPT^{+}04, INH^{+}09, JRP01, JBA02, JPNM^{+}06, KKTN05, KSSH02, KR04, KOY^{+}05, KCSPT09, KYYM00, KGS^{+}08, LB08a, LRK09, Lap01a, Lap01b, Lee00, LNPB04, LG06, Les05, LNH00, LLW09, MHW06, MTV06, MBR^{+}05, MJG01, MS05, MJKH03, MK03c, MIL^{+}00, MIH01, NJPI03, NW05a, OHN01, OKH^{+}07, PC02a, PCBL05, PI07a, RH03, RKLK07, RZ00, RMZC07, RCB05, SRJ09a, SH06b, TK01, TNHL02, TNT08, TDTR03, TDP^{+}00, TLNH01, TBH06, TBL^{+}09, War05, WAB^{+}02, WAMP08, WLJ08, WRW03, Zab09, BCBQ01, CLS00, DSM01, DGdG^{+}09, FS02, GGR06, HDCA06, JN09, JH04, KKL^{+}03, KSHP00, LDL^{+}04, MB07b, MP06, MHHS06, NDV^{+}00a, NDV^{+}00b, Pee02]. bay [SBJ02, WYFX02, ZYZ^{+}06, AI04, APM^{+}01, BSBD^{+}04, BMV02, BCCM03, BIR00, BTC^{+}04, BLHB04, CN04, CJRT00, CD05, CS08, CT09, CDML06, DBH^{+}07, DTV09, DCM^{+}06, DDSR07, DCRT07, DBvOM06, FBN^{+}07, GHH^{+}04, GHL02, HTW^{+}05, HK03b, HAP^{+}08, HK04a, HS07, IT06b, JSMB05, JWR05, JTR05, JH04, KHB00a, KBA^{+}05, KSHP00, LHD^{+}01, LBS^{+}05, LLH06, Lap01a, Lap01b, LRR^{+}09, LGD04, LSdM^{+}03, LS02, MFS07, MDBC03, MGA^{+}01, MPS^{+}08, MJ07, MIH^{+}00, MILH00, NLG05, NA04a, Niz07, NSH^{+}08, NGMT01, OFHM03, PCJF06, PJJ^{+}05, PSRB00, PD01, PAMJ03, RHS01b, SSH^{+}04, SAT^{+}03, SSGJ00, SMHC03, SAH09, TAM03, TAM04, Tor09, ULL03, VMN^{+}03, WGD05, XZUK05, ZIUG07]. bay-ocean [DSM01]. Bay-scale [WGD05]. bay/open [MP06]. Bayer [Cai09]. Bayesian [MPR^{+}06]. Bays [JBT^{+}07, DCO^{+}08, MSA07, NKK^{+}01, SMH^{+}03].

BDM [WLL03]. be [FI09, GSM05, GPHC09, GC02, POH^{+}02, Sim00, TS04]. Beach [DFT^{+}07, HGS^{+}02, BVVM02, CD07a, DGL01, DM05b, GY00, ILL06, JSN08, MZJ06, TK04, WCRB03]. beach-cast [CD07a, MZJ06]. beachcast [NCN^{+}09]. Beached [PBN^{+}07]. beaches [Bra01, CM08, CHA06, DC02, DG05, FDDS09]. bead [EB05]. beaked [MPS^{+}08]. beam [MHDS09]. bearded [GNAD^{+}03, TBBI04, WLS^{+}09b]. bearing [HKR^{+}08]. bears [GNAD^{+}03, TBBI04]. beat [RL07, TJSG03, YA09]. Beaufort [BNP^{+}09, CMJ04, MRB^{+}08, XZUK05]. becomes [RSR02]. bed [BFK^{+}05b, BSED08, BH02, DGM02, DFV01, DH01, JBS08, JYS06, NM07, RWBS03, SM08a, SKR09, SWL07, SCD^{+}09, WPB08a, WPB^{+}08b]. beds [AM02a, BS00, BvK07, CH09, DGC^{+}05b, DVNvdV07, DKFV05, ENAT07, GA08a, GDKN08, HSWG^{+}06, HFM^{+}02, JNMB08, LJ09, MR05, MJS09, MCCT07a, MCCT07b, MSUK01, MS01b, MJKH03, NKK^{+}01, NRvdV^{+}02, NvdV04a, NvdV04b, NBN^{+}08, O’D08, PPCU07, PA06, RPO00, SS09a, SASR03, SG09, SCZ^{+}06, TM02a, TTD^{+}06, TÁCTR06, TJA08, VT00, Van02, Van03, WKD^{+}04, WPB^{+}08b]. been [Nug02, RGVC07]. before [GSC^{+}01]. begets [GHP02, PMGE01]. Behavior [KTK^{+}08a, MBBW08, Ada01, AM02b, AY03, AAY^{+}07, BBHT02, BGC06, BBDW09, BHH07, BJJ01, BLH02, CCdAR00, CCK03, CR04, CCL^{+}06, DCP^{+}08, DJG^{+}07, DSM01, EDB^{+}04, FPB02, FBD06, FHN^{+}09, FAB^{+}07, GYD04, GHJ02, Goc04, GCS06, HPS07, HMBD01, HSSS05, IAKS03, JP04, JNH^{+}09, KRKS09, KBB09, KKW^{+}03, KB05, LHJD^{+}03, MR04, MY02, MY04, MAH^{+}08, MGC^{+}09, MAW06, Mil07a, MWTB00, MWS^{+}04, MW01, MHR03, NSH^{+}08, OJ02, Par06, PE05, Pea00, PTB03, PKLD09, RFLC07, RSIS09, SHT^{+}04a, SICT03, SHK05, SAH^{+}05, SAM^{+}06, TE00, TC03a, TCD01, Tit01, TT07a, TT06, VFM06, VPG04, WB07, WLS^{+}09b, WS09, WOL^{+}07a, YOM^{+}09, ZBDE07, ZZT05, SKMS06b]. behavior-based [FBD06]. Behavioral [BEW03a, BEW03b, KCTG06, LBF^{+}07, Lev08, MB08, OCB^{+}07, RW07, RST04, CHL^{+}05, Goc04, KSR^{+}07, LCHG03, RBG04, SO01a, TSC^{+}07, Tya09]. behavioral-physical [CHL^{+}05]. behaviorally [Rye02]. behaviors [IMS05, LS09b, LDT07, PBS^{+}00, TSI02]. Behaviour [CTB07, LGS07, Lei07, Trn02, ULL03, AFHS08, ABY00, BLD^{+}05, BCHS09, BBG^{+}06, BKH^{+}09, BDCvB^{+}09, BH01, BGL^{+}07, BM05b, CRLP09, CSTC01, CYC^{+}09, CVT09, DPS^{+}03, DMA03, DGH08, EWG^{+}08, EEWBdN08, FBD^{+}02, FA06, FCCP06, FJK^{+}07, FB03a, GMZvdL07, GL04b, GLA06, GDR^{+}04a, GML02, Gru03, GBD^{+}06, HHF^{+}07, HHG^{+}07, HHB^{+}00, HLPHA^{+}01, HDG^{+}08, HBBH07, HDRC^{+}08, HBG^{+}02, ICH04, JSIT02, JT08, JdSM09, JMJ07, JBS01, LKDO07, LTH^{+}06, LAG08, MVN^{+}06, MSH08, MC07, MEHD07, MSF^{+}09, NW06, NZM08a, OST07, PMG05b, Pea03, PC05, PPS^{+}07, PWC^{+}09, PRvdL^{+}07, PCW05, PEGG^{+}08, PK05b, QMM02, RM00, RWJ02, SSN04, SDD^{+}07, SG05, SJM06, SK07, SAS^{+}06, SCKC07, SSR^{+}03, SRB04, SRJ07b, TT05a, Tal01, TJEM02, TML03, TR05, TC00, TC03c, VJKF07, WBS^{+}07, WST^{+}09, WWV08, WGCM^{+}03, WDLM05]. behaviour-based [WGCM^{+}03]. Behavioural [CT04, GSM^{+}08, GBW^{+}05, KCH^{+}03, Kro05, LGB^{+}04, NZM07, SB07, WBDR06, BB09, BSP03, FDDS09, FJK^{+}07, HJH^{+}02, Lec05, RHSB08, SP03a, SLB^{+}08]. behaviours [LHH09, SV08]. behind [VJ04]. Behrens [Ano04c]. belcheri [CBDH02]. Belgian [RLDL02]. Belgium [SVG^{+}05]. belief [WDB^{+}05]. Belize [GR07b, HGKJ01, MBH^{+}00, MSS^{+}03, PCBS05, RPS^{+}04, SLHT05, WPH01]. bellied [KWHD08]. bellus [HS05]. below [KD00, KHGA02, KGB^{+}01, RHG05]. below-ground [KD00]. belowground [CC00]. beluga [GRRH07]. beneath [PHO^{+}07]. benedicti [EWSH08]. beneficial [LKW03]. benefit [BDR06, RFLC07]. Benefits [BSVM^{+}05, Enb05, CHW06, EB04b, PL03, SZW^{+}07, WDLM05]. Bengal [HAP^{+}08]. Benguela [BAB^{+}01, DRF08, FPB^{+}07, GB01, HFMP03, LVvdLH07, PRvdL^{+}07, PRL^{+}09, RVMI^{+}03]. Benthic [ATTX00, BD11, FAPA^{+}09, FAPOF09, GFP^{+}05, GRPT^{+}00, GKWR02, HKH04, HS08b, HGT01, IOSPT^{+}04, KH00a, KTF^{+}01, LRR^{+}09, LFF^{+}09, LWB^{+}00, LBH01, LBC^{+}09, MTR^{+}07, MDBC03, MAJ^{+}06, MVD^{+}08a, NLG05, NOO^{+}08, PAMJ03, RBMB00, SBFP08, SCS^{+}07, SYCG05, TC00, WBG09, WRJ^{+}05, WKPK05, AGS^{+}09, Ano07g, AWA^{+}05, Arm01, AHW03, ABD^{+}02, ARFP09, BWMA02, BCM^{+}09, BLHB04, BRF03, BM07b, BS04a, BYH^{+}06, BK06c, BSIS07, BDC00b, CM04a, CMD07, CC02, CJRT00, CLH^{+}07, CS03c, CM05, CB05c, CAA07, DP02, DCT^{+}03, DSPHV04, DB02, DJ04, DSMBG03, DR01, DGW01, DEvdZ03, DdNEvdZ05, Ell02, EEA01, EF02, FGLM06, FA08, FWF02, FH08a, FCC^{+}08, FWC02, FTO^{+}09, Fro00, FCH^{+}08, FR08b, FEO07, GSMD00, GBSQ09, GRP^{+}06, GF04, GMd^{+}02, GCHC00, GBR01, GGTKG09, HDP02, HLVU00, HSPH05, HCV03b, HDDR08]. benthic [HTL^{+}02, HLTT09, HCCY02, HDCA06, HAD07, HAA00, JBSB01, JPPW01, JPPW02, JPC^{+}06, JTB^{+}03, Jos09, KCH^{+}06, KH07a, KNTK05, KKL^{+}03, Kar07, KGHW05, KBL^{+}07, KKH07, KGHL^{+}01, KVV^{+}00, KT03, KP01b, KST^{+}08, KGS^{+}08, LHQ03, LHQ05a, LHQ05b, LDL^{+}04, LPK^{+}03, Les05, LCD^{+}05, LBCS05, MCFM02, MCCV08, MTV06, MEHD07, MM03, MLC08, MAT01, MSY02, MVK^{+}09, MC08a, MBM^{+}00, MMB^{+}02, MRB^{+}08, NReZW09, NR00, NDOS03, NWFV06, NHN^{+}06, NAB09, NRH^{+}07, ORD00, OIP04, OW02, PGCC^{+}06, PTLK04, PDT09, PvDMB08, PFZS^{+}06, Phi05, PD01, PSGP04, PCS04b, PHO^{+}07, PDFD09, RGS^{+}09, RMM04, RWB^{+}09, RATD07, RWP07, RC08, RT00b, RAH^{+}02, RGD^{+}08, RS05, RST04, Saf03, SR09, SLL02, SF02a, SABK03, SE04, SWL07, Sma07, SW09, SS04b, SAF^{+}09, SPRK02, SMSD07, SSdWM00, Sta08, SG01, SJA^{+}09, SPGS03, SMH^{+}03]. benthic [SCH^{+}07, TDT^{+}03, TRH^{+}06, TGA02, TGGA04, TdMF^{+}08, Thu07, THJK06, TGM^{+}03, TNSM00a, TPS^{+}05, TSJ05, VCA^{+}09, VV04, Vii07, VFP^{+}07, WWAC00, WE04, WTS09, WWL04, WDSA09, WAMP08, Win09, WHB09, WASW03, WKK07, WdBS02, WS02, YAF^{+}07, YI03b, YLL^{+}04, YH02, ZS03, ZYZ^{+}06, dJTD07, vDS05, CHW02, CBE04]. Benthic-pelagic [LRR^{+}09, LWB^{+}00, LBH01, CJRT00, DP02, KBL^{+}07, LDL^{+}04, PSGP04, PCS04b, Saf03, ZYZ^{+}06]. benthic-planktonic [FCH^{+}08]. benthivores [AL08]. bentho [HPB^{+}04]. bentho-pelagic [HPB^{+}04]. benthopelagic [ACAG07, CMMN^{+}01, JTB^{+}03, KST^{+}08]. benthos [BA01, BGK09, BMF06, Bow05, BCPB06, BTF^{+}06, Cla08, GLG00, GP03, HBK^{+}05, ISBP06, Lar03, Lin04, LCB^{+}05, REM^{+}06, SAV09b, SAV^{+}09c, TGGA04, WG01, WAA^{+}09]. benzo [ABB^{+}00, LP01a, WW06b]. bergi [SJS^{+}01]. Bering [CGL^{+}02, JCZ^{+}05, Tyn04, BHPC06, BSH02, CB05a, CGC^{+}09, HLD^{+}07, LSM^{+}02, LCB^{+}05, NEK^{+}07, OP08, RL03, RZS05, SBI07, YM06]. Berkelmans [Dun02]. Bermuda [CSMvE04, LDR00, MGO^{+}07]. berndti [CEBR07]. Beroe [KFA^{+}04]. Bertalanffy [HLR04]. Besiktepe [Ano02c]. Best [Ano01a, Hou09]. Bet [Kru01]. Bet-hedging [Kru01]. beta [DFV01, dVCH^{+}06]. Betaeus [LMC08]. betaproteobacterial [BCM^{+}09]. better [EM07b, JM01, MMCF07, MVH^{+}06, MB07e, TJM03]. between [Abd07, AG02a, AAPO00, AI04, AMP^{+}02, ABHE09, ANF^{+}00, BLD^{+}05, BMK07, BCC^{+}03, Bay02, BDH01, BW04, BC06b, BCH00, BGS02a, Bra01, BBT02, BOZ03, BJJ01, BM02d, Bul05, CED^{+}07, CÁ04, CA00, COPB06, CC07, CDW^{+}09, DEAI06, DPS^{+}03, DRDD09, DFSS08, DFV01, DPD^{+}03a, DVNvdV07, DBT02b, DOML^{+}07, EJM01, EG03, EGJ07, EPM06, FPDS^{+}02, FTP^{+}09, Fau09, FSW09, FSB07, FH08c, FZT^{+}05, FH03, FD02, FBM07, GMA^{+}09, GGHA02, GRP^{+}06, Gil02a, GAB^{+}03, GB02a, GHT^{+}01, GSCW00, GSFN00, GCU07, Gor03, GC07, GMd^{+}02, GPSG06, GGDB09, GRLML09, Hal00, HAMW09, HGSH08, HHC00, HK04a, HH05b, HTL^{+}02, HLY05, HST^{+}09, HK04b, HFM^{+}04, JC08, JMA02, JSIT06, JCK03, JLJ01, JNH^{+}09, KS08, KWP07, KH00b, KH08a, Kri02, KSWP04, KCA^{+}08, KKI^{+}06, LHJD^{+}03, Lar03]. between [LKR^{+}06, Lei08, Les05, LTNH00, LGZ^{+}09, LD03, LVZW06, LMF^{+}01, MR06a, MVHV06, MHM^{+}02, MJS09, MSK00a, MFS07, MAM^{+}01, MSS^{+}03, McI02, MW04b, MWTB00, Mis04, MMMB^{+}05, MB04, NDV^{+}00a, NKK^{+}01, NBN^{+}08, NWÓ03, NTGK09, NFN^{+}09, NOO^{+}08, NB06b, OC01, PSW^{+}08, PDM^{+}04, PSPZ09, PBAA04, PGK^{+}08, PH01b, PC02b, PRL^{+}09, PJ08, PVV^{+}06, RMM04, RWE04, RTB06, RFN02, RB04, RFG04, RCGA^{+}04, RZEACGF00, RC05, SG06b, SH07a, SEH07a, ST06a, SJE^{+}06, SICT03, SN02, SH07c, SYA02, SSL^{+}07, SSBG02, Ste04b, SB05b, SMGCW00, SJ09b, SH00, Tak02, TMJ^{+}05, Thu07, TCS^{+}09, TMFF00, TBAT02, TA01b, TBC^{+}00, TBH06, TBSJ^{+}04, Uth01, UB01, VHK09, VLW09, VBB06, VAJ^{+}03, Vli05, WH02, WHP00, WBS^{+}04, WIDM09, WFG^{+}08, WÓ02, WL05, Wie05, WGM02, WHB09, YNM^{+}07]. beyond [ARP09, CCR08, NNLF09]. Bias [MJG08, MUT03]. biased [BGRD00, PPJ^{+}05]. bibliometric [GDDM05]. bichique [HLFG^{+}07]. bifasciatum [GCS06, SGCP06]. bifilosa [CW03b, Gor03, KSKL^{+}03, SKK^{+}06]. Big [GPS07, Joh04]. BIG-STf [Joh04]. bigeye [NMOO03]. Bigger [MVH^{+}06, SPM03]. Bight [Har06b, SSSV00, SAHK06, BAA04, BAA07, CH06b, Hob00, HH07a, Joh07, KSG03, LL03, LSG^{+}05a, SRJ09b, SP01b, SC00b, TCC04, TCS^{+}09, VFP^{+}07]. bilinearis [SC00b]. billed [CBDH02, FBD^{+}02, QMM^{+}08]. billfish [BTW08]. billfishes [LC08]. bimodal [SL05d]. Binary [Mon07]. Binding [BG04c, SW04]. binned [TÅ07b]. Bio [BBC^{+}00, BPK^{+}00, BPS^{+}00, RSK04, SSH^{+}00, VMN^{+}03]. Bio-mineralogy [BBC^{+}00]. Bio-optical [BPK^{+}00, BPS^{+}00, RSK04, SSH^{+}00, VMN^{+}03]. Bioaccumulation [LTMDB02, WW03, WW06b, BGQR01, BMM^{+}04, CR02, FSSW^{+}00, GGHA02, GFL02, LLY^{+}01, LW05, PW08, RSL09b, WKYL00, YLL^{+}04]. Bioactivity [LMCQ02, JQ05, PCL^{+}09]. bioassay [SSL00b]. bioassays [TA07a, XYH^{+}09]. Bioavailability [DDF01, NN06, TA03, WAM^{+}08, WG00, WYFX02, ABBN09, AHL^{+}01, PDFD09]. bioavailable [DFMD00]. Biocean [FGLM06]. Biochemical [DTV08, DTV09, WdBS02, AMR00b, AMR00c, CDFRL03, CDT^{+}03, HDM04, LFRG^{+}05, LCHG03, PBCB03, PEdlP^{+}02, RHG04, Sar06, SJ05, She08, SKMS06a, Ven06, WJWL05]. biochemical-based [PBCB03]. Biochemically [FMWGB08]. biochemistry [TSSQ06]. biocides [DBD^{+}05]. bioconcentration [BF01a]. bioconstruction [CFMB00]. biodeposit [CWMD06]. biodeposition [BGKN00, ZYZ^{+}06]. biodeposits [GP04a, MNP02]. biodiffusor [ODG^{+}04]. Biodiversity [BRIP00, CAA07, Duf06, Ell01, FCS^{+}06, Raf06, ABD^{+}02, BB03a, BWRP06, Cha03, CW01, CBEL06, CVB06b, CV06, DS06b, Ell02, EHM^{+}09, GLB^{+}04, GS09, HSM^{+}09, HDDR08, ISBP06, MTF^{+}08, Nae06, NRH^{+}07, Pri02, PVV^{+}06, RDH^{+}09, SRP^{+}06, SvAHH09b, SvAHH09a, SYA02, THF^{+}01a, THF^{+}01b, TLIL^{+}09, WAB^{+}02, WAK^{+}04]. biodiversity-based [Nae06]. Biodiversity-ecosystem [FCS^{+}06, BWRP06]. biodiversity-productivity [EHM^{+}09]. Biodynamic [RSL09a]. biodynamics [PW09]. Bioeconomics [SBC04]. Bioenergetic [Kne03, BDGS07, WTR02]. Bioenergetics [WO05, APG^{+}09, MR07a, NRC07]. bioeroding [SL05a]. bioerosion [LCCTFO03]. biofilm [HGRB04, HBP^{+}05, OTBR00]. biofilms [DQ05, HLQ02, HH03b, HTWQ05, HTZ^{+}07, JQ05, KF02, LMCQ02, MBW03, PA05, POH^{+}02, PBF07, TMJ^{+}05, Uth06]. biofouling [DQW05, MGDP01]. Biogenic [SVG^{+}05, FFML02, KSS^{+}08, LAW06, LZL05, MGRP07, RWR^{+}00, RDG^{+}07, RGD^{+}08, SES01, VPP04]. Biogeochemical [CHW02, CHSW03, GRCdLvB08, HCA06, KHD07, AB04, DP02, FHL^{+}09, FDD^{+}03, GÁSPR03, HTJ00, IAT^{+}07, NCÁSGP05, SVC^{+}01, TK08, TMFF00, TBC^{+}00, War05, ZLGC04]. Biogeochemistry [TSW^{+}09, AC08, BCM^{+}09, EKA03, GDDM05, GK02, GKK03, SGCC06, TPS^{+}05]. Biogeographic [FGLM06, HMK06, PPS00, Smi04, AVE^{+}06, BS08a, BBG^{+}04, CC07, CDS^{+}04, HRB^{+}06, LBH^{+}04, LMC08, POH08, RC09, SC03, TMFB06, ZBNW09]. biogeographical [Kad03]. Biogeography [UST^{+}04, ABD^{+}02, BIL^{+}02, GSG08, HF06, Lei07, MG05b, SKW^{+}06]. bioindicators [GSSR02]. bioirrigation [JGH05]. Biokinetics [BTF^{+}02, NW05b]. Biological [ASR^{+}09, BL00b, KJMP05, LOT^{+}00, LCL^{+}00, LH08, LHGC^{+}09, MM06a, MMCF07, MLR02, SPAMG00, WG01, WGV^{+}02, XZUK05, ARAMPW02, ACR07, AC02, Bel08b, Ber00, CCL^{+}05, CMSS09, CS08, FC04, FFML02, FZT^{+}05, GNPB07, GNP07, Han07, KPA^{+}09, LRR^{+}09, Las03, Mil07b, MHT07, MK02, MR01, MWD^{+}09, PPCU07, Pee02, PNCS^{+}08, PNCT^{+}08, RCB05, SDFM00, SM08b, SK07, She05, TKGS04, WBH^{+}04, ZLAK04, dCLLT^{+}02]. biologically [WLL03]. Biology [CRB07, HFW^{+}05, AYSL09, Car05, CTVP03, CY06, CV06, DM03, DM05a, DBS05, DB07b, HCP^{+}06, Kri02, PCS09, TK04, WB05a, WB05b]. bioluminescence [NOP02]. bioluminescent [RB05b]. Biomagnification [ZWYL01]. Biomarker [WS04a, GPB02, HPBT03, HBTD04, LBCS05, NBDG05, SZ09, SPGS03]. Biomarkers [LLSL06, BSIS07, CPG^{+}09, DBOR09, EOL^{+}05, FPDS^{+}02, KPA^{+}09, MGBB06, MRB^{+}08, OGMC06]. Biomass [RKGH01, AGDP^{+}06, BBBSPH03, BKM08, BB03a, BA00, BM09, BL00b, BFLÁ01, BSH02, BOZ03, BK06c, BNP^{+}09, BD01, CTA^{+}03, CSL^{+}05, CHT09, DT08, DAG^{+}00, FMMS03, FPF01, FPWM03, FH08a, FD02, FBN^{+}07, GHV02, GSMD00, GVKF03, Gis05, GSFN00, GUL^{+}07, GBG^{+}00b, GCM01, HKH04, HCGC09, HBP^{+}05, HRK^{+}08, JHJS04, KD00, KMR01, LOT^{+}00, LRHV07, LD00, LSE06, LV05, MCFM02, Mag02, MBG^{+}03, MDLS01, MSD05, MIL^{+}00, MVB^{+}05, MUP06, NPN06, NGMT01, NPM01, OSF08, OBB00, PCSS01, RHHH09, SNH04, SDS^{+}03, SO08, SdSC01, SS07, SFL^{+}09, SV01, SLM07, STRB06, TKVL04, TNHL02, VSL^{+}04, VLP^{+}04, VSVV03, WH08, WHL^{+}06, WAW^{+}09, WHB09, YFH^{+}02, YQC^{+}00, Yin02, ZHWCS08, Zhu06, vOSF^{+}09]. biomaterials [Dis07]. biomechanical [DESK03, RBG04]. biomineral [LG00b]. biomixing [LKR^{+}06]. biomonitor [LMW^{+}01]. Biomonitoring [RBG05, MCVAR08]. Biophysical [HTW^{+}05, BMT07, LJK01a, PCCL05]. biopsies [HBW^{+}05]. biopsy [Eva08, HYR^{+}09, Rob06]. biorepositories [Dis07]. biosilica [BTC^{+}04]. biosynthesis [BSVK05]. biosynthetic [VHV^{+}04]. biota [CFMB00, GGC^{+}05, GCU07, GKB09, KCH^{+}06, RWPJ04, SPRK02, WBSL00, WBP^{+}06]. Biotic [Rue07, ABB02, BPBH09, GLV^{+}09, HH07b, LSH00, MUT04, PBF08, Tom08]. biotope [NDV^{+}00b, VBV^{+}00]. biotopes [NDV^{+}00a, NDV^{+}00b]. biotoxin [BDGS07]. biotransformation [LK02]. bioturbated [ODA^{+}09]. bioturbating [Kar07]. Bioturbation [SFH^{+}00a, TR05, BRP00, CWR06, GK02, GGM^{+}00, HBT^{+}08, NTH^{+}02, OKP06, OSS^{+}04, Orv05, SWH^{+}04b, Vii07]. bioturbators [WAK^{+}04]. bipinnata [BSVK05]. Bipolar [BK05c]. bird [CBM^{+}03, DMA03, MS04a, NCC^{+}07, Tor09, GSCW00]. birds [BD06, Hen09, KMB^{+}08, PBN^{+}07, QWY07, SP01b]. Birth [FGT^{+}07]. Birth-date [FGT^{+}07]. birthday [ZK09]. Biscay [BMF06, AI04, BSBD^{+}04, BIR00, BLHB04, DTV09, LHD^{+}01, LBS^{+}05, LGD04, MPS^{+}08, ULL03, ZIUG07]. bispinosa [BSS05]. Bivalva [KTK08b]. Bivalve [Wid01, ASP^{+}03, ABB02, Bay04, BDDL03, BDH01, BDEM01, BDJ09, Bla09, BdVD^{+}01, CTHF01, DSS04, DB07a, FK08, FON03, GZ02, GFP^{+}05, GLCL04, GW06, HSW^{+}08, HW02, Hid03, HWvdM03, HDW04, HM06b, HMFC07, KNTK05, KJL03, Kim06, KSS^{+}05, KCB06b, KB04, LCC07, LC07b, MMTT00, MDBG07, MPDC03, NPTW05, NT06, Pet04, PPSP07, RAAT^{+}07, Rii01a, Rii04a, SB05a, SSBG02, TKN00, Tom08, TYD07, YI03b]. bivalves [BDLB02, BMS04, CDK^{+}07, CTvdM^{+}08, CTD^{+}09, Cra01, CT04, CVT09, ISSK03, KLC^{+}06, KH05, LLSL06, MCS09, NCT^{+}01, NPTW05, Pen00, PH01a, PH01b, Rii01b, Rii01c, RJH02, Sar06, Str02, Urb02, VN02, VN05, WPG08, WLL03, WL03, dZCVC01]. bivalves-the [Rii01b]. Bivalvia [NSSI06, PTH08, BDR04, DF06, MTS07, OJPM01, SSJ^{+}02, SBR09]. Black [GWWC00, AS06b, Bar07, BBC^{+}09b, CWT00, DFPG^{+}03, DBH^{+}02, EG03, FEMW07, HJW08, LWW^{+}01, MARM06, NHS^{+}08, PBFB08, PCB08, RWA^{+}02, SG05, SKR08, SIK^{+}02, ABD^{+}02, AVE^{+}06, Das02, Das03, GF04, Ogu07, OSF08, PE09, SNY^{+}05, SYH^{+}09, VMVG^{+}07, YVB^{+}02, YMC05]. black-backed [SG05]. black-band [AS06b]. Black-browed [GWWC00, CWT00]. black-legged [Bar07, DFPG^{+}03, DBH^{+}02, FEMW07, LWW^{+}01, SIK^{+}02]. black-lip [PBFB08]. black-tail [NHS^{+}08]. blade [RC04b]. Blake [HGC^{+}07]. blame [BN04]. Blanes [PSRB00, KSSH02]. blasted [Fox04]. bleach [WS04b]. bleached [DPM02, RGL01]. Bleaching [FOL02, BM02a, Ber02a, Ber02b, BMM^{+}09, BB02a, BDW^{+}00, BD08, Dov04, Dun02, FIRL02, FHGJB04, GYLÖ06, HR07, Jon04, KHIA02, KJ03, LAK^{+}08, MMMPB05, MBA^{+}05, MAG^{+}07a, MW02a, NvW01, NYvW03, RSY^{+}06, SFL05, SP08b, UBRvO06, Woo09, YH04, vWNYS05]. bleekeri [IMS05, IS07]. blennies [DGCM09]. blenny [WBC01]. Blicher [Ano07c]. Blind [BK05a, BK05b, KBT05]. block [Via07]. blocks [Lam00a]. blood [ACJ^{+}04, KG07]. blooded [SWHSA05]. bloom [AD00, ABWK06, BGC06, CVF^{+}03, CLS^{+}01, EO04, FHL^{+}09, GVM^{+}05, GGK^{+}07, GLS^{+}00, GHH^{+}04, GSW01a, GHL02, GCHC00, HJ03, HHC00, HGAS^{+}01, HMKT02, HCJ^{+}07, HK03c, JKNJ07, JPNM^{+}06, KWAT08, KBSWG09, Kre01, KFN01, KGRW06, KYYM00, KB04, KB05, LPK^{+}07, LRN^{+}09, LWR^{+}02, LTD^{+}02, LW00, MNR^{+}02, MLS^{+}08, MJ07, ONP02, OPSM04, PFW^{+}02, PCYC02, PTD05, RYF01, RJBW03, RFG^{+}09, SV08, SRS05, TGSW06, TNT08, VGS^{+}01, VSV04, WHD^{+}05, WG01, WPTS08, ZGV00, Zam02, ZLAK04, vdBÓME01, RMC^{+}05, WL06]. bloom-forming [Kre01, LWR^{+}02, LTD^{+}02, LW00, PCYC02]. blooms [AGB03, BGJ^{+}07, CCL^{+}04, CS08, CvBGR09, CTFC08, DCC^{+}03, FTK^{+}04, GMC^{+}09, GSW01b, GEHZ07, JNS04, JERS03, KLRS00, KNT^{+}07, LHD^{+}01, LSG^{+}05a, LRK09, LGZ^{+}09, MKL^{+}03, MGZ^{+}03, MSH^{+}03a, NHNB01, PDT09, PCJF06, PUL07, RGB02, RHGB02, SAT^{+}03, SBH^{+}02, SMH^{+}03, WP07, WW04, ZTW08]. Blubber [BRI^{+}07, NFN^{+}09, GNHB^{+}05, HBW^{+}05, HMY^{+}08, HYR^{+}09, MMC^{+}08b, RBGM06, WHP00, WP03]. Blue [DDV^{+}08, EP01, AML^{+}03, Alt06, ABY00, BFK06, BEW03a, BEW03b, BJN^{+}05, BPLD^{+}05, CJFM09, CBTW02, CMT^{+}08, DRDM09, DES^{+}00, EE00, FCI^{+}04, GRPE08, GKS^{+}07, GLP^{+}08, HCM01, HD07b, HF05, HLH^{+}05, JSMB05, JKK^{+}06, JSH01, KPW03, KWWH02, KHvW02, KSH^{+}09, KH05, LLH06, LS02, LW02, LVK05, MNB^{+}07, MH06a, MGH00, NM07, ODF09, OCB^{+}07, OvM02, PMS^{+}08, RBHS06, RW08, SMHC03, SJPC02a, SJPC02b, SCM^{+}09a, SAH^{+}05, TW00, TE00, Thi06, WST^{+}09, WS09, WJ06, WLM06, ZBDE07, vMRO03]. blue-footed [WST^{+}09, ZBDE07]. bluefin [GDdS^{+}07, KNK^{+}00, KNKT01, LFF^{+}09, MAA^{+}07, RSZI01, RSD^{+}08, RFG04, RF06, SDD^{+}07, TSIY06]. bluefish [BM02d, HBNA04, MBL07, SBJ02, SLB07]. bluehead [GCS06]. boat [Lus05, SDD^{+}07, SB07]. boat-based [SB07]. boats [Lus03]. Bodden [HP06, KMR01]. bodies [PK01]. Body [JMSM08, JHL04, MB01, APN06, AVNN05a, BLH01, BPM05, BH04, DFH05, DJ04, EBG^{+}00, FCCP06, GF06, HGS^{+}02, HM01, HS01, HW00, JWM02, LSG05b, Mad02, MSK00a, MHSS08, MRJ05, ÖBFO01, PEG^{+}00, REM^{+}06, Roy02, RBGM06, SJR00, SBJ02, SN02, SJ09b, SHW09, TBH08, WWV08, YA09, ZBDE07]. Body-size [JMSM08]. Bolinao [SESP03]. bonasus [NRC07]. bonefish [HALL05]. bonga [PDM^{+}04]. Bonnemaisonia [NCHP05, NCP^{+}08]. boobies [WLB08, WST^{+}09, ZBDE07]. booby [WLJM05]. Bookhout [Cos03]. Boom [Kav05]. Boom-or-bust [Kav05]. boosted [LEF^{+}06]. booster [DBD^{+}05]. border [HSC^{+}09]. boreal [BAA04]. borealis [FLUF^{+}04, HM03, MH08c, OSL07]. boreoatlantic [ZKHP07]. Boreogadus [BF08b]. Boreomysis [CM01]. Boring [CBGLM08, BB02b, BBST07, SZ09, TYD07, WNWH02]. borne [MHB01a, MY04]. Bornholm [HVW^{+}07]. Bosmina [VV04]. both [BD06, GA03, THBSJ05]. Bothrocara [KSHO01]. Botryllus [PDM^{+}03, SBSRW02]. bottle [NNS01]. bottleneck [NW05a]. bottlenecks [Hal04b]. Bottlenose [FGA^{+}04, ARGS01, FFN^{+}08, HR08, HWGB02, IR02, Lus03, Lus05, LW06, MTLT02, SMP^{+}09, SB07, WGH^{+}00]. Bottom [CMLG07, OÁSG^{+}08, WLJ08, AWS^{+}09, BMF06, BS07c, BLP^{+}02, CSM02, DF02, DH09b, Fro00, HKK^{+}06, HBBD09, HHM^{+}06, HJG06, JFD09, KRKS09, KLN06, KKTN05, KHB00a, KOC^{+}07, LEB07, MAE03, ML00, Niz07, PCS04b, PPC^{+}05, RBRH02, RGP^{+}09, SKMS06b, Som00, TYT^{+}04, THJK06, TLPG06, VHT^{+}05, WDT^{+}08b, WAW^{+}09, Wie05, WMCF08, YM06, ZSKA01]. bottom-feeding [WDT^{+}08b]. bottom-layer [KHB00a]. bottom-mounted [KRKS09]. bottom-trawl [YM06]. Bottom-up [CMLG07, OÁSG^{+}08, WLJ08, BS07c, HBBD09, HHM^{+}06, JFD09, ML00, Som00, TLPG06, WAW^{+}09, Wie05, WMCF08, ZSKA01]. bottom-water [BLP^{+}02]. bottoms [BF04, DF02, MUT03, NWÓ03]. boulder [Bar06, CM05, LAK07, MB01]. boulders [SC07]. bound [AHL^{+}01, OSA05, WG00]. boundaries [FMT09, HF09, PSPZ09, SBA01, TMFB06, VHB^{+}07, ZSR03]. Boundary [CS01, BAM^{+}04, BBABL01, CC02, CMMN^{+}01, GGTKG09, HSHA06, tAHSHA07, HSC^{+}09, ID02, KGHW05, MBG^{+}08, PSGP04, WWL04]. boundary-layer [PSGP04]. Boundary-shear [CS01]. bounds [WG02c]. Bourgneuf [MBR^{+}05]. Bouvetøya [BKH^{+}09]. bowhead [LHJN07, LSMR05]. box [BPC09, TK08]. box-jellyfishes [BPC09]. box-modeling [TK08]. boyeri [ADR^{+}05, CRC02]. Brachidontes [TM02a]. brachiopod [Tom08]. brachydactyla [SMFP08]. Brachyura [BM05b, FTB09, KMF02, RMAY04]. brachyuran [BF04, DSM01, FPB02, FCCP06, Fis06, FCP02a, LW02, PPV^{+}06, PDQ01]. brackish [JHJ^{+}03, JHJS04, KTNK07, KTK08b, KLN06, Mal02, RC00, SGB09, SGCC06, WWAC00]. brackish-marsh [SGCC06, WWAC00]. brackish-water [Mal02]. Bradbury [SRJ07a]. brain [CSMvE04, LWCF04]. branching [Lir00, ZAZL08]. Bransfield [KKKK03, VAP^{+}07]. Branta [RS07a]. brasiliensis [DGL01, DC02]. Bray [Som08, Yos08a, Yos08b]. Brazil [BF04, KW02a, LFSR06, WGR07]. Brazilian [BBBSPH03, GLÅ^{+}03, KWD05, PCT00, SYCG05, WLdG^{+}04]. breadcrumb [TL08]. breadth [SS00a, SJR00]. break [CWC00]. breakage [LMC08]. Breaking [SGCD01, TDL05]. bream [EG03, HJW08]. breathing [CDW^{+}09]. breeder [MGC^{+}09]. breeding [BDG^{+}04, BCH00, CDF^{+}08, CEO06, CMT^{+}08, CCCD09, FBD^{+}02, FWH^{+}05, FEMW07, GSCW00, HPS07, HGB01, HM06a, Hip08, HFA02, JHL04, KHF^{+}00, KS00a, KKL07, LWW^{+}01, LIM^{+}06, MH09, MSJ^{+}09, MRQ07, MND^{+}09, MH06b, NGSV07, NMBH02, OBG^{+}07, PSKC04, RMA^{+}02, RHC^{+}08, SWS^{+}09, SPH^{+}09, SAB^{+}02, SKE05, WHP00, WISH09, WHRS05, WHL^{+}08, WWR09]. bresilioid [PGS^{+}00]. Brest [BTC^{+}04, LRR^{+}09, CJRT00]. breve [LJK01a, LJK01b, LJK02, SLT^{+}01]. Brevetoxin [HR08, KLR^{+}06]. brevetoxins [FFN^{+}08]. brevicornis [DHCH04, Joh01]. brevifissura [KPN03]. brevior [MC05b]. brevipes [SAWG05, SAJG06]. brevirostris [WGR07]. brevis [BMV02, GMC^{+}09, HR08, JK06, SKMS06a, SKMS06b, TDvdW^{+}01]. Brevoortia [HD06a]. Briareum [HBC09]. briareus [BL09]. Bridging [SB05b]. brine [SW04]. Britain [CWCP^{+}06]. British [DER01, DD01, ESDKM09, EBG^{+}00, HRCB02, MZJ06, FFA03, FE06, FGT^{+}07, LWH^{+}04]. Brittany [HPC04]. brittle [BHFC08, GTPC04, LS09b, McG02, MW09, SCC05, SLCC09]. brittlestar [DHT^{+}08, HSD03]. broad [BP08, Lin04, THN^{+}03]. broad-scale [BP08, THN^{+}03]. broadcast [BMWH08, MYH07, NWHH01, NKS03]. broadcast-spawning [MYH07, NKS03]. broadnose [Bra08, LME05]. brominated [NCP^{+}08, SCH^{+}07]. bromophenolosus [TKR06]. Brood [CSM^{+}09, GGFSP06, SIK^{+}02, TGT^{+}06]. brood-guard [TGT^{+}06]. brood-rearing [SIK^{+}02]. Brooding [FCCP06, AFHA05, AM04, AM06, CSTC01, FLK^{+}06, FPB02, MCCP09, NKS03, PV05, RHG04, Ver05, WWDH09, WEEBdN08]. broods [JR01, MS08, Too04]. browed [CWT00, GWWC00]. brown [BMS04, BM07b, BNP07, CCL^{+}04, CC05, CHV^{+}09, GSW01a, GSW01b, GDS02, GLCL04, HTM01, KLFH04, MS08, NB06b, PB00a, PCT00, PCS^{+}07, PVGO01, RK08, RMKR00, RRRM08, SLT^{+}01, TGSW06, TP00a, WST^{+}09, WGP^{+}00]. brown-tide [SLT^{+}01]. Browns [Bri03]. Bryan [EM07b]. Bryde [Ano01a]. Bryopsidales [jMFVPL05]. Bryozoa [LTL06c, OGL09b]. bryozoan [CFMB00, GQ00, IH02, Lid08, LP01b, MHB01a, NK03, Pra05, QRS00, SM07b, SM08a, SG09]. bryozoans [BK05c, BWL06, OHCFJ04, PJ06b]. bubbler [SW08a]. Buccinum [SJ02, WHB^{+}06]. bucktooth [GHV05]. budget [BM06, GF04, KW02a, LMD^{+}02, MIH^{+}00, MILH00, MFP^{+}05, SAJW04, SMSD07, SLB^{+}08, TAÁO^{+}03, WKPS06, WVL01]. budgets [BFB01, BKH^{+}09, CHSW03, GXL^{+}08, MABB03, MIH01, TDUN07, YLK00]. buffer [RB08]. Bugula [GQ00, LTL06c]. Building [Lei08, BVP^{+}06, BS04c, CNW07a, Cum02, DDV^{+}08, FFHW07, FBM08, GNC07, MDW^{+}09, NvWY05, OYNO08, VYM06, ZDA^{+}06, dBKSV00]. built [MIH^{+}00]. bulk [OHL^{+}06, SC06b, dGvdBvO^{+}08]. bulky [ABB^{+}00]. bull [CWKE05, TS05]. bullet [MOP^{+}08]. buoy [FHK^{+}08, RDS^{+}09]. buoy-based [RDS^{+}09]. buoyancy [CD03a, MVHV06, SSL00b]. burden [CC03b, PEG^{+}00]. burial [ARB04, MF03, PKR07, RU02]. Burrow [CCdAR00, OJ02, GDL04, KM03b, MGM05]. burrow-dwelling [KM03b]. Burrowing [Tal01, BVI^{+}05, CTD^{+}09, Gri09, GWF07, MR06c, PNK^{+}02, PZA09, SW08a, WE04]. burrows [BIG^{+}06, GHS05, HAA00, KWKF03, MLW^{+}02]. Burry [WGCM^{+}03]. burtoni [KAPTK02]. business [Lei08]. bust [Kav05]. butcheri [EG03, HJW08]. butterflyfishes [BPM05, Goc04]. butyltins [SJPC02b]. Buzzards [CT01, PCBL05, TBL^{+}09]. by-catch [BRGS^{+}07]. bycatch [Rye02, SAH^{+}05]. byssus [CEB^{+}08, CYC^{+}09]. Bythograea [PEDT03]. bythograeid [HVCT09].

C [FFYS^{+}07, SKFPS02, ZER^{+}01, AW03a, AHC^{+}06, AHSPG08, AZKJAM06, BW01, BVI^{+}05, BGS^{+}02b, CF02, DLLB03, Fre03, GXL^{+}08, HH00a, HGMG^{+}04, HD07b, HCCY02, KPM00, KP01a, KECJ06, KW05, KKL^{+}03, KSK07, LCC01, LHB09, MSF^{+}07, MH08a, MRHP01, MOOB06, MD07, MBM^{+}00, NEMF09, NHN^{+}06, Pah05, PCL08, RGL01, RSN02, RH03, RW08, RLKW04, Rol00, RCGA^{+}04, SAT^{+}03, SJE^{+}06, SR03, STF^{+}09, STH^{+}06, SKO^{+}08, TKR06, TPSW00, TBI07, WPRS05, ZER^{+}01]. C-depleted [TPSW00]. C-labeled [NHN^{+}06]. C-labelled [AW03a, KW05]. C-SOLAS [FFYS^{+}07]. C-tracer [KECJ06]. C. [DMM^{+}03, HB07a, HFM^{+}04, MNTS08, MTN03, PL07, RRIHC02, SKFPS02, TDvdW^{+}01]. C8 [ZRG00]. CA [BP03, CIT04, HYLT03, KSK07, LIT05, MC05c, SCIT09, TSI02, TCW^{+}07]. Ca2 [dBKSV00]. Cádiz [FAPA^{+}09, FAPOF09, PD01, SLP^{+}09]. Cadmium [PMG^{+}06, RGT01, RGT06, WDX01, BGD^{+}03, BTF^{+}02, LBS^{+}05, MW04b, PW08, PW09, RB05a, RWJ02, SSG^{+}00]. cadmium-inducible [SSG^{+}00]. caerulea [CÁ04, CBTW02]. caesia [Bur02]. caesium [TH02]. caespitosa [RMRAFP08]. caespitula [PVGO01]. Caeté [BBBSPH03, KW02a]. Cafeteria [JSK^{+}07]. cage [DSJBS^{+}02, DFJSJ^{+}05, HJK01, NLB08]. cages [LFF^{+}09]. caging [MJKH04]. Cairns [PHS^{+}07]. calamary [PTSJ06, SMG02, SPM03, TA01a]. Calanipeda [BCB^{+}08b]. Calanoid [SVL^{+}04, AMR01b, BRIP00, BIL01, BIL^{+}02, BI02, BI04, BD02, CT05, HH01, JKS^{+}01, JTLG09, NMVH06, Reb01, SSS^{+}05, SAHK06, SKK^{+}06, SKU^{+}08, WWWY07]. Calanoida [BC01b, BCB^{+}08b, LQRGT06, MRZJ05, NOP02, PIY04, STA^{+}05, TSW^{+}04, TSW07, WJWL05]. Calanoides [STA^{+}05, TSW^{+}04]. Calanus [BT06, BCC^{+}03, BSPK03, CWT^{+}01, EO04, GEM02, Gis05, HHC00, HB07a, HGHB00, HBN01, HBH07, ITH^{+}03, IP07, JBT^{+}07, Joh03, KP00, KH08a, MNTS08, MMC^{+}07, MCM05, MGZ^{+}03, MTN03, NH05, Pea03, PTS00, PRM^{+}09b, PJR^{+}03, PJ08, PWL^{+}06, RRIHC02, RRIH^{+}02a, RRIH^{+}02b, RHI^{+}01, RVMI^{+}03, RJBW03, RBRSP00, SHH09, SKFPS02, SAHK06, SGH^{+}06, SYH^{+}09, TJEM02, TJM03, TBVJ08, TBH06, WCBD01, WLS^{+}09a, XW01]. calcareous [BG07a, IF09, SBB09, SEM01]. calcarifer [MC05c]. Calcidiscus [LNJ07]. Calcification [CWN^{+}07, BG07a, CGT^{+}00, GPM08, Lou08, MBLA01, MEB^{+}07, MJG^{+}00, RFPB^{+}04, SBB09, TAS04]. calcified [EG05]. calcifiers [AMB08]. calcifying [BF08a, HOT08]. calcite [AMB08]. calcitrans [TDvdW^{+}01]. calcium [IAKS03]. Calculating [Kir02a, Kir02b, Kir02c, Pet02a]. calculation [SMSD07, WVL01]. calculations [WKPS06]. Caledonia [FPKC05, GDSB04, JDTB06, VMRL04]. Caledonian [BBS07]. calico [SCM09b]. Calidris [LPM^{+}06]. California [BHN09, CH06b, DJG^{+}07, GES^{+}06, GGDHRA00, Hen09, Hob00, HB08b, IOSPT^{+}04, Joh07, KGS^{+}08, LL03, Les05, NW02, SRN02, SH06b, War05, ADF^{+}09, ABWK06, BKM08, BDGS07, BMG02, BH06, CHW02, CHSW03, CSR^{+}00, CS09, CL04, Dav00a, DTB07, DBS05, FS02, GGR06, HLGCO06, HWBZ05, HM09, HGZ^{+}00, JPL04, KBB09, KAAH05, KM03a, KB07, LTPL05, LRK09, LH09, LJM^{+}00, LZM^{+}03, LBC^{+}09, LMB^{+}09, LW04, Mad02, MHABGC04, NW05a, OAS01, PLCG^{+}05, PBN^{+}07, PCB08, RWA^{+}02, Reb01, RCB05, SSS^{+}09, SVG07, SFR^{+}09, SB08, TKGS04, TLC06, WST^{+}09, WH08, WG02a, ZM06]. Californian [BL04]. californianus [HMH04, POH08]. californica [MC08b]. californicus [Mad02]. Caligus [TWMW06]. call [HM01, OCB^{+}07, SCM^{+}09a]. Callianassa [BRP00, PBF07]. Callianira [SDPT06]. Callinectes [BPLD^{+}05, CHL^{+}05, DRDM09, DES^{+}00, FCI^{+}04, HCM01, JSMB05, JSH01, KWWH02, KH05, LS02, ODF09, OvM02, RAA02, SMHC03, TW00, TE00, vMRO03]. Calliostoma [HSCD01]. Callorhinus [KW02b]. Callyspongia [DDQ04, QDDP06, TLPG06]. Calonectris [GDGSP09, NGSV07, NGS09]. calves [BKLGN05]. Calvi [GHL02]. calving [HST^{+}09]. Calycophorae [PGR^{+}01]. Calyptogena [FAB^{+}07]. calyptraeid [BFV08, Col03, CMGBJ05]. Calyptraeidae [MCCP09]. Camargue [BM06]. camera [FTDK07, WDT^{+}08b]. Cameras [BKP07, JTB^{+}03, WBL^{+}06]. camouflage [RDD^{+}02]. Campbell [Ano04d]. campechanus [WCF08, WPPS04]. Campeche [AWA^{+}05]. campylocercoides [TVG00]. camtschatica [TT06]. Can [BDLB02, GSM05, GPHC09, GC02, GQ00, MC01b, POH^{+}02, Sim00, SES01, SAH^{+}05, Arm01, CT05, DSJT^{+}06, DB07b, KRK09, MBW03, MM07a, TB08, TS04]. Canada [EBG^{+}00, FLR05, GSM06, RATD07, SG06b, BRB^{+}00, CWMD06, CSMBH08, CAA07, DER01, DHRG05, DCM^{+}06, DHR04b, DHR06, DD01, GJT09, HS08b, LGBWH06, LHK01, LWH^{+}04, LFSR06, MHCB08, MZJ06, PDRT02, PT04, RBHS06, RS07a, SM07b, SM08a, SH07b, SF07b, WGD05, WAMP08]. canadensis [RS07a]. Canadian [JPMGS08, CMJ04]. canaliculus [GP04a, SAG03]. Canarian [TBSJ^{+}04, TH06]. Canary [BA00, AAHDS06, BBV^{+}01, FVAB00]. Canary-NW [BA00]. cancellata [Iye02]. Cancer [BFE^{+}03, FTB09, HE08, HMA06, MH08c, MS08, SP08a, SO01b, THM^{+}03, TT06]. candel [CDT^{+}01, SKH06]. candelabrum [DUW08]. Cannibalism [BT06, KP00, KM00, SML02a, TOK^{+}04]. Cannibalistic [FSC05]. cannot [FI09]. canonical [SGCD01]. canopies [Abd03a, BJH04, BVN^{+}03, CEF04, DJM07, MC08b, PLMS04]. Canopy [SKH06, BCPB^{+}01, CÅJ05, CCB^{+}09b, EDS03, FWC02, SG06b, SL07, VAB09]. canopy-benthic [FWC02]. canopy-forming [CCB^{+}09b, SG06b]. cantonensis [NWT00]. canyon [OGL^{+}09a, DS09a]. canyons [DBG^{+}09, SSHW^{+}07]. Cap [OGL^{+}09a]. capability [BC06a, SWM03]. capacity [CLSM09, GR07b, JT08, KF02, MMB03, RLBB04, SO02, TJSG03]. Cape [JBT^{+}07, CRV03, CDML06, GDR^{+}04b, MND^{+}09, PRvdL^{+}07, SHA^{+}05]. capelin [Car05, CB05a, DM03, DM05a, HSO04, PWFC08, RFCM03, WJDA06]. capensis [MND^{+}09, NPVW06, PRvdL^{+}07]. capillare [GPB02]. Capitella [LGFM00, LGVF00, LT00, MW04a, RF08a, TSSQ06, TWM^{+}01]. Caprella [OAM02]. caprellid [OAM02]. capricorni [HUP^{+}00]. capsules [BHMR01]. Captain [SLP^{+}09]. captive [KKK^{+}05, NWIT08]. capture [DBCB05, GDSB04, RL05a, WDT^{+}08b]. capture-recapture [GDSB04]. captured [JBSB01]. captures [GZY06]. Caranx [MHP07]. carbaryl [LBF^{+}07]. carbohydrate [AMR01a, DFMD00, PUB^{+}01, WSL06]. carbohydrates [GKM02, HHT^{+}09, NCÁSBP04, dBS01]. Carbon [CHO09, CVK00, CBE04, CRBE04, CRCV04, CVSM09, DES^{+}00, DM09, EP01, GCL04, JBS08, Lee00, MHM^{+}02, NSJ02, NMA00, RBDK00, SSN^{+}09, VNJ06, VHV^{+}04, vDSD02, AMR01a, AJ03, ARFP09, BSR09, BWSH00, BWCW02, BJR04, BGS^{+}02b, BVN^{+}03, CMG04, CCMK03, CMRF05, CH07c, CNW^{+}07b, CNPS05, CvBGR09, DRG^{+}07, DZ06, DÁSG^{+}01, EF02, GPW^{+}02, GRPT^{+}00, GM08, GSW01a, GGRC^{+}08, GC06b, GTSS03, HARC05, HH03a, HLR07, HKY^{+}02, HO02, HBK^{+}05, IT06a, JFA02, JCK03, KTNK07, KKIV01, Kin01, KKB^{+}06, Koe05, KMR01, KKSR05, KW02b, KG07, LR02, LG00b, LGM^{+}09, MBW03, MCP05, ML06a, MLG^{+}03a, MDLS01, MP05, MG00, MLM^{+}02, MFP^{+}05, MGL^{+}00, MLG03b, NOO^{+}08, OAGV07, PW04, PTK^{+}09, PMNP^{+}01, RFPS^{+}02, SN06, SAJW04, SS02a, SEVSP02, SHS00, SH07d, SHT^{+}04b, SFR^{+}09, TYT^{+}04, TRH^{+}06, TSF01, TAÁO^{+}03]. carbon [TPQ^{+}03, TCS^{+}05, TSC^{+}07, TPSW00, TKTJ07, TDC^{+}03, TBI07, Tyn04, VLP^{+}04, VFJ06, VTS^{+}02, WWWR09, WdBS02, WLT^{+}03, XZUK05, XZLP09, XW03, dBGLT08, dGvdBvO^{+}08, vDHHM08, MCCT07a]. carbon- [Lee00]. Carbon-isotopic [VHV^{+}04]. carbon-specific [CMRF05]. Carbon-to-chlorophyll [SSN^{+}09]. Carbonate [AED^{+}03, JWG01, MCCT07a, MBLA01, MSUK01, MAKMK09, NSEL00, RZWG08, SGHG07, WRJ^{+}05, WWH05]. carcasses [BKP07]. carcharhinid [PWLC06]. Carcharhinus [CPBS07, HCD^{+}09]. Carcharias [BB09, DNL08a, LKDO07, WOL^{+}07a]. Carcharodon [DNL08a, LKDO07, WOL^{+}07a]. carcinophaga [SKE05]. Carcinus [AQSP07, BH01, GD07, HJH^{+}02, JMA02, LTMDB02, LBCS05, Mis04, MHR03, Mok04, QMM02, SvNvdMvdV09, Smi04, SRLA04]. cardinal [ÖNCNG05]. cardinalfish [BBH06]. cardinalfishes [GJ05, MB02a]. care [CC03b]. Caretta [BBM00, CLP^{+}06, CAF^{+}08, HTM^{+}02, HGS^{+}02, HBG^{+}02]. Caribbean [AP06, ALPK00, BMK07, BHH05, BSVK05, BM07a, BS04c, BH09b, BIC^{+}03, BI03, BL09, CPBS07, CSS00, CRAP03, CH05, DVNvdV07, DGdG^{+}09, Edm04, EP08, ET08, GPK^{+}04, GS05, GEC08, IE06, KPEF00, KKSR05, Les00, LVZW06, LÁMBF09, NDV^{+}00a, NDV^{+}00b, NKK^{+}01, NvdV04a, NvdV04b, NBN^{+}08, NB06b, OP02, PB08a, PHP^{+}03a, PCBS05, PFP03, SKR08, SC06a, SFH00b, SSBZ07, SPT04, TSMC09, Ver05, VFM06, VFJ^{+}07, WP00a, WP00b]. Caridea [CY06]. caring [VMdBV06]. cariosus [Nod04]. carnivore [KLHO02, VRJ02]. Carnivore/non [VRJ02]. Carnivore/non-carnivore [VRJ02]. carnivorous [HB04, SK03, TNT06]. carnosa [NP05]. Carolina [DRDM09, GB07, HGC^{+}07, PPHT00, BBS05, FCI^{+}04, HFM^{+}02, MBL07, NRM05]. carotenoids [ZRG00]. Carpentaria [JRP01]. carponotatus [Kri02]. carrion [ITN^{+}06]. carrying [MMB03]. Carscadden [DM05a]. carterae [JKS^{+}01]. cascade [Das02, KVS08b, NRJB04, SB03a]. cascades [Bas06, ÖCOB06]. Cascadia [SPRK02, HSPH05, SRL^{+}02]. Cascading [JHJS04, SLL07, SLL09, MVK^{+}09]. Case [RSK04, AAHDS06, Arr05, BDR06, BER^{+}03, BDÁM04, DSDS04, GR07b, GDR^{+}04b, GBI^{+}06, JMB^{+}05, MSY02, MGM05, MB00b, PBPLB06, PB08b, PBJ^{+}05, QMF05, QWY07, RAM03, SAMN04, SWGC^{+}05, Sto03, TBCB^{+}05]. cases [SF05]. Casey [Sta08]. Casitas [BFLÁ01]. Caspian [BN04, FKA^{+}06, KFA^{+}04]. Cassin [AS06a, HRCB02]. cast [CD07a, MZJ06]. catadromous [AWBO09]. catadromy [TDC^{+}07, TA01b]. Catalan [CBS^{+}08, KSSH02, CM01]. Catastrophe [EE06, GLS^{+}00]. catastrophic [San00]. catastrophically [HP09]. catch [BRGS^{+}07, CWM^{+}07, RNL00]. catch-up [RNL00]. catchability [ASA^{+}03]. catches [RF06]. categorical [dM01a]. catenella [CCM01, SPP^{+}09, VGMC01]. Catharacta [BPT^{+}00]. cathodoluminescence [LABR^{+}06]. Catostylus [PK00]. caudatus [KRP^{+}06]. caught [BCL^{+}01, CES^{+}03, FOFG09, MCP^{+}06]. Caulerpa [BBC08, CMJ00, DDHQ06, JCMR^{+}03, MGD^{+}09, MPB^{+}09, MS03, PP02, PCC01, PC02b, RLBB04, WG02a, WS04b, YBGP06, ZM06, jMFVPL05]. Caulerpaceae [ZM06]. Caulerpales [PP02]. caupo [OJ02]. causal [FH03]. causative [Ryp08]. cause [KJ03, Nod04, WHRS05, WÓ02]. caused [BOM^{+}08, GP03, HWG06, JTVC08, KKSR05, LMC08, LPK^{+}07, LKW03, RMPC08]. causes [BDEM01, FMMS03, GY00, HGS^{+}02, HGAS^{+}01, KL03, MB07d, Mou04, PUL07, TKVL04]. causing [Bul05]. cauta [HGB01]. caution [AG04]. cavernosa [LFRR^{+}07]. caves [BTF^{+}06]. cavities [SBvD05]. cavity [KNKT01, dGvdBvO^{+}08]. Cazlyn [Cos03]. Cd [GFL02, KSS^{+}05, LLY^{+}01, LW05, NW04, NWT00, RW01, RSL09a, SW04, WLL03, WL03, WG00, WR00a, Wan02a, XW02]. cDNA [SSG^{+}00]. CDOM [SNCP04, URFA06]. cecilia [OTB^{+}00]. celebesensis [AMT03b]. celebrates [Ano00d, ZK09]. Cell [FHGJB04, RGP^{+}03, TAS04, BVH02, BBRJ04, CF08, GM01, LWR^{+}02, MDLS01, MW04b, NGdN^{+}07, OHL^{+}06, PBVP09, PEGG^{+}08, RG09, SH01, VFF09]. cell-based [RG09]. cell-size [CF08]. cell-specific [MW04b]. Cellana [BCU04, WDLM05]. Celleporella [MHB01a]. cells [AKRA02, CMRF05, COB^{+}03, FCH^{+}09, IRM08, SLT^{+}01]. Cellular [GKM02, KLR^{+}06, OGMC06, SH03]. Celtic [BIR00, DFR08, TPD^{+}05]. census [MVdMFF08, WMB00]. censuses [GTD^{+}08, SB00]. Centimetre [WM02a, WMS03b]. Centimetre-scale [WM02a, WMS03b]. central [ANCC07, BSR09, BBC^{+}09b, BH06, CC04, CAF^{+}08, CSH00, CL05, DFSS08, DD01, FC00, GSP^{+}02, HCM01, HQDH06, HK03c, KAAH05, KJMP05, KHP09, KCEB09, LNV^{+}05, LTNC07, LBF^{+}02, LC02, LCB^{+}05, LBC^{+}09, LMB^{+}09, MC01a, MDBC03, NNLV06, NNW04, PCL^{+}09, PPL^{+}02, RT00a, Rod03, SMH09, SHS08, SAB^{+}02, SSL^{+}07, VDA^{+}03, WKIK04, AAPO00, CAD00, DMV07, FCP02a, GGM^{+}00, HK04b, LSHL03, MC05b, WR06, WCH09, YCSS06]. central-southern [HQDH06]. centre [TWT08]. centric [KKKK03]. Centropages [BK05a, BC01b, CID^{+}06, WJWL05]. Centropomus [AWBO09]. century [CED^{+}07, CLB^{+}09, EOK^{+}08, STF^{+}09, TLIL^{+}09]. cepedianus [LME05]. cephalic [PL00]. Cephalophanes [NOP02]. Cephalopod [CDG04, GPM08]. Cephalopoda [DGC^{+}05a, GMS02, IAKS03]. cephalus [CIT04]. Cerastoderma [BD06, CWR06, DdMB04, KJL03, KvKB06, UNII01]. Ceratium [ONP02, SCS03, SH07d, TLIL^{+}09]. cercariae [TdMF^{+}08]. Cercopagis [LG08a]. Cerithidea [SR04b]. Cerorhinca [TS07]. certain [KNT^{+}07, WA00]. cervicornis [RR05, WM05]. cessation [FRJ09]. Cestode [BDRM06]. Cetacea [MZD^{+}05]. Cetacean [BKM08, Tyn04, RFB^{+}06, dSHCR07]. cetacean-habitat [RFB^{+}06]. cetaceans [dSCV^{+}08]. Cetorhinus [NWJ^{+}08]. Ceuta [GGGVGG03]. cf. [PPB^{+}01]. CH [KFU^{+}08]. Chaceon [FVAB00]. Chaenocephalus [PSPZ09]. Chaetoceros [TDvdW^{+}01]. Chaetodontidae [BPM05]. Chaetogammarus [LP01a]. chaetognath [CDT^{+}03, Øre00]. Chaetognatha [TSMC09]. chaetognaths [TT05b]. Chaetomorpha [BRBPL06]. Chagos [SHS08]. Chain [PO09, MF08, ODA^{+}09, PAB^{+}02, RGJ^{+}04, SM00b, WKYL00, WW06b, XW02]. chains [Wan02b, WL05]. chalcogramma [BBC^{+}09a, Bai00, DABC02, PCH^{+}05, WJDA06, YHSH02]. challenge [BSS05]. challenged [LW05]. challenges [HP07, Nae06]. chamber [CSM^{+}09, HAA00, ZLGC04]. chambers [MRLK06, MCCP09, OIP04]. Champsocephalus [RHDC05]. Change [CH05, Bai00, BGK09, Cal06, CCJT09, CCL^{+}08, CS02b, CMGBJ05, DFH05, Den08, DCA^{+}09, Fie02, FBBV09, FH03, Gou06, GB09, Hue08b, Joh03, Kra07, KYYM00, MS04a, MSJ^{+}09, MHT07, MK02, OAS01, PZ07, PBF08, RC08, SSL00b, SHW09, TT05a, TdLvdVR08, THN^{+}03, TSF04, WM04, WAN^{+}01, YMC05]. changed [MF07]. Changes [BVMBC07, BB02a, CTA^{+}03, CAS00, DBJ^{+}00, DLD^{+}00, EMB^{+}09, FF02, GHL02, GTFB03, HW02, IHW^{+}01, JPL04, KSSH02, KK02, KWTR08, MG00, PCSS01, RV05, RBG^{+}07, RFH^{+}06, STAS06, SMVF03, VSV04, VHMB09, WBP^{+}06, YA09, ACAG07, AVNN05b, ABL08, AP00b, BCB^{+}06, BBBSPH03, BIL01, BI04, BDG^{+}04, Boo02, BTB^{+}09, BPA^{+}08, BES^{+}08, BD01, BMJ02, CBD04, Car05, CPBS07, CR08, CSPA09, CMT^{+}08, CWN^{+}07, Das03, DM05a, DPD^{+}03a, EKA03, EH04, FC08, FHN^{+}09, FR08b, GXL^{+}08, GRWB02, GSCW00, GAM02, HS08a, HNYI07, HSM^{+}09, HGJM09, HM01, HHT^{+}09, HSB08, HB07b, IMTW09, JSN08, JHL04, JCL09, JF02, JPMGS08, KHIH08, KKTN05, KYML02, KHH08, LBBM06, LG01, LG08b, LP04, MMC08a, MBVBC09, MGA^{+}01, MBR^{+}05, MRC^{+}08, MSK^{+}00b, MKR07, MRQ07, NLB08, Pea00, PMT^{+}02, QTB02]. changes [RL09a, RB02, RHS01a, RWH00, RGCC^{+}08, RZWG08, STPA04, STV^{+}09, SB07, TTM^{+}05, TBID09, TBH08, TLIL^{+}09, Tyn04, VSVV03, VBMBC08a, VML05, WAB^{+}02, WPB08a, WDR^{+}05, WPTS07, XW01, YOM^{+}09, Yap09, Zam02, dJTD07, dlMPN^{+}03]. Changing [HGKK06, NLHO09, WCCA09, WGH^{+}00, BTC^{+}04, CCK^{+}09, Dav00b, DM03, GHL02, HF08, KF02, Mon07, NCC^{+}07, Tho07, WWDH09, WCM08]. Channel [ABWK06, LW04, NW02, Øre00, PD01, Zam02, JS08b, MDBC03, VBS06, AMS^{+}06, BIR00, BER^{+}03, DD04, GPSG06, HBH07, JHB06, SML^{+}09, SK07, WLJ^{+}04]. Chao [UG04]. Character [DS09b]. characterisation [LVK05, WMS03b]. characterise [TBM^{+}08]. characteristic [OIP04]. Characteristics [BBC^{+}09b, HGC^{+}07, LC07a, MAB^{+}03, ROCB09, VJ04, AFPS08, BDG^{+}04, BEU^{+}07, CMB^{+}09, CLMR07, CLRM08, CTH^{+}03, DC06, DE08, DD09b, EH00, FDD^{+}03, GDS02, HM05a, HM05b, HGKK06, IS07, JSMM07, JMM^{+}09, KNTK05, Kel02, KFA^{+}04, KG00b, LSHL03, MRPL07, MBH07, MD08b, MZD^{+}05, Nem05, OBB00, PZA09, QRS00, RS03, SMM^{+}09, SSH^{+}00, TFSJ00, VSV04, VMH04, WGP^{+}00, WA03]. Characterization [BBSO03, CB00b, EKN08, HHH02, KMR01, SRS05, SLKF03, UYNO01, DMK^{+}08, DHZ02, FTU^{+}00, HCS07b, NEMF09, RFCM03, RK09, SN06, YFH^{+}02]. characterized [WGR07]. characterizes [CTvdM^{+}08]. Characterizing [BS07b]. chase [AW03a]. Chasmagnathus [BIG^{+}06, CGCDA02, GA03]. Chatham [Chi09, GGS09]. Chattonella [WL06]. check [ST02, WM04]. Checking [BWO04, MdMR04]. Cheilinus [CDAM06, Tup07]. Cheilodactylus [LS04]. Cheilodipterus [ÖNCNG05]. cheilostome [CFMB00]. cheiragonus [KMF02]. Chelonia [ABL08, BBC03, BHN09, GSB^{+}06, KCEB09, SRN02, SJE^{+}06, TTD^{+}06, TBBB06, YA09]. Chemical [ALPK00, BMWH08, BNP07, CWM^{+}09, DDTQ05, GP01, GWBdN09, KFH04, LS09a, NCHP05, NP05, PdNS06a, AKRA02, AIM^{+}05, BFK^{+}05b, Bro06, CTA^{+}03, CH00a, COK02, DKAH02, FPB02, GA08b, HJG03, JSN08, Jon04, Kaa04, KZ00, KLFH04, LH06, LTL06c, NATR^{+}05, O’C07, OP02, Paw00b, PAC^{+}00, PHM^{+}08, PAM^{+}09, PLC^{+}02, PEdlP^{+}02, PPS00, PPBS02, PFP03, RW06c, SR08a, SJ06, UP01, VH03, VBAR07, VJ04, YI03a, BIC^{+}03]. Chemically [DdNS06, CR04, GZG^{+}07, GLG^{+}05, KPW03]. chemicals [BBM^{+}02a, DMC09, HLY05, LBC07]. chemistry [AAJ07, BS07b, BTKG^{+}08, BML^{+}08, CSMvE04, GB03, HJG03, HCP^{+}06, KHP09, KVV^{+}00, MC01b, NGdN^{+}07, RSZI01, RZWG08, SRJ09b, Tak02]. chemistry-toxicity [BML^{+}08]. chemoautotrophic [BS03, VPSO01]. chemodetection [JK04]. Chemoreceptive [ITN^{+}06]. chemosynthetic [MCFM02, TWM^{+}01, YFH^{+}02]. chemotaxonomic [DK06, GVKF03]. chemotaxonomy [ZJW^{+}04]. CHEMTAX [EDBvdL08, Lat07]. Chenier [BA04]. Chernobyl [PE09]. Chesapeake [CS08, KHB00a, BMV02, CN04, DBH^{+}07, DDSR07, DBvOM06, FBN^{+}07, HTW^{+}05, JSMB05, JH04, KBA^{+}05, KR04, KSHP00, LLH06, LS02, MHW06, MDBC03, MJG01, MJ07, NSH^{+}08, OFHM03, PAMJ03, RHS01b, SRJ09a, SMHC03]. chick [AFHA05, Bar02b, DGH08, GRIR06, GBW^{+}07, HH00a, HDRC^{+}08, PC05, PCW05, TC03c]. chick-brooding [AFHA05]. chick-rearing [DGH08]. chicks [CWC00, GCLC^{+}02, MSH08, ÖBFO01]. Chikugo [ST07, SKO^{+}08]. Childress [Ike08]. Childs [Ano03c]. Chile [GEM02, APM^{+}01, BH06, CDO08, CC04, CGF04, CSH00, DDQ^{+}00, FMCB05, FC00, GLFU^{+}04, GSFN00, GGM^{+}00, HQDH06, HK04b, LNV^{+}05, LTNC07, LC02, MC01a, MFE^{+}05, MVLM^{+}06, NNLV06, OW02, PCL^{+}09, PPL^{+}02, RWE04, RT00a, SSGJ00, SAB^{+}02, TDC^{+}03, VHT^{+}05, YCSS06]. Chilean [ANCC07, AO01, FH08c, Rod03, WKN^{+}03, ZBNW09]. chilensis [CSTC01, GEM02, KKKK03, MC05a, MC07, VN02]. chimerism [BGGR02, BSDR01]. China [CCL03, JCL09, LGZ^{+}09, LZL05, MOM^{+}02, NLCS05, SAW^{+}01, TKVL04, ZTW08, HGG^{+}03, Lam00a, Lam00b, Lee00, SHZ^{+}05, ZYZ^{+}06]. Chincoteague [LLW09]. chinensis [KS08]. Chinese [HCB02]. chinned [CWC00]. Chinook [HWBZ05, HFWW05]. chinstrap [IBB^{+}07, TDT^{+}03]. Chionoecetes [DSMBG03, RGT01, SML02a]. chirodropid [BPC09]. Chironomidae [GJT09]. chitobiase [OBB00, SD09]. chitobiase-based [SD09]. chiton [NW06]. chl [JHJ^{+}03, ZER^{+}01]. chl-a [JHJ^{+}03]. Chlamydomonas [EKN08]. Chlamys [BG04c, MBH^{+}09, PW08, PW09, ZYZ^{+}06]. chlorinated [SCH^{+}07]. chlorine [WS04b]. Chlorophyceae [LS05, Ste04b]. Chlorophyll [ZER^{+}01, BH04, CSD04, FR06, FK04, FBHB03, GBHO^{+}08, HST^{+}09, HHAH00, HHKH05, KKA^{+}06, KW00b, LG01, MFS07, MTCU08, Pah05, PAK^{+}00, POH^{+}02, SSN^{+}09, SC01, SdSC01, SMF00, WKN^{+}03, WPTS07, YH00, YVB^{+}02, YMC05, ZT07, MGBB06]. chlorophyll- [BH04, CSD04, HHAH00, HHKH05, KKA^{+}06, MFS07, SMHL00, SdSC01, WKN^{+}03, YVB^{+}02, YMC05]. Chlorophyll-c [ZER^{+}01]. ‘Chlorophyll-like’ [MGBB06]. chlorophyll-nutrient [Pah05]. chlorophylls [ZRG00].

Chlorophyta [CMJ00, PP02, TMM05, ZM06, jMFVPL05]. chloroplast [JHL00]. chloroplast-retaining [JHL00]. chloroticus [LB01, PS06, Sew05, Wal07, WGL03]. choice [BLDG01, EÖM08, FP08, GW06, NLK00, SHNV09]. cholesterol [GWBdN09]. Chondrilla [UST^{+}04]. choose [UÓ03]. chordate [ATT06a, TBAT02]. Chorus [CGT^{+}00]. choruses [RJTM08]. Chromis [ACH07, BTKG^{+}08, Gla07, MAW06]. Chromophoric [RNDF^{+}04, SZBZ04, SNCP04, URFA06]. Chronic [APW^{+}04, PHB02, THJK06, CNW07a, Gil02b, JNDL02, LGFM00, PMG^{+}06, SW03, TGM07, WKPK05]. chronographic [PTCH01]. chronologies [Bla09]. chronology [FVAB00, HGAS^{+}01]. Chrysaora [DBH^{+}07]. Chrysoblephus [GKAS08]. Chrysochromulina [GLS^{+}00, SH01, ZER^{+}01]. chrysocome [RTPS07, SR04a, PISL02, RTPS07, SR04a]. chrysolophus [BC02, DGH08, TGT^{+}06]. chrysostoma [CPP^{+}04, NLP^{+}01]. chrysurus [WMG02]. chthamalid [DMM^{+}03]. Chthamalus [BCABC00, CCB^{+}09a, DMM^{+}03, JC08, LTPL05, TCW^{+}08]. Chukchi [EKN08]. chum [FWD^{+}01]. cilia [BDDL03, RL07]. ciliary [RNL00, RGA^{+}02, RL05a]. Ciliate [VPSO01, BKR04, BDÁM04, GT02, GJ03, Kam00, KS06, MR04, MRB04, OHS^{+}04, VF09]. Ciliate-generated [VPSO01]. ciliated [EF07]. ciliates [BJJ01, LDWC05, NOE^{+}04, PG00, PBW^{+}07, SJK^{+}06]. ciliatum [AM02a, SBR09]. cinctipes [Don04, Too04]. cinereus [LIT05]. Ciona [MK03a]. Circatidal [LDT07]. circuit [SM00b]. Circulation [ABWK06, ZZT05, BWP^{+}06, DHRG05, NW02, SVG07, SR01, TMFF00]. Circumpolar [AAJ07, PWFC08, RHHH09, WAN^{+}01, ASP^{+}08, BPT07]. cirri [Mil07a]. cirripede [dSSC07]. Cirripedia [DP09]. clade [Gou07, PBS^{+}05, WBZ^{+}09, RLKW04]. clades [LFP^{+}04, LWB09]. Cladocera [VV04]. cladoceran [ASC06, LG08a, Vii07]. cladocerans [KSSH02]. Cladocora [RMRAFP08]. Cladophora [SRS05]. clam [ASS05, BM07b, CAS05, CDFRL03, DD06, FKM^{+}01, GDL04, GFL02, Hun04, Hun05, LEB07, LBD00, Mar04, MM05, MDG04, NW04, NWFV06, ÓUA05, PTP01, RL03, SMHC03, Tal01, TK01, TE00]. clams [ARB04, CGP08, DSS^{+}09, GB02a, HGC^{+}07, LBC07, NTGK09, SMT08]. clan [MWR07]. clarity [Zam07]. class [BRI^{+}07, BM00, DUW08, EDBvdL08, Lat07, LG00b]. classes [AML^{+}03, OGL09b, PTD05, VGW00, WÓ02]. classical [MRHP01]. Classification [DFW^{+}03, Cos09, EYH^{+}06, HD06b, RC08]. classified [SDH01]. clathratus [FA02, LTCP03, SMF^{+}02]. clausi [BGF^{+}06, CT05, CAT^{+}06, FGM^{+}00, GFC^{+}02, LQRGT06, MGZ^{+}03, TT05a]. Clausidium [CS03b]. Clavelina [dCLLT^{+}02]. clavigera [BW04]. clavus [GLB^{+}03]. claw [Smi04]. claws [OP08]. clay [AGB03, NTH^{+}02, SLT^{+}01]. Clayoquot [DD01]. Cleaner [GL02, BG02, CC03a, CC03b, Gru00, SHK05]. cleaner-seeking [CC03b, SHK05]. cleaning [GML02, Gru03, SFH00b]. clear [HSP^{+}08]. Clearance [HMKR05, Bay04, LP01b, PBS^{+}04, Pet04, PK05b, Rii01a, SJPC02a, Wid01]. Clearwater [ARGS01]. cleavage [VWF^{+}01]. client [BG02]. Climate [Bla09, Den08, KBL^{+}07, LBPH06, MJ07, MSJ^{+}09, PT04, SHW09, WIDM09, AS04, AS06a, BBC^{+}09a, BGK09, BPB07, BDJ09, BCP^{+}08, BPA^{+}08, CH06a, CCL^{+}08, CSMvE04, EBR^{+}02, FBBV09, FH03, GHL02, GB09, GCVR01, HANM^{+}05, HSM^{+}09, HRCB02, HBRJ06, Hip08, Hip09, Hue08b, ILL06, Jen05, KAAH05, KS09, MHSS08, MS04a, MHT07, NLHO09, OAS01, PZ07, RCL^{+}09, SHB03, SGCD01, SOB^{+}06, SB09, Syd09, TdLvdVR08, TDTR03, TSF04, WH08, WFT^{+}08, WCM08, WDB^{+}05, YMC05]. Climate-driven [Bla09, HSM^{+}09]. Climate-forced [WIDM09]. climate-induced [SHB03]. climate-oceanic [MHSS08]. climate-related [AS06a]. Climatic [FFML02, AM03, AR02b, Bai00, BI04, BRS07, CCL^{+}05, HRR^{+}07, HSO04, LGC^{+}06, RRH09, SAB^{+}02]. cline [BA01, BFE^{+}03]. clines [MC06]. Clinocardium [SBR09]. Clinocottus [DL02]. Cliona [CU06, SL05a]. clip [KDAK01]. clipping [KDAK01]. clocking [ATT06a]. Clonal [BVP^{+}06, AAHDS06, AM04, DDD08, HR02a, SZ08a, SZ08b]. Clonality [KDM05]. clones [TLLKC04]. cloning [SSG^{+}00]. close [CWC00]. closed [BSVM^{+}05, FF02, MJS07, SH06c]. Closely [LBH^{+}04, FSHB09, GB02a, VSBB03]. closterium [Rij03]. closure [MGW^{+}09, NYK03, WWMF06]. closures [LHT^{+}02]. cloud [KBT05]. Cloudy [MCE^{+}01b]. clump [CYC^{+}09]. Clupea [Bar07, CMB^{+}09, CKTE01, Enb05, EH04, FFG03, JVT01, MST03, TFSJ00, VMCG09, VFV01]. clupeid [MVH^{+}06]. Clupeidae [PDM^{+}04]. clupeoid [SZGS08]. clutch [LF02, RPC^{+}06]. clutches [MH00]. Clyde [AM03]. Cnemidocarpa [STE04a]. Cnidaria [BPC09, CLR^{+}06, FGO09, PCS09, RCR03, RZEACGF00, STRD02]. cnidarian [CSG01, DPGN08, RLKW04]. cnidarian-algal [CSG01]. cnidarian/dinoflagellate [RLKW04]. Co [Hen09, PPBS02, TDvdW^{+}01, ADF^{+}09, CAT^{+}06, CCK03, CTvdM^{+}08, CKS05, EPM06, HD05, MSK00a, PBVP09, PC05, PSC05, RSN02, SJK^{+}06, VPS08, WG02b, KES00, CGCMPA01, CCB^{+}09a, DHT^{+}08, FKSW09, HLD^{+}07, IHK08, KFU^{+}08, KS04c, Kur08, MMC^{+}07, PZ07, RR05, SHL^{+}03, TDSM02, WDM^{+}09, GFL02]. ’Co-consumption’ [KES00]. co-cultured [PBVP09]. co-existence [PSC05]. co-existing [EPM06]. Co-limitation [TDvdW^{+}01]. Co-occurrence [Hen09, PPBS02, ADF^{+}09, CTvdM^{+}08, HD05]. co-occurring [CAT^{+}06, CCK03, RSN02, SJK^{+}06, VPS08, WG02b]. co-ordinated [CKS05, PC05]. co-variation [MSK00a]. coalescence [SAHF03]. coalitus [Sha00]. Coast [BA04, AP06, ANCC07, AO01, AB06, BRGS^{+}07, BBBSPH03, BMB02, BRP07, BAB04, BLHB04, BAMJ09, CL05, CK09b, CK05, CRT^{+}08, DP01, DCAL^{+}02, DCF^{+}05, FCP02a, FS02, FZT^{+}05, GTE^{+}09, GRPE08, GLS^{+}00, GSRH02, HKZ04, HMK06, ILL06, JNMB08, Joh02, JNFL04, KJMP05, KMWH04, KUN02, KCEB09, LFRG^{+}05, LNV^{+}05, LTNC07, LNB^{+}03, LLSL06, LÅCN01, MTL^{+}03, MVBGM04, MP06, McQ05, MPE03, PCSS01, PDM^{+}03, PMG^{+}06, PPL^{+}02, QAA^{+}06, RWA^{+}02, RS03, Rod03, SHH09, SPP^{+}09, SSW08, SYCG05, TSE^{+}08, UB01, VHT^{+}05, VGMC01, WDT^{+}08a, WGV^{+}02, WCRB03, XBW07, YW04, dSSC07, NAB^{+}06, RBG01]. Coastal [ADS^{+}09, DSJT^{+}06, DUSJ^{+}09, NZM^{+}08b, PPV^{+}06, AD00, AHC^{+}06, ÁSG04, Ant06, Arr05, ADR^{+}05, BCL03, BDR06, BA00, Ber00, BF04, BR08, BGKN00, BRG^{+}02, Bra08, BPA^{+}08, BK06c, BDÁM04, CHLV07, CCMK03, CBAd00, CMRF05, CLL^{+}09, CPH^{+}02, CWM^{+}09, Clo01, CAS00, CABS02, CRC02, COSM03, CHT09, CSD^{+}07, CHV^{+}09, CMB^{+}05b, CVT09, CRZ04, DP02, DSPHV04, DGBS08, DR00, DRB^{+}01, DES^{+}00, DL01, DAG^{+}00, EB04b, EvBB^{+}09, EOK^{+}08, FD00, FH08b, FNDP08, FOH^{+}08, GÁSPR03, GN04, GVKF03, GDDM05, GMK^{+}03, GEM02, GB03, GLÅ^{+}03, GMTN00, GT02, GWO02, GK03, GBW02, GPM^{+}09, GGRC^{+}08, HDC^{+}08, HKM00, HMH^{+}08, HHD^{+}09, HGG^{+}03, Hin02, HWBZ05, HFWW05, HCBP05, HCH09, INH^{+}09, IG04, JHTV02, JK06, JLW04, JS08b, JLJ01, JKES07, KD00, KKL^{+}03, KHD07, KAR04, KOC^{+}07, KHK^{+}06, KMR01, KM03b]. coastal [LBvdH03, LC02, LMC08, LHB09, LWH^{+}04, LHT06, MH08a, MFS07, Maz04, MAT01, MSA07, MCD^{+}05, MTLT02, MGTK00, MW04b, MHY^{+}06, MAW06, Min02, MHABGC04, MSF04, NE01, NEG01, NLB08, NCÁSBP04, NCÁSGP05, NW02, OAGV07, OLWB06, OÁSG^{+}09, PZ07, PTD05, PDT09, PBBC06, PNCS^{+}08, PNCT^{+}08, PEdlP^{+}02, PW07, QRWJF07, RSK04, RJTM08, RLBB04, RWE04, RGCC^{+}08, Rol00, RGT06, RLDL02, RMPC08, RHGB02, Saf03, SSN04, SM01a, SHB03, SLC05, SMHL00, SRGS01, SRZZ01, SRN02, SGL02a, SSM05, SP01a, She09, SHZ^{+}05, SWL^{+}06, SABK03, SS06, SDRH01, SZBZ04, SMSJ04, SOMM06, SCS^{+}07, SLHT05, TFA^{+}08, TSF01, TAÁO^{+}03, TL05, TBID09, TMFF00, VSRGW09, VMH^{+}08, VPS^{+}02, VVTH07, WDX01, WPC^{+}03, WDSA09, WRH05, YYK^{+}08, YYL^{+}00, YCSS06, Yin02, YF04, dSSC^{+}08]. Coastline [JMM^{+}07, vdMPEM08]. coastlines [BHR08]. coasts [AMT03b, FNB07, GhGD^{+}02, MMGC06, SAC^{+}00, SBSRW02]. coating [KES00]. cobalt [BTD^{+}04]. cobble [HKK^{+}06, RT00b, SKR09]. coccolithophorid [CLS^{+}01, CWN^{+}07, SHL^{+}03]. coccoliths [LNJ07]. cochlear [PB00b]. Cochlodinium [JTLG09]. Cockle [DdMB04, BD06, CWR06, Mou04, SBR09, UNII01]. cockles [KvKB06]. Cocos [KP01b]. Cod [CRV03, CDML06, JBT^{+}07, SOB^{+}06, AP00a, BM00, BM02b, BM02c, BF08b, BCLS06, CSB03, CHH^{+}05, CSMBH08, CDH^{+}09, CCL^{+}05, CB05a, CCCH01, CPG^{+}09, COSM03, EOK^{+}08, FRvdMF02, FBSG02, GBR^{+}04, GJB04, Hel00, HVW^{+}07, HRMH07, HJG06, HRC^{+}07, HMHB03a, HMHB03b, HNMC09, HCR01, JTN09, JLCA01, JCG^{+}02, JKES07, KFL09, KOC^{+}07, KLW^{+}09, LGB03a, LGB03b, LGB^{+}04, LAK07, LG02, LHB^{+}06, MHP^{+}00, MB02b, MSF^{+}09, MBS03, MBS07, MGSJ08, Mun07, NB06a, NM04, NWHH^{+}09, PMS^{+}08, PTM08, PBCB03, PKLD09, RR06, RGL06, RW08, RHSB08, SSN04, SHNV09, SSL00b, SGCD01, SMM^{+}09, SRJ07b, TKK05, UP01, VJKF07, WPIB09, YW04, vdMM03, vHGH01]. cod-capelin [CB05a]. Codium [SG06b, SS05]. coelestis [PKM04]. coexistence [CL05, DH09a, TLTB03]. coexisting [CNPS05, RvdPHW04, SZGS08, vdBÓME01]. coffeus [PD05]. cohort [CCC09, HSO04, MBL07, SLB07, WCM08]. Cohorts [RT00a]. coioides [APW^{+}04]. Cold [DFR08, OGL^{+}09a, SWHSA05, AMB08, ATLC03, ACPS09, AWP^{+}01, BD03, BDJ09, BEKT09, CLO^{+}09, CY06, DBOR09, DLB04, FA08, FH08c, FKM^{+}01, GN04, GGDB09, HAMW09, HGC^{+}07, HDG^{+}08, ISSK03, JHB06, JNFL04, PTD05, PDT09, PBL^{+}09, RDH^{+}09, SDHG03, SP02, SYH^{+}09, TNHL02, TDTR03, WKMK06, WMW^{+}08, YFH^{+}02]. Cold-blooded [SWHSA05]. cold-ocean [TDTR03]. cold-seep [FKM^{+}01, ISSK03]. Cold-water [DFR08, OGL^{+}09a, AMB08, ACPS09, BEKT09, CLO^{+}09, DBOR09, DLB04, FH08c, GGDB09, HAMW09, HDG^{+}08, JHB06, JNFL04, PBL^{+}09, RDH^{+}09]. ’coldspot [Pri02]. coldwater [CPG^{+}09]. colias [RW08]. colimitation [PO09]. collapse [BVC^{+}02, CDH^{+}09, JCG^{+}02, WGF09]. collected [BHMR01, LLSL06, WWL04]. collecting [Eva08, RNL00]. collection [GK00b, KST^{+}08, Rob06]. collections [MS05]. colloid [WG00]. colloid-bound [WG00]. colloidal [Bur01]. Colloquium [PIM^{+}05]. Colne [DTNU00, ABBN09, KUN02, TDUN07]. colonial [BSDR01, HKK^{+}06, HH05b, MPK03, SNCP04, TLLCT02, VHMB09, ZPP09]. coloniality [HF03, Pra05]. colonies [BL00a, GDR^{+}04b, HPH^{+}01, Lid08, MAVO^{+}06, WHP00, ZE01]. Colonisation [Chi09, GGS09, GF02, PZMG05, SDM06, BCPB06, EPM06, GMVF08, NCP^{+}08, PH05, PHO^{+}07, Sta08]. colonising [HAC06a]. colonists [LTH^{+}06]. Colonization [LGQ^{+}00, TAS07, TVD02, BS08b, BP02b, CUB08, KAPTK02, Lir00, NCHP05, THP^{+}05, TCW^{+}08, WH02, WR00b, ZPP09]. colonizing [KT01]. Colony [BPG08, LGC^{+}08, MHK^{+}01, PC05, BPT^{+}00, Cum02, DBH^{+}02, FBD^{+}02, FFHW07, GAMB05, HAMW09, KS08, LWW^{+}01, Lir00, Mar00, MG08, Pra05, SAB^{+}02, WLB08]. Colony-based [LGC^{+}08]. Colony-specific [BPG08, PC05]. color [SCD^{+}05, SAH^{+}05]. colour [DSP07, FFYS^{+}07, SWD^{+}04, SDR01]. Columbia [DER01, DD01, EBG^{+}00, LWH^{+}04, MZJ06, CB00b, DMS^{+}05, ESDKM09, FE06, HRCB02, MDZ05]. columbiana [Iye02]. column [CS02b, Lar03, LDHC04, MGDP01, PC02a, PSGP04, RMWO09, SVC^{+}01, UÓ03, VJ04, ZRG00]. comb [BN04]. combination [MMA03a, MH06b]. Combined [Ano00a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b, PSGP04, Prz05, GFN^{+}09, HL05, MGC03, PEdlP^{+}02, TPGQ07]. Combining [FGLM06, LBR^{+}09, RB05b, FMW07, HMW^{+}00, SMP01]. come [CB08a]. commensal [GDL04, HS05]. Comment [BG09, BR07, Ber02a, BD11, CST08, Dun02, EM07b, GHP02, Kir02b, MH09, Pea03, She08, Som08, TBBI04, TW10, vWNYS05, BL07, RN06, Car05, GNAD^{+}04]. commercial [BT01, BD06, CWT00, GSM06, HDM04, HK04a, HK04b, JLCA01, Kro05, NSBK05, RJH02, SS04b, SAH^{+}05, TK01, VBFF08]. commercially [BLT03, NM09, QFJ^{+}07, SJ09b, Tup07, UB02, WHB^{+}06]. common [BBM03, BR07, Bar02b, BCH00, BPT07, BTD^{+}04, CD08, CU08, DFPG^{+}03, DC06, DMA02, GC02, IBB^{+}07, IAKS03, JRA02, JSK^{+}07, KR01, KS04a, LBS^{+}05, LG00a, LGD04, LC03, McI02, MPS^{+}08, MR07a, MR07b, NMBH02, OHN01, ÖBFO01, OÁSG^{+}08, Pet04, PAMJ03, RMT^{+}05, RE00a, SLC05, SLB^{+}08, SJ02, TVD02, TBFA04, VGB^{+}05, WWAC00, ZPP09]. commons [HRB^{+}06]. communal [Swa04]. communication [BK05b, JBW^{+}09]. communites [BHSR00]. communities [AP05a, AR02a, AAPO00, AAGS02, AI04, AMP^{+}02, APH^{+}04, ACPS09, AJ03, BS03, BAA07, BBC^{+}00, BL04, BFK^{+}05b, BF04, BR08, BSJ02, BM05a, BS04a, BNSG01, BMJ02, CGH08, CMD07, CDS07, CH03, CF08, CWE^{+}08, CK04, CLH^{+}07, CHP^{+}05, CS03c, DSPHV04, DB02, DTD03, DJ04, DDTQ05, DQW05, DAG^{+}00, ESS09, ESQ03, ENC^{+}04, FD00, FAJM07, FTO^{+}09, Fro00, GhGD^{+}02, GSMD00, GMC^{+}09, GJG^{+}03, GGC^{+}05, GZG^{+}07, GGG^{+}04, GT02, GBR01, GBC^{+}04, HF08, HF06, JDTB06, JPPW01, JNDL02, JHB06, JSMM06, KVS08b, KKH07, KT03, KSWP04, LBH^{+}04, LSH00, LSG08, LMW04, LBB08, LTFH06, LWB09, LBPH06, LTH^{+}04, LBC^{+}09, MFCM05, MKG^{+}05, MUT04, MGDP01, MBC^{+}07, MSS^{+}03, MGM^{+}07, MAG^{+}07a, MGW^{+}09, MdL06, MC08a, MHDS09, NLG05, NBB^{+}04, Nug02, OAGV07, OGA03, OPSB07, PKR07, PFZS^{+}06]. communities [PMGE01, PCJF06, PAMJ03, RS03, RGS^{+}09, RCAK03, RS06, RGP^{+}09, RW06b, SGB09, SAB04, SFW01, SC03, SFF^{+}07, SF07b, Sma07, SYA02, SAF^{+}09, SLL07, SLL09, SHFP^{+}03, SPCG09, SWGC^{+}05, Tal09, TMD^{+}07, TSF01, TGA02, TSSA03, TC03b, THJK06, TSJH05, UB03, UYNO01, VHT^{+}05, VSV04, VAP^{+}07, VD00, WBR07, WAMP08, WDB^{+}05, YFH^{+}02]. Community [BG07a, GS07, HAMW09, MCCT07a, MCCT07b, MHTJ07, MRB09, Van02, VBS06, YM06, AGCF^{+}08, APM^{+}09, AL09, AB06, BBABL01, BBA06a, BEU^{+}07, BRP00, BLHB04, BB00, BPS^{+}03, BSH^{+}09, BFE^{+}05, BH09b, BHR08, CHSW03, Cal06, CH07a, CC02, CSMK09, CMMN^{+}01, CRJ^{+}09, CÅJ05, CLJ^{+}08, CIHF01, CS02b, CZW^{+}08, CD05, CT06b, CT09, CHA06, CV05, DDQ^{+}00, DHHY09, DDL^{+}03, DB05, DHRG05, DBD^{+}05, DFT^{+}07, DDSR07, DKFV05, DAG^{+}00, DAA00, DH01, DLB04, DVLH^{+}00, Edm00, Ell01, EB08, EB05, EÖKS^{+}02, FGPS04, FHL^{+}09, FKA^{+}06, FF02, FLG03, FD04, GL00a, Gim04, GSRH02, GDF03, GK07, GGMGR07, GAM02, GLG^{+}05, GG09, HLD^{+}07, Har06a, HGS^{+}07, HCB^{+}07, HHIS06, HS08b, Her06, HWL00, HLTT09, HSB08, HB07b, HAD07, HJ07, JZL09, JHTV02, JHJS04, JLW04, JPPW02]. community [JCL09, JQ05, JFR02, KOTL07, KMP^{+}06, Kas00, KSSH02, KK02, KKIV01, KGHL^{+}01, LSH^{+}07a, LOT^{+}00, LBBM06, LYL^{+}07, LPK^{+}03, LHK01, LJM^{+}00, LCD^{+}05, LSM^{+}02, LTL^{+}06b, LUIAH03, LWMW02, MB05a, MRB04, MLW^{+}02, MTR^{+}07, MTH^{+}08, MB00a, MB01, MAJ^{+}06, MRC^{+}08, MARM06, MLW03, MASFS00, MVD^{+}08a, MFP04, MPE03, MDZ05, NDV^{+}00b, NvdVVD06, NLB08, NTGK09, NIF00, NTN08, NS07, OMP^{+}05, PH03a, PP00, PWHW09, PWB^{+}03, PDRT02, PK06, PCRR08, QASA08, RBJB^{+}05, RSN02, RT00b, RT00c, RBG^{+}07, RGDN02, SRL^{+}02, SFH^{+}00a, SD09, SEL02, SKR09, SL07, SH07b, SKR08, SSM05, SW03, SH06b, SBFP08, SS04b, SGHG07, SMSD07, SCD^{+}09, SPT^{+}00, SWG^{+}09, SH06c, SBC^{+}06, SOMM06, SMHG09, SFG05, SW08c, TS05, TPD^{+}05, Van03, WSA^{+}06, WAB^{+}02, WPC^{+}03, WBN^{+}00, WKL01, WH03, WHB09, WASW03, WWWC02, WDR^{+}05]. community [Yap09, dBMSL02, dJTD07, dVBC09]. community-mediated [MVD^{+}08a]. Community-wide [GS07]. comosa [CK09b]. compaction [HCH09]. Comparability [CJFM09]. Comparative [BGD^{+}03, CCM01, CRJ^{+}09, DSDS04, FGO09, HS03, KWD05, MCW^{+}07, NLCS05, OSP01, RW01, RGA^{+}02, SFW01, WJDA06, YLK00, CDK^{+}07, MGHA05, PGCC^{+}06, RRS^{+}09]. compare [BH08, BBDW09]. compared [CWN^{+}07, GDR^{+}04a, HTM01, MRHP01, MGD^{+}09, NHNB01, PKS00, TKR06]. Comparing [BSB03, SRB04, dSHCR07, TNTH04]. Comparison [BMG00, BDG^{+}07, BH06, BIC^{+}03, BI03, CSD04, CWT06, EF02, GLV^{+}09, HPVL05, LCHG03, MAM^{+}01, MHBB05, MJG01, OC01, PNK^{+}02, PTCH01, PJ08, RWR^{+}00, SPI^{+}07, TC01b, WNS^{+}05, ZTW08, Bay02, BRF03, CED^{+}07, CU08, CNT^{+}04, CWR^{+}03, DF02, FBD^{+}02, FS07, GVKF03, JBSB01, JBS01, KDAK01, LV02, MCC^{+}00, MP06, MA00, MWM^{+}05, NKK^{+}01, NLG05, PDM^{+}04, RHC^{+}08, SBF08, SP02, SSL^{+}07, STH^{+}06, SG02a, VFV01, VAJ^{+}03, VPG04, WMB00, XYH^{+}09, MC01b]. Comparisons [Bay04, GRM00, MOOB06, FBM07, KHHM07, KR01, Las03, PDS^{+}07, PPS00, SRDG05]. compartmentalization [CW05, RPS^{+}06b, RAAT^{+}07, WLL03, WL03]. compartments [DPD^{+}03a, KVV^{+}00]. compatibility [LC07a, RZLS09]. compensate [CT05]. compensates [Mil07a]. Compensation [AL09, SF07b, CRH01, KR05]. Compensatory [DSH03, HSSS05, SHK05, EH04, SO02]. competent [MVLM^{+}06, PPL^{+}02, RGL01, TT07a]. competing [BKB^{+}09]. Competition [BM02d, JK03a, LVZW06, NWÓ03, SJ09b, ZE01, ADF^{+}09, BK03, BK05c, Bar06, BM07a, BM05b, DDD08, Don04, EEA01, GSM^{+}08, GDR^{+}04b, HM04, JM03b, LC06, LJM03, LOB03, Mok04, Pet02b, PSC05, PC02b, RS07b, SML02a, SM09, SvNvdMvdV09, SB06, Ste04b, WH00, WÓ02, vWCAB07]. Competitive [FBM08, SG06b, BK04, DTAF00, JMA02, MP05, MN04, NB06b, PWLC06, PBFB08, SMM^{+}09, ZE01]. competitively [MN04, ZA07]. competitor [DDD08, JNH^{+}09]. competitors [BWO04, LWCF04, SF02b]. complementary [LW00]. Complete [BCH00]. Complex [RFLC07, BDF^{+}04, BHFC08, CML08, CTHF01, Gim04, HEN^{+}08, KMWH04, Kne09, LFP^{+}04, PB08a, RDH^{+}09, SS04a, SB00, TSRR05, TA01a]. Complexity [OLG^{+}09, AO01, BDC00a, BBE03, Bib07, BB00, CH09, CV05, DTD03, DHR04a, FSW09, FSJ03, GP04b, GS05, GDW06, GBNÖ08, HAC06a, HJK01, HMJ^{+}09, HFM^{+}02, PTH08, QS08, Raf06, SH06b, SMGCW00, TNTH04]. complexity-the [BB00]. compliance [MGW^{+}09]. complicated [RGVC07]. Complications [BCS^{+}06]. component [GAB^{+}03, Kev01, SS07, SGCD01]. components [BCM^{+}07, JQ05, Kaa04, MUT03, PMGE02]. Composition [GVKF03, GSFN00, HLGCO06, HH03b, MGDP01, AM03, AMR00b, AMR00c, AVNN05a, AVNN05b, Bel07, BVH02, BRP00, BD11, BCB08a, BFE^{+}05, Bro06, CTA^{+}03, CDFRL03, CH07a, CSMK09, CKO^{+}07, CDT^{+}03, CMB^{+}05a, CTFC08, DHBZ00, DTV08, DTV09, DK06, DPGN09, EL05, ESDKM09, EBG^{+}00, GPB02, GKWK02, GNAD^{+}03, GNHB^{+}05, GRM00, GBNÖ08, GTSS03, HKZ04, HAMW09, HNYI07, HK03b, HS08b, HMY^{+}08, HPB^{+}04, IFL02, JHJS04, KF02, KKTN05, KTK^{+}02, KSK07, KWL^{+}05, LPK^{+}03, LLJ05, LTFH06, LPBD00, LSHL03, Mag02, MLW^{+}02, McQ05, MBR^{+}05, MO05, MRG08, MLW03, MDZ05, MC08b, NM07, NGMT01, OTB^{+}00, PDO01, PDT09, PWB^{+}03, PGS^{+}00, RBJB^{+}05, RHG04, Reb01, RFPS^{+}02, RMH^{+}09, RGF^{+}03, RBGM06, RMC^{+}07, RW06c, SMM^{+}04, SLCB00, SKR08, She06, SHS^{+}05, SABK03, SDH01, SVC^{+}01, SCD^{+}09, SHH08, SJ04, TSW^{+}04, TGGA04]. composition [TBBI04, TBH08, TSEH08, THJK06, TFLA03, TDSM02, TGNT05, Ven06, Ver05, VBS06, WR00a, WJWL05, WHD^{+}05, WA05, WHL^{+}06, WDR^{+}05, YI03a, dVBC09, vBWSS06, vDvHT^{+}09]. compositional [HHT^{+}09, Thr09]. compositions [EdPP^{+}00, HYR^{+}09]. compounds [LSdM^{+}03, LELC02, MGBB06, TL08, Zam03b, Zam07]. Comprehensive [AIM^{+}05]. compressa [MBLA01]. compression [Zab09]. comprised [BGGR02]. compromise [FKSW09, PHP^{+}03a]. Compton [BD11]. Computation [FFYS^{+}07]. Computer [SDW07b, RDG^{+}07, RGD^{+}08]. computer-aided [RDG^{+}07, RGD^{+}08]. Concentration [BMS04, APP^{+}09, BCB^{+}06, CSD04, DCD^{+}03, EG05, Fin00, FR06, GMZvdL07, HSGS00, HL05, KTK^{+}02, KMAC^{+}02, LZM^{+}03, LT00, LP01b, LSG05b, MBLA01, PPB^{+}09a, PTP01, RSR02, SH01, SFPGA05, SH07c, Ste04b, WKN^{+}03]. Concentration-dependent [BMS04]. concentrations [AHK^{+}02, BPT^{+}00, BCB02, CAMC07, DSF^{+}04, EG03, GSRH02, GHV05, GLCL04, HST^{+}09, HHAH00, HHKH05, KWTR08, KB05, LSdM^{+}03, LLL^{+}09, MCP^{+}06, MG00, MLR02, PW08, RWJ02, SH03, SCS03, SMB^{+}03, SH07d, SMVF03, TGSW06, TP00a, VN02, VN05, WS09]. concentrica [RDC06]. concentricus [MB07b]. Concepción [SSGJ00]. Conception [BMG02, CDT^{+}03, TDTR03]. concepts [BSS05, Cos09, DKA^{+}06, HCS07b]. Conceptual [PL03, Clo01, Woo09]. concerning [CST08]. concerns [BN04]. conch [GL04b, Sto03]. Concholepas [MC01a, PPL^{+}02]. concinna [FCP02b, PTP01]. conclusions [PMGE01, GHP02]. Concordance [PVV^{+}06]. concrete [Lam00a]. Concurrent [LS02, SA06, KG00a]. Condition [LP03, AMG^{+}07, ALE^{+}09, BLH01, BPM05, COPB06, CS02a, DTV08, FHVMDS07, FFHW07, GCS06, GNC07, HW00, HCD^{+}09, JHL04, KKMH08, KHvW02, MCW^{+}07, MB02b, MSS^{+}03, MWTB00, MSP^{+}06, NPTW05, NT06, OA02, PDS^{+}07, PTM08, PWAT01, SSL00b, SJ05, SNY^{+}05, SB03b, SJ09b, TOBR07, VMCG09, VCBH06]. Conditional [HH05b, TDC^{+}07]. conditioning [SSS^{+}09]. conditions [AP05a, AMP^{+}02, AMR01b, AVNN05a, AVNN05b, ABWK06, BTL05, BPVL03, BPE^{+}06, Ber00, BFV08, BFE^{+}05, BPA^{+}08, CSA02, CCK^{+}09, CVÁM06, CES^{+}03, CNPS05, CEO06, DDL^{+}03, GVM^{+}05, GMP07, Gil02b, GJ03, HGA07, HNMC09, JNS04, KWV^{+}05, KCB06b, LHQ05a, LWJC04, LB05b, LJK01b, LS06, MNB^{+}07, MSF^{+}07, MTH04, MM03, MF03, Mon07, OSL07, PPV^{+}06, PW04, PJR^{+}03, PVGO01, PHS09, RRIH^{+}02a, RTPS07, RJBW03, RAH^{+}02, RC07, SSR^{+}09, SW08a, STF^{+}09, TFLA03, ULL^{+}05, VN05, VFM06, VHMB09, WCBD01, WW07, WLB08, WKOE09, Woo09, dZCVC01]. conductance [ML06a]. conducted [MJG08]. configuration [DVNvdV07, GC02, GC05, GBNÖ08, JMM^{+}07]. confirm [GDSB04, PTAR07]. confirmation [MFWD08]. confirmed [CMRB08, PPB^{+}01, WHCP^{+}04]. conflicting [BS07a, JS08a]. conflicts [AM04]. Confluence [HRK^{+}08]. confounding [BCO07, LMW^{+}01]. congener [DCT^{+}03]. congeneric [LYL^{+}07, LS09b, MHT07, QDDP06, SRDG05]. congeners [TG03]. congruence [KMP^{+}06]. conifera [ALPK00]. Connecticut [COK02]. connections [CC02, Chi09, GGS09, LPD09]. Connectivity [CK09b, Gil02a, Bel08a, Chi09, CEBR07, DVNvdV07, FSO08, FCS^{+}09, GAB^{+}03, GGS09, LHH09, PCC07, PBTB08, RRDHG08, SVG07, She05, SJ08, She09, SSW08, TCS^{+}09, TCW^{+}08, UDB^{+}09, USG^{+}08, WT09]. consecutive [ZIUG07]. consequence [GA03, vdVBW00]. Consequences [CLR^{+}06, DABC02, HLD^{+}07, HSM^{+}09, MSC04, PCL^{+}09, PvDMB08, Pra05, SHB03, She09, Tan02, Ark09, BCPB^{+}01, BDEM01, CC03b, CK09a, CMT^{+}08, CC05, CB08c, DRDD09, DCF^{+}05, DH07b, FMT09, GEM02, Gru00, JRA02, JTLG09, JVT01, Kap06, LGB^{+}04, MDLS01, Mou04, PMGE02, PUL07, Rod03, She05, Tar03, VBMBC08a, WH00]. Conservation [CLB^{+}09, GPSG06, AFPS08, BRGS^{+}07, BT09, BP08, BRSM07, DSJT^{+}06, FF04, GFB^{+}07, GB09, LIM^{+}06, MPR^{+}06, MNT09, OHB^{+}00, Pri02, PISL02, SDFM00, SP01c, SMCE07]. Conserving [GL04b, LAV04]. consider [BHZW08]. considerations [DFMD00, Kne03, PM00]. considered [FI09]. consistency [FWC02, HHG^{+}07, QMM^{+}08, SM07a]. Consistent [HWC^{+}06, MEHD07]. consortia [UHvO^{+}08, vDHHM08]. conspecific [CMGBJ05, DMA03, Fer04, NPLZ00, OKP06, Swa04]. conspecifics [FSC05]. conspicillata [BDF09]. constituents [CHW02]. constitutes [Sto03]. Constraining [MFP^{+}05]. constrains [PSKC04]. constraint [ECK05, GSB^{+}06, HHG^{+}07]. constraints [BS08a, CMHP06, CEO06, DMA02, EBT^{+}02, GSM05, GMD08, MCH^{+}06, Mil07a, PBB^{+}02, RCWYK07, TN04]. constructed [Cha03, CS03c, PPGL09]. construction [CHT06]. Consumer [HGJM09, DDSR07]. consumers [BCT^{+}08, CH03, CRGH01, CHV^{+}09, DSP00, KFH04, LWMW02, NFN03, SR08a]. Consumption [KJB^{+}06, Rod03, TKN00, BKM08, CM07, GHV05, Hyr01, KYML02, KVH02, KKP08, KYYM00, LM09, LDWC05, MSLP06, MH08c, MNC^{+}02, OYNO08, Paf06, PBCB03, PD07, PKA06, RAA02, SLS^{+}00, WH08, XZUK05, ZLAK04, vBSSH02, WJB09, KES00]. contact [GMA^{+}09, NL05]. container [CK04]. containing [JWTP06, TCD01, ZRG00]. Contaminant [BML^{+}08, PEG^{+}00]. contaminants [AHL^{+}01, FHG^{+}03, HBT^{+}08, KR01, SCH^{+}07]. contaminated [ABB^{+}00, CTLR03, GGHA02, RSL09b, SRDPT^{+}07, WYFX02, YLL^{+}04]. contamination [ARBF02, BMB02, BGS02a, DL03b, HVH^{+}07, KGS^{+}08, LMW^{+}01, MTL^{+}03, RCC01, RGEB00]. Contemporary [WBC08]. content [BGLL09, CF09, EP01, FKBB01, HNYI07, HGJM09, HBK^{+}05, IFL02, JF02, KWWH02, LG00b, PDT09, RHG04, SN06, SNY^{+}05, SHS00, SP00, THN^{+}03, TJ02, WKOE09, dlMPN^{+}03]. contents [CBD04, CvBGR09, DJM^{+}05, FC08, MGA^{+}01, PJN01, PJN03]. Context [RGMH08, OCB^{+}07, Orb08, dSCV^{+}08]. Context-dependent [RGMH08]. contiguous [BBT02]. Continental [BM07c, RWB^{+}09, BN09, BHD04, CSP^{+}03, CAA07, DFR08, ESS09, EHWS07, Ell01, Ell02, GBdS^{+}05, GL00b, GL00a, GMd^{+}02, HGT01, KH07a, Kor01, LG02, LF06, MLC08, MBB^{+}04, NBB^{+}04, NLHO09, OJPM01, SSHW^{+}07, ST06a, SC00b, SCAF00, SCAF03, ULL03, VCA^{+}09, WKL01]. Continental-scale [RWB^{+}09, ESS09]. contingents [KS04b, SRZZ01]. Continued [RWA^{+}02]. Continuing [SB03a]. Continuity [MCP05]. Continuous [AG00, SRR06, FPC05, KJRG09, vWCAB07]. continuous-time [KJRG09]. continuum [APA07, KH00c, dlMPN^{+}03]. contradictions [RZWG08]. Contrasted [CH07d, HAD07]. Contrasting [Bel07, FPB^{+}07, HPH^{+}01, JM03a, KAM^{+}06, MMA^{+}04, SFT^{+}05, VK05, dCLLT^{+}02, BPE^{+}06, BMO^{+}04, CCK^{+}09, CLL^{+}09, DHR04a, DM09, EFL^{+}09, FDDS09, FSHB09, FSV^{+}08, GSMD00, HTZ^{+}07, LBPS03, LB05b, LP07, LTL^{+}06b, MM06a, MHB^{+}01b, MB00b, OC01, PPV^{+}06, SPAMG00, TRT04, WDX01, vOSF^{+}09]. Contrasts [FD02, SO02]. contribute [DCO^{+}08, DGC^{+}05b, KJL03]. contributes [Gil02b, WT06]. contributing [IFL02]. Contribution [GYD04, HH03a, Mil07b, MHDS09, PFW^{+}02, SFKT09, BDGK^{+}06, CGCMPA01, CLP08, CCB^{+}09a, FMMS03, FH08b, HSCD01, LSE06, MdL06, SGL^{+}02b]. Contributions [KCP^{+}07, BAOV06, BSP03, CMG04, GÁSPR03]. contributors [VLK04]. Control [HRK^{+}08, KGHW05, MIL^{+}00, TNHL02, Bai00, BVI^{+}05, BH09b, COK02, CMLG07, CKO^{+}07, DDTQ05, DDSR07, EDE05, EFL^{+}09, HBBD09, KN00, LR02, MLW^{+}02, MPDS07, MHHS06, OÁSG^{+}08, PMH07, QAA^{+}06, QA01, ROCB09, SP01a, Som00, Sot03, TGY^{+}06, TSJH05, VSL^{+}04, WHD^{+}05, Wie05, WLJ08, ZHWCS08]. controlled [AVNN05a, AVNN05b, Fen08, GH08, JNS04, LSK08, SEH07b]. controlling [BFB^{+}09, CA06, EAF^{+}05, HDCA06, MSS^{+}03, PDQ01, PCJF06, SW02a, VLP^{+}04, WAW^{+}09, dBMSL02]. Controls [Bri03, ABF07, AW03b, CMHP06, CMSS09, DP02, HM04, HHM^{+}06, JFD09, MCH^{+}06, MDBC03, MTS07, MSH^{+}03b, MVB^{+}05, NW05b, OSF08, RL07, SBB09, Seu05, SW02b, TKGS04, TSJ05]. conundrum [HCD^{+}09].

Convection [BHW^{+}03]. conventional [RRS^{+}09]. Convergence [KH00c, TBI08, WRW06]. Convergent [SPRK02, HSPH05, SRL^{+}02]. conversion [JWM09]. Cook [AP05a, Gar00, GRRH07, RHC^{+}08]. cookei [DS09a]. cookii [RHC^{+}08]. cool [GF02]. cooler [JM01]. Coos [MS05]. cope [MMCF07]. Copepod [CLS00, CSA02, CVÁM06, ITE^{+}04, Lei01, OLWB06, SDS^{+}03, SWW02, WB06, WB07, AMR00b, AMR00c, AMR01b, BGF^{+}06, BI02, BD02, BK06b, BGS^{+}02b, BKR04, BSR05, CT05, CWT^{+}01, CBD04, CA06, CK09a, CD03c, CRB07, CS03b, DHCH04, DD09a, DD09b, EC07a, ESDKM09, EBG^{+}00, FGM^{+}00, GT02, GRM00, GMR^{+}02, GFC^{+}02, HK03b, HM01, HPR05, HW08, HB04, HDP08, IHW^{+}01, ISG04, JWG01, JB07a, JDCB09, JP04, JP08, JTLG09, Joh03, JFA02, JNH^{+}09, KHHM07, Kiø03, KVS08b, KB03, KKP08, LNJ07, LSdM^{+}03, LG08a, LWH06, MB05a, MdMR04, MPE03, NMS^{+}02, NMVH06, OSA05, OBB00, PIM^{+}05, PPB^{+}09a, PDRT02, PWL^{+}06, PK06, PI08, RRS05, RBRSP00, SSB09, SHT^{+}04a, SK03, SAHK06, SKU^{+}08, SGH^{+}06, SNGR00, TD01a, Thi03, TKTJ07, Tit01, TK03a, TSRR05, TNT06, TIM^{+}01, TLNH01, TBH06]. copepod [WTP08, XW01, XW03, Zam02]. Copepoda [BC01b, KI01, LQRGT06, MRZJ05, PIY04, RVMI^{+}03, SHT^{+}04a, STA^{+}05, TSW^{+}04, TSW07, TBVJ08, TCC04, WJWL05]. copepodite [TT05a]. copepodites [BCB^{+}08b]. copepods [Ano02i, AG04, BK05a, BK05b, BRIP00, BIL01, BIL^{+}02, BI04, BP03, BJJ01, BH04, BLH02, CAT^{+}06, CID^{+}06, CR02, DQ05, DFV01, FPDS^{+}02, FR06, FSSW^{+}00, FI09, GK08, Gor03, GEHZ07, GGTKG09, HH01, HKM01, HK02b, ISYM06, ISY07, Ike08, IP07, JKS^{+}01, JKS^{+}07, JP04, JPL04, Joh07, JPNM^{+}06, KSSH02, KBT05, KH08b, KSKL^{+}03, KBSWG09, LEÖK^{+}02, LTNH00, LSBJ08, LVA07, MGZ^{+}03, NOP02, OHS^{+}04, PL00, PK05a, Paf06, PKLD09, PLC^{+}02, PK05b, PK06, Reb01, RFB07, ROBFR08, SNH04, SSS^{+}05, SKK^{+}06, SVL^{+}04, TS04, TNT08, Tit01, TF04, VSL^{+}04, VTS^{+}02, WB06, WWWC02, WWWY07]. Coplanar [DCT^{+}04]. Copper [HJM07, CCB06, CR02, DL03a, FMCB05, LMC08, LP01a, PJP06, PJ06b, RGJ^{+}04, TAW^{+}09, TP00a, XWQQ05]. copper-tolerant [RGJ^{+}04]. coprochaly [IP07]. Coprophagy [PK05b, IP07]. Coprorhexy [IP07, PK05b]. Copulation [KMF02]. Coral [ACR07, BMM^{+}09, DY05, FAJM07, Fox04, JBF08, JM03b, Lou08, NReZW09, NDB04, WYB^{+}04, WM05, AR02a, AE02, ACT09, AAGS02, AS06b, APA07, AH08, ATR^{+}08, AWS^{+}09, ACH07, APH^{+}04, ACPS09, AC08, AM04, AM06, BBM03, BCLRL07, BF07, BGGR02, BG04a, BHH05, BJC02, BP08, BL00a, Bel07, BPM05, BMWH08, BBC^{+}09b, BSL02, BB02a, BES^{+}08, BCKP09, BHY09, BDW^{+}00, BDDG02, BD08, BH02, BH09b, BHSR00, CBGLM08, CH07a, CM07, CLO^{+}09, CHM06, CWE^{+}08, CSM02, Chi02, CCC05, CSvO09, CCB^{+}09b, DSL07, DFY00, DB03, DB05, DBT^{+}02a, DC07a, DH07b, DBOR09, DGNvdV05, DGC^{+}05b, DVNvdV07, DWL09, DLB04, DFR08, DS07, EJM01, Edm00, Edm04, EMB^{+}09, EFL^{+}09, EEWBdN08, FPBAT02, FOL02, FIRL02, FB03a, Fis05, FBM08, FB07, FSHB09, FHGJB04, FH08d, GJ05, GYLÖ06, GPSH01, GPLAG04]. coral [GLB^{+}03, Goc04, GAZ02, GEC08, GDKN08, GCS06, GL02, GGDB09, GB06, GNC07, Hal04a, HWC^{+}06, HW09, HCH07, HAMW09, HNYI07, HKY^{+}02, HSS09, HSP^{+}08, HM05a, HM05b, HF06, HSRF02, HR07, HM04, HGMG^{+}04, HSB08, HOVG07, Hou09, Hue08a, HWG06, JS01, JK03b, JBHGB04, Jon04, JSMM06, JNFL04, KAMR00, KKG06, Kav05, KHIA02, KJ03, KKB^{+}06, KSR^{+}07, KN00, KWHS^{+}06, Kuf01, KWB^{+}06, KKSR05, LBH^{+}04, Lam00a, Lam00b, LWCF04, LC06, LG06, LCEC02, LCE03, LHH09, LAK^{+}08, LB02, Les00, LFRR^{+}07, LDS^{+}04, LWK^{+}07, Lir00, LC09b, LLCHG01, LVZW06, LRM00, LKW03, LSS^{+}09, MRN07, MHO07, MSG07, MH08b, MB07a, MBHN07, MBLA01, MG08, MSS^{+}03, MMMPB05, MBA^{+}05, MG05a, MGM^{+}07, MAG^{+}07a, MGW^{+}09, MAOP09, MAW01, MHK^{+}01, MDW^{+}09, MVK^{+}09, MNT09, MSUK01, MCN09, MAG07b, NDV^{+}00b, NvdV02]. coral [NRvdV^{+}02, NBN^{+}08, NYvW03, NvWY05, NHS^{+}08, NMDS00, NNLF09, NTN08, Nug02, NB06b, OYNO08, OP09, OM04, PS08, PR09, PBAA04, PFZS^{+}06, PKH00, PV05, Pra01, PKS07, PR05, PBL^{+}09, RBRH02, RR05, RFPS^{+}02, RFPB^{+}04, RMH^{+}09, RCAK03, RV05, Rit06, RDH^{+}09, RFB07, RM06, RSAK02, RL05b, RL08, SAB04, SSL00a, SPL00, San00, SCMB03, SDHG03, SBvD05, SS00b, SKR08, SPBM05, SC06a, SFL05, SBL08, SHS08, SFPGA05, SYWM05, SP08b, SRS05, SWD^{+}02, SP04, SGCP06, SLHR09, Sta09, SJ04, TAS07, TAA04, TAS04, TTN01, TCWR08, UHR05, UBRvO06, USG^{+}08, Uth01, VHB^{+}07, VKH08, VHK09, VB02, VSBB03, Ver05, VFM06, VFJ^{+}07, VNdG^{+}06, VYM06, WPRS05, WKH07, WH00, WBW^{+}06, Whi06, Wie03, WRW^{+}04, WTH04, WWH05, WPH01, WWMF06, WM01, WBC01, WM02b, WH03, WS08]. coral [WCH09, WDB^{+}05, YBY^{+}09, Yap09, ZDA^{+}06, dBKSV00, dGvdBvO^{+}08, dlMPN^{+}03, vDHHM08, vWNYS05, RS00]. Coral-algal [JM03b]. Coral-associated [WYB^{+}04, RSAK02]. Coral-dwelling [BMM^{+}09, HM04]. coral-feeding [BPM05, Goc04]. coral-macroalgal [NNLF09]. coral-reef [AH08, LCEC02, LCE03, LHH09, MSUK01, NRvdV^{+}02, OM04, Sta09, VKH08, WH00]. corallanid [GL02]. Corallimorpharia [MC08b]. corallimorpharians [FGO09, KCA^{+}08]. Coralline [SCM09b, BP02b, BBM02b, Bul06, DEAI06, ENAT07, HJ07, KSS^{+}08, Kel02, KCS04, LG08b, SHA^{+}05, WW06a]. corallines [DB02]. Corallium [CFA07, GPSH01, SCMB03]. corallivorous [RWSP03]. corals [ARAMPW02, ASLD09, Ant06, BM02a, BG09, BR07, BMK07, BWPB04, BSAR03, BMM^{+}09, BM07a, BEKT09, BB05, BM07c, BS04c, CNW07a, CSMvE04, CD07b, Cum02, DDV^{+}08, DPM02, Dov04, DKAH02, DS09b, Edm00, EBC04, Edm07, EG08, FD00, FLK^{+}06, FHPW07, FFHW07, FH08c, Gil02b, GC06a, GA08b, Gou06, Gou07, HAC06b, HTRFP04, IE06, JM03b, JM03a, JMHS03, KS08, Kel04, LG01, LWB09, MFK07, MH09, MACS07, Mar00, MGBB06, MND^{+}02, MCE^{+}01b, MB00b, NvW01, NWHH01, NSFW09, NKS03, NM09, NL05, OP02, OGL^{+}09a, PGM05, PB08a, PPBS02, RGL01, RGDI06, RPS^{+}04, RMRAFP08, RSY^{+}06, RK09, SGV^{+}02, SHS^{+}05, STF^{+}09, SPCG09, SPT04, SHH08, SHSF08, Thr09, TLLKC04, UHvO^{+}08, VYM08, WMW^{+}08, WH03, Woo09, YH04, ZAZL08]. Corbula [HM06b, Kim06, TK01]. cordgrass [PAMJ03, WWAC00]. core [CGF04, HLH^{+}07, WKMK06]. coregonid [MCR03]. cores [MRLK06, RHH^{+}05, TRPNE06]. coriacea [Eck02, GF06, LV02, MH06b, RPC^{+}06, VL00]. Coris [FCS^{+}09]. cormorant [QWY07]. cornigera [HTS06]. cornuta [SWH04a]. Corophium [HDP02, MS00a, MR06c]. corrected [SLL09]. Correcting [NNS01, KVS08b]. correction [HD06b]. corrections [KHHM07]. correlate [DE08, RBG04, RLKW04, Zam07]. correlated [MAG07b]. correlates [BTG^{+}08, BCJ05, Min02, RHSB08, SCM^{+}09a, Thr09]. correlation [BRJ^{+}03, DBT^{+}02a, GBHO^{+}08, KWL^{+}05, KKI^{+}06, MUT04, SGCD01, Tak02]. correlations [HST^{+}09]. correspondence [GKS^{+}07]. corroborate [PHM^{+}08]. Corticium [DUW08]. Coruña [TAÁO^{+}03]. Cory [GDGSP09, MSH08, NGSV07, NGS09]. Corynactis [MC08b]. Coscinasterias [SBM02]. Coscinoderma [DWL09]. Coscinodiscus [Zam02]. cosmopolitan [BHFC08, LFP^{+}04]. cost [BDR06, FCCP06, HAC06b, JHH01, KB06, RRIHC02, SNP03, vDS05]. cost-benefit [BDR06]. Costa [TBBB06, IMW^{+}02]. costatum [GKM02, YNM^{+}07]. Costs [SR08b, EM07a, GML02, HHF^{+}07, MND^{+}09, SZW^{+}07, vHGH01]. Coudert [Ano07f]. could [AKG02, HG08]. Counter [EZTZ02]. counteract [Fen08]. counteracts [LM09]. countercurrent [IMW^{+}02]. countergradient [Tru02, WPIB09]. countermeasures [Rii03]. counts [HLBS07, SMH09, TM02b, WWTH06]. couple [Kap09]. Coupled [HH05a, CHL^{+}05, FSSW^{+}00, LRR^{+}09, Mil07b, MKK^{+}09, MGPAE01, NYK03, NRC07, NSH^{+}08, OSF08]. couples [PGS^{+}00]. Coupling [BD02, DRG^{+}07, GRLML09, MTV06, Abd07, BWSH00, CJRT00, DP02, FCH^{+}08, HPB^{+}04, JCK03, KBL^{+}07, LRR^{+}09, LDL^{+}04, LTNH00, MM06a, MS04b, MRB^{+}08, NReZW09, PNCS^{+}08, PNCT^{+}08, PSGP04, PCS04b, RCB05, TRH^{+}06, TMFF00, VCBH06, WDB^{+}03, ZYZ^{+}06]. Cove [STE04a]. cover [Bar02c, GEC08, GGDB09, HSB08, LRM00, MAG07b, NDOS03, PS08, RM06, SBR09, WPH01, WHB09]. covering [Ada01]. Cr [NWT00, RW01, WG00, WR00a]. crab [ATLC03, BF04, BIG^{+}06, BFE^{+}03, BM05b, BG07b, BPLD^{+}05, BBST07, CTA^{+}03, CRV03, CSMBH08, CGCDA02, CACDC07, DRDD09, DGL01, DC02, DFCH05, DTB07, DSMBG03, Don04, EE00, FVAB00, FCP02a, FCI^{+}04, GA03, GD07, GGDHRA00, HCB02, HS05, HCM01, HE08, HVCT09, HH07a, HH07b, HMA06, HF05, JSN08, JMA02, JSMB05, JSH01, KMF02, KWWH02, LLH06, LTMDB02, LWJC04, LAW06, LS02, LDT07, LBCS05, MM06a, MCE01a, MADPQ06, MH08c, MS08, MdL06, MS04b, MH06a, MSP^{+}06, MFM^{+}09b, NE01, NEG01, OJ04, O’C07, ODF09, OvM02, PB07, PEDT03, PJJ^{+}05, PPJ^{+}05, QMM02, QAA^{+}06, RB04, RAA02, RE00a, RAS07, RGT01, SML02a, SW08a, SAWG05, SG06a, SAJG06, SMHC03, SP08a, Smi04, SMFP08, THM^{+}03, TE00, Too04, TL03, vMRO03].

Crabeater [WBS^{+}07, SKE05]. crabs [AED^{+}07, AF06, BD00, Bar02a, Bar03, BEW03a, BEW03b, BKB^{+}09, BVI^{+}05, BLT03, BKB05, DRDM09, DES^{+}00, FPB02, FCCP06, FSC05, GKK03, Hal04b, Har06b, HSST01, HD07b, HCCY02, KPW03, KWD05, KH05, MHR03, Mok04, OWBD07, RB02, RB05a, She08, SvNvdMvdV09, SO01b, SRLA04, TMAN08, TW00, TDL05, VHW^{+}07, WP00a, WS09, WL05]. Crangon [OHN01, PGBA04, TC03a, Tay03, Tay05]. Crangonidae [PGBA04]. crassipes [MSP^{+}06]. Crassostrea [Bay02, BCJ05, BC06a, BDG^{+}07, BGD^{+}03, BH09a, CBMB03, DBR^{+}07, DVLH^{+}00, HFM^{+}04, KLR^{+}06, LNV^{+}03, LABR^{+}06, MSF^{+}07, MWMH05, MMGC06, NMCT03, PCS04b, RRS00, RSN02, YIY08]. crawling [Mou04]. create [HC02, RJTM08]. created [Vie09]. creation [BGGR02]. creek [FH08a, HH03a, PDQ01, TD01b]. creeks [AHBM^{+}07, BBBSPH03, CLSM09, GPHC09, NA04a]. creels [ASA^{+}08]. Crepidula [BCB^{+}06, BDG^{+}07, CTP02, CNW^{+}07b, DBR^{+}07, DJV03, DEV07, LGD04, MTR^{+}07, PL07, RSN02, Thi05, UP07]. Crepipatella [MCCP09]. crested [GRIR06]. Creus [OGL^{+}09a]. crinoids [CBP^{+}01]. crispata [Lam00a, Lam00b]. criteria [Ley05, PPC09]. Critical [HHO03, SD00a, FOFG09, GT08, GDBF07, HDDR08, IR02, SW08a, Sve02]. critically [LIM^{+}06]. critique [Kin02, LKW03, RASB03]. Crittercam [HMBD01]. croaker [CC05, JBCL03, MFA03, SRJ09a, SRJ09b, SEH07a]. crocea [RGMJM07]. Croix [Chi02, CS03a]. Cross [CBM^{+}07, JS08b, KMP^{+}06, MFM^{+}09b, Mun07, SOMM06, WHB09, AI04, FMT09, JSMB05, KHP09, SB05a, VHB^{+}07, SS09b]. Cross-channel [JS08b]. Cross-community [KMP^{+}06]. Cross-frontal [Mun07]. cross-habitat [VHB^{+}07]. Cross-shelf [CBM^{+}07, MFM^{+}09b, SOMM06, WHB09, AI04, KHP09, SB05a, SS09b]. cross-validation [JSMB05]. crown [GHdN07, ITN09, LRM00, Pra01]. crown-of-thorns [GHdN07, LRM00, Pra01]. Crozet [BLD^{+}05, FDD^{+}03]. crucial [RK08]. crude [SMC^{+}05]. Crustacea [FFR^{+}03, AY04a, AY04b, BF04, CGCDA02, GLB^{+}04, Gru03, GLBF07, JR01, KSSH02, KI01, PD01, RMAY04, TD01b, Zup00]. Crustacean [YCSS06, BRB^{+}00, GSFN00, OHS^{+}07, OBB00, QS05, RPS^{+}06b, RPS^{+}06a, SD09, SEVSP02, TFSP09, Ven06]. crustaceans [CSP^{+}03, DPGN08, GPHC09, HSJ08, HLR04, JLJ01, PGCC^{+}06, TMFB06]. crustose [BP02b, ENAT07, NWHH01]. crusts [DS01]. crustulenta [LP01b]. cryptic [CLABRO^{+}03, DB05, DRV^{+}05, SB00, TAS07, VGB^{+}05, WA03]. cryptobenthic [DB03]. cryptomonad [LN05]. Cryptomonas [SLCB00]. Cryptophyceae [AG04]. cryptophyte [KNT^{+}07]. cryptophytes [HP06]. crystallina [KSH^{+}09]. Ctenophora [LF06]. ctenophore [CS08, DBP04, FKA^{+}06, GT02, KFA^{+}04, SDPT06]. ctenophore-copepod-ciliate-diatom-phytoflagellate [GT02]. ctenophores [FPDS^{+}02, GB01, WB06]. Cu [KSS^{+}05, LW05, NK03, RCC01]. Cuba [BF07]. Cuban [PCCL05]. Cubozoa [BPC09]. Cue [WWWY07, KZ00, MDW^{+}09, SdNH^{+}06, TJM00]. cues [BBSO03, BK06a, DK03, EEWBdN08, FPB02, HLQ02, IH02, LHQ03, LHQ05a, OJ04, O’C07, RWP07, RE00a, SVN04, Swa04]. cul [BCT^{+}08]. cul-de-sac [BCT^{+}08]. cultivated [DOML^{+}07]. Cultivation [GTE^{+}09, CBM^{+}03]. Cultural [Orb08]. culture [CMD07, FPC05, HSRF02, HLH^{+}08, LGQ^{+}00, LNPB04, LZW^{+}08, Pep06, SLT^{+}01, WGD05]. Cultured [PPI04, AGCF^{+}08, FB03b, GDS02, LCHG03, LMF^{+}01, PBVP09, ZYZ^{+}06]. cultures [CPV^{+}03, MGDP01, Maz04, MFCC02, RATD07, SML05, WPD02]. Cumacea [AY04a, AY04b, GLB^{+}04]. cumingi [KHK^{+}04]. cunner [NMJ02]. cup [AY03]. cupuladriid [OHCFJ04]. Curacao [DBT^{+}02a, NvdV02]. Current [BER^{+}03, DDQ^{+}00, EDS03, HM09, IG04, PGR^{+}01, TDC^{+}03, AB04, BKP07, BFK^{+}09, CH09, CWR06, DESK01, FC00, GGR06, LJK02, MWT^{+}06, MBG^{+}08, PBPLB06, RHS01a, RZWG08, SvAHH09b, SvAHH09a, SRR06, Whi02, ADF^{+}09, AAJ07, BKM08, CCL03, GSFN00, KBB09, KB07, KCG07, MOM^{+}02, MHABGC04, OAS01, PBN^{+}07, RGMJM07, RGCL^{+}05, RGC^{+}06, SSS^{+}09, SB08, WAN^{+}01]. currents [BSED08, CBM^{+}07, DTB07, DCF^{+}05, FK00, GMA^{+}09, GSB^{+}06, GBD^{+}06, JPG05, LLM^{+}08, NRJB04, RWBS03, Zam03a]. Curtis [Som08, Yos08a, Yos08b]. curve [BBNK02]. curves [CS02a, HLR04, Uth06]. cutthroat [LBF^{+}07]. cuttlefish [BTF^{+}02, BTD^{+}04]. cuvier [MCP^{+}09]. Cyanea [Pur03, ACH07]. Cyanobacteria [KLRS00, EÖKS^{+}02, ET08, Hen07, JNS04, KOTL07, KSKK^{+}07, KP01b, KWB^{+}06, LFRR^{+}07, MPDS07, NReZW09, RGB02, SGK04, SBC^{+}06, WVL01]. cyanobacteria-dominated [KOTL07]. cyanobacteria-driven [Hen07]. cyanobacterial [BP04, CMHP06, FPC05, GHH^{+}04, HMKT02, LEÖK^{+}02, MSH^{+}03a, RHGB02, SFG05, WK00]. cyanobacterium [APP^{+}09, EC07a, KSKL^{+}03, RMC^{+}05, RAE07, SNCP04]. cycle [AY04a, ACLF09, BVC^{+}02, BTF^{+}02, BTD^{+}04, CBD04, CID^{+}06, CB01, DL03a, FC08, GCD^{+}03, GGRC^{+}08, HPVL05, Joh03, KI01, KWL^{+}05, LRR^{+}09, MSF^{+}07, MTLT02, NPM01, OMT06, PIY04, PMM^{+}00, PEGG^{+}08, PJR^{+}03, RGP^{+}03, STPA04, TSW^{+}04, TSW07]. cycle-related [STPA04]. cycles [AS06a, BPT07, CWR06, DAM^{+}08, DRG^{+}07, GKS^{+}07, Gim04, GTPC04, HH01, JC08, LSG^{+}05a, LLRDT05, MGREA07, PDHH06, SNGR00, TAW^{+}09, WHL^{+}06, ZLAK04, dM01a]. cyclic [SP04]. Cycling [NCÁSBP04, AJ03, BWSH00, BAE^{+}08, CGF04, CBE04, CRBE04, CRCV04, CVSM09, FDRR05, GM08, GCHC00, HH03a, HO02, HCA06, IAT^{+}07, MFE^{+}05, NSJ02, NMA00, NWFV06, RPO00, SCB00, TC01b, TACM01, TT07b, WK00]. Cycliophora [Ned04]. Cyclope [BSBD^{+}04, CMRB08]. cyclopoid [BK06b, HW08, NMS^{+}02]. cydippid [SDPT06]. cylindrica [MFK07]. cylindrical [GK00b]. Cylindrotheca [Rij03]. Cymbastela [RDC06]. Cymodocea [AMS01, AAHDS06, CPKV00, MD01, OEDSJ02]. Cynoscion [HH00b, Kup03, LBHL^{+}09, TT07b]. CYP1A [DCT^{+}04]. CYP1A-IPP [DCT^{+}04]. cyprid [THC^{+}05, TOBR07, dSSC07]. cyprids [OTBR00, TPGQ07]. Cyprinodontidae [Mal02]. cyprinoides [CT06a, SCIT09]. Cyprus [GBM01, HBG^{+}02]. cyst [KH00b, Kre01, SEM01]. cyst-forming [SEM01]. cysteine [SSG^{+}00]. cysteine-rich [SSG^{+}00]. Cystodytes [LLRDT05]. Cystoseira [VAB09]. cysts [ASH03, KSA03, ONA04, PCBL05, RMC^{+}07]. Cytochrome [CHMR05, SRLA04, DCT^{+}04]. cytometric [BMW^{+}00]. cytometry [SML05].

D [Lei01, NH03a, GV01, TBI07, AHC^{+}06, GLS02, SMM^{+}04]. dab [MCW^{+}07, SF05]. dabneyi [WCZ^{+}09]. daedalea [LWCF04, MSG07]. Daily [CT06a, DHR06, EdPP^{+}00, SS03, ASC06, HSCD01, MTS07, PEGG^{+}08, RP06, SSM^{+}01]. Dall [OKAM00]. dalli [OKAM00]. Damage [BDW^{+}00, HSS09, BdBK^{+}00, BKSS07, BD08, BWvdE^{+}06, JBSB01, Mar00, VPS^{+}02, YBY^{+}09]. damaging [Kev01]. damicornis [AM04, HR07, Kuf01, PKH00, UHR05, VYM06, Whi06]. dampens [Kim06]. damselfish [ACH07, Boo02, CHM06, CC03b, EM05, HK02a, HK03a, KHG06, LCEC02, McI02, MAW06, SHK05]. damselfishes [BKDB00, JSMM06, MAW01]. Danger [KWHD08, FSC05]. dangerous [CSM^{+}09]. danilevskii [OAM02]. Danish [ORS02]. dark [CGCMPA01]. darkening [ADS^{+}09]. Darss [HP06]. Darwinian [BHZW08]. Dascyllus [KKG06, KHG06]. Dasyatis [SR08b]. data [BP08, BMW^{+}00, BHFC08, CLB^{+}09, CSD04, CST08, CNT^{+}04, Cos09, CMRB08, CG05, CWN^{+}07, CAA07, DLLB03, DEV07, EOK^{+}08, EGPT03, FGLM06, FTDK07, FSSW^{+}00, FMW07, FWD^{+}01, GAMB05, GBHO^{+}08, GNAD^{+}04, GBI^{+}06, HRB^{+}06, HNMC09, Kir02a, KJRG09, LTMF01, Las03, LBR^{+}09, MWT^{+}06, MWM^{+}05, MH06b, PMNP^{+}01, PR05, Rii01b, SWD^{+}04, SMH09, SAS^{+}06, SW09, SLL09, TBH08, TW10, WPM09, WMS^{+}03a, WAZ07, vOSF^{+}09]. data-logging [MH06b]. database [ASR^{+}09, EHM^{+}09, FGLM06, GLV^{+}09, RWB^{+}09, SAV09b, SAV^{+}09c, SRR06, VCA^{+}09, WAA^{+}09]. databases [SAV^{+}09a]. Date [RMH07, FGT^{+}07, Lap01b, LB05a, RMM04]. dates [SL05d]. dating [ASLD09, BK05a, BK05b, KBT05]. Davies [AGDP^{+}06]. davisae [HSCH07]. Davoren [Car05]. Day [EC07b, NDV^{+}00a, STRB06, OMT06, PPS^{+}07, WBS^{+}04]. Day-night [NDV^{+}00a, OMT06]. Day-to-day [EC07b]. daylength [SEM01]. days [SR04a]. Daytime [FLR05]. DBT [SJPC02a]. DCMs [HHKH05]. DDT [MLR02]. dead [BD07, FCH^{+}09, TJM03, TRS^{+}06]. death [CPB04b, FHGJB04, HBG^{+}04, YBY^{+}09]. Decadal [TPG09, Win09, WRW03, DJM07, Edm07, Fie02, Tyn04]. Decadal-scale [Win09, WRW03, Edm07]. decades [CSPA09, GHL02, SAC^{+}03]. decadienal [CBO04]. decapod [BSH^{+}09, CSP^{+}03, DFCH05, MW01, PGCC^{+}06, RPS^{+}06b, RPS^{+}06a, RT00c, SEVSP02, TSSA03, dSSC^{+}08]. Decapoda [ASA^{+}03, BF04, BO00, CGCDA02, CACDC07, CCdAR00, FTB09, KMF02, LWJC04, MNC^{+}02, NPVW06, PGBA04, RGVC07, Zup00]. decapods [JH06, PPV^{+}06, RT00b]. decay [EÖKS^{+}02, MMBT03]. decaying [VWH00]. deciduous [DZ06]. decisions [SNP03]. decline [Cum02, Edm07, GEC08, HGS^{+}02, LS02, LRM00, MGO^{+}07, NEK^{+}07, OKAM00, WL06]. Declines [CPG08, CWE^{+}08, CST08, FBD06, RWA^{+}02, RM06]. declining [GCSSJV07, HBBD09, RCL^{+}09]. decomposing [MMBT03]. decomposition [ATTX00, CD07a, DDS^{+}06, EF02, GRCdLvB08, HO02, MZJ06, SYVC00]. Deconstructing [MRJ05]. decoration [HTS06]. Decoupling [TIM^{+}01]. decrease [HW02, HMJ^{+}09, ÖBFO01, WWHE04]. Decreased [LIT05]. decreasing [KdLCM09, MG00]. decussata [SHH08]. decussatus [CDFRL03]. Deducing [MBN^{+}09]. Deep [AW03a, AHW03, CPBS07, FKI^{+}09, HHKH05, LHJD^{+}03, LSG08, RHHH09, TCS^{+}05, TPBF09, AWS^{+}09, AHK^{+}02, BS03, BKP07, BD07, BWP^{+}06, BHY09, BM07c, Cai09, CLO^{+}09, CM02, CC02, CM01, DBG^{+}09, DFMD00, DZ06, DS09b, EM07b, FGLM06, FC08, FLG02, FVAB00, FMZvW02, GHV02, GLB^{+}04, GMVF08, GB02b, GGTKG09, HDDR08, Hou09, HPBT03, HBTD04, HPB^{+}04, Joh07, JTB^{+}03, KCB06a, KBP06, Koe05, KSHO01, KVV^{+}00, KT03, KW00b, KST^{+}08, LTFJ00, LTG^{+}01, LTMF01, LBF^{+}02, LMS04, LG06, LLJ05, LSBJ08, LAK07, MBC^{+}07, MRJ05, MC06, MB07e, MNT09, MMB^{+}02, NNG03, NSSI06, NOP02, NHN^{+}06, PB07, PCS09, PEG^{+}00, PKA06, QASA08, RW07, RSE01, REM^{+}06, RGDI06, RB05b, SSRG05, SHS^{+}05, STF^{+}09, SJS^{+}01, SFL^{+}09, SV01, STRB06, TSC^{+}07, Thr09, TFL^{+}04, TSW^{+}09, TYC00, TYD07, VT00]. deep [Van02, VTY06, WASW03, vDHHM08, Lee00]. deep-coral [AWS^{+}09]. Deep-diving [CPBS07, LHJD^{+}03]. Deep-ocean [TCS^{+}05]. Deep-sea [AW03a, AHW03, TPBF09, AHK^{+}02, BS03, BD07, BHY09, CM02, CC02, DBG^{+}09, DFMD00, DS09b, FGLM06, FVAB00, FMZvW02, GHV02, GMVF08, HPBT03, HBTD04, HPB^{+}04, KCB06a, KBP06, KSHO01, LTG^{+}01, LTMF01, LBF^{+}02, LMS04, LLJ05, LSBJ08, MRJ05, MC06, MNT09, MMB^{+}02, NNG03, NSSI06, NOP02, NHN^{+}06, PB07, PCS09, PEG^{+}00, PKA06, RW07, RSE01, REM^{+}06, RGDI06, RB05b, SSRG05, SHS^{+}05, SJS^{+}01, SV01, TSC^{+}07, TSW^{+}09, TYC00, TYD07, VT00, Van02, VTY06, WASW03]. deep-water [BKP07, BM07c, Cai09, CM01, DZ06, GB02b, Hou09, KST^{+}08, Thr09, TFL^{+}04]. deepwater [Bra08, CMD^{+}09, SJ05]. defecation [BD02]. Defence [TLLCT02, CNPS05, FI09, GHdN07, NDB04, NCP^{+}08, PdNS06a, PLC^{+}02, Rij03, TJ02]. defences [CC04, GWBdN09, MKKK08, PEG^{+}00]. defend [KPEF00]. defense [BLS03, BIC^{+}03, BI03, DDHQ06, FD08, HLY05, KKBH03, KFH04, LS09a, LH06, LTL06c, PCL^{+}09, PHM^{+}08, RWP07, VWF^{+}01, VH03]. Defenses [WP00a, WP00b, AIM^{+}05, ALPK00, BIC^{+}03, BI03, CH00a, Fre07, GP01, Goc04, GA08b, HLY05, KLFH04, OP02, PAM^{+}09, PPS00, PPBS02, PFP03, WLdG^{+}04]. defensive [IH02]. deferred [EM07a]. deficiency [BSDR01, BHFC08, BOM^{+}08, DHBZ00, NR00, VDA^{+}03]. Defining [BL07, FBM07, HM09, LBHL^{+}09, SRJ07a]. definition [HCS07b]. defoliation [PTAR07, SM08a]. deformation [ZAZL08]. degeneration [FHGJB04]. deglaciated [THM^{+}03]. degradability [DMH01]. Degradation [PI08, WRW^{+}04, BLP^{+}02, ECM^{+}05, EWBH04, FAJM07, HHT^{+}09, HCJ^{+}07, KM03b, NFN03, OSA05, PK06, PD05, VWH00, WRJ^{+}05]. Degraded [KFL09]. DEHP [SFHS00]. del [SR04a]. delagoae [GTGM06]. Delaware [NA04a, PJJ^{+}05, XZUK05]. Delayed [NK03, GL04b, JPC00, LCD^{+}05, MPK03, SAM^{+}06, TPGQ07]. Deleterious [JTLG09]. delicatissima [OPSM04]. Delineation [DCT^{+}03, DSP07, MBH^{+}09]. Delisea [DdNS06, WdNS00]. delivery [PMR06, Sha09a, Sha09b, SMF^{+}02]. Delphinapterus [BKLGN05, GRRH07, LMKG01]. delphinid [JBW^{+}09]. Delphinidae [MZD^{+}05]. Delphinus [DC06, LBS^{+}05, MPS^{+}08, MWD^{+}09, SLB^{+}08, TBFA04]. delphis [DC06, LBS^{+}05, MPS^{+}08, MWD^{+}09, TBFA04]. delta [ATTX00, DSPHV04, ARBF02, AHDT^{+}09]. deltaic [APN06]. deltoidalis [KSS^{+}05]. demand [BSR09, PSE^{+}03]. Demersal [Mag02, AHSPG08, AC07, BVVM02, ECM^{+}05, EL05, GBdS^{+}05, HCB^{+}07, HS03, HMHB03a, KKK^{+}05, KBP06, KSHO01, Kor01, LEF^{+}06, MKG^{+}05, MHP03, MSSP06, PPC^{+}05, RHSN07, RM09a, RGT06, ULL03, WKL01]. Demise [HK03a]. demissa [CWT06]. democratica [TL05]. demographic [AR02b, CRAP03, CSO08, CMT^{+}08, DT06, LBHL^{+}09, MAVO^{+}06, MGC03, MHBB05, MDG04, PCBS05, TPG09, TCWR08]. Demography [MAW01, PCMB06, DNL08b, FF03, OEDSJ02, RAC^{+}05, SP03b, TKGS04, TR06]. demonstration [HMH04, PSS06]. demosponge [QASA08, RCC^{+}04, SL05a]. Demospongiae [UST^{+}04]. Denaturing [CBD^{+}05]. Dendrophylliidae [GAZ02]. Dendropoma [KAMR00, RHG05, ZAZL08]. Denitrification [DTNU00, MRLK06, ORPND01, WBN^{+}00, AG02b, CB00a, CRTP02, EF02, FE07, HH03a, HH05a, HMKT02, KGTH01, LJ06, SMG00, TNSM00a, TNSM00b, TRPNE06, WE04]. Denmark [MR07b, SPT^{+}00, AI02, GBB^{+}05, GHP03, HMKR05, HCA06, JHTV02, NRPS^{+}01]. dense [CS05, PBS^{+}05]. densities [BS06, CU08, DFH05, DFCH05, DE08, FGS00, HBNA04, HH07a, HFM^{+}02, ML05, NvdV02]. Density [BBC^{+}09a, CSM02, CHT09, CRCN07, FSV^{+}08, KH05, LSH07b, MHSS08, MDG04, ST07, WDR^{+}05, APA07, ABBF01, ABB02, BBBSPH03, BSJ02, BvK07, BKK^{+}05, BB00, BT04, CCL^{+}05, CBL^{+}09, DBJ^{+}00, DGC^{+}05b, FCC^{+}08, FMP05, FD02, HH03b, HFA09, ISBP06, KEB07, LGB^{+}04, LGS07, LBD00, MK03a, MAWH03, MvdBÓ00, Mok04, NMJ02, Nod04, OGL^{+}09a, OFHM03, OM04, PvDMB08, PHS09, RCF09, RF06, Sha09a, SD00b, SH06b, Ste04b, SR01, SMGCW00, SRC00, Tan02, TBBB06, VHW^{+}07, VÅ01, WH02, WJWL05, WWV08, WBP09, WKL01, WMB00, WA03, Wil09, WR00b, ZSR03, vDvHT^{+}09]. Density-dependence [ST07, LBD00]. Density-dependent [BBC^{+}09a, CRCN07, KH05, LSH07b, MHSS08, WDR^{+}05, CCL^{+}05, LGB^{+}04, LGS07, Mok04, NMJ02, RF06, SRC00, Tan02, TBBB06]. density-driven [SR01]. density-independent [CCL^{+}05]. dentata [SG06a]. dentatus [NRM05, SAG08]. dentex [AFHS09, MSK^{+}00b]. denticulatum [TL05]. Dentition [BKSS07]. depalmai [TYD07]. depend [HBP^{+}05, VJ05, Zam03a]. Dependance [IMS05]. Dependence [MBLA01, NKK^{+}01, BI02, BT04, LBD00, MDG04, MHHS06, OM04, RBB^{+}08, RBDK00, ST07, vDvHT^{+}09]. dependency [BCB08a, BSPK03]. dependent [AA04, BBC^{+}09a, BK04, BB03b, BHH09, BB08a, BTB^{+}02, BMS04, CCL^{+}05, CT09, CRCN07, DHCH04, DMA02, Don04, FE02, FMZvW02, GBGH02, GDL^{+}01, GP03, HT07, HDGM01, IT06a, JMSM08, JS08b, KHG06, KH05, LGB^{+}04, LSH07b, LGS07, LQRGT06, Les00, MVHV06, MHB01a, MHSS08, Mok04, NM04, NMJ02, NTH^{+}02, PA07, RGMH08, RF06, SBJ02, She06, SFPGA05, SD00a, SRC00, SWHSA05, SNGR00, TW04b, TSIY06, Tan02, Tar03, Tay03, TBBB06, WCM08, WBC08, WIB08, WDR^{+}05]. depends [BRGS^{+}07, CYC^{+}09, MKKK08, NBN^{+}08]. depleted [CPG08, SKMS06b, TPSW00, WLJ08]. Depletion [FD00, Her06, WWGR07, ZS03, GKM02, HSHA06, tAHSHA07, JRA02, MR06b]. deployment [CUB08]. deposit [AHL^{+}01, BEE06, EEA01, HSD03, KW00a, KSA03, MDBG07, STJ03, TA03, WÓ02]. deposit-feeder [KSA03]. deposit-feeding [AHL^{+}01, BEE06, EEA01, HSD03, MDBG07, STJ03, TA03, WÓ02]. deposited [WT06]. Deposition [LTL^{+}06b, AH07, BIG^{+}06, FSV^{+}08, VSV04]. depositional [SKDP02]. deposits [CM08, CTH^{+}03, CT04, CVT09, LTH^{+}04, SRL^{+}02]. depression [OYT^{+}02, SP01c]. Depth [ACPS09, DSPHV04, KBP06, Les00, OEDSJ02, TLPG06, AL03b, ARB04, BBM03, BP04, BL00b, BES^{+}08, CST08, CLMR07, CLRM08, DFT^{+}07, EBC04, EMB^{+}09, FGO09, FB07, GRWB02, GK03, HHB^{+}00, HHD^{+}09, HRK^{+}08, HKSW01, Hue08a, JS02, JS08b, KHB00a, MRS00, MW02a, PGM05, QWY07, RSE01, RMRAFP08, SJM06, SMH02, TC01a, VPS^{+}02, VPWE09, WGP^{+}00]. Depth-acclimation [OEDSJ02]. Depth-dependent [Les00, JS08b]. depth-related [CST08, CLMR07, CLRM08, SMH02]. Depth-specific [TLPG06]. depths [JT05, TAFBC07, WMS^{+}03a]. depuration [BGD^{+}03, BTF^{+}02, SSO^{+}01]. depurator [RMAY04]. Derawan [dVBC09]. derived [BCC08, DK06, GBHO^{+}08, HRR^{+}01, HJH^{+}02, HCV03a, KPKD06, LV05, LTH^{+}04, MTV06, SEL02, SW09, SBC^{+}06, TBI08, VWH00, dSHCR07]. Dermochelys [Eck02, GF06, MH06b, RPC^{+}06]. Descending [BES^{+}08, UÓ03]. descends [TJM03]. describe [ONA04]. described [BJR04, GLÅ^{+}03, JKK^{+}05, SFH^{+}00a]. describing [BSW^{+}09]. Description [GMS02, Joh04, HLH^{+}05, Bea04]. descriptions [GT08, RRS00, SKGG03]. descriptors [VSVV03]. Desiccation [HDW02, Prz05, SC01]. Design [FDM^{+}09, HSTS^{+}09, AFHS08, AFHS09, CM04a, GHP02, MM06b, MHP07, MKK^{+}09, NZ08, PMGE01, SNP03, TMAN08, dSHCR07]. design-based [dSHCR07]. Designing [WBT^{+}08, CH06b]. designs [MA00, PMGE02]. Desmodus [CD08]. desolata [CBDH02]. Desperate [EM07a, MK03b, TP01b, BK06a]. despite [FSO08, TSRR05, VMH^{+}08, WLM06]. destinations [CSR^{+}00]. destroyers [DSP00]. Destruction [OKP06, LPK^{+}07]. destructive [GWF07]. detached [CHA06]. detachment [THYBS09]. detect [GPK^{+}04, KS04a, MC01b, MK02, OWBD07, RFB07, SN02, WP03, WBL^{+}06]. detectability [MTF^{+}08]. detected [AGCF^{+}08]. Detecting [RC08, dM01a]. Detection [EOL^{+}05, PPC09, SGK04, WMB00, AKG02, FS07, GHD^{+}02, KT01, LCEC02, NOP02, PMGE01, PBL^{+}09, WB06, WHCP^{+}04]. detections [SCM^{+}09a]. detectors [CHT06]. deter [TL08]. deteriorating [SBL08]. determinant [JMM^{+}07, YH04]. Determinants [BCBVM05, WDT^{+}08a, AHBM^{+}07, GBNÖ08, HCM01, MRS00, NGS09, OvM02, PM03, TGSW06]. Determination [DSS^{+}09, FBJ00b, HDM04, Jef03, JBH00, KVS08b, ASLD09, BBE03, BGKN00, EDBvdL08, GBGH02, HSJ08, HYR^{+}09, MG08, MdMR04, PEdlP^{+}02, She08]. determine [ACLF09, BEU^{+}07, CC03a, CEO06, DPSD08b, GMP07, GM05, GRM00, HH03b, LGB03a, NB06a, SVL^{+}04, Ver05]. determined [AKM09, BCM^{+}07, CDS07, COSM03, HTM^{+}02, HALL05, KPM00, KKMH08, KMB^{+}08, LBS^{+}05, LMD^{+}02, MTCU08, OHL^{+}06, RTH06, SCD^{+}05, TTD^{+}06, WNS^{+}05]. determines [AFHS08, BC06b, BG02, DKR09, FJK^{+}07, KLC^{+}06, MCM^{+}03, TOBR07]. Determining [BLH01, BBM^{+}02a, LSP08, AHL^{+}01, FPB02, HMHB03a, LF02, MC07, WWR^{+}01, ZE01]. Determinism [BBG^{+}06, JCD^{+}05]. deterrence [Fre07, PdNS06b, SP08b, VBAR07]. deterrents [KKBH03]. detours [GSFF^{+}09]. Detoxification [NW04, BGD^{+}03]. detoxified [WLL03]. detrimental [VYM06]. detrital [KL03]. detritivore [PMH08]. detritivorous [Wil04]. detritus [HO02, KM03b, PTLK04, SRGS01, Tho04, VGS^{+}01, VWH00, WWAC00]. develop [BVP^{+}06]. developed [UG04]. Developing [HDDH06, HRB^{+}06, KHK^{+}04, VHV^{+}04]. Development [BAB^{+}06, BC01b, CS08, DKFV05, ECK05, FBJ00a, GJG^{+}03, Gla01, Kup03, LRK09, SYH^{+}09, WPPS04, ATR^{+}08, ATLC03, BMG00, BR08, BSAR03, BPvdM07, BDGK^{+}06, BFV08, BSR05, BH09b, CTA^{+}03, CDFRL03, CWT^{+}01, CCL^{+}04, CHW06, Col03, DHT^{+}08, EÖKS^{+}02, FMWGB08, FSP06, FPB02, FKSW09, GHL02, GKB09, HGS^{+}07, HS05, HHC00, HLU06, HK02c, HFH09, JSN08, JCMR^{+}03, JR01, KR01, KH08b, KN00, KSKK^{+}07, KS04c, LWCF04, LC06, LQRGT06, LC09a, LGZ^{+}09, LGFM00, MHP^{+}00, MFY02, MW01, MCCP09, MNC^{+}02, NFV^{+}06, NPVW06, NMCT03, NS07, PC05, PCC06, PEDT03, PV05, PDB04, Prz05, RMM04, RRIHC02, RRIH^{+}02b, RGB02, RGP^{+}09, SA03a, SAC^{+}00, Sew05, SWW02, Sta08, SMG02, TT05a, WCCA09, vDS05, SW02b]. Developmental [MSK^{+}00b, TW04a, TSSA03, ATT06a, ATT06b, CID^{+}06, FSP06, GMS02, HKSW01, JVT01, Kev01, Kur08, SZGS08, Tho07, Urb02, Wen00]. deviations [BGM09]. devices [DK03, GTD^{+}08, HG08, JLW04, MH06b, RCWYK07]. dextrilobatus [HB04]. DGGE [CBD^{+}05, MCS09, MRQM06]. Diadema [MRN07, MAG07b, TBSJ^{+}04, THBSJ05]. diadromous [BS07b, HS08a]. diagenetic [BPN^{+}09]. diagnosis [LIM^{+}06]. diameter [SN02]. Diamond [WWMF06]. dianthus [TP01a, TP01b]. diapause [AY04a, CW03b, Joh03, SHH09, WJWL05]. diapausing [AY04b, DBvOM06, TBVJ08]. Diaphus [MKWO00, SK04, SK05]. Diarrhetic [KSG03, BCBQ01, GSRH02, MFG^{+}00]. Diatom [PBVP09, RLDL02, SP03a, TMFF00, BVH02, BSR05, CBO04, CS02b, CPV^{+}03, CG04, FI09, GT02, GKM02, GHP03, HDP02, HLQ02, HSHA06, tAHSHA07, KNTK05, KKKK03, LSG^{+}05a, LN05, LELC02, LHK05, MGREA07, MDLS01, MLS^{+}08, MW04b, MGZ^{+}03, MGRP07, NHNB01, OPSM04, PIM^{+}05, PA05, RYF01, RBDK00, Rij03, RRH09, SDRH01, SP00, SSH^{+}00, TIM^{+}01, XW03, vBSSH02, vLvSGS05, vdBÓME01]. diatom- [NHNB01]. diatom-copepod [PIM^{+}05]. diatom-dominated [vBSSH02]. diatoms [Ano02i, DHBZ00, EWBH04, FCS^{+}06, GKWR02, HJ03, INH^{+}09, LHQ03, LHQ05a, LHQ05b, LSE06, LG00b, LHK04, MLC08, OLWB06, PK05a, PLC^{+}02, PWL^{+}06, SWD^{+}04, SS07, SSdWM00, SSH09, WdBS02, WS02]. diazotroph [HF06]. diazotrophic [Fen08, RMC^{+}05, RAE07, SHS06]. diazotrophy [GRS07]. dichaeta [TDvdW^{+}01]. Dictyota [PCT00]. Did [Car05, DM05a]. didactylus [MLGG03]. dieffenbachii [AKL^{+}04, JT02, JT05]. Diego [SH06b]. Diel [ASA^{+}03, AAY^{+}07, BBABL01, BBZM08, CL04, CLSM09, CM05, FLK^{+}06, FMZvW02, FHN^{+}09, HLPHA^{+}01, KBB09, KvNVH00, MS00b, Øre00, Rin03, SAS^{+}06, SS02b, TTM^{+}05, VTDS09, WKMK06, dSSC^{+}08, BCC^{+}03, BIL01, BBA06a, BSKR^{+}06, BZLD02, BCVC07, BT04, CPBS07, GKM02, ICH04, JC08, KN00, LWR^{+}02, MBBW08, MYH07, MHP07, MG09b, MPDS07, Pea00, SSL00a, SKMS06a, SH05, SML05, Tar03, TAW^{+}09, TT07b]. diel-cycling [TT07b]. Diet [BTB^{+}09, CBTW02, CD03b, CB08b, DHCH04, GBHO^{+}08, HM06a, JBG00, KRP^{+}06, LWW^{+}01, MB02a, QMF05, TT05b, YIY08, AS04, AHC^{+}06, ALE^{+}09, ABL08, ASC06, BHPC06, Bar02b, BPT^{+}00, BRI^{+}07, BPM05, BD02, BGS02a, BZLD02, CDF^{+}08, DSS^{+}09, DGH08, DCA^{+}09, DPGN09, EMB^{+}09, FHVMDS07, GRIR06, GP04a, GP01, GBW02, Gru00, HARC05, HRCB02, HBBD09, Hip09, HTRFP04, IMTW09, ISK07, JRT^{+}06, KKK^{+}05, KCV^{+}06, KP00, KJL03, KHH08, KSSB04, KB03, KCSPT09, KB05, LBvdH03, LCGN02, LAH07, LW06, MSF05, MRQM06, MW08, MNP02, MS04a, NHNB01, NA04a, NB06a, OGA03, PWLC06, PJN03, PCW05, RGJ^{+}04, RHDC05, RHI^{+}01, RHSN07, RdNS02, SSB09, SJE^{+}06, SAH09, STAS06, TKTJ07, TSEH08, TC03c, TBI07, TBI08, VBS06, VBFF08, WP03, WR00a, WISH09, WH08, WMCF08, WS06, YTH^{+}05]. Diet-dependent [DHCH04]. Diet-tissue [YIY08, YTH^{+}05]. Dietary [GL00a, PMG05a, SR09, WR00a, AL03a, BFK05a, CC02, CMT^{+}08, EOL^{+}05, GFF08, GFL02, HH00a, HEN^{+}08, IT06b, KFK^{+}07, KTNK07, KTK08b, KFS^{+}09, KW02b, LBS^{+}05, LWH06, LW05, MFJT08, MPS^{+}08, Mon07, NFN^{+}09, PS01, TGNT05, WHP00, WBC01, WAZ07, dlMPN^{+}03]. Diets [CM02, CL05, GL00b, Thu07, YHSH02, BGF^{+}06, BHH09, CF02, HPW^{+}00, HHG^{+}07, ITE^{+}04, KHIH08, KSIG08, NEMF09, NMBH02, RW01, SR08a, TBI^{+}09, TIM^{+}01]. differ [LBH^{+}04, UP07, WBS^{+}04]. difference [MB04, RCWG^{+}06]. Differences [BLD^{+}05, BGS02a, DFSS08, GCU07, HK04a, MGD^{+}09, RSL09b, SFR^{+}09, TMJ^{+}05, TJS08, WJWL05, AE06, BF01a, BCHS09, BGK^{+}06, BRI^{+}07, BPM05, BC06b, BOZ03, Bul05, CC03b, CWM^{+}09, DPD^{+}03a, DBOR09, GSFN00, HTM^{+}02, JCD^{+}05, JVT01, Kne09, KG07, LHJD^{+}03, LRC06, LMF^{+}01, MH08a, MM03, MMC^{+}08b, NMBH02, NT06, PMG05b, RHG04, SKFPS02, ST06a, SFH00b, Smi04, TH02, TLPG06, VFJ^{+}07, WHP00, WST^{+}09, WC03, WPTS08, vLvSGS05]. Different [KVV^{+}00, NHN^{+}06, Air00, ABD^{+}03b, ATLC03, AKL^{+}04, Arm00, Arr05, BC01a, BCBVM05, BCBVM06, BC01b, BFV08, BHY09, BTF^{+}02, BTD^{+}04, CAT^{+}06, CEB^{+}08, CPKV00, CW05, CNPS05, CHA06, DB02, DBT02b, DHR06, FBD^{+}02, FB00, FGSH^{+}05, GLS02, GMP07, GY00, GZG^{+}07, GLG^{+}05, GJ03, HDP02, HGS^{+}07, HKSW01, IBB^{+}07, ISG04, JM03b, KPA^{+}09, KW05, KNTK05, KH00a, KKMH08, KSWP04, LV05, LB08b, LGFM00, LWB09, LG08b, LDT07, LW00, Lus03, MNB^{+}07, MTL^{+}03, MHTJ07, MST03, MS00b, MHDS09, NZM08a, OHS^{+}07, OTB^{+}00, OMP^{+}05, PW04, PP02, PBTB08, PVV^{+}06, RB02, RB05a, RGA^{+}02, RNDF^{+}04, Rod03, RvdPHW04, RU02, SSN04, SG05, SSSV00, SHSV04, SML02b, TAFBC07, TFLA03, THBSJ05, Urb02, VD00, WCBD01, Wan02b, WNS^{+}05, WdBS02, YYL^{+}00, YM09, ZTW08, Zhu06, vdMM03]. Differential [BKB^{+}09, Bye00, FBBH07, FWF02, GRIR06, GLBF07, LSE06, NB06b, PS06, PJ06b, QS05, TBVJ08, TÁCTR06, YH04, DUW08, LL03, NvW01, NDB04, PMR06, RPS^{+}04, WDR^{+}05]. differentially [HRR^{+}07]. differentiated [HFM^{+}04]. Differentiating [MRMA07]. differentiation [AAPO00, CTvdM^{+}08, DFY00, GRPE08, HWvdM03, HDW04, LPHH^{+}09, MB07b, ORS02, PZB^{+}05, RGCL^{+}07, TCW^{+}08, VMVG^{+}07, WGM02]. differing [FKBB01, KG00b, PLMS04, RST04, SABK03]. differs [EPM06, VLW09]. diffuse [UBM^{+}07]. diffusion [NvW01, SWD^{+}02]. diffusive [KGHW05]. digenean [DdMB04]. Digestion [PK05a, RAAT^{+}07]. Digestive [JRTJ04, Pen00, AHL^{+}01, AVNN05a, AVNN05b, IRM08]. Digital [DDGK^{+}08, BHR08, CLABRO^{+}03]. digitata [BER^{+}03, SDHG03]. digitifera [NvWY05]. dilatata [MCCP09]. Dilsea [NP05]. Dilution [DGM00, GRLML09, HRK^{+}08, RSR02, TUK09]. dimensional [DFW^{+}03, EHWS07, FJ01, GHR^{+}07, GBD^{+}06, HHWT02, LMF^{+}01, NBDG05, SDW07a, SWD^{+}02]. dimensionless [BDC00a]. Dimensions [TBI07, SPM03]. Dimethylsulfide [VH03, SLS^{+}00, SLM^{+}02, SLM^{+}00, ZHS^{+}05, ZLAK04, vBSSH02]. Dimethylsulfoniopropionate [KWL^{+}05, SLHT05, YM09, HD06a, HD07a, SLM^{+}02, SLM^{+}00, SH03, ZHS^{+}05, vBSSH02, vBWSS06, LRN^{+}09]. dimethylsulfoxide [LSdM^{+}03, ZHS^{+}05]. dimethylsulphoniopropionate [KSS^{+}08]. dimidiatus [Gru00, GL02]. dimorphism [GBF^{+}02, IS07, Kru01]. DIN [TGM^{+}03]. dinner [FSC05]. dinoflagellate [ASH03, AGB03, BWPB04, BGF^{+}06, BCBQ01, BSVK05, CVF^{+}03, CBM00, CD03c, CTFC08, DCAL^{+}02, FGM^{+}00, GFM^{+}02, HBC09, HSGS00, HMA04, HCJ^{+}07, ITE^{+}04, JKS^{+}01, JKK^{+}05, JTLG09, JERS03, KS06, KJ04, KSA03, LGQ^{+}00, LGZ^{+}09, LWR^{+}02, LTD^{+}02, LWB09, LJK01a, LCW06, MGREA07, MDLS01, ONP02, PCYC02, PPHT00, PCBL05, RH09, RLKW04, SdV06, SBH^{+}02, SHS00, SCS03, SH07d, SSBG02, SG02b, TLIL^{+}09, TIM^{+}01, VSRGW09, VGMC01, WTP08]. dinoflagellates [AC02, CCM01, CWR^{+}03, FHGJB04, GVM^{+}05, HLR07, HGMG^{+}04, INH^{+}09, JHL00, JSK^{+}07, JKS^{+}07, JF07, JF02, KJ04, Kre01, LBR04, LSE06, LLCHG01, PCYC02, Pep06, RGP^{+}03, SGV^{+}02, SL05a, SSLM01, SEM01, SS07, TJ02, YH04]. Dinophyceae [KH00b, LJK01b, LJK02]. Dinophysis [GSRH02, KSG03, KNT^{+}07, PEGG^{+}08, RGP^{+}03, RH09, RJBW03, SPP^{+}09, VSRGW09]. dioecious [LBPS03]. dioica [ATT06a, FLUF^{+}04, LSG05b, LUA99, SVL^{+}04, TCG03, TBAT02, TGNT05, URFA06]. Diomedea [NRP^{+}02, GDGSP09, NGSV07, NGS09]. dioxide [TCS^{+}05, TSC^{+}07]. dioxide-rich [TSC^{+}07]. diphenyl [MAE03]. Diploastrea [TLLKC04]. Diplodus [GWPRC^{+}07, LP03]. diploid [HPVL05, MSF^{+}07]. diploids [TG03]. diplomacy [GDR^{+}04b]. Diploria [CSMvE04]. Diplosoma [MPK03]. Diptera [GJT09]. Direct [BS04a, BN02, PBPLB06, TRPNE06, EH00, GSM^{+}02, Kin02, KHK^{+}04, PdNS06b, PJP06, PTAR07, PUL07, SMC09, VHV^{+}04, WS00]. direct- [VHV^{+}04]. direct-developing [KHK^{+}04]. direction [KJMP05]. directions [EP00b, Han07, Law07, RZWG08, SRR06]. Disappearance [AP06, BVC^{+}02, PD01]. discarded [BWMA02]. discards [ARBF02, AO02, CFG06, FEO07, GZY06, LIM^{+}06, PRvdL^{+}07]. Discernment [GDBF07, MBW07]. discharge [CAS00, DM08, PCSS01, RSNG09]. discharged [Kev01]. discontinuities [BER^{+}03]. discontinuity [DCF^{+}05, PQRAC08]. discounting [MGC03]. Discovery [LLG^{+}06, LG06]. discrepancies [FE07]. Discriminant [WR07]. Discriminating [VRST^{+}07, KGTH01, PJ08, RCF07]. Discrimination [HSST01, RSZI01, SWD^{+}04, APH^{+}04, CHO09, LPK^{+}03, ML06a, NNG03, SJE^{+}06]. Disease [SPT04, AS06b, BL04, BB04, CWCP^{+}06, GA08b, JBHGB04, MT09, NCC^{+}07, Nug02, RS02, RV05, Ryp08, SKR08, SCBA05, SBLF09, THYBS09, WGF09, WKH07, WNWH02, WM05]. disentangle [Kra07]. Disentangling [MB07d, SB06]. Disko [JPNM^{+}06, LNH00, TNHL02, TNT08, TLNH01]. dislodgement [Kaw01]. dislodgment [HS01]. Disparities [SICT03]. Dispersal [EBLE07, FFT^{+}07, GKS^{+}07, GBD^{+}06, HRC^{+}07, STCJ01, ANCC07, Arr05, BTG^{+}08, BUCC09, BBA06b, BHH07, CH06b, CB05b, CFA07, DKR09, DD06, DH06, DEV07, EHWS07, EvBO^{+}08, EE00, FJK^{+}07, Fis05, GLS02, HS03, Hob00, HDW04, HMFC07, JM07, KSHP00, Kru01, Kum06, LVRN02, Lei07, LHH09, LYMCG06, MRN07, MK03b, MSC04, MBBP08, MP00, MNP02, MKK^{+}09, Mok04, MJKH04, NJPI03, NLK00, NCT^{+}01, NSH^{+}08, PLG09, PISL02, SSMM06, SWM03, SP01b, TP07, TÅ07b, UP07, WT09, ZMP^{+}03, Zac05a, Zac05b, vDvTJD^{+}09]. dispersal-limited [DKR09]. dispersers [JM07]. dispersing [KHP09]. dispersion [ACR07, CWMD06, SKGG03, SC07, Und04]. displacement [GWH03, SM09]. display [RCC^{+}09]. disposal [CM04b, FMCB05, OWBD07, VFP^{+}07]. disrupting [LBC07]. Disruption [KSKK^{+}07, FMCB05, LBCS05, NMCT03]. Dissimilatory [AG02b, KGTH01]. dissimile [CS03b]. dissipation [CMSS09]. Dissodactylus [DRDD09]. Dissolution [LNJ07, AED^{+}03, BTC^{+}04, JWG01, MGRP07]. Dissolved [BD09, BVCG01, DÁSG^{+}01, LLL^{+}09, MGPAE01, MIH01, SA03b, vDHHM08, ABBN09, BLS06, CB00a, CR02, DBP04, DR00, GHH^{+}04, GFL02, GB05b, GTSS03, GGM^{+}00, HKM00, HHKH05, JHR05, KSR^{+}07, LSdM^{+}03, MCP05, MN00b, ME01, NN06, NCÁSBP04, Pas00, RWJ02, RNDF^{+}04, SSB09, SLM^{+}00, SZBZ04, SNCP04, TRDV00, TGSW06, TSF01, TPQ^{+}03, TT07b, URFA06, VGS^{+}01, VLP^{+}04, WAM^{+}08, dGvdBvO^{+}08]. Distance [Kum06, BRGS^{+}07, BR09, DD05, GB02b, JM07, KWTR08, LLM^{+}08, THP^{+}05, TC01a]. distances [SDR01]. distant [BCC^{+}04, CWC00, WFG^{+}08]. Disteria [WH09]. Distinct [RLKW04, vWCAB07, BBT02, KCEB09, LJM^{+}00, NSSI06]. distinctness [CW01]. distinguish [GK00a, WHP00]. Distinguishing [Cos09, WGF09, DS06a]. distributed [BDR06, TCWR08]. Distribution [Ano01a, AB06, BMV02, BBS05, BO00, Boo02, BLT03, BSM02, CCR08, CCL03, CV05, DFCH05, DSMBG03, DGNvdV05, DCRT07, Eck02, EHM^{+}09, FVS^{+}08, GDG02, HM05c, JCZ^{+}05, JBCL03, KvKB06, KKL07, MAE03, MMS09, MOM^{+}02, MZD^{+}05, MBS03, OLGG^{+}02, PGCC^{+}06, PWLC06, RWJ02, RK09, TL05, VLP^{+}04, VHH^{+}02, VAP^{+}07, WWS^{+}05, ZLGC04, vdWHF^{+}08, Alt06, AVA^{+}09, BBC^{+}09a, BP02a, BBM03, BA00, BG04a, BJC02, BIL01, BP08, BM02b, BJH04, BCJ05, BF04, BBRJ04, BL00b, Bri03, BG02, BTD^{+}04, CPA^{+}03, CH07a, CS09, CRLP09, CF08, CMD^{+}06, CML08, CSM02, CDAM06, CBA02, CC07, CMD^{+}09, DGVK^{+}04, DKR09, DFV01, DBH^{+}07, DMM^{+}03, DDL^{+}03, DSM01, DM01b, DGCM09, EJM01, Edm00, EWBH04, EB04b, EAF^{+}05, FA06, FLG02, FGO09, FDD^{+}03, FF04, FGA^{+}04, FSV^{+}08, FTBC05, FHQ^{+}06, FHN^{+}09, GVF^{+}03]. distribution [GRWB02, GSS09a, GY00, GEM02, GPB02, GDL^{+}01, GBI^{+}06, GCM01, HLH^{+}07, HCGC09, HSP^{+}08, HW02, HLTT09, HST^{+}09, HM04, HLH04, HH00b, HOVG07, HWGB02, HKSW01, Hyr01, IBB^{+}07, JRP01, Jef03, JH01, JSMB05, JF07, JWR05, JS08b, JTB^{+}03, JGH05, KMWH04, KBP06, KPT01, KT01, KNK^{+}00, KCB06b, KOC^{+}07, KZ00, LL03, Lap01b, LTV^{+}01, LGB03a, LGB03b, LWH^{+}04, LAW06, LPM^{+}06, LTH^{+}06, MFG^{+}04, MR06a, Mag02, MHB^{+}09, MAB^{+}03, MGF^{+}03, MTLT02, MS05, MC08a, MH06a, MVLM^{+}06, MMK^{+}08, MDZ05, MC08b, MB00b, NDV^{+}00b, NA04a, NM07, NLHO09, OSP01, OGL^{+}09a, PIY04, PSW^{+}08, PNK^{+}02, PCD04, PJJ^{+}05, PCS^{+}07, PBBC06, PSRB00, PG00, PCBL05, PDG^{+}02, PPL^{+}02, PKA06, PBS^{+}00, QTB02, Ran02, RFLC07, RSAK02, RZS05, Rue01, SML^{+}09, SZGS08, SSR^{+}09, SGV^{+}02, SLM^{+}02, SHNV09, SSHW^{+}07, SG05, SK07]. distribution [SK03, SvNvdMvdV09, SAHK06, SKE05, SGH^{+}06, SDW07b, SO01b, SGCC06, SM01b, THM^{+}03, TN04, TRDV00, TP07, TCC04, TF04, TT05b, TBFA04, TW10, TT07b, VSRGW09, VAB09, VB02, Ver05, VGMC01, Vli05, WHM^{+}01, WMS^{+}03a, WK02, WJ06, WBR07, WKOE09, Wil09, WKK07, WCH09, WPTS07, YCSS06, YLv^{+}05, ZH02, ZIUG07, dBS01, dSCV^{+}08, dSSC07, vdMPEM08, DMS^{+}05, MBS07]. Distributional [CTvdM^{+}08, DMA03, JNFL04, JDCB09, RHPB08]. Distributions [BRG^{+}02, OP09, ALE^{+}09, BM00, BDDL03, BBZM08, BPC09, BKB05, CTB07, CGC^{+}09, DHG^{+}04, Dav00b, DFH05, EYH^{+}06, ENAT07, Fin00, FJ01, FTU^{+}00, GSM05, Gar00, GSC^{+}01, GK03, HDM04, HMY^{+}08, IHW^{+}01, KWTP06, KHB00a, LBPS03, LB05a, MVN^{+}06, MHP03, MJS09, MHK^{+}01, MA00, MFM^{+}09b, MIH01, MMS^{+}05, RC04a, SS09b, WLM06, ZM06]. Disturbance [EP08, HW03, MW04a, NSBK05, Air00, AVA^{+}09, AF06, Bel08b, BCPB^{+}01, BB00, CÅJ05, CT04, DY05, EPM06, FBT^{+}01, GA08a, GC05, HSS09, IC06, JMdG^{+}06, JPPW01, JNDL02, JHR05, JK03a, JHB06, KH00a, LTG^{+}01, LMW04, MS09, MYE03, MvdBÓ00, MJKH03, NMJ02, NTH^{+}02, RH06, RGS^{+}09, RS00, RU02, SL07, SCZ^{+}06, Sma07, SCAF03, SJ04, TGGA04, TGM07, THF^{+}01a, THF^{+}01b, Tom08, TPS^{+}05, VHT^{+}05, VPWE09, WAK^{+}04, WDRH07, WKPK05, dJTD07]. Disturbance-induced [EP08]. disturbances [AAGS02, BLH02, CBP^{+}05, Fro00, HKWM07, JKQ02, KRVB05, NIF00, PKS07]. disturbed [GMVF08, HP09, SBFP08, WKD^{+}04]. Diterpene [BSVK05]. Ditrema [OAM06]. ditropis [WFG^{+}08]. Ditrupa [MJG^{+}00, RGA^{+}02]. Diurnal [Kar07, LG01, LDS^{+}04, NS02, VPS^{+}02, BVVM02, DCP^{+}08, EdPP^{+}00, KWHD08]. diurnally [EvdKB^{+}08]. diuron [JMHS03]. divaricata [SHH08]. Dive [QWY07, WMS^{+}03a, FFT^{+}07, HHF^{+}07, HHB^{+}00, SJM06, SWHSA05]. diver [PNK^{+}02]. divergence [EGPT03, HEN^{+}08, PC05, RZLS09, WRW06]. Divergent [BCM^{+}07, PGPL03]. divers [EJKN03, SWHSA05]. Diverse [CW09, DS01, KCS04, SLHT05]. diversicolor [CVK00, CBA02, FGS^{+}02, KPA^{+}09, KM03b, OKH^{+}07, RPS^{+}06a, RSL09a, RSL09b, VBA06, WBP09]. diversification [WBC07]. diversion [PI07a]. Diversity [AR02a, ARP09, BWPB04, BIL^{+}02, CBM00, Goo09, HCCY02, LDP^{+}08, LCW06, RC04a, RSAK02, RMC^{+}07, VT00, WKL01, AAHDS06, AR00, AHDT^{+}09, AM04, BBBSPH03, BIL01, BI02, BJH04, BSZ05, BGM09, BD11, BLHB04, BRF03, BSH^{+}09, BPA^{+}08, CLS^{+}01, DBG^{+}09, DDS^{+}06, DFV01, DH09a, DGC^{+}05b, DS06b, DJV03, EWR08, ET08, EHM^{+}09, FSO08, GTD^{+}08, GHV02, GBSQ09, GSP^{+}02, GP08, GB06, GP03, HF08, HAC06a, HGJM09, HKR^{+}08, IP01, JHJS04, JMdG^{+}06, JWTP06, Jos09, KHT^{+}02, Kas00, Kel04, KRVB05, LTFJ00, LTG^{+}01, LTMF01, LBF^{+}02, LCS^{+}01, LH09, LPS08, Mag02, MA02, MLW03, MG05b, NAB09, NK07, OPSB07, PDO01, PCC01, PTH08, PPI04, QDDP06, RL09a, RH06, RWB^{+}09, RSE01, RFN02, RKJ01, RGC05, SAB^{+}05, SGB09, SGV^{+}02, SHB03, SSGB^{+}08, SH07b, SPI^{+}07, SdV06, SE04]. diversity [SMH02, SB06, SG02a, TSSA03, VHT^{+}05, VLK04, WDM^{+}09, WKK07, WWWC02, Yap09, ZM06, Zhu06, dVCH^{+}06]. dives [DFW^{+}03, MFWD08, TC00]. Diving [BBHT02, DJG^{+}07, GLA06, HHB^{+}00, HBG^{+}02, SAM^{+}06, WLS^{+}09b, ZBDE07, AY03, AAY^{+}07, BPM^{+}06, BCH00, BZLD02, CPBS07, GBG00a, HHF^{+}07, HHWT02, KRCGC06, LHJD^{+}03, LAG08, MBH07, MWS^{+}04, MWM^{+}05, PMG05b, PCC07, RBMB00, SAS^{+}06, WBS^{+}07]. division [LWR^{+}02, RGP^{+}03, VSRGW09]. DMS [SLS^{+}00, SPAMG00, ZLAK04, vBSSH02]. DMSO [SPAMG00]. DMSP [ACLF09, HD06a, MLS^{+}08, SPAMG00, SH03, SLHT05, TRDV00, VWF^{+}01, YM09, vBSSH02]. DNA [ABB^{+}00, BAB^{+}06, BdBK^{+}00, BCC08, BWvdE^{+}06, BHMR01, CSB03, Cal05, CAMC07, CRC02, DRV^{+}05, GTGM06, HCB02, HBdBV01, KKMH08, MND^{+}02, MWTB00, NGM^{+}05, RP02, RZLS09, RFCM03, RvdPHW04, SAG03, TGGM05, VBA06, VPS^{+}02, WGP^{+}00, BS04c]. DNA-barcoding [BAB^{+}06]. DNRA [AG02b]. Do [CC03a, CWR^{+}03, DL01, FBD06, KMP^{+}06, KJL03, LJM^{+}00, MH06b, NvdV02, RSE01, WR00b, AM04, BSKR^{+}06, BWO04, CB08a, Gil02a, GDKN08, HC02, HRDA05, LH02, MHR03, MCU06, RZLS09, RBG04, SDR01, WRW06]. DOC [KECJ06, MTN03, dGvdBvO^{+}08]. documenting [BF08a]. Does [DDS^{+}06, Hel00, HBP^{+}05, HM01, KS00b, Kra07, Lin02, PB00b, SBJ02, SJ06, HMJ^{+}09, KAPTK02, Kne09, MHM^{+}08, MB04, RGCC^{+}08, RB08, VJ05]. Dogger [SF05]. dogwhelk [CH07d, LMW^{+}01]. doliolids [KSSH02, PK05a]. Doliolum [TL05]. dolphin [AG02a, CPG08, DCP^{+}08, FGA^{+}04, GF05, MPS^{+}08, OPSB07, RBB^{+}08, SMP^{+}09, SLB^{+}08]. dolphin-habitat [RBB^{+}08]. dolphins [ARGS01, DC06, FFN^{+}08, HR08, IR02, LBS^{+}05, Lus03, Lus05, LW06, MTLT02, MZD^{+}05, SB07, TBFA04, WWGR07, WGH^{+}00]. DOM [KSKK^{+}07, NCÁSGP05, PPCJ^{+}06]. DOM-stimulated [KSKK^{+}07]. domain [BGM09, CCR08]. domains [MFP04]. Dome [IMW^{+}02]. Dominance [BKB05, BTC^{+}04, HR02a, HSHA06, tAHSHA07, JPL04, JKQ02, LBH^{+}04, RL03, RC05, SZ08a]. dominans [KJ04]. dominant [BCT^{+}08, BSR09, DH09a, EHWS07, HH01, MKWO00, RF08b, RGCL^{+}05, SB08, WOH05, ZA07]. dominate [DP01]. dominated [BFB01, BG07a, CÅJ05, CV05, Eyr00, EF02, FD00, FDM^{+}09, IC02, KKKK03, KOTL07, LW04, MIL^{+}00, MILH00, NHNB01, PvDMB08, PK06, RTB06, RSN02, RB04, RCDC06, SPP^{+}09, SH05, YBGP06, vBSSH02]. Domoic [CG04, KGS^{+}08, BPC^{+}02, BLS06, BDGS07, LHK04, LHK05, MTH04, PPB^{+}01]. domuncula [SSG^{+}00]. DON [ABBN09, WB01]. donacium [RCL^{+}09]. dormancy [HBBH07, NH05, OHB^{+}00]. dormant [Joh07]. Dorvilleidae [SP03b]. Doryteuthis [WHS^{+}09]. Dose [BSPK03]. dose-rate [BSPK03]. doses [VTDS09]. Dosidicus [GES^{+}06, GMB^{+}06, IMW^{+}02, WR06]. Doto [YHSH02]. Doubtful [Lus03, RW08, RW06b]. Douro [MBW02]. Dover [EGPT03]. Dower [Ano04d]. Down [LKR^{+}06, BS07c, BES^{+}08, FD02, GT02, HBBD09, HHM^{+}06, JFD09, KH07b, MHHS06, SGCD01, Som00, Wie05, WLJ08, ZSKA01]. Down-mixing [LKR^{+}06]. Downs [DCBvBE09, Pea00]. Downstream [RNL00]. downward [CCL03, PFW^{+}02]. ’downwelling [PTB03, GGRC^{+}08, SB05a, SS09b]. downwelling-upwelling [GGRC^{+}08]. drag [GPS07, ÖNCNG05]. dragging [NSBK05]. Drake [AML^{+}03]. dramatic [ACM^{+}02]. dredge [CM04b]. dredged [WBP^{+}06]. dredging [JBSB01]. Drepanopus [AMR01b]. dressing [RGVC07]. Drift [IOB04, MBBP08, VJKF07, DJV03, Rod03, VK05]. Drifting [HG08, Bib07, GTD^{+}08, Hob00]. drive [GMA^{+}09, SH07a]. driven [Bla09, BD01, DHRG05, DTB07, DHT^{+}08, FAK07, GBD^{+}06, HSM^{+}09, Hen07, HBBD09, JTR05, KSR^{+}07, Kur08, MP00, NB06a, SR01, TAÁO^{+}03, Win09]. driver [BJLH04]. drivers [GPM^{+}09, SJ09a]. drives [AQHTL02, Das02, DS07, NZM^{+}08b]. driving [SSdWM00]. droebachiensis [Ada01, BRS07, DHR04b, DHR06, KSIG08, LGS07, MH08c, MS01a, NA04b, SS05]. Droop [SSH09]. drought [AP00b]. drought-induced [AP00b]. drum [FOFG09, HHH02, PDHH06, SML02b]. Dual [CKS05, FKM^{+}01, MSH08, HD07b, HD06b, KKL^{+}03, PWC^{+}09]. Dual-foraging [CKS05, MSH08]. Dubois [RN06]. duck [EBT^{+}02, KEB07]. ducks [ALE^{+}09, LEB07]. due [KH00a, Kiø03, RRIH^{+}02a, RC04b, SH07d, SR01, VSVV03, vWCAB07]. duebeni [RJ03]. dugon [AL03a]. dugong [AL03a, AL03a]. dump [SRS04]. Dungeness [HE08, HMA06, SO01b, THM^{+}03]. dunlin [LPM^{+}06]. Duration [Bul06, CSP^{+}03, AMT03b, BM09, BBM00, Fis06, GBM01, HPH^{+}01, HHB^{+}00, HHJ04, LMMT09, MR06a, RPC^{+}06, VW00, ZH02, vdVBGW09]. during [ABWK06, AFHA05, ANF^{+}00, BP02a, BCCM03, BKLGN05, BSKR^{+}06, BCH00, BVCG01, BA04, BCC^{+}04, BSP03, BDW^{+}00, CTA^{+}03, CLB^{+}09, Car05, CLS^{+}01, CVÁM06, CSTC01, CSM^{+}09, CT06a, CACDC07, CD07b, DPD^{+}03a, DPD^{+}03b, DR00, DTV08, DRG^{+}02, DCC^{+}03, EO04, EÖKS^{+}02, FHM02, FWF02, FRvdMF02, FFT^{+}07, FHGJB04, FTK^{+}04, GNP07, GVM^{+}05, GGK^{+}07, GJG^{+}03, GLA06, GFB^{+}07, GCD^{+}03, GHH^{+}04, GSW01a, GR07b, GJB04, GBW^{+}05, GBW^{+}07, GCHC00, GG09, GGRC^{+}08, GGM^{+}00, HJ03, HM06a, HLGCO06, HW02, HDG^{+}08, HHT^{+}09, HOVG07, HDRC^{+}08, HBG^{+}02, HDCA06, HFA02, HK03c, IMW^{+}02, IG04, IT06b, JTVC08, JKNJ07, JMM^{+}09, JHL04, JPL04, Joh03, KJMP05, KCSPT09, KYYM00, LGRB^{+}08, LBB08, LIT05, LSM^{+}02, LP04, LLW09, MNR^{+}02, MNTS08, MWG02, MFY02, MGF^{+}03, MLS^{+}08, Min02, MGZ^{+}03, MSH^{+}03a, MCCP09, MNC^{+}02, MGPAE01, MPE03, MGL^{+}00]. during [MCE^{+}01b, MH06b, NvW01, NSJ02, NNW04, NGSV07, NWHC03, PFW^{+}02, PS07, PB07, PTK^{+}09, PTD05, PEGG^{+}08, RHC^{+}08, RMKR00, RJBW03, RLKW04, RCWG^{+}06, RGDN02, RGP^{+}09, SAB^{+}05, SAT^{+}03, SHNV09, SO08, Sew05, SBBS^{+}02, SKE05, SMSD07, SMVF03, SRLA04, SMH^{+}03, SYH^{+}09, TW04a, TT05a, TRH^{+}06, TND^{+}05, TPQ^{+}03, TFSJ00, TDRM08, TFLA03, TGT^{+}06, Tyn04, VH03, VGS^{+}01, VSV04, VGW00, WTP08, WW07, WCCA09, WWL04, WdBS02, Woo09, YOM^{+}09, YQC^{+}00, YQW^{+}01, ZIUG07, ZLAK04, vdVBW00]. Durvillaea [TS05]. Dusky [DCP^{+}08, HDP^{+}06, HCD^{+}09]. dust [EDKT06, Ryp08]. Dutch [EvBO^{+}08, vKH00, vdVBW00, vdVBGW09]. DVM [BSKR^{+}06]. dwarf [CDS^{+}04, ZCR^{+}09]. dwarfism [SSL^{+}07]. dwelling [BMM^{+}09, HM04, HD07b, KM03b, LS09a, SJ04, TCS^{+}05, YFH^{+}02, Zam07]. Dynamic [EKA03, GBG^{+}00b, SN06, ARP02, ANF^{+}00, BCL03, Bar02a, BB03b, BdVD^{+}01, GEHZ07, GK03, KWKF03, LSH07b, NRPS^{+}01, WLJM05]. Dynamics [APN06, CC00, DPD^{+}03a, DZ06, FBSG02, GCHC00, GGTKG09, HHT^{+}09, HvW08, JBHGB04, KS04b, MPSJ06, NGBG07, PJJ^{+}05, SAT^{+}03, SLM^{+}00, YQC^{+}00, AI02, AQHTL02, AMR00b, ABB05, BAA07, BPVL03, BPE^{+}06, BLS04, BP03, BB04, BCVC07, BMCH09, CÁ04, CLB^{+}09, CLS^{+}01, CA06, CMHP06, CCL^{+}05, CSO08, CB05a, CDML06, CSD^{+}07, CCB^{+}09b, DL02, DUW08, DPD^{+}03b, DAM^{+}08, DC02, DPM02, DMV07, DHZ02, DÁSG^{+}01, DPD^{+}08, DdNEvdZ05, EDS03, Edm04, EZTZ02, EDKT06, EB05, EKBM05, FCP^{+}07, GPW^{+}02, GB03, GSS09c, GCSSJV07, GPSG06, GS02, GHP03, GPM^{+}09, GNC07, HP06, HKWM07, HCH07, Har08, HIM05, HGRB04, HTM01, HBN01, HBH07, HL00, HH07a, HGK02, HDW04, HH00b, HDP08, HGZ^{+}00, ITN09, JRP01, JÅC^{+}01, Joh01, KVH02, KWD05, Kor01, KFN01, LCS^{+}01, Lee00, LNH00, LBD00, LCD^{+}05]. dynamics [LJ09, LJK01a, LJK01b, LJK02, LUA99, MBW03, MPM^{+}02, MPM^{+}03, MTH04, MSF^{+}09, MJG^{+}00, MR07b, MASFS00, MVD^{+}08a, MNC^{+}02, MJKH04, MBL07, MAKMK09, MFCC02, MR01, NNW04, NTN08, OSF08, OHS^{+}07, OÁSG^{+}09, Pah05, PS07, PMM^{+}00, PBH^{+}03, PD01, PR05, QASA08, RF08a, RCDC06, RWBS03, SAG08, SG09, SMHC03, SSM05, SPBM05, SS04a, SJ08, SFL05, SBL08, SPH^{+}09, SW02b, SW02a, ST08, SA03b, SC06b, SAHK06, SGH^{+}06, SSBZ07, SVL^{+}04, SBHE05, TM02a, TGY^{+}06, Tar03, TSW07, TMFF00, TT02, TGT^{+}06, TCT^{+}00, THBSJ05, UB03, VGS^{+}01, VHM09, WBW^{+}06, WBH^{+}04, Win09, WMD^{+}07, ZAH03, dBS01, vKH00, vdWHF^{+}08, LCL^{+}00]. Dysidea [BP04].

ear [LP04]. Early [CSG01, DRB^{+}01, FTB09, KCSPT09, Lam00a, RCF07, ACR07, ATLC03, BP02b, BBC04, BF08b, BDRM06, BSP03, BPN^{+}09, CID^{+}06, CGM00, CS02b, DE01, DSMBG03, DQW05, FRvdMF02, FDRR05, FGT^{+}07, GNPB07, GNP07, GB07, GQ00, GDKN08, GS02, GBW02, Hel00, HDRC^{+}08, HRC^{+}07, JTVC08, Jar00, KHK^{+}06, KS04c, Kur08, LJ08, LC02, LWJC04, LOB03, MPSJ06, MK03a, MT05, MAM^{+}01, NH03a, NH03b, OSP01, OvM02, PTK^{+}09, PH05, PGBA04, PBCB03, PSPD07, RTL05, RLFEF03, SZGS08, Sew05, SL05b, ST06b, SBBS^{+}02, SGCP06, SCM09b, TKK05, WHS^{+}09, WMG02, WDRH07, vdVBW00, vHGH01]. early-juvenile [GBW02]. early-life [NH03a, NH03b]. earthworm [CM08]. ease [VSL^{+}04]. East [ABD^{+}03b, CKF^{+}03, INH^{+}09, KHT^{+}02, MHB^{+}09, PHO^{+}07, WRW03, BBBSPH03, BND02, CBGLM08, CIHF01, DGC^{+}05a, Eyr00, KMWH04, KUN02, KCEB09, LBvdH03, PH05, QA01, SKE05, SBSRW02, TK01, TLFS06, UB01, BHN09, CNC00, DKFV05, GZG^{+}07, JMA02, JCL09, KUOI06, LGZ^{+}09, LZL05, MOM^{+}02, PZMG05, SAW^{+}01, THS^{+}02, Van03]. Eastern [RC09, WR06, AVE^{+}06, BSBD^{+}04, BAM^{+}04, BL00b, Bra08, BGRD00, DC06, DD04, DNL08a, DPGN09, DHR04b, DHR06, Fie02, FGT^{+}07, GF05, GUL^{+}07, HANM^{+}05, HBW^{+}05, HW00, JTB^{+}03, JLCA01, JCG^{+}02, KKKK03, KLR^{+}06, LNV^{+}03, LBBH08, LOT^{+}00, LCL^{+}00, Lee00, LCCTFO03, MWMH05, MMPD09, MAW01, MBG^{+}08, NMOO03, NEK^{+}07, ORD00, OBG^{+}07, OCB^{+}07, PAK^{+}00, RGMJM07, RBB^{+}08, RHS01a, ROCB09, RGEB00, RZS05, SH07b, SK07, SV07, SMA^{+}03, TPQ^{+}03, TR06, TSJH05, VTDS09, VW00, VBF06, WWGR07, WBS^{+}07, WHL^{+}08, WOL^{+}07a, WH03, YM06, YBGP06, ZH02, AHW03, HGT01, KvKB06, PG00, WKPS06]. Eastwood [Ano02d]. Eat [DER01, JK04]. Eating [JNH^{+}09, FE06]. ebb [KKA^{+}06, YH00]. eblanae [DGC^{+}05a]. Ebro [ARBF02, CGCMPA01]. ecdysteroid [SRLA04]. Ecdysteroids [Joh03]. echinoderm [MMS09, WMCF08]. Echinodermata [CBP^{+}01, CL05, FB00, HSD03, HPBT03, HBTD04, KAPTK02, LFP^{+}04, SCC05, SDR01, TYC00, UB01, VTY06]. echinoderms [BS00, HDG^{+}08]. Echinogammarus [DJM^{+}05]. echinoid [BP01, BPT07, DRDD09, HCK^{+}05, MY02, SPB09]. Echinoidea [FB00, TYC00]. echinoids [KR01]. echinoplutei [SPB09]. Echinopluteus [SPB09]. Echinorhinus [DS09a]. echo [KRKS09]. Echolocation [CR04, CHT06]. echosounder [CDW^{+}09]. Eco [ZK09, CMLG07, MDS^{+}07]. Eco-Ethics [ZK09]. eco-physiological [MDS^{+}07]. eco-regions [CMLG07]. Ecol [tAHSHA07]. Ecological [BTG^{+}08, BIG^{+}06, CMHP06, Cla08, DSF^{+}04, ESS09, GFM^{+}02, LZL05, MMC08a, MCH^{+}06, MBM^{+}00, SJ09a, SPRK02, WBC07, ABD^{+}03b, BF01a, BVC^{+}02, BH08, BPC09, BN04, CNT^{+}04, Cos09, DSP07, DH07b, DWO^{+}07, ED02, EP00b, FVS^{+}08, GEM02, Gru00, HG08, HRB^{+}06, Kad03, KWV^{+}05, KBA^{+}05, KW02a, KWHS^{+}06, LBS^{+}05, LABR^{+}06, LSK08, MC06, MGHA05, NLCS05, OGL09b, PB08a, PCD04, PL03, RWE04, Rod03, She09, TBR^{+}03, TH02, VWF^{+}01, VMN^{+}03, WWWY07, ZS04]. Ecologically [CCB06, KCEB09, LJ09, RZS05]. ecologies [RB02, RB05a, WB05a, WB05b]. Ecology [BD00, Bar02a, BK06b, GPB02, Kin05b, MC08a, PBJ^{+}05, PBS^{+}05, PI07a, Wil09, WJB09, AGDP^{+}06, ASC06, Bar03, BM03, BDG^{+}04, BCABC00, BBC04, Bra08, BBS07, Car00, CAF^{+}08, CRJ^{+}09, CWT00, CBTW02, CBDH02, CD05, CCC05, DLLB03, DMS^{+}05, DMM^{+}03, DMV07, DRG^{+}02, DD06, DWO^{+}07, EDS03, EBR^{+}02, FJK^{+}07, FHG^{+}03, GJT09, GPLAG04, Gim04, GDGSP09, GSS09b, GTPC04, GB09, Gru03, HBDD02, HFT^{+}09, HHC00, HBW^{+}05, HP07, HM06b, HPBT03, HBTD04, JB07b, JdSM09, KD00, KET^{+}03, KMB^{+}08, KFS^{+}09, KLW^{+}09, LMS04, LCL^{+}03, LSBJ08, LG02, Lóp07, LGM^{+}09, MCS09, MHCB08, MOP^{+}08, NGSV07, NCT^{+}01, OHN01, ORD00, OLGG^{+}02, Paw00b, PAC^{+}00, PMH07, RRS^{+}09, RHC^{+}08, RF08b, RH03, RGCL^{+}05, RTH06, RCB05, STE04a, SR04a, SDPT06, SV07, SCT^{+}09, SW09, SN09, SB05b, SB08, SCH^{+}07]. ecology [TS07, TEB^{+}07, TKGS04, TSEH08, TLNH01, VBF06, WKY09, WBG09, WTS09, WCF08, WR07, WJDA06, YSS03, dM01a]. ecomorphological [FB02a]. ecomorphology [BBH06, BH06]. economics [BS05b]. Ecophysiological [MS00a, STM06]. Ecophysiology [CBD04, Mad02, NT06, PH03c]. ecoregions [WFG^{+}08]. Ecosystem [LHT^{+}02, Pör08, SW02b, SW02a, AC07, AMS^{+}06, AGM06, AWP^{+}01, Bai00, BAA04, BAA07, BKM08, Bea04, BG07a, BGKN00, BSR09, BMSD00, BSH02, BK06c, BS04b, BS05b, BWRP06, Car00, CCMK03, CLC04, CMRF05, CJRT00, CBS^{+}08, DEAI06, DUSJ^{+}09, DS06b, EHK06, FCS^{+}06, FP08, FSHB09, FH03, FSJ03, GN04, GL00b, GL00a, GBB^{+}05, GB01, GKWR02, GS09, GRRH07, Gri09, GPM^{+}09, HMKR05, Har06a, HSM^{+}09, HKK^{+}06, ISBP06, ISK07, JTR05, KCP^{+}07, KMP^{+}06, KWD05, LGSB09, LNB^{+}03, LCT^{+}08, LF06, MDHL08, MT00, MR01, NReZW09, NMA00, NCC^{+}07, NPN06, NRH^{+}07, NK07, PvDMB08, PL03, PKM03, PMG^{+}06, PE09, QRWJF07, Raf06, RHS01a, RGMH08, RFH^{+}06, SAT^{+}03, SLC05, She09, SWL^{+}06, SNY^{+}05, SRP^{+}06, SBI07, SB06, TDTR03, TMFF00, Tyn04, WFT^{+}08, dSSC^{+}08]. ecosystem-based [BS04b, BS05b]. ecosystems [AWS^{+}09, AG00, Bea04, BI04, BRF03, BVHG08a, BVHG08b, CRS^{+}07, CPK00, CK04, CES^{+}09, CS01, Duf06, EBA^{+}03, EKBM05, Hou09, JBS^{+}09, KBH01, KBPB07, LHB09, LB02, LVvdLH07, LTNH00, Lin02, MMB03, MRMA07, MNT09, PPC09, PPB09b, PSS^{+}07, PFT^{+}08, SB09, Syd09, WLZ05]. ecotone [TACM01, WMCF08]. ecotourism [LKDO07]. ecotoxicological [BBM^{+}02a]. ecotoxicology [QHD04]. ectoenzymatic [CAA00]. ectoenzyme [AG00]. Ectohydrolase [TWYS03]. ectoparasite [CC03b, GML02, Gru03, KSH^{+}09, SFH00b]. ectoparasites [CC03a, FF03]. Ectoparasitic [TWMW06, TSRR05]. ectosymbionts [VPSO01]. Ectosymbiosis [HDDR08]. Ectosymbiotic [TT06]. Ectotherms [IF09]. ectothermy [CGH08]. eddies [HVWH06, SML^{+}09]. eddy [BRJ^{+}03, HGZ^{+}00, KKI^{+}06, MTH04, NW02]. eddy-correlation [BRJ^{+}03, KKI^{+}06]. eddy-jet [HGZ^{+}00]. edeni [Ano01a]. Edge [JRC06, SHJC08, TM02a, DGNvdV05, GMD08, LB08a, LAH07, MHBB05, NGM^{+}05, RBG04, SFR^{+}09, THM^{+}03]. edible [PK00, SP08a]. edule [BD06, CWR06, DdMB04, KJL03, KvKB06, UNII01]. edulis [Alt06, ABBF01, BFK06, BJN^{+}05, BJN^{+}06, BGKN00, BH01, DSP00, DM01b, DLD^{+}00, Fre07, HD07a, KR05, LGBWH06, LLSL06, LK02, LVK05, MNB^{+}07, MMGC06, MGH00, NSBK05, NM07, NK07, PBS^{+}04, QTB02, RL07, SS04b, SJPC02a, SJPC02b, SAP02, TBR^{+}03, Thi06, VN02, WGD05, WJ06, WLT^{+}03]. Edward [GFPM01, KPM00, LGBWH06, WGD05]. edwardsii [CB08a, FMT09, GFN^{+}09, JWM09, JRTJ04]. eel [AMT03b, BCP^{+}08, DLT^{+}06, DT06, EDE05, EBLE07, HYLT03, HT07, KOY^{+}05, OYT^{+}02, PMV06, STCJ01, SICT03, SSM^{+}01, TTL06, TA01b, TSI02, YOM^{+}09]. Eelgrass [EvBO^{+}08, HCV03a, Abd03a, Abd07, AW03b, BN01, BvK07, BN02, CGP08, CPG^{+}09, CDS^{+}04, DDSR07, DH01, EWR08, FSO08, HB08b, JBS08, MT09, MF03, MSTS^{+}06, NSBK05, OFHM03, PZ07, PABM08, RS02, RH06, RS07a, SDK^{+}02, SH06b, SCD^{+}09, VK02, WDRH07, ZSKA01, ZCR^{+}09]. eelgrass-bed [DH01]. eelpout [ORS02]. eels [ALOT01, AKL^{+}04, EDB^{+}04, JT02, JT05, JSIT02, JSIT04, JSIT06, JSIT07, KYML02, LIT05, MAM^{+}01, RLFEF03, SAW^{+}01, SAM^{+}01, TDC^{+}07, TCW^{+}07, Wes03]. Effect [Abd03a, AAF^{+}04, ARB04, BPH07, BLS06, BBRJ04, BM07a, BH01, BFV08, BSR05, BSED08, BG07b, BVN^{+}03, CM04a, CH09, CMB^{+}09, CCL^{+}04, CAMC07, CMJ00, DMH01, DQW05, DAG^{+}00, DEvdZ03, EWBH04, ENC^{+}04, ERR06, FCH^{+}09, FB03b, FSJ03, GBR^{+}04, GHJ02, GMK^{+}03, GB07, GCVR01, HUP^{+}00, HANM^{+}05, HN03, HTWQ05, ILL06, JQ04, JFA02, KS08, KNK^{+}00, KR05, LF02, LRHV07, LGD04, LP07, LWR^{+}02, LAH07, LHK04, MSF^{+}07, MJS07, MW06, ML00, MDLS01, MW02a, NLB08, NIF00, PDT09, PBFB08, PCW05, PSRB00, PCS04b, PTP01, PBW^{+}07, RMWO09, RBRH02, RM05, RR05, RTPS07, RFPS^{+}02, RFPB^{+}04, RMH^{+}09, RS07a, SDD^{+}07, SML05, SLEH01, Swa04, TC03a, Thi06, WW06a, WE04, WLdG^{+}04, WPB08a, WBP09, YAF^{+}07, Zup00, AO01, BCO07, BGPL08, CLN00, CCR08, CABS02, CTHF01, DCP^{+}08, DUSJ^{+}09]. effect [DvD02, DDTQ05, EFL^{+}09, Enr05, FPB02, FOL02, FLG03, GMZvdL07, HGSH08, HCV03a, HSO04, HMD^{+}07, HFMP03, IMSO09, JHL00, JHJS04, LP03, LHB^{+}06, MK04b, MR07c, MRH^{+}08, MCCP09, MVD^{+}08b, MFP^{+}05, RMZC07, SLL02, SH01, SAMN04, SLB07, SOB^{+}06, SJ09b, SIK^{+}02, TPGQ07, Whi06, WG02b, WS02]. effective [CFA07, DKA^{+}06, EFL^{+}09, GHdN07, JWM09, LDK^{+}06, PB07, RS06, SBJ06]. Effectiveness [PLCG^{+}05, LLH06, MTLG09, RC08]. Effects [AMP^{+}02, AM02a, ASS05, AC07, AM02b, AGS^{+}09, AL03b, AVA^{+}09, APP^{+}09, AP00b, BFK06, BGC06, BPE^{+}06, BP04, BL04, BP02b, BSZ05, BVH02, BR08, BCM^{+}09, BH09a, BPvdM07, BF08b, BFK^{+}09, BTW08, BSH^{+}09, BM07b, BFLÁ01, BVHG08a, BVHG08b, BT04, CBM^{+}03, CGT^{+}00, CSMK09, CUB08, CLH^{+}07, CPH^{+}02, CBA02, CEF04, CCU07, CD03c, CMGBJ05, CvBGR09, CRZ04, DT08, DD05, DB02, DHBZ00, DAM^{+}08, DBP04, DDD08, EBC04, ESQ03, EHK06, EB04a, EÖKS^{+}02, EMD^{+}01, EWSH08, FGPS04, FHL^{+}09, FJV^{+}02, FDRR05, FVB^{+}09, Fro00, FBN^{+}07, FH08d, GN04, GMC^{+}09, GBMHD08, GMHA09, GP04a, GSW01a, GHV05, GCHA05, GKAS08, GJ03, GS05, GWO02, GLCL04, GK02, Gri09, GBW02, GGM^{+}00, HLVU00, Har06a, Har06b, HAC06a, HVP^{+}06, HBRJ06, HQDH06, Hin02, HLU06, HK02c, HGK02, HFM^{+}02, HCH09, Hun04]. Effects [Hun05, HB08b, HW00, JHTV02, JNDL02, JSIT07, JPC^{+}06, JH06, JWR05, JMHS03, JHB06, KP01a, KVS08a, KT06, KNTK05, KHB00a, KJ03, KYML02, Kim02, KR01, KHK^{+}06, KG00a, Kre01, KSA03, KP01b, Kuf01, KS04c, Kur08, Lar03, LKDO07, LD00, LRN^{+}09, LEB07, LÅCN01, LGFM00, LGVF00, LWMW02, MR05, MCL07, MBVBC09, MDR^{+}06, ML06a, Mar04, MK03a, MKL^{+}03, MRP09, MY04, MMMPB05, MGW^{+}09, MGC^{+}09, MMA^{+}09, MCD^{+}05, MO05, MOPP05, Mis04, MN04, MVLM^{+}06, MRGC09, MRQ07, MGL^{+}00, NM07, NPTW05, OAM02, OGMC06, ODG^{+}04, PGM05, PH05, PCYC02, PBCB03, PH03a, PH03b, PCT00, PM00, Pet02b, PC02b, PMT^{+}02, PPJ^{+}05, PBF07, PPC^{+}05, PDB04, PZA09, RBJB^{+}05, RB02, RBW03, RGR^{+}01, RGS^{+}09, RL03, RTL05, Rij03, RWH00, RCF09, RdNS02, RRH09, RFG^{+}09, RGDN02, RR01, RW06b]. Effects [RSIS09, SSB09, SR08a, SGB09, SBMAE06, SAWG05, SEL02, SZ08b, SMHC03, SVG07, SFPGA05, SCZ^{+}06, SWL07, SH03, SS04b, SG01, SCC06, SFG05, SCAF03, TAA04, TE00, TdLvdVR08, TTR05a, UH03, VHT^{+}05, VP00, VBMBC08b, VMN^{+}03, VSB04, VFJ^{+}07, VÅ01, Vii07, VMM^{+}04, VVTH07, Wal07, WAK^{+}09, WJ06, WKOE09, WDM^{+}09, WPB^{+}08b, WDSA09, WWMF06, vKH00, vdBÓME01, AR02a, AH07, ARAMPW02, Ano02i, ARP09, AP00a, AR02b, AC08, APW^{+}04, BBC^{+}09a, BSBD^{+}04, BDC00a, Bar02c, BMV02, BG04a, BM02b, BM02c, BJH04, BCPB^{+}01, BCBAC01, BCO07, BBDW09, BM05a, BMWH08, BLHB04, BA04, BW00, BOR08, BMS04, BS04a, BJJ01, BCLS06, CAS05, CSA02, CCB06, CMD^{+}06, Cha03, CK04, CCL^{+}05, CCL^{+}08, CNW^{+}07b, CZW^{+}08, CBS^{+}08, Con05, CWBR05, DEAI06, DCT^{+}04, DBN^{+}09, DDGK^{+}08, DJM^{+}05, DGM00, DGNvdV05]. effects [DBS05, DRG^{+}02, DHR04a, DdNEvdZ05, DJV03, EH00, ESS09, EG05, EvBB^{+}09, EAF^{+}05, EM04, FTB09, FI09, FAK07, FK04, FBM08, FBHB03, FGM^{+}00, FD02, GWH03, GRCdLvB08, GYLÖ06, GL04a, GL04b, Gim04, GC01, GSM^{+}02, GC05, GKM02, GM05, GC07, GDL04, GWBdN09, GFN^{+}09, GS07, GB06, HLR07, HSBD08, HL05, HJK01, HOT08, HJM07, HAA00, HBRW04, HMFC07, ISBP06, JMdG^{+}06, JRC06, JCD^{+}05, JBW^{+}09, JM03b, JM03a, KF02, KF01, KHT^{+}02, KES00, Kar07, KLR^{+}06, Kim06, KBL^{+}07, KVS08b, KHvW02, Kon00, KGRW06, Kum06, LSH^{+}07a, LWCF04, LC06, LMMT09, Law08, LLSL06, LHGC^{+}09, LHT^{+}02, LCT^{+}08, LCD^{+}05, LTFH06, Lir00, LP01b, LN05, LRC06, LHT06, LW00, LHK05, MNB^{+}07, MCR03, MACS07, MK04a, MM07a, MT05, MHTJ07, ML05, MRC^{+}08, MVK^{+}09, MvdBÓ00, Mok04, MMMB^{+}05, MLW03]. effects [MK02, MC08b, Mou04, MCN09, NRC07, NK03, NW04, NMCT03, Nod04, NK07, NDB04, NS07, OFHM03, OSS^{+}04, PPCU07, PCC07, PDO01, PE05, PBPLB06, PDHH06, PFZS^{+}06, PJP06, PS06, PCC01, PJ08, PHH^{+}05, Pör08, PLG09, Prz05, Pur07, RL09a, RB05a, RGL^{+}00, RGMV04, RCF07, RCP^{+}05, RFB07, RvdPHW04, RGMH08, RH08, Sar06, SMC^{+}05, SJR00, SR08b, SC00a, SB03a, SWL^{+}06, SN02, SMC09, SGCD01, SH06b, SP08b, SH05, SHJC08, SBF08, SYA02, SLL07, SLL09, STH^{+}06, Sta09, SF02b, Ste04b, STM06, SWG^{+}09, SRC00, SMT08, SAP02, Tal01, TM02a, THC^{+}05, TC03b, TKTJ07, TJ02, TDSM02, TLPG06, TDL05, UP07, VHB^{+}07, VBMBC08a, VML05, VBF06, VMH04, Vli05, VPS08, VPWE09, WGF09, WG01, Wen00, WWV08, WAK^{+}04, WK00, WOH05, WS00, YLK00, ZSR03, ZS04]. effects [ZSKA01, BDCvB^{+}09, DCBvBE09, HSGS00, Sha09a, SHS00, vWNYS05]. efficiencies [HMJ^{+}09, JWM02]. Efficiency [dBGLT08, AGB03, DGM02, Enr05, FLUF^{+}04, FK04, FCH^{+}09, JFA02, KBH01, LP01a, LP01b, MYNY08, STJ03, Uth06, WH02]. efficient [BMT07, NReZW09]. effluent [Kev01, PWHW09]. effluents [CPH^{+}02]. efflux [LWH06, TSW^{+}09]. effort [AM04, BRGS^{+}07, GDDM05, GP08, GPSG06, RHG04, RB08, WZ04]. egested [IAT^{+}07]. Egestion [OSA05, SKK^{+}06]. Egg [AMR00b, AMR00c, AMR01b, BL00a, CCK^{+}09, HK03b, KOC^{+}07, KM00, LZW^{+}08, MRZJ05, MPR02, NMS^{+}02, RRIH^{+}02a, RGF^{+}03, AKM09, BDH01, BC01b, BFE^{+}03, BHMR01, CSA02, CIHL05, CA06, CVÁM06, CW03b, FTB09, Gis05, Gor03, HM01, HLU06, HCBP05, ITE^{+}04, ISG04, KP00, MNTS08, MSK00a, MH00, MS01a, MHW07, NJPI03, Pee02, PJ08, PWL^{+}06, RTL05, SKK^{+}06, SRC00, TD01a, TIM^{+}01, VHMB09]. egg-hatching [TIM^{+}01]. egg-laying [VHMB09]. Egg-production [AMR00b]. eggs [BDCvB^{+}09, DCBvBE09, EvBB^{+}09, FWF02, GRM00, HFMP03, HLH^{+}05, JSN08, JVJ09, KCG07, MHP^{+}00, MLGG03, MBS03, MBS07, MGSJ08, NMVH06, PEDT03, VV04, Vii07, WJWL05, vdVBGW09, vDS05]. Egregia [BMG02]. Egypt [WK00]. eider [DFPG^{+}03]. eiders [LG00a, OP08, RL03]. eight [BAA07, CCK03, KGS^{+}08, VFJ^{+}07]. Eilat [LKW03, GC06a, Wie03]. Eisenia [RC04b]. either [OHCFJ04]. ejaculate [IS07, SAWG05]. ejaculates [KWWH02]. elaboration [OW02]. elasmobranch [MFJT08]. elasmobranchs [PCBS05]. elassodon [RZS05]. Electra [LP01b]. Electrocardiogram [RCWG^{+}06]. Electron [MG09b, BK00, POH^{+}02, YH04]. electronic [GMB^{+}06, TBB^{+}04]. electrophoresis [CBD^{+}05, LMF^{+}01]. elegans [CDT^{+}03, DQ05, GQ00, HLQ02, HH03b, HTWQ05, JQ04, JQ05, KWHD08, LHQ03, LHQ05b, LMCQ02, MHW07, MNC^{+}02, PABM08, TT05b, XWQQ05]. Eleginus [JTN09]. element [BMM^{+}04, MW06, TÁCTR06, XW01]. Elemental [GK00a, HKZ04, PKM04, BCB08a, CF09, DHBZ00, DPSD08a, DPSD08b, EG03, LZW^{+}08, SRZZ01, SSW08, SMT08, VRST^{+}07, Ven06, WHS^{+}09, ZMP^{+}03]. elements [FS02, RHH^{+}05, XW03, Zac05a, Zac05b]. elephant [CDF^{+}08, CCL^{+}06, FBBH07, GAMB05, MHBB05, NFN^{+}09, SCKC07, TBH08, VSB04, RHHH09]. elephantotus [DFY00]. elephas [GQR06]. elevate [PH01a]. Elevated [DS07, RHH^{+}05, BOM^{+}08, CPG^{+}09, DGW01, GPM08, KKB^{+}06, MDW^{+}09, MCN09, RR05, SHL^{+}03, SO01a]. elevation [HS08b]. elevational [RGCL^{+}07]. elimination [ABB^{+}00, QHD04]. elisabethae [LGRB^{+}08]. Elkhorn [CS09]. elliptica [ASP^{+}03, PTP01]. Elminius [RW01]. elongatus [EO04, IP07, KBT05, TWMW06]. else [HMR00]. elucidated [HD07b, LS06, PMNP^{+}01]. elucidation [AHSPG08]. elutriate [GGHA02]. Elysia [AKM09]. emarginatus [LMC08]. embayment [CHLV07, Cam01, CSL^{+}05, CRK06, CSD^{+}07, GSW01b, HDC^{+}08, HMH^{+}08, MSTS^{+}06, PBF08]. embayments [FH08b, PCBL05, She06]. embryo [BFE^{+}03, FPB02, OYNO08]. embryo/planula [OYNO08]. embryogenesis [LP01a]. Embryonic [VAJ^{+}03, AKM09, BS07b, CGT^{+}00, FKSW09, GA03, MNC^{+}02, OAM02, SYWM05, SMG02]. embryos [ATT06b, BSPK03, CHMR05, PDB04, Prz05, TYC00, VP00, VTY06]. Emergence [GLG00, ST06a, TN04, PBFB08, SL05d, Thi03]. Emergent [BDF^{+}04, TS04, KDM05, VPG04]. emerging [Hou09, MNT09]. Emerita [DGL01, DC02, DTB07]. emersion [LMW04, PKS00, PUB^{+}01, SMSD07, Vie09]. emigrate [MHR03]. emigration [BWW^{+}06, SMP^{+}09]. Emiliania [CLS^{+}01, CWN^{+}07, HPVL05, SHL^{+}03, SH03]. Emission [KFU^{+}08]. emissions [Law08, SLP^{+}09]. emperor [KHA^{+}00, KSSB04, MHB^{+}09, RBMB00]. emphasis [AVS03, BVVM02, HJ05, JPNM^{+}06, LNH00, LDWC05, NDV^{+}00a, NWFV06]. empirical [BBNK02, JP04, KS00b, SRP^{+}06]. empirical-modeling [JP04]. employed [LB08b]. Employing [HMBD01]. Empress [DLD^{+}00]. emptive [LOB03]. enable [BJN^{+}05]. enabled [HP07]. enabling [TD01b]. encapsulated [PDB04, Prz05]. encapsulation [KSA03]. enclosed [AB00, Car00, CSL^{+}05, JN09]. enclosure [LCS^{+}01, RHGB02]. enclosures [NFV^{+}06]. encounter [TK03a]. encountered [AHK^{+}02]. encountering [KCTG06]. encrasicolus [BTB^{+}09, CGH^{+}03, DTV08, HFMP03, SZGS08]. encrusting [BA01, BK04, BUCC09, BBM02b, Bul06, DB02, Fer04, LVZW06, TPG09, dGvdBvO^{+}08]. Encystment [ONA04, AHZ03, Pep06, SEM01]. end [CTFC08, LFP^{+}04]. endangered [BPB07, DSH03, LIM^{+}06]. endemic [SV07, TWT08, WG07]. Endemism [VW00, DF04, TPBF09]. Endocrine [LBCS05, LBC07, NMCT03]. endocrine-disrupting [LBC07]. endocuticular [KCB06a]. endofauna [SRDG05]. endogenous [HMY^{+}08, KYML02]. endolithic [MFK07]. endolymph [EdPP^{+}00, Tak02, TTM^{+}05]. endosymbionts [FTU^{+}00, ISSK03, YFH^{+}02]. endosymbiotic [PLM03]. Endothermy [CGH08]. endpoints [MK02]. endurance [GDdS^{+}07]. endurance-exhaustion [GDdS^{+}07]. endures [BP09]. Energetic [HAC06b, KB06, MND^{+}09, RRIHC02, RH04, CC05, EL05, HSCD01, WG02c]. energetics [ASP^{+}03, AGS^{+}09, GBR^{+}04, HHF^{+}07, KHWH03, PRvdL^{+}07, RRS00]. Energy [BAA04, CLN00, CDFRL03, KW02a, WKPS06, AMR01b, Ant06, BK06a, CAT^{+}06, DGM02, DB05, FSB07, HSSS05, JRS^{+}06, MVHV06, MCE01a, MABB03, MBL07, QFJ^{+}07, RHG04, RL03, RG09, TdMF^{+}08, THC^{+}05, VHB^{+}07, WHRS05, WS06, YLK00]. energy-rich [WS06]. enflata [Øre00]. engineer [DEAI06, Gri09]. engineering [CLC04, Har06a, PvDMB08]. England [Min02, DTNU00, DPD^{+}08, HH03a, HFH09, OWBD07, PH05]. English [PD01, AMS^{+}06, BIR00, BER^{+}03, Bro06, CFG06, DD04, GPSG06, HBH07, SK07]. Engraulis [BTB^{+}09, CSH00, CCK^{+}09, CGH^{+}03, DTV08, HFMP03, LRC06, SZGS08, TW04a, TW04b, TAM03, TAM04, TAO07, WHM^{+}01]. Enhalus [LVRN02]. enhance [CABS02, GDKN08, KAPTK02, LWH08, SCM09b, TJA08]. Enhanced [Ant06, RGJ^{+}04, RS00, BP02a, BW00, CR04, HLY05, LFSR06, NBDG05, PKM03, SMC^{+}05, SBvD05, Woo09, dBKSV00]. Enhancement [BT00, HSRF02, ASES07, CCJT09, DAMJ08, KET^{+}03, MP05, OLG^{+}09, PGP03, PGPL03, SPD03, Uth01]. enhances [ACT09, CT06b, GW06, HHW^{+}07, LS08b, MYNY08, PPSP07, RWPJ04]. enhancing [Dis07]. enigma [TSSA03]. enough [GM08]. enriched [Cam01, CSD^{+}07, KHGA02, OAGV07, PTLK04]. Enrichment [RMGL07, APP^{+}09, BSAR03, FH08d, GWO02, HDC^{+}08, HVP^{+}06, HFD08, KF01, KL03, KBH01, LD00, LV05, MMCF07, MMC08a, MBC^{+}07, MM08, MSF04, PZ07, PTK^{+}09, PPHT00, RBRH02, RR05, SAMN04, SVL^{+}04, SCC06, TFA^{+}08, VP00, WHCP^{+}04, WASW03]. ENSO [BP02a, Fie02, MCE^{+}01b, ZHWCS08]. enter [Kin03]. entering [DBvOM06]. Enteromorpha [BR03, CRZ04, KF01, LS05, RWBS03, SAC^{+}00, Ste04b, TNSM00b]. entities [BGGR02]. entrainment [ONKK03]. Entrée [RATD07]. entry [ARFP09, CWMD06]. envelope [KS09]. environment [APN06, AM03, BRSM07, CS02b, CHV^{+}09, GCLC^{+}02, GS09, GSS09c, GWO02, HK03b, HGC^{+}07, HVW^{+}07, HGG^{+}05, JLW04, JS01, KKW^{+}03, KPT01, KWAT08, LHT06, MCW^{+}07, MdMR04, MK02, NLB08, PTD05, PDT09, PDHH06, PFZS^{+}06, PKM03, PHP^{+}03b, PMP04, PB08b, SBL08, SVN04, SRDG05, SvNvdMvdV09, STAS06, TK01, THHY05, TLLKC04, TBH^{+}09, VHK09, WLJM05, Wil09, WJB09, YH02, vdWHF^{+}08]. Environmental [ARAMPW02, BP08, BM02b, BM02c, CSH00, CEO06, CDML06, Fie02, GB06, LNV^{+}03, MTH04, MG08, MdMR04, MTS07, NGS09, PDHH06, PCH^{+}05, Pur07, SSL00a, SEH07a, SCM^{+}09a, Thr09, WM01, ABWK06, AC02, BCO07, BFK^{+}05b, BUA02, CM04a, CSA02, CMB^{+}09, CCR08, CCK^{+}09, COPB06, CIT04, CR07b, CGC^{+}09, Das03, DDL^{+}03, Ell02, FC04, FZT^{+}05, FMW07, FR08b, GMP07, GR07a, GBI^{+}06, GMD08, HAP^{+}08, HDP^{+}06, HNMC09, JBCL03, KTK08b, KWTP06, KSK07, LBH^{+}04, LABR^{+}06, LVvdLH07, LPHH^{+}09, LJK01b, LS06, MHP^{+}00, ML06a, MM07a, MCVAR08, MM03, MW02b, MCS05, MG05b, MTLG09, NPTW05, NT06, OP09, PCS^{+}04a, PCW05, POH08, PJR^{+}03, QMM^{+}08, RWJ02, RGCC^{+}08, RGP^{+}09, SSR^{+}09, SH07b, SP01a, SKMS06a, SHFP^{+}03, Sta08, SBLF09, TBR^{+}03, TBCB^{+}05, TKGS04, VHMB09, WAB^{+}02, WLB08, WGS^{+}02]. environmental [YH02, Zam07, ZKHP07]. Environmentally [NB06a, ABY00, PQRAC08]. environments [BP06, DL01, DM09, GTSS03, HTZ^{+}07, KG00b, MHB^{+}01b, MMVB07, MYNY08, PJB03, RU02, SN04, SJH^{+}04, TSHT02, WRW^{+}04]. Enzymatically [DFMD00]. enzyme [JRTJ04, MLG03b, OBB00]. enzymes [AJ03, BG04c, PEG^{+}00, RL09b]. eperlanus [PA07]. ephemeral [Cam01, DEAI06, DJV03, MRP09, SVG07, SRS05]. epi [KCB06a]. epi- [KCB06a]. epibenthic [BBC^{+}00, DAMJ08, GGR06, Hun04, LS09a, SG02a, SH06c, Str02, WP07, YM06]. epibenthos [CED^{+}07, KW02a]. epibionts [GF02]. epibiosis [DDTQ05, DDHQ06]. epibiota [CCU07, HC02]. epibiotic [GC01]. epidermal [WGH^{+}00]. epifauna [BMO^{+}04, HKH04, HH05b]. Epifaunal [GLG^{+}05, JHR05, SCD^{+}09, Bat09, HKK^{+}06, HK02c, IAT^{+}07, LTFH06, PDO01, RH06, SH06b, SCZ^{+}06, VN02]. epilithic [HK04a]. Epinephelus [APW^{+}04, DE01, HDP^{+}06, KAR04, PCMB06, SSBZ07]. epineustonic [MVLM^{+}06, PPL^{+}02]. epipelagic [IT06a, LUAIH03, SSH^{+}04, TNT08]. epiphyte [CHLV07, IOB04, KES00, PAR08, WS00]. epiphytes [BN02, HVP^{+}06, KES00, TTR05a, WdNS00, YLv^{+}05]. epiphytic [CHG05, ID02, JZL09, LV02, LRHV07, MRP09, MS01b, SASR03, VLK04, VL00]. Episodic [GPM^{+}09, QR03, BEW03a, BEW03b]. episymbionts [Bat07]. epizoic [HK04a]. epizootic [RV05, Thi05]. epizootics [PHH^{+}05]. Eppley [BBNK02]. EPS [HSD03, WS02]. Equatorial [IT06a, RGMJM07, SBA01, BL00b, LBF^{+}02, LOT^{+}00, LCL^{+}00, MH08b, MH09, TDSM02]. equilibrium [SN06]. equina [TLP^{+}03]. equivalence [BLF06, CM04b]. era [CV06]. Eradication [WS04b]. erect [IC06]. Eretmochelys [SWHSA05]. Erin [OHN01]. Eriocheir [CACDC07]. Eritrea [ABB05]. erodability [RWPJ04, WBSL00, WBP^{+}06, WPB08a, WBP09]. erodibility [OSS^{+}04]. erosion [GP04a]. Errata [Ano03b, CFMB00]. Erratum [Ano00c, Ano00b, Ano01d, Ano01e, Ano01f, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03c, Ano03d, Ano03e, Ano03f, Ano04c, Ano04d, Ano04e, Ano05c, Ano05d, Ano06, Ano07c, Ano07d, Ano07e, Ano07f, Ano10a, Ano10b, tAHSHA07, KH07c, MBS07, NH03a, PCS04b]. Errina [WCZ^{+}09]. error [MDHL08, SDH01, SWD^{+}02]. Errors [HPR05, Kir02b, Kir02c, MM07a, Pet02a]. Erylus [KPEF00]. erythraea [LWH06]. erythraeum [RMC^{+}05]. erythrogramma [SC07]. Escape [BLH02, Jak01, BKB^{+}09, JNH^{+}09, Tit01, VFV01, WB07]. escapee [Rye02]. escapees [Rye02]. Eschrichtius [BWW^{+}06]. especially [SSH^{+}04]. Essential [MB02b, AMG^{+}07, BKR04, BSR05, CLP08, LBPH06, Sto03]. established [MHCB08]. Establishment [BUF02, DW09, BB00, DDGK^{+}08, OFHM03, SM09]. ester [ZER^{+}01]. esterification [SAHK06]. esters [SKFPS02]. estimate [FTDK07, GP08, HSHA06, tAHSHA07, POH^{+}02, Sim00]. Estimated [PGP03, BAOV06, KCG07, MOOB06, NRPS^{+}01, TBI^{+}09, WMS^{+}03a]. Estimates [PL07, Whi02, BGDS05, BBC03, BWW^{+}06, BSED08, CM04a, CJFM09, CDT^{+}01, DFMD00, HPR05, MTF^{+}08, MGM^{+}07, MMA03a, MWM^{+}05, MAKMK09, MMBF09, PBAD05, PRM^{+}09b, PRM^{+}09a, PSC^{+}04, RGP^{+}03, TBB^{+}04, TBI08, UB03, dSHCR07]. Estimating [BK00, BWCW02, BLP^{+}02, Cal05, CH06b, FAFR03, FSHB09, GAMB05, GF06, GBM01, KHHM07, MSY02, NDOS03, PGPL03, RPC^{+}06, RKJ01, SMH09, SMP^{+}09, SDH01, Bra05, CSD04, DC07b, GHV02, HHAH00, MW01, OBB00, SBMAE06, SMF^{+}02, WTR02]. Estimation [GMR^{+}02, OIP04, UG04, BBA^{+}05, DR00, FCC^{+}08, FON03, GM01, Gim09, HLR04, KDAK01, LALB08, MB02b, SO08, TUK09, YFH^{+}02]. estimations [Lat07, MUT03, SN04]. estrogenic [DMC09]. estrogens [TAA04]. estuaries [BVI^{+}05, Bro06, CGP08, CVF07, CMHP06, Eyr00, FTO^{+}09, FEFT08, Gil02a, HCV03a, LBC07, LHT^{+}02, Ley05, LPM^{+}06, MTL^{+}03, MCH^{+}06, MN00b, PDM^{+}04, RSL09b, SJ09a, SCC06, TWYS03, Tay05, THN^{+}03, WBP^{+}06]. Estuarine [CLL^{+}09, NRM05, AP04a, AP04b, AFFH04, AP00b, ARFP09, BPBH09, BS05a, BS06, BBBSPH03, BK00, BMG00, Bay02, BFB01, BR08, BMSD00, BDRM06, BCB08a, Bye00, CGCMPA01, CSM^{+}09, CLSM09, CPK00, CK09a, DKR09, DS06a, ECK05, EYH^{+}06, FFA03, FAK07, GGHA02, GFP^{+}05, GK00a, GA03, GDL04, GCL04, GC06b, HKZ04, HHH02, HLTT09, HHT^{+}09, HS07, HM05c, JSN08, Jos09, KGTH01, KSK07, Kim02, KS05, LC07a, LF08, LN05, LDP^{+}08, LS08b, MBW03, MPM^{+}02, MPSJ06, MTV06, MRMA07, MHV^{+}00, MLM^{+}02, MCCP09, MGL^{+}00, MHHS06, MLG03b, MTCU08, NvdV02, NZ08, NTH^{+}02, ODF09, PMM^{+}00, PCD04, PBF08, PCS04b, PCBL05, PI07b, QAA^{+}06, RGJ^{+}04, RSC08, RMKR00, Rue01, RT05, SN06, SR08a, SEL02, SRZZ01, She01, She06, SJ08, SD00a, SD00b, SP01b, SLM07, SSBG02, SMSJ04, SHM^{+}01]. estuarine [STM06, SG02b, SMP01, SML02b, TE00, TLFS06, TMFB06, Trn02, TT07b, VMV^{+}08, VRST^{+}07, WM02a, WL05, WBR07, WBP09, WBC08, WDB^{+}03, WMD^{+}07, XZLP09, YQW^{+}01, YTH^{+}05, YMHV02, YH02, dBMSL02, vBWSS06]. estuarine-dependent [WBC08]. Estuary [AQSP07, BBBSPH03, GB07, PPHT00, PDRT02, RZ00, SAH09, SHM^{+}01, TDUN07, WBH^{+}04, AI02, AED^{+}03, ABBN09, AG02b, BS07a, BBS05, BP03, BMF06, BK06b, BdVD^{+}01, BAE^{+}08, BRB^{+}00, BLP^{+}02, CB00a, CHW02, CHSW03, CIFN00, CS08, CBE04, CRBE04, CRCV04, CVSM09, CRB07, CRCN07, CB00b, CRZ04, CAA00, DSM01, DTNU00, DGW01, FLR05, FCS^{+}06, FCI^{+}04, FR08b, GBB^{+}05, HP06, HJW08, HMA06, HHM^{+}06, HCCY02, JBCL03, JSIT06, KWV^{+}05, KPN03, KSHP00, KW02a, KNU02, KUN02, KS04b, KWL^{+}05, Lar03, Lee00, MBW02, MPM^{+}03, MdL06, MN00a, MS04b, MIL^{+}00, MIH^{+}00, MILH00, NRM05, NRPS^{+}01, NPM01, OWBD07, PMPI03, PMM^{+}00, PA07, RF08b, RMKR00, RGB02, RAS07, RFH^{+}06, SAG08, SSGB^{+}08, SLFM03, SEVSP02, SSLM01, SZZ08, SW02b, SW02a, SBH^{+}02, SSS06, SR01, SWGC^{+}05, SO01b, SVG^{+}05, SAM^{+}01, SKO^{+}08, TRDV00]. estuary [TLFS06, TB08, TK08, TGM^{+}03, TNSM00a, Trn02, VHH^{+}02, WAB^{+}02, WHL^{+}06, WLJ08, WAZ07, YQC^{+}00, Yin02, ZS03, CGCMPA01, CLH^{+}07, CB08b, CS05, GRBB07, HB04, Kim06, KUN02, LHK01, PJR^{+}03, ZGV00]. estuary-dependent [CRCN07]. ether [MAE03]. Ethics [ZK09]. Ethmalosa [PDM^{+}04]. ethylhexyl [SFHS00]. etiological [WNWH02]. ETM [MAKMK09, NH03a, NH03b]. Etnoyer [MB07e]. ETS [MG09b]. Eubalaena [DCM^{+}06, EB04b]. Euchone [RGA^{+}02]. Eudyptes [BC02, DGH08, PISL02, RTPS07, SR04a, TGT^{+}06]. Eudyptula [CCCD09, HDRC^{+}08]. euendolithic [CSMK09]. Eugenia [GGDHRA00]. eulachon [MWG02]. Eumetopias [GT08, KG07, WTR02, WWS^{+}05, WS06]. Eunicella [CLR^{+}06]. Euphausia [AS06a, AVNN05a, AVNN05b, BLS06, DBS05, FHM02, FMWGB08, FSP06, GGFSP06, HRQ^{+}02, HL00, JT08, MRQM06, MABB03, MO05, NOE^{+}04, QR03, SN02, STA^{+}05, TCRT^{+}07]. Euphausiacea [HB03]. Euphausiid [BSA^{+}08, AM03, GGR06, GFPM01, KCG07, NOE^{+}04, PZB^{+}05, Tor01]. Euphausiids [Hip09, AS04, DBS05, ZZT05]. euphotic [KvNVH00]. Euplacella [DDQ04]. Eurasian [KVV^{+}00]. europaea [GAZ02]. Europe [SEH07b, VCA^{+}09]. European [JKNJ07, ASR^{+}09, BMG00, CH06c, CMLG07, CTD^{+}09, CBEL06, DT06, DFR08, EGPT03, FHVMDS07, FEFT08, HSM^{+}09, HB08a, JAA^{+}01, JÅC^{+}01, KTT04, LPHH^{+}09, LP04, MTL^{+}03, MAVO^{+}06, MCC^{+}00, PMV06, RHS01a, RGMV04, SJCM01, SAV09b, SAV^{+}09c, Str02, TBR^{+}03, Thi05, VCA^{+}09, VML05, WDT^{+}08b, WAA^{+}09]. European-scale [JAA^{+}01]. euryhaline [BMV02]. Eurytemora [KSKL^{+}03, LG08a, SKK^{+}06, SKU^{+}08]. eurythermal [DD07]. eutrophic [AED^{+}03, BAE^{+}08, CHLV07, CMRF05, FTO^{+}09, HLH^{+}08, JHJ^{+}03, JHJS04, KUN02, MCP05, PC02a, QRWJF07, RC07, SCC06, ZYZ^{+}06]. Eutrophication [CVF07, KBA^{+}05, MSA07, Sve02, VLRJ07, Clo01, DT08, DSDS04, LKW03, PMM^{+}00, PPC^{+}05, SRGS01, TA07a, YMC05]. euxinus [SYH^{+}09]. evade [RFB07]. evaluate [ALE^{+}09, Bro06, FLL09, KJRG09, SSW08]. evaluated [FHB^{+}02, vdMM03]. Evaluating [CR07a, Cra01, DBN^{+}09, NWIT08, TSEH08, CH09, HJ05]. Evaluation [AB04, KMAC^{+}02, Rob06, SJ05, Sha00, TBBB06, AB00, AIM^{+}05, BAB^{+}06, BBC08, Eyr00, FAFR03, GBB^{+}05, HHO03, HSBD08, JCMR^{+}03, JSIT07, KET^{+}03, Kup03, MA02, MGHA05, NRM05, SWD^{+}02, Sve02]. Evans [SHA^{+}05]. Evechinus [LB01, PS06, Sew05, Wal07, WGL03]. evenness [SH07b]. event [AW03a, BB02a, DdNEvdZ05, FAK07, GYLÖ06, GPSH01, LCD^{+}05, MW02a, MCE^{+}01b, NvW01, RVMI^{+}03, RSY^{+}06]. event-driven [FAK07]. event-scale [RVMI^{+}03]. events [BLH01, BP02a, BUCC09, CPP^{+}04, DBS05, ESB^{+}04, GB01, GPM^{+}09, LS08a, LGRB^{+}08, LDL^{+}04, MG09a, MM05, OKH^{+}07, OÁSG^{+}09, RWA^{+}02, RMA^{+}02, SP04, SYCG05, VBMBC08b, VSB04, WAB^{+}02]. Everglades [DCD^{+}03]. Evidence [AKL^{+}04, Edm07, GAB^{+}03, GHH^{+}04, HBBD09, KB07, LHHB05, MVD^{+}08b, MGRP07, MKH^{+}02, NHS^{+}08, NMJ02, RWE04, RSD^{+}08, SRC00, SHSF08, TAM03, Tru02, WWHE04, WVL01, ZCR^{+}09, AZKJAM06, AYSL09, APM^{+}01, ARP09, BS07a, BKR04, CBO04, CAF^{+}08, CR08, CHW06, CFA07, DBH^{+}02, DMK^{+}08, DDQ04, ESB^{+}04, EKBM05, FAK07, FTBC05, GMVF08, GSL^{+}06, GGDHRA00, GAAB^{+}08, HPW^{+}00, HYLT03, HF03, HLY05, HEN^{+}08, HPB^{+}04, IF09, JB07a, JMdG^{+}06, JCK03, KHG06, KVH02, KR05, LC07a, LHK01, MWR07, McI02, MP00, MSD05, MS01b, NW05a, NOO^{+}08, OKP06, PWLC06, PM00, PWC^{+}09, PBTB08, QS08, RJBW03, SESP03, SBSPS04, SSRG05, STF^{+}09, SSR^{+}03, SAF^{+}09, SPGS03, TK01, THP^{+}05, TPBF09, Too04, TTL06, VWF^{+}01, VBB06, VBAR07, Whi06, YQC^{+}00, YI03b, YTK^{+}05, CC03a, DL01]. evidenced [CVF07, RH03]. evokes [Ada01]. ’Evolution [BHC^{+}00, WG09, ALOT01, DS07, GWF07, HLU06, HWG06, Law07, OGL09b, PKS00, SdV06, SRR06, SKW^{+}06, Wri07]. evolutionary [BHZW08, EH04, MKR07]. evolving [Clo01]. exacerbates [HJM07]. examination [BR03, BPCvdM06, CW03a, GDGSP09, LB08b, ROCB09, SM00a]. examinations [JB07a]. examining [BBE03, CES^{+}03]. example [Bar03, Kup03, SNP03]. excavating [LVZW06]. exceeds [MBBW08]. Exceptional [EF07]. Exceptionally [VGW00]. excess [Koe05]. Excessive [SSLM01]. exchange [DSM01, DRG^{+}02, HTBT07, KWHD08, KKI^{+}06, LWB^{+}00, LBH01, PQRAC08, TMD^{+}07, TDUN07, VHB^{+}07]. exclusion [DGL01, HK03a, LV05, MN04, PWLC06]. excreta [SZL^{+}05]. excretion [BP01, MMTT00]. exemplified [BBG^{+}06, WCZ^{+}09]. Exhaustion [KYYM00, GDdS^{+}07]. exhibited [PGPL03]. exigua [JMU09]. existence [PSC05]. existing [EPM06]. exoculata [PGS^{+}00, ZLGC04]. exogenous [KYML02]. exopolymer [BPVL03, EP01, LWH08, MWMH05, MGF^{+}03, Pas00, Pas02, RYF01]. exopolymeric [RMGL07, Tho04]. exopolymers [PA05]. exopolysaccharide [SSdWM00]. exoskeleton [IF09]. Exotic [PDCH06, LW02, SMC09, TK01, VJ05]. exotica [TD01b]. exotoxins [SKU^{+}08]. Expanding [Nae06, VGMC01]. expands [SN09]. Expansion [FLG02, SAB04, Arr05, BOR08, CH07d, ESB^{+}04, TGGM05, WHM^{+}01, MVBGM04]. expectations [Kne03]. expected [FR08b]. expelled [HR07, RGL01]. expenditure [JRS^{+}06]. experience [BK06a, GA03, WWTH06]. experienced [KSK07]. experiment [BDW^{+}00, DAA00, GRCdLvB08, HVP^{+}06, HS07, MK02, RM00, RGDN02, RHGB02, SCD^{+}09, SVL^{+}04, VNdG^{+}06, WC03]. Experimental [AD00, HMH04, HBNA04, HSBD08, LABR^{+}06, MRN07, MLW^{+}02, RDC06, SDFM00, TTN01, VBAR07, ADA00, AMR01b, ABBF01, BSZ05, BBC08, CPK00, CK04, CRCN07, CS01, DAG^{+}00, DTAF00, EÖM08, FH08d, GG09, HKWM07, HJG06, HGMH06, JRC06, KET^{+}03, KSI09, LMMT09, LV05, LTFH06, MCL07, MF03, PDO01, PTB03, PM00, PH01a, PSRB00, QRS00, RNDF^{+}04, RU02, RR01, SYVC00, STM06, TFA^{+}08, VD00, WASW03]. experimentally [KGRW06]. experimenting [CTHF01]. experiments [AW03a, AHV01, BS07a, DGM00, FDDS09, GRLML09, HM03, KECJ06, LMD^{+}02, LH02, LFSR06, MBA^{+}05, MJG08, MBM^{+}00, NNS01, NB06b, OTBR00, OS01, PS07, SLT^{+}01, SDRH01, Tya09, UB03, WB01, WBT^{+}08, vKH00]. experts [CWR^{+}03]. explain [BS08a, CPSC01, GSRH02, Kne09, NvW01, PSRB00, PCHM07]. explained [GSM05, Som00]. Explaining [MB00a]. explains [SM07b, SYPO09]. explicit [BM00, SW02b, SMP01]. exploit [JBF08, JTR05, KRK09]. Exploitation [BCC^{+}04, CWC00, DBR^{+}07, NLP^{+}01, TWM^{+}01, DSPHV04, Ogu07, Tom08]. exploited [BLT03, FF04, KdLCM09, MB07d, MGC03, PMG^{+}06, SF07b, SWG^{+}09, TP00b, WHB^{+}06]. exploration [MWTB00]. explore [HGS^{+}02, HOT08, JPG05, MS03]. Exploring [BDDG02, GDdS^{+}07, KRKS09]. explosion [KUOI06]. Export [GM08, HCH07, BJR04, BAE^{+}08, CGCDA02, FR07, HSHA06, tAHSHA07, JPMGS08, KVS08a, PBT^{+}09, RWR^{+}00, RSC08, RRH09]. exportation [MdL06, VMM^{+}04]. exposed [AW03a, ABB^{+}00, BVVM02, BWB^{+}06, BSPK03, CHMR05, CPB04a, CK05, DP01, GGHA02, GB05b, GTFB03, HGG^{+}03, LBF^{+}07, LÅCN01, NZM07, ORPND01, PKS07, RCDC06, TYT^{+}04, TMJ^{+}05, WWV08, YLL^{+}04]. Exposure [ACT09, BDGS07, DQ05, HHW^{+}07, HJG06, TSC^{+}07, XW02, AC02, BVMBC07, BvK07, BHR08, CNW07a, CBP^{+}01, DD05, EG05, FNB07, FNDP08, GR07a, HS08b, HR08, HFD08, ISG04, JQ05, KWTR08, LGFM00, MB08, MFJT08, ML00, ML05, NK03, NW04, NMCT03, OAM02, PJP06, PJ06b, RBW03, RBG04, RWJ02, RBSP00, SKMS06a, SB03b, TPGQ07, TH06, Wan02a, WW03, WJ06]. exposures [KLR^{+}06, NMBH02]. expression [BGL^{+}07, CSG01, CSvO09, HKH09, LMF^{+}01, MMGC06, POH08, RCC^{+}09, RPS^{+}04, SSG^{+}00, SRLA04, TTM^{+}05, TBVJ08]. extended [PJ06b, MH09]. extender [LJ09]. Extending [RSR02, MLC08, RFSC08]. extension [Arm01]. Extensive [BWH09, PTAR07]. extent [BML^{+}08, BLP^{+}02, JCMR^{+}03, JKES07, WAN^{+}01]. external [CBAd00, PV05, WB05a, WB05b]. externally [RCWYK07]. externally-attached [RCWYK07]. extinction [JLCA01, MAG^{+}07a]. extra [AKM09]. extra-embryonic [AKM09]. Extracellular [MLG03b, PB07, AJ03, GKM02, HHT^{+}09, HSD03, LHQ05a, LHQ05b, SLKF03, WS02, dBS01]. extractable [HSJ08]. extracting [TML03]. Extraction [BDC00b, FRJ09, She08]. extracts [DDTQ05, EP00a, NGdN^{+}07]. Extraordinarily [CM08]. extraordinary [QWY07]. extrapolative [FAFR03]. Extreme [BBA06a, BHFC08, BK03, Ber00, Mil07a]. extremes [MEB^{+}07, WGR07]. Extrinsic [PB00a, HWB^{+}09, SB06]. exudate [TD01a]. exudates [PBVP09]. Exxon [BBD^{+}02, DBJ^{+}00, DBF^{+}02, EBT^{+}02, GSM^{+}02, PMGE01]. eye [SN02]. eyed [MEHD07]. Ezo [TSE^{+}08].

f [SWH^{+}04b]. f-SPI [SWH^{+}04b]. Fabia [HS05]. fabricii [ZKHP07]. facilitated [WHM^{+}01]. facilitates [CU06, HDW02, NYvW03, ZA07]. facilitating [RK08]. Facilitation [BMC^{+}09, BBC08, HLTT09, WPG08, ADF^{+}09, APM^{+}01, BCS^{+}06, KKG06, KT04, LGS07, MN04, Mou04, Rue07, SB06, vWCAB07]. facilitators [IE06]. facilitatory [UM08]. facing [GB09]. factor [BBC^{+}00, Gar00, HDDR08, LMW^{+}01, XW02]. Factors [AS06b, CA06, CGM00, DCD^{+}03, HDCA06, HMHB03a, MMBF09, PBBC06, SFL^{+}09, ZPP09, dBMSL02, AAPO00, ABB02, BPBH09, BMV02, BP08, Bra01, CDML06, DSF^{+}04, EAF^{+}05, FC04, GMHA09, HRR^{+}07, HCM01, HDP^{+}06, IFL02, JLW04, JBCL03, LSH00, MUT04, NA04a, PB00a, PMH07, POH08, PCH^{+}05, Ree03, RGP^{+}09, SAB^{+}02, SW02a, SB06, VLP^{+}04, VMN^{+}03, War05, WAW^{+}09]. Facultative [SCIT09, TDC^{+}07, TA01b, Orv05, PH01b]. FADs [DK03, GTD^{+}08, JLW04]. faecal [BND02, BGS^{+}02b, IAT^{+}07, JB07a, LEÖK^{+}02, RRS05, RJBW03]. faeces [CSTC01]. failed [MRMA07]. fails [BD06]. failure [MARM06, SHZ^{+}05, WHRS05, Wes03]. fair [RASB03, RASB03]. Falkland [PCS^{+}07, PISL02]. fall [KJB^{+}06, TSW^{+}09]. fallow [LBB08]. falls [BKP07, PKA06]. families [MCC^{+}00]. family [BJC02, NNG03, ORD00, WKH02, ZK01]. famine [FCP02b]. fan [BMK07, DKAH02, Ryp08]. Fanafjord [HB07b]. fans [PHP^{+}03a, THYBS09]. far [DB07b, KWV^{+}05, LHT06, RW07]. far-field [LHT06]. far-red [RW07]. Farfantepenaeus [RRRM08]. farm [BSAR03, CWMD06, CHT06, GSL^{+}06, GSRH02, HN03, KHGA02, VMM^{+}04, WHCP^{+}04]. farmed [ABS^{+}09, MSF^{+}09, SMM^{+}09]. farming [EHK06, GXL^{+}08, KH00a, KTSR02, NWFV06, SVC^{+}01, VYM06]. farms [DSJBS^{+}02, DFJSJ^{+}05, DUSJ^{+}09, DCO^{+}08, FHDA^{+}07, GSS09a, HSWG^{+}06, HK02a, HK03a, LKW03, MKG^{+}05, PTA^{+}09, SSMM06, UDB^{+}09, WW05]. Faroe [BMM^{+}04, JHB06, Mag02]. farreri [ZYZ^{+}06]. fasciatum [DNL08b]. fasciatus [Mal02]. Fast [MHB^{+}09, Cam01, DS01, FHK^{+}08, GHD^{+}02, GKB09, HGG^{+}03, INH^{+}09, KFJ^{+}07, KKL07, LG01, MVH^{+}06, MOOB06, MSH^{+}03b, RSK04, RCF07, SMHG09]. fast-growing [Cam01, MVH^{+}06]. fast-ice [KKL07]. fast-paced [DS01]. fasting [GBW^{+}07]. Fat [IFL02, SNY^{+}05, MPS^{+}08]. Fate [CPKV00, GFC^{+}02, HKM01, HBT^{+}08, IAT^{+}07, JFR02, LEÖK^{+}02, MdL06, RRS05, SSH^{+}04, SMS06, AP05b, BCBQ01, CBA02, CS08, DPM02, GMTN00, HRR^{+}01, INH^{+}09, TCRT^{+}07, TDRM08, vdBÓME01]. Fates [XW03]. ’Fatty [TBBI04, BSIS07, CPG^{+}09, GKWK02, GNAD^{+}03, GNHB^{+}05, KKK^{+}05, KCV^{+}06, KFK^{+}07, KSIG08, KSI09, MT00, MMBT03, RBGM06, BN09, BRI^{+}07, BKLGN05, CLP08, DK06, DPGN08, ESDKM09, GSL^{+}06, GNAD^{+}04, GFN^{+}09, HRR^{+}01, HK03b, HBW^{+}05, HMY^{+}08, HYR^{+}09, HPBT03, HPB^{+}04, HCK^{+}05, IFL02, ISK07, JBS08, KKIV01, LCGN02, LYL^{+}07, LBPH06, MMC^{+}08b, NFN^{+}09, NWIT08, ODA^{+}09, OTB^{+}00, PTD05, PJN01, PJN03, PGS^{+}00, PAB^{+}02, RF08b, RTH06, RMC^{+}07, SKFPS02, SZ09, SKDP02, SPGS03, TGNT05, TBI08, TBI^{+}09, WHP00, WP03, WISH09, WIB08]. fatty-acid [HPBT03]. fauna [AP05b, BSBD^{+}04, CDG04, CNF09, DD05, EH00, GAB^{+}03, JFD09, MFCM05, NR00, NRH^{+}07, PJP06, RAH^{+}02, She06, SJ09a, WBL^{+}06, WP07, WMW^{+}08, WKL01]. fauna-microbe [WMW^{+}08]. Faunal [AR00, BB00, DR01, PCSS01, SFHS00]. faunas [NKK^{+}01]. Favella [KS06]. faveolata [AS06b, GB06, Les00, VFM06]. Favia [TLLKC04, dBKSV00]. Faviidae [Lam00a, Lam00b]. Favites [KS08]. favors [BBZM08]. FDA [GMP07]. Fe [VNJ06, VFJ06, VF09]. feasibility [RH04]. feasible [MBW07]. Feast [FCP02b]. Feather [BGSBC02]. feathers [HM06a, OP08]. features [ADF^{+}09, BCL03, BIR00, CBAd00, GGDB09, HHKH05, JWR05, MFP^{+}05, MG05b, NLP^{+}01, RFG04, Und04]. February [WG01]. Fecal [SS09a, CMG04, HRR^{+}01, Kiø03, KSA03, LCGN02, MTN03, OSA05, PK06, PI08, SSH^{+}04]. fecund [MS08]. fecunda [CTP02]. Fecundity [BBT02, BH04, MM03, BSS05, FBM08, FAB^{+}07, LT00, TCRT^{+}07, TIM^{+}01]. Fecundity-time [MM03]. fed [BC01b, FB00, HM03, OTB^{+}00, SSO^{+}01]. feed [ABB^{+}00, CSR^{+}00, HGKJ01]. feedback [Hen07, SW02b]. feeder [GSM^{+}08, KSA03, LHB09, Orv05]. feeders [BDG^{+}07, DBR^{+}07, GLS02, LWH08, RNL00, RKLK07, TSJ05, Uth01]. Feeding [ASC06, BCB^{+}06, BGF^{+}06, Bat07, BMdNP09, BP03, BZLD02, Bra08, BCB^{+}08b, CT05, CIHL05, CSTC01, CTP02, DMV07, DBCB06, DVLH^{+}00, FC04, FK00, GM04, GBW^{+}07, GDR^{+}04a, Gru03, HB03, HSCH07, HBW^{+}05, HPBT03, HBTD04, ITH^{+}03, JKK^{+}05, JSK^{+}07, JKS^{+}07, JHH01, KJ04, KFS^{+}09, KSKL^{+}03, LT00, LP01b, MHCB08, MABB03, Mil07a, MKWO00, MR07a, OHN01, OST07, Par06, PP00, PMNP^{+}01, PCRR08, RRS^{+}09, RN06, RGCL^{+}05, RGVC07, SDPT06, SCT^{+}09, SCS03, SDK^{+}08, TNT06, VN02, VN05, WKH02, WKIK04, WKY09, WCF08, YI03b, YSS03, AHL^{+}01, AMR01b, AL08, ABY00, Ark09, APP^{+}09, AL09, Bay02, BPM05, BGPL08, BKSS07, BTB^{+}09, BMSD00, BH06, BH01, BEE06, CEB^{+}08, CPP^{+}04, CWT00, CBTW02, CBDH02, CVK00, CCK03, CDK^{+}07, CS05, CRH01, CTHF01, DCP^{+}08, DLLB03, DER01, DMS^{+}05]. feeding [DvD02, DOML^{+}07, EEA01, FM02, FKA^{+}06, FHN^{+}09, GMZvdL07, GBHO^{+}08, GHJ02, Goc04, GPS07, GJ03, GLCL04, HPW^{+}00, HHG^{+}07, HS05, HM03, HTM^{+}02, Hel00, HL05, HLPHA^{+}01, HH00b, HSD03, JS01, KLHO02, KT06, Kam00, KS06, KNTK05, KAMR00, KHG06, KPN03, KKBH03, KW00a, KHvW02, KLW^{+}09, LKR^{+}06, LCL^{+}03, LSMR05, LB02, LC08, LP03, LUIAH03, LVA07, LPBD00, LME05, MVN^{+}06, MNTS08, MSH08, MDBG07, MCS09, MWR07, MMVB07, MK03b, MAW06, MTN03, MR06c, MOP^{+}08, NDV^{+}00a, NvdV04a, NvdV04b, NBN^{+}08, NGSV07, NTGK09, NATR^{+}05, NWIT08, OLWB06, Øre00, OJ02, PGM05, PdNS06b, PWM^{+}08, Pen00, PH01a, PH01b, Phi05, PRvdL^{+}07, PCH^{+}05, PK05b, Pra01, Pra05, Ran02, RW06a, RRS00, RRIH^{+}02a, RFPS^{+}02, RMH^{+}09, RHG05, RCR03, RGA^{+}02, RL05a, Rin03, RS00, RWPJ04, RCF09, RGMH08]. feeding [RH08, SMMM07, SK04, SK05, STPA04, SG05, SV08, SWH04a, SK03, Sot03, SSS06, STJ03, SPGS03, SCH^{+}07, SAH^{+}05, TDT^{+}03, Tan02, TCD01, TJS08, TA03, TLNH01, WPG08, WDT^{+}08b, WÓ02, WJDA06, WWWR09, WC03, YLK00, Zam03a, vdMM03, BEW03b]. feeding-related [MOP^{+}08]. Feldmannia [PVGO01]. Female [EB04b, MH00, SKP^{+}07, APG^{+}09, BGQR01, BMB02, BT06, BKH^{+}09, BSM02, BM05b, BGRD00, CS09, CDF^{+}08, CH00b, CCL^{+}06, DC06, DRDM09, FPB02, FTB09, GGFSP06, GDL^{+}01, HBDD02, HJH^{+}02, HTM^{+}02, HHWT02, HM01, LAH07, Lus03, MCM05, NH05, OAM06, SAWG05, SR04a, SHZ^{+}05, SHK05, SB07, YA09]. female-derived [HJH^{+}02]. Female-female [SKP^{+}07]. females [GRM00, RRIH^{+}02a]. Feminisation [LBC07]. femorata [BEE06, WS01, vdBÓME01]. fencing [MAH^{+}08]. Fernando [BKSS07]. Ferritin [DGB^{+}02]. fertilisation [HJM07, MSK00a, PTP01, SS04a]. fertility [MHB01a]. Fertilization [BT01, LBPS03, BCB02, BL00b, BB08b, CB08c, Den08, FR08a, Hue08b, LOT^{+}00, LCL^{+}00, Law08, Lei08, MC05a, Mar02, MSY02, Orb08, WBT^{+}08, dBGLT08]. fertilize [PB00b]. fertilized [DPD^{+}08, HLH^{+}05]. few [She01]. FFA [GHD^{+}02]. ficiformis [RCC^{+}04]. fiddler [GKK03, KWD05, LDT07, OJ04, PJJ^{+}05]. Fidelity [AWBO09, BCLRL07, BHPC06, BLDG01, BCB08a, HPH^{+}01, HALL05, KAR04, LAK07, MG05b, TLC06]. Field [BD01, SESP03, SAHF03, SH06a, VBB06, Ark09, BLF06, BDW^{+}00, BKB05, BHMR01, CSD04, CTVP03, FDDS09, FS06, FSSW^{+}00, GSM05, GMR^{+}02, HVP^{+}06, HMW^{+}00, HJM07, HAA00, IF09, Kam00, LHT06, MK03a, MK04a, MBA^{+}05, MSH^{+}03b, MK03c, MTCU08, NPLZ00, OTBR00, PSRB00, RBW03, RFLC07, SCD^{+}09, SMG02, TF04, VN05, VNdG^{+}06, WG09, WS04a, Wri07, PWL^{+}06]. field-based [MTCU08]. field-collected [BHMR01]. field-exposure [RBW03]. field-spectrometer [CSD04]. fields [Bar06, CM05, MCP^{+}06, MB01, WB06, YFH^{+}02]. Fiji [MC08a, PPI04]. filamentous [BRBPL06, CNPS05, PdNS06b, SSGJ00]. filaments [NDB04]. filholi [BRP00]. filibranch [BDR04]. filiforme [HKWM07, SZ08a, SZ08b]. filiformis [MLR02, OKP06]. films [HLQ02]. Filter [MR06c, GLS02, LKR^{+}06, MSA07, NATR^{+}05, OJ02, Phi05, RGA^{+}02, TSJ05]. filter-feeders [GLS02]. filter-feeding [LKR^{+}06, NATR^{+}05, OJ02, Phi05, RGA^{+}02]. filtered [LCC01]. filtering [LSG05b, ZYZ^{+}06]. filtration [CNW^{+}07b, Cra01, KR05, MPDC03, Rii01b, Rii01c, Rii04a, RL07]. fimbria [SO01a]. fimbriata [PDM^{+}04]. Fin [JTR05, GPS07, RBGM06, SCM^{+}09a]. final [OYT^{+}02]. find [JK04]. finding [BK05a, BK05b, KBT05, RB05b, TLFS06]. Fine [ARGS01, CFA07, HHF^{+}07, HFT^{+}09, HHWT02, JN09, McI03, BCJ05, GYD04, JWR05, KW05, LBvdH03, NJPI03, ST08, TML03, Wal07, WBP^{+}06, WRH05]. fine-grained [KW05, WRH05]. Fine-scale [ARGS01, CFA07, HHF^{+}07, HFT^{+}09, HHWT02, JN09, McI03, BCJ05, GYD04, JWR05, LBvdH03, ST08]. finfish [FF02, HDM04]. fingerprint [UB02]. fingerprinting [ZMP^{+}03]. fingerprints [GK00a, SRZZ01, VRST^{+}07]. finite [TÅ07b]. Finland [KLRS00, LLSL06]. finmarchicus [ITH^{+}03, BT06, BCC^{+}03, BSPK03, CWT^{+}01, EO04, Gis05, HHC00, HB07a, HGHB00, HBN01, JBT^{+}07, KH08a, MNTS08, MMC^{+}07, MCM05, MTN03, Pea03, PTS00, PRM^{+}09b, PJ08, RRIHC02, RBRSP00, SHH09, SKFPS02, SGH^{+}06, TJEM02, TJM03, TBVJ08, TBH06, WCBD01]. finned [BBHT02]. Finnish [RSK04]. Fiordland [JWM09, Win09]. First [CBO04, DMK^{+}08, HF03, HGAS^{+}01, JT02, SF05, TND^{+}05, JPMGS08, LVA07, OGMC06, RMH07, UB03, vMRO03]. first-year [JPMGS08]. Fish [BGM09, BFB^{+}09, GA08a, HMD^{+}07, Kap09, KTSR02, LH02, MCC^{+}00, ÖCOB06, TC01a, YYK^{+}08, YBGP06, AB00, AR02a, AWBO09, AFHS09, APA07, AP04a, AP04b, AMG^{+}07, AO01, ARBF02, AGDP^{+}06, Arm00, AKG02, AL09, BPBH09, BF07, BS07b, Bar02b, BF01b, BSZ05, BBG^{+}06, BVVM02, BLHB04, BGPL08, BDCvB^{+}09, BSAR03, BB02a, BH06, BL07, BGL^{+}07, BPA^{+}08, BES^{+}08, BG02, BCC08, CH06a, CC02, CWE^{+}08, CLSM09, CC03a, CDF^{+}08, CWM^{+}07, Chi02, CS03a, CDJ^{+}07, CWM^{+}09, CCU07, CHPS04, CBL^{+}09, CRCN07, CEBR07, CHA06, DSL07, Das03, DF04, DFSS08, DSJBS^{+}02, DK03, DFJSJ^{+}05, DUSJ^{+}09, DBN^{+}09, DTV09, DCBvBE09, DCO^{+}08, DGC^{+}05b, DVNvdV07, DGdG^{+}09, DE08, DGB^{+}02, DS07, ESDG08, EB08, Enb05, EÖM08, EvBB^{+}09, EOK^{+}08, Eva08, FGPS04, FA06, FTDK07, FS06, FAJM07, FPKC05]. fish [FBBV09, FM02, FA02, FF03, FFN^{+}08, FBJ00a, FB03a, FH08b, FE06, FHDA^{+}07, FBM07, FCH^{+}08, FOH^{+}08, GBdS^{+}05, GTD^{+}08, GM04, GNPB07, GMC^{+}09, GSL^{+}06, GXL^{+}08, GYLÖ06, GL00a, GMK^{+}03, GK00a, Gil02a, GS05, GK03, GM05, GB05a, GDKN08, GCS06, Gru00, GL02, Gru03, GTFB03, GBNÖ08, GCM01, HS08a, HP09, HSWG^{+}06, HG08, HF08, Hal00, HWC^{+}06, HW09, HMKR05, HTW^{+}05, HCB^{+}07, HCS^{+}07a, HCGC09, HLBS07, HBRJ06, HGAS^{+}01, HGA07, HGKJ01, HJK01, HJ05, HSB08, HM06c, Hue08a, HRC^{+}07, HLH^{+}05, IFL02, JLW04, JRC06, Jen05, JCL09, JS08a, JS08b, JBS01, KRF^{+}05, KTK^{+}08a, KKK^{+}05, KH00a, KKG06, KHGA02, KFT08, KSK07, KKBH03, KHP09, KRP^{+}06, KSHO01, KN00, KS05, Kro05, KNF^{+}03, LC02, LSH00, LEF^{+}06, Lec05, Lei07, LB08b, Ley05, LPHH^{+}09, LPS08, Lin04, LBPH06]. fish [LC09b, LW05, LIM^{+}06, LLW09, LS04, LKW03, MMB03, MD08a, MTF^{+}08, MKG^{+}05, Mag02, MGL^{+}07, MCR03, MG09a, MACS07, MHP03, MMVB07, MB07d, MCP^{+}06, MG05a, MGM^{+}07, MGW^{+}09, MS06, MCM^{+}03, MSSP06, MPS^{+}08, MAW06, MS04b, MS05, Mil07b, MVdMFF08, MJKH03, MJKH04, MBB^{+}04, NKK^{+}01, NvdV02, NRvdV^{+}02, NvdV04a, NvdV04b, NvdVVD06, NA04a, NW02, NMJ02, NH03a, NH03b, NLHO09, Ogu07, OMP^{+}05, OT08, ÖNCNG05, OM04, PDS^{+}07, PGK^{+}08, PP00, PGP03, PTA^{+}09, PMP04, PCHM07, PB08b, PEG^{+}00, PGPL03, PFT^{+}08, RM05, RW07, RDS^{+}09, RFLC07, RWP07, RP06, RGF^{+}03, RHSN07, RGT06, Rue01, RW06c, RB05b, RSIS09, SS00a, SESP03, San00, SSMM06, SJR00, SLFM03, SS00b, SVG07, SF07b, She06, SJ08, SJ09a, SR03, SF02a, SS09c, SNY^{+}05, SFH00b, SD00a, SA03b, SH05, SHJC08]. fish [SSL^{+}07, SA06, SLM07, SP04, SGCP06, SLHR09, SSW08, SF02b, SBLF09, SMS06, SB00, SWG^{+}09, SIK^{+}02, SRJ07a, SH00, TOK08, Tor09, TFL^{+}04, TPD^{+}05, TCWR08, TBSJ^{+}04, UDB^{+}09, ULL03, USG^{+}08, VJ03, VLW09, VRST^{+}07, VAB09, VNdG^{+}06, VYM06, VPG04, VMM^{+}04, WW05, WM04, WT06, WW03, WHRS05, WAK^{+}09, WH00, WCRB03, WKL01, WA03, WM02b, Wil04, WWS^{+}05, XW02, ZWYL01, ZK01, vdVBGW09]. fish-assemblage [DFSS08]. fish-eating [FE06]. fish-farm [BSAR03]. Fish-habitat [GA08a]. fish-killing [HGAS^{+}01]. Fish-seaweed [LH02]. fished [CSS00, DFSS08, RGS^{+}09, SJ09b, UB02]. Fisheries [GRB^{+}03, Law07, MKR07, AMS^{+}06, AO02, BRGS^{+}07, BHC^{+}00, BHZW08, CJFM09, CMLG07, CWT00, CCL^{+}08, CHG05, FVB^{+}09, FEO07, GAAB^{+}08, GWWC00, KH07a, KWP07, MGW^{+}09, PPC09, PRL^{+}09, RHDC05, RSC08, SMA^{+}03, SAH^{+}05, VBFF08, WLB08, WH08, Wri07]. fisheries-independent [SMA^{+}03]. Fisheries-induced [GRB^{+}03, Law07, MKR07, Wri07]. Fishery [FCC^{+}08, WWMF06, BWMA02, BD06, BDF09, CFG06, CPB04a, CPG08, EFL^{+}09, FF02, FT00, GHT^{+}01, GQR06, GZY06, HGSH08, Hou09, HK04b, HW00, Jen05, JCG^{+}02, LTL06a, MRGC09, PCMB06, PRvdL^{+}07, PPJ^{+}05, SH06c, SRJ07a, WGF09, WBR07]. Fishery-independent [FCC^{+}08]. Fishes [IHK08, LH06, AP05a, AE02, AH08, AA04, BBBSPH03, BBG^{+}04, BP08, BTG^{+}08, BMM^{+}09, BSP03, CM02, CDAM06, CLHT05, CCC05, Dav00a, DB03, DB05, DGNvdV05, ECM^{+}05, FWF02, FF04, Fis05, FVS^{+}08, FMZvW02, FEFT08, FD02, FB02a, FB02b, GNP07, GPHC09, Gri02, Hal04a, Her06, HLR04, HS03, HR08, HOVG07, HMJ^{+}09, JS01, JBF08, JH06, JBA02, KRKS09, KK02, KBP06, KMB^{+}08, KPEF00, LCE03, LHL^{+}07, LHH09, LYMCG06, MR06a, MCBT07, MB07a, MFJT08, MBH^{+}00, MKWO00, MBG^{+}09, MCN09, NDV^{+}00a, NDV^{+}00b, NKK^{+}01, NWT00, OP02, PD07, PA06, PPC^{+}05, PPSP07, PS01, RS00, RC09, RTH06, RM09a, SSL00a, SPL00, SBSPS04, She01, SCT^{+}09, SYWM05, SO02, Sta09, SMF^{+}02, SCAF00, TM02b, TJM00, Trn02, TWT08, VKH08, VHK09, VHW^{+}07, VW00, VBS06, VMM^{+}04, WKH02]. fishes [WBC07, WPH01, WWMF06, WWTH06, WMB00, WM01, WJDA06, Wil09, WJB09, Zam07, ZSR03, dlMPN^{+}03]. Fishing [BLHB04, EvdKB^{+}08, THF^{+}01a, THF^{+}01b, ASA^{+}08, BRGS^{+}07, BHZW08, DBN^{+}09, FD02, FBM07, GKAS08, GFN^{+}09, HBG^{+}03, HCV03b, JCL09, JLCA01, KCH^{+}06, KS04a, LHT^{+}02, LCT^{+}08, LS04, PE05, RGS^{+}09, SG06a, SD01b, SLL07, SLL09, Sta09, SMCE07, SMVF03, SAM^{+}06, TSJH05, WAK^{+}09, WAK^{+}04, dJTD07, RCP^{+}05]. fishing-induced [HBG^{+}03]. fission [BUCC09]. fissiparous [SBM02]. Fitness [JRA02, CRH01, FLL09, MM03]. fittest [Bra05]. five [BF03, FLK^{+}06, GL00b, ITN09, VRST^{+}07]. fixation [BWK^{+}08, CGCMPA01, CMHP06, DGBS08, EB04a, Fen08, FB03b, GRS07, GPW^{+}02, LJ06, LFRR^{+}07, MCH^{+}06, MP05, RAE07, SFKT09, WVL01, WNS^{+}05, WBN^{+}00, vDHHM08]. fixed [Gla01, KKLL02, PFZS^{+}06, WdBS02]. fixing [FPC05, MSH^{+}03a, MPDS07, SBC^{+}06]. Fjord [AI02, GBB^{+}05, NRPS^{+}01, SHS06, ACM^{+}02, ASES07, BND02, BRS07, BSR09, DFPG^{+}03, Fis06, GKWR02, JGH05, KFE^{+}03, MD08b, ORS02, RTB06, RDD^{+}02, RAH^{+}02, RKGH01, SSJ^{+}02, THM^{+}03, TDRM08, VFJ06, VTS^{+}02, WL06, WGL03, WKPK05]. fjordic [SWM03]. fjords [ADS^{+}09, FH08c, KTK^{+}08a, RWR^{+}00, ZZT05]. Fladen [EO04]. flagellate [NHNB01, TGY^{+}06, WL06, vLvSGS05]. flagellate-microzooplankton-dominated [NHNB01]. flagellated [AHK^{+}02]. Flagging [SAC^{+}00]. flanking [HFM^{+}04]. flapping [RCWG^{+}06]. flat [ABB05, BWB^{+}06, BWP^{+}06, BdVD^{+}01, CTD^{+}09, KT06, KWKF03, LWB^{+}00, LBH01, MMTT00, MBR^{+}05, MT00, MLM^{+}02, SN06, SR09, SDW07a, SHM^{+}01, TBR^{+}03, VMV^{+}09, WBSL00, WBH^{+}04, YH00]. flatfish [BP02a, Ber00, LSH07b, LR06, NGBG07, RSSA07, SF05, SCAF03, WP07]. flatfishes [GRWB02, RST04]. flathead [RZS05]. flats [CDK^{+}07, DB07a, ENAT07, ORPND01, SR08a, SAC^{+}00]. flavescens [CD08, CKB^{+}09, HK04b, UP01]. flavimanus [NW05a]. flavolineatus [KFS^{+}09]. flectosiphonata [KT06]. fledging [CKF^{+}03, ÖBFO01]. flemingeri [KI01]. fleshy [BP02b, CU06]. flesus [MTL^{+}03]. flexibility [BCLRL07, GBW^{+}05, HHG^{+}07, RB08]. Flexible [ICH04, MBG^{+}09, FBN^{+}07]. flexuosus [WG02b]. flight [BLH01, RCWG^{+}06]. flightlessness [DMA02]. flipper [GLÅ^{+}03, YA09]. Floating [PFZS^{+}06, EvBO^{+}08, HC02]. flocculation [SLT^{+}01]. flocs [JKNJ07]. flood [BSvdH^{+}02, HFA09, KKA^{+}06, YH00]. flood-ebb [KKA^{+}06]. flooded [SMS06]. flooding [Zam03a]. floor [KTF^{+}01]. Florida [CSO08, MIL^{+}00, ARGS01, BYH^{+}06, CF09, CSS00, CSM02, CL05, CT09, CH05, DSL07, DE08, EBC04, FFN^{+}08, GMC^{+}09, GHH^{+}04, HHV06, HKZ04, HK03b, HALL05, KVH02, KH07b, KCEB09, LDL^{+}04, LSG08, LC08, McG02, MMPD09, MIH01, Niz07, PMH07, PCJF06, PBBC06, PI07a, ROCB09, RV05, SC06a, SGCP06, SZBZ04, SCS^{+}07, Tor09, VHKW00, WPRS05, WLJ08]. flotsam [BF03]. flounder [BWCW02, KET^{+}03, KMAC^{+}02, MPM^{+}02, MPM^{+}03, MPSJ06, MTL^{+}03, MGTK00, NRM05, SAG08, SK07, STM06, SMP01, TC03a, Tay03, Tay05]. Flow [Ark09, GSC^{+}01, WG02c, Abd03a, Abd07, BAA04, Bib07, BSR09, BMW^{+}00, BP06, CT06b, CT09, DESK01, DEV07, ERR06, FMCB05, FHPW07, FGS00, GLJ07, HSBD08, HL05, Hun04, Hun05, HMFC07, IT06a, JBF08, KPKD06, Kim02, KSR^{+}07, LKR^{+}06, LWJC04, LZM^{+}03, MB07b, MMMPB05, MSTS^{+}06, NvW01, NYvW03, NvWY05, NKS03, PSPZ09, PvDMB08, PLMS04, PVGO01, PSGP04, PW07, PLG09, QRS00, RL05a, RS00, RFB07, SL05c, SML05, TdMF^{+}08, TC03b, TDC^{+}03, Tyn04, UBM^{+}07, UB01, VPWE09, WB06, ZBNW09, vDvTJD^{+}09, vWNYS05]. flow-dependent [CT09]. Flow-induced [WG02c]. Flow-mediated [Ark09]. Flower [SC06a]. flowering [RHBW09]. flows [HCCY02, Mil07a, OAGV07]. fluctuating [Enb05, MM03, SBM02, vLvSGS05]. fluctuation [SHH08]. Fluctuations [HWGB02, BIR00, CS02a, GSRH02, HGS^{+}02, HGA07, KHF^{+}00, MR01, RF06, SIK^{+}02, WÓ02]. Fluid [KPT01, KT01, Jak01, LZM^{+}03, SMB^{+}03]. fluids [PPI04]. flume [Fin00, PvDMB08, RWBS03]. fluoranthene [LGFM00, LGVF00]. Fluorescence [LCF02, FK04, HHAH00, JPC^{+}06, LG01, MOOB06, NCÁSGP05, POH^{+}02, RMWO09, SC01, SO08, SdSC01, SMHG09, WM02a, WMS03b, YF04]. fluorescent [SWH^{+}04b]. fluorescently [CNC00]. fluorimetric [GHD^{+}02]. fluorometer [FJ01, Joh04, KGB^{+}01, LG01, MSH^{+}03b]. fluorometrically [BCC08]. fluorometry [FHK^{+}08, RSK04]. flushing [ZGV00]. flux [BND02, CMG04, CCL03, DCD^{+}03, DL01, DLB04, FHPW07, FFML02, GK00b, HKY^{+}02, KVS08a, KT06, KKI^{+}06, NMA00, NRPS^{+}01, OSA05, PFW^{+}02, PK06, SRL^{+}02, SES01, Sve02, TDRM08, VPP04, WBSL00]. Fluxes [LR02, CB00a, CHW02, CCMK03, CVK00, CIFN00, CCB^{+}09a, CS01, DP02, EF02, GFP^{+}05, GCHC00, HAA00, Kar07, KGHW05, KPT01, LBF^{+}02, LTMDB02, Lar03, LFF^{+}09, MBW02, MCCT07a, MCCT07b, MFS07, MLG^{+}03a, RATD07, SMG00, TSF01, TGM^{+}03, VTS^{+}02, WE04, WBN^{+}00, WDM^{+}09, WWH05, YH00]. flying [CCC09, DGC^{+}05a, IMW^{+}02, IMSO09, PSC^{+}04, WKIK04]. foliose [MRB09]. follow [HGA07, IF09]. following [ABU01, BM02a, BMF06, BHSR00, CCB^{+}09b, DBJ^{+}00, DBF^{+}02, DLD^{+}00, EPM06, EBT^{+}02, GB01, HSSS05, ISG04, JPNM^{+}06, KNT^{+}07, KGRW06, LGB03b, LPK^{+}07, MFJT08, MARM06, Nem05, PWHW09, RV05, RWJ02, RBSP00, RAH^{+}02, SHK05, SP04, WBP^{+}06]. follows [BVC^{+}02]. Food [AND05, Bar07, BGPL08, CWT00, CBDH02, CSPA09, DBF^{+}02, ENAT07, FHB^{+}02, Fre03, Kaa04, KH07a, KW00a, LJM03, LUAIH03, LGME06, RGCC^{+}08, SAH^{+}05, VHB^{+}07, AG02a, AHC^{+}06, AHSPG08, AZKJAM06, AW03b, ABBF01, AVS03, ATLC03, AMS^{+}06, ARFP09, BKP07, BPC^{+}02, Bay02, BEU^{+}07, BCL^{+}01, BAOV06, BC01b, BPvdM07, BVI^{+}05, BJR04, BKR04, BSIS07, CC02, CVÁM06, CKO^{+}07, CNW^{+}07b, CR08, CLP08, CMB^{+}05a, CGH^{+}03, CPG^{+}09, CHV^{+}09, DBR^{+}07, DJM^{+}05, DBOR09, DCO^{+}08, DPGN08, DOML^{+}07, DF06, DLB04, DWM04, EOL^{+}05, ECK05, EEA01, FB00, FGSH^{+}05, GCLC^{+}02, Gla07, GRM00, GFPM01, GLBF07, HPS07, HKM01, HK03b, HL05, HMW^{+}00, HMK06, HR08, HVH^{+}07, HS07, HW00, IT06a, ILL06, ISK07, JMB^{+}05, JWM02, JMSM08, JRS^{+}06, JVJ09, JFA02, KPM00, KSSH02, KSI09, KJB^{+}06, KWHD08, KKIV01]. food [KVH02, KHF^{+}00, KPW07, KKH07, KSKK^{+}07, KMK01, KGS^{+}08, LG00a, LTV^{+}01, LGS07, LLY^{+}01, LR02, LR08, LNB^{+}03, LNPB04, LEÖK^{+}02, Les05, Lin02, LSG05b, LS08b, MNR^{+}02, MNB^{+}07, MSH08, MDR^{+}06, MFG^{+}00, MR04, MRB04, MTV06, MF08, MCM^{+}03, MBH^{+}09, MT00, MdL06, MSD05, Mon07, NM07, NOP02, NFN03, NAB09, NPTW05, PRB^{+}08, PMH08, PBCB03, PD07, PHS^{+}07, PTA^{+}09, PKA06, PI07b, RBW03, RGJ^{+}04, RGR^{+}01, RTL05, RH09, RE00b, ROBFR08, SRDPT^{+}07, SLGD00, SAJW04, SNH04, SZZ08, SM00b, SWW02, SNY^{+}05, SD00b, SBM02, SO01a, STH^{+}06, SG01, TYT^{+}04, Thu07, TK03a, TWM^{+}01, TSJ05, VLRJ07, VN05, VNdG^{+}06, VCBH06, WCBD01, WWAC00, WT06, WKYL00, Wan02b, WW06b, WÓ02, WL05, WDSA09, WAZ07, WTR02, WLT^{+}03, XW01, XW02, YTK^{+}05, YI07, Zup00]. food-chain [RGJ^{+}04]. food-mediated [LGS07]. food-regulated [ABBF01]. Food-web [CSPA09, ECK05, HR08, LR02, WLT^{+}03]. foodweb [Kim06]. foodwebs [MH08a]. foot [KRCGC06, Und04]. foot-propelled [KRCGC06]. footed [WLJM05, WST^{+}09, ZBDE07]. forage [AP05a, IFL02, KS00a, MBG^{+}09, MH06b, SIK^{+}02, WJDA06, WRW06]. forage- [MBG^{+}09]. forager [CWT00, SR08b, SMH09]. Foraging [AFHA05, BDG^{+}04, BBC04, BR09, BLDG01, BBS07, CAF^{+}08, CPP^{+}04, DBH^{+}02, DD01, EÖM08, FBD^{+}02, GBG00a, GSCW00, HPW^{+}00, HBDD02, HGB01, HC05, HDRC^{+}08, KKW^{+}03, LB05b, NRP^{+}02, PPS^{+}07, PSC05, PRvdL^{+}07, RHC^{+}08, TDT^{+}03, TTD^{+}06, TML03, TGT^{+}06, WBS^{+}04, WLJ^{+}04, WLJM05, WLB08, ARGS01, APM^{+}09, ALE^{+}09, BHPC06, BLD^{+}05, BCHS09, BC02, BDC00a, Bar02c, BM03, BPG08, BN09, BPM^{+}06, BTB^{+}02, BSM02, BHN09, BG02, CRJ^{+}09, CH07c, CWDG05, CW09, CEO06, CCK06, CKS05, CCL^{+}06, Dav00b, DGH08, DNL08a, EWG^{+}08, EC07b, FE06, FB02b, GBR^{+}04, GSM05, Goc04, GBW^{+}05, GBW^{+}07, GPSG06, GDL^{+}01, HHF^{+}07, HPH^{+}01, HHG^{+}07, Har00, HFT^{+}09, HFW^{+}05, HMJ^{+}09, ITN^{+}06, JCZ^{+}05, JBT^{+}07, JMJ07, KRCGC06, KB06, KCEB09, KG07, KB05, LHJD^{+}03, LCGN02, LGC^{+}08, Lei01, LB08b, LS08b, LGM^{+}09, LGME06]. foraging [MFWD08, MSH08, MB02a, MEHD07, MGC^{+}09, MBG^{+}09, MB04, MND^{+}09, NGS09, NW06, PMS^{+}06, PSKC04, PC05, PD07, PWC^{+}09, PJB03, QMF05, QMM^{+}08, RCCK09, RWE04, RBMB00, RB08, SBC04, SR04a, SRN02, STV^{+}09, SVN04, SCKC07, Sim00, SSS^{+}05, SRB04, TT05a, TE00, TC03a, TSEH08, TBM^{+}08, TC00, TC03c, VBF06, VACT^{+}08, WST^{+}09, WIB08, WBJ^{+}07, WWWY07, dBTO^{+}09]. Foraminifera [CR07b, DEvdZ03, DdNEvdZ05, HSPH05, Les05, MBM^{+}00, MMB^{+}02, NHN^{+}06, NOO^{+}08, SPGS03]. Foraminiferal [SdV06]. Foraminiferan [GPB02]. forced [GGMGR07, WIDM09]. forces [Kin03, O’D08, SW03]. forcing [CS05, FFML02, FH03, HWB^{+}09, HSO04, MBBW08, MJ07, PBN^{+}07, RRH09, Saf03, SK07, TDTR03, WMCF08]. fore [LG06]. fore-reef [LG06]. foreshore [JSN08]. forest [Fre03, HO02, IAT^{+}07, KKH07, MM08, NFN03, PRB^{+}08, Ros01, SFR^{+}09, SKH06, WMCF08]. forests [ATR^{+}08, BMO^{+}04, CRT^{+}08, CDT^{+}01, EE06, Kon00, KFU^{+}08, VHH^{+}02]. Foreword [Bel04]. Form [PAK^{+}00, SL05b, SL05c, BMG02, NATR^{+}05, Pra05, PK01, SR03, Sto03, WCCA09, dlSPLV09]. Formation [ASH03, CS05, Pas00, ST02, Tho04, VWH00, ACM^{+}02, BGC06, CW03a, CPKV00, CT06a, DQW05, EdPP^{+}00, FFG03, GYD04, GJG^{+}03, HWG06, MW02a, MSK^{+}00b, ONA04, RMH07, Tak02, TVG00, WM04]. formations [GLB^{+}03, JS08a]. formed [RMGL07]. formerly [WHS^{+}09]. forming [CK09b, CCB^{+}09b, DW09, FGS^{+}02, Kre01, LWR^{+}02, LTD^{+}02, LW00, MBVBC09, PCYC02, SG06b, SEM01]. Formosa [BRAS03]. formosus [KPEF00]. forms [FS07, GLS02, LGZ^{+}09, TGGA04, WGM02]. Formulation [BMT07]. fornicata [BCB^{+}06, BDG^{+}07, DBR^{+}07, DJV03, DEV07, LGD04, MTR^{+}07, PL07, RSN02, Thi05]. forsteri [MHB^{+}09, RBMB00]. Forsterygion [MS06]. fortii [KNT^{+}07]. foul [RASB03]. fouling [GHdN07, LGQ^{+}00, LC09a, LMW04, NS07, NP05]. found [SL05a]. foundation [CHP^{+}05]. founded [JS08a]. founder [ESB^{+}04]. four [BMG00, BBC04, BS04c, CLHT05, DVBLS07, KWD05, Kro05, LBvdH03, MPR^{+}06, MF08, ML07, NA04a, PCL^{+}09, PWLC06, PPS^{+}07, PJ06b, PB08b, RWSP03, RFH^{+}06, SICT03, SHM02, SAC^{+}03, SCAF00, SM01b, TG03, YI03a, YI06]. four-step [MF08]. fraction [CB00b, DFMD00]. fractionated [HKM00, MFP04, RGC^{+}06, Rol00]. fractionation [BGS^{+}02b, NWHC03, RBDK00, SMM^{+}04, WB01, YTH^{+}05, YIY08]. fractions [MLG^{+}03a, SSSV00, TDP^{+}00, WLL03]. fragile [SG06b, SS05]. Fragilidium [HSGS00, SHS00]. fragilis [DHT^{+}08, TCT^{+}00]. fragment [MS03, UYNO01]. fragmentation [GCHA05, JH06, RH06, RRRM08]. fragmented [CK09b, GC02, GCU07]. fragmenting [BSL02]. Fragments [HHV06, Goo09, JWTP06, SZ08b]. Fram [AP05b, MGF^{+}03]. framework [RBRH02, SBvD05]. frameworks [HF09]. France [BTC^{+}04, BM06, GPSH01, LABR^{+}06, LRR^{+}09, LNPB04, RH03, BLHB04, DT06, DAM^{+}08, DCRT07, HGRB04, HDCA06, HAD07, LNB^{+}03, MTL^{+}03, MVBGM04, Maz04, MBR^{+}05, PMG^{+}06, SAT^{+}03]. franciscana [SW04]. franciscanus [BJ08, NA04b]. Francisco [BP03, CD05, CB08b, GRBB07, GCHC00, HB04, HS07, Kim06, Les05, MFS07, NW05a, NLG05]. Fratercula [Bar02b, HANM^{+}05, SH00]. Frederick [Cai09]. Free [UÓ03, BEW03a, BEW03b, DRV^{+}05, EBA^{+}03, HS05, HMO^{+}04, KNKT01, KCH^{+}03, KCTG06, MFWD08, MSK00a, Mar02, MA02, RW06a, RCWG^{+}06, Tya09]. Free-living [UÓ03, DRV^{+}05, HS05, MA02]. free-ranging [BEW03a, BEW03b, KCH^{+}03, KCTG06, MFWD08, RW06a, RCWG^{+}06, Tya09]. free-spawning [MSK00a, Mar02]. free-swimming [HMO^{+}04, KNKT01]. French [AAPO00, AAGS02, LPBD00, MACS07, SLGD00, FPF01, FFT^{+}07, GhGD^{+}02, GPSG06, LYMCG06, MLGG03, MMGC06]. frenulate [SLP^{+}09]. frequencies [CHH^{+}05]. frequency [BHMR01, HBN01, HFA09, LSH^{+}07a, LCF02, MHK^{+}01, MAW06, MS04a, OÁSG^{+}09, PW07, RL07, RPC^{+}06, SAWG05, YA09]. frequent [BUCC09]. frequently [SWD^{+}02]. fresh [GGM^{+}00, KM03b]. freshets [LNV^{+}03]. freshly [CVÁM06]. Freshwater [WHD^{+}05, AKL^{+}04, CHW02, CAS00, JT02, JSIT06, KR04, Kim02, LF08, NMA00, RW06b, SJ08, SVG^{+}05, TA01b, Ven06]. friction [OIP04]. frigatebirds [WLJ^{+}04]. fringe [CWT06]. fringing [EB08, ERR06, KM01, McI03, TACM01]. fringing-reef [ERR06]. Fritillaria [FLUF^{+}04]. Front [SBA01, JCZ^{+}05, JK06, MTLT02, MSH^{+}03b, VLP^{+}04, PWB^{+}03]. Frontal [LCGN02, ADF^{+}09, AI04, AAJ07, FDD^{+}03, Mun07, ONKK03, Pee02, RMPC08, dBTO^{+}09]. fronts [BRG^{+}02, DMS^{+}05, DVBLS07, GN04, GGMGR07, LGS07, MDZ05, SP01b]. FRR [SMHG09]. Fuca [CTVP03, LJM03, UBM^{+}07, MTH04]. fucalean [BMC^{+}09]. Fucales [CK09b, SEH07b, Ste04b]. fucata [TKN00]. fucoid [ENAT07, LBPS03, LJ08, MCCV08]. Fucus [Arr05, BSJ02, BNP07, HDW02, HHJ04, ID02, KLFH04, PKS00, RBJB^{+}05, RK08, Ste04b, VP00]. Fuego [SR04a]. Fuel [FRvdMF02, SRDPT^{+}07]. Full [vDS05]. fullness [SLHR09]. fulmar [DFPG^{+}03]. fulmarine [HH00a]. fulmars [WISH09]. Fulmarus [WISH09]. Fulvia [RJH02]. fulvus [NHS^{+}08]. fumatus [TK01]. function [ADK^{+}06, BD02, BMW^{+}00, BWK^{+}08, EHK06, FLL09, FCS^{+}06, FSHB09, GD07, GLJ07, HF05, KMHS04, KP00, KCSPT09, MPDC03, PdNS06a, RJ03, RAA02, SSMM06, SL05b, SL05c, SRP^{+}06, SJPC02b, Sto03, TB08, TVD02, TJSG03, TLP^{+}03, VBS06, WM06, WCCA09, WR07, YH04]. Functional [ED02, FGS^{+}02, HF08, HVH^{+}07, HB04, JVT01, NOP02, RRS00, SAB^{+}05, dJTD07, Bat09, Bel07, BCJ05, BB08a, BRF03, BVP^{+}06, EJKN03, EGJ07, FVAB00, FEFT08, GLS02, GL04b, GDR^{+}04a, HDGM01, HOT08, JNH^{+}09, JFR02, KFT08, KT03, KSWP04, MMC08a, MCC^{+}00, NRH^{+}07, NK07, RM00, STRD02, TC03a, THJK06, TSF04, TSJH05, VSVV03, WBDR06, dlSPLV09]. functional-form [dlSPLV09]. functional-group [Bat09]. functional-structural [BVP^{+}06]. functioning [AC07, AGM06, BWRP06, CH06c, Duf06, HSM^{+}09, ISBP06, MHTJ07, NK07, Raf06, RGMH08, SB06]. functions [LS08a, NRH^{+}07, Saf03, SKFPS02, SJPC02a, THBSJ05]. Fundamental [RGCL^{+}07]. Fundulus [GV01, HFA09, Kne09, WBC08]. Fundy [DCM^{+}06, WAMP08, BCCM03, HK04a, JWR05, JTR05]. fundyense [AKRA02, CD03c]. fungal [DKAH02]. fungicide [MBHN07]. Funka [KYYM00]. furanones [WdNS00]. furca [SCS03]. furnieri [JBCL03]. further [CW01]. fuscata [EC07b]. fuscescens [PBAA04]. fuscoguttatus [PCMB06]. fuscus [LS04, SG05]. fusiformis [GBD^{+}06, PSRB00]. fusion [BUCC09, JY08]. fusion-rejection [JY08]. Future [FKSW09, Han07, Mar01, BDH01, CWB^{+}06, DGVK^{+}04, DB07b, DHT^{+}08, EP00b, GNP07, Law07, RZWG08, SRP^{+}06, SvAHH09b, SvAHH09a, SRR06].

G. [KJ04, LGB03b, LGB^{+}04]. GA [Ano02i]. gadoid [Joh02]. Gadus [AP00a, BM00, BM02b, BM02c, CSB03, CHH^{+}05, CDH^{+}09, CCL^{+}05, CCCH01, CPG^{+}09, FRvdMF02, GBR^{+}04, Hel00, HJG06, HMHB03a, HMHB03b, HNMC09, HCR01, JKES07, LGB03a, LGB03b, LGB^{+}04, LAK07, LG02, MHP^{+}00, MB02b, MSF^{+}09, MGSJ08, Mun07, NB06a, NM04, NWHH^{+}09, PTM08, PBCB03, RR06, RHSB08, SHNV09, SSL00b, SRJ07b, TKK05, UP01, WPIB09, YW04, vdMM03]. gag [HKZ04]. gain [ODA^{+}09]. Galapagos [SLL07, SLL09, VACT^{+}08, RW06c]. Galaxias [BS07b, HS08a]. galbana [OMT06]. Galeocerdo [MCP^{+}09]. Galeorhinus [LGME06]. galeus [LGME06]. Galicia [DLB04, LFRG^{+}05]. Galician [BAOV06, NAB^{+}06, OÁSG^{+}08, VPP04]. Gallager [Ano06]. gallery [FGS^{+}02, ODG^{+}04]. gallery-forming [FGS^{+}02]. galloprovincialis [ABB^{+}00, BOR08, BWH09, CLN00, GGHA02, JDVG08, LMF^{+}01, MP00, MOPP05, MMGC06, NZM07, RBG^{+}07, SSB02, SB03b, XBW07, ZMN07, ZMTB07]. Galveston [RZ00, RMZC07]. GAM [EOK^{+}08]. Gambia [PDM^{+}04]. Gamete [BCB02, MSC04, MYH07, Pep06]. gameteogenesis [MHW06]. gametes [MY04]. ’gametic [GDSB04, BGDS05, PBAD05]. Gametogenesis [BPT07, MY02]. gametogenic [GTPC04, MSF^{+}07]. ‘Gamma’ [GR07b]. gammarid [PBJ^{+}05]. gammarid-trematode [PBJ^{+}05]. Gammarus [CNT^{+}04, Pol00, RJ03, DL03a, SJCM01]. Gamo [AK02]. gannets [GDR^{+}04b, GPSG06, HPW^{+}00, HPH^{+}01, HHG^{+}07, KGF^{+}09, Mon07, MND^{+}09, PRvdL^{+}07, RCWG^{+}06, WHL^{+}08]. gap [KTH^{+}04, PTK^{+}09, SB05b]. gaps [CBEL06]. Garden [SC06a, PB00b]. gardneri [DZ06, HDW02, VP00]. garmani [SK04]. gas [FGPS04, HTJ00, HSPH05, Law08, NMA00, SRL^{+}02, SAB04, SPRK02, SGP07, TBK^{+}03]. gas-tight [HTJ00]. Gasterosteus [JHJ^{+}03, JHJS04]. Gastropod [Fer04, AE06, BSBD^{+}04, BT01, BCB^{+}06, Bar03, Bel08a, BK06a, BSH^{+}09, BBST07, CTP02, CU08, CW05, CNW^{+}07b, ED02, FMP05, KAMR00, KSH^{+}09, KHK^{+}04, MC01a, MVBGM04, MCCP09, PD05, RAAT^{+}07, RHG05, SRC00, UP07, WHB^{+}06, ZMP^{+}03, ZAZL08]. Gastropoda [DJV03, CR07a, CMRB08, Iye02, LCW06, RWSP03]. gastropods [BTL05, Bat07, CCR08, CKO^{+}07, Col03, CMGBJ05, DD07, KT06, LP07, MRJ05, MC06, PL07, PCC06, PM03, Prz05, RGEB00, Roy02, RBG05, SHZ^{+}05, SKW^{+}06, TDL05, Und04, VRJ02, WWV08]. Gastrosaccus [KPN03]. gaudichaudii [BCL^{+}01]. Gaussia [FSW02]. Gaussian [PLG09]. gazella [BKH^{+}09, BLDG01, GDL^{+}01, SRB04]. Gazi [DGVK^{+}04]. gear [GLP^{+}08, HCV03b, MAA^{+}07, SH05, SAM^{+}06]. gears [SMCE07]. geese [RS07a]. gel [CBD^{+}05, LMF^{+}01]. Gelatinous [BBG^{+}08, PGR^{+}01, GB05b, HB07b, OSF08]. Gelidiella [GTE^{+}09]. Gene [JWTP06, POH08, CSvO09, DEV07, FMCB05, HKH09, MB07b, NKS03, PSPZ09, PQRAC08, RCC^{+}09, TBVJ08, UB01, ZBNW09, vDvTJD^{+}09]. genera [LHK04]. General [TW10, BSS05, CSP^{+}03]. generalisation [Orv05]. Generalised [DRF08]. generalist [HBDD02]. generality [KCS04]. Generalized [SHZ^{+}05, JSMB05, SMP01]. generated [SV01, VGW00, VPSO01]. generating [MR01]. generation [WBT^{+}08]. generations [PLG09]. generative [EvBO^{+}08]. generator [KCH^{+}03]. genes [LK02, TBR^{+}03]. Genetic [AMS01, AAHDS06, ABD^{+}03b, AHDT^{+}09, CCML07, DGC^{+}05a, FSO08, KSH^{+}09, LPHH^{+}09, LLCHG01, Mal02, MDW^{+}09, MSTS^{+}06, NL05, PZB^{+}05, PDM^{+}03, PBTB08, PQRAC08, PMV06, RRDHG08, SC06a, SWM03, SMFP08, Sot03, SAG03, SBSRW02, Too04, TS06, WHB^{+}06, Whi06, BCM^{+}07, BRAS03, BER^{+}03, BUCC09, BGH06, BT04, CB05b, CH07d, CFA07, CEBR07, DFY00, DS06b, DJV03, EWR08, ESB^{+}04, GRPE08, GPK^{+}04, GWPRC^{+}07, HDW04, HCR01, HKR^{+}08, JKES07, KTT04, Kra07, MM07a, MDSP05, MC01b, NGM^{+}05, NW05a, NZM^{+}08b, OPSB07, PPI04, PWFC08, RGMJM07, RFCM03, SPI^{+}07, SDR01, TTL06, TLP^{+}03, UB02, VMVG^{+}07, VBA06, WTF^{+}05, WBC08, WG07, ZMTB07, vDvTJD^{+}09, vdVBW00, GGS09]. Genetically [MST03, BBT02, DBT02b, GTGM06, HFM^{+}04]. genetics [CDS^{+}04, DCF^{+}05, Gil02b, GDGSP09, HM06b, MSG07, NCMC09, Reu01, RMT^{+}05, WT09]. genets [HR02a]. geniculate [SHA^{+}05]. Gennady [ZK09]. genome [JWTP06]. Genomics [HP07, DWO^{+}07, HOT08, JB07b, TEB^{+}07]. Genomics-enabled [HP07]. Genotype [TLLKC04, SHH08]. Genotypic [Kne09, AM04, HWvdM03]. genotyping [GDSB04]. gentoo [LB05b]. genus [ALOT01, DBT^{+}02a, DH07b, KTT04, KWD05, MP05, MZD^{+}05, NOP02, PFP03, RM05, RGP^{+}03, SGV^{+}02, SH07d, TLIL^{+}09, UP07]. geochemical [WCZ^{+}09]. geochemistry [CMD07, CVT09]. geoenvironmental [SW08a]. Geographic [BGK09, BMG02, BNSG01, HTS06, KS09, LT07, PK00, TC03c, WHS^{+}09, WdNS00, AED^{+}07, BPC09, CHH^{+}05, DKAH02, GSSR02, GDL^{+}01, HSC^{+}09, JKES07, MR06a, RMT^{+}05, SBR09, SMP01, WHB^{+}06, ZM06, Zam03b, ZH02, Wil09]. Geographical [BIL01, CH07c, BRAS03, CPA^{+}03, DD05, DBOR09]. geography [ASR^{+}09, BGS02a, MMMPB05, RAC^{+}05, vWNYS05]. geolocation [GNL08, LALB08, PSC^{+}04, TBB^{+}04]. geology [KTF^{+}01]. geometry [PPV^{+}06]. Geomorphological [AHBM^{+}07]. geomorphology [VBS06]. George [ASP^{+}03, KKLL02, VSB04, FBJ00b, JBH00, JH01, Jen05]. Georges [JDCB09, ML06b, ML07, BCLS06, GYD04, GJB04, HCV03b, KLW^{+}09, LB05a, LAV04, LHB^{+}06, MJS07, MJS09, MBS03, MBS07, MGSJ08, SBL^{+}00, SH06c, TCS^{+}09, TT02]. Georgia [AWP^{+}01, ESDKM09, EBG^{+}00, GPW^{+}02, BC02, BGSBC02, CTB07, GSCW00, GKK03, KWAT08, MR01, PSC05, PWC^{+}09, RHDC05, SD09, SWW02, STA^{+}05, TCRT^{+}07, TGT^{+}06, WAW^{+}09]. Georgian [BCH00]. Geostatistical [WMS03b, HK04b, RT00a, Rue01, RMAY04]. geostrophic [VLP^{+}04]. geotaxis [MFY02]. Gerardia [PR09]. Gerlache [HLPHA^{+}01]. German [GSS09a, KSG03, SP01b, SAHK06, ZCR^{+}09]. germination [KH00b, Kre01, KSA03, McQ05, OHB^{+}00, OFHM03]. germling [GLJ07]. germlings [Ste04b]. gets [MK04b]. Geukensia [CWT06]. Ghana [Kor01]. ghost [BRP00, CS03b, PTA^{+}09, ASA^{+}08]. giant [FGSH^{+}05, GSCW00, GSSR02, HH05b, HHW^{+}07, KUOI06, MHP07, SFR^{+}09]. gibba [HM06b, TK01]. gibbesi [DFCH05]. gibbosus [WW03]. Gibraltar [dSCV^{+}08]. gigantea [KM03a]. giganteus [CGT^{+}00]. Gigartinales [RvdPHW04]. gigas [Bay02, BCJ05, BC06a, BDG^{+}07, BGD^{+}03, CBMB03, DBR^{+}07, DVLH^{+}00, GL04b, GES^{+}06, GMB^{+}06, HFM^{+}04, IMW^{+}02, LABR^{+}06, MSF^{+}07, MMGC06, NMCT03, RRS00, RSN02, TSW07, WR06, YIY08]. gilchristi [GB02b, TGGM05]. gill [APW^{+}04, BDDL03, BCJ05, CDK^{+}07, CR04, RJ03, ZLGC04]. gillnets [KCTG06]. gills [LTMDB02]. Gironde [DT06, LHD^{+}01]. GIS [EMG01]. Gitanopsis [YSS03]. glacial [CH07d, SO01b, WKPK05]. glacialis [DCM^{+}06, MNTS08, MTN03, NH05, SKFPS02, VTY06, WISH09]. Glacier [HE08]. gladius [WKY09]. gland [AVNN05a, AVNN05b]. glandula [GSK^{+}08, Kad03]. glass [EDB^{+}04, EDE05, FD08, KYML02, LWH^{+}04, STCJ01]. glassy [NWT00]. gliding [RCWG^{+}06]. Glimpse [SMMM07]. Global [CB05c, KCH^{+}06, REM^{+}06, Bar03, BTW08, BH04, CGH08, CWM^{+}07, CCL^{+}08, EWR08, GB09, HK02b, ISY07, KWP07, LHGC^{+}09, Pri02, Whi02]. global-bathymetric [ISY07]. globally [GCSSJV07]. globosa [SW08a, SV08, SLKF03]. glomerata [Bay02]. Glover [MSS^{+}03]. Glover’s [PCBS05, MBH^{+}00]. Glutathione [AWOO02]. Glycogen [SGL^{+}02b]. glycolytic [BG04c]. glycosides [KPEF00]. Gnathia [Gru03]. gnathiid [GLG00, PGK^{+}08]. go [AM04, CB08a, DB07b]. goatfish [KFS^{+}09]. gobies [CB08b, DH07b, HM05a, HM05b, UP01]. gobiid [HLFG^{+}07]. Gobiidae [MPR02]. Gobiodon [DH07b, HM04]. Gobiusculus [UP01]. goby [HM04, JRA02, MPR02, NW05a, Sha00, ST02, SJ04, WS08]. godwits [ZS03]. Going [BP06, PW07, PLG09, WWR^{+}01, MK04b]. gold [PR09]. goldband [NSEL00]. Golfo [GZY06]. Gompertz [HLR04]. Gonad [HCK^{+}05, BSAR03, BPT07, CDFRL03, CHW06, RRIHC02]. gonadal [CCCH01, KM03a]. gonads [KSIG08]. Gonatus [ZKHP07]. Goniastrea [BDW^{+}00, BDDG02, KS08]. goose [KT04]. gorbuscha [CHMR05]. Gorgonacea [Thr09]. Gorgonia [DKAH02, THYBS09]. gorgonian [BL00a, BSVK05, BSL02, BM07c, CCB^{+}09b, LCD^{+}05, OP02, PPBS02, RCR03, SHS^{+}05]. gorgonians [EM07b, MB07e, PPS00, WPRS05]. gouldi [JSWH03]. Goulet [BR07]. governs [BWP^{+}06, GUT04]. Gowing [Ano03b]. GPS [GDR^{+}04b, KJRG09, RDS^{+}09]. grab [NR00]. Gracilaria [TMM05]. Gracilariopsis [HB08b]. gracilis [EC07a, JTN09, MVH^{+}06, TT06]. gracillima [HSD03]. gradient [BA01, BJ08, BRS07, BES^{+}08, CML08, CLMR07, CLRM08, CBD^{+}05, CAA00, CV05, DFT^{+}07, DGNvdV05, DESK03, EMB^{+}09, FB07, GFP^{+}05, HM05c, HDCA06, Joh04, LPHH^{+}09, MCCV08, MCS05, MFE^{+}05, MLW03, MLG03b, PMM^{+}00, RMWO09, RGCL^{+}07, RSE01, SRGS01, SS03, SCS^{+}07, TJS08, UBRvO06, Uth06, Vie09, ZSR03, vWCAB07]. gradient-single [Joh04]. Gradients [SL07, ZHS^{+}05, BFK^{+}05b, BRBPL06, CCR08, EYH^{+}06, FS07, HWC^{+}06, HS08b, Hil04, KTK08b, LBH^{+}04, LTFJ00, MND^{+}02, PCD04, RWJ02, SH07b, SLM07, SOMM06, TA07a, Win09, YMHV02]. Grahl [GNAD^{+}04, TBBI04]. Grahl-Nielsen [TBBI04]. grained [KW05, WRH05]. Granata [Ano01d]. Grande [RATD07]. grandis [WBC08]. grani [HSCH07]. granulata [CGCDA02, GA03]. granulatus [BIG^{+}06]. granulometric [HDCA06]. granulosa [ATLC03, BP04, CTA^{+}03]. Grapsid [WL05, CGCDA02]. Grapsidae [BM05b, DFCH05]. Grapsoidea [CACDC07]. grass [KD00, KS00b, MRP09, PABM08, RGR^{+}01]. grasses [ML06a]. grata [WDLM05]. gratilla [VPS08]. gravid [FFT^{+}07]. gravity [NRJB04]. gray [BWW^{+}06, DD01, LSS^{+}09, MW06, AL08, KHGA02, SAV^{+}09a, WFHS08]. Grazer [RK08, SEH07b, APP^{+}09, LV05, WS00]. Grazer-controlled [SEH07b]. grazers [DB02, FH08d, HBRW04, PTA^{+}09, RC05, RC07, SS07]. Grazing [AK02, CU06, DUW08, FH02, HRQ^{+}02, KPN03, MNB^{+}07, ONP02, PMNL07, PD05, RJBW03, SLC05, SSSV00, SBBS^{+}02, URFA06, AH07, AED^{+}07, AM02a, AAF^{+}04, BLS06, BL00b, BS07c, CLS00, CVF^{+}03, CLL^{+}09, DJM07, DGM00, DH01, FTO^{+}09, GSFN00, GRLML09, HVP^{+}06, HLGCO06, HWL00, JMU09, JHL00, JAA^{+}01, JCD^{+}05, JKK^{+}05, KSSH02, Kiø03, KVH02, KVS08b, LCL^{+}00, LGS07, LSdM^{+}03, LÅCN01, LD03, LUIAH03, MR05, MRN07, MB07a, MB05b, NNS01, NGMT01, PFW^{+}02, RSR02, RS07a, RFG^{+}09, SLGD00, SWL07, SHSV04, SKK^{+}06, SKU^{+}08, SWEBS00, SG02b, SS05, TUK09, TGY^{+}06, TIM^{+}01, TLNH01, UM08, Uth01, VSL^{+}04, VHKW00, VJ05, VH03, VPS08, WTP08, WH02, WW06a, WWV08, WPH01, WDR^{+}05, XW03, ZSKA01, CMHP06, MR04]. greasy [KAR04]. Great [Ber02a, Boo02, Dun02, MB02a, BPT^{+}00, FS08, KCV^{+}06, MR06b, WLJ^{+}04, BM02a, Ber02b, BBA06b, BB08a, CWMD06, CDS07, CWE^{+}08, CMD^{+}09, EC07b, FMZvW02, Gru03, JMHS03, KHP09, KRP^{+}06, Kri02, LRM00, NMDS00, PKM04, PCMB06, SRDG05, UBRvO06, Uth06, WNWH02, WHB09]. Great-Entry [CWMD06]. greatly [PD05]. ’Green [PAK^{+}00, Cos03, HMA04, ABL08, APW^{+}04, ARB04, BR03, BRBPL06, BBC03, BGRD00, BHN09, BBC08, CD07a, CYC^{+}09, CNPS05, CvBGR09, CRZ04, DEAI06, DHR04b, GSL^{+}06, GXL^{+}08, GSB^{+}06, GBF^{+}02, GBGH02, GLÅ^{+}03, HHB^{+}00, JMA02, JHL04, KF01, KCEB09, MH08c, MB05b, MS03, NW05b, NDB04, PW04, PCC01, SG06b, SAC^{+}00, SRN02, SJE^{+}06, SJM06, SRS05, TTD^{+}06, TBBB06, Wan02a, WC03, YA09, vLvSGS05]. green-lipped [CYC^{+}09, GSL^{+}06, GXL^{+}08]. greenhouse [HMA04]. Greenland [ABD^{+}03b, GRPT^{+}00, GKWR02, JPNM^{+}06, LHJN07, LNH00, MRG08, RKGH01, SSJ^{+}02, SBR09, TNHL02, TNT08, TLNH01, WHD^{+}05]. greens [SAC^{+}00]. greenshell [SAG03]. gregarious [Don04, TP01a, TP01b, TP01a]. grenadier [BD07]. grey [BMM^{+}04, CPP^{+}04, CIT04, DPD^{+}03a, DPD^{+}03b, NLP^{+}01, PHH^{+}05, TBI07, TBI08, WHP00, WP03]. grey-headed [CPP^{+}04, NLP^{+}01]. grid [BOZ03]. griffithii [LV02, LMMT09, MCL07]. griseus [LSS^{+}09, MW06, SWS^{+}09]. groenlandica [CS02a, WHL00]. groenlandicus [CDH^{+}09]. gross [HHIS06]. Ground [EO04, BLDG01, GRWB02, HFW^{+}05, KD00, MPSJ06, dJTD07]. groundfish [CGC^{+}09, GP08, ML06b, ML07]. grounding [GP03]. grounds [AWBO09, BGDS05, DvD02, GDSB04, HK04a, KMHS04, MC01a, NvdV04a, NGBG07, NMBH02, OBG^{+}07, PA06, SSS06, WWWR09]. groundtruthing [WHCP^{+}04]. Groundwater [KHT^{+}02, CTLR03, DM08, GSW01b, GB03]. group [AP00a, Bat09, DDS^{+}06, KHG06, MWR07, SR08b, TdLvdVR08, TSJH05, vdVBW00]. group-living [SR08b]. group-size [KHG06]. grouper [AP06, APW^{+}04, DE01, HDP^{+}06, KAR04, MRGC09, PCMB06, SSBZ07]. Groupers [CSS00, Tup07]. groups [ACJ^{+}04, BBT02, HLH^{+}07, HSS^{+}04, KKL07, Kri02, Lid08, NSSI06, RGC^{+}06, SMHL00, TSF04]. grow [KFJ^{+}07]. growing [Cam01, MVH^{+}06]. Growth [ASP^{+}03, ATR^{+}08, AMG^{+}07, ASLD09, BRS07, CWT^{+}01, CS02a, DS01, DL03a, FPC05, GSM06, GB05b, GPM08, HLR07, HL05, HLR04, ITN09, JHL00, JSH01, KCEB09, KdLCM09, LHQ05a, LMW^{+}01, MD01, MGSJ08, MHHS06, NM04, PR09, Phi05, PEGG^{+}08, Pol00, RRIH^{+}02b, SRDPT^{+}07, SML02a, SSJ^{+}02, ST06b, SH07d, SO01a, SB03b, SML02b, SAP02, TW04b, TOK^{+}04, TSIY06, THC^{+}05, UBM^{+}07, WM02b, Wil04, vdMM03, AS04, ADA00, ATT06a, AW03b, AM01, AM02b, AHZ03, AGS^{+}09, APP^{+}09, APW^{+}04, BM02a, BA01, BWL06, Bar02b, BGJ^{+}07, BGK^{+}06, Bay02, BDG^{+}04, BJH04, BVH02, BBM00, Bla09, BSAR03, BH09a, BF08b, BCPB06, BM07a, BWK^{+}08, BMS04, BBT02, BY09, BHV^{+}02, BVN^{+}03, BVP^{+}06, BM02d, BCLS06, BCC08, BH09b, CSB03, Cal05, CGP08, CHO09, COPB06, CU06, CÅJ05, COK02, CR07a]. growth [CT06a, CLL^{+}09, CMJ00, CNW^{+}07b, CDAM06, CK09a, CSMvE04, CB08b, CLMR07, CRCN07, DJM08, DTV08, DTV09, DvD02, DC07a, DFR08, DH06, EKN08, EDE05, EBC04, Edm07, EZTZ02, EB04a, EÖM08, Eyr00, FMWGB08, FJK^{+}07, FR06, FFG03, FB03b, FON03, GM01, GVM^{+}05, GC06a, GSW01a, Gor03, GQ00, GJB04, GM05, GLCL04, GDKN08, GBW02, GRLML09, Hal00, HWB^{+}09, HN03, HRCB02, Hel00, HSP^{+}08, HHW^{+}07, HM01, HPR05, HL00, HGK02, HGKK06, HSSS05, HMHB03b, JM01, JSWH03, JSIT04, JSIT07, JF07, JFA02, KS08, KLC^{+}06, Kav05, KCK^{+}06, KM03a, KWTR08, KHHM07, Kir02a, Kir02b, Kir02c, KHWH03, Kne09, KP01b, KAM^{+}06, KMAC^{+}02, Lam00a, LCL^{+}00, LC06, Lap01a, Lap01b, LGBWH06, LGQ^{+}00, LQRGT06, LD00, LG06, LSP08, LT00, LWR^{+}02, LN05, LD03, LLRDT05, LUAIH03, LHB^{+}06, LDR00, LHK04, MCR03]. growth [MPM^{+}02, Mar04, MHSS08, MFA03, ML00, MCM^{+}03, MW02a, MGTK00, MW04b, MHY^{+}06, MOPP05, MAWH03, MGZ^{+}03, MTS07, Mok04, MR07a, MR07b, MR07c, MBG^{+}08, Mun07, NWJ^{+}08, NRM05, NWHC03, NRC07, NHNB01, ÓUA05, OA02, Pah05, PCYC02, PBVP09, PHM^{+}08, PBCB03, PD07, PH03b, Pep06, PBPLB06, PDHH06, Pet02a, Pet02b, PCW05, PA07, PPSP07, QA01, RK08, Ran02, RM05, RPN^{+}01, RGR^{+}01, RR05, RHI^{+}01, RVMI^{+}03, RBDK00, RMKR00, RH09, Rij03, Rit02, RGDI06, RCF07, RCF09, RvdPHW04, RM09b, RH08, SM01a, SBMAE06, SSN^{+}09, SBvD05, SH01, SZ08a, SZ08b, SLCB00, SS00b, SEH07a, SBR09, SM00a, SJ05, SSRG05, STCJ01, SH04, SF02a, SWW02, SD00a, SD00b, SML05, SO02, SPC^{+}00, SLHR09, STJ03, SF02b, SCM09b, SC00b, SFR^{+}09, SOMM06, SSH09, TW04a, TAM03, TAM04, TAO07]. growth [TK01, TN04, TAA04, TdLvdVR08, TFA^{+}08, TGGA04, TFSJ00, TAS04, Thi05, Thi06, TPGQ07, Thr09, THYBS09, TCG03, TC03c, TBAT02, Tru02, TLPG06, TL03, TYD07, Urb02, ULL^{+}05, UB02, VMCG09, VP00, VBB06, VJKF07, WCBD01, WGD05, WM04, WW07, WKH07, WWL04, WKOE09, WCM08, WPIB09, WPPS04, WLJ08, WG02b, WG02c, WO05, XBW07, YQW^{+}01, ZS04, ZSKA01, dlSPLV09, jMFVPL05, vdVBW00, HM05a, KCK^{+}06]. Growth-dependent [TSIY06, SD00a]. Growth-history [KdLCM09]. growth-increment [Bla09]. growth-rate-limiting [RBDK00]. Growth-related [WM02b, TAM04]. Growth-selective [ST06b, TOK^{+}04, TAM03, TAO07]. grunts [FS08, NBN^{+}08]. Gryllotalpidae [MLGG03]. grypus [BMM^{+}04, DPD^{+}03a, DPD^{+}03b, PHH^{+}05, TBI07, WHP00, WP03]. GTT [WGM02]. Guadalupe [DNL08a]. Guadiana [RGB02]. Guam [KP01b, PPBS02]. guard [TGT^{+}06]. guarding [TSRR05]. Guiana [FFT^{+}07, MLGG03]. guild [BJC02, BH06, Dav00a, FEFT08, GJ05, GL00a]. guilds [AL09, MCC^{+}00]. guillemot [Bar02b, ÖBFO01, RMT^{+}05]. guillemots [GSM^{+}02]. Guinea [CWCP^{+}06]. Gulf [GES^{+}06, JRP01, LHK01, MGH00, OA02, SAB04, SLB^{+}08, TS06, CL05, SSS06, AQHTL02, ARAMPW02, AL03b, AWA^{+}05, BP02a, BTKG^{+}08, BUF02, BA04, BY09, BHY09, BRB^{+}00, CSMBH08, CDH^{+}09, CRTP02, CC05, DJG^{+}07, DR00, DRB^{+}01, DH07b, DVBLS07, DABC02, DCC^{+}03, FHVMDS07, FAPA^{+}09, FAPOF09, GMd^{+}02, GS02, HCGC09, HCM01, HTM01, HST^{+}09, HCBP05, HRC^{+}07, JDCB09, JWTP06, JPG05, JH06, KLRS00, KHGA02, KK02, LB05a, LAK07, LKW03, MCFM02, MFCM05, MD08a, MB05a, MRZJ05, ML06b, ML07, MA07, MIH^{+}00, MILH00, NEK^{+}07, OPSM04, OSL07, PPB^{+}01, PAK^{+}00, PAW00a, PRM^{+}09b, PRM^{+}09a, PT04, PDG^{+}02, PCB08, PR05, QS08, RBHS06, Ran02, RBRH02, RCF09, RTH06, RRRM08, SRN02, SBBS^{+}02, SLP^{+}09, SOMM06, UB03, WPC^{+}03, WST^{+}09, WCF08, WBC08, Wil09, WJB09]. gull [SG05]. Gullmarsfjord [QMM02]. Gullmarsfjorden [HLVU00]. gulls [HW00, KKK^{+}05, RRS^{+}09]. Gully [HKH04]. gulps [GPS07]. gunnari [RHDC05]. Gut [LUA99, BLF06, CSTC01, DJM^{+}05, DC07b, FC08, KSA03, Pen00, SMB^{+}03, SLHR09, STJ03, dlMPN^{+}03]. gut-content [dlMPN^{+}03]. guts [JWG01, LNJ07, SMMM07]. Gutt [Ano02g]. Gymnodinium [HGAS^{+}01, LJK01a, LJK01b, LJK02, SLT^{+}01]. Gymnothorax [LIT05]. Gyre [LSK08, EC07a, Fen08, TPQ^{+}03]. Gyrodinium [KJ04].

H [COB^{+}03]. Håkon [VMdBV06]. Habitat [AA04, BBM03, BP09, BL09, CS09, CKB^{+}09, CH00b, DC06, DvD02, DH06, ECM^{+}05, FDDS09, FSW09, FP08, FMT09, GJ05, GRRH07, GP04b, Hal04b, Hal00, HJW08, LPS08, MD08a, MGE^{+}09, MR03, OvM02, PCBS05, PBF08, SAG08, SFH00b, SJ04, TNTH04, TSHT02, THN^{+}03, AR02a, AFHS08, AFHS09, ARGS01, AP04a, AP04b, AMG^{+}07, AO01, AKL^{+}04, AGDP^{+}06, ABL08, AFHA05, BPBH09, BT09, BC02, BDC00a, BTL05, BCCM03, Bay02, BB03b, BBE03, BRP07, BDLB02, BER^{+}03, BGS02a, BB00, BFLÁ01, BM07c, BDF09, BV08, BLP^{+}02, CH09, CH06c, CPBS07, CUB08, CB01, CNF09, CK09b, CTvdM^{+}08, Con05, Cos09, CC05, CTH^{+}03, CV05, DLT^{+}06, Dav00a, DMA03, DMK^{+}08, DG05, DBT^{+}02a, DCF^{+}05, DSMBG03, DVBLS07, DGNvdV05, DVNvdV07, DDSR07, DGCM09, DH09b, EJM01, EMG01, EDE05, EM07a, EÖM08]. habitat [EM07b, FS06, FHG^{+}03, FMZvW02, FVB^{+}09, FWC02, FKI^{+}09, FBM07, FB02a, GA08a, GBG00a, GPHC09, GD07, Gla07, GC02, GC05, GLP^{+}08, GS05, GT08, GW06, GDW06, GBI^{+}06, GF02, HT07, HTM^{+}02, HAC06a, HSS09, HMBD01, HHH02, HJK01, HWBZ05, HFWW05, HGG^{+}05, HMA06, Hor06, HMJ^{+}09, HDRC^{+}08, HF05, HS01, IBB^{+}07, IR02, JWM09, JSIT02, JSIT07, JH06, JFD09, JHH01, KLN06, KvKB06, KBH01, KHA^{+}00, KSSB04, KTF^{+}01, KKL07, KCSPT09, KCA^{+}08, Kup03, LGB03a, LGB03b, LGB^{+}04, LSH07b, LGD04, Lec05, LAW06, LP03, LRGdS00, LTCP03, LÁMBF09, MHP^{+}00, MR05, MBVBC09, MPM^{+}02, MHM^{+}08, MGTK00, MB07e, ML06b, ML07, MH06a, MHT07, MZD^{+}05, MSP^{+}06, NBN^{+}08, NRM05, NSBK05, OLG^{+}09, OT08, OJ02, PS08, PGCC^{+}06, PCD04, PRM^{+}09a, PGP03, PCW05, PPC^{+}05, PPSP07, PZMG05]. habitat [PCB08, RFB^{+}06, RBB^{+}08, RH06, RWE04, RBG04, RDH^{+}09, RT00b, RC00, RRRM08, RMZC07, RM09b, RGC05, RST04, RSSA07, SHNV09, SDW^{+}01, SH07a, SKBP09, SWS^{+}09, SCIT09, SICT03, SJCM01, SV01, SLP^{+}09, SC00b, SMP01, Sto03, SML02b, SHSF08, TLFS06, THC^{+}05, TSSQ06, THF^{+}01a, THF^{+}01b, Tor09, TOBR07, ULL03, UÓ03, Und04, VHB^{+}07, VFJ^{+}07, VBF06, WPM09, WBC07, WOL^{+}07a, WA03, YAF^{+}07, vWCAB07]. Habitat- [CKB^{+}09]. Habitat-based [BL09, DC06]. Habitat-dependent [AA04, FMZvW02]. habitat-forming [CK09b, MBVBC09]. habitat-limited [BFLÁ01]. habitat-related [VFJ^{+}07]. habitat-survival [HF05]. Habitat/sex [SFH00b]. habitats [AE02, ABC^{+}05, ASP^{+}08, BS05a, BBS05, Bib07, BBM02b, CWKE05, CCK^{+}09, CSM02, CHG05, CHT09, CVT09, CB05c, DKA^{+}06, DGdG^{+}09, DHR06, EE00, FC00, FBBV09, FRJ09, FLL09, FS02, GPLAG04, GK00a, GAB^{+}03, GCU07, Goo09, GLG^{+}05, GRG^{+}02, GBW02, GCL04, GC06b, HKZ04, HHD^{+}09, HC02, HFA09, HFA02, ISSK03, IC02, JH01, JBF08, JTN09, KTH^{+}04, KEB07, KvdMDP09, KCEB09, LB08a, LDK^{+}06, LJ06, Lei01, LEB07, LAV04, LSS^{+}09, MGD^{+}09, MH07, MA00, MFR01, MHR03, Mok04, NKK^{+}01, NvdV02, NvdV04b, NLG05, NR00, OC01, OKM06, PMS^{+}06, PDS^{+}07, PZ07, PFZS^{+}06, PTCH01, PBL^{+}09, RGA^{+}02, RFLC07, RT00c, RAH^{+}02, RC05, SH06a, She01, SBJ06, SW08b, SB00, SDM06, TNTH04, TRT04, TA01b, Tup07, USG^{+}08, UBM^{+}07, VNdG^{+}06, WBJ^{+}07, YBGP06]. habits [FC04, KNTK05, LME05, LGME06, MKWO00, NOP02, RW06a, SK04, SK05, WKH02, WKIK04]. habituation [BSW^{+}09]. hadal [FKM^{+}01]. hadalis [FKM^{+}01]. haddock [BW01, Bri03, BCLS06, Cal05, GJB04, LB05a, LHB^{+}06, MBS03, MBS07, MGSJ08, PBCB03, PKLD09, RTL05]. haemastoma [RBG01]. haemocyanin [SS02b]. haemolymph [DL03a]. Haemolytic [ULL^{+}05]. Haemulidae [NDV^{+}00a, NBN^{+}08]. hairs [FSW02]. hake [FHVMDS07, GL00b, GZY06, GBW02, HGSH08, LP04, SC00b]. halibut [FS02, HSSS05, LS06, Mad02]. Halichoerus [BMM^{+}04, DPD^{+}03a, DPD^{+}03b, PHH^{+}05, TBI07, WHP00, WP03]. Halichondria [TL08]. Haliclona [CÁ04]. Halimeda [NDB04]. haline [GGMGR07]. Haliotis [BT01, BT00, CMS^{+}01, ESB^{+}04, MBBP08, OKH^{+}07, PBFB08, PS06, SM08b]. Halipteris [MS09]. Halobaena [CBTW02]. Halocladius [GJT09]. halocline [MMS09]. Halodule [BH09a, BD01, HHV06, HVP^{+}06, KDAK01, TS06]. halogenated [EWSH08, WdNS00]. Halophila [HHV06, KWTR08, VHM09]. halotolerant [EKN08]. hamifera [NCHP05, NCP^{+}08]. hammerhead [DH06]. handling [AP00a, BH01]. Hapalogaster [SG06a]. haploid [HPVL05]. haploids [TG03]. haplotype [VGB^{+}05]. haplotypes [LW04]. Haptophyta [CLS^{+}01, ZER^{+}01, ZJW^{+}04]. haptophyte [GJ03, SKK^{+}06, ULL^{+}05]. harbor [CR04, CKTE01, FBD06, HBBD09, KCTG06, ZS04]. harbors [LL03]. Harbour [BML^{+}08, WG01, CHT06, DSF^{+}04, GSS09a, GGGVGG03, JWR05, MGE^{+}09, NvdV02, NWIT08, RFLK03, SAH09, ST08, TH02, TR06, VMVG^{+}07]. harbours [dCLLT^{+}02]. hard [BKSS07, BM07b, CGP08, CSM02, DF02, DH09b, FBT^{+}01, Goc04, NTGK09, TYT^{+}04, VHT^{+}05]. hard-bottom [CSM02, DH09b, VHT^{+}05]. hard-bottoms [DF02]. hardbottom [CT09]. hare [GP01, RdNS02]. harengus [VMCG09, Bar07, CMB^{+}09, CKTE01, Enb05, EH04, FFG03, JVT01, MST03, TFSJ00, VFV01]. Harlequin [EBT^{+}02]. Harmful [KB04, AGB03, BGJ^{+}07, BGC06, CTFC08, GMC^{+}09, GVM^{+}05, HSW^{+}08, JKK^{+}05, Kam00, KGRW06, KB05, NSFW09]. harp [CS02a, CDH^{+}09, GNAD^{+}03, TBBI04, TBI^{+}09, WHL00]. Harpacticoid [Thi03, CLABRO^{+}03, DHCH04, DFV01, TS04]. Harpacticoida [Joh01]. Harrison [BG09]. harvest [NM09, PPGL09, TS07]. harvested [SRC00]. harvesting [ABC^{+}05, ASS05, CAS05, Enb05, NSBK05, RLBB04]. Hatch [LB05a, Lap01b]. hatch-date [Lap01b]. hatched [KB06]. hatchery [WPPS04]. hatchery-reared [WPPS04]. hatching [AMR00b, BF08b, GR07a, ITE^{+}04, JVJ09, MRZJ05, MMC^{+}07, NMS^{+}02, RM05, RBRSP00, TD01a, TIM^{+}01, WIDM09]. hatchling [BGRD00, GBM01, TBBB06]. hatchlings [SPM03]. haul [GAMB05]. haul-out [GAMB05]. Hauraki [SLB^{+}08]. having [BBG^{+}04]. hawaiensis [PCT00]. Hawai’i [SRS05, CPB04a, CKB^{+}09, FBM07, MCP^{+}09, Par06, WWMF06, Zab09]. Hawaii-based [CPB04a]. Hawaiian [DF04, BPH07, BBABL01, BBZM08, CH00b, DP01, DFH05, FD02, GA08b, LBBM06, MBBW08, MHP07, PWLC06, PR09, RGDI06, ZA07]. hawksbill [DvD02, LB02, SWHSA05]. Hays [CPB04b]. HCl [SZL^{+}05]. HCl-soluble [SZL^{+}05]. Head [WWMF06]. headed [CPP^{+}04, NLP^{+}01, WH09]. headland [MM06a, MM06b]. health [KPA^{+}09, NCC^{+}07]. Heard [PJN01]. Hearing [WW05]. Heart [FD08, LC03, RCWG^{+}06, VVTH07]. heat [BFB01, BCKP09, HMH04, RCC^{+}09, RPS^{+}04, WDLM05]. heat-inducible [RCC^{+}09]. heavy [DL03b, GdLB^{+}09, LMW^{+}01, MBH^{+}09]. heavy-metal [DL03b]. hedging [Kru01]. Hediste [KTK08b, VBA06]. height [AL03b, DD05, FS07, KDAK01, MHTJ07, ML05, NW06]. Heightened [SW08b]. Helgoland [KSG03]. helgolandicus [HB07a, HBH07, IP07, KP00, MGZ^{+}03, PWL^{+}06, RRIHC02, RRIH^{+}02a, RRIH^{+}02b, RHI^{+}01, RJBW03, SAHK06]. Helicoradomenia [KCB06a]. Heliocidaris [SC07]. heliopora [TLLKC04]. help [WBJ^{+}07]. Hemicentrotus [ENAT07]. Hemigrapsus [BM05b, BG07b, JMA02, MCE01a, O’C07]. Hemigymnus [GL02]. Hemisphere [BCP^{+}08]. hemocyanin [DGB^{+}02]. Hemocyte [SJPC02a, SJPC02b]. hemprichii [GNC07, LVRN02]. hence [Hel00]. heptanone [NCP^{+}08]. herbicide [CNW07a]. herbicides [JMHS03, JK03b]. Herbivore [KF02, AL03a, BH09b, CDO08, Kru01, PCL^{+}09, PdNS06b, RFN02, SWL07, TJS08, WHB09]. Herbivore-mediated [KF02]. herbivores [BMdNP09, LW00, MKKK08, PI07b, TTR05a, VBAR07, WdNS00]. herbivorous [AED^{+}07, CCC05, CRH01, FSHB09, HK02a, JSMM06, MB07a, MACS07, Sot03, WPH01]. Herbivory [SKR09, BP02b, BB04, FB07, Har06a, HHJ04, LT07, MSS^{+}03, MKKK08, PTAR07, PFT^{+}08, SDS^{+}03, TTR05b, VLW09, VAB09, WDN00]. herbstii [RE00a]. Heritability [MHW07]. Heritage [CMD^{+}09]. hermaphroditic [LBPS03, RZLS09]. hermatypic [BSAR03, HNYI07, MBLA01, dBKSV00]. hermit [BD00, Bar02a, Bar03, Hal04b, RB04, TL03, TDL05, WP00a]. Heron [CMS^{+}01]. herring [ALE^{+}09, ANF^{+}00, Bar07, BCM^{+}07, CMB^{+}09, CKTE01, DCBvBE09, Enb05, EH04, FFG03, JVT01, KKK^{+}05, KNF^{+}03, LHHB05, MRS00, MHH^{+}05, MST03, TFSJ00, VMCG09, WDT^{+}08a, Zam03a]. Hesiolyra [SJS^{+}01]. Heterocapsa [AGB03, LWR^{+}02, LTD^{+}02]. heteroclitus [GV01, HFA09, Kne09]. heterocystous [WVL01]. heterogeneity [BWW^{+}06, Con05, DJM^{+}05, FGA^{+}04, GBR01, HTL^{+}02, HVW^{+}07, IC02, KPT01, LZM^{+}03, MHP^{+}00, MD08a, MR05, MdMR04, MA07, PR01, RBGM06, SV01, WDN00]. heterogeneous [KKW^{+}03, LTPL05, UHvO^{+}08, WJ06]. Heteromastus [MLR02]. heterorhabdic [BDR04]. Heterosigma [BGC06, JKK^{+}05, YNM^{+}07]. Heterospecific [GK08]. Heterotrophic [SBL^{+}00, SS07, AHW03, AJ03, CNC00, CPH^{+}02, CLP08, CAA00, INH^{+}09, JKS^{+}01, JKK^{+}05, JSK^{+}07, JKS^{+}07, KJ04, KW00b, MN00a, ONP02, PC02a, PH03b, SLGD00, SJK^{+}06, SPBM05, SWL^{+}06, Tan02, TJ02, VTDS09, WLJ08, YYL^{+}00]. heterotrophy [VHV^{+}04]. Heteroxenia [PBAA04]. heterozygote [BHFC08, BSDR01]. Hexactinellida [LWH^{+}04]. hexadecanoic [GWBdN09]. Hexaplex [PCB08]. Hg [PW04]. Hiatella [SSJ^{+}02]. hibernation [SSR^{+}03]. hidden [KvdMDP09, RDD^{+}02, vDS05]. hierarchical [PMP04, WEEBdN08]. hierarchies [BK03, BKB05]. hierarchy [EEWBdN08, VL00, WWWY07]. High [AJ03, BS03, BK03, BSR09, BL01, CEBR07, DJV03, Kiø03, MBG^{+}03, MH07, OÁSG^{+}09, PBAD05, RBSP00, SM08a, UDB^{+}09, USG^{+}08, vDvTJD^{+}09, ABF07, AQSP07, AMB08, ATLC03, Ant06, AHK^{+}02, BS06, BWL06, BRS07, BKK^{+}05, Bra08, BHMR01, BLH02, CHW02, CSD04, CC00, CM08, DHCH04, Dis07, FA06, FSO08, FKSW09, Fis06, FP08, FBHB03, GRPT^{+}00, GKWR02, GPS07, GTSS03, HSTS^{+}09, HBG^{+}03, HBN01, IHK08, JSMM07, JBF08, LCF02, MD08a, MAC^{+}07, MP05, NPN06, NTN08, PPC09, PAK^{+}00, PH03a, PH03b, Ran02, RGEB00, RKGH01, SSJ^{+}02, SWM03, SH07d, SP01c, SDW07b, SMF00, SZW^{+}07, TP07, TSSA03, TTN01, TYC00, VMH^{+}08, WH03, WS08, YBY^{+}09, MSJ^{+}09]. high-Arctic [BRS07, GRPT^{+}00, RKGH01, SSJ^{+}02, MSJ^{+}09]. high-CO [IHK08]. high-density [BKK^{+}05]. high-flow [JBF08]. High-frequency [OÁSG^{+}09, HBN01, LCF02]. high-latitude [Fis06, FP08, MAC^{+}07, NPN06, NTN08]. high-molecular-weight [GTSS03]. high-nitrate [SMF00]. high-nutrient [FBHB03]. high-order [Bra08]. high-seas [Ran02]. high-throughput [Dis07]. high-turbidity [Ant06]. higher [CMRF05, GH08, JRS^{+}06, NvdV02, Sta08]. higher-order [GH08]. highlighted [HFM^{+}04]. highly [AFHS09, BB03b, BK06b, BdVD^{+}01, BGRD00, CVÁM06, FBBH07, GDGSP09, KTH^{+}04, MCW^{+}07, MS08, PTSJ06, SDK^{+}08]. hilsa [MC01b]. Himasthla [DdMB04]. hind [Nem05]. hinumensis [AY04a, AY04b]. Hippoglossoides [RZS05, SM01b]. Hippoglossus [HSSS05, LS06]. Hippolyte [Zup00]. hirundo [NMBH02]. hispida [GNHB^{+}05, KKL07, WHL00]. hispidus [CSO08]. histochemical [LTMDB02]. Histological [PKH00]. Historic [DCA^{+}09, NEK^{+}07]. Historical [GC05, TSF04, BKP07, CLB^{+}09, CWN^{+}07, CAA07, KBA^{+}05, RJS04, SZL^{+}05, TR06, Whi02, WBC08]. histories [Air00, HWB^{+}09, Kap09, MAOP09, NH03a, NH03b, RLFEF03, SS09c]. History [PHO^{+}07, AY04a, AY04b, ATT06a, ABL08, BDRM06, BSP03, CAF^{+}08, CIT04, CKB^{+}09, CEO06, DC06, DL02, DC02, DG05, DH07a, FOH^{+}08, GM04, GNPB07, GKAS08, GV01, GLBF07, HM05a, HM05b, IMSO09, JRP01, JSIT04, JDCB09, KWWH02, KHK^{+}06, KOY^{+}05, KdLCM09, MAM^{+}01, MGW^{+}09, MMA^{+}09, MVH^{+}06, OHCFJ04, OTB^{+}00, PDM^{+}04, RBW03, RF08a, RCL^{+}09, RCC01, RTL05, SSN04, Sha00, SP03b, SR04b, SGCP06, TKK05, TD01b, TCW^{+}07, WA00, WTF^{+}05, WTS09, WDRH07, Wol07b, ZKHP07]. histrio [HM04]. HNLC [FBHB03]. hoc [SNP03]. Hokkaido [BWH09, OKAM00, WDT^{+}08a]. holdfasts [HAC06a]. holding [BKB05]. hollandi [KSHO01]. Holling [KFT08]. Holm [INH^{+}09]. holobenthic [DPSD08a]. Holocene [EBA^{+}03, AHC^{+}06]. Holocentridae [CEBR07]. Holopneustes [SdNH^{+}06]. Holothuria [UB01, UB02]. holothurian [MYH07]. holothurians [HPB^{+}04, UB02]. Holothuroidea [UB01]. Holt [Ano10a]. Homalaspis [FC00]. Homarus [BHH07, DL03a, GSM06, JPC00, OA02, RBSP00, SJCM01, WA00]. Home [SRN02, TAW^{+}09, JKK^{+}06, LTCP03, PBH^{+}03, TLFS06]. Homing [DK03, GSB^{+}06, KAR04, SSB02, BL07, RSD^{+}08, SRJ07a, SRJ07b]. homogeneity [AMS01, FLG02]. homogeneous [UHvO^{+}08]. homogenized [KvdMDP09]. homothermy [SYH^{+}09]. Hong [JKQ02, CLL^{+}09, DDTQ05, DQW05, GXL^{+}08, Lam00a, Lam00b, LMCQ02, LYL^{+}07, MHY^{+}06, QDDP06, SE04, XYH^{+}09, YQC^{+}00, Yin02, Yin03]. Honshu [ENAT07]. hooded [BCHS09, TBI^{+}09]. hookeri [CWDG05, MMC^{+}08b]. hope [HR02b, HR02c, Pet04]. Horizontal [dSSC07, BBABL01, BBA06a, BBZM08, DSM01, GMB^{+}06, MBBW08, RKGH01]. hormonal [PMT^{+}02]. hormones [CCCH01, KPW07, SH04]. Hormosira [BMC^{+}09, Kev01]. horse [MNP02, RCC01]. horseshoe [BLT03, CRV03, JSN08]. hospite [UHR05]. Host [Bat09, CS03b, RWSP03, BMdNP09, BMM^{+}09, CBM00, CD08, DDV^{+}08, DRDD09, Dov04, FMP05, FTU^{+}00, Gou07, KES00, LBH^{+}04, MW09, OHS^{+}04, OHS^{+}07, PWdN^{+}00, SKR08, SGL^{+}02b, SKW^{+}06, TLTB03, TT06]. host-specificity [OHS^{+}04]. host-virus [TLTB03]. hosting [SGP07]. hosts [GBI^{+}06, HGMG^{+}04, JHH01, KES00, SdV06, TFSP09]. Hot [SBI07, CS05]. ’hotspot [Pri02, SBSPS04]. hour [MAB^{+}03]. house [STI03]. Houtman [WAK^{+}09]. Howe [LPS08]. HPLC [BDC00b, EDBvdL08, LPBD00, ZRG00]. HPLC-CHEMTAX [EDBvdL08]. HR [COB^{+}03]. hsp70 [HKH09, RCC^{+}09]. hub [LR08]. Hudson [SLFM03, SRZZ01, TWYS03]. hudsonica [CD03c]. hull [LC09a]. hull-fouling [LC09a]. Human [DGVK^{+}04, DKR09, AVA^{+}09, CWR^{+}03, FBT^{+}01, FDM^{+}09, RC08, VSVV03]. human-dominated [FDM^{+}09]. Human-impacted [DGVK^{+}04]. human-induced [RC08]. Human-mediated [DKR09]. humans [PUL07]. Humber [FR08b, SWGC^{+}05]. Humboldt [DDQ^{+}00, GSFN00, HC05, IG04, KCG07, LJC00, PGR^{+}01, RCL^{+}09, RGCL^{+}05, SAB^{+}02, TDC^{+}03]. humboldti [SAB^{+}02]. humilis [PV05]. humpback [BGDS05, CSR^{+}00, CH00b, FHN^{+}09, GDSB04, HFT^{+}09, HYR^{+}09, OBG^{+}07, PBAD05, SAC^{+}03, WWWR09, WWR09, ZAHJ^{+}06]. humphead [Tup07]. hurricane [BHSR00]. Hurricanes [PWHW09]. huxleyi [CLS^{+}01, CWN^{+}07, HPVL05, SHL^{+}03, SH03]. hyalina [MHB01a]. hybrid [BDLB02, GKS^{+}07, NLG05]. hybridisation [PSPD07]. hybridization [BWH09, HFM^{+}04, LCC07]. hydrate [SRL^{+}02, SPRK02, TBK^{+}03]. hydrates [HSPH05, SPRK02, SGP07, TBK^{+}03]. hydration [FWF02]. hydro [BI04]. hydro-climatic [BI04]. Hydroacoustic [SHNV09]. Hydroacoustical [GGDHRA00]. Hydrobia [CBP^{+}05, HDP02, OSS^{+}04]. hydrobiid [SAC^{+}00]. hydrocarbon [CHP^{+}05, GGGVGG03, KFH04, MFCM05]. hydrocarbons [CWR06, NAB^{+}06, SCH^{+}07]. Hydrocharitaceae [vTWMG08]. hydroclimatic [BIR00]. hydrocoral [WCZ^{+}09]. Hydrodynamic [CDJ^{+}07, JP08, NJPI03, Sar06, Seu05, BvK07, BLH02, DESK03, FBJ00b, HFM^{+}02, JBH00, MKK^{+}09, NSH^{+}08, PvDMB08, SW03, Thi03, WB07, WKOE09, vWCAB07]. hydrodynamics [BDCvB^{+}09, CFMB00, DCBvBE09, GBR01, Lei07, LTH^{+}06, SSMM06, SA03a, SW02b, WPB08a, WPB^{+}08b]. Hydrographic [JMM^{+}09, GMHA09, JNS04, NW05b, NIF00, RRH09, TCC04, TFLA03]. Hydrography [NBB^{+}04, DFR08, Lei07, RDH^{+}09, MDZ05]. hydroid [Fer04, HK04a]. Hydroides [DQ05, GQ00, HLQ02, HH03b, HTWQ05, JQ04, JQ05, LHQ03, LHQ05b, LMCQ02, MHW07, TP01a, TP01b, XWQQ05]. Hydrologic [BA04, COK02, DM08]. hydrological [BPVL03]. hydrolysed [DFMD00]. Hydromechanical [Vis01, BK05b]. hydromedusa [HB07b]. hydromedusae [CCK03, JN09, MR07a, MR07b, MR07c, RCCK09, WCCA09]. Hydrophis [KWHD08]. hydrophobic [AHL^{+}01]. Hydrostatic [GG09]. Hydrothermal [RCC^{+}09, AHK^{+}02, BTL05, Bat07, BDDL03, BEU^{+}07, FTU^{+}00, GZG^{+}07, GLG^{+}05, HVCT09, KCB06a, KFH04, LJM03, LLJ05, LSBJ08, MCP^{+}06, MC05b, MMS^{+}05, PPI04, PGS^{+}00, PAB^{+}02, SKW^{+}06, VT00, Van02, Van03, WTF^{+}05, YFH^{+}02]. hydrothermal-vent [SKW^{+}06]. hydrothermica [PAB^{+}02]. hyperbenthic [CH07a, CDT^{+}03]. hyperborea [NFN03, STJ03]. hyperboreus [SKFPS02, MTN03, PJR^{+}03]. hypercapnia [MOPP05, PB07]. hypereutrophic [BL01]. Hyperia [TT06]. hyperiid [LH09]. hypernutrified [KNU02]. Hyperoxic [ID02]. hypersaline [WK00]. hypersensitive [Dov04]. hypo [KJ03]. hypo-osmosis [KJ03]. Hypoplectrus [HEN^{+}08]. hyposaline [QTB02]. hypotheses [BGDS05, MAWH03, RL09a, Thi03, WWGR07]. Hypothesis [FDDS09, BPCvdM06, CS02b, CST08, DGL01, FMZvW02, GDdS^{+}07, HHO03, JRT^{+}06, KS05, LMW04, LBPH06, MK03b, NRvdV^{+}02, PBF07, SB05a, TOK^{+}04, TD01a, TDC^{+}07, TP01b, TL03, TWT08, ZGV00]. hypothesized [Sha09b]. Hypoxia [CC05, Alt06, BEW03a, BEW03b, BH01, BPLD^{+}05, BLP^{+}02, Bye00, CLSM09, CRTP02, DP09, ECM^{+}05, GWO02, KHB00a, KHK^{+}06, LS08b, Mis04, NPTW05, PMT^{+}02, RT05, SMHC03, STM06, Tal01, TW00, TE00, UHR05, VMV^{+}08, VMV^{+}09, WS01, WS04a]. Hypoxia-induced [CC05]. hypoxic [HCGC09, LS08a, PC02a]. hypsilepis [Gla07]. hysteresis [LDS^{+}04]. hystrix [LLCHG01].

IA [PAK^{+}00]. Ianthella [CWCP^{+}06]. ianthina [LCW06]. Iberia [CTFC08, RGB02]. Iberian [CMRB08, FAPA^{+}09, FAPOF09, ILL06, MMV05, NCÁSBP04, NCÁSGP05, TAÁO^{+}03, VPP04, DCF^{+}05, GCVR01, GUT04, OÁSG^{+}09, RGC^{+}06, TSF01, TFLA03]. Ice [PHO^{+}07, Sma07, TRH^{+}06, AP05b, BCS09, BGS02a, BCC^{+}04, CK05, EKN08, EBA^{+}03, FFML02, FKI^{+}09, GJG^{+}03, GGC^{+}05, GRG^{+}02, GGG^{+}04, GKB09, HS08b, INH^{+}09, JPMGS08, Kaa04, KTH^{+}04, KTK^{+}02, KKL07, KGB^{+}01, LCB^{+}05, MHB^{+}09, MAJ^{+}06, MGF^{+}03, MRG08, MG00, NGMT01, PTK^{+}09, RMGL07, RGP^{+}09, RKGH01, SH06a, SBR09, SHFP^{+}03, SNGR00, WBL^{+}06, WA05, WAN^{+}01]. ice-dependent [SNGR00]. ice-free [EBA^{+}03]. ice-independent [SNGR00]. iceberg [SBFP08, TGGA04]. icebergs [GP03]. icefish [RHDC05]. Iceland [BG04c, Gis05]. Icelandic [PTM08]. Ichthyofauna [LMB^{+}09]. ichthyofaunas [GDW06]. ichthyoplankton [AB06, BRB^{+}00, LVvdLH07, RSC08, ST06b, STRB06, CT01]. ICP [RCC01]. Identification [BMW^{+}00, CWCP^{+}06, KKIV01, KWHS^{+}06, LFP^{+}04, LK02, LVK05, MFWD08, SRZZ01, Som08, Tup07, BAB^{+}06, CBD^{+}05, CWR^{+}03, DNL08a, FT00, GZ02, HLH^{+}05, LTL06a, OPSB07, PSPD07, SC06a, SHSF08]. identified [SAS^{+}06]. identify [BT09, EHWS07, Kev01, LBR^{+}09, WBJ^{+}07]. Identifying [CBL^{+}09, GMD08, HGG^{+}05, IR02, JMJ07, BUA02, HDDH06, LPD09, SMHL00]. identity [DDD08, ISBP06, MBVBC09, MKKK08, RK08, SL05a, Sta09]. Idotea [GLBF07]. If [TT07a]. Ifremeria [SKW^{+}06]. ignobilis [MHP07]. II [BL00b, LOT^{+}00, LCL^{+}00, PW04, AY04b, AVNN05a, BI04, BEW03b, BL00b, BPS^{+}00, BDW^{+}00, BI03, CMHP06, Chi09, CRBE04, DMS^{+}05, DCBvBE09, FBJ00b, FCH^{+}09, HSGS00, HM05b, HHKH05, JK03b, KT01, KBT05, Lap01b, Lee00, MRB04, MVBGM04, MCCT07b, MDBC03, PNCT^{+}08, RCP^{+}05, RFG^{+}09, Sha09b, SL05c, SW02a, SKMS06b, TP01b, WP00b, WL03, WJB09]. III [VNJ06, VFJ06, AVNN05b, BK05b, BDCvB^{+}09, CRCV04, KFT08, LCL^{+}00, LRN^{+}09]. Ikeda [CST08]. Iles [BCH00, CBTW02, CBDH02]. ilicifolium [ABB05]. ilisha [MC01b]. image [CWB^{+}06, GDR^{+}04a, HD05, MDBG07, MWM^{+}05, WMS^{+}03a]. imagery [KTSR02, OKP06, SK02]. images [KS04a, NDOS03]. imaging [FJ01, MUP06, NR00, SWH^{+}04b, UM08]. imbricata [SWHSA05]. immaculatus [JBG00]. immature [APG^{+}09, BBC03]. Immediate [LCD^{+}05, BDH01, Wie03]. immigrating [vdVBGW09]. Immigration [MYNY08]. immune [DLD^{+}00]. immunity [SPT04]. immunoassays [KCG07]. immunochemical [HRQ^{+}02]. immunolocalization [CACDC07]. Immunomagnetic [AKRA02]. immunomodulation [MM05]. immunopositive [DCT^{+}04]. Impact [ABC^{+}05, BSW^{+}09, BCP^{+}08, BVI^{+}05, BRB^{+}00, CCL^{+}06, DBD^{+}05, FF03, FMP05, GKK03, HSWG^{+}06, JHJ^{+}03, JBSB01, KM03b, LTG^{+}01, LP01a, NWFV06, Nug02, PTLK04, PMM^{+}00, PJP06, PPCJ^{+}06, RS02, RBRSP00, RWBS03, RSY^{+}06, RJH02, RL08, SA03a, SP01b, SH06c, TK01, Thi05, THYBS09, TPS^{+}05, WGP^{+}00, WWTH06, AWA^{+}05, BF08a, BCO07, BPN^{+}09, CVF^{+}03, DFT^{+}07, FR07, FPKC05, FKA^{+}06, GSFN00, GHS05, GP03, HMKR05, HBdBV01, HLGCO06, HB04, HJG06, JKK^{+}05, JSMM06, Kad03, LM09, LNPB04, LUIAH03, MHCB08, MR07b, NT06, NGMT01, ORD00, OGL^{+}09a, PGR^{+}01, PP00, RMT^{+}05, RGC^{+}06, SFW01, SPT^{+}00, SCS^{+}07, TBCB^{+}05, UHR05, URFA06, VSVV03, WPH01, ZSKA01]. impacted [DGVK^{+}04, LRM00, MTR^{+}07, SZL^{+}05]. Impacts [AWS^{+}09, CHT06, CHM06, GB03, HKK^{+}06, JPPW01, LGC^{+}06, RKLK07, SG06a, TMFB06, AF06, BUA02, BHSR00, CM04b, DH01, ECM^{+}05, FSO08, Fis05, FFR^{+}03, Hou09, KTSR02, KGHL^{+}01, LW02, LHHB05, MAKMK09, NSBK05, OWBD07, PZ07, PPC09, PHB02, SF02a, SA03b, Syd09, WGF09]. Impaired [WS01, Rye02]. impairment [LBF^{+}07]. impairs [CLSM09]. imperial [QWY07]. implant [Gri02]. implementation [DWB04, Fen08, LSK08]. implemented [PPB09b]. implementing [HF09]. implication [CES^{+}09, ONKK03, OSA05, SMSD07, WMW^{+}08]. Implications [BB08b, AFHS08, AFHS09, ADS^{+}09, AMB08, ALOT01, AGB03, AJ03, BLH01, BRGS^{+}07, BF01a, BPE^{+}06, BO00, BPS^{+}03, BM07b, BSH02, CW03a, CH06b, CÅJ05, CKO^{+}07, CM04b, CMB^{+}05a, CDML06, CT04, DB02, DFJSJ^{+}05, DSM01, DR00, DMV07, DH09a, DPGN08, EDS03, EB08, EB04a, FC00, FBBV09, GFB^{+}07, Goc04, Gor03, GPSG06, Gru03, GFPM01, HR08, HCA06, Hyr01, IAT^{+}07, JWM09, JSN08, JM07, JBT^{+}07, JS01, KEB07, KS09, KW02b, LABR^{+}06, LB05a, MSLP06, MM06b, MWT^{+}06, MSH08, MAVO^{+}06, MHB^{+}09, MH00, MS06, MHP07, MKK^{+}09, NReZW09, NCT^{+}01, NT06, OLG^{+}09, ÓUA05, OHB^{+}00, PCMB06, PRL^{+}09, PPGL09, Pri02, PI07a, PISL02, QTB02, QHD04, RBJB^{+}05, Ran02, RHDC05, RVMI^{+}03, RvdPHW04, SMM^{+}04, SN04, STA^{+}05, SYA02, SPRK02, SMCE07, TMAN08, Thi03, TBID09, Tit01, TK03a, TÁCTR06, TBC^{+}00]. implications [Tyn04, VBV^{+}00, VMN^{+}03, VSRGW09, WH08, WGP^{+}00, WWWY07, ZJW^{+}04, dSSC^{+}08, GC02, KT01, SL05b]. implied [WB01]. import [HCH07]. Importance [BSJ02, CFG06, DL02, EWR08, HTRFP04, Kin02, LW06, MNTS08, MDHL08, MWD^{+}09, NDV^{+}00b, NOE^{+}04, NRH^{+}07, PKR07, RGL^{+}00, SW04, SMH^{+}03, VTS^{+}02, WAK^{+}04, WDRH07, AMR00a, AMR01a, BB00, BIG^{+}06, BSvdH^{+}02, CVÁM06, CR02, CSO08, DDD08, EKBM05, GFF08, Gim04, GP01, GKWR02, GS09, GAM02, GHS05, GRM00, HW01, JLW04, JFR02, KPKD06, KW05, KKL^{+}03, KVH02, KDAK01, LSH00, MD08a, MRPL07, MPS^{+}08, MW01, MH06a, MS01b, NvdV04b, NFN03, PRB^{+}08, PTAR07, RH03, SAJW04, SDFM00, SZ08a, SZZ08, SH07c, SF02b, SJH^{+}04, TDT^{+}03, TA03, TTR05a, VBA06, WWAC00, WB01, YI06, ZE01]. important [AM02b, BRAS03, DS06b, GAB^{+}03, LB02, LJ09, MUT03, MAC^{+}07, NRvdV^{+}02, NvdV04a, QFJ^{+}07, SEVSP02]. imported [BMO^{+}04]. impose [RGCC^{+}08]. imposed [BRP00, DMA02, HCV03a]. Imposex [BGQR01, BMB02, HPC04, SJ02, BMG00, MHW06, RGEB00, SHZ^{+}05]. impoundments [MMPD09]. Improved [CBS^{+}08, Bur01, GT08, MVdMFF08, RAH^{+}02, SP01c, YF04]. improvement [Kin04]. improves [SB00]. Improving [Lat07, GNL08, LIM^{+}06]. improvisus [BJLH04, TL08]. IMS [BWK^{+}08]. IMS-101 [BWK^{+}08]. Inaccurate [Rii01a, Wid01]. Inachus [RGVC07]. inbreeding [CFA07, LP07]. incident [GLS^{+}00]. incidental [GZY06]. incipient [RC04b]. including [OLWB06]. income [MGC^{+}09]. incongruous [BGGR02]. inconsistencies [CC07]. incorporating [BWW^{+}06, MB02b]. incorporation [DM09, EG05, MT05, MW06]. increase [AL08, Bar07, CLS00, FBD06, GM01, Hel00, KF02, LF06, PMS^{+}08, PTH08, PAR08, SAC^{+}03, TKVL04, VSL^{+}04, WS09, YBY^{+}09]. Increased [BOM^{+}08, DL03b, JRS^{+}06, JERS03, MHM^{+}08, SV08, DAA00, FHL^{+}09, HLD^{+}07, IHW^{+}01, JDTB06, KS04c, LBF^{+}07, LRN^{+}09, MAG07b, RB08, RFG^{+}09, RU02, TP00a, WS01]. Increases [BSH02, BWL06, BBG^{+}04, DJM08, MRN07, MAW06, PD05, SHB03, SW08b, ZT07]. Increasing [GEC08, LM09, AHDT^{+}09, Min02, YG09]. Increment [WFC03, Bla09, CW03a, CT06a, FFG03, LSP08, RMH07]. increments [STCJ01, SSM^{+}01]. Incubation [BGRD00, GBM01, HTJ00, MCCP09, NNS01, PS07, dZCVC01]. incubations [WdBS02, dZCVC01]. Independent [OGL09b, BRGS^{+}07, BGL^{+}07, CCL^{+}05, FCC^{+}08, SMA^{+}03, SNGR00, TJ02, WHCP^{+}04]. independently [GC02, NRPS^{+}01]. Index [Ano00a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b, BBE03, CHO09, CW01, IP01, LCT^{+}08, RGP^{+}03, Som08, Yos08a, Yos08b]. India [GTE^{+}09]. Indian [CRJ^{+}09, CBTW02, FDD^{+}03, KFS^{+}09, MAG^{+}07a, MC05b, MBB^{+}04, NBB^{+}04, Øre00, RW03, SHS08, BCL^{+}01, Bur02, CRLP09, CEBR07, GMA^{+}09, GTGM06, HHV06, HVWH06, JMM^{+}09, LLM^{+}08, MGW^{+}09, MRQ07, NLP^{+}01, PMNL07, PVV^{+}06, RLFEF03, RRH09, SC03, TTD^{+}06, TCW^{+}08]. indicate [BPM^{+}06, BHMR01, Car05, DM03, DM05a, DPSD08a, EGPT03, FKBB01, FS02, MC05c, Mon07, NFN^{+}09, STAS06, VDA^{+}03, WKPS06]. indicated [LSMR05]. indicates [BSR09, EWG^{+}08, PPB^{+}09a]. indication [NH05]. indicator [AC02, BPS^{+}03, GKWK02, GNC07, HPS07, HBK^{+}05, KHIH08, KMAC^{+}02, MSP^{+}06, NCC^{+}07, SNY^{+}05]. indicators [Bar02b, Bea04, BCS09, BPA^{+}08, BPLD^{+}05, BS04c, CTH^{+}03, FFHW07, FMW07, IAKS03, ISK07, KKK^{+}05, KCV^{+}06, KFK^{+}07, KJMP05, LP03, MCVAR08, PWAT01, PBCB03, PSW07, PSS^{+}07, PHS^{+}07, RGF^{+}03, Sha09a, Sha09b, TMD^{+}07, dJTD07]. indices [AMG^{+}07, BDC00a, BHR08, CH06a, Das03, GLV^{+}09, HM09, HCD^{+}09, LCHG03, MRQ07, RC08, SJ05, TGA02]. Indies [RC04a]. indigenous [EPM06, JPG05, NZM07, NZM08a, SPT^{+}00, ZMN07, ZMTB07, vdMPEM08]. Indirect [CC03b, HSO04, PE05, RH08, SP08b, BS04a, GSM^{+}02, KF02, Mou04, PJP06, SMC09, Sta09, VML05, VHV^{+}04, WAK^{+}04, WS00]. indirect-developing [VHV^{+}04]. indirectly [HGJM09]. Individual [KSR^{+}07, KLW^{+}09, KGF^{+}09, MMVB07, SHT^{+}04a, TCG03, VPG04, XW01, AML^{+}03, BPM^{+}06, BWW^{+}06, BCKP09, BHMR01, CBM00, CDJ^{+}07, DGCM09, HWGB02, JT08, KTK^{+}08a, Lei01, LGM^{+}09, MPR^{+}06, MGC^{+}09, Mil07b, NRC07, NMBH02, OPSB07, PW09, PD07, PKLD09, QMM^{+}08, SRB04, SM05, TBI08, WPM09, WH00]. Individual-based [KSR^{+}07, KLW^{+}09, MMVB07, CDJ^{+}07, Lei01, Mil07b, PD07, PKLD09]. individuals [BHH07, BOZ03, NWHH^{+}09, SDH01]. Indo [AH08, DGC^{+}05b, LLCHG01, LCW06, SB07, SPT04, THP^{+}05, TCWR08, USG^{+}08, VS08, WGV^{+}02]. Indo-West [LLCHG01]. Indonesia [ATR^{+}08, CR07b, Fox04, dVCH^{+}06, dVBC09]. Indonesian [AMT03b, BSK06, KAM^{+}06]. induce [IMTW09, PHP^{+}03a]. induced [AP00b, BGK^{+}06, BdBK^{+}00, BWvdE^{+}06, CEB^{+}08, CBAd00, CSL^{+}05, CC05, EP08, FGPS04, FI09, Fre07, FGS00, Goc04, GRB^{+}03, HDW02, HBG^{+}03, HGMG^{+}04, HSSS05, Kaw01, KFT08, KBSWG09, KGRW06, LWCF04, Law07, LT07, MMC^{+}07, MYE03, MGM05, MKR07, NH05, PCSS01, PQRAC08, RS00, RC08, SHB03, VPS^{+}02, WWL04, WDM^{+}09, WG02c, Wri07, ZT07, ZLAK04]. inducer [KSR^{+}07]. induces [MM05]. Inducible [Alt06, HHJ04, CC04, MKKK08, RCC^{+}09, SSG^{+}00]. inducing [IH02, SCC05]. Induction [ABB^{+}00, HLQ02, LHQ03, LHQ05b, AD00, CHMR05, LBC07, LMCQ02, LW05, NLK00, RE00a, SCBA05, TP00a, WLdG^{+}04]. industrial [BMF06, KK02]. ineffective [PB07]. inermis [Zup00]. inexpensive [AL03a]. infauna [AFFH04, CS03c, CTHF01, FAK07, GFF08, PH05, RL09a]. infaunal [DM08, DJ04, EWSH08, ISBP06, JNDL02, KKBH03, KH05, LBBM06, LBB08, MK03c, PZA09, RW06b, SCZ^{+}06, SYA02, TE00, VN02, WDM^{+}09]. infected [LWB09]. infecting [OHS^{+}04]. infection [AML^{+}03, ESQ03, LNV^{+}03, OST07, PR01, Thi06, TFSP09]. infections [PM03]. infer [TSEH08]. inference [FBSG02, MM07a, SRJ07a]. inferences [CNT^{+}04, DPGN09, KSHP00]. inferior [MN04]. inferred [ACJ^{+}04, BPT^{+}00, BM00, BDDL03, GRPE08, GF04, HM06a, HH00a, KNKT01, MC05c, QMF05, RW08, TR06, WTF^{+}05, XYH^{+}09]. Inferring [ÁSG04, BPC09, RSA05, TH02, DK06]. infestations [TWMW06]. infested [MS00a]. inflatus [GRPE08]. inflow [DSM01, KLRS00]. Influence [AML^{+}03, AM03, APM^{+}09, ABY00, BFK05a, Cai09, CCC09, CW05, CWR06, CSD^{+}07, CHA06, DTD03, DD04, DVBLS07, DVNvdV07, FHG^{+}03, FBKB03, FEO07, GGMGR07, GSFF^{+}09, GCS06, HDP02, HDC^{+}08, HAP^{+}08, Hir06, HF05, HFA09, HZ04, JPC00, Jen05, JT08, Kel02, KWWH02, KB03, LLM^{+}08, LG00a, Las03, LLY^{+}01, LSdM^{+}03, LC03, LAW06, LJK01b, LJK02, LSG05b, LPM^{+}06, MADPQ06, Maz04, MYE03, MC06, MGM^{+}07, McQ05, MAA^{+}07, MTLT02, MNP02, NPVW06, PDM^{+}04, PJ06a, PA05, PWL^{+}06, Pra01, RMA^{+}02, RGMJM07, RHHH09, RHI^{+}01, RGP^{+}09, SFHS00, SLM^{+}02, SKBP09, SJ08, SR03, SW03, SD00b, SVC^{+}01, SRS04, SP00, TGM07, TSJH05, WBSL00, WWH05, Yin03, ZYZ^{+}06, dZCVC01, BPBH09, BPT^{+}00, BGSBC02, BP08, BGS02a, BOZ03, CFMB00, EM07a, EB08, FC04, FC08, Fer04, FHB^{+}02, FTU^{+}00, GHV02]. influence [GFP^{+}05, GR07a, GMTN00, GJ03, GC06b, GBR01, Har00, JRT^{+}06, KR04, LH09, LR06, LZM^{+}03, LTH^{+}06, MB02b, MBG^{+}08, NW06, OTBR00, Orv05, OÁSG^{+}09, PBVP09, Pra05, RATD07, RRRM08, SML^{+}09, SESP03, SMHL00, SB06, VK05, WR00a, WAMP08, Yin02, dBTO^{+}09, DB07b, MRB04]. influenced [Bib07, CCL03, GB03, HSPH05, SCS03]. Influences [SWH04a, YH00, AP04a, AP04b, BBM03, BS08b, BM00, CSH00, CVT09, EHWS07, EEWBdN08, FNB07, GM04, GC05, HM06c, JDTB06, MAH^{+}08, NSH^{+}08, Pee02, PBF08, RH06, RAS07, RU02, RCB05, SSL00a, SEH07a, SRB04, TKK05, VHW^{+}07, Wan02a, WBC08, WM01, WS06, SKMS06a]. influencing [MMBF09, NA04a, PB00a, PCH^{+}05, SAB^{+}02, ZPP09]. informatics [CVB06b, CV06]. Information [FGLM06, GDF03, AVE^{+}06, CDG04, CLJ^{+}08, GSB^{+}06, GDL^{+}01, LTL06a, MDHL08, PKS07, SVN04, SMP01]. Information-theoretic [GDF03]. infralittoral [TK04]. infrared [AL03a]. ingens [CD03b, JBG00, PJN01, PJN03, SJ05, TR05]. ingested [CVÁM06, GFC^{+}02, HD07a, TKTJ07]. ingestion [AG04, BD02, CNC00, CD03c, CG04, DC07b, HSST01, NHN^{+}06, RHI^{+}01, SD00b, STJ03, TPSW00]. ingress [HTW^{+}05, LLW09, ODF09, SRJ09a]. inhabited [CVK00, ID02]. inhabiting [BAB04, FVAB00, JH06, RJ03]. Inhaca [PDQ01]. Inhibition [KWB^{+}06, MGZ^{+}03, BNP07, CBO04, FNDP08, KSKK^{+}07, MP05, NCHP05, NP05, TD01a, VPS^{+}02]. inhibitors [GDBF06, Jon04]. inhibits [NCP^{+}08]. Initial [GR07a, FBKB03, FJ01, OFHM03]. Initiation [LHD^{+}01, PBAA04, EBA^{+}03, SMH^{+}03, TGSW06]. injured [FSC05]. injury [Cum02, DSH03, PMGE01]. Inland [AY04a, AY04b, MMTT00, OHS^{+}04, TYT^{+}04, TD01b]. Inlet [AP05a, DER01, GB07, GRRH07, KST^{+}08]. inner [JCZ^{+}05, LP04, TNSM00a, WBG09]. innocent [GLS^{+}00]. innovation [SN09]. inoculated [HGS^{+}07]. inopinus [CRB07]. Inorganic [PTK^{+}09, CSMK09, GSW01a, HLR07, JHTV02, LTMDB02, LS05, LLL^{+}09, MN00b, SSB09, SH07d, TDUN07, WW03, WBN^{+}00]. inorganic-N [WBN^{+}00]. inornatus [MVBGM04]. input [CPG^{+}09, KHD07, KR04, Lei07, MP06, RW06b, vOSF^{+}09]. inputs [ADA00, BMO^{+}04, CHW02, CRK06, DAA00, DGW01, PMG^{+}06, PWHW09]. insect [GJT09]. Insecticides [MBHN07, NSFW09]. Inshore [SAM^{+}01, BGH06, GGDHRA00, MASFS00, PDQ01, PTSJ06, VLW09]. insidious [FI09]. insight [AGM06]. insights [HH00a, LRR^{+}09, MH06b, PCS^{+}07, RWB^{+}09, WISH09, WAA^{+}09, vOSF^{+}09]. Inspiration [WZ04]. installation [FGPS04]. instantaneous [DRTM09, MGM^{+}07]. instar [GSM06]. instars [OvM02, vMRO03]. insular [OPSB07]. intact [KHD07, TRPNE06]. intake [AG02a, EOL^{+}05]. integer [RWJ02, WG02b]. Integrated [MM07a, Ber02b, CG05, DRTM09, Dun02, Sar06, VCA^{+}09, Ber02a]. Integrating [ATT06a, HTL^{+}02, LTL06a, GPLAG04, Lin04, PKS07]. Integration [TEB^{+}07, CPP^{+}04, CSMvE04, SZ08a, SZ08b, WAZ07]. integrative [MDS^{+}07]. intended [CB08c]. Intense [EJKN03]. intensely [SBFP08]. intensities [ZTW08]. Intensity [VLW09, AED^{+}07, BCS09, BVMBC07, BWK^{+}08, GBHO^{+}08, LMMT09, LMW04, MRN07, MMMPB05, MSK^{+}00b, Phi05, SH06a, Sma07, VHW^{+}07, VBMBC08b, vWNYS05]. Intensive [VYM06]. intensively [RGS^{+}09]. Inter [CDS07, HRCB02, JNS04, Kin05a, LRC06, PMG05b, PW09, TCWR08, WAB^{+}02, AS04, CMD^{+}09, FBD^{+}02, FDRR05, IMTW09, JSIT07, KHHM07, LCL^{+}03, LCGN02, SAHF03, SBR09, SM05, TM02b, WP03]. inter- [SAHF03]. Inter-annual [HRCB02, JNS04, WAB^{+}02, AS04, IMTW09, LCL^{+}03, LCGN02, SBR09, TM02b, WP03]. inter-colony [FBD^{+}02]. inter-comparisons [KHHM07]. inter-habitat [JSIT07]. Inter-individual [PW09, SM05]. Inter-oceanic [TCWR08]. Inter-population [LRC06]. Inter-reef [CDS07, CMD^{+}09]. Inter-Research [Kin05a]. Inter-sexual [PMG05b]. inter-tidal [FDRR05]. interact [IC02, JFD09]. Interacting [ML05, MCN09, OSS^{+}04]. Interaction [DD07, HK04b, KCG07, LSH^{+}07a, MSS^{+}03, PDQ01, SH07c, BJJ01, CÁ04, CST08, DCP^{+}08, Fau09, GLP^{+}08, JC08, KH00c, QRS00, RBGM06, RC05, SKP^{+}07, TE00, WM06, WMW^{+}08]. Interactions [AG02a, FZT^{+}05, JKS^{+}01, LD03, PGK^{+}08, SSBG02, Uth01, WH02, Wan02b, AMS^{+}06, Bar07, CB05a, CDW^{+}09, CBM^{+}07, DM08, DEAI06, DTAF00, FTP^{+}09, FE02, FSW09, Fer04, FD04, FT00, GNPB07, GNP07, GHJ02, GHT^{+}01, Han07, HGSH08, HBRJ06, HTL^{+}02, HLY05, HCK^{+}05, IOB04, JMA02, JNMB08, JY08, KES00, KBA^{+}05, LS08b, Lus03, MACS07, MGDP01, Mis04, PIM^{+}05, PdNS06b, PBFB08, PH01b, PBF07, PBF08, RWPJ04, SGB09, SG06b, SCBA05, Tit01, TLLKC04, TÁCTR06, TBH^{+}09, VHB^{+}07, WTP08, WB01, Wie05, YNM^{+}07, dBTO^{+}09]. Interactive [DESK01, LMMT09, JMdG^{+}06, SO01a, THC^{+}05]. Interannual [CMT^{+}08, DCC^{+}03, ESDKM09, PHS09, SD09, AAGS02, BC02, BKDB00, GTPC04, HST^{+}09, MGE^{+}09, MHB^{+}09, MMA03a, MRH^{+}08, OA02, vdVBGW09]. Intercalibration [PBS^{+}04, Rii04a]. interception [FMZvW02]. interchange [OPSB07]. intercohort [SML02a]. intercomparison [KFE^{+}03]. Intercontinental [KCS04]. interday [LWR^{+}02]. Interdecadal [CS02b, HBRJ06]. interdisciplinary [RDH^{+}09]. interface [BBT02, CIFN00, DBT02b, HL05]. interface-feeding [HL05]. interfaces [CGCMPA01]. Interference [Mok04, GSM^{+}08, GSS09a, HLTT09, TE00]. interhabitat [HKH09]. Interlinkage [NBN^{+}08]. interlinked [NvdV04b]. intermedia [OYNO08]. intermediate [LMW04]. Intermittent [LG06, ECM^{+}05]. Internal [LCCTFO03, CBM^{+}07, JKK^{+}06, LCF02, PG01, TFA^{+}08, TVG00, WB05a, WB05b]. International [ZK09, FR08a, GWWC00]. internesting [HBG^{+}02]. internode [KWTR08]. Interplay [BBM02b, DPS^{+}03, LKR^{+}06, vWCAB07]. interpopulation [RZLS09]. Interpretation [SMHG09, Rii01b, SMM^{+}04]. interpretations [TM02b]. intersex [BMG00, SF05]. Intersexual [BCHS09]. Interspecific [BF01a, CF09, HKH09, Ste04b, TN07, GHJ02, MHK^{+}01, SS00a, SF07a, SJR00]. interstices [TAS07]. interstitium [PVGO01].

Intertidal [BCS09, BAMJ09, DGCM09, HMA06, Hor06, MBW03, Mou04, PA06, TB08, AHBM^{+}07, Alt06, APM^{+}01, AFFH04, ABU01, ARP09, ABB02, BAA04, BAA07, BWB^{+}06, BWP^{+}06, BvK07, BH06, BM05b, BNSG01, Bul05, CB00a, CBM^{+}03, CDO08, CH03, CC04, CEB^{+}08, CSD04, CLC04, CH06c, CML08, Chi09, CZW^{+}08, CCB^{+}09a, CDK^{+}07, CTvdM^{+}08, CHG05, CBE04, CRBE04, CRCV04, CVSM09, CvBGR09, CTH^{+}03, DM08, DMH01, DD05, DMM^{+}03, DD04, DS01, ED02, EB08, EPM06, EF07, FA06, FMCB05, FC00, FKSW09, FNB07, FH08a, FH02, FD04, FS07, FMP05, FR08b, GJT09, GRCdLvB08, GH08, GGS09, GLFU^{+}04, GBG^{+}00b, GBR01, HDP02, Hal00, HMH04, HDW02, Har06a, HS08b, HSC^{+}09, HGRB04, HBP^{+}05, Hid03, HMK06, HDCA06, HAD07, HJ07, JC08, JMU09, Jar00, Jef03, JÅC^{+}01, JCD^{+}05, Joh01, KT06, KLC^{+}06]. intertidal [KCP^{+}07, KT04, Kne03, KHK^{+}04, KvdMDP09, KS09, KKI^{+}06, LNV^{+}05, LTNC07, LNB^{+}03, LEB07, LÅCN01, LLL^{+}09, LMF^{+}01, MBW02, Mar02, MHTJ07, MTH^{+}08, MB00a, MB01, ML00, MP00, MP06, MT00, MA00, MMMB^{+}05, MMA03a, MK03c, MUP06, OC01, OGMC06, ORPND01, PMPI03, PKS00, PH03c, PMG^{+}06, PTH08, POH08, PM03, PD05, RS06, RP06, RFN02, RS07a, RBG^{+}07, Rod03, RU02, RGMH08, ROBFR08, SF06, SAHF03, SA03a, SL07, SH06a, SH07b, SC00a, SdSC01, SS09b, SM09, SCZ^{+}06, SvNvdMvdV09, SP01c, SL05d, SMSD07, Str02, SHW09, TK04, Tal09, TN04, Thi06, TJS08, THN^{+}03, Too04, TNSM00b, Tru02, Und04, VT00, Vie09, WBN^{+}00, WBP^{+}06, WPB08a, WG02c, Wol07b, YG09, YH00, YTK^{+}05, YH02, dBS01, vBSSH02, vBWSS06, vKH00, vdWHF^{+}08]. intertidal-subtidal [Alt06]. interval [HBG^{+}02]. intestinalis [BR03, BAB04, CRZ04, KF01, MK03a, RWBS03]. Intra [LG00b, SS00a, SF07a, TM02b, GHJ02, GDR^{+}04b, RC00, SML02a]. Intra- [SS00a, SF07a, TM02b, GHJ02, SML02a]. Intra-class [LG00b]. intra-habitat [RC00]. intra-specific [GDR^{+}04b]. intracapsular [BFV08, CGT^{+}00]. intracellular [SCS03]. Intracolony [FHPW07]. intraguild [BL09, CDO08]. intrapopulation [CC04]. Intraspecific [EEA01, HM04, LMC08, MSK00a, MAH^{+}08, MPS^{+}08, NW06, WÓ02, BK05c, Pet02b, SAHF03, WH00]. Intrinsic [CWM^{+}07, HWB^{+}09, MG08]. Introduced [CMRB08, Arm01, BDG^{+}07, BKB^{+}09, BK06b, BS04a, BBC08, CRB07, DJV03, HK02c, HB04, HRDA05, JMA02, LG08a, NWÓ03, PCC01, PC02b, PJ06b, RJH02, RKLK07, RFH^{+}06, SM07b, SM08a, SEH07b, Smi04, VGB^{+}05, WOH05]. Introducing [JB07b]. Introduction [Boy08, CVB06a, GNP07, GSS09b, JB07c, KBP07, MB07c, PSW07, Sol09, VHG08, BCS^{+}06, SB09]. introductions [KRVB05]. introgression [BWH09]. intrusion [KKTN05, MTLT02]. intrusions [CAS00, CABS02]. intuition [FBSG02]. intuitions [CES^{+}09]. inundation [FDRR05, RRRM08]. invaded [MGD^{+}09, NLG05, SS09a, SSB02, SZZ08]. invader [DD06, IOTL09, LC07a, Thi05, WOH05]. invaders [MRMA07, WWR^{+}01]. Invading [HCS^{+}07a, HH07a, MHW06, Zab09]. Invariant [WLM06]. invasibility [SF06]. Invasion [BSBD^{+}04, DBvOM06, JTN09, Kad03, MVBGM04, SMC09, SPT^{+}00, Bye00, CLH^{+}07, FHDA^{+}07, HCB02, NW05a, RBG^{+}07, SBSPS04, SM09, SB06, TD01b, TYC00, VGB^{+}05, VTY06, WGV^{+}02, WG02a, WWR^{+}01]. invasions [BP06, GSK^{+}08, SBSRW02]. Invasive [AH08, FKA^{+}06, GW06, AW03b, BRJ^{+}03, BG04b, BG07b, CD07b, DBR^{+}07, DFCH05, EPM06, Gri09, HKH09, HH07b, JHR05, KSI09, KMWH04, Kim06, Lar03, LGD04, LC07b, MTR^{+}07, MP00, MS03, NZM07, NZM08a, PG01, RLBB04, SG06b, SG09, SH07a, SM09, SJ09b, SDM06, SS05, TKGS04, TJA08, WS04b, ZA07, ZMN07, ZMTB07, vdMPEM08]. inventory [CBEL06]. inverse [ÁSG04, BJR04, CVSM09]. invertebrate [Ark09, AP00b, Bat09, BLHB04, BWO04, DKR09, DP01, HJ07, IC02, JPPW01, JS03, JKQ02, KB05, LBC^{+}09, Mar02, MK03b, NNLV06, OJ02, PDCH06, PH05, PJP06, PSPD07, RWB^{+}09, RBG^{+}07, RGC05, SM08b, SS09b, THJK06, WP00a, WP00b, WBL^{+}06]. invertebrates [BJH04, BC01a, BWMA02, BGS02a, CM04a, CD07a, CZW^{+}08, CHT09, DBvOM06, FA08, FHG^{+}03, GC05, GCL04, IE06, IFL02, KMHS04, KKIV01, KvdMDP09, Kur08, LTNC07, LBCS05, MSK00a, MM03, MSY02, MK03c, NATR^{+}05, PDT09, Phi05, PPSP07, PCB08, RMM04, RWP07, SLL02, UB02, VCA^{+}09, YFH^{+}02, YM06, YLL^{+}04, ZS03]. investigate [ABS^{+}09, CH07c, DDGK^{+}08, MCS09, NMBH02, SSMM06]. Investigating [EGJ07, FCS^{+}09, SW02a, TBH^{+}09]. investigation [HLQ02, MLW^{+}02]. Investigations [HHC00, GMZvdL07, GNAD^{+}04, RJBW03, Wri07]. investment [BJ08, BFE^{+}03, YW04]. involve [SP01c]. involved [NA04b]. iodate [WPD02]. iodide [WPD02]. ion [MBLA01]. Ionian [CLO^{+}09]. IPP [DCT^{+}04]. Ireland [GGDB09]. iris [PS06]. Irish [OHN01, FGT^{+}07, GMTN00, HDM04, ICH04, JKNJ07]. Irminger [HSHA06, tAHSHA07, ITH^{+}03, WW07]. Iron [FBHB03, BL00b, BB08b, CB08c, DHBZ00, DAM^{+}08, Den08, FR08a, GSW01a, HHKH05, KB07, KHWH03, KWAT08, KSKK^{+}07, KBSWG09, KP01b, LOT^{+}00, LCL^{+}00, Law08, Lei08, MHY^{+}06, SDRH01, SBC^{+}06, SP00, TDvdW^{+}01, WBT^{+}08, dBGLT08]. iron- [KSKK^{+}07]. iron-induced [KBSWG09]. iron-replete [KWAT08]. IronEx [BL00b, LOT^{+}00, LCL^{+}00]. irradiance [ACLF09, BRBPL06, JHL00, Joh04, MG00, STRD02, SP01a, SHS00, SH07c, SHM02, SP00, VMH04, WS02, jMFVPL05, vLvSGS05]. irradiances [WdBS02]. irradians [MB07b, BCBQ01, HD07a]. irregularities [BBG^{+}08]. irreversible [vdVBW00]. Irrigation [MGM05, OJ02, SFH^{+}00a, WE04]. irrorata [AFHA05]. Ischyroceridae [JYS06]. Isefjord [JHTV02]. Ishigaki [HKY^{+}02]. Isididae [DS09b, Thr09]. Island [ASP^{+}03, BD00, BB03a, Boo02, BG07b, CGL^{+}02, DNL08a, GK07, GSCW00, HKY^{+}02, HGB01, HRCB02, KHT^{+}02, KKLL02, KFS^{+}09, LGBWH06, LPS08, LW02, OKM06, PDQ01, RHHH09, RTPS07, RAM03, SBFP08, SKH06, TDT^{+}03, TRDV00, TWYS03, VBV^{+}00, WGD05, DWB04, HWC^{+}06, JTR05, Kri02, NRvdV^{+}02, NvdV04b, RC04a, RS00, STRB06, BDG^{+}04, BKDB00, BHD04, BT00, CWT06, FVAB00, GBG00a, GSW01b, GBGH02, GHT^{+}01, KMB^{+}08, Lap01a, Lap01b, MGTK00, MCE^{+}01b, PJN03, RLFEF03, RW06c, SAM^{+}01, TTD^{+}06, VSB04, dBTO^{+}09]. island-induced [RS00]. islandica [BG04c, HWvdM03]. Islands [CMMN^{+}01, DFSS08, DvD02, GSM06, GFPM01, KPM00, Kon00, MGH00, OA02, RDD^{+}02, STE04a, SH00, WAK^{+}09, WG07, WCH09, Zam02, dVBC09, CDG04, FD02, MTH^{+}08, AAHDS06, BLD^{+}05, BBV^{+}01, BZLD02, BMM^{+}04, CD03b, CDF^{+}08, Chi09, Chi02, CS03a, CMT^{+}08, CH00b, DMK^{+}08, DF04, Edm00, GGS09, GDBF06, KHK^{+}04, LW04, Nem05, PWLC06, PJN01, PISL02, RM06, UB01, WH03]. Isle [OHN01, MFG^{+}04]. Isles [FGT^{+}07]. Isochrysis [OMT06]. isolated [BAB04, EKN08, HFA09, LMCQ02, LW06, NWHH01]. isolates [BMS04, FPC05]. Isolation [MCCP09, OPSB07, RFCM03, AKRA02, BSZ05, CSM^{+}09, DFY00, NJPI03, VFJ06]. isopod [BB02b, JHH01, ÖNCNG05]. Isopoda [Gru03, GLBF07, TD01b]. isopods [GLG00, GL02, PGK^{+}08]. isotope [AND05, ARAMPW02, AYSL09, BPT^{+}00, BGS^{+}02b, CF09, CF02, CHO09, CH07c, CKO^{+}07, CR08, ECK05, FE07, FHB^{+}02, Fre03, FCH^{+}08, GFF08, GSL^{+}06, GV01, GFN^{+}09, HARC05, HTM^{+}02, HM06a, HBW^{+}05, HHH02, HMK06, HH00a, HD07b, HEN^{+}08, JMB^{+}05, JBS08, JWM02, JMSM08, JCK03, KTNK07, KKL^{+}03, KSK07, KKIV01, KMB^{+}08, KMK01, KW02b, KG07, LSMR05, LHK01, LGM^{+}09, ML06a, MRHP01, MD07, MS01b, MPB^{+}09, MSF04, NHS^{+}08, NWHC03, NSEL00, PMNP^{+}01, QFJ^{+}07, QMM^{+}08, RBDK00, RSNG09, ROBFR08, SJE^{+}06, SSS^{+}05, STH^{+}06, TRH^{+}06, TSEH08, WAZ07, WWWR09, WWR09, YI07, ZS04, vDSD02, vOSF^{+}09]. isotopes [BPM^{+}06, BW01, BEU^{+}07, BAOV06, BWCW02, BMSD00, CLJ^{+}08, CDF^{+}08, DFPG^{+}03, DSS^{+}09, DOML^{+}07, FC08, FGSH^{+}05, FHDA^{+}07, FOH^{+}08, HWB^{+}09, HMW^{+}00, KFK^{+}07, LF08, Lee00, MSF05, MWR07, MLM^{+}02, NMBH02, OP08, PCS^{+}07, QMF05, QRWJF07, RTH06, RSD^{+}08, RCGA^{+}04, SSMM06, TBI07, TBI08, WIB08, dlMPN^{+}03]. Isotopic [YTK^{+}05, AKRA02, DHHY09, DCA^{+}09, GTSS03, KTK^{+}02, NEK^{+}07, NOO^{+}08, PTK^{+}09, RRS^{+}09, RW06a, RFPS^{+}02, RMH^{+}09, Sar06, SHS^{+}05, TFA^{+}08, VHV^{+}04, YI03b, YTH^{+}05, YIY08]. Israel [PDM^{+}03, RGEB00]. Israeli [RBG01]. ISSR [SPI^{+}07]. issues [OHB^{+}00, Raf06]. Isthmus [OHCFJ04]. Italian [CRC02]. Italy [SCB00]. ITS-types [OPSM04]. ITS2 [PG01]. IV [VNJ06, CVSM09, vdVBGW09]. iwatensis [HIM05].

J [Ano00c, Ano00b]. jack [BBG^{+}06, HK04b, MSK^{+}00b]. Jamaica [LG06]. Jania [CÁ04]. Japan [AK02, AY04a, AY04b, AAY^{+}07, BWH09, CS02b, ENAT07, FTK^{+}04, FKM^{+}01, GSK^{+}08, HHIS06, HKY^{+}02, HTM^{+}02, HGS^{+}02, HTS06, KKTN05, KS04a, KKA^{+}06, KSHO01, KOY^{+}05, KYYM00, KST^{+}08, MMTT00, MT00, MTS07, NTN08, OKAM00, OHS^{+}04, OKH^{+}07, RYF01, ST07, SKH06, TK04, TYT^{+}04, TSE^{+}08, WTF^{+}05, WDT^{+}08a, YHSH02, YTK^{+}05]. Japanese [HYLT03, HT07, HN03, IAKS03, IT06b, Kad03, KS04a, KHK^{+}04, KOY^{+}05, OKAM00, OYT^{+}02, ST06b, ST07, SKO^{+}08, TW04a, TW04b, TAM03, TAM04, TAO07, TKN00, TTL06, TSI02, YOM^{+}09]. japonica [CMD^{+}06, DLT^{+}06, HYLT03, HT07, Kaw01, KYML02, KOY^{+}05, Lar03, NW06, OYT^{+}02, SAW^{+}01, SICT03, SP03b, TK04, TTL06, TA01b, TSI02, YOM^{+}09, YSS03]. japonicum [MTS07]. japonicus [ST02, ST07, SKO^{+}08, TW04a, TW04b, TAM03, TAM04, TAO07, TLFS06]. jarbua [LW05]. Jassa [JYS06]. Jasus [BO00, CB08a, FMT09, GFN^{+}09, JWM09, JRTJ04]. Jeddah [RSNG09]. jelly [BN04]. Jellyfish [PKRK07, BG04b, BAB^{+}01, BBA^{+}05, BSH02, GMM03, HMKR05, HDDH06, JPG05, KUOI06, LHHB05, MR07a, MR07b, MR07c, PK00, PBS^{+}00, PS01, Pur03, PUL07, RTB06, TRS^{+}06]. jellyfishes [BPC09]. Jersey [WCRB03]. jet [HGZ^{+}00]. John [Edm00, GDBF06, RM06, SAV^{+}09a, WFHS08]. johnsonii [HHV06, VHM09]. Johnston [SPCG09]. joint [PMGE02]. Jonah [MH08c]. Jorge [GZY06]. journal [RBL^{+}04, Rii04b]. Ju [She08]. Juan [RDD^{+}02, Zam02, CTVP03, LJM03, MTH04, UBM^{+}07]. jubatus [GT08, KG07, WTR02, WWS^{+}05, WS06]. julis [FCS^{+}09]. July [MLG^{+}03a]. jumbo [DJG^{+}07, GES^{+}06, GMB^{+}06, IMW^{+}02, RCGA^{+}04]. Juvenile [DE08, Edm04, MRH^{+}08, NA04b, TPGQ07, TT07b, APA07, AMG^{+}07, ATLC03, APW^{+}04, BM00, BBS05, BVVM02, BBM00, BM07a, BMS04, Bri03, BFLÁ01, BM02d, BL09, CLP^{+}06, CNW^{+}07b, COSM03, CRCN07, CT04, DE01, DKA^{+}06, DMS^{+}05, DES^{+}00, DGdG^{+}09, DH06, Eck02, EBLE07, Edm00, EBC04, Edm07, EM05, EG03, FHVMDS07, FBBH07, FH08b, FLL09, FCS^{+}09, FS02, GSM06, GRWB02, Gil02a, GAB^{+}03, GC06a, GR07a, GBW02, HJG03, HKZ04, HCM01, Hel00, HD06a, HPR05, HF05, HMFC07, HSSS05, HMHB03a, HMHB03b, IT06b, JTVC08, JSIT06, JS01, KMHS04, KHHM07, KCEB09, Kup03, LGB03a, LGB03b, LGB^{+}04, LDK^{+}06, LGD04, LR06, LP07, LWB09, Mad02, MPSJ06, MT05, MW06, MS04b, MS05, Mis04, Mok04, NDV^{+}00b, NvdV02, NvdV04a, NvdV04b, NGBG07, NRM05, NFN^{+}09, NM04, NW02, NCT^{+}01, NL05, OvM02]. juvenile [PBCB03, PGK^{+}08, PBF07, PPSP07, PS01, RMKR00, RS00, RJH02, RK09, Rye02, SRDPT^{+}07, SBMAE06, SBJ02, SEH07a, SBJ06, SL05b, SL05c, SO01a, STM06, SMP01, SMT08, TNTH04, TW00, TC03a, THC^{+}05, TSSQ06, TSHT02, Trn02, Und04, VHH^{+}02, VGW00, VNdG^{+}06, WKH02, WMG02, WK02, WOL^{+}07a, WGR07, WPIB09, WBDR06, vMRO03, vdVBW00]. Juvenile-adult [NA04b]. juveniles [GQ00, GDKN08, GLG00, Hun04, JP04, KKG06, Kro05, LWJC04, MH06a, NA04a, SKO^{+}08, TOK^{+}04, WBR07]. juxtaposition [Win09].

kaleidoscope [Tor09]. Kalimantan [dVBC09]. Kandelia [CDT^{+}01, SKH06]. Kaneohe [Zab09]. Kangerluarsunnguaq [MRG08]. Kara [FA08]. Karakassis [UG04]. Karenia [GMC^{+}09, HR08, JK06, SKMS06a, SKMS06b]. Karlodinium [WTP08]. Kasyanov [Pud05]. Katrina [PWHW09]. Kattegat [PWB^{+}03, AM03, CLB^{+}09, CCMK03, SJ02]. Kelletia [LZW^{+}08]. kelletii [LZW^{+}08]. Kelp [Tal09, BS00, BER^{+}03, BSED08, CWKE05, CHO09, CEF04, CRT^{+}08, DBJ^{+}00, DZ06, EE06, FMCB05, FWC02, Fre03, HAC06a, HKH09, HH05b, HHW^{+}07, Hob00, IC06, KPKD06, Kaw01, Kon00, LJ09, LTCP03, LW04, MH08c, MC08b, NFN03, PRB^{+}08, PK01, RC04b, Rod03, SM08a, SG09, SMF^{+}02, SFR^{+}09, TS05, TKGS04, VJ05, VLW09]. kelp-derived [KPKD06]. kelps [DESK01, DESK03, EDS03]. kelts [HHD^{+}09]. Kenya [DGVK^{+}04, KHT^{+}02, KAR04, MHO07, MSG07, MH08b, MH09]. Kenyan [AM02a, MG05a]. Keratoisidinae [DS09b]. Kerguelen [CBTW02, AMR00c, BCH00, BCL^{+}01, BZLD02, CWT00, CBDH02, CD03b, CDF^{+}08, CMT^{+}08, LGC^{+}08, LB05b]. Kermadec [WG07]. kernel [EOK^{+}08]. kernels [ANCC07, EHWS07, PLG09]. Kerr [Ano10b]. keto [SGK04]. Key [TLFS06, BHH09, BVP^{+}06, FT00, KCG07, LBR^{+}09, MMB^{+}02, NMCT03, PSGP04, PI08, TSF04]. Keys [CH05, KVH02, VHKW00, CSS00, CSM02, CSO08, DSL07, DE08, EBC04, HALL05, KH07b, LSG08, McG02, RV05, SC06a, SGCP06]. keystone [GH08, Kon00, MAH^{+}08, SHW09]. Khalaf [Ano02e]. Kiel [SSSV00, SHS06]. killed [JM03a]. Killer [GDdS^{+}07, FE06, HLH^{+}07, HBW^{+}05, HMY^{+}08, MAC^{+}07, MR06b, NEMF09]. killifish [WBC08]. killing [HGAS^{+}01]. Kimmerer [Ano07d]. kin [JY08]. kin-structured [JY08]. kinematics [Pen00]. Kinetic [GFL02]. Kinetics [FPBAT02, SAMN04, DCT^{+}04, FPC05, FSSW^{+}00, FBHB03, LW05, MSK00a, NW04, RSL09b, SSO^{+}01, SYPO09, WM06, WW03]. king [BCC^{+}04, CB01, HHF^{+}07, OP08, SAWG05, SAJG06, TMAN08, ASP^{+}03, BZLD02, FBJ00b, GCLC^{+}02, JBH00, JH01, Jen05, KKLL02, VSB04]. Kinne [Ano00d]. kittiwake [DFPG^{+}03, DBH^{+}02, FWH^{+}05, FEMW07, SIK^{+}02]. kittiwakes [Bar07, LWW^{+}01, SPH^{+}09]. kleptoparasitism [Iye02]. knife [MHBB05]. knife-edge [MHBB05]. knowledge [CBEL06, MWT^{+}06, RZWG08]. Kobbefjord [MRG08]. Kochzius [Ano02e]. Kodiak [MW08]. kokuboi [TK04]. Kong [Lam00a, Lam00b, CLL^{+}09, DDTQ05, DQW05, GXL^{+}08, JKQ02, LMCQ02, LYL^{+}07, MHY^{+}06, QDDP06, SE04, XYH^{+}09, YQC^{+}00, Yin02, Yin03]. Kongsfjorden [KKL07]. Korea [CNC00, JYS06, KKL^{+}03, PC02a]. Korsfjord [HB07b]. kraits [BBS07]. Kramer [KH07c]. Kraus [Ano05c]. kraussi [PBF07]. Krill [BPC^{+}02, SSS06, AM01, AM02b, AVNN05a, AVNN05b, ASP^{+}08, BLS06, BT04, CS05, CDW^{+}09, CTB07, FHM02, FMWGB08, HRQ^{+}02, HRR^{+}01, HLPHA^{+}01, HL00, JT08, KLHO02, KRF^{+}05, KCK^{+}06, KFJ^{+}07, LTV^{+}01, MRQM06, MABB03, MO05, MR01, Pea03, QR03, Rit02, STPA04, SN02, SWW02, STA^{+}05, SS02b, TDT^{+}03, TJEM02, TJM03, Tar03, TCRT^{+}07, TJSG03, WAW^{+}09, WCM08, WJB09]. krill-better [TJM03]. Kuroshio [TW04a, TW04b, CCL03, CCJT09, MOM^{+}02, ONKK03, WKH02]. Kuroshio-Oyashio [TW04a, TW04b]. Kuwait [AZKJAM06, HGAS^{+}01].

Kvamme [Ano03d]. Kyucho [KKTN05].

L [Ano00c, Ano00b, SH03]. L. [BDJ09, HW00]. labeled [CNC00, NHN^{+}06]. labelled [AW03a, KW05]. labelling [HMW^{+}00, MBM^{+}00]. labiata [PBS^{+}00, Pur07, PHS09]. labile [MSD05]. lability [DS09b]. Laboratory [GMZvdL07, HM03, WG02b, AMR00b, BLF06, BC01b, BPCvdM06, CWT^{+}01, CSvO09, DPH02, EEA01, FOFG09, FSP06, FSSW^{+}00, KB03, MCR03, MMA03a, NPVW06, PWL^{+}06, RGR^{+}01, RRIH^{+}02a, RRIH^{+}02b, SML02a, VN05, WCBD01, ZLAK04, SLT^{+}01]. laboratory-induced [ZLAK04]. laboratory-reared [FOFG09, MCR03]. Labrador [HHC00, LSHL03, SLM^{+}00, SSH^{+}00, WPTS08]. labrid [CDAM06, FB02a, FB02b, RM05]. Labridae [FWF02, TLC06]. Labroides [Gru00, GL02]. labroids [DFH05]. labyrinthiformis [CSMvE04]. Labyrinthula [MT09, RS02]. Lack [GRPE08, KS06, LHK05, TP00a, WWGR07, RM06]. lacking [Bel08a]. lactating [BPG08, BKH^{+}09, CWDG05, CW09, GBG00a, HHWT02, dBTO^{+}09]. lactation [BKLGN05, DPD^{+}03a, DPD^{+}03b, GLA06]. lactuca [MMMB^{+}05, TMM05]. laevigata [KSH^{+}09]. Lagoon [AK02, CWMD06, HHV06, KP01b, AE02, AM02a, AMT03a, BKDB00, Boo02, BFLÁ01, BDÁM04, CH07a, CM07, CMD^{+}09, CAS00, CABS02, DRG^{+}07, ERR06, HLH^{+}08, JDTB06, JCK03, KD00, KTNK07, KM01, LABR^{+}06, LPBD00, MHO07, MACS07, Maz04, MAT01, MSUK01, MHABGC04, MPB^{+}09, RATD07, RGCC^{+}08, Rue01, SKBP09, SVC^{+}01, SLHT05, SJ04, TSMC09, BSP^{+}08, BBS07, DVLH^{+}00, HvW08, SCB00]. lagoonal [KHT^{+}02, MFR01]. lagoons [AAPO00, ADR^{+}05, CRC02, EFL^{+}09, EF02, FPF01, JHJS04, KTK08b, Kav05, MSA07, MGTK00, SPC^{+}00, SCS^{+}07]. Lagrangian [FFYS^{+}07, NYK03, PFW^{+}02, SKGG03, SWD^{+}02]. Laguna [AG02b]. Lake [WK00, RC00]. lakes [BDR06, JHJ^{+}03, PBS^{+}05, SHB03]. Lamellibrachia [CHP^{+}05]. Laminaria [AL03b, BER^{+}03, GCD^{+}03, Kaw01, NFN03, PK01]. Laminariales [DZ06]. laminarian [FS07]. Lamna [WFG^{+}08]. Lampris [JBG00]. lance [FHN^{+}09]. Land [MHB^{+}09, HCV03a, LV05, MTV06, PRB^{+}08, SBC^{+}06, YI07]. land-derived [HCV03a, LV05, MTV06, SBC^{+}06]. Landing [Und04, BWO04]. Landsat [MAKMK09]. Landscape [CB05a, BBC^{+}09a, BHSR00, DW09, DE08, GBNÖ08, HFM^{+}02, HF05, Kne03, KvdMDP09, LAV04, PMP04, RB04, TGA02, TGM07, WK02, WBR07]. landscapes [GC02, JRC06, KDM05, Kne09, SESP03, WBJ^{+}07]. lane [LDR00]. lapidosa [MFY02, MYE03]. lapillum [MS06]. lapillus [SJ06, CH07d, DD07, HPC04, LMW^{+}01]. lapponica [ZS03]. lapse [OKP06, SK02, SWH^{+}04b]. Laptev [BF08b]. Large [CMRF05, EE00, GBdS^{+}05, GLB^{+}04, GAM02, GRG^{+}02, HCMR09, JFD09, LYMCG06, MGO^{+}07, Nod04, PM03, SAV09b, SAV^{+}09c, SDR01, TFSP09, WWWC02, AD00, BS06, BCPB^{+}01, BP03, BB08b, CPSC01, CDAM06, CBMB03, CB08c, CTHF01, DB07a, DUSJ^{+}09, DSDS04, EE06, GC01, GPB02, GMM03, HMBD01, HGKK06, HW00, KJL03, KJB^{+}06, LTG^{+}01, LGB03b, Law08, LH02, LAV04, LS04, MCFM02, MBG^{+}09, NZ08, NMDS00, OS01, PS07, PCMB06, PKA06, Pur03, RMA^{+}02, SDS^{+}03, SBJ02, SYVC00, SRS05, Tup07, VCA^{+}09, VDA^{+}03, Vli05, WKL^{+}03, WTS09]. Large-scale [EE00, GLB^{+}04, GAM02, HCMR09, JFD09, LYMCG06, MGO^{+}07, Nod04, PM03, SAV09b, SAV^{+}09c, TFSP09, WWWC02, BCPB^{+}01, BB08b, CB08c, DB07a, DSDS04, EE06, LTG^{+}01, LGB03b, Law08, LH02, OS01, RMA^{+}02, SRS05, VCA^{+}09, Vli05, WKL^{+}03, WTS09]. Large-sized [CMRF05]. larger [RSSA07, TDP^{+}00]. largest [FPF01, WLB08]. Larus [HW00, RRS^{+}09, SG05]. larva [IT06b, MK03b, TP01b]. larvacean [URFA06]. larvae [ACT09, AYSL09, ACR07, AM04, BBM03, Bel08a, BF01b, BDCvB^{+}09, BPvdM07, BK06a, BM07b, BPA^{+}08, BCLS06, BWO04, CSB03, Cal05, COPB06, CSTC01, CES^{+}03, CB08a, CDJ^{+}07, CLHT05, CHPS04, CGH^{+}03, CBM^{+}07, DSM01, DTV08, DTV09, DCBvBE09, DMV07, DBT02b, DCRT07, DS07, EM07a, EÖM08, EvBB^{+}09, FOFG09, FBJ00a, FB03a, Fis06, FFG03, GBR^{+}04, GZ02, GGHA02, GK03, GM05, GBD^{+}06, HMKR05, HNYI07, HHH02, HSRF02, HS03, HLFG^{+}07, Hue08a, HFMP03, HTZ^{+}07, HLH^{+}05, JS03, JRTJ04, KHvW02, KCB06b, Kru01, LTPL05, LCC07, Lec05, LCEC02, LCE03, Lei07, LHL^{+}07, LHH09, LB08b, LRC06, LC09b, LYMCG06, LDT07, MM06a, MK03b, MADPQ06, MFA03, MP00, MCM^{+}03, MMS09, MVLM^{+}06, MOP^{+}08, MBS03, MBS07, MGSJ08, MBB^{+}04, NE01, NSH^{+}08, ONKK03, PCC06, PTB03, PJJ^{+}05, PPL^{+}02, PSPD07, QRS00]. larvae [QMM02, RC04a, RBSP00, RE00a, SS00a, SMMM07, SRJ09b, SLFM03, SEVSP02, SK07, SB05a, SS09b, ST06b, SD00a, SPB09, TP07, TLLCT02, TCRT^{+}07, TJM00, TKN00, TT07a, TOBR07, Trn02, TYC00, VTY06, VCBH06, WPPS04, YBY^{+}09, YOM^{+}09, YCSS06, dSSC07, dSSC^{+}08, vdMM03, vdVBGW09]. Larval [ABF07, AKM09, ATLC03, AMT03b, AKG02, BP02a, BBA06b, Car00, DM01b, DBT02b, FSP06, GA03, KZ00, MM06b, MCE01a, MPK03, MFM09a, NKS03, PCCL05, Ros01, SK04, SK05, WEEBdN08, ANCC07, Arm00, BMWH08, BRG^{+}02, BUCC09, BGPL08, BDCvB^{+}09, BO00, BFV08, BNP07, BCC08, CTA^{+}03, CSTC01, CGCDA02, CWBS03, CB08a, CNW^{+}07b, CFA07, CBM^{+}07, DSL07, DQ05, DRTM09, DABC02, DB07b, DEV07, DHT^{+}08, DS07, EHWS07, EWSH08, EEWBdN08, FLK^{+}06, FM02, FJK^{+}07, Fis05, FSB07, FBJ00b, GJB04, HLQ02, HTW^{+}05, HSST01, HLFG^{+}07, HH00b, HH03b, HTWQ05, JRP01, JBH00, JMM^{+}07, JM07, JQ04, JQ05, JS01, KCB06b, KSR^{+}07, KLW^{+}09, Kuf01, LHQ03, LHQ05a, LHQ05b, LMCQ02, LS02, LC08, LZW^{+}08, LHB^{+}06, LLW09, MMV05, MR06a, MCR03, MCW^{+}07, MHO07, MCBT07, MC07, MMVB07, MK03b, MK04a, MK04b, MT05]. larval [MFY02, McI02, McI03, MP06, MABB03, MO05, MS05, MWTB00, MKK^{+}09, MW01, MBG^{+}08, MMS^{+}05, Mun07, NJPI03, NNLV06, NEG01, NPVW06, NK03, NM04, OTBR00, OvM02, OÁSG^{+}09, OA02, PGBA04, PL07, PD07, PBT^{+}09, PP00, PDHH06, PKLD09, PJJ^{+}05, PS06, PMR06, PCH^{+}05, RM05, RMM04, RGL^{+}00, RGF^{+}03, RCF09, RGCL^{+}05, RAS07, SRJ09a, STPA04, SSL00b, SS00b, SEH07a, Sha09a, Sha09b, SF02a, SS09c, SKGG03, SD00b, SLHR09, SMF^{+}02, SCM09b, SMT08, TAM03, TOK^{+}04, TAM04, TAO07, TSSQ06, VFM06, VW00, VJKF07, VCBH06, WGF09, WHS^{+}09, WT09, WM01, Win09, ZMP^{+}03, Zac05a, Zac05b, ZH02, vdVBGW09, vHGH01]. Lasaea [OJPM01]. laser [MDBG07, RCC01]. last [CLB^{+}09, CSPA09, GHL02]. lasting [AWS^{+}09]. Late [EBA^{+}03, ZGV00, ACR07, Arm00, CVÁM06, GGK^{+}07, GRG^{+}02, KNF^{+}03, LCB^{+}05, SAB^{+}05, SHFP^{+}03, Tyn04]. Late-Holocene [EBA^{+}03]. Latent [CNW^{+}07b]. Lateolabrax [ST07, SKO^{+}08]. Later [WHL^{+}08]. lateral [MMS^{+}05]. Laternula [ASP^{+}03, KTK08b, PTP01]. Lates [MC05c]. lathosterol [GWBdN09]. Laticauda [BBS07]. laticeps [GKAS08]. latirostris [CGM00]. latitude [AMB08, BK03, CC00, Fis06, FP08, GUL^{+}07, MAC^{+}07, NPN06, NTN08, RJS04, TTN01, WH03]. latitudes [ATLC03, JSMM07, LFSR06, TSSA03, TYC00]. Latitudinal [CAD00, CML08, LTFJ00, LBF^{+}02, RFN02, BA01, BHH09, BFE^{+}03, Hil04, KBP06, LBH^{+}04, LTMF01, RSE01, SGV^{+}02, UBRvO06]. Laurel [SRJ07a]. Lavalli [WA00]. law [FR08a]. lawn [FGS00]. Lawrence [BRB^{+}00, CSMBH08, CGL^{+}02, CS05, FLR05, LF08, LHK01, PDRT02, PJR^{+}03, PT04, RBHS06, WDB^{+}03, ZGV00, CDH^{+}09, DHRG05, DVBLS07, MRZJ05, MHCB08, MGH00, OA02, OSL07, QS08, RCF09, SSS06, XZLP09]. Lawton [WA00]. lay [SS04b, vDS05]. layer [ACM^{+}02, CC02, CMMN^{+}01, CMSS09, FLG02, FMZvW02, GGTKG09, GB06, HRK^{+}08, JN09, KHB00a, KGHW05, KW00b, MBBW08, MFP^{+}05, MLR02, PSE^{+}03, PSGP04, WG09, WWL04]. layer-interception [FMZvW02]. layers [DDS^{+}01, FK04, GYD04, HMY^{+}08, ID02, KTK^{+}08a, KTH^{+}04, Lei01, MAB^{+}03, MCD^{+}05, MD08b, PTK^{+}09, RDD^{+}02, RMPC08]. laying [VHMB09]. Lazarev [FVS^{+}08, PFW^{+}02]. Lea [Ano03e]. Leach [HM06a]. lead [SFT^{+}05]. lead-210 [SFT^{+}05]. leads [FTO^{+}09]. leaf [CDT^{+}01, ITN09, KDAK01, LRHV07, LD00, PD05, RC00]. leaf-clipping [KDAK01]. leaf-marking [KDAK01]. leakage [MTN03]. learn [LH06]. learning [PBL^{+}09]. least [GRIR06]. leatherback [Eck02, FFT^{+}07, GF06, GFB^{+}07, JSMM07, JMJ07, LLM^{+}08, MLGG03, MH06b, WKPS06, WBJ^{+}07]. leaves [BSvdH^{+}02, Enr05, LRHV07, MMBT03]. lecithotrophic [CTA^{+}03]. lee [MM06a, MM06b]. legacy [SAV^{+}09a]. legged [Bar07, DFPG^{+}03, DBH^{+}02, FEMW07, LWW^{+}01, RRS^{+}09, SIK^{+}02]. leidyi [BN04, DBP04, FKA^{+}06, HCS^{+}07a]. Leiostomus [CRCN07, MT05]. lemon [WGR07]. length [ABBF01, CLN00, GGFSP06, KCK^{+}06, MS03, SN02, SH06b, SFL^{+}09, UYNO01, WDT^{+}08a]. length-mass-density [ABBF01]. lengths [WAK^{+}09]. lenorostralum [YSS03]. leopard [CS09, HARC05]. lepadiformis [dCLLT^{+}02]. Lepeophtheirus [SBHE05, TSRR05, TWMW06]. lepetodrilid [Bat07]. Lepidochelys [MBH07, SAM^{+}06]. Lepidopsetta [RSSA07, RH08]. Lepidopus [KRP^{+}06]. Leptasterias [RHG04]. leptocephali [KAM^{+}06, MOM^{+}02, MM07b, MMA^{+}04]. leptocephalus [RC04a]. Leptogorgia [RCR03]. Leptonychotes [HHWT02, MFWD08]. leptoporus [LNJ07]. Lesion [FFHW07, Lir00]. lesions [SCBA05, WGH^{+}00]. Lessepsian [BTG^{+}08]. lesser [DGC^{+}05a, MCW^{+}07, SG05, WWHE04, vDvHT^{+}09]. lesson [MTLG09]. Lessonia [FMCB05]. lessoniana [SM00a]. lessons [ABD^{+}02]. lethal [AFPS08, FSW09, TPGQ07]. Leuca [CLO^{+}09]. leucas [BKLGN05, GRRH07, LMKG01]. Leuconidae [AY04a, AY04b]. leucopsarus [MKWO00]. leucorhoa [HM06a]. level [AP00a, BPB07, CAS05, DDGK^{+}08, DHT^{+}08, FR08b, HGK02, MSLP06, OTBR00, RHDC05, SMM^{+}04, SW03, TBB^{+}04, TC01b, VBMBC08a, WFG^{+}08, WWTH06]. levels [APM^{+}09, ARBF02, BGSBC02, BOM^{+}08, CCCH01, DCT^{+}04, GGHA02, Gor03, GEHZ07, HYLT03, HHB^{+}00, HBG^{+}03, JBSB01, JHR05, KPA^{+}09, KH00a, KKB^{+}06, LMW04, LGVF00, MSF^{+}07, MPK03, MR03, NNG03, NAB^{+}06, PBAD05, PR01, RGEB00, RNDF^{+}04, Rod03, RHH^{+}05, SJPC02b, SFT^{+}05, SRLA04, TFSP09, WLdG^{+}04, vDvTJD^{+}09]. Levu [MC08a]. lewini [DH06]. libellula [MHCB08]. library [NGM^{+}05]. lice [SBHE05, TWMW06]. Life [AY04a, AY04b, AMB08, FEFT08, HM05a, HM05b, IMSO09, JDCB09, KI01, LLRDT05, OHCFJ04, OTB^{+}00, PIY04, PJR^{+}03, RF08a, RBG04, SP03b, TSW^{+}04, TSW07, TD01b, WTS09, Air00, ATT06a, ABL08, BBC04, BDRM06, BSP03, BTF^{+}02, BTD^{+}04, CAF^{+}08, CKB^{+}09, CEO06, CCB^{+}09a, DC06, DL02, DC02, DG05, FRvdMF02, FGT^{+}07, FS07, FOH^{+}08, GNPB07, GNP07, Gim04, GB07, GKAS08, GV01, GLBF07, HWB^{+}09, HGS^{+}07, HLFG^{+}07, HPVL05, HRC^{+}07, JRP01, Kap09, KHK^{+}06, Kro05, LW00, MBHN07, MAM^{+}01, MMA^{+}09, MVH^{+}06, MW09, NGM^{+}05, NH03a, NH03b, PDM^{+}04, PMM^{+}00, RCL^{+}09, RTL05, RLFEF03, RFB07, RvdPHW04, SSN04, Sha00, SR04b, SP01c, SGCP06, SNGR00, TKK05, VAJ^{+}03, WA00, WDRH07, Wol07b, XWQQ05, vdVBW00]. Life-cycle [TSW^{+}04, TSW07, HPVL05]. Life-history [HM05a, HM05b, ATT06a, BDRM06, DL02, DG05, JRP01, RCL^{+}09, SSN04, Wol07b]. life-span [HM05a]. life-stages [FS07]. lifespan [WS08]. Lifetime [DRDM09, BDH01]. Light [Enr05, MFK07, OMT06, WWL04, BP04, BMSH02, BN01, BSJ02, BWK^{+}08, BOM^{+}08, BN02, BD01, CGP08, CLMR07, CLRM08, Con05, DDV^{+}08, FR06, HCV03a, IC02, JS01, JS02, KUN02, LWR^{+}02, LLL^{+}09, MCL07, MBLA01, MEB^{+}07, MND^{+}02, MS05, MSK^{+}00b, NPM01, PO09, PCL^{+}09, PCYC02, PUB^{+}01, PKR07, PSC05, PDB04, PHS09, RMWO09, RLBB04, RW07, RFPB^{+}04, RMH^{+}09, SMHL00, SLCB00, SA06, SMSJ04, TN07, TBB^{+}04, TDvdW^{+}01, Uth06, VB02, ZSKA01, dBKSV00]. light-driven [BD01]. Light-induced [WWL04]. light-mantled [PSC05]. light-saturated [BMSH02]. light-trap [MS05]. lighting [RSIS09]. lights [Pea03]. Ligia [TD01b]. lignin [KHB00b]. Ligurian [AM03, BBHT02]. like [BDJ09, CCM01, LBR04, PPHT00, SGK04, SSG^{+}00, SPBM05, SFPGA05, MGBB06]. likely [RJS04, SPM03]. lima [BCBQ01, LGQ^{+}00]. Limanda [MCW^{+}07, SF05]. Limfjorden [HMKR05, MR07b, SPT^{+}00]. limit [AMS01, AAHDS06, AVA^{+}09, LBPS03, MARM06, RS06, SSR^{+}09, WR00b]. Limitation [SCGC07, AW03b, ATLC03, BGPL08, BPvdM07, BSR05, CGP08, CNW^{+}07b, DBF^{+}02, DRB^{+}01, EE06, FIS04, FTK^{+}04, GFM^{+}02, GLBF07, HLR07, HN03, HCV03a, KOTL07, KB07, LBPH06, LUAIH03, MS06, MHY^{+}06, MSH^{+}03a, MPDS07, MFM09a, NNW04, NPM01, RGB02, RHGB02, SG06a, SHL^{+}03, SH07d, SMF^{+}02, SG01, TA07a, TDvdW^{+}01, WKPS06, Wol07b, XYH^{+}09, YQC^{+}00, YQW^{+}01]. limitations [DH07a]. Limited [BWSH00, JMdG^{+}06, JKK^{+}06, Kon00, MP00, RB08, BFLÁ01, CFA07, DKR09, JM07, MSC04, PWC^{+}09, SH06c, SJH^{+}04, TP07, ULL^{+}05]. limiting [Gar00, Hal04a, RBDK00]. Limits [DEV07, WPH01, BRAS03, BP06, DB07b, KS09, MLC08, PD07, PWM^{+}08, SHNV09, SMF00]. Limosa [ZS03]. Limpet [DJM07, AH07, BCU04, CCB06, CHW06, DEV07, FCP02b, JAA^{+}01, KM03a, LGD04, PB00b, PTP01, SSB02, Thi05, WDLM05, WG07, ZA07, ZSKA01]. limpets [AAF^{+}04, BCBAC01, Chi09, GGS09, MCU06, SBC04]. Limulus [BLT03, CRV03, JSN08]. Linckia [KSH^{+}09]. line [PLC^{+}02, dSHCR07, DFSS08]. lineage [THP^{+}05]. lineages [LC07a, OJPM01, SKW^{+}06]. Linear [TBM^{+}08, BLS04, DHZ02, RMAY04]. linearis [SD01a]. linearisation [Kir02b, Kir02c, Pet02a]. lined [MSP^{+}06]. lingcod [TAW^{+}09]. Link [COPB06, McI02, AG02b, EGJ07, GRP^{+}06, HNMC09, KPW07, WL05, Wie05, WLT^{+}03]. linkages [FPDS^{+}02, HS07, Kim02, RS06]. linked [ABF07, RHS01a]. Linking [DRTM09, FJK^{+}07, JPPW02, Ley05, MDBG07, NH03b, Pah05, PMP04, SCKC07, SSS^{+}05, Ven06, AHC^{+}06, EE00, PdNS06a, RFLC07, TL03, WB05a, WB05b, dlSPLV09, NH03a]. Links [RTB06, CMB^{+}05b, FH03, MNT09, NOO^{+}08, PDS^{+}07, PMH08, RWE04, Saf03, Thu07, VBFF08]. linum [BRBPL06]. Liocarcinus [RMAY04]. lion [CD08, DCA^{+}09, GT08, MMC^{+}08b, STV^{+}09, VACT^{+}08, WH08, WS06]. lionfish [AH08, KMWH04, WGV^{+}02]. lions [BRI^{+}07, BDGS07, CWDG05, CW09, HK04b, KG07, MW08, STAS06, WTR02, WWS^{+}05, DR00, DRB^{+}01, FHVMDS07, GMd^{+}02]. lip [PBFB08]. Lipid [CMB^{+}05a, DBOR09, DPGN08, DPGN09, FPDS^{+}02, LHK06, MNC^{+}02, PTD05, AP05a, AM03, AVNN05b, CBD04, CVÁM06, EBG^{+}00, GPB02, HNYI07, PDT09, SAP02, SNGR00, TPSW00, WA05, AVNN05a]. lipid-poor [AP05a]. lipid-rich [AP05a]. Lipids [AVNN05a, AVNN05b, DFPG^{+}03, BKR04, BSR05, CD03a, FHG^{+}03, LMS04, PGS^{+}00, Sew05, WBC01, vDSD02]. lipofuscin [HSJ08, JSH01, She08]. lipped [CYC^{+}09, GSL^{+}06, GXL^{+}08]. listerianum [MPK03]. lists [CW01]. lithodid [ATLC03]. Litopenaeus [WK02]. Litophyton [TAS04]. litter [PD05, VS08, WL05, WMCF08]. little [CCCD09, GPK^{+}04, Gil02b, HDRC^{+}08, KKW^{+}03, Mar04]. little-neck [Mar04]. littoral [BAOV06, DJM^{+}05, MR06a, TD01b]. littoralis [RBJB^{+}05]. littorea [DD07, ED02, OST07, Pet02b]. Littorina [DD07, ED02, OST07, PJ06a, Pet02b, PQRAC08, SP01c, WGM02]. littorinids [SDW07b]. Live [MLC08, AKRA02, HJ03, HKR^{+}08, PRvdL^{+}07]. live-bearing [HKR^{+}08]. lived [CHP^{+}05, TPG09]. lividus [FTP^{+}09, FB00, RM00, TRT04, TÁCTR06]. Living [GBC^{+}04, HSPH05, SRDG05, THM^{+}03, DRV^{+}05, HS05, MA02, QHD04, SR08b, SJS^{+}01, UÓ03]. load [BGS02a, GPM^{+}09, jMFVPL05]. loading [CLL^{+}09, HWL00, LHT06, ZHWCS08]. loads [CWT06, HCV03a, LV05, SFH00b]. Lobodon [SKE05]. lobster [ASA^{+}08, BHH07, BL09, DL03a, DGB^{+}02, EJKN03, EP08, FCC^{+}08, FMT09, GRPE08, GSM06, GTGM06, GB02b, GFN^{+}09, JWM09, JRTJ04, MJG08, NPLZ00, Niz07, OA02, PMS^{+}08, PWAT01, PE05, RT00a, RBSP00, TGGM05, WA00, WGF09]. lobsters [BLS03, BFLÁ01, CWBS03, CH05, GQR06, JPC00, LÁMBF09, MGC03, OWBD07, SJCM01]. Local [AED^{+}07, BVC^{+}02, Bar03, BHSR00, DAMJ08, HR02a, NCT^{+}01, UP07, Yap09, BEU^{+}07, BCS^{+}06, BMO^{+}04, CC04, CTB07, EM04, GSP^{+}02, HDDR08, HW01, JKES07, LS09a, LW06, McI02, MYNY08, NMDS00, OP09, PZB^{+}05, PJR^{+}03, PKS07, RMA^{+}02, RFN02, RC04b, SD01b, WAB^{+}02, Zam03b]. local-scale [CTB07]. localised [BB09, HCR01]. Localization [BDR04, LTMDB02]. located [dJTD07]. Location [HM06c, ITN^{+}06, KHA^{+}00, MWR07, WBZ^{+}09]. locations [BN09, HPW^{+}00, HHG^{+}07, JÅC^{+}01, KJRG09, LBvdH03, MSH08, Saf03]. loch [BPN^{+}09]. loci [BHH05, NWHH^{+}09]. Locomotion [KRCGC06]. locus [HFM^{+}04]. locusta [CNT^{+}04]. Lofoten [HMD^{+}07]. Loggerhead [BRGS^{+}07, BBM00, CLP^{+}06, CAF^{+}08, CPB04a, GBM01, HTM^{+}02, HGS^{+}02, HBG^{+}02, SBMAE06, SKP^{+}07]. loggers [PCS^{+}07, WDT^{+}08b]. logging [MH06b]. logistic [BRP07, EYH^{+}06, LRK09, YMHV02, HLR04]. Loire [AED^{+}03]. loliginid [SM00a, VAJ^{+}03]. Loligo [BHMR01, BGH06, BMCH09, IMS05, IS07, MH00, SS04a, VAJ^{+}03, WHS^{+}09]. Loliolus [JM01]. Lolliguncula [BMV02]. Long [BBG^{+}04, BI04, BMF06, BOR08, CJRT00, CS03c, Das03, DB07a, FMCB05, FAB^{+}07, GYLÖ06, GSM^{+}02, GTPC04, GB02b, KR04, KST^{+}08, LBS^{+}05, LC06, MYH07, MCP^{+}09, NMCT03, Pea00, PD01, Reb01, SR04a, SWG^{+}09, TBID09, YF04, YVB^{+}02, YMC05, AWS^{+}09, BBHT02, BPM^{+}06, BIR00, CR08, CHP^{+}05, CCB^{+}09b, DL01, FGM^{+}00, GLS^{+}00, HWC^{+}06, JM07, LLM^{+}08, Lin04, LPM^{+}06, MMC08a, MS08, MOPP05, MGL^{+}00, NB06b, ODA^{+}09, PAB^{+}02, RF06, RMZC07, SN02, TPG09, THP^{+}05, TM02b, TPS^{+}05, BG07b, GSW01b, GK07, LW02, TRDV00, TWYS03]. long-chain [ODA^{+}09, PAB^{+}02]. Long-distance [GB02b, LLM^{+}08, THP^{+}05]. long-finned [BBHT02]. long-lasting [AWS^{+}09]. long-lived [CHP^{+}05, TPG09]. Long-term [BBG^{+}04, BI04, BMF06, BOR08, CJRT00, CS03c, Das03, DB07a, FMCB05, FAB^{+}07, GYLÖ06, GSM^{+}02, GTPC04, KR04, KST^{+}08, LBS^{+}05, LC06, MYH07, MCP^{+}09, NMCT03, Pea00, PD01, Reb01, SWG^{+}09, TBID09, YF04, YVB^{+}02, YMC05, BPM^{+}06, BIR00, CR08, CCB^{+}09b, DL01, FGM^{+}00, GLS^{+}00, HWC^{+}06, Lin04, LPM^{+}06, MMC08a, MS08, MOPP05, NB06b, RF06, RMZC07, SN02, TM02b, TPS^{+}05]. longa [PJ08]. longevity [RCP^{+}05, UBM^{+}07]. longfin [BGH06, CC03b, JT05]. longicornis [BGS^{+}02b, CA06, GK08, JVJ09, MRZJ05, OBB00, PK05b, SV08]. longicyathus [LLCHG01]. longimana [Sot03]. longirostris [KN00]. longispina [VV04]. longitude [JMM^{+}09]. Longitudinal [Bur02, SSGB^{+}08]. longline [BDF09, CJFM09, CPB04a, GLP^{+}08, SAM^{+}06]. Longshore [GY00]. look [BGK09, MWTB00]. loon [SPD03]. loop [OW02, Thu07]. López [Ano00c, Ano00b]. López-Urrutia [Ano00c]. Lophelia [BY09, BHY09, DBOR09, JNFL04]. Lophopanopeus [HS05]. lophyropoda [BUCC09]. Lord [LPS08]. Loss [MCCV08, ADA00, BCPB^{+}01, CHLV07, DJM07, GC02, HCV03a, JKG^{+}05, MHB01a, RH06, RC00, RMZC07, VJ03]. losses [BFE^{+}03, Uth01]. lost [CRT^{+}08, MR03, MPDC03, Nug02, SS09b]. Lottia [KM03a]. Louisiana [LB08a, BUF02, CY06, JBA02]. ’low [KJ03, JS03, RBG01, TPSW00, WHRS05, AWOO02, AJ03, BGS^{+}02b, BOM^{+}08, CMJ00, DBP04, DSM01, FBHB03, FGS00, GDSB04, GL04b, GDS02, GB05b, HDCA06, JHR05, KRK09, KG00a, Lam00b, MSC04, MP06, MS04a, MJG08, NAB09, PS08, PAK^{+}00, Pea03, PPI04, RCC^{+}09, RPN^{+}01, SLGD00, SW03, SH07d, SJPC02b, SMF00, TSSA03, TPGQ07, WW06a, WPH01, WO05, vDvTJD^{+}09, Gar00]. low-chlorophyll [FBHB03, SMF00]. low-diversity [NAB09]. low-frequency [MS04a]. low-inflow [DSM01]. low-level [SW03]. low-predation [MJG08]. ’low-salinity [KJ03, TPGQ07]. low-shore [WW06a]. low-temperature [Lam00b, PPI04]. lower [AP05a, AQSP07, BBBSPH03, BCCM03, FTO^{+}09, HSHA06, tAHSHA07, PDRT02, PJR^{+}03, SMF00, WKK07, ZGV00]. lower-boundary [HSHA06, tAHSHA07]. Loxothylacus [BBSO03, RAA02]. LTER [QR03]. lubrica [YI03b]. Lucia [Nug02]. Luciella [JKS^{+}07]. luetkeana [BSED08, CWKE05]. lugworm [ODA^{+}09, SP02]. luminophore [MDBG07, OKP06, SWH^{+}04b]. Lunar [SP04, ZDA^{+}06, BBDW09, JC08, MYH07, dM01a]. lunge [GPS07]. Lurefjord [NH05]. Lutjanidae [LDR00, NDV^{+}00a, NSEL00, PCCL05, ZH02]. Lutjanus [Kri02, LDR00, LSS^{+}09, MW06, NHS^{+}08, WCF08, WPPS04, XW02]. Lützow [INH^{+}09]. Lützow-Holm [INH^{+}09]. luymesi [CHP^{+}05]. lyase [YM09]. Lyngbya [APP^{+}09, EB04a]. lysianassid [ITN^{+}06]. lysis [GM01, TUK09]. Lysmata [RZLS09]. lysosomes [IRM08]. Lytechinus [CL05, MY02, MY04, VHKW00].

M [Cai09, BES^{+}08, FLUF^{+}04, GG09, HH07a, HBN01]. m-2 [HH07a]. M. [BWH09, LRGdS00, MMGC06, NH03a, NH03b, QTB02, SD01a, VV04]. MAAs [SFPGA05]. macaroni [BC02, DGH08, GBW^{+}05, GBW^{+}07, TGT^{+}06]. maccoyii [LFF^{+}09]. Machine [HD05, CWR^{+}03, PBL^{+}09]. machine-learning [PBL^{+}09]. Mackenzie [JPMGS08, RMGL07]. Mackerel [RHDC05, BBG^{+}06, HK04b, NS02, RCF07, RCF09, ST06b]. Macoma [ABB02, BDH01, BD03, BDJ09, BPCvdM06, BPvdM07, BdVD^{+}01, DB07a, GFL02, HW02, Hid03, LLSL06, ÓUA05, SMHC03, Tal01]. Macquarie [GHT^{+}01, PJN01, PJN03]. Macro [CHH^{+}05, HNMC09, JNFL04, Zhu06]. Macro- [CHH^{+}05, JNFL04]. macro-algal [Zhu06]. macro-structure [HNMC09]. macroalga [BRBPL06, BAB04, CRZ04, DDHQ06, KF01, SDM06]. macroalgae [AIM^{+}05, APH^{+}04, BG07a, Bib07, BSJ02, BM05a, BMWH08, Cam01, CKO^{+}07, EF02, FKBB01, FBM08, FS07, GLS02, GLFU^{+}04, GKWK02, HKWM07, HBNA04, HKSW01, JM03b, KSI09, Kin01, KP01b, KWB^{+}06, LGQ^{+}00, Lev08, LWK^{+}07, LW00, MBBP08, MKKK08, NDB04, PVV^{+}06, RH03, SBB09, TBSJ^{+}04, TMM05, VWF^{+}01, VLK04, VL00, VK05, WLdG^{+}04, YLv^{+}05, dlSPLV09]. macroalgae-dominated [EF02]. Macroalgal [BJH04, CRH01, JS02, MB07a, PPSP07, TFA^{+}08, ARP09, ARB04, BM07a, BS07c, BH09b, CÅJ05, CvBGR09, GRCdLvB08, Hir06, HB08b, KES00, KM03b, LV02, LSG08, LPM^{+}06, MRN07, MH07, MMMB^{+}05, MC08b, MAG07b, NNLF09, PTLK04, PCC01, SF06, SPT^{+}00, SMH^{+}03, Tho04, WAM^{+}08, WP07, WPH01, WHB09, YLv^{+}05]. MacroBen [ASR^{+}09, EHM^{+}09, GLV^{+}09, RWB^{+}09, SAV09b, SAV^{+}09c, VCA^{+}09, WAA^{+}09]. Macrobenthic [NPN06, VMV^{+}08, YMHV02, AAPO00, CRZ04, DFT^{+}07, ENC^{+}04, FGS^{+}02, GGGVGG03, JMdG^{+}06, KvdMDP09, LTH^{+}04, LWMW02, MGD^{+}09, MLW03, RU02, RKJ01, SFW01, SEL02, VHT^{+}05]. macrobenthos [Ell01, GHV02, HMW^{+}00, LTL^{+}06b, MFR01, SR09, WAB^{+}02, YTK^{+}05, YTH^{+}05, YI07, vdWHF^{+}08]. macroborers [LCCTFO03]. macrochisma [Tom08]. macrocrustaceans [BVVM02]. Macrocystis [HH05b, HHW^{+}07, Hob00, SFR^{+}09]. macrodetritus [CM08]. macrodiversity [PAW00a]. macroecological [BDF^{+}04, CCL^{+}08]. Macroecology [DC02, HB07a, WAA^{+}09]. macrofauna [AW03a, BRP00, CTH^{+}03, DAMJ08, DM05b, FH02, FR08b, GGM^{+}00, HLVU00, HAC06a, HTL^{+}02, HJK01, HGMH06, ILL06, KW05, KGHL^{+}01, LZM^{+}03, MAE03, PBF07, RF08b, SG01, THN^{+}03, VMV^{+}09, VFP^{+}07, YAF^{+}07, YH02]. macrofauna-mediated [VMV^{+}09]. Macrofaunal [LG08b, SRL^{+}02, WM06, APN06, BS08b, CMD07, DKFV05, EYH^{+}06, EHM^{+}09, GHS05, HFM^{+}02, JFR02, Kel02, KSWP04, Las03, LJM^{+}00, LTH^{+}06, MLW^{+}02, NLG05, NTH^{+}02, RS03, SW08c, UH03]. Macrofouling [QRS00, DDTQ05]. macroinfauna [AWA^{+}05, GY00, HCH09]. macroinvertebrate [CB05c, DL03b, Jos09, KCP^{+}07, KGRW06, VH03]. macroinvertebrates [ABHE09, FFA03, GBSQ09, JBA02]. macromolecules [HDQ04]. Macronectes [GSCW00, GSSR02]. Macronutrient [MHY^{+}06]. macrophyte [BAOV06, BWSH00, BSvdH^{+}02, JH01, KRVB05, TGM07]. Macrophytes [CNF09, CVF07, CHA06, ERR06, PvDMB08]. Macrosetella [EC07a]. macrotidal [BAE^{+}08, CC00, MBR^{+}05, SAT^{+}03, TGM^{+}03]. macrourid [CM02]. macrourids [DPGN09]. macrozoobenthic [BSR09, Vii07]. Macrozoobenthos [DDL^{+}03, DD04, KTNK07, KKL^{+}03].

Macrozooplankton [DHRG05, SHFP^{+}03]. mactroides [LBD00]. maculatus [BS07b, HS08a]. Madagascar [Bar02a, MAOP09]. made [BF03, SWD^{+}02]. Madeleine [RATD07]. Madracis [DBT^{+}02a, LG06, VB02]. Madre [AG02b]. Madre/Baffin [AG02b]. Madrepora [HAMW09]. maenas [AQSP07, BH01, GD07, HJH^{+}02, JMA02, LTMDB02, LBCS05, MHR03, Mok04, QMM02, SvNvdMvdV09, Smi04, SRLA04]. maerl [HSWG^{+}06, KMHS04, MCCT07a, MCCT07b]. Magdalen [GSM06, MGH00, OA02]. Magdalena [GGMGR07]. Magellan [MGHA05]. Magellanic [BR09, FHB^{+}02, WZ04]. magellanicus [Har06b, LCHG03, MJS09, WBDR06]. magister [HE08, HMA06, SO01b, THM^{+}03]. magnesium [MT05]. Magnitude [Koe05, KWAT08, MAT01]. main [BDEM01, FD02, PLC^{+}02]. Maine [HST^{+}09, JDCB09, LAK07, ML06b, ML07, AL03b, DCC^{+}03, HRC^{+}07, LB05a, MD08a, MB05a, NSBK05, PRM^{+}09b, PRM^{+}09a]. Maine [?]GeorgesLapolla:2005:HDD. mainland [SHZ^{+}05]. maintain [BDLB02, IC02]. maintained [MB00b, RNDF^{+}04]. maintaining [BBM02b]. maintains [BCT^{+}08]. Maintenance [DH09a, JKQ02, CD03a, CMSS09, Con05, HKR^{+}08, KWAT08, PW07, WAK^{+}04, Zam02]. Maja [SMFP08]. Majidae [RGVC07]. Major [dGvdBvO^{+}08, Bai00, DJG^{+}07, DLD^{+}00, EM05, EPM06, FSO08, MBW03, OGL09b, PI07b, RCAK03, SS07, VJ03, WHRS05, YYK^{+}08, YTK^{+}05, WH09]. majority [WT06]. majuscula [EB04a]. Makaira [GLP^{+}08]. make [CWR^{+}03, MB04]. Malathion [WS09]. malayensis [TCW^{+}08]. Male [KFJ^{+}07, LGRB^{+}08, Lus03, BM05b, CS02a, FKI^{+}09, HWB^{+}09, JHL04, KWWH02, LBC07, MCM05, PBAD05, RBMB00, SAWG05, SG06a, SAJG06, SRLA04]. male-only [SG06a]. male-type [MCM05]. males [CGM00]. Mallotus [PWFC08, RFCM03, WJDA06]. mammal [BT09, BBDW09, BTB^{+}02, EOL^{+}05, KWTP06, KAAH05, Tor09]. mammals [CH03, DLLB03, HRB^{+}06, Hen09, JdSM09, LHK01, MWT^{+}06, GNAD^{+}04]. Man [OHN01, BF03]. man-made [BF03]. managed [HFH09, PH05, ROCB09]. Management [VCSL^{+}09, WWMF06, ARP02, AFPS08, BRGS^{+}07, BP08, BA04, BS04b, BS05b, CV06, Cos09, DFJSJ^{+}05, Dis07, EFL^{+}09, EB04a, FEO07, HF09, JWM09, Kap06, LNV^{+}03, MNT09, OW02, PCMB06, PTSJ06, RHDC05, SMCE07, WH08, WGCM^{+}03, FC00]. Managing [Hou09, PPC09]. manatees [RW06a]. manganese [MT05]. Mangrove [BB02b, APA07, ATTX00, ATR^{+}08, AHDT^{+}09, BBBSPH03, BKK^{+}05, BDGK^{+}06, BMSD00, BMO^{+}04, CDT^{+}01, DCD^{+}03, DDGK^{+}08, ENC^{+}04, EB08, FS08, GF02, HO02, HCCY02, ITN09, KW02a, KWD05, KKH07, KFU^{+}08, LJ06, LHT^{+}02, Ley05, LSS^{+}09, MT00, MdL06, MMBT03, MM08, PDQ01, PD05, ROCB09, Ros01, SM00b, She05, SR04b, SH05, SR01, SKH06, TVD02, USG^{+}08, VHH^{+}02, VNdG^{+}06, WL05, dlMPN^{+}03]. mangrove-seagrass-coral [APA07]. mangrove-seagrass-reef [dlMPN^{+}03]. mangrove/intertidal [MT00]. Mangroves [FS06, GDKN08, BMC^{+}09, DGVK^{+}04, DL01, DGC^{+}05b, DVNvdV07, DE08, GR07b, Hal04a, HJ05, KM01, NKK^{+}01, NvdV02, NRvdV^{+}02, NvdV04a, NvdV04b, NHS^{+}08, PMH07, PCHM07, SEVSP02]. mangroves/seagrass [NKK^{+}01]. Mangubhai [BG09]. Manicina [RZEACGF00]. Manila [DSS^{+}09, MDG04]. manipulated [SG06a]. manipulating [ZSR03]. manipulation [BVMBC07, LGB03b, MK04b, RDC06]. manipulations [JFD09, MK03c]. manipulative [BDW^{+}00, HMW^{+}00, HS07]. mantle [BC06a]. mantled [PSC05]. Mantua [Ano10a]. manuscripts [Kin04]. Maorithyas [FKM^{+}01]. maorum [DPSD08b]. Mapping [KWTP06, LWK^{+}07, RC00, ZIUG07, CG05, KWP07, KvKB06, KTF^{+}01, RT00a]. Mar [tAHSHA07, PGS^{+}00]. MarBEF [SAV^{+}09a]. marbled [BB03b, BPB07, MPR02]. March [WG01]. Marennes [LNPB04, PMG^{+}06]. Marenzelleria [BBRJ04, NWÓ03]. Margalef [Est04, SSLM01]. margaritifera [LPBD00, YLK00]. Margin [SPRK02, AMB08, CSP^{+}03, DFR08, HSPH05, RHS01a, SRL^{+}02, SSHW^{+}07, SMH02]. Marginal [TRH^{+}06, AVA^{+}09, BGS02a, BCC^{+}04, SHFP^{+}03]. marginatus [HDP^{+}06, KKG06, KHG06]. Margins [BM07c, DBG^{+}09, SGV^{+}02]. Maria [CLO^{+}09]. mariculture [LBBM06]. marilina [KTK08b]. marina [Abd03a, Abd07, AW03b, BRAS03, BN01, BN02, CBA02, DRV^{+}05, EWR08, FSO08, HR02a, HCV03a, HCA06, HB08b, JBS08, JKG^{+}05, JKS^{+}01, JYS06, JNMB08, KKIV01, LPK^{+}07, MT09, MGM05, MF03, MSTS^{+}06, NSBK05, ODA^{+}09, OFHM03, PZ07, RS02, RH06, RPO00, RS07a, SASR03, SDK^{+}02, SH06b, SMB^{+}03, SP02, SCD^{+}09, TA03, TJA08, WL06, WPG08, WKOE09, WPB^{+}08b, ZSKA01, vKH00]. marinas [BCO07]. Marine [BDF09, CDS07, CH06a, CLC04, CABS02, DKA^{+}06, DLLB03, DR01, DWO^{+}07, EP00b, FC00, Gim04, HWL00, KAAH05, Kin05b, KVH02, LDK^{+}06, LHK01, LPS08, MHO07, NT07, OT08, PMS^{+}08, Paw00b, PAC^{+}00, PFT^{+}08, RRDHG08, RWP07, SBJ06, SRP^{+}06, SvAHH09b, SvAHH09a, SHMN07, SB09, Syd09, TP00b, AG02a, AFHS08, AFHS09, AWOO02, AHC^{+}06, AMP^{+}02, ABHE09, ACM^{+}02, AMB08, AL03a, Ano07g, Ark09, ASC06, BT09, BBM^{+}02a, BPC^{+}02, Bar06, Bas06, BMdNP09, BBC^{+}00, Bea04, BI04, BL04, BDF^{+}04, BLS04, BCBVM05, BBDW09, BJ08, BVH02, BDLB02, BCM^{+}09, BdBK^{+}00, BRSM07, BLDG01, BPS^{+}03, Bow05, BCPB06, BTB^{+}02, BSP03, BL07, Bra05, BRF03, BGS^{+}02b, BKR04, BRB^{+}00, BS04b, BS05b, BVHG08a, BVHG08b, BWvdE^{+}06, BH04, BTF^{+}06, CGH08, Cam01, CH03, CH06b, CLS^{+}01, CFG06, CWCP^{+}06]. marine [CR02, COB^{+}03, CK04, CH07c, CAMC07, CSO08, CLP08, CLHT05, CES^{+}09, Cla08, CWM^{+}09, CAS00, CBD^{+}05, CBEL06, CV06, Cos09, CM08, CH05, CVT09, CB05c, DSPHV04, DPS^{+}03, DMA03, DTD03, DWB04, DRV^{+}05, DHRG05, DBD^{+}05, DSDS04, DdNEvdZ05, DWM04, DB07b, DGB^{+}02, ESS09, EP08, EOL^{+}05, ESQ03, EB05, ERR06, EAF^{+}05, FPC05, FJV^{+}02, FR06, FWF02, FSSW^{+}00, FF04, FHG^{+}03, FAFR03, FKBB01, FBKB03, FVB^{+}09, FFR^{+}03, FP08, FTBC05, FDM^{+}09, FMT09, FBM07, FSJ03, GSM05, GJT09, GMP07, GL04a, GAB^{+}03, Gim09, Gla01, GB03, GRRH07, GK08, GCHA05, GQR06, GAAB^{+}08, GSS09c, GCSSJV07, Goo09, GNAD^{+}04, GNHB^{+}05, GHS05, GKM02, GS05, GWO02, Gra02, GG09, GS07, GTSS03, HRB^{+}06, HR02a, HLR07, HLQ02, HKM01, HBG^{+}03, HBRJ06, HHD^{+}09, HBdBV01, Hen09, HGA07, Hil04, Hin02]. marine [HK02b, HVH^{+}07, HP07, HOT08, HGG^{+}05, HBK^{+}05, ISBP06, ISY07, ISSK03, JWM09, JTVC08, JBS^{+}09, JSK^{+}07, JM07, JS03, JB07b, JdSM09, JKQ02, JHH01, JFR02, KPM00, KP01a, KF02, KSS^{+}08, Kap06, Kap09, KMP^{+}06, KWTP06, KAR04, Kin01, KBPB07, KHWH03, KRVB05, KOC^{+}07, KMB^{+}08, KH07b, KM03b, Kru01, KJRG09, Lam00a, LHQ03, LTMF01, LBF^{+}02, LMCQ02, LSH07b, Law08, LLY^{+}01, LHK06, LHL^{+}07, LPK^{+}03, LH08, LHGC^{+}09, LNH00, LTNH00, Ley05, LP07, LWH08, Lid08, LAV04, LIT05, Lin02, LP01b, LWH06, LTL06a, LTH^{+}04, LHT06, LTL^{+}06b, LW05, Lóp07, LWMW02, LBR^{+}09, LELC02, LTCP03, LHK05, LW04, LBCS05, Mar01, MM06b, MWT^{+}06, MGL^{+}07, MCS09, MC01a, MHCB08, MSK00a, MVBGM04, MF08, MG05a, MM03, MLC08, MBH07, MSY02, MOPP05, MS04a, Mil07b, MA02, MHDS09, MKH^{+}02, MHK^{+}04, Nae06]. marine [NWHC03, NZ08, NZM^{+}08b, NMVH06, NLK00, NR00, NRH^{+}07, OAS01, OTB^{+}00, OJ02, PL00, Paf06, PRB^{+}08, PJ06a, PDO01, PPC09, PLCG^{+}05, PBH^{+}03, PL07, PH03a, PH03b, PBT^{+}09, PGK^{+}08, PPB09b, PKM03, PHP^{+}03b, PSS^{+}07, PHS^{+}07, PRL^{+}09, PBTB08, PPJ^{+}05, PQRAC08, PSPD07, Pri02, QS08, RMC^{+}05, RMM04, RBDK00, Rij03, RGEB00, RBRSP00, RB08, RAH^{+}02, RDG^{+}07, RGD^{+}08, Roy02, RAM03, RSA05, RC05, RH04, SM08b, SFW01, SJR00, SSG^{+}00, SN04, SBSPS04, SF07b, Sha00, SB03a, SJ08, SR03, SKGG03, SWL07, SS06, SMB^{+}03, SO02, SAF^{+}09, SPRK02, SSS^{+}05, STH^{+}06, SDRH01, STAS06, SN09, SHM02, SRS04, Sta08, SNP03, SFT^{+}05, SWG^{+}09, SJA^{+}09, SRC00, Sto03, SP00, SZW^{+}07, SSH09, SLEH01, SRJ07a, SAH^{+}05, TMAN08, TMFB06, TdMF^{+}08, TFSP09, TVG00, TEB^{+}07, TR05, TMJ^{+}05]. marine [TKTJ07, TG03, THF^{+}01a, THF^{+}01b, TLTB03, Tya09, UÓ03, UP07, UYNO01, VRJ02, VWF^{+}01, Ven06, VACT^{+}08, VF09, VBFF08, WKL^{+}03, WT06, WG00, WKYL00, Wan02b, WW03, WW06b, WTS09, WT09, WHB^{+}06, WLZ05, WLJM05, WB05a, WB05b, WEEBdN08, WR07, WDSA09, WGS^{+}02, WS06, WPD02, XW01, XW02, XW03, YYK^{+}08, ZRG00, ZSR03, ZWYL01, ZK01, dM01a, dZCVC01, vDvTJD^{+}09, vTWMG08, HSTS^{+}09, MHP07, VB02]. marine-phytoplankton [BLS04]. marine-terrestrial [BJ08]. marinus [DJM^{+}05, EvdKB^{+}08, HGG^{+}05, HW00, LNV^{+}03, LP01a, MCW^{+}07, WWHE04, vDvHT^{+}09]. Marion [dBTO^{+}09]. maritima [KLN06, MRP09, VV04]. Maritime [CH03, SZL^{+}05]. mark [GAMB05]. mark-recapture [GAMB05]. Marked [HCR01]. marker [FH02, FON03, GMR^{+}02, MLR02, SM08b, ZER^{+}01]. markers [AVNN05a, AVNN05b, BHMR01, DK06, KZ00, PEdlP^{+}02, Reu01, SPI^{+}07, SHSF08]. markers-old [Reu01]. market [WHS^{+}09]. marking [KDAK01]. marlin [GLP^{+}08, HLH^{+}05]. marmorata [GMA^{+}09, SICT03, SSM^{+}01]. marmoratus [MPR02]. marsh [AHBM^{+}07, BA04, CF02, CLSM09, DT08, FCH^{+}08, GPHC09, GK02, Hal00, HH03a, HH05a, HGK02, HM05c, HFA09, KCP^{+}07, KLN06, Kne09, KWL^{+}05, KMK01, LB08a, LSH00, MD08a, NA04a, NMA00, PMH08, PAMJ03, RGCL^{+}07, RZ00, RM09b, SZZ08, SA03b, SGCC06, SVG^{+}05, TC01b, TC01a, TACM01, WWAC00, WK02, WHL^{+}06]. marsh-aquifer [TACM01]. marsh-pond [LSH00]. marshes [AED^{+}07, BSvdH^{+}02, BCB^{+}08b, CH06c, COK02, CWT06, CS03c, DPD^{+}08, GBSQ09, HS07, Kne03, MD08a, MCC^{+}00, MJG01, Min02, MAWH03, MA07, MSP^{+}06, OJ04, PAMJ03, QFJ^{+}07, RRRM08, VGB^{+}05]. martensii [TKN00]. Marthasterias [VTY06]. martia [MNC^{+}02, MNC^{+}02]. Mártir [GES^{+}06]. Maryland [KS04b, LLW09]. MAS [COB^{+}03]. Masan [PC02a]. masanensis [JKS^{+}07]. masked [WLB08]. masking [CES^{+}09, RM00, RGVC07]. masks [OM04, WWDH09]. Mass [GPSH01, TCT^{+}00, AMR00a, ABBF01, ACPS09, BPT07, CLR^{+}06, CCB^{+}09b, DJ04, GF06, Har08, HSRF02, HDG^{+}08, JWM02, JS08a, LCD^{+}05, LSG05b, MARM06, ÖBFO01, RWA^{+}02, RC00, SHMN07, VWH00, WHM^{+}01, WTH04, ZAH03]. mass-spawning [WTH04]. Massachusetts [CT01, HGT01, MTV06, PCBL05, TBL^{+}09, JBT^{+}07, TDP^{+}00, TBH06]. masses [CBAd00, FPB02, PGR^{+}01, WKMK06]. massive [GLB^{+}03, HSP^{+}08, JM03a, TLLKC04]. mat [BM06, WK00]. match [BPCvdM06, OP09]. match/mismatch [BPCvdM06]. Matches [Hip08]. mate [BK05a, BK05b, KBT05, RB05b, TSRR05]. mate-guarding [TSRR05]. material [BSL02, JPMGS08, KT06, KKKK03, OSA05, Pas00, SRGS01, WBSL00, WBP^{+}06]. materials [CGL^{+}02, DCD^{+}03, KWKF03]. Maternal [GM05, GC07, SBF08, TKK05, BLD^{+}05, BDG^{+}04, BJ08, CCK^{+}09, ITE^{+}04, MK04b, MGC^{+}09, STAS06, UP07]. Mating [BM05b, CK09a, AM06, BMCH09, DRDD09, GK08, HM05b, KWWH02, MPR02, SAWG05, SRC00, Thi03]. matrices [ABY00, HD05]. matrix [NRC07, TTM^{+}05]. matrotrophy [OGL09b]. mats [AZKJAM06, ARB04, BRBPL06, LPM^{+}06, ÓUA05, RWBS03, SAC^{+}00]. matter [BD09, BIG^{+}06, BMSD00, BMO^{+}04, BEE06, CRS^{+}07, CSMK09, CC00, CRCV04, DDF01, DSPHV04, DMH01, DDS^{+}06, EKBM05, FD00, GHS05, GTSS03, GGM^{+}00, HRR^{+}01, HBT^{+}08, HCCY02, IAT^{+}07, KPKD06, KCP^{+}07, KHD07, KTK^{+}02, LLJ05, LCB^{+}05, MGA^{+}01, MT00, MdL06, MSD05, MLM^{+}02, PRB^{+}08, PTD05, PDFD09, RF08a, RWR^{+}00, RMKR00, RNDF^{+}04, RE00b, SR09, Sar06, SAT^{+}03, SCD^{+}09, SZBZ04, SNCP04, SBC^{+}06, SES01, TRDV00, TDRM08, TNSM00a, URFA06, VK02, VGS^{+}01, WAM^{+}08, WRJ^{+}05, WMW^{+}08, YYK^{+}08, vOSF^{+}09]. matters [BWPB04, SFL05]. matthewsi [DWL09]. Maturation [Tho07, DH07a, FVAB00, GRB^{+}03, GHR^{+}07, HM05b, Kra07, KdLCM09, MB07d, MM07a, MKR07, VMCG09, Wri07]. mature [BM02b, WFC03]. maturing [CGM00]. Maturity [EH04, FVAB00, GL04b, JSWH03, MF07, NL05, RMH07]. Maui [SRS05]. mauretanicus [AO02]. Mauritiu [vWNYS05]. Mauritius [MMMPB05]. Maurolicus [LC02]. ‘Maury [RJS04]. Mawson [CKF^{+}03]. maxima [HHKH05, BSK06, EdPP^{+}00, KAMR00, SP08b, YLK00, ZAZL08]. Maximal [HS02]. maximisation [MB04]. maximising [GML02]. Maximizing [BDH01]. maximum [DBH^{+}02, KW00b, LF08, RHS01b, SD00b, RHG05]. maximus [BDR04, NWJ^{+}08, PMT^{+}02]. may [AL08, FHVMDS07, FKSW09, FKBB01, GK00a, Gou06, MP05, MCE^{+}01b, OST07, RFLC07, SH07a, SS09c, VSL^{+}04]. Mayotte [TTD^{+}06]. Mazzaella [SD01a, TG03]. McClanahan [vWNYS05]. Meadow [MDSP05, BD09, BBC^{+}09b, CPKV00, CLMR07, DAM^{+}08, FTP^{+}09, GSC^{+}01, IOB04, PFT^{+}08, RS07a, SBB09, TTR05a, VK02, WBN^{+}00]. Meadow-scale [MDSP05]. Meadows [CRS^{+}07, BWSH00, CPG^{+}09, GCSSJV07, HHO03, HCA06, JH06, MSTS^{+}06, MGO^{+}07, PDHH06, PTAR07, SKBP09, VHKW00, VLK04, VS08]. mean [BVMBC07, PPB^{+}09a, VBMBC08b, WDT^{+}08a]. means [AL03a, SGK04]. measure [Kam00, KS00b, LPS08, MUP06, NPTW05, She08]. measured [AMG^{+}07, BRJ^{+}03, FE07, GCD^{+}03, HH05a, KGB^{+}01, KKI^{+}06, LG01, NRM05, ORPND01, TNSM00a]. Measurement [MRHP01, AG00, BY09, GWF07, HHAH00, KUN02, LCC01, Osb07, PCJ06, Pet04, Rii01b, Rii04a, TRPNE06]. measurements [Bay04, BLF06, BN02, Cra01, DLLB03, DL01, FK04, GF04, HCBP05, HAA00, JPC^{+}06, KKP08, LALB08, MRLK06, NRPS^{+}01, PBS^{+}04, PTAR07, Rii01a, Rii01c, TRH^{+}06, Wid01]. measures [AFFH04, CBL^{+}09, Mar02, PPB09b, PVV^{+}06]. Measuring [AHV01, FFR^{+}03, IP01, EZTZ02, MD07, Pen00, RSA05, SHW09, UB02, WGS^{+}02]. Mechanical [FSW02, Wol07b, Jak01, Kaw01, Mil07a, NP05]. mechanics [SWH04a]. Mechanism [CTLR03, DBCB06, FI09, FK00, Lev08, MBBP08, PH01a, RCCK09, RN06, SL05d, SRJ07b, TJ02, YYK^{+}08, ZSKA01]. Mechanisms [KCB06b, KCA^{+}08, ODF09, SLFM03, SM09, WH00, ACM^{+}02, APM^{+}01, ACH07, Ano07g, ALOT01, BDDL03, BGD^{+}03, BBC08, CTP02, CD03c, DM05b, DBCB05, GD07, HTW^{+}05, Hen07, KNKT01, KBSWG09, MM06a, RST04, SN06, Sha09b, SO02, TAM04, TLLCT02, TNHL02, WBDR06]. mechanistic [AB04]. mechanosensory [FSW02]. mediate [FD04, KHH08, MC08b, WL05]. mediated [Ark09, DKR09, DdNS06, FK08, Gim04, GEC08, GBR01, Har08, KF02, KPW03, LGS07, MVD^{+}08a, OAGV07, SG09, SGCP06, VMV^{+}09]. mediates [KWHD08, WTS09]. Mediating [MACS07, GS09]. mediation [AMT03a]. Mediterranean [AHW03, BPVL03, BCVC07, CM01, CMMN^{+}01, CLR^{+}06, GBdS^{+}05, GPSH01, GHL02, GWPRC^{+}07, GGTKG09, KSSH02, LCD^{+}05, MGA^{+}01, ME01, OEDSJ02, OSP01, PSRB00, RGEB00, RBG01, RMRAFP08, SMA^{+}03, TLIL^{+}09, VLP^{+}04, dJTD07, ADA00, ASA^{+}03, ARBF02, AO02, ABD^{+}02, AVE^{+}06, ADR^{+}05, BCL03, BGLL09, BGM09, BCABC00, BCBAC01, BG07a, BBC^{+}09b, BFB^{+}09, BTB^{+}09, BGRD00, BCB^{+}08b, BTF^{+}06, CSA02, Car00, CPA^{+}03, CLO^{+}09, CM02, CC02, CLP^{+}06, CAF^{+}08, COPB06, CBS^{+}08, CSP^{+}03, DGM02, DBG^{+}09, DSPHV04, DUW08, DRG^{+}07, DSJBS^{+}02, DFCH05, DBN^{+}09, DR00, DRB^{+}01, DGC^{+}05a, DAG^{+}00, DAA00, DVLH^{+}00, EDKT06, FC04, FC08, FHVMDS07, FIRL02, FGA^{+}04, FHDA^{+}07, GDG02, GBSQ09, GLA06, GAZ02, GAAB^{+}08, GCSSJV07, GMd^{+}02, GTFB03, GBC^{+}04, GS07, Her06, JCMR^{+}03, JTB^{+}03, KTSR02, KAPTK02, KRVB05, LP03, LLRDT05]. Mediterranean [MR06a, Mal02, MHP03, MUT04, MJG^{+}00, MAA^{+}07, MG09b, MG05b, Ned04, OMP^{+}05, OGL^{+}09a, PDM^{+}03, PP02, PCC01, PG00, PMNP^{+}01, PPCJ^{+}06, RRS^{+}09, RLBB04, RCC^{+}04, RC00, RFG04, RBGM06, RMAY04, RR01, SSR^{+}09, SCMB03, SDD^{+}07, SYVC00, SSL^{+}07, TBCB^{+}05, TAFBC07, TRT04, TBC^{+}00, VGS^{+}01, VAB09, VD00, VGMC01, VAJ^{+}03, VMM^{+}04, dCLLT^{+}02]. medium [Som00]. medium-sized [Som00]. MedOBIS [AVE^{+}06]. medusae [CCK06, GB01, RT05]. medusarum [TT06]. MEERC [CS01]. meet [Lei07, PKM03]. meets [DWO^{+}07, JMA02]. Mega [SG02a, SWL07]. Mega-epibenthic [SG02a]. mega-herbivore [SWL07]. megabenthic [CCR08]. megabenthos [SSHW^{+}07]. megafauna [BLHB04, JBSB01, JNFL04, PNK^{+}02]. megafaunal [HCV03b, HAA00, JHB06]. megalopae [AQSP07, FCP02a, FCI^{+}04, HS05, HE08, LWJC04, MS04b, MH06a, OJ04, O’C07, ODF09, PDQ01, vMRO03]. megalopal [QAA^{+}06]. Megalops [CT06a, SCIT09]. Meganyctiphanes [AM03, KLHO02, LTV^{+}01, PZB^{+}05, STPA04, SS02b, Tar03, TJSG03, Tor01]. Megaptera [CH00b, FHN^{+}09, HFT^{+}09, HYR^{+}09, ZAHJ^{+}06]. meiobenthic [BBM^{+}02a, GZG^{+}07]. Meiobenthos [VMdBV06, FSV^{+}08]. meiofauna [Bur01, DF02, GMd^{+}02, GGTKG09, GF02, ID02, NGMT01, SGP07]. Meiofaunal [DGM02, DF02, CF02, DQ05, JH01, LPD09, ÓUA05, VBV^{+}00]. Mejillones [GEM02]. Mekong [ATTX00, AHDT^{+}09]. Melampus [PD05]. melanization [PHP^{+}03a]. Melanogrammus [BW01, Cal05, LB05a, MGSJ08, PBCB03, RTL05]. melanostictus [OKAM00]. melapterus [GL02]. Melita [KSS^{+}05]. membranacea [IH02, Pra05, SM07b, SM08a, SG09]. membrane [IRM08]. Membranipora [IH02, Pra05, SM07b, SM08a, SG09]. membras [VMCG09]. Memoriam [Cos03, Est04, Pud05, WFHS08]. Menai [TSJ05]. Mendelian [BHH05]. menhaden [HD06a]. Menidia [CWM^{+}09, DMC09, GV01]. Menorca [CLR^{+}06]. menstrualis [PCT00]. menziesii [BMG02]. Meoma [DRDD09]. MEPS [Ano00d, Bel04, Kin02, Kin03, Kin04, Kin05b, Wie03]. Mer [MGA^{+}01]. Mercenaria [BM07b, CCL^{+}04, GLCL04, MW04a, NTGK09]. Mercury [ARBF02, MCP^{+}06, APM^{+}09, BG04a, BPT^{+}00, BGSBC02, LTMDB02, NMBH02, WW03]. merguiensis [VHH^{+}02]. meristematic [GBMHD08]. Merlangius [CCML07]. merlangus [CCML07]. Merluccius [GBW02, LP04, SC00b]. meroplankton [TSSA03, VAP^{+}07]. meroplanktonic [MUT03]. mesenterial [NDB04]. mesh [KCTG06]. Meso [LNV^{+}05]. Meso-scale [LNV^{+}05]. Mesocosm [TGY^{+}06, DHZ02, DAA00, GT02, GK02, LFSR06, MGL^{+}00, ODA^{+}09, OS01, SO08, SKMS06b, SHSV04, SSS^{+}05, SHS06, SVL^{+}04, SES01]. mesocosms [DTD03, HGS^{+}07, MHHS06, PSGP04, RNDF^{+}04, SG01, vdMM03]. Mesodesma [LBD00, RCL^{+}09]. Mesodinium [SH07c]. mesograzer [KES00]. mesograzer-epiphyte-host [KES00]. mesograzers [DESK01, IAT^{+}07, WLdG^{+}04]. mesopelagic [ADS^{+}09, BBABL01, CDF^{+}08, CNC00, ISYM06, Ike08, KTK^{+}08a, KRKS09, LC02]. Mesoscale [BA00, CPA^{+}03, CBAd00, FDD^{+}03, MP06, SML^{+}09, HGZ^{+}00, HVWH06, LH09, MB00a, MFP^{+}05, NLP^{+}01, WR06]. mesoscales [SCKC07, Wie05]. mesotidal [GD07]. mesotrophic [DGW01, HGKK06]. Mesozooplankton [GRBB07, LDWC05, CVF^{+}03, DSP00, FPWM03, GK03, KR04, LD03, MBB^{+}04, NNS01, SSSV00, SHSV04, SHS06, BL00b]. Meta [GFB^{+}07, BCC08, HBRW04, MAWH03, SFW01]. Meta-analysis [GFB^{+}07, BCC08, HBRW04, MAWH03, SFW01]. Metabolic [CKO^{+}07, HHIS06, JDVG08, LJC00, MBC^{+}07, RT05, SP02, vHGH01, AMR01a, BS04c, CST08, FIS04, FRvdMF02, LC03, MNC^{+}02, Sim00, SML05, SP01c, SM05]. metabolising [PEG^{+}00]. Metabolism [HD06a, ISYM06, Ike08, PCC06, YI03a, BR03, BVH02, BYH^{+}06, CMJ00, CBE04, CvBGR09, GBB^{+}05, GAM02, HGT01, HDCA06, HAD07, IOSPT^{+}04, LGVF00, MCE01a, MCCT07a, MCCT07b, MEB^{+}07, MAT01, MLS^{+}08, MO05, NWFV06, PTLK04, PBVP09, PWM^{+}08, PMT^{+}02, SDPT06, SCC06, SCS^{+}07, SAP02, WE04, WRJ^{+}05, YI06]. metabolite [APP^{+}09]. metabolites [BMdNP09, BP04, DDQ04, EWSH08, GP01, MAE03, NGdN^{+}07, PCT00]. Metabolomics [Via07, JB07b]. metacercarial [Thi06]. metacommunity [MYNY08]. metal [AWOO02, AHK^{+}02, BFK05a, BW04, CW05, DSF^{+}04, DL03b, GGHA02, KR01, LMW^{+}01, MBH^{+}09, NW04, RSL09b, RCC01, SW04, WLL03, WL03]. metal-binding [SW04]. metal-contaminated [RSL09b]. metal-sensitive [WLL03]. Metallothionein [LW05, TBR^{+}03, GGHA02, SSG^{+}00]. metallothionein-like [SSG^{+}00]. metallothioneins [BW04]. Metals [GSSR02, BFK06, CW05, FSSW^{+}00, GdLB^{+}09, MFJT08, MF08, NATR^{+}05, RBW03, RPS^{+}06b, RPS^{+}06a, RAAT^{+}07, WYFX02, Wan02b]. metamorphic [MK04a]. metamorphosed [BWCW02]. metamorphosing [EBLE07]. Metamorphosis [NWHH01, CT06a, JQ04, LIT05, MPK03, PA05, RE00a, ST02, TW04a, TPGQ07, YOM^{+}09]. Metapopulation [Joh01]. metapopulations [SJ08]. Metaxas [EM07b]. metazoan [FSV^{+}08, KRP^{+}06, OAGV07, SGP07, Thu07]. metazoans [NOO^{+}08]. metazooplankton [KSKK^{+}07]. Meteor [KRP^{+}06, FMZvW02]. meteorological [GMHA09, GPM^{+}09, NIF00]. Methane [AI02, LJM^{+}00, LZM^{+}03, SLP^{+}09, TBK^{+}03]. methanogenesis [TSW^{+}09]. method [ÁSG04, CH07b, CUB08, DHG^{+}04, DC07b, HTJ00, HDM04, HPR05, HDDH06, Kam00, KKI^{+}06, LCC01, LCC07, LPD09, PTS00, RSR02, SMF^{+}02, TUK09, YF04, ZRG00]. Methodological [MUT03, DFMD00, PJ08]. methodologies [Dis07]. methodology [LLH06]. Methods [MW01, Pet04, BKP07, CF02, CLABRO^{+}03, DRTM09, GBB^{+}05, MRHP01, RRS^{+}09, Rii04a, STH^{+}06, UG04, WNS^{+}05, WHCP^{+}04, dSHCR07]. methylamines [FDRR05]. methylated [LTMDB02]. methylmercury [PW04, WW03]. metres [GCL04]. metric [PKS07, RM09b]. Metridia [HCBP05, PIY04, PJ08]. metropolitan [CRT^{+}08]. Mexican [AP06]. Mexico [AWA^{+}05, BUF02, BY09, BHY09, CRTP02, CC05, DJG^{+}07, DNL08a, GAMB05, GRPE08, GES^{+}06, GGDHRA00, GGMGR07, HCGC09, HCM01, HTM01, IOSPT^{+}04, JWTP06, JPG05, JH06, MCFM02, MFCM05, MA07, MIH^{+}00, MILH00, PPB^{+}01, PAK^{+}00, PAW00a, PCD04, RTH06, RRRM08, SAB04, SRN02, TS06, WPC^{+}03, WCF08, BHN09, GGR06]. Mg [AMB08]. Mg-calcite [AMB08]. MGDG [ZER^{+}01]. michahellis [RRS^{+}09]. Michel [DCRT07]. Micippa [HTS06]. Micro [PAW00a, CHH^{+}05, DDTQ05, FFG03, GBB^{+}05, HTRFP04, LD03, MD07]. Micro- [PAW00a, DDTQ05, LD03]. micro-diet [HTRFP04]. micro-geographic [CHH^{+}05]. micro-increment [FFG03]. micro-optode [MD07]. micro-tidal [GBB^{+}05]. microadriaticum [YM09]. Microalgae [HGMH06, AK02, BWvdE^{+}06, DCO^{+}08, EF02, JPC^{+}06, KKL^{+}03, LWB^{+}00, LBH01, SR09, SG01, Uth01, WWAC00, YI03b]. microalgae- [EF02]. Microalgal [MRG08, Saf03, BDC00b, COB^{+}03, DB02, DBD^{+}05, JAA^{+}01, KF02, KKLL02, NPM01, ÓUA05, TGM^{+}03, Zup00]. microbe [WMW^{+}08, vDHHM08]. microbenthic [HWL00, RGDN02]. microbenthos [GRCdLvB08]. microbes [RTB06]. Microbial [AMT03a, CLS^{+}01, DRG^{+}02, GHP03, SKR08, SYVC00, XZLP09, AZKJAM06, AG00, AJ03, BM06, CSMK09, CGCMPA01, DTD03, DM09, FH08a, GJG^{+}03, GGC^{+}05, GHS05, HUP^{+}00, JHTV02, KSSH02, Kir02a, KKB^{+}06, KMR01, KSKK^{+}07, KSWP04, LR08, MR04, MRB04, MMBT03, MSD05, NLB08, ONA04, PBF07, QDDP06, Rit06, SSM05, SGHG07, TSF01, TEB^{+}07, Thu07, TRS^{+}06, UB03, UYNO01, VLP^{+}04, WMCF08, dBMSL02]. microbioerosion [CSMK09]. Microbiological [MSUK01]. microchemical [HS08a]. microchemistry [ABS^{+}09, AKL^{+}04, FH08b, JSIT02, JSIT06, MAM^{+}01, ZKHP07]. Microclimate [PVGO01, MFK07]. microcopepod [BSS05]. microcosm [GRCdLvB08, RGDN02]. microcosms [PTLK04]. microendolithic [FH08c]. microenvironment [QASA08]. Microgadus [GC07]. microgastropod [CUB08]. microgastropods [OC01]. Microgeographical [NWHH^{+}09]. Microhabitat [AP04a, AP04b, COSM03, DH07b, WDT^{+}08b, BJC02, DE01, EBC04, MC07, MS06, NLK00, OHS^{+}07, RK09, SH05]. Microhabitat-scale [AP04a, AP04b]. microhabitats [KKH07, LJ08, RH03, ST06a, WH09]. microincrement [HMHB03a]. microlayer [Kel04]. microlepis [HKZ04]. Micromesistius [AML^{+}03]. micronekton [BBA06a, BBZM08, SML^{+}09]. microorganisms [RMGL07]. microphallid [MS00a]. microphytobenthic [DPH02, FK04, FCS^{+}06, GBG^{+}00b, HGRB04, HBP^{+}05, KKA^{+}06, LV05, MS00b, POH^{+}02]. Microphytobenthos [KLC^{+}06, BK00, FH02, GFF08, LS08a, LLG^{+}06, LLL^{+}09, MBR^{+}05, MMA03a, OSS^{+}04, PUB^{+}01, SC00a, SdSC01, SMG00, TYT^{+}04, dBS01, vBWSS06, vdWHF^{+}08]. microphytobenthos-originating [TYT^{+}04]. microphytoplankton [BTC^{+}04, GhGD^{+}02]. Microplankton [CTFC08, ESQ03, LDWC05, McQ05, MASFS00, NIF00, OLWB06, LOT^{+}00]. Micropogonias [JBCL03, MFA03, SEH07a]. micropredation [PGK^{+}08]. microprobe [KSHP00]. microps [JRA02]. Microsatellite [BSK06, BHMR01, MPR^{+}06, BHH05, DEV07, GDSB04, HFM^{+}04, MHH^{+}05, WGM02]. microsatellites [HCR01, RFCM03]. Microscale [FJ01, SSM05]. microscopic [EDBvdL08, FS07, FON03, GVKF03, JB07a]. microscopy [HSS^{+}04, LPBD00]. microsensor [dBKSV00]. microsensors [KGB^{+}01]. Microsetella [MVN^{+}06]. Microstructural [AE06, HS08a]. Microstructure [HGRB04, FBJ00b, HMHB03b, JBH00, MAM^{+}01, vdMM03]. microtidal [QMM02, SMG00]. Microtiter [SM05]. Microtopographic [SGCC06]. microzones [JKG^{+}05]. Microzooplankton [PI07b, SWEBS00, BND02, CVF^{+}03, CLL^{+}09, DGM00, GRLML09, LCL^{+}00, NNS01, NHNB01, RSR02, SS07, RFG^{+}09]. mid [AAF^{+}04, BGM09, CCR08, CLC04, HLBS07, HTBT07, HW03, LÅCN01, MSSP06, OJPM01, PTS00, RJS04, SAHF03, Har06b, KBP06, MCP^{+}06, SRJ09b, MG09a]. mid-Atlantic [MSSP06, Har06b, KBP06, MCP^{+}06, SRJ09b, MG09a]. mid-domain [BGM09, CCR08]. mid-intertidal [CLC04, SAHF03]. mid-latitude [RJS04]. mid-ocean [HTBT07]. mid-oceanic [OJPM01]. mid-shelf [PTS00]. mid-shore [AAF^{+}04, HW03, LÅCN01]. mid-water [HLBS07]. midae [ESB^{+}04]. Middle [TCS^{+}09]. midge [SL05d]. Midnight [TJEM02, BSKR^{+}06, Hel00, Pea03]. midsummer [CLS00]. midwater [BBZM08, MBBW08]. migrating [EB08, EvdKB^{+}08, JT02, JT05, SAW^{+}01, SS02a]. Migration [BdVD^{+}01, DNL08a, HHD^{+}09, JSIT06, KNF^{+}03, NLK00, SKO^{+}08, WFG^{+}08, Wes03, ALOT01, BBABL01, BBZM08, BSKR^{+}06, BZLD02, BT04, CHL^{+}05, CKF^{+}03, DSM01, DEvdZ03, FWD^{+}01, GSFF^{+}09, GB02b, HMH^{+}08, HW02, Hid03, HMA06, HFA09, ICH04, JSIT04, JSIT07, JF07, KS04a, KAM^{+}06, LS06, LLG^{+}06, LLL^{+}09, MADPQ06, MBBW08, NHS^{+}08, NS02, OYT^{+}02, Pea00, QMM02, Ran02, RMKR00, SS02b, SHM^{+}01, SAM^{+}01, TND^{+}05, Tar03, TCC04, WKMK06, WEEBdN08, dSSC^{+}08]. migrations [BBA06a, CCCH01, CBM^{+}07, FLR05, GMB^{+}06, JSMM07, JPC^{+}06, NBN^{+}08]. Migratory [CSR^{+}00, CIT04, JSIT02, SRJ07b, ACAG07, BLH01, EDB^{+}04, IAKS03, KOY^{+}05, Kro05, KGF^{+}09, KW02b, SC00a, SICT03, TSI02, TCW^{+}07, WKH02, WWR09]. Miguel [BT00]. Mikawa [KOY^{+}05]. mild [BD03]. miles [KMWH04]. Milford [Lus05]. miliaris [HCK^{+}05]. Milk [BN09]. millepora [BMWH08, CSvO09, MDW^{+}09]. mimic [GF02]. mimics [Fin00]. minax [GB07, OJ04]. mine [FMCB05, LMC08]. mineral [VK02]. mineralisation [TNSM00a, TNSM00b]. mineralization [GRPT^{+}00, GHS05, VNJ06, WBW^{+}06, WRW^{+}04, WRH05]. mineralogy [BBC^{+}00, GBC^{+}04]. minerals [ZLGC04]. Mingulay [RDH^{+}09]. miniature [QRS00]. Minidiscus [KKKK03]. minimal [BWH09]. minimising [GML02]. Minimum [MFE^{+}05, CGF04, RCWG^{+}06, SDW^{+}01, KJ04]. mining [SFW01]. minke [ABD^{+}03b, JTR05, MFG^{+}04]. minor [CCCD09, HDRC^{+}08]. minutum [BGF^{+}06, CVF^{+}03, CCM01, FGM^{+}00, GFM^{+}02]. Mira [AQSP07]. mirabilis [LG06]. MIRO [LSG^{+}05a]. Mirounga [GAMB05]. Misidentification [TCW^{+}07, Yos08b]. misidentified [Yos08a]. mismatch [BPCvdM06, GPSG06, WIDM09]. mismatched [SPH^{+}09]. mismatches [Hip08]. misrepresentation [Wid01]. missing [CRT^{+}08, SJM06]. Mississippi [BJR04, CB08b, LD03, LDWC05]. mistakes [CWR^{+}03]. Misuse [Rii03, BSW^{+}09]. mitchilli [JH04, KSHP00, Lap01a, Lap01b, MHHS06, Pee02, SBJ02]. mitigate [GDR^{+}04b]. mitigates [GP04b]. mitigation [AGB03, BCO07, DSH03, Hue08b, MR03]. Mitochondrial [ADR^{+}05, DRV^{+}05, GTGM06, TGGM05, VBA06, CMRB08]. mitten [CACDC07, HCB02]. Mixed [Zam03a, BGF^{+}06, DY05, FLG02, HRK^{+}08, JM03a, PCC06, PSE^{+}03, SGL02a, TIM^{+}01, Whi06]. mixed-species [DY05, JM03a]. mixing [CGH^{+}03, CRH01, HGG^{+}03, KHvW02, LKR^{+}06, MLR02, PSGP04, SFHS00, TKTJ07]. mixotrophic [ASH03, HSGS00, RH09, SBL^{+}00, SHS00, SCS03, ZAH03]. Mixotrophy [HP06, PG00]. mixta [JR01, HDGM01, VFV01, VV04]. mixture [BCM^{+}07]. Mn [VNJ06]. Mnemiopsis [BN04, DBP04, FKA^{+}06, HCS^{+}07a]. Mobile [HCV03b, PCB08, AFHS09, ABHE09, Cha03, CHT09, GPHC09, GAB^{+}03, Gim09, GC05, KS00b, MCFM02, PTSJ06, PJP06, PPSP07, TAS07, YAF^{+}07, ZRG00]. mobilisation [RG09]. mode [BBM^{+}02a, BWPB04, CCB^{+}09a, Col03, MC05a, VFJ^{+}07]. Model [LGBWH06, PRM^{+}09b, PRM^{+}09a, ANCC07, ATT06a, ATT06b, AHZ03, ARP02, Arr05, BMSH02, BDGS07, BGM09, BBM00, BMT07, BCC08, BWRP06, BD01, BT04, CHL^{+}05, CAF^{+}08, Clo01, CVSM09, DTD03, DPH02, DRF08, DEV07, EvBO^{+}08, EM07b, FSSW^{+}00, FBD06, GMVF08, GBG^{+}00b, GBD^{+}06, Hal04b, Hen07, HLR04, HDP08, ISY07, JWG01, JSMB05, JF07, JPG05, Jon04, JMJ07, KWTP06, KGTH01, KGHW05, KSR^{+}07, KBSWG09, KJRG09, LSG^{+}05a, LRK09, LRR^{+}09, Lei01, LJK01a, Mal02, MGC03, MDG04, MB07e, MYNY08, MKK^{+}09, NYK03, NSH^{+}08, OSF08, Orv05, Pah05, PO09, PKLD09, QHD04, RM09a, RG09, SDK^{+}02, SW02b, SW02a, SWH^{+}04b, SMGCW00, TF04, TT07a, TCG03, VFF09, VHM09, WGCM^{+}03, WBR07, WCM08, WTR02, Woo09]. Model-based [PRM^{+}09b, PRM^{+}09a]. Modeled [JK06]. Modeling [Abd07, BFB01, BRP07, BBNK02, FHM02, GBI^{+}06, HL00, JZL09, LNPB04, Ran02, TDTR03, TCS^{+}09, Urb02, WKYL00, WSL06, ZAH03, AB04, BP08, BPC09, CBGLM08, CN04, DH09b, FMWGB08, FGS^{+}02, FBN^{+}07, GPLAG04, GF04, GFL02, HSBD08, HRC^{+}07, JP04, KECJ06, Kap06, KLW^{+}09, NRC07, PCCL05, PD07, Pen00, RFB^{+}06, RMM04, SW02b, SW02a, STM06, TK08, WWGR07, WCH09].

Modelling [APG^{+}09, AMS^{+}06, Arm00, BPN^{+}09, CPB04a, EMG01, FK04, Hid03, HFMP03, JF02, KWP07, KCK^{+}06, LSG^{+}05a, LGSB09, MR01, Pep06, SK07, SC00a, SPC^{+}00, AC07, BT09, CDJ^{+}07, DDGK^{+}08, FBJ00b, GNPB07, GNP07, Han07, JBH00, Lei07, NJPI03, NRPS^{+}01, RSK04, RSL09a, SGH^{+}06, TSW^{+}04, VSRGW09, WPTS07, EvBB^{+}09]. Models [SJH^{+}04, BRSM07, BPS^{+}00, BVP^{+}06, CHPS04, ECK05, FM02, FSJ03, GDF03, KBH01, KS09, Kup03, LDHC04, MBA^{+}05, MM03, MDHL08, Mil07b, NZ08, PB08b, QRS00, RBB^{+}08, SWD^{+}02, SMP01, SSH09, TW10, dSHCR07]. Moderate [VSL^{+}04, MOPP05, WS01]. moderately [LÅCN01, TMJ^{+}05, WO05]. modern [PSW07, PCBL05]. modes [BFV08, FA08, KT04, LGFM00, LGME06, Whi06]. modesta [Kru01]. modestus [RW01]. modification [BS07c, BB02b, MPM^{+}03, SL07, SDM06, TS05, WM06]. modified [PBJ^{+}05, TUK09]. modifies [JK03a, RSSA07]. modify [PKRK07, WS09]. Modiolus [RCC01]. modular [VÅ01]. modulate [MG09a, SJPC02b]. modulated [KGB^{+}01]. modulates [BD08, CAT^{+}06, JQ05]. Modulation [BGJ^{+}07]. Mola [CL04]. molding [TL03]. mole [DGL01, DC02]. Molecular [BPLD^{+}05, Dis07, MRQM06, SGV^{+}02, SL05a, THP^{+}05, AWOO02, BBS05, CLABRO^{+}03, DGB^{+}02, GTSS03, NCMC09, STRD02, SHSF08]. molecule [NBB06]. Molenplaat [SHM^{+}01]. mollis [PPS00]. mollissima [LG00a]. mollusc [CDK^{+}07, DSS04, DJV03, HWvdM03, HDW04, LGD04, PBTB08, TYD07]. Mollusca [GMS02, Iye02, KCB06a, LCW06, NSSI06, OJPM01, PTH08, RGMJM07]. molluscan [BMC^{+}09, BAMJ09, KCS04, MRMA07, PDB04]. molluscs [Bay04, FK08, HSW^{+}08, HLR04, Niz07, RSN02, SSBG02]. mollusk [CSM^{+}09, MTR^{+}07]. molt [AY04a, Joh03]. Molting [OJ04, O’C07]. Mombasa [MHO07]. Mona [DvD02]. Monacanthidae [KN00]. Monachus [BPH07]. monasteriata [MFK07]. Mondego [PMM^{+}00]. Monito [DvD02]. monitored [AY03, SJM06, ZAHJ^{+}06]. Monitoring [Bea04, BI04, BMG00, BBB^{+}06, BOR08, CM04a, CHT06, FCH^{+}08, FAB^{+}07, HDDH06, Law08, LCC07, Mal02, MTLG09, PR05, Sta08, SBLF09, TBR^{+}03, TM02b, TLC06, YF04]. monk [BPH07, GLA06]. monocerata [TS07]. monoceros [LHJD^{+}03]. monocultural [HK02a, HK03a]. Monod [SSH09]. Monodon [LHJD^{+}03, LF02, RT00a]. monogalactosyldiacylglyceride [ZER^{+}01]. Monoporeia [BEE06, EEA01, Kar07, NWÓ03, WÓ02, WS01, vdBÓME01]. monopterygius [NS02]. monosaccharides [vDSD02]. monospecific [HLQ02]. monoxide [Kin01, XZUK05, XZLP09]. monsoon [MMA03b, NSJ02]. Monsoonal [Yin02]. monsoons [Yin03]. Mont [DCRT07]. Mont-Saint-Michel [DCRT07]. montagui [CCB^{+}09a, DMM^{+}03]. Montastraea [AS06b, FFHW07, FBM08, GB06, Les00, LFRR^{+}07, MW02a, MW02b, PB08a, VFM06]. Monterey [DS09a, Hen09, KGS^{+}08, LRK09, RCB05, War05]. Montevecchi [Car05]. month [KJB^{+}06]. monthly [MIH01]. Montipora [MFK07, SDHG03]. Moodley [Ano02f]. Moorea [AAGS02, MACS07]. moratorium [JCL09]. Morbillivirus [PHH^{+}05]. mordax [BDRM06, SD00a, SD00b]. Moreton [EB04a, KCSPT09]. Morgan [MB07e]. morhua [AP00a, BM00, BM02b, BM02c, CSB03, CHH^{+}05, CDH^{+}09, CCL^{+}05, CCCH01, CPG^{+}09, FRvdMF02, GBR^{+}04, Hel00, HJG06, HMHB03a, HMHB03b, HNMC09, HCR01, JKES07, LGB03b, LGB^{+}04, LAK07, LG02, MHP^{+}00, MB02b, MSF^{+}09, MGSJ08, Mun07, NB06a, NM04, NWHH^{+}09, PTM08, PBCB03, RR06, RHSB08, SHNV09, SSL00b, SRJ07b, TKK05, UP01, WPIB09, YW04, vdMM03]. Morlaix [PD01]. Morone [KS04b, NH03a, NH03b]. Moroteuthis [CD03b, JBG00, PJN01, PJN03, SJ05]. morphodynamic [DFT^{+}07]. Morphodynamics [DG05, Bra01, FDDS09]. Morphological [BBB^{+}06, BCBVM06, BG04b, CLMR07, DD09a, RHSB08, VMVG^{+}07, BCL03, BLS03, Bat07, CLABRO^{+}03, Goc04, LNJ07, MC01b, SMH02, TN04, WPB08a, WGM02]. morphologically [DB02, TS04]. morphologies [FKBB01, SL05b]. Morphology [KEB07, OHS^{+}04, PB08a, BS08a, BND02, BSL02, BB00, CCdAR00, CDK^{+}07, EMB^{+}09, Fer04, GDGSP09, JS02, LD00, Mar00, MCS05, NOP02, OEDSJ02, PGM05, RC04b, SA03a, SvNvdMvdV09, SFR^{+}09, TL03, jMFVPL05]. Morphometric [FVAB00, DESK03, LCHG03, SM08b, VSV04]. Morphometrics [RJ03]. morphometry [MSG07, PBPLB06]. morphotypes [BHY09, CWN^{+}07, LLRDT05]. mortalities [Har08, RWA^{+}02, WHM^{+}01]. Mortality [BM02a, HK02b, MF03, TNT08, BMM^{+}09, BPvdM07, Bra05, BSPK03, CJFM09, CPB04a, CMD^{+}06, CLR^{+}06, CGH^{+}03, CRCN07, CCB^{+}09b, DMM^{+}03, DdMB04, DS07, EZTZ02, EO04, EvdKB^{+}08, FJK^{+}07, GVM^{+}05, GPSH01, GC06a, GBD^{+}06, HDW02, HBG^{+}03, HBH07, HW08, Hun04, JTVC08, Jar00, Jef03, JLCA01, KKB^{+}06, KKSR05, LBF^{+}07, Lap01a, LCD^{+}05, MAVO^{+}06, ML00, MVH^{+}06, MARM06, MBS03, MBS07, MGSJ08, MMBF09, MGH00, NCC^{+}07, NCN^{+}09, NZM08a, NM04, NMJ02, OLG^{+}09, PL07, Prz05, SSL00b, Sha09a, SMG02, SG02b, SHMN07, SAM^{+}06, TUK09, TSE^{+}08, TP07, TSW^{+}04, WMG02, WS09, Wil04, WS08, vdMM03, vdVBW00, HM05b]. Morus [HPW^{+}00, HPH^{+}01, MND^{+}09, PRvdL^{+}07]. mosaic [BDLB02, GMVF08, She09]. mosaicus [PK00]. Mosby [VMdBV06]. mosquito [MMPD09, ROCB09]. Most [Gou06, Gou07]. mother [DPD^{+}03a, DPD^{+}03b]. mothers [BKLGN05]. motif [WGM02]. motile [HSCH07]. Motility [TK03a, CBO04, RS07b, RB05b]. motion [GWH03, Kaw01, KPT01, KT01, Paf06, PJ06a, PFZS^{+}06, RC04b]. motivation [JS08a]. motor [HKWM07]. Moult [KHA^{+}00, KSSB04, STPA04, DL03a, HM06a, SRLA04, TJSG03, HPR05]. moulting [WG02b]. mound [GWF07]. mounds [vDHHM08]. mounted [KRKS09, WBL^{+}06]. move [LC07b]. Movement [GLÅ^{+}03, GC06b, LTCP03, LSS^{+}09, MC05c, NZM08a, SJCM01, WGR07, BCU04, CL04, CHPS04, DHR04b, DHR06, DB07b, DS07, EOK^{+}08, FH08b, GWH03, GCL04, HJW08, JKK^{+}06, KNK^{+}00, KJRG09, LLM^{+}08, LGB^{+}04, LRHV07, LAK07, MCP^{+}09, MSP^{+}06, PJ06a, RW08, RSD^{+}08, SSB02, SC00b, TVD02, TLC06, WPM09, WMG02, BEW03a]. Movements [CRLP09, DS09a, HALL05, PCS^{+}07, WOL^{+}07a, WCJL08, ZSR03, BB09, BHN09, CPP^{+}04, DFW^{+}03, GFB^{+}07, HRMH07, JWR05, LMKG01, MHP07, RTPS07, SSR^{+}03, SJL05, SO01b, TMAN08, UDB^{+}09, WWR09, ZAHJ^{+}06, dBTO^{+}09]. moving [Gla01, WPTS07]. Mozambique [PDQ01, BD00, SML^{+}09, WLJ^{+}04]. MPA [PBFB08]. MPAs [EFL^{+}09]. mRNA [HYLT03, LK02, LVK05, TTM^{+}05]. MS [RCC01]. MSF [WLL03]. mt [HCB02]. mt-DNA [HCB02]. mtDNA [GRPE08, LW04, MB07b, NW05a, TR06]. much [Nug02]. mucilage [CPV^{+}03]. mucocyte [BDDL03, BCJ05]. mucoid [SRGS01]. Mucopolysaccharide [CPV^{+}03]. Mucus [HWG06, KAMR00, RHG05, BB05, ESDG08, ED02, GB06, HSCD01, NReZW09, Rit06, UNII01, WRW^{+}04, WWH05, Woo09, Zam03b, Zam07]. mucus-associated [Rit06]. Mucus-net [KAMR00, RHG05]. Mud [KWKF03, VMdBV06, CCdAR00, EKA03, GDL04, Gri09, ORPND01, PPJ^{+}05, RE00a, SLP^{+}09, THN^{+}03]. mud-flats [ORPND01]. muddy [HCA06]. mudflat [AZKJAM06, AF06, DAM^{+}08, FH08a, HGRB04, KKA^{+}06, LNB^{+}03, LNPB04, OSS^{+}04, VMV^{+}08, WPB08a, WOH05, dBS01]. mudflats [CBE04, CRBE04, CRCV04, CVSM09, GBG^{+}00b, HBP^{+}05, MTCU08]. mudskipper [NWT00]. Mugil [CIT04]. Mulinia [VN02]. Mull [MFG^{+}04]. mullet [CIT04, SSL^{+}07]. mullets [KHGA02]. Mullidae [KFS^{+}09, LRGdS00]. Mulloidichthys [KFS^{+}09]. mulloway [TLFS06]. Mullus [LRGdS00, PEdlP^{+}02, SSL^{+}07]. Multi [AFHS09, BCC08, CMD07, STF^{+}09, TOK08, THHY05, BCL03, DM05b, EJM01, Gar00, GCM01, HDG^{+}08, LPD09, OT08, TH06]. Multi-century [STF^{+}09]. multi-method [LPD09]. Multi-scale [AFHS09, CMD07, THHY05, BCL03, DM05b, EJM01, GCM01]. multi-scaled [OT08, TH06]. Multi-species [BCC08, TOK08, Gar00, HDG^{+}08]. Multibeam [CDW^{+}09, KTF^{+}01]. multidens [NSEL00]. multidentata [ACAG07]. multidisciplinary [CLABRO^{+}03]. multidrug [LVK05]. multimodal [RWP07]. Multiple [BR07, BGH06, HKR^{+}08, JY08, OPSM04, RE00a, SS04a, VACT^{+}08, AA04, BHMR01, CU08, CEF04, CC07, EE00, HCD^{+}09, Jos09, KHIH08, KW02b, LPK^{+}03, MSF05, MBHN07, MCVAR08, MH00, MR03, MS04a, MS01b, NvdVVD06, NDOS03, NPTW05, PKM04, PMGE02, PLG09, RFSC08, SAG08, Too04, VBAR07]. multiple-indicator [KHIH08]. multiplex [HLH^{+}05]. Multiscale [TSMC09]. multiseries [LHK05]. Multitrophic [FD04, BSP^{+}08]. Multivariate [MB05a, SO08, TBCB^{+}05, GGGVGG03, Hir06, KHD07, Las03].

Multixenobiotic [MMGC06, LK02]. mummichogs [KS00b, Kne09]. Muraenesox [LIT05]. murex [PCB08, Swa04]. muricata [SBM02]. muricid [CGT^{+}00]. murphyi [BBG^{+}06]. murrayi [PCS09]. murre [ZPP09]. murrelet [BB03b, BPB07]. murres [DMA02, FBD^{+}02, SPH^{+}09, SH00]. muscle [BG04c, JVT01, TFSJ00]. Musculista [Mis04]. musculus [RBHS06]. mushroom [Gil02b]. mushrooming [WDLM05]. Mussel [MDS^{+}07, AW03b, Alt06, ABY00, BFK^{+}05b, BvK07, BOR08, CBM^{+}03, CWMD06, CMD07, CC04, CYC^{+}09, CSD^{+}07, DM01b, DKFV05, DLD^{+}00, EM04, GP04a, GKS^{+}07, GSRH02, HMH04, HDC^{+}08, HTM01, IRM08, KR05, LFRG^{+}05, LW02, LVK05, MNB^{+}07, MB00a, MC05b, ML00, MP00, ML05, MNP02, NSBK05, NSSI06, NW05b, NM07, NWFV06, O’D08, PPCU07, PW04, PTB03, PBS^{+}04, POH08, RATD07, RCC01, RS06, RBG^{+}07, SH07a, SSB02, SS04b, SJPC02a, SJPC02b, SAG03, SMGCW00, TM02a, Thi06, VT00, Van02, Van03, WC03, ZMTB07, ZHWCS08]. Mussels [DSP00, ABB^{+}00, BFK05a, BFK06, BJH04, BJN^{+}05, CLN00, CEB^{+}08, EPM06, FTU^{+}00, GSL^{+}06, GXL^{+}08, Gar00, HMK06, HD07a, HS01, JNH^{+}09, KT04, KEB07, LKR^{+}06, LGBWH06, LGQ^{+}00, LMF^{+}01, MP06, MOPP05, MN04, MGH00, NZM07, NZM^{+}08b, NZM08a, OGMC06, PCL^{+}09, PMR06, QTB02, RL07, RBG05, SM09, Thi05, WGD05, Wan02a, WJ06, WLM06, WLT^{+}03, ZMN07, vdMPEM08]. muticum [BS04a, IOTL09, SEH07b, SPT^{+}00, SDM06, TJA08]. mutual [TE00]. mutualism [HH05b, PH01b]. mutualistic [CÁ04]. MW [KCH^{+}03]. MXR [LK02]. Mya [ARB04, Hun04, MB08, SMT08, TE00]. Mycedium [DFY00]. mycosporine [CCM01, LBR04, SFPGA05]. mycosporine-like [CCM01, LBR04, SFPGA05]. Mycteroperca [HKZ04]. myctophid [MKWO00, RS00, SCT^{+}09, SB08, WKH02]. Myctophidae [SK04, SK05, WKH02]. myctophids [PJN01, RGCL^{+}05]. myctophiform [SS00a]. Myctophum [SK04]. mydas [ABL08, BBC03, BHN09, GSB^{+}06, KCEB09, SRN02, SJE^{+}06, TTD^{+}06, TBBB06, YA09]. Myrionecta [BDÁM04]. Myripristis [CEBR07]. mysid [CM01, KPN03, RWPJ04, VV04, WMD^{+}07]. Mysidacea [JR01, RWJ02]. mysids [CM07, JR01, TK04, Vii07, WG02b]. Mysis [HDGM01, VFV01, VV04]. mysticetus [LSMR05]. mystinus [JKK^{+}06]. Mytilus [ABB^{+}00, Alt06, ABBF01, ABY00, BFK06, BJN^{+}05, BJN^{+}06, BGKN00, BOR08, BWH09, BH01, CLN00, DSP00, DM01b, DLD^{+}00, Fre07, GGHA02, HMH04, HD07a, JDVG08, KR05, LGBWH06, LC07b, LLSL06, LMF^{+}01, LK02, LVK05, MNB^{+}07, MP00, MOPP05, MMGC06, MGH00, NSBK05, NZM07, NM07, NK07, NF06, PTB03, PBS^{+}04, POH08, QTB02, RL07, RBG^{+}07, SSB02, SS04b, SJPC02a, SJPC02b, SB03b, SAP02, Thi06, VN02, WGD05, WJ06, WLT^{+}03, XBW07, ZMN07, ZMTB07]. myxoxanthophyll [SGK04].

N [WBN^{+}00, CLP08, AHSPG08, BMK07, BRS07, BVI^{+}05, CF02, CGF04, CVF07, CST08, DLLB03, DGBS08, DPD^{+}08, EKBM05, FPC05, Fen08, FE07, Fre03, GXL^{+}08, GLS02, GKM02, GV01, HH00a, HGMG^{+}04, HD07b, KPM00, KP01a, KKL^{+}03, KBP06, LSH^{+}07a, LHB09, LWK^{+}07, MSF^{+}07, MH08a, MSF04, NEMF09, ORPND01, Pah05, PO09, PSW^{+}08, PH03c, PCL08, RSN02, RH03, RW08, Rol00, RCGA^{+}04, SAT^{+}03, SMM^{+}04, SJE^{+}06, STF^{+}09, SFKT09, SCGC07, SHSV04, STH^{+}06, THS^{+}02, TBI07, VS08, WPRS05, YQW^{+}01]. n-3 [CLP08]. N-ammonium [ORPND01]. N-specific [CST08]. N. [CVK00, PCSS01, RPS^{+}06a]. N2 [FB03b]. Nacella [FCP02b, PTP01]. naïve [UP01]. Namibian [BAB^{+}01]. Nanaura [KKA^{+}06]. Nanism [SSL^{+}07]. nannochir [MKWO00]. Nannochloris [SML05]. nanoflagellate [JSK^{+}07]. nanoflagellates [AHW03, CNC00, CPH^{+}02, FK00, KW00b, PC02a, SJK^{+}06]. Nanophyetus [JTVC08]. nanoplankton [CLS00, FPF01, GRS07, HB03, SBL^{+}00]. Naples [OPSM04]. Narragansett [Lap01a, Lap01b]. narwhals [LHJD^{+}03]. nasa [WS08]. Naso [RAM03]. nassariid [BSBD^{+}04]. Nassariidae [CR07a, CMRB08]. Nassarius [BGQR01, BMB02, CR07a, EKA03]. Nassau [AP06, DE01, SSBZ07]. Natal [FCS^{+}09, BS07b, BL07, RSD^{+}08, RW06c, SSW08, SRJ07a, SRJ07b, WHS^{+}09]. National [CH05, KVH02, MHO07]. Native [AW03b, BCS^{+}06, CB08b, DKR09, GW06, HKH09, HH07a, HH07b, JMA02, JHR05, KSI09, LC07a, LG08a, NWÓ03, Rue07, SS09a, SG06b, SH07a, SBSPS04, SMC09, SM09, WOH05, YBGP06]. Natural [AAGS02, HFM^{+}04, IRM08, RCR03, TKR06, ASC06, BGGR02, BR03, BdBK^{+}00, CBP^{+}05, CF02, CPV^{+}03, CB08b, DBR^{+}07, DPH02, DS06a, DBD^{+}05, Dis07, DPD^{+}08, DVLH^{+}00, DLD^{+}00, EvdKB^{+}08, FLG03, GHD^{+}02, Goo09, GMS02, GWBdN09, HJG03, HM03, HKM00, HF09, Hil09, JLW04, KSIG08, KEB07, KVS08b, KH05, LTG^{+}01, LABR^{+}06, MFCC02, OJ02, OM04, PDS^{+}07, PH03a, PKR07, RHH^{+}05, SFH^{+}00a, SFHS00, SMC^{+}05, SLC05, SJCM01, SSSV00, SCD^{+}09, SSW08, SFG05, SCAF03, SLEH01, SdNH^{+}06, TCD01, TFSJ00, TS06, UP01, VSVV03, WKL^{+}03, WE04, WRJ^{+}05, WKPK05, WKK07, YF04, Zac05a, Zac05b]. natural-abundance [CF02]. naturally [BW00, GCU07, KWAT08]. Nature [She05, DJM^{+}05, KJ03, RCC^{+}09]. naupliar [BT06]. nauplii [BCB^{+}08b, HSCH07, KKP08, RHI^{+}01, SM05, Tit01, TK03a, TK03b, TLNH01, ITH^{+}03]. nautilei [SKW^{+}06]. navigation [TJM00]. Neanthes [LLY^{+}01, ODG^{+}04]. neap [KKA^{+}06]. Near [AL03a, DHT^{+}08, AAHDS06, ARGS01, BBV^{+}01, BWO04, DCF^{+}05, Gil02a, GGMGR07, GKB09, ISG04, JM01, KSG03, LBPS03, LW04, MB05a, MUT03, NRJB04, Pee02, RWBS03, SSL00b, SP01a, SCBA05, SH00, VMRL04, WPB08a, WPB^{+}08b, ZHS^{+}05]. near-bed [RWBS03, WPB08a, WPB^{+}08b]. Near-future [DHT^{+}08]. near-reef [NRJB04]. near-shore [DCF^{+}05, GKB09, JM01, MB05a, MUT03, SP01a]. near-surface [GGMGR07]. near-term [SSL00b]. Nearest [SVN04]. nearly [HK02a]. Nearshore [PBB^{+}02, BGJ^{+}07, BB03b, BBA06a, BR08, BRG^{+}02, CWKE05, GMC^{+}09, GSC^{+}01, HMK06, JTN09, LYMCG06, SVG07, Sma07, SW09, SRS04, TP07, VLRJ07, WKN^{+}03, Zam03a, ZT07]. nearshore-offshore [HMK06]. nebouxii [ZBDE07]. nebulosus [HH00b, Kup03, LBHL^{+}09]. necessary [MHP^{+}00, SH04]. necessity [BUA02]. neck [Mar04]. Necora [LWJC04]. necto [Vii07]. necto-benthic [Vii07]. need [SRJ07a]. needs [BHZW08, EM07b, GBB^{+}01, MWT^{+}06, MB07e, SRP^{+}06]. Neer [Ano07e]. Negaprion [WGR07]. Negative [AH07, DEAI06, Ada01, CABS02, KF01, UB02]. negatively [Bar07]. negligible [FAK07]. neighbors [SVN04]. neighbouring [GDR^{+}04b, RGC05]. Nekton [KLN06, RMZC07, AHBM^{+}07, BR08, BFE^{+}05, EB08, Kne03, LB08a, MJG01, MAWH03, RZ00, RM09b, VMH^{+}08, WBR07]. nektonic [QFJ^{+}07]. Nematode [VSVV03, CLH^{+}07, DRV^{+}05, LTMF01, MHV^{+}00, SFL^{+}09, TVG00, VMV^{+}09, VMdBV06, VSV04]. nematodes [BAB^{+}06, CBD^{+}05, DBG^{+}09, DDS^{+}06, GSM05, GMVF08, LTFJ00, LTG^{+}01, LBF^{+}02, MLM^{+}02, MA02, RSE01, RH03, ROBFR08, SMH02, SDW07a, SHM^{+}01, UÓ03]. Nematostella [DKR09]. Nemopilema [KUOI06]. Neo [BBS07]. Neo-Caledonian [BBS07]. Neocalanus [BM09, CBD04, ESDKM09, EBG^{+}00, KI01]. neogastropods [SJ02]. Neogoniolithon [RZEACGF00]. Neomysis [RWJ02, WG02b]. neon [CCC09, IMSO09, WKIK04]. neoverrucid [WTF^{+}05]. Nephropidae [ASA^{+}03]. Nephrops [ASA^{+}08, ASA^{+}03, BWMA02, CFG06, FCC^{+}08, PWAT01, SMA^{+}03]. Neptunea [SJ02]. Nereididae [VBA06]. Nereis [CVK00, CBA02, FGS^{+}02, KPA^{+}09, KM03b, LOB03, RPS^{+}06a, RSL09a, RSL09b, WBP09]. Nereocystis [BSED08, CWKE05]. neritea [BSBD^{+}04, CMRB08]. neritic [CWT00, HMKR05, Wil09, WJB09]. neritic-slope [CWT00]. neritina [GQ00, LTL06c]. Nervión [BMF06]. nest [DGCM09]. Nested [FS07, MVdMFF08]. nesting [AS04, FFT^{+}07, GFB^{+}07, HGS^{+}02, HBG^{+}02]. nestling [BHH09]. nests [MAW06]. Net [GBB^{+}05, ODA^{+}09, WB01, CvBGR09, CDT^{+}01, GAM02, KAMR00, MFP04, RHG05]. Netherlands [BdVD^{+}01, NvdV02, RSN02, SHM^{+}01, WBSL00, WBH^{+}04]. nets [BCL^{+}01, CR04]. netted [CR07a]. netting [PPSP07]. Network [DWM04, BMW^{+}00, LDP^{+}08, OT08, VLRJ07, WDB^{+}05]. networks [MHP03, Win09]. neumayeri [BP01, BPT07, KR01, TYC00]. neural [BMW^{+}00, FSW02, MHP03]. neurotoxic [ITE^{+}04]. Neuse [PPHT00]. neustonic [RBSP00]. neutral [CD03a]. neutralising [PHH^{+}05]. never [RGVC07]. newcomers [Zab09]. Newfoundland [CDT^{+}03, COSM03, RS03, TDRM08, TDTR03]. newly [Gri09, GCS06, HM06c, JPC00, JKK^{+}05, KB06, MPM^{+}02, PS08, PZMG05, RMGL07, SML02b]. newly-opened [PZMG05]. next [WBT^{+}08]. NH [HH05a]. niche [BD00, Bar02a, BH08, BPC09, BSIS07, DD05, FMZvW02, GMR^{+}02, HEN^{+}08, LTD^{+}02, OS01, PCRR08, RGCL^{+}07, RGF^{+}03, SS00a, SJR00, TLFS06, WHL00, Zab09]. niches [NRP^{+}02, WIB08]. Nielsen [GNAD^{+}04, DBCB06, TBBI04]. niger [Rod03]. night [BBHT02, HOVG07, NDV^{+}00a, OMT06, PPS^{+}07, STRB06, WBS^{+}04]. night-time [BBHT02]. nighttime [FLR05, KS04a]. nigrescens [FMCB05]. nigricans [GLP^{+}08, HK02a]. nigritus [PCB08]. Niña [HLGCO06, IMW^{+}02, JPL04, VSB04]. nine [Eyr00]. Ningaloo [MBP^{+}06]. Niño [CCL^{+}06, GSFN00, GBW02, GGM^{+}00, HLGCO06, IMW^{+}02, IG04, JPL04, KJMP05, Min02, RWA^{+}02, VSB04]. niphonius [ST06b]. Nippoleucon [AY04a, AY04b]. nitida [GDR^{+}04a, MDR^{+}06]. Nitrate [GDS02, PSE^{+}03, SKMS06a, SKMS06b, YLv^{+}05, AG02b, CT06b, DGW01, FBHB03, KGTH01, KYYM00, KP01b, LS05, LFSR06, MRC^{+}08, MN00a, MIH01, NWHC03, NWFV06, SYPO09, SMF00, YMC05]. nitrate-depleted [SKMS06b]. nitrate-rich [MN00a]. nitrates [PT04]. nitrification [ÁSG04, HH05a, MKH^{+}02, MHK^{+}04, PBW^{+}07, War05, CHSW03]. nitrification-denitrification [HH05a]. nitrifying [PBW^{+}07]. nitrite [AQHTL02]. Nitrogen [BR03, BM06, BWK^{+}08, BAE^{+}08, DR00, FR07, GF04, GTSS03, HIM05, KF01, KG00b, LFRR^{+}07, LGM^{+}09, MN00a, MIH^{+}00, NWHC03, NRPS^{+}01, PH03c, RPO00, RAE07, SPP^{+}09, SMG00, SCB00, TACM01, jMFVPL05, ABBN09, AG02b, AMT03a, BVH02, BCM^{+}09, BVCG01, BWCW02, BSvdH^{+}02, BS07c, CB00a, CHSW03, CMHP06, CKO^{+}07, CWT06, CVK00, CBE04, CRBE04, CRCV04, CVSM09, CvBGR09, CSD^{+}07, DRG^{+}07, DES^{+}00, DPD^{+}08, EB04a, EP01, FBKB03, GRS07, GHH^{+}04, GSW01a, GHV05, HARC05, HH03a, HKM00, HKM01, HCV03a, HHW^{+}07, HWL00, HO02, JMSM08, JF02, JFA02, KTNK07, KHWH03, KG00a, KW02b, KG07, LS05, LFF^{+}09, Lee00, LD00, LJ06, LMD^{+}02, LV05, LGZ^{+}09, LWK^{+}07, LLL^{+}09, MHM^{+}02, MCH^{+}06, MTV06, Maz04, MN00b, MSUK01, MSH^{+}03a, MPDS07, MIH01, MGL^{+}00, MLG03b, NNW04, NWFV06, NOO^{+}08, PLM03, PMNP^{+}01]. nitrogen [PK01, RMH^{+}09, RCAK03, SS02a, SLCB00, SHL^{+}03, SSS^{+}05, SHS06, SFR^{+}09, SBC^{+}06, TRH^{+}06, TGSW06, TC01b, TDUN07, TGM^{+}03, TBI07, VS08, WDX01, WPRS05, WB01, WVL01, WNS^{+}05, WBN^{+}00, WWWR09, CRBE04]. nitrogen-fixing [MSH^{+}03a, MPDS07, SBC^{+}06]. nitrogen-stable [Lee00]. nitrous [dBMSL02]. nitzschia [ABWK06, CG04, FDBT06, LRK09, LHK04, LHK05, MTH04, MP05, OLWB06, OPSM04, PPB^{+}01, RDD^{+}02, SPP^{+}09]. nm [KFE^{+}03, RBSP00]. NMR [COB^{+}03]. NO [KYYM00, DBT^{+}02a, HF09, HMKT02, BKR04, CHW06, DWB04, HEN^{+}08, KR05, LH08, LHGC^{+}09, LTCP03, PWC^{+}09, PPGL09, RSA05, SB03a, SSR^{+}03, SJA^{+}09, TPBF09, vDS05]. no-harvest [PPGL09]. no-take [DWB04, LH08, LHGC^{+}09, LTCP03, RSA05, SB03a, SJA^{+}09]. nobilis [PW08, PW09, UB02]. Noctiluca [DCAL^{+}02, HMA04, Kiø03, JM01]. nocturnal [FB03a, HOVG07, MB02a, NDV^{+}00a]. nodes [LR02]. nodosa [AMS01, AAHDS06, CPKV00, MD01, OEDSJ02]. nodosum [AVA^{+}09, DJM07, DJM08, PB00a, TP00a]. Nodularia [EÖKS^{+}02, KSKL^{+}03, SGK04]. noise [HF09, Hil09, JBW^{+}09, KCH^{+}03, MWT^{+}06, SDD^{+}07, SYWM05]. noltii [ASS05, BHV^{+}02, BVN^{+}03, BOM^{+}08, CAS05, CDS^{+}04, DAM^{+}08, DCF^{+}05, PBPLB06, PA06, WBN^{+}00, WPB^{+}08b, ZCR^{+}09]. nominally [CCC05]. nomurai [KUOI06]. Non [AFPS08, BLS04, GF05, GWF07, SMC^{+}05, AI02, ASA^{+}08, BGF^{+}06, BRJ^{+}03, BCS^{+}06, BMS04, CDH^{+}09, CVK00, CB08b, DKR09, DHZ02, EJKN03, FSW09, GSFN00, HSCH07, HH07a, HW01, ITE^{+}04, JBSB01, JPG05, JBS01, KLN06, KHP09, LB08a, LC07a, MK03b, MC05c, MRGC09, NvdV02, OSA05, PLG09, Rue07, RMAY04, SHA^{+}05, SBSPS04, SKDP02, TS04, TTL06, WB07, WVL01, WAK^{+}09, vdVBW00, AY04b]. non-captured [JBSB01]. non-carnivore [VRJ02]. non-depositional [SKDP02]. Non-destructive [GWF07]. Non-diapausing [AY04b]. non-dispersing [KHP09]. non-El [GSFN00]. non-emergent [TS04]. non-estuarine [NvdV02]. non-feeding [MK03b]. non-fish [JBS01]. non-Gaussian [PLG09]. non-genetic [vdVBW00]. non-geniculate [SHA^{+}05]. non-heterocystous [WVL01]. non-indigenous [JPG05]. non-invasive [BRJ^{+}03]. Non-lethal [AFPS08, FSW09]. Non-linear [BLS04, DHZ02, RMAY04]. non-local [HW01]. non-motile [HSCH07]. non-native [BCS^{+}06, CB08b, DKR09, HH07a, LC07a, Rue07, SBSPS04]. non-neurotoxic [ITE^{+}04]. non-panmictic [TTL06]. non-participation [MC05c]. non-pellet-bound [OSA05]. Non-recovery [GF05, CDH^{+}09]. non-saturating [EJKN03]. non-spawning [MRGC09]. non-suspension-feeding [CVK00]. Non-synergistic [SMC^{+}05]. non-target [ASA^{+}08]. non-targeted [WAK^{+}09]. non-tidal [AI02]. non-toxic [BGF^{+}06, BMS04]. non-turbulent [WB07]. non-vegetated [KLN06, LB08a]. nongregarious [TP01a]. nonindigenous [HTM01, LL03, Lar03]. Nonlinear [Ogu07, OSF08]. nonvegetated [RZ00]. Nonylphenol [MM05, NMCT03]. Nordic [DFR08]. Nordrügensche [KMR01]. norm [Kra07]. normally [SR08b]. Normanby [SR01]. norms [DH07a, GRB^{+}03, GHR^{+}07, MKR07, VMCG09]. Noronha [BKSS07]. North [EO04, GSS09a, HSHA06, tAHSHA07, KRP^{+}06, KW02a, LSHL03, NOE^{+}04, PPHT00, RLDL02, SAHK06, SF05, WW07, ADR^{+}05, HCM01, LCB^{+}05, LME05, OSL07, QA01, RHS01a, SC06b, SAM^{+}01, UB01, AC07, ABD^{+}03b, Ano07g, ARP02, AR02b, BCC^{+}03, BCCM03, BM03, BBG^{+}04, BRIP00, BIL01, BIL^{+}02, BI02, BI04, BPK^{+}00, BPS^{+}00, BEKT09, BM07c, CED^{+}07, CMG04, CIHL05, CFG06, CCL^{+}05, CR08, DRDM09, DLLB03, DSF^{+}04, DBH^{+}02, DMH01, DCM^{+}06, DPGN09, DSDS04, EC07a, EBR^{+}02, EWBH04, FBD^{+}02, FHL^{+}09, Fen08, FCI^{+}04, FBD06, FT00, GMHA09, GB07, GGFSP06, GP08, GRB^{+}03, GHR^{+}07, HH01, HPW^{+}00, HPH^{+}01, HHG^{+}07, HJ03, HSM^{+}09, HB07a, HBW^{+}05, HLY05, Hip09, HRMH07, HFM^{+}02, Hum00, HMO^{+}04, HCMR09, HW00, HCR01, HFA02, IMSO09, ISYM06, KFK^{+}07, KW05, KHIH08, KFT08, KH08a, KBL^{+}07, KWD05, KSWP04]. North [KW00b, LTFJ00, LTG^{+}01, LTMF01, LSG^{+}05a, LSdM^{+}03, LRN^{+}09, LR06, LSK08, LSE06, LWW^{+}01, LHHB05, MPSJ06, MRS00, MHH^{+}05, MLG^{+}03a, MR06b, MKWO00, MKR07, MBL07, Mun07, NRM05, NM04, NWHH^{+}09, NPN06, NGMT01, OCB^{+}07, PGBA04, PKS00, PPB^{+}09a, PPI04, RGMJM07, RJS04, RHS01a, RCC01, RMT^{+}05, Rin03, RGL06, RJBW03, RMRAFP08, RFG^{+}09, SSH^{+}04, SK04, SK05, SLS^{+}00, SAC^{+}00, SP01a, SFKT09, SP02, SCM^{+}09a, SAC^{+}03, SOB^{+}06, SKDP02, TSE^{+}08, TOK08, TKO08, TB08, TPQ^{+}03, TR06, TPS^{+}05, TBH06, VGB^{+}05, VMN^{+}03, WWHE04, WHRS05, WKH02, WKIK04, WKMK06, WKY09, WLZ05, WGV^{+}02, WWWR09, WWR09, YW04, vOSF^{+}09]. north-central [HCM01, LCB^{+}05]. north-east [QA01, UB01]. north-western [OSL07]. Northeast [SSJ^{+}02, Bla09, CCML07, CRB07, DPGN08, GL00b, LSBJ08, LG02, NSJ02, OÁSG^{+}08, VNJ06, BM09, DBG^{+}09, DD06, GSMD00, GL00a, NLHO09, PCRR08, VJKF07]. northeastern [ASLD09, CWM^{+}09, DDF01, EDKT06, FAPA^{+}09, FAPOF09, GSK^{+}08, MIH^{+}00, MILH00, OJPM01, PCB08, SDW^{+}01, SDK^{+}02, SMFP08, VRJ02, Wal07, vDHHM08]. Northern [BCP^{+}08, TJSG03, UB03, ARAMPW02, AMS01, AO01, AAF^{+}04, Arr05, BAA07, BBBSPH03, BVCG01, BTB^{+}09, BY09, BT04, CDO08, CMG04, CH07b, CGF04, CCL^{+}04, CDH^{+}09, CSS00, CES^{+}03, CGH^{+}03, CCL^{+}06, CGC^{+}09, DFPG^{+}03, DFSS08, DH07b, DBS05, DKFV05, ENAT07, FMCB05, FZT^{+}05, GAMB05, GEM02, GC06a, GSFN00, GEHZ07, GPSG06, GS02, HM03, HWGB02, HALL05, IG04, IRL^{+}09, JN09, JCMR^{+}03, JCL09, JH06, KWV^{+}05, KGF^{+}09, KW02b, LTPL05, LSH^{+}07a, LTV^{+}01, LSdM^{+}03, LG08a, LVvdLH07, LJM^{+}00, LZM^{+}03, LKW03, MFE^{+}05, Mon07, MA07, NSJ02, NSEL00, NEK^{+}07, NLCS05, NGMT01, OW02, OSL07, PPB^{+}01, PCL^{+}09, RWR^{+}00, RWE04, RV05, Rol00, RSZI01, RRRM08, RSSA07, RH08, SSR^{+}09, SESP03, SSH^{+}04, SAJW04, SCT^{+}09, SCKC07, SEH07b, Som00, SBBS^{+}02, SHT^{+}04b, SMVF03, SB08, Tar03, VMN^{+}03, VV04, WISH09]. northern [WHL^{+}08, WCF08, WAW^{+}09, Wie03, WCRB03, WRJ^{+}05, YHSH02, YW04, ZBDE07, BYH^{+}06, SAB04, SPC^{+}00]. northern-central [PCL^{+}09]. Northwest [BRP07, OLWB06, Tal09, TFLA03, AL09, BW01, BCABC00, BCBAC01, HB08a, HK04a, HKK^{+}06, MGE^{+}09, SLM^{+}02, TTL06, WPIB09, BCHS09, CS02a, KFL09, MLS^{+}08, RMC^{+}05]. northwestern [BCL03, BGLL09, COPB06, DGBS08, FFYS^{+}07, FD02, GF04, KRVB05, LP03, MGA^{+}01, OGL^{+}09a, SYVC00, TSIY06, VGS^{+}01, VYM08, DF04]. norvegica [AM03, KLHO02, LTV^{+}01, PZB^{+}05, RJBW03, STPA04, SK03, SS02b, Tar03, TJSG03, TNT06, Tor01]. norvegicus [ASA^{+}08, ASA^{+}03, CFG06, FCC^{+}08, PWAT01, SMA^{+}03]. Norway [NH05, ASA^{+}08, FCC^{+}08, FZT^{+}05, HB07b, JN09, LMKG01, NPN06, PWAT01, SHH09, WBG09]. Norwegian [ADS^{+}09, ASES07, Enb05, EH04, EAF^{+}05, Fre03, FBSG02, GLS^{+}00, Hel00, HBN01, Joh02, KTK^{+}08a, KFE^{+}03, KNF^{+}03, PTS00, RWR^{+}00, ZZT05]. norwegica [MVN^{+}06]. nosed [PCW05]. notatus [FOH^{+}08]. notes [KI01, Ano03g]. Notorynchus [LME05]. Nototodarus [JSWH03]. nourishment [WCRB03]. Nova [BS00, GJT09, SM07b, SM08a, SG06b, SG09, SS05]. novaeangliae [CH00b, FHN^{+}09, HFT^{+}09, HYR^{+}09, ZAHJ^{+}06]. novaezelandiae [TNTH04]. Novel [HJH^{+}02, LCT^{+}08, BRJ^{+}03, FCC^{+}08, FK00, GMM03, GT08, HC02, LCC07, MGBB06, PCS^{+}04a, PBF07, PBS^{+}05, RC05, ZER^{+}01]. nowcast [JPG05]. Nucella [CH07d, DD07, GR07a, HPC04, LMW^{+}01]. nuclear [PG01, PE09, SWL^{+}06]. Nucleic [GEHZ07, PWAT01, WCBD01, Gor03, MCR03]. nucleotide [PG01]. nudibranch [LH06]. nudibranchs [RWSP03]. nudus [ENAT07]. number [AG02a, MBVBC09]. numbers [GM01, HW02, HW00, SAJG06]. Numerical [LNB^{+}03, ANCC07, AHZ03, DR01, JWG01, LJK01b, LJK02]. nurseries [BS06, CBL^{+}09, DKA^{+}06, GAB^{+}03, LDK^{+}06, MAWH03, SBJ06, VRST^{+}07]. Nursery [ADK^{+}06, KMHS04, BS05a, Bro06, CSM^{+}09, DE01, DMK^{+}08, DH06, EE00, FH08b, FLL09, FS02, GRWB02, GK00a, GBW02, HHO03, HHH02, HCS07b, Hid03, KS05, KCSPT09, LC02, LR06, MPM^{+}02, MPSJ06, MHR03, Mok04, NKK^{+}01, NRvdV^{+}02, NGBG07, PDHH06, PPSP07, RSZI01, SC00b, STM06, Sto03, Tup07, WGR07]. Nursery-area [KMHS04]. nursery-role [KS05]. nursery-specific [LR06]. nursing [WLS^{+}09b]. Nutrient [AMR00a, BWB^{+}06, BH09b, CPK00, CZW^{+}08, DDSR07, EKBM05, FKBB01, HGMG^{+}04, KBH01, MM08, RHGB02, VD00, XYH^{+}09, ADA00, ACLF09, APP^{+}09, BGLL09, BPE^{+}06, BMSH02, BWP^{+}06, BSAR03, BSR05, CHW02, Cam01, CGP08, CH07b, CLL^{+}09, CRK06, CSD^{+}07, DP02, DY05, DAG^{+}00, DAA00, Eyr00, EF02, FAPA^{+}09, FBKB03, FBHB03, FH08d, GFP^{+}05, GM08, GJ03, GCHC00, GPM^{+}09, HDP02, HDC^{+}08, HN03, HVP^{+}06, HHJ04, HSHA06, tAHSHA07, HGJM09, HFD08, HGT01, IOTL09, IRL^{+}09, JHTV02, JK06, JMdG^{+}06, KOTL07, Kar07, KWTR08, KM01, KUN02, KG00b, KFN01, KKSR05, LSH^{+}07a, LSG^{+}05a, Lar03, LV05, LJK01b, LLL^{+}09, MBW02, MBW03, MMTT00, MBC^{+}07, MRP09, MG00, MK03c, MLG03b, NPM01, OAGV07, Pah05, PPHT00, PKRK07, PTA^{+}09, PAR08, PPCJ^{+}06, RBRH02, RPN^{+}01, Ree03, RR05, RL03, RC07, SGL02a, SW02a]. nutrient [SKMS06b, SCS03, SCD^{+}09, SHM02, Ste04b, SVL^{+}04, SMVF03, SOMM06, SCC06, SMH^{+}03, TA07a, TFA^{+}08, TDUN07, TDSM02, TT02, ULL^{+}05, VP00, WE04, WLdG^{+}04, WDM^{+}09, WWH05, WR00b, WS00, YH00, ZHWCS08, MCH^{+}06, MCCT07b]. nutrient-based [BMSH02]. nutrient-enriched [Cam01, OAGV07]. Nutrient-induced [HGMG^{+}04]. nutrient-limited [ULL^{+}05]. nutrient-load [GPM^{+}09]. nutrient-rich [HDC^{+}08]. nutrients [AF06, BP02b, CSMK09, DL01, FKBB01, GXL^{+}08, GSW01a, GMTN00, HBRW04, KPT01, LD03, MSS^{+}03, MPB^{+}09, NLB08, NFV^{+}06, RMGL07, RDC06, RS07b, RU02, RC05, SSB09, SEL02, SMHL00, SSLM01, YQC^{+}00, Yin02]. Nutrition [FB00, BKR04, CHG05, ET08, GPB02, LJM^{+}00, NA04b, SPP^{+}09]. Nutritional [CCC05, KKMH08, MFCM05, NGSV07, BGLL09, CES^{+}03, HK03b, JRT^{+}06, LLJ05, MCW^{+}07, MMBT03, PWAT01, SWH04a, TBAT02]. nutshell [HR02b, HR02c]. Nuttallia [DD06]. NW [BMG00, CLR^{+}06, DLB04, GPSH01, GHL02, HKH04, JKNJ07, KSSH02, KHIA02, NIF00, PSRB00, TAÁO^{+}03, VPP04, ÁSG04, ARBF02, BGQR01, BA00, BPVL03, BG07a, BGH06, CSA02, CCC09, CBS^{+}08, CSP^{+}03, CMRB08, CTFC08, DSPHV04, DFH05, DR00, DRB^{+}01, FHVMDS07, FGA^{+}04, GBdS^{+}05, GMd^{+}02, GGTKG09, HLH04, HPC04, ILL06, JERS03, LCD^{+}05, LAV04, MMV05, MPM^{+}03, NOE^{+}04, NCÁSBP04, NCÁSGP05, OSP01, OÁSG^{+}09, PIY04, PNCS^{+}08, PNCT^{+}08, PPCJ^{+}06, RGC^{+}06, RHGB02, SAMN04, SCS^{+}07, Tay05, TSF01, TPBF09, TS06, TLIL^{+}09, VGMC01, XZUK05, dJTD07]. Ny [MAJ^{+}06]. Nyctiphanes [GGR06, HB03]. nylon [KCTG06].

O [BW01, CGF04, KSK07, KGB^{+}01, MRHP01, MD07, RSY^{+}06, VVTH07]. Oahu [WWMF06, Zab09]. oases [SGP07]. OAWRS [JBS^{+}09]. OBI [CVB06b]. OBIS [HRB^{+}06]. OBIS-SEAMAP [HRB^{+}06]. obligate [RH04]. obscurata [DD06]. obscurus [HCD^{+}09, KKMH08]. Observation [SK02, DHG^{+}04, FHK^{+}08, TFL^{+}04, VPG04]. observations [BAB^{+}01, CDW^{+}09, DFW^{+}03, GGDB09, MGL^{+}07, MHBB05, PNK^{+}02, PWL^{+}06, TF04, Trn02, LJM^{+}00]. observatory [MG09a]. observed [BBABL01, EB08]. observer [WWTH06]. observers [NDOS03]. obtain [RSR02]. obtusus [DER01]. occupancy [FFA03, OHS^{+}07]. occupation [MD01]. Occurrence [ACM^{+}02, IMW^{+}02, MHW06, Ned04, SAW^{+}01, ADF^{+}09, AHW03, BP08, CGM00, CTvdM^{+}08, DDS^{+}01, DTAF00, Hen09, HD05, KKTN05, KAAH05, MGE^{+}09, PPBS02, RJS04, SS06, WGH^{+}00]. occurrences [GBI^{+}06]. occurring [CAT^{+}06, CCK03, PBS^{+}05, RSN02, SJK^{+}06, VPS08, WG02b]. Ocean [AS04, BF08a, BPB07, BJN^{+}05, CVB06b, EBR^{+}02, FGLM06, FR08a, GMA^{+}09, GFPM01, HWBZ05, Hue08b, JBS^{+}09, KPM00, MS04b, MBB^{+}04, NBB^{+}04, PFW^{+}02, RAC^{+}05, RCP^{+}05, SGH^{+}06, ANCC07, AMB08, AC08, BD07, BPA^{+}08, BVHG08a, BVHG08b, CL04, CB08c, CCCD09, Den08, DSP07, DSM01, DÁSG^{+}01, DRG^{+}02, DHT^{+}08, FFYS^{+}07, FWD^{+}01, GTD^{+}08, HRCB02, Hil09, HFWW05, Hip08, Hip09, HOT08, HGKK06, HTBT07, KAAH05, KB06, Koe05, Kur08, LLM^{+}08, Law08, LR02, Lei08, LCE03, LBB08, MHB^{+}01b, MMA^{+}09, MCD^{+}05, MS04a, MRQ07, MAKMK09, MCN09, NEG01, Orb08, PTD05, PDT09, PSE^{+}03, Pör08, PW07, RZWG08, SWD^{+}04, SYPO09, SMH02, SCD^{+}05, SN09, SNCP04, SMF00, STRB06, TMD^{+}07, TCS^{+}05, TDTR03, TSF04, WBT^{+}08, WH08, WFT^{+}08, dBGLT08, ASLD09, BAM^{+}04, BF03, BI02, BW01, BSA^{+}08, BCL^{+}01, BdBK^{+}00, BKSS07]. Ocean [Bur02, CCC09, CBGLM08, CRJ^{+}09, CRLP09, CBTW02, CH07c, CAS00, CV06, CWN^{+}07, DGJ^{+}03, DÁSG^{+}01, DPGN08, DD06, EL05, FR07, FDD^{+}03, Fie02, FBBH07, FPWM03, FVS^{+}08, GMA^{+}09, GVF^{+}03, GF05, GAM02, GTGM06, GUL^{+}07, GMD08, HKH04, HB07a, HBN01, HVWH06, IMSO09, ISYM06, JMM^{+}09, KFS^{+}09, KVV^{+}00, KvNVH00, LLM^{+}08, LGC^{+}06, LSE06, LSM^{+}02, LGM^{+}09, MG09a, MBG^{+}03, MFS07, MRPL07, MAG^{+}07a, MGW^{+}09, MGREA07, MAW01, MRQ07, MZD^{+}05, MVB^{+}05, MR01, NMOO03, NLP^{+}01, NCMC09, Øre00, OÁSG^{+}08, PIY04, PSW^{+}08, PPC09, PPB09b, PMNL07, PVV^{+}06, QMF05, RW03, RPN^{+}01, RBB^{+}08, RLFEF03, RRH09, RSZI01, SC03, SFF^{+}07, SHS08, SFKT09, SCGC07, SCM^{+}09a, TSIY06, TTD^{+}06, TBK^{+}03, TTL06, VMRL04, VW00, VDA^{+}03, WRW06, WWWC02, YI03a, ZAHJ^{+}06, vDHHM08]. ocean-colour [SWD^{+}04]. Ocean-estuary [MS04b]. Ocean-scale [SGH^{+}06]. oceana [SL05d]. Oceanic [ASP^{+}08, HLFG^{+}07, ALOT01, BPH07, BN09, BMSH02, BBV^{+}01, CB01, DDQ^{+}00, FCS^{+}09, GSB^{+}06, Joh03, LCL^{+}03, LJ06, LC08, LC09b, LYMCG06, MTH^{+}08, MHSS08, OJPM01, PCRR08, RFG04, SP04, SMSJ04, TND^{+}05, TCWR08]. oceanica [AMR00a, AMR01a, BCL03, BGLL09, BD09, CRS^{+}07, DGM02, FJV^{+}02, FHDA^{+}07, GBMHD08, GCSSJV07, HSBD08, LMD^{+}02, MDSP05, OEDSJ02, PTAR07, PAR08, RP02, RR01, SPI^{+}07, TTR05a, TTR05b, VBAR07]. Oceanodroma [HM06a]. Oceanographic [AP05a, GSRH02, HRR^{+}07, HFA02, SAB^{+}02, BSH^{+}09, BFE^{+}05, CCC09, FHQ^{+}06, GDL^{+}01, IMSO09, IMTW09, JWR05, LB05b, Mon07, NLP^{+}01, OSL07, RTPS07, STF^{+}09, TKVL04, WWWY07, Yin03, ZBNW09, dSCV^{+}08, Chi09]. oceanography [AML^{+}03, DS06b, FJK^{+}07, RS03, SCKC07, SBI07, SIK^{+}08, WR06, ZJW^{+}04]. Oceans [BRSM07, BB08b, IHK08, LEF^{+}06, CEBR07, TCW^{+}08]. ocellatus [FOFG09, HHH02, HH00b, PDHH06, SML02b]. ochraceus [BV08, SM07a]. Ocinebrellus [MVBGM04]. octocoral [ACPS09, Cai09, EMPF07, GC03, HBC09, Kum06, LGRB^{+}08, MW09, OLGG^{+}02, RCR03, SSRG05]. octocoral-algal [GC03]. octocoral-associated [Kum06]. octocoral-gorgonian [RCR03]. Octocorallia [CLR^{+}06, DS09b, PCS09, RCR03]. octocorals [BR07, Gou07, GSG08, OGA03, TPBF09]. octopus [APG^{+}09, BL09, DPSD08a, LSP08, OÁSG^{+}08, BL09, DPSD08a, DPSD08b, OÁSG^{+}08, OÁSG^{+}09]. octopuses [AGS^{+}09]. oculata [HAMW09]. Oculina [FOL02, FIRL02, RMRAFP08, SFL05]. Ocypodidae [KWD05]. Ocyurus [WMG02]. odd [Kap09]. Odontaster [MAH^{+}08, PWM^{+}08]. odor [BD07, Fin00]. odorabile [WNWH02]. odors [NPLZ00]. odour [AKG02]. off [AP06, Ano01a, AAY^{+}07, AB06, ANF^{+}00, BF04, BDH01, BRP07, BFE^{+}05, BHD04, CSR^{+}00, CL04, CGF04, CSH00, CIHF01, CES^{+}03, CTFC08, DE01, DDQ^{+}00, DSPHV04, DBS05, EB04b, GSM06, GEM02, GSFN00, GGM^{+}00, HGG^{+}03, HLGCO06, HQDH06, HK04b, HGZ^{+}00, IMW^{+}02, JPL04, KAAH05, KJMP05, Kor01, KGHL^{+}01, LC02, LH09, LF06, LBC^{+}09, LMB^{+}09, MFG^{+}04, MVHV06, MGE^{+}09, MSSP06, MC08a, MFE^{+}05, MPE03, MBL07, MBG^{+}08, NNLV06, NAB^{+}06, Par06, PHM^{+}08, Reb01, RHS01a, RWE04, RSNG09, RT00a, SHH09, SSHW^{+}07, SDW^{+}01, SSGJ00, SPP^{+}09, SKE05, SMVF03, SYCG05, TAÁO^{+}03, TDC^{+}03, VMH04, WDT^{+}08a, WWL04, WKL01, YCSS06, dSSC07]. off-reef [DE01]. Offense [BLS03]. officinalis [BTF^{+}02, BTD^{+}04, GPM08]. offs [FSB07, Hal00, HP07, TSSA03]. Offshore [GDR^{+}04b, ONKK03, TKVL04, ZT07, BP02a, BGH06, CHT06, DGBS08, DWB04, DNL08a, FF02, GSS09a, HMK06, HLH04, JT05, KST^{+}08, MGPAE01, MPE03, MGO^{+}07, NCN^{+}09, PDCH06, PDS^{+}07, PDQ01, PGPL03, RJS04, SAB04, SFW01, SIK^{+}08, TOK^{+}04, VLW09, WW05, dSSC^{+}08]. offspring [FHB^{+}02, OAM06]. ogac [LGB03b, LGB^{+}04]. Oikopleura [ATT06a, FLUF^{+}04, LSG05b, LUA99, SVL^{+}04, TCG03, TBAT02, TGNT05, URFA06]. Oil [Rui04, AGCF^{+}08, AWA^{+}05, BBD^{+}02, CTLR03, CHMR05, CMB^{+}05a, DBJ^{+}00, DBF^{+}02, DLD^{+}00, EBT^{+}02, GSM^{+}02, HCP^{+}06, HMD^{+}07, LFRG^{+}05, MAVO^{+}06, NAB^{+}06, OGMC06, PDCH06, PDS^{+}07, PMGE01, PHB02, PBB^{+}02, PD01, RMT^{+}05, SRDPT^{+}07, SAB04, SMC^{+}05, VML05]. oil-contaminated [CTLR03]. oil-exposed [CHMR05]. oil-related [AWA^{+}05]. Oithona [CIHL05, CRSL05, DD09a, DD09b, EO04, HSCH07, HW08, IP07, JP08, MVN^{+}06, NMS^{+}02, RRS05]. Okinawa [KST^{+}08, MT00, RSY^{+}06, SKH06]. old [Reu01, RATD07, SP08a]. Oléron [LNPB04, PMG^{+}06]. olfactory [FSW09, KPW03]. oligochaete [NLK00]. Oligochaetes [ARFP09]. oligotrophic [FMMS03, GAM02, HSCF06, HGKK06, KHGA02, MBG^{+}03, MCP05, MBC^{+}07, MFP04, PG00, RF08b, RC07, SLGD00, SN04, SHT^{+}04b, TBC^{+}00]. oligotrophy [VGS^{+}01]. olivacea [MBH07, SAM^{+}06]. Olive [WH09, MBH07, SAM^{+}06]. Olive-headed [WH09]. omic [Via07]. omissa [CCdAR00]. Ommastrephes [CCC09, Par06, WKIK04]. ommastrephid [Par06]. Ommastrephidae [DGC^{+}05a, IAKS03]. omnivorous [JP04, KLHO02, PABM08, VSL^{+}04]. omnivory [CDO08, FTO^{+}09, GRBB07]. onagawae [ITN^{+}06]. Oncaea [BSS05, SHT^{+}04a]. Oncholaimus [TVG00]. Oncorhynchus [CHMR05, HFWW05]. one [TWMW06, Und04, BKDB00, Boo02, CS03a, MB02a]. onitis [KMAC^{+}02]. online [FGLM06]. only [BAB04, GT08, SG06a, VGW00]. onset [CCCD09, PFW^{+}02, RMH07, RLKW04, WHD^{+}05]. Onshore [MPE03, NNLV06, BP02a, DRTM09]. Onshore-offshore [MPE03]. onto [CD07b, TC01a]. Ontogenetic [ACAG07, ABL08, BCLS06, FR06, FB02a, GRWB02, Gru00, IT06b, LÁMBF09, MS06, ML07, NWJ^{+}08, OT08, QTB02, TT05a, TF04, YOM^{+}09, dlMPN^{+}03, BGPL08, BTB^{+}09, CSG01, KFS^{+}09, NHS^{+}08, NEMF09, YHSH02, SL05c]. Ontogeny [CGCDA02, FOFG09, LHL^{+}07, LHH09, LTL06c, MFY02, BGGR02, BC06a, CACDC07, CLHT05, LSH07b, Lei07, MOP^{+}08, SJR00]. Onychoteuthidae [CD03b]. onyx [CNW^{+}07b]. oocyte [FWF02]. Oosterschelde [RSN02]. opah [JBG00]. opalescens [WHS^{+}09]. Open [FWD^{+}01, DBG^{+}09, DRG^{+}02, EHK06, FS02, FBM07, GTD^{+}08, GTE^{+}09, GD07, KPN03, KFU^{+}08, LBB08, LCB^{+}05, MHB^{+}01b, MP06, NPM01, SPAMG00, SSW08, SNCP04, SMF00, SJH^{+}04, TMD^{+}07, YVB^{+}02, YMC05]. open-coast [FS02, SSW08]. Open-ocean [FWD^{+}01, DRG^{+}02, GTD^{+}08, LBB08, MHB^{+}01b]. open-sea [SPAMG00]. open-water [EHK06]. opened [PZMG05]. operated [PNK^{+}02, TFL^{+}04]. operation [LBBM06]. operations [WCRB03]. opercularis [KMHS04]. Ophiactis [McG02]. Ophionotus [GTPC04]. Ophiopholis [DHR04a]. Ophiothrix [DHT^{+}08, TCT^{+}00]. ophiuroid [DHR04a, LFP^{+}04]. Ophiuroida [OKP06]. Ophiuroidea [SCC05, SDR01, TCT^{+}00].

Ophryotrocha [SP03b]. opilio [DSMBG03, RGT01, SML02a]. Opisthobranchia [RWSP03]. opportunist [AP04a, AP04b]. Opportunistic [MLS^{+}06, KT06, LHB09, MBG^{+}09, RF08a, TBAT02]. opportunists [CRGH01]. opportunities [AL08, Tho07]. Opportunity [SNP03]. opposed [MPDC03]. Opposing [FWC02]. Opposite [YG09]. Optical [EZTZ02, HTBT07, BPK^{+}00, BPS^{+}00, BPS^{+}03, DHG^{+}04, KFE^{+}03, LPBD00, LSHL03, RSK04, SLCB00, SSH^{+}00, VMN^{+}03]. optima [AJ03, TOK08]. Optimal [KH08b, PJB03, SYPO09, VFF09, Pah05, SDK^{+}02, TT07a]. optimise [TLFS06]. Optimising [GML02]. optimized [BMT07]. Optimizing [GNL08, FSB07]. optimum [SW08a]. optional [HBBH07]. optode [MD07, SM05]. opuntia [NDB04]. Oran [VLP^{+}04]. Orange [AHDT^{+}09]. oratoria [KHK^{+}06]. Oratosquilla [KHK^{+}06]. orca [FE06, HLH^{+}07, HBW^{+}05, HMY^{+}08, MAC^{+}07]. Orchomene [DER01]. Orcinus [FE06, HLH^{+}07, HBW^{+}05, HMY^{+}08, MAC^{+}07]. order [Bra08, GH08, NYK03]. ordinated [CKS05, PC05]. Oreaster [MSY02]. Oregon [AB06, BFE^{+}05, BPA^{+}08, CSR^{+}00, HWBZ05, KJMP05, MS04b, MS05, MPE03, RAS07, WWL04]. Øresund [GHP03]. organ [Tak02, TTM^{+}05]. Organic [HKY^{+}02, HBK^{+}05, LCB^{+}05, MK02, WMW^{+}08, vOSF^{+}09, ABBN09, AHL^{+}01, ARFP09, BD09, BCM^{+}09, BVCG01, BIG^{+}06, BMSD00, BMO^{+}04, BEE06, CRS^{+}07, CMG04, CSMK09, CC00, CRCV04, CGL^{+}02, DDF01, DSPHV04, DMH01, DR00, DÁSG^{+}01, EKBM05, GSL^{+}06, GHH^{+}04, GSW01a, GHS05, GWO02, GGRC^{+}08, GTSS03, GGM^{+}00, HRR^{+}01, HDC^{+}08, HKM00, HBT^{+}08, HCCY02, IAT^{+}07, JHTV02, KCP^{+}07, KHD07, KTK^{+}02, KKB^{+}06, KKSR05, LBF^{+}02, LT00, MMCF07, MMC08a, MCP05, MGA^{+}01, MT00, MdL06, MSD05, MLM^{+}02, MIH01, NN06, PRB^{+}08, PW04, PDFD09, RF08a, RHG05, RMKR00, RNDF^{+}04, RE00b, SSB09, SR09, Sar06, SFW01, SAT^{+}03, SS02a, SYVC00, SCD^{+}09, SHT^{+}04b, SZBZ04, SNCP04, SBC^{+}06, SCH^{+}07, TYT^{+}04, TSF01, TPQ^{+}03, TKTJ07, TNSM00a, TNSM00b, URFA06, VK02, VGS^{+}01, VLP^{+}04, WAM^{+}08, WKOE09, WRJ^{+}05, WHCP^{+}04, YYK^{+}08, dGvdBvO^{+}08]. organically [KHGA02]. organics [GMd^{+}02]. organisation [HM04, KPA^{+}09, Kin02]. organism [BVMBC07, GBI^{+}06, LTH^{+}06, SV01]. organism-generated [SV01]. Organismal [GDSB04]. organisms [BBM^{+}02a, CH06c, FNB07, FH08c, FGS^{+}02, GRS07, Gim09, HOT08, JBSB01, JK03a, KP01a, KHB00a, KFH04, Kim02, KST^{+}08, LVZW06, NDOS03, QFJ^{+}07, QHD04, Via07, WB05a, WB05b]. organization [AO01, GDF03, JWTP06, NW06, SKR09]. organizing [BF07]. organochlorine [FHG^{+}03, HBW^{+}05]. organotin [BGQR01, BMB02, RBG05]. organs [LP04, MOP^{+}08]. oRia [BMG00]. orientalis [KNK^{+}00, KNKT01, SL05a, TSIY06, TL05]. Orientation [GC01, LCE03, AKG02, FDDS09, FWD^{+}01, Har08, KPW03]. Origin [AP05b, LLJ05, ABS^{+}09, BSBD^{+}04, OBG^{+}07, PKH00, Sar06, SLB07]. originating [HTZ^{+}07, TYT^{+}04]. Origins [PAB^{+}02, BM00, BM02c, BCM^{+}07, SKFPS02]. Orkney [PSPZ09]. ornata [TCC04]. Orthoptera [MLGG03]. Oscillation [KH08a, SOB^{+}06, TBH06]. oscillations [BJN^{+}05]. oscillatory [TC03b]. Oslofjord [WBG09, KFN01]. Osmerus [BDRM06, PA07, SD00a, SD00b]. osmolality [RB02, RB05a]. osmolytes [EOL^{+}05]. Osmoregulation [CACDC07, APW^{+}04, CGCDA02]. osmosis [KJ03]. Osmotic [SSL00b]. Osmundea [OST07]. OSST [FJ01]. Osteichthyes [ORD00]. ostracod [MvdBÓ00]. Ostrea [CSTC01, TBR^{+}03]. ostrina [GR07a]. Otaria [CD08, HK04b]. other [AWOO02, GGTKG09, HTM01, Hue08b, IE06, KMHS04, LVZW06, PP02, SWD^{+}04, SPD03]. Otolith [AAJ07, BCB08a, FFG03, HS08a, HJG03, HW09, Hum00, KHP09, Tak02, WM04, AKL^{+}04, BS07b, BTKG^{+}08, CIT04, EdPP^{+}00, EG03, FH08b, FBJ00b, GM04, HYLT03, HMHB03a, HMHB03b, HNMC09, JBH00, JSIT02, JSIT06, MW06, MAM^{+}01, MC01b, NSEL00, RMH07, RSZI01, SRJ09b, SRZZ01, Sha00, SCIT09, SR03, STCJ01, SH04, SSW08, SSM^{+}01, TTM^{+}05, TSI02, VRST^{+}07, vdMM03]. Otolith-based [Hum00]. Otolith-check [WM04]. otolithic [LP04]. otoliths [BW01, Ber00, Bro06, CW03a, CT06a, CWM^{+}09, FS02, GK00a, HKZ04, KSK07, KSHP00, LIT05, MT05, MC05c, PKM04, RSD^{+}08, RHH^{+}05, RW06c, ST02, SJ06, TCW^{+}07, WT06, WFC03]. otter [BBD^{+}02, DBJ^{+}00, HKK^{+}06, KB04, RL09a]. otters [DBF^{+}02, Kon00]. Otto [Ano00d]. oualaniensis [Par06]. Oulastrea [Lam00a, Lam00b]. Our [Clo01]. outbreak [JBHGB04, Nug02, SG09, WM05]. outcome [SKP^{+}07, ZE01]. outer [AAGS02, HMY^{+}08, SFR^{+}09, TNSM00a]. output [CNW07a, CPG08, FFR^{+}03, MH00, McI02, RWH00, SM07a, TBAT02, ZMN07]. outputs [BDH01]. ovalis [KWTR08]. ovata [GDR^{+}04a, KFA^{+}04, MDR^{+}06, MDBG07]. over-exploitation [Ogu07]. over-harvested [SRC00]. over-wintering [FKI^{+}09]. overall [GEC08, JRP01]. Overcoming [Kir02c, Pet02a, Kir02b]. overexploited [SCMB03]. Overfishing [Das02, CHM06, VLRJ07]. overgrazed [JMU09]. overgrowth [Bul06]. Overlap [CDK^{+}07, PRL^{+}09, Bar02a, BDF09, CC02, CTvdM^{+}08, FMZvW02, GJ05, HH00a, KWP07, KFT08, PS01, TKO08, WHL00]. overlaps [NB06a]. overlooked [TdMF^{+}08]. overlying [ATR^{+}08]. oversights [WR07]. Overview [SPB09, Hue08b, PM00, SMMM07]. overwhelms [HH07b]. Overwinter [TSE^{+}08]. Overwintering [HGHB00, HBBH07, SLB07, CBD04, EBG^{+}00, KRK09]. ovigerous [CHL^{+}05]. oviposition [Swa04]. ovotestis [SF05]. Owenia [GBD^{+}06, PSRB00]. owl [KM03a]. own [HMA04, PB00b]. Oxic [JKG^{+}05, CGCMPA01]. oxic/anoxic [CGCMPA01]. Oxidation [VPWE09, DSS04, Kin01, TBK^{+}03, TRPNE06, VFJ06]. oxidative [YBY^{+}09]. oxide [dBMSL02]. oxidising [ISSK03]. oxidizers [BCM^{+}09]. oxidizing [PPI04]. oxodGuo [ABB^{+}00]. Oxygen [BRJ^{+}03, JGH05, KKI^{+}06, OYNO08, Paf06, RAA02, SDW^{+}01, ARAMPW02, BFB01, BFV08, CGT^{+}00, CGF04, DCT^{+}04, DMH01, DD07, DBP04, DRG^{+}07, FPB02, FHPW07, GCHC00, GB05b, GGM^{+}00, HH03a, HGK02, JKG^{+}05, JHR05, KRK09, Kar07, KYML02, KKP08, LF02, LGVF00, LSG05b, MBW02, MFE^{+}05, NR00, PKR07, QRWJF07, RFPS^{+}02, RWJ02, RAH^{+}02, SA03b, TSF01, TNSM00a, TT07b, WBN^{+}00, WBW^{+}06, WK00, WO05, MFE^{+}05]. Oxygen-minimum [SDW^{+}01]. oxygenation [LC03]. Oxygenic [SSdWM00]. Oxymonacanthus [KN00]. Oxyrrhis [JKS^{+}01]. Oyashio [TW04a, TW04b, KI01, PIY04, YI03a, YI06]. oyster [BCJ05, BC06a, BSK06, BCS^{+}06, BGD^{+}03, BH09a, CSTC01, CBMB03, DCO^{+}08, DVLH^{+}00, DGB^{+}02, FBN^{+}07, GPHC09, GP04b, HLH^{+}08, KLR^{+}06, LNPB04, LPBD00, MGDP01, Maz04, MWMH05, NSH^{+}08, PGP03, PPGL09, RRS00, Rue07, TBR^{+}03, TB08, TKN00, WDRH07, YIY08]. oyster-culture [HLH^{+}08]. oystercatchers [WGCM^{+}03]. oysters [Bay02, DOML^{+}07, LNV^{+}03, MSF^{+}07, MM08, PCS04b, YLK00]. Ozone [HCP^{+}06].

P [ABB^{+}00, GXL^{+}08, LSH^{+}07a, LCC01, PO09, XYH^{+}09, YQW^{+}01, SRLA04]. P-contaminated [ABB^{+}00]. P. [ACAG07, CBDH02, GNHB^{+}05, HRR^{+}01, MNC^{+}02, SLKF03, SHH08, WHL00, YLK00]. P450 [DCT^{+}04]. P4501A [CHMR05]. paced [DS01]. Pachygrapsus [MSP^{+}06]. Pachyptila [CBDH02, GLG^{+}05, SGL^{+}02b, THS^{+}02]. Pacific [CCC09, DGJ^{+}03, DKFV05, DD06, EC07a, Fen08, GZG^{+}07, GUL^{+}07, IMSO09, ISYM06, MVB^{+}05, PIY04, PPC09, PPB09b, PZMG05, SFKT09, SPRK02, SCM^{+}09a, THS^{+}02, TBK^{+}03, VMRL04, Van03, ABF07, AH08, ACM^{+}02, ASLD09, BM09, Bay02, BBG^{+}06, BRP07, Bla09, BL00b, BNSG01, BHN09, BM07c, CJFM09, CBGLM08, CT06a, CRB07, CEBR07, DC06, DHHY09, DFSS08, DNL08a, DGC^{+}05b, DPGN08, DPGN09, DM09, EE06, Fie02, FBD06, FKM^{+}01, GRPE08, GRS07, GSK^{+}08, GF05, GSP^{+}02, GGFSP06, GBW02, HGSH08, HSPH05, HBW^{+}05, HVCT09, Hip09, HMA06, Hum00, HSSS05, HFA02, IT06a, JTVC08, Kad03, KHIA02, KNK^{+}00, KNKT01, KI01, KAM^{+}06, LBF^{+}02, LOT^{+}00, LCL^{+}00, Lar03, LR06, LSK08, LLJ05, LSBJ08, LSM^{+}02, LLCHG01, LCW06, LS06, LCCTFO03, MSF^{+}07, MFS07, MAM^{+}01, MRPL07, MAW01, MMA^{+}04, MR06b, MKWO00]. Pacific [MF07, NOE^{+}04, NMOO03, NCMC09, OBG^{+}07, OCB^{+}07, OLWB06, OPSB07, Pra01, PPS00, RBB^{+}08, RRS00, RC09, RSZI01, SRL^{+}02, SK04, SK05, SV07, SCIT09, SP08b, SBA01, SCM09b, SB07, SPT04, TSE^{+}08, TOK08, TKO08, Tal09, TND^{+}05, TSIY06, TMD^{+}07, THP^{+}05, TDSM02, TCWR08, TCW^{+}08, TTL06, USG^{+}08, VRJ02, VTDS09, VW00, VBF06, VS08, WWGR07, WKPS06, WR06, WKH02, WKIK04, WKMK06, WKY09, WOL^{+}07a, WCJL08, WGV^{+}02, WAN^{+}01, WS08, WWWR09, WWR09, YI03a, YI06, YIY08, Zam03b, ZH02]. pacifica [AS06a, BLS06, DBS05, FSP06, GGFSP06, NOE^{+}04]. pacificus [IAKS03, Joh03, KS04a, MWG02, PC05]. pack [GRG^{+}02, GGG^{+}04, LCB^{+}05]. package [Enr05]. Packaging [LBR04]. Paffenhöfer [Ano02i]. Pagophilus [CDH^{+}09]. Pagrus [DWB04, HJG03, NJPI03, PBH^{+}03, TSHT02]. pagurus [MS08, SP08a]. PAH [HFD08]. PAHs [NAB^{+}06]. paints [RGEB00]. paired [DPD^{+}08]. pairing [FE07]. Pakistan [SDW^{+}01]. Palaemon [MNC^{+}02, PABM08]. Palaemonetes [RPS^{+}06a, RGR^{+}01]. Palaemonidae [MNC^{+}02]. Palatability [KKBH03, PAM^{+}09, AIM^{+}05, TLLCT02]. palatable [Her06]. Palau [HCH07, Tup07]. palifera [AM06]. Palinuridae [BO00]. Palinurus [GQR06, GTGM06, GB02b, TGGM05]. pallidus [DPSD08a]. palmata [BHH05, GDBF06, GDBF07, Lir00, MBW07]. Palmer [QR03]. palmiformis [CTVP03]. Palmones [CIFN00]. Palmyra [RK09]. palp [CDK^{+}07]. Palythoa [Bur02]. PAM [KGB^{+}01]. Pan [CHH^{+}05]. Panama [OHCFJ04]. Panamanian [PGM05]. pandalid [CPA^{+}03, CGM00, FC04]. Pandalidae [MNC^{+}02]. pandalids [FC08]. Pandalus [CGM00, HM03, OSL07]. panicea [TL08]. panmictic [TTL06]. panmixia [TGGM05]. Panopeus [RE00a]. pantophysin [CHH^{+}05]. Panulirus [BFLÁ01, BL09, CH05, EJKN03, GRPE08, NPLZ00, PE05]. papillosa [RGA^{+}02]. Papua [CWCP^{+}06]. Paracentrotus [FTP^{+}09, FB00, RM00, TRT04, TÁCTR06]. Paradigm [SS09b, BS07a, CH09]. paradox [HCBP05, LFP^{+}04]. Paraeuchaeta [AMR00b, AMR00c, SK03]. Paralabidocera [SNGR00]. Paralabrax [FA02, LTCP03, SMF^{+}02]. Paralichthys [Mad02, NRM05, SAG08]. Paralithodes [SAWG05, SAJG06]. Paralomis [ATLC03, CTA^{+}03]. Paralvinella [CTVP03]. paralvinellid [RL09b]. Paralytic [DCM^{+}06, CCM01, GFC^{+}02, JF02, KS06, SSO^{+}01]. Paramesopodopsis [JR01]. parameters [AHZ03, BA00, DLD^{+}00, GK07, JM01, KCA^{+}08, MAA^{+}07, ME01, MWD^{+}09, NPTW05, RMA^{+}02, SFH^{+}00a, SP01a]. Paramorea [DL03b]. Parapercis [RW08]. Parasagitta [CDT^{+}03]. Parasite [BG02, JTVC08, BCS^{+}06, DdMB04, Gri09]. Parasite-associated [JTVC08]. parasites [AGM06, BKB^{+}09, KRP^{+}06, OHS^{+}07, SD00b, SS06]. Parasitic [ÖNCNG05, AML^{+}03, GLG00, GL02, MC06, MdMR04, PCYC02]. parasitically [PBJ^{+}05]. parasitism [BDRM06, LM09, MRH^{+}08, SDFM00]. parasitized [RAA02]. parasitoid [TT06]. Parental [RGMV04, CC03b, KN00, OAM06, RGL^{+}00, VMdBV06]. parental-caring [VMdBV06]. Pareuchaeta [TNT06]. Parhyale [PCT00]. Park [LPS08, MHO07, Lam00a, CDS07]. parks [MGL^{+}07]. Parrotfish [KVH02, AFHS08, BB08a, CRAP03, GPK^{+}04, GHV05]. parrotfishes [BKSS07, MR05, RL05b]. parsimony [TW10]. Partial [LBB08, DBN^{+}09, HJH^{+}02, RR05, SHL^{+}03, Tan02]. partially [LH08, NvW01, TK08]. participation [MC05c]. Particle [BDDL03, BCJ05, DBCB05, DLB04, GDL04, RCAK03, BPVL03, BDR04, BC06a, BDG^{+}07, CB00b, FK00, GSC^{+}01, GK00b, HSBD08, HWG06, LWH08, LP01b, MKK^{+}09, NSH^{+}08, Pen00, PSRB00, RL05a, RGD^{+}08, SFHS00, TGY^{+}06, TCD01, VPP04, VJ04]. particle-associated [TGY^{+}06]. particle-attached [FK00]. particle-processing [Pen00]. particle-tracking [MKK^{+}09, NSH^{+}08]. particles [EP01, FLUF^{+}04, FFML02, FK00, JK04, KJL03, LWH08, MWMH05, MGF^{+}03, MYNY08, PS07, Pas00, Pas02, RYF01, RHG05, SSH^{+}04, SRGS01, SYVC00, SPBM05, Tho04]. particular [VHH^{+}02]. Particulate [KTK^{+}02, BVCG01, DSPHV04, DF06, EKBM05, FD00, GGRC^{+}08, JPMGS08, KPKD06, LCC01, LLJ05, MGA^{+}01, MHDS09, MGPAE01, ME01, NCÁSBP04, SAT^{+}03, SS02a, TSF01, TDRM08, VGS^{+}01, WRJ^{+}05]. partition [CC02]. Partitioning [GÁSPR03, MRC^{+}08, QRWJF07, BM02d, CPK00, CLJ^{+}08, Dav00a, DMK^{+}08, DOML^{+}07, GSCW00, Jos09, LJM03, NFN^{+}09, OHS^{+}07, PMG05a, PWLC06, PLM03, RBG04, RGF^{+}03, SW04, WdBS02]. partitus [WM02b]. Partners [MW09]. partnerships [BR07]. parts [BGGR02, KMP^{+}06]. parvipinnis [LC02]. parvula [GP01, RdNS02]. parvum [FLG03, GJ03, SKK^{+}06, SKU^{+}08, ULL^{+}05]. Pasiphaea [ACAG07]. passage [BLF06, GN04, KSA03, STJ03, AML^{+}03]. Passamaquoddy [WAMP08]. Passive [BSP03, BIG^{+}06, NvW01, VCSL^{+}09]. past [RRIH^{+}02a]. pastures [MB05b]. PAT [LS06, LALB08]. Patagonia [DCA^{+}09, LME05, VMH04]. Patagonian [GHT^{+}01, GWWC00, JBG00]. patagonica [FOL02, FIRL02, GMD08, RMRAFP08, SFL05]. patagonicus [HHF^{+}07]. Patapsco [SSLM01]. Patch [APM^{+}01, CS03a, GMVF08, GC06b, JRC06, KH07b, LGB03a]. patch-reef [CS03a]. patches [CS05, DW09, FMT09, HLTT09, KW00a, LCF02, NE01, NEG01, O’D08, PJJ^{+}05, SGL02a, SHJC08, VHM09, YAF^{+}07]. patchily [BDR06]. Patchiness [KvdMDP09, VL00, AR02a, BSP03, BDÁM04, FE02, Fra05, GB05a, PMV06, RKGH01, Seu05]. patchy [GCU07, Lei01, MA00, PJB03, SVN04]. Patella [AH07, CCB06, CHW06, DJM08, PB00b]. paternal [GM05, TKK05]. paternity [BHMR01, BMCH09, HKR^{+}08, IMS05, JY08, MS08, SS04a, Too04]. path [TdMF^{+}08]. pathogen [DKAH02, MT09, RS02, SCBA05, WKH07]. Pathogens [PHP^{+}03a]. pathological [DSF^{+}04]. Pathologies [KKSR05]. pathway [SBSPS04, WWR^{+}01, YLL^{+}04]. Pathways [RL09b, VFJ06, ATTX00, BP02a, BTKG^{+}08, DCT^{+}03, FSP06, FBJ00b, HCB02, JBH00, LCB^{+}05, MBH^{+}09, NWFV06, PCCL05, PMG^{+}06, TYT^{+}04, TBM^{+}08, VNJ06, VHV^{+}04, WW03]. Patinopecten [SSO^{+}01]. Patiriella [JMU09]. patrol [KRF^{+}05]. Pattern [CMS^{+}01, AGS^{+}09, BCBVM05, BSP03, CSP^{+}03, GLB^{+}04, JSWH03, LMC08, LWR^{+}02, LHB^{+}06, Lus05, MADPQ06, NM04, PTB03, Rin03, RvdPHW04, STCJ01, SYPO09, TGA02, WM05]. Patterns [ACH07, BHPC06, BCU04, CW03a, CFMB00, CC07, CAA00, DSL07, DM05b, DWL09, Edm00, FA06, FPB02, FTBC05, FB02b, GBSQ09, GDDM05, GLFU^{+}04, HP09, JBA02, KM01, LWH^{+}04, MHO07, MHB^{+}01b, McG02, NW02, PHM^{+}08, PWdN^{+}00, RHBW09, SRJ09a, Sta08, VJ03, AFHS09, ASA^{+}03, ABD^{+}02, BS08a, Bar03, BK05c, Bel07, Bel08a, BGM09, BCBVM06, BJ08, BER^{+}03, BKDB00, BMO^{+}04, BTW08, BG02, BH02, BH04, CPA^{+}03, CHL^{+}05, CL04, CKB^{+}09, Col03, CH07d, CB05c, DHHY09, Dav00a, DMA03, DF04, DCAL^{+}02, DMM^{+}03, DHRG05, DS06a, DPM02, DNL08a, DR01, DH06, DD01, DS07, EJM01, EHM^{+}09, FLK^{+}06, FPB^{+}07, FA02, FCP02a, FFA03, FFT^{+}07, FWH^{+}05, FBM07, GhGD^{+}02, GVF^{+}03, GLS02, GY00, GSP^{+}02, GLÅ^{+}03, HWB^{+}09, HLH^{+}07, HSCH07, HDP^{+}06, HHH02, HJ05, HK02b, HWGB02, HFMP03]. patterns [HMO^{+}04, HCMR09, JF07, JDCB09, JNFL04, KMP^{+}06, KCS04, KHP09, Kum06, KAM^{+}06, KNF^{+}03, LB08a, LTMF01, LBF^{+}02, LV02, Lec05, LJ06, LH02, LP07, LLW09, LTCP03, MR05, MBW02, MPM^{+}03, MH08b, MHW06, MB07b, MYE03, MRJ05, MC06, McI03, MGREA07, MCP^{+}09, MGM05, MC05c, MA02, MMA03a, MB00b, NDV^{+}00b, NWJ^{+}08, NA04a, NMDS00, NHN^{+}06, NAB09, OC01, OSP01, OP09, OKM06, OT08, OGL09b, PGCC^{+}06, PEDT03, PCS^{+}07, PCHM07, RS03, RGP^{+}03, RWB^{+}09, RHPB08, REM^{+}06, RRDHG08, RP06, RB04, RLFEF03, RHSN07, RHS01b, RF06, RZ00, SML^{+}09, SM07b, SH07a, SSB02, SHS08, SE04, SS03, SAB^{+}02, SH05, SML05, SMH02, SSS^{+}05, SA06, SP04, SDW07b, SDK^{+}08, TMAN08, TAS07, TA07a, TLLCT02, TMFB06, TFSP09, TF04, TAW^{+}09, TLC06, Tor09, TCT^{+}00, TH06, UM08, VKH08]. patterns [Ver05, VFM06, VBA06, WFC03, WBW^{+}06, WGR07, WBR07, WM01, WH03, WKK07, WWWC02, ZS04, ZLAK04, vdMM03]. Patuxent [KHB00a, KS04b]. patzneri [WBC01]. Paucity [Cha03]. Pavona [GLB^{+}03, SHH08]. pay [HP07]. pay-offs [HP07]. PB [Ano01a, ASLD09, RCC01]. PCB [BGKN00, CBP^{+}01, DCT^{+}03, DCT^{+}04, DPD^{+}03a, DPD^{+}03b]. pCO [FHL^{+}09, GÁSPR03, GPM08, LRN^{+}09, RNDF^{+}04, RFG^{+}09, SFF^{+}07]. PCR [FS07, HLH^{+}05, MCS09, MRQM06]. PCR-DGGE [MCS09, MRQM06]. PDV [PHH^{+}05]. pea [DRDD09]. peak [BM09]. pealeii [BHMR01, BGH06, BMCH09, MH00]. Pearl [Lee00, BSK06, DCO^{+}08, HN03, LPBD00, TKN00, YLK00, YQC^{+}00, YQW^{+}01, Yin02]. Pechora [DDL^{+}03]. Pecten [BDR04, TK01, TNTH04]. pectinatus [AMR01b]. Pectinidae [BDR04, RL05a]. pedal [HSCD01]. Pedicellariae [GHdN07]. Pedro [GES^{+}06]. pedunculate [BWW^{+}04]. Peer [RBL^{+}04, Rii04b, Rii03, RQF^{+}00]. peer-review [Rii03, RQF^{+}00]. Pelagic [BBV^{+}01, HLBS07, SAJW04, SdV06, SF02a, VFM06, VBF06, ZH02, AHW03, AWP^{+}01, ANF^{+}00, BM00, Bel08a, BRP07, BBM00, BGS02a, BGL^{+}07, BFE^{+}05, BK06c, CRLP09, CPB04a, CJRT00, CLP08, DP02, DK03, DTV09, DS06b, EL05, FWF02, FVS^{+}08, FNDP08, GMK^{+}03, GRP^{+}06, GDGSP09, GGDHRA00, GSFN00, GLP^{+}08, GFPM01, HS03, HHKH05, HPB^{+}04, Hue08a, ISY07, KKK^{+}05, KBH01, KPT01, KH08b, KBL^{+}07, KMB^{+}08, KCG07, LRR^{+}09, LDL^{+}04, LEÖK^{+}02, LCE03, LHL^{+}07, LNH00, LPHH^{+}09, LYMCG06, LIM^{+}06, LWB^{+}00, LBH01, LW06, MNB^{+}07, MR06a, MVHV06, MR04, MTV06, MAT01, MBG^{+}09, MRB^{+}08, NReZW09, NLB08, NM04, NM07, NW02, Ogu07, OHS^{+}04, PCC06, PSC05, PSGP04, PCS04b, RM05, RWE04, RS00, RGF^{+}03, Saf03, SNH04, SAS^{+}06, SO02, TOK08, TRH^{+}06, TPSW00, Thu07, TJM00, TC00, TBAT02, TGNT05]. pelagic [VGW00, VLRJ07, VW00, WPM09, WDSA09, WJDA06, YI03a, YI06, ZYZ^{+}06, vdVBW00]. pelagic-benthic [MRB^{+}08, NReZW09, TRH^{+}06]. pelagic-spawning [LHL^{+}07]. pelagica [OHS^{+}04]. pellet [KSA03, OSA05]. pelletization [WRH05]. pellets [BND02, BGS^{+}02b, CMG04, IAT^{+}07, JB07a, Kiø03, LEÖK^{+}02, MTN03, PK06, PI08, RRS05, RJBW03, SRDPT^{+}07, SSH^{+}04]. Pellioditis [DRV^{+}05]. Pelorus [ZHWCS08]. penaeid [BLF06, PCD04, VHH^{+}02]. Penaeus [JRP01, OLG^{+}09, RMKR00, VHH^{+}02]. Penetration [NCMC09, CAS00, Con05, IC02, WBW^{+}06]. pengoi [LG08a]. penguin [BRSM07, BZLD02, FHB^{+}02, HDRC^{+}08, KHA^{+}00, LB05b, MB04, SZL^{+}05, TBM^{+}08, WAN^{+}01]. Penguins [Wil03, BC02, BWW^{+}04, BR09, BCC^{+}04, CB01, CKF^{+}03, CEO06, CCCD09, DGH08, GCLC^{+}02, GBW^{+}05, GBW^{+}07, HHF^{+}07, HC05, IBB^{+}07, KSSB04, LGM^{+}09, LJC00, MHB^{+}09, MEHD07, PISL02, RTPS07, RBMB00, RCWYK07, SR04a, SAB^{+}02, TDT^{+}03, TSEH08, TGT^{+}06, TC00, TC03c, WZ04]. Penilia [ASC06]. Peninsula [FAPA^{+}09, FAPOF09, GEM02, TAÁO^{+}03, VPP04, GGR06, AIM^{+}05, CMRB08, FHQ^{+}06, GVF^{+}03, ILL06, MMV05, NCÁSBP04, NCÁSGP05, PAM^{+}09, SPGS03]. pennatulacean [PCS09]. pentachlorophenol [LP01a]. Peracarida [GLB^{+}04, RWJ02, RJ03]. perception [JP08, NOP02]. perch [KS04b, NH03a, NH03b]. Percnon [DFCH05]. Percoidei [JH01]. perennial [RK08]. perennis [RGCL^{+}07]. perezi [CPBS07]. perform [BSKR^{+}06]. Performance [FBBV09, Ark09, BM05a, EGJ07, GA03, GM05, HRR^{+}07, HCB^{+}07, HC05, JHH01, KGB^{+}01, LWW^{+}01, LP07, LIM^{+}06, LÁMBF09, MPK03, MK04a, MMMB^{+}05, MRQ07, MCN09, RdNS02, RCWYK07, RT05, SL05b, SPH^{+}09, Smi04, SWHSA05, TSSQ06, Wil03]. performances [GLV^{+}09]. peri [BDDL03]. peri-hydrothermal [BDDL03]. period [AFHA05, DPD^{+}03a, LBB08, LP04, MAB^{+}03, PUB^{+}01, SH05, TNT08, WWHE04, WWDH09]. periodic [JHR05, Sha00, dM01a]. periodicities [MYH07]. periodicity [BB09, BVVM02, Joh02, KvNVH00, ZDA^{+}06]. periods [BSKR^{+}06, BGRD00, GKS^{+}07, SD00a]. Periophthalmus [NWT00]. Periphyton [CK04, HLH^{+}08]. peritoneal [KNKT01]. periwinkle [LM09, Pet02b, WH02]. Perkinsus [LNV^{+}03]. perlemoen [ESB^{+}04]. Permanent [RM06, GZG^{+}07]. permeable [EWBH04, HCJ^{+}07, MGM05, WTH04]. Perna [CYC^{+}09, EM04, GSL^{+}06, GXL^{+}08, GP04a, HTM01, ML00, ML05, NW05b, NZM07, PW04, PMR06, SAG03, Wan02a, WLT^{+}03, WC03, ZMN07, ZMTB07]. peroxidation [SAP02]. persist [TB08]. Persistence [LL03, MSD05, SHSV04, VJ05, BUF02, Bib07, CBL^{+}09, DW09, DS01, MAB^{+}03, RCL^{+}09, WW06a, ZPP09]. Persistent [HSC^{+}09, HFWW05, BKDB00, DUSJ^{+}09, LF06, MFM09a, RL09a, ZCR^{+}09]. perspective [BDF^{+}04, DRG^{+}07, EE06, FH03, FB02a, GLS^{+}00, KdLCM09, LRM00, Mar01, NB06b, Roy02, VLRJ07]. Perspectives [BS04b, BB05, MMPD09, CWB^{+}06, EOK^{+}08, MM07a]. perturbation [BCM^{+}09]. perturbations [HGMG^{+}04]. pertusa [BY09, BHY09, DBOR09, JNFL04]. Peru [ZBDE07]. Pestarella [PTLK04]. Peters [Kir02c]. Peterson [GHP02]. petrel [AHC^{+}06, FGSH^{+}05, HM06a, QMF05, RHC^{+}08]. petrels [BCH00, BDF09, CWC00, CBTW02, CMT^{+}08, GSCW00, GSSR02, HH00a]. Petrolisthes [Don04, HH07a, Too04]. Petrosia [RCC^{+}04]. PFA [Lam00a]. PFA-concrete [Lam00a]. Pfiesteria [ASH03, AHZ03, JKS^{+}07, PPHT00, SSBG02, SG02b, ZAH03]. Pfiesteria-like [PPHT00]. pH [HLR07, Hin02, ISG04, LHK04, MH07, MCCP09, PH03a, PH03b, SH01, SH07c, SH07d, Tak02]. Phacellophora [TT06]. Phacosoma [MTS07]. Phaennidae [NOP02]. Phaeocystis [HRQ^{+}02, HRR^{+}01, JKNJ07, LSG^{+}05a, NFV^{+}06, SV08, SLKF03]. Phaeodactylum [SDRH01]. Phaeophyceae [CK09b, SEH07b, Ste04b]. Phaeophyta [ABB05, Kev01]. Phaethon [LCL^{+}03]. phagotrophic [HP06]. phagotrophy [SHS00]. phalangium [RGVC07]. phase [LGSB09, NNLF09, RM06, SH01, Trn02, ZRG00]. phases [HPVL05, WA00, ZRG00]. phenolic [LELC02]. phenolics [SCBA05]. phenological [MSJ^{+}09]. phenology [AS04, Hip08, SSS^{+}09, SB09, Syd09]. Phenotypic [CEB^{+}08, DLT^{+}06, HWvdM03, MSG07, SR04b, BS08a, BGL^{+}07, Kra07, SPB09, TSW^{+}04]. pheophorbide [FH02]. pheophorbide-a [FH02]. pheromone [BK05a, HJH^{+}02, KMF02, KBT05]. pheromones [SCC05]. Phestilla [RWSP03]. philippinarum [MM05, MDG04, NW04, SPC^{+}00]. Philippine [AMT03b, RGMJM07]. Philippines [RAM03, SESP03, SJA^{+}09, VYM06, VYM08]. Phillip [Cam01, MJKH03, MK03c, RKLK07, TK01]. philopatric [BUCC09]. phlorotannin [PB00a, TP00a, VP00]. phlorotannins [KLFH04]. Phoca [CS02a, GNHB^{+}05, HBBD09, NWIT08, WHL00, ZS04]. Phocarctos [CWDG05, MMC^{+}08b]. Phocoena [CR04, CKTE01, DSF^{+}04, JWR05, KCTG06, MGE^{+}09, ST08, TH02, VMVG^{+}07]. Phocoenoides [OKAM00]. Phoebastria [AFHA05]. Pholadidae [PTH08]. phoronid [LS09a]. phosphatase [AC02, FB03b, HM05c, IRL^{+}09, OHL^{+}06, SN04, SAMN04, VDA^{+}03]. Phosphate [MVB^{+}05, AC02, BOM^{+}08, CAD00, CIFN00, DM09, GDS02, KP01b, NNW04, SAMN04, VMRL04]. phosphatic [AGB03]. Phospholipid [DK06, SKDP02]. Phospholipid-derived [DK06]. phospholipids [SKFPS02]. Phosphorus [FTK^{+}04, LWH06, MPDS07, MILH00, MFCC02, NNW04, OHL^{+}06, SC06b, DAM^{+}08, DRB^{+}01, EB04a, FPC05, FBKB03, GHH^{+}04, GFM^{+}02, HCA06, JF02, LFF^{+}09, LGZ^{+}09, MSH^{+}03a, NN06, NRPS^{+}01, VDA^{+}03, WLJ08, YQC^{+}00]. phosphorus-depleted [WLJ08]. photic [MFP^{+}05, PGS^{+}00]. photo [RG09, UHvO^{+}08]. photo-acclimation [RG09]. photo-physiology [UHvO^{+}08]. Photoacclimation [BMSH02, HSP^{+}08, MEB^{+}07, SHM02, WBZ^{+}09, vLvSGS05]. photoacclimatization [Les00]. photobiology [RS02]. photochemical [PKR07]. photographs [KTF^{+}01]. Photoinhibition [CNW07a, NvWY05, GCD^{+}03, HAC06b, Jon04, LDHC04]. Photoperiod [CW03b, BCLS06]. photophysiology [JPC^{+}06, SHH08]. Photoprotective [HKSW01, Dov04]. photoresponse [NYK03]. photosymbiosis [SdV06]. Photosynthesis [CNPS05, VMH04, dlSPLV09, BMSH02, BW00, CMRF05, ERR06, FJV^{+}02, FHPW07, FNDP08, GCD^{+}03, GLFU^{+}04, HBdBV01, LDHC04, LDS^{+}04, LN05, NvWY05, OEDSJ02, PCYC02, RH09, RvdPHW04, RG09, SBB09, SP01a, SFPGA05, SS03, SHS00, SSdWM00, SKH06, TN07, TFLA03, dBKSV00, jMFVPL05]. photosynthesis-irradiance [SP01a]. Photosynthetic [KGB^{+}01, MRPL07, RLBB04, SDHG03, SC01, UHR05, Uth06, ZJW^{+}04, BK00, BA00, BSJ02, BM05a, CABS02, FK04, GK07, HBP^{+}05, HPVL05, JS02, KF02, KCA^{+}08, MCP05, MG00, ME01, MUP06, POH^{+}02, PKR07, STRD02, SSGB^{+}08, TMJ^{+}05, WGP^{+}00]. photosynthetically [RGL01, VMN^{+}03, WdBS02]. photosyntheticperformance [HSGS00]. Photosystem [BDW^{+}00, JK03b, FCH^{+}09]. photosythesis [MD07]. phototactic [YOM^{+}09]. phototaxis [Ada01, MFY02]. Phototrophic [ET08, FH08c]. phototrophs [MFK07]. Phragmatopoma [MFY02, MYE03]. Phragmites [COK02, LC07a, MJG01, Min02, PAMJ03, VGB^{+}05, WWAC00]. phthalate [SFHS00]. phyla [BF03]. phylloclimate [BVP^{+}06]. Phyllonotus [Swa04]. phyllosoma [JRTJ04]. phyllosomas [BO00]. Phyllospadix [HIM05, RHBW09]. Phyllospora [CK09b]. Phylogenetic [FTU^{+}00, OJPM01, YFH^{+}02, GKWK02]. Phylogenetically [FK08, NSSI06]. phylogeny [ISSK03, Lei07, UST^{+}04]. phylogeographic [TMFB06]. Phylogeography [KSHO01, KHK^{+}04, PG01]. phylum [Ned04]. physalus [JTR05, RBGM06]. Physical [AAPO00, CMSS09, DP01, IC06, JSN08, LB08a, LSK08, MSH^{+}03b, PNCS^{+}08, PNCT^{+}08, RCB05, TSJ05, AF06, BFK^{+}05b, BB00, CHL^{+}05, DDS^{+}01, FD08, FH03, GÁSPR03, GNPB07, GNP07, GCHA05, Han07, Har06a, HH00b, JHB06, Kaa04, Kel02, Kim02, LTG^{+}01, LRR^{+}09, MM06a, MCD^{+}05, MBBW08, Mil07b, MJKH03, O’C07, OP02, PPCU07, PBN^{+}07, Pee02, SHH09, SK07, SCZ^{+}06, SWM03, TK08, WBH^{+}04, Wil09, WJB09, WKPK05, BI03]. Physical-biological [PNCS^{+}08, PNCT^{+}08, RCB05, GNP07, Han07, LRR^{+}09, Mil07b]. physically [WKD^{+}04]. Physico [DP02, NATR^{+}05]. Physico-biogeochemical [DP02]. Physico-chemical [NATR^{+}05]. Physicochemistry [RB05a]. physics [DPS^{+}03, NH03a, NH03b]. physiography [MFG^{+}04]. Physiological [BRBPL06, CLRM08, DP09, KFA^{+}04, MWTB00, NMVH06, PBH^{+}06, SP01c, TW00, TOBR07, WDLM05, AKRA02, Bay02, CGCDA02, JHL04, MB08, MCE01a, MDS^{+}07, MMMB^{+}05, POH08, SP03a, SZ08b, SYPO09, SMHG09, TAS04, XW01]. Physiologically [PD07]. physiologist [Pör08]. Physiology [HSGS00, SHS00, ENC^{+}04, FCP02b, GB02a, GG09, KVS08a, Lei07, LUIAH03, MSH^{+}03b, RB05a, SKMS06b, UHvO^{+}08, VN02, VN05, WC03]. phyto [Pas02]. phyto- [Pas02]. phytodetrital [SW08c]. phytodetritus [AW03a, DD04, JFR02, MvdBÓ00, MMB^{+}02, NOO^{+}08, SPGS03]. phytoflagellate [GT02]. Phytoplankton [AQHTL02, ACLF09, BAOV06, BNP^{+}09, CMJ04, CH07b, CSL^{+}05, CCJT09, CD05, DGJ^{+}03, EDBvdL08, GM01, GVF^{+}03, IG04, LPBD00, MLG^{+}03a, PWHW09, RS07b, RMPC08, SSGB^{+}08, SP01a, TD01a, TSF01, TFLA03, TLIL^{+}09, WW07, AD00, AWOO02, AVS03, BF01a, BF08a, BAM^{+}04, BA00, BLS04, BMW^{+}00, BBV^{+}01, BPS^{+}03, BW00, BOZ03, BBNK02, CMG04, CBAd00, CMRF05, CF08, CLL^{+}09, CMSS09, CABS02, DDS^{+}01, DBD^{+}05, DK06, DAA00, DCC^{+}03, ESQ03, EDKT06, EB05, EAF^{+}05, Eyr00, EF02, FPB^{+}07, FHL^{+}09, FPWM03, FIS04, FSV^{+}08, FCH^{+}09, FJ01, FNDP08, FHK^{+}08, FBN^{+}07, GFF08, GVKF03, GMP07, GHD^{+}02, GSW01a, GB03, GHL02, GEHZ07, GCHC00, GDL04, GRLML09, HLD^{+}07, HGS^{+}07, HN03, HSS^{+}04, HHC00, HBdBV01, HGG^{+}03, HRK^{+}08, Hin02, HLH04, HLH^{+}08, HS02, HK03c, JDTB06, JHJS04, KVS08a, KW05, KOTL07, KSSH02]. phytoplankton [KB07, KNU02, KWAT08, KBSWG09, KYYM00, KWL^{+}05, LHD^{+}01, LCL^{+}00, LF08, LCS^{+}01, LKR^{+}06, Lat07, LSE06, LD03, LFSR06, LSHL03, MHB^{+}01b, MBG^{+}03, MTH04, MKL^{+}03, MAJ^{+}06, MOOB06, MRC^{+}08, MHY^{+}06, MN00a, MSH^{+}03a, MHABGC04, MSH^{+}03b, MIL^{+}00, MGL^{+}00, NWHC03, NTGK09, NATR^{+}05, NRJB04, NLCS05, OS01, OHL^{+}06, Pah05, PO09, PFW^{+}02, PDT09, PAW00a, PCJF06, PPHT00, PTA^{+}09, PPCJ^{+}06, QA01, RMWO09, RW01, RPN^{+}01, RSR02, RHHH09, RGC^{+}06, RG09, RZWG08, RCB05, SWD^{+}04, SSN^{+}09, SDS^{+}03, SAT^{+}03, SLM^{+}02, SLL02, SMHL00, SO08, Seu05, SCGC07, SW02b, SW02a, SYPO09, Som00, SHSV04, SDRH01, SHM02, SMSJ04, SMF00, SOMM06, SMHG09, TA07a, TKVL04, TBID09, TDSM02, TT02, TSF04, TLNH01, VSV04, VFF09, VWH00, VMH04, WWAC00, WG09, WPC^{+}03, WWL04, WW04, WHL^{+}06, WAW^{+}09, WPD02, YQC^{+}00, YQW^{+}01, Yin02]. phytoplankton [YTK^{+}05, ZGV00, ZRG00, ZHWCS08, ZTW08, ZLAK04, TBL^{+}09]. phytoplanktonic [KUN02]. Phytotoxicity [JK03b]. Pico [FPF01]. Pico- [FPF01]. picophytoplankton [BCVC07, FMMS03, KKTN05]. Picoplankton [GUL^{+}07, LSM^{+}02, BdBK^{+}00, GRS07, RCAK03, SBL^{+}00, TKN00]. picoplanktonic [NReZW09]. Piddocks [PTH08]. pigeon [GSM^{+}02]. piggy [BWW^{+}04]. piggy-back [BWW^{+}04]. Pigment [AVS03, BAM^{+}04, SMSJ04, AG04, CCM01, DDV^{+}08, DC07b, HSS^{+}04, HBTD04, LG00b, LPBD00, LSHL03, MBR^{+}05, RGC^{+}06, SMHL00, SGK04, SLCB00, WHL^{+}06, ZER^{+}01, LWB^{+}00]. pigment/chlorophyll [SMHL00]. pigment/chlorophyll- [SMHL00]. pigmentation [FB03b, WBZ^{+}09]. pigments [BDC00b, Dov04, GMR^{+}02, HJ03, Rij03, SMHL00, SLHT05, ZJW^{+}04]. Pilayella [RBJB^{+}05]. pilchard [COPB06, OKAM00, OSP01]. pilchardus [COPB06, CG04, GMZvdL07, GBHO^{+}08, GCVR01, GUT04, RGL^{+}00, RGMV04, SMVF03]. Pilot [KET^{+}03, BBHT02, PBL^{+}09]. Pinctada [BSK06, LPBD00, TKN00, YLK00]. pink [CTLR03, CHMR05, CWBR05, CBM^{+}07]. pinnatifida [CHO09, TKGS04, VJ05]. pinnid [GFP^{+}05]. pinniped [BWW^{+}04]. pioneer [Lam00a, PZMG05]. pipes [QRS00]. piping [DSH03]. Pisaster [BV08, SM07a]. Pisces [CSS00, Gla07, LDR00, MPR02, NSEL00, ORD00, PDM^{+}04, SK04, SK05, Tup07, ZH02]. piscesae [PAB^{+}02, UBM^{+}07]. piscicida [ASH03, AHZ03, JKS^{+}07, SSBG02, SG02b, ZAH03]. Piscivore [DGdG^{+}09, BS05a, JBS01, VHB^{+}07]. piscivores [BS06]. Piscivorous [KRF^{+}05, HMJ^{+}09, IMTW09, She01, ST06b, SP01b, TS07, ZWYL01]. pistillata [AYSL09, EMB^{+}09, FPBAT02, FHGJB04, KJ03, MEB^{+}07, NYvW03, NKS03, RFPS^{+}02, SFPGA05, WBZ^{+}09, ZDA^{+}06]. pituitary [HYLT03]. Pivotal [GKB09, BVC^{+}02]. place [SMH09, vDS05]. placement [WBP^{+}06]. Placopecten [Har06b, LCHG03, MJS09, WBDR06]. plaice [BVVM02, BDCvB^{+}09, FGT^{+}07, GRB^{+}03, GHR^{+}07, HMO^{+}04, HCMR09, SM01b, TdLvdVR08, VGW00, Wen00, vdVBW00, vdVBGW09]. Plain [BA04, DDF01]. plana [FC00, LBC07, Orv05, PL07]. planci [GHdN07, LRM00, Pra01, San00]. planipes [GGDHRA00]. planktivorous [AS04, HR08, MHHS06, SBI07]. Plankton [BH02, CT01, EZTZ02, Fra05, GGRC^{+}08, HFD08, NFV^{+}06, SRR06, TAÁO^{+}03, TBL^{+}09, WSA^{+}06, AMS^{+}06, BSP^{+}08, BIR00, Bea04, BI04, BSB03, BBG^{+}08, CH07a, CWB^{+}06, DHG^{+}04, EÖKS^{+}02, FLG03, GYD04, GSRH02, HVH^{+}07, HD05, HD06b, JN09, KVS08b, LSH^{+}07a, LC09b, MBC^{+}07, MP00, MCM^{+}03, MD08b, NHNB01, OSF08, Osb07, PTD05, PM00, Rol00, RE00b, SMC^{+}05, SFF^{+}07, SW02b, SW02a, SC06b, SHSV04, SFT^{+}05, SBC^{+}06, SFG05, SLEH01, TCD01, UB03, Vis01, WKL^{+}03, WG00, WDX01, ZIUG07]. plankton-anchovy [OSF08]. plankton-rich [MD08b]. Planktonic [AQSP07, KKKK03, LR08, PG00, ATT06a, ATT06b, AMP^{+}02, Ano02i, BK05a, BK05b, BH04, BLH02, CBAd00, CMHP06, CRK06, DCO^{+}08, DAG^{+}00, DVLH^{+}00, EE00, FCH^{+}08, GK08, GAM02, GPM^{+}09, HJ03, HKM00, HK02b, Jak01, JS03, KBT05, LDWC05, MCH^{+}06, ONA04, OAGV07, OGA03, PL00, PK05a, Paf06, PL07, PH03a, PKRK07, PI07b, SLGD00, SSM05, SdV06, TMD^{+}07, VSL^{+}04, VD00, VF09, WL05]. Plant [DPD^{+}08, PDO01, GK02, GKK03, HR02a, PH05, SGB09, SC03, SWL^{+}06, SGCC06, WR00b]. plantation [BKK^{+}05]. plantations [ATTX00]. plantigrades [ML05]. Planting [BvK07]. plants [MSA07, SZZ08, vDSD02]. planula [NKS03, OYNO08, ZDA^{+}06]. planula-brooding [NKS03]. planulae [ACR07, PKH00, VSBB03]. planular [PV05]. planulation [VYM08]. plasma [EdPP^{+}00, KKK^{+}05]. plastic [HTJ00, JBG00]. Plasticity [WMCF08, BS08a, BCBVM06, BBG^{+}06, CGCDA02, DLT^{+}06, DD09a, EDE05, HMH04, JVT01, Kra07, RLBB04, SR04b, SPB09, TD01b]. plastics [BWW^{+}04]. Plata [JBCL03]. plate [SM05]. platessa [BVVM02, FGT^{+}07, VGW00, Wen00, vdVBW00]. platessoides [SM01b]. platform [FA06, MUP06, PDS^{+}07]. platforms [FGPS04, PDCH06, SAB04]. Platichthys [MTL^{+}03]. platipes [HTS06]. Platygyra [LWCF04, MSG07]. plausibility [FBD06]. plausible [BGDS05]. play [MBW03]. playback [Tya09]. players [MMB^{+}02]. plays [BV08]. Pleasant [CRV03]. plebejus [OLG^{+}09]. Plectorhinchus [WW03]. pleopod [TJSG03]. Plesionika [CPA^{+}03, MNC^{+}02]. Pleuragramma [MVHV06]. Pleurogrammus [NS02]. Pleuroncodes [GGDHRA00, RT00a]. Pleuronectes [BVVM02, Bro06, FGT^{+}07, VGW00, Wen00, vdVBW00]. Pleurophycus [DZ06]. plicae [BCJ05]. plots [DY05]. plover [DSH03]. plumage [QWY07]. plumchrus [BM09, CBD04, ESDKM09, EBG^{+}00]. plume [BJR04, DMS^{+}05, HLBS07, LHD^{+}01, LD03, LDWC05, MDZ05, PAK^{+}00, Pee02, PPCJ^{+}06, VJ04, WPC^{+}03, YQW^{+}01]. plumes [Fin00]. plumifera [JP08]. plumulosa [KSS^{+}05]. pneumatophores [GF02]. Poaceae [VGB^{+}05]. Pocillopora [AM04, HR07, Kuf01, LCCTFO03, MG08, PKH00, RRDHG08, UHR05, VYM06, Whi06]. pocilloporid [VYM08]. Pockmarks [WAMP08, WBG09]. poculata [DC07a]. Pododesmus [Tom08]. PODs [CHT06]. Poecilostomatoida [SHT^{+}04a]. Poigar [SAM^{+}01]. point [LWK^{+}07, Rii04a, WM02b, BMG02]. poisoning [BCBQ01, DCM^{+}06, GFC^{+}02, KS06, MFG^{+}00]. Polar [BWW^{+}04, LCGN02, BK03, BF08b, GNAD^{+}03, HDQ04, HW08, KG00a, KG00b, PBH^{+}06, SG02a, SCH^{+}07, TBBI04]. polaris [RHG04]. Poleward [RGC^{+}06]. policy [WGCM^{+}03]. Polikarpov [ZK09]. Politics [BS05b]. pollen [vTWMG08]. pollock [BBC^{+}09a, Bai00, DABC02, PCH^{+}05, PBS^{+}00, Rye02, WJDA06, YHSH02]. pollutant [QHD04]. polluted [BGQR01, PCSS01, PCBL05]. pollution [FFR^{+}03, GGGVGG03, GTFB03, JDVG08, KWV^{+}05, LFRG^{+}05, LLSL06, MK02, PHB02, PBB^{+}02, PEdlP^{+}02, RBG05, SFW01]. polonium [SFT^{+}05]. polonium-210 [SFT^{+}05]. poly [NCP^{+}08]. poly-brominated [NCP^{+}08]. polyacanthus [KHP09]. Polyandry [TSRR05]. Polyaromatic [SCH^{+}07]. Polychaeta [BBRJ04, CTVP03, DQ05, DBCB05, DCRT07, FCP^{+}07, Iye02, KTK08b, LGVF00, MJG^{+}00, MG05b, PZMG05, SP03b, TP01a, TP01b, VBA06]. Polychaete [GSP^{+}02, MGHA05, CH06b, DBCB06, EWSH08, FGS^{+}02, FGS00, GQ00, HLQ02, HK02c, HH03b, JQ04, KW00a, KM03b, LHQ03, LHQ05b, LMCQ02, LLY^{+}01, LBBM06, LOB03, MFY02, MYE03, MHW07, NWÓ03, ODG^{+}04, PG01, QRS00, RPS^{+}06b, RPS^{+}06a, RSL09a, RSL09b, RF08a, RN06, SWH04a, SZ09, SMB^{+}03, TSSQ06, TA03, TWM^{+}01, XWQQ05]. polychaete-scallop [SZ09]. polychaetes [AHL^{+}01, ABD^{+}02, BBST07, CVK00, CBA02, DM08, DPGN08, EWSH08, HL05, HCCY02, LJM03, MTLG09, PMPI03, RGA^{+}02]. polychaetous [SJS^{+}01]. polyculture [GSL^{+}06]. polycyclic [CWR06, NAB^{+}06]. Polydora [SWH04a]. polykrikoides [JTLG09]. polylepis [SH01]. polymeric [HSD03, LHQ05a, WS02]. polymers [LHQ05b, SLKF03]. polymorphic [CRC02, SAG03]. polymorphism [FOH^{+}08, MS03, UYNO01, VBB06]. Polynesia [AAPO00, AAGS02, LPBD00, FPF01, LYMCG06, MACS07, SLGD00]. Polynesian [OBG^{+}07]. Polynya [VMN^{+}03, CGL^{+}02, FBD^{+}02, GGK^{+}07, SSH^{+}04, SDS^{+}03, KHIH08, MLG^{+}03a]. polynyas [BF08b]. polyp [MW02a, PGM05, PHS09]. polyphemus [BLT03, CRV03, JSN08]. polyps [BCKP09]. polyunsaturated [ODA^{+}09, PAB^{+}02]. polyxystra [RSSA07, RH08]. POM [WASW03]. Pomacanthidae [EJM01]. Pomacentridae [Boo02, Gla07, LCEC02, SYWM05, WM02b]. pomacentrids [CW03a]. Pomacentrus [PKM04]. Pomatomus [HBNA04, MBL07, SLB07]. Pomatoschistus [JRA02, MPR02]. ’Pompeii [CFC^{+}00, FCP^{+}07]. pompejana [FCP^{+}07, PZMG05]. pomum [Swa04]. pond [BFB01, HFA09, LSH00, SG01]. ponds [BA04, BL01, KLN06]. Pontellidae [TCC04]. Pontomyia [SL05d]. Pontoporeia [BEE06, WS01, vdBÓME01]. pool [MHTJ07]. pools [DFMD00, EF07, Joh01, KMR01, MD08a, MHTJ07, SA03b, SAF^{+}09, TFA^{+}08, WW06a]. poor [AP05a, ABF07, CDK^{+}07, DW09]. pop [LS06, LALB08]. pop-up [LS06, LALB08]. Population [ACJ^{+}04, BCABC00, BSS05, CAS05, CÁ04, CCL^{+}05, CSO08, CRC02, CDS^{+}04, DCF^{+}05, DBS05, DESK03, DdNEvdZ05, ESB^{+}04, EGPT03, GJT09, GDGSP09, GNC07, HTM01, HH07a, HDW04, JSMM07, JÅC^{+}01, LCS^{+}01, LBD00, LJ09, LS04, MJG^{+}00, MBP^{+}06, MR07b, Nem05, OBG^{+}07, ORS02, RMT^{+}05, RPS^{+}04, SSGJ00, SBL08, SBHE05, TKGS04, TR06, TCW^{+}08, TTL06, WTF^{+}05, WT09, WWV08, WWWR09, AMR00c, ABD^{+}03b, Ano01a, ARP02, BS07b, BVC^{+}02, BCM^{+}07, BLS04, BJN^{+}06, BDH01, BSVM^{+}05, BER^{+}03, BBD^{+}02, BLT03, BDRM06, BHH07, BMJ02, CW03a, Car00, CRV03, CLS^{+}01, CMHP06, CK09a, CB05b, CY06, CCB^{+}09b, DL02, DC02, DdMB04, DZ06, DPSD08a, DPSD08b, DB07b, Edm04, EZTZ02, EBT^{+}02, FLL09, FMP05, FGS00, FBN^{+}07, GAMB05, GPK^{+}04, Gil02b, GCSSJV07, GSCW00, GKAS08, GMS02, GPSG06]. population [GDBF07, HWC^{+}06, HP06, HGSH08, HGS^{+}02, HBBD09, HBN01, HWvdM03, HM06b, HDP08, Hum00, HFA09, JY08, JVT01, KUOI06, KM03a, KH00b, KPW07, KOC^{+}07, KWD05, KSH^{+}09, KS04b, Kri02, LHH09, LPHH^{+}09, LCD^{+}05, LJK01a, LJK01b, LJK02, LRC06, LBHL^{+}09, LW06, MHH^{+}05, MHSS08, MS06, McI02, MHBB05, MP06, MC01b, MGZ^{+}03, MRH^{+}08, MR01, NRC07, NCMC09, NWHH^{+}09, Nod04, NLHO09, OSF08, OKAM00, PIY04, PZB^{+}05, PMM^{+}00, PCC07, PTS00, PCBS05, PWFC08, PKA06, RF08a, RGMJM07, Reu01, RCL^{+}09, RV05, RBG01, RT00a, RFCM03, RMAY04, SCMB03, SBMAE06, SMP^{+}09, SP02, SP04, SDR01, SMF^{+}02, SVL^{+}04, SMGCW00, SZL^{+}05, Tar03, TSW07, TFSJ00, TGGM05, TA01a, VFJ^{+}07, VHMB09, WTS09, WHB^{+}06, WÓ02, Whi06, Whi02, WAN^{+}01, WGL03, WMCF08, Win09, WG07, ZAH03, vDvTJD^{+}09, Lap01a]. Population-level [CAS05]. populations [AVA^{+}09, Arm01, ADR^{+}05, AP00b, BR03, BSDR01, BCBVM05, BCBVM06, BSK06, BHH09, BDJ09, BG04b, BFB^{+}09, BCP^{+}08, BL07, Bro06, BT04, CMB^{+}09, CBP^{+}05, CLSM09, CWT06, CK09b, CLR^{+}06, CS03b, CNT^{+}04, CPG08, DF04, DGL01, DBT02b, DLD^{+}00, FSO08, GMA^{+}09, GPSH01, GL04b, GF05, GMK^{+}03, Gil02a, GGFSP06, GWPRC^{+}07, GTGM06, GC07, GDBF06, HKH09, HMY^{+}08, HTM01, HMD^{+}07, HCR01, HFM^{+}04, Hyr01, JDCB09, JKES07, KPA^{+}09, KHIA02, KH05, KB04, LQRGT06, LW02, MB08, Mal02, MTL^{+}03, MJS07, MB07d, MGM^{+}07, MHBB05, ML00, MHK^{+}01, Mil07b, MR06b, MRGC09, MCU06, MFCC02, MAG07b, ONA04, OM04, PSPZ09, PAW00a, PDM^{+}03, PKS00, PP02, PD01, PBJ^{+}05, PBS^{+}05, QAA^{+}06, RSL09b, RHC^{+}08, RDD^{+}02, Rit06, Rod03, RMZC07, RMC^{+}07, Saf03, SSN04, SM07a, SG06a]. populations [SVG07, SP03b, SPD03, SSW08, SAG03, SSH^{+}00, SLEH01, SRJ07a, SJH^{+}04, SNGR00, SM01b, TG03, TCS^{+}09, TS06, TTL06, Tyn04, UB01, VSRGW09, VB02, VAJ^{+}03, WKL^{+}03, WE04, ZMTB07, dCLLT^{+}02]. porcelain [Don04]. Porcellanidae [Don04]. Porcupine [DDF01, KHvW02]. pore [BWP^{+}06, CHW02, EWBH04]. porewater [WM06]. Porifera [LWH^{+}04]. Porites [BG04a, GEC08, JM03a, MFK07, MBLA01, RL05b]. porpoise [CHT06, MGE^{+}09, OKAM00, ST08, VMVG^{+}07]. porpoises [CHT06, CR04, CKTE01, DSF^{+}04, GSS09a, JWR05, KCH^{+}03, KCTG06, RFLK03, TH02, TR06, TW10]. Port [Cam01, JKQ02, MJKH03, MK03c, OHN01, RKLK07, TK01]. Porter [Ano04e]. Portugal [AQSP07, BMG00, BRAS03, CB00a, CES^{+}03, FCP02a, JC08, MBW02, PMM^{+}00, SMVF03]. Portuguese [BMB02, CVF07, QAA^{+}06, dSSC07, dSSC^{+}08]. Portunidae [LWJC04]. Posidonia [AMR00a, AMR01a, BCL03, BGLL09, BD09, CRS^{+}07, CLMR07, CLRM08, DGM02, FJV^{+}02, FHDA^{+}07, GBMHD08, GCSSJV07, HSBD08, LV02, LMD^{+}02, MDSP05, OEDSJ02, OHT02, OKM06, PTAR07, PAR08, PCL08, RP02, RR01, SC01, SPI^{+}07, TTR05a, TTR05b, VL00, VBAR07]. position [APM^{+}09, BGSBC02, CDF^{+}08, CYC^{+}09, GC01, GV01, HM06a, NBDG05, PH03c, ROBFR08]. positions [GFPM01]. Positive [MMMB^{+}05, PH01b, WOH05, RSE01, Wie05]. Possible [Ano07g, AC08, WPIB09, ARFP09, CHO09, EAF^{+}05, Gar00, GSS09a, GA08b, HCB02, JLCA01, Koe05, KWHS^{+}06, LR06, PZ07, Rod03, SPT^{+}00, YH04, NMCT03]. Post [CKF^{+}03, HR07, CPB04a, CH07d, CT04, CVT09, DMM^{+}03, DM01b, EM05, EE00, GLA06, HCM01, HLFG^{+}07, Jef03, JH01, Kin02, KSSB04, KCSPT09, MPM^{+}03, MK03a, MPK03, MK04a, MB00b, MMBF09, NW05a, OvM02, RM05, Sha09a, ST07, Tay03, Tay05, TNT08, TCT^{+}00]. Post-bleaching [HR07]. post-disturbance [CT04]. Post-fledging [CKF^{+}03]. post-glacial [CH07d]. post-invasion [NW05a]. post-larvae [HLFG^{+}07]. post-larval [DM01b, MPK03, OvM02]. post-metamorphic [MK04a]. post-moult [KSSB04]. post-publication [Kin02]. post-recruit [ST07]. post-release [CPB04a, MMBF09]. post-settlement [CVT09, DMM^{+}03, EM05, EE00, HCM01, Jef03, JH01, KCSPT09, MPM^{+}03, MK03a, MB00b, RM05, Sha09a, Tay03, Tay05, TCT^{+}00]. post-spring [TNT08]. postbreeding [BCHS09]. postcopulatory [TSRR05]. postglacial [TCW^{+}08]. postlarvae [CWBR05, MHR03, MFM^{+}09b]. postlarval [EE00, HS05, OLG^{+}09, TW00]. postmitotic [RGP^{+}03]. Potamocorbula [Pen00]. Potential [CDO08, FMMS03, GDBF06, HSW^{+}08, MSF04, PCJF06, SZW^{+}07, BP02a, BPC^{+}02, BM00, BJLH04, BGPL08, BGD^{+}03, BKB05, CHMR05, CH06b, CSTC01, CCL^{+}08, DRDM09, DRDD09, DD06, DS06b, FSB07, Fox04, GM01, GY00, Gla07, GSFN00, GCS06, GRLML09, GGM^{+}00, HHV06, HMKR05, HCS^{+}07a, HGS^{+}02, HGAS^{+}01, HLGCO06, HB04, HTBT07, JP04, JDVG08, KET^{+}03, KMWH04, KS04a, KS00b, LNV^{+}03, LP03, LTH^{+}06, MFG^{+}00, MK03b, MM07a, MB02b, MBBP08, MAKMK09, NBB06, OBB00, PGR^{+}01, PZ07, RSE01, RHH^{+}05, SR09, SAJW04, SHA^{+}05, SN04, SD01a, SA03b, SDK^{+}08, TBR^{+}03, TP07, TPSW00, TYC00, UP07, VTY06, WGS^{+}02, WBJ^{+}07, WLT^{+}03, WWR^{+}01, XW02, vDvTJD^{+}09]. potentially [AHK^{+}02, DFMD00, LHK04, SRDG05, TdMF^{+}08]. Potter [STE04a]. pouchetii [HRR^{+}01, NFV^{+}06]. pouchetii-derived [HRR^{+}01]. pout [KB06]. Power [CLABRO^{+}03, BGDS05, CPH^{+}02, SWL^{+}06, SMGCW00, dlSPLV09]. power-scaling [dlSPLV09]. Practical [BLF06]. practices [FEO07, Hou09]. praeputialis [APM^{+}01, CLC04]. Praunus [WG02b]. prawn [JRP01, Kro05]. prawns [PMP04, VHH^{+}02]. Pre [HCM01, LOB03, PV05, DL02, KSSB04, MB00b, NW04, SSS^{+}09]. Pre- [HCM01, KSSB04, MB00b]. pre-conditioning [SSS^{+}09]. Pre-emptive [LOB03]. pre-exposure [NW04]. Pre-planular [PV05]. pre-recruitment [DL02]. precautionary [ARP02, CM04b]. precious [NM09]. Precipitation [Min02]. Precocious [LC08]. precopulatory [TSRR05]. precursor [Pas00]. Precursors [WDB^{+}05]. predacious [GSM05, MHV^{+}00]. Predation [AF06, BKK^{+}05, Goc04, LG08a, Lid08, MHV^{+}00, Niz07, OHT02, OKM06, PJN01, Pur03, Tay05, TDL05, VV04, AHV01, AE06, BLS03, BSZ05, BCT^{+}08, BFB^{+}09, BJJ01, BG07b, BH02, BBST07, BL09, CDO08, CH09, CDH^{+}09, DBP04, DGdG^{+}09, DHR04a, EvdKB^{+}08, FBD06, GRWB02, GR07a, GW06, GDdS^{+}07, Hal00, HMKR05, HBNA04, HJK01, HLTT09, HM06c, JP04, JS03, JNH^{+}09, KB06, KS00b, KCG07, KH05, KB04, LBF^{+}07, LGB03a, LR06, LEB07, LB08b, MS09, MPSJ06, MHCB08, MHM^{+}08, MLS^{+}06, MAW06, MJG08, MH06a, MR07b, NZM07, OLG^{+}09, PMPI03, Pea03, PGK^{+}08, PTCH01, RdNS02, RL05b, Rye02, SBL^{+}00, San00, SBJ02, SW08b, SH05, Str02, SDK^{+}08, TAM03, TOK^{+}04, TAO07, TNTH04, TJEM02, Tay03, Tor01, TLNH01, VHW^{+}07, Vii07, Wen00, WH09, vTWMG08, EFL^{+}09]. Predation-induced [Goc04]. Predator [CC04, KKG06, KFT08, LGB03a, MWG02, NS07, SPL00, SJR00, Sta09, TAO07, TK03b, ALE^{+}09, BPH07, BS08a, BDC00a, Bar02c, BBDW09, BLDG01, CSMBH08, CYC^{+}09, DFH05, FK08, FE02, FH03, Fre07, HBDD02, HLY05, JWM02, KH00c, LHJN07, LGC^{+}06, LB02, MLGG03, Mis04, NB06a, Nod04, PBFB08, RAAT^{+}07, RB08, RST04, RH08, SKBP09, SMHC03, SP08b, Smi04, SHW09, Tal01, TE00, Tit01, TPD^{+}05, UP01, VACT^{+}08, VBFF08, WTP08, WDT^{+}08b, WLJ^{+}04, WFG^{+}08, WA03, WS06, YG09]. Predator-inducible [CC04]. Predator-specific [TAO07]. predators [AG02a, ADF^{+}09, AM01, BZLD02, Bra08, CEB^{+}08, CDG04, CH07c, CDW^{+}09, DMA02, DHR04a, FTP^{+}09, FSW09, FD02, Hun04, KEB07, LS09a, MCFM02, OKM06, PHB02, PBB^{+}02, PSC05, PRL^{+}09, PCRR08, QS05, RHDC05, RS06, Rit02, RFB07, RL08, SJR00, SMC09, SF02b, Syd09, Tor09, WP00a, WP00b, WFT^{+}08]. Predatory [DJM^{+}05, JBS01, BIC^{+}03, BI03, CW05, EGJ07, FKA^{+}06, GH08, HB04, HRDA05, KPEF00, OP02, PGR^{+}01, RBG04, VHW^{+}07, WB06]. predict [BN09, CCL^{+}08, JWM02, LH02, RZS05, SDW07b, Wil03]. predictability [CMS^{+}01, FWC02]. predictable [MYH07]. predicted [BDC00a, Bar02c, vdWHF^{+}08]. Predicting [BCPB^{+}01, BM07c, CB08c, CCCD09, DBH^{+}07, DJ04, EYH^{+}06, HS01, Law08, MHP03, RW03, RM09a, SM08b, SKE05, SMGCW00, SWGC^{+}05, SG02b, DGVK^{+}04, PL03, SHW09, THN^{+}03]. Prediction [CT09, BHR08, GPLAG04, LRK09, RSL09a, YMHV02]. predictions [GNAD^{+}04, MAOP09, MDLS01, TF04, WCM08]. Predictive [EM07b, MB07e, WCH09, SMGCW00]. predictors [Bat09]. predicts [CHMR05, Cum02]. preference [CD07b, EDE05, HYLT03, MS06, OLG^{+}09, RSSA07, Sot03, WHL00]. preferences [BMdNP09, CH00b, HGG^{+}05, IR02, JHH01, KPN03, MGE^{+}09, Pra01, TJS08, WPM09, WOL^{+}07a]. preferential [TGM^{+}03]. preferred [RCF09]. pregnant [BB09]. Preingestive [UNII01]. Preliminary [KHG06, WG02a, KWP07, KFH04, MWTB00, SMF^{+}02, YQC^{+}00]. preparation [STH^{+}06]. Presence [MH08c, NSSI06, RSSA07, SCC05, Bat07, DL03b, Fin00, GT08, HSCH07, KSKL^{+}03, KSKK^{+}07, MKG^{+}05, RM09a, Tal01]. presence-only [GT08]. present [Law07, RRIH^{+}02a]. preservation [KP01a, MDLS01]. preserved [AKRA02]. President [ZK09]. prespawning [MRS00]. pressure [AK02, DGdG^{+}09, GG09, HWL00, HH07b, KSSH02, LOB03, PTCH01, RCC^{+}04, RR05, SHL^{+}03, SW08b, SHSV04, TTR05b, TYC00, VTY06, WZ04]. pressures [BHZW08, LR06]. Prestige [AGCF^{+}08, LFRG^{+}05, MAVO^{+}06, NAB^{+}06, OGMC06, SRDPT^{+}07, VML05]. Prevalence [FFN^{+}08, BBG^{+}04, HAP^{+}08, WM05]. prevent [OST07]. prevents [FIRL02, ÓUA05]. previous [MB08, RBW03]. Prey [AS06a, Bas06, BTB^{+}02, JWM09, OKAM00, PKLD09, PS01, RCCK09, RGL06, WHL00, WBJ^{+}07, ADF^{+}09, ARBF02, AP00a, ASC06, BKM08, BS08b, BCC^{+}03, BPB07, BBDW09, BHH09, BZLD02, BH01, BJJ01, BHD04, CSMBH08, CW05, DJG^{+}07, Dav00b, DMA02, DMA03, DFH05, DCM^{+}06, EL05, EWG^{+}08, EGJ07, FE02, Fau09, FSW09, FM02, FFN^{+}08, FTO^{+}09, FHQ^{+}06, GWH03, GMZvdL07, GNAD^{+}03, HSGS00, HFT^{+}09, HBRJ06, HC05, HSCH07, HGJM09, HLTT09, HSST01, Hip08, Hyr01, IBB^{+}07, JRA02, JHL00, JWM02, JBS01, KFT08, KS00b, KH00c, KLW^{+}09, KH05, KW00b, KB05, MSLP06, MFG^{+}04, MAH^{+}08, MAC^{+}07, Mis04, MWS^{+}04, MR06b, MHV^{+}00, MWM^{+}05, Mun07, NOE^{+}04, NB06a, NWT00, NH03a, NH03b, OvM02, PBFB08, PRvdL^{+}07, PLM03, PPC^{+}05, QMF05, RAAT^{+}07, RFLC07, RCF09, RB08, SS00a, SPL00, SJR00, SLC05]. prey [SKBP09, SG05, SMHC03, SMC09, ST06b, SPH^{+}09, STV^{+}09, SvNvdMvdV09, SW08b, SH07c, STAS06, SIK^{+}02, TE00, TS07, TBBI04, Tit01, TT05b, Tor09, TPD^{+}05, VML05, VFV01, Vli05, WTP08, WMS^{+}03a, WDT^{+}08b, WFT^{+}08, Wen00, WLM06, Wil09, WS06, ZBDE07].

Prey-dependent [BTB^{+}02]. prey-handling [BH01]. Prey-switching [AS06a]. preying [MBG^{+}09, TC03a, WBDR06]. Pribilof [SH00]. prickly [DS09a]. primarily [KLHO02]. Primary [ASES07, CGCMPA01, DDQ^{+}00, GPW^{+}02, SHA^{+}05, SCD^{+}05, WDN00, ACT09, BHW^{+}03, BKM08, BK00, BMSH02, BG07a, BPS^{+}00, CN04, CPSC01, CHV^{+}09, DRF08, DGJ^{+}03, DRB^{+}01, EAF^{+}05, FFYS^{+}07, FTK^{+}04, GRS07, GKWR02, HHIS06, HB08a, JBS01, KCP^{+}07, KM01, KNU02, KUN02, MBG^{+}03, MLC08, MOOB06, MRG08, MRQ07, MGPAE01, ME01, MFP04, NEK^{+}07, PCJ06, PUB^{+}01, RSK04, SR08a, SGB09, SC00a, SCC06, TFLA03, TBC^{+}00, WBN^{+}00, YF04, YVB^{+}02]. prime [SRJ07b]. primitivus [DRDD09]. Primnoa [SSRG05, SHS^{+}05]. Primnoella [OGA03]. Primnoisis [OGA03]. Prince [GSM^{+}02, GFPM01, IFL02, JTN09, KPM00, LGBWH06, PBS^{+}00, PS01, Pur03, WGD05]. princeps [FSW02]. principal [SGCD01, VNJ06]. principle [RNL00]. principles [CD05]. prions [CBDH02, QMM^{+}08]. prior [SKMS06a]. priority [BT09]. prison [CSM^{+}09]. pristine [EF07, RFH^{+}06]. Pristipomoides [NSEL00]. Probabilistic [DH07a, Kra07]. probability [BBC03, ONA04]. probable [WHRS05]. probably [ABD^{+}03b]. probe [Eva08]. probes [Rob06]. problem [Clo01]. Problems [WA00, Bra05, BBNK02, HPR05]. procedures [ARP02, BH08, EMG01]. proceedings [GBB^{+}01]. Procellaria [BDF09]. Procellariiform [CKS05, CMB^{+}05a, NBB06]. Procellariiformes [TPSW00]. Process [CN04, FM02, BBD^{+}02, BWRP06, EBT^{+}02, GS09, HTJ00, Kin03, SD01b, SSdWM00, Wri07]. Process-based [CN04, FM02]. processes [BPE^{+}06, BS07c, BPN^{+}09, CHW02, CHSW03, DSL07, DP01, DDS^{+}06, DM05b, DDS^{+}01, Den08, DBT02b, DS06b, FHQ^{+}06, HUP^{+}00, HBBD09, HH00b, HW01, KPW07, Kum06, LBD00, MPM^{+}03, MCD^{+}05, MKK^{+}09, MVLM^{+}06, MB00b, MR01, NCÁSGP05, NAB09, ODG^{+}04, PDQ01, SP03a, SHH09, SHS^{+}05, SWL^{+}06, TAS04, TK08, TMFF00, TBBB06, TKK05, UNII01, WBH^{+}04, Wie05, Yin03, ZLGC04]. Processing [HD07a, KW05, BDDL03, BCJ05, BC06a, BDG^{+}07, CBMB03, DBCB05, FWF02, Pen00, RGA^{+}02, SAS^{+}06, vOSF^{+}09]. processors [KWKF03]. Prochlorophytes [KW00b]. produce [IBB^{+}07]. produced [LTL06c]. producer [BK06c, PPB^{+}01]. producers [TBC^{+}00]. producing [LHK05]. Production [BTC^{+}04, GMTN00, JH04, MWMH05, MTN03, Pas02, SLS^{+}00, SNCP04, TdMF^{+}08, WAZ07, WS02, YI06, vBSSH02, AMP^{+}02, AMR00b, AMR00c, AMR01b, AVS03, AMS^{+}06, ASES07, BHW^{+}03, BKM08, BK00, BMSH02, BL00a, BG07a, BDH01, BD06, BSJ02, BRS07, BC01b, BWSH00, BPS^{+}00, BFE^{+}03, BGRD00, BOZ03, BNP^{+}09, BBNK02, BL01, CSA02, CWMD06, CM07, CEB^{+}08, CGCMPA01, CIHL05, CA06, CVÁM06, CPKV00, CN04, CMLG07, CSL^{+}05, CYC^{+}09, CW03b, CPSC01, CDT^{+}01, CB05c, DDQ^{+}00, DCT^{+}04, DGM02, DHBZ00, DRF08, DR00, EAF^{+}05, EF02, FR07, FPF01, FFYS^{+}07, GRS07, GGK^{+}07, GPW^{+}02, Gis05, GKWR02, Gor03, GKM02, GGRC^{+}08, GFM^{+}02, HHIS06, HK03b, HB08a, HSHA06, tAHSHA07, HCV03b, HQDH06, HM01, HCBP05, HK03c, ITE^{+}04, JZL09, JNDL02, JWM02, JYS06, JVJ09, KD00, KCP^{+}07, KM03a, Kin01]. production [Kir02a, Kne03, KM01, KNU02, KUN02, KZ00, LCT^{+}08, LHK04, LW06, MD08a, MNTS08, MG09a, MRZJ05, MBG^{+}03, MCP05, MRHP01, MR03, MPDC03, MLC08, MJG^{+}00, MLG^{+}03a, MOOB06, MRG08, MASFS00, MMA03a, MHABGC04, MSH^{+}03b, MGPAE01, ME01, MFP04, MIL^{+}00, NPN06, OCB^{+}07, OBB00, PMM^{+}00, PB00a, PCJ06, Pep06, PUB^{+}01, PGP03, PJ08, Pol00, PWL^{+}06, PGPL03, PPCJ^{+}06, RSK04, RGS^{+}09, RRIH^{+}02a, RSC08, RRH09, RRRM08, RKGH01, SGB09, SD09, SNH04, SS09a, SLM^{+}00, SHA^{+}05, SSJ^{+}02, SZZ08, SFKT09, SCGC07, STA^{+}05, SCD^{+}05, SKK^{+}06, SMSD07, SSdWM00, SMS06, SRC00, SCC06, SKH06, TSF01, TPQ^{+}03, TNHL02, TFLA03, TBBB06, TNT06, UNII01, Uth01, VP00, VLP^{+}04, VTDS09, VPS^{+}02, WL05, YF04, YVB^{+}02]. productive [AWP^{+}01, CVÁM06, KTH^{+}04, SdSC01, SDK^{+}08]. Productivity [KDAK01, BPH07, CMJ04, CCJT09, CHT09, DB07a, DGJ^{+}03, DRB^{+}01, DCC^{+}03, EHM^{+}09, FMMS03, FEMW07, FTK^{+}04, FHK^{+}08, GSP^{+}02, JMM^{+}09, JRT^{+}06, MHB^{+}01b, MS00b, MRB09, MF03, MRQ07, MB05b, MAKMK09, NEK^{+}07, NPN06, PH01a, Pol00, RFH^{+}06, STI03, SC00a, WPG08, WBN^{+}00, WDSA09, Win09]. products [GWBdN09, WTH04]. productus [GBW02]. Professor [Cos03]. profile [DPD^{+}03a, GPB02, LYL^{+}07, NR00, OKP06, SLM^{+}02, SK02, SWH^{+}04b]. profile-imaging [NR00]. profiler [GMM03]. profiles [BRI^{+}07, BSB03, CCM01, HW09, JRTJ04, KSIG08, MMC^{+}08b, MWM^{+}05, NFN^{+}09, SJM06, SBLF09, SKDP02, VPS^{+}02, WHP00, WP03, WM02a, YMC05]. profiling [FHK^{+}08, JDVG08, KTSR02, UYNO01]. profitability [BTB^{+}02]. Prog [tAHSHA07]. progeny [BHFC08, KN00]. progeny-array [BHFC08]. Programme [HR02b, HR02c]. Progress [Kin05b, SP08a, EBT^{+}02, PL03, WGM02]. projection [GSM06]. prokaryotes [MYNY08]. Prokaryotic [QASA08]. prolactin [HYLT03]. prolifera [MPB^{+}09, jMFVPL05]. prolonged [KSSH02]. promote [Cal06]. promotes [SM09]. promoting [PBS^{+}00, SLFM03]. proof [CM04b]. prooxidant [RCC^{+}04]. propagating [Gim04]. Propagule [HH07b, LVRN02, DDGK^{+}08]. propagules [BKK^{+}05, BFK^{+}09, SR01]. propelled [KRCGC06]. propeller [UH03]. Proper [RAE07, WNS^{+}05, GEHZ07, HLH04, Rol00, RHGB02]. properties [ACR07, BDF^{+}04, BSL02, BPK^{+}00, BPS^{+}00, BPS^{+}03, JM03b, Kaa04, KDM05, KFE^{+}03, SR08a, SLCB00, VMV^{+}09, VPG04, WTP08]. proposals [RBL^{+}04, Rii04b]. proposed [KES00, WA00]. Prorocentrum [BCBQ01, KJ04, LGQ^{+}00]. prosobranch [BMG00, BMB02, ED02]. prospects [HFH09, MARM06]. protandrous [CGM00]. protect [KB04]. Protected [FC00, HSTS^{+}09, MHP07, OT08, RRDHG08, SHMN07, BT09, CSS00, ESS09, FVB^{+}09, FDM^{+}09, FBM07, GAAB^{+}08, GCSSJV07, HF09, HCH09, Kap09, KAR04, KH07b, LH08, LAV04, LW04, MC01a, MG05a, NZ08, PRL^{+}09, SWG^{+}09, TMAN08, TP00b]. Protection [PPB09b, BSVM^{+}05, BD08, GA08b, LP03, Nem05, PPJ^{+}05, PFT^{+}08, SB03a, SAC^{+}03, SJA^{+}09, WAK^{+}09, Wie03]. Protective [GR07b, CSM^{+}09, KES00]. protects [BEKT09]. Protein [QHD04, CBD04, DFMD00, HDM04, LMF^{+}01, LVK05, MMGC06, PK01, RPS^{+}04, SSG^{+}00, VBB06]. protein-based [HDM04]. proteins [BCKP09, DCT^{+}04, FWF02, SW04, TTM^{+}05]. proteobacterium [WNWH02]. proteomics [JB07b, Lóp07, NT07, TEB^{+}07]. Protis [PAB^{+}02]. Protist [QA01, KWHS^{+}06, NBB^{+}04, RGP^{+}09]. protist-coral [KWHS^{+}06]. Protistan [FH08a, HSST01, SS06]. protists [AHK^{+}02, CLP08, Jak01, JKS^{+}07, LTNH00, PH03b, SABK03]. protocol [Bur01]. protoconch [MCCP09, Zac05a, Zac05b]. protogynous [DFH05, WM04]. Protomyctophum [SK05]. Protoperidinium [ONP02]. Protozoa [CGT^{+}00, EF07]. Protozoan [OAGV07, TUK09]. protozoans [LNH00, TLNH01]. protozooplankton [PI08]. Provannidae [SKW^{+}06]. Provence [GPSH01]. provendering [HSCD01]. provide [DL01, GCLC^{+}02, GA08b, MHM^{+}08, MP05, MH06b, SBJ02]. provides [LFRR^{+}07]. provinces [CC07, DSP07, NLCS05]. Provisioning [IBB^{+}07, AKM09, CKS05, HRCB02, HH00a, LKDO07, LGC^{+}06, MSH08, MGC^{+}09, PSC05, PWC^{+}09, PCW05, SR08b]. proxies [KSK07]. Proximate [EL05]. Proximity [RGC05, VHW^{+}07, LAH07]. proxy [SW09, SZL^{+}05]. Prydz [WRW03]. Prymnesiophyceae [HPVL05, OMT06]. prymnesiophyte [SH01, SSH^{+}00]. prymnesiophytes [LG00b]. Prymnesium [FLG03, GJ03, SKK^{+}06, SKU^{+}08, ULL^{+}05]. Psammechinus [HCK^{+}05]. Psetta [EdPP^{+}00]. pseudo [SCBA05, ABWK06, CG04, FDBT06, LRK09, LHK04, LHK05, MTH04, MP05, OLWB06, OPSM04, PPB^{+}01, RDD^{+}02, SPP^{+}09]. pseudo-induction [SCBA05]. Pseudo-nitzschia [ABWK06, CG04, FDBT06, LRK09, LHK04, LHK05, MTH04, MP05, OLWB06, OPSM04, PPB^{+}01, RDD^{+}02, SPP^{+}09]. Pseudocalanus [EO04, IP07, KBT05]. Pseudocaranx [AFHS09, MSK^{+}00b]. pseudodelicatissima [PPB^{+}01]. Pseudodiaptomus [CRB07]. pseudofaeces [UNII01]. pseudofeces [BC06a]. pseudonana [BVH02]. Pseudopentaceros [Hum00]. Pseudopleuronectes [BWCW02, KMAC^{+}02, MPM^{+}02, TC03a, Tay03, Tay05]. Pseudopterogorgia [BSVK05, EMPF07, LGRB^{+}08]. PSII [JK03b, YH04]. PSP [DCM^{+}06, KS06, MB08, TJ02]. PSP-toxin [TJ02]. Psychrochlamys [GMD08]. psychrophilic [EKN08, VFJ06]. psychrotolerant [VFJ06]. Pteraeolidia [LCW06]. Pterodroma [RHC^{+}08]. Pterois [AH08, KMWH04, WGV^{+}02]. Ptychoramphus [AS04, AS06a]. puber [LWJC04]. publication [GDDM05, Kin02]. publishing [BKB^{+}04]. Puerto [APA07, DvD02, PCHM07]. PUFA [ODA^{+}09]. pufferfish [KKMH08]. puffin [Bar02b]. puffins [HANM^{+}05, SH00, WIB08]. Puffinus [AO02, CMB^{+}05a, PC05, SWS^{+}09]. pugettensis [GDL04]. pugilator [BBS05, GB07, LDT07]. pugio [RGR^{+}01]. pugnax [BBS05, GB07]. pulcher [TLC06]. pulcherrimus [ENAT07]. pulchra [DdNS06, WdNS00]. Pulsating [VHM09]. pulse [AW03a, CCB06, JKQ02, KGB^{+}01, LS05, SW08c, WS00]. pulse-amplitude-modulated [KGB^{+}01]. pulse-chase [AW03a]. Pulses [RHS01a, BHV^{+}02, PKRK07, RC07]. pulvinata [DDQ04]. pumping [RL07]. punctata [Jen05]. punctulata [CL05]. Punta [GGDHRA00]. pup [BDG^{+}04, DPD^{+}03a, DPD^{+}03b]. pups [GLA06, WLS^{+}09b]. Purification [SSG^{+}00, HJH^{+}02]. purposes [LTL06a]. purpurascens [SdNH^{+}06]. purse [CPG08, SMVF03]. purse-seine [CPG08]. pursuit [vHGH01]. Pusa [KKL07]. Putative [RJS04, CH07d]. Py [BS08b]. pycnogona [BWW^{+}04]. Pygoscelis [CKF^{+}03, TSEH08]. pyrene [ABB^{+}00, CBA02, GHS05, HVH^{+}07, LP01a, TA03, WW06b]. pyrene-related [ABB^{+}00]. pyridine [ZRG00]. pyridine-containing [ZRG00]. pyrifera [HH05b, HHW^{+}07, Hob00, SFR^{+}09]. Pyrodinium [PBBC06]. Pyrosoma [PMNL07]. Pyura [APM^{+}01, CLC04, MC05a, MC07].

QFASA [NWIT08, TBI^{+}09]. quahog [CCL^{+}04]. qualitative [AHW03, BDR04, ISK07]. qualities [VN02]. Quality [BKB^{+}04, Kin04, AMG^{+}07, AP00b, BMK07, Bar02c, BKR04, CT05, FHB^{+}02, Gar00, GHS05, GRM00, JRT^{+}06, JVJ09, JFA02, LHQ05a, LLJ05, LP03, MK04b, MGTK00, MMBT03, NFN03, OTBR00, OvM02, PCS^{+}04a, PCS04b, Sar06, STM06, SWGC^{+}05, TGNT05, TBL^{+}09, Uth06]. Quantification [AHZ03, BCKP09, DHZ02, GMS02, GMM03, HSS^{+}04, HD06b, LCT^{+}08, RGD^{+}08, SK02, SWH^{+}04b, SdNH^{+}06, TGA02, VPS08]. quantified [PR05]. quantify [RDG^{+}07, SH06a, SdSC01]. Quantifying [FB07, GGDB09, MHP^{+}00, PBT^{+}09, PCJ06, WPM09, BMSH02, FT00, GFL02, SMHL00]. Quantile [PB08b]. quantiles [EMG01]. Quantitative [AFFH04, CLP08, DPD^{+}03b, LP04, MUP06, AHW03, BMdNP09, BEU^{+}07, GÁSPR03, ISK07, NWIT08, PL03, PGP03, PKM03, SPD03, TBI^{+}09, UM08]. quantitatively [ONA04]. Quantity [PDFD09, CT05, GRM00, JRT^{+}06, LHQ05a, TGNT05, Wan02a, XW01]. quantum [PCYC02]. Quebec [GSM06, RATD07]. queen [GL04b, KMHS04, Sto03]. Queensland [BB09, EB04a, TC01a]. question [SP08a]. Questioning [She08]. questions [Reu01]. quillback [TAW^{+}09]. quinone [UYNO01]. quinquecirrha [DBH^{+}07]. quinquelineatus [ÖNCNG05]. Quintin [IOSPT^{+}04]. Quirimba [BD00]. quissetensis [DdMB04]. quota [PEGG^{+}08]. quoyana [PR01].

rabbitfishes [FSHB09]. racemosa [DDHQ06, BBC08, PCC01, RLBB04]. racks [HLH^{+}08]. radial [BMW^{+}00, JKG^{+}05]. radians [GHV05]. Radiation [MLW03, Ada01, AL03b, BW00, BRB^{+}00, BSPK03, BD08, BV08, CABS02, DQW05, DL03b, FJV^{+}02, FIRL02, HBdBV01, HGG^{+}03, HQDH06, HTWQ05, Kuf01, LDHC04, Les00, LWMW02, MGL^{+}00, OMT06, PDB04, Prz05, Rij03, RBSP00, RBRSP00, RGDN02, SMC^{+}05, SH03, VMN^{+}03, WGP^{+}00]. radically [DHT^{+}08]. radioactive [TH02]. Radiocarbon [RGDI06, SSRG05]. Radiocarbon-based [RGDI06]. radiocesium [WKYL00, ZWYL01]. radiolabelled [DCT^{+}03]. radiometers [KFE^{+}03]. radiometry [DSP07]. radiotracer [BGD^{+}03, MBH^{+}09]. Rafting [BF03]. rafts [Hob00]. ragworm [KPA^{+}09]. rainbow [BDRM06, BCB08a, SD00a, SD00b]. rainfall [FAK07]. rainwater [ZT07]. ramets [MHM^{+}02, SZ08b]. Ramon [Est04]. ramosissima [OST07]. Randers [AI02, GBB^{+}05, NRPS^{+}01]. Random [AM06, CHPS04, CRC02, SAF^{+}09, TÅ07b]. randomly [SAG03]. range [AAHDS06, APM^{+}09, Arr05, BC02, BGK09, BP06, CEO06, CHT09, CDS^{+}04, DBH^{+}02, DFV01, DB07b, EE06, ESB^{+}04, GMD08, HWGB02, JKK^{+}06, LJ09, LTCP03, PBH^{+}03, SSR^{+}09, SRN02, TLFS06, TAW^{+}09, VN02, WHB^{+}06]. range-extender [LJ09]. ranges [BLH01, HPW^{+}00, TBM^{+}08]. ranging [BEW03a, BEW03b, KCH^{+}03, KCTG06, LLCHG01, MFWD08, RW06a, RCWG^{+}06, Tya09]. Ranong [ITN09]. rapa [MHW06]. Rapana [MHW06]. RAPD [AMS01, CNT^{+}04, EGPT03, HWvdM03]. RAPDs [GPK^{+}04]. raphidophyte [BGC06, KLR^{+}06, YNM^{+}07]. Rapid [DPS^{+}03, DWB04, HGA07, WTH04, WASW03, WS00, AL03a, BAB^{+}06, HDDH06, Uth06]. Rapid-response [DPS^{+}03]. rare [CU08, GC02]. rate [ASLD09, BLS06, BGJ^{+}07, BVH02, BBT02, BSPK03, BM02d, CRSL05, CNW^{+}07b, Cra01, DJM08, DVLH^{+}00, EPM06, FKSW09, FR06, FFG03, FHK^{+}08, GM01, Hel00, HHW^{+}07, HM05b, HPR05, JPC00, JSIT07, JSH01, KR05, KYYM00, LC03, LG01, LMW^{+}01, LP01b, MOOB06, MW04b, MOPP05, MSH^{+}03b, MB04, MBG^{+}08, Mun07, NMS^{+}02, OAM02, ONA04, PBCB03, POH^{+}02, PBS^{+}04, Pet04, PJ08, RSK04, RB02, RR05, RBDK00, Rii01a, Rii01b, Rii01c, RL07, RCWG^{+}06, RC00, SSN^{+}09, SAJG06, SDS^{+}03, SLCB00, SF02a, Sim00, SP01c, STJ03, SB03b, SAC^{+}03, SOMM06, SMHG09, TTM^{+}05, TW04b, TJSG03, Thr09, WJWL05, YH04, HPR05]. rate-dependent [TW04b]. rate-hope [Pet04]. Rates [DSS04, LRM00, TPQ^{+}03, ADA00, ATTX00, ÁSG04, BK00, Bay04, BHFC08, BH09a, BDGK^{+}06, BWCW02, BWK^{+}08, BCLS06, BH04, CWMD06, CWT^{+}01, CSMK09, CPK00, CST08, CD03c, CABS02, DvD02, DM09, DH06, Edm07, EZTZ02, ESQ03, EB04a, FIS04, FFHW07, FRJ09, FBD06, FON03, GVM^{+}05, GP04a, Gim09, Gor03, GDL04, GLG00, GK00b, HMKR05, HL05, HLTT09, HK02b, HBH07, HCBP05, IMS05, JM01, JS03, JSH01, Kam00, KPN03, KM03a, KHHM07, Kiø03, Kir02a, KCEB09, KKSR05, LBBH08, LGB03a, LWR^{+}02, LLRDT05, LVA07, LJC00, MGC^{+}09, ML00, MDG04, MGTK00, MZJ06, MHV^{+}00, MGRP07, MGL^{+}00, NvW01, NKS03, Pet02a, PK05b, PK06, PKA06, PD05, Pur07, RGP^{+}03, RRIH^{+}02a, Rii04a, RGDI06, RSA05, SD09, STPA04, SBB09, SL05c, SS09c, SWH04a, SM05, TW04a, TP07, TNT08, TIM^{+}01]. rates [VSRGW09, VD00, VÅ01, War05, WNS^{+}05, WBW^{+}06, Wid01, Wil04, Kir02b, Kir02c]. rathbunae [RAA02]. rather [CTvdM^{+}08, DL01, GSB^{+}06, RL09a]. Ratio [HDGM01, CSB03, CPSC01, KKIV01, MWTB00, NSEL00, OAM02, Pah05, SSN^{+}09, SZL^{+}05, VSB04]. Ratio-dependent [HDGM01]. ration [ASC06, SLB07]. ratios [BGRD00, BS04c, BCC08, Cal05, CLP^{+}06, CIT04, CAMC07, CDK^{+}07, GBM01, HYLT03, HMK06, IAKS03, JWM02, JMSM08, KTNK07, KKMH08, KW02b, KG07, LSH^{+}07a, LSMR05, LHB09, LGZ^{+}09, LIT05, MND^{+}02, MC05c, MPB^{+}09, SMHL00, SCIT09, SHSV04, TSI02, TCW^{+}07, VRJ02, VWH00, WWWR09, XYH^{+}09, ZS04]. rbcL [JWTP06, PAK^{+}00, PAW00a]. rbcL-containing [JWTP06]. rDNA [ISSK03, OPSM04, PG01, SdV06]. rDNA-based [ISSK03]. Re [EPM06, JCMR^{+}03, MA02, AB00, CH09, ROCB09, RQF^{+}00, WVL01]. re-assessment [RQF^{+}00]. re-calculation [WVL01]. Re-colonisation [EPM06]. re-evaluating [CH09]. Re-evaluation [JCMR^{+}03, MA02, AB00]. re-examination [ROCB09]. reaction [DH07a, GRB^{+}03, GHR^{+}07, Kra07, MKR07, VMCG09]. Reactions [CKTE01, KCH^{+}03, KCTG06, NL05, VFV01, WW05]. reactivation [PKR07]. ’reactive [AMT03a, DCT^{+}04, LLY^{+}01]. Real [DHG^{+}04, GF02, VCSL^{+}09]. Real-time [DHG^{+}04, VCSL^{+}09]. realignment [HFH09, PH05]. realistic [BHZW08, PSGP04]. reallocation [RH08]. really [BN04, She01]. reanalysis [SLL09]. reappearance [BJN^{+}05]. reappraisal [OP02]. reared [ATLC03, CWT^{+}01, FOFG09, LGBWH06, MCR03, RRIH^{+}02a, RRIH^{+}02b, WPPS04, YOM^{+}09, vdMM03]. rearing [CWC00, DGH08, GBW^{+}07, HDRC^{+}08, SIK^{+}02]. reassessing [BT00]. Reassessment [DC07b]. rebuilding [HGSH08]. recapture [ACJ^{+}04, BGDS05, GAMB05, GDSB04, JPC00, PBAD05, UB02]. recently [BWCW02, HB04, MHCB08, SML02a, SCAF03, THM^{+}03]. receptivity [BM05b]. reciprocal [HGMH06, WC03]. reclamation [EvBB^{+}09]. recognise [HRDA05]. recognition [GK08, HD05, KKG06]. recognized [TS04]. recolonisation [CT04, DWB04, LTH^{+}06]. recolonising [CTH^{+}03]. Recolonization [CK05, LPK^{+}07, APM^{+}01, AHDT^{+}09, SG01]. recommendations [KHHM07]. Reconsidering [Cor09]. reconstructing [CLB^{+}09]. Reconstruction [WPRS05]. reconstructions [KW02b]. record [CSMvE04, HGAS^{+}01, RSNG09, TND^{+}05]. recorded [DFW^{+}03]. Recorder [SRR06, JN09]. recorders [DPS^{+}03, SJM06]. recording [EB08, JdSM09]. recordings [RCWG^{+}06]. records [ARAMPW02, Lou08, MFWD08, NM09, RSNG09, RSY^{+}06, ZS04]. Recovering [CRT^{+}08, LRM00, MRN07, Nem05]. Recovery [FRJ09, KHIA02, KGRW06, MG05a, MAG07b, RAH^{+}02, TGGA04, VFP^{+}07, AL09, BBD^{+}02, BMF06, BDW^{+}00, CDH^{+}09, DER01, DBF^{+}02, EE06, EBT^{+}02, FNDP08, GF05, GDBF06, GDBF07, HP09, HKWM07, HW03, KCH^{+}06, KWV^{+}05, LS08a, LBB08, LRM00, MMCF07, MMC08a, MAVO^{+}06, MARM06, NYvW03, NSBK05, NTH^{+}02, PMGE01, PBB^{+}02, PMT^{+}02, RPS^{+}04, RM06, RSA05, San00, SAJG06, SHS08, VMV^{+}08, WWGR07, YH04]. Recreational [MAW06, ABC^{+}05, EJKN03, PE05]. recruit [EM04, MMB03, Sha09a, ST07]. recruited [BDGK^{+}06, Lam00a]. recruiting [BG07b, FBBV09, RLFEF03, SAM^{+}01]. Recruitment [ALOT01, BC06b, BKDB00, BPCvdM06, FC00, GC06a, LTFH06, MARM06, PS08, TCRT^{+}07, WKD^{+}04, vDvHT^{+}09, ABF07, AH08, ACH07, AA04, AR02b, Bai00, BS00, BM00, BM02c, BJH04, BTKG^{+}08, BP09, BDCvB^{+}09, BO00, BB00, Bow05, BM07b, Bul05, CA00, CH06a, CMB^{+}09, CID^{+}06, CÅJ05, CHPS04, CD07b, DL02, DPM02, DTB07, EM04, EE00, Fis05, FD04, Fox04, FWH^{+}05, Gil02a, Gil02b, GW06, GCVR01, Har06b, HSRF02, Hum00, HW01, Jen05, Joh02, KFL09, KCSPT09, KWB^{+}06, LBPS03, LNV^{+}05, LTNC07, LB01, LJ08, LB05a, LPK^{+}07, LBD00, LAW06, LS02, LWMW02, LHB^{+}06, LW04, LHHB05, MW04a, MP06, Mil07b, MM08, MFM^{+}09b, MFM09a, NMJ02, Nod04, NH03a, NH03b, OSL07, PPV^{+}06, PHM^{+}08, PBF07, PSS06, PSRB00, PB08b, QR03, RBJB^{+}05, RHBW09, RS06, RCF07, RLFEF03, RHSN07, RT00b]. recruitment [RT00c, RAS07, RvdPHW04, RS05, SHH09, SM08a, SVG07, Sha09a, Sha09b, SS09c, SGCD01, SEH07b, SGCP06, Sta09, SRS04, SOB^{+}06, SCAF00, SJH^{+}04, SdNH^{+}06, TW04b, TSIY06, THS^{+}02, TTN01, TRT04, TCT^{+}00, VKH08, VHK09, Ver05, WR06, XBW07, ZA07]. recruitment-environment [PB08b]. recruits [DB02, GW06, GDBF07, HHV06, LFRG^{+}05, MBW07, PMV06, SC06a, SHSF08]. Recurrence [DHZ02]. Recurrent [RF06]. recycling [BWP^{+}06, DP02, FAPA^{+}09, RMH^{+}09, TGM^{+}03, WTH04, WMCF08]. Red [AQHTL02, ARAMPW02, DH07b, GC06a, KSS^{+}08, KHGA02, KK02, PDG^{+}02, RBRH02, BCBVM06, BT00, CÁ04, DCAL^{+}02, DdNS06, FOFG09, FVAB00, GPSH01, GGDHRA00, HLR07, HHH02, HB08b, JWM09, JTLG09, JM03a, KJ04, LCL^{+}03, LPK^{+}07, MB08, Nem05, NCHP05, NP05, NCP^{+}08, OST07, PdNS06a, PDHH06, RW07, SCMB03, SAHF03, SLT^{+}01, SJK^{+}06, SSL^{+}07, SJ02, SML02b, TMAN08, TDP^{+}00, WLJM05, WCF08, WPPS04, WdNS00, Yin03, ABB05, BWPB04, BSS05, BIC^{+}03, BI03, GS02, HR02b, HR02c, KAPTK02, NRJB04, Som00, SBBS^{+}02, Wie03, WRJ^{+}05]. red- [SLT^{+}01]. red-footed [WLJM05]. red-tailed [LCL^{+}03]. red-tide [HLR07, KJ04, SJK^{+}06]. redds [CTLR03]. Redefining [BS05a]. Redistribution [BD03]. redox [EMD^{+}01]. RedTeld [Pah05]. reduce [AH08, AF06, CD03c, GL02, MM07a, SAH^{+}05]. Reduced [Str02, BN01, CHMR05, ITE^{+}04, KF01, KB06, Mar04, MCCP09, MGRP07, MAG07b, NvWY05, PKS00, RGDN02, WZ04]. reduces [BS07c, CNW07a, EC07b, HCV03b, HJM07, MRN07, MH08c, Thi05]. reducing [LNV^{+}03, VFJ06]. reductase [LS05, YLv^{+}05]. Reduction [O’D08, AG02b, DBJ^{+}00, HSSS05, KGTH01, MCL07, MDBC03, MBH^{+}00, NWFV06, PCL^{+}09, SR09, TSW^{+}09, VNJ06]. reductionist [SM00a]. reductions [FT00, VHB^{+}07]. Redundancy [MFR01]. redundant [GPHC09]. reed [PAMJ03, VGB^{+}05, WWAC00]. Reef [AB00, AGDP^{+}06, BB08a, CMS^{+}01, EC07b, JMHS03, LAK^{+}08, MB02a, MBH^{+}00, MSS^{+}03, MBP^{+}06, PKM04, PCBS05, Uth06, WPRS05, AR02a, AE02, ADK^{+}06, AAGS02, AFHS09, APA07, AH08, ABHE09, ACH07, APH^{+}04, AA04, Arm00, AKG02, ABB05, AC08, BF07, BJC02, BP08, BL04, Bel07, BF01b, BPM05, BFLÁ01, BYH^{+}06, BES^{+}08, BDDG02, BB05, BG02, BS04c, BH02, BH09b, BL09, BHSR00, CNW07a, CDS07, CH07a, CM07, CHM06, CPBS07, CWE^{+}08, CSM02, Chi02, CS03a, CHPS04, CMD^{+}09, CEBR07, Cum02, DDV^{+}08, DE01, DP01, DF04, DB03, DB05, DBN^{+}09, DBT^{+}02a, DGNvdV05, DGdG^{+}09, DWL09, EJM01, ESDG08, ESS09, Edm00, Edm04, EG08, EFL^{+}09, ERR06, EEWBdN08, FD00, FAJM07, FD08, FA02, FF03, FBJ00a, FB03a, Fis05, FFHW07, FBM08, FB07, FD02]. reef [FBM07, GJ05, GYLÖ06, GPLAG04, GPHC09, GM05, GB05a, GDKN08, GCS06, GL02, GTFB03, GB06, GCM01, GNC07, HP09, Hal04a, HF08, HWC^{+}06, HW09, HCH07, HKY^{+}02, HM05a, HM05b, HF06, Hir06, HSB08, HM06c, HOVG07, Hue08a, IE06, JS01, JBF08, JSMM06, JNFL04, KKG06, Kav05, KHP09, KN00, KH07b, KPEF00, Kuf01, LBH^{+}04, Lec05, LG06, LSG08, LCEC02, LCE03, LHH09, LB02, LG01, LPS08, LC09b, LVZW06, LS04, LSS^{+}09, MTF^{+}08, MRN07, MGL^{+}07, MH09, MCBT07, MB07a, Mar00, MG05a, MGM^{+}07, MGW^{+}09, MS06, McI03, MAW01, MCM^{+}03, MDW^{+}09, MVK^{+}09, MVdMFF08, MSUK01, MAKMK09, MCE^{+}01b, MCN09, NDV^{+}00b, NKK^{+}01, NvdV02, NRvdV^{+}02, NvdV04a, NvdV04b, NvWY05, NReZW09, NRJB04, NMJ02, OYNO08, OMP^{+}05, OT08, OM04, PDS^{+}07, PB08a, PCC07, PFZS^{+}06]. reef [PGP03, PCB08, RBRH02, RCAK03, RFLC07, RWP07, RDC06, RDH^{+}09, RFB07, RL05b, RW06c, SSL00a, SPL00, San00, SBvD05, SS00b, SPBM05, SBL08, SF02a, SS09c, SFH00b, SYWM05, SLL07, SLL09, SWD^{+}02, SP04, SGCP06, SLHR09, Sta09, SMF^{+}02, SF02b, SLHT05, SJ04, TM02b, TJM00, TSMC09, TCWR08, TLPG06, TBSJ^{+}04, TWT08, USG^{+}08, VHB^{+}07, VKH08, VHK09, VHW^{+}07, VNdG^{+}06, VW00, VYM06, WH00, WBW^{+}06, Wie03, WRW^{+}04, WTH04, WRJ^{+}05, WPH01, WWMF06, WA03, WM01, WBC01, WM02b, WS08, Woo09, ZDA^{+}06, Zam07, ZSR03, dBKSV00, dGvdBvO^{+}08, dlMPN^{+}03, Ber02a, Dun02, BM02a, Ber02b, BBA06b, Boo02, CWE^{+}08, CMD^{+}09, FPKC05, Gru03, KHIA02, KHP09, Kri02, LRM00, NMDS00, PCMB06, SRDG05, UBRvO06, WNWH02, WHB09]. reef-associated [DP01]. reef-building [BS04c, CNW07a, Cum02, DDV^{+}08, FFHW07, FBM08, GNC07, MDW^{+}09, NvWY05, OYNO08, VYM06, ZDA^{+}06, dBKSV00]. reef-fish [MTF^{+}08, OT08]. reef-seagrass [DGNvdV05]. Reefs [VLK04, Air00, ATR^{+}08, Ant06, AKG02, BSZ05, Ber02a, Ber02b, CBGLM08, CCC05, DSL07, DGC^{+}05b, DVNvdV07, Dun02, DS07, FSHB09, FH08d, GP04b, GEC08, HWG06, KAMR00, KSR^{+}07, Lev08, LH02, LWK^{+}07, LAK07, LRM00, LKW03, MHO07, MH08b, MAOP09, NBN^{+}08, NHS^{+}08, NMJ02, NNLF09, OM04, PS08, PGPL03, PPGL09, PKS07, RV05, RCDC06, RM06, RL08, SRS05, TB08, THBSJ05, Uth01, VLW09, WPRS05, WWV08]. reefscape [AGDP^{+}06]. reference [KWV^{+}05, LTFJ00, TL05, VHH^{+}02, VFV01]. referenced [OWBD07]. refined [ASLD09]. Refining [MBA^{+}05]. reflect [BHH09, CHW02, FSB07, KMP^{+}06, SPH^{+}09]. reflectance [APH^{+}04]. reflected [TOK08]. reflects [ABD^{+}03b, AAJ07, RP06, ZBNW09]. reforested [BKK^{+}05, BDGK^{+}06]. refuge [AP04a, AP04b, BS07a, Bar02a, GW06, Lev08, MHM^{+}08, RST04, SBJ02]. refuge-use [Bar02a]. refuges [DHR04a, JBF08]. refugia [Bas06]. refugium [MKK^{+}09]. regalis [TT07b]. regenerated [FR07, GGK^{+}07]. Regeneration [VS08, BKK^{+}05, CBP^{+}01, DR00, FR07, FOL02, FFHW07, HGT01, Lir00, Maz04, SMH^{+}03]. Regime [WLZ05, ADS^{+}09, Ano07g, HFM^{+}02, Hun05, HMFC07, IMSO09, LBPH06, PBF08, TOK08]. regimes [FNB07, HDP02, LB08b, LDT07, MHSS08, MND^{+}02, VB02, WB07, WDX01, WO05, vdMM03]. region [BRG^{+}02, CCL^{+}05, CKF^{+}03, CGL^{+}02, GPSH01, HKH04, JDCB09, KCK^{+}06, KI01, LB05a, MM06a, ML06b, ML07, MS04b, MFM^{+}09b, MFM09a, ONKK03, OFHM03, PIY04, Phi05, RS00, SK04, SK05, SAMN04, SDK^{+}08, SB08, TCRT^{+}07, WKH02, WKIK04, WKMK06, WKY09, YI03a, YI06, QR03]. Regional [Eyr00, FWH^{+}05, FEMW07, JCD^{+}05, KG07, LBvdH03, LSMR05, MAKMK09, PPB^{+}09a, WPTS08, Bar03, BGK^{+}06, BRI^{+}07, BBA06b, DH09b, FBT^{+}01, GSP^{+}02, HF09, HM09, KI01, Ley05, LRM00, OP09, PKS07, RC04a, RFN02, SAF^{+}09, TH02, WAB^{+}02]. Regional-scale [PPB^{+}09a]. regions [ABD^{+}03b, CMLG07, FBHB03, GMTN00, HFM^{+}04, MAC^{+}07, OP08, SKR08, SMF00]. regression [BRP07, EMG01, EYH^{+}06, LRK09, LEF^{+}06, PB08b, YMHV02, dM01a]. regulate [SEM01, SES01]. regulated [ABBF01, HWL00]. regulates [BKK^{+}05, CMLG07, LAK^{+}08, SLM07]. regulating [ODF09, SS05, VHKW00]. Regulation [KHWH03, KNU02, RPN^{+}01, Rit06, BBRJ04, CW03b, CS08, DDV^{+}08, GT02, Hue08a, LR02, ML00, PB07, PWC^{+}09]. Regulative [Hen07]. regulator [BDRM06]. regulatory [DM05b, QS05]. rehabilitation [Fox04]. reigns [SSLM01]. reinforcement [AE06]. reiterative [MBA^{+}05]. Rejection [BJLH04, BC06a, JY08, UNII01]. related [AS06a, AG02a, ABB^{+}00, AWA^{+}05, CST08, CLMR07, CLRM08, DSF^{+}04, DL03a, DF06, EDB^{+}04, FSHB09, FWH^{+}05, FEMW07, GB02a, GCHC00, HTM^{+}02, JBHGB04, LBH^{+}04, LSdM^{+}03, LK02, LVK05, Lus05, MOP^{+}08, PD01, PWL^{+}06, RWA^{+}02, STPA04, SMH02, TAM04, TC01b, UNII01, VSBB03, VFJ^{+}07, War05, WM02b, WC03, ZLGC04, vBWSS06, vDvTJD^{+}09]. relatedness [Bat09, TLP^{+}03]. Relating [CR07b, AGDP^{+}06, Mar00, TNHL02]. Relation [JP04, AMR01b, AL03b, ABB02, BIR00, BI04, BDG^{+}04, BPVL03, BH04, CHLV07, CCCH01, CGC^{+}09, DJG^{+}07, DFV01, DGH08, DDS^{+}01, DDL^{+}03, DMV07, DTB07, DFR08, Ell02, EP01, FPC05, FFG03, FHQ^{+}06, GBG00a, GRWB02, GDL^{+}01, HC05, HDP^{+}06, HJK01, IRL^{+}09, IP01, JBCL03, JH01, JS02, KKTN05, KSG03, LTV^{+}01, LT00, LRGdS00, MB08, MFG^{+}04, Mad02, MHSS08, MND^{+}02, MNC^{+}02, NLHO09, Paf06, PH03c, PHS09, RS03, RF08a, RMH07, SML02a, SWM03, SvNvdMvdV09, SLHT05, TSE^{+}08, TN07, TAW^{+}09, TT05b, TT07b, VHMB09, WMS^{+}03a, WP07, WAN^{+}01, Wil09, WJB09, WWS^{+}05, XBW07, ZBDE07, dSCV^{+}08, vDvHT^{+}09, HHKH05]. relations [SPL00, SK02]. Relationship [BMK07, Bra01, CSB03, EG03, KSWP04, Les05, MR06a, TBC^{+}00, ABBF01, AFFH04, CA00, CSMBH08, CG04, EdPP^{+}00, FCS^{+}06, GSCW00, HCV03a, HGMH06, KH08a, KVV^{+}00, Lee00, MW02b, MWTB00, NL05, PJR^{+}03, PEG^{+}00, RFN02, SDS^{+}03, SEH07a, SN02, SMGCW00, TNSM00b, WLB08, HHC00]. Relationships [AI04, ABHE09, BJC02, BW04, GGHA02, Gor03, GMd^{+}02, LGZ^{+}09, MW04b, NTGK09, RMM04, RB04, RZEACGF00, SYA02, TBSJ^{+}04, BML^{+}08, CLO^{+}09, CC02, CWT00, Chi02, DJ04, EJM01, EHM^{+}09, FF04, FH08c, Gim09, GSP^{+}02, GKWK02, GK03, GGDB09, HH00a, KH00b, Lei08, MMB03, MB02b, MSF04, OJPM01, PB08b, PMNP^{+}01, RZS05, RCGA^{+}04, SJR00, SWW02, SZ09, SIK^{+}08, TRH^{+}06, TC01a, THHY05, TT06, TPD^{+}05, TBH06, VHK09, VGS^{+}01, WFT^{+}08, WA03, WHB09, YH02]. Relative [BF01b, BSvdH^{+}02, CVF^{+}03, CR02, GS09, GHS05, HBRW04, LR06, LW02, MSLP06, MH06a, NvdV04b, RRRM08, SZZ08, STA^{+}05, WWAC00, AF06, BS08b, BOR08, DGdG^{+}09, GDG02, GNHB^{+}05, GEC08, GRM00, JCZ^{+}05, JRT^{+}06, KW05, KWTP06, KS00b, LBH^{+}04, MRC^{+}08, MMS09, MHT07, OPSM04, PPB^{+}09a, RC05, SF02b, SJH^{+}04, TBH08, WDN00, WMB00]. relaxation [NNLV06]. relaxed [CAF^{+}08]. Release [HKM00, OBB00, BD09, BWB^{+}06, CPB04a, DR00, FLK^{+}06, JR01, MMV05, MYH07, MJG08, MMBF09, SSB09, SAM^{+}06, VH03, VSBB03, WB01, WMW^{+}08, Woo09, ZDA^{+}06]. released [ACR07, INH^{+}09]. relevance [FVS^{+}08, MMTT00, RC04a, SDRH01, BPS^{+}00]. relevant [AM02b, ABY00, BUA02, CCB06, CST08, FTBC05, LHH09]. ’reliability [Cra01, HHAH00]. reliable [Rii01b, Rii01c]. relicta [VV04]. relocation [HSW^{+}08]. remain [FSO08, Gil02a]. remains [Yos08a]. remediation [PZ07]. remineralization [GCHA05, MYNY08, YYK^{+}08]. Remote [JBS^{+}09, WGS^{+}02, BSDR01, BPS^{+}03, BYH^{+}06, CDS07, DGVK^{+}04, FFYS^{+}07, GPLAG04, HCB^{+}07, KT01, MHP07, MTCU08, OT08, PR05, SCD^{+}05, vdWHF^{+}08]. Remote-sensing [WGS^{+}02, PR05]. Remotely [WR06, PNK^{+}02, TFL^{+}04]. Removal [AGB03, SLT^{+}01, WKL^{+}03, ABU01, CK04, EJKN03, GXL^{+}08, GPW^{+}02, GLJ07, GDL04, GML02, HLH^{+}08, LCT^{+}08, PTK^{+}09, RCAK03, WDR^{+}05, dBGLT08, dGvdBvO^{+}08]. removing [FPKC05]. renewal [BGJ^{+}07]. Renicola [Thi06]. reorganisation [WCZ^{+}09]. reorganization [LBPH06]. repair [BB02b]. repeat [WGM02]. Repeatable [CTD^{+}09]. repeated [BHSR00, UDB^{+}09]. repetition [FHK^{+}08, LG01, MOOB06, MSH^{+}03b, RSK04, SMHG09]. repetition-rate [MSH^{+}03b]. replacement [TS05]. replacing [MPDC03]. replenishment [HWC^{+}06, McI02, SP04, WM01, WM02b]. replete [KWAT08]. replicates [RKJ01]. Replication [BCO07, BSB03]. Reply [Ber02a, GNAD^{+}04, HSJ08, Ike08, Kir02b, Kir02c, MH09, MMPD09, MB07e, Yos08a, CPB04b, DBCB06, LTMF01, Rii01a, Rii01c, SRJ07a, DM05a]. reported [SF05]. reports [WG02a]. represent [HM01]. representative [KGS^{+}08, TW10]. Reproduction [LME05, MGH00, NH05, SCMB03, AM03, ASS05, AW03b, ANF^{+}00, BM02a, BRAS03, CT05, CSP^{+}03, DDD08, FCP^{+}07, FA08, FKA^{+}06, GCLC^{+}02, GAZ02, GB05b, GRM00, HRR^{+}07, HHC00, KS08, KLC^{+}06, KAPTK02, KSKL^{+}03, Lam00b, LOB03, LP07, LDR00, MC05a, McG02, MP06, MW02a, MW02b, NTGK09, PHP^{+}03a, PWL^{+}06, PLG09, Pur07, RGR^{+}01, RR06, SD01a, SBM02, SJ09b, SRLA04, TAA04, TYD07, WLS^{+}09a, Whi06, WS01, WG02b, WO05, ZMN07]. Reproductive [BB09, BFE^{+}03, DFY00, DD06, FA08, GBF^{+}02, HH01, HVCT09, KP00, PCS09, STE04a, SM07a, TK04, WGL03, ABF07, AMR01b, AFPS08, AM04, BPB07, BDH01, BDEM01, BR09, BG04c, CAT^{+}06, CNW07a, CLSM09, CHW06, CTVP03, CY06, CH00b, CPG08, DRDM09, DD09b, EB04b, FC08, FFR^{+}03, GDSB04, GL04b, GKS^{+}07, HC05, IMS05, JM01, KD00, LBBH08, LGRB^{+}08, LP07, LP03, MB02b, MH00, MY02, MM03, McI02, MSF^{+}09, MAA^{+}07, NJPI03, NHNB01, OLGG^{+}02, OGL09b, PJ06a, PCC06, RMA^{+}02, RHG04, RZLS09, RWH00, RHSB08, SAWG05, SV07, SHZ^{+}05, SPH^{+}09, TS07, TBAT02, TD01b, VFJ^{+}07, YW04, ZMN07]. reptile [MBH07]. require [GPS07, TW10]. required [GP08, MB02b]. requirements [BKM08, Cam01, EM07a, HSCD01, WTR02, ZSKA01]. requires [Wie03]. RES [KWTP06]. Research [Kin05a, AC08, BSW^{+}09, Cai09, CV06, EP00b, HRB^{+}06, HP07, Mar01, Nae06, RBL^{+}04, Rii04b, RZWG08, SRP^{+}06, SvAHH09b, SvAHH09a, VCSL^{+}09]. resedaeformis [SSRG05]. Reserve [MHO07, AFHS08, AFHS09, ABHE09, CH05, DWB04, FPKC05, FMT09, GQR06, LP03, LTCP03, PLCG^{+}05, PBH^{+}03, PPJ^{+}05, RSC08, SB03a, SNP03, SJA^{+}09, THC^{+}05, ZSR03]. reserves [AMR01a, Bas06, BL04, CH06b, EP08, FSB07, GS07, JWM09, JM07, Kap06, LH08, LHGC^{+}09, Ley05, MM06b, MVHV06, PMS^{+}08, PBT^{+}09, RG09, RAM03, RSA05, SZW^{+}07]. residence [GGDHRA00, JTVC08, JSIT07, UDB^{+}09]. Residency [Lus05, CS09]. Resident [BHH07, CM07, BSBD^{+}04, BBA06a, BBC08, HLH^{+}07]. residual [DBvOM06]. Resilience [CBP^{+}05, Vie09, EWR08, HSB08, MP05, NZM08a, WDB^{+}05]. Resistance [HSB08, LELC02, DHCH04, DKAH02, HHJ04, HH07b, LK02, LVK05, Mar00, MMGC06, NZM08a, PGBA04, PJ06b, Rue07, SJS^{+}01, SMC09, SM09, Tom08, WKH07]. resistance-related [LVK05]. resistant [BMM^{+}09, OP09]. resolution [AQSP07, CSD04, CCMK03, EM07b, Hir06, KH00a, Las03, MB07e, Pee02, SHNV09, WKK07]. resolve [MSF04]. Resolving [ST08]. Resonating [RJTM08]. Resource [BMO^{+}04, CGP08, CLJ^{+}08, FIS04, JSMM06, KHF^{+}00, SF06, TBAT02, BD00, Bar03, BKB05, BM02d, FBBH07, FOL02, FMZvW02, FHB^{+}02, FBD06, Hal04a, Hor06, JBA02, KS08, KWP07, LJM03, NvdVVD06, NFN^{+}09, PMG05a, PWLC06, PHM^{+}08, RBDK00, RGF^{+}03, RB04, SB06, TK01, TWM^{+}01, WKPS06]. Resource-restricted [CGP08, TK01]. resources [AP04a, AP04b, BS04b, BS05b, CC02, CBEL06, HMW^{+}00, Les05, PPC^{+}05]. respect [HKSW01, MWS^{+}04, RBGM06]. Respiration [ISY07, MR07c, SHT^{+}04b, AMP^{+}02, BP01, CRSL05, DDQ^{+}00, ESQ03, FAPOF09, FKA^{+}06, GK07, HKH04, HHIS06, MTR^{+}07, MASFS00, SKH06, TN07, VLP^{+}04, WA05, ZK01]. respiratory [MG09b, VNJ06]. respirometry [SM05]. respond [Bar07, FSC05]. Response [ADA00, BD07, Ber00, DAA00, FNDP08, HKWM07, IOTL09, JDTB06, Kim06, KBSWG09, LFRG^{+}05, LS05, LV05, MKG^{+}05, MS09, MBH^{+}00, MMA03b, PZ07, PE09, RC07, SHL^{+}03, SYWM05, SBA01, SW08c, UP01, WW04, WCRB03, Arm00, BVMBC07, BL00b, BB02b, BSPK03, BLH02, CCK^{+}09, CT04, DPS^{+}03, Dav00b, DMC09, EH00, EKN08, EKA03, FE02, FHL^{+}09, FBKB03, FH03, FHN^{+}09, GSM^{+}08, GRIR06, GHL02, GEHZ07, GB05a, GDR^{+}04a, HMH04, HB03, HB04, HHM^{+}06, Hyr01, JWM09, Jak01, JHR05, JRS^{+}06, JF02, JK03a, KCH^{+}06, KHD07, KFT08, LOT^{+}00, LCL^{+}00, LMW04, LTH^{+}04, LTL^{+}06b, LFSR06, LS04, MBC^{+}07, MAJ^{+}06, MS04a, MR06b, MMB^{+}02, MK03c, NWHH01, NZM07, NR00, Ogu07, Pea03, PKRK07, POH08, PAR08, RASB03, RWJ02, SRDPT^{+}07, STRD02, SC01, STV^{+}09, SBFP08, SV01, SB07, SYCG05, TJEM02]. response [TJM03, TC03a, TP00a, TBAT02, TGNT05, Tru02, TH06, VN05, WWDH09, WDRH07, WASW03, WKPK05, WBDR06, WPTS07, YMHV02, vdWHF^{+}08, GPHC09]. Responses [Air00, BN01, CWE^{+}08, HLH^{+}08, MTL^{+}03, PUB^{+}01, PPHT00, SLCB00, XWQQ05, BCB^{+}06, BGK09, BLS03, BSW^{+}09, BEW03a, BEW03b, CLS00, CH07b, CRK06, DP09, DY05, EJKN03, EH04, FK08, FSW02, FR06, FSV^{+}08, FR08b, GLBF07, HF08, HDGM01, HRCB02, HVH^{+}07, HFD08, HPVL05, IBB^{+}07, JHL00, JS02, KOTL07, KHF^{+}00, KSR^{+}07, LCHG03, LH08, LT07, MB08, MAG^{+}07a, MDW^{+}09, MSJ^{+}09, Mon07, MHT07, MLR02, OAS01, PWHW09, RW07, RC08, RMRAFP08, SR09, San00, SDHG03, Sim00, SW08b, SO01a, SHSV04, SBI07, SB03b, STM06, SJA^{+}09, TW00, TFA^{+}08, Tho07, THN^{+}03, Tya09, VMV^{+}09, WDLM05, XW01, YG09]. resting [HW00, KSA03, McQ05, ONA04]. Restoration [CWKE05, HS07, MPDC03, PHP^{+}03b, DSH03, FBN^{+}07, GPHC09, HFH09, MR03, MGC03, MVK^{+}09, OHB^{+}00, PKM03, PPGL09, RMZC07, RM09b, SPD03]. restorations [PL03]. restored [CB08b, TS06]. restoring [PGP03]. Restricted [UB01, ZBNW09, CGP08, MSTS^{+}06, TK01, TBM^{+}08]. restriction [UYNO01]. restrictions [DBN^{+}09]. Restructuring [FTO^{+}09]. result [DBJ^{+}00, TAS04, TVG00, WS01]. resulting [CB05a, JCL09, PGP03]. results [ASR^{+}09, BRB^{+}00, BCC08, CHL^{+}05, FJ01, HCM01, HVP^{+}06, HWvdM03, MC01b, OGMC06]. resuspended [HJG06]. Resuspension [SABK03, CWR06, KKA^{+}06, Kre01, WPB^{+}08b, YYK^{+}08]. retaining [BCU04, JHL00]. retardation [TRS^{+}06]. Retention [FLUF^{+}04, Gri02, Joh07, KHB00b, AKG02, BSH^{+}09, BAE^{+}08, BP06, CSTC01, DCRT07, GKS^{+}07, JM07, JBT^{+}07, LP01b, LHB^{+}06, MADPQ06, MFJT08, NRPS^{+}01]. reticularis [LIT05]. reticulata [HDQ04, PPS00]. reticulatus [BGQR01, BMB02, CR07a, MSY02]. retinoids [TBFA04]. Retracted [SLL07]. Retraction [Ano07h]. retranslocation [AMR00a]. Retrospective [CHLV07, CDT^{+}01, NEMF09, DGVK^{+}04, Sha00]. return [FWD^{+}01, LLH06]. returned [KHB00b]. returning [TWMW06]. Reunion [KMB^{+}08, KFS^{+}09, HLFG^{+}07, RLFEF03]. reveal [BHPC06, BS06, BRI^{+}07, BCB08a, BSIS07, CDF^{+}08, CDW^{+}09, DPS^{+}03, DBOR09, FGSH^{+}05, FOH^{+}08, GLS02, HRR^{+}01, JFD09, MSF05, MH08a, MPB^{+}09, RCWG^{+}06, SBI07, WIB08]. revealed [AMS01, AKL^{+}04, CIT04, DCA^{+}09, DOML^{+}07, DEV07, GMB^{+}06, HCR01, JN09, JSIT02, JSIT06, KTNK07, LF08, MHH^{+}05, MAM^{+}01, MS04a, MRB^{+}08, OPSB07, RHSN07, SCIT09, STCJ01, SAG03, SKO^{+}08, TRH^{+}06, TA07a, TBI07, UDB^{+}09, UM08, WWWR09]. reveals [AS06a, BBA06b, GFF08, HARC05, HEN^{+}08, MBH07, QMM^{+}08, TA01a]. reversal [Zup00]. Reverse [TCC04]. reversed [ZRG00]. reversible [RM06]. review [Rii03, CH06c, CDAM06, Fau09, Fie02, GB09, HMD^{+}07, JdSM09, OHB^{+}00, PUL07, QMF05, RQF^{+}00, Rii01b, RBL^{+}04, Rii04b, TBH^{+}09, WPM09]. reviews [Wid01]. revisited [AR02a, BK06a, CJRT00, DBCB06, Lou08, MG05b, PHS^{+}07, RN06, SSH09, TOK^{+}04, ZGV00]. reworking [FGS^{+}02, MDR^{+}06, MDBG07, ODG^{+}04]. Rex [LTMF01]. Rey [MMPD09]. reynaudii [SS04a]. RFLP [UYNO01]. Rhône [PPCJ^{+}06]. Rhincalanus [TSW07]. Rhincodon [HGKJ01, MBP^{+}06]. rhinoceros [Dav00b, HBRJ06, WIDM09]. Rhinoptera [NRC07]. rhizocephalan [BBSO03, RAA02]. rhizome [DT08]. rhizomes [KECJ06]. Rhizophora [ATTX00]. rhizosphere [HGK02, HCA06]. Rhizostomeae [KUOI06, PK00]. Rhode [Lap01a, Lap01b, CWT06, MGTK00]. rhodoliths [SCM09b]. Rhodophyta [Bib07, RZEACGF00, SD01a, TMM05]. Rhone [DSPHV04]. Rhopaloeides [WNWH02]. rhythm [EdPP^{+}00]. rhythms [ACAG07, KYML02, SC00a, SAS^{+}06]. ria [RYF01, VPP04, BRAS03, ÁSG04, NIF00, PNCS^{+}08, PNCT^{+}08]. ribosomal [PSPD07]. Rica [IMW^{+}02, TBBB06]. rich [AP05a, CDK^{+}07, DW09, HDC^{+}08, MD08b, MN00a, SSG^{+}00, TSC^{+}07, VMdBV06, WS06]. Richards [HLR04]. richardsi [HBBD09, NWIT08]. Richness [SSHW^{+}07, ARP09, BS03, BAMJ09, BTW08, Bra01, CF08, EF07, FAFR03, Gra02, HF08, ISBP06, LEF^{+}06, MWM^{+}05, RFN02, RKJ01, SH07b, SLM07, TFSP09, UG04, Zhu06]. Rico [APA07, DvD02, PCHM07]. Ridge [SPRK02, TBK^{+}03, IMW^{+}02, OJPM01, CTVP03, HGC^{+}07, HVCT09, KBP06, LJM03, MCP^{+}06, MC05b, UBM^{+}07]. Ridgea [PAB^{+}02]. Ridgeia [UBM^{+}07]. ridley [MBH07, SAM^{+}06]. Riftia [GLG^{+}05, SGL^{+}02b, THS^{+}02]. right [BCC^{+}03, BCCM03, BM03, CDML06, DCM^{+}06, EB04b, HST^{+}09, JBT^{+}07, PPB^{+}09a, PRM^{+}09a, RJS04]. right-whale [HST^{+}09]. rigida [BVN^{+}03]. Riisgård [DBCB06]. Rimicaris [PGS^{+}00, ZLGC04]. ring [WKMK06]. ringed [GNAD^{+}03, GNHB^{+}05, KKL07, TBBI04, WHL00]. ringens [CSH00, CCK^{+}09, LRC06]. rings [CR07a, SSRG05]. Río [JBCL03]. rise [DDGK^{+}08, FR08b, HGK02, KWHD08, TC01b, VBMBC08a, DKFV05, GZG^{+}07, PZMG05, THS^{+}02, Van03]. rising [Edm04]. Risk [JCG^{+}02, BJJ01, BH02, DHR04a, DBvOM06, GRWB02, HFW^{+}05, MPSJ06, NZM07, SW08b]. risks [HP07, PHB02]. risky [SW08b]. Rissa [Bar07, DBH^{+}02, FWH^{+}05, LWW^{+}01, SIK^{+}02]. Rissoella [BCBVM05, BCBVM06]. Rita [PWHW09]. River [CGCMPA01, CIFN00, CS08, HHV06, KHB00a, KS04b, Lee00, MBW02, PPHT00, ST07, SW02b, SR01, Eyr00, KKMH08, MIL^{+}00, MILH00, PCSS01, Pee02, PI07a, BJR04, CB00b, DSPHV04, DMS^{+}05, DL01, DTNU00, LF08, LD03, LDWC05, MDZ05, PAK^{+}00, PPCJ^{+}06, SLFM03, SRZZ01, SSLM01, SZZ08, SW02a, SAM^{+}01, SKO^{+}08, TWYS03, TNSM00a, WDB^{+}03, YQC^{+}00, YQW^{+}01, Yin02]. river-dominated [Eyr00, MIL^{+}00, MILH00]. river-plume [Pee02]. riverine [RE00b, SRZZ01, TSI02, ZHWCS08]. rivers [DL01, SR08a, SICT03, Tal09]. rivulatus [BB08a]. RNA [BS04c, BCC08, CSB03, Cal05, CAMC07, KKMH08, KMAC^{+}02, MND^{+}02, PSPD07]. RNA/DNA [BCC08, CSB03, KKMH08, MND^{+}02, BS04c]. roach [WLM06]. road [Rii01b, Rii01c]. robustness [DWM04, PKS07]. robustus [BWW^{+}06]. Rocas [WGR07]. rochei [MOP^{+}08]. Rock [BB09, Bay02, CWBS03, EF07, GB02b, GFN^{+}09, JWM09, Joh01, MHTJ07, MUP06, PMS^{+}08, RSSA07, RH08, WW06a]. rock-platform [MUP06]. Rockall [GGDB09, HSTS^{+}09]. Rockfish [MS04a, Bla09, FSB07, HGSH08, JKK^{+}06, TAW^{+}09]. rockfishes [BRG^{+}02, HKR^{+}08, SBF08]. rockhopper [PISL02, RTPS07, SR04a, TC00, TC03c]. rockpool [CB05b, Gri02, GDW06, MSC04]. rockpools [BAB04]. rocks [GBC^{+}04]. Rocky [MTH^{+}08, ABC^{+}05, APM^{+}01, BL04, BCPB^{+}01, BCBAC01, BC01a, BCO07, BCB02, BVMBC07, BUA02, BFB^{+}09, BCU04, Bul05, BMJ02, BHR08, CDO08, CWKE05, CLC04, CD07b, DBN^{+}09, ESS09, FA06, FCP02a, FD04, FTBC05, GJT09, GTFB03, GBC^{+}04, Har08, HSM^{+}09, HW01, HW03, JCD^{+}05, KM03a, Las03, LGSB09, LÅCN01, MBW03, MBVBC09, MA00, MMMB^{+}05, OMP^{+}05, PCSS01, Prz05, RS06, RFN02, RBG^{+}07, Rue07, SH06a, SH07b, SMC09, SLL07, SLL09, SDW07b, SHW09, Tal09, TMJ^{+}05, TJS08, TBSJ^{+}04, VBMBC08a, VBMBC08b, Vie09, Wie05, WDN00]. rocky-reef [DBN^{+}09, GTFB03, TBSJ^{+}04]. rocky-shore [BCO07]. Role [AZKJAM06, BTL05, BKR04, Bul05, CD03a, CD07a, CDH^{+}09, DDS^{+}06, DMM^{+}03, EDE05, EDKT06, GDW06, HO02, HAD07, INH^{+}09, JRP01, KPW03, KBPB07, KKB^{+}06, KKH07, LHJN07, LS08a, LAV04, MR04, MRB04, MT09, MHT07, MJKH03, RSC08, RH09, RGB02, SFF^{+}07, SZBZ04, SS05, SNGR00, VHKW00, VAB09, WAW^{+}09, WBH^{+}04, WRH05, AHL^{+}01, ARBF02, BF07, BSR09, BB08a, BSR05, Bro06, BV08, CHSW03, CM07, DW09, DK03, DB03, DdMB04, DGW01, EH00, FBBV09, FHPW07, GA08a, GKB09, HK03b, HHO03, HBT^{+}08, Hip09, HH00b, ITN^{+}06, JVJ09, JPNM^{+}06, KW02a, Koe05, KWHS^{+}06, KS05, KLFH04, LJ08, LGB^{+}04, LSH07b, LPK^{+}07, LC09b, MBW03, MC06, MSA07, NBB06, NWHH^{+}09, NTH^{+}02, OGA03, PABM08, PCJF06, PSW07, PI08, RRS05, RSE01, RMH^{+}09, SHH09, SNH04, SBM02, SS06]. role [SO01a, SMG00, THF^{+}01a, THF^{+}01b, TLPG06, UBRvO06, VWF^{+}01, VH03, WM06, WWR^{+}01]. Roles [MC07, BKB05, CRH01, DS06b, DESK01, FSHB09, GW06, ISBP06, QS05, WDN00]. roman [GKAS08]. Rømø [BAA04, BAA07]. root [BB02b, DT08, EKBM05]. roots [GK02, GKK03, JKG^{+}05, MD01]. Ropert [Ano07f]. Ropert-Coudert [Ano07f]. rorqual [DVBLS07]. ROS [DCT^{+}04].

Roscoff [HDCA06, HAD07, RH03]. roscovita [Thi06]. roseate [MRQ07, MRH^{+}08, RMA^{+}02]. Ross [SHA^{+}05, EBA^{+}03, GJG^{+}03, GGC^{+}05, GRG^{+}02, GGG^{+}04, KSSB04]. rostrata [DLT^{+}06, JSIT02, JSIT04, JSIT06, JSIT07]. rotational [WWMF06]. rotenone [AB00]. rotifer [OTB^{+}00, WLT^{+}03]. Rottnest [OKM06]. rotundus [CD08]. rough [MK04b]. round [GBW^{+}05]. route [ARFP09, BLDG01, SBC04]. route-choice [BLDG01]. routes [CH07d, FWD^{+}01, KS04a, MEHD07, TA03]. Routine [HSS^{+}04]. ROV [TFL^{+}04]. rowing [WCCA09]. rubble [TAS07]. rubens [DCT^{+}04, VTY06]. rubra [BDÁM04, PBFB08, SM08b]. rubricauda [LCL^{+}03]. rubrum [CFA07, GPSH01, SCMB03, SH07c]. Ruditapes [CDFRL03, NW04]. rufa [JR01]. rufescens [BT00]. Rugosity [DH09b, WH02]. Rugosity-based [DH09b]. rule [BBG^{+}08, IF09, NGBG07]. rules [BVP^{+}06, VJKF07]. run [DER01]. runoff [HHM^{+}06, ZT07]. runs [MWG02]. Ruppia [KLN06, MRP09]. Russian [FA08]. Ryukyu [NM09]. Ryukyus [KHIA02].

S [GUL^{+}07, HCP^{+}06, HCCY02, MFE^{+}05, PSW^{+}08, WKL01]. S. [MKWO00, NA04b, TT05b, YSS03]. S/T [HCP^{+}06]. Saanich [DER01]. Sabella [HK02c, PG01]. Sabellaria [DBCB05, DBCB06, DCRT07, RN06]. sabellariid [MFY02, MYE03]. Sabellariidae [DBCB05, DCRT07]. Sabellidae [PG01]. sablefish [SO01a]. sac [BCT^{+}08, LRC06]. saccharina [AL03b, GCD^{+}03]. Saccoglossus [TKR06]. Saccostrea [Bay02]. saccular [EdPP^{+}00]. Saco [PDQ01]. Safety [FDDS09, Ree03, CM04b, DG05, DGdG^{+}09, KWHD08]. saffron [JTN09]. Saga [KKA^{+}06]. Sagami [FTK^{+}04, HHIS06, OKH^{+}07, TAM03, TAM04]. sagax [MBG^{+}08, RVMI^{+}03, WHM^{+}01]. sagitta [SJ06, Øre00, TT05b]. sagittal [NSEL00]. Saharan [EDKT06]. saida [BF08b]. sailing [HR02b, HR02c]. Saint [DCRT07]. Saipan [HvW08]. salar [ABS^{+}09, HMH^{+}08, HHD^{+}09, TWMW06, WFC03]. Salarias [WBC01]. saline [CMHP06, MCH^{+}06, PBS^{+}05]. salinities [KKMH08]. Salinity [CAT^{+}06, HYLT03, LALB08, BGC06, BBRJ04, CSM^{+}09, CAA00, EKN08, EDE05, FB03b, GDS02, HM05c, KF01, KSK07, KJ03, Mad02, ML06a, Mar04, MT05, MW06, MT09, McQ05, MLG03b, PDM^{+}04, PAK^{+}00, PHS09, RB02, RB05a, RDC06, RAA02, SHB03, SMT08, TA07a, TPGQ07]. salinity/high [PAK^{+}00]. salmincola [JTVC08]. Salmo [ABS^{+}09, HMH^{+}08, HHD^{+}09, TWMW06, WFC03]. salmon [BRP07, CTLR03, CHMR05, DMS^{+}05, DUSJ^{+}09, FWD^{+}01, HMH^{+}08, HHD^{+}09, HWBZ05, HFWW05, JTVC08, KdLCM09, MHSS08, MF07, NLB08, Ran02, SAH09, SBHE05, TND^{+}05, TWMW06, UDB^{+}09, WFC03, WFG^{+}08, WHCP^{+}04]. salmonids [Tho07]. salmonis [SBHE05, TSRR05, TWMW06]. Saloum [PDM^{+}04]. salp [HLGCO06]. salps [NSJ02]. Salt [CGCMPA01, MAWH03, VGB^{+}05, AED^{+}07, AHBM^{+}07, BSvdH^{+}02, BCB^{+}08b, CWT06, DT08, DPD^{+}08, FCH^{+}08, GBSQ09, GK02, Hal00, HH03a, HGK02, KCP^{+}07, KWL^{+}05, KMK01, MD08a, MMS09, MA07, MSP^{+}06, OJ04, PMH08, RGCL^{+}07, Rij03, RRRM08, SZZ08, SA03b, TC01b, WHL^{+}06]. salt-marsh [Hal00]. saltatory [LB08b]. saltatrix [HBNA04, MBL07, SLB07]. saltmarsh [CC00, GFF08, GKK03, GC06b, HD07b, HFH09, JFD09, PH05, SMS06, SG01]. saltmarsh-pond [SG01]. saltmarshes [TC01a]. saltwater [KLRS00]. same [LCC01]. Sample [STH^{+}06, CF02, JMB^{+}05, LCC01, SDH01]. samplers [Lin04]. samples [Eva08, GHV02, GHD^{+}02, HYR^{+}09, NR00, Rob06]. ’Sampling [GHP02, BDR06, MA00, PMGE01, BZLD02, CM04a, DHG^{+}04, FAK07, GNAD^{+}04, GP08, HMW^{+}00, HJ05, MAA^{+}07, MVdMFF08, NWHH^{+}09, PTS00, PMGE02]. San [GZY06, GCHC00, Les05, Zam02, BP03, BT00, CD05, CB08b, GRBB07, GES^{+}06, HB04, HS07, IOSPT^{+}04, Kim06, MFS07, NW05a, NLG05, RDD^{+}02, SH06b, Zam02]. sanctuaries [PPGL09]. Sanctuary [CH05, LLH06, KVH02]. Sand [HJ07, NGM^{+}05, ADR^{+}05, BWB^{+}06, BWP^{+}06, CRC02, DTB07, FHN^{+}09, LAV04, PKR07, SW08a, SHJC08, SJ04, TC03a, Tay03, Tay05]. sand-dwelling [SJ04]. Sandeel [AR02b, FWH^{+}05, FT00, HGG^{+}05, MCW^{+}07, RWH00, vDvHT^{+}09]. sandeels [ARP02, EvdKB^{+}08, SAH09, WWHE04]. sandflat [KKI^{+}06, NCT^{+}01, YTK^{+}05]. sandflats [CTH^{+}03, NTH^{+}02]. sandlance [Zam03a]. sands [WRJ^{+}05, WWH05]. sandy [AR00, BVVM02, Bra01, CHA06, DGL01, DC02, DG05, DM05b, FDDS09, GY00, HCA06, ILL06, KW05, She06, SDW07a, SGHG07, TK04, VK02, YSS03]. sandy-shelf [KW05]. sanguineus [BG07b, MCE01a, O’C07]. Santa [ABWK06, CLO^{+}09, NW02]. sapidus [BPLD^{+}05, CHL^{+}05, DRDM09, DES^{+}00, FCI^{+}04, HCM01, JSMB05, JSH01, KWWH02, KH05, LS02, ODF09, OvM02, SMHC03, TW00, TE00, vMRO03]. saplings [ITN09]. Sarasota [FFN^{+}08]. Sarcocornia [RGCL^{+}07]. Sardina [COPB06, CG04, GMZvdL07, GBHO^{+}08, GCVR01, GUT04, RGL^{+}00, RGMV04, SMVF03]. sardine [ABF07, CES^{+}03, CG04, GBHO^{+}08, GCVR01, GUT04, KBB09, RVMI^{+}03, RGMV04, SMVF03, TKO08, WHM^{+}01]. Sardinella [SZGS08, SSR^{+}09]. Sardinops [MBG^{+}08, OKAM00, RVMI^{+}03, WHM^{+}01]. Sargasso [MM07b, SBL^{+}00, SS02a]. Sargassum [RTH06, ABB05, BS04a, IOTL09, IAT^{+}07, SEH07b, SPT^{+}00, SDM06, TJA08]. Sargassum-associated [RTH06]. sargus [GWPRC^{+}07, LP03]. sarmentosa [RCR03]. sarsi [DSS04]. Sarsia [MR07a, MR07b, MR07c]. Satellite [CG05, HBG^{+}03, MBH07, SW09, ZAHJ^{+}06, GBHO^{+}08, GLÅ^{+}03, HPW^{+}00, HTM^{+}02, JMJ07, KJRG09, TML03]. Satellite-derived [SW09, GBHO^{+}08]. Satellite-monitored [ZAHJ^{+}06]. satiation [AP00a, Pea03, TJEM02]. saturated [BMSH02, RSR02]. saturating [EJKN03]. saturation [Tak02, TTM^{+}05]. Saudi [RSNG09]. saved [MCE^{+}01b]. savignyi [McG02]. saxatilis [EM05, NH03a, NH03b, PJ06a, PQRAC08, SP01c, WGM02]. saxitoxin [GHD^{+}02]. scabbard [KRP^{+}06]. scabra [UB01]. scalariformis [SR04b]. Scale [ABS^{+}09, BK04, BB03b, DMA02, FE02, GP03, She06, AFHS09, ARGS01, AP04a, AP04b, BCL03, BP08, BSZ05, BCPB^{+}01, BCJ05, BM05a, BUCC09, BOZ03, BB08b, CMD07, CÅJ05, CMLG07, CFA07, CTB07, CB08c, DM05b, DB07a, DBS05, DSDS04, DOML^{+}07, EJM01, ESS09, Edm07, EE06, EE00, GBdS^{+}05, GLB^{+}04, GYD04, GAM02, GCL04, GDL^{+}01, GCM01, HHF^{+}07, HSGS00, HFT^{+}09, HHWT02, HLTT09, HGKK06, HLH^{+}08, HCMR09, IP01, JN09, JAA^{+}01, JCD^{+}05, JFD09, JWR05, JKES07, KWP07, KPT01, KOC^{+}07, KvdMDP09, LNV^{+}05, LBvdH03, LTG^{+}01, LGB03b, Law08, LAK^{+}08, LH02, LYMCG06, MB07b, McI03, MdL06, MDSP05, MASFS00, MJKH04, MHHS06, MGO^{+}07, NRvdV^{+}02, NJPI03, Nod04, OS01, Osb07, PdNS06b, PPB^{+}09a, POH08, PR01, PM03, Raf06, RMA^{+}02, RBB^{+}08, RWB^{+}09, RVMI^{+}03, RFN02, RU02, RZ00, RMAY04, SHNV09]. scale [SGL02a, Seu05, SSM05, ST08, SRS05, SC07, SAV09b, SAV^{+}09c, SGH^{+}06, SDW07b, SMCE07, TFSP09, TPBF09, TML03, THN^{+}03, THHY05, THJK06, TTR05b, Und04, VCA^{+}09, VBA06, Vli05, WKL^{+}03, WGD05, WM02a, WMS03b, WTS09, Win09, WRW03, WG07, WWWC02, dBMSL02]. scale-dependence [MHHS06]. Scale-dependent [BK04, BB03b, DMA02, FE02, GP03, She06, GDL^{+}01]. scaled [BYH^{+}06, OT08, TH06]. Scales [BCBAC01, CU08, TAFBC07, AA04, BF08a, BCC^{+}03, BC01a, BCBVM05, BWP^{+}06, BUA02, BK06c, BHZW08, CC03a, CC07, DJM07, DWL09, EM04, FTBC05, GY00, HTL^{+}02, HSCF06, JÅC^{+}01, Jos09, LJ08, Lin04, LG08b, LBHL^{+}09, MTH^{+}08, MS04a, Nae06, NE01, NZM08a, NMDS00, OP09, OMP^{+}05, PKM04, PBTB08, SC00a, STA^{+}05, VCSL^{+}09, VL00, WHB^{+}06, WFC03, Zam03b]. Scaling [GBR01, MGC03, PKM03, SPD03, APH^{+}04, BSS05, CPK00, CK04, CRK06, DSH03, FIS04, FRvdMF02, Kne03, MMTT00, PHP^{+}03b, Sha00, SL05c, SCAF00, dlSPLV09]. scaling-up [APH^{+}04]. scallop [BCBQ01, BG04c, GMD08, Har06b, HK04a, JBSB01, MJS09, MB07b, MBH^{+}09, PW08, PW09, PBH^{+}06, SSO^{+}01, SZ09, SCM09b, SH06c, SHMN07, TK01, TCS^{+}09, ZYZ^{+}06]. scalloped [DH06]. scallops [BSVM^{+}05, HD07a, HRDA05, KMHS04, LCHG03, RL05a, TNTH04, WBDR06]. Scapteriscus [MLGG03]. scarce [RB08]. Scarcity [Som00]. scarring [UH03]. Scarus [BB08a]. scattering [FMZvW02, KTK^{+}08a, MBBW08]. scavenger [FTDK07, RH04]. scavengers [BD07, BWMA02, CFG06, FEO07, KH07a, KBPB07, PKA06]. scavenging [FEO07, ITN^{+}06, KBP06, Pet02b]. scenarios [FCS^{+}09, PGPL03]. scene [Kin03]. scent [FSC05]. schauinslandi [BPH07]. schedules [KGF^{+}09]. Schelde [SHM^{+}01, SVG^{+}05]. Scheldt [KvKB06]. Schizoporella [CFMB00]. schlosseri [PDM^{+}03, SBSRW02]. Schmaljohann [Ano01e]. schomburgkii [CHG05]. school [ANF^{+}00, LGME06, MSK^{+}00b]. Schooling [BGL^{+}07, BBG^{+}06]. schools [KBB09, RFG04, VPG04, Zam03a]. sciaenids [HH00b]. Sciaenops [FOFG09, HHH02, HH00b, PDHH06, SML02b]. science [BHC^{+}00, BKB^{+}04, BHZW08, Cor09, MNT09, SN09, SRJ07a]. sciences [Dis07]. scientific [Cor09, Kin03]. scientists [Lei08]. scintillans [DCAL^{+}02, HMA04, Kiø03]. Scirpus [KCP^{+}07]. Scleractinia [DBOR09, GAZ02, Lam00a, Lam00b, RZEACGF00]. Scleractinian [Dov04, IE06, BM02a, BG09, BR07, CSvO09, DFY00, DC07a, Edm07, FLK^{+}06, FPBAT02, Gou07, HR07, HTRFP04, LG06, LFRR^{+}07, MFK07, MHK^{+}01, MND^{+}02, MB00b, NTN08, RR05, RFPB^{+}04, SC06a, TTN01, UHR05, UHvO^{+}08, WH03]. scleractinians [FGO09]. Scomber [RCF07]. Scomberomorus [ST06b]. scombrus [RCF07]. Scopalina [BUCC09]. scope [BGPL08, LC03, STJ03]. Scopelocheirus [ITN^{+}06]. Scophthalmus [PMT^{+}02]. Scopimera [SW08a]. Scorpaenidae [ZK01]. Scotia [GJT09, HRK^{+}08, SM07b, SM08a, SG06b, BS00, FHM02, SCT^{+}09, SS05, TSW^{+}04, TSW07, WSA^{+}06]. Scotian [BCCM03, FF02, HKH04, KTF^{+}01, SG09, SF07b]. Scotland [BWMA02, MFG^{+}04, MGE^{+}09, NWHH^{+}09, SAH09, TWMW06]. Scottish [FDBT06, PHH^{+}05, YW04]. scour [BCS09, HS08b]. scouring [BFK^{+}09, SH06a, SBFP08, VBMBC08b]. scours [CK05]. screening [CH07b]. Scrippsiella [KH00b]. Scrobicularia [LBC07, Orv05]. sculpin [DL02]. sculpins [Zam03b]. Scutellastra [SSB02]. Scylla [PPJ^{+}05]. scyphomedusa [DBH^{+}07]. scyphomedusae [SDK^{+}08]. Scyphozoa [KUOI06]. scyphozoan [MR07c, Pur07, PHS09]. SE [DDL^{+}03, JH06, BF07, DTAF00, HGG^{+}03, HFH09, JCZ^{+}05, LVRN02, SH05, SYCG05, Tyn04, WWS^{+}05, RW01, Wan02a, XW02]. Sea [AY04a, AY04b, ABB05, BBD^{+}02, CM01, CD08, CNC00, CPSC01, DDL^{+}03, EBA^{+}03, FA08, FVS^{+}08, FR08b, GGC^{+}05, JKNJ07, KSSH02, KHGA02, KHH08, KKA^{+}06, LLSL06, MMTT00, ME01, MBB^{+}04, NBB^{+}04, OHS^{+}04, PFW^{+}02, PSKC04, PSRB00, SHA^{+}05, SBR09, SHMN07, TYT^{+}04, THBSJ05, WW07, AW03a, Ada01, AM02a, ALE^{+}09, AHW03, ABL08, AHK^{+}02, BS03, BMK07, BD07, BRI^{+}07, BDGS07, BCS09, BC01a, BKH^{+}09, BBM00, BRS07, BFB^{+}09, BPN^{+}09, BBS07, BOZ03, BHY09, BBM02b, BV08, CM02, CC02, CLP^{+}06, CAF^{+}08, CRLP09, CU06, CPB04a, CSM02, CWDG05, CW09, CL05, DCT^{+}03, DCT^{+}04, DBG^{+}09, DBJ^{+}00, DBF^{+}02, DFMD00, DSJBS^{+}02, DFJSJ^{+}05, DDGK^{+}08, DCA^{+}09, DKAH02, DS09b, DHR04b, DHR06, Eck02, ENAT07, EC07b, FGLM06, FTP^{+}09, FVAB00, FB00, FMZvW02, FFML02, FEMW07, GHV02, GWH03, GMVF08, GNL08, GTE^{+}09]. sea [GJG^{+}03, GHJ02, GH08, GBG00a, GP01, GSS09b, GSFF^{+}09, GKB09, GT08, GHdN07, GWBdN09, GBI^{+}06, HRB^{+}06, Har06b, HF03, HFW^{+}05, HGA07, Her06, HDDR08, HBBH07, HGK02, HPBT03, HBTD04, HK04b, HPB^{+}04, HZ04, IT06b, JSMM07, JT02, JPMGS08, KTH^{+}04, KCB06a, KSIG08, KTK^{+}02, KWHD08, KR01, KBP06, KEB07, KHB00b, KHK^{+}06, KSHO01, Kon00, KTF^{+}01, KGB^{+}01, KS04c, KG07, KB04, LCHG03, LB01, LLM^{+}08, LTFJ00, LTG^{+}01, LTMF01, LBF^{+}02, LFF^{+}09, LMS04, LGS07, LLJ05, LEB07, LSBJ08, LJ09, LCW06, LVA07, MSH08, MS09, MACS07, MJS07, MJS09, MRJ05, MC06, MAH^{+}08, MH08c, MW08, MS01a, MGF^{+}03, MSY02, MMC^{+}08b, MRG08, MNT09, MG00, MMB^{+}02, MSH^{+}03b, MAG07b, NNG03, NSSI06, NCN^{+}09, NOP02, NA04b, NHN^{+}06, NGMT01, OBB00, PB07, PTK^{+}09, PHB02, PHP^{+}03a]. sea [PCS09, PEG^{+}00, PKA06, RJTM08, RW03, RHG04, RW07, RSE01, REM^{+}06, RM00, RMGL07, RPC^{+}06, RGDI06, Rod03, RdNS02, RB05b, Ryp08, RKGH01, SM07a, SF07a, SSN^{+}09, SS09a, SGV^{+}02, SFF^{+}07, SKP^{+}07, SG05, SH06a, SJE^{+}06, Sew05, SS07, SSRG05, SHS^{+}05, SJS^{+}01, STV^{+}09, SPAMG00, SWM03, SC07, SFL^{+}09, SV01, STAS06, SBHE05, SH06c, SS05, SdNH^{+}06, SAM^{+}06, TND^{+}05, TTD^{+}06, TBB^{+}04, TSC^{+}07, TPBF09, TC01b, TCS^{+}09, THJK06, TWMW06, THYBS09, TRT04, TÁCTR06, TSW^{+}09, TBC^{+}00, TLP^{+}03, TBSJ^{+}04, TSJH05, TYC00, TYD07, VHKW00, VJ05, VT00, Van02, VK05, VLW09, VBMBC08a, VTY06, VACT^{+}08, VHV^{+}04, WP00b, Wal07, WH08, WAN^{+}01, WGL03, Win09, WTR02, WH09, WASW03, WG02c, WWS^{+}05, WS06, WBDR06, WDR^{+}05, YI07, ZHS^{+}05, AHSPG08, AQHTL02, ARAMPW02, AM03, AC07, ABD^{+}03b, AVS03]. Sea [Ano07g, AHW03, ARP02, AR02b, ABD^{+}02, AVE^{+}06, BHPC06, BBHT02, BAA07, BGLL09, BWPB04, Bar02b, Bar07, BBG^{+}04, BRIP00, BIR00, BIL01, BI04, BPVL03, BGM09, BDH01, BDEM01, BD03, BD06, BN04, BBC^{+}09b, BTB^{+}09, BSS05, BF08b, BSH02, BNP^{+}09, BCVC07, BIC^{+}03, BI03, CED^{+}07, CLO^{+}09, CMJ04, CH07b, CLP^{+}06, CAF^{+}08, CFG06, CCL03, CCL^{+}05, CS02b, CWBS03, CPSC01, CR08, CB05a, CBS^{+}08, CSPA09, CGH^{+}03, CGL^{+}02, CGC^{+}09, DLLB03, DSF^{+}04, Das02, Das03, DBH^{+}02, DMH01, DGBS08, DB07a, DSJBS^{+}02, DH07b, DM01b, DSDS04, DD09b, DLD^{+}00, EKN08, EBR^{+}02, EWBH04, EO04, EDBvdL08, EÖM08, EvBO^{+}08, FGPS04, FHM02, FC04, FSO08, FKA^{+}06, FGT^{+}07, FT00, GBdS^{+}05, GDG02, GJG^{+}03, GGC^{+}05, GMHA09, GSS09a, GC06a, GCSSJV07, GWPRC^{+}07, GMTN00, GRG^{+}02, GGG^{+}04, GP08, GF04]. Sea [GRB^{+}03, GHR^{+}07, GS02, GBC^{+}04, GS07, GGTKG09, HH01, HPW^{+}00, HHG^{+}07, HLD^{+}07, HHC00, HBT^{+}08, HDM04, HGAS^{+}01, Hel00, HR02b, HR02c, HMKT02, HVW^{+}07, HBN01, HSO04, HMD^{+}07, HRMH07, HLH04, HMO^{+}04, HCMR09, HW00, HCR01, HK03c, ITH^{+}03, ICH04, IRL^{+}09, JCZ^{+}05, JNS04, JCL09, JGH05, Kaa04, KLRS00, KFK^{+}07, KW05, KOTL07, KAPTK02, KWV^{+}05, KFT08, KK02, KFA^{+}04, KBL^{+}07, KS04a, KRVB05, KSHO01, KSSB04, KMR01, KSKK^{+}07, Kre01, KSWP04, LSH^{+}07a, LSG^{+}05a, LG08a, LSE06, LWW^{+}01, LGZ^{+}09, LCD^{+}05, LSM^{+}02, LZL05, LCB^{+}05, LSHL03, LHHB05, MHP^{+}00, MHP03, MHH^{+}05, MOM^{+}02, MM07b, MG09b, MFR01, MvdBÓ00, MSH^{+}03a, MPDS07, MKR07, MGHA05, MMA03b, MRB^{+}08, Mun07, NSJ02, NNW04, NN06, NB06a, NEK^{+}07, NM04, NWHH^{+}09, NRJB04, NLCS05, Ogu07, OSF08, OHN01, ORD00, OEDSJ02, OSP01, OP08]. Sea [OMP^{+}05, OLGG^{+}02, ÖBFO01, ÖCOB06, PGBA04, PKS00, PE09, PA06, PDG^{+}02, PBW^{+}07, RSK04, RBRH02, RHS01a, RCC01, RL03, Rin03, RGL06, RJBW03, RS00, RMRAFP08, RE00b, RZS05, RFLK03, RLDL02, RKJ01, SAJW04, SBL^{+}00, SDD^{+}07, SAW^{+}01, SNH04, SA03a, SGK04, SDW^{+}01, SS02a, SLM^{+}00, SRGS01, SYVC00, SO08, SP01a, SCT^{+}09, SNY^{+}05, SMA^{+}03, SSSV00, Som00, SBBS^{+}02, SP02, SAHK06, SHS06, SCD^{+}05, SHFP^{+}03, SBI07, SHT^{+}04b, SF05, SKDP02, SBC^{+}06, Str02, SSH^{+}00, SYH^{+}09, TRH^{+}06, TA07a, TKVL04, TSW^{+}04, TSW07, TPD^{+}05, TPS^{+}05, TLIL^{+}09, Tyn04, UB03, VGS^{+}01, VLP^{+}04, VAB09, VMVG^{+}07, VFV01, VV04, WLS^{+}09a, WWHE04, WHRS05, WSA^{+}06, WVL01, WHL00, WLZ05, Wie03, WRJ^{+}05, WR00b, WPTS08, XZUK05, YM06, YW04, YVB^{+}02, YMC05, ZTW08, ZCR^{+}09, dJTD07, vKH00, vOSF^{+}09, vdVBW00]. Sea [vdVBGW09]. sea-air [SFF^{+}07]. sea-basin [THJK06]. sea-cage [DSJBS^{+}02, DFJSJ^{+}05]. sea-cages [LFF^{+}09]. Sea-ice [GGC^{+}05, NGMT01, WAN^{+}01]. Sea-level [FR08b, TC01b]. Sea-surface [CPSC01, EC07b, GBG00a]. sea-winter [TWMW06]. Seabed [MRS00, SLP^{+}09, FRJ09]. Seabird [BHH09, DHHY09, EWG^{+}08, FT00, GCLC^{+}02, GZY06, HPS07, HVWH06, JLW04, SD01b, AS04, APM^{+}09, AO02, BB03b, BPB07, CRJ^{+}09, CRLP09, CLJ^{+}08, CMB^{+}05a, DPS^{+}03, EGJ07, GMHA09, GDGSP09, Hip08, Hip09, Hyr01, IMTW09, JRT^{+}06, JRS^{+}06, LCL^{+}03, MSJ^{+}09, MBG^{+}09, RHSN07, SSS^{+}09, SG05, SPD03, TS07, TBH^{+}09, Vli05, VHMB09, WHRS05, WLJM05, WRW03, ZPP09]. seabird-environment [TBH^{+}09]. Seabird-fishery [FT00]. Seabirds [FMW07, ISK07, PSS^{+}07, PHS^{+}07, ARBF02, BPM^{+}06, BGSBC02, BCL^{+}01, BHD04, Car05, DM03, DM05a, FE02, Fau09, FT00, FEO07, GHT^{+}01, GSS09b, GSS09c, GDR^{+}04b, GB09, HRB^{+}06, KFK^{+}07, KHIH08, KWP07, KRCGC06, KHF^{+}00, KPW07, LBR^{+}09, NBB06, NCN^{+}09, OAS01, ÖCOB06, PSC^{+}04, PSW07, QMF05, RW03, RHPB08, RWH00, SBA01, SBI07, TPSW00, VBF06, Vli05, WPM09, WGS^{+}02]. seaboard [RGMJM07]. seabream [LP03]. seaduck [SPD03]. seafloor [EP00b, GGDB09]. Seagrass [CHG05, FTP^{+}09, GBNÖ08, HMJ^{+}09, RH06, SBB09, SCBA05, APA07, AMS01, AMR00a, AMR01a, AM02a, ASS05, Bib07, BRAS03, BH09a, BB00, BVN^{+}03, BVP^{+}06, BH02, BD01, CAS05, CHLV07, CH09, CRS^{+}07, CPKV00, CMD^{+}09, CLMR07, CLRM08, CT06b, CT09, CRZ04, CV05, DGM02, DFV01, DGNvdV05, DGC^{+}05b, DVNvdV07, Duf06, EH00, EHK06, EFL^{+}09, EMD^{+}01, Enr05, EKBM05, EF02, FHDA^{+}07, GBMHD08, GSC^{+}01, GDKN08, HUP^{+}00, HBNA04, HHO03, HSBD08, HO02, HFM^{+}02, HF05, IOB04, JRC06, JH01, Jen05, JYS06, JH06, KECJ06, KDM05, KKIV01, KWTR08, KVH02, Lar03, LRHV07, LD00, LMD^{+}02, LSS^{+}09, MR05, MCL07, MD01, MHM^{+}02, MCVAR08, MRHP01, MSUK01, MS01b, MJKH03, MJKH04, MB05b, MGO^{+}07, NKK^{+}01, NRvdV^{+}02, NvdV04a, NvdV04b, NvdVVD06, NBN^{+}08, OHB^{+}00, OHT02, OKM06, PTLK04]. seagrass [PDHH06, PTCH01, PH01a, PH01b, PLMS04, PA06, PTAR07, PFT^{+}08, RP02, RB04, RR01, SESP03, SASR03, SKBP09, SZ08a, SZ08b, SPI^{+}07, SS03, SCZ^{+}06, SWL07, SHJC08, SYA02, SZBZ04, TTD^{+}06, TC03b, TTR05a, TTR05b, TÁCTR06, TJA08, UH03, USG^{+}08, VHB^{+}07, VLK04, VJ03, VHW^{+}07, VBAR07, VNdG^{+}06, VHM09, VPS08, VS08, WPG08, WKD^{+}04, WPB^{+}08b, WH09, WR00b, dlMPN^{+}03]. seagrass-associated [Jen05]. seagrass-coral [EFL^{+}09, VHB^{+}07]. seagrass-dominated [CV05, RB04]. seagrass-hardbottom [CT09]. Seagrass-pathogen [SCBA05]. seagrasses [CF09, DTAF00, HHV06, HKWM07, HBRW04, JCK03, KHT^{+}02, KDM05, KHB00b, LVRN02, LV02, PBPLB06, Reu01, SA03a, TN04, TN07, YBGP06, YLv^{+}05]. seahorse [CV05, THP^{+}05]. Seal [WBL^{+}06, BLD^{+}05, BCHS09, BPH07, BN09, BLDG01, DFW^{+}03, GLA06, GdLB^{+}09, GNHB^{+}05, HWB^{+}09, HBDD02, HBBD09, KW02b, LBvdH03, MGC^{+}09, MHBB05, NEK^{+}07, PMG05b, SCKC07, SZL^{+}05, WBS^{+}07, WLS^{+}09b, ZS04]. Seal-mounted [WBL^{+}06]. Seals [SAH09, BWW^{+}04, BPG08, BDG^{+}04, BKH^{+}09, BSM02, BMM^{+}04, CS02a, CDH^{+}09, CDF^{+}08, CCL^{+}06, DPD^{+}03a, DPD^{+}03b, FBBH07, FBD06, GAMB05, GBG00a, GHT^{+}01, GNAD^{+}03, GDL^{+}01, HARC05, HHWT02, IBB^{+}07, KHH08, KCH^{+}03, KKL07, LBBH08, LCGN02, LGC^{+}08, LAH07, LAG08, MFWD08, MWS^{+}04, MWM^{+}05, NFN^{+}09, NWIT08, PMG05a, PMS^{+}06, PHH^{+}05, SKE05, SRB04, TBBI04, TML03, TBH08, TBI07, TBI08, TBI^{+}09, VSB04, WHP00, WP03, WMS^{+}03a, WBS^{+}04, WHL00, dBTO^{+}09, TBBI04]. SEAMAP [HRB^{+}06]. Seamount [KRP^{+}06, KGHL^{+}01, DLB04, FMZvW02, Hum00, MVD^{+}08b, TPBF09]. seamount-scale [TPBF09]. seamounts [AWS^{+}09, BSH^{+}09, LBC^{+}09, LMB^{+}09, MMK^{+}08, STF^{+}09]. seaperch [ST07]. seas [ACLF09, ASR^{+}09, BIL^{+}02, CMLG07, HSTS^{+}09, HB08a, HB07a, MH08a, PPC09, Ran02, SZW^{+}07, DD09a, HJ03, KAM^{+}06]. Seascape [ABD^{+}02, FDM^{+}09, PCHM07, SWM03]. seascapes [RS06]. seashore [EP00b]. seashores [MBVBC09, VBMBC08a]. season [ACR07, BF08b, CTFC08, FFT^{+}07, GFB^{+}07, HPS07, HS01, HFA02, KCK^{+}06, Lap01b, MPE03, MH06b, PH03c, RHC^{+}08, SKE05, TFSJ00]. Seasonal [ABB05, BBBSPH03, BC02, CB00a, CMG04, CDT^{+}03, CIFN00, CGL^{+}02, DGBS08, FE07, Fis06, GXL^{+}08, GRS07, GSS09a, Gis05, HBH07, HST^{+}09, HB07b, IMTW09, JPMGS08, KKLL02, KHIH08, KKTN05, KH00b, KN00, KOY^{+}05, LH09, LCL^{+}03, LJ06, LTD^{+}02, LS06, LVA07, MFG^{+}04, MUT04, MFS07, ML06b, MHP07, MCM05, MASFS00, MMA03a, OA02, PGBA04, PAR08, RCC^{+}04, RHPB08, RVMI^{+}03, RT00b, Rol00, SSR^{+}03, SS06, SO01b, SHH08, TA07a, TRDV00, TTR05b, TNSM00a, TNSM00b, UHvO^{+}08, UB03, VMRL04, VK02, WA05, WS06, YYL^{+}00, vBWSS06, AS06a, ASA^{+}03, ASP^{+}03, ALE^{+}09, ACLF09, BIR00, BIL01, BDG^{+}04, BRI^{+}07, BCLS06, CBD04, CS09, CS02a, CCCH01, CWBR05, DHHY09, Dav00a, DAM^{+}08, DRG^{+}07, DNL08b, FR07, FBBH07, FFML02, FCP02b, GMHA09, GKB09, HARC05, Hip09, HCMR09, HW01, HW03, IRM08]. seasonal [JSMM07, JAA^{+}01, LRR^{+}09, LSMR05, MSF^{+}07, MHW06, MADPQ06, MYE03, MGREA07, McQ05, MRG08, MRB^{+}08, NLB08, NT06, RLBB04, RPO00, SDS^{+}03, SMH09, SPH^{+}09, STV^{+}09, SGH^{+}06, TLC06, WW06a, WIDM09, WDN00, YHSH02, Yin02, dBTO^{+}09]. Seasonality [Bow05, BP01, CY06, FDBT06, GBF^{+}02, KLC^{+}06, ZMN07]. Seasonally [LSH00, SMS06, VBV^{+}00]. seasons [MH09, PSKC04]. seastar [HRDA05, SBM02]. seastars [HPBT03, HBTD04, RJH02, VTY06]. seatrout [Kup03, LBHL^{+}09]. Seawalls [MCU06, BC06b, BCU04, Bul05, Cha03]. seaward [OYT^{+}02]. seawater [Edm04, GPM08, HTBT07, LCS^{+}01, LZW^{+}08, MAOP09, MBH^{+}09, NFV^{+}06, RL07, SV08, SMVF03, TSE^{+}08, TSC^{+}07, TA01b, VPSO01, VPWE09, WAM^{+}08, WDM^{+}09, YYL^{+}00]. seawaters [SAB^{+}05]. Seaweed [NCP^{+}08, PdNS06b, BR03, Bat09, BMdNP09, BCB02, EHK06, GW06, HDQ04, JHH01, LH02, LT07, MUT04, MKKK08, NGdN^{+}07, OST07, PCL^{+}09, TP00a, VÅ01, WS04b, Wol07b, WS00]. seaweed-epiphyte-grazer [WS00]. Seaweed-herbivore [PdNS06b]. seaweed-invertebrate [Bat09]. seaweeds [CU06, CH00a, KG00a, KG00b, LBPS03, MD07, PH03c, Ree03, SG06b]. Sebastes [BRG^{+}02, HKR^{+}08, JKK^{+}06, SBF08]. Sebastolobus [RM09a]. Secondary [JYS06, APP^{+}09, BMdNP09, BP04, CM07, GMA^{+}09, GP01, Kev01, KW00b, LCT^{+}08, MJG^{+}00, MJKH04, PCT00, RGS^{+}09, SR01]. secosterols [EMPF07]. secretion [BB05]. secretions [HSD03]. Section [Rii03, Ano03h, BHC^{+}00, BS05b, BVHG08b, GBB^{+}01, GSS09c, HR02c, Mar01, PAC^{+}00, RQF^{+}00, Rii04b, SvAHH09b, SvAHH09a, SAV^{+}09c, SB05b, Syd09, WB05b]. sector [DÁSG^{+}01, WAN^{+}01]. sectors [FBD^{+}02]. sedentary [SM08b]. Sediment [BHY09, CRTP02, RWPJ04, TDUN07, TGM^{+}03, ACT09, AHL^{+}01, AH07, AFFH04, BD11, BWP^{+}06, BCM^{+}09, BFK^{+}09, CMD07, CVK00, CWR06, CIFN00, CTvdM^{+}08, CTD^{+}09, CK05, CTH^{+}03, CT04, CVT09, DSS04, DR01, EKA03, Ell01, Ell02, EMD^{+}01, FSV^{+}08, FGS^{+}02, GSM05, GBMHD08, GGHA02, GPB02, GKK03, HTJ00, HUP^{+}00, HDC^{+}08, HF06, HGG^{+}05, HO02, HCH09, HCJ^{+}07, HAA00, HJG06, Hun05, HMFC07, HGMH06, KECJ06, KW05, KNTK05, KKKK03, KTSR02, KHD07, KL03, KGTH01, KWTR08, KSS^{+}05, KM03b, KGRW06, KVV^{+}00, KKI^{+}06, LS08a, LS09a, LT00, LTH^{+}04, LHT06, LTL^{+}06b, MBW02, MAE03, MDR^{+}06, MDBG07, MD01, MGD^{+}09, MT00, MF03, MSD05, MC08a, MVD^{+}08a, MK02, NR00, NRH^{+}07, OWBD07, OKP06, Orv05, ODG^{+}04, PTLK04, PSRB00, PZA09, PDFD09, QS05, RSL09a, RS03, RF08a, RBRH02]. sediment [RWBS03, SN06, SR08a, SABK03, SE04, SS04b, SK02, SWH^{+}04b, SGHG07, SCD^{+}09, SRS04, SCC06, Tal01, TCS^{+}05, TSSQ06, THF^{+}01a, THF^{+}01b, TSHT02, THN^{+}03, TA03, TPS^{+}05, TRPNE06, Uth01, Uth06, VMV^{+}09, VFJ06, VBMBC08b, VVTH07, VPWE09, WE04, WAA^{+}09, WKOE09, WDM^{+}09, WBSL00, WBH^{+}04, WBP^{+}06, WPB08a, WPB^{+}08b, WBP09, YYK^{+}08, YH02, vBSSH02]. sediment-biofilms [Uth06]. Sediment-biota [RWPJ04]. sediment-bound [AHL^{+}01]. sediment-dwelling [TCS^{+}05]. sediment-feeders [Uth01]. sediment-profile [OKP06]. Sediment-water [TDUN07, HAA00, MBW02]. sedimentary [BEE06, GMd^{+}02, JCK03, MRB^{+}08, QASA08, RS03, VK02, WYFX02]. Sedimentation [BJR04, CWMD06, IC02, JPNM^{+}06, AW03a, AFFH04, BPE^{+}06, Con05, CGL^{+}02, DD04, ENC^{+}04, Gil02b, HRR^{+}01, WKPK05]. sedimenting [MGA^{+}01]. Sediments [SGP07, AHC^{+}06, AGCF^{+}08, AJ03, BRJ^{+}03, BML^{+}08, BWSH00, BVI^{+}05, BPN^{+}09, BL01, Bur01, CB00a, CBA02, CvBGR09, CM05, DDF01, DMH01, DP01, DRG^{+}07, DFMD00, DTNU00, EWBH04, EHM^{+}09, FDRR05, GMVF08, GKB09, GHS05, Gra02, GGM^{+}00, HH03a, HJ03, HBT^{+}08, HSPH05, HHT^{+}09, HGK02, HCA06, HM05c, HBK^{+}05, JGH05, JFR02, KHGA02, KWV^{+}05, KMR01, KFU^{+}08, MRLK06, MGM05, MSD05, MSUK01, MKH^{+}02, MHK^{+}04, MLR02, OSS^{+}04, PS06, PCBL05, PBW^{+}07, QS08, RDG^{+}07, SP03a, SFH^{+}00a, SFHS00, SDW^{+}01, SV01, SHT^{+}04b, SVG^{+}05, SZL^{+}05, SMG00, SCB00, TKR06, TR05, TNSM00b, TJA08, UYNO01, VMdBV06, VNJ06, Wal07, WBW^{+}06, WTH04, WRH05, WHCP^{+}04, YYK^{+}08, YLL^{+}04, dBS01, vBWSS06, vOSF^{+}09]. See [BBG^{+}08]. Seed [BGLL09, OFHM03, EvBO^{+}08, LPK^{+}07, OHB^{+}00, RHBW09, ZCR^{+}09]. Seed-density [OFHM03]. seeding [Kre01, MC01a]. seedling [OFHM03, RHBW09]. seedlings [BGLL09, BN01, BFK^{+}09, WKD^{+}04]. seeds [OHT02, OKM06]. seeking [CC03b, SHK05]. seep [BUF02, BFK^{+}05b, CY06, CHP^{+}05, FKM^{+}01, HGC^{+}07, ISSK03, KFH04, MFCM05, YFH^{+}02]. seepage [GSW01b, GB03]. seeps [LJM^{+}00, LZM^{+}03]. segregate [BGH06]. Segregation [MB04, NvdVVD06, BB09, BCH00, DGNvdV05, FHB^{+}02, FGSH^{+}05, GSSR02, HLH^{+}07, KCA^{+}08, LBS^{+}05, LGC^{+}08, LRGdS00, PCRR08, SKBP09, SLM07, TBI07, WISH09]. Seine [SAT^{+}03, CPG08, SMVF03]. select [SSLM01]. selected [BMV02, KP01a, WKK07]. selecting [MJG08]. Selection [AP00a, CBMB03, Gla07, MCVAR08, ARGS01, AP04a, AP04b, BBM03, BTL05, BB03b, BDR04, BDG^{+}07, BP09, BV08, CS03b, DMA03, DSMBG03, DVBLS07, EDE05, EM07a, FM02, FGT^{+}07, FKI^{+}09, GDF03, HDRC^{+}08, KHA^{+}00, KLW^{+}09, LSH07b, MSLP06, MH06a, NPLZ00, PS08, PKLD09, PCW05, PS01, RCCK09, RCF07, RGVC07, SBC04, SH07a, SCIT09, ST06b, SDK^{+}02, SIK^{+}02, TCD01, TSSQ06, TOBR07, TL03, UNII01, VBF06, WLM06, vMRO03]. Selective [FS08, FE06, LB02, RL05b, SPGS03, vdVBW00, ABY00, BWMA02, BMSD00, BS07c, Enb05, HM06c, LOB03, LC08, LPBD00, MVH^{+}06, ST06b, TAM03, TOK^{+}04, TAO07, TSF04]. selectively [KLHO02]. selectivity [ASC06, BTB^{+}09, CBS^{+}08, CRH01, HTS06, JBS01, LGME06, MB07a, MHV^{+}00, OLWB06, SH05, SHSF08, VFV01, WEEBdN08]. selenite [BF01a]. selenium [GSSR02, SLL02]. Self [MC05a, TS05, APM^{+}01, Fis05, LC07a, NGBG07]. self-compatibility [LC07a]. self-facilitation [APM^{+}01]. Self-fertilization [MC05a]. self-recruitment [Fis05]. Self-replacement [TS05]. self-thinning [NGBG07]. selfing [BHFC08]. semi [CSL^{+}05, CM08, HFA09, JC08, JN09, dZCVC01]. semi-arid [CM08]. semi-enclosed [CSL^{+}05, JN09]. semi-isolated [HFA09]. semi-lunar [JC08]. semi-static [dZCVC01]. Semibalanus [BBT02, BG07b, DBT02b, FKSW09, JÅC^{+}01, Nod04, PSS06, SJH^{+}04, VBB06]. Semichi [Kon00]. Semicossyphus [TLC06]. semifasciata [CS09]. Semimytilus [CC04]. semisuberites [QASA08]. semisulcatus [JRP01]. senescence [MLS^{+}08]. senhousia [Mis04]. Senri [HGS^{+}02]. sensation [HMR00]. sensed [WR06]. Sensing [JBS^{+}09, BPS^{+}03, BYH^{+}06, DGVK^{+}04, FFYS^{+}07, GPLAG04, JCMR^{+}03, MTCU08, OT08, PR05, SCD^{+}05, WGS^{+}02, vdWHF^{+}08]. sensitisation [BSW^{+}09]. sensitive [FT00, TCS^{+}05, WLL03]. Sensitivity [RZWG08, BK06a, DL03b, EM05, FT00, HPS07, HRC^{+}07, JS01, KR01, SK07, SW08b, UBRvO06, WB06, WS04a]. sensors [VCSL^{+}09]. Sensory [WB05a, WB05b, DK03, JdSM09]. Separating [PPCU07, Bur01, EG05, SRJ07b]. Separation [ZRG00, Ano01a, BSIS07, OS01, PMS^{+}06, PK00, TH02]. Sepia [BTF^{+}02, BTD^{+}04, GPM08]. Sepioteuthis [GMS02, PTSJ06, SM00a, SMG02, SPM03, TA01a]. septemspinosa [TC03a, Tay03, Tay05]. sequence [CMRB08, HFM^{+}04, PG01]. sequences [GRPE08, HCB02, LK02, PAW00a, TR06]. sequential [CS02a]. sequestered [TCS^{+}05]. sequestering [SHT^{+}04b]. sequestration [GP01, GM08, Koe05]. Ser [tAHSHA07]. sergeant [EM05]. Seriatopora [LLCHG01, NL05]. sericea [SRS05]. Series [Kin05b, BP09, GPM^{+}09, HBN01, KKP08, MS05, RFSC08, Sha09b, STF^{+}09]. Serpula [Iye02]. serpulid [MHW07]. Serpulidae [TP01a, TP01b]. Serranidae [CSS00, KAR04, Tup07]. serrata [PPJ^{+}05]. serratus [Arr05, BSJ02, ID02, MNC^{+}02]. serum [CCCH01]. services [PKM03]. sessile [Ark09, Bow05, BCPB06, BTF^{+}06, CGT^{+}00, GSM^{+}08, Gla01, JKQ02, Sta08, TBCB^{+}05, TAFBC07, VBMBC08a]. Seston [SRGS01, ABY00, BCB^{+}06, CRGH01, Gar00, KJL03, Mar04, SSSV00, TRDV00, VN02, Wan02a]. setae [PL00]. setation [MCM05]. setiferus [WK02]. Seto [AY04a, AY04b, MMTT00, OHS^{+}04, TYT^{+}04]. setosa [TT05b]. setosus [BFE^{+}03, FTB09]. sets [FMW07, SMF00]. setting [DMM^{+}03]. settings [GSRH02, HAP^{+}08]. settled [GCS06, HM06c, JPC00, MPM^{+}02, PS08, SML02a, SML02b, SCAF03]. Settlement [FCI^{+}04, GD07, LB01, LGB03b, LWJC04, MHR03, PTB03, SC00b, TRT04, TP01a, TP01b, TYD07, WMG02, AQSP07, ACH07, BS00, BJLH04, BMWH08, BP09, BBSO03, Boo02, BK06a, BNP07, CB08a, CVT09, DQ05, DMM^{+}03, DRTM09, EM05, EWSH08, EE00, EEWBdN08, FCP^{+}07, FA02, FCP02a, GKS^{+}07, Gim09, GLJ07, HS08a, HW09, HLQ02, HNYI07, HCM01, HHH02, HH03b, HTWQ05, JC08, JPC00, Jef03, JH01, JQ05, KSR^{+}07, KCSPT09, KZ00, Kru01, Kuf01, LTPL05, LNV^{+}05, LHQ03, LHQ05a, LHQ05b, LMCQ02, Lec05, LP04, MPM^{+}03, MHO07, MK03a, MDW^{+}09, MVK^{+}09, MW01, MB00b, NNLV06, NKS03, OTBR00, PPV^{+}06, PDQ01, PS06, PSS06, PMR06, RM05, RP06, Ros01, SM07b, Sha09a, Sha09b, SF02a, Ste04b, SCM09b, SdNH^{+}06, Tay03, Tay05, TT07a, Trn02, TCT^{+}00, VFM06, WW06a, Wal07, WEEBdN08]. settlement-stage [Trn02]. settler [MK03a]. settling [BWO04, GG09, HBT^{+}08, HLH04, JKNJ07, PTD05, Wen00]. seven [MF08, ULL03, WS04a]. sevengill [Bra08, LME05]. several [WLdG^{+}04]. severe [Str02]. severity [GR07a]. Sewage [BUA02, ABU01, BMF06, GLS02, GTFB03, Kev01, RSNG09, RJ03, SEL02, VFP^{+}07]. sewage-derived [SEL02]. Sex [CLP^{+}06, HT07, MAVO^{+}06, RHG04, SM01b, Tar03, BPM^{+}06, BRI^{+}07, BGSBC02, BGRD00, CC03b, CAMC07, CKB^{+}09, CMGBJ05, DFH05, FHB^{+}02, FGSH^{+}05, GBM01, GBGH02, HJH^{+}02, JSIT04, KCK^{+}06, KKL07, MMC^{+}08b, MdMR04, OAM02, PPJ^{+}05, SFH00b, TJSG03, VSB04, WM04, WST^{+}09, ZBDE07, Zup00]. sex-biased [PPJ^{+}05]. Sex-dependent [HT07, Tar03]. Sex-specific [MAVO^{+}06, SM01b, BPM^{+}06, CKB^{+}09, WST^{+}09]. sexdentatus [BM05b]. Sexual [GAZ02, Lam00b, ASS05, DDD08, FVAB00, GBF^{+}02, GDBF07, McG02, MBW07, PMG05b, QMM^{+}08, SCMB03, SBM02]. Seychelles [CRJ^{+}09]. shad [MC01b, PDM^{+}04]. shade [CEF04, RDC06, VNdG^{+}06]. shaded [BC06b]. shading [BVN^{+}03, GBR01, LMMT09, RR01, SCC06]. shag [WDT^{+}08b]. shags [MAVO^{+}06, VML05]. Shallow [BS07a, KTT04, MC08b, AI02, ACM^{+}02, AG02b, ABB05, BS05a, BS06, BD08, BTF^{+}06, BLP^{+}02, CCMK03, CIFN00, CAS00, CD07b, CRZ04, CAA00, DF04, DQW05, DdNEvdZ05, EDS03, FBT^{+}01, GRWB02, GGR06, HvW08, ITN^{+}06, JHJ^{+}03, JHJS04, KRCGC06, KFH04, KMR01, LS08a, LDL^{+}04, MSA07, MD08b, MH07, MLW03, NDV^{+}00a, NDV^{+}00b, NvdV02, NGBG07, OMP^{+}05, PGCC^{+}06, PB07, RGCC^{+}08, Roy02, RZ00, RW06b, SAG08, SSGB^{+}08, SPBM05, She01, SBFP08, SLCC09, SCS^{+}07, SMG00, SMH^{+}03, TFA^{+}08, TGM^{+}03, TH06, Zam07]. shallow-diving [KRCGC06]. shallow-dwelling [Zam07]. shallow-subtidal [CD07b]. Shallow-water [BS07a, DF04, DdNEvdZ05, GRWB02, HvW08, ITN^{+}06, KFH04, LS08a, MH07, MLW03, NDV^{+}00a, NDV^{+}00b, NvdV02, NGBG07, PB07, RW06b, SLCC09, SMG00]. shape [GM04, Lir00, SS09c, SMH02, SB03b]. shaped [BRSM07]. shaping [DP01, RCL^{+}09]. shared [Bay02, DNL08a]. Shark [HCS07b, BBC04, CJFM09, DMK^{+}08, HCD^{+}09, LME05, LGME06, NMOO03, NWJ^{+}08, PWLC06, Rob06, WFG^{+}08, WH09, NJPI03]. sharks [AG02a, AFPS08, Bra08, CS09, CPBS07, DS09a, DNL08a, DNL08b, DH06, HMBD01, HFW^{+}05, HGKJ01, LKDO07, MSF05, MBP^{+}06, MCP^{+}09, SAS^{+}06, Sim00, SSR^{+}03, WOL^{+}07a, WGR07]. Sharp [GdLB^{+}09]. sharpnose [Bra08]. shear [CS01, LJK02, PCS04b, RMPC08]. shearwater [AO02, CRLP09, CKS05, GDGSP09, GSFF^{+}09, LIM^{+}06, NGSV07, PSKC04, PC05]. shearwaters [BLH01, BHPC06, JCZ^{+}05, KS00a, MSH08, NGS09, SWS^{+}09, WRW06]. Sheehy [HSJ08]. sheephead [TLC06]. Shelf [HKH04, JPMGS08, PHO^{+}07, SHH09, SF07b, AI04, AWA^{+}05, ACLF09, AL09, BN09, BVCG01, BOZ03, BHD04, CWC00, CC05, CBM^{+}07, CAA07, Ell01, Ell02, FAPA^{+}09, FAPOF09, FHQ^{+}06, GBdS^{+}05, GL00b, GL00a, GRP^{+}06, GF04, GMd^{+}02, HCGC09, HB08a, HGT01, KH07a, KW05, KHP09, Kor01, LJM^{+}00, LG02, LF06, LAH07, MSD05, MSH^{+}03b, MFM^{+}09b, MBB^{+}04, NBB^{+}04, NLHO09, OÁSG^{+}09, PTM08, PTS00, RHS01a, RMGL07, ST06a, SB05a, SS09b, SFL^{+}09, SHT^{+}04b, SC00b, SOMM06, SPGS03, SCAF00, SCAF03, Thi03, Tyn04, WHD^{+}05, WBL^{+}06, WHB09, BCCM03, CMJ04, FF02, GWWC00, JBG00, KTF^{+}01, JKNJ07]. shelf-break [CWC00]. shelf-sea [BOZ03]. shell [AE06, BBST07, CGT^{+}00, Fer04, Hun04, LABR^{+}06, MTS07, RB04, SB03b, TL03, TDL05, ABBF01, BBST07, HSCD01]. Shell-boring [BBST07]. shell-breaking [TDL05]. shell-species [TL03]. shelled [ARB04, PTP01]. shellfish [BCBQ01, CCM01, DCM^{+}06, GSRH02, GFC^{+}02, JF02, KS06, KSG03, MFG^{+}00, SSO^{+}01, SVC^{+}01]. shellfishery [WGCM^{+}03]. shells [Bar03, RCC01, SMT08]. Shelter [JKQ02, PJ06a, TVD02]. sheltered [RCDC06, TMJ^{+}05]. shelters [BFLÁ01, LÁMBF09]. shelves [EHWS07, MGHA05, VCA^{+}09]. Shetland [JHB06, STE04a, BMJ02, KCV^{+}06]. Shetlands [PSPZ09]. Shift [YQW^{+}01, Ano07g, BGK09, DH09a, IT06b, KTNK07, KFS^{+}09, PBF08, RM06, WH09]. Shifting [Bai00, Rui04, SP04]. shifts [ADS^{+}09, CC05, FTO^{+}09, HGA07, HD07b, JWM09, KH07b, KCSPT09, LGSB09, LCGN02, LBPH06, LÁMBF09, MS06, MARM06, NDV^{+}00a, NEMF09, NNLF09, TOK08, VHV^{+}04, WLZ05, WDSA09]. Ship [HBN01]. Shipboard [HLH^{+}05, TMD^{+}07]. ships [DBvOM06, MBN^{+}09]. shipwreck [PHM^{+}08]. Shiraho [HKY^{+}02, KHIA02]. shoalgrass [HVP^{+}06, TS06]. shoaling [GB05a]. shoals [JS08a, SH00]. shock [BCKP09, HMH04, RPS^{+}04]. shoot [BB00, BFK^{+}09, LPK^{+}07, PvDMB08, SH06b]. shoot-density [BB00]. shoots [EvBO^{+}08, PCL08]. Shore [RC09, ANCC07, AR00, AAF^{+}04, BCO07, BCB02, BVMBC07, BMJ02, BHR08, CLC04, DHCH04, DD05, DCF^{+}05, GD07, GKB09, HW01, HW03, JM01, JMA02, KK02, LTMDB02, LÅCN01, LBCS05, MB05a, MUT03, MCE01a, MHTJ07, MHR03, Mok04, MSP^{+}06, NW06, O’C07, PH03c, Prz05, QMM02, SP01a, SvNvdMvdV09, SDW07b, SRLA04, TYT^{+}04, Vie09, WW06a, WDN00, WDLM05, YSS03]. Shorebird [KB05, SWGC^{+}05]. shorefishes [SMMM07]. shoreline [DS06a, DL03b, PMGE01]. shorelines [FS08]. shores [BCPB^{+}01, BCBAC01, BC01a, BUA02, BFB^{+}09, BCU04, Bul05, CMD^{+}06, CD07b, FTBC05, Gar00, Har08, HSM^{+}09, HW03, Kad03, KM03a, LÅCN01, Rue07, SMC09, SDW07b, TMJ^{+}05, VBMBC08b, WBP^{+}06, Wie05]. Short [FGM^{+}00, GB01, KSS^{+}05, KS00a, MG09a, ME01, SM00b, SMSD07, SIK^{+}02, VML05, WK00, WS08, WPTS07, dBS01, BLH01, BHPC06, BBM^{+}02a, DSJBS^{+}02, DBS05, GÁSPR03, GBG^{+}00b, JM07, Joh02, KLR^{+}06, MO05, MPS^{+}08, MASFS00, MMB^{+}02, MK02, PB07, RBSP00, SdSC01, SH06c, TAM04, WRW06, WdBS02]. short-beaked [MPS^{+}08]. Short-circuit [SM00b]. Short-tailed [KS00a, BLH01, BHPC06, WRW06]. Short-term [FGM^{+}00, GB01, KSS^{+}05, MG09a, ME01, SMSD07, SIK^{+}02, VML05, WK00, WPTS07, dBS01, DSJBS^{+}02, GÁSPR03, GBG^{+}00b, Joh02, KLR^{+}06, MO05, MMB^{+}02, MK02, PB07, RBSP00, SdSC01, SH06c, TAM04, WdBS02]. short-time-scale [MASFS00]. shortbill [HLH^{+}05]. Shortened [BM09]. shortspine [RM09a]. should [BGK09]. show [RSE01, TPBF09]. shown [GFN^{+}09, LCGN02, NS02, WDT^{+}08b]. shows [CHW06]. shrimp [BRP00, BLF06, BL01, CGM00, CSO08, CCdAR00, CY06, CS03b, CC05, CWBR05, CBM^{+}07, GDL04, Gri09, GWF07, HLVU00, HM03, KPN03, KWKF03, KS00b, LMC08, MRP09, OHN01, OSL07, PGBA04, PABM08, PGS^{+}00, RGR^{+}01, RZLS09, RMKR00, RRRM08, SW04, TC03a, Tay03, Tay05, WK02, WE04, ZLGC04]. shrimp-trawling [HLVU00]. shrimps [AND05, ACAG07, CPA^{+}03, FC04, PCD04, RCC^{+}09, VV04]. shy [HGB01]. Si [SCGC07, SHSV04]. sibirica [GNHB^{+}05]. sibling [BBRJ04, Kru01, LGFM00, LGVF00, LW04, SDR01]. sibling-dominated [LW04]. Sicily [BCVC07]. Sicyopterus [ST02]. side [DMM^{+}03, OHCFJ04]. Siderastrea [SKR08]. siderea [SKR08]. Siderophores [SDRH01]. sides [THBSJ05]. Sieve [GHV02]. Siganidae [FSHB09]. sightings [Eck02]. sign [HMKT02]. signal [JP08, NBB06, SAS^{+}06]. signalling [RB05b]. Signals [DM03, Car05, DM05a, HCCY02, Jak01, Vis01, WCZ^{+}09]. signature [NWIT08, TBI^{+}09]. signatures [BS07b, BPT^{+}00, CH07c, CWM^{+}09, DPSD08a, DPSD08b, FCS^{+}09, GLS02, HTBT07, KP01a, KKK^{+}05, KCV^{+}06, KFK^{+}07, LZW^{+}08, PKM04, PDFD09, SSS^{+}05, SSW08, SMCE07, SMHG09, TFA^{+}08, TPSW00, WISH09, WHS^{+}09, WIB08]. Significance [AO02, HS05, MC01a, SRGS01, CLP08, DGBS08, JBG00, LNV^{+}03, LZL05, MRPL07, MBM^{+}00, NSJ02, RM00, RFSC08, STRD02, ST06b, WLL03, WL03]. significant [CRS^{+}07, DSP00, MH08a, SS07]. significantly [MH08c]. Signy [TDT^{+}03, VBV^{+}00]. silica [Bur01, DHBZ00, LRR^{+}09, LZL05, MGRP07, SVG^{+}05]. silicate [SES01, Sve02, WRJ^{+}05]. silicic [FBHB03, RLDL02]. silicification [RLDL02]. silicon [RGB02, YQC^{+}00]. sill [MVLM^{+}06]. Sillaginidae [JH01]. Sillaginodes [Jen05]. Sillago [CHG05, KCSPT09]. silt [RDC06]. silver [BTD^{+}04, JSIT04, KRP^{+}06, NW05b, SAW^{+}01, SC00b]. silverfish [MVHV06]. silvering [HYLT03]. silversides [CB08b]. similar [NMJ02]. similarities [WWDH09]. Similarity [Bel08a, IP01, Som08, Yos08a, Yos08b]. similis [CIHL05, CRSL05, DD09a, DD09b, EO04, HW08, IP07, MVN^{+}06, NMS^{+}02]. Simple [BH08, BVP^{+}06, CSD04, HDDH06, KBH01, Kne09, SM08b, SMF^{+}02, THF^{+}01a, THF^{+}01b, TT07a, VJKF07]. simplex [GGR06]. simplification [MDHL08]. simplified [HDP08, WW06b]. Simulated [BB04, MB05b, SH06b, AW03a, BOZ03, CHM06, DdNEvdZ05, FGS00, KCH^{+}03, MS09, NSH^{+}08, RSIS09, WW04, vdBÓME01]. Simulating [MKK^{+}09, NRC07, TÅ07b]. Simulation [BW00, RB05b, SWH^{+}04b]. simulations [DTD03, MMVB07, RAAT^{+}07, SDW07b]. Simultaneous [KUN02, Pri02, TUK09, MGL^{+}00, NHNB01, RZLS09]. since [RSNG09, SHS08]. sinensis [CACDC07]. Single [MS08, SKW^{+}06, BGF^{+}06, DBD^{+}05, DY05, Gou07, Joh04, OYNO08, OHL^{+}06, SN04, WS00]. single- [DY05]. single-cell [OHL^{+}06]. single-substrate [SN04]. singletons [NW05a]. singularis [CLR^{+}06]. sinicus [WLS^{+}09a, XW01]. sink [AG02b, DGW01]. sinking [CMG04, GP04a, GK00b, HS02, JPMGS08, KPT01, KT01, MYNY08, PFW^{+}02, PDT09, Pea03, SSH^{+}04, TJEM02, VJ04, WJWL05]. sinking-particle [GK00b]. sinks [BBA06b, CB08a, CSO08, QRWJF07, WGL03]. Sinularia [SP08b]. sinuosa [CLMR07, CLRM08]. sinusigerus [EKA03]. SiO [KYYM00]. Siphonophora [PGR^{+}01]. siphonostomatoid [LSBJ08]. Site [LAK07, SDK^{+}02, SBLF09, TLC06, WC03, BHPC06, BCB08a, FFYS^{+}07, GN04, Gla07, HALL05, KAR04, KK02, NTH^{+}02, SPL00, SBFP08, SWGC^{+}05, Thi03]. site-dependent [NTH^{+}02]. Site-related [WC03]. Site-selection [SDK^{+}02]. Site-specific [SBLF09]. sites [ALE^{+}09, BDR04, CLL^{+}09, Fox04, GSMD00, GZG^{+}07, HAD07, MJG08, MK03c, NLG05, PGS^{+}00, PZMG05, SABK03, SKDP02]. sitosterol [GWBdN09]. situ [SdNH^{+}06, WASW03, AW03a, APH^{+}04, BGJ^{+}07, BGKN00, BSAR03, BY09, CCK03, DC07b, EZTZ02, EB08, FHK^{+}08, FAB^{+}07, GCD^{+}03, Gor03, GMM03, GF04, GK00b, HH05a, HBdBV01, HSRF02, HHAH00, KKI^{+}06, LCC07, LCEC02, LHH09, LMD^{+}02, LG01, LWR^{+}02, LUIAH03, LPBD00, LKW03, Mar02, MGDP01, MP00, OKP06, ÓUA05, OT08, PNCS^{+}08, PJ08, PSPD07, PHS09, PR05, RGP^{+}03, Rob06, RR01, SK02, SWH^{+}04b, SG01, TSC^{+}07, Trn02, VPS08, ZSKA01]. situation [ABBF01]. sivado [ACAG07]. six [BBM03, Cla08, CTD^{+}09, FAFR03, GAAB^{+}08, KJB^{+}06, LK02, PH03b]. Size [BCT^{+}08, BB08a, DFH05, Don04, DKAH02, DHR04b, HTM^{+}02, HD07b, IT06a, JKNJ07, MFP04, RGC^{+}06, SBJ02, SD01a, SFL05, SMH02, Tay03, APN06, AHSPG08, AKM09, APG^{+}09, AP00a, BS08b, Bas06, BLHB04, BMG02, BFK^{+}09, BSED08, CF08, CS02a, CSMBH08, CSM02, CDAM06, CMGBJ05, CH05, DE01, DUW08, DFV01, DCF^{+}05, DWL09, EM04, FCCP06, Fer04, FIS04, FTB09, FSB07, FD02, GHV02, GAMB05, GMZvdL07, GBSQ09, GSM06, GGFSP06, GR07a, GM05, GC06b, GCM01, Har00, HKM00, HGS^{+}02, HLBS07, HBP^{+}05, HM06c, HS01, IG04, IF09, IRM08, JHJS04, JPC00, JRC06, JPPW02, JWM02, JMSM08, JH04, KS08, KWWH02, KHG06, KVV^{+}00, KdLCM09, LF02, LS02, Lir00, MHP^{+}00, MSLP06, Mad02, MHB^{+}01b, MBG^{+}03, MSK00a, MK03a, MK03b, MK04a, MK04b, MB02b, MHTJ07]. size [MHSS08, MPR02, MRJ05, MC06, MGW^{+}09, MB01, MVH^{+}06, MBP^{+}06, MHK^{+}01, MLG^{+}03a, MLS^{+}06, MHW07, MR07c, MF07, NLHO09, OvM02, OA02, PGM05, PMS^{+}08, PBFB08, PSRB00, PA07, PFT^{+}08, Pra05, RM05, REM^{+}06, RTL05, RCP^{+}05, Rol00, Roy02, RMAY04, RGC05, SS00a, SAWG05, SAJG06, SJR00, SW02b, SW02a, Smi04, SSSV00, SAP02, SZL^{+}05, STRB06, TSF01, TJSG03, TAW^{+}09, TDP^{+}00, VFF09, VHMB09, WWHE04, WB01, WWV08, WLM06, Whi02, WBZ^{+}09, Wol07b, ZBDE07, ZAH03, ZE01]. size-at-age [APG^{+}09]. size-at-hatching [RM05]. Size-based [HD07b, JPPW02, JWM02, JH04, MLS^{+}06]. size-class [DUW08]. size-compartments [KVV^{+}00]. Size-dependent [BB08a, Don04, IT06a, SBJ02, Tay03]. Size-fractionated [MFP04, RGC^{+}06, HKM00, Rol00]. size-fractions [MLG^{+}03a]. size-frequency [MHK^{+}01]. size-prey [SJR00]. Size-related [HTM^{+}02]. size-selective [HM06c, MVH^{+}06]. Size-specific [BCT^{+}08, DHR04b]. size-structure [TSF01]. size-structured [AHSPG08, SW02b, SW02a]. sized [BD03, CMRF05, SSSV00, Som00, SPRK02]. sizes [HB03, Zhu06]. Skagerrak [CLB^{+}09, FBSG02, GLS^{+}00, Joh02, MNR^{+}02, PWB^{+}03, SJ02, TNT06]. Skeletal [BSL02, GLB^{+}03, Mar00, WCZ^{+}09, CSMvE04, Lou08, MW02a]. skeleton [BS03, RFPS^{+}02]. Skeletonema [GKM02, YNM^{+}07]. skeletons [RJTM08]. skewed [RBG01]. skills [MFA03]. Skimming [FGS00]. skipjack [TOK^{+}04]. Skov [TW10]. skua [BPT^{+}00, KCV^{+}06]. skuas [BPT^{+}00, HRR^{+}07, KCV^{+}06, PCS^{+}07]. slatteryi [JYS06]. slipper [DEV07, LGD04, Thi05]. slope [AAGS02, BUF02, CWT00, CD03b, GMd^{+}02, Hil04, Kor01, LG06, LJM^{+}00, MLC08, MBB^{+}04, NBB^{+}04, ULL03, VBMBC08a, WKL01, CY06]. slopes [DBG^{+}09]. sloppy [MTN03]. Slough [CS09]. Slow [BWL06, DS01]. sludge [VFP^{+}07]. slug [LCW06]. sluggish [ZK01]. Small [BUCC09, CF02, CÅJ05, DOML^{+}07, LIM^{+}06, MJKH04, PR01, RU02, RZ00, RMAY04, SGL02a, SC07, WG07, BSZ05, BGL^{+}07, CHW02, CHSW03, CT05, CIFN00, CHT09, DB05, DCF^{+}05, DD09a, DD09b, FCCP06, FF03, GMK^{+}03, Gri02, JM01, JS08b, KOC^{+}07, KBSWG09, LAK^{+}08, LH02, LPHH^{+}09, MM06a, MM06b, MdL06, OGA03, Osb07, PLCG^{+}05, PdNS06b, PK06, RK09, SR08a, SHNV09, Seu05, SSM05, SPRK02, SDW07b, SMCE07, SIK^{+}08, TOK08, TF04, TTR05b, Und04, VHW^{+}07, VBA06, Vli05, Wil04, Wil09, WJB09, ZE01, KPT01]. Small-sample [CF02]. Small-scale [BUCC09, CÅJ05, DOML^{+}07, MJKH04, PR01, RU02, RZ00, RMAY04, SGL02a, SC07, BSZ05, KOC^{+}07, LAK^{+}08, LH02, MdL06, Osb07, SHNV09, Seu05, SSM05, SDW07b, SMCE07, TTR05b, Und04, Vli05, KPT01]. small-sized [SPRK02]. smaller [CMRF05]. smart [TT07a]. Smear’ [RJS04]. smelt [ADR^{+}05, BDRM06, BCB08a, CRC02, PA07, SD00a, SD00b]. smoker [SJS^{+}01]. smolt [HMH^{+}08]. smooth [WWAC00]. smoothed [GNL08]. smoothers [EOK^{+}08]. smothered [ATR^{+}08]. Smothering [NTH^{+}02]. snail [EKA03, KS09, PJ06a, PQRAC08, SR04b, Swa04, Tru02]. snails [Iye02, KMK01, PCB08, SAC^{+}00]. snake [KWHD08]. snakes [WH09]. Snapper [PBH^{+}03, DWB04, HJG03, LDR00, LSS^{+}09, MW06, NJPI03, NHS^{+}08, NSEL00, PCCL05, TSHT02, WMG02, WCF08, WPPS04]. snappers [FS08, ZH02]. snapping [LMC08]. Sneaky [Iye02]. snow [AHC^{+}06, ACM^{+}02, CSMBH08, DSMBG03, DdNEvdZ05, GCHA05, RGT01, SML02a]. Social [AFHS08, AM02b, DGVK^{+}04, HLH^{+}07, HM04, KKG06, MWR07]. Society [MCE^{+}01b]. socio [BS05b]. socio-economics [BS05b]. sockeye [MHSS08, Ran02]. Socotra [SC03]. Soft [BGGR02, Ell02, AFFH04, ARB04, BWPB04, BF04, BMF06, CTvdM^{+}08, CK05, DF02, Ell01, EHM^{+}09, FD00, Fro00, Gra02, Hun04, KL03, KGRW06, LS09a, LEB07, MAE03, MGD^{+}09, MC08a, MK02, NWÓ03, Niz07, PBAA04, PTP01, QS05, SP03a, SJE^{+}06, SE04, SP08b, SS04b, SRS04, SDM06, TAS04, THF^{+}01a, THF^{+}01b, TSHT02, TJA08, WAA^{+}09, YH02]. soft-bottom [BMF06, DF02, Fro00, LEB07, MAE03, Niz07]. Soft-coral [BGGR02]. Soft-sediment [Ell02, AFFH04, CK05, Ell01, KL03, KGRW06, LS09a, MGD^{+}09, MC08a, QS05, SE04, SRS04, THF^{+}01a, THF^{+}01b, TSHT02, YH02]. soft-shelled [ARB04, PTP01]. softshell [SMT08]. Solar [BD08, BV08, GUT04, BRBPL06, BDW^{+}00, CABS02, DQW05, DL03b, FJV^{+}02, FNDP08, HBdBV01, HGG^{+}03, HQDH06, Les00, Prz05, RBRSP00, VMH04, WK00]. SOLAS [FFYS^{+}07]. sole [AMG^{+}07, Bro06, EMG01, EGPT03, LGD04, LC03, MKR07, RZS05, RSSA07, RH08, TdLvdVR08]. Solea [AMG^{+}07, EMG01, EGPT03, LGD04, LC03, MKR07]. Solemyidae [NSSI06]. Solenogastres [KCB06a]. solids [APW^{+}04]. solidungula [PK01]. solitary [GAZ02, MC05a, SR08b, WH03]. Solomon [UB01]. sols [Bur01]. soluble [SMC^{+}05, SZL^{+}05, TL08]. solute [HAA00, KPT01, KT01, MW04a]. solution [RSL09a]. solutions [Nae06, WR07]. Somateria [LG00a]. Somatic [BBM00, DTV09, LQRGT06, Sha00, TBAT02]. somatic-otolith [Sha00]. Some [BDJ09, DGdG^{+}09, JTB^{+}03, KKBH03, MMCF07, PM00, TSC^{+}07, WR07]. Somerfield [Yos08a]. something [HMR00]. sometimes [BBG^{+}08]. Sonnenholzner [Ano07h]. sooty [EC07b, PSC05, SWS^{+}09]. sorenseni [MBBP08]. sorting [TPS^{+}05]. Sound [BG07b, DD01, GKWR02, GSM^{+}02, IFL02, JTN09, LCEC02, Lus03, Lus05, MZJ06, PBS^{+}00, PS01, Pur03, RW08, RW06b, RKGH01, SSJ^{+}02, YTK^{+}05, ZHWCS08, FMZvW02, GRPT^{+}00, KCTG06, MG09a, MBBW08, SN09, TJM00, Tya09, GK07, LW02, TRDV00, TWYS03]. sound-scattering [FMZvW02, MBBW08]. sounders [KRKS09]. sounds [MCBT07, WW05]. soup [BEKT09]. Source [RFLK03, ARBF02, BS07b, CRS^{+}07, CWBS03, CHV^{+}09, CTFC08, GSB^{+}06, GC07, Hun05, KSS^{+}08, KHWH03, LFRR^{+}07, LZW^{+}08, RCAK03, Ryp08, ROBFR08, SM07a, SEVSP02, SKMS06b, SZBZ04, TPSW00, WWR^{+}01, YTK^{+}05]. Sources [ABBN09, BMSD00, CB08a, GSK^{+}08, TBI^{+}09, WBC01, AND05, BBA06b, BVI^{+}05, DBR^{+}07, DES^{+}00, FKBB01, GFL02, Hil09, HMK06, JBS08, KCP^{+}07, KKIV01, KMK01, LWK^{+}07, MTV06, MLM^{+}02, MBN^{+}09, PRB^{+}08, QFJ^{+}07, QRWJF07, RHBW09, WPRS05, WAZ07, WGL03, YI07, CRCV04]. South [BDR06, GPW^{+}02, GB07, GSCW00, MGL^{+}07, PDQ01, PK00, STE04a, TCC04, Bra08, CMMN^{+}01, DSJBS^{+}02, DL01, HW00, PH05, PMNL07, RS03, RGB02, SZBZ04, TK01, TLFS06, TWT08, YBGP06, AWP^{+}01, BC02, BGSBC02, BBS05, BBG^{+}06, BCH00, CF09, CCL03, CBS^{+}08, CTB07, DC02, DGJ^{+}03, DCA^{+}09, EB04b, ESB^{+}04, FDD^{+}03, FBJ00b, GdLB^{+}09, GCHC00, GUL^{+}07, GB02b, HGC^{+}07, HH07a, JBCL03, KPN03, KWAT08, Les05, MP00, MR01, NLCS05, NPM01, OBG^{+}07, OPSB07, PSPZ09, PPC09, PBT^{+}09, PPB09b, PSC05, PWC^{+}09, RHDC05, RF08b, RMKR00, RBG^{+}07, SSB02, SPP^{+}09, SWW02, STA^{+}05, STAS06, SNP03, TKVL04, TCRT^{+}07, TMFB06, TML03, TGT^{+}06, VTDS09, WHM^{+}01, WAW^{+}09, XBW07, ZMTB07, ZTW08]. south-east [PH05, TK01, TLFS06]. south-eastern [Bra08, HW00, YBGP06]. south-western [DSJBS^{+}02, RGB02, TWT08]. Southampton [CW03b]. southeast [BB09, DCAL^{+}02, MMA03b, SAH09, DM09]. southeastern [BHPC06, BF04, BNSG01, BNP^{+}09, CB05a, FBJ00b, GTE^{+}09, JBH00, MRB^{+}08, MBN^{+}09, PAW00a, RRH09, SSHW^{+}07, SSL^{+}07, SO01b]. Southern [CBTW02, FVS^{+}08, GFPM01, KPM00, PFW^{+}02, AML^{+}03, AP06, Ano01a, AWA^{+}05, AVA^{+}09, ABB05, BMG00, BBG^{+}04, BL04, BRBPL06, BCL^{+}01, BM06, BO00, BMF06, CTA^{+}03, CL04, CDF^{+}08, CIFN00, CNT^{+}04, DLLB03, Dav00a, DCD^{+}03, DRF08, EDBvdL08, EB04b, EBA^{+}03, FPB^{+}07, FJV^{+}02, FGSH^{+}05, GMHA09, GB01, GLFU^{+}04, GFN^{+}09, HP06, HLH^{+}07, HQDH06, HMK06, HK04b, HFMP03, HMO^{+}04, JBG00, JYS06, KHIA02, KM03a, KB07, KGHL^{+}01, KMR01, LBPS03, LFF^{+}09, LLSL06, LVvdLH07, LBC^{+}09, LMB^{+}09, MRZJ05, MHBB05, MHABGC04, MGH00, NLP^{+}01, NFN^{+}09, OA02, PLCG^{+}05, PTSJ06, Reb01, RRDHG08, RCF09, SR08b, SMG02, TS05, TSSA03, TBH08, TA01a, TPS^{+}05, VSB04, WG02a, ZM06, Zam02, BF03, CH06b, CH07c, CAS00, CWN^{+}07, DÁSG^{+}01, FBBH07, GVF^{+}03, Hob00, JMM^{+}09, Joh07, LL03, LSG^{+}05a, LGC^{+}06, LGM^{+}09, MVLM^{+}06, MR01]. Southern [QMF05, RW03, RPN^{+}01, RLDL02, SRGS01, SVG07, WRW06]. Southwest [GRS07, LLM^{+}08, LBC07, BKSS07, GMC^{+}09, MRPL07, MZD^{+}05, ZAHJ^{+}06]. southwestern [ACPS09, Bar07, CMMN^{+}01, GTGM06, JDTB06, JMM^{+}09, MM07b, MBG^{+}08, NTN08]. sp [APP^{+}09, BK06a, CD07a, EKN08, FB03b, Gru03, HGAS^{+}01, HB08b, JSK^{+}07, LS05, LT00, LELC02, LW06, MW04a, NWÓ03, OGA03, PPB^{+}01, PR09, RF08a, RAE07, RZEACGF00, SLCB00, SLB^{+}08, SM05, TSSQ06, TNSM00b, TWM^{+}01, dBKSV00, BS08b]. Sp/Py [BS08b]. Space [BS08b, AM01, BF07, BPC09, GC03, JSWH03, JMJ07, KH00c, LVZW06, MD01, Mok04, Rit02, RT00a, SJH^{+}04, TAW^{+}09]. space-limited [SJH^{+}04]. space-what [JSWH03]. spacer [PG01]. spacing [BBE03]. Spain [NIF00, ÁSG04, AAF^{+}04, Arr05, BGQR01, BRBPL06, BVCG01, CIFN00, DLB04, FJV^{+}02, PCSS01, PNCS^{+}08, PNCT^{+}08]. spallanzanii [HK02c, PG01]. span [HM05a]. Spanish [OEDSJ02, ST06b]. spanning [SVG07]. Sparidae [DWB04, GKAS08, HJW08, PBH^{+}03]. Sparisoma [CRAP03, GPK^{+}04, GHV05, PS08]. Spartina [CLH^{+}07, CS03c, HH05a, HGK02, ML06a, MJG01, NLG05, PMH08, PAMJ03, SZZ08, WWAC00, WPB08a]. Spartina-invaded [SZZ08]. spat [BD03]. spat-sized [BD03]. spatangoid [TR05, VVTH07]. Spatial [ABB02, BCL03, BIR00, BI02, BJ08, BK06c, BHD04, BMJ02, BTF^{+}06, CHL^{+}05, CCMK03, CB05b, CMD^{+}09, CTVP03, CGC^{+}09, Dav00a, DF04, DTB07, DD09b, EM04, Fau09, FCP^{+}07, FBT^{+}01, GPLAG04, GEM02, GB09, GDL^{+}01, GCM01, HF06, HJK01, HVW^{+}07, HMFC07, JH01, JKES07, KTNK07, KTK08b, LTNC07, Lec05, LZM^{+}03, LRGdS00, LTH^{+}06, MNR^{+}02, MGL^{+}07, MPM^{+}02, MW08, MGREA07, MD08b, MCS05, MRB^{+}08, MSP^{+}06, MTCU08, NE01, NA04a, PMS^{+}06, PSPZ09, PTM08, PCD04, PEDT03, PWB^{+}03, PDRT02, PCBL05, PDG^{+}02, PR05, RS03, RP06, RHSN07, RT00c, RW08, Rue01, RS05, RW06c, SM01a, SASR03, SRJ09b, SC03, SH07a, SH07b, SPBM05, SE04, SBM02, SvNvdMvdV09, SGHG07, SLHR09, SSBZ07, SCAF00, SH00, TMAN08, TGA02, THS^{+}02, TG03, TPD^{+}05, TH06, VMCG09, Vli05, WISH09]. Spatial [WBW^{+}06, WBZ^{+}09, YH02, Zac05a, Zac05b, Air00, AA04, Arm01, BS00, BK03, Bar06, BM02b, BC01a, BCBVM05, BCBVM06, BRP00, BWP^{+}06, BML^{+}08, BUA02, BBRJ04, BUCC09, BKDB00, BSP03, BDÁM04, BH02, BWO04, BLP^{+}02, CMD07, CU08, CC03a, CSM02, CWM^{+}09, CC07, CRK06, DSS^{+}09, DC06, DDS^{+}01, DSJBS^{+}02, DDL^{+}03, DB05, DS06a, DOML^{+}07, DWL09, EMG01, FA06, FHG^{+}03, FCP02a, Fra05, GBdS^{+}05, GVF^{+}03, GLS02, GMK^{+}03, GY00, Gis05, GDF03, GK07, GSS09b, HWC^{+}06, HLH^{+}07, HTL^{+}02, HSCF06, HHWT02, HLTT09, HM04, HGKK06, IP01, JWM09, JÅC^{+}01, Jos09, KCS04, KvdMDP09, LNV^{+}05, LBvdH03, LJ08, LV02, LAK^{+}08, LG08b, LBHL^{+}09, MBW02, MC07, MRS00, MUT04, MTH^{+}08, MKL^{+}03, McI03, MB01, ML05, MHK^{+}01, MG09b, MIH01, NDV^{+}00b, NJPI03, NZM08a, NMDS00]. spatial [NT06, NLHO09, OC01, OSP01, OMP^{+}05, OGL^{+}09a, OM04, PMM^{+}00, PNK^{+}02, PKM04, PTSJ06, PTB03, PJJ^{+}05, PG00, PCHM07, PR01, PPL^{+}02, PKS07, RFN02, RB04, RGCC^{+}08, RHS01b, SDS^{+}03, SV07, SJ09a, SDW07a, SDW07b, SRB04, TAFBC07, TTR05b, TW10, VKH08, VCSL^{+}09, VL00, VHM09, VBFF08, WGD05, WG09, WM02a, WMS03b, WBSL00, WDN00, WMB00, WG07, YW04, dSHCR07, dSCV^{+}08]. Spatially [LTPL05, OWBD07, SS09c, SMP01, BM00, TP00b]. Spatio [CIHF01, DE01, FC08, FHVMDS07, HW01, LLW09, MJS09, MBR^{+}05, PMR06, TS07, TSSQ06, HGRB04, HFMP03, KL03, MTLT02, SNH04, ST08, WW06a]. Spatio-temporal [CIHF01, DE01, FC08, FHVMDS07, HW01, LLW09, MJS09, MBR^{+}05, PMR06, TS07, TSSQ06, HGRB04, HFMP03, KL03, MTLT02, SNH04, ST08, WW06a]. Spatiotemporal [GSS09c, SWS^{+}09, SLM07, JDCB09]. spawn [ALE^{+}09, FWF02, HGKJ01, SCC05]. spawn-inducing [SCC05]. Spawner [MMB03, CMB^{+}09]. Spawner-recruit [MMB03]. Spawning [BG09, BM00, GES^{+}06, HDG^{+}08, LS09b, LHB^{+}06, MBG^{+}08, SSR^{+}09, SLCC09, TKO08, WH03, AWBO09, AP06, ANF^{+}00, BBC^{+}09a, BP02a, BM02c, BMWH08, BCB08a, CHL^{+}05, CSH00, CCK^{+}09, CMS^{+}01, EMG01, Enb05, EH04, FBJ00b, FAB^{+}07, Gla07, GMS02, HNYI07, HDM04, HDP^{+}06, HS03, HVW^{+}07, HFMP03, HCR01, JBH00, KN00, KNF^{+}03, LLH06, LC02, LGRB^{+}08, LHL^{+}07, LS02, LBHL^{+}09, MH08b, MH09, MSK00a, Mar02, MSC04, MWG02, MH00, MYE03, MST03, MS01a, MYH07, MC05c, MRGC09, NWHH01, Nem05, NKS03, OKH^{+}07, PCCL05, PA06, RMM04, RVMI^{+}03, RTL05, SSL00a, SPL00, SHNV09, SHK05, TOK08, TFSJ00, WDT^{+}08a, WTH04, WWS^{+}05, Lap01b]. spawnings [HDG^{+}08]. spear [LS04]. spearfish [HLH^{+}05]. special [BVVM02, JPNM^{+}06, LTFJ00, LDWC05, TL05]. specialisation [Bat07, BJC02, BPM^{+}06, BDLB02, DH07b, GJ05, MSLP06]. specialist [APP^{+}09, BMdNP09, Kru01, PMH08]. specialization [JHH01]. Speciation [TGSW06, DRV^{+}05, RC04b, WGM02]. specie [DFV01, GMD08]. Species [BF07, BFE^{+}05, CF08, Chi02, DH01, GB02a, Gra02, KOTL07, Kas00, KRVB05, MZJ06, MM07b, PSPD07, STI03, SKFPS02, SB06, WST^{+}09, Zhu06, ASA^{+}08, AKRA02, APH^{+}04, ACPS09, ARP09, Arm01, AP00a, AAF^{+}04, Arr05, BS03, BLD^{+}05, BBM03, BVC^{+}02, BK04, BMG00, BIL^{+}02, BH08, Bel07, Bel08a, BBDW09, BLF06, BBC04, BDEM01, BD11, BBRJ04, BSR09, BAMJ09, BMW^{+}00, BRSM07, BHFC08, Boo02, BDGK^{+}06, BFV08, Bra01, BS04c, BCC08, CLS00, CF02, CH06b, CLABRO^{+}03, CML08, CSM^{+}09, Cha03, CU08, CAMC07, CB05a, CW01, Cla08, CR07b, CMGBJ05, CDK^{+}07, CHG05, CHP^{+}05, CV05, DCT^{+}04, DLT^{+}06, DJG^{+}07, DW09, DMK^{+}08, DBD^{+}05, DTV09, DH09a, DY05, DVBLS07, DSH03, DPSD08a, DOML^{+}07, DS06b, DB07b, EL05, EYH^{+}06, EHM^{+}09, EF07, FAFR03, FCC^{+}08, Gar00, GL00b, GBSQ09]. species [GKS^{+}07, GSP^{+}02, GK08, GLFU^{+}04, GDS02, GHS05, GP08, GFPM01, GLBF07, HG08, Hal04b, HF08, HDM04, HSP^{+}08, HSC^{+}09, HVCT09, HDG^{+}08, HEN^{+}08, ISBP06, ITN09, IFL02, JHL00, JHR05, JKQ02, JM03a, JTB^{+}03, KW05, KNTK05, KTT04, KWTP06, KNT^{+}07, Kon00, KCG07, Kro05, KJRG09, KKSR05, KGS^{+}08, LFP^{+}04, LEF^{+}06, LR06, LTFH06, LGFM00, LGVF00, LG08b, LTL06a, LLCHG01, LW02, LRGdS00, MPR^{+}06, MBVBC09, Mag02, MAE03, MHP03, MUT04, MCP^{+}06, MCC^{+}00, MGC03, MC06, McI02, MS08, MP00, ML07, MN04, MR06b, MvdBÓ00, MSJ^{+}09, MHV^{+}00, MA02, MMMB^{+}05, MLW03, MHT07, MB04, MB00b, MG05b, NNG03, NWHC03, NA04a, NRH^{+}07, NK07, NB06b, PDCH06, PGM05, PB08a, PCL^{+}09, PWLC06, PDO01, PBVP09, PTSJ06, PKLD09, PPS^{+}07, PP02, PTH08, PJ08, PZMG05, PCL08, RW07, Reb01]. species [RSE01, RL03, RGEB00, RFN02, RGF^{+}03, Rue01, RFH^{+}06, RKJ01, SZGS08, SGB09, SLL02, SH07a, SLFM03, ST06a, SMC09, SM09, SH07d, SBF08, SAF^{+}09, SLM07, SDR01, SSBG02, SHM02, SF05, Sto03, SB08, SHSF08, TOK08, TN04, TVD02, TS04, TSC^{+}07, THHY05, TMFF00, TL03, UG04, ULL03, UBRvO06, UP07, UST^{+}04, VRST^{+}07, VSBB03, VFJ^{+}07, Vii07, VMH^{+}08, VPS08, WM06, WHD^{+}05, WAK^{+}09, WBR07, WMCF08, WMD^{+}07, WOH05, WG07, WDR^{+}05, YMHV02, YSS03, Zab09, ZM06, Zam03a, ZER^{+}01, ZJW^{+}04, ZAH03]. Species- [WST^{+}09]. Species-area [Chi02]. species-environment [THHY05]. species-poor [CDK^{+}07, DW09]. species-rich [CDK^{+}07, DW09]. Species-specific [DH01, GB02a, KOTL07, MZJ06, STI03, AKRA02, CLS00, GKS^{+}07, HDM04, NRH^{+}07, PKLD09, PJ08, SHSF08]. specific [AKRA02, BBM^{+}02a, BPG08, BPM^{+}06, BCT^{+}08, CLS00, CMRF05, CST08, CKB^{+}09, DHG^{+}04, DH01, DHR04b, EO04, GVM^{+}05, GKS^{+}07, GB02a, GDR^{+}04b, HDM04, HD06b, KOTL07, LR06, MAVO^{+}06, MZJ06, MW04b, NRH^{+}07, OJ04, PC05, PKLD09, PJ08, STI03, SMM^{+}04, SMSJ04, SBLF09, SHSF08, SM01b, TAO07, TP07, TLPG06, WST^{+}09, vLvSGS05]. Specificity [Fre07, IH02, MKKK08, SPCG09, CSG01, MW04a, OHS^{+}04, RWSP03, RLKW04]. speciosa [TLLKC04]. spectacled [BDF09, RL03]. spectra [Fra05, VSVV03]. Spectral [JS01, APH^{+}04, CSD04, SO08]. spectral-resolution [CSD04]. Spectrally [LDHC04]. spectrographic [JCMR^{+}03]. spectrometer [CSD04]. spectroscopy [AL03a, COB^{+}03]. spectrum [Mag02, STRB06]. speed [BLH02, FOFG09, GLJ07, HL05, LHL^{+}07, TND^{+}05, TBH08, VPWE09, YA09]. speeds [BF01b, Fis05, ZTW08]. Sperm [AY03, MWR07, RWE04, SMM^{+}09, AAY^{+}07, CBO04, DJG^{+}07, GDG02, KWWH02, MHB01a, MS08, PTP01, RCGA^{+}04, SG06a, SAJG06, Whi02, WCJL08]. Spermonde [CR07b, dVCH^{+}06]. Spheniscus [SAB^{+}02]. Sphyrna [DH06]. SPI [SK02, SWH^{+}04b]. spicule [FD08]. spicules [HLY05]. spider [SMFP08]. spill [AGCF^{+}08, BBD^{+}02, DBJ^{+}00, DBF^{+}02, DLD^{+}00, EP08, EFL^{+}09, EBT^{+}02, GSM^{+}02, LFRG^{+}05, MAVO^{+}06, NAB^{+}06, OGMC06, PMGE01, PD01, VML05]. spill-in [EP08, EFL^{+}09]. Spillover [GQR06, GAAB^{+}08, RAM03, FVB^{+}09, SWG^{+}09]. spills [HMD^{+}07, RMT^{+}05, Rui04]. spined [JHJ^{+}03, JHJS04]. spines [IH02]. spinifera [AS06a]. spinner [GF05, OPSB07, WWGR07]. spinning [PCC07]. spiny [BFLÁ01, BL09, CH05, EJKN03, EP08, GRPE08, GQR06, GTGM06, JRTJ04, LÁMBF09, NPLZ00, Niz07, PE05, SAWG05, SAJG06, SMFP08, TGGM05]. spionid [EWSH08, KW00a, SWH04a]. Spionidae [BBRJ04]. spirale [KJ04]. splendens [SD01a]. splitting [ANF^{+}00]. spoil [CM04b]. Sponge [dVCH^{+}06, dVBC09, BP04, BBB^{+}06, Bel07, BUCC09, CÁ04, CU06, CWCP^{+}06, DUW08, DDQ04, DDTQ05, DWL09, EP00a, ET08, EEWBdN08, HSS09, KPEF00, LWH^{+}04, PMH07, PHM^{+}08, PCJF06, PWdN^{+}00, RDC06, RCDC06, SSHW^{+}07, SSG^{+}00, TPG09, TL08, TLPG06, WNWH02, WEEBdN08, vDHHM08]. sponge-associated [PWdN^{+}00]. sponge-cyanobacteria [ET08]. sponge-dominated [RCDC06]. sponge-microbe [vDHHM08]. Sponges [Bel08b, ALPK00, BDR06, BIC^{+}03, BI03, CBGLM08, DDTQ05, FD08, HLY05, LYL^{+}07, LVZW06, NSFW09, PAM^{+}09, PFP03, QDDP06, SRDG05, Thu07, TCT^{+}00, WP00a, WP00b, dGvdBvO^{+}08]. spongin [WNWH02]. spongin-boring [WNWH02]. spontaneous [ZK01]. spore [GLJ07]. spot [CRCN07, MT05]. spots [CS05]. spotted [GF05, Kup03, LBHL^{+}09, UP01, WWGR07]. spp [ALOT01, BRG^{+}02, BP03, BO00, BWSH00, BBS07, CPA^{+}03, EC07a, FDBT06, FFHW07, FHN^{+}09, FAB^{+}07, GSRH02, GSCW00, GSSR02, HHC00, HAP^{+}08, HCBP05, HKR^{+}08, JC08, JMA02, JKS^{+}01, JM03a, KTK08b, KCB06a, KSG03, KCSPT09, LBH^{+}04, LTPL05, LGB03a, LCCTFO03, Lus03, Lus05, MJS07, MJS09, MVB^{+}05, MFCC02, OJPM01, OLWB06, PIY04, PTB03, PDG^{+}02, QDDP06, RRS05, RDD^{+}02, RLFEF03, RMC^{+}07, STPA04, SAC^{+}00, SSGJ00, SPP^{+}09, SHS^{+}05, SYWM05, SBF08, SC06b, SNCP04, SAM^{+}01, SH00, TCD01, TJ02, VSRGW09, VAB09, WSL06]. Sprat [KRK09, BGK^{+}06, DMV07, KM00, LPHH^{+}09, RMH07, SNY^{+}05, VCBH06]. Spratelloides [MVH^{+}06]. Sprattus [BGK^{+}06, DMV07, KRK09, KM00, LPHH^{+}09, RMH07]. spray [ORPND01]. spread [MP00, Min02]. spreading [HHJ04]. Spring [HW08, KFN01, LC02, OSL07, WLS^{+}09a, BVCG01, CCC09, CVÁM06, CSL^{+}05, CS02b, DD04, DR00, DRB^{+}01, EO04, Enb05, EH04, FHL^{+}09, FTK^{+}04, GHL02, GCHC00, GS02, HJ03, HHC00, HLH04, HM09, HK03c, JPNM^{+}06, KKA^{+}06, KNF^{+}03, LRN^{+}09, LLW09, MNR^{+}02, MNTS08, MLS^{+}08, MJ07, OPSM04, PTD05, PDT09, RYF01, RFG^{+}09, RGP^{+}09, SAB^{+}05, SAT^{+}03, SV08, SBBS^{+}02, SHFP^{+}03, SYH^{+}09, TDRM08, TNT08, VGS^{+}01, VSV04, WHD^{+}05, WDT^{+}08a, WWS^{+}05, WPTS08, ZGV00, ZIUG07, vdBÓME01]. spring-neap [KKA^{+}06]. spring-spawning [EH04, KNF^{+}03, WWS^{+}05]. spring-summer [FTK^{+}04]. Springtime [JBT^{+}07, TT02, RWR^{+}00]. spumigena [EÖKS^{+}02, KSKL^{+}03, SGK04]. squamata [BHFC08, LFP^{+}04, SDR01]. squat [RT00a]. squid [BMV02, BHMR01, BGH06, BMCH09, CCC09, CD03b, DJG^{+}07, DGC^{+}05a, GES^{+}06, GMB^{+}06, IMW^{+}02, IMSO09, IAKS03, IMS05, IS07, JBG00, JM01, JSWH03, KS04a, MH00, PTSJ06, PJN01, PJN03, RCGA^{+}04, SM00a, SJ05, SS04a, SJL05, VAJ^{+}03, WHS^{+}09, WKIK04, WKMK06, ZKHP07]. squids [Par06]. Sr [MC05c, SZL^{+}05, CIT04, HYLT03, KSK07, LIT05, MC05c, SCIT09, SZL^{+}05, TSI02, TCW^{+}07]. Sr/ [MC05c, SZL^{+}05]. Sr/Ca [KSK07, LIT05, MC05c, TCW^{+}07]. ssp. [SG06b, SS05]. SSR [SPI^{+}07]. SST [SW09]. St [BRB^{+}00, CSMBH08, CDH^{+}09, DVBLS07, LHK01, MRZJ05, MGH00, OA02, OSL07, PT04, QS08, RBHS06, RCF09, ZGV00]. St. [Chi02, CS03a, CGL^{+}02, CS05, DHRG05, Edm00, FLR05, GDBF06, LF08, MHCB08, Nug02, PDRT02, PJR^{+}03, RM06, SAH09, SSS06, WDB^{+}03, XZLP09]. Stability [GC03, HBC09, San00, Bar06, BHSR00, HGMH06, IRM08, LDP^{+}08, PZB^{+}05, Reb01, RL09b]. Stable [BPM^{+}06, BW01, BGS^{+}02b, CDF^{+}08, CR08, FGSH^{+}05, FCH^{+}08, FOH^{+}08, GFF08, GSL^{+}06, HARC05, HMW^{+}00, HEN^{+}08, JMB^{+}05, JCK03, KSK07, KMK01, KW02b, MSF05, MD07, QMM^{+}08, ROBFR08, SJE^{+}06, SHS^{+}05, TLTB03, VMH^{+}08, WIB08, ZS04, AND05, ARAMPW02, AYSL09, BPT^{+}00, BEU^{+}07, BAOV06, BWCW02, BMSD00, CF09, CHO09, CÅJ05, CH07c, CLJ^{+}08, DFPG^{+}03, DSS^{+}09, DOML^{+}07, ECK05, FC08, FHB^{+}02, FHDA^{+}07, Fre03, GFN^{+}09, HWB^{+}09, HTM^{+}02, HM06a, HBW^{+}05, HHH02, HH00a, HD07b, JBS08, JWM02, JMSM08, KFK^{+}07, KTNK07, KKL^{+}03, KTK^{+}02, KKIV01, KMB^{+}08, KG07, LF08, Lee00, LSMR05, LHK01, MWR07, MRHP01, MLM^{+}02, MS01b, MPB^{+}09, NHS^{+}08, NSEL00, NMBH02, OP08, PMNP^{+}01, QFJ^{+}07, QMF05, QRWJF07, RSNG09, RTH06, RSD^{+}08, RCGA^{+}04, SSMM06, STF^{+}09, SSS^{+}05, STH^{+}06, TRH^{+}06, TSEH08, TBI07, TBI08]. stable [WAZ07, WWWR09, WWR09, YI07, dlMPN^{+}03, vOSF^{+}09]. stable-isotope [AND05, ECK05, FHB^{+}02, HH00a, PMNP^{+}01, TSEH08]. stactophila [CY06]. Stage [EO04, TP07, ABF07, Arm00, BBM^{+}02a, BBM00, BSED08, CID^{+}06, Hal04b, HDP08, JSMB05, MW01, NMCT03, OAM02, STCJ01, TJSG03, Trn02, VGW00, Wen00, WIB08, vdVBGW09]. stage-dependent [WIB08]. Stage-specific [EO04, TP07, BBM^{+}02a]. stage-structured [Hal04b]. stages [BT06, BBC04, BTF^{+}02, BTD^{+}04, CPKV00, DL02, DSMBG03, FRvdMF02, FGT^{+}07, FS07, GB07, HGS^{+}07, HKSW01, HRC^{+}07, Kev01, Kro05, Kur08, LVA07, LW00, MBHN07, MHTJ07, McQ05, OSP01, PSPD07, RvdPHW04, SZGS08, Urb02, WDRH07, XWQQ05]. staining [Kam00]. standard [KCTG06]. standardising [HLBS07]. standardized [JMB^{+}05]. standing [CLS00, Kir02a, REM^{+}06, WJB09]. stands [BDGK^{+}06]. star [BHFC08, BV08, DCT^{+}03, DCT^{+}04, GWH03, GH08, GTPC04, GHdN07, MJS07, MJS09, MAH^{+}08, McG02, MSY02, MW09, RHG04, SM07a, SWM03]. starfish [BFB^{+}09, JMU09, KSH^{+}09, LRM00, PWM^{+}08, Pra01]. Starmans [Ano02g]. stars [GHJ02, GWBdN09, Har06b, LS09b, SCC05, SLCC09, WP00b, WBDR06]. starvation [BLH01, MO05, PGBA04, SFPGA05, SN02]. starving [CES^{+}03]. STAT [CG05]. State [BS04a, MVHV06, AB04, CBEL06, JT08, JMJ07, KWHD08, NL05, SWH04a, SKMS06a, SMHG09, Tak02]. State-dependent [MVHV06]. state-space [JMJ07]. Staten [RTPS07]. States [DMK^{+}08, FWC02, NNLF09, WDR^{+}05, CWM^{+}09, GL00b, GL00a, KMWH04, NLHO09]. static [dZCVC01, dZCVC01]. Station [HBN01, Sta08, CPH^{+}02, HBH07, KKLL02, Rol00, TAÁO^{+}03]. stations [AB00, CDS07, HCB^{+}07, SFH00b]. Statistical [BGDS05, Kup03, PBAD05, PMGE01, PMGE02]. Statolith [ZKHP07, CR07a, SPM03, Zac05a, Zac05b]. statoliths [IAKS03, LZW^{+}08, VAJ^{+}03, WHS^{+}09, ZMP^{+}03]. status [AFHS08, Ant06, BGLL09, BR03, BPT^{+}00, BBD^{+}02, BG04c, CH07b, CMRB08, CH00b, DWB04, EDB^{+}04, FPKC05, FBKB03, GNP07, HHW^{+}07, IRL^{+}09, KPA^{+}09, Law07, MAOP09, MWD^{+}09, PDFD09]. steep [LPHH^{+}09]. Stegastes [HK02a, WM02b]. Stegostoma [DNL08b]. steinii [ONP02]. stellatus [BCABC00, DMM^{+}03]. Steller [BRI^{+}07, GT08, KG07, MW08, STV^{+}09, WTR02, WWS^{+}05, WS06]. stems [WPB08a]. Stenella [MZD^{+}05]. Stenobrachius [MKWO00]. stenolepis [HSSS05, LS06]. Stenopus [CSO08]. step [MF08]. steps [BVP^{+}06]. Stercorarius [KCV^{+}06]. Sterechinus [BP01, BPT07, KR01, TYC00]. sterility [BGQR01, BMB02]. Sterna [EC07b, NMBH02]. steroidal [TAA04]. sterols [DPGN08, HRR^{+}01]. STf [Joh04]. Sthenoteuthis [Par06]. stickleback [JHJ^{+}03, JHJS04]. sticklebacks [WP07]. stiffness [BFK^{+}09, PvDMB08]. still [AHDT^{+}09]. stimulated [KSKK^{+}07]. Stimulation [O’C07, SBC^{+}06, TRS^{+}06]. stimulatory [OJ04]. stimuli [Arm00, BSW^{+}09, BGL^{+}07, UP01]. stingrays [SR08b]. stirred [JKNJ07, OIP04]. Stochastic [ARP02, GBD^{+}06, LBD00, MDG04, SJH^{+}04]. Stock [NSEL00, BM02b, BM02c, CLS00, CA00, CLB^{+}09, Enb05, KET^{+}03, LLH06, LS02, MMB03, OLG^{+}09, REM^{+}06, SP08a, SMA^{+}03, SRJ07b, WJB09, DÁSG^{+}01]. stock-separating [SRJ07b]. stocked [Wes03]. stocking [TLFS06]. stocks [Bar02b, BGH06, BT00, DM03, EGPT03, KFL09, Kir02a, NLHO09, Ogu07, PK00, SMF00, TP00b]. Stoeckeria [JKK^{+}05, JKS^{+}07]. stoichiometry [Cla08]. stomach [CMB^{+}05a, PJN01, PJN03]. stomachs [Lin04]. stomatal [ML06a]. stomatopod [KHK^{+}06]. stone [CTA^{+}03]. stony [ACPS09, Rii01b, Rii01c]. stoplight [CRAP03, GPK^{+}04]. stopover [RPC^{+}06]. Storage [PK01, AMR01b, BGD^{+}03, CVÁM06, EBG^{+}00, FWD^{+}01, HSSS05, HNMC09, KP01a, LS05, LHK06, MS08, MBL07, PH03c, PGS^{+}00, SGL^{+}02b]. storm [GSC^{+}01, HM06a, HHM^{+}06, QMF05, WPTS07]. storm-petrel [HM06a, QMF05]. storms [GB07, GR07b]. story [BT00]. strain [AGCF^{+}08, CMJ00, ITE^{+}04, SH03]. strains [BGF^{+}06, LMCQ02, PCYC02, YM09, ZJW^{+}04]. Strait [Gar00, HLPHA^{+}01, HDRC^{+}08, KKKK03, Kev01, TSJ05, VAP^{+}07, AP05b, BCVC07, ESDKM09, EBG^{+}00, KHH08, MGF^{+}03, OHL^{+}06, SD09, dSCV^{+}08]. Straits [LC08]. Stramonita [RBG01]. stranded [RH03]. strategies [AM01, Arr05, BGF^{+}06, CEB^{+}08, CPP^{+}04, CW09, DBH^{+}02, Enb05, FA08, FBBH07, FSB07, FEFT08, FBN^{+}07, HPH^{+}01, HGB01, HGHB00, HH00a, Hue08b, IBB^{+}07, JF07, LGC^{+}06, LB08b, Lus03, MGC^{+}09, MM03, NGS09, NZM08a, OW02, PCC06, PSC05, Rit02, RST04, SSN04, SSS^{+}05, TDT^{+}03, VACT^{+}08]. strategy [AMR01b, BLD^{+}05, Bat07, CGCDA02, KRCGC06, Kru01, LTD^{+}02, MH00, NOP02, TDC^{+}07, TD01b, WLJ^{+}04, WLJM05, WLB08, Wol07b]. stratified [MVLM^{+}06, PDG^{+}02, QRWJF07, SSLM01, TK08]. Streblospio [EWSH08]. Strength [Hil04, BM00, BBST07, GK03, HS01, SF07a, ZMN07]. strengths [DH07a]. Stress [KPW07, AWOO02, BOM^{+}08, CGP08, CSM^{+}09, CNPS05, CSvO09, HFD08, HBK^{+}05, IHW^{+}01, JRT^{+}06, LS08b, MGBB06, OP09, OHL^{+}06, Rij03, Rye02, SDHG03, SH07b, TN04, TPGQ07, Vie09, WDLM05, XWQQ05, vWCAB07]. stress-resistant [OP09]. stresses [PKS00, SLCB00]. stressful [TSC^{+}07]. stressors [KKSR05, LPK^{+}03, PZA09]. striatus [APP^{+}09, DE01, SSBZ07]. strings [HRR^{+}01]. striped [BM02d, Har00, MSK^{+}00b, NH03b, SRZZ01]. stripey [Kri02]. strobilation [PHS09]. stromatolite [PKR07]. Strombus [GL04b]. strong [MFE^{+}05, MFM^{+}09b, MFM09a, Som00, Tak02, VGW00]. strongly [Bib07, HSWG^{+}06]. Strongylocentrotus [Ada01, BJ08, BRS07, DHR04b, DHR06, ENAT07, KSIG08, LGS07, MH08c, MS01a, NA04b, SS05]. Strontium [EG05, FON03, IAKS03, FPBAT02, KSHP00, RFPB^{+}04, WT06]. Structural [BBE03, KT03, NK07, ZAZL08, BCL03, BDC00a, BA04, BDGK^{+}06, BVP^{+}06, GBNÖ08, HVH^{+}07, HFM^{+}02, KSWP04, PPS00, PPBS02, SH06b]. Structure [AO01, CS03a, Kor01, MSSP06, NTN08, RCDC06, SLGD00, VNdG^{+}06, WA03, AP05a, ABHE09, AMR00c, ABD^{+}03b, ABU01, AGDP^{+}06, ACJ^{+}04, APW^{+}04, AB06, BGC06, BER^{+}03, BSJ02, BLHB04, BUCC09, BSS05, BPS^{+}03, BH06, BJR04, BFE^{+}05, BHR08, CHSW03, CGH08, CW03a, CRV03, CMMN^{+}01, CF08, CCML07, CSM02, CIHF01, CS02b, CDAM06, CR08, CSMvE04, CB05b, CY06, CHA06, DW09, DFSS08, DDL^{+}03, DB05, DGJ^{+}03, DPSD08a, DPSD08b, DDSR07, DKFV05, DAG^{+}00, DKAH02, DWM04, Edm00, EHK06, EM04, ESB^{+}04, EGPT03, FTP^{+}09, FAJM07, FF02, GBdS^{+}05, GSMD00, GYD04, GL00a, GMK^{+}03, Gla07, GDGSP09, GGMGR07, GKAS08, GLG^{+}05, GG09, GGRC^{+}08, GBC^{+}04, GFPM01, HLD^{+}07, Har06a, HCB^{+}07, HWL00, HFM^{+}02, HF05, HNMC09, IG04, ISK07, JLW04, JBS08, JH01, JPPW01, JPPW02, JMM^{+}07, JKK^{+}06]. structure [KPM00, KTT04, KK02, KEB07, KOC^{+}07, KGHL^{+}01, KvdMDP09, Lap01a, LAK^{+}08, LHK01, LJM^{+}00, LPHH^{+}09, LCD^{+}05, LÅCN01, LSM^{+}02, LG08b, LS04, LAG08, MNR^{+}02, MCCV08, MHB^{+}01b, MBG^{+}03, MR04, MRB04, MHH^{+}05, MHTJ07, MTH^{+}08, MHM^{+}08, MPDC03, MB00a, MB01, MCD^{+}05, MP06, MBP^{+}06, MBR^{+}05, MRC^{+}08, MDSP05, MCS05, MRB09, MC01b, MARM06, MB05b, MPE03, MBL07, MMS^{+}05, MSTS^{+}06, MTLG09, NDV^{+}00b, NTGK09, NSEL00, NZM^{+}08b, NWHH^{+}09, NIF00, OBG^{+}07, OM04, PIY04, PZB^{+}05, PdNS06a, PDM^{+}03, PCD04, PLMS04, PCBS05, PWHW09, PDRT02, PWFC08, PKA06, RGMJM07, RBG01, RP06, RT00a, RT00b, RT00c, RBG^{+}07, RW08, RGCC^{+}08, RMAY04, RST04, SRL^{+}02, SAJW04, SCMB03, SKP^{+}07, Sha00, SW03, SDW07a, SGHG07, SMGCW00, SOMM06, SMHG09, SJ04, TNTH04, TSF01, THF^{+}01a, THF^{+}01b, TSHT02, TR06]. structure [TS06, TPD^{+}05, TA01a, TTL06, THBSJ05, Tyn04, Van02, Van03, VBV^{+}00, VFJ^{+}07, VMH^{+}08, WG09, WSA^{+}06, WAB^{+}02, WTF^{+}05, WM02a, WMS03b, WPC^{+}03, WHB^{+}06, WP07, WWV08, WBR07, Whi06, WDSA09, WKL01, WBC08, WAZ07, WMCF08, WHB09, WWWR09, WG07, WWWC02, WWWY07, YM06, ZMTB07, vDvTJD^{+}09]. structure-forming [DW09]. structured [AHSPG08, GTGM06, Hal04b, JY08, SW02b, SW02a, TP00b, VB02]. Structures [SMB^{+}03, AI04, BS08b, CCU07, EG05, FZT^{+}05, HC02, KM03a, KSH^{+}09, LH09, MAW01, MVK^{+}09, OKP06, RDG^{+}07, RGD^{+}08, Sma07]. structuring [BBC^{+}00, CFA07, GDW06, HCR01, JKES07, MJKH03, SAG03, VAB09, VBA06]. Stuart [WFHS08]. studied [KTSR02]. studies [AMR00b, CM04a, CCM01, CKO^{+}07, CT01, CMB^{+}05a, DSDS04, DESK03, EB08, FHM02, FS06, Fre03, GF04, GV01, HRB^{+}06, HTJ00, HBNA04, Hyr01, JMB^{+}05, MBA^{+}05, MS00a, MC01b, MMBF09, NNG03, NWIT08, NPTW05, OGMC06, Osb07, PBH^{+}03, PKH00, PB08b, PSGP04, RFCM03, RWBS03, SFW01, SYA02, SAV09b, SAV^{+}09c, STH^{+}06, TM02b, TBL^{+}09, WG02b]. study [AND05, AC07, AMR01b, BKP07, BDR06, BRAS03, BGD^{+}03, BC01b, BDÁM04, CLABRO^{+}03, DFPG^{+}03, DPD^{+}08, EJM01, EEA01, EÖM08, FMWGB08, Fin00, GWH03, GR07b, GT02, GDR^{+}04b, GK02, GBI^{+}06, HMBD01, Hen07, HJG06, JRP01, JP04, KGHW05, KJB^{+}06, LNJ07, LMS04, LJK01b, LJK02, LPBD00, LGM^{+}09, MGDP01, MYH07, MSY02, MGM05, MRG08, MSH^{+}03b, MJKH04, MGL^{+}00, MB00b, NMA00, OSP01, PFW^{+}02, PBPLB06, PvDMB08, PBJ^{+}05, PBL^{+}09, QR03, QMF05, RGR^{+}01, RGA^{+}02, RDH^{+}09, RB05b, SML02a, Sar06, SSGJ00, SAMN04, SO08, SHSV04, SSS^{+}05, SHS06, SPM03, SWGC^{+}05, Sto03, SMG00, SES01, TGY^{+}06, TBCB^{+}05, TSC^{+}07, TLC06, Tya09, TSI02, WNS^{+}05, WS04a, WASW03, WPTS07, YI07, ZMP^{+}03, dBKSV00]. Studying [JdSM09, MVdMFF08, QRS00]. stutchburyi [Mar04]. stylasterid [WCZ^{+}09]. style [WLS^{+}09b]. Stylet [DPSD08a, DPSD08b, LSP08]. stylifera [AKM09, CID^{+}06, GK08]. Stylocheilus [APP^{+}09]. Stylophora [AYSL09, EMB^{+}09, FPBAT02, FHGJB04, KJ03, MEB^{+}07, NYvW03, NKS03, RFPS^{+}02, SFPGA05, WBZ^{+}09, ZDA^{+}06]. Sub [BOZ03, SSN04, Eyr00, JCK03, KPKD06, KW00b, QS05, RT00b, RT00c, SHT^{+}04a, SWEBS00, TPGQ07, TDRM08, ZZT05, Chi09, GGS09, BSR09]. sub-Antarctic [KPKD06, Chi09, GGS09]. sub-arctic [QS05, ZZT05, BSR09]. Sub-grid-scale [BOZ03]. sub-lethal [TPGQ07]. Sub-populations [SSN04]. sub-tidal [RT00b, RT00c]. sub-tropical [Eyr00, JCK03, SHT^{+}04a, SWEBS00]. sub-zero [TDRM08]. subadult [DS09a, HMA06]. Subantarctic [GBG00a, TSSA03, AMR00b, AMR00c, AMR01b, BDG^{+}04, CDG04, LAG08, dBTO^{+}09]. subarctic [EBG^{+}00, KI01, LSM^{+}02, MB01, MKWO00, RYF01, YI03a, YI06, YYL^{+}00]. subcarinatus [FMP05]. Subcellular [NW05b, WLL03, WL03, CW05, RPS^{+}06b, RAAT^{+}07]. subcolony [ZPP09]. subdivision [SWM03]. subdivisions [RC09]. Suberites [SSG^{+}00]. subglobosum [HSGS00, SHS00]. subitaneous [WJWL05]. subjected [BSAR03]. Sublethal [BG04a, KLR^{+}06, FFR^{+}03, MLR02, SF02b]. sublittoral [GGM^{+}00, HCJ^{+}07, PCSS01, SS04b, SJ02, WJ06]. submesoscale [AFHA05]. submicron [FK00]. suboxic [MKH^{+}02]. subpopulation [AY04b]. subpopulations [MS01a]. subquadrata [HS05]. subsequent [DCT^{+}04, DER01, HTWQ05, HW03, JY08, MS09, PMT^{+}02, VMV^{+}09, ZT07]. subsidies [KH07a, LW06]. subsidize [Tal09]. subsidized [LG06]. Subsidy [DJM08]. subspecies [RGCL^{+}07]. substances [HKSW01, RMGL07, WS02]. Substantial [Gil02b, RGCC^{+}08]. Substituting [SJ06]. substrata [BJLH04, BCPB06, Gla01, GC01, RS05, SAC^{+}00]. Substrate [Tom08, Ver05, vMRO03, DJM^{+}05, EH00, FBT^{+}01, GBC^{+}04, MRS00, MW04a, SN04, Swa04]. Substratum [CD07b, Bib07, DHR04a, MBW03, ML05, MWTB00, PFZS^{+}06, SDM06, VBMBC08a, WH02]. substructuring [GPK^{+}04]. subsurface [FK04]. subtidal [Air00, Alt06, AIM^{+}05, AO01, BBB^{+}06, BS04a, Con05, CD07b, DMH01, DW09, DFT^{+}07, DGCM09, DHR06, EDS03, EHM^{+}09, FWC02, FBT^{+}01, GHJ02, GH08, GBC^{+}04, Her06, HC02, Hor06, IC02, KGRW06, LGSB09, LAW06, LTL^{+}06b, LLG^{+}06, MRB09, RS06, SKR09, SC03, SABK03, SW03, TK01, TBCB^{+}05, TAFBC07, WBC07, WWV08]. Subtle [FAK07]. subtorquata [NK03]. Subtropical [Fen08, LSK08, TPQ^{+}03, Bar02a, BPK^{+}00, BPS^{+}00, CLS00, CMJ00, CRK06, FPWM03, FMZvW02, GGR06, GAM02, HVWH06, IOB04, JS08b, KD00, KvNVH00, LDL^{+}04, LSS^{+}09, MdL06, MS01b, MFP04, MFP^{+}05, Niz07, PTCH01, PPJ^{+}05, QA01, SJ08, SE04, SCZ^{+}06, SWL07, SC06b, SKH06, TN04, TC01a, VHKW00, WKY09, WLJ08, ZS03]. Success [BDGK^{+}06, PPGL09, ABF07, AMR00b, BT01, BDC00a, Bar02c, BPB07, BTG^{+}08, BCB02, BDEM01, BR09, BT00, CAT^{+}06, CH09, CK09a, CCCD09, EC07b, FWH^{+}05, GSM05, GBW^{+}07, GCVR01, HC05, Hip08, HST^{+}09, HJM07, HFMP03, ITE^{+}04, KP00, KHG06, KHvW02, LBPS03, LGRB^{+}08, MRZJ05, Mar02, MMC^{+}07, MSY02, NHNB01, OSL07, PSKC04, PTB03, PTP01, Pra05, RWE04, RTL05, RCF09, RHSB08, SAWG05, SGCP06, SB06, TVD02, TOBR07, TIM^{+}01]. successful [DJV03, LC07a, LC08, MHP^{+}00, MRMA07, PSS06]. Succession [CHP^{+}05, NR00, CS03c, LNH00, LWMW02, RLDL02, TM02a, TMFF00, WDN00]. Successional [HFH09, DQW05, MHTJ07]. suction [AY03]. suction-cup-attached [AY03]. sudden [RC07]. suffer [BDJ09]. sufficiency [DS06a]. Sugar [PS07]. suggest [MND^{+}02]. suggests [GTGM06, HBG^{+}03, TGGM05]. suitability [CTH^{+}03, EMG01, KWTP06, Kup03, LGD04, MPM^{+}02]. suitable [BM07c]. Sula [Mon07, ZBDE07]. Sulaibikhat [AZKJAM06]. Sulawesi [dVCH^{+}06, ATR^{+}08, SAM^{+}01]. sulfamates [KKBH03]. sulfate [MDBC03, TSW^{+}09]. Sulfide [FHDA^{+}07, AHK^{+}02, CPV^{+}03, GB02a, ISG04, LLY^{+}01, LZM^{+}03, SRL^{+}02, TSW^{+}09, WK00]. sulfides [GBMHD08]. sulfur [FHDA^{+}07, KKIV01, LF08, LSdM^{+}03, PPI04, SSGJ00]. sulfur-oxidizing [PPI04]. sulfur-stable [LF08]. sulfuris [SV07]. Sulphide [TVG00, DSS04, DF06, KWTR08, NMVH06, VMdBV06]. sulphide-rich [VMdBV06]. sulphur [ISSK03, TVG00]. sulphur-oxidising [ISSK03]. Sumatra [ATR^{+}08]. Summer [CMD^{+}06, CWDG05, HLH^{+}07, LMKG01, RATD07, dSCV^{+}08, BBV^{+}01, BCVC07, BGH06, CS02b, FTK^{+}04, GGK^{+}07, GGC^{+}05, GFP^{+}05, GB07, GGG^{+}04, JMM^{+}09, JCL09, KET^{+}03, KTK^{+}02, KNF^{+}03, LSM^{+}02, LCB^{+}05, MABB03, MGPAE01, MPE03, MGH00, NRM05, PTK^{+}09, SAG08, SSGB^{+}08, SWEBS00, STM06, TRH^{+}06, TPQ^{+}03, Thi03, WSA^{+}06, YM06, YQC^{+}00, YQW^{+}01, Zam02, AY04a]. Summertime [BM03]. sun [BSKR^{+}06, Hel00]. sunfish [CL04]. sunscreens [ESDG08]. superabundance [DABC02]. superba [AVNN05a, AVNN05b, FHM02, FMWGB08, HRQ^{+}02, HL00, JT08, MRQM06, MABB03, MO05, QR03, SN02, STA^{+}05, TCRT^{+}07]. superciliosus [NMOO03]. superior [BWO04]. Supplemental [TKTJ07]. supplementary [LLY^{+}01]. supplies [HPS07, PHS^{+}07, VPSO01]. ’supply [DMM^{+}03, DSL07, DBOR09, DAG^{+}00, DRTM09, DLB04, FPB02, IOTL09, JF02, KHF^{+}00, KG00a, LSH^{+}07a, LG00a, LC09a, McI02, McI03, MW01, OTBR00, PSE^{+}03, QAA^{+}06, RHBW09, RTL05, Ros01, ST06b, SNY^{+}05, SSS06, SMF^{+}02, TSJ05, VD00, VVTH07, WGF09]. ’supply-side [DMM^{+}03]. Support [HD05, CHG05, FMZvW02, LJM^{+}00, PWC^{+}09, SDR01]. Suppression [RR06, RH08]. suprabenthic [CM01, PDT09]. supralittoral [DG05]. supratidal [CD07a]. surf [CHA06, RHG05, WCRB03, YSS03]. Surface [KTH^{+}04, LYL^{+}07, NGdN^{+}07, BAM^{+}04, CCC09, CPSC01, DR01, EDS03, EC07b, FDD^{+}03, FEMW07, GNL08, GBG00a, GGMGR07, GSFF^{+}09, GB06, HSCF06, Kel04, KW00a, KGB^{+}01, LHQ03, MP05, Orv05, PSKC04, QRS00, RS03, Rit06, SG05, SHS^{+}05, SH06b, TWYS03, TBB^{+}04, UYNO01, VPWE09, WB01, WKH02, WWWC02, YVB^{+}02, ZHS^{+}05, dBS01]. surface-associated [KGB^{+}01, LHQ03]. surface-canopy [EDS03]. surface-feeding [SG05]. surfaces [BKSS07, MD08a, VK02, YMHV02]. surfactants [AHL^{+}01, SMB^{+}03]. surfgrass [GA08a, HIM05, Hor06, RHBW09]. Surficial [BWP^{+}06, KTF^{+}01, SPRK02]. Surfing [PCC07]. surfperch [OAM06]. surgeonfish [RAC^{+}05, RCP^{+}05]. surmuletus [LRGdS00]. surplus [SKH06]. Surrogacy [WKK07]. surrogate [LPS08]. surrogates [MTLG09]. surrounding [LEF^{+}06, LBBM06, MH08a, TKR06]. survey [GDDM05, HPC04, Hyr01, IF09]. surveys [BS06, DGVK^{+}04, LBR^{+}09, NCN^{+}09, WWTH06, YM06, ZIUG07]. Survival [AM01, BBC03, BSAR03, BWW^{+}06, Bra05, ISG04, RBHS06, Rit02, AH07, AM02b, ANF^{+}00, BPH07, BMG02, BGPL08, BvK07, BF08b, BDRM06, BM07a, BY09, BHV^{+}02, Bul06, CHMR05, CID^{+}06, DUW08, DS01, DMV07, DESK01, DHT^{+}08, FHM02, FKSW09, FSB07, FMP05, GRIR06, GQ00, GJ03, GCS06, HANM^{+}05, HF05, KB03, LGBWH06, LGVF00, MK03a, MFA03, MAWH03, MRH^{+}08, MMBF09, NvW01, OAM02, ONKK03, OvM02, PH03b, PDHH06, PS06, RGR^{+}01, RH09, RGL^{+}00, RBSP00, Ros01, RT05, SML02a, SBMAE06, SZ08b, SD01a, SF02a, ST06b, SMP^{+}09, SD00a, SLB07, SKK^{+}06, SJH^{+}04, TAM04, TNTH04, Thi05, THC^{+}05, VFM06, VJKF07, VCBH06, Wal07, WM02b, ZMN07, ZSKA01, dZCVC01]. survive [SPM03]. survivorship [BLT03, BBT02, DJM08, EBC04, RM05, RJH02, SJ09b, SLEH01, THYBS09]. susceptibility [AS06b, Bye00, DMC09, MAG^{+}07a, MT09, RdNS02, YH04]. suspended [APW^{+}04, BMSD00, CWMD06, CMD07, CBMB03, DÁSG^{+}01, FD00, HL05, JKNJ07, KPKD06, LGQ^{+}00, MGDP01, Maz04, PS06, WGD05]. suspension [BDG^{+}07, CVK00, CTHF01, DBR^{+}07, DOML^{+}07, GSM^{+}08, LHB09, LWH08, MCS09, NTGK09, Orv05, Pen00, PH01a, PH01b, RNL00, RL05a, RKLK07, RGMH08, SWH04a, STJ03, WPG08, YLK00]. suspension- [STJ03]. suspension-feeder [Orv05]. suspension-feeders [DBR^{+}07]. suspension-feeding [CVK00, CTHF01, DOML^{+}07, NTGK09, Pen00, PH01b, RL05a, SWH04a, WPG08]. sustain [BD06, CMRF05, GQ00, MCU06]. sustainable [HS02, OW02]. Svalbard [AP05b, BJN^{+}05, BJN^{+}06, FPDS^{+}02, JGH05, KKL07, LMKG01, MAJ^{+}06, VNJ06, VFJ06]. Svedäng [BL07]. SW [AQSP07, FAPA^{+}09, FAPOF09, KFS^{+}09, ME01, SHM^{+}01, VLP^{+}04, WBH^{+}04, dVCH^{+}06, APA07, AED^{+}07, Bar02a, BCCM03, BVI^{+}05, BIG^{+}06, BHD04, COK02, DT06, EL05, FC04, GWPRC^{+}07, GMD08, JC08, MVB^{+}05, PMPI03, PCHM07, RSK04, RGCC^{+}08, SFF^{+}07, TL05, VMRL04]. swamps [ROCB09]. swarming [JR01]. swarms [GGR06, KRF^{+}05]. swash [DGL01, ILL06]. Sweden [HLVU00, QMM02]. Swedish [BAB04, GSRH02, JNMB08, JNFL04, LÅCN01, McQ05, PABM08, RAH^{+}02, VTS^{+}02, WL06]. Sweeper [LWCF04, LC06]. sweetlips [WW03]. swim [KNF^{+}03, TBH08]. Swimming [CLHT05, Fis05, JT05, TJSG03, Tit01, Arm00, BGC06, BF01b, BJJ01, CCK03, FOFG09, FBJ00a, FB03a, GYD04, GBD^{+}06, HHD^{+}09, HSCH07, HMO^{+}04, HCMR09, JT08, KNKT01, LP01a, LWJC04, LHL^{+}07, LDT07, LJC00, NSH^{+}08, PBH^{+}06, RWJ02, SHT^{+}04a, SV08, SR03, TND^{+}05, WLS^{+}09b, WZ04, YA09]. switched [MR06b]. switching [AS06a, IMTW09, JMJ07, MSF05, OKAM00, RGL06]. swordfish [WKY09]. Sydney [Bay02, BML^{+}08]. sydowii [Ryp08]. Syllidae [MG05b]. Sylt [BAA04, BAA07]. Symbiodinium [ACT09, GSG08, KCA^{+}08, LBH^{+}04, LWB09, LCW06, RLKW04, STRD02, SGV^{+}02, SPCG09, UHvO^{+}08, YM09]. Symbiont [DF06, BWPB04, BSVK05, ET08, LTL06c, SKW^{+}06]. symbiont-produced [LTL06c]. symbionts [BWPB04, BP04, Bur02, CBM00, Gou06, LCW06, OP09, PR01, Pra01, SdV06, SGL^{+}02b, YBY^{+}09]. symbioses [BCLRL07, ET08]. Symbiosis [DRDD09, KCB06a, CSG01, DC07a, FKM^{+}01, GC03, MC05b, PBAA04, RCC^{+}04, RLKW04]. Symbiotic [JLJ01, AC02, BR07, CS03b, FH08c, FHGJB04, Gou07, HBC09, HGMG^{+}04, LFRR^{+}07, LLCHG01, STRD02, SGV^{+}02, SL05a, Woo09, YH04, YM09]. symmetricus [HK04b]. symmetry [MG08]. sympagic [BGS02a, Pol00]. sympatric [BLD^{+}05, BCM^{+}07, BBC04, CBDH02, CS03b, CV05, DBR^{+}07, HFA02, LAG08, MB04, PMG05a, PSC05, WMD^{+}07]. Symphodus [RM05]. symposium [GBB^{+}01]. synagris [LDR00]. Synchaeta [OTB^{+}00]. Synchelidium [YSS03]. Synchronized [BDEM01, OKH^{+}07, vTWMG08]. Synchronous [JR01, LTPL05]. synchrony [BG09, Bla09, BMJ02, CHW06, CMS^{+}01, HDG^{+}08, KFL09, LTNC07, Mar02]. Synechococcus [ADA00, JWTP06, PBS^{+}05]. synergistic [Con05, SMC^{+}05, TAM04]. synoptic [HMD^{+}07, PTS00]. synthesis [Boy08, BRB^{+}00, CVSM09, Fau09, GSS09b, LHGC^{+}09]. Syringodium [HKWM07, SZ08a, SZ08b]. System [FGLM06, KBB09, KB07, MG09b, MHABGC04, PBN^{+}07, ÁSG04, AFFH04, Rii03, AVE^{+}06, BMG00, BB03b, BOZ03, CB05a, CG05, CMB^{+}05b, CTHF01, DE01, DP02, DRDD09, DRF08, DHRG05, ECK05, FPB^{+}07, FHK^{+}08, GÁSPR03, GSL^{+}06, GB03, Hal04b, HM05b, HQDH06, HMD^{+}07, HLH^{+}08, HGZ^{+}00, JMU09, KPKD06, KFE^{+}03, KOC^{+}07, LS08a, MHCB08, MADPQ06, MPR02, MVLM^{+}06, MASFS00, NCÁSBP04, NCÁSGP05, NAB09, NL05, ORD00, OÁSG^{+}08, PABM08, PRvdL^{+}07, PRL^{+}09, PNCS^{+}08, PNCT^{+}08, PBJ^{+}05, QMM02, QS05, RVMI^{+}03, RQF^{+}00, RGC^{+}06, RMPC08, SM01a, SPBM05, SHZ^{+}05, SH05, SSS06, SJL05, SGCC06, SLHT05, SLEH01, TS07, TFLA03, TSMC09, WKL^{+}03, WDSA09, WS00, XZLP09, dBTO^{+}09, dSSC07, DDQ^{+}00, HM09, IG04, PGR^{+}01, RCL^{+}09, TDC^{+}03]. system-scale [HLH^{+}08]. systems [ADK^{+}06, AAJ07, BDF^{+}04, BLS04, BWRP06, CTvdM^{+}08, CTD^{+}09, DPH02, EYH^{+}06, EP00b, EKBM05, FKBB01, GL04a, GRP^{+}06, GDL^{+}01, Hen07, KKL^{+}03, MMCF07, Nae06, NZ08, Niz07, OW02, PTB03, RWE04, She05, SMP01, TdMF^{+}08, THF^{+}01a, THF^{+}01b, TLTB03, VWF^{+}01, WT09, WGS^{+}02].

T [CHT06, HCP^{+}06, UYNO01, SK02]. T-PODs [CHT06]. T-RFLP [UYNO01]. t-SPI [SK02]. T. [GK08, TL05]. Tactics [DT06, BLDG01, BTB^{+}02, MBG^{+}09, WBS^{+}04]. taeniatus [GB05a]. tag [AY03, ACJ^{+}04, GNL08, LLH06, MWTB00]. tag-recapture [ACJ^{+}04]. tag-return [LLH06]. TAG/DNA [MWTB00]. tagged [DNL08a]. tagging [BCB08a, EOK^{+}08, GMB^{+}06, GLÅ^{+}03, PBH^{+}03, SAS^{+}06, SSR^{+}03]. tags [Ber00, FWD^{+}01, Gri02, HJG03, HNMC09, JdSM09, LS06, LALB08, NS02, RHH^{+}05, SSW08, TBB^{+}04, Zac05a, Zac05b]. Tagus [CB00a]. tail [NHS^{+}08]. tailed [BLH01, BHPC06, CRLP09, CKS05, KS00a, LCL^{+}03, PSKC04, PC05, WRW06, ZS03]. tailings [LMC08]. Taiwan [HT07, HCH09, OHL^{+}06, TL05]. Taiwanese [SICT03]. Takapoto [LPBD00, SLGD00]. take [DWB04, LH08, LHGC^{+}09, LTCP03, RSA05, SB03a, SJA^{+}09]. Takifugu [KKMH08]. taking [SMP^{+}09]. Talcahuano [LC02]. Talitridae [PR01]. Talorchestia [PR01]. TAM [SW04, WL03]. tamarense [CCM01, FLR05, ITE^{+}04, KS06, SSO^{+}01]. tang [CKB^{+}09]. tank [GWH03]. tanker [HCP^{+}06, SRDPT^{+}07]. Tanner [TMAN08]. Tanzanian [KFU^{+}08]. Tapes [MM05, MDG04, SPC^{+}00]. taraikaensis [KS06]. target [ASA^{+}08]. targeted [WAK^{+}09]. targets [MPDC03, PKM03]. Tarletonbeania [SK05]. Tarling [Pea03]. tarpon [CT06a, LALB08, SCIT09]. Tasman [CWBS03]. Tasmania [HGB01, KGHL^{+}01]. tasmaniae [JR01]. Tasmanian [PBFB08, STF^{+}09]. tasmanica [MCS05]. taurus [BB09]. tautog [KMAC^{+}02]. Tautoga [KMAC^{+}02]. Tautogolabrus [NMJ02]. tauvina [KAR04]. taxa [ACLF09, BDR06, DHG^{+}04, GC02, HB03, HD06b, Lid08, MMMPB05, WTS09, WKK07, vWNYS05]. taxa-specific [DHG^{+}04, HD06b]. taxifolia [CMJ00, JCMR^{+}03, MGD^{+}09, MS03, PP02, WG02a, WS04b, YBGP06]. Taxon [GVM^{+}05, vLvSGS05, Sta08]. Taxon-specific [GVM^{+}05, vLvSGS05]. Taxonomic [DS06a, MTLG09, ZM06, Bat09, CW01, CC07, CBEL06, Hir06, IP01, JS08b, KH00a, Las03, NNG03, RC08, SLM^{+}02, SDH01, TDSM02, WHL^{+}06, WKK07, LBH01]. taxonomic- [JS08b]. taxonomy [RLFEF03]. Tay [SAH09]. taylori [SK05]. TBT [BMG00, RGEB00, SJPC02a]. TBT-based [RGEB00]. TDR [AY03]. technique [BRJ^{+}03, GT08, GDdS^{+}07, HCBP05, MVdMFF08, MLR02, ORPND01, SH06a, SB00, UB02]. Techniques [GZ02, RFB^{+}06, BBS05, CF02, CSD04, FE07, FAFR03, KDAK01, MBH^{+}09]. technology [Bra05, EB05, Osb07, SN09]. technology-problems [Bra05]. Ted [Cai09]. Tees [WAB^{+}02, WAB^{+}02]. teeth [HWB^{+}09, NEK^{+}07, NEMF09]. teleconnections [RHHH09]. telemetry [COSM03, HPW^{+}00, HTM^{+}02, HBG^{+}03, HALL05, JMJ07, MDBG07, PTSJ06, RDS^{+}09, SAG08, SMH09, SJL05]. teleost [CDAM06, Eva08, MFJT08]. teleostean [SH04]. Teleostei [CRC02]. television [SMA^{+}03]. tell [Arm01]. Tellina [KSS^{+}05]. tellinid [BDH01, BDJ09, DB07a]. Telmessus [KMF02]. temmincki [OAM06]. Temora [BGS^{+}02b, CID^{+}06, CA06, GK08, JVJ09, MRZJ05, OBB00, PK05b, SV08]. temperate [AFHS08, ABHE09, AMT03a, AA04, ACLF09, BFK05a, BFK06, BRSM07, BP03, BK06b, BAE^{+}08, CH00a, CLHT05, CBE04, CRBE04, CRCV04, CVSM09, CHV^{+}09, DLT^{+}06, DC07a, EF02, FD08, FBBV09, FA02, FNB07, FR06, FWC02, GB05a, GDW06, GBG^{+}00b, GF02, HCV03a, HBdBV01, HJ05, HAD07, Jen05, JLJ01, KGTH01, KR01, KG00a, Lev08, LH02, LUAIH03, LUIAH03, LVA07, LTCP03, LS04, MCCT07a, MCCT07b, MAM^{+}01, MS06, MAT01, MD08b, NLB08, NMS^{+}02, NMJ02, PDS^{+}07, PTCH01, PBH^{+}06, PTAR07, RM05, RCR03, RF08b, RPO00, RDC06, RCDC06, RMRAFP08, Ros01, SICT03, SF02a, SS09c, SBH^{+}02, SH07d, SF02b, SWEBS00, SKO^{+}08, TB08, TPG09, TBID09, TTR05b, Trn02, VLRJ07, WAZ07, Zam03b]. Temperature [BS08a, BGK^{+}06, BPS^{+}03, CRSL05, GBGH02, LQRGT06, MCR03, MT05, PWM^{+}08, PA07, SM07b, SG09, SEM01, SJS^{+}01, SGCP06, SMT08, TYC00, VTY06, WKH07, WCM08, AGS^{+}09, AJ03, BFK05a, BFK06, BJN^{+}05, BVH02, BHH09, BTW08, BCLS06, BCC08, BH04, CSB03, Cal05, CW03b, CMJ00, CPSC01, EKN08, Edm04, EC07b, FPC05, FHL^{+}09, FBBV09, FTB09, FEMW07, FWD^{+}01, GNL08, GRWB02, HDP02, HLD^{+}07, Har08, HFA09, IF09, JVT01, JBHGB04, KYML02, KHH08, KNK^{+}00, KNKT01, KR05, KG00a, Lam00b, LM09, LSH07b, LRN^{+}09, LC03, LP01b, LRC06, LZW^{+}08, LSG05b, Mad02, MW06, MMMPB05, MAOP09, MCM^{+}03, MDG04, MDW^{+}09, MR07c, MCN09, NRC07, NPVW06, NH05, ODG^{+}04, PBCB03, PSKC04, PPI04, PHS09, RPN^{+}01, RFPB^{+}04, RCF09, RFG^{+}09, SDHG03, SH07a, SC01, SWH04a, SWW02, SBM02, SW09, Ste04b]. temperature [SWHSA05, SP00, SM01b, SHW09, TSE^{+}08, TOK08, TC03a, TBB^{+}04, TBM^{+}08, WCBD01, WK00, WS01, WS04a, WG02b, WS02, YLK00, vWNYS05]. temperature-dependence [MDG04]. Temperature-dependent [GBGH02, LQRGT06, PA07, WCM08, BHH09, SWHSA05]. Temperature-induced [BGK^{+}06, NH05]. Temperature-mediated [SG09, SGCP06, Har08]. temperature-related [JBHGB04]. temperature-size [IF09]. temperatures [ATLC03, CCCD09, GBG00a, SO01a, TDRM08, WWDH09, YBY^{+}09, YG09]. Temporal [ABU01, BS00, Bar06, BRP00, BMCH09, CRK06, DDS^{+}01, DCAL^{+}02, DC07a, EBG^{+}00, FA02, FCP02a, GSW01b, GK07, GWPRC^{+}07, HNYI07, HDP^{+}06, HSCF06, HPB^{+}04, JM01, Jar00, JCL09, LJ08, LBBM06, LC09a, LAG08, MBW02, MMTT00, MGA^{+}01, MMC^{+}08b, NAB^{+}06, NMDS00, NAB09, OHS^{+}07, Pee02, PJN03, PPL^{+}02, RYF01, RMRAFP08, RHS01b, SV07, SMG02, VKH08, VBFF08, War05, WH08, YW04, ZLAK04, Air00, AQSP07, BGK^{+}06, BVMBC07, BWP^{+}06, BDÁM04, BTF^{+}06, CCMK03, CA06, CIHF01, CWM^{+}09, CB05b, DE01, DC06, DSJBS^{+}02, FC08, FCP^{+}07, FHVMDS07, GLS02, GDF03, GAM02, HJG03, HGRB04, HJK01, HVW^{+}07, HWG06, HFMP03, HMFC07, HW01, JMSM08, KL03, LBvdH03, LV02, LLW09, LBHL^{+}09, MNR^{+}02, MGL^{+}07, MJS09, MKL^{+}03, McI03, MW08, MBR^{+}05, MD08b, MTLT02, MSP^{+}06, MMS^{+}05, NE01]. temporal [NW06, PZB^{+}05, PSPZ09, PTM08, PEDT03, PDRT02, PMR06, PDG^{+}02, PA07, PKA06, PKS07, PR05, RH08, SM01a, SNH04, SSM05, SP01a, ST08, SA06, SDW07b, TS07, THS^{+}02, TSSQ06, TG03, TPD^{+}05, VHK09, VCSL^{+}09, VBMBC08b, VMH04, VHM09, WW06a, WISH09, WBSL00, WKN^{+}03, WBZ^{+}09, YH02, Zac05a, Zac05b, AVNN05b, MDBC03]. temporarily [KPN03, NPM01]. temporarily-open [NPM01]. ’temporary [GL02, BPvdM07, BWW^{+}06, RL09a, SMP^{+}09]. ten [HPC04]. Tenagomysis [JR01]. tension [CGT^{+}00]. tentacle [LC06]. tentacles [LWCF04]. Tentorium [QASA08]. Tenualosa [MC01b]. tenue [JM03a]. tenuicostata [RJH02]. tenuifolia [RPS^{+}04]. tenuiremis [WJWL05]. tenuirostris [CMB^{+}05a]. tenuis [HNYI07, HDW04, NKS03]. TEP [BPVL03, BPE^{+}06, EP01, Koe05, LWH08, MR04, MRB04, MWMH05, Pas00, Pas02, RYF01, Tho04]. Terapon [LW05]. Terebratalia [Tom08]. term [BBG^{+}04, BPM^{+}06, BIR00, BI04, BMF06, BOR08, CJRT00, CR08, CS03c, CCB^{+}09b, Das03, DB07a, DSJBS^{+}02, DL01, FMCB05, FGM^{+}00, FAB^{+}07, GÁSPR03, GYLÖ06, GB01, GLS^{+}00, GSM^{+}02, GTPC04, GBG^{+}00b, HWC^{+}06, Joh02, KLR^{+}06, KR04, KSS^{+}05, KST^{+}08, LBS^{+}05, LC06, Lin04, LPM^{+}06, MMC08a, MG09a, MS08, MYH07, MO05, MCP^{+}09, MOPP05, MMB^{+}02, ME01, MK02, NMCT03, NB06b, PB07, Pea00, PD01, Reb01, RBSP00, RF06, RMZC07, SSL00b, SdSC01, SN02, SMSD07, SWG^{+}09, SH06c, SIK^{+}02, TAM04, TM02b, TBID09, TPS^{+}05, VML05, WK00, WdBS02, WPTS07, YF04, YVB^{+}02, YMC05, dBS01]. terminal [UYNO01]. terminal-restriction [UYNO01]. terminates [KYYM00]. termination [SHH09]. terminology [WA00]. terms [ONA04]. tern [EC07b, RMA^{+}02]. terns [MRQ07, MRH^{+}08, NMBH02]. terracing [RM09b]. terrain [DDGK^{+}08]. Terrestrial [CTH^{+}03, BJ08, CH03, CPG^{+}09, CT04, DSPHV04, SB05b, ZT07, vDSD02]. Terrestrially [LTH^{+}04]. Terrigenous [CVT09, ARFP09, ENC^{+}04, LTH^{+}04, LHT06, LTL^{+}06b, NTH^{+}02, SAJW04]. territorial [HSTS^{+}09, HK03a]. territoriality [CHM06]. test [CD03c, DAG^{+}00, DTAF00, FSSW^{+}00, HAA00, HGMH06, JRC06, KCS04, KS00b, LBF^{+}02, LTFH06, MW04a, PH01a, SR09, SB05a, SMF^{+}02, TD01a]. tested [NRvdV^{+}02]. Testing [CHW06, DGL01, FDDS09, Jon04, LMW04, RL09a, TWT08, MAWH03, NPLZ00, RFSC08, SAHF03, SMHL00, TDC^{+}07, TP01b]. tests [BBM^{+}02a, CM04b, HDGM01, Kev01, LW04, NGdN^{+}07, PdNS06b]. testudinum [BB04, EMD^{+}01, Enr05, GHV05, HKWM07, IOB04, KD00, KVH02, LD00, RZEACGF00, SZBZ04, UH03, VHKW00, WKD^{+}04, vTWMG08]. tethering [AHV01, BS07a, BH02, KS00b, MJG08, PTCH01]. tetrabrominated [MAE03]. Tetraclita [CMD^{+}06, CML08]. Tetrapygus [Rod03].

Teuthoidea [CD03b]. texanus [BBSO03, RAA02]. Texas [RZ00, RMZC07, AG02b, RMKR00]. Thai [TVD02]. Thailand [ITN09]. Thais [BW04]. Thalassarche [CPP^{+}04, HGB01, NLP^{+}01]. Thalassia [BB04, EMD^{+}01, Enr05, GHV05, HKWM07, IOB04, KD00, KVH02, LVRN02, LD00, RZEACGF00, SZBZ04, UH03, VHKW00, WKD^{+}04, vTWMG08]. Thalassinidae [NPVW06]. Thalassinidea [CCdAR00]. thalassinidean [AND05, WE04]. Thalassiosira [BVH02, LELC02, SP00]. Thalassodendron [AM02a]. Thalassoma [GCS06, SGCP06]. Thaleichthys [MWG02]. Thalia [TL05]. Thaliacea [PMNL07, TL05]. thallus [BM05a, JS02]. Thames [TNSM00a, MN00a, PA07]. Thau [DVLH^{+}00, LABR^{+}06, Maz04, SVC^{+}01]. their [AFHS09, ARBF02, AHW03, BF01a, BCC^{+}03, Ber02a, Ber02b, BKLGN05, BSR09, BDGK^{+}06, BHD04, CH06c, CPB04b, CH07c, DH07a, DCM^{+}06, DGCM09, Dun02, EOL^{+}05, EB04b, EP01, FNB07, GSM05, GCLC^{+}02, Gil02a, GY00, Gou06, GNAD^{+}03, GBI^{+}06, HJG03, HMKR05, HBG^{+}04, HBRJ06, HMY^{+}08, HGC^{+}07, HMK06, HGMG^{+}04, HKSW01, KES00, KHF^{+}00, KVV^{+}00, LBPS03, LALB08, MGDP01, MMMB^{+}05, NATR^{+}05, ONKK03, OPSM04, OGL09b, PPCU07, PLM03, PMP04, PD01, PUL07, QMF05, RFB07, RHH^{+}05, SGV^{+}02, SC03, Sim00, SvNvdMvdV09, SDRH01, Sta08, SBLF09, SAM^{+}06, TBBI04, Tor09, TNSM00b, UP01, VHW^{+}07, WWGR07, WW05, WR07, WGS^{+}02, Wil03, WWWR09]. them [vDS05]. Theme [Rii03, Ano03h, BHC^{+}00, BS05b, BVHG08b, GBB^{+}01, GSS09c, HR02c, Mar01, PAC^{+}00, RQF^{+}00, Rii04b, SvAHH09b, SvAHH09a, SAV^{+}09c, SB05b, Syd09, WB05b]. Themisto [BCL^{+}01, MHCB08]. Theora [YI03b]. theoretic [GDF03]. theoretical [MS01a, PM00, TNHL02]. theory [CCL^{+}08, JH04, Lin02, VPG04]. Theragra [BBC^{+}09a, Bai00, DABC02, PCH^{+}05, WJDA06, YHSH02]. there [CSP^{+}03, EB04b, Joh02, PFP03, Rye02, She01]. Thermal [KMWH04, SWL^{+}06, BTL05, Ber02a, Ber02b, CPH^{+}02, CSvO09, DD05, DD07, DVBLS07, Dov04, Dun02, KS09, RL09b, SBA01]. thermally [HSSS05]. thermo [GGMGR07]. thermo-haline [GGMGR07]. thermocline [GK03, RS07b, SBA01]. Thermoconservation [KNKT01]. thermophilic [RL09b]. thermophily [NF06]. Thermotolerance [CFC^{+}00, BDDG02]. thermydron [PEDT03]. theta [MKWO00, SK05]. thick [FBD^{+}02]. thick-billed [FBD^{+}02]. Thin [RDD^{+}02, ACM^{+}02, CBDH02, CMSS09, DDS^{+}01, Lei01, LTH^{+}04, MAB^{+}03, MCD^{+}05, QMM^{+}08, RMPC08, WG09]. thin-billed [CBDH02, QMM^{+}08]. thinning [NGBG07]. thiols [AWOO02]. Thioploca [SDW^{+}01, SSGJ00]. thompsoni [SK05]. Thomsen [TW10]. Thoracica [DP09]. thorns [GHdN07, LRM00, Pra01]. thornyhead [RM09a]. thousands [HH07a]. thread [CEB^{+}08, CYC^{+}09]. Threat [PCC01, HCS^{+}07a, KB06]. threatened [AO02, DSH03, VHM09]. Three [GHR^{+}07, SDW07a, TAM04, BR03, BLS03, BCPB06, CAMC07, CSPA09, CMGBJ05, DLT^{+}06, DTV09, DESK03, EJM01, ERR06, FAK07, FBHB03, FTU^{+}00, GSMD00, GGFSP06, GDS02, GBD^{+}06, GFPM01, HHWT02, JHJ^{+}03, JHJS04, JÅC^{+}01, JSMM06, KNTK05, KCEB09, KKSR05, LR06, LGVF00, LFSR06, LBC^{+}09, LMB^{+}09, MMV05, MvdBÓ00, MKWO00, MB00b, NLG05, NRP^{+}02, NLCS05, PGM05, PD01, RB02, RB05a, RWR^{+}00, Saf03, SLL02, SKR08, SMC09, SP03b, SFT^{+}05, SCS^{+}07, SB08, TN07, TMFB06, Vii07, VPS08, YLL^{+}04, YSS03, ZIUG07]. Three-dimensional [GHR^{+}07, SDW07a, GBD^{+}06, HHWT02]. three-spined [JHJ^{+}03, JHJS04]. thresher [NMOO03]. Threshold [IS07, SW08a, Enb05, MSK^{+}00b, Sim00]. thresholds [Ber02a, Ber02b, Dun02, GP04a]. Thriving [RCL^{+}09, VMdBV06]. through-flow [KPKD06]. throughout [CDS^{+}04, DL03a, DTB07, FPB02, MVH^{+}06, TBSJ^{+}04]. throughput [Dis07, LUA99]. Thrush [Ano02h]. Thunnus [ACJ^{+}04, KNK^{+}00, KNKT01, LFF^{+}09, SDD^{+}07, TSIY06]. Thyasira [DSS04]. thyasirid [FKM^{+}01]. thyasirids [DF06]. Thyca [KSH^{+}09]. thynnus [KNK^{+}00, KNKT01, SDD^{+}07]. Thyroid [EDB^{+}04, SH04]. thyroidal [CCCH01]. Thysanoessa [AS06a, STPA04]. Tiahura [AAGS02]. Tiarinia [HTS06]. tick [MRH^{+}08]. Tidal [BVVM02, BHN09, CWBR05, EB08, FNB07, GGR06, Har08, JC08, KKA^{+}06, QFJ^{+}07, SHM^{+}01, VMV^{+}09, Zam02, AI02, AL03b, AF06, BBBSPH03, BCC^{+}03, BdVD^{+}01, BSvdH^{+}02, BSED08, COK02, CWR06, CTD^{+}09, CS05, CBM^{+}07, CM05, CAA00, DAM^{+}08, DB07a, FDRR05, GBB^{+}05, GPHC09, GCD^{+}03, HSWG^{+}06, HH03a, IRM08, KT06, KWHD08, KWKF03, Kne09, KCB06b, KWL^{+}05, LDT07, LWB^{+}00, LBH01, MMTT00, MCC^{+}00, ML05, MTLT02, MN00b, MS00b, MLM^{+}02, MPB^{+}09, NMA00, OTBR00, RT00b, RT00c, RAS07, SN06, SR08a, SR09, SAC^{+}00, Seu05, SAS^{+}06, SDW07a, SVG^{+}05, TAW^{+}09, VMV^{+}08, VBS06, WK02, WGR07, WHL^{+}06, WBSL00, WBH^{+}04, WAZ07, Zam03a]. tidal-current [CWR06]. tidal-flat [KWKF03]. tidally [BFB01, EB08, JKNJ07, JTR05, SGL02a]. Tide [QAA^{+}06, BMS04, BM07b, CCL^{+}04, CSL^{+}05, DCAL^{+}02, FA06, GSW01a, GSW01b, GLCL04, HLR07, HDCA06, JTLG09, KJ04, LPK^{+}07, MSC04, SLT^{+}01, SJK^{+}06, TGSW06, TDP^{+}00]. tidepool [Dav00a, Zam03b]. tidepools [KCSPT09, Zhu06]. tides [CBM^{+}07, GGMGR07, MB08, YH00, Yin03]. Tierra [SR04a]. tiger [HFW^{+}05, MCP^{+}09]. tight [HTJ00]. Tigriopus [DHCH04, Joh01]. Tilstone [Ano03f]. Time [Ber02b, BKH^{+}09, Dun02, KKP08, ATT06b, AM01, Rii03, BBHT02, BCC^{+}03, BLF06, BP09, BHZW08, CSTC01, DJM07, DHG^{+}04, DBS05, EG05, GC03, GWF07, GPM^{+}09, HSGS00, HBN01, HLU06, JSWH03, KH08b, KH00c, KN00, KJRG09, MM03, MS04a, MS05, MASFS00, OKP06, RMM04, RTL05, RQF^{+}00, Rit02, RFSC08, SJM06, SC00a, STF^{+}09, SFH00b, SK02, SWH^{+}04b, SLCC09, STJ03, VCSL^{+}09, Sha09b, Ber02a]. Time-integrated [Ber02b, Dun02, Ber02a]. time-lapse [OKP06, SK02, SWH^{+}04b]. time-scale [DBS05]. Time-series [KKP08, GPM^{+}09, MS05, STF^{+}09, Sha09b]. time-space [KH00c]. times [AP00a, HGSH08, Pör08, Vie09, XYH^{+}09]. Timing [CCCH01, HSHA06, MMV05, MW02b, PSS06, SPH^{+}09, VYM08, BCB02, CSP^{+}03, Fis06, HW09, HM05b, KJMP05, LMMT09, RMA^{+}02, SHH09, SSL00a, VHMB09, WPTS08, tAHSHA07]. tintinnid [KS06]. Tisbe [BBM^{+}02a]. Tissue [Cum02, TBFA04, Eva08, FBKB03, GNAD^{+}03, MW02a, Rob06, SMM^{+}04, SFR^{+}09, TBBI04, YTH^{+}05, YIY08]. tissue-specific [SMM^{+}04]. tissues [BTD^{+}04, DL03a, HKSW01, KW02b, MSF05, RFPS^{+}02, SJE^{+}06]. Title [Ano00a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b]. Tobago [MH08a]. Todarodes [IAKS03, KS04a]. Todaropsis [DGC^{+}05a]. together [DSJT^{+}06]. Tolerance [AHK^{+}02, Alt06, ATLC03, BSW^{+}09, BHY09, DD07, Kaw01, KMWH04, NMVH06, PKS00, QTB02, TVG00, VGB^{+}05, VTY06]. tolerances [TYC00]. tolerant [Lam00b, RGJ^{+}04]. Tomales [HB08b]. tomcod [GC07, GC07]. tomentosoides [SG06b, SS05]. tomography [RDG^{+}07, RGD^{+}08]. tonsa [BK05b, BD02, CAT^{+}06, HK03b, ISG04, JTLG09, JFA02, OSA05, PK05b, PI07a, SSB09, TKTJ07, WTP08]. Tonsina [HCP^{+}06]. Tool [CG05, ABS^{+}09, CH07c, CBD^{+}05, DHZ02, LGM^{+}09, MS03, OBB00, TBR^{+}03, VCA^{+}09]. toothcarp [Mal02]. toothed [MSLP06]. toothfish [GHT^{+}01]. Top [BS07c, GT02, KH07b, ZSKA01, HBBD09, HHM^{+}06, JFD09, MHHS06, PRL^{+}09, PCRR08, Som00, Syd09, Tor09, VBFF08, WDT^{+}08b, WLJ^{+}04, WFT^{+}08, Wie05, WLJ08]. Top-down [BS07c, GT02, KH07b, ZSKA01, HBBD09, HHM^{+}06, JFD09, MHHS06, Som00, Wie05, WLJ08]. Topical [SZ09]. topics [AC08]. topographic [GBR01, Har08, RCB05, Und04]. topographical [PTH08]. topography [GMK^{+}03, GWF07, KOC^{+}07, NZM^{+}08b, vdMPEM08]. torreyi [RHBW09]. Tortanus [HB04]. Tortuguero [TBBB06]. tostriped [NH03a]. Total [Bar02c, DBN^{+}09, FMMS03]. tour [Lus03]. Tourism [SLB^{+}08, SR08b, SB07]. tourist [DCP^{+}08]. Toxic [SKK^{+}06, TJ02, BBM^{+}02a, BGF^{+}06, BGC06, BMS04, CVF^{+}03, CCM01, CD03c, EÖKS^{+}02, FGM^{+}00, GJ03, JKS^{+}01, JK03a, KS06, KSKL^{+}03, LGQ^{+}00, LHK04, SGK04, SBH^{+}02, SRDG05, SSBG02, TCD01, VGMC01, WTP08, WL06, WG01]. Toxicity [Kev01, BML^{+}08, BHV^{+}02, HCP^{+}06, KSG03, LELC02, MUT04, TLLCT02]. toxigenic [ABWK06, FPB^{+}07, LRK09, OLWB06]. toxin [GSRH02, GFM^{+}02, JF02, KSKL^{+}03, KB05, PEGG^{+}08, TJ02]. toxins [BCBQ01, CCM01, DCM^{+}06, GFC^{+}02, KS06, KB04, LEÖK^{+}02, MB08, MFG^{+}00, SSO^{+}01, TDP^{+}00]. Trace [BMM^{+}04, FSSW^{+}00, FS02, ZMP^{+}03, DSF^{+}04, HKZ04, Law08, MF08, NATR^{+}05, RBW03, RPS^{+}06b, RPS^{+}06a, RAAT^{+}07, RSL09b, RCC01, RHH^{+}05, Wan02b, WHS^{+}09, XW01, XW03, Zac05a, Zac05b]. trace-element [XW01]. trace-elemental [WHS^{+}09]. traced [BMSD00]. tracer [CF02, DPD^{+}08, HH05a, KECJ06, LBS^{+}05, NCÁSGP05]. tracer-addition [CF02]. tracers [HMW^{+}00, KSI09, MDBG07, MT00, OKP06, SWH^{+}04b, SCH^{+}07]. traces [ZKHP07]. Trachinops [GB05a]. Trachurus [BBG^{+}06, HK04b, KTT04]. Tracing [KWV^{+}05, MLM^{+}02, WWR09, KZ00]. track [Edm04, JT02, KJRG09, Zac05a, Zac05b]. tracked [NCMC09]. Tracking [BTKG^{+}08, CPB04b, CG05, FH08b, HBG^{+}04, RDS^{+}09, RSNG09, SJL05, TBH08, DR01, GLÅ^{+}03, GDR^{+}04b, KJRG09, LBR^{+}09, MBH07, MKK^{+}09, NMOO03, NSH^{+}08, ZAH03, BK05a]. tracks [TML03, TBM^{+}08]. tract [BYH^{+}06, LSG08, WPRS05]. tracts [BC06a]. Trade [FSB07, ANF^{+}00, BDH01, Hal00, MVHV06, PHM^{+}08, SBSPS04, TSSA03]. trade-off [ANF^{+}00, BDH01, MVHV06, PHM^{+}08]. Trade-offs [FSB07, Hal00, TSSA03]. traded [ZM06]. traffic [Lus05]. trail [BK05a]. trails [ED02]. trait [CCK^{+}09, Gim04]. trait-mediated [Gim04]. traits [CR07b, CRH01, DC02, FTB09, JDCB09, Lid08, MC07, MGC^{+}09, NRH^{+}07, PDM^{+}04, SGCP06, TKK05, dCLLT^{+}02]. trajectories [GBR01, HS07, MG05a, ZMP^{+}03]. trajectory [Whi02]. tramoserica [BCU04]. Trampling [EH00, AVA^{+}09]. Trans [CWBS03, TMD^{+}07, RSD^{+}08]. trans-Atlantic [RSD^{+}08]. Trans-Pacific [TMD^{+}07]. Trans-Tasman [CWBS03]. transcribed [PG01]. transcripts [PAK^{+}00]. transect [AI04, SSGB^{+}08, TRDV00, TNSM00a, YI07, dSHCR07]. transepithelial [LTMDB02]. Transfer [BKLGN05, NWT00, OAM06, SHS06, AVS03, BTD^{+}04, CW05, DGM02, DPD^{+}03a, DPD^{+}03b, GRS07, GP04b, HR08, JWM02, KBH01, LEÖK^{+}02, MF08, NFV^{+}06, RGJ^{+}04, RPS^{+}06b, RPS^{+}06a, RAAT^{+}07, SVN04, SFT^{+}05, VF09, WDX01, WW06b, XW02]. transferred [PTA^{+}09]. transformation [CPV^{+}03, Eyr00, MSUK01, RFH^{+}06, WPD02]. transformations [AMT03a, BVI^{+}05, Las03]. transgenerational [NMCT03]. transience [SMP^{+}09]. transient [PA06, RGB02]. transients [RMWO09, SBMAE06]. transition [BA00, HJ03, HM09, IT06b, NNW04, RGP^{+}09, SK04, SK05, TSF01, VGS^{+}01, WKMK06, WKY09, WDB^{+}03, WMD^{+}07, ZBNW09]. transitional [MKWO00, TW04a, TW04b, WKIK04]. transitions [OGL09b]. translation [ZTW08]. translocation [FOL02, HS07, MHM^{+}02, PCL08]. transmission [EAF^{+}05, WM05]. transmitters [JT02, TTD^{+}06]. transmitting [LS06, LALB08]. transoceanic [DRG^{+}02]. transparency [LDHC04]. Transparent [BPVL03, LWH08, EP01, MWMH05, Pas00, Pas02, RYF01, SRGS01, Tho04]. transplant [HR02a, MK02, Sar06]. transplantation [SDK^{+}02, WC03, Yap09, vKH00]. transplanted [BvK07, TLLKC04]. transplants [SD01a]. Transport [Bib07, HE08, HCJ^{+}07, MG09b, NEG01, TYT^{+}04, Abd03a, BP02a, BK00, BSA^{+}08, BDCvB^{+}09, CTLR03, CCL03, CWBS03, CWBR05, CBM^{+}07, DKR09, DCBvBE09, DBT02b, EWBH04, EvBB^{+}09, EE00, FHM02, Fra05, GD07, Gla07, GGR06, HSW^{+}08, HFMP03, Hun04, Hun05, HMFC07, JBT^{+}07, JPG05, KCB06b, KSR^{+}07, KZ00, LC07b, LVvdLH07, MM06a, MCD^{+}05, MS04b, MGL^{+}00, Orv05, PCCL05, POH^{+}02, RGD^{+}08, SS02a, SLFM03, SB05a, SWD^{+}02, TK08, VPSO01, Win09, YH04, dSSC^{+}08, vdVBGW09]. transversa [Tom08]. transversus [SPB09]. trap [HG08, Hor06, HWG06, KKKK03, MS05, NReZW09, OWBD07, RF06]. Trapping [SR01, HSBD08]. traps [DGB^{+}02, GK00b, JTB^{+}03, KWKF03]. travel [KHA^{+}00, KSSB04]. Trawl [Rye02, AHSPG08, BZLD02, CBS^{+}08, GZY06, MS09, PPC09, TPS^{+}05, YM06]. trawler [ARBF02]. trawling [AC07, AWS^{+}09, CED^{+}07, HLVU00, HKK^{+}06, HJG06, JPPW01, JNDL02, KGHL^{+}01, LIM^{+}06, RL09a, THJK06, TFL^{+}04, dJTD07]. treated [GC02, Kev01]. treatment [BMF06, GNAD^{+}04, HCP^{+}06, JMB^{+}05, RJ03, SLEH01, WKL^{+}03]. Tree [BKDB00, Boo02, CS03a, MB02a, Bla09, BDGK^{+}06, DGVK^{+}04, ITN09]. trees [LEF^{+}06]. trematode [JTVC08, OST07, PBJ^{+}05, PM03, Thi06, TdMF^{+}08, TFSP09]. trematodes [FMP05, MS00a]. Trench [FKM^{+}01]. trends [AS04, BCLS06, FHG^{+}03, GRB^{+}03, HMK06, KCK^{+}06, KBA^{+}05, KR04, Kon00, MB07d, PTM08, RFN02, SH07b, TGGA04, TPG09]. trevally [AFHS09, MHP07]. Triakis [CS09]. trials [TMD^{+}07]. triangle [JNH^{+}09, HRCB02]. tributary [TT07b]. tributyltin [OAM02, OAM06]. Trichodesmium [BWK^{+}08, CCL03, EC07a, FB03b, HAP^{+}08, MRPL07, MVB^{+}05, MFCC02, PDG^{+}02, RMC^{+}05, SC06b, SNCP04, WSL06]. Trichotropis [Iye02]. tricornutum [SDRH01]. Tridacna [RGMJM07]. Tridacnidae [RGMJM07]. tridactyla [Bar07, DBH^{+}02, FWH^{+}05, LWW^{+}01, SIK^{+}02]. Trieste [UB03]. trigger [OYT^{+}02]. triggered [KLRS00, OKH^{+}07, ZTW08]. triggerfish [BJC02]. trigonus [WR00a]. triiodothyronine [APW^{+}04]. Trimma [WS08]. Trinidad [MH08a]. trioostegitum [YSS03]. trip [HPH^{+}01]. triplefin [WBC07]. triploid [MSF^{+}07]. Tripneustes [VPS08]. trips [SR04a]. Tripterygiidae [WBC07]. triqueter [KCP^{+}07]. triquetra [AGB03, LWR^{+}02, LTD^{+}02]. Triterpene [KPEF00]. trochoidea [KH00b]. Trophic [AYSL09, AGM06, BAA07, BBH06, BJJ01, CLO^{+}09, CC02, CMMN^{+}01, GHT^{+}01, GFN^{+}09, GFPM01, HGSH08, HH00a, JNMB08, KPM00, KPKD06, KKL^{+}03, KMB^{+}08, Lee00, LHB09, LSBJ08, LG02, LS08b, MF08, MS01b, MA07, MOP^{+}08, NFN03, NBDG05, ORD00, PDS^{+}07, PABM08, QS08, RPS^{+}06b, RPS^{+}06a, RAAT^{+}07, RF08b, RH03, RS06, RTH06, RCGA^{+}04, SMM^{+}04, SMCE07, SIK^{+}08, SJA^{+}09, SB08, TRH^{+}06, VF09, WDB^{+}03, WMD^{+}07, AVNN05a, AVNN05b, APM^{+}09, AO01, AGDP^{+}06, AR02b, BPH07, BS08a, Bas06, BPT^{+}00, BPVL03, BGSBC02, BPB07, BCT^{+}08, BCH00, BKR04, CM07, CR02, CDF^{+}08, CMB^{+}05b, DFPG^{+}03, Das02, DDS^{+}06, DS06b, FPDS^{+}02, FHG^{+}03, FGSH^{+}05, FTO^{+}09, FSV^{+}08, GRP^{+}06, GSSR02, GP04b, GKWK02, GV01, GGRC^{+}08, GMR^{+}02, HRR^{+}01, HM06a, HBRJ06, HMW^{+}00, HD07b, HCK^{+}05, JBS08, JPPW01, JPPW02, JWM02]. trophic [JS08b, KBH01, KK02, Kim02, KVS08b, KH07b, LMS04, LPD09, LTNH00, Lid08, LC09b, LDP^{+}08, MSF^{+}07, MCS09, MCH^{+}06, MR03, MDHL08, MSF04, NFV^{+}06, NOO^{+}08, ÖCOB06, PDFD09, Raf06, RHDC05, RGCC^{+}08, RBGM06, ROBFR08, SJR00, SB03a, SZ09, SFT^{+}05, TÁCTR06, TT06, VMH^{+}08, WDX01, WB01, WFG^{+}08, YI06]. trophic- [JS08b]. Trophic-level [SMM^{+}04]. trophic-niche [SJR00]. trophically [SW04, WL03]. tropho [AM03]. tropho-climatic [AM03]. Trophodynamics [CM01, DHHY09, DB03]. trophosome [SGL^{+}02b]. Tropical [HGS^{+}07, Kel04, RLFEF03, AND05, ADK^{+}06, ALOT01, AMT03b, BS07a, BD00, BBA06a, Bib07, BdBK^{+}00, BFLÁ01, CW03a, Car00, CRJ^{+}09, CRLP09, CH00a, CMD^{+}06, CLJ^{+}08, CKS05, CMS^{+}01, CD07b, DC06, DFV01, EG08, EHK06, Eyr00, FAJM07, Fie02, FBBV09, FPWM03, FKBB01, FBKB03, Fro00, GM04, GSMD00, GF05, GSG08, GR07b, GM05, GBNÖ08, HG08, HO02, HHM^{+}06, HLH^{+}08, HW01, HW03, JM01, JLW04, JS08b, JCK03, KKBH03, KR01, KW00b, KAM^{+}06, LCL^{+}03, LHL^{+}07, LHT^{+}02, Ley05, LYMCG06, LCCTFO03, MAM^{+}01, MAW01, MCM^{+}03, MSY02, MRQ07, MRH^{+}08, MFP04, NvdVVD06, PCT00, PCD04, RMA^{+}02, RBB^{+}08, RRH09, RL08, Rue01, SMMM07, SLC05, SEVSP02, SBB09, SHT^{+}04a, Sha00, She06, SJ09a, SICT03, SP08b, SC06b, SBA01, SWEBS00, SAM^{+}01, SSM^{+}01, VHH^{+}02, VBF06, VPS^{+}02, VPS08, VS08]. tropical [WWGR07, WW06a, WM04, WLJ^{+}04, WLJM05, WCJL08, WDN00, WDLM05, Zam07, ZH02, ZK01, vDvTJD^{+}09, RC09]. tropicalis [BDG^{+}04]. tropicbird [LCL^{+}03]. trossulus [ABY00, BWH09, QTB02]. Trough [GGDB09]. trout [LBF^{+}07, SBHE05, Tak02, TTM^{+}05]. truncatus [ARGS01, IR02]. tshawytscha [HFWW05]. Tuamotu [AAPO00, LPBD00]. Tube [Cal06, FGS00, THS^{+}02, TP01a, TP01b]. tuberculata [BT01]. Tubeworm [GZG^{+}07, UBM^{+}07]. Tubeworm-associated [GZG^{+}07]. tubeworms [SLP^{+}09]. tubulosa [MR07a, MR07b, MR07c]. tufted [WIB08]. tumor [GLB^{+}03]. tuna [BTW08, CPG08, GDdS^{+}07, HG08, KNK^{+}00, KNKT01, LFF^{+}09, MAA^{+}07, MLS^{+}06, MOP^{+}08, RSZI01, RSD^{+}08, RFG04, RF06, SDD^{+}07, TSIY06]. tunas [TOK^{+}04]. Tunicata [KSSH02, LUAIH03, PDM^{+}03, PMNL07]. tunicate [APM^{+}01, DDD08, MC05a]. tunicates [GSFN00]. turbid [AED^{+}03, EÖM08, HSP^{+}08, KUN02, MN00b]. turbidity [Ant06, LF08, MKL^{+}03, MND^{+}02, RWPJ04, RHS01b, SD00b]. turbinata [DDHQ06]. turbine [MWT^{+}06]. turbinid [WWV08]. turbot [CLB^{+}09, EdPP^{+}00, PMT^{+}02, SRDPT^{+}07]. Turbulence [MVN^{+}06, AMP^{+}02, BPE^{+}06, GBR^{+}04, GBD^{+}06, HZ04, IHW^{+}01, JNH^{+}09, MNB^{+}07, NYK03, Osb07, PTB03, PM00, SES01, TSF04]. turbulent [Fin00, Fra05, KSR^{+}07, LB08b, MMVB07, PJB03, RFB07, TÅ07b, VJ04, WB07, WG09]. turf [DEAI06, HJ07, JM03a, Kel02, KCS04, LG08b, MRB09]. turfs [BBC08, JM03a]. turning [Rii04a]. Turnover [TRS^{+}06, ABBN09, BWCW02, CSL^{+}05, Joh04, Kiø03, LG08b, SPAMG00, VMH^{+}08, Wil04, XW01, XYH^{+}09]. Tursiops [ARGS01, IR02, Lus03, Lus05, LW06]. turtle [ABL08, BRGS^{+}07, Bra05, BHN09, DvD02, Eck02, HGS^{+}02, JMJ07, KD00, LB02, MLGG03, MBH07, MB05b, RPC^{+}06, SJE^{+}06, SWHSA05, SAH^{+}05, TTD^{+}06, TBBB06]. turtlegrass [GHV05, KVH02, SCBA05, VHKW00]. turtles [BBM00, BBC03, BGRD00, CLP^{+}06, CAF^{+}08, CPB04a, CPB04b, FFT^{+}07, GF06, GFB^{+}07, GSB^{+}06, GBM01, GBF^{+}02, GBGH02, GLÅ^{+}03, HRB^{+}06, HTM^{+}02, HHB^{+}00, HBG^{+}03, HBG^{+}04, HFW^{+}05, HBBH07, HBG^{+}02, JSMM07, JHL04, KCEB09, LLM^{+}08, MH06b, SBMAE06, SKP^{+}07, SRN02, SJM06, SAM^{+}06, WKPS06, WBJ^{+}07, YA09]. twilight [BBC^{+}09b, BES^{+}08, NRJB04]. twilight-zone [BES^{+}08]. Two [LMF^{+}01, ASH03, AA04, BLD^{+}05, Bel08a, BCBVM05, BDG^{+}07, BBDW09, BLF06, BKB^{+}09, BBRJ04, BRSM07, BSAR03, Bra08, BFV08, BBT02, BHY09, CSM^{+}09, CLL^{+}09, CC03a, CNPS05, CC07, CH07d, CS03b, CM08, CPG08, CD07b, DCD^{+}03, DMK^{+}08, DBR^{+}07, DHR04a, DBT02b, ET08, FGPS04, FC08, FPF01, FGSH^{+}05, FJ01, GHJ02, GF05, GHL02, GZG^{+}07, GMTN00, GLV^{+}09, GLBF07, Hal04a, HPH^{+}01, HVCT09, HR08, HBH07, HH00b, HAD07, HFH09, HFA02, JHL00, JSMB05, KTK08b, KES00, KCG07, Kre01, KFU^{+}08, Kri02, LQRGT06, LYL^{+}07, LP07, LS09b, LIT05, LW02, LTL^{+}06b, LRGdS00, LLRDT05, MD08a, MF03, MSJ^{+}09, MHV^{+}00, MR07a, MR07b, MR07c, MB04, MK03c, NSSI06, NW05b, OHS^{+}07, OGA03, PCYC02, Par06, PC02b, PLM03, PCL08, QDDP06, RM05, RW07, RHC^{+}08, RZLS09, RGA^{+}02, RGEB00, RPO00, Rob06]. two [RMRAFP08, RM06, RKLK07, RST04, SZGS08, SLC05, SLFM03, SRDG05, SO02, SCC05, SLCC09, SWD^{+}02, SSBG02, SJ02, SHSF08, TRT04, TSF04, UBRvO06, UHvO^{+}08, UP07, UP01, VN05, WLL03, WL03, WHP00, WJDA06, WMCF08, WMD^{+}07, WG07, YYL^{+}00, ZTW08, ZZT05, vOSF^{+}09]. Two-dimensional [LMF^{+}01, FJ01, SWD^{+}02]. two-spotted [UP01]. two-stage [JSMB05]. type [EWG^{+}08, HBC09, KFT08, MCM05, SFHS00, SPB09, Tal01, VHW^{+}07, ZBDE07]. types [ASH03, BLS03, BC01b, CHA06, FSV^{+}08, KLN06, KCA^{+}08, LTL^{+}06b, MF03, OPSM04, PCHM07, RLKW04, SML02b, vWCAB07]. typhoon [OKH^{+}07, ZT07]. typhoons [CCJT09, ZTW08]. typical [SPB09]. typicus [BK05a, BC01b, CID^{+}06]. typology [LHB09]. typus [HGKJ01, MBP^{+}06]. tyrannus [HD06a]. tyrrhena [PTLK04].

U. [BBS05, GB07, NPVW06]. ubiquitous [MC05b]. Uca [BBS05, GB07, KWD05, LDT07, OJ04]. Uchiumi [KKTN05]. UK [ABBN09, BAMJ09, DPD^{+}03a, DPD^{+}03b, DLD^{+}00, FRJ09, FR08b, KUN02, LBC07, SP01a, SWGC^{+}05, TDUN07, TNSM00a, TSJ05, WHP00, WAB^{+}02]. ultimate [CRGH01]. Ultrastructure [BND02, PL00]. Ultraviolet [ESDG08, FIRL02, Zam03b, Zam07, AL03b, BRB^{+}00, BSPK03, DQW05, HBdBV01, HTWQ05, Kuf01, LDHC04, Les00, MGL^{+}00, OMT06, PDB04, RBSP00, RBRSP00, SMC^{+}05, SH03, VMN^{+}03]. Ultraviolet-absorbing [Zam03b, Zam07]. ultraviolet-B [MGL^{+}00, SMC^{+}05]. Ulva [BAB04, BVN^{+}03, CD07a, HDQ04, MMMB^{+}05, TNSM00b, TMM05]. ulvae [CBP^{+}05, HDP02, OSS^{+}04]. Ulvales [Ste04b]. unaffected [JM03a, LRM00, SS09b]. unblasted [Fox04]. uncertainties [LSK08, PCJ06]. Uncertainty [Gim09, PBAD05, PGPL03]. uncoupled [MGPAE01, TAS04]. uncoupling [MBG^{+}03]. uncultivated [JWTP06]. Undaria [CHO09, TKGS04, VJ05]. undatum [SJ02, WHB^{+}06]. undecimalis [AWBO09]. under-ice [WA05]. underestimation [NNS01]. underlies [VGB^{+}05]. underlying [Bra01, BBC08, RST04, SO02, WBDR06]. underside [WBL^{+}06]. Understanding [Wri07, Mil07b, Pet04, SvAHH09b, SvAHH09a]. understorey [IC06]. understory [CEF04, DESK03, EDS03]. Underwater [SJM06, SMA^{+}03, CDS07, CCB^{+}09a, Eva08, FHK^{+}08, HCB^{+}07, KFE^{+}03, MWT^{+}06, NDOS03, Rob06, RSIS09, VTDS09, WWTH06, WMB00]. underway [AG00]. Undisturbed [FB03a]. undulatus [CDAM06, MFA03, SEH07a, Tup07]. unequal [FMMS03]. Unexpected [CCB^{+}09b, UM08, ZMTB07]. unexpectedly [BGS^{+}02b]. unfed [TJM03]. unfished [DFSS08, SG06a]. unicellular [FPC05]. unidirectional [QRS00, TC03b]. unintended [CB08c]. uninvaded [MGD^{+}09, NLG05]. Union [ZK09]. unique [HK03c, SL05a]. United [GL00a, NLHO09, CWM^{+}09, DMK^{+}08, GL00b, KMWH04]. units [CUB08]. univariate [TBCB^{+}05]. unlikely [Ryp08]. ’unmarkable [UB02]. unobserved [Rye02]. unpredictability [MdMR04]. unpredictable [GCLC^{+}02]. Unregulated [RGEB00]. unsaturated [PLC^{+}02]. unshaded [BC06b]. unsuitable [BJLH04]. Untangling [WFT^{+}08]. Unusual [FH08c, KUOI06, PGS^{+}00, MBH07]. unvegetated [JH01]. Up-estuary [KSHP00]. Update [KH08a]. upgrading [BKR04]. Upogebia [CCdAR00, GDL04, NPVW06]. Upogebiidae [CCdAR00, NPVW06]. upon [GEM02, SRZZ01, Zam03a]. upper [BPH07, BUF02, CD03b, CS02b, DSL07, DÁSG^{+}01, GRBB07, GMd^{+}02, KT04, Kor01, KS09, LR02, MLC08, RHDC05, WFG^{+}08, WBP^{+}06, CSO08, RHHH09]. upriver [SLFM03]. ups [Pea00]. Upstream [CCK06, CCJT09]. Uptake [BCBQ01, BTD^{+}04, BEE06, MN00b, MvdBÓ00, PW04, TMM05, YLL^{+}04, BFK05a, BFK06, BRJ^{+}03, BGD^{+}03, BSvdH^{+}02, BS07c, CAD00, CPK00, CNPS05, CT06b, CT09, DCT^{+}03, DCT^{+}04, FPBAT02, FSSW^{+}00, FBHB03, GSL^{+}06, GPW^{+}02, GHH^{+}04, GDS02, LS05, LMD^{+}02, LSG05b, LLL^{+}09, MOOB06, MRC^{+}08, MBH^{+}09, MW04b, MN00a, MLG03b, NW04, PSW^{+}08, PH03c, PPCJ^{+}06, QHD04, RB02, RBW03, RB05a, RPN^{+}01, Ree03, RFPB^{+}04, RH09, SKMS06a, SKMS06b, SYPO09, SDRH01, SMT08, SSH09, TC03b, TA03, TNSM00a, WR00a, WDX01, Wan02a, WB01, XZLP09, dBKSV00]. Upwelling [RMC^{+}05, RCL^{+}09, SB05a, Wie05, ÁSG04, ASES07, BB03b, BRG^{+}02, BAOV06, BPA^{+}08, BK06c, CSL^{+}05, CTFC08, CMB^{+}05b, DP02, DRF08, FPB^{+}07, Fen08, GÁSPR03, GB01, GEM02, GGRC^{+}08, HQDH06, IG04, LC02, LG06, LSK08, MM06a, MG09a, MADPQ06, MS04b, MASFS00, MPE03, MFM^{+}09b, MFM09a, NNLV06, NCÁSBP04, NCÁSGP05, OÁSG^{+}08, OÁSG^{+}09, PGM05, Phi05, PRvdL^{+}07, PRL^{+}09, PNCS^{+}08, PNCT^{+}08, RVMI^{+}03, RGCL^{+}05, RGC^{+}06, RMPC08, SM01a, SAMN04, SS09b, SDK^{+}08, SYCG05, TAÁO^{+}03, TS07, TMFF00, TFLA03, WW04, XBW07, YCSS06, dSSC07, dSSC^{+}08]. urban [BMF06, CCU07, KWV^{+}05, PWHW09]. urbanisation [MCCV08]. urbanised [CK09b]. urbanization [Cha03]. urchin [Ada01, AM02a, BL04, BRS07, BFB^{+}09, DHR04b, DHR06, FTP^{+}09, FB00, KSIG08, KR01, KS04c, LB01, LGS07, LJ09, MH08c, MAG07b, RJTM08, Rod03, Sew05, SC07, SJ09b, SdNH^{+}06, TR05, TRT04, TBSJ^{+}04, TSJH05, THBSJ05, TYC00, VHKW00, VJ05, Wal07, WGL03, Win09, WDR^{+}05]. urchins [BBM02b, CU06, CSM02, CL05, DBJ^{+}00, DHR04b, ENAT07, Her06, MACS07, MS01a, NA04b, PS06, RM00, SF07a, SS09a, SS05, TÁCTR06, VK05, VLW09, VHV^{+}04, VVTH07]. Urea [PSW^{+}08, TMM05]. Urechis [OJ02]. Uria [Bar02b, ÖBFO01, RMT^{+}05, SH00, ZPP09]. urotaenia [NWT00]. Urrutia [Ano00b, Ano00c]. ursinus [KW02b]. Uruguay [DGL01]. Uruguayan [GY00]. USA [BRP07, GCHC00, Har06b, PCBL05, RZ00, SW02b, SW02a, Tal09, VHKW00, AP05a, AB06, BA04, BS04a, BG07b, COK02, CT01, CKB^{+}09, CH05, DRDM09, GB07, Hen09, HGC^{+}07, HST^{+}09, HGT01, Hou09, HB08b, JBT^{+}07, JTN09, KJMP05, KR04, KS04b, KCEB09, KGS^{+}08, LB08a, Les05, LAK07, LF06, LLW09, LBC^{+}09, LMB^{+}09, MHW06, MTV06, MMPD09, MJG01, Min02, MIL^{+}00, NW02, PPHT00, PBS^{+}00, ROCB09, RDD^{+}02, RMZC07, SAG08, SRJ09a, SH06b, SFR^{+}09, SBSRW02, TDP^{+}00, TBH06, TBL^{+}09, War05, WWL04, WCRB03, WLJ08, WTR02, WCH09, ZM06, Zam02]. usage [AFHS08]. Use [AE02, APH^{+}04, DHR04a, FLL09, GWH03, GLS02, JWM02, KRKS09, KS04a, LMS04, MCFM02, MCS09, MS03, NMBH02, PTSJ06, PBL^{+}09, RE00b, SSMM06, SdSC01, SSW08, TSI02, WHP00, WP03, AFHS09, AHBM^{+}07, AKL^{+}04, ABL08, AKG02, AFHA05, BPBH09, Bar02a, Bar03, BSW^{+}09, BDF09, CS09, CLABRO^{+}03, CSTC01, CPBS07, COSM03, DLT^{+}06, DPH02, DH06, EJM01, FS08, FC00, FS02, FHK^{+}08, FB02a, GRRH07, GZY06, GLP^{+}08, Hal00, HT07, HSJ08, HTM^{+}02, HMBD01, Hid03, HHWT02, HWBZ05, HFWW05, HSCD01, HMA06, HWGB02, HTBT07, JK04, JT02, JH01, JSIT02, JBA02, JSMM06, KLN06, LB08a, LGB^{+}04, LMW^{+}01, MJG01, OT08, PGCC^{+}06, PJ06a, PRM^{+}09a, PCBS05, PWdN^{+}00, RWE04, RGEB00, RWP07, RZ00, RHH^{+}05, SR09, SMHL00, Sha00, SWS^{+}09, She08, SICT03, SH05, SN09, Sta08, SBLF09, TC01a, TAW^{+}09]. use [Tor09, TL03, UH03, WPM09, WDT^{+}08b, WBC07, WGS^{+}02, SW02a]. used [BBSO03, HKZ04, POH^{+}02, Sim00]. useful [FH02]. uses [SRR06]. Using [ALE^{+}09, BT09, BBDW09, BEU^{+}07, Bro06, BWRP06, DPSD08b, EHWS07, HGS^{+}02, HOT08, HMHB03b, HNMC09, JPG05, Lin04, MGTK00, PCHM07, RZS05, RM09b, SN02, WBR07, AW03a, AB00, AY03, AHZ03, ASLD09, AG04, BKP07, BS07b, BDR06, BCM^{+}07, BTKG^{+}08, BPC09, BRP07, Bur01, CF02, CSD04, CHT06, CS02a, CR07a, CM04b, CBL^{+}09, CDK^{+}07, DCT^{+}03, DSP07, DRTM09, EMG01, EZTZ02, EOL^{+}05, EYH^{+}06, FTDK07, FBJ00b, FHDA^{+}07, FCH^{+}08, GDSB04, GMM03, GDL^{+}01, GBI^{+}06, GMR^{+}02, HLBS07, HDM04, HD06b, ISK07, IP01, JCMR^{+}03, JRC06, JBH00, JdSM09, Jon04, JMJ07, KPA^{+}09, KWTP06, KWV^{+}05, KET^{+}03, Kev01, KKIV01, KFH04, KKMH08, KS00b, KJRG09, LCHG03, LBS^{+}05, LLH06, Lat07, LEF^{+}06, Lei01, LSP08, LWK^{+}07, LPBD00, MRLK06, MFWD08, MHP03, MW04a, MRHP01, MBH^{+}09, MKK^{+}09, MWM^{+}05, MAKMK09]. using [MLR02, NWIT08, NPTW05, OP08, PL07, QRWJF07, RRS^{+}09, RDS^{+}09, RSNG09, RPC^{+}06, RSZI01, RCGA^{+}04, SAG08, SWD^{+}04, SM08b, SC01, SLT^{+}01, SMH09, SK02, SWH^{+}04b, SLL09, TUK09, TGA02, TBH08, TSEH08, WPM09, WGCM^{+}03, WWR09, ZRG00]. usually [BD06]. Utilisation [BWMA02, CB01, BSJ02, FMZvW02, FHB^{+}02, HJW08, PH03c]. utility [MSP^{+}06]. Utilization [HGG^{+}03, HSD03, KECJ06, RMKR00, Sew05, BMO^{+}04, FBM07, JFD09, LTCP03, SOMM06, TDSM02, ULL03, WPPS04]. utilizing [RGD^{+}08]. UV [Ada01, BW00, CAS00, CABS02, DL03b, EAF^{+}05, FJV^{+}02, HGG^{+}03, KFE^{+}03, Kuf01, LN05, LFSR06, LWMW02, MP05, MLW03, RGDN02, SFPGA05, SML05, VTDS09, VPS^{+}02, WGP^{+}00]. UV-A [SML05]. UV-B [LFSR06, MLW03, SML05]. UV-dependent [SFPGA05]. UV-induced [VPS^{+}02]. UV-radiation [WGP^{+}00]. UVA [MP05]. UVAR [DHCH04]. UVB [Rij03]. UVBR [BdBK^{+}00, BWvdE^{+}06, DHCH04, RvdPHW04]. UVBR-induced [BdBK^{+}00, BWvdE^{+}06]. UVR [FNDP08, VMH04].

V [CT01]. vaginalis [TLPG06]. Valdez [BBD^{+}02, DBJ^{+}00, DBF^{+}02, EBT^{+}02, GSM^{+}02, PMGE01]. valentina [OTB^{+}00]. valid [CA00]. valida [PMM^{+}00]. Validation [SSM^{+}01, TBB^{+}04, CT09, GT08, JSMB05, PR09]. validus [MAH^{+}08, PWM^{+}08]. valuable [NM09, Tup07]. valuation [PGP03]. value [BGS^{+}02b, CLB^{+}09, CH06c, Dis07, EL05, Hid03, NPTW05, RST04]. values [ISG04, LGM^{+}09, SMM^{+}04, STH^{+}06, WHRS05]. vampire [CD08]. Vampyrophrya [OHS^{+}04]. Vancouver [BHD04]. var. [DDHQ06, PBBC06]. varia [MBH^{+}09]. variabilis [GJT09]. Variability [BPT^{+}00, BC01a, BDCvB^{+}09, Dav00b, DGH08, DCBvBE09, EvBB^{+}09, FPWM03, GhGD^{+}02, GGFSP06, GK03, LC09b, LSHL03, MSH^{+}03a, TFSJ00, TK08, WHL^{+}06, YTH^{+}05, vdVBGW09, AAGS02, ABF07, Air00, ADR^{+}05, BS00, BA00, BI04, BVMBC07, BMG02, BPK^{+}00, BCVC07, BTF^{+}06, BHSR00, CAD00, CCC09, CH06a, CMB^{+}09, CC04, CEB^{+}08, CIHF01, CPSC01, CB05b, DE01, DC06, DB07a, DSJBS^{+}02, DFJSJ^{+}05, DB05, DTB07, DOML^{+}07, ESDKM09, Ell02, FSP06, Fie02, FHPW07, FD04, FBT^{+}01, FWH^{+}05, GÁSPR03, GB01, Gis05, GSW01b, GK07, GAM02, GBG^{+}00b, GCM01, HAMW09, HBRJ06, Hil04, HJK01, HMFC07, IOTL09, IFL02, ISK07, IRM08, JNS04, JAA^{+}01, JÅC^{+}01, JSH01, KBH01, LBD00, LG00b, LLCHG01, LHB^{+}06, LTH^{+}06, MNR^{+}02, MBG^{+}03, MCE01a, MTH^{+}08, MGE^{+}09, MHB^{+}09, MKL^{+}03, MPR02, MAT01, MG09b, ME01, NDOS03]. variability [Nod04, NAB09, NT06, OA02, PDHH06, PCW05, PWB^{+}03, PT04, PMR06, QMM^{+}08, RCC^{+}04, RRIH^{+}02a, RCL^{+}09, RT00c, RGCC^{+}08, SS00a, SD09, SASR03, SH07a, SC00a, SHT^{+}04a, SJ09a, SBM02, SC07, SMSD07, SMG02, SBSRW02, SGCC06, TS07, TTR05b, TSMC09, VTDS09, VMH04, WR06, War05, WSA^{+}06, WH08, WA05, WKN^{+}03, WMB00, YMC05, vDSD02]. Variable [CTHF01, DMC09, JLCA01, KT04, LFSR06, LW00, MK04a, OAS01, SS00b, TGNT05, VHK09, ACH07, BPH07, BCBVM06, BFE^{+}05, FE02, JS01, MCW^{+}07, Phi05, SM00a, SS09c, UBM^{+}07, VN05, WLM06]. variables [ABB02, BCO07, CR07b, GDL^{+}01, KvKB06, MB02b, MW02b, SHFP^{+}03, WGS^{+}02, YH02]. Variance [OM04, BVMBC07, SM05, VBMBC08b]. varians [RPS^{+}06a]. Variation [APA07, CSvO09, GSMD00, HMR00, JSWH03, JSIT04, KL03, KM03a, Kri02, LEF^{+}06, LPK^{+}03, MK03b, MS01a, MK03c, RC04b, SAJG06, SF02b, UBRvO06, Van03, VSBB03, WWDH09, YI07, AS04, AHC^{+}06, AED^{+}07, Alt06, AMS^{+}06, ABU01, ABB02, BCL03, BC02, BK04, BMdNP09, BCBAC01, BBZM08, BSK06, Ber02a, Ber02b, BRP00, BHH09, BGPL08, BG04b, BB08a, BPCvdM06, BNSG01, Bur02, CB00a, CF09, CHH^{+}05, CCK^{+}09, CU08, CJRT00, CH07c, CSO08, CDT^{+}03, CW01, CH07d, CGL^{+}02, CTVP03, CGC^{+}09, DSPHV04, DRV^{+}05, DGC^{+}05a, DC07a, DKAH02, Dun02, DCC^{+}03, EC07b, EBG^{+}00, FHVMDS07, FTBC05, FEMW07, GBdS^{+}05, GB02a, GWPRC^{+}07, GTGM06, Gru00, HARC05, HJG03, HHG^{+}07, HRCB02, HHJ04, HKH09, HSCF06, Hip09, HKSW01, HTS06, Hum00, HS01, HCMR09, HW01, HFM^{+}04, JM01, Jar00, JH01]. variation [Joh03, KKLL02, Kne09, KN00, KOY^{+}05, KS09, LB08a, LNV^{+}05, LBvdH03, LJ08, LG00a, LH09, LCL^{+}03, LAK^{+}08, LT07, LVA07, MGL^{+}07, MSF^{+}07, MSG07, MWR07, MSK00a, MUT04, MMMPB05, MGM^{+}07, MW08, MB00a, MB01, ML05, MP06, MHK^{+}01, ML06b, ML07, MDW^{+}09, MPS^{+}08, MS00b, MCM05, MCS05, MRH^{+}08, Mun07, MTCU08, NW06, NAB^{+}06, OHCFJ04, OC01, PGBA04, PPL^{+}02, PA07, PW07, PHS09, RHBW09, RVMI^{+}03, RPO00, RP06, RT00b, Rol00, RC00, RS05, RW06c, SM01a, SF07a, SDS^{+}03, SM08b, STRD02, SRJ09b, SJR00, SC03, SBR09, SP01a, SV07, She06, SS03, SMFP08, SGCP06, SLHR09, TN07, TAFBC07, TSSQ06, TM02b, TG03, TC03c, TCWR08, TBAT02, Tru02, TBI^{+}09, UHvO^{+}08, VMCG09, VK02, VBFF08, WW06a, WGD05, WHS^{+}09, WBSL00, WPIB09, WdNS00, YYL^{+}00, YW04, YH02]. variation [Zac05a, Zac05b, Zam03b, vBWSS06, vWNYS05]. Variations [BFK^{+}05b, OMP^{+}05, SRLA04, CMD07, CM05, DD09b, EMG01, EdPP^{+}00, GKM02, HPB^{+}04, HMHB03a, JMSM08, KTK08b, KHH08, KI01, KKA^{+}06, LWR^{+}02, MJS09, MRB^{+}08, MPE03, PW09, PJN03, PDRT02, RYF01, SNH04, SFF^{+}07, SdSC01, THS^{+}02, UB03, VMRL04, VPS^{+}02, WP03, YHSH02, Yin02, YVB^{+}02]. variegated [MBH^{+}09]. variegatus [CL05, MY02, MY04, VHKW00]. varies [KCB06b]. varieties [SDR01]. variety [LB08b, RLBB04]. various [BDJ09, HB03, LMW04, RH03]. variously [BWMA02]. varnish [DD06]. vary [GC07, KB05]. varying [CWT06, DWL09, HBNA04, LSH00, PTB03, TNTH04, VCSL^{+}09, VBV^{+}00]. vascular [SZZ08]. vaterite [TCW^{+}07]. vectensis [DKR09]. Vector [HD05, BPC^{+}02, MFG^{+}00]. vegetated [KLN06, LB08a]. vegetation [DGVK^{+}04, DGW01, Fin00, SVG^{+}05, TC01a]. vegetative [KH00b, Pep06]. vehicle [PNK^{+}02, RSIS09, TFL^{+}04]. vehicles [ACR07]. veined [MHW06]. velocities [CS01]. Velocity [Fin00, Bib07, BFK^{+}09, HS02, JKNJ07, Kaw01, OIP04, PBPLB06, PTB03, PCS04b, Tru02]. velvet [LWJC04]. veneficum [WTP08]. Veneridae [MTS07]. Venice [SCB00, BSP^{+}08]. venosa [MHW06]. vent [AHK^{+}02, BTL05, BDDL03, BEU^{+}07, CTVP03, FTU^{+}00, GZG^{+}07, GLG^{+}05, KFH04, MCP^{+}06, PEDT03, PGS^{+}00, PAB^{+}02, PZMG05, RCC^{+}09, RL09b, SV07, SJS^{+}01, SKW^{+}06, VT00, Van03, WTF^{+}05, YFH^{+}02, ZLGC04]. vent-endemic [SV07]. ventalina [DKAH02, THYBS09]. Ventiella [SV07]. ventilation [BFE^{+}03, MGM05]. ventricosa [DRDD09]. ventricosus [SCM09b]. vents [Bat07, HVCT09, KCB06a, LJM03, LLJ05, LSBJ08, MC05b, MMS^{+}05, Van02]. venusta [SHT^{+}04a]. verification [BD01, HTBT07]. verify [HW09]. verifying [CB08c]. vermetid [KAMR00, RHG05, ZAZL08]. vermiculophylla [TMM05]. verrucosa [BCBVM06, MG08, RRDHG08, STE04a]. verruculosa [BCBVM05]. versus [AG02b, BLS03, BN09, BFB^{+}09, BZLD02, BMO^{+}04, BB02b, BH09b, CH00a, CR02, DDSR07, DRTM09, EvdKB^{+}08, FS02, GM01, Gla01, HLR07, HS03, HGKK06, HW01, IC02, LH08, LJ09, LW06, MG08, MRP09, NPM01, NP05, NGdN^{+}07, OM04, PGK^{+}08, RSN02, RBL^{+}04, Rii04b, RCWG^{+}06, Rui04, RSA05, RC07, SFW01, Sha09a, Sha09b, SC06b, SBA01, SMHG09, TOK^{+}04, TC03b, TM02b, TA03, Uth01, VTS^{+}02, VMH04, WAW^{+}09, WKK07, dM01a]. vertebral [NWJ^{+}08]. vertebrate [CGH08, CDS07, PHB02, PBB^{+}02]. Vertical [BCVC07, DFJSJ^{+}05, DSM01, GMB^{+}06, GLP^{+}08, HCGC09, HRMH07, HLH04, HH00b, HMO^{+}04, JF07, MSF^{+}09, MGF^{+}03, MMS^{+}05, NSH^{+}08, OS01, PK06, QMM02, SZGS08, SK03, TDRM08, VPP04, VSRGW09, WPC^{+}03, ACAG07, BSB03, BSKR^{+}06, BZLD02, BT04, CPBS07, CBM^{+}07, DF02, DM01b, DEvdZ03, ENAT07, EAF^{+}05, FLG02, FLR05, FFML02, FSV^{+}08, GYD04, GK03, HF06, ICH04, JPC^{+}06, KNK^{+}00, KCB06b, LTV^{+}01, MVN^{+}06, MADPQ06, MB00b, NS02, NM07, OSA05, PIY04, Pea00, PG00, PT04, RWR^{+}00, SHNV09, SS03, SAHK06, SS02b, SES01, Sve02, Tar03, TCC04, TJS08, TÅ07b, TF04, TT05b, VTS^{+}02, VAB09, WMS^{+}03a, WKMK06, WEEBdN08, YLv^{+}05, YMC05, dSSC07, dSSC^{+}08, BPK^{+}00]. vertically [ANF^{+}00, SS02a]. verweyi [RFPB^{+}04]. Very [SJPC02b, FLG02]. Vesicomyid [GB02a, HGC^{+}07]. vesiculosus [BNP07, HHJ04, KLFH04, PKS00, RBJB^{+}05, RK08]. Vessel [JBW^{+}09, HKWM07, LC09a, LBR^{+}09]. vessel-based [LBR^{+}09]. vessels [Hyr01, SD01b]. vestimentiferan [BUF02]. vetulus [Bro06]. VI [TBL^{+}09]. via [EvBO^{+}08, FH08b, HRR^{+}01, HSW^{+}08, NEMF09, SAB04]. viability [GMP07, HR07, JSN08, JB07a, RvdPHW04, SKU^{+}08, WGP^{+}00]. viable [MCU06]. vibrissae [ZS04]. Vice [ZK09]. vicinity [DBT02b, GPW^{+}02, JK06, NMJ02, Yin02]. Victoria [Kev01, JBH00]. victoriae [GTPC04]. Victorian [HJG03]. video [BLH02, CDS07, DFW^{+}03, DR01, EB08, GMM03, GGDB09, HCB^{+}07, HLBS07, JN09, MFWD08, MGL^{+}07, NDOS03, SJM06, WMB00]. video-recorded [DFW^{+}03]. video-time-depth [SJM06]. Vietnam [ATTX00]. view [Pör08, SBJ06]. Vigo [NIF00, ÁSG04, PNCS^{+}08, PNCT^{+}08]. VII [HRR^{+}01]. villosus [PWFC08, RFCM03, WJDA06]. viral [GGG^{+}04, TUK09]. virens [CVK00, LOB03, ODG^{+}04, RPS^{+}06a]. Virgin [Chi02, CS03a, DMK^{+}08, Edm00, GDBF06, Nem05, RM06]. Virginia [SW02b, SW02a]. virginica [BH09a, KLR^{+}06, LNV^{+}03, MWMH05, PCS04b]. viride [CRAP03, GPK^{+}04, PS08]. viridis [BTKG^{+}08, CU06, CYC^{+}09, GSL^{+}06, GXL^{+}08, NW05b, PW04, Wan02a, WLT^{+}03, WC03]. virus [ESQ03, SPBM05, TUK09, TLTB03]. virus-dilution [TUK09]. virus-like [SPBM05]. viruses [GRG^{+}02, LCS^{+}01]. Viscosity [RL07, SV08]. visible [Gri02, KS04a, PDB04, SH03]. visitors [MVD^{+}08b]. Visual [BS06, GTD^{+}08, MBW07, Tor01, CST08, MVdMFF08, RFLC07, TFL^{+}04, UP01, WWTH06, WMB00]. visual-census [MVdMFF08]. visualization [RGD^{+}08]. visualize [RDG^{+}07]. vital [Kam00, MDG04, VÅ01]. Viti [MC08a]. vitrina [MR07a, MR07b, MR07c]. vitro [ODG^{+}04, PS07, PNCT^{+}08, RAAT^{+}07]. vitulina [HBBD09, NWIT08, ZS04]. viviparous [OAM06]. viviparus [ORS02]. vivo [ABB^{+}00, SO08, SdSC01, SJPC02a, SMSJ04, WM02a, WMS03b]. Vladimir [Pud05]. vlamingii [RAM03]. volatile [PLC^{+}02]. Volcano [LLJ05, SLP^{+}09, VMdBV06]. volitans [AH08, KMWH04, WGV^{+}02]. volitans/miles [KMWH04]. Volume [Ano00d, BF07, BBRJ04, MDLS01, VWH00]. Volumes [Ano00a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a, Ano04a, Ano04b, Ano05a, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b, CHA06, Kin05b, TÅ07b, Ano05b]. volumetric [SL05c]. volutator [HDP02, MS00a, MR06c]. voyage [DRG^{+}02]. VPR [JN09]. vs [FBSG02, GKS^{+}07, JHR05, NNG03, SPB09]. vulcani [BDDL03]. vulgaris [SS04a, GWH03, OÁSG^{+}08, OÁSG^{+}09, VAJ^{+}03, WBDR06]. vulgata [AH07, CCB06, CHW06, DJM08]. Vulnerability [Wen00, BOM^{+}08, CWM^{+}07, LR06, MS09, Rye02, SBJ02]. vulnerable [PPC09, PPB09b, PRL^{+}09]. vulpes [HALL05].

W [NH03a, DGJ^{+}03]. Wadden [EvBO^{+}08, ZCR^{+}09, vKH00, vdVBW00, vdVBGW09, BAA07, BDH01, BDEM01, BD03, BD06, DB07a, FSO08, PA06, SA03a, Str02, WLZ05]. wahoo [HLH^{+}05]. Waikiki [WWMF06]. Waite [Ano05d]. wake [JTR05, STRB06]. Wales [BDR06, MGL^{+}07, PK00, DLD^{+}00]. walk [CHPS04]. walkeri [DL03b]. walks [TÅ07b]. walleye [BBC^{+}09a, Bai00, DABC02, PCH^{+}05, PBS^{+}00, Rye02, WJDA06, YHSH02]. walls [CK04, SJS^{+}01]. walruses [FKI^{+}09]. Waquoit [MTV06]. warm [EF02, KJMP05, NNLV06, Trn02, WKMK06]. warm-temperate [EF02, Trn02]. warm-water [KJMP05]. warmer [RHS01a]. warming [BDJ09, EWR08, Pör08, SBI07, TLIL^{+}09, WDSA09, WDB^{+}05]. warning [KCTG06]. Washington [RDD^{+}02, BS04a, BFE^{+}05, CSR^{+}00]. waste [FMCB05, GSL^{+}06, SSMM06, SRS04, VMM^{+}04]. wasted [Iye02]. wasting [MT09, RS02, SCBA05]. Water [CMG04, CT06b, FBD^{+}02, JM07, KHIH08, LSdM^{+}03, MLG^{+}03a, NvW01, NYvW03, NGMT01, RL05a, SSH^{+}04, TL08, VMN^{+}03, WT06, Abd07, ADS^{+}09, AMB08, ACPS09, AWP^{+}01, AP00b, BKP07, BMK07, BS07a, BGJ^{+}07, BFB01, BWP^{+}06, BSvdH^{+}02, BEKT09, BD08, BM07c, BLP^{+}02, CHW02, Cai09, Cal05, CLO^{+}09, CM01, CBAd00, CS02b, CMJ00, CIFN00, DF04, DHRG05, DBOR09, DZ06, DRG^{+}02, DdNEvdZ05, DLB04, DFR08, EWBH04, EHK06, EG03, EG05, EÖM08, ERR06, FA08, FHPW07, FH08c, GWH03, GRWB02, GB03, GB02b, GGDB09, HAMW09, HLBS07, HCP^{+}06, HDG^{+}08, HvW08, Hou09, HAA00, HTBT07, INH^{+}09, ITN^{+}06, JHB06, JNFL04, Kaw01, KJMP05, KFH04, KFA^{+}04, KYML02, KSS^{+}05, KMR01, Kro05, KST^{+}08, KKI^{+}06, Lar03, LS08a, LKR^{+}06, LDHC04, LSG08, LCB^{+}05, Mad02, MBW02, Mal02, MHB01a, MRS00, MGDP01]. water [MBC^{+}07, MY04, MMMPB05, MH07, MLW03, MSTS^{+}06, MBN^{+}09, NDV^{+}00a, NDV^{+}00b, NvdV02, NvWY05, NNLV06, NGBG07, NGM^{+}05, OGL^{+}09a, PGR^{+}01, PB07, PJ06a, PC02a, PLMS04, PSGP04, PCS04b, PBL^{+}09, PI07a, RMWO09, RBW03, RHS01a, RGA^{+}02, RL07, RWPJ04, RC04b, RDH^{+}09, RFB07, Roy02, RW06b, SGB09, SMC^{+}05, SBFP08, SW09, SP01c, SLCC09, SVC^{+}01, SMG00, SLEH01, TMD^{+}07, TL05, TVD02, TC01a, TDRM08, TDUN07, Thr09, TA03, TFL^{+}04, Tru02, TBL^{+}09, TH06, UÓ03, Uth06, VJ04, WKL^{+}03, WKMK06, WMW^{+}08, WWWC02, Zam07, vDHHM08, vKH00, vWNYS05, CW03b, RHHH09]. water-borne [MHB01a, MY04]. water-column [LDHC04, PC02a, PSGP04]. Water-flow [NvW01]. water-processing [RGA^{+}02]. water-pumping [RL07]. Water-soluble [TL08, SMC^{+}05]. Waterborne [HDQ04, DDQ04, KZ00, LW05, SJPC02b]. waters [AD00, BFK06, BA01, BN09, BBZM08, BCL^{+}01, BVCG01, BBV^{+}01, BO00, CWC00, CML08, CCML07, CLL^{+}09, CWT00, CD03b, CPH^{+}02, DGBS08, DFCH05, DGC^{+}05a, DÁSG^{+}01, DL03b, EAF^{+}05, FPDS^{+}02, FDBT06, FVAB00, FMMS03, FHQ^{+}06, GTD^{+}08, GVKF03, GLÅ^{+}03, GAM02, GT02, GK03, HSTS^{+}09, HJG03, HKM00, HCGC09, HR02b, HR02c, HSP^{+}08, HSCF06, HWBZ05, HW08, HHAH00, HHKH05, HCR01, KRK09, KTT04, KUOI06, KS00a, KFU^{+}08, LHD^{+}01, LMCQ02, LWH^{+}04, LYMCG06, LUIAH03, MB05a, MBG^{+}03, MCP05, MBBW08, MP05, MHY^{+}06, MKWO00, MGPAE01, NE01, NLB08, NSEL00, NMS^{+}02, PGCC^{+}06, PJJ^{+}05, PBBC06, PCH^{+}05, QASA08, RSK04, RJS04, RLDL02, SLS^{+}00, SRGS01, SKFPS02, SGL02a, SP01a, SHZ^{+}05, SE04, SPAMG00, Sma07, SDRH01, SWD^{+}02, SMSJ04, SLHT05, TW04a, TW04b, TOK^{+}04, TP07, TWYS03, TFA^{+}08, TNHL02]. waters [TBID09, VPS^{+}02, WB01, WLJ^{+}04, WAW^{+}09, WKN^{+}03, XYH^{+}09, YQC^{+}00, Yin02, Yin03, YF04]. watershed [DT06]. Watersipora [NK03]. Wave [BHR08, DB05, BOR08, DD05, EDS03, GWH03, GR07a, HDW02, HS08b, HS01, Kaw01, KSR^{+}07, MYE03, ML00, ML05, NTH^{+}02, O’D08, RBG04, SB03b, TH06, WWV08, WJ06, ZMN07]. wave-driven [KSR^{+}07]. wave-exposed [WWV08]. wave-exposure [RBG04]. wave-induced [HDW02, Kaw01, MYE03]. waved [AFHA05]. Waveguide [JBS^{+}09]. wavelet [RFSC08]. waves [HHW^{+}07, LCF02, MST03]. wax [SKFPS02]. ways [BDJ09, CAT^{+}06]. weak [Som00]. weakfish [TT07b]. weaning [GLA06, GdLB^{+}09]. Weather [HBN01, MCE^{+}01b, RMA^{+}02]. web [AHC^{+}06, AVS03, AMS^{+}06, Bar07, BEU^{+}07, BCL^{+}01, BJR04, BSIS07, CKO^{+}07, CR08, CSPA09, CMB^{+}05a, CPG^{+}09, DPGN08, ECK05, Fre03, GFPM01, HR08, HVH^{+}07, HS07, IT06a, ISK07, JMB^{+}05, Kaa04, KPM00, KSSH02, KSI09, KVH02, KSKK^{+}07, KGS^{+}08, LR02, LNB^{+}03, LNPB04, LEÖK^{+}02, Lin02, LS08b, MNR^{+}02, MNB^{+}07, MFG^{+}00, MR04, MRB04, MT00, Mon07, NM07, NAB09, PTA^{+}09, PI07b, RGCC^{+}08, SLGD00, SAJW04, SNH04, SZZ08, STH^{+}06, TYT^{+}04, Thu07, VHB^{+}07, WDSA09, WAZ07, WLT^{+}03]. webs [AHSPG08, AZKJAM06, ARFP09, BPC^{+}02, BAOV06, BVI^{+}05, CLP08, DCO^{+}08, DWM04, FGSH^{+}05, JWM02, JMSM08, KKH07, LR08, MTV06, PRB^{+}08, PMH08, RE00b, VLRJ07, WWAC00]. Weddell [DFW^{+}03, HRK^{+}08, HHWT02, LBvdH03, LBBH08, MFWD08, MWS^{+}04, MGHA05, MWM^{+}05, ORD00, OLGG^{+}02, WMS^{+}03a]. weddellii [HHWT02, MFWD08]. Wedge [CGCMPA01, CRLP09, CKS05, MMS09, PSKC04, PC05]. wedge-tailed [CRLP09, CKS05, PSKC04, PC05]. Weeding [HK02a]. week [MGL^{+}00]. week-long [MGL^{+}00]. weight [AWOO02, BH04, GTSS03, HM01]. weighted [LDHC04]. weissflogii [SP00]. Wellington [WG01]. were [TT07a]. West [JPNM^{+}06, Kor01, LNH00, MPSJ06, MRG08, RC04a, TLNH01, BWMA02, DCF^{+}05, GSRH02, GGDB09, HKZ04, JMA02, JNFL04, LÅCN01, McQ05, NWHH^{+}09, QAA^{+}06, RHS01a, SHH09, SPP^{+}09, SBSRW02, XBW07, YW04, LHJN07, LLCHG01, SRS05]. Western [MBP^{+}06, NJPI03, OKM06, SHS06, TNHL02, WAK^{+}09, ASA^{+}03, AMS^{+}06, ASES07, BWW^{+}06, BG07b, CW03a, CCC09, CM02, CM01, CRJ^{+}09, CRLP09, DGM02, DSJBS^{+}02, DFCH05, FDBT06, FKM^{+}01, GVF^{+}03, GAAB^{+}08, HRK^{+}08, HLY05, HB07b, HRC^{+}07, ISYM06, JPMGS08, KI01, KAM^{+}06, LAK07, LLRDT05, MB05a, MGW^{+}09, MAA^{+}07, MMA^{+}04, MKWO00, MLW03, MRQ07, NCMC09, NSEL00, NEK^{+}07, NH05, OMP^{+}05, Øre00, OSL07, PAM^{+}09, PD01, Pra01, RRS^{+}09, RGB02, RFG04, RMAY04, SSR^{+}09, SAB^{+}05, SK04, SK05, SFKT09, SCGC07, SPGS03, TOK08, TKO08, TWYS03, TNT08, TWT08, VPP04, WKH02, WKMK06, WKY09, WKL01, Wil09, WJB09, WS08, YI03a, YI06, dCLLT^{+}02, AHSPG08, BSA^{+}08, CPA^{+}03, FHQ^{+}06, HKK^{+}06, IT06a, MAG^{+}07a, McI03, OHT02, PCL08, RKJ01]. Westerschelde [BdVD^{+}01, FCS^{+}06, WBSL00, WBH^{+}04]. wetland [HCH09, JS08b, RMZC07, SJ08]. wetlands [CB08b, DCD^{+}03, DES^{+}00, JS08b]. wettable [QWY07]. Whale [FHQ^{+}06, HGKJ01, AY03, Ano01a, BS03, BCCM03, CDML06, CS05, DVBLS07, FHN^{+}09, GPS07, GDdS^{+}07, HMY^{+}08, HST^{+}09, JBT^{+}07, Kas00, LHJN07, MWR07, MBP^{+}06, MR06b, NEMF09, PRM^{+}09a, RWE04, RBGM06, SSS06, SCM^{+}09a, SAC^{+}03, TSW^{+}09]. whales [ABD^{+}03b, AL08, AAY^{+}07, BBHT02, BGDS05, BCC^{+}03, BM03, BKLGN05, BWW^{+}06, CSR^{+}00, CH00b, CMB^{+}05b, DJG^{+}07, DCM^{+}06, DD01, EB04b, FE06, GDG02, GDSB04, GRRH07, HLH^{+}07, HFT^{+}09, HBW^{+}05, HYR^{+}09, HWGB02, JTR05, LSMR05, LMKG01, MSLP06, MFG^{+}04, MAC^{+}07, MR06b, OBG^{+}07, OCB^{+}07, PBAD05, PPB^{+}09a, RBHS06, RJS04, RCGA^{+}04, Whi02, WCJL08, WWWR09, WWR09, ZAHJ^{+}06]. whaling [Cor09, SAC^{+}03]. wheeleri [Hum00]. whelk [BW04, CR07a, MHW06, RBG01, SJ02]. whelks [RBG04]. Where [AM04, CB08a, Gar00, BGK09, WRW06]. which [PH01a, RSR02]. while [VHW^{+}07]. whip [MS09]. White [DD09a, AFHS09, BKLGN05, CWC00, DNL08a, KS04b, LKDO07, LP03, LMKG01, MBBP08, NH03a, NH03b, RW07, WK02, WOL^{+}07a, DM01b, SP02]. white-chinned [CWC00]. whitemouth [JBCL03]. whiting [AML^{+}03, CCML07, FBJ00b, JBH00, JH01, Jen05, KHvW02, Rin03, RGL06]. Who [GBB^{+}01]. whole [COB^{+}03, KMP^{+}06, LCC07, NGdN^{+}07, PSPD07, SM00a]. whole-cell [NGdN^{+}07]. whole-larvae [PSPD07]. wide [CHT09, GS07, HTM01, KWTP06, LLCHG01, SHS08, VN02]. widely [TCWR08]. wider [SKR08, STA^{+}05]. Widespread [Fen08, LF06, Gri09, HPR05, LSK08]. widgeongrass [MRP09]. width [BD00, BS08b, DD05, HMHB03a, SFL^{+}09]. wiedenmayeri [ALPK00]. wild [ABS^{+}09, APG^{+}09, DSJBS^{+}02, DFJSJ^{+}05, DUSJ^{+}09, FOFG09, HMH^{+}08, LCHG03, LSP08, MSF^{+}09, SMM^{+}09, SJCM01, SWHSA05, TWMW06, UDB^{+}09, VMM^{+}04, WGH^{+}00]. wild-caught [FOFG09]. wildlife [BSW^{+}09]. wilkitzkii [Pol00]. will [RZLS09]. willemoesi [MS09]. William [GSM^{+}02, IFL02, JTN09, PBS^{+}00, PS01, Pur03]. Williams [Ano01f]. ‘Wilma’ [GR07b]. Wilson [QMF05]. Wind [CGH^{+}03, KHvW02, LDL^{+}04, MWT^{+}06, RAS07, CHT06, CMB^{+}05b, DTB07, DBS05, FPB^{+}07, GSS09a, GBD^{+}06, IHW^{+}01, MP00, NTH^{+}02, QAA^{+}06, TAÁO^{+}03, WW05]. wind-driven [DTB07, GBD^{+}06, MP00, TAÁO^{+}03]. Wind-mixing [KHvW02]. wind-wave [NTH^{+}02]. window [BGGR02, PSS06]. windpower [KCH^{+}03]. winds [GSFF^{+}09, ZT07]. wing [KRCGC06]. wing-propelled [KRCGC06]. Winter [JSMB05, MPM^{+}03, MBL07, PISL02, SSS^{+}09, BHW^{+}03, BA04, BWCW02, BCC^{+}04, BGH06, CSA02, CCC09, CSH00, CES^{+}03, CKF^{+}03, DCC^{+}03, GHL02, HGA07, HW00, KMAC^{+}02, LHD^{+}01, LNH00, LLW09, LCB^{+}05, MVHV06, MPM^{+}02, MPSJ06, MGTK00, PCS^{+}07, RTPS07, RBMB00, RGP^{+}09, SSR^{+}03, SHFP^{+}03, STM06, SMP01, Str02, TC03a, Tay03, Tay05, Thi03, TWMW06, WW07, WWL04]. winter-spring [CCC09, GHL02, RGP^{+}09]. wintering [BGDS05, BD06, BDF09, FKI^{+}09, GDSB04, HANM^{+}05, KGF^{+}09, KNF^{+}03, LHJD^{+}03, LG00a, NMBH02, OP08]. winters [BD03]. wintertime [JNS04]. wiping [HSCD01]. Within [BPM05, APM^{+}09, ANF^{+}00, BT09, BA01, BRBPL06, BHH07, BHMR01, CC02, CLJ^{+}08, DDS^{+}06, DWB04, DRV^{+}05, FH08b, GB03, GSRH02, GPB02, HAMW09, IFL02, JB07a, KLN06, KAR04, Kri02, KGS^{+}08, LHD^{+}01, LHGC^{+}09, MG05a, MS08, MK03c, MPB^{+}09, O’D08, OPSM04, PSKC04, POH^{+}02, PLM03, PCL08, RGS^{+}09, RSN02, RS00, RB04, SP03a, SAG08, SAB04, SSH^{+}04, SC03, SPBM05, SDR01, SFR^{+}09, SKO^{+}08, TBI08, Van03, WK02, WÓ02, WGR07, WGL03]. Within-reef [BPM05]. within-trophic [DDS^{+}06]. without [NKK^{+}01]. Wolf [BB09]. wollebaeki [VACT^{+}08]. wood [TYD07]. wood-boring [TYD07]. woolly [DL02]. work [DSJT^{+}06, Lin02]. works [RJ03]. World [CMD^{+}09, BBG^{+}08, DS01, HTM01, KWTP06, WLB08]. world-wide [HTM01, KWTP06]. Worldwide [Col03, GSG08]. worm [FCP^{+}07, RPS^{+}06a, THS^{+}02, TP01a, TP01b]. worms [Cal06, Iye02, KKBH03, PAB^{+}02, RGJ^{+}04, RL09b, CFC^{+}00]. would [Fen08]. wrack [MZJ06, RU02]. wrasse [GCS06, Tup07]. wrightii [BH09a, BD01, HHV06, HVP^{+}06, KDAK01, TS06]. wurdemanni [RZLS09].

xanthid [BKB05, LAW06]. xanthophyll [GCD^{+}03]. xanthurus [CRCN07, MT05]. Xenobiotic [PEG^{+}00]. xenophyophores [LMS04]. Xiphias [WKY09]. Xylophaga [TYD07].

Yangtze [CLH^{+}07, SZZ08]. year [AML^{+}03, BGK^{+}06, BM00, BJN^{+}05, BLT03, GBW^{+}05, HPC04, JPMGS08, KSHP00, KG07, KMAC^{+}02, Lap01b, LB05a, MWR07, MBL07, NB06b, OGMC06, SLB07, TSE^{+}08, VGW00, WÓ02, WS04a]. year-class [BM00]. year-classes [AML^{+}03]. year-round [GBW^{+}05]. Years [Kin05a, AHDT^{+}09, BWL06, BCPB06, FS06, FAK07, GSFN00, HBH07, SB03a, VJ03]. yellow [CKB^{+}09, LBD00, MEHD07, PCW05, RRS^{+}09, CPSC01, HK03c, SCD^{+}05, WLS^{+}09a]. yellow-eyed [MEHD07]. yellow-legged [RRS^{+}09]. yellow-nosed [PCW05]. yellowfin [CB08b, CPG08, NW05a]. yellowstripe [KFS^{+}09]. yellowtail [SHK05, WMG02]. yessoensis [SSO^{+}01]. yield [PCYC02, ZHWCS08]. yields [LG01]. Yoldia [STJ03]. yolk [FWF02, LRC06, WPPS04]. yolk-sac [LRC06]. York [SC00b, VFP^{+}07, CGP08, CS08, SW02b, SW02a]. Yoshioka [Som08]. young [BGK^{+}06, BRI^{+}07, BLT03, GCLC^{+}02, KFJ^{+}07, KSHP00, KG07, KMAC^{+}02, KNF^{+}03, Lap01b, LB05a, MRQM06, MH06a, MBL07, SLB07, TSE^{+}08, GRPT^{+}00, GKWR02, RKGH01, SSJ^{+}02]. young-of-the-year [BGK^{+}06, BLT03, KSHP00, KG07, KMAC^{+}02, Lap01b, MBL07, SLB07, TSE^{+}08]. young-of-year [LB05a].

Z. [WPB^{+}08b]. Zalophus [VACT^{+}08]. Zanzibar [KHT^{+}02, Øre00]. Zeacumantus [FMP05]. Zealand [Gar00, AKL^{+}04, BPG08, BN09, BSDR01, CWDG05, CW09, CWBS03, CB08a, Chi09, GGS09, HBDD02, LB01, LEF^{+}06, Lus03, Lus05, Mar04, MAM^{+}01, MDS^{+}07, MMC^{+}08b, Mou04, MBN^{+}09, PMG05a, PMG05b, PMS^{+}06, PPB09b, PH03c, PS06, RRS00, RW08, RW06b, Saf03, SLB^{+}08, TA01a, Wal07, WG01, WGL03, Win09, WG07, ZHWCS08]. zebra [DNL08b]. Zebrasoma [CKB^{+}09]. zelandica [GFP^{+}05, HTL^{+}02, MNP02]. zero [PVGO01, TDRM08]. zero-flow [PVGO01]. zinc [BTF^{+}02, DHBZ00, FBHB03, MW04b, PW08, PW09, RB02]. Zingst [HP06]. zizyphinum [HSCD01]. Zn [LLY^{+}01, NW04, NWT00, RW01, RSL09a, RCC01, SW04, VF09, WLL03, WL03, WG00, WR00a, XW02]. zoanthid [Bur02]. Zoarces [ORS02]. zoea [HM03, RBSP00]. zonalis [KT06]. zonation [ACPS09, CY06, DF02, HJ05, IRM08, JS02, KBP06, MSSP06, MDS^{+}07, PB08a, RvdPHW04, TGM07, WGP^{+}00]. Zone [LCGN02, TRH^{+}06, BA00, BDLB02, BBC^{+}09b, BGS02a, BPA^{+}08, BES^{+}08, BWO04, CSD04, CGF04, CRTP02, CDT^{+}03, CCB^{+}09a, DR00, DRB^{+}01, FKM^{+}01, GJT09, GH08, GKS^{+}07, HJ03, HHD^{+}09, KvNVH00, LC02, LLG^{+}06, MMMB^{+}05, NRJB04, Pee02, PGS^{+}00, RSK04, RHG05, RHGB02, SSH^{+}04, SDW^{+}01, SHFP^{+}03, SHW09, TSF01, TD01b, WKY09, WCRB03, WDB^{+}03, WMD^{+}07, YCSS06, YSS03, ZBNW09, MFE^{+}05]. zones [BT09, BCC^{+}04, CHA06, FDD^{+}03, NW05b, RMPC08, TC01b, WST^{+}09]. zoning [LPS08]. Zoobenthic [BB03a, BB00]. zooplankters [KT01]. zooplanktivore [HDGM01]. zooplanktivorous [Hip08, PS01]. Zooplanktivory [VFV01]. Zooplankton [EZTZ02, GB05a, HHM^{+}06, HGZ^{+}00, JK04, MFG^{+}00, MG09b, NHNB01, TCD01, TLNH01, TBH06, ABF07, AI04, AVS03, BBZM08, BHH09, BSKR^{+}06, BMSD00, BRB^{+}00, CD03a, CIHF01, CDML06, DBP04, DMV07, DCM^{+}06, DCC^{+}03, DC07b, FLG02, FKA^{+}06, FP08, GGMGR07, GB05b, HCH07, HMKR05, HM03, HB07b, ICH04, JHJ^{+}03, LHK06, LCB^{+}05, MRS00, MCD^{+}05, MDZ05, NCMC09, NWT00, NH03a, NH03b, OAGV07, PGR^{+}01, PFW^{+}02, PP00, PK06, Pur03, RW01, RF08b, RHS01b, SD09, SHB03, SS02a, SS06, SNCP04, STRB06, TRDV00, TNHL02, TDP^{+}00, Ven06, vDvHT^{+}09]. zooplankton-mediated [OAGV07]. zoospores [PPHT00, WGP^{+}00]. zooxanthella [Gou07, WBZ^{+}09]. Zooxanthellae [RGL01, AC02, Bur02, CBM00, DBT^{+}02a, HR07, PBAA04, PLM03, UHR05, UBRvO06]. zooxanthellar [SHH08]. zooxanthellate [FD00, RFPS^{+}02, RMH^{+}09, SPT04]. Zostera [Abd03a, Abd07, ASS05, AW03b, BRAS03, BN01, BWSH00, BHV^{+}02, BVN^{+}03, BOM^{+}08, BN02, CAS05, CDS^{+}04, DAM^{+}08, DCF^{+}05, EWR08, FSO08, HR02a, HUP^{+}00, HCV03a, HCA06, HB08b, JBS08, JKG^{+}05, JYS06, JNMB08, KKIV01, Lar03, LPK^{+}07, MT09, MCS05, MF03, MSTS^{+}06, NSBK05, OFHM03, PZ07, PBPLB06, PA06, RS02, RH06, RPO00, RS07a, SASR03, SDK^{+}02, SH06b, SCD^{+}09, TJA08, WPG08, WBN^{+}00, WKOE09, WPB^{+}08b, ZSKA01, ZCR^{+}09, vKH00]. zosterae [MT09, RS02].

References

Andrews:2004:HDR

[AA04] K. S. Andrews and T. W. Anderson. Habitat-dependent recruitment of two temperate reef fishes at multiple spatial scales. Marine Ecology Progress Series, 277:231–244, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/277/m277p231.pdf.

Arrontes:2004:EGL

[AAF^{+}04] J. Arrontes, F. Arenas, C. Fernández, J. M. Rico, J. Oliveros, B. Martínez, R. M. Viejo, and D. Alvarez. Effect of grazing by limpets on mid-shore species assemblages in northern Spain. Marine Ecology Progress Series, 277:117–133, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p117.pdf.

Adjeroud:2002:NDI

[AAGS02] M. Adjeroud, D. Augustin, R. Galzin, and B. Salvat. Natural disturbances and interannual variability of coral reef communities on the outer slope of Tiahura (Moorea, French Polynesia): 1991 to 1997. Marine Ecology Progress Series, 237:121–131, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p121.pdf.

Alberto:2006:GDC

[AAHDS06] F. Alberto, S. Arnaud-Haond, C. M. Duarte, and E. A. Serrão. Genetic diversity of a clonal angiosperm near its range limit: the case of Cymodocea nodosa at the Canary Islands. Marine Ecology Progress Series, 309:117–129, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p117.pdf.

Ashford:2007:OCR

[AAJ07] J. R. Ashford, A. I. Arkhipkin, and C. M. Jones. Otolith chemistry reflects frontal systems in the Antarctic Circumpolar Current. Marine Ecology Progress Series, 351:249– 260, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/351/m351p249.pdf.

Adjeroud:2000:PFD

[AAPO00] M. Adjeroud, S. Andréfouët, C. Payri, and J. Orempüller. Physical factors of differentiation in macrobenthic communities between atoll lagoons in the Central Tuamotu Archipelago (French Polynesia). Marine Ecology Progress Series, 196: 25–38, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/196/m196p025.pdf.

Aoki:2007:DDB

[AAY^{+}07] K. Aoki, M. Amano, M. Yoshioka, K. Mori, D. Tokuda, and N. Miyazaki. Diel diving behavior of sperm whales off Japan. Marine Ecology Progress Series, 349:277–287, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p277.pdf.

Ackerman:2000:RFA

[AB00] J. L. Ackerman and D. R. Bellwood. Reef fish assemblages: a re-evaluation using enclosed rotenone stations. Marine Ecology Progress Series, 206:227–237, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/206/m206p227.pdf.

Arhonditsis:2004:ECS

[AB04] G. B. Arhonditsis and M. T. Brett. Evaluation of the current state of mechanistic aquatic biogeochemical modeling. Marine Ecology Progress Series, 271:13–26, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p013.pdf.

Auth:2006:DCS

[AB06] T. D. Auth and R. D. Brodeur. Distribution and community structure of ichthyoplankton off the coast of Oregon, USA, in 2000 and 2002. Marine Ecology Progress Series, 319:199–213, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p199.pdf.

Akcha:2000:IEB

[ABB^{+}00] F. Akcha, T. Burgeot, H. Budzinski, A. Pfohl-Leszkowicz, and J. F. Narbonne. Induction and elimination of bulky benzo[a]pyrene-related DNA adducts and 8-oxodGuo in mussels Mytilus galloprovincialis exposed in vivo to B[a]Pcontaminated feed. Marine Ecology Progress Series, 205:

195–206, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p195.pdf.

Azouzi:2002:SVI

[ABB02] L. Azouzi, E. Bourget, and D. Borcard. Spatial variation in the intertidal bivalve Macoma balthica: biotic variables in relation to density and abiotic factors. Marine Ecology Progress Series, 234:159–170, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p159.pdf.

Ateweberhan:2005:SDS

[ABB05] M. Ateweberhan, J. H. Bruggemann, and A. M. Breeman. Seasonal dynamics of Sargassum ilicifolium (Phaeophyta) on a shallow reef flat in the southern Red Sea (Eritrea). Marine Ecology Progress Series, 292:159–171, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p159.pdf.

Alunno-Bruscia:2001:SAL

[ABBF01] M. Alunno-Bruscia, E. Bourget, and M. Fréchette. Shell allometry and length-mass-density relationship for Mytilus edulis in an experimental food-regulated situation. Marine Ecology Progress Series, 219:177–188, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p177.pdf.

Agedah:2009:STB

[ABBN09] E. C. Agedah, H. E. Binalaiyifa, A. S. Ball, and D. B. Nedwell. Sources, turnover and bioavailability of dissolved organic nitrogen (DON) in the Colne estuary, UK. Marine Ecology Progress Series, 382:23–33, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/382/m382p023.pdf.

Airoldi:2005:IRH

[ABC^{+}05] L. Airoldi, F. Bacchiocchi, C. Cagliola, F. Bulleri, and M. Abbiati. Impact of recreational harvesting on assemblages in artificial rocky habitats. Marine Ecology Progress Series, 299: 55–66, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p055.pdf.

Arvanitidis:2002:SBP

[ABD^{+}02] C. Arvanitidis, G. Bellan, P. Drakopoulos, V. Valavanis, C. Dounas, A. Koukouras, and A. Eleftheriou. Seascape biodiversity patterns along the Mediterranean and the Black Sea: lessons from the biogeography of benthic polychaetes. Marine Ecology Progress Series, 244:139–152, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p139.pdf.

Abdelrhman:2003:EEZ

[Abd03a] M. A. Abdelrhman. Effect of eelgrass Zostera marina canopies on flow and transport. Marine Ecology Progress Series, 248: 67–83, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p067.pdf.

Andersen:2003:GPS

[ABD^{+}03b] L. W. Andersen, E. W. Born, R. Dietz, T. Haug, N. Øien, and C. Bendixen. Genetic population structure of minke whales Balaenoptera acutorostrata from Greenland, the North East Atlantic and the North Sea probably reflects different ecological regions. Marine Ecology Progress Series, 247:263– 280, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p263.pdf.

Abdelrhman:2007:MCB

[Abd07] M. A. Abdelrhman. Modeling coupling between eelgrass Zostera marina and water flow. Marine Ecology Progress Series, 338:81–96, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p081.pdf.

Agostini:2007:LSC

[ABF07] V. N. Agostini, A. Bakun, and R. C. Francis. Larval stage controls on Pacific sardine recruitment variability: high zooplankton abundance linked to poor reproductive success. Marine Ecology Progress Series, 345:237–244, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p237.pdf.

Alexander:2009:RBM

[ABHE09] T. J. Alexander, N. Barrett, M. Haddon, and G. Edgar. Relationships between mobile macroinvertebrates and reef structure in a temperate marine reserve. Marine Ecology Progress Series, 389:31–44, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p031.pdf.

Arthur:2008:OCD

[ABL08] K. E. Arthur, M. C. Boyle, and C. J. Limpus. Ontogenetic changes in diet and habitat use in green sea turtle (Chelonia mydas) life history. Marine Ecology Progress Series, 362:303–311, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/362/m362p303.pdf.

Adey:2009:SMT

[ABS^{+}09] E. A. Adey, K. D. Black, T. Sawyer, T. M. Shimmield, and C. N. Trueman. Scale microchemistry as a tool to investigate the origin of wild and farmed Salmo salar. Marine Ecology Progress Series, 390:225–235, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/390/m390p225.pdf.

Archambault:2001:TVS

[ABU01] P. Archambault, K. Banwell, and A. J. Underwood. Temporal variation in the structure of intertidal assemblages following the removal of sewage. Marine Ecology Progress Series, 222:51–62, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p051.pdf.

Anderson:2006:CEC

[ABWK06] C. R. Anderson, M. A. Brzezinski, L. Washburn, and R. Kudela. Circulation and environmental conditions during a toxigenic Pseudo-nitzschia australis bloom in the Santa Barbara Channel, California. Marine Ecology Progress Series, 327:119–133, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/327/m327p119.pdf.

Arifin:2000:ISF

[ABY00] Z. Arifin and L. I. Bendell-Young. Influence of a selective feeding behaviour by the blue mussel Mytilus trossulus on the assimilation of 109Cd from environmentally relevant seston matrices. Marine Ecology Progress Series, 192:181–193, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p181.pdf.

Annis:2002:APA

[AC02] E. R. Annis and C. B. Cook. Alkaline phosphatase activity in symbiotic dinoflagellates (zooxanthellae) as a biological indicator of environmental phosphate exposure. Marine Ecology Progress Series, 245:11–20, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p011.pdf.

Allen:2007:EDT

[AC07] J. I. Allen and K. R. Clarke. Effects of demersal trawling on ecosystem functioning in the North Sea: a modelling study. Marine Ecology Progress Series, 336:63–75, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p063.pdf.

Atkinson:2008:PEO

[AC08] M. J. Atkinson and P. Cuet. Possible effects of ocean acidification on coral reef biogeochemistry: topics for research. Marine Ecology Progress Series, 373:249–256, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p249.pdf.

Aguzzi:2007:OCV

[ACAG07] J. Aguzzi, J. B. Company, P. Abelló, and J. A. García. Ontogenetic changes in vertical migratory rhythms of benthopelagic shrimps Pasiphaea multidentata and P. sivado. Marine Ecology Progress Series, 335:167–174, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p167.pdf.

Anderson:2007:PMV

[ACH07] T. W. Anderson, M. H. Carr, and M. A. Hixon. Patterns and mechanisms of variable settlement and recruitment of a coral reef damselfish, Chromis cyanea. Marine Ecology Progress Series, 350:109–116, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/350/m350p109.pdf.

Arrizabalaga:2004:PSA

[ACJ^{+}04] H. Arrizabalaga, E. Costas, J. Juste, A. González-Garcés, B. Nieto, and V. López-Rodas. Population structure of albacore Thunnus alalunga inferred from blood groups and tagrecapture analyses. Marine Ecology Progress Series, 282: 245–252, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p245.pdf.

Archer:2009:PTI

[ACLF09] S. D. Archer, D. G. Cummings, C. A. Llewellyn, and J. R. Fishwick. Phytoplankton taxa, irradiance and nutrient availability determine the seasonal cycle of DMSP in temperate shelf seas. Marine Ecology Progress Series, 394:111– 124, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/394/m394p111.pdf.

Alldredge:2002:OMF

[ACM^{+}02] A. L. Alldredge, T. J. Cowles, S. MacIntyre, J. E. B. Rines, P. L. Donaghay, C. F. Greenlaw, D. V. Holliday, M. M. Dekshenieks, J. M. Sullivan, and J. R. V. Zaneveld. Occurrence and mechanisms of formation of a dramatic thin layer of marine snow in a shallow Pacific fjord. Marine Ecology Progress Series, 233:1–12, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p001.pdf.

Arantes:2009:DWM

[ACPS09] R. C. M. Arantes, C. B. Castro, D. O. Pires, and J. C. S. Seoane. Depth and water mass zonation and species associations of cold-water octocoral and stony coral communities in the southwestern Atlantic. Marine Ecology Progress Series, 397:71–79, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p071.pdf.

Amar:2007:CPD

[ACR07] K. O. Amar, N. E. Chadwick, and B. Rinkevich. Coral planulae as dispersion vehicles: biological properties of larvae released early and late in the season. Marine Ecology Progress Series, 350:71–78, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/350/m350p071.pdf.

Adams:2009:ESE

[ACT09] L. M. Adams, V. R. Cumbo, and M. Takabayashi. Exposure to sediment enhances primary acquisition of Symbiodinium by asymbiotic coral larvae. Marine Ecology Progress Series, 377:149–156, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/377/m377p149.pdf.

Agusti:2000:EIL

[AD00] S. Agustí and C. M. Duarte. Experimental induction of a large phytoplankton bloom in Antarctic coastal waters. Marine Ecology Progress Series, 206:73–85, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/206/m206p073.pdf.

Agawin:2000:RMS

[ADA00] N. S. R. Agawin, C. M. Duarte, and S. Agustí. Response of Mediterranean Synechococcus growth and loss rates to experimental nutrient inputs. Marine Ecology Progress Series, 206:97–106, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/206/m206p097.pdf.

Adams:2001:URE

[Ada01] N. L. Adams. UV radiation evokes negative phototaxis and covering behavior in the sea urchin Strongylocentrotus droebachiensis. Marine Ecology Progress Series, 213:87–95, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
213/m213p087.pdf.

Ainley:2009:APP

[ADF^{+}09] D. G. Ainley, K. D. Dugger, R. G. Ford, S. D. Pierce, D. C. Reese, R. D. Brodeur, C. T. Tynan, and J. A. Barth. Association of predators and prey at frontal features in the California Current: competition, facilitation, and co-occurrence. Marine Ecology Progress Series, 389:271–294, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p271.pdf.

Adams:2006:NFT

[ADK^{+}06] A. J. Adams, C. P. Dahlgren, G. T. Kellison, M. S. Kendall, C. A. Layman, J. A. Ley, I. Nagelkerken, and J. E. Serafy. Nursery function of tropical back-reef systems. Marine Ecology Progress Series, 318:287–301, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p287.pdf.

Astolfi:2005:MVS

[ADR^{+}05] L. Astolfi, I. Dupanloup, R. Rossi, P. M. Bisol, E. Faure, and L. Congiu. Mitochondrial variability of sand smelt Atherina boyeri populations from north Mediterranean coastal lagoons. Marine Ecology Progress Series, 297:233–243, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p233.pdf.

Aksnes:2009:CWD

[ADS^{+}09] D. L. Aksnes, N. Dupont, A. Staby, Ó. Fiksen, S. Kaartvedt, and J. Aure. Coastal water darkening and implications for mesopelagic regime shifts in Norwegian fjords. Marine Ecology Progress Series, 387:39–49, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m387p039.pdf.

Adams:2002:UBR

[AE02] A. J. Adams and J. P. Ebersole. Use of back-reef and lagoon habitats by coral reef fishes. Marine Ecology Progress Series, 228:213–226, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p213.pdf.

Avery:2006:MDR

[AE06] R. Avery and R. J. Etter. Microstructural differences in the reinforcement of a gastropod shell against predation. Marine Ecology Progress Series, 323:159–170, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p159.pdf.

Abril:2003:CDT

[AED^{+}03] G. Abril, H. Etcheber, B. Delille, M. Frankignoulle, and A. V. Borges. Carbonate dissolution in the turbid and eutrophic Loire estuary. Marine Ecology Progress Series, 259: 129–138, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/259/m259p129.pdf.

Alberti:2007:LGV

[AED^{+}07] J. Alberti, M. Escapa, P. Daleo, O. Iribarne, B. R. Silliman, and M. Bertness. Local and geographic variation in grazing intensity by herbivorous crabs in SW Atlantic salt marshes. Marine Ecology Progress Series, 349:235–243, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p235.pdf.

Armitage:2006:PPD

[AF06] A. R. Armitage and P. Fong. Predation and physical disturbance by crabs reduce the relative impacts of nutrients in a tidal mudflat. Marine Ecology Progress Series, 313: 205–213, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/313/m313p205.pdf.

Anderson:2004:QMS

[AFFH04] M. J. Anderson, R. B. Ford, D. A. Feary, and C. Honeywill. Quantitative measures of sedimentation in an estuarine system and its relationship with intertidal soft-sediment infauna. Marine Ecology Progress Series, 272:33–48, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p033.pdf.

Awkerman:2005:FAS

[AFHA05] J. A. Awkerman, A. Fukuda, H. Higuchi, and D. J. Anderson. Foraging activity and submesoscale habitat use of waved albatrosses Phoebastria irrorata during chick-brooding period. Marine Ecology Progress Series, 291:289–300, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p289.pdf.

Afonso:2008:SSD

[AFHS08] P. Afonso, J. Fontes, K. N. Holland, and R. S. Santos. Social status determines behaviour and habitat usage in a temperate parrotfish: implications for marine reserve design. Marine Ecology Progress Series, 359:215–227, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/359/
m359p215.pdf.

Afonso:2009:MSP

[AFHS09] P. Afonso, J. Fontes, K. N. Holland, and R. S. Santos. Multiscale patterns of habitat use in a highly mobile reef fish, the white trevally Pseudocaranx dentex, and their implications for marine reserve design. Marine Ecology Progress Series, 381:273–286, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p273.pdf.

Awruch:2008:NLA

[AFPS08] C. A. Awruch, S. D. Frusher, N. W. Pankhurst, and J. D. Stevens. Non-lethal assessment of reproductive characteristics for management and conservation of sharks. Marine Ecology Progress Series, 355:277–285, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/355/m355p277.pdf.

Ammerman:2000:CUM

[AG00] J. W. Ammerman and W. B. Glover. Continuous underway measurement of microbial ectoenzyme activities in aquatic ecosystems. Marine Ecology Progress Series, 201:1–12, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p001.pdf.

Acevedo-Gutierrez:2002:IBM

[AG02a] A. Acevedo-Gutiérrez. Interactions between marine predators: dolphin food intake is related to number of sharks. Marine Ecology Progress Series, 240:267–271, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p267.pdf.

An:2002:DNR

[AG02b] S. An and W. S. Gardner. Dissimilatory nitrate reduction to ammonium (DNRA) as a nitrogen link, versus denitrification as a sink in a shallow estuary (Laguna Madre/ Baffin Bay, Texas). Marine Ecology Progress Series, 237: 41–50, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p041.pdf.

Antajan:2004:ACI

[AG04] E. Antajan and S. Gasparini. Assessment of cryptophyceae ingestion by copepods using alloxanthin pigment: a caution. Marine Ecology Progress Series, 274:191–198, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p191.pdf.

Archambault:2003:RED

[AGB03] M. C. Archambault, J. Grant, and V. M. Bricelj. Removal efficiency of the dinoflagellate Heterocapsa triquetra by phosphatic clay , and implications for the mitigation of harmful algal blooms. Marine Ecology Progress Series, 253:97– 109, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p097.pdf.

Alonso-Gutierrez:2008:ASD

[AGCF^{+}08] J. Alonso-Gutiérrez, M. M. Costa, A. Figueras, J. Albaigés, M. Viñas, A. M. Solanas, and B. Novoa. Alcanivorax strain detected among the cultured bacterial community from sediments affected by the ‘Prestige’ oil spill. Marine Ecology Progress Series, 362:25–36, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p025.pdf.

Arias-Gonzalez:2006:TRE

[AGDP^{+}06] J. E. Arias-González, T. J. Done, C. A. Page, A. Cheal, S. Kininmonth, and J. R. Garza-Pérez. Towards a reefscape ecology: relating biomass and trophic structure of fish assemblages to habitat at Davies Reef, Australia. Marine Ecology Progress Series, 320:29–41, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/320/m320p029.pdf.

Arias-Gonzalez:2006:TFP

[AGM06] J. E. Arias-González and S. Morand. Trophic functioning with parasites: a new insight for ecosystem analysis. Marine Ecology Progress Series, 320:43–53, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p043.pdf.

Andre:2009:ETE

[AGS^{+}09] J. André, E. P. M. Grist, J. M. Semmens, G. T. Pecl, and S. Segawa. Effects of temperature on energetics and the growth pattern of benthic octopuses. Marine Ecology Progress Series, 374:167–179, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/374/m374p167.pdf.

Airoldi:2007:NES

[AH07] L. Airoldi and S. J. Hawkins. Negative effects of sediment deposition on grazing activity and survival of the limpet Patella vulgata. Marine Ecology Progress Series, 332:235– 240, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/332/m332p235.pdf.

Albins:2008:IIP

[AH08] M. A. Albins and M. A. Hixon. Invasive Indo–Pacific lionfish Pterois volitans reduce recruitment of Atlantic coralreef fishes. Marine Ecology Progress Series, 367:233–238, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/367/m367p233.pdf.

Allen:2007:GDN

[AHBM^{+}07] D. M. Allen, S. S. Haertel-Borer, B. J. Milan, D. Bushek, and R. F. Dame. Geomorphological determinants of nekton use of intertidal salt marsh creeks. Marine Ecology Progress Series, 329:57–71, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/329/m329p057.pdf.

Ainley:2006:HVA

[AHC^{+}06] D. G. Ainley, K. A. Hobson, X. Crosta, G. H. Rau, L. I. Wassenaar, and P. C. Augustinus. Holocene variation in the Antarctic coastal food web: linking δ D and δ ^{13}C in snow petrel diet and marine sediments. Marine Ecology Progress Series, 306:31–40, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p031.pdf.

Arnaud-Haond:2009:GRM

[AHDT^{+}09] S. Arnaud-Haond, C. M. Duarte, S. Teixeira, S. I. Massa, J. Terrados, N. H. Tri, P. N. Hong, and E. A. Serrão. Genetic recolonization of mangrove: genetic diversity still increasing in the Mekong Delta 30 years after Agent Orange. Marine Ecology Progress Series, 390:129–135, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p129.pdf.

Atkins:2002:TFP

[AHK^{+}02] M. S. Atkins, M. A. Hanna, E. A. Kupetsky, M. A. Saito, C. D. Taylor, and C. O. Wirsen. Tolerance of flagellated protists to high sulfide and metal concentrations potentially encountered at deep-sea hydrothermal vent. Marine Ecology Progress Series, 226:63–75, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p063.pdf.

Ahrens:2001:RDS

[AHL^{+}01] M. J. Ahrens, J. Hertz, E. M. Lamoureux, G. R. Lopez, A. E. McElroy, and B. J. Brownawell. The role of digestive surfactants in determining bioavailability of sediment-bound hydrophobic organic contaminants to 2 deposit-feeding polychaetes. Marine Ecology Progress Series, 212:145–157, 2001.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p145.pdf.

Al-Habsi:2008:ESS

[AHSPG08] S. H. Al-Habsi, C. J. Sweeting, N. V. C. Polunin, and N. A. J. Graham. δ ^{15}N and δ ^{13}C elucidation of size-structured food webs in a Western Arabian Sea demersal trawl assemblage. Marine Ecology Progress Series, 353:55–63, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p055.pdf.

Aronson:2001:MPT

[AHV01] R. B. Aronson, K. L. Heck, Jr., and J. F. Valentine. Measuring predation with tethering experiments. Marine Ecology Progress Series, 214:311–312, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p311.pdf.

Arndt:2003:DSH

[AHW03] H. Arndt, K. Hausmann, and M. Wolf. Deep-sea heterotrophic nanoflagellates of the Eastern Mediterranean Sea: qualitative and quantitative aspects of their pelagic and benthic occurrence. Marine Ecology Progress Series, 256:45– 56, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/256/m256p045.pdf.

Anderson:2003:QPP

[AHZ03] J. T. Anderson, R. R. Hood, and X. Zhang. Quantification of Pfiesteria piscicida growth and encystment parameters using a numerical model. Marine Ecology Progress Series, 246:105–113, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/246/m246p105.pdf.

Abril:2002:MDS

[AI02] G. Abril and N. Iversen. Methane dynamics in a shallow non-tidal estuary (Randers Fjord, Denmark). Marine Ecology Progress Series, 230:171–181, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/230/m230p171.pdf.

Albaina:2004:RBF

[AI04] A. Albaina and X. Irigoien. Relationships between frontal structures and zooplankton communities along a cross-shelf transect in the Bay of Biscay (1995 to 2003). Marine Ecology Progress Series, 284:65–75, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/284/m284p065.pdf.

Amsler:2005:CEP

[AIM^{+}05] C. D. Amsler, K. Iken, J. B. McClintock, M. O. Amsler, K. J. Peters, J. M. Hubbard, F. B. Furrow, and B. J. Baker. Comprehensive evaluation of the palatability and chemical defenses of subtidal macroalgae from the Antarctic Peninsula. Marine Ecology Progress Series, 294:141–159, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p141.pdf.

Airoldi:2000:RAD

[Air00] L. Airoldi. Responses of algae with different life histories to temporal and spatial variability of disturbance in subtidal reefs. Marine Ecology Progress Series, 195:81–92, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p081.pdf.

Arnosti:2003:HAL

[AJ03] C. Arnosti and B. B. Jørgensen. High activity and low temperature optima of extracellular enzymes in Arctic sediments: implications for carbon cycling by heterotrophic microbial communities. Marine Ecology Progress Series, 249: 15–24, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/249/m249p015.pdf.

Aikins:2002:GPA

[AK02] S. Aikins and E. Kikuchi. Grazing pressure by amphipods on microalgae in Gamo Lagoon, Japan. Marine Ecology Progress Series, 245:171–179, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/245/m245p171.pdf.

Atema:2002:LRF

[AKG02] J. Atema, M. J. Kingsford, and G. Gerlach. Larval reef fish could use odour for detection, retention and orientation to reefs. Marine Ecology Progress Series, 241:151– 160, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/241/m241p151.pdf.

Arai:2004:EDH

[AKL^{+}04] T. Arai, A. Kotake, P. M. Lokman, M. J. Miller, and K. Tsukamoto. Evidence of different habitat use by New Zealand freshwater eels Anguilla australis and A. dieffenbachii, as revealed by otolith microchemistry. Marine Ecology Progress Series, 266:213–225, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/266/m266p213.pdf.

Allen:2009:LSE

[AKM09] R. M. Allen, P. J. Krug, and D. J. Marshall. Larval size in Elysia stylifera is determined by extra-embryonic provisioning but not egg size. Marine Ecology Progress Series, 389:127–137, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p127.pdf.

Aguilera:2002:IIL

[AKRA02] A. Aguilera, B. A. Keafer, G. H. Rau, and D. M. Anderson. Immunomagnetic isolation of live and preserved Alexandrium fundyense cells: species-specific physiological, chemical, and isotopic analyses. Marine Ecology Progress Series, 237:65–78, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p065.pdf.

Andre:2003:NIS

[AL03a] J. André and I. R. Lawler. Near infrared spectroscopy as a rapid and inexpensive means of dietary analysis for the marine herbivore, dugong Dugong dugon. Marine Ecology Progress Series, 257:259–266, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/257/m257p259.pdf.

Apprill:2003:EUR

[AL03b] A. M. Apprill and M. P. Lesser. Effects of ultraviolet radiation on Laminaria saccharina in relation to depth and tidal height in the Gulf of Maine. Marine Ecology Progress Series, 256:75–85, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p075.pdf.

Anderson:2008:GWM

[AL08] E. M. Anderson and J. R. Lovvorn. Gray whales may increase feeding opportunities for avian benthivores. Marine Ecology Progress Series, 360:291–296, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p291.pdf.

Auster:2009:CRF

[AL09] P. J. Auster and J. S. Link. Compensation and recovery of feeding guilds in a northwest Atlantic shelf fish community. Marine Ecology Progress Series, 382:163–172, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/382/
m382p163.pdf.

Anderson:2009:UPD

[ALE^{+}09] E. M. Anderson, J. R. Lovvorn, D. Esler, W. S. Boyd, and K. C. Stick. Using predator distributions, diet, and condition to evaluate seasonal foraging sites: sea ducks and herring spawn. Marine Ecology Progress Series, 386:287– 302, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/386/m386p287.pdf.

Arai:2001:RMT

[ALOT01] T. Arai, D. Limbong, T. Otake, and K. Tsukamoto. Recruitment mechanisms of tropical eels Anguilla spp. and implications for the evolution of oceanic migration in the genus Anguilla. Marine Ecology Progress Series, 216:253–264, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p253.pdf.

Assmann:2000:CDC

[ALPK00] M. Assmann, E. Lichte, J. R. Pawlik, and M. Köck. Chemical defenses of the Caribbean sponges Agelas wiedenmayeri and Agelas conifera. Marine Ecology Progress Series, 207:255–262, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p255.pdf.

Altieri:2006:IVH

[Alt06] A. H. Altieri. Inducible variation in hypoxia tolerance across the intertidal-subtidal distribution of the blue mussel Mytilus edulis. Marine Ecology Progress Series, 325:295– 300, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/325/m325p295.pdf.

Alonzo:2001:SSG

[AM01] S. H. Alonzo and M. Mangel. Survival strategies and growth of krill: avoiding predators in space and time. Marine Ecology Progress Series, 209:203–217, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p203.pdf.

Alcoverro:2002:ESU

[AM02a] T. Alcoverro and S. Mariani. Effects of sea urchin grazing on seagrass (Thalassodendron ciliatum) beds of a Kenyan lagoon. Marine Ecology Progress Series, 226:255–263, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p255.pdf.

Alonzo:2002:ESB

[AM02b] S. H. Alonzo and M. Mangel. Effects of social behavior on survival and growth of krill: important, but how relevant? Marine Ecology Progress Series, 244:309–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p309.pdf.

Albessard:2003:ITC

[AM03] E. Albessard and P. Mayzaud. Influence of tropho-climatic environment and reproduction on lipid composition of the euphausiid Meganyctiphanes norvegica in the Ligurian Sea, the Clyde Sea and the Kattegat. Marine Ecology Progress Series, 253:217–232, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p217.pdf.

Ayre:2004:WDC

[AM04] D. J. Ayre and K. J. Miller. Where do clonal coral larvae go? Adult genotypic diversity conflicts with reproductive effort in the brooding coral Pocillopora damicornis. Marine Ecology Progress Series, 277:95–105, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p095.pdf.

Ayre:2006:RMB

[AM06] D. J. Ayre and K. Miller. Random mating in the brooding coral Acropora palifera. Marine Ecology Progress Series, 307:155–160, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p155.pdf.

Andersson:2008:LMI

[AMB08] A. J. Andersson, F. T. Mackenzie, and N. R. Bates. Life on the margin: implications of ocean acidification on Mg-calcite, high latitude and cold-water marine calcifiers. Marine Ecology Progress Series, 373:265–273, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m373p265.pdf.

Amara:2007:GCI

[AMG^{+}07] R. Amara, T. Meziane, C. Gilliers, G. Hermel, and P. Laffargue. Growth and condition indices in juvenile sole Solea solea measured to assess the quality of essential fish habitat. Marine Ecology Progress Series, 351:201–208, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p201.pdf.

Agnew:2003:IDP

[AML^{+}03] D. J. Agnew, T. R. Marlow, K. Lorenzen, J. Pompert, R. C. Wakeford, and G. A. Tingley. Influence of Drake Passage oceanography on the parasitic infection of individual yearclasses of southern blue whiting Micromesistius australis. Marine Ecology Progress Series, 254:281–291, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p281.pdf.

Alcaraz:2002:ETC

[AMP^{+}02] M. Alcaraz, C. Marrasé, F. Peters, L. Arin, and A. Malits. Effects of turbulence conditions on the balance between production and respiration in marine planktonic communities. Marine Ecology Progress Series, 242:63–71, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/242/
m242p063.pdf.

Alcoverro:2000:NMB

[AMR00a] T. Alcoverro, M. Manzanera, and J. Romero. Nutrient mass balance of the seagrass Posidonia oceanica: the importance of nutrient retranslocation. Marine Ecology Progress Series, 194:13–21, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p013.pdf.

Alonzo:2000:EPD

[AMR00b] F. Alonzo, P. Mayzaud, and S. Razouls. Egg-production dynamics, biochemical composition and hatching success of the subantarctic copepod Paraeuchaeta antarctica: laboratory studies. Marine Ecology Progress Series, 205:219–227, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p219.pdf.

Alonzo:2000:EPP

[AMR00c] F. Alonzo, P. Mayzaud, and S. Razouls. Egg production, population structure and biochemical composition of the subantarctic copepod Paraeuchaeta antarctica in the Kerguelen Archipelago. Marine Ecology Progress Series, 205:207– 217, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p207.pdf.

Alcoverro:2001:AMC

[AMR01a] T. Alcoverro, M. Manzanera, and J. Romero. Annual metabolic carbon balance of the seagrass Posidonia oceanica: the importance of carbohydrate reserves. Marine Ecology Progress Series, 211:105–116, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/211/m211p105.pdf.

Alonzo:2001:EPE

[AMR01b] F. Alonzo, P. Mayzaud, and S. Razouls. Egg production and energy storage in relation to feeding conditions in the subantarctic calanoid copepod Drepanopus pectinatus: an experimental study of reproductive strategy. Marine Ecology Progress Series, 209:231–242, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p231.pdf.

Alberto:2001:GHS

[AMS01] F. Alberto, L. Mata, and R. Santos. Genetic homogeneity in the seagrass Cymodocea nodosa at its northern Atlantic limit revealed through RAPD. Marine Ecology Progress Series, 221:299–301, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/221/m221p299.pdf.

Araujo:2006:MFW

[AMS^{+}06] J. N. Araújo, S. Mackinson, R. J. Stanford, D. W. Sims, A. J. Southward, S. J. Hawkins, J. R. Ellis, and P. J. B. Hart. Modelling food web interactions, variation in plankton production, and fisheries in the western English Channel ecosystem. Marine Ecology Progress Series, 309:175–187, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p175.pdf.

Anderson:2003:MMR

[AMT03a] I. C. Anderson, K. J. McGlathery, and A. C. Tyler. Microbial mediation of ’reactive’ nitrogen transformations in a temperate lagoon. Marine Ecology Progress Series, 246: 73–84, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p073.pdf.

Arai:2003:LDT

[AMT03b] T. Arai, M. J. Miller, and K. Tsukamoto. Larval duration of the tropical eel Anguilla celebesensis from Indonesian and Philippine coasts. Marine Ecology Progress Series, 251:255–261, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p255.pdf.

Aiken:2007:ASL

[ANCC07] C. M. Aiken, S. A. Navarrete, M. I. Castillo, and J. C. Castilla. Along-shore larval dispersal kernels in a numerical ocean model of the central Chilean coast. Marine Ecology Progress Series, 339:13–24, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/339/m339p013.pdf.

Abed-Navandi:2005:FST

[AND05] D. Abed-Navandi and P. C. Dworschak. Food sources of tropical thalassinidean shrimps: a stable-isotope study. Marine Ecology Progress Series, 291:159–168, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p159.pdf.

Axelsen:2000:APD

[ANF^{+}00] B. E. Axelsen, L. Nøttestad, A. Fernö, A. Johannessen, and O. A. Misund. ’Await’ in the pelagic: dynamic trade-off between reproduction and survival within a herring school splitting vertically during spawning. Marine Ecology Progress Series, 205:259–269, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p259.pdf.

Anonymous:2000:CAT

[Ano00a] Anonymous. Combined author and title index: Volumes 191 to 200 (1999–2000). Marine Ecology Progress Series, 200:303– 315, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2000:ELUb

[Ano00b] Anonymous. Erratum: López–Urrutia, Á., Acuña, J. L. 191:195–205 (1999). Marine Ecology Progress Series, 193:310, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See [LUA99].

Anonymous:2000:ELUa

[Ano00c] Anonymous. Erratum: López-Urrutia, Á, Acuña, J. L., 191:195–205 (1999). Marine Ecology Progress Series, ??:??, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See [LUA99].

Anonymous:2000:OKM

[Ano00d] Anonymous. Otto Kinne MEPS celebrates: Volume 200 in 2000 MEPS 200:1. Marine Ecology Progress Series, ??: ??, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/200/m200p001.pdf.

Anonymous:2001:BPD

[Ano01a] Anonymous. Best PB distribution and population separation of Bryde’s whale Balaenoptera edeni off southern Africa. Marine Ecology Progress Series, 220:277–289, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p277.pdf.

Anonymous:2001:CATa

[Ano01b] Anonymous. Combined author and title index: Volumes 201 to 210 (2000–2001). Marine Ecology Progress Series, 210:303–314, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/authorindex/m210p303.pdf.

Anonymous:2001:CATb

[Ano01c] Anonymous. Combined author and title index: Volumes 211 to 220 (2001). Marine Ecology Progress Series, 220: 297–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m220p297.pdf.

Anonymous:2001:EGA

[Ano01d] Anonymous. Erratum: Granata et al., 218:95–106. Marine Ecology Progress Series, 222:309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:/

/www.int-res.com/articles/meps/222/m222p309.pdf. See

[GSC^{+}01].

Anonymous:2001:ESA

[Ano01e] Anonymous. Erratum: Schmaljohann et al., 211:27–42 (2001). Marine Ecology Progress Series, 220:295, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p295.pdf. See [SDW^{+}01].

Anonymous:2001:EWA

[Ano01f] Anonymous. Erratum: Williams et al., 203:81–94 (2000). Marine Ecology Progress Series, 209:311, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
209/m209p311.pdf. See [WDN00].

Anonymous:2002:CATa

[Ano02a] Anonymous. Combined author and title index: Volumes 221 to 230 (2001–2002). Marine Ecology Progress Series, 230:303–314, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p303.pdf.

Anonymous:2002:CATb

[Ano02b] Anonymous. Combined author and title index: Volumes 231 to 240 (2002). Marine Ecology Progress Series, 240: 297–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m240p297.pdf.

Anonymous:2002:EBA

[Ano02c] Anonymous. Erratum: Besiktepe et al., 224:151–164. (2001). Marine Ecology Progress Series, 234:310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p310.pdf. See [BD02].

Anonymous:2002:EEA

[Ano02d] Anonymous. Erratum: Eastwood et al., 224:251–266 (2001). Marine Ecology Progress Series, 230:301, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p301.pdf. See [EMG01].

Anonymous:2002:EKK

[Ano02e] Anonymous. Erratum: Khalaf & Kochzius, 239:287–299 (2002). Marine Ecology Progress Series, 241:306–307, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/241/m241p306.pdf. See [KK02].

Anonymous:2002:EMA

[Ano02f] Anonymous. Erratum: Moodley et al., 236:23–29. Marine Ecology Progress Series, 239:230, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/

m239p230.pdf. See [MMB^{+}02].

Anonymous:2002:ESG

[Ano02g] Anonymous. Erratum: Starmans & Gutt, 225:45–52, 2002. Marine Ecology Progress Series, 229?:313, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p313.pdf. See [SG02a].

Anonymous:2002:ETA

[Ano02h] Anonymous. Erratum: Thrush et al., 221:255–264, 2001 and 223:277–286, 2001. Marine Ecology Progress Series, 240:296, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/240/m240p296.pdf. See [THF^{+}01a, THF^{+}01b].

Anonymous:2002:PGA

[Ano02i] Anonymous. Paffenhöfer GA: An assessment of the effects of diatoms on planktonic copepods. Marine Ecology Progress Series, 227:305–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/227/m227p305.pdf.

Anonymous:2003:CAT

[Ano03a] Anonymous. Combined author and title index: Volumes 251 to 260 (2003). Marine Ecology Progress Series, 260: 297–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m260p297.pdf.

Anonymous:2003:EGA

[Ano03b] Anonymous. Errata: Gowing et al., 241:1–11, (2002). Marine Ecology Progress Series, 254:313, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/

m254p313.pdf. See [GRG^{+}02].

Anonymous:2003:ECA

[Ano03c] Anonymous. Erratum: Childs et al., 240:285–290. Marine Ecology Progress Series, 247:310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/

m247p310.pdf. See [CRTP02].

Anonymous:2003:EKA

[Ano03d] Anonymous. Erratum: Kvamme et al., 247:197–210 (2003). Marine Ecology Progress Series, 263:309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/

m263p309.pdf. See [KNF^{+}03].

Anonymous:2003:ELA

[Ano03e] Anonymous. Erratum: Lea et al., 245:281–297 (2002). Marine Ecology Progress Series, 253:310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/

m253p310.pdf. See [LCGN02].

Anonymous:2003:ETA

[Ano03f] Anonymous. Erratum: Tilstone et al., 252:89–104 (2003). Marine Ecology Progress Series, 254:313, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/

m254p313.pdf. See [TFLA03].

Anonymous:2003:N

[Ano03g] Anonymous. NOTES. Marine Ecology Progress Series, ??:??, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2003:TS

[Ano03h] Anonymous. Theme section. Marine Ecology Progress Series, ??:??, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2004:CATa

[Ano04a] Anonymous. Combined author and title index: Volumes 261 to 270 (2003/2004). Marine Ecology Progress Series, 270:289–301, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/authorindex/m270p289.pdf.

Anonymous:2004:CATb

[Ano04b] Anonymous. Combined author and title index; volumes 271 to 280 (2004). Marine Ecology Progress Series, 280:295– 306, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m280p295.pdf.

Anonymous:2004:EBA

[Ano04c] Anonymous. Erratum: Behrens et al., 279:129–139 (2004). Marine Ecology Progress Series, 281:307, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/

m281p307.pdf. See erratum [BL04].

Anonymous:2004:ECD

[Ano04d] Anonymous. Erratum: Campbell & Dower, 263:93–99, 2003. Marine Ecology Progress Series, 274:304, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p304.pdf. See [CD03a].

Anonymous:2004:EPA

[Ano04e] Anonymous. Erratum: Porter et al., 271:61–75 (2004). Marine Ecology Progress Series, 278:309, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/278/

m278p309.pdf. See [PCS04b].

Anonymous:2005:CATa

[Ano05a] Anonymous. Combined author and title index: Volumes 281 to 290 (2005). Marine Ecology Progress Series, 290: 297–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m290p297.pdf.

Anonymous:2005:CATb

[Ano05b] Anonymous. Combined author and title index Volumes 291 to 300 (2005). Marine Ecology Progress Series, 300:297– 310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/authorindex/m300p297.pdf.

Anonymous:2005:EKA

[Ano05c] Anonymous. Erratum: Kraus et al., 291:301–305 (2005). Marine Ecology Progress Series, 294:310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p310.pdf. See [KS05].

Anonymous:2005:EWA

[Ano05d] Anonymous. Erratum: Waite et al., 297:157–167 (2005). Marine Ecology Progress Series, 305:305, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/305/
m305p305.pdf. See [WGD05].

Anonymous:2006:EGA

[Ano06] Anonymous. Erratum: Gallager et al., 267:27–43 (2005). Marine Ecology Progress Series, 318:309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p309.pdf. See [GYD04].

Anonymous:2007:CATa

[Ano07a] Anonymous. Combined author and title index: Volumes 321 to 330 (2006–2007). Marine Ecology Progress Series, 330:297–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p297.pdf.

Anonymous:2007:CATb

[Ano07b] Anonymous. Combined author and title index: Volumes 341 to 350 (2007). Marine Ecology Progress Series, 350: 299–312, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/350/m350p299.pdf.

Anonymous:2007:EBA

[Ano07c] Anonymous. Erratum: Blicher et al., 341:89–102 (2007). Marine Ecology Progress Series, 346:314, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p314.pdf. See [BRS07].

Anonymous:2007:EK

[Ano07d] Anonymous. Erratum: Kimmerer, 324:207–218 (2006). Marine Ecology Progress Series, 343:308, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/343/
m343p308.pdf.

Anonymous:2007:ENA

[Ano07e] Anonymous. Erratum: Neer et al., 329:211–223 (2007). Marine Ecology Progress Series, 335:311, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p311.pdf. See [NRC07].

Anonymous:2007:ERC

[Ano07f] Anonymous. Erratum: Ropert-Coudert et al. 328:275–284, 2006. Marine Ecology Progress Series, 331:309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/

m331p309.pdf. See [RCWG^{+}06].

Anonymous:2007:PMM

[Ano07g] Anonymous. Possible mechanisms for a marine benthic regime shift in the North Sea. Marine Ecology Progress Series, 330: 39–47, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/330/m330p039.pdf.

Anonymous:2007:RSA

[Ano07h] Anonymous. Retraction: Sonnenholzner et al., 343:77– 85. Marine Ecology Progress Series, ??:??, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See [SLL07].

Anonymous:2008:CATa

[Ano08a] Anonymous. Combined author and title index: Volumes 351 to 360 (2007–2008). Marine Ecology Progress Series, 360:297–310, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p297.pdf.

Anonymous:2008:CATb

[Ano08b] Anonymous. Combined author and title index: Volumes 361 to 370 (2008). Marine Ecology Progress Series, 370: 295–307, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p295.pdf.

Anonymous:2009:CATa

[Ano09a] Anonymous. Combined author and title index: Volumes 371 to 380 (2009). Marine Ecology Progress Series, 370: 295–308, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/380/m380p295.pdf.

Anonymous:2009:CATb

[Ano09b] Anonymous. Combined author and title index: Volumes 381 to 390 (2009). Marine Ecology Progress Series, 390: 291–304, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p291.pdf.

Anonymous:2010:EHM

[Ano10a] Anonymous. Erratum: Holt & Mantua, 393:285–299 (2009). Marine Ecology Progress Series, 398:305–306, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/398/
m398p305.pdf. See [HM09].

Anonymous:2010:EKA

[Ano10b] Anonymous. Erratum: Kerr et al., 349:245–253, (2007). Marine Ecology Progress Series, 417:301, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2010/417/
m417p301.pdf. See [KSK07].

Anthony:2006:EES

[Ant06] K. R. N. Anthony. Enhanced energy status of corals on coastal, high-turbidity reefs. Marine Ecology Progress Series, 319:111–116, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p111.pdf.

Angel:2001:STO

[AO01] A. Angel and F. P. Ojeda. Structure and trophic organization of subtidal fish assemblages on the northern Chilean coast: the effect of habitat complexity. Marine Ecology Progress Series, 217:81–91, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/217/m217p081.pdf.

Arcos:2002:SFD

[AO02] J. M. Arcos and D. Oro. Significance of fisheries discards for a threatened Mediterranean seabird, the Balearic shearwater Puffinus mauretanicus. Marine Ecology Progress Series, 239:209–220, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p209.pdf.

Arnott:2000:SPS

[AP00a] S. A. Arnott and L. Pihl. Selection of prey size and prey species by 1-group cod Gadus morhua: effects of satiation level and prey handling times. Marine Ecology Progress Series, 198:225–238, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p225.pdf.

Attrill:2000:EIP

[AP00b] M. J. Attrill and M. Power. Effects on invertebrate populations of drought-induced changes in estuarine water quality. Marine Ecology Progress Series, 203:133–143, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p133.pdf.

Alofs:2004:MSIa

[AP04a] K. M. Alofs and K. M. Polivka. Microhabitat-scale influences of resources and refuge on habitat selection by an estuarine opportunist fish. Marine Ecology Progress Series, 271:297–306, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/271/m271p297.pdf.

Alofs:2004:MSIb

[AP04b] K. M. Alofs and K. M. Polivka. Microhabitat-scale influences of resources and refuge on habitat selection by an estuarine opportunist fish. Marine Ecology Progress Series, 271:297– 306, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Abookire:2005:OCS

[AP05a] A. A. Abookire and J. F. Piatt. Oceanographic conditions structure forage fishes into lipid-rich and lipid-poor communities in lower Cook Inlet, Alaska, USA. Marine Ecology Progress Series, 287:229–240, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/287/m287p229.pdf.

Arndt:2005:OFI

[AP05b] C. E. Arndt and O. Pavlova. Origin and fate of ice fauna in the Fram Strait and Svalbard area. Marine Ecology Progress Series, 301:55–66, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p055.pdf.

Aguilar-Perera:2006:DNG

[AP06] A. Aguilar-Perera. Disappearance of a Nassau grouper spawning aggregation off the southern Mexican Caribbean coast. Marine Ecology Progress Series, 327:289–296, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p289.pdf.

Aguilar-Perera:2007:VJF

[APA07] A. Aguilar-Perera and R. S. Appeldoorn. Variation in juvenile fish density along the mangrove-seagrass-coral reef continuum in SW Puerto Rico. Marine Ecology Progress Series, 348:139–148, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p139.pdf.

Andre:2009:MSA

[APG^{+}09] J. André, G. T. Pecl, E. P. M. Grist, J. M. Semmens, M. Haddon, and S. C. Leporati. Modelling size-at-age in wild immature female octopus: a bioenergetics approach. Marine Ecology Progress Series, 384:159–174, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/384/
m384p159.pdf.

Andrefouet:2004:ULS

[APH^{+}04] S. Andréfouët, C. Payri, E. J. Hochberg, C. Hu, M. J. Atkinson, and F. E. Muller-Karger. Use of in situ and airborne reflectance for scaling-up spectral discrimination of coral reef macroalgae from species to communities. Marine Ecology Progress Series, 283:161–177, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p161.pdf.

Alvarado:2001:PRT

[APM^{+}01] J. L. Alvarado, R. Pinto, P. Marquet, C. Pacheco, R. Guiñez, and J. C. Castilla. Patch recolonization by the tunicate Pyura praeputialis in the rocky intertidal of the Bay of Antofagasta, Chile: evidence for self-facilitation mechanisms. Marine Ecology Progress Series, 224:93–101, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/224/m224p093.pdf.

Anderson:2009:ITP

[APM^{+}09] O. R. J. Anderson, R. A. Phillips, R. A. McDonald, R. F. Shore, R. A. R. McGill, and S. Bearhop. Influence of trophic position and foraging range on mercury levels within a seabird community. Marine Ecology Progress Series, 375: 277–288, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/375/m375p277.pdf.

Akoumianaki:2006:DMB

[APN06] I. Akoumianaki, S. Papaspyrou, and A. Nicolaidou. Dynamics of macrofaunal body size in a deltaic environment. Marine Ecology Progress Series, 321:55–66, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p055.pdf.

Arthur:2009:ENE

[APP^{+}09] K. E. Arthur, V. J. Paul, H. W. Paerl, J. M. O’Neil, J. Joyner, and T. Meickle. Effects of nutrient enrichment of the cyanobacterium Lyngbya sp. on growth, secondary metabolite concentration and feeding by the specialist grazer Stylocheilus striatus. Marine Ecology Progress Series, 394: 101–110, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/394/m394p101.pdf.

Au:2004:CES

[APW^{+}04] D. W. T. Au, C. A. Pollino, R. S. S. Wu, P. K. S. Shin, S. T. F. Lau, and J. Y. M. Tang. Chronic effects of suspended solids on gill structure, osmoregulation, growth, and triiodothyronine in juvenile green grouper Epinephelus coioides. Marine Ecology Progress Series, 266:255–264, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/266/m266p255.pdf.

Al-Qutob:2002:PDN

[AQHTL02] M. Al-Qutob, C. Häse, M. M. Tilzer, and B. Lazar. Phytoplankton drives nitrite dynamics in the Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 239:233–239, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/239/m239p233.pdf.

Amaral:2007:PAS

[AQSP07] V. Amaral, H. Queiroga, M. Skov, and J. Paula. Planktonic availability and settlement of Carcinus maenas megalopae at high temporal resolution in the lower Mira Estuary (SW Portugal). Marine Ecology Progress Series, 348:

239–248, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/348/m348p239.pdf.

Armonies:2000:FDA

[AR00] W. Armonies and K. Reise. Faunal diversity across a sandy shore. Marine Ecology Progress Series, 196:49–57, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p049.pdf.

Acosta:2002:DCR

[AR02a] C. A. Acosta and D. N. Robertson. Diversity in coral reef fish communities: the effects of habitat patchiness revisited. Marine Ecology Progress Series, 227:87–96, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p087.pdf.

Arnott:2002:SRN

[AR02b] S. A. Arnott and G. D. Ruxton. Sandeel recruitment in the North Sea: demographic, climatic and trophic effects. Marine Ecology Progress Series, 238:199–210, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p199.pdf.

Al-Rousan:2002:EBE

[ARAMPW02] S. Al-Rousan, S. Al-Moghrabi, J. Pätzold, and G. Wefer. Environmental and biological effects on the stable oxygen isotope records of corals in the northern Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 239:301–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p301.pdf.

Auffrey:2004:EGM

[ARB04] L. M. Auffrey, S. M. C. Robinson, and M. A. Barbeau. Effect of green macroalgal mats on burial depth of soft-shelled clams Mya arenaria. Marine Ecology Progress Series, 278: 193–203, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/278/m278p193.pdf.

Arcos:2002:MLS

[ARBF02] J. M. Arcos, X. Ruiz, S. Bearhop, and R. W. Furness. Mercury levels in seabirds and their fish prey at the Ebro Delta (NW Mediterranean): the role of trawler discards as a source of contamination. Marine Ecology Progress Series, 232:281– 290, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/232/m232p281.pdf.

Attrill:2009:OPE

[ARFP09] M. J. Attrill, S. D. Rundle, A. Fraser, and M. Power. Oligochaetes as a possible entry route for terrigenous organic carbon into estuarine benthic food webs. Marine Ecology Progress Series, 384:147–157, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/384/m384p147.pdf.

Allen:2001:FSH

[ARGS01] M. C. Allen, A. J. Read, J. Gaudet, and L. S. Sayigh. Finescale habitat selection of foraging bottlenose dolphins Tursiops truncatus near Clearwater, Florida. Marine Ecology Progress Series, 222:253–264, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/222/m222p253.pdf.

Arkema:2009:FMF

[Ark09] K. K. Arkema. Flow-mediated feeding in the field: consequences for the performance and abundance of a sessile marine invertebrate. Marine Ecology Progress Series, 388:207– 220, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/388/m388p207.pdf.

Armsworth:2000:MSR

[Arm00] P. R. Armsworth. Modelling the swimming response of late stage larval reef fish to different stimuli. Marine Ecology Progress Series, 195:231–247, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p231.pdf.

Armonies:2001:WIS

[Arm01] W. Armonies. What an introduced species can tell us about the spatial extension of benthic populations. Marine Ecology Progress Series, 209:289–294, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p289.pdf.

Arnott:2002:SDP

[ARP02] S. A. Arnott, G. D. Ruxton, and E. S. Poloczanska. Stochastic dynamic population model of North Sea sandeels, and its application to precautionary management procedures. Marine Ecology Progress Series, 235:223–234, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p223.pdf.

Arenas:2009:DEB

[ARP09] F. Arenas, F. Rey, and I. Sousa Pinto. Diversity effects beyond species richness: evidence from intertidal macroalgal assemblages. Marine Ecology Progress Series, 381:99– 108, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/381/m381p099.pdf.

Arrontes:2005:MRE

[Arr05] J. Arrontes. A model for range expansion of coastal algal species with different dispersal strategies: the case of Fucus serratus in northern Spain. Marine Ecology Progress Series, 295:57–68, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p057.pdf.

Abraham:2004:OCE

[AS04] C. L. Abraham and W. J. Sydeman. Ocean climate, euphausiids and auklet nesting: inter-annual trends and variation in phenology, diet and growth of a planktivorous seabird, Ptychoramphus aleuticus. Marine Ecology Progress Series, 274:235–250, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p235.pdf.

Abraham:2006:PSC

[AS06a] C. L. Abraham and W. J. Sydeman. Prey-switching by Cassin’s auklet Ptychoramphus aleuticus reveals seasonal climate-related cycles of Euphausia pacifica and Thysanoessa spinifera. Marine Ecology Progress Series, 313:271–283, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/313/m313p271.pdf.

Aeby:2006:FAS

[AS06b] G. S. Aeby and D. L. Santavy. Factors affecting susceptibility of the coral Montastraea faveolata to black-band disease. Marine Ecology Progress Series, 318:103–110, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p103.pdf.

Aguzzi:2003:DSP

[ASA^{+}03] J. Aguzzi, F. Sardà, P. Abelló, J. B. Company, and G. Rotllant. Diel and seasonal patterns of Nephrops norvegicus (Decapoda: Nephropidae) catchability in the western Mediterranean. Marine Ecology Progress Series, 258:201–211, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/258/m258p201.pdf.

Adey:2008:GFT

[ASA^{+}08] J. M. Adey, I. P. Smith, R. J. A. Atkinson, I. D. Tuck, and A. C. Taylor. ‘Ghost fishing’ of target and non-target species by Norway lobster Nephrops norvegicus creels. Marine Ecology Progress Series, 366:119–127, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/366/m366p119.pdf.

Atienza:2006:FEM

[ASC06] D. Atienza, E. Saiz, and A. Calbet. Feeding ecology of the marine cladoceran Penilia avirostris: natural diet, prey selectivity and daily ration. Marine Ecology Progress Series, 315:211–220, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p211.pdf.

Aure:2007:PPE

[ASES07] J. Aure, Ó. Strand, S. R. Erga, and T. Strohmeier. Primary production enhancement by artificial upwelling in a western Norwegian fjord. Marine Ecology Progress Series, 352: 39–52, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/352/m352p039.pdf.

Alvarez-Salgado:2004:INR

[ÁSG04] X. A. Álvarez-Salgado and M. Gilcoto. Inferring nitrification rates with an inverse method in a coastal upwelling system, Ría de Vigo (NW Spain). Marine Ecology Progress Series, 276:3–17, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p003.pdf.

Anderson:2003:FTT

[ASH03] J. T. Anderson, D. K. Stoecker, and R. R. Hood. Formation of two types of cysts by a mixotrophic dinoflagellate, Pfiesteria piscicida. Marine Ecology Progress Series, 246: 95–104, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p095.pdf.

Andrews:2009:GRA

[ASLD09] A. H. Andrews, R. P. Stone, C. C. Lundstrom, and A. P. DeVogelaere. Growth rate and age determination of bamboo corals from the northeastern Pacific Ocean using refined^{210}Pb dating. Marine Ecology Progress Series, 397:173–185, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p173.pdf.

Ahn:2003:GSE

[ASP^{+}03] I. Y. Ahn, J. Surh, Y. G. Park, H. Kwon, K. S. Choi, S. H. Kang, H. J. Choi, K. W. Kim, and H. Chung. Growth and seasonal energetics of the Antarctic bivalve Laternula elliptica from King George Island, Antarctica. Marine Ecology Progress Series, 257:99–110, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p099.pdf.

Atkinson:2008:OCH

[ASP^{+}08] A. Atkinson, V. Siegel, E. A. Pakhomov, P. Rothery, V. Loeb, R. M. Ross, L. B. Quetin, K. Schmidt, P. Fretwell, E. J. Murphy, G. A. Tarling, and A. H. Fleming. Oceanic circumpolar habitats of Antarctic krill. Marine Ecology Progress Series, 362:1–23, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Arvanitidis:2009:BGE

[ASR^{+}09] C. Arvanitidis, P. J. Somerfield, H. Rumohr, S. Faulwetter, V. Valavanis, A. Vasileiadou, G. Chatzigeorgiou, E. Vanden Berghe, J. Vanaverbeke, C. Labrune, A. Grémare, M. L. Zettler, M. Kedra, M. Włodarska-Kowalczuk, I. F. Aleffi, J. M. Amouroux, N. Anisimova, G. Bachelet, M. Büntzow, S. J. Cochrane, M. J. Costello, J. Craeymeersch, S. Dahle, S. Degraer, S. Denisenko, C. Dounas, G. Duineveld, C. Emblow, V. Escavarage, M. C. Fabri, D. Fleischer, J. S. Gray, C. H. R. Heip, M. Herrmann, H. Hummel, U. Janas, I. Karakassis, M. A. Kendall, P. Kingston, L. Kotwicki, J. Laudien, A. S. Y. Mackie, E. L. Nevrova, A. Occhipinti-Ambrogi, P. G. Oliver, F. Olsgard, R. Palerud, A. Petrov, E. Rachor, N. K. Revkov, A. Rose, R. Sardá, W. C. H. Sistermans, J. Speybroeck, G. Van Hoey, M. Vincx, P. Whomersley, W. Willems, and A. Zenetos. Biological geography of the European seas: results from the MacroBen database. Marine Ecology Progress Series, 382:265–278, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Alexandre:2005:ECH

[ASS05] A. Alexandre, R. Santos, and E. Serrão. Effects of clam harvesting on sexual reproduction of the seagrass Zostera noltii. Marine Ecology Progress Series, 298:115–122, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p115.pdf.

Anger:2003:LEJ

[ATLC03] K. Anger, S. Thatje, G. Lovrich, and J. Calcagno. Larval and early juvenile development of Paralomis granulosa reared at different temperatures: tolerance of cold and food limitation in a lithodid crab from high latitudes. Marine Ecology Progress Series, 253:243–251, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/253/m253p243.pdf.

Alongi:2008:GDM

[ATR^{+}08] D. M. Alongi, L. A. Trott, Rachmansyah, F. Tirendi, A. D. McKinnon, and M. C. Undu. Growth and development of mangrove forests overlying smothered coral reefs, Sulawesi and Sumatra, Indonesia. Marine Ecology Progress Series, 370: 97–109, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p097.pdf.

Aksnes:2006:IDC

[ATT06a] D. L. Aksnes, C. Troedsson, and E. M. Thompson. Integrating developmental clocking and growth in a life-history model for the planktonic chordate Oikopleura dioica. Marine Ecology Progress Series, 318:81–88, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m318p081.pdf.

Aksnes:2006:MDT

[ATT06b] D. L. Aksnes, C. Troedsson, and E. M. Thompson. A model of developmental time applied planktonic embryos. Marine Ecology Progress Series, 318:75–80, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m318p075.pdf.

Alongi:2000:BDR

[ATTX00] D. M. Alongi, F. Tirendi, L. A. Trott, and T. T. Xuan. Benthic decomposition rates and pathways in plantations of the mangrove Rhizophora apiculata in the Mekong delta, Vietnam. Marine Ecology Progress Series, 194:87–101, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p087.pdf.

Araujo:2009:EDM

[AVA^{+}09] R. Araújo, S. Vaselli, M. Almeida, E. Serrão, and I. Sousa-Pinto. Effects of disturbance on marginal populations: human trampling on Ascophyllum nodosum assemblages at its southern distribution limit. Marine Ecology Progress Series, 378:81–92, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p081.pdf.

Arvanitidis:2006:MBI

[AVE^{+}06] C. Arvanitidis, V. D. Valavanis, A. Eleftheriou, M. J. Costello, S. Faulwetter, P. Gotsis, M. S. Kitsos, I. Kirmtzoglou, A. Zenetos, A. Petrov, B. Galil, and N. Papageorgiou. MedOBIS: biogeographic information system for the eastern Mediterranean and Black Sea. Marine Ecology Progress Series, 316: 225–230, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m316p225.pdf.

Alonzo:2005:LTMa

[AVNN05a] F. Alonzo, P. Virtue, S. Nicol, and P. D. Nichols. Lipids as trophic markers in Antarctic krill. II. Lipid composition of the body and digestive gland of Euphausia superba in controlled conditions. Marine Ecology Progress Series, 296: 65–79, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/296/m296p065.pdf.

Alonzo:2005:LTMb

[AVNN05b] F. Alonzo, P. Virtue, S. Nicol, and P. D. Nichols. Lipids as trophic markers in Antarctic krill. III. Temporal changes in digestive gland lipid composition of Euphausia superba in controlled conditions. Marine Ecology Progress Series, 296: 81–91, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/296/m296p081.pdf.

Andersson:2003:PTP

[AVS03] M. Andersson, L. Van Nieuwerburgh, and P. Snoeijs. Pigment transfer from phytoplankton to zooplankton with emphasis on astaxanthin production in the Baltic Sea food web. Marine Ecology Progress Series, 254:213–224, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p213.pdf.

Aberle:2003:DSM

[AW03a] N. Aberle and U. Witte. Deep-sea macrofauna exposed to a simulated sedimentation event in the abyssal NE Atlantic: in situ pulse-chase experiments using^{13}C-labelled phytodetritus. Marine Ecology Progress Series, 251:37–47, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/251/
m251p037.pdf.

Allen:2003:NEZ

[AW03b] B. J. Allen and S. L. Williams. Native eelgrass Zostera marina controls growth and reproduction of an invasive mussel through food limitation. Marine Ecology Progress Series, 254:57–67, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p057.pdf.

Arana:2005:AIO

[AWA^{+}05] H. A. Hernández Arana, R. M. Warwick, M. J. Attrill, A. A. Rowden, and G. Gold-Bouchot. Assessing the impact of oilrelated activities on benthic macroinfauna assemblages of the Campeche shelf, southern Gulf of Mexico. Marine Ecology Progress Series, 289:89–107, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/289/m289p089.pdf.

Adams:2009:FSG

[AWBO09] A. Adams, R. K. Wolfe, N. Barkowski, and D. Overcash. Fidelity to spawning grounds by a catadromous fish, Centropomus undecimalis. Marine Ecology Progress Series, 389: 213–222, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/389/m389p213.pdf.

Ahner:2002:GOL

[AWOO02] B. A. Ahner, L. Wei, J. R. Oleson, and N. Ogura. Glutathione and other low molecular weight thiols in marine phytoplankton under metal stress. Marine Ecology Progress Series, 232:93–103, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/232/m232p093.pdf.

Atkinson:2001:SGA

[AWP^{+}01] A. Atkinson, M. J. Whitehouse, J. Priddle, G. C. Cripps, P. Ward, and M. A. Brandon. South Georgia, Antarctica: a productive, cold water, pelagic ecosystem. Marine Ecology Progress Series, 216:279–308, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/216/m216p279.pdf.

Althaus:2009:IBT

[AWS^{+}09] F. Althaus, A. Williams, T. A. Schlacher, R. J. Kloser, M. A. Green, B. A. Barker, N. J. Bax, P. Brodie, and M. A. Schlacher-Hoenlinger. Impacts of bottom trawling on deepcoral ecosystems of seamounts are long-lasting. Marine Ecology Progress Series, 397:279–294, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m397p279.pdf.

Amano:2003:SWD

[AY03] M. Amano and M. Yoshioka. Sperm whale diving behavior monitored using a suction-cup-attached TDR tag. Marine Ecology Progress Series, 258:291–295, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p291.pdf.

Akiyama:2004:LHNa

[AY04a] T. Akiyama and M. Yamamoto. Life history of Nippoleucon hinumensis (Crustacea: Cumacea: Leuconidae) in Seto Inland Sea of Japan. I. Summer diapause and molt cycle. Marine Ecology Progress Series, 284:211–225, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p211.pdf.

Akiyama:2004:LHNb

[AY04b] T. Akiyama and M. Yamamoto. Life history of Nippoleucon hinumensis (Crustacea: Cumacea: Leuconidae) in Seto Inland Sea of Japan. II. Non-diapausing subpopulation. Marine Ecology Progress Series, 284:227–235, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p227.pdf.

Alamaru:2009:TBS

[AYSL09] A. Alamaru, R. Yam, A. Shemesh, and Y. Loya. Trophic biology of Stylophora pistillata larvae: evidence from stable isotope analysis. Marine Ecology Progress Series, 383: 85–94, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/383/m383p085.pdf.

Al-Zaidan:2006:RMM

[AZKJAM06] A. S. Y. Al-Zaidan, H. Kennedy, D. A. Jones, and S. Y. Al-Mohanna. Role of microbial mats in Sulaibikhat Bay (Kuwait) mudflat food webs: evidence from? ^{13}C analysis. Marine Ecology Progress Series, 308:27–36, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p027.pdf.

Basterretxea:2000:MVP

[BA00] G. Basterretxea and J. Arístegui. Mesoscale variability in phytoplankton biomass distribution and photosynthetic parameters in the canary-NW African coastal transition zone. Marine Ecology Progress Series, 197:27–40, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/197/m197p027.pdf.

Barnes:2001:GCE

[BA01] D. K. A. Barnes and R. Arnold. A growth cline in encrusting benthos along a latitudinal gradient within Antarctic waters. Marine Ecology Progress Series, 210:85–91, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
210/m210p085.pdf.

Bolduc:2004:HAM

[BA04] F. Bolduc and A. D. Afton. Hydrologic aspects of marsh ponds during winter on the Gulf Coast Chenier Plain, USA: effects of structural marsh management. Marine Ecology Progress Series, 266:35–42, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p035.pdf.

Baird:2004:EFB

[BAA04] D. Baird, H. Asmus, and R. Asmus. Energy flow of a boreal intertidal ecosystem, the Sylt–Rømø Bight. Marine Ecology Progress Series, 279:45–61, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/279/m279p045.pdf.

Baird:2007:TDE

[BAA07] D. Baird, H. Asmus, and R. Asmus. Trophic dynamics of eight intertidal communities of the Sylt–Rømø Bight ecosystem, northern Wadden Sea. Marine Ecology Progress Series, 351:25–41, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p025.pdf.

Brierley:2001:AOJ

[BAB^{+}01] A. S. Brierley, B. E. Axelsen, E. Buecher, C. A. J. Sparks, H. Boyer, and M. J. Gibbons. Acoustic observations of jellyfish in the Namibian Benguela. Marine Ecology Progress Series, 210:55–66, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p055.pdf.

Bjork:2004:WUI

[BAB04] M. Björk, L. Axelsson, and S. Beer. Why is Ulva intestinalis the only macroalga inhabiting isolated rockpools along the Swedish Atlantic coast? Marine Ecology Progress Series, 284:109–116, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p109.pdf.

Bhadury:2006:DED

[BAB^{+}06] P. Bhadury, M. C. Austen, D. T. Bilton, P. J. D. Lambshead, A. D. Rogers, and G. R. Smerdon. Development and evaluation of a DNA-barcoding approach for the rapid identification of nematodes. Marine Ecology Progress Series, 320:1–9, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brion:2008:NCR

[BAE^{+}08] N. Brion, M. G. I. Andersson, M. Elskens, C. Diaconu, W. Baeyens, F. Dehairs, and J. J. Middelburg. Nitrogen cycling, retention and export in a eutrophic temperate macrotidal estuary. Marine Ecology Progress Series, 357: 87–99, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/357/m357p087.pdf.

Bailey:2000:SCR

[Bai00] K. M. Bailey. Shifting control of recruitment of walleye pollock Theragra chalcogramma after a major climatic and ecosystem change. Marine Ecology Progress Series, 198:215–224, 2000.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p215.pdf.

Barlow:2004:PAS

[BAM^{+}04] R. G. Barlow, J. Aiken, G. F. Moore, P. M. Holligan, and S. Lavender. Pigment adaptations in surface phytoplankton along the eastern boundary of the Atlantic Ocean. Marine Ecology Progress Series, 281:13–26, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p013.pdf.

Blight:2009:IMA

[BAMJ09] A. J. Blight, A. L. Allcock, C. A. Maggs, and M. P. Johnson. Intertidal molluscan and algal species richness around the UK coast. Marine Ecology Progress Series, 396:235–243, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p235.pdf.

Bode:2006:PMC

[BAOV06] A. Bode, M. T. Alvarez-Ossorio, and M. Varela. Phytoplankton and macrophyte contributions to littoral food webs in the Galician upwelling estimated from stable isotopes. Marine Ecology Progress Series, 318:89–102, 2006. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m318p089.pdf.

Barnes:2002:ESH

[Bar02a] D. K. A. Barnes. Ecology of subtropical hermit crabs in SW Madagascar: refuge-use and dynamic niche overlap. Marine Ecology Progress Series, 238:163–172, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p163.pdf.

Barrett:2002:APF

[Bar02b] R. T. Barrett. Atlantic puffin Fratercula arctica and common guillemot Uria aalge chick diet and growth as indicators of fish stocks in the Barents Sea. Marine Ecology Progress Series, 230:275–287, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p275.pdf.

Bartholomew:2002:TCC

[Bar02c] A. Bartholomew. Total cover and cover quality: predicted and actual effects on a predator’s foraging success. Marine Ecology Progress Series, 227:1–9, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p001.pdf.

Barnes:2003:LRG

[Bar03] D. K. A. Barnes. Local, regional and global patterns of resource use in ecology: hermit crabs and gastropod shells as an example. Marine Ecology Progress Series, 246:211– 223, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p211.pdf.

Barnes:2006:TSS

[Bar06] D. K. A. Barnes. Temporal–spatial stability of competition in marine boulder fields. Marine Ecology Progress Series, 314:15–23, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p015.pdf.

Barrett:2007:FWI

[Bar07] R. T. Barrett. Food web interactions in the southwestern Barents Sea: black-legged kittiwakes Rissa tridactyla respond negatively to an increase in herring Clupea harengus. Marine Ecology Progress Series, 349:269–276, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p269.pdf.

Baskett:2006:PSR

[Bas06] M. L. Baskett. Prey size refugia and trophic cascades in marine reserves. Marine Ecology Progress Series, 328:285– 293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p285.pdf.

Bates:2007:FSM

[Bat07] A. E. Bates. Feeding strategy, morphological specialisation and presence of bacterial episymbionts in lepetodrilid gastropods from hydrothermal vents. Marine Ecology Progress Series, 347:87–99, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/347/m347p087.pdf.

Bates:2009:HTR

[Bat09] C. R. Bates. Host taxonomic relatedness and functionalgroup affiliation as predictors of seaweed-invertebrate epifaunal associations. Marine Ecology Progress Series, 387:125– 136, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/387/m387p125.pdf.

Bayne:2002:PCB

[Bay02] B. L. Bayne. A physiological comparison between Pacific oysters Crassostrea gigas and Sydney rock oysters Saccostrea glomerata: food, feeding and growth in a shared estuarine habitat. Marine Ecology Progress Series, 232:163–178, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/232/m232p163.pdf.

Bayne:2004:CMC

[Bay04] B. L. Bayne. Comparisons of measurements of clearance rates in bivalve molluscs. Marine Ecology Progress Series, 276:305–306, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p305.pdf.

Bostrom:2000:ZCE

[BB00] C. Boström and E. Bonsdorff. Zoobenthic community establishment and habitat complexity-the importance of seagrass shoot-density, morphology and physical disturbance for faunal recruitment. Marine Ecology Progress Series, 205:123– 138, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p123.pdf.

Booth:2002:CFA

[BB02a] D. J. Booth and G. A. Beretta. Changes in a fish assemblage after a coral bleaching event. Marine Ecology Progress Series, 245:205–212, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p205.pdf.

Brooks:2002:MRA

[BB02b] R. A. Brooks and S. S. Bell. Mangrove response to attack by a root boring isopod: root repair versus architectural modification. Marine Ecology Progress Series, 231: 85–90, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p085.pdf.

Barnes:2003:ZBB

[BB03a] D. K. A. Barnes and S. Brockington. Zoobenthic biodiversity, biomass and abundance at Adelaide Island, Antarctica. Marine Ecology Progress Series, 249:145–155, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p145.pdf.

Becker:2003:SDH

[BB03b] B. H. Becker and S. R. Beissinger. Scale-dependent habitat selection by a nearshore seabird, the marbled murrelet, in a highly dynamic upwelling system. Marine Ecology Progress Series, 256:243–255, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/256/m256p243.pdf.

Bowles:2004:SHD

[BB04] J. W. Bowles and S. S. Bell. Simulated herbivory and the dynamics of disease in Thalassia testudinum. Marine Ecology Progress Series, 283:127–132, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p127.pdf.

Brown:2005:PMS

[BB05] B. E. Brown and J. C. Bythell. Perspectives on mucus secretion in reef corals. Marine Ecology Progress Series, 296: 291–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/296/m296p291.pdf.

Bonaldo:2008:SDV

[BB08a] R. M. Bonaldo and D. R. Bellwood. Size-dependent variation in the functional role of the parrotfish Scarus rivulatus on the Great Barrier Reef, Australia. Marine Ecology Progress Series, 360:237–244, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/360/m360p237.pdf.

Browman:2008:ILS

[BB08b] Howard Browman and Philip W. Boyd. Implications of largescale iron fertilization of the oceans. Marine Ecology Progress Series, 364:213–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m364ThemeSection.pdf.

Bansemer:2009:RPL

[BB09] C. S. Bansemer and M. B. Bennett. Reproductive periodicity, localised movements and behavioural segregation of pregnant Carcharias taurus at Wolf Rock, southeast Queensland, Australia. Marine Ecology Progress Series, 374:215– 227, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/374/m374p215.pdf.

Brierley:2005:TAE

[BBA^{+}05] A. S. Brierley, D. C. Boyer, B. E. Axelsen, C. P. Lynam, C. A. J. Sparks, H. J. Boyer, and M. J. Gibbons. Towards the acoustic estimation of jellyfish abundance. Marine Ecology Progress Series, 295:105–111, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/295/m295p105.pdf.

Benoit-Bird:2006:EDH

[BBA06a] K. J. Benoit-Bird and W. W. L. Au. Extreme diel horizontal migrations by a tropical nearshore resident micronekton community. Marine Ecology Progress Series, 319:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bode:2006:LDR

[BBA06b] M. Bode, L. Bode, and P. R. Armsworth. Larval dispersal reveals regional sources and sinks in the Great Barrier Reef. Marine Ecology Progress Series, 308:17–25, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p017.pdf.

Benoit-Bird:2001:DHM

[BBABL01] K. J. Benoit-Bird, W. W. L. Au, R. E. Brainard, and M. O. Lammers. Diel horizontal migration of the Hawaiian mesopelagic boundary community observed acoustically. Marine Ecology Progress Series, 217:1–14, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p001.pdf.

Bell:2006:MMS

[BBB^{+}06] J. J. Bell, M. Burton, B. Bullimore, P. B. Newman, and K. Lock. Morphological monitoring of subtidal sponge assemblages. Marine Ecology Progress Series, 311:79–91, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/311/m311p079.pdf.

Barletta:2003:SCD

[BBBSPH03] M. Barletta, A. Barletta-Bergan, U. Saint-Paul, and G. Hubold. Seasonal changes in density, biomass, and diversity of estuarine fishes in tidal mangrove creeks of the lower Caeté Estuary (northern Brazilian coast, east Amazon). Marine Ecology Progress Series, 256:217–228, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p217.pdf.

Bavestrello:2000:BMS

[BBC^{+}00] G. Bavestrello, C. N. Bianchi, B. Calcinai, R. Cattaneo-Vietti, C. Cerrano, C. Morri, S. Puce, and M. Sarà. Bio-mineralogy as a structuring factor for marine epibenthic communities. Marine Ecology Progress Series, 193:241–249, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p241.pdf.

Bjorndal:2003:SPE

[BBC03] K. A. Bjorndal, A. B. Bolten, and M. Y. Chaloupka. Survival probability estimates for immature green turtles Chelonia mydas in the Bahamas. Marine Ecology Progress Series, 252:273–281, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p273.pdf.

Bethea:2004:FEE

[BBC04] D. M. Bethea, J. A. Buckel, and J. K. Carlson. Foraging ecology of the early life stages of four sympatric shark species. Marine Ecology Progress Series, 268:245–264, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p245.pdf.

Bulleri:2008:FIG

[BBC08] F. Bulleri and L. Benedetti-Cecchi. Facilitation of the introduced green alga Caulerpa racemosa by resident algal turfs: experimental evaluation of underlying mechanisms. Marine Ecology Progress Series, 364:77–86, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p077.pdf.

Bacheler:2009:DDL

[BBC^{+}09a] N. M. Bacheler, K. M. Bailey, L. Ciannelli, V. Bartolino, and K. S. Chan. Density-dependent, landscape, and climate effects on spawning distribution of walleye pollock Theragra chalcogramma. Marine Ecology Progress Series, 391:1–12, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bo:2009:CBC

[BBC^{+}09b] M. Bo, G. Bavestrello, S. Canese, M. Giusti, E. Salvati, M. Angiolillo, and S. Greco. Characteristics of a black coral meadow in the twilight zone of the central Mediterranean Sea. Marine Ecology Progress Series, 397:53–61, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m397p053.pdf.

Bodkin:2002:SOP

[BBD^{+}02] J. L. Bodkin, B. E. Ballachey, T. A. Dean, A. K. Fukuyama, S. C. Jewett, L. McDonald, D. H. Monson, C. E. O’Clair, and G. R. VanBlaricom. Sea otter population status and the process of recovery from the 1989 ‘Exxon Valdez’ oil spill. Marine Ecology Progress Series, 241:237–253, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p237.pdf.

Benoit-Bird:2009:UAA

[BBDW09] K. J. Benoit-Bird, A. D. Dahood, and B. Würsig. Using active acoustics to compare lunar effects on predator–prey behavior in two marine mammal species. Marine Ecology Progress Series, 395:119–135, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p119.pdf.

Bell:2003:SSD

[BBE03] S. S. Bell, R. A. Brooks, and W. E. Ellis. Structural spacing and the determination of habitat complexity: examining the Bartholomew et al. (2000) index. Marine Ecology Progress Series, 248:293–295, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/248/m248p293.pdf.

Beare:2004:LTI

[BBG^{+}04] D. J. Beare, F. Burns, A. Greig, E. G. Jones, K. Peach, M. Kienzle, E. McKenzie, and D. G. Reid. Long-term increases in prevalence of North Sea fishes having southern biogeographic affinities. Marine Ecology Progress Series, 284: 269–278, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p269.pdf.

Bertrand:2006:DPF

[BBG^{+}06] A. Bertrand, M. A. Barbieri, F. Gerlotto, F. Leiva, and J. Córdova. Determinism and plasticity of fish schooling behaviour as exemplified by the South Pacific jack mackerel Trachurus murphyi. Marine Ecology Progress Series, 311:145–156, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/311/m311p145.pdf.

Boero:2008:WSI

[BBG^{+}08] F. Boero, J. Bouillon, C. Gravili, M. P. Miglietta, T. Parsons, and S. Piraino. As we see it: Gelatinous plankton: irregularities rule the world (sometimes). Marine Ecology Progress Series, 356:299–310, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p299.pdf.

Barnett:2006:TEC

[BBH06] A. Barnett, D. R. Bellwood, and A. S. Hoey. Trophic ecomorphology of cardinalfish. Marine Ecology Progress Series, 322:249–257, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p249.pdf.

Baird:2002:DNT

[BBHT02] R. W. Baird, J. F. Borsani, M. B. Hanson, and P. L. Tyack. Diving and night-time behavior of long-finned pilot whales in the Ligurian Sea. Marine Ecology Progress Series, 237:301–305, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p301.pdf.

Bjorndal:2000:SGM

[BBM00] K. A. Bjorndal, A. B. Bolten, and H. R. Martins. Somatic growth model of juvenile loggerhead sea turtles Caretta caretta: duration of pelagic stage. Marine Ecology Progress Series, 202:265–272, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/202/m202p265.pdf.

Barata:2002:DEM

[BBM^{+}02a] C. Barata, D. J. Baird, M. Medina, A. Albalat, and A. M. V. M. Soares. Determining the ecotoxicological mode of action of toxic chemicals in meiobenthic marine organisms: stagespecific short tests with Tisbe battagliai. Marine Ecology Progress Series, 230:183–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/230/m230p183.pdf.

Bulleri:2002:IEC

[BBM02b] F. Bulleri, I. Bertocci, and F. Micheli. Interplay of encrusting coralline algae and sea urchins in maintaining alternative habitats. Marine Ecology Progress Series, 243:101– 109, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p101.pdf.

Baird:2003:HSL

[BBM03] A. H. Baird, R. C. Babcock, and C. P. Mundy. Habitat selection by larvae influences the depth distribution of six common coral species. Marine Ecology Progress Series, 252:289–293, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p289.pdf.

Brush:2002:MPP

[BBNK02] M. J. Brush, J. W. Brawley, S. W. Nixon, and J. N. Kremer. Modeling phytoplankton production: problems with the Eppley curve and an empirical alternative. Marine Ecology Progress Series, 238:31–45, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/238/m238p031.pdf.

Blank:2004:ESS

[BBRJ04] M. Blank, R. Bastrop, M. Röhner, and K. Jürss. Effect of salinity on spatial distribution and cell volume regulation in two sibling species of Marenzelleria (Polychaeta: Spionidae). Marine Ecology Progress Series, 271:193–205, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p193.pdf.

Behum:2005:DJU

[BBS05] M. E. Behum, R. J. Brodie, and J. L. Staton. Distribution of juvenile Uca pugnax and U. pugilator across habitats in a South Carolina estuary, assessed by molecular techniques. Marine Ecology Progress Series, 288:211–220, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p211.pdf.

Brischoux:2007:FES

[BBS07] F. Brischoux, X. Bonnet, and R. Shine. Foraging ecology of sea kraits Laticauda spp. in the Neo-Caledonian Lagoon. Marine Ecology Progress Series, 350:145–151, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/350/
m350p145.pdf.

Boone:2003:CSC

[BBSO03] E. J. Boone, A. A. Boettcher, T. D. Sherman, and J. J. O’Brien. Characterization of settlement cues used by the rhizocephalan barnacle Loxothylacus texanus. Marine Ecology Progress Series, 252:187–197, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p187.pdf.

Buschbaum:2007:SBP

[BBST07] C. Buschbaum, G. Buschbaum, I. Schrey, and D. W. Thieltges. Shell-boring polychaetes affect gastropod shell strength and crab predation. Marine Ecology Progress Series, 329: 123–130, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p123.pdf.

BrindAmour:2002:FGR

[BBT02] A. Brind’Amour, E. Bourget, and R. Tremblay. Fecundity, growth rate and survivorship at the interface between two contiguous genetically distinct groups of Semibalanus balanoides. Marine Ecology Progress Series, 229:173–184, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p173.pdf.

Bode:2001:PBP

[BBV^{+}01] A. Bode, S. Barquero, M. Varela, J. G. Braun, and D. de Armas. Pelagic bacteria and phytoplankton in oceanic waters near the Canary Islands in summer. Marine Ecology Progress Series, 209:1–17, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/209/m209p001.pdf.

Benoit-Bird:2008:DVZ

[BBZM08] K. J. Benoit-Bird, M. J. Zirbel, and M. A. McManus. Diel variation of zooplankton distributions in Hawaiian waters favors horizontal diel migration by midwater micronekton. Marine Ecology Progress Series, 367:109–123, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/367/
m367p109.pdf.

Benedetti-Cecchi:2001:VAA

[BC01a] L. Benedetti-Cecchi. Variability in abundance of algae and invertebrates at different spatial scales on rocky sea shores. Marine Ecology Progress Series, 215:79–92, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p079.pdf.

Bonnet:2001:DEP

[BC01b] D. Bonnet and F. Carlotti. Development and egg production in Centropages typicus (Copepoda: Calanoida) fed different food types: a laboratory study. Marine Ecology Progress Series, 224:133–148, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p133.pdf.

Barlow:2002:SIV

[BC02] K. E. Barlow and J. P. Croxall. Seasonal and interannual variation in foraging range and habitat of macaroni penguins Eudyptes chrysolophus at South Georgia. Marine Ecology Progress Series, 232:291–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/232/m232p291.pdf.

Beninger:2006:APP

[BC06a] P. G. Beninger and R. Cannuel. Acquisition of particle processing capability in the oyster Crassostrea gigas: ontogeny of the mantle pseudofeces rejection tracts. Marine Ecology Progress Series, 325:153–163, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p153.pdf.

Blockley:2006:RDD

[BC06b] D. J. Blockley and M. G. Chapman. Recruitment determines differences between assemblages on shaded or unshaded seawalls. Marine Ecology Progress Series, 327:27– 36, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p027.pdf.

Benedetti-Cecchi:2000:PEB

[BCABC00] L. Benedetti-Cecchi, S. Acunto, F. Bulleri, and F. Cinelli. Population ecology of the barnacle Chthamalus stellatus in the northwest Mediterranean. Marine Ecology Progress Series, 198:157–170, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p157.pdf.

Berndt:2002:GCT

[BCB02] M. L. Berndt, J. A. Callow, and S. H. Brawley. Gamete concentrations and timing and success of fertilization in a rocky shore seaweed. Marine Ecology Progress Series, 226: 273–285, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p273.pdf.

Barille:2006:FRG

[BCB^{+}06] L. Barillé, B. Cognie, P. Beninger, P. Decottignies, and Y. Rincé. Feeding responses of the gastropod Crepidula fornicata to changes in seston concentration. Marine Ecology Progress Series, 322:169–178, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/322/m322p169.pdf.

Bradbury:2008:OEC

[BCB08a] I. R. Bradbury, S. E. Campana, and P. Bentzen. Otolith elemental composition and adult tagging reveal spawning site fidelity and estuarine dependency in rainbow smelt. Marine Ecology Progress Series, 368:255–268, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p255.pdf.

Brucet:2008:FNC

[BCB^{+}08b] S. Brucet, J. Compte, D. Boix, R. López-Flores, and X. D. Quintana. Feeding of nauplii, copepodites and adults of Calanipeda aquaedulcis (Calanoida) in Mediterranean salt marshes. Marine Ecology Progress Series, 355:183–191, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/355/m355p183.pdf.

Benedetti-Cecchi:2001:SVE

[BCBAC01] L. Benedetti-Cecchi, F. Bulleri, S. Acunto, and F. Cinelli. Scales of variation in the effects of limpets on rocky shores in the northwest Mediterranean. Marine Ecology Progress Series, 209:131–141, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/209/m209p131.pdf.

Bauder:2001:UFD

[BCBQ01] A. G. Bauder, A. D. Cembella, V. M. Bricelj, and M. A. Quilliam. Uptake and fate of diarrhetic shellfish poisoning toxins from the dinoflagellate Prorocentrum lima in the bay scallop Argopecten irradians. Marine Ecology Progress Series, 213:39–52, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p039.pdf.

Benedetti-Cecchi:2005:DSP

[BCBVM05] L. Benedetti-Cecchi, I. Bertocci, S. Vaselli, and E. Maggi. Determinants of spatial pattern at different scales in two populations of the marine alga Rissoella verruculosa. Marine Ecology Progress Series, 293:37–47, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/293/m293p037.pdf.

Benedetti-Cecchi:2006:MPV

[BCBVM06] L. Benedetti-Cecchi, I. Bertocci, S. Vaselli, and E. Maggi. Morphological plasticity and variable spatial patterns in different populations of the red alga Rissoella verrucosa. Marine Ecology Progress Series, 315:87–98, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p087.pdf.

Baumgartner:2003:ABN

[BCC^{+}03] M. F. Baumgartner, T. V. N. Cole, R. G. Campbell, G. J. Teegarden, and E. G. Durbin. Associations between North Atlantic right whales and their prey, Calanus finmarchicus, over diel and tidal time scales. Marine Ecology Progress Series, 264:155–166, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/264/m264p155.pdf.

Bost:2004:EDM

[BCC^{+}04] C. A. Bost, J. B. Charrassin, Y. Clerquin, Y. Ropert-Coudert, and Y. Le Maho. Exploitation of distant marginal ice zones by king penguins during winter. Marine Ecology Progress Series, 283:293–297, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/283/m283p293.pdf.

Buckley:2008:MSL

[BCC08] L. J. Buckley, E. M. Caldarone, and C. Clemmesen. Multispecies larval fish growth model based on temperature and fluorometrically derived RNA/DNA ratios: results from a meta-analysis. Marine Ecology Progress Series, 371:221– 232, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/371/m371p221.pdf.

Baumgartner:2003:NAR

[BCCM03] M. F. Baumgartner, T. V. N. Cole, P. J. Clapham, and B. R. Mate. North Atlantic right whale habitat in the lower Bay of Fundy and on the SW Scotian Shelf during 1999–2001. Marine Ecology Progress Series, 264:137–154, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p137.pdf.

Bocher:2000:CTS

[BCH00] P. Bocher, Y. Cherel, and K. A. Hobson. Complete trophic segregation between South Georgian and common diving petrels during breeding at Iles Kerguelen. Marine Ecology Progress Series, 208:249–264, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/208/m208p249.pdf.

Bajzak:2009:IDP

[BCHS09] C. E. Bajzak, S. D. Côté, M. O. Hammill, and G. Stenson. Intersexual differences in the postbreeding foraging behaviour of the Northwest Atlantic hooded seal. Marine Ecology Progress Series, 385:285–294, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/385/m385p285.pdf.

Beninger:2005:PPG

[BCJ05] P. G. Beninger, R. Cannuel, and S. Jaunet. Particle processing on the gill plicae of the oyster Crassostrea gigas: fine-scale mucocyte distribution and functional correlates. Marine Ecology Progress Series, 295:191–199, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/295/m295p191.pdf.

Bromage:2009:QCH

[BCKP09] E. Bromage, L. Carpenter, S. Kaattari, and M. Patterson. Quantification of coral heat shock proteins from individual coral polyps. Marine Ecology Progress Series, 376:123– 132, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p123.pdf.

Bocher:2001:ABF

[BCL^{+}01] P. Bocher, Y. Cherel, J. P. Labat, P. Mayzaud, S. Razouls, and P. Jouventin. Amphipod-based food web: Themisto gaudichaudii caught in nets and by seabirds in Kerguelen waters, southern Indian Ocean. Marine Ecology Progress Series, 223:261–276, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p261.pdf.

Balestri:2003:SVP

[BCL03] E. Balestri, F. Cinelli, and C. Lardicci. Spatial variation in Posidonia oceanica structural, morphological and dynamic features in a northwestern Mediterranean coastal area: a multi-scale analysis. Marine Ecology Progress Series, 250: 51–60, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p051.pdf.

Baird:2007:FFC

[BCLRL07] A. H. Baird, V. R. Cumbo, W. Leggat, and M. Rodriguez-Lanetty. Fidelity and flexibility in coral symbioses. Marine Ecology Progress Series, 347:307–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/347/
m347p307.pdf.

Buckley:2006:OST

[BCLS06] L. J. Buckley, E. M. Caldarone, R. G. Lough, and J. M. St. Onge-Burns. Ontogenetic and seasonal trends in recent growth rates of Atlantic cod and haddock larvae on Georges Bank: effects of photoperiod and temperature. Marine Ecology Progress Series, 325:205–226, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p205.pdf.

Bekkevold:2007:DOS

[BCM^{+}07] D. Bekkevold, L. A. W. Clausen, S. Mariani, C. André, T. B. Christensen, and H. Mosegaard. Divergent origins of sympatric herring population components determined using genetic mixture analysis. Marine Ecology Progress Series, 337:187–196, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p187.pdf.

Bissett:2009:EOP

[BCM^{+}09] A. Bissett, P. L. M. Cook, C. Macleod, J. P. Bowman, and C. Burke. Effects of organic perturbation on marine sediment betaproteobacterial ammonia oxidizers and on benthic nitrogen biogeochemistry. Marine Ecology Progress Series, 392:17–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p017.pdf.

Benedetti-Cecchi:2007:RME

[BCO07] L. Benedetti-Cecchi and G. C. Osio. Replication and mitigation of effects of confounding variables in environmental impact assessment: effect of marinas on rocky-shore assemblages. Marine Ecology Progress Series, 334:21–35, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p021.pdf.

Bonhommeau:2008:ICE

[BCP^{+}08] S. Bonhommeau, E. Chassot, B. Planque, E. Rivot, A. H. Knap, and O. Le Pape. Impact of climate on eel populations of the northern hemisphere. Marine Ecology Progress Series, 373:71–80, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/373/m373p071.pdf.

Benedetti-Cecchi:2001:PCA

[BCPB^{+}01] L. Benedetti-Cecchi, F. Pannacciulli, F. Bulleri, P. S. Moschella, L. Airoldi, G. Relini, and F. Cinelli. Predicting the consequences of anthropogenic disturbance: large-scale effects of loss of canopy algae on rocky shores. Marine Ecology Progress Series, 214:137–150, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p137.pdf.

Bowden:2006:ASM

[BCPB06] D. A. Bowden, A. Clarke, L. S. Peck, and D. K. A. Barnes. Antarctic sessile marine benthos: colonisation and growth on artificial substrata over three years. Marine Ecology Progress Series, 316:1–16, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bishop:2006:CNN

[BCS^{+}06] M. J. Bishop, R. B. Carnegie, N. A. Stokes, C. H. Peterson, and E. M. Burreson. Complications of a non-native oyster introduction: facilitation of a local parasite. Marine Ecology Progress Series, 325:145–152, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p145.pdf.

Belt:2009:IBI

[BCS09] K. M. Belt, S. W. B. Cole, and R. A. Scrosati. Intertidal barnacles as indicators of the intensity of scour by sea ice. Marine Ecology Progress Series, 381:183–187, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/381/
m381p183.pdf.

Bishop:2008:SSP

[BCT^{+}08] M. J. Bishop, M. R. Cole, S. L. Taylor, E. M. Wilkie, and B. P. Kelaher. Size-specific predation by dominant consumers maintains a ‘trophic cul-de-sac’. Marine Ecology Progress Series, 354:75–83, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p075.pdf.

Bulleri:2004:PML

[BCU04] F. Bulleri, M. G. Chapman, and A. J. Underwood. Patterns of movement of the limpet Cellana tramoserica on rocky shores and retaining seawalls. Marine Ecology Progress Series, 281:121–129, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p121.pdf.

Brunet:2007:VVD

[BCVC07] C. Brunet, R. Casotti, V. Vantrepotte, and F. Conversano. Vertical variability and diel dynamics of picophytoplankton in the Strait of Sicily, Mediterranean Sea, in summer. Marine Ecology Progress Series, 346:15–26, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p015.pdf.

Barnes:2000:ETH

[BD00] D. K. A. Barnes and S. De Grave. Ecology of tropical hermit crabs at Quirimba Island, Mozambique: niche width and resource allocation. Marine Ecology Progress Series, 206: 171–179, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/206/m206p171.pdf.

Burd:2001:FVL

[BD01] A. B. Burd and K. H. Dunton. Field verification of a lightdriven model of biomass changes in the seagrass Halodule wrightii. Marine Ecology Progress Series, 209:85–98, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
209/m209p085.pdf.

Besiktepe:2002:CID

[BD02] S. Besiktepe and H. G. Dam. Coupling of ingestion and defecation as a function of diet in the calanoid copepod Acartia tonsa. Marine Ecology Progress Series, 229:151–164, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/229/m229p151.pdf. See erratum [Ano02c].

Beukema:2003:RSS

[BD03] J. J. Beukema and R. Dekker. Redistribution of spat-sized Macoma balthica in the Wadden Sea in cold and mild winters. Marine Ecology Progress Series, 265:117–122, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p117.pdf.

Beukema:2006:ACC

[BD06] J. J. Beukema and R. Dekker. Annual cockle Cerastoderma edule production in the Wadden Sea usually fails to sustain both wintering birds and a commercial fishery. Marine Ecology Progress Series, 309:189–204, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/309/m309p189.pdf.

Barry:2007:RDS

[BD07] J. P. Barry and J. C. Drazen. Response of deep-sea scavengers to ocean acidification and the odor from a dead grenadier. Marine Ecology Progress Series, 350:193–207, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m350p193.pdf.

Brown:2008:SRM

[BD08] B. E. Brown and R. P. Dunne. Solar radiation modulates bleaching and damage protection in a shallow water coral. Marine Ecology Progress Series, 362:99–107, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p099.pdf.

Barron:2009:DOM

[BD09] C. Barrón and C. M. Duarte. Dissolved organic matter release in a Posidonia oceanica meadow. Marine Ecology Progress Series, 374:75–84, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/374/m374p075.pdf.

Beukema:2011:BSD

[BD11] J. J. Beukema and R. Dekker. Benthic species diversity and sediment composition: Comment on Compton et al. (2008).

Marine Ecology Progress Series, 440:281–283, 2011. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2011/440/

m440p281.pdf. See [CTvdM^{+}08].

Bulit:2004:AST

[BDÁM04] C. Bulit, C. Díaz-Ávalos, and D. J. S. Montagnes. Assessing spatial and temporal patchiness of the autotrophic ciliate Myrionecta rubra: a case study in a coastal lagoon. Marine Ecology Progress Series, 268:55–67, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p055.pdf.

Boelen:2000:UID

[BdBK^{+}00] P. Boelen, M. K. de Boer, G. W. Kraay, M. J. W. Veldhuis, and A. G. J. Buma. UVBR-induced DNA damage in natural marine picoplankton assemblages in the tropical Atlantic Ocean. Marine Ecology Progress Series, 193:1–9, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p001.pdf.

Bartholomew:2000:NDI

[BDC00a] A. Bartholomew, R. J. Diaz, and G. Cicchetti. New dimensionless indices of structural habitat complexity: predicted and actual effects on a predator’s foraging success. Marine Ecology Progress Series, 206:45–58, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/206/m206p045.pdf.

Buffan-Dubau:2000:EBM

[BDC00b] E. Buffan-Dubau and K. R. Carman. Extraction of benthic microalgal pigments for HPLC analyses. Marine Ecology Progress Series, 204:293–297, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/204/m204p293.pdf.

Bolle:2009:VTF

[BDCvB^{+}09] L. J. Bolle, M. Dickey-Collas, J. K. L. van Beek, P. L. A. Erftemeijer, J. I. J. Witte, H. W. van der Veer, and A. D. Rijnsdorp. Variability in transport of fish eggs and larvae. III. Effects of hydrodynamics and larval behaviour on recruitment in plaice.

Marine Ecology Progress Series, 390:195–211, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p195.pdf.

Brown:2002:EBT

[BDDG02] B. E. Brown, C. A. Downs, R. P. Dunne, and S. W. Gibb. Exploring the basis of thermotolerance in the reef coral Goniastrea aspera. Marine Ecology Progress Series, 242:119– 129, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p119.pdf.

Beninger:2003:PPM

[BDDL03] P. G. Beninger, S. C. Dufour, P. Decottignies, and M. Le Pennec. Particle processing mechanisms in the archaic, perihydrothermal vent bivalve Bathypecten vulcani, inferred from cilia and mucocyte distributions on the gill. Marine Ecology Progress Series, 246:183–195, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/246/m246p183.pdf.

Beukema:2001:SRS

[BDEM01] J. J. Beukema, R. Dekker, K. Essink, and H. Michaelis. Synchronized reproductive success in the main bivalve species in the Wadden Sea: causes and consequences. Marine Ecology Progress Series, 211:143–155, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/211/m211p143.pdf.

Belgrano:2004:EPC

[BDF^{+}04] A. Belgrano, J. A. Dunne, Z. V. Finkel, D. Grünbaum, H. Hillebrand, D. Ø. Hjermann, A. J. Irwin, M. Lima, N. D. Martinez, G. Ottersen, J. K. Parrish, O. Schofield, N. C. Stenseth, S. V. Viscido, and R. J. William. Emergent properties of complex marine systems: a macroecological perspective. Marine Ecology Progress Series, 273:227–302, 2004. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m273_TS.pdf.

Bugoni:2009:MHU

[BDF09] L. Bugoni, L. D’Alba, and R. W. Furness. Marine habitat use of wintering spectacled petrels Procellaria conspicillata, and overlap with longline fishery. Marine Ecology Progress Series, 374:273–285, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/374/m374p273.pdf.

Beauplet:2004:FES

[BDG^{+}04] G. Beauplet, L. Dubroca, C. Guinet, Y. Cherel, W. Dabin, C. Gagne, and M. Hindell. Foraging ecology of subantarctic fur seals Arctocephalus tropicalis breeding on Amsterdam Island: seasonal changes in relation to maternal characteristics and pup growth. Marine Ecology Progress Series, 273:211–225, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p211.pdf.

Beninger:2007:CPP

[BDG^{+}07] P. G. Beninger, P. Decottignies, F. Guiheneuf, L. Barillé, and Y. Rincé. Comparison of particle processing by two introduced suspension feeders: selection in Crepidula fornicata and Crassostrea gigas. Marine Ecology Progress Series, 334:165–177, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p165.pdf.

Bosire:2006:SRR

[BDGK^{+}06] J. O. Bosire, F. Dahdouh-Guebas, J. G. Kairo, S. Wartel, J. Kazungu, and N. Koedam. Success rates of recruited tree species and their contribution to the structural development of reforested mangrove stands. Marine Ecology Progress Series, 325:85–91, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/325/m325p085.pdf.

Bejarano:2007:EAB

[BDGS07] A. C. Bejarano, F. M. V. Dolah, F. M. Gulland, and L. Schwacke. Exposure assessment of the biotoxin domoic acid in California sea lions: application of a bioenergetic model. Marine Ecology Progress Series, 345:293–304, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p293.pdf.

Beukema:2001:MLE

[BDH01] J. J. Beukema, J. Drent, and P. J. C. Honkoop. Maximizing lifetime egg production in a Wadden Sea population of the tellinid bivalve Macoma balthica: a trade-off between immediate and future reproductive outputs. Marine Ecology Progress Series, 209:119–129, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/209/m209p119.pdf.

Beukema:2009:SLI

[BDJ09] J. J. Beukema, R. Dekker, and J. M. Jansen. Some like it cold: populations of the tellinid bivalve Macoma balthica (L.) suffer in various ways from a warming climate. Marine Ecology Progress Series, 384:135–145, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m384p135.pdf.

Bierne:2002:CHS

[BDLB02] N. Bierne, P. David, A. Langlade, and F. Bonhomme. Can habitat specialisation maintain a mosaic hybrid zone in marine bivalves? Marine Ecology Progress Series, 245:157– 170, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p157.pdf.

Beninger:2004:LQP

[BDR04] P. G. Beninger, P. Decottignies, and Y. Rincé. Localization of qualitative particle selection sites in the heterorhabdic filibranch Pecten maximus (Bivalvia: Pectinidae). Marine Ecology Progress Series, 275:163–173, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p163.pdf.

Barnes:2006:SPD

[BDR06] P. B. Barnes, A. R. Davis, and D. E. Roberts. Sampling patchily distributed taxa: a case study using cost-benefit analyses for sponges and ascidians in coastal lakes of New South Wales, Australia. Marine Ecology Progress Series, 319: 55–64, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p055.pdf.

Bourque:2006:CPR

[BDRM06] J. F. Bourque, J. J. Dodson, D. A. J. Ryan, and D. J. Marcogliese. Cestode parasitism as a regulator of early lifehistory survival in an estuarine population of rainbow smelt Osmerus mordax. Marine Ecology Progress Series, 314:295– 307, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p295.pdf.

Bouma:2001:MBM

[BdVD^{+}01] H. Bouma, P. P. de Vries, J. M. C. Duiker, P. M. J. Herman, and W. J. Wolff. Migration of the bivalve Macoma balthica on a highly dynamic tidal flat in the Westerschelde estuary, The Netherlands. Marine Ecology Progress Series, 224:157–170, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p157.pdf.

Brown:2000:DRP

[BDW^{+}00] B. E. Brown, R. P. Dunne, M. E. Warner, I. Ambarsari, W. K. Fitt, S. W. Gibb, and D. G. Cummings. Damage and recovery of photosystem II during a manipulative field experiment on solar bleaching in the coral Goniastrea aspera. Marine Ecology Progress Series, 195:117–124, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p117.pdf.

Beaugrand:2004:MMPa

[Bea04] G. Beaugrand. Monitoring marine plankton ecosystems. I: Description of an ecosystem approach based on plankton indicators. Marine Ecology Progress Series, 269:69–81, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/269/m269p069.pdf.

Byren:2006:USO

[BEE06] L. Byrén, G. Ejdung, and R. Elmgren. Uptake of sedimentary organic matter by the deposit-feeding Baltic amphipods Monoporeia affinis and Pontoporeia femorata. Marine Ecology Progress Series, 313:135–143, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/313/m313p135.pdf.

Brock:2009:ASP

[BEKT09] R. Brock, E. English, E. Kenchington, and M. Tasker. The alphabet soup that protects cold-water corals in the North Atlantic. Marine Ecology Progress Series, 397:355–360, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p355.pdf.

Belgrano:2004:FM

[Bel04] A. Belgrano. Foreword MEPS 273:227. Marine Ecology Progress Series, ??:??, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/273/m273p227.pdf.

Bell:2007:CPS

[Bel07] J. J. Bell. Contrasting patterns of species and functional composition of coral reef sponge assemblages. Marine Ecology Progress Series, 339:73–81, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/339/m339p073.pdf.

Bell:2008:SCP

[Bel08a] J. J. Bell. Similarity in connectivity patterns for two gastropod species lacking pelagic larvae. Marine Ecology Progress Series, 357:185–194, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/357/m357p185.pdf.

Bell:2008:SAB

[Bel08b] J. J. Bell. Sponges as agents of biological disturbance. Marine Ecology Progress Series, 368:127–135, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p127.pdf.

Berghahn:2000:REC

[Ber00] R. Berghahn. Response to extreme conditions in coastal areas: biological tags in flatfish otoliths. Marine Ecology Progress Series, 192:277–285, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/192/m192p277.pdf.

Berkelmans:2002:RCT

[Ber02a] R. Berkelmans. Reply comment: ‘Time-integrated thermal bleaching thresholds of reefs and their variation on the Great Barrier Reef’. Marine Ecology Progress Series, 237: 309–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p309.pdf. See [Ber02b].

Berkelmans:2002:TIT

[Ber02b] R. Berkelmans. Time-integrated thermal bleaching thresholds of reefs and their variation on the Great Barrier Reef. Marine Ecology Progress Series, 229:73–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/

m229p073.pdf. See comment [Dun02] and reply [Ber02a].

Billot:2003:CPH

[BER^{+}03] C. Billot, C. R. Engel, S. Rousvoal, B. Kloareg, and M. Valero. Current patterns, habitat discontinuities and population genetic structure: the case of the kelp Laminaria digitata in the English Channel. Marine Ecology Progress Series, 253: 111–121, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p111.pdf.

Brokovich:2008:DTZ

[BES^{+}08] E. Brokovich, S. Einbinder, N. Shashar, M. Kiflawi, and S. Kark. Descending to the twilight-zone: changes in coral reef fish assemblages along a depth gradient down to 65 m. Marine Ecology Progress Series, 371:253–262, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/371/
m371p253.pdf.

Bergquist:2007:USI

[BEU^{+}07] D. C. Bergquist, J. T. Eckner, I. A. Urcuyo, E. E. Cordes, S. Hourdez, S. A. Macko, and C. R. Fisher. Using stable isotopes and quantitative community characteristics to determine a local hydrothermal vent food web. Marine Ecology Progress Series, 330:49–65, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/330/m330p049.pdf.

Bell:2003:BRFa

[BEW03a] G. W. Bell, D. B. Eggleston, and T. G. Wolcott. Behavioral responses of free-ranging blue crabs to episodic hypoxia. I. Movement. Marine Ecology Progress Series, 259: 215–225, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/259/m259p215.pdf.

Bell:2003:BRFb

[BEW03b] G. W. Bell, D. B. Eggleston, and T. G. Wolcott. Behavioral responses of free-ranging blue crabs to episodic hypoxia. II. Feeding. Marine Ecology Progress Series, 259:227– 235, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/259/m259p227.pdf.

Baines:2001:IDB

[BF01a] S. B. Baines and N. S. Fisher. Interspecific differences in the bioconcentration of selenite by phytoplankton and their ecological implications. Marine Ecology Progress Series, 213: 1–12, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/213/m213p001.pdf.

Bellwood:2001:RSS

[BF01b] D. R. Bellwood and R. Fisher. Relative swimming speeds in reef fish larvae. Marine Ecology Progress Series, 211:299– 303, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/211/m211p299.pdf.

Barnes:2003:RFP

[BF03] D. K. A. Barnes and K. P. P. Fraser. Rafting by five phyla on man-made flotsam in the Southern Ocean. Marine Ecology Progress Series, 262:289–291, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p289.pdf.

Bertini:2004:BDB

[BF04] G. Bertini and A. Fransozo. Bathymetric distribution of brachyuran crab (Crustacea, Decapoda) communities on coastal soft bottoms off southeastern Brazil. Marine Ecology Progress Series, 279:193–200, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/279/m279p193.pdf.

Balasubramanian:2007:SSR

[BF07] H. Balasubramanian and R. S. Foster. Species and space: role of volume in organizing coral reef fish assemblages in SE Cuba. Marine Ecology Progress Series, 345:229–236, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p229.pdf.

Balch:2008:OAD

[BF08a] W. M. Balch and V. J. Fabry. Ocean acidification: documenting its impact on calcifying phytoplankton at basin scales. Marine Ecology Progress Series, 373:239–247, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p239.pdf.

Bouchard:2008:EPH

[BF08b] C. Bouchard and L. Fortier. Effects of polynyas on the hatching season, early growth and survival of polar cod Boreogadus saida in the Laptev Sea. Marine Ecology Progress Series, 355:247–256, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m355p247.pdf.

Beck:2001:MWH

[BFB01] N. G. Beck, A. T. Fisher, and K. W. Bruland. Modeling water, heat, and oxygen budgets in a tidally dominated estuarine pond. Marine Ecology Progress Series, 217:43–58, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p043.pdf.

Bonaviri:2009:FVS

[BFB^{+}09] C. Bonaviri, T. Vega Fernández, F. Badalamenti, P. Gianguzza, M. Di Lorenzo, and S. Riggio. Fish versus starfish predation in controlling sea urchin populations in Mediterranean rocky shores. Marine Ecology Progress Series, 382: 129–138, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/382/m382p129.pdf.

Brante:2003:RIC

[BFE^{+}03] A. Brante, M. Fernández, L. Eckerle, F. Mark, H. O. Pörtner, and W. Arntz. Reproductive investment in the crab Cancer setosus along a latitudinal cline: egg production, embryo losses and embryo ventilation. Marine Ecology Progress Series, 251:221–232, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p221.pdf.

Brodeur:2005:SCC

[BFE^{+}05] R. D. Brodeur, J. P. Fisher, R. L. Emmett, C. A. Morgan, and E. Casillas. Species composition and community structure of pelagic nekton off Oregon and Washington under variable oceanographic conditions. Marine Ecology Progress Series, 298:41–57, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p041.pdf.

Baines:2005:ITD

[BFK05a] S. B. Baines, N. S. Fisher, and E. L. Kinney. Influence of temperature on dietary metal uptake in Arctic and temperate mussels. Marine Ecology Progress Series, 289:201– 213, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/289/m289p201.pdf.

Bergquist:2005:VSM

[BFK^{+}05b] D. C. Bergquist, C. Fleckenstein, J. Knisel, B. Begley, I. R. MacDonald, and C. R. Fisher. Variations in seep mussel bed communities along physical and chemical environmental gradients. Marine Ecology Progress Series, 293:99– 108, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p099.pdf.

Baines:2006:ETU

[BFK06] S. B. Baines, N. S. Fisher, and E. L. Kinney. Effects of temperature on uptake of aqueous metals by blue mussels Mytilus edulis from Arctic and temperate waters. Marine Ecology Progress Series, 308:117–128, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/308/m308p117.pdf.

Bouma:2009:ESS

[BFK^{+}09] T. J. Bouma, M. Friedrichs, P. Klaassen, B. K. van Wesenbeeck, F. G. Brun, S. Temmerman, M. M. van Katwijk, G. Graf, and P. M. J. Herman. Effects of shoot stiffness, shoot size and current velocity on scouring sediment from around seedlings and propagules. Marine Ecology Progress Series, 388:293–297, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m388p293.pdf.

Briones-Fourzan:2001:EAS

[BFLÁ01] P. Briones-Fourzán and E. Lozano-Álvarez. Effects of artificial shelters (Casitas) on the abundance and biomass of juvenile spiny lobsters Panulirus argus in a habitat-limited tropical reef lagoon. Marine Ecology Progress Series, 221: 221–232, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/221/m221p221.pdf.

Brante:2008:EOC

[BFV08] A. Brante, M. Fernández, and F. Viard. Effect of oxygen conditions on intracapsular development in two calyptraeid species with different modes of larval development. Marine Ecology Progress Series, 368:197–207, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p197.pdf.

Bshary:2002:PDC

[BG02] R. Bshary and A. S. Grutter. Parasite distribution on client reef fish determines cleaner fish foraging patterns. Marine Ecology Progress Series, 235:217–222, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p217.pdf.

Bastidas:2004:SEM

[BG04a] C. Bastidas and E. M. García. Sublethal effects of mercury and its distribution in the coral Porites astreoides. Marine Ecology Progress Series, 267:133–143, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/267/m267p133.pdf.

Bolton:2004:MVA

[BG04b] T. F. Bolton and W. M. Graham. Morphological variation among populations of an invasive jellyfish. Marine Ecology Progress Series, 278:125–139, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p125.pdf.

Brokordt:2004:BGE

[BG04c] K. Brokordt and H. Guderley. Binding of glycolytic enzymes in adductor muscle of Iceland scallop Chlamys islandica is altered by reproductive status. Marine Ecology Progress Series, 268:141–149, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/268/m268p141.pdf.

Bensoussan:2007:CPP

[BG07a] N. Bensoussan and J. P. Gattuso. Community primary production and calcification in a NW Mediterranean ecosystem dominated by calcareous macroalgae. Marine Ecology Progress Series, 334:37–45, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p037.pdf.

Brousseau:2007:EPI

[BG07b] D. J. Brousseau and R. Goldberg. Effect of predation by the invasive crab Hemigrapsus sanguineus on recruiting barnacles Semibalanus balanoides in western Long Island Sound, USA. Marine Ecology Progress Series, 339:221–228, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p221.pdf.

Baird:2009:SSS

[BG09] A. H. Baird and J. R. Guest. Spawning synchrony in scleractinian corals: Comment on Mangubhai & Harrison (2008). Marine Ecology Progress Series, 374:301–304, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/

m374p301.pdf. See [MH08b] and reply [MH09].

Bearon:2006:ESS

[BGC06] R. N. Bearon, D. Grünbaum, and R. A. Cattolico. Effects of salinity structure on swimming behavior and harmful algal bloom formation in Heterosigma akashiwo, a toxic raphidophyte. Marine Ecology Progress Series, 306:153–163, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/306/m306p153.pdf.

Boisson:2003:CRS

[BGD^{+}03] F. Boisson, F. Goudard, J. P. Durand, C. Barbot, J. Pieri, J. C. Amiard, and S. W. Fowler. Comparative radiotracer study of cadmium uptake, storage, detoxification and depuration in the oyster Crassostrea gigas: potential adaptive mechanisms. Marine Ecology Progress Series, 254:177–186, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/254/m254p177.pdf.

Baker:2005:SPP

[BGDS05] C. S. Baker, C. Garrigue, R. Dodemont, and D. Steel. Statistical power and plausible alternate hypotheses of gametic recapture estimates from wintering grounds of humpback whales. Marine Ecology Progress Series, 295:309–311, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p309.pdf.

Barreiro:2006:FSC

[BGF^{+}06] A. Barreiro, C. Guisande, M. Frangópulos, A. González-Fernández, S. Muñoz, D. Pérez, S. Magadán, I. Maneiro, I. Riveiro, and P. Iglesias. Feeding strategies of the copepod Acartia clausi on single and mixed diets of toxic and non-toxic strains of the dinoflagellate Alexandrium minutum.

Marine Ecology Progress Series, 316:115–125, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/316/
m316p115.pdf.

Barki:2002:SCN

[BGGR02] Yael Barki, Daniel Gateño, Dan Graur, and Baruch Rinkevich. Soft-coral natural chimerism: a window in ontogeny allows the creation of entities comprised of incongruous parts. Marine Ecology Progress Series, 231:91–99, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p091.pdf.

Buresch:2006:MGS

[BGH06] K. C. Buresch, G. Gerlach, and R. T. Hanlon. Multiple genetic stocks of longfin squid Loligo pealeii in the NW Atlantic: stocks segregate inshore in summer, but aggregate offshore in winter. Marine Ecology Progress Series, 310:263– 270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/310/m310p263.pdf.

Basterretxea:2007:MNH

[BGJ^{+}07] G. Basterretxea, E. Garcés, A. Jordi, S. Anglès, and M. Masó. Modulation of nearshore harmful algal blooms by in situ growth rate and water renewal. Marine Ecology Progress Series, 352:53–65, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m352p053.pdf.

Baumann:2006:TIR

[BGK^{+}06] H. Baumann, T. Gröhsler, G. Kornilovs, A. Makarchouk, V. Feldmann, and A. Temming. Temperature-induced regional and temporal growth differences in Baltic young-of-the-year sprat Sprattus sprattus. Marine Ecology Progress Series, 317:225–236, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p225.pdf.

Barnes:2009:GRS

[BGK09] D. K. A. Barnes, H. J. Griffiths, and S. Kaiser. Geographic range shift responses to climate change by Antarctic benthos: where we should look. Marine Ecology Progress Series, 393:13–26, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/393/m393p013.pdf.

Bjork:2000:LSD

[BGKN00] M. Björk, M. Gilek, N. Kautsky, and C. Näf. In situ determination of PCB biodeposition by Mytilus edulis in a Baltic coastal ecosystem. Marine Ecology Progress Series, 194:193–201, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p193.pdf.

Brehmer:2007:SBS

[BGL^{+}07] P. Brehmer, F. Gerlotto, C. Laurent, P. Cotel, A. Achury, and B. Samb. Schooling behaviour of small pelagic fish: phenotypic expression of independent stimuli. Marine Ecology Progress Series, 334:263–272, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/334/m334p263.pdf.

Balestri:2009:SNC

[BGLL09] E. Balestri, S. Gobert, G. Lepoint, and C. Lardicci. Seed nutrient content and nutritional status of Posidonia oceanica seedlings in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 388:99–109, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/388/
m388p099.pdf.

BenRaisLasram:2009:FDP

[BGM09] F. Ben Rais Lasram, F. Guilhaumon, and D. Mouillot. Fish diversity patterns in the Mediterranean Sea: deviations from a mid-domain model. Marine Ecology Progress Series, 376: 253–267, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p253.pdf.

Bochdansky:2008:FLL

[BGPL08] A. B. Bochdansky, P. Grønkjær, P. Pepin, and W. C. Leggett. Food limitation in larval fish: ontogenetic variation in feeding scope and its potential effect on survival. Marine Ecology Progress Series, 367:239–248, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p239.pdf.

Barreiro:2001:IOB

[BGQR01] R. Barreiro, R. González, M. Quintela, and J. M. Ruiz. Imposex, organotin bioaccumulation and sterility of female Nassarius reticulatus in polluted areas of NW Spain. Marine Ecology Progress Series, 218:203–212, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/218/m218p203.pdf.

Broderick:2000:IPS

[BGRD00] A. C. Broderick, B. J. Godley, S. Reece, and J. R. Downie. Incubation periods and sex ratios of green turtles: highly female biased hatchling production in the eastern Mediterranean. Marine Ecology Progress Series, 202:273–281, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p273.pdf.

Borgaa:2002:DCL

[BGS02a] K. Borgå, G. W. Gabrielsen, and J. U. Skaare. Differences in contamination load between pelagic and sympagic invertebrates in the Arctic marginal ice zone: influence of habitat, diet and geography. Marine Ecology Progress Series, 235:157–169, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p157.pdf.

Breteler:2002:SCI

[BGS^{+}02b] W. C. M. Klein Breteler, K. Grice, S. Schouten, H. T. Kloosterhuis, and J. S. S. Damsté. Stable carbon isotope fractionation in the marine copepod Temora longicornis: unexpectedly low δ ^{13}C value of faecal pellets. Marine Ecology Progress Series, 240:195–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p195.pdf.

Becker:2002:FML

[BGSBC02] P. H. Becker, J. González-Solís, B. Behrends, and J. Croxall. Feather mercury levels in seabirds at South Georgia: influence of trophic position, sex and age. Marine Ecology Progress Series, 243:261–269, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/243/m243p261.pdf.

Brante:2001:EHP

[BH01] A. Brante and R. N. Hughes. Effect of hypoxia on the preyhandling behaviour of Carcinus maenas feeding on Mytilus edulis. Marine Ecology Progress Series, 209:301–305, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
209/m209p301.pdf.

Bullard:2002:PTA

[BH02] S. G. Bullard and M. E. Hay. Plankton tethering to assess spatial patterns of predation risk over a coral reef and seagrass bed. Marine Ecology Progress Series, 225:17–28, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p017.pdf.

Bunker:2004:FMP

[BH04] A. J. Bunker and A. G. Hirst. Fecundity of marine planktonic copepods: global rates and patterns in relation to chlorophylla, temperature and body weight. Marine Ecology Progress Series, 279:161–181, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/279/m279p161.pdf.

Boyle:2006:CFG

[BH06] K. S. Boyle and M. H. Horn. Comparison of feeding guild structure and ecomorphology of intertidal fish assemblages from central California and central Chile. Marine Ecology Progress Series, 319:65–84, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/319/m319p065.pdf.

Beaugrand:2008:SPA

[BH08] G. Beaugrand and P. Helaouët. Simple procedures to assess and compare the ecological niche of species. Marine Ecology Progress Series, 363:29–37, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/363/m363p029.pdf.

Booth:2009:EAO

[BH09a] D. M. Booth and K. L. Heck, Jr. Effects of the American oyster Crassostrea virginica on growth rates of the seagrass Halodule wrightii. Marine Ecology Progress Series, 389:117–126, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m389p117.pdf.

Burkepile:2009:NVH

[BH09b] D. E. Burkepile and M. E. Hay. Nutrient versus herbivore control of macroalgal community development and coral growth on a Caribbean reef. Marine Ecology Progress Series, 389:71–84, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p071.pdf.

Browman:2000:TSE

[BHC^{+}00] Howard I. Browman, Jeffrey A. Hutchings, David O. Conover, Kevin Stokes, Richard Law, and Carl Walters. Theme section: ’Evolution’ of fisheries science. Marine Ecology Progress Series, 208:299–313, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p299.pdf.

Burger:2004:SAS

[BHD04] A. E. Burger, C. L. Hitchcock, and G. K. Davoren. Spatial aggregations of seabirds and their prey on the continental shelf off SW Vancouver Island. Marine Ecology Progress Series, 283:279–292, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/283/m283p279.pdf.

Boissin:2008:ESR

[BHFC08] E. Boissin, T. B. Hoareau, J. P. Féral, and A. Chenuil. Extreme selfing rates in the cosmopolitan brittle star species complex Amphipholis squamata: data from progeny-array and heterozygote deficiency. Marine Ecology Progress Series, 361:151–159, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p151.pdf.

Baums:2005:MML

[BHH05] I. B. Baums, C. R. Hughes, and M. E. Hellberg. Mendelian microsatellite loci for the Caribbean coral Acropora palmata. Marine Ecology Progress Series, 288:115–127, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p115.pdf.

Bowlby:2007:RDB

[BHH07] H. D. Bowlby, J. M. Hanson, and J. A. Hutchings. Resident and dispersal behavior among individuals within a population of American lobster Homarus americanus. Marine Ecology Progress Series, 331:207–218, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/331/m331p207.pdf. See

[Kim06].

Bertram:2009:SND

[BHH09] D. F. Bertram, A. Harfenist, and A. Hedd. Seabird nestling diets reflect latitudinal temperature-dependent variation in availability of key zooplankton prey populations. Marine Ecology Progress Series, 393:199–210, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m393p199.pdf.

Buresch:2001:MDM

[BHMR01] K. M. Buresch, R. T. Hanlon, M. R. Maxwell, and S. Ring. Microsatellite DNA markers indicate a high frequency of multiple paternity within individual field-collected egg capsules of the squid Loligo pealeii. Marine Ecology Progress Series, 210:161–165, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p161.pdf.

Brooks:2009:TME

[BHN09] L. B. Brooks, J. T. Harvey, and W. J. Nichols. Tidal movements of East Pacific green turtle Chelonia mydas at a foraging area in Baja California Sur, México. Marine Ecology Progress Series, 386:263–274, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/386/m386p263.pdf.

Baduini:2006:PDR

[BHPC06] C. L. Baduini, G. L. Hunt, Jr., A. I. Pinchuk, and K. O. Coyle. Patterns in diet reveal foraging site fidelity of short-tailed shearwaters in the southeastern Bering Sea. Marine Ecology Progress Series, 320:279–292, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p279.pdf.

Burrows:2008:WEI

[BHR08] M. T. Burrows, R. Harvey, and L. Robb. Wave exposure indices from digital coastlines and the prediction of rocky shore community structure. Marine Ecology Progress Series, 353:1– 12, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bythell:2000:LVL

[BHSR00] J. C. Bythell, Z. M. Hillis-Starr, and C. S. Rogers. Local variability but landscape stability in coral reef communites following repeated hurricane impacts. Marine Ecology Progress Series, 204:93–100, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p093.pdf.

Brun:2002:ATA

[BHV^{+}02] F. G. Brun, I. Hernández, J. J. Vergara, G. Peralta, and J. L. Pérez-Lloréns. Assessing the toxicity of ammonium pulses to the survival and growth of Zostera noltii. Marine Ecology Progress Series, 225:177–187, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/225/m225p177.pdf.

Backhaus:2003:CPP

[BHW^{+}03] J. O. Backhaus, E. N. Hegseth, H. Wehde, X. Irigoien, K. Hatten, and K. Logemann. Convection and primary production in winter. Marine Ecology Progress Series, 251: 1–14, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/251/m251p001.pdf.

Brooke:2009:STT

[BHY09] S. D. Brooke, M. W. Holmes, and C. M. Young. Sediment tolerance of two different morphotypes of the deep-sea coral Lophelia pertusa from the Gulf of Mexico. Marine Ecology Progress Series, 390:137–144, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/390/m390p137.pdf.

Brown:2008:DFS

[BHZW08] C. J. Brown, A. J. Hobday, P. E. Ziegler, and D. C. Welsford. Darwinian fisheries science needs to consider realistic fishing pressures over evolutionary time scales. Marine Ecology Progress Series, 369:257–266, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/369/m369p257.pdf.

Beaugrand:2002:SDC

[BI02] G. Beaugrand and F. Ibañez. Spatial dependence of calanoid copepod diversity in the North Atlantic Ocean. Marine Ecology Progress Series, 232:197–211, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p197.pdf.

Burns:2003:CAPb

[BI03] E. Burns and M. Ilan. Comparison of anti-predatory defenses of Red Sea and Caribbean sponges. II. Physical defense. Marine Ecology Progress Series, 252:115–123, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p115.pdf.

Beaugrand:2004:MMPb

[BI04] G. Beaugrand and F. Ibanez. Monitoring marine plankton ecosystems. II: Long-term changes in North Sea calanoid copepods in relation to hydro-climatic variability. Marine Ecology Progress Series, 284:35–47, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/284/m284p035.pdf.

Biber:2007:TPD

[Bib07] P. D. Biber. Transport and persistence of drifting macroalgae (Rhodophyta) are strongly influenced by flow velocity and substratum complexity in tropical seagrass habitats. Marine Ecology Progress Series, 343:115–122, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/343/
m343p115.pdf.

Burns:2003:CAPa

[BIC^{+}03] E. Burns, I. Ifrach, S. Carmeli, J. R. Pawlik, and M. Ilan. Comparison of anti-predatory defenses of Red Sea and Caribbean sponges. I. Chemical defense. Marine Ecology Progress Series, 252:105–114, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p105.pdf.

Botto:2006:EIP

[BIG^{+}06] F. Botto, O. Iribarne, J. Gutierrez, J. Bava, A. Gagliardini, and I. Valiela. Ecological importance of passive deposition of organic matter into burrows of the SW Atlantic crab Chasmagnathus granulatus. Marine Ecology Progress Series, 312:201–210, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/312/m312p201.pdf.

Beaugrand:2001:GDS

[BIL01] G. Beaugrand, F. Ibañez, and J. A. Lindley. Geographical distribution and seasonal and diel changes in the diversity of calanoid copepods in the North Atlantic and North Sea. Marine Ecology Progress Series, 219:189–203, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p189.pdf.

Beaugrand:2002:DCC

[BIL^{+}02] G. Beaugrand, F. Ibañez, J. A. Lindley, C. Philip, and P. C. Reid. Diversity of calanoid copepods in the North Atlantic and adjacent seas: species associations and biogeography. Marine Ecology Progress Series, 232:179–195, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p179.pdf.

Beaugrand:2000:SSL

[BIR00] G. Beaugrand, F. Ibañez, and P. C. Reid. Spatial, seasonal and long-term fluctuations of plankton in relation to hydroclimatic features in the English Channel, Celtic Sea and Bay of Biscay. Marine Ecology Progress Series, 200:93–102, 2000.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p093.pdf.

Berger:2008:SPM

[BJ08] M. S. Berger and D. E. Jelinski. Spatial patterns of maternal investment in Strongylocentrotus franciscanus along a marine-terrestrial gradient. Marine Ecology Progress Series, 364:119–127, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p119.pdf.

Bean:2002:RAD

[BJC02] K. Bean, G. P. Jones, and M. J. Caley. Relationships among distribution, abundance and microhabitat specialisation in a guild of coral reef triggerfish (family balistidae). Marine Ecology Progress Series, 233:263–272, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p263.pdf.

Begin:2004:MCD

[BJH04] C. Bégin, L. E. Johnson, and J. H. Himmelman. Macroalgal canopies: distribution and diversity of associated invertebrates and effects on the recruitment and growth of mussels. Marine Ecology Progress Series, 271:121–132, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p121.pdf.

Broglio:2001:TIB

[BJJ01] E. Broglio, M. Johansson, and P. R. Jonsson. Trophic interaction between copepods and ciliates: effects of prey swimming behavior on predation risk. Marine Ecology Progress Series, 220:179–186, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p179.pdf.

Berntsson:2004:RUS

[BJLH04] K. M. Berntsson, P. R. Jonsson, A. I. Larsson, and S. Holdt. Rejection of unsuitable substrata as a potential driver of aggregated settlement in the barnacle Balanus improvisus. Marine Ecology Progress Series, 275:199–210, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p199.pdf.

Berge:2005:OTO

[BJN^{+}05] J. Berge, G. Johnsen, F. Nilsen, B. Gulliksen, and D. Slagstad. Ocean temperature oscillations enable reappearance of blue mussels Mytilus edulis in Svalbard after a 1000 year absence. Marine Ecology Progress Series, 303:167–175, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m303p167.pdf.

Berge:2006:MEP

[BJN^{+}06] J. Berge, G. Johnsen, F. Nilsen, B. Gulliksen, D. Slagstad, and D. M. Pampanin. The Mytilus edulis population in Svalbard: how and why. Marine Ecology Progress Series, 309:305–306, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/309/m309p305.pdf.

Breed:2004:SCE

[BJR04] G. A. Breed, G. A. Jackson, and T. L. Richardson. Sedimentation, carbon export and food web structure in the Mississippi River plume described by inverse analysis. Marine Ecology Progress Series, 278:35–51, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p035.pdf.

Barranguet:2000:EPP

[BK00] C. Barranguet and J. Kromkamp. Estimating primary production rates from photosynthetic electron transport in estuarine microphytobenthos. Marine Ecology Progress Series, 204:39–52, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p039.pdf.

Barnes:2003:HPS

[BK03] D. K. A. Barnes and P. Kuklinski. High polar spatial competition: extreme hierarchies at extreme latitude. Marine Ecology Progress Series, 259:17–28, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/259/m259p017.pdf.

Barnes:2004:SDV

[BK04] D. K. A. Barnes and P. Kuklinski. Scale-dependent variation in competitive ability among encrusting Arctic species. Marine Ecology Progress Series, 275:21–32, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p021.pdf.

Bagoien:2005:BDMa

[BK05a] E. Bagøien and T. Kiørboe. Blind dating — mate finding in planktonic copepods. I. Tracking the pheromone trail of Centropages typicus. Marine Ecology Progress Series, 300: 105–115, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/300/m300p105.pdf.

Bagoien:2005:BDMb

[BK05b] E. Bagøien and T. Kiørboe. Blind dating — mate finding in planktonic copepods. III. Hydromechanical communication in Acartia tonsa. Marine Ecology Progress Series, 300:129–133, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p129.pdf.

Barnes:2005:BPI

[BK05c] D. K. A. Barnes and P. Kuklinski. Bipolar patterns of intraspecific competition in bryozoans. Marine Ecology Progress Series, 285:75–87, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/285/m285p075.pdf.

Botello:2006:DLR

[BK06a] G. Botello and P. J. Krug. ‘Desperate larvae’ revisited: age, energy and experience affect sensitivity to settlement cues in larvae of the gastropod Alderia sp. Marine Ecology Progress Series, 312:149–159, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/312/m312p149.pdf.

Bouley:2006:EHA

[BK06b] P. Bouley and W. J. Kimmerer. Ecology of a highly abundant, introduced cyclopoid copepod in a temperate estuary.

Marine Ecology Progress Series, 324:219–228, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p219.pdf.

Broitman:2006:SSB

[BK06c] B. R. Broitman and B. P. Kinlan. Spatial scales of benthic and pelagic producer biomass in a coastal upwelling ecosystem. Marine Ecology Progress Series, 327:15–25, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p015.pdf.

Browman:2004:QSP

[BKB^{+}04] Howard I. Browman, D. S. Kirby, H. H. Bauer, B. P. Dancik, G. Dresser, J. S. Gray, P. J. B. Hart, D. S. Kirby, J. C. Roff, and A. J. Underwood. Quality in science publishing. Marine Ecology Progress Series, 270:265–287, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m270p265.pdf.

Brown:2005:DHX

[BKB05] K. M. Brown, S. F. Keenan, and P. D. Banks. Dominance hierarchies in xanthid crabs: roles in resource holding potential and field distributions. Marine Ecology Progress Series, 291:189–196, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/291/m291p189.pdf.

Blakeslee:2009:DEP

[BKB^{+}09] A. M. H. Blakeslee, C. L. Keogh, J. E. Byers, A. M. Kuris, K. D. Lafferty, and M. E. Torchin. Differential escape from parasites by two competing introduced crabs. Marine Ecology Progress Series, 393:83–96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m393p083.pdf.

Booth:2000:RDO

[BKDB00] D. J. Booth, M. J. Kingsford, P. J. Doherty, and G. A. Beretta. Recruitment of damselfishes in One Tree Island lagoon: persistent interannual spatial patterns. Marine Ecology Progress Series, 202:219–230, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/202/m202p219.pdf.

Biuw:2009:TBS

[BKH^{+}09] M. Biuw, B. A. Krafft, G. J. G. Hofmeyr, C. Lydersen, and K. M. Kovacs. Time budgets and at-sea behaviour of lactating female Antarctic fur seals Arctocephalus gazella at Bouvetøya. Marine Ecology Progress Series, 385:271–284, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p271.pdf.

Bosire:2005:PPR

[BKK^{+}05] J. O. Bosire, J. G. Kairo, J. Kazungu, N. Koedam, and F. Dahdouh-Guebas. Predation on propagules regulates regeneration in a high-density reforested mangrove plantation. Marine Ecology Progress Series, 299:149–155, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p149.pdf.

Birkeland:2005:TFA

[BKLGN05] A. Birkeland, K. M. Kovacs, C. Lydersen, and O. Grahl-Nielsen. Transfer of fatty acids from mothers to their calves during lactation in white whales Delphinapterus leucas. Marine Ecology Progress Series, 298:287–294, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p287.pdf.

Barlow:2008:CBP

[BKM08] J. Barlow, M. Kahru, and B. G. Mitchell. Cetacean biomass, prey consumption, and primary production requirements in the California Current ecosystem. Marine Ecology Progress Series, 371:285–295, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/371/m371p285.pdf.

Bailey:2007:CCH

[BKP07] D. M. Bailey, N. J. King, and I. G. Priede. Cameras and carcasses: historical and current methods for using artificial food falls to study deep-water animals. Marine Ecology Progress Series, 350:179–191, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m350p179.pdf.

Breteler:2004:REL

[BKR04] W. C. M. Klein Breteler, M. Koski, and S. Rampen. Role of essential lipids in copepod nutrition: no evidence for trophic upgrading of food quality by a marine ciliate. Marine Ecology Progress Series, 274:199–208, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p199.pdf.

Bonaldo:2007:DDP

[BKSS07] R. M. Bonaldo, J. P. Krajewski, C. Sazima, and I. Sazima. Dentition damage in parrotfishes feeding on hard surfaces at Fernando de Noronha Archipelago, southwest Atlantic Ocean. Marine Ecology Progress Series, 342:249–254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p249.pdf.

Beiring:2000:EPC

[BL00a] E. A. Beiring and H. R. Lasker. Egg production by colonies of a gorgonian coral. Marine Ecology Progress Series, 196: 169–177, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/196/m196p169.pdf.

Bollens:2000:BRI

[BL00b] G. C. Rollwagen Bollens and M. R. Landry. Biological response to iron fertilization in the eastern equatorial Pacific (IronEx II). II. Mesozooplankton abundance, biomass, depth distribution and grazing. Marine Ecology Progress Series, 201: 43–56, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/201/m201p043.pdf.

Burford:2001:HAP

[BL01] M. A. Burford and A. R. Longmore. High ammonium production from sediments in hypereutrophic shrimp ponds. Marine Ecology Progress Series, 224:187–195, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p187.pdf.

Behrens:2004:EMR

[BL04] M. D. Behrens and K. D. Lafferty. Effects of marine reserves and urchin disease on southern Californian rocky reef communities. Marine Ecology Progress Series, 279:129–139, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/279/m279p129.pdf. See erratum [Ano04c].

Bradbury:2007:DNH

[BL07] I. R. Bradbury and B. J. Laurel. Defining ‘natal homing’ in marine fish populations: comment on Svedäng et al. (2007). Marine Ecology Progress Series, 349:307–308, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/

m349p307.pdf. See [SRJ07b].

Butler:2009:HBI

[BL09] M. J. Butler IV and J. A. Lear. Habitat-based intraguild predation by Caribbean reef octopus Octopus briareus on juvenile Caribbean spiny lobster Panulirus argus. Marine Ecology Progress Series, 386:115–122, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/386/m386p115.pdf.

Black:2009:CDS

[Bla09] B. A. Black. Climate-driven synchrony across tree, bivalve, and rockfish growth-increment chronologies of the northeast Pacific. Marine Ecology Progress Series, 378:37–46, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/378/m378p037.pdf.

Bailleul:2005:DFS

[BLD^{+}05] F. Bailleul, S. Luque, L. Dubroca, J. P. Y. Arnould, and C. Guinet. Differences in foraging strategy and maternal behaviour between two sympatric fur seal species at the Crozet Islands. Marine Ecology Progress Series, 293:273– 282, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p273.pdf.

Bonadonna:2001:FGF

[BLDG01] F. Bonadonna, M. A. Lea, O. Dehorter, and C. Guinet. Foraging ground fidelity and route-choice tactics of a marine predator: the Antarctic fur seal Arctocephalus gazella. Marine Ecology Progress Series, 223:287–297, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/223/m223p287.pdf.

Beseres:2006:PEL

[BLF06] J. J. Beseres, A. L. Lawrence, and R. J. Feller. Practical equivalence of laboratory and field measurements of gut passage time in two penaeid shrimp species. Marine Ecology Progress Series, 309:221–231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/309/m309p221.pdf.

Baduini:2001:DBC

[BLH01] C. L. Baduini, J. R. Lovvorn, and G. L. Hunt, Jr. Determining the body condition of short-tailed shearwaters: implications for migratory flight ranges and starvation events. Marine Ecology Progress Series, 222:265–277, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/222/m222p265.pdf.

Buskey:2002:EBP

[BLH02] E. J. Buskey, P. H. Lenz, and D. K. Hartline. Escape behavior of planktonic copepods in response to hydrodynamic disturbances: high speed video analysis. Marine Ecology Progress Series, 235:135–146, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/235/m235p135.pdf.

Blanchard:2004:FED

[BLHB04] F. Blanchard, F. LeLoc’h, C. Hily, and J. Boucher. Fishing effects on diversity, size and community structure of the benthic invertebrate and fish megafauna on the Bay of Biscay coast of France. Marine Ecology Progress Series, 280: 249–260, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/280/m280p249.pdf.

Buzzelli:2002:ESE

[BLP^{+}02] C. P. Buzzelli, R. A. Luettich, Jr., S. P. Powers, C. H. Peterson, J. E. McNinch, J. L. Pinckney, and H. W. Paerl. Estimating the spatial extent of bottom-water hypoxia and habitat degradation in a shallow estuary. Marine Ecology Progress Series, 230:103–112, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p103.pdf.

Barshaw:2003:OVD

[BLS03] D. E. Barshaw, K. L. Lavalli, and E. Spanier. Offense versus defense: responses of three morphological types of lobsters to predation. Marine Ecology Progress Series, 256:171– 182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/256/m256p171.pdf.

Belgrano:2004:NLD

[BLS04] A. Belgrano, M. Lima, and N. C. Stenseth. Non-linear dynamics in marine-phytoplankton population systems. Marine Ecology Progress Series, 273:281–289, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/273/
m273p281.pdf.

Bargu:2006:EDD

[BLS06] S. Bargu, K. Lefebvre, and M. W. Silver. Effect of dissolved domoic acid on the grazing rate of krill Euphausia pacifica. Marine Ecology Progress Series, 312:169–175, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p169.pdf.

Botton:2003:DAS

[BLT03] M. L. Botton, R. E. Loveland, and A. Tiwari. Distribution, abundance, and survivorship of young-of-the-year in a commercially exploited population of horseshoe crabs Limulus polyphemus. Marine Ecology Progress Series, 265:175– 184, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p175.pdf.

Begg:2000:SOP

[BM00] G. A. Begg and G. Marteinsdottir. Spawning origins of pelagic juvenile cod Gadus morhua inferred from spatially explicit age distributions: potential influences on year-class strength and recruitment. Marine Ecology Progress Series, 202:193–217, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p193.pdf.

Baird:2002:MGR

[BM02a] A. H. Baird and P. A. Marshall. Mortality, growth and reproduction in scleractinian corals following bleaching on the Great Barrier Reef. Marine Ecology Progress Series, 237: 133–141, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p133.pdf.

Begg:2002:ESEa

[BM02b] G. A. Begg and G. Marteinsdottir. Environmental and stock effects on spatial distribution and abundance of mature cod Gadus morhua. Marine Ecology Progress Series, 229:245– 262, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/229/m229p245.pdf.

Begg:2002:ESEb

[BM02c] G. A. Begg and G. Marteinsdottir. Environmental and stock effects on spawning origins and recruitment of cod Gadus morhua. Marine Ecology Progress Series, 229:263– 277, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/229/m229p263.pdf.

Buckel:2002:CBJ

[BM02d] J. A. Buckel and K. A. McKown. Competition between juvenile striped bass and bluefish: resource partitioning and growth rate. Marine Ecology Progress Series, 234:191–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p191.pdf.

Baumgartner:2003:SFE

[BM03] M. F. Baumgartner and B. R. Mate. Summertime foraging ecology of North Atlantic right whales. Marine Ecology Progress Series, 264:123–135, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p123.pdf.

Binzer:2005:TCS

[BM05a] T. Binzer and A. L. Middelboe. From thallus to communities: scale effects and photosynthetic performance in macroalgae communities. Marine Ecology Progress Series, 287:65– 75, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/287/m287p065.pdf.

Brockerhoff:2005:MBF

[BM05b] A. M. Brockerhoff and C. L. McLay. Mating behaviour, female receptivity and male–male competition in the intertidal crab Hemigrapsus sexdentatus (Brachyura: Grapsidae). Marine Ecology Progress Series, 290:179–191, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p179.pdf.

Bonin:2006:NBM

[BM06] P. C. Bonin and V. D. Michotey. Nitrogen budget in a microbial mat in the Camargue (southern France). Marine Ecology Progress Series, 322:75–84, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/322/m322p075.pdf.

Box:2007:EMC

[BM07a] S. J. Box and P. J. Mumby. Effect of macroalgal competition on growth and survival of juvenile Caribbean corals. Marine Ecology Progress Series, 342:139–149, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p139.pdf.

Bricelj:2007:EBT

[BM07b] V. M. Bricelj and S. P. MacQuarrie. Effects of brown tide (Aureococcus anophagefferens) on hard clam Mercenaria mercenaria larvae and implications for benthic recruitment. Marine Ecology Progress Series, 331:147–159, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p147.pdf.

Bryan:2007:PSH

[BM07c] T. L. Bryan and A. Metaxas. Predicting suitable habitat for deep-water gorgonian corals on the Atlantic and Pacific continental margins of North America. Marine Ecology Progress Series, 330:113–126, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/330/m330p113.pdf. See com-

ment [EM07b].

Batten:2009:SDA

[BM09] S. D. Batten and D. L. Mackas. Shortened duration of the annual Neocalanus plumchrus biomass peak in the Northeast Pacific. Marine Ecology Progress Series, 393:189–198, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m393p189.pdf.

Barroso:2002:IFS

[BMB02] C. M. Barroso, M. H. Moreira, and M. J. Bebianno. Imposex, female sterility and organotin contamination of the prosobranch Nassarius reticulatus from the Portuguese coast. Marine Ecology Progress Series, 230:127–135, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p127.pdf.

Bishop:2009:FMA

[BMC^{+}09] M. J. Bishop, T. Morgan, M. A. Coleman, B. P. Kelaher, L. K. Hardstaff, and R. W. Evenden. Facilitation of molluscan assemblages in mangroves by the fucalean alga Hormosira banksii. Marine Ecology Progress Series, 392:111– 122, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p111.pdf.

Buresch:2009:TDM

[BMCH09] K. C. Buresch, M. R. Maxwell, M. R. Cox, and R. T. Hanlon. Temporal dynamics of mating and paternity in the squid Loligo pealeii. Marine Ecology Progress Series, 387: 197–203, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/387/m387p197.pdf.

Baumgartner:2009:FPH

[BMdNP09] F. A. Baumgartner, C. A. Motti, R. de Nys, and N. A. Paul. Feeding preferences and host associations of specialist marine herbivores align with quantitative variation in seaweed secondary metabolites. Marine Ecology Progress Series, 396:1– 12, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Borja:2006:LTR

[BMF06] Á. Borja, I. Muxika, and J. Franco. Long-term recovery of soft-bottom benthos following urban and industrial sewage treatment in the Nervión estuary (southern Bay of Biscay). Marine Ecology Progress Series, 313:43–55, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p043.pdf.

Barroso:2000:CII

[BMG00] C. M. Barroso, M. H. Moreira, and P. E. Gibbs. Comparison of imposex and intersex development in four prosobranch species for TBT monitoring of a southern European estuarine system (oRia de Aveiro, NW Portugal). Marine Ecology Progress Series, 201:221–232, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/201/m201p221.pdf.

Blanchette:2002:GVF

[BMG02] C. A. Blanchette, B. G. Miner, and S. D. Gaines. Geographic variability in form, size and survival of Egregia menziesii around Point Conception, California. Marine Ecology Progress Series, 239:69–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/239/m239p069.pdf.

Burrows:2002:SSP

[BMJ02] M. T. Burrows, J. J. Moore, and B. James. Spatial synchrony of population changes in rocky shore communities in Shetland. Marine Ecology Progress Series, 240:39–48, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p039.pdf.

Baker:2007:RBW

[BMK07] D. M. Baker, S. E. MacAvoy, and K. Kim. Relationship between water quality, δ ^{1}5N, and aspergillosis of Caribbean sea fan corals. Marine Ecology Progress Series, 343:123– 130, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p123.pdf.

Birch:2008:CCT

[BML^{+}08] G. F. Birch, S. McCready, E. R. Long, S. S. Taylor, and G. Spyrakis. Contaminant chemistry and toxicity of sediments in Sydney Harbour, Australia: spatial extent and chemistry-toxicity relationships. Marine Ecology Progress Series, 363:71–87, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/363/m363p071.pdf.

Bustamante:2004:TEB

[BMM^{+}04] P. Bustamante, C. F. Morales, B. Mikkelsen, M. Dam, and F. Caurant. Trace element bioaccumulation in grey seals Halichoerus grypus from the Faroe Islands. Marine Ecology Progress Series, 267:291–301, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/267/m267p291.pdf.

Bonin:2009:CDF

[BMM^{+}09] M. C. Bonin, P. L. Munday, M. I. McCormick, M. Srinivasan, and G. P. Jones. Coral-dwelling fishes resistant to bleaching but not to mortality of host corals. Marine Ecology Progress Series, 394:215–222, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/394/m394p215.pdf.

Bouillon:2004:RUP

[BMO^{+}04] S. Bouillon, T. Moens, I. Overmeer, N. Koedam, and F. Dehairs. Resource utilization patterns of epifauna from mangrove forests with contrasting inputs of local versus imported organic matter. Marine Ecology Progress Series, 278:77– 88, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/278/m278p077.pdf.

Bricelj:2004:CDE

[BMS04] V. M. Bricelj, S. P. MacQuarrie, and R. Smolowitz. Concentration-dependent effects of toxic and non-toxic isolates of the brown tide alga Aureococcus anophagefferens on growth of juvenile bivalves. Marine Ecology Progress Series, 282:101–114, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p101.pdf.

Bouillon:2000:SSO

[BMSD00] S. Bouillon, P. C. Mohan, N. Sreenivas, and F. Dehairs. Sources of suspended organic matter and selective feeding by zooplankton in an estuarine mangrove ecosystem as traced by stable isotopes. Marine Ecology Progress Series, 208: 79–92, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/208/m208p079.pdf.

Behrenfeld:2002:PNB

[BMSH02] M. J. Behrenfeld, E. Marañón, D. A. Siegel, and S. B. Hooker. Photoacclimation and nutrient-based model of light-saturated photosynthesis for quantifying oceanic primary production. Marine Ecology Progress Series, 228:103–117, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p103.pdf.

Brickman:2007:FAE

[BMT07] D. Brickman, G. Marteinsdottir, and L. Taylor. Formulation and application of an efficient optimized biophysical model. Marine Ecology Progress Series, 347:275–284, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p275.pdf.

Bartol:2002:DES

[BMV02] I. K. Bartol, R. Mann, and M. Vecchione. Distribution of the euryhaline squid Lolliguncula brevis in Chesapeake Bay: effects of selected abiotic factors. Marine Ecology Progress Series, 226:235–247, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/226/m226p235.pdf.

Boddy:2000:IPS

[BMW^{+}00] L. Boddy, C. W. Morris, M. F. Wilkins, L. Al-Haddad, G. A. Tarran, R. R. Jonker, and P. H. Burkill. Identification of 72 phytoplankton species by radial basis function neural network analysis of flow cytometric data. Marine Ecology Progress Series, 195:47–59, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/195/m195p047.pdf.

Birrell:2008:CEM

[BMWH08] C. L. Birrell, L. J. McCook, B. L. Willis, and L. Harrington. Chemical effects of macroalgae on larval settlement of the broadcast spawning coral Acropora millepora. Marine Ecology Progress Series, 362:129–137, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p129.pdf.

Bintz:2001:REZ

[BN01] J. C. Bintz and S. W. Nixon. Responses of eelgrass Zostera marina seedlings to reduced light. Marine Ecology Progress Series, 223:133–141, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/223/m223p133.pdf.

Brush:2002:DML

[BN02] M. J. Brush and S. W. Nixon. Direct measurements of light attenuation by epiphytes on eelgrass Zostera marina. Marine Ecology Progress Series, 238:73–79, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p073.pdf.

Bilio:2004:CJR

[BN04] M. Bilio and U. Niermann. Is the comb jelly really to blame for it all? Mnemiopsis leidyi and the ecological concerns about the Caspian Sea. Marine Ecology Progress Series, 269:173–183, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/269/m269p173.pdf.

Baylis:2009:MFA

[BN09] A. M. M. Baylis and P. D. Nichols. Milk fatty acids predict the foraging locations of the New Zealand fur seal: continental shelf versus oceanic waters. Marine Ecology Progress Series, 380:271–286, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p271.pdf.

Beaumont:2002:UMF

[BND02] K. L. Beaumont, G. V. Nash, and A. T. Davidson. Ultrastructure, morphology and flux of microzooplankton faecal pellets in an east Antarctic fjord. Marine Ecology Progress Series, 245:133–148, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p133.pdf.

Brock:2007:CIB

[BNP07] E. Brock, G. M. Nylund, and H. Pavia. Chemical inhibition of barnacle larval settlement by the brown alga Fucus vesiculosus. Marine Ecology Progress Series, 337:165– 174, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p165.pdf.

Brugel:2009:PBP

[BNP^{+}09] S. Brugel, C. Nozais, M. Poulin, J. E. Tremblay, L. A. Miller, K. G. Simpson, Y. Gratton, and S. Demers. Phytoplankton biomass and production in the southeastern Beaufort Sea in autumn 2002 and 2003. Marine Ecology Progress Series, 377:63–77, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/377/m377p063.pdf.

Broitman:2001:GVS

[BNSG01] B. R. Broitman, S. A. Navarrete, F. Smith, and S. D. Gaines. Geographic variation of southeastern Pacific intertidal communities. Marine Ecology Progress Series, 224:21–34, 2001.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p021.pdf.

Booth:2000:DJS

[BO00] J. D. Booth and J. R. Ovenden. Distribution of Jasus spp. (Decapoda: Palinuridae) phyllosomas in southern waters: implications for larval recruitment. Marine Ecology Progress Series, 200:241–255, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p241.pdf.

Brun:2008:IVZ

[BOM^{+}08] F. G. Brun, I. Olivé, E. Malta, J. J. Vergara, I. Hernández, and J. L. Pérez-Lloréns. Increased vulnerability of Zostera noltii to stress caused by low light and elevated ammonium levels under phosphate deficiency. Marine Ecology Progress Series, 365:67–75, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p067.pdf.

Booth:2002:DCA

[Boo02] D. J. Booth. Distribution changes after settlement in 6 species of damselfish (Pomacentridae) in One Tree Island lagoon, Great Barrier Reef. Marine Ecology Progress Series, 226:157–164, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p157.pdf.

Branch:2008:LTM

[BOR08] G. M. Branch, F. Odendaal, and T. B. Robinson. Long-term monitoring of the arrival, expansion and effects of the alien mussel Mytilus galloprovincialis relative to wave action. Marine Ecology Progress Series, 370:171–183, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/370/
m370p171.pdf.

Bowden:2005:SRA

[Bow05] D. A. Bowden. Seasonality of recruitment in Antarctic sessile marine benthos. Marine Ecology Progress Series, 297: 101–118, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/297/m297p101.pdf.

Boyd:2008:IS

[Boy08] P. W. Boyd. Introduction and synthesis. Marine Ecology Progress Series, 364:213–218, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m364p213.pdf.

Broekhuizen:2003:SGS

[BOZ03] N. Broekhuizen, J. Oldman, and J. Zeldis. Sub-grid-scale differences between individuals influence simulated phytoplankton production and biomass in a shelf-sea system. Marine Ecology Progress Series, 252:61–76, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p061.pdf.

Brockington:2001:SRA

[BP01] S. Brockington and L. S. Peck. Seasonality of respiration and ammonium excretion in the Antarctic echinoid Sterechinus neumayeri. Marine Ecology Progress Series, 219:159– 168, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/219/m219p159.pdf.

Bailey:2002:LDO

[BP02a] K. M. Bailey and S. J. Picquelle. Larval distribution of offshore spawning flatfish in the Gulf of Alaska: potential transport pathways and enhanced onshore transport during ENSO events. Marine Ecology Progress Series, 236:205– 217, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/236/m236p205.pdf.

Belliveau:2002:EHN

[BP02b] S. A. Belliveau and V. J. Paul. Effects of herbivory and nutrients on the early colonization of crustose coralline and fleshy algae. Marine Ecology Progress Series, 232:105–114, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/232/m232p105.pdf.

Bollens:2003:FDA

[BP03] G. C. Rollwagen Bollens and D. L. Penry. Feeding dynamics of Acartia spp. copepods in a large, temperate estuary (San Francisco Bay, CA). Marine Ecology Progress Series, 257:139–158, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p139.pdf.

Becerro:2004:EDL

[BP04] M. A. Becerro and V. J. Paul. Effects of depth and light on secondary metabolites and cyanobacterial symbionts of the sponge Dysidea granulosa. Marine Ecology Progress Series, 280:115–128, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p115.pdf.

Byers:2006:GAF

[BP06] J. E. Byers and J. M. Pringle. Going against the flow: retention, range limits and invasions in advective environments. Marine Ecology Progress Series, 313:27–41, 2006. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m313p027.pdf.

Beger:2008:EFI

[BP08] M. Beger and H. P. Possingham. Environmental factors that influence the distribution of coral reef fishes: modeling occurrence data for broad-scale conservation and management. Marine Ecology Progress Series, 361:1–13, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Blythe:2009:HSS

[BP09] J. N. Blythe and J. Pineda. Habitat selection at settlement endures in recruitment time series. Marine Ecology Progress Series, 396:77–84, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/396/m396p077.pdf.

Brodeur:2008:ADC

[BPA^{+}08] R. D. Brodeur, W. T. Peterson, T. D. Auth, H. L. Soulen, M. M. Parnel, and A. A. Emerson. Abundance and diversity of coastal fish larvae as indicators of recent changes in ocean and climate conditions in the Oregon upwelling zone. Marine Ecology Progress Series, 366:187–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/366/
m366p187.pdf.

Becker:2007:OCP

[BPB07] B. H. Becker, M. Z. Peery, and S. R. Beissinger. Ocean climate and prey availability affect the trophic level and reproductive success of the marbled murrelet, an endangered seabird. Marine Ecology Progress Series, 329:267–279, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p267.pdf.

Bacheler:2009:ABF

[BPBH09] N. M. Bacheler, L. M. Paramore, J. A. Buckel, and J. E. Hightower. Abiotic and biotic factors influence the habitat use of an estuarine fish. Marine Ecology Progress Series, 377:263–277, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/377/m377p263.pdf.

Bargu:2002:KPV

[BPC^{+}02] S. Bargu, C. L. Powell, S. L. Coale, M. Busman, G. J. Doucette, and M. W. Silver. Krill: a potential vector for domoic acid in marine food webs. Marine Ecology Progress Series, 237:209–216, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p209.pdf.

Bentlage:2009:IDC

[BPC09] B. Bentlage, A. T. Peterson, and P. Cartwright. Inferring distributions of chirodropid box-jellyfishes (Cnidaria: Cubozoa) in geographic and ecological space using ecological niche modeling. Marine Ecology Progress Series, 384: 121–133, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p121.pdf.

Bos:2006:RVM

[BPCvdM06] O. G. Bos, C. J. M. Philippart, G. C. Cadée, and J. van der Meer. Recruitment variation in Macoma balthica: a laboratory examination of the match/mismatch hypothesis. Marine Ecology Progress Series, 320:207–214, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p207.pdf.

Beauvais:2006:ETT

[BPE^{+}06] S. Beauvais, M. L. Pedrotti, J. Egge, K. Iversen, and C. Marrasé. Effects of turbulence on TEP dynamics under contrasting nutrient conditions: implications for aggregation and sedimentation processes. Marine Ecology Progress Series, 323: 47–57, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m323p047.pdf.

Baylis:2008:CSF

[BPG08] A. M. M. Baylis, B. Page, and S. D. Goldsworthy. Colonyspecific foraging areas of lactating New Zealand fur seals. Marine Ecology Progress Series, 361:279–290, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p279.pdf.

Baker:2007:EVO

[BPH07] J. D. Baker, J. J. Polovina, and E. A. Howell. Effect of variable oceanic productivity on the survival of an upper trophic predator, the Hawaiian monk seal Monachus schauinslandi. Marine Ecology Progress Series, 346:277–283, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p277.pdf.

Bouman:2000:BOPa

[BPK^{+}00] H. A. Bouman, T. Platt, G. W. Kraay, S. Sathyendranath, and B. D. Irwin. Bio-optical properties of the subtropical North Atlantic. I. Vertical variability. Marine Ecology Progress Series, 200:3–18, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p003.pdf.

Brown-Peterson:2005:MIH

[BPLD^{+}05] N. J. Brown-Peterson, P. Larkin, N. Denslow, C. King, S. Manning, and M. Brouwer. Molecular indicators of hypoxia in the blue crab Callinectes sapidus. Marine Ecology Progress Series, 286:203–215, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/286/m286p203.pdf.

Berumen:2005:WRD

[BPM05] M. L. Berumen, M. S. Pratchett, and M. I. McCormick. Within-reef differences in diet and body condition of coralfeeding butterflyfishes (Chaetodontidae). Marine Ecology Progress Series, 287:217–227, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/287/m287p217.pdf.

Bearhop:2006:SII

[BPM^{+}06] S. Bearhop, R. A. Phillips, R. McGill, Y. Cherel, D. A. Dawson, and J. P. Croxall. Stable isotopes indicate sex-specific and long-term individual foraging specialisation in diving seabirds. Marine Ecology Progress Series, 311:157–164, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/311/
m311p157.pdf.

Brigolin:2009:MIA

[BPN^{+}09] D. Brigolin, R. Pastres, T. D. Nickell, C. J. Cromey, D. R. Aguilera, and P. Regnier. Modelling the impact of aquaculture on early diagenetic processes in sea loch sediments. Marine Ecology Progress Series, 388:63–80, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m388p063.pdf.

Bouman:2000:BOPb

[BPS^{+}00] H. A. Bouman, T. Platt, S. Sathyendranath, B. D. Irwin, M. R. Wernand, and G. W. Kraay. Bio-optical properties of the subtropical North Atlantic. II. Relevance to models of primary production. Marine Ecology Progress Series, 200: 19–34, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/200/m200p019.pdf.

Bouman:2003:TIO

[BPS^{+}03] H. A. Bouman, T. Platt, S. Sathyendranath, W. K. W. Li, V. Stuart, C. Fuentes-Yaco, H. Maass, E. P. W. Horne, O. Ulloa, V. Lutz, and M. Kyewalyanga. Temperature as indicator of optical properties and community structure of marine phytoplankton: implications for remote sensing. Marine Ecology Progress Series, 258:19–30, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/258/m258p019.pdf.

Bearhop:2000:VMC

[BPT^{+}00] S. Bearhop, R. A. Phillips, D. R. Thompson, S. Waldron, and R. W. Furness. Variability in mercury concentrations of great skuas Catharacta skua: the influence of colony, diet and trophic status inferred from stable isotope signatures. Marine Ecology Progress Series, 195:261–268, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p261.pdf.

Brockington:2007:GGM

[BPT07] S. Brockington, L. S. Peck, and P. A. Tyler. Gametogenesis and gonad mass cycles in the common circumpolar Antarctic echinoid Sterechinus neumayeri. Marine Ecology Progress Series, 330:139–147, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p139.pdf.

Bos:2007:ETF

[BPvdM07] O. G. Bos, C. J. M. Philippart, and J. van der Meer. Effects of temporary food limitation on development and mortality of Macoma balthica larvae. Marine Ecology Progress Series, 330:155–162, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p155.pdf.

Beauvais:2003:TEP

[BPVL03] S. Beauvais, M. L. Pedrotti, E. Villa, and R. Lemée. Transparent exopolymer particle (TEP) dynamics in relation to trophic and hydrological conditions in the NW Mediterranean Sea. Marine Ecology Progress Series, 262:97–109, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/262/
m262p097.pdf.

Barr:2003:NSM

[BR03] N. G. Barr and T. A. V. Rees. Nitrogen status and metabolism in the green seaweed Enteromorpha intestinalis: an examination of three natural populations. Marine Ecology Progress Series, 249:133–144, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p133.pdf.

Baker:2007:MSP

[BR07] A. C. Baker and A. M. Romanski. Multiple symbiotic partnerships are common in scleractinian corals, but not in octocorals: Comment on Goulet (2006). Marine Ecology Progress Series, 335:237–242, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m335p237.pdf. See [Gou06].

Bilkovic:2008:ECD

[BR08] D. M. Bilkovic and M. M. Roggero. Effects of coastal development on nearshore estuarine nekton communities. Marine Ecology Progress Series, 358:27–39, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p027.pdf.

Boersma:2009:FDA

[BR09] P. D. Boersma and G. A. Rebstock. Foraging distance affects reproductive success in Magellanic penguins. Marine Ecology Progress Series, 375:263–275, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/375/m375p263.pdf.

Brazeiro:2001:RBS

[Bra01] A. Brazeiro. Relationship between species richness and morphodynamics in sandy beaches: what are the underlying factors? Marine Ecology Progress Series, 224:35–44, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p035.pdf.

Bradshaw:2005:SFT

[Bra05] C. J. A. Bradshaw. Survival of the fittest technology-problems estimating marine turtle mortality. Marine Ecology Progress Series, 287:261–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/287/m287p261.pdf.

Braccini:2008:FET

[Bra08] J. M. Braccini. Feeding ecology of two high-order predators from south-eastern Australia: the coastal broadnose and the deepwater sharpnose sevengill sharks. Marine Ecology Progress Series, 371:273–284, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/371/m371p273.pdf.

Billingham:2003:ARM

[BRAS03] M. R. Billingham, T. B. H. Reusch, F. Alberto, and E. A. Serrão. Is asexual reproduction more important at geographical limits? A genetic study of the seagrass Zostera marina in the Ria Formosa, Portugal. Marine Ecology Progress Series, 265:77–83, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p077.pdf.

Browman:2000:IUR

[BRB^{+}00] Howard I. Browman, C. A. Rodriguez, F. Béland, J. J. Cullen, R. F. Davis, J. H. M. Kouwenberg, P. S. Kuhn, B. McArthur, J. A. Runge, J. F. St-Pierre, and R. D. Vetter. Impact of ultraviolet radiation on marine crustacean zooplankton and ichthyoplankton: a synthesis of results from the estuary and Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 199:293–311, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p293.pdf.

Bischof:2006:PAG

[BRBPL06] K. Bischof, R. Rautenberger, L. Brey, and J. L. Pérez-Lloréns. Physiological acclimation to gradients of solar irradiance within mats of the filamentous green macroalga Chaetomorpha linum from southern Spain. Marine Ecology Progress Series, 306:165–175, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/306/m306p165.pdf.

Bremner:2003:AFD

[BRF03] J. Bremner, S. I. Rogers, and C. L. J. Frid. Assessing functional diversity in marine benthic ecosystems: a comparison of approaches. Marine Ecology Progress Series, 254: 11–25, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p011.pdf.

Bjorkstedt:2002:DLR

[BRG^{+}02] E. P. Bjorkstedt, L. K. Rosenfeld, B. A. Grantham, Y. Shkedy, and J. Roughgarden. Distributions of larval rockfishes Sebastes spp. across nearshore fronts in a coastal upwelling region. Marine Ecology Progress Series, 242:215–228, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/242/m242p215.pdf.

Baez:2007:LTC

[BRGS^{+}07] J. C. Báez, R. Real, C. García-Soto, J. M. de la Serna, D. Macías, and J. A. Camiñas. Loggerhead turtle by-catch depends on distance to the coast, independent of fishing effort: implications for conservation and fisheries management. Marine Ecology Progress Series, 338:249–256, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/338/
m338p249.pdf.

Brickman:2003:CDB

[Bri03] D. Brickman. Controls on the distribution of Browns Bank juvenile haddock. Marine Ecology Progress Series, 263:235– 246, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/263/m263p235.pdf.

Beck:2007:BFA

[BRI^{+}07] C. A. Beck, L. D. Rea, S. J. Iverson, J. M. Kennish, K. W. Pitcher, and B. S. Fadely. Blubber fatty acid profiles reveal regional, seasonal, age-class and sex differences in the diet of young Steller sea lions in Alaska. Marine Ecology Progress Series, 338:269–280, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/338/m338p269.pdf.

Beaugrand:2000:BNA

[BRIP00] G. Beaugrand, P. C. Reid, F. Ibañez, and B. Planque. Biodiversity of North Atlantic and North Sea calanoid copepods. Marine Ecology Progress Series, 204:299–303, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p299.pdf.

Berg:2003:OUA

[BRJ^{+}03] P. Berg, H. Røy, F. Janssen, V. Meyer, B. B. Jørgensen, M. Huettel, and D. de Beer. Oxygen uptake by aquatic sediments measured with a novel non-invasive eddy-correlation technique. Marine Ecology Progress Series, 261:75–83, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/261/m261p075.pdf.

Brown:2006:UCC

[Bro06] J. A. Brown. Using the chemical composition of otoliths to evaluate the nursery role of estuaries for English sole Pleuronectes vetulus populations. Marine Ecology Progress Series, 306:269–281, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p269.pdf.

Berkenbusch:2000:TSV

[BRP00] K. Berkenbusch, A. A. Rowden, and P. K. Probert. Temporal and spatial variation in macrofauna community composition imposed by ghost shrimp Callianassa filholi bioturbation. Marine Ecology Progress Series, 192:249–257, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p249.pdf.

Bi:2007:MPH

[BRP07] H. Bi, R. E. Ruppel, and W. T. Peterson. Modeling the pelagic habitat of salmon off the Pacific Northwest (USA) coast using logistic regression. Marine Ecology Progress Series, 336:249–265, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p249.pdf.

Blicher:2007:GPS

[BRS07] M. E. Blicher, S. Rysgaard, and M. K. Sejr. Growth and production of sea urchin Strongylocentrotus droebachiensis in a high-Arctic fjord, and growth along a climatic gradient (64 to 77^{o}N). Marine Ecology Progress Series, 341:89–102, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/341/m341p089.pdf. See erratum [Ano07c].

Boersma:2007:OAC

[BRSM07] P. D. Boersma, G. A. Rebstock, D. L. Stokes, and P. Majluf. Oceans apart: conservation models for two temperate penguin species shaped by the marine environment. Marine Ecology Progress Series, 335:217–225, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/335/m335p217.pdf.

Balch:2000:TSV

[BS00] T. Balch and R. E. Scheibling. Temporal and spatial variability in settlement and recruitment of echinoderms in kelp beds and barrens in Nova Scotia. Marine Ecology Progress Series, 205:139–154, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p139.pdf.

Baco:2003:HSR

[BS03] A. R. Baco and C. R. Smith. High species richness in deep-sea chemoautotrophic whale skeleton communities. Marine Ecology Progress Series, 260:109–114, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p109.pdf.

Britton-Simmons:2004:DIE

[BS04a] K. H. Britton-Simmons. Direct and indirect effects of the introduced alga Sargassum muticum on benthic, subtidal communities of Washington State, USA. Marine Ecology Progress Series, 277:61–78, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/277/m277p061.pdf.

Browman:2004:PEB

[BS04b] Howard I. Browman and Konstantinos I. Stergiou. Perspectives on ecosystem-based approaches to the management of marine resources. Marine Ecology Progress Series, 274:269– 303, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/274/m274p269.pdf.

Buckley:2004:RDR

[BS04c] Bradley A. Buckley and Alina M. Szmant. RNA/dnA ratios as indicators of metabolic activity in four species of Caribbean reef-building corals. Marine Ecology Progress Series, 282:143–149, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p143.pdf.

Baker:2005:RPA

[BS05a] R. Baker and M. Sheaves. Redefining the piscivore assemblage of shallow estuarine nursery habitats. Marine Ecology Progress Series, 291:197–213, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/291/m291p197.pdf.

Browman:2005:TSP

[BS05b] Howard I. Browman and Konstantinos I. Stergiou. Theme section: Politics and socio-economics of ecosystem-based management of marine resources. Marine Ecology Progress Series, 300:241–296, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m300p241.pdf.

Baker:2006:VSR

[BS06] R. Baker and M. Sheaves. Visual surveys reveal high densities of large piscivores in shallow estuarine nurseries. Marine Ecology Progress Series, 323:75–82, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p075.pdf.

Baker:2007:SWR

[BS07a] R. Baker and M. Sheaves. Shallow-water refuge paradigm: conflicting evidence from tethering experiments in a tropical estuary. Marine Ecology Progress Series, 349:13–22, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/349/m349p013.pdf.

Barbee:2007:CNS

[BS07b] N. C. Barbee and S. E. Swearer. Characterizing natal source population signatures in the diadromous fish Galaxias maculatus, using embryonic otolith chemistry. Marine Ecology Progress Series, 343:273–282, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/343/m343p273.pdf.

Bracken:2007:TMB

[BS07c] M. E. S. Bracken and J. J. Stachowicz. Top-down modification of bottom-up processes: selective grazing reduces macroalgal nitrogen uptake. Marine Ecology Progress Series, 330:75–82, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p075.pdf.

Baldridge:2008:TCP

[BS08a] A. K. Baldridge and L. D. Smith. Temperature constraints on phenotypic plasticity explain biogeographic patterns in predator trophic morphology. Marine Ecology Progress Series, 365:25–34, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p025.pdf.

Bartholomew:2008:SSR

[BS08b] A. Bartholomew and R. L. Shine. Space size relative to prey width (Sp/Py) influences macrofaunal colonization of artificial structures. Marine Ecology Progress Series, 358: 95–102, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/358/m358p095.pdf.

Berline:2008:ETW

[BSA^{+}08] L. Berline, Y. H. Spitz, C. J. Ashjian, R. G. Campbell, W. Maslowski, and S. E. Moore. Euphausiid transport in the Western Arctic Ocean. Marine Ecology Progress Series, 360:163–178, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p163.pdf.

Bongiorni:2003:SGG

[BSAR03] L. Bongiorni, S. Shafir, D. Angel, and B. Rinkevich. Survival, growth and gonad development of two hermatypic corals subjected to in situ fish-farm nutrient enrichment. Marine Ecology Progress Series, 253:137–144, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p137.pdf.

Beet:2003:CVP

[BSB03] A. Beet, A. R. Solow, and S. M. Bollens. Comparing vertical plankton profiles with replication. Marine Ecology Progress Series, 262:285–287, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/262/m262p285.pdf.

Bachelet:2004:IEB

[BSBD^{+}04] G. Bachelet, B. Simon-Bouhet, C. Desclaux, P. Garcia-Meunier, G. Mairesse, X. de Montaudouin, H. Raigné, K. Randriambao, P. G. Sauriau, and F. Viard. Invasion of the eastern Bay of Biscay by the nassariid gastropod Cyclope neritea: origin and effects on resident fauna. Marine Ecology Progress Series, 276:147–159, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p147.pdf.

Ben-Shlomo:2001:HDC

[BSDR01] R. Ben-Shlomo, J. Douek, and B. Rinkevich. Heterozygote deficiency and chimerism in remote populations of a colonial ascidian from New Zealand. Marine Ecology Progress Series, 209:109–117, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/209/m209p109.pdf.

Britton-Simmons:2008:ETC

[BSED08] K. Britton-Simmons, J. E. Eckman, and D. O. Duggins. Effect of tidal currents and tidal stage on estimates of bed size in the kelp Nereocystis luetkeana. Marine Ecology Progress Series, 355:95–105, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/355/m355p095.pdf.

Brodeur:2002:IJB

[BSH02] R. D. Brodeur, H. Sugisaki, and G. L. Hunt, Jr. Increases in jellyfish biomass in the Bering Sea: implications for the ecosystem. Marine Ecology Progress Series, 233:89– 103, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p089.pdf.

Brewin:2009:EOR

[BSH^{+}09] P. E. Brewin, K. I. Stocks, D. B. Haidvogel, C. Condit, and A. Gupta. Effects of oceanographic retention on decapod and gastropod community diversity on seamounts. Marine Ecology Progress Series, 383:225–237, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m383p225.pdf.

Budge:2007:FAB

[BSIS07] S. M. Budge, A. M. Springer, S. J. Iverson, and G. Sheffield. Fatty acid biomarkers reveal niche separation in an Arctic benthic food web. Marine Ecology Progress Series, 336:305– 309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/336/m336p305.pdf.

Binzer:2002:ISD

[BSJ02] T. Binzer and K. Sand-Jensen. Importance of structure and density of macroalgae communities (Fucus serratus) for photosynthetic production and light utilisation. Marine Ecology Progress Series, 235:53–62, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/235/m235p053.pdf.

Benzie:2006:MVA

[BSK06] J. A. H. Benzie and C. Smith-Keune. Microsatellite variation in Australian and Indonesian pearl oyster Pinctada maxima populations. Marine Ecology Progress Series, 314:197– 211, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p197.pdf.

Blachowiak-Samolyk:2006:AZD

[BSKR^{+}06] K. Blachowiak-Samolyk, S. Kwasniewski, K. Richardson, K. Dmoch, E. Hansen, H. Hop, S. Falk-Petersen, and L. T.

Mouritsen. Arctic zooplankton do not perform diel vertical migration (DVM) during periods of midnight sun. Marine Ecology Progress Series, 308:101–116, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p101.pdf.

Boller:2002:SMM

[BSL02] M. L. Boller, T. D. Swain, and H. R. Lasker. Skeletal morphology and material properties of a fragmenting gorgonian coral. Marine Ecology Progress Series, 228:131–141, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p131.pdf.

Boyd:2002:DFF

[BSM02] I. L. Boyd, I. J. Staniland, and A. R. Martin. Distribution of foraging by female Antarctic fur seals. Marine Ecology Progress Series, 242:285–294, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/242/m242p285.pdf.

Bradbury:2003:PAB

[BSP03] I. R. Bradbury, P. V. R. Snelgrove, and P. Pepin. Passive and active behavioural contributions to patchiness and spatial pattern during the early life history of marine fishes. Marine Ecology Progress Series, 257:233–245, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/257/
m257p233.pdf.

Bandelj:2008:AMP

[BSP^{+}08] Vinko Bandelj, Giorgio Socal, Young-Seuk Park, Sovan Lek, Joan Coppola, Elisa Camatti, Elisa Capuzzo, Luisella Milani, and Cosimo Solidoro. Analysis of multitrophic plankton assemblages in the Lagoon of Venice. Marine Ecology Progress Series, 368:23–40, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m368p023.pdf.

Browman:2003:DDR

[BSPK03] Howard I. Browman, J. F. St-Pierre, and P. Kuhn. Dose and dose-rate dependency in the mortality response of Calanus finmarchicus embryos exposed to ultraviolet radiation. Marine Ecology Progress Series, 247:297–302, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p297.pdf.

Breteler:2005:EDN

[BSR05] W. C. M. Klein Breteler, N. Schogt, and S. Rampen. Effect of diatom nutrient limitation on copepod development: the role of essential lipids. Marine Ecology Progress Series, 291:125–133, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/291/m291p125.pdf.

Blicher:2009:HCD

[BSR09] M. E. Blicher, M. K. Sejr, and S. Rysgaard. High carbon demand of dominant macrozoobenthic species indicates their central role in ecosystem carbon flow in a sub-Arctic fjord. Marine Ecology Progress Series, 383:127–140, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p127.pdf.

Bottger-Schnack:2005:PSF

[BSS05] R. Böttger-Schnack and D. Schnack. Population structure and fecundity of the microcopepod Oncaea bispinosa in the Red Sea — a challenge to general concepts for the scaling of fecundity. Marine Ecology Progress Series, 302:159– 175, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p159.pdf.

Bouma:2002:RIM

[BSvdH^{+}02] T. J. Bouma, J. Stapel, J. van der Heiden, B. Koutstaal, J. van Soelen, and L. van IJzerloo. Relative importance of macrophyte leaves for nitrogen uptake from flood water in tidal salt marshes. Marine Ecology Progress Series, 240: 93–104, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/240/m240p093.pdf.

Boehnlein:2005:DBD

[BSVK05] J. M. Boehnlein, L. Z. Santiago-Vázquez, and R. G. Kerr. Diterpene biosynthesis by the dinoflagellate symbiont of the Caribbean gorgonian Pseudopterogorgia bipinnata. Marine Ecology Progress Series, 303:105–111, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p105.pdf.

Beukers-Stewart:2005:BCA

[BSVM^{+}05] B. D. Beukers-Stewart, B. J. Vause, M. W. J. Mosley, H. L. Rossetti, and A. R. Brand. Benefits of closed area protection for a population of scallops. Marine Ecology Progress Series, 298:189–204, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p189.pdf.

Bejder:2009:IAR

[BSW^{+}09] L. Bejder, A. Samuels, H. Whitehead, H. Finn, and S. Allen. Impact assessment research: use and misuse of habituation, sensitisation and tolerance in describing wildlife responses to anthropogenic stimuli. Marine Ecology Progress Series, 395:177–185, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p177.pdf.

Belmaker:2005:ESS

[BSZ05] J. Belmaker, N. Shashar, and Y. Ziv. Effects of small-scale isolation and predation on fish diversity on experimental reefs. Marine Ecology Progress Series, 289:273–283, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p273.pdf.

Burton:2000:ERA

[BT00] R. S. Burton and M. J. Tegner. Enhancement of red abalone Haliotis rufescens stocks at San Miguel Island: reassessing a success story. Marine Ecology Progress Series, 202:303– 308, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/202/m202p303.pdf.

Baker:2001:FSC

[BT01] M. C. Baker and P. A. Tyler. Fertilization success in the commercial gastropod Haliotis tuberculata. Marine Ecology Progress Series, 211:205–213, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/211/m211p205.pdf.

Burrows:2004:EDD

[BT04] M. T. Burrows and G. Tarling. Effects of density dependence on diel vertical migration of populations of northern krill: a genetic algorithm model. Marine Ecology Progress Series, 277:209–220, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/277/m277p209.pdf.

Basedow:2006:CFC

[BT06] S. L. Basedow and K. S. Tande. Cannibalism by female Calanus finmarchicus on naupliar stages. Marine Ecology Progress Series, 327:247–255, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/327/m327p247.pdf.

Bailey:2009:UMM

[BT09] H. Bailey and P. M. Thompson. Using marine mammal habitat modelling to identify priority conservation zones within a marine protected area. Marine Ecology Progress Series, 378:279–287, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p279.pdf.

Bowen:2002:PDF

[BTB^{+}02] W. D. Bowen, D. Tully, D. J. Boness, B. M. Bulheier, and G. J. Marshall. Prey-dependent foraging tactics and prey profitability in a marine mammal. Marine Ecology Progress Series, 244:235–245, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p235.pdf.

Borme:2009:DEE

[BTB^{+}09] D. Borme, V. Tirelli, S. B. Brandt, S. Fonda Umani, and E. Arneri. Diet of Engraulis encrasicolus in the northern Adriatic Sea (Mediterranean): ontogenetic changes and feeding selectivity. Marine Ecology Progress Series, 392:193– 209, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p193.pdf.

Beucher:2004:PDB

[BTC^{+}04] C. Beucher, P. Tréguer, R. Corvaisier, A. M. Hapette, and M. Elskens. Production and dissolution of biosilica, and changing microphytoplankton dominance in the Bay of Brest (France). Marine Ecology Progress Series, 267:57– 69, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/267/m267p057.pdf.

Bustamante:2004:UTD

[BTD^{+}04] P. Bustamante, J. L. Teyssié, B. Danis, S. W. Fowler, P. Miramand, O. Cotret, and M. Warnau. Uptake, transfer and distribution of silver and cobalt in tissues of the common cuttlefish Sepia officinalis at different stages of its life cycle. Marine Ecology Progress Series, 269:185–195, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p185.pdf.

Bustamante:2002:BZC

[BTF^{+}02] P. Bustamante, J. L. Teyssié, S. W. Fowler, O. Cotret, B. Danis, P. Miramand, and M. Warnau. Biokinetics of zinc and cadmium accumulation and depuration at different stages in the life cycle of the cuttlefish Sepia officinalis. Marine Ecology Progress Series, 231:167–177, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/231/m231p167.pdf.

Bussotti:2006:STV

[BTF^{+}06] S. Bussotti, A. Terlizzi, S. Fraschetti, G. Belmonte, and F. Boero. Spatial and temporal variability of sessile benthos in shallow Mediterranean marine caves. Marine Ecology Progress Series, 325:109–119, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p109.pdf.

BenRaisLasram:2008:ECD

[BTG^{+}08] F. Ben Rais Lasram, J. A. Tomasini, F. Guilhaumon, M. S. Romdhane, T. Do Chi, and D. Mouillot. Ecological correlates of dispersal success of Lessepsian fishes. Marine Ecology Progress Series, 363:273–286, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/363/m363p273.pdf.

Ben-Tzvi:2008:TRP

[BTKG^{+}08] O. Ben-Tzvi, M. Kiflawi, S. D. Gaines, M. Al-Zibdah, M. S. Sheehy, G. L. Paradis, and A. Abelson. Tracking recruitment pathways of Chromis viridis in the Gulf of Aqaba using otolith chemistry. Marine Ecology Progress Series, 359: 229–238, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/359/m359p229.pdf.

Bates:2005:RTC

[BTL05] A. E. Bates, V. Tunnicliffe, and R. W. Lee. Role of thermal conditions in habitat selection by hydrothermal vent gastropods. Marine Ecology Progress Series, 305:1–15, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Boyce:2008:ETG

[BTW08] D. G. Boyce, D. P. Tittensor, and B. Worm. Effects of temperature on global patterns of tuna and billfish richness. Marine Ecology Progress Series, 355:267–276, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p267.pdf.

Bishop:2002:SEI

[BUA02] M. J. Bishop, A. J. Underwood, and P. Archambault. Sewage and environmental impacts on rocky shores: necessity of identifying relevant spatial scales. Marine Ecology Progress Series, 236:121–128, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p121.pdf.

Blanquer:2009:SSS

[BUCC09] A. Blanquer, M. J. Uriz, and J. Caujapé-Castells. Smallscale spatial genetic structure in Scopalina lophyropoda, an encrusting sponge with philopatric larval dispersal and frequent fission and fusion events. Marine Ecology Progress Series, 380:95–102, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p095.pdf.

Bergquist:2002:EPS

[BUF02] D. C. Bergquist, I. A. Urcuyo, and C. R. Fisher. Establishment and persistence of seep vestimentiferan aggregations on the upper Louisiana slope of the Gulf of Mexico. Marine Ecology Progress Series, 241:89–98, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/241/m241p089.pdf.

Bulleri:2005:RRC

[Bul05] F. Bulleri. Role of recruitment in causing differences between intertidal assemblages on seawalls and rocky shores. Marine Ecology Progress Series, 287:53–65, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p053.pdf.

Bulleri:2006:DOA

[Bul06] F. Bulleri. Duration of overgrowth affects survival of encrusting coralline algae. Marine Ecology Progress Series, 321: 79–85, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p079.pdf.

Burgess:2001:IPS

[Bur01] R. Burgess. An improved protocol for separating meiofauna from sediments using colloidal silica sols. Marine Ecology Progress Series, 214:161–165, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p161.pdf.

Burnett:2002:LVA

[Bur02] W. J. Burnett. Longitudinal variation in algal symbionts (zooxanthellae) from the Indian Ocean zoanthid Palythoa caesia. Marine Ecology Progress Series, 234:105–109, 2002.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p105.pdf.

Burnaford:2008:SRP

[BV08] J. L. Burnaford and M. Vasquez. Solar radiation plays a role in habitat selection by the sea star Pisaster ochraceus. Marine Ecology Progress Series, 368:177–187, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p177.pdf.

Barnes:2002:LPD

[BVC^{+}02] D. K. A. Barnes, E. Verling, A. Crook, I. Davidson, and M. O’Mahoney. Local population disappearance follows (20 yr after) cycle collapse in a pivotal ecological species. Marine Ecology Progress Series, 226:311–313, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p311.pdf.

Bode:2001:DPO

[BVCG01] A. Bode, M. Varela, M. Canle, and N. González. Dissolved and particulate organic nitrogen in shelf waters of northern Spain during spring. Marine Ecology Progress Series, 214: 43–54, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p043.pdf.

Berges:2002:ETG

[BVH02] J. A. Berges, D. E. Varela, and P. J. Harrison. Effects of temperature on growth rate, cell composition and nitrogen metabolism in the marine diatom Thalassiosira pseudonana (Bacillariophyceae). Marine Ecology Progress Series, 225:139–146, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p139.pdf.

Browman:2008:EOA

[BVHG08a] Howard I. Browman, Alain F. Vézina, and Ova Hoegh-Guldberg. Effects of ocean acidification on marine ecosystems. Marine Ecology Progress Series, ??:??, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Browman:2008:TSE

[BVHG08b] Howard I. Browman, Alain F. Vézina, and Ove Hoegh-Guldberg. Theme section: Effects of ocean acidification on marine ecosystems. Marine Ecology Progress Series, 373: 199–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m373_ThemeSection.pdf.

Botto:2005:IBC

[BVI^{+}05] F. Botto, I. Valiela, O. Iribarne, P. Martinetto, and J. Alberti. Impact of burrowing crabs on C and N sources, control, and transformations in sediments and food webs of SW Atlantic estuaries. Marine Ecology Progress Series, 293:155– 164, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p155.pdf.

Bos:2007:PDH

[BvK07] A. R. Bos and M. M. van Katwijk. Planting density, hydrodynamic exposure and mussel beds affect survival of transplanted intertidal eelgrass. Marine Ecology Progress Series, 336:121–129, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p121.pdf.

Bertocci:2007:CTV

[BVMBC07] I. Bertocci, S. Vaselli, E. Maggi, and L. Benedetti-Cecchi. Changes in temporal variance of rocky shore organism abundances in response to manipulation of mean intensity and temporal variability of aerial exposure. Marine Ecology Progress Series, 338:11–20, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p011.pdf.

Brun:2003:ESU

[BVN^{+}03] F. G. Brun, J. J. Vergara, G. Navarro, I. Hernández, and J. L. Pérez-Lloréns. Effect of shading by Ulva rigida canopies on growth and carbon balance of the seagrass Zostera noltii. Marine Ecology Progress Series, 265:85–96, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p085.pdf.

Brun:2006:CBS

[BVP^{+}06] F. G. Brun, J. J. Vergara, G. Peralta, M. P. García-Sánchez, I. Hernández, and J. L. Pérez-Lloréns. Clonal building, simple growth rules and phylloclimate as key steps to develop functional-structural seagrass models. Marine Ecology Progress Series, 323:133–148, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/323/m323p133.pdf.

Beyst:2002:TDP

[BVVM02] B. Beyst, J. Vanaverbeke, M. Vincx, and J. Mees. Tidal and diurnal periodicity in macrocrustaceans and demersal fish of an exposed sandy beach, with special emphasis on juvenile plaice Pleuronectes platessa. Marine Ecology Progress Series, 225:263–274, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p263.pdf.

Bracher:2000:SEN

[BW00] A. U. Bracher and C. Wiencke. Simulation of the effects of naturally enhanced UV radiation on photosynthesis of Antarctic phytoplankton. Marine Ecology Progress Series, 196:127–141, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p127.pdf.

Begg:2001:SIO

[BW01] G. A. Begg and C. R. Weidman. Stable δ ^{13}C and δ ^{1}8O isotopes in otoliths of haddock Melanogrammus aeglefinus from the northwest Atlantic Ocean. Marine Ecology Progress Series, 216:223–233, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/216/m216p223.pdf.

Blackmore:2004:RBM

[BW04] G. Blackmore and W. X. Wang. Relationships between metallothioneins and metal accumulation in the whelk Thais clavigera. Marine Ecology Progress Series, 277:135–145, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/277/m277p135.pdf.

Billerbeck:2006:NRE

[BWB^{+}06] M. Billerbeck, U. Werner, K. Bosselmann, E. Walpersdorf, and M. Huettel. Nutrient release from an exposed intertidal sand flat. Marine Ecology Progress Series, 316:35– 51, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/316/m316p035.pdf.

Bosley:2002:ETR

[BWCW02] K. L. Bosley, D. A. Witting, R. C. Chambers, and S. C. Wainright. Estimating turnover rates of carbon and nitrogen in recently metamorphosed winter flounder Pseudopleuronectes americanus with stable isotopes. Marine Ecology Progress Series, 236:233–240, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p233.pdf.

Brannock:2009:EHM

[BWH09] P. M. Brannock, D. S. Wethey, and T. J. Hilbish. Extensive hybridization with minimal introgression in Mytilus galloprovincialis and M. trossulus in Hokkaido, Japan. Marine Ecology Progress Series, 383:161–171, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p161.pdf.

Breitbarth:2008:NFG

[BWK^{+}08] E. Breitbarth, J. Wohlers, J. Kläs, J. LaRoche, and I. Peeken. Nitrogen fixation and growth rates of Trichodesmium IMS-101 as a function of light intensity. Marine Ecology Progress Series, 359:25–36, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p025.pdf.

Barnes:2006:SGA

[BWL06] D. K. A. Barnes, K. Webb, and K. Linse. Slow growth of Antarctic bryozoans increases over 20 years and is anomalously high in 2003. Marine Ecology Progress Series, 314: 187–195, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p187.pdf.

Bergmann:2002:UID

[BWMA02] M. Bergmann, S. K. Wieczorek, P. G. Moore, and R. J. A. Atkinson. Utilisation of invertebrates discarded from the Nephrops fishery by variously selective benthic scavengers in the west of Scotland. Marine Ecology Progress Series, 233:185–198, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p185.pdf.

Bullard:2004:CLZ

[BWO04] S. G. Bullard, R. B. Whitlatch, and R. W. Osman. Checking the landing zone: do invertebrate larvae avoid settling near superior spatial competitors? Marine Ecology Progress Series, 280:239–247, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p239.pdf.

Billerbeck:2006:SDP

[BWP^{+}06] M. Billerbeck, U. Werner, L. Polerecky, E. Walpersdorf, D. de-Beer, and M. Huettel. Surficial and deep pore water circulation governs spatial and temporal scales of nutrient recycling in intertidal sand flat sediment. Marine Ecology Progress Series, 326:61–76, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p061.pdf.

Barneah:2004:DDS

[BWPB04] O. Barneah, V. M. Weis, S. Perez, and Y. Benayahu. Diversity of dinoflagellate symbionts in Red Sea soft corals: mode of symbiont acquisition matters. Marine Ecology Progress Series, 275:89–95, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p089.pdf.

Bulling:2006:UMS

[BWRP06] M. T. Bulling, P. C. L. White, D. Raffaelli, and G. J. Pierce. Using model systems to address the biodiversity-ecosystem functioning process. Marine Ecology Progress Series, 311: 295–309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m311p295.pdf.

Boschker:2000:LCM

[BWSH00] H. T. S. Boschker, A. Wielemaker, B. E. M. Schaub, and M. Holmer. Limited coupling of macrophyte production and bacterial carbon cycling in the sediments of Zostera spp. meadows. Marine Ecology Progress Series, 203:181–189, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p181.pdf.

Buma:2006:PAU

[BWvdE^{+}06] A. G. J. Buma, S. W. Wright, R. van den Enden, W. H. van de Poll, and A. T. Davidson. PAR acclimation and UVBR-induced DNA damage in Antarctic marine microalgae. Marine Ecology Progress Series, 315:33–42, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p033.pdf.

Barnes:2004:PPB

[BWW^{+}04] D. K. A. Barnes, N. L. Warren, K. Webb, B. Phalan, and K. Reid. Polar pedunculate barnacles piggy-back on pycnogona, penguins, pinniped seals and plastics. Marine Ecology Progress Series, 284:305–310, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/284/m284p305.pdf.

Bradford:2006:SEW

[BWW^{+}06] A. L. Bradford, P. R. Wade, D. W. Weller, A. M. Burdin, Y. V. Ivashchenko, G. A. Tsidulko, G. R. VanBlaricom, and R. L. Brownell, Jr. Survival estimates of western gray whales Eschrichtius robustus incorporating individual heterogeneity and temporary emigration. Marine Ecology Progress Series, 315:293–307, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p293.pdf.

Brooke:2009:LSM

[BY09] S. Brooke and C. M. Young. In situ measurement of survival and growth of Lophelia pertusa in the northern Gulf of Mexico. Marine Ecology Progress Series, 397:153–161, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p153.pdf.

Byers:2000:DSH

[Bye00] J. E. Byers. Differential susceptibility to hypoxia aids estuarine invasion. Marine Ecology Progress Series, 203:123– 132, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/203/m203p123.pdf.

Brock:2006:NFR

[BYH^{+}06] J. C. Brock, K. K. Yates, R. B. Halley, I. B. Kuffner, C. W. Wright, and B. G. Hatcher. Northern Florida reef tract benthic metabolism scaled by remote sensing. Marine Ecology Progress Series, 312:123–139, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/312/m312p123.pdf.

Bost:2002:FDP

[BZLD02] C. A. Bost, T. Zorn, Y. Le Maho, and G. Duhamel. Feeding of diving predators and diel vertical migration of prey: King penguin diet versus trawl sampling at Kerguelen Islands. Marine Ecology Progress Series, 227:51–61, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p051.pdf.

Cardinale:2000:RBS

[CA00] M. Cardinale and F. Arrhenius. The relationship between stock and recruitment: are the assumptions valid? Marine Ecology Progress Series, 196:305–309, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p305.pdf.

Carballo:2004:PDM

[CÁ04] J. L. Carballo and E. Ávila. Population dynamics of a mutualistic interaction between the sponge Haliclona caerulea and the red alga Jania adherens. Marine Ecology Progress Series, 279:93–104, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p093.pdf.

Castellani:2006:FCT

[CA06] C. Castellani and Y. Altunbas. Factors controlling the temporal dynamics of egg production in the copepod Temora longicornis. Marine Ecology Progress Series, 308:143–153, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/308/m308p143.pdf.

Cunha:2000:PEH

[CAA00] M. A. Cunha, M. A. Almeida, and F. Alcântara. Patterns of ectoenzymatic and heterotrophic bacterial activities along a salinity gradient in a shallow tidal estuary. Marine Ecology Progress Series, 204:1–12, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/204/m204p001.pdf.

Cusson:2007:BBA

[CAA07] M. Cusson, P. Archambault, and A. Aitken. Biodiversity of benthic assemblages on the Arctic continental shelf: historical data from Canada. Marine Ecology Progress Series, 331:291–304, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/331/m331p291.pdf.

Conde:2002:MIC

[CABS02] D. Conde, L. Aubriot, S. Bonilla, and R. Sommaruga. Marine intrusions in a coastal lagoon enhance the negative effect of solar UV radiation on phytoplankton photosynthetic rates. Marine Ecology Progress Series, 240:57–70, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p057.pdf.

Cieluch:2007:OIK

[CACDC07] U. Cieluch, K. Anger, M. Charmantier-Daures, and G. Charmantier. Osmoregulation and immunolocalization of Na^{+} / K^{+}-ATPase during the ontogeny of the mitten crab Eriocheir sinensis (Decapoda, Grapsoidea). Marine Ecology Progress Series, 329:169–178, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/329/m329p169.pdf.

Canellas:2000:LVP

[CAD00] M. Cañellas, S. Agustí, and C. M. Duarte. Latitudinal variability in phosphate uptake in the Central Atlantic. Marine Ecology Progress Series, 194:283–294, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p283.pdf.

Casale:2008:FEL

[CAF^{+}08] P. Casale, G. Abbate, D. Freggi, N. Conte, M. Oliverio, and R. Argano. Foraging ecology of loggerhead sea turtles Caretta caretta in the central Mediterranean Sea: evidence for a relaxed life history model. Marine Ecology Progress Series, 372:265–276, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/372/m372p265.pdf.

Cairns:2009:IFT

[Cai09] S. D. Cairns. Influence of Frederick (Ted) M. Bayer on deep-water octocoral research. Marine Ecology Progress Series, 397:7–10, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p007.pdf.

Cervin:2005:SSD

[CÅJ05] G. Cervin, P. Åberg, and S. R. Jenkins. Small-scale disturbance in a stable canopy dominated community: implications for macroalgal recruitment and growth. Marine Ecology Progress Series, 305:31–40, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p031.pdf.

Caldarone:2005:EGH

[Cal05] E. M. Caldarone. Estimating growth in haddock larvae Melanogrammus aeglefinus from RNA: DNA ratios and water temperature. Marine Ecology Progress Series, 293:241– 252, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p241.pdf.

Callaway:2006:TWP

[Cal06] R. Callaway. Tube worms promote community change. Marine Ecology Progress Series, 308:49–60, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p049.pdf.

Campbell:2001:ARF

[Cam01] S. Campbell. Ammonium requirements of fast-growing ephemeral macroalgae in a nutrient-enriched marine embayment (Port Phillip Bay, Australia). Marine Ecology Progress Series, 209:99–107, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/209/m209p099.pdf.

Chicharo:2007:ESR

[CAMC07] M. A. Chícharo, A. Amaral, P. Morais, and L. Chícharo. Effect of sex on ratios and concentrations of DNA and RNA in three marine species. Marine Ecology Progress Series, 332:241–245, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/332/m332p241.pdf.

Carballo:2000:LEA

[Car00] J. L. Carballo. Larval ecology of an ascidian tropical population in a Mediterranean enclosed ecosystem. Marine Ecology Progress Series, 195:159–167, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p159.pdf.

Carscadden:2005:DSS

[Car05] J. E. Carscadden. Did signals from seabirds indicate changes in capelin biology during the 1990s? Comment on Davoren & Montevecchi (2003). Marine Ecology Progress Series, 285: 289–297, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p289.pdf. See [DM03] and re-

ply [DM05a].

Conde:2000:CUP

[CAS00] D. Conde, L. Aubriot, and R. Sommaruga. Changes in UV penetration associated with marine intrusions and freshwater discharge in a shallow coastal lagoon of the Southern Atlantic Ocean. Marine Ecology Progress Series, 207:19–31, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p019.pdf.

Cabaco:2005:PLE

[CAS05] S. Cabaço, A. Alexandre, and R. Santos. Population-level effects of clam harvesting on the seagrass Zostera noltii. Marine Ecology Progress Series, 298:123–129, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p123.pdf.

Calliari:2006:SME

[CAT^{+}06] D. Calliari, C. M. Andersen, P. Thor, E. Gorokhova, and P. Tiselius. Salinity modulates the energy balance and reproductive success of co-occurring copepods Acartia tonsa and A. clausi in different ways. Marine Ecology Progress Series, 312:177–188, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/312/m312p177.pdf.

Cabrita:2000:SVD

[CB00a] M. T. Cabrita and V. Brotas. Seasonal variation in denitrification and dissolved nitrogen fluxes in intertidal sediments of the Tagus estuary, Portugal. Marine Ecology Progress Series, 202:51–65, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p051.pdf.

Crump:2000:CBA

[CB00b] B. C. Crump and J. A. Baross. Characterization of the bacterially-active particle fraction in the Columbia River estuary. Marine Ecology Progress Series, 206:13–22, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p013.pdf.

Charrassin:2001:UOH

[CB01] J. B. Charrassin and C. A. Bost. Utilisation of the oceanic habitat by king penguins over the annual cycle. Marine Ecology Progress Series, 221:285–298, 2001. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/221/m221p285.pdf.

Ciannelli:2005:LDR

[CB05a] L. Ciannelli and K. M. Bailey. Landscape dynamics and resulting species interactions: the cod-capelin system in the southeastern Bering Sea. Marine Ecology Progress Series, 291:227–236, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/291/m291p227.pdf.

Coleman:2005:STV

[CB05b] M. A. Coleman and S. H. Brawley. Spatial and temporal variability in dispersal and population genetic structure of a rockpool alga. Marine Ecology Progress Series, 300: 63–77, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/300/m300p063.pdf.

Cusson:2005:GPM

[CB05c] M. Cusson and E. Bourget. Global patterns of macroinvertebrate production in marine benthic habitats. Marine Ecology Progress Series, 297:1–14, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chiswell:2008:SSL

[CB08a] S. M. Chiswell and J. D. Booth. Sources and sinks of larval settlement in Jasus edwardsii around New Zealand: Where do larvae come from and where do they go? Marine Ecology Progress Series, 354:201–217, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/354/m354p201.pdf.

Cohen:2008:DGN

[CB08b] S. E. Cohen and S. M. Bollens. Diet and growth of non-native Mississippi silversides and yellowfin gobies in restored and natural wetlands in the San Francisco Estuary. Marine Ecology Progress Series, 368:241–254, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/368/m368p241.pdf.

Cullen:2008:PVI

[CB08c] J. J. Cullen and P. W. Boyd. Predicting and verifying the intended and unintended consequences of large-scale ocean iron fertilization. Marine Ecology Progress Series, 364:295– 301, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m364p295.pdf.

Christensen:2002:EPN

[CBA02] M. Christensen, G. T. Banta, and O. Andersen. Effects of the polychaetes Nereis diversicolor and Arenicola marina on the fate and distribution of pyrene in sediments. Marine Ecology Progress Series, 237:159–172, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/237/m237p159.pdf.

Casotti:2000:MFP

[CBAd00] R. Casotti, C. Brunet, B. Aronne, and M. R. d’Alcala. Mesoscale features of phytoplankton and planktonic bacteria in a coastal area as induced by external water masses. Marine Ecology Progress Series, 195:15–27, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/195/m195p015.pdf.

Campbell:2004:EOC

[CBD04] R. W. Campbell, P. Boutillier, and J. F. Dower. Ecophysiology of overwintering in the copepod Neocalanus plumchrus: changes in lipid and protein contents over a seasonal cycle. Marine Ecology Progress Series, 280:211–226, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p211.pdf.

Cook:2005:DGG

[CBD^{+}05] A. A. Cook, P. Bhadury, N. J. Debenham, B. H. M. Meldal, M. L. Blaxter, G. R. Smerdon, M. C. Austen, P. J. D. Lambshead, and A. D. Rogers. Denaturing gradient gel electrophoresis (DGGE) as a tool for identification of marine nematodes. Marine Ecology Progress Series, 291:103– 113, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p103.pdf.

Cherel:2002:FFE

[CBDH02] Y. Cherel, P. Bocher, C. De Broyer, and K. A. Hobson. Food and feeding ecology of the sympatric thin-billed Pachyptila belcheri and Antarctic P. desolata prions at Iles Kerguelen. Marine Ecology Progress Series, 228:263–281, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p263.pdf.

Cook:2004:CNCa

[CBE04] P. L. M. Cook, E. C. V. Butler, and B. D. Eyre. Carbon and nitrogen cycling on intertidal mudflats of a temperate Australian estuary. I. Benthic metabolism. Marine Ecology Progress Series, 280:25–38, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/280/m280p025.pdf.

Costello:2006:EMB

[CBEL06] M. J. Costello, P. Bouchet, C. S. Emblow, and A. Legakis. European marine biodiversity inventory and taxonomic resources: state of the art and gaps in knowledge. Marine Ecology Progress Series, 316:257–268, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m316p257.pdf.

Carballo:2008:BSM

[CBGLM08] J. L. Carballo, E. Bautista-Guerrero, and G. E. Leyte-Morales. Boring sponges and the modeling of coral reefs in the east Pacific Ocean. Marine Ecology Progress Series, 356:113–122, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p113.pdf.

Colloca:2009:IFN

[CBL^{+}09] F. Colloca, V. Bartolino, G. J. Lasinio, L. Maiorano, P. Sartor, and G. Ardizzone. Identifying fish nurseries using density and persistence measures. Marine Ecology Progress Series, 381:287–296, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p287.pdf.

Carlos:2000:DDS

[CBM00] A. A. Carlos, B. K. Baillie, and T. Maruyama. Diversity of dinoflagellate symbionts (zooxanthellae) in a host individual. Marine Ecology Progress Series, 195:93–100, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p093.pdf.

Caldow:2003:EIM

[CBM^{+}03] R. W. G. Caldow, H. A. Beadman, S. McGrorty, M. J. Kaiser, J. D. Goss-Custard, K. Mould, and A. Wilson. Effects of intertidal mussel cultivation on bird assemblages. Marine Ecology Progress Series, 259:173–183, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/259/m259p173.pdf.

Criales:2007:CST

[CBM^{+}07] M. M. Criales, J. A. Browder, C. N. K. Mooers, M. B. Robblee, H. Cardenas, and T. L. Jackson. Cross-shelf transport of pink shrimp larvae: interactions of tidal currents, larval vertical migrations and internal tides. Marine Ecology Progress Series, 345:167–184, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/345/m345p167.pdf.

Cognie:2003:SPL

[CBMB03] B. Cognie, L. Barillé, G. Massé, and P. G. Beninger. Selection and processing of large suspended algae in the oyster Crassostrea gigas. Marine Ecology Progress Series, 250: 145–152, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p145.pdf.

Caldwell:2004:FES

[CBO04] Gary S. Caldwell, Matthew G. Bentley, and Peter J. W. Olive. First evidence of sperm motility inhibition by the diatom aldehyde 2E, 4E-decadienal. Marine Ecology Progress Series, 273:97–108, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p097.pdf.

Carnevali:2001:PER

[CBP^{+}01] M. D. C. Carnevali, F. Bonasoro, M. Patruno, M. C. Thorndyke, and S. Galassi. PCB exposure and regeneration in crinoids (Echinodermata). Marine Ecology Progress Series, 215:155–167, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p155.pdf.

Cardoso:2005:RHU

[CBP^{+}05] P. G. Cardoso, A. Brandão, M. A. Pardal, D. Raffaelli, and J. C. Marques. Resilience of Hydrobia ulvae populations to anthropogenic and natural disturbances. Marine Ecology Progress Series, 289:191–199, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/289/m289p191.pdf.

Coll:2008:ITS

[CBS^{+}08] M. Coll, N. Bahamon, F. Sardà, I. Palomera, S. Tudela, and P. Suuronen. Improved trawl selectivity: effects on the ecosystem in the South Catalan Sea (NW Mediterranean). Marine Ecology Progress Series, 355:131–147, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p131.pdf.

Cherel:2002:DFE

[CBTW02] Y. Cherel, P. Bocher, C. Trouvé, and H. Weimerskirch. Diet and feeding ecology of blue petrels Halobaena caerulea at Iles Kerguelen, Southern Indian Ocean. Marine Ecology Progress Series, 228:283–299, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/228/m228p283.pdf.

Connor:2000:DAB

[CC00] R. Connor and G. L. Chmura. Dynamics of above- and belowground organic matter in a high latitude macrotidal saltmarsh. Marine Ecology Progress Series, 204:101–110, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p101.pdf.

Carrasson:2002:TRM

[CC02] M. Carrassón and J. E. Cartes. Trophic relationships in a Mediterranean deep-sea fish community: partition of food resources, dietary overlap and connections within the benthic boundary layer. Marine Ecology Progress Series, 241: 41–55, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/241/m241p041.pdf.

Cheney:2003:DED

[CC03a] K. L. Cheney and I. M. Côté. Do ectoparasites determine cleaner fish abundance? Evidence on two spatial scales. Marine Ecology Progress Series, 263:189–196, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p189.pdf.

Cheney:2003:ICP

[CC03b] K. L. Cheney and I. M. Côté. Indirect consequences of parental care: sex differences in ectoparasite burden and cleaner-seeking activity in longfin damselfish. Marine Ecology Progress Series, 262:267–275, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p267.pdf.

Caro:2004:PID

[CC04] A. U. Caro and J. C. Castilla. Predator-inducible defences and local intrapopulation variability of the intertidal mussel Semimytilus algosus in central Chile. Marine Ecology Progress Series, 276:115–123, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/276/m276p115.pdf.

Craig:2005:HIH

[CC05] J. K. Craig and L. B. Crowder. Hypoxia-induced habitat shifts and energetic consequences in Atlantic croaker and brown shrimp on the Gulf of Mexico shelf. Marine Ecology Progress Series, 294:79–94, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/294/m294p079.pdf.

Cole:2007:PDA

[CC07] V. J. Cole and M. G. Chapman. Patterns of distribution of annelids: taxonomic and spatial inconsistencies between two biogeographic provinces and across multiple spatial scales. Marine Ecology Progress Series, 346:235–241, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p235.pdf.

Cartwright:2006:ERE

[CCB06] S. R. Cartwright, R. A. Coleman, and M. A. Browne. Ecologically relevant effects of pulse application of copper on the limpet Patella vulgata. Marine Ecology Progress Series, 326:187–194, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p187.pdf.

Clavier:2009:AML

[CCB^{+}09a] J. Clavier, M. D. Castets, T. Bastian, C. Hily, G. Boucher, and L. Chauvaud. An amphibious mode of life in the intertidal zone: aerial and underwater contribution of Chthamalus montagui to CO_{2} fluxes. Marine Ecology Progress Series, 375:185–194, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p185.pdf.

Cupido:2009:ULT

[CCB^{+}09b] R. Cupido, S. Cocito, M. Barsanti, S. Sgorbini, A. Peirano, and G. Santangelo. Unexpected long-term population dynamics in a canopy-forming gorgonian coral following mass mortality. Marine Ecology Progress Series, 394:195–200, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/394/m394p195.pdf.

Crossman:2005:NEN

[CCC05] D. J. Crossman, J. H. Choat, and K. D. Clements. Nutritional ecology of nominally herbivorous fishes on coral reefs. Marine Ecology Progress Series, 296:129–142, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/296/
m296p129.pdf.

Cao:2009:ISO

[CCC09] J. Cao, X. Chen, and Y. Chen. Influence of surface oceanographic variability on abundance of the western winter-spring cohort of neon flying squid Ommastrephes bartramii in the NW Pacific Ocean. Marine Ecology Progress Series, 381: 119–127, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/381/m381p119.pdf.

Cullen:2009:POS

[CCCD09] J. M. Cullen, L. E. Chambers, P. C. Coutin, and P. Dann. Predicting onset and success of breeding in little penguins Eudyptula minor from ocean temperatures. Marine Ecology Progress Series, 378:269–278, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/378/m378p269.pdf.

Comeau:2001:TAC

[CCCH01] L. A. Comeau, S. E. Campana, G. A. Chouinard, and J. M. Hanson. Timing of Atlantic cod Gadus morhua seasonal migrations in relation to serum levels of gonadal and thyroidal hormones. Marine Ecology Progress Series, 221:245–253, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p245.pdf.

Coelho:2000:BMB

[CCdAR00] V. R. Coelho, R. A. Cooper, and S. de Almeida Rodrigues. Burrow morphology and behavior of the mud shrimp Upogebia omissa (Decapoda: Thalassinidea: Upogebiidae). Marine Ecology Progress Series, 200:229–240, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p229.pdf.

Chen:2009:PPE

[CCJT09] Y. L. L. Chen, H. Y. Chen, S. Jan, and S. H. Tuo. Phytoplankton productivity enhancement and assemblage change in the upstream Kuroshio after typhoons. Marine Ecology Progress Series, 385:111–126, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m385p111.pdf.

Colin:2003:LSS

[CCK03] S. P. Colin, J. H. Costello, and E. Klos. In situ swimming and feeding behavior of eight co-occurring hydromedusae. Marine Ecology Progress Series, 253:305–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p305.pdf.

Colin:2006:UFM

[CCK06] S. P. Colin, J. H. Costello, and H. Kordula. Upstream foraging by medusae. Marine Ecology Progress Series, 327:143– 155, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p143.pdf.

Castro:2009:ETV

[CCK^{+}09] L. R. Castro, G. Claramunt, M. C. Krautz, A. Llanos-Rivera, and P. Moreno. Egg trait variation in anchoveta Engraulis ringens: a maternal response to changing environmental conditions in contrasting spawning habitats. Marine Ecology Progress Series, 381:237–248, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/381/m381p237.pdf.

Chen:2003:DDF

[CCL03] Y. L. L. Chen, H. Y. Chen, and Y. H. Lin. Distribution and downward flux of Trichodesmium in the South China Sea as influenced by the transport from the Kuroshio Current. Marine Ecology Progress Series, 259:47–57, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/259/
m259p047.pdf.

Cerrato:2004:ENQ

[CCL^{+}04] R. M. Cerrato, D. A. Caron, D. J. Lonsdale, J. M. Rose, and R. A. Schaffner. Effect of the northern quahog Mercenaria mercenaria on the development of blooms of the brown tide alga Aureococcus anophagefferens. Marine Ecology Progress Series, 281:93–108, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p093.pdf.

Chen:2005:PDC

[CCL^{+}05] Q. Chen, K. S. Chan, K. Lekve, E. Torstensen, J. Gjøsæter, G. Ottersen, and N. C. Stenseth. Population dynamics of cod Gadus morhua in the North Sea region: biological densitydependent and climatic density-independent effects. Marine Ecology Progress Series, 302:219–232, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p219.pdf.

Crocker:2006:INN

[CCL^{+}06] D. E. Crocker, D. P. Costa, B. J. Le Boeuf, P. M. Webb, and D. S. Houser. Impact of El Niño on the foraging behavior of female northern elephant seals. Marine Ecology Progress Series, 309:1–10, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cheung:2008:AMT

[CCL^{+}08] W. W. L. Cheung, C. Close, V. Lam, R. Watson, and D. Pauly. Application of macroecological theory to predict effects of climate change on global fisheries potential. Marine Ecology Progress Series, 365:187–197, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/365/m365p187.pdf.

Carreto:2001:CSM

[CCM01] J. I. Carreto, M. O. Carignan, and N. G. Montoya. Comparative studies on mycosporine-like amino acids, paralytic shellfish toxins and pigment profiles of the toxic dinoflagellates Alexandrium tamarense, A. catenella and A. minutum. Marine Ecology Progress Series, 223:49–60, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p049.pdf.

Carstensen:2003:STR

[CCMK03] J. Carstensen, D. Conley, and B. Müller-Karulis. Spatial and temporal resolution of carbon fluxes in a shallow coastal ecosystem, the Kattegat. Marine Ecology Progress Series, 252: 35–50, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/252/m252p035.pdf.

Charrier:2007:GSW

[CCML07] G. Charrier, S. H. Coombs, I. H. McQuinn, and J. Laroche. Genetic structure of whiting Merlangius merlangus in the northeast Atlantic and adjacent waters. Marine Ecology Progress Series, 330:201–211, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/330/m330p201.pdf.

Carranza:2008:DMG

[CCR08] A. Carranza, R. K. Colwell, and T. F. L. V. B. Rangel. Distribution of megabenthic gastropods along environmental gradients: the mid-domain effect and beyond. Marine Ecology Progress Series, 367:193–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p193.pdf.

Clynick:2007:EEA

[CCU07] B. G. Clynick, M. G. Chapman, and A. J. Underwood. Effects of epibiota on assemblages of fish associated with urban structures. Marine Ecology Progress Series, 332:201– 210, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/332/m332p201.pdf.

Campbell:2003:RLM

[CD03a] R. W. Campbell and J. F. Dower. Role of lipids in the maintenance of neutral buoyancy by zooplankton. Marine Ecology Progress Series, 263:93–99, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/263/m263p093.pdf. See

erratum [Ano04d].

Cherel:2003:DSM

[CD03b] Y. Cherel and G. Duhamel. Diet of the squid Moroteuthis ingens (Teuthoidea: Onychoteuthidae) in the upper slope waters of the Kerguelen Islands. Marine Ecology Progress Series, 250:197–203, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p197.pdf.

Colin:2003:ETD

[CD03c] S. P. Colin and H. G. Dam. Effects of the toxic dinoflagellate Alexandrium fundyense on the copepod Acartia hudsonica: a test of the mechanisms that reduce ingestion rates. Marine Ecology Progress Series, 248:55–65, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/248/
m248p055.pdf.

Cloern:2005:PCE

[CD05] J. E. Cloern and R. Dufford. Phytoplankton community ecology: principles applied in San Francisco Bay. Marine Ecology Progress Series, 285:11–28, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/285/m285p011.pdf.

Catenazzi:2007:RSI

[CD07a] A. Catenazzi and M. A. Donnelly. Role of supratidal invertebrates in the decomposition of beach-cast green algae Ulva sp. Marine Ecology Progress Series, 349:33–42, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p033.pdf.

Creed:2007:SPD

[CD07b] J. C. Creed and A. F. De Paula. Substratum preference during recruitment of two invasive alien corals onto shallowsubtidal tropical rocky shores. Marine Ecology Progress Series, 330:101–111, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p101.pdf.

Catenazzi:2008:SLO

[CD08] A. Catenazzi and M. A. Donnelly. Sea lion Otaria flavescens as host of the common vampire bat Desmodus rotundus. Marine Ecology Progress Series, 360:285–289, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/360/
m360p285.pdf.

Choat:2006:ASG

[CDAM06] J. H. Choat, C. R. Davies, J. L. Ackerman, and B. D. Mapstone. Age structure and growth in a large teleost, Cheilinus undulatus, with a review of size distribution in labrid fishes. Marine Ecology Progress Series, 318:237–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p237.pdf.

Cherel:2008:SIR

[CDF^{+}08] Y. Cherel, S. Ducatez, C. Fontaine, P. Richard, and C. Guinet. Stable isotopes reveal the trophic position and mesopelagic fish diet of female southern elephant seals breeding on the Kerguelen Islands. Marine Ecology Progress Series, 370:239– 247, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p239.pdf.

Camacho:2003:EBG

[CDFRL03] A. Pérez Camacho, M. Delgado, M. J. Fernández-Reiriz, and U. Labarta. Energy balance, gonad development and biochemical composition in the clam Ruditapes decussatus. Marine Ecology Progress Series, 258:133–145, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p133.pdf.

Cherel:2004:CFS

[CDG04] Y. Cherel, G. Duhamel, and N. Gasco. Cephalopod fauna of subantarctic islands: new information from predators. Marine Ecology Progress Series, 266:143–156, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p143.pdf.

Chassot:2009:RPH

[CDH^{+}09] E. Chassot, D. Duplisea, M. Hammill, A. Caskenette, N. Bousquet, Y. Lambert, and G. Stenson. Role of predation by harp seals Pagophilus groenlandicus in the collapse and nonrecovery of northern Gulf of St. Lawrence cod Gadus morhua. Marine Ecology Progress Series, 379:279–297, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m379p279.pdf.

Christensen:2007:HBF

[CDJ^{+}07] A. Christensen, U. Daewel, H. Jensen, H. Mosegaard, M. St John, and C. Schrum. Hydrodynamic backtracking of fish larvae by individual-based modelling. Marine Ecology Progress Series, 347:221–232, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m347p221.pdf.

Compton:2007:OFM

[CDK^{+}07] T. J. Compton, J. Drent, R. Kentie, G. B. Pearson, J. van der Meer, and T. Piersma. Overlap in the feeding morphology of bivalves from species-rich and species-poor intertidal flats using gill: palp ratios for comparative analyses of mollusc assemblages. Marine Ecology Progress Series, 348:213– 220, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/348/m348p213.pdf.

Costa:2006:EFA

[CDML06] A. DeLorenzo Costa, E. G. Durbin, C. A. Mayo, and E. G. Lyman. Environmental factors affecting zooplankton in Cape Cod Bay: implications for right whale dynamics. Marine Ecology Progress Series, 323:281–298, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p281.pdf.

Camus:2008:POA

[CDO08] P. A. Camus, K. Daroch, and L. F. Opazo. Potential for omnivory and apparent intraguild predation in rocky intertidal herbivore assemblages from northern Chile. Marine Ecology Progress Series, 361:35–45, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/361/m361p035.pdf.

Coyer:2004:PGD

[CDS^{+}04] J. A. Coyer, O. E. Diekmann, E. A. Serrão, G. Procaccini, N. Milchakova, G. A. Pearson, W. T. Stam, and J. L. Olsen. Population genetics of dwarf eelgrass Zostera noltii throughout its biogeographic range. Marine Ecology Progress Series, 281:51–62, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p051.pdf.

Cappo:2007:IRV

[CDS07] M. Cappo, G. De’ath, and P. Speare. Inter-reef vertebrate communities of the Great Barrier Reef Marine Park determined by baited remote underwater video stations. Marine Ecology Progress Series, 350:209–221, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m350p209.pdf.

Coulter:2001:REN

[CDT^{+}01] S. C. Coulter, C. M. Duarte, M. S. Tuan, N. H. Tri, H. T. Ha, L. H. Giang, and P. N. Hong. Retrospective estimates of net leaf production in Kandelia candel mangrove forests. Marine Ecology Progress Series, 221:117–124, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p117.pdf.

Choe:2003:SVB

[CDT^{+}03] N. Choe, D. Deibel, R. J. Thompson, S. H. Lee, and V. K. Bushell. Seasonal variation in the biochemical composition of the chaetognath Parasagitta elegans from the hyperbenthic zone of Conception Bay, Newfoundland. Marine Ecology Progress Series, 251:191–200, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/251/m251p191.pdf.

Cox:2009:MEO

[CDW^{+}09] M. J. Cox, D. A. Demer, J. D. Warren, G. R. Cutter, and A. S. Brierley. Multibeam echosounder observations reveal interactions between Antarctic krill and air-breathing predators. Marine Ecology Progress Series, 378:199–209, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m378p199.pdf.

Caro:2008:PVB

[CEB^{+}08] A. U. Caro, J. Escobar, F. Bozinovic, S. A. Navarrete, and J. C. Castilla. Phenotypic variability in byssus thread production of intertidal mussels induced by predators with different feeding strategies. Marine Ecology Progress Series, 372:127–134, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/372/m372p127.pdf.

Craig:2007:HGC

[CEBR07] M. T. Craig, J. A. Eble, B. W. Bowen, and D. R. Robertson. High genetic connectivity across the Indian and Pacific Oceans in the reef fish Myripristis berndti (Holocentridae). Marine Ecology Progress Series, 334:245–254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p245.pdf.

Callaway:2007:CNS

[CED^{+}07] R. Callaway, G. H. Engelhard, J. Dann, J. Cotter, and H. Rumohr. A century of North Sea epibenthos and trawling: comparison between 1902–1912, 1982–1985 and 2000. Marine Ecology Progress Series, 346:27–43, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m346p027.pdf.

Clark:2004:ESM

[CEF04] R. P. Clark, M. S. Edwards, and M. S. Foster. Effects of shade from multiple kelp canopies on an understory algal assemblage. Marine Ecology Progress Series, 267:107– 119, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/267/m267p107.pdf.

Clarke:2006:ECL

[CEO06] J. Clarke, L. M. Emmerson, and P. Otahal. Environmental conditions and life history constraints determine foraging range in breeding Adélie penguins. Marine Ecology Progress Series, 310:247–261, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/310/m310p247.pdf.

Chicharo:2003:SLC

[CES^{+}03] M. A. Chícharo, E. Esteves, A. M. P. Santos, A. dos Santos, A. Peliz, and P. Ré. Are sardine larvae caught off northern Portugal in winter starving? An approach examining nutritional conditions. Marine Ecology Progress Series, 257:303–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p303.pdf.

Clark:2009:AMM

[CES^{+}09] C. W. Clark, W. T. Ellison, B. L. Southall, L. Hatch, S. M. Van Parijs, A. Frankel, and D. Ponirakis. Acoustic masking in marine ecosystems: intuitions, analysis, and implication. Marine Ecology Progress Series, 395:201–222, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p201.pdf.

Carman:2002:SSM

[CF02] K. R. Carman and B. Fry. Small-sample methods for δ ^{13}C and δ ^{15}N analysis of the diets of marsh meiofaunal species using natural-abundance and tracer-addition isotope techniques. Marine Ecology Progress Series, 240:85–92, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p085.pdf.

Cermeno:2008:SRC

[CF08] P. Cermeño and F. G. Figueiras. Species richness and cell-size distribution: size structure of phytoplankton communities. Marine Ecology Progress Series, 357:79–85, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p079.pdf.

Campbell:2009:IVE

[CF09] J. E. Campbell and J. W. Fourqurean. Interspecific variation in the elemental and stable isotope content of seagrasses in South Florida. Marine Ecology Progress Series, 387:109–123, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/387/m387p109.pdf.

Costantini:2007:FSG

[CFA07] F. Costantini, C. Fauvelot, and M. Abbiati. Fine-scale genetic structuring in Corallium rubrum: evidence of inbreeding and limited effective larval dispersal. Marine Ecology Progress Series, 340:109–119, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p109.pdf.

Chevaldonne:2000:TPW

[CFC^{+}00] P. Chevaldonné, C. R. Fisher, J. J. Childress, D. Desbruyères, D. Jollivet, F. Zal, and A. Toulmond. Thermotolerance and the ‘’Pompeii worms’. Marine Ecology Progress Series, 208: 293–295, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/208/m208p293.pdf.

Catchpole:2006:IDE

[CFG06] T. L. Catchpole, C. L. J. Frid, and T. S. Gray. Importance of discards from the English Nephrops norvegicus fishery in the North Sea to marine scavengers. Marine Ecology Progress Series, 313:215–226, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/313/m313p215.pdf.

Cocito:2000:PBC

[CFMB00] S. Cocito, F. Ferdeghini, C. Morri, and C. N. Bianchi. Patterns of bioconstruction in the cheilostome bryozoan Schizoporella errata: the influence of hydrodynamics and associated biota. Marine Ecology Progress Series, 192:153–161, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p153.pdf.

Costa:2004:DAA

[CG04] P. R. Costa and S. Garrido. Domoic acid accumulation in the sardine Sardina pilchardus and its relationship to Pseudonitzschia diatom ingestion. Marine Ecology Progress Series, 284:261–268, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p261.pdf.

Coyne:2005:STA

[CG05] M. S. Coyne and B. J. Godley. Satellite tracking and analysis tool (STAT): an integrated system for archiving, analyzing and mapping animal tracking data. Marine Ecology Progress Series, 301:1–7, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cui:2009:SDG

[CGC^{+}09] X. Cui, J. M. Grebmeier, L. W. Cooper, J. R. Lovvorn, C. A. North, W. L. Seaver, and J. M. Kolts. Spatial distributions of groundfish in the northern Bering Sea in relation to environmental variation. Marine Ecology Progress Series, 393:147–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/393/m393p147.pdf.

Charmantier:2002:OOP

[CGCDA02] G. Charmantier, L. Giménez, M. Charmantier-Daures, and K. Anger. Ontogeny of osmoregulation, physiological plasticity and larval export strategy in the grapsid crab Chasmagnathus granulata (Crustacea, Decapoda). Marine Ecology Progress Series, 229:185–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/229/m229p185.pdf.

Casamayor:2001:PPE

[CGCMPA01] E. O. Casamayor, J. García-Cantizano, J. Mas, and C. Pedrós-Alió. Primary production in estuarine oxic/anoxic interfaces: contribution of microbial dark CO_{2} fixation in the Ebro River Salt Wedge Estuary. Marine Ecology Progress Series, 215:49–56, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p049.pdf.

Castro-Gonzalez:2004:CCO

[CGF04] M. Castro-González and L. Farías. N_{2}O cycling at the core of the oxygen minimum zone off northern Chile. Marine Ecology Progress Series, 280:1–11, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/280/m280p001.pdf.

Coombs:2003:WMF

[CGH^{+}03] S. H. Coombs, O. Giovanardi, N. C. Halliday, G. Franceschini, D. V. P. Conway, L. Manzueto, C. D. Barrett, and I. R. B. McFadzen. Wind mixing, food availability and mortality of anchovy larvae Engraulis encrasicolus in the northern Adriatic Sea. Marine Ecology Progress Series, 248:221– 235, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p221.pdf.

Cairns:2008:EEG

[CGH08] D. K. Cairns, A. J. Gaston, and F. Huettmann. Endothermy, ectothermy and the global structure of marine vertebrate communities. Marine Ecology Progress Series, 356:239–250, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/356/m356p239.pdf.

Cooper:2002:SVS

[CGL^{+}02] L. W. Cooper, J. M. Grebmeier, I. L. Larsen, V. G. Egorov, C. Theodorakis, H. P. Kelly, and J. R. Lovvorn. Seasonal variation in sedimentation of organic materials in the St. Lawrence Island polynya region, Bering Sea. Marine Ecology Progress Series, 226:13–26, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p013.pdf.

Chiba:2000:FAO

[CGM00] S. Chiba, S. Goshima, and T. Mizushima. Factors affecting the occurrence of early maturing males in the protandrous pandalid shrimp Pandalus latirostris. Marine Ecology Progress Series, 203:215–224, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/203/m203p215.pdf.

Carroll:2008:RRG

[CGP08] J. Carroll, C. J. Gobler, and B. J. Peterson. Resourcerestricted growth of eelgrass in New York estuaries: light limitation, and alleviation of nutrient stress by hard clams. Marine Ecology Progress Series, 369:51–62, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/369/
m369p051.pdf.

Cancino:2000:ESP

[CGT^{+}00] J. M. Cancino, J. A. Gallardo, F. Torres, G. Leiva, and J. M. Navarro. Effects of sessile protozoa on intracapsular oxygen tension and embryonic shell calcification in the muricid Chorus giganteus. Marine Ecology Progress Series, 200:141–148, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p141.pdf.

Cetrulo:2000:ACD

[CH00a] G. L. Cetrulo and M. E. Hay. Activated chemical defenses in tropical versus temperate seaweeds. Marine Ecology Progress Series, 207:243–253, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/207/m207p243.pdf.

Craig:2000:HPF

[CH00b] A. S. Craig and L. M. Herman. Habitat preferences of female humpback whales Megaptera novaeangliae in the Hawaiian Islands are associated with reproductive status. Marine Ecology Progress Series, 193:209–216, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p209.pdf.

Carlton:2003:MMT

[CH03] J. T. Carlton and J. Hodder. Maritime mammals: terrestrial mammals as consumers in marine intertidal communities. Marine Ecology Progress Series, 256:271–286, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p271.pdf.

Cox:2005:CSA

[CH05] C. Cox and J. H. Hunt. Change in size and abundance of Caribbean spiny lobsters Panulirus argus in a marine reserve in the Florida Keys National Marine Sanctuary, USA. Marine Ecology Progress Series, 294:227–239, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p227.pdf.

Cardinale:2006:MFR

[CH06a] M. Cardinale and J. Hjelm. Marine fish recruitment variability and climate indices. Marine Ecology Progress Series, 309:307–309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/309/m309p307.pdf.

Carson:2006:EDP

[CH06b] H. S. Carson and B. T. Hentschel. Estimating the dispersal potential of polychaete species in the Southern California Bight: implications for designing marine reserves. Marine Ecology Progress Series, 316:105–113, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/316/m316p105.pdf.

Cattrijsse:2006:EIM

[CH06c] A. Cattrijsse and H. Hampel. European intertidal marshes: a review of their habitat functioning and value for aquatic organisms. Marine Ecology Progress Series, 324:293–307, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/324/m324p293.pdf.

Carleton:2007:HPC

[CH07a] J. H. Carleton and W. M. Hamner. The hyperbenthic plankton community: composition, distribution, and abundance in a coral reef lagoon. Marine Ecology Progress Series, 336: 77–88, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/336/m336p077.pdf.

Carstensen:2007:PRN

[CH07b] J. Carstensen and A. S. Heiskanen. Phytoplankton responses to nutrient status: application of a screening method to the northern Baltic Sea. Marine Ecology Progress Series, 336: 29–42, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m336p029.pdf.

Cherel:2007:GVC

[CH07c] Y. Cherel and K. A. Hobson. Geographical variation in carbon stable isotope signatures of marine predators: a tool to investigate their foraging areas in the Southern Ocean. Marine Ecology Progress Series, 329:281–287, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p281.pdf.

Colson:2007:CPG

[CH07d] I. Colson and R. N. Hughes. Contrasted patterns of genetic variation in the dogwhelk Nucella lapillus along two putative post-glacial expansion routes. Marine Ecology Progress Series, 343:183–191, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p183.pdf.

Canion:2009:EHC

[CH09] C. R. Canion and K. L. Heck, Jr. Effect of habitat complexity on predation success: re-evaluating the current paradigm in seagrass beds. Marine Ecology Progress Series, 393:37– 46, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m393p037.pdf.

Chapman:2003:PMS

[Cha03] M. G. Chapman. Paucity of mobile species on constructed seawalls: effects of urbanization on biodiversity. Marine Ecology Progress Series, 264:21–29, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p021.pdf.

Crawley:2006:IDV

[CHA06] K. R. Crawley, G. A. Hyndes, and S. G. Ayvazian. Influence of different volumes and types of detached macrophytes on fish community structure in surf zones of sandy beaches. Marine Ecology Progress Series, 307:233–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/307/
m307p233.pdf.

Connolly:2005:SEA

[CHG05] R. M. Connolly, J. S. Hindell, and D. Gorman. Seagrass and epiphytic algae support nutrition of a fisheries species, Sillago schomburgkii, in adjacent intertidal habitats. Marine Ecology Progress Series, 286:69–79, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/286/m286p069.pdf.

Case:2005:MMG

[CHH^{+}05] R. A. J. Case, W. F. Hutchinson, L. Hauser, C. Van Oosterhout, and G. R. Carvalho. Macro- and micro-geographic variation in pantophysin (Pan I) allele frequencies in NE Atlantic cod Gadus morhua. Marine Ecology Progress Series, 301:267–278, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p267.pdf.

Chittaro:2002:SAR

[Chi02] P. M. Chittaro. Species-area relationships for coral reef fish assemblages of St. Croix, US Virgin Islands. Marine Ecology Progress Series, 233:253–261, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/233/m233p253.pdf.

Chiswell:2009:CCI

[Chi09] S. M. Chiswell. Colonisation and connectivity by intertidal limpets among New Zealand, Chatham and Sub-Antarctic Islands. II. Oceanographic connections. Marine Ecology Progress Series, 388:121–135, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p121.pdf.

Carr:2005:SPO

[CHL^{+}05] S. D. Carr, J. L. Hench, R. A. Luettich, Jr., R. B. Forward, Jr., and R. A. Tankersley. Spatial patterns in the ovigerous Callinectes sapidus spawning migration: results from a coupled behavioral-physical model. Marine Ecology Progress Series, 294:213–226, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/294/m294p213.pdf.

Cambridge:2007:RAE

[CHLV07] M. L. Cambridge, J. R. How, P. S. Lavery, and M. A. Vanderklift. Retrospective analysis of epiphyte assemblages in relation to seagrass loss in a eutrophic coastal embayment. Marine Ecology Progress Series, 346:97–107, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p097.pdf.

Ceccarelli:2006:ISO

[CHM06] D. M. Ceccarelli, T. P. Hughes, and L. J. McCook. Impacts of simulated overfishing on the territoriality of coral reef damselfish. Marine Ecology Progress Series, 309:255– 262, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/309/m309p255.pdf.

Carls:2005:CPI

[CHMR05] M. G. Carls, R. A. Heintz, G. D. Marty, and S. D. Rice. Cytochrome P4501A induction in oil-exposed pink salmon Oncorhynchus gorbuscha embryos predicts reduced survival potential. Marine Ecology Progress Series, 301:253–265, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/301/m301p253.pdf.

Carvalho:2009:CSI

[CHO09] M. C. Carvalho, K. Hayashizaki, and H. Ogawa. Carbon stable isotope discrimination: a possible growth index for the kelp Undaria pinnatifida. Marine Ecology Progress Series, 381:71–82, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m381p071.pdf.

Cordes:2005:SHS

[CHP^{+}05] E. E. Cordes, S. Hourdez, B. L. Predmore, M. L. Redding, and C. R. Fisher. Succession of hydrocarbon seep communities associated with the long-lived foundation species Lamellibrachia luymesi. Marine Ecology Progress Series, 305: 17–29, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/305/m305p017.pdf.

Codling:2004:RWM

[CHPS04] E. A. Codling, N. A. Hill, J. W. Pitchford, and S. D. Simpson. Random walk models for the movement and recruitment of reef fish larvae. Marine Ecology Progress Series, 279:215–224, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p215.pdf.

Caffrey:2003:BPS

[CHSW03] J. M. Caffrey, N. Harrington, I. Solem, and B. B. Ward. Biogeochemical processes in a small California estuary. 2. Nitrification activity, community structure and role in nitrogen budgets. Marine Ecology Progress Series, 248:27– 40, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p027.pdf.

Carstensen:2006:IOW

[CHT06] J. Carstensen, O. D. Henriksen, and J. Teilmann. Impacts of offshore wind farm construction on harbour porpoises: acoustic monitoring of echolocation activity using porpoise detectors (T-PODs). Marine Ecology Progress Series, 321:295– 308, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m321p295.pdf.

Cowles:2009:DBP

[CHT09] A. Cowles, J. E. Hewitt, and R. B. Taylor. Density, biomass and productivity of small mobile invertebrates in a wide range of coastal habitats. Marine Ecology Progress Series, 384:175–185, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/384/m384p175.pdf.

Crawley:2009:ABA

[CHV^{+}09] K. R. Crawley, G. A. Hyndes, M. A. Vanderklift, A. T. Revill, and P. D. Nichols. Allochthonous brown algae are the primary food source for consumers in a temperate, coastal environment. Marine Ecology Progress Series, 376:33–44, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/376/m376p033.pdf.

Caffrey:2002:BPS

[CHW02] J. M. Caffrey, N. Harrington, and B. Ward. Biogeochemical processes in a small California estuary. 1. Benthic fluxes and pore water constituents reflect high nutrient freshwater inputs. Marine Ecology Progress Series, 233:39–53, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p039.pdf.

Coleman:2006:TRB

[CHW06] R. A. Coleman, S. J. Hawkins, and H. L. Wood. Testing the reproductive benefits of aggregation: the limpet Patella vulgata shows no evidence of synchrony in gonad development. Marine Ecology Progress Series, 306:201–207, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p201.pdf.

Carotenuto:2006:ACE

[CID^{+}06] Y. Carotenuto, A. Ianora, M. Di Pinto, D. Sarno, and A. Miralto. Annual cycle of early developmental stage survival and recruitment in the copepods Temora stylifera and Centropages typicus. Marine Ecology Progress Series, 314:227– 238, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p227.pdf.

Clavero:2000:SFP

[CIFN00] V. Clavero, J. J. Izquierdo, J. A. Fernández, and F. X. Niell. Seasonal fluxes of phosphate and ammonium across the sediment–water interface in a shallow small estuary (Palmones River, southern Spain). Marine Ecology Progress Series, 198:51–60, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p051.pdf.

Chiba:2001:STV

[CIHF01] S. Chiba, T. Ishimaru, G. W. Hosie, and M. Fukuchi. Spatiotemporal variability of zooplankton community structure off east Antarctica (90 to 160^{o}e). Marine Ecology Progress Series, 216:95–108, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/216/m216p095.pdf.

Castellani:2005:FEP

[CIHL05] C. Castellani, X. Irigoien, R. P. Harris, and R. S. Lampitt. Feeding and egg production of Oithona similis in the North Atlantic. Marine Ecology Progress Series, 288:173–182, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/288/m288p173.pdf.

Chang:2004:MEH

[CIT04] C. W. Chang, Y. Iizuka, and W. N. Tzeng. Migratory environmental history of the grey mullet Mugil cephalus as revealed by otolith Sr:Ca ratios. Marine Ecology Progress Series, 269:277–288, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/269/m269p277.pdf.

Campana:2009:CBS

[CJFM09] S. E. Campana, W. Joyce, M. P. Francis, and M. J. Manning. Comparability of blue shark mortality estimates for the Atlantic and Pacific longline fisheries. Marine Ecology Progress Series, 396:161–164, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/396/m396p161.pdf.

Chauvaud:2000:LTV

[CJRT00] L. Chauvaud, F. Jean, O. Ragueneau, and G. Thouzeau. Long-term variation of the Bay of Brest ecosystem: benthicpelagic coupling revisited. Marine Ecology Progress Series, 200:35–48, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/200/m200p035.pdf.

Chen:2004:PCE

[CK04] C. C. Chen and W. M. Kemp. Periphyton communities in experimental marine ecosystems: scaling the effects of removal from container walls. Marine Ecology Progress Series, 271:27–41, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/271/m271p027.pdf.

Conlan:2005:RSS

[CK05] K. E. Conlan and R. G. Kvitek. Recolonization of softsediment ice scours on an exposed Arctic coast. Marine Ecology Progress Series, 286:21–42, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p021.pdf.

Choi:2009:MSC

[CK09a] K. H. Choi and W. Kimmerer. Mating success and its consequences for population growth in an estuarine copepod. Marine Ecology Progress Series, 377:183–191, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/377/
m377p183.pdf.

Coleman:2009:CAF

[CK09b] M. A. Coleman and B. P. Kelaher. Connectivity among fragmented populations of a habitat-forming alga, Phyllospora comosa (Phaeophyceae, Fucales) on an urbanised coast. Marine Ecology Progress Series, 381:63–70, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/381/
m381p063.pdf.

Claisse:2009:HSS

[CKB^{+}09] J. T. Claisse, M. Kienzle, M. E. Bushnell, D. J. Shafer, and J. D. Parrish. Habitat- and sex-specific life history patterns of yellow tang Zebrasoma flavescens in Hawaii, USA. Marine Ecology Progress Series, 389:245–255, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p245.pdf.

Clarke:2003:PFW

[CKF^{+}03] J. Clarke, K. Kerry, C. Fowler, R. Lawless, S. Eberhard, and R. Murphy. Post-fledging and winter migration of Adélie penguins Pygoscelis adeliae in the Mawson region of East Antarctica. Marine Ecology Progress Series, 248:267–278, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/248/m248p267.pdf.

Chikaraishi:2007:MCN

[CKO^{+}07] Y. Chikaraishi, Y. Kashiyama, N. O. Ogawa, H. Kitazato, and N. Ohkouchi. Metabolic control of nitrogen isotope composition of amino acids in macroalgae and gastropods: implications for aquatic food web studies. Marine Ecology Progress Series, 342:85–90, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/342/m342p085.pdf.

Congdon:2005:DFC

[CKS05] B. C. Congdon, A. K. Krockenberger, and B. V. Smithers. Dual-foraging and co-ordinated provisioning in a tropical procellariiform, the wedge-tailed shearwater. Marine Ecology Progress Series, 301:293–301, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/301/m301p293.pdf.

Culik:2001:RHP

[CKTE01] B. M. Culik, S. Koschinski, N. Tregenza, and G. M. Ellis. Reactions of harbor porpoises Phocoena phocoena and herring Clupea harengus to acoustic alarms. Marine Ecology Progress Series, 211:255–260, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/211/m211p255.pdf.

Cartamil:2004:DMP

[CL04] D. P. Cartamil and C. G. Lowe. Diel movement patterns of ocean sunfish Mola mola off southern California. Marine Ecology Progress Series, 266:245–253, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p245.pdf.

Cobb:2005:DCS

[CL05] J. Cobb and J. M. Lawrence. Diets and coexistence of the sea urchins Lytechinus variegatus and Arbacia punctulata (Echinodermata) along the central Florida gulf coast. Marine Ecology Progress Series, 295:171–182, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p171.pdf.

Clarke:2008:ESS

[Cla08] A. Clarke. Ecological stoichiometry in six species of Antarctic marine benthos. Marine Ecology Progress Series, 369: 25–37, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m369p025.pdf.

Castro-Longoria:2003:PMA

[CLABRO^{+}03] E. Castro-Longoria, J. Alvarez-Borrego, A. Rocha-Olivares, S. Gomez, and V. Kober. Power of a multidisciplinary approach: use of morphological, molecular and digital methods in the study of harpacticoid cryptic species. Marine Ecology Progress Series, 249:297–303, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/249/m249p297.pdf.

Cardinale:2009:CVH

[CLB^{+}09] M. Cardinale, M. Linder, V. Bartolino, L. Maiorano, and M. Casini. Conservation value of historical data: reconstructing stock dynamics of turbot during the last century in the Kattegat–Skagerrak. Marine Ecology Progress Series, 386:197–206, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/386/m386p197.pdf.

Castilla:2004:MEE

[CLC04] J. C. Castilla, N. A. Lagos, and M. Cerda. Marine ecosystem engineering by the alien ascidian Pyura praeputialis on a mid-intertidal rocky shore. Marine Ecology Progress Series, 268:119–130, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/268/m268p119.pdf.

Chen:2007:ESA

[CLH^{+}07] H. Chen, B. Li, J. Hu, J. Chen, and J. Wu. Effects of Spartina alterniflora invasion on benthic nematode communities in the Yangtze Estuary. Marine Ecology Progress Series, 336:99–110, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p099.pdf.

Clark:2005:SOL

[CLHT05] D. L. Clark, J. M. Leis, A. C. Hay, and T. Trnski. Swimming ontogeny of larvae of four temperate marine fishes. Marine Ecology Progress Series, 292:287–300, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p287.pdf.

Cherel:2008:RPW

[CLJ^{+}08] Y. Cherel, M. Le Corre, S. Jaquemet, F. Ménard, P. Richard, and H. Weimerskirch. Resource partitioning within a tropical seabird community: new information from stable isotopes. Marine Ecology Progress Series, 366:281–291, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/366/
m366p281.pdf.

Chen:2009:ENL

[CLL^{+}09] B. Chen, H. Liu, M. R. Landry, M. Chen, J. Sun, L. Shek, X. Chen, and P. J. Harrison. Estuarine nutrient loading affects phytoplankton growth and microzooplankton grazing at two contrasting sites in Hong Kong coastal waters. Marine Ecology Progress Series, 379:77–90, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/379/m379p077.pdf.

Collier:2007:MGM

[CLMR07] C. J. Collier, P. S. Lavery, R. J. Masini, and P. J. Ralph. Morphological, growth and meadow characteristics of the seagrass Posidonia sinuosa along a depth-related gradient of light availability. Marine Ecology Progress Series, 337:103– 115, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p103.pdf.

Camacho:2000:EBM

[CLN00] A. P. Camacho, U. Labarta, and E. Navarro. Energy balance of mussels Mytilus galloprovincialis: the effect of length and age. Marine Ecology Progress Series, 199:149–158, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p149.pdf.

Cloern:2001:OEC

[Clo01] J. E. Cloern. Our evolving conceptual model of the coastal eutrophication problem. Marine Ecology Progress Series, 210: 223–253, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/210/m210p223.pdf.

Carlier:2009:TRD

[CLO^{+}09] A. Carlier, E. Le Guilloux, K. Olu, J. Sarrazin, F. Mastrototaro, M. Taviani, and J. Clavier. Trophic relationships in a deep Mediterranean cold-water coral bank (Santa Maria di Leuca, Ionian Sea). Marine Ecology Progress Series, 397:125–137, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p125.pdf.

Casale:2006:SRJ

[CLP^{+}06] Paolo Casale, Bojan Lazar, Sara Pont, Jesús Tomás, Nicola Zizzo, Ferran Alegre, Javier Badillo, Aldo Di Summa, Daniela Freggi, Gordana Lackovic, Juan Antonio Raga, Lucio Rositani, and Nikola Tvrtković. Sex ratios of juvenile loggerhead sea turtles Caretta caretta in the Mediterranean Sea. Marine Ecology Progress Series, 324:281–285, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p281.pdf.

Chu:2008:QSE

[CLP08] F. L. E. Chu, E. D. Lund, and J. A. Podbesek. Quantitative significance of n-3 essential fatty acid contribution by heterotrophic protists in marine pelagic food webs. Marine Ecology Progress Series, 354:85–95, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/354/
m354p085.pdf.

Coma:2006:CMM

[CLR^{+}06] R. Coma, C. Linares, M. Ribes, D. Diaz, J. Garrabou, and E. Ballesteros. Consequences of a mass mortality in populations of Eunicella singularis (Cnidaria: Octocorallia) in Menorca (NW Mediterranean). Marine Ecology Progress Series, 327:51–60, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/327/m327p051.pdf.

Collier:2008:PCS

[CLRM08] C. J. Collier, P. S. Lavery, P. J. Ralph, and R. J. Masini. Physiological characteristics of the seagrass Posidonia sinuosa along a depth-related gradient of light availability. Marine Ecology Progress Series, 353:65–79, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/353/m353p065.pdf.

Calbet:2000:CGS

[CLS00] A. Calbet, M. R. Landry, and R. D. Scheinberg. Copepod grazing in a subtropical bay: species-specific responses to a midsummer increase in nanoplankton standing stock. Marine Ecology Progress Series, 193:75–84, 2000. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/193/m193p075.pdf.

Castberg:2001:MPD

[CLS^{+}01] T. Castberg, A. Larsen, R. A. Sandaa, C. P. D. Brussaard, J. K. Egge, M. Heldal, R. Thyrhaug, E. J. van Hannen, and G. Bratbak. Microbial population dynamics and diversity during a bloom of the marine coccolithophorid Emiliania huxleyi (Haptophyta). Marine Ecology Progress Series, 221:39–46, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/221/m221p039.pdf.

Cheek:2009:DHM

[CLSM09] A. O. Cheek, C. A. Landry, S. L. Steele, and S. Manning. Diel hypoxia in marsh creeks impairs the reproductive capacity of estuarine fish populations. Marine Ecology Progress Series, 392:211–221, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p211.pdf.

Cartes:2001:TDW

[CM01] J. E. Cartes and F. Maynou. Trophodynamics of the deepwater suprabenthic mysid Boreomysis arctica in the Catalan Sea (western Mediterranean). Marine Ecology Progress Series, 211:225–234, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/211/m211p225.pdf.

Carrasson:2002:DDS

[CM02] M. Carrassón and J. Matallanas. Diets of deep-sea macrourid fishes in the western Mediterranean. Marine Ecology Progress Series, 234:215–228, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/234/m234p215.pdf.

Cabral:2004:ESD

[CM04a] H. N. Cabral and A. G. Murta. Effect of sampling design on abundance estimates of benthic invertebrates in environmental monitoring studies. Marine Ecology Progress Series, 276:19–24, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p019.pdf.

Cole:2004:AID

[CM04b] R. Cole and G. McBride. Assessing impacts of dredge spoil disposal using equivalence tests: implications of a precautionary (proof of safety) approach. Marine Ecology Progress Series, 279:63–72, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p063.pdf.

Cruz-Motta:2005:DTV

[CM05] J. J. Cruz-Motta. Diel and tidal variations of benthic assemblages in sediments associated with boulder fields. Marine Ecology Progress Series, 290:97–107, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p097.pdf.

Carleton:2007:RMS

[CM07] J. H. Carleton and A. D. McKinnon. Resident mysids: secondary production, consumption, and trophic role in a coral reef lagoon. Marine Ecology Progress Series, 336:89– 98, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/336/m336p089.pdf.

Coupland:2008:EHE

[CM08] G. T. Coupland and J. I. McDonald. Extraordinarily high earthworm abundance in deposits of marine macrodetritus along two semi-arid beaches. Marine Ecology Progress Series, 361:181–189, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p181.pdf.

Connan:2005:LCS

[CMB^{+}05a] M. Connan, P. Mayzaud, M. Boutoute, H. Weimerskirch, and Y. Cherel. Lipid composition of stomach oil in a procellariiform seabird Puffinus tenuirostris: implications for food web studies. Marine Ecology Progress Series, 290:277– 290, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/290/m290p277.pdf.

Croll:2005:WWT

[CMB^{+}05b] D. A. Croll, B. Marinovic, S. Benson, F. P. Chavez, N. Black, R. Ternullo, and B. R. Tershy. From wind to whales: trophic links in a coastal upwelling system. Marine Ecology Progress Series, 289:117–130, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/289/m289p117.pdf.

Cardinale:2009:EEV

[CMB^{+}09] M. Cardinale, C. Möllmann, V. Bartolino, M. Casini, G. Kornilovs, T. Raid, P. Margonski, A. Grzyb, J. Raitaniemi, T. Gröhsler, and J. Flinkman. Effect of environmental variability and spawner characteristics on the recruitment of Baltic herring Clupea harengus populations. Marine Ecology Progress Series, 388:221–234, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p221.pdf.

Chan:2006:SME

[CMD^{+}06] B. K. K. Chan, D. Morritt, M. De Pirro, K. M. Y. Leung, and G. A. Williams. Summer mortality: effects on the distribution and abundance of the acorn barnacle Tetraclita japonica on tropical shores. Marine Ecology Progress Series, 328:195–204, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p195.pdf.

Callier:2007:MSS

[CMD07] M. D. Callier, C. W. McKindsey, and G. Desrosiers. Multiscale spatial variations in benthic sediment geochemistry and macrofaunal communities under a suspended mussel culture. Marine Ecology Progress Series, 348:103–115, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p103.pdf.

Coles:2009:SDD

[CMD^{+}09] R. Coles, L. McKenzie, G. De’ath, A. Roelofs, and W. Lee Long. Spatial distribution of deepwater seagrass in the interreef lagoon of the Great Barrier Reef world heritage area. Marine Ecology Progress Series, 392:57–68, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/392/
m392p057.pdf.

Caron:2004:SCP

[CMG04] G. Caron, C. Michel, and M. Gosselin. Seasonal contributions of phytoplankton and fecal pellets to the organic carbon sinking flux in the North Water (northern Baffin Bay). Marine Ecology Progress Series, 283:1–13, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p001.pdf.

Collin:2005:ECA

[CMGBJ05] R. Collin, M. McLellan, K. Gruber, and C. Bailey-Jourdain. Effects of conspecific associations on size at sex change in three species of calyptraeid gastropods. Marine Ecology Progress Series, 293:89–97, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p089.pdf.

Chan:2006:ECP

[CMHP06] F. Chan, R. L. Marino, R. W. Howarth, and M. L. Pace. Ecological constraints on planktonic nitrogen fixation in saline estuaries. II. Grazing controls on cyanobacterial population dynamics. Marine Ecology Progress Series, 309:41–53, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m309p041.pdf.

Chisholm:2000:ELW

[CMJ00] J. R. M. Chisholm, M. Marchioretti, and J. M. Jaubert. Effect of low water temperature on metabolism and growth of a subtropical strain of Caulerpa taxifolia (Chlorophyta). Marine Ecology Progress Series, 201:189–198, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p189.pdf.

Carmack:2004:PPC

[CMJ04] E. C. Carmack, R. W. Macdonald, and S. Jasper. Phytoplankton productivity on the Canadian Shelf of the Beaufort Sea. Marine Ecology Progress Series, 277:37–50, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p037.pdf.

Chan:2008:LGD

[CML08] B. K. K. Chan, A. Murata, and P. F. Lee. Latitudinal gradient in the distribution of intertidal barnacles of the Tetraclita species complex in Asian waters. Marine Ecology Progress Series, 362:201–210, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/362/m362p201.pdf.

Chassot:2007:BCR

[CMLG07] E. Chassot, F. Mélin, O. Le Pape, and D. Gascuel. Bottomup control regulates fisheries production at the scale of ecoregions in European seas. Marine Ecology Progress Series, 343:45–55, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p045.pdf.

Cartes:2001:TSB

[CMMN^{+}01] J. E. Cartes, F. Maynou, B. Morales-Nin, E. Massutí, and J. Moranta. Trophic structure of a bathyal benthopelagic boundary layer community south of the Balearic Islands (southwestern Mediterranean). Marine Ecology Progress Series, 215:23–35, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p023.pdf.

Couceiro:2008:ISC

[CMRB08] L. Couceiro, A. Míguez, J. M. Ruiz, and R. Barreiro. Introduced status of Cyclope neritea (Gastropoda, Nassariidae) in the NW Iberian Peninsula confirmed by mitochondrial sequence data. Marine Ecology Progress Series, 354: 141–146, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p141.pdf.

Cermeno:2005:LSP

[CMRF05] P. Cermeño, E. Marañón, J. Rodríguez, and E. Fernández. Large-sized phytoplankton sustain higher carbon-specific photosynthesis than smaller cells in a coastal eutrophic ecosystem. Marine Ecology Progress Series, 297:51–60, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p051.pdf.

Counihan:2001:PSP

[CMS^{+}01] R. T. Counihan, D. C. McNamara, D. C. Souter, E. J. Jebreen, N. P. Preston, C. R. Johnson, and B. M. Degnan. Pattern, synchrony and predictability of spawning of the tropical abalone Haliotis asinina from Heron Reef, Australia. Marine Ecology Progress Series, 213:193–202, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/213/m213p193.pdf.

Cheriton:2009:PBC

[CMSS09] O. M. Cheriton, M. A. McManus, M. T. Stacey, and J. V. Steinbuck. Physical and biological controls on the maintenance and dissipation of a thin phytoplankton layer. Marine Ecology Progress Series, 378:55–69, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m378p055.pdf.

Connan:2008:IDC

[CMT^{+}08] M. Connan, P. Mayzaud, C. Trouvé, C. Barbraud, and Y. Cherel. Interannual dietary changes and demographic consequences in breeding blue petrels from Kerguelen Islands. Marine Ecology Progress Series, 373:123–135, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/373/
m373p123.pdf.

Cerco:2004:PBP

[CN04] C. F. Cerco and M. R. Noel. Process-based primary production modeling in Chesapeake Bay. Marine Ecology Progress Series, 282:45–58, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p045.pdf.

Cho:2000:AIF

[CNC00] B. C. Cho, S. C. Na, and D. H. Choi. Active ingestion of fluorescently labeled bacteria by mesopelagic heterotrophic nanoflagellates in the East Sea, Korea. Marine Ecology Progress Series, 206:23–32, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p023.pdf.

Christie:2009:MHF

[CNF09] H. Christie, K. M. Norderhaug, and S. Fredriksen. Macrophytes as habitat for fauna. Marine Ecology Progress Series, 396:221–233, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m396p221.pdf.

Choo:2005:PCU

[CNPS05] K. S. Choo, J. Nilsson, M. Pedersén, and P. Snoeijs. Photosynthesis, carbon uptake and antioxidant defence in two coexisting filamentous green algae under different stress conditions. Marine Ecology Progress Series, 292:127–138, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p127.pdf.

Costa:2004:RAS

[CNT^{+}04] F. O. Costa, T. Neuparth, C. W. Theodorakis, M. H. Costa, and L. R. Shugart. RAPD analysis of southern populations of Gammarus locusta: comparison with allozyme data and ecological inferences. Marine Ecology Progress Series, 277: 197–207, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/277/m277p197.pdf.

Cantin:2007:PCH

[CNW07a] N. E. Cantin, A. P. Negri, and B. L. Willis. Photoinhibition from chronic herbicide exposure reduces reproductive output of reef-building corals. Marine Ecology Progress Series, 344:81–93, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m344p081.pdf.

Chiu:2007:LEL

[CNW^{+}07b] J. M. Y. Chiu, T. Y. T. Ng, W. X. Wang, V. Thiyagarajan, and P. Y. Qian. Latent effects of larval food limitation on filtration rate, carbon assimilation and growth in juvenile gastropod Crepidula onyx. Marine Ecology Progress Series, 343:173–182, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p173.pdf.

Chauton:2003:HNS

[COB^{+}03] M. S. Chauton, O. I. Optun, T. F. Bathen, Z. Volent, I. S. Gribbestad, and G. Johnsen. HR MAS ^{1}H NMR spectroscopy analysis of marine microalgal whole cells. Marine Ecology Progress Series, 256:57–62, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p057.pdf.

Chambers:2002:HCC

[COK02] R. M. Chambers, D. T. Osgood, and N. Kalapasev. Hydrologic and chemical control of Phragmites growth in tidal marshes of SW Connecticut, USA. Marine Ecology Progress Series, 239:83–91, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p083.pdf.

Collin:2003:WPM

[Col03] R. Collin. Worldwide patterns in mode of development in calyptraeid gastropods. Marine Ecology Progress Series, 247: 103–122, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p103.pdf.

Connell:2005:AMS

[Con05] S. D. Connell. Assembly and maintenance of subtidal habitat heterogeneity: synergistic effects of light penetration and sedimentation. Marine Ecology Progress Series, 289:53–61, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/289/m289p053.pdf.

Catalan:2006:LBE

[COPB06] I. A. Catalán, M. P. Olivar, I. Palomera, and E. Berdalet. Link between environmental anomalies, growth and condition of pilchard Sardina pilchardus larvae in the northwestern Mediterranean. Marine Ecology Progress Series, 307: 219–231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p219.pdf.

Corkeron:2009:RSS

[Cor09] P. J. Corkeron. Reconsidering the science of scientific whaling. Marine Ecology Progress Series, 375:305–309, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p305.pdf.

Costlow:2003:MPC

[Cos03] John D. Costlow. In memoriam: Professor Cazlyn Green Bookhout (1907–2003). Marine Ecology Progress Series, 258:1, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/misc/m258p001.pdf.

Costello:2009:DMH

[Cos09] M. J. Costello. Distinguishing marine habitat classification concepts for ecological data management. Marine Ecology Progress Series, 397:253–268, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m397p253.pdf.

Cote:2003:MUJ

[COSM03] D. Cote, L. M. N. Ollerhead, D. A. Scruton, and R. S. McKinley. Microhabitat use of juvenile Atlantic cod in a coastal area of Newfoundland determined by 2D telemetry. Marine Ecology Progress Series, 265:227–234, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/265/m265p227.pdf.

Carbonell:2003:MGP

[CPA^{+}03] A. Carbonell, M. Palmer, P. Abelló, P. Torres, R. Alemany, and L. G. de Sola. Mesoscale geographical patterns in the distribution of pandalid shrimps Plesionika spp. in the Western Mediterranean. Marine Ecology Progress Series, 247: 151–158, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p151.pdf.

Chaloupka:2004:MPR

[CPB04a] M. Chaloupka, D. Parker, and G. Balazs. Modelling postrelease mortality of loggerhead sea turtles exposed to the Hawaii-based pelagic longline fishery. Marine Ecology Progress Series, 280:285–293, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/280/m280p285.pdf.

Chaloupka:2004:TTT

[CPB04b] M. Chaloupka, D. Parker, and G. Balazs. Tracking turtles to their death — reply to Hays et al. Marine Ecology Progress Series, 283:301–302, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p301.pdf. See

[HBG^{+}04].

Chapman:2007:DDD

[CPBS07] D. D. Chapman, E. K. Pikitch, E. A. Babcock, and M. S. Shivji. Deep-diving and diel changes in vertical habitat use by Caribbean reef sharks Carcharhinus perezi. Marine Ecology Progress Series, 344:271–275, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/344/m344p271.pdf.

Cramer:2008:DRO

[CPG08] K. L. Cramer, W. L. Perryman, and T. Gerrodette. Declines in reproductive output in two dolphin populations depleted by the yellowfin tuna purse-seine fishery. Marine Ecology Progress Series, 369:273–285, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/369/m369p273.pdf.

Copeman:2009:FAB

[CPG^{+}09] L. A. Copeman, C. C. Parrish, R. S. Gregory, R. E. Jamieson, J. Wells, and M. J. Whiticar. Fatty acid biomarkers in coldwater eelgrass meadows: elevated terrestrial input to the food web of age-0 Atlantic cod Gadus morhua. Marine Ecology Progress Series, 386:237–251, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/386/m386p237.pdf.

Choi:2002:ETE

[CPH^{+}02] D. H. Choi, J. S. Park, C. Y. Hwang, S. H. Huh, and B. C. Cho. Effects of thermal effluents from a power station on bacteria and heterotrophic nanoflagellates in coastal waters. Marine Ecology Progress Series, 229:1–10, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p001.pdf.

Chen:2000:NUE

[CPK00] C. C. Chen, J. E. Petersen, and W. M. Kemp. Nutrient uptake in experimental estuarine ecosystems: scaling and partitioning rates. Marine Ecology Progress Series, 200:103–116, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p103.pdf.

Cebrian:2000:FPS

[CPKV00] J. Cebrián, M. F. Pedersen, K. D. Kroeger, and I. Valiela. Fate of production of the seagrass Cymodocea nodosa in different stages of meadow formation. Marine Ecology Progress Series, 204:119–130, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p119.pdf.

Catry:2004:FSG

[CPP^{+}04] P. Catry, R. A. Phillips, B. Phalan, J. R. D. Silk, and J. P. Croxall. Foraging strategies of grey-headed albatrosses Thalassarche chrysostoma: integration of movements, activity and feeding events. Marine Ecology Progress Series, 280: 261–273, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/280/m280p261.pdf.

Cho:2001:SST

[CPSC01] B. C. Cho, M. G. Park, J. H. Shim, and D. H. Choi. Seasurface temperature and f-ratio explain large variability in the ratio of bacterial production to primary production in the Yellow Sea. Marine Ecology Progress Series, 216:31–41, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p031.pdf.

Ciglenecki:2003:MTS

[CPV^{+}03] I. Ciglenecki, M. Plavsic, V. Vojvodic, B. Cosovic, M. Pepi, and F. Baldi. Mucopolysaccharide transformation by sulfide in diatom cultures and natural mucilage. Marine Ecology Progress Series, 263:17–27, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/263/m263p017.pdf.

Chang:2002:RID

[CR02] S. Chang II and J. R. Reinfelder. Relative importance of dissolved versus trophic bioaccumulation of copper in marine copepods. Marine Ecology Progress Series, 231:179– 186, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p179.pdf.

Cox:2004:EBH

[CR04] T. M. Cox and A. J. Read. Echolocation behavior of harbor porpoises Phocoena phocoena around chemically enhanced gill nets. Marine Ecology Progress Series, 279:275– 282, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p275.pdf.

Chatzinikolaou:2007:EGA

[CR07a] E. Chatzinikolaou and C. A. Richardson. Evaluating growth and age of netted whelk Nassarius reticulatus (Gastropoda: Nassariidae) using statolith growth rings. Marine Ecology Progress Series, 342:163–176, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p163.pdf.

Cleary:2007:RST

[CR07b] D. F. R. Cleary and W. Renema. Relating species traits of Foraminifera to environmental variables in the Spermonde Archipelago, Indonesia. Marine Ecology Progress Series, 334: 73–82, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/334/m334p073.pdf.

Christensen:2008:SIE

[CR08] J. T. Christensen and K. Richardson. Stable isotope evidence of long-term changes in the North Sea food web structure. Marine Ecology Progress Series, 368:1–8, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cranford:2001:ERF

[Cra01] P. J. Cranford. Evaluating the ’reliability’ of filtration rate measurements in bivalves. Marine Ecology Progress Series, 215:303–305, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p303.pdf. See reply [Rii01c].

Choat:2003:ABD

[CRAP03] J. H. Choat, D. R. Robertson, J. L. Ackerman, and J. M. Posada. An age-based demographic analysis of the Caribbean stoplight parrotfish Sparisoma viride. Marine Ecology Progress Series, 246:265–277, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p265.pdf.

Cordell:2007:BIC

[CRB07] J. R. Cordell, M. Rasmussen, and S. M. Bollens. Biology of the introduced copepod Pseudodiaptomus inopinus in a northeast Pacific estuary. Marine Ecology Progress Series, 333:213–227, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p213.pdf.

Cook:2004:CNCb

[CRBE04] P. L. M. Cook, A. T. Revill, E. C. V. Butler, and B. D. Eyre. Carbon and nitrogen cycling on intertidal mudflats of a temperate Australian estuary. II. Nitrogen cycling. Marine Ecology Progress Series, 280:39–54, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p039.pdf.

Congiu:2002:PAS

[CRC02] L. Congiu, R. Rossi, and G. Colombo. Population analysis of the sand smelt Atherina boyeri (Teleostei Atherinidae), from Italian coastal lagoons by random amplified polymorphic DNA. Marine Ecology Progress Series, 229:279–289, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/229/m229p279.pdf.

Craig:2007:DDG

[CRCN07] J. K. Craig, J. A. Rice, L. B. Crowder, and D. A. Nadeau. Density-dependent growth and mortality in an estuarydependent fish: an experimental approach with juvenile spot Leiostomus xanthurus. Marine Ecology Progress Series, 343:

251–262, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p251.pdf.

Cook:2004:CNCc

[CRCV04] P. L. M. Cook, A. T. Revill, L. A. Clementson, and J. K. Volkman. Carbon and nitrogen cycling on intertidal mudflats of a temperate Australian estuary. III. Sources of organic matter. Marine Ecology Progress Series, 280:55–72, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p055.pdf.

Coma:2001:UOC

[CRGH01] R. Coma, M. Ribes, J. M. Gili, and R. N. Hughes. The ultimate opportunists: consumers of seston. Marine Ecology Progress Series, 219:305–308, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p305.pdf.

Cruz-Rivera:2001:MTF

[CRH01] E. Cruz-Rivera and M. E. Hay. Macroalgal traits and the feeding and fitness of an herbivorous amphipod: the roles of selectivity, mixing, and compensation. Marine Ecology Progress Series, 218:249–266, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/218/m218p249.pdf.

Catry:2009:CFE

[CRJ^{+}09] T. Catry, J. A. Ramos, S. Jaquemet, L. Faulquier, M. Berlincourt, A. Hauselmann, P. Pinet, and M. Le Corre. Comparative foraging ecology of a tropical seabird community of the Seychelles, western Indian Ocean. Marine Ecology Progress Series, 374:259–272, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/374/m374p259.pdf.

Cox:2006:TSS

[CRK06] E. F. Cox, M. Ribes, and R. A. Kinzie III. Temporal and spatial scaling of planktonic responses to nutrient inputs into a subtropical embayment. Marine Ecology Progress Series, 324: 19–35, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/324/m324p019.pdf.

Catry:2009:MSD

[CRLP09] T. Catry, J. A. Ramos, M. Le Corre, and R. A. Phillips. Movements, at-sea distribution and behaviour of a tropical pelagic seabird: the wedge-tailed shearwater in the western Indian Ocean. Marine Ecology Progress Series, 391:231–242, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m391p231.pdf.

Cardona:2007:MSP

[CRS^{+}07] L. Cardona, M. Revelles, M. Sales, A. Aguilar, and A. Borrell. Meadows of the seagrass Posidonia oceanica are a significant source of organic matter for adjoining ecosystems. Marine Ecology Progress Series, 335:123–131, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p123.pdf.

Castellani:2005:TAR

[CRSL05] C. Castellani, C. Robinson, T. Smith, and R. S. Lampitt. Temperature affects respiration rate of Oithona similis. Marine Ecology Progress Series, 285:129–135, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p129.pdf.

Connell:2008:RLB

[CRT^{+}08] S. D. Connell, B. D. Russell, D. J. Turner, S. A. Shepherd, T. Kildea, D. Miller, L. Airoldi, and A. Cheshire. Recovering a lost baseline: missing kelp forests from a metropolitan coast. Marine Ecology Progress Series, 360:63–72, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m360p063.pdf.

Childs:2002:SDG

[CRTP02] C. R. Childs, N. N. Rabalais, R. E. Turner, and L. M. Proctor. Sediment denitrification in the Gulf of Mexico zone of hypoxia. Marine Ecology Progress Series, 240:285–290, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/

m240p285.pdf. See erratum [Ano03c].

Carmichael:2003:APS

[CRV03] R. H. Carmichael, D. Rutecki, and I. Valiela. Abundance and population structure of the Atlantic horseshoe crab Limulus polyphemus in Pleasant Bay, Cape cod. Marine Ecology Progress Series, 246:225–239, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/246/m246p225.pdf.

Cummins:2004:EGM

[CRZ04] S. P. Cummins, D. E. Roberts, and K. D. Zimmerman. Effects of the green macroalga Enteromorpha intestinalis on macrobenthic and seagrass assemblages in a shallow coastal estuary. Marine Ecology Progress Series, 266:77–87, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/266/m266p077.pdf.

Crawford:2001:BSV

[CS01] S. M. Crawford and L. P. Sanford. Boundary-shear velocities and fluxes in the MEERC experimental ecosystems. Marine Ecology Progress Series, 210:1–12, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/210/m210p001.pdf.

Chabot:2002:GSF

[CS02a] D. Chabot and G. B. Stenson. Growth and seasonal fluctuations in size and condition of male Northwest Atlantic harp seals Phoca groenlandica: an analysis using sequential growth curves. Marine Ecology Progress Series, 227: 25–42, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/227/m227p025.pdf.

Chiba:2002:ICU

[CS02b] S. Chiba and T. Saino. Interdecadal change in the upper water column environment and spring diatom community structure in the Japan Sea: an early summer hypothesis. Marine Ecology Progress Series, 231:23–35, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p023.pdf.

Chittaro:2003:SPR

[CS03a] P. M. Chittaro and P. F. Sale. Structure of patch-reef fish assemblages at St. Croix, US Virgin Islands, and One Tree Reef, Australia. Marine Ecology Progress Series, 249:277– 287, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/249/m249p277.pdf.

Corsetti:2003:HSS

[CS03b] J. L. Corsetti and K. M. Strasser. Host selection of the symbiotic copepod Clausidium dissimile in two sympatric populations of ghost shrimp. Marine Ecology Progress Series, 256:151–159, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p151.pdf.

Craft:2003:LTS

[CS03c] C. Craft and J. Sacco. Long-term succession of benthic infauna communities on constructed Spartina alterniflora marshes. Marine Ecology Progress Series, 257:45–58, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/257/
m257p045.pdf.

Cotte:2005:FDK

[CS05] C. Cotté and Y. Simard. Formation of dense krill patches under tidal forcing at whale feeding hot spots in the St. Lawrence Estuary. Marine Ecology Progress Series, 288:199– 210, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/288/m288p199.pdf.

Condon:2008:DBR

[CS08] R. H. Condon and D. K. Steinberg. Development, biological regulation, and fate of ctenophore blooms in the York River estuary, Chesapeake Bay. Marine Ecology Progress Series, 369:153–168, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p153.pdf.

Carlisle:2009:HUR

[CS09] A. B. Carlisle and R. M. Starr. Habitat use, residency, and seasonal distribution of female leopard sharks Triakis semifasciata in Elkhorn Slough, California. Marine Ecology Progress Series, 380:213–228, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m380p213.pdf.

Calbet:2002:CEP

[CSA02] A. Calbet, E. Saiz, and M. Alcaraz. Copepod egg production in the NW Mediterranean: effects of winter environmental conditions. Marine Ecology Progress Series, 237:173– 184, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p173.pdf.

Caldarone:2003:RRD

[CSB03] E. M. Caldarone, J. M. St. Onge-Burns, and L. J. Buckley. Relationship of RNA/DNA ratio and temperature to growth in larvae of Atlantic cod Gadus morhua. Marine Ecology Progress Series, 262:229–240, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p229.pdf.

Carrere:2004:CST

[CSD04] V. Carrère, N. Spilmont, and D. Davoult. Comparison of simple techniques for estimating chlorophyll-a concentration in the intertidal zone using high spectral-resolution fieldspectrometer data. Marine Ecology Progress Series, 274: 31–40, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/274/m274p031.pdf.

Cranford:2007:IMA

[CSD^{+}07] P. J. Cranford, P. M. Strain, M. Dowd, B. T. Hargrave, J. Grant, and M. Archambault. Influence of mussel aquaculture on nitrogen dynamics in a nutrient enriched coastal embayment. Marine Ecology Progress Series, 347:61–78, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/347/m347p061.pdf.

Coffroth:2001:EOE

[CSG01] M. A. Coffroth, S. R. Santos, and T. L. Goulet. Early ontogenetic expression of specificity in a cnidarian-algal symbiosis. Marine Ecology Progress Series, 222:85–96, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p085.pdf.

Castro:2000:EIW

[CSH00] L. R. Castro, G. R. Salinas, and E. H. Hernández. Environmental influences on winter spawning of the anchoveta Engraulis ringens off central Chile. Marine Ecology Progress Series, 197:247–258, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p247.pdf.

Chen:2005:PBB

[CSL^{+}05] C. C. Chen, F. K. Shiah, H. J. Lee, K. Y. Li, P. J. Meng, S. J. Kao, Y. F. Tseng, and C. L. Chung. Phytoplankton and bacterioplankton biomass, production and turnover in a semi-enclosed embayment with spring tide induced upwelling. Marine Ecology Progress Series, 304:91–100, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p091.pdf.

Chiappone:2002:DSD

[CSM02] M. Chiappone, D. W. Swanson, and S. L. Miller. Density, spatial distribution and size structure of sea urchins in Florida Keys coral reef and hard-bottom habitats. Marine Ecology Progress Series, 235:117–126, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/235/m235p117.pdf.

Chaparro:2009:BCI

[CSM^{+}09] O. R. Chaparro, C. J. Segura, J. A. Montory, J. M. Navarro, and J. A. Pechenik. Brood chamber isolation during salinity stress in two estuarine mollusk species: from a protective nursery to a dangerous prison. Marine Ecology Progress Series, 374:145–155, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/374/m374p145.pdf.

Chabot:2008:ACS

[CSMBH08] D. Chabot, B. Sainte-Marie, K. Briand, and J. M. Hanson. Atlantic cod and snow crab predator–prey size relationship in the Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 363:227–240, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/363/m363p227.pdf.

Carreiro-Silva:2009:EIN

[CSMK09] M. Carreiro-Silva, T. R. McClanahan, and W. E. Kiene. Effects of inorganic nutrients and organic matter on microbial euendolithic community composition and microbioerosion rates. Marine Ecology Progress Series, 392:1–15, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cohen:2004:HBC

[CSMvE04] A. L. Cohen, S. R. Smith, M. S. McCartney, and J. van Etten. How brain corals record climate: an integration of skeletal structure, growth and chemistry of Diploria labyrinthiformis from Bermuda. Marine Ecology Progress Series, 271: 147–158, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/271/m271p147.pdf.

Chockley:2008:PSU

[CSO08] B. R. Chockley, C. M. St. Mary, and C. W. Osenberg. Population sinks in the Upper Florida Keys: the importance of demographic variation in population dynamics of the marine shrimp Stenopus hispidus. Marine Ecology Progress Series, 360:135–145, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p135.pdf.

Company:2003:DTR

[CSP^{+}03] J. B. Company, F. Sardà, P. Puig, J. E. Cartes, and A. Palanques. Duration and timing of reproduction in decapod crustaceans of the NW Mediterranean continental margin: is there a general pattern? Marine Ecology Progress Series, 261:201–216, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p201.pdf.

Coll:2009:FWC

[CSPA09] M. Coll, A. Santojanni, I. Palomera, and E. Arneri. Food-web changes in the Adriatic Sea over the last three decades. Marine Ecology Progress Series, 381:17–37, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m381p017.pdf.

Calambokidis:2000:MDH

[CSR^{+}00] John Calambokidis, Gretchen H. Steiger, Kristin Rasmussen, Jorge Urbán R., Kenneth C. Balcomb, Paloma Ladrón de Guevara P., Mario Salinas Z., Jeff K. Jacobsen, C. Scott Baker, Louis M. Herman, Salvatore Cerchio, and James D. Darling. Migratory destinations of humpback whales that feed off California, Oregon and Washington. Marine Ecology Progress Series, 192:295–304, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p295.pdf.

Chiappone:2000:GPS

[CSS00] M. Chiappone, R. Sluka, and K. S. Sealey. Groupers (Pisces: Serranidae) in fished and protected areas of the Florida Keys, Bahamas and northern Caribbean. Marine Ecology Progress Series, 198:261–272, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/198/m198p261.pdf.

Childress:2008:SMD

[CST08] J. J. Childress, B. A. Seibel, and E. V. Thuesen. N-specific metabolic data are not relevant to the ‘visual interactions’ hypothesis concerning the depth-related declines in metabolic rates: Comment on Ikeda et al. (2006). Marine Ecology Progress Series, 373:187–191, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/373/m373p187.pdf. See

[ISYM06] and reply [Ike08].

Chaparro:2001:FBD

[CSTC01] O. R. Chaparro, C. J. Soto, R. J. Thompson, and I. I. Concha. Feeding behaviour during brooding in the oyster Ostrea chilensis: gut retention time in adults and larvae and potential use of larval faeces by adults. Marine Ecology Progress Series, 222:119–129, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p119.pdf.

Csaszar:2009:VAG

[CSvO09] N. B. M. Császár, F. O. Seneca, and M. J. H. van Oppen. Variation in antioxidant gene expression in the scleractinian coral Acropora millepora under laboratory thermal stress. Marine Ecology Progress Series, 392:93–102, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/392/
m392p093.pdf.

Chute:2001:PSB

[CT01] A. S. Chute and J. T. Turner. Plankton studies in Buzzards Bay, Massachusetts, USA. V. Ichthyoplankton, 1987 to 1993. Marine Ecology Progress Series, 224:45–54, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p045.pdf.

Cummings:2004:BRJ

[CT04] V. J. Cummings and S. F. Thrush. Behavioural response of juvenile bivalves to terrestrial sediment deposits: implications for post-disturbance recolonisation. Marine Ecology Progress Series, 278:179–191, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/278/m278p179.pdf.

Calliari:2005:FRS

[CT05] D. Calliari and P. Tiselius. Feeding and reproduction in a small calanoid copepod: Acartia clausi can compensate quality with quantity. Marine Ecology Progress Series, 298:241– 250, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p241.pdf.

Chen:2006:DGI

[CT06a] H. L. Chen and W. N. Tzeng. Daily growth increment formation in otoliths of Pacific tarpon Megalops cyprinoides during metamorphosis. Marine Ecology Progress Series, 312: 255–263, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/312/m312p255.pdf.

Cornelisen:2006:WFE

[CT06b] C. D. Cornelisen and F. I. M. Thomas. Water flow enhances ammonium and nitrate uptake in a seagrass community. Marine Ecology Progress Series, 312:1–13, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cornelisen:2009:PVF

[CT09] C. D. Cornelisen and F. I. M. Thomas. Prediction and validation of flow-dependent uptake of ammonium over a seagrasshardbottom community in Florida Bay. Marine Ecology Progress Series, 386:71–81, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m386p071.pdf.

Calcagno:2003:CBC

[CTA^{+}03] J. A. Calcagno, S. Thatje, K. Anger, G. A. Lovrich, and A. Kaffenberger. Changes in biomass and chemical composition during lecithotrophic larval development of the southern stone crab Paralomis granulosa. Marine Ecology Progress Series, 257:189–196, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p189.pdf.

Cresswell:2007:BAL

[CTB07] K. A. Cresswell, G. A. Tarling, and M. T. Burrows. Behaviour affects local-scale distributions of Antarctic krill around South Georgia. Marine Ecology Progress Series, 343: 193–206, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p193.pdf.

Compton:2009:RSA

[CTD^{+}09] T. J. Compton, T. A. Troost, J. Drent, C. Kraan, P. Bocher, J. Leyrer, A. Dekinga, and T. Piersma. Repeatable sediment associations of burrowing bivalves across six European tidal flat systems. Marine Ecology Progress Series, 382: 87–98, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/382/m382p087.pdf.

Crespo:2008:MCN

[CTFC08] B. G. Crespo, I. G. Teixeira, F. G. Figueiras, and C. G. Castro. Microplankton composition off NW Iberia at the end of the upwelling season: source areas of harmful dinoflagellate blooms. Marine Ecology Progress Series, 355:31–43, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p031.pdf.

Cummings:2003:TDI

[CTH^{+}03] V. Cummings, S. Thrush, J. Hewitt, A. Norkko, and S. Pickmere. Terrestrial deposits on intertidal sandflats: sediment characteristics as indicators of habitat suitability for recolonising macrofauna. Marine Ecology Progress Series, 253:39– 54, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p039.pdf.

Cummings:2001:VEL

[CTHF01] V. J. Cummings, S. F. Thrush, J. E. Hewitt, and G. A. Funnell. Variable effect of a large suspension-feeding bivalve on infauna: experimenting in a complex system. Marine Ecology Progress Series, 209:159–175, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/209/m209p159.pdf.

Carls:2003:MTO

[CTLR03] M. G. Carls, R. E. Thomas, M. R. Lilly, and S. D. Rice. Mechanism for transport of oil-contaminated groundwater into pink salmon redds. Marine Ecology Progress Series, 248:245–255, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/248/m248p245.pdf.

Chaparro:2002:FMG

[CTP02] O. R. Chaparro, R. J. Thompson, and S. V. Pereda. Feeding mechanisms in the gastropod Crepidula fecunda. Marine Ecology Progress Series, 234:171–181, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p171.pdf.

Compton:2008:DOR

[CTvdM^{+}08] Tanya J. Compton, Tineke A. Troost, Jaap van der Meer, Casper Kraan, Pieter J. C. Honkoop, Danny I. Rogers, Grant B. Pearson, Petra de Goeij, Pierrick Bocher, Marc S. S. Lavaleye, Jutta Leyrer, Mick G. Yates, Anne Dekinga, and Theunis Piersma. Distributional overlap rather than habitat differentiation characterizes co-occurrence of bivalves in intertidal soft sediment systems. Marine Ecology Progress Series, 373:25–35, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/373/m373p025.pdf. See comment

[BD11].

Copley:2003:SVR

[CTVP03] J. T. P. Copley, P. A. Tyler, C. L. Van Dover, and S. J. Philip. Spatial variation in the reproductive biology of Paralvinella palmiformis (Polychaeta: Alvinellidae) from a vent field on the Juan de Fuca Ridge. Marine Ecology Progress Series, 255:171–181, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p171.pdf.

Cebrian:2006:GFS

[CU06] E. Cebrian and M. J. Uriz. Grazing on fleshy seaweeds by sea urchins facilitates sponge Cliona viridis growth. Marine Ecology Progress Series, 323:83–89, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p083.pdf.

Chapman:2008:SVG

[CU08] M. G. Chapman and A. J. Underwood. Scales of variation of gastropod densities over multiple spatial scales: comparison of common and rare species. Marine Ecology Progress Series, 354:147–160, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p147.pdf.

Chapman:2008:EMD

[CUB08] M. G. Chapman, A. J. Underwood, and D. Blockley. Effects of method of deployment of artificial units of habitat on microgastropod colonization. Marine Ecology Progress Series, 366:49–57, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p049.pdf.

Cumming:2002:TIP

[Cum02] R. L. Cumming. Tissue injury predicts colony decline in reef-building corals. Marine Ecology Progress Series, 242: 131–141, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p131.pdf.

Curtis:2005:DSS

[CV05] J. M. R. Curtis and A. C. J. Vincent. Distribution of sympatric seahorse species along a gradient of habitat complexity in a seagrass-dominated community. Marine Ecology Progress Series, 291:81–91, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/291/m291p081.pdf.

Costello:2006:OBIb

[CV06] M. J. Costello and E. Vanden Berghe. ‘Ocean biodiversity informatics’: a new era in marine biology research and management. Marine Ecology Progress Series, 316:203–214, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m316p203.pdf.

Ceballos:2006:CEP

[CVÁM06] S. Ceballos, L. Viesca, and F. Álvarez-Marqués. Copepod egg production during highly productive late spring conditions: importance of freshly ingested food and lipid storage. Marine Ecology Progress Series, 317:171–186, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p171.pdf.

Costello:2006:I

[CVB06a] M. J. Costello, E. Vanden Berghe, and Howard I. Browman. Introduction. Marine Ecology Progress Series, 316: 201–202, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m316p201.pdf.

Costello:2006:OBIa

[CVB06b] M. J. Costello, E. Vanden Berghe, and Howard I. Browman. Ocean biodiversity informatics (OBI). Marine Ecology Progress Series, 316:201–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m316_TS.pdf.

Corzo:2009:EGM

[CvBGR09] A. Corzo, S. A. van Bergeijk, and E. García-Robledo. Effects of green macroalgal blooms on intertidal sediments: net metabolism and carbon and nitrogen contents. Marine Ecology Progress Series, 380:81–93, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m380p081.pdf.

Calbet:2003:RGI

[CVF^{+}03] A. Calbet, D. Vaqué, J. Felipe, M. Vila, M. Montserrat Sala, M. Alcaraz, and M. Estrada. Relative grazing impact of microzooplankton and mesozooplankton on a bloom of the toxic dinoflagellate Alexandrium minutum. Marine Ecology Progress Series, 259:303–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/259/m259p303.pdf.

Castro:2007:EPE

[CVF07] P. Castro, I. Valiela, and H. Freitas. Eutrophication in Portuguese estuaries evidenced by δ ^{15}N of macrophytes. Marine Ecology Progress Series, 351:43–51, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p043.pdf.

Christensen:2000:CNF

[CVK00] B. Christensen, A. Vedel, and E. Kristensen. Carbon and nitrogen fluxes in sediment inhabited by suspension-feeding (Nereis diversicolor) and non-suspension-feeding (N. virens) polychaetes. Marine Ecology Progress Series, 192:203–217, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/192/m192p203.pdf.

Cook:2009:CNC

[CVSM09] P. L. M. Cook, D. Van Oevelen, K. Soetaert, and J. J. Middelburg. Carbon and nitrogen cycling on intertidal mudflats of a temperate Australian estuary. IV. Inverse model analysis and synthesis. Marine Ecology Progress Series, 394:35–48, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m394p035.pdf.

Cummings:2009:TDC

[CVT09] V. Cummings, K. Vopel, and S. Thrush. Terrigenous deposits in coastal marine habitats: influences on sediment geochemistry and behaviour of post-settlement bivalves. Marine Ecology Progress Series, 383:173–185, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p173.pdf.

Clarke:2001:FBI

[CW01] K. R. Clarke and R. M. Warwick. A further biodiversity index applicable to species lists: variation in taxonomic distinctness. Marine Ecology Progress Series, 216:265–278, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p265.pdf.

Caldow:2003:PAI

[CW03a] C. Caldow and G. M. Wellington. Patterns of annual increment formation in otoliths of pomacentrids in the tropical western Atlantic: implications for population age-structure examination. Marine Ecology Progress Series, 265:185–195, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/265/m265p185.pdf.

Chinnery:2003:PTR

[CW03b] F. E. Chinnery and J. A. Williams. Photoperiod and temperature regulation of diapause egg production in Acartia bifilosa from Southampton Water. Marine Ecology Progress Series, 263:149–157, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p149.pdf.

Cheung:2005:ISM

[CW05] M. Cheung and W. X. Wang. Influence of subcellular metal compartmentalization in different prey on the transfer of metals to a predatory gastropod. Marine Ecology Progress Series, 286:155–166, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p155.pdf.

Chilvers:2009:DFS

[CW09] B. L. Chilvers and I. S. Wilkinson. Diverse foraging strategies in lactating New Zealand sea lions. Marine Ecology Progress Series, 378:299–308, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/378/m378p299.pdf.

Culverhouse:2006:AIA

[CWB^{+}06] P. F. Culverhouse, R. Williams, M. Benfield, P. R. Flood, A. F. Sell, M. G. Mazzocchi, I. Buttino, and M. Sieracki. Automatic image analysis of plankton: future perspectives. Marine Ecology Progress Series, 312:297–309, 2006. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m312p297.pdf.

Criales:2005:TSE

[CWBR05] M. M. Criales, J. Wang, J. A. Browder, and M. B. Robblee. Tidal and seasonal effects on transport of pink shrimp postlarvae. Marine Ecology Progress Series, 286:231–238, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/286/m286p231.pdf.

Chiswell:2003:TTS

[CWBS03] S. M. Chiswell, J. Wilkin, J. D. Booth, and B. Stanton. Trans-Tasman Sea larval transport: Is Australia a source for New Zealand rock lobsters? Marine Ecology Progress Series, 247:173–182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p173.pdf.

Catard:2000:EDA

[CWC00] A. Catard, H. Weimerskirch, and Y. Cherel. Exploitation of distant Antarctic waters and close shelf-break waters by white-chinned petrels rearing chicks. Marine Ecology Progress Series, 194:249–261, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/194/m194p249.pdf.

Cervino:2006:IBA

[CWCP^{+}06] J. M. Cervino, K. Winiarski-Cervino, S. W. Polson, T. Goreau, and G. W. Smith. Identification of bacteria associated with a disease affecting the marine sponge Ianthella basta in New Britain, Papua New Guinea. Marine Ecology Progress Series, 324:139–150, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/324/m324p139.pdf.

Chilvers:2005:SFA

[CWDG05] B. L. Chilvers, I. S. Wilkinson, P. J. Duignan, and N. J. Gemmell. Summer foraging areas for lactating New Zealand sea lions Phocarctos hookeri. Marine Ecology Progress Series, 304:235–247, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p235.pdf.

Cheal:2008:RRF

[CWE^{+}08] A. J. Cheal, S. K. Wilson, M. J. Emslie, A. M. Dolman, and H. Sweatman. Responses of reef fish communities to coral declines on the Great Barrier Reef. Marine Ecology Progress Series, 372:211–223, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/372/m372p211.pdf.

Carney:2005:RBK

[CWKE05] L. T. Carney, J. R. Waaland, T. Klinger, and K. Ewing. Restoration of the bull kelp Nereocystis luetkeana in nearshore rocky habitats. Marine Ecology Progress Series, 302:49–61, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p049.pdf.

Cheung:2007:IVG

[CWM^{+}07] W. W. L. Cheung, R. Watson, T. Morato, T. J. Pitcher, and D. Pauly. Intrinsic vulnerability in the global fish catch. Marine Ecology Progress Series, 333:1–12, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Clarke:2009:CSO

[CWM^{+}09] L. M. Clarke, B. D. Walther, S. B. Munch, S. R. Thorrold, and D. O. Conover. Chemical signatures in the otoliths of a coastal marine fish, Menidia menidia, from the northeastern United States: spatial and temporal differences. Marine Ecology Progress Series, 384:261–271, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/384/m384p261.pdf.

Callier:2006:SRS

[CWMD06] M. D. Callier, A. M. Weise, C. W. McKindsey, and G. Desrosiers. Sedimentation rates in a suspended mussel farm (Great-Entry Lagoon, Canada): biodeposit production and dispersion. Marine Ecology Progress Series, 322:129– 141, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p129.pdf.

Cubillos:2007:CMC

[CWN^{+}07] J. C. Cubillos, S. W. Wright, G. Nash, M. F. de Salas, B. Griffiths, B. Tilbrook, A. Poisson, and G. M. Hallegraeff. Calcification morphotypes of the coccolithophorid Emiliania huxleyi in the Southern Ocean: changes in 2001 to 2006 compared to historical data. Marine Ecology Progress Series, 348:47–54, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p047.pdf.

Culverhouse:2003:DEM

[CWR^{+}03] P. F. Culverhouse, R. Williams, B. Reguera, V. Herry, and S. González-Gil. Do experts make mistakes? A comparison of human and machine identification of dinoflagellates. Marine Ecology Progress Series, 247:17–25, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p017.pdf.

Ciutat:2006:ICC

[CWR06] A. Ciutat, J. Widdows, and J. W. Readman. Influence of cockle Cerastoderma edule bioturbation and tidal-current cycles on resuspension of sediment and polycyclic aromatic hydrocarbons. Marine Ecology Progress Series, 328:51–64, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/328/m328p051.pdf.

Cherel:2000:FFE

[CWT00] Yves Cherel, Henri Weimerskirch, and Colette Trouvé. Food and feeding ecology of the neritic-slope forager black-browed albatross and its relationships with commercial fisheries in Kerguelen waters. Marine Ecology Progress Series, 207:183– 199, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/207/m207p183.pdf.

Campbell:2001:GDR

[CWT^{+}01] R. G. Campbell, M. M. Wagner, G. J. Teegarden, C. A. Boudreau, and E. G. Durbin. Growth and development rates of the copepod Calanus finmarchicus reared in the laboratory. Marine Ecology Progress Series, 221:161–183, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p161.pdf.

Chintala:2006:CGD

[CWT06] M. M. Chintala, C. Wigand, and G. Thursby. Comparison of Geukensia demissa populations in Rhode Island fringe salt marshes with varying nitrogen loads. Marine Ecology Progress Series, 320:101–108, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/320/m320p101.pdf.

Copley:2006:SZR

[CY06] J. T. P. Copley and C. M. Young. Seasonality and zonation in the reproductive biology and population structure of the shrimp Alvinocaris stactophila (Caridea: Alvinocarididae) at a Louisiana Slope cold seep. Marine Ecology Progress Series, 315:199–209, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p199.pdf.

Cheung:2009:APB

[CYC^{+}09] S. G. Cheung, F. Y. Yang, J. M. Y. Chiu, C. C. Liu, and P. K. S. Shin. Anti-predator behaviour in the green-lipped mussel Perna viridis: byssus thread production depends on the mussel’s position in clump. Marine Ecology Progress Series, 378:145–151, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p145.pdf.

Chiu:2008:NEI

[CZW^{+}08] J. M. Y. Chiu, R. Zhang, H. Wang, V. Thiyagarajan, and P. Y. Qian. Nutrient effects on intertidal community: from bacteria to invertebrates. Marine Ecology Progress Series, 358: 41–50, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/358/m358p041.pdf.

Duarte:2000:RMP

[DAA00] C. M. Duarte, S. Agustí, and N. S. R. Agawin. Response of a Mediterranean phytoplankton community to increased nutrient inputs: a mesocosm experiment. Marine Ecology Progress Series, 195:61–70, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/195/m195p061.pdf.

Duffy-Anderson:2002:CSL

[DABC02] J. T. Duffy-Anderson, K. M. Bailey, and L. Ciannelli. Consequences of a superabundance of larval walleye pollock Theragra chalcogramma in the Gulf of Alaska in 1981. Marine Ecology Progress Series, 243:179–190, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p179.pdf.

Duarte:2000:ENS

[DAG^{+}00] C. M. Duarte, S. Agustí, J. M. Gasol, D. Vaqué, and E. Vazquez-Dominguez. Effect of nutrient supply on the biomass structure of planktonic communities: an experimental test on a Mediterranean coastal community. Marine Ecology Progress Series, 206:87–95, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p087.pdf.

Deborde:2008:ESD

[DAM^{+}08] J. Deborde, G. Abril, A. Mouret, D. Jézéquel, G. Thouzeau, J. Clavier, G. Bachelet, and P. Anschutz. Effects of seasonal dynamics in a Zostera noltii meadow on phosphorus and iron cycles in a tidal mudflat (Arcachon Bay, France). Marine Ecology Progress Series, 355:59–71, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p059.pdf.

DAmours:2008:LEE

[DAMJ08] O. D’Amours, P. Archambault, C. W. McKindsey, and L. E. Johnson. Local enhancement of epibenthic macrofauna by aquaculture activities. Marine Ecology Progress Series, 371: 73–84, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/371/m371p073.pdf.

Daskalov:2002:ODT

[Das02] G. M. Daskalov. Overfishing drives a trophic cascade in the Black Sea. Marine Ecology Progress Series, 225:53–63, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p053.pdf.

Daskalov:2003:LTC

[Das03] G. M. Daskalov. Long-term changes in fish abundance and environmental indices in the Black Sea. Marine Ecology Progress Series, 255:259–270, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/255/m255p259.pdf.

Doval:2001:DSO

[DÁSG^{+}01] M. D. Doval, X. A. Álvarez-Salgado, J. M. Gasol, L. M. Lorenzo, I. Mirón, F. G. Figueiras, and C. Pedrós-Alió. Dissolved and suspended organic carbon in the Atlantic sector of the Southern Ocean. Stock dynamics in upper ocean waters. Marine Ecology Progress Series, 223:27–38, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p027.pdf.

Davis:2000:SSP

[Dav00a] J. L. D. Davis. Spatial and seasonal patterns of habitat partitioning in a guild of southern California tidepool fishes. Marine Ecology Progress Series, 196:253–268, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p253.pdf.

Davoren:2000:VFR

[Dav00b] G. K. Davoren. Variability in foraging in response to changing prey distributions in rhinoceros auklets. Marine Ecology Progress Series, 198:283–291, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/198/m198p283.pdf.

Day:2002:EBG

[DB02] E. Day and G. M. Branch. Effects of benthic grazers on microalgal communities of morphologically different encrusting corallines: implications for abalone recruits. Marine Ecology Progress Series, 244:95–103, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/244/m244p095.pdf.

Depczynski:2003:RCR

[DB03] M. Depczynski and D. R. Bellwood. The role of cryptobenthic reef fishes in coral reef trophodynamics. Marine Ecology Progress Series, 256:183–191, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p183.pdf.

Depczynski:2005:WES

[DB05] M. Depczynski and D. R. Bellwood. Wave energy and spatial variability in community structure of small cryptic coral reef fishes. Marine Ecology Progress Series, 303:283– 293, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/303/m303p283.pdf.

Dekker:2007:LTL

[DB07a] R. Dekker and J. J. Beukema. Long-term and large-scale variability in productivity of the tellinid bivalve Macoma balthica on Wadden Sea tidal flats. Marine Ecology Progress Series, 337:117–134, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p117.pdf.

Dunstan:2007:HFC

[DB07b] P. K. Dunstan and N. J. Bax. How far can marine species go? Influence of population biology and larval movement on future range limits. Marine Ecology Progress Series, 344: 15–28, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/344/m344p015.pdf.

Dubois:2005:PCP

[DBCB05] S. Dubois, L. Barillé, B. Cognie, and P. G. Beninger. Particle capture and processing mechanisms in Sabellaria alveolata (Polychaeta: Sabellariidae). Marine Ecology Progress Series, 301:159–171, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p159.pdf. See comment

[RN06].

Dubois:2006:FMP

[DBCB06] S. Dubois, L. Barillé, B. Cognie, and P. Beninger. Feeding mechanism of the polychaete Sabellaria alveolata revisited: reply to Riisgård & Nielsen (2006). Marine Ecology Progress Series, 328:307–311, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p307.pdf.

Devilla:2005:IAB

[DBD^{+}05] R. A. Devilla, M. T. Brown, M. Donkin, G. A. Tarran, J. Aiken, and J. W. Readman. Impact of antifouling booster biocides on single microalgal species and on a natural marine phytoplankton community. Marine Ecology Progress Series, 286:1–12, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p001.pdf.

Dean:2002:FLR

[DBF^{+}02] T. A. Dean, J. L. Bodkin, A. K. Fukuyama, S. C. Jewett, D. H. Monson, C. E. O’Clair, and G. R. VanBlaricom. Food limitation and the recovery of sea otters following the ‘Exxon Valdez’ oil spill. Marine Ecology Progress Series, 241:255–270, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p255.pdf.

Danovaro:2009:DDS

[DBG^{+}09] R. Danovaro, S. Bianchelli, C. Gambi, M. Mea, and D. Zeppilli. α-, β-, γ-, δ- and ϵ-diversity of deep-sea nematodes in canyons and open slopes of Northeast Atlantic and Mediterranean margins. Marine Ecology Progress Series, 396:197– 209, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m396p197.pdf.

deBaar:2008:ECR

[dBGLT08] H. J. W. de Baar, L. J. A. Gerringa, P. Laan, and K. R. Timmermans. Efficiency of carbon removal per added iron in ocean iron fertilization. Marine Ecology Progress Series, 364:269–282, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m364p269.pdf.

Daunt:2002:FSB

[DBH^{+}02] F. Daunt, S. Benvenuti, M. P. Harris, L. Dall’Antonia, D. A. Elston, and S. Wanless. Foraging strategies of the black-legged kittiwake Rissa tridactyla at a North Sea colony: evidence for a maximum foraging range. Marine Ecology Progress Series, 245:239–247, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p239.pdf.

Decker:2007:PDS

[DBH^{+}07] M. B. Decker, C. W. Brown, R. R. Hood, J. E. Purcell, T. F. Gross, J. C. Matanoski, R. O. Bannon, and E. M. Setzler-Hamilton. Predicting the distribution of the scyphomedusa Chrysaora quinquecirrha in Chesapeake Bay. Marine Ecology Progress Series, 329:99–113, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/329/m329p099.pdf.

Dean:2000:CSU

[DBJ^{+}00] T. A. Dean, J. L. Bodkin, S. C. Jewett, D. H. Monson, and D. Jung. Changes in sea urchins and kelp following a reduction in sea otter density as a result of the Exxon Valdez oil spill. Marine Ecology Progress Series, 199:281–291, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p281.pdf.

deBeer:2000:MSL

[dBKSV00] D. de Beer, M. Kühl, N. Stambler, and L. Vaki. A microsensor study of light enhanced ca2+ uptake and photosynthesis in the reef-building hermatypic coral Favia sp. Marine Ecology Progress Series, 194:75–85, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/194/m194p075.pdf.

deBie:2002:FCN

[dBMSL02] M. J. M. de Bie, J. J. Middelburg, M. Starink, and H. J. Laanbroek. Factors controlling nitrous oxide at the microbial community and estuarine scale. Marine Ecology Progress Series, 240:1–9, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p001.pdf.

DiFranco:2009:EET

[DBN^{+}09] A. Di Franco, S. Bussotti, A. Navone, P. Panzalis, and P. Guidetti. Evaluating effects of total and partial restrictions to fishing on Mediterranean rocky-reef fish assemblages. Marine Ecology Progress Series, 387:275–285, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p275.pdf.

Dodds:2009:LBR

[DBOR09] L. A. Dodds, K. D. Black, H. Orr, and J. M. Roberts. Lipid biomarkers reveal geographical differences in food supply to the cold-water coral Lophelia pertusa (Scleractinia). Marine Ecology Progress Series, 397:113–124, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p113.pdf.

Decker:2004:ELD

[DBP04] M. B. Decker, D. L. Breitburg, and J. E. Purcell. Effects of low dissolved oxygen on zooplankton predation by the ctenophore Mnemiopsis leidyi. Marine Ecology Progress Series, 280:163–172, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p163.pdf.

Decottignies:2007:ENF

[DBR^{+}07] P. Decottignies, P. G. Beninger, Y. Rincé, R. J. Robins, and P. Riera. Exploitation of natural food sources by two sympatric, invasive suspension-feeders: Crassostrea gigas and Crepidula fornicata. Marine Ecology Progress Series, 334: 179–192, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/334/m334p179.pdf.

deBrouwer:2001:STD

[dBS01] J. F. C. de Brouwer and L. J. Stal. Short-term dynamics in microphytobenthos distribution and associated extracellular carbohydrates in surface sediments of an intertidal mudflat. Marine Ecology Progress Series, 218:33–44, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p033.pdf.

Dorman:2005:PBE

[DBS05] J. G. Dorman, S. M. Bollens, and A. M. Slaughter. Population biology of euphausiids off northern California and effects of short time-scale wind events on Euphausia pacifica. Marine Ecology Progress Series, 288:183–198, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p183.pdf.

Diekmann:2002:NHC

[DBT^{+}02a] O. E. Diekmann, R. P. M. Bak, L. Tonk, W. T. Stam, and J. L. Olsen. No habitat correlation of zooxanthellae in the coral genus Madracis on a Curaçao reef. Marine Ecology Progress Series, 227:221–232, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/227/m227p221.pdf.

Drouin:2002:LTP

[DBT02b] C. A. Drouin, E. Bourget, and R. Tremblay. Larval transport processes of barnacle larvae in the vicinity of the interface between two genetically different populations of Semibalanus balanoides. Marine Ecology Progress Series, 229:165– 172, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/229/m229p165.pdf.

deBruyn:2009:BFS

[dBTO^{+}09] P. J. N. de Bruyn, C. A. Tosh, W. C. Oosthuizen, M. N. Bester, and J. P. Y. Arnould. Bathymetry and frontal system interactions influence seasonal foraging movements of lactating subantarctic fur seals from Marion Island. Marine Ecology Progress Series, 394:263–276, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p263.pdf.

Duggan:2006:IRA

[DBvOM06] I. C. Duggan, S. A. Bailey, C. D. A. van Overdijk, and H. J. MacIsaac. Invasion risk of active and diapausing invertebrates from residual ballast in ships entering Chesapeake Bay. Marine Ecology Progress Series, 324:57–66, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p057.pdf.

Defeo:2002:MPD

[DC02] O. Defeo and R. S. Cardoso. Macroecology of population dynamics and life history traits of the mole crab Emerita brasiliensis in Atlantic sandy beaches of South America. Marine Ecology Progress Series, 239:169–179, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p169.pdf.

Danil:2006:HBS

[DC06] K. Danil and S. J. Chivers. Habitat-based spatial and temporal variability in life history characteristics of female common dolphins Delphinus delphis in the eastern tropical Pacific. Marine Ecology Progress Series, 318:277–286, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p277.pdf.

Dimond:2007:TVS

[DC07a] J. Dimond and E. Carrington. Temporal variation in the symbiosis and growth of the temperate scleractinian coral Astrangia poculata. Marine Ecology Progress Series, 348:161– 172, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/348/m348p161.pdf.

Durbin:2007:RGP

[DC07b] E. G. Durbin and R. G. Campbell. Reassessment of the gut pigment method for estimating in situ zooplankton ingestion. Marine Ecology Progress Series, 331:305–307, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p305.pdf.

Drago:2009:HDC

[DCA^{+}09] M. Drago, E. A. Crespo, A. Aguilar, L. Cardona, N. García, S. L. Dans, and N. Goodall. Historic diet change of the South American sea lion in Patagonia as revealed by isotopic analysis. Marine Ecology Progress Series, 384:273– 286, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p273.pdf.

Dela-Cruz:2002:TAP

[DCAL^{+}02] J. Dela-Cruz, P. Ajani, R. Lee, T. Pritchard, and I. Suthers. Temporal abundance patterns of the red tide dinoflagellate Noctiluca scintillans along the southeast coast of Australia. Marine Ecology Progress Series, 236:75–88, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p075.pdf.

Dickey-Collas:2009:VTF

[DCBvBE09] M. Dickey-Collas, L. J. Bolle, J. K. L. van Beek, and P. L. A. Erftemeijer. Variability in transport of fish eggs and larvae. II. Effects of hydrodynamics on the transport of downs herring larvae. Marine Ecology Progress Series, 390:183– 194, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p183.pdf.

Durbin:2003:IVP

[DCC^{+}03] E. G. Durbin, R. G. Campbell, M. C. Casas, M. D. Ohman, B. Niehoff, J. Runge, and M. Wagner. Interannual variation in phytoplankton blooms and zooplankton productivity and abundance in the Gulf of Maine during winter. Marine Ecology Progress Series, 254:81–100, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/254/m254p081.pdf.

Davis:2003:FAC

[DCD^{+}03] S. E. Davis III, D. L. Childers, J. W. Day, Jr., D. T. Rudnick, and F. H. Sklar. Factors affecting the concentration and flux of materials in two southern Everglades mangrove wetlands. Marine Ecology Progress Series, 253:85–96, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p085.pdf.

Diekmann:2005:PGZ

[DCF^{+}05] O. E. Diekmann, J. A. Coyer, J. Ferreira, J. L. Olsen, W. T. Stam, G. A. Pearson, and E. A. Serrão. Population genetics of Zostera noltii along the west Iberian coast: consequences of small population size, habitat discontinuity and near-shore currents. Marine Ecology Progress Series, 290: 89–96, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/290/m290p089.pdf.

deCaralt:2002:CBT

[dCLLT^{+}02] S. de Caralt, S. López-Legentil, I. Tarjuelo, M. J. Uriz, and X. Turon. Contrasting biological traits of Clavelina lepadiformis (Ascidiacea) populations from inside and outside harbours in the western Mediterranean. Marine Ecology Progress Series, 244:125–137, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/244/m244p125.pdf.

Doucette:2006:PSP

[DCM^{+}06] G. J. Doucette, A. D. Cembella, J. L. Martin, J. Michaud, T. V. N. Cole, and R. M. Rolland. Paralytic shellfish poisoning (PSP) toxins in North Atlantic right whales Eubalaena glacialis and their zooplankton prey in the Bay of Fundy, Canada. Marine Ecology Progress Series, 306:303– 313, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/306/m306p303.pdf.

Doi:2008:AMC

[DCO^{+}08] H. Doi, K. H. Chang, Y. Obayashi, M. Yoshihara, M. Shime, T. Yamamoto, Y. Nishibe, and S. Nakano. Attached microalgae contribute to planktonic food webs in bays with fish and pearl oyster farms. Marine Ecology Progress Series, 353:107–113, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/353/m353p107.pdf.

Dans:2008:DDT

[DCP^{+}08] S. L. Dans, E. A. Crespo, S. N. Pedraza, M. Degrati, and G. V. Garaffo. Dusky dolphin and tourist interaction: effect on diurnal feeding behavior. Marine Ecology Progress Series, 369:287–296, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p287.pdf.

Dubois:2007:DRS

[DCRT07] S. Dubois, T. Comtet, C. Retière, and E. Thiébaut. Distribution and retention of Sabellaria alveolata larvae (Polychaeta: Sabellariidae) in the Bay of Mont-Saint-Michel, France. Marine Ecology Progress Series, 346:243–254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p243.pdf.

Danis:2003:DPU

[DCT^{+}03] B. Danis, O. Cotret, J. L. Teyssié, S. W. Fowler, P. Bustamante, and M. Warnau. Delineation of PCB uptake pathways in a benthic sea star using a radiolabelled congener. Marine Ecology Progress Series, 253:155–163, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p155.pdf.

Danis:2004:CPU

[DCT^{+}04] B. Danis, O. Cotret, J. L. Teyssié, S. W. Fowler, and M. Warnau. Coplanar PCB 77 uptake kinetics in the sea star Asterias rubens and subsequent effects on reactive oxygen species (ROS) production and levels of cytochrome P450 immunopositive proteins (CYP1A-IPP). Marine Ecology Progress Series, 279:117–128, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p117.pdf.

Dunham:2001:FPG

[DD01] J. S. Dunham and D. A. Duffus. Foraging patterns of gray whales in central Clayoquot Sound, British Columbia, Canada. Marine Ecology Progress Series, 223:299–310, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p299.pdf.

Desroy:2004:ISP

[DD04] N. Desroy and L. Denis. Influence of spring phytodetritus sedimentation on intertidal macrozoobenthos in the eastern English Channel. Marine Ecology Progress Series, 270: 41–53, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/270/m270p041.pdf.

Davenport:2005:ESH

[DD05] J. Davenport and J. L. Davenport. Effects of shore height, wave exposure and geographical distance on thermal niche width of intertidal fauna. Marine Ecology Progress Series, 292:41–50, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p041.pdf.

Dudas:2006:RED

[DD06] S. E. Dudas and J. F. Dower. Reproductive ecology and dispersal potential of varnish clam Nuttallia obscurata, a recent invader in the Northeast Pacific Ocean. Marine Ecology Progress Series, 320:195–205, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p195.pdf.

Davenport:2007:ITT

[DD07] J. Davenport and J. L. Davenport. Interaction of thermal tolerance and oxygen availability in the eurythermal gastropods Littorina littorea and Nucella lapillus. Marine Ecology Progress Series, 332:167–170, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p167.pdf.

Dvoretsky:2009:MPS

[DD09a] V. G. Dvoretsky and A. G. Dvoretsky. Morphological plasticity in the small copepod Oithona similis in the Barents and White Seas. Marine Ecology Progress Series, 385:165– 178, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/385/m385p165.pdf.

Dvoretsky:2009:SVR

[DD09b] V. G. Dvoretsky and A. G. Dvoretsky. Spatial variations in reproductive characteristics of the small copepod Oithona similis in the Barents Sea. Marine Ecology Progress Series, 386:133–146, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/386/m386p133.pdf.

Dias:2008:ECS

[DDD08] G. M. Dias, C. G. M. Delboni, and L. F. L. Duarte. Effects of competition on sexual and clonal reproduction of a tunicate: the importance of competitor identity. Marine Ecology Progress Series, 362:149–156, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p149.pdf.

Danovaro:2001:BOM

[DDF01] R. Danovaro, A. Dell’Anno, and M. Fabiano. Bioavailability of organic matter in the sediments of the Porcupine Abyssal Plain, northeastern Atlantic. Marine Ecology Progress Series, 220:25–32, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p025.pdf.

DiNitto:2008:DTM

[DDGK^{+}08] D. Di Nitto, F. Dahdouh-Guebas, J. G. Kairo, H. Decleir, and N. Koedam. Digital terrain modelling to investigate the effects of sea level rise on mangrove propagule establishment. Marine Ecology Progress Series, 356:175–188, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/356/
m356p175.pdf.

Dobretsov:2006:ADA

[DDHQ06] S. Dobretsov, H. U. Dahms, T. Harder, and P. Y. Qian. Allelochemical defense against epibiosis in the macroalga Caulerpa racemosa var. turbinata. Marine Ecology Progress Series, 318:165–175, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/318/m318p165.pdf.

Denisenko:2003:MPS

[DDL^{+}03] S. G. Denisenko, N. V. Denisenko, K. K. Lehtonen, A. B. Andersin, and A. O. Laine. Macrozoobenthos of the Pechora Sea (SE Barents Sea): community structure and spatial distribution in relation to environmental conditions. Marine Ecology Progress Series, 258:109–123, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/258/m258p109.pdf.

Desclaux:2004:CCE

[DdMB04] C. Desclaux, X. de Montaudouin, and G. Bachelet. Cockle Cerastoderma edule population mortality: role of the digenean parasite Himasthla quissetensis. Marine Ecology Progress Series, 279:141–150, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p141.pdf.

Duijnstee:2005:PDB

[DdNEvdZ05] I. A. P. Duijnstee, L. J. de Nooijer, S. R. Ernst, and G. J. van der Zwaan. Population dynamics of benthic shallowwater Foraminifera: effects of a simulated marine snow event. Marine Ecology Progress Series, 285:29–42, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p029.pdf.

Dworjanyn:2006:CMA

[DdNS06] S. A. Dworjanyn, R. de Nys, and P. D. Steinberg. Chemically mediated antifouling in the red alga Delisea pulchra. Marine Ecology Progress Series, 318:153–163, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p153.pdf.

Daneri:2000:PPC

[DDQ^{+}00] G. Daneri, V. Dellarossa, R. Quiñones, B. Jacob, P. Montero, and O. Ulloa. Primary production and community respiration in the Humboldt Current System off Chile and associated oceanic areas. Marine Ecology Progress Series, 197: 41–49, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/197/m197p041.pdf.

Dobretsov:2004:AAA

[DDQ04] S. Dobretsov, H. U. Dahms, and P. Y. Qian. Antilarval and antimicrobial activity of waterborne metabolites of the sponge Callyspongia (Euplacella) pulvinata: evidence of allelopathy. Marine Ecology Progress Series, 271:133–146, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p133.pdf.

Dekshenieks:2001:TSO

[DDS^{+}01] M. M. Dekshenieks, P. L. Donaghay, J. M. Sullivan, J. E. B. Rines, T. R. Osborn, and M. S. Twardowski. Temporal and spatial occurrence of thin phytoplankton layers in relation to physical processes. Marine Ecology Progress Series, 223:61–71, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p061.pdf.

DeMesel:2006:RND

[DDS^{+}06] I. De Mesel, S. Derycke, J. Swings, M. Vincx, and T. Moens. Role of nematodes in decomposition processes: Does withintrophic group diversity matter? Marine Ecology Progress Series, 321:157–166, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p157.pdf.

Douglass:2007:NVC

[DDSR07] J. G. Douglass, J. E. Duffy, A. C. Spivak, and J. P. Richardson. Nutrient versus consumer control of community structure in a Chesapeake Bay eelgrass habitat. Marine Ecology Progress Series, 348:71–83, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p071.pdf.

Dobretsov:2005:CCE

[DDTQ05] S. Dobretsov, H. U. Dahms, M. Y. Tsoi, and P. Y. Qian. Chemical control of epibiosis by Hong Kong sponges: the effect of sponge extracts on micro- and macrofouling communities. Marine Ecology Progress Series, 297:119–129, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p119.pdf.

DAngelo:2008:BLR

[DDV^{+}08] C. D’Angelo, A. Denzel, A. Vogt, M. V. Matz, F. Oswald, A. Salih, G. U. Nienhaus, and J. Wiedenmann. Blue light regulation of host pigment in reef-building corals. Marine Ecology Progress Series, 364:97–106, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m364p097.pdf.

Dahlgren:2001:STV

[DE01] C. P. Dahlgren and D. B. Eggleston. Spatio-temporal variability in abundance, size and microhabitat associations of early juvenile Nassau grouper Epinephelus striatus in an offreef nursery system. Marine Ecology Progress Series, 217: 145–156, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/217/m217p145.pdf.

Drew:2008:JFD

[DE08] C. A. Drew and D. B. Eggleston. Juvenile fish densities in Florida Keys mangroves correlate with landscape characteristics. Marine Ecology Progress Series, 362:233–243, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/362/m362p233.pdf.

Daleo:2006:NEA

[DEAI06] P. Daleo, M. Escapa, J. Alberti, and O. Iribarne. Negative effects of an autogenic ecosystem engineer: interactions between coralline turf and an ephemeral green alga. Marine Ecology Progress Series, 315:67–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/315/m315p067.pdf.

Denman:2008:CCO

[Den08] K. L. Denman. Climate change, ocean processes and ocean iron fertilization. Marine Ecology Progress Series, 364:219– 225, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m364p219.pdf.

DeRobertis:2001:ERA

[DER01] A. De Robertis, K. Eiane, and G. H. Rau. Eat and run: anoxic feeding and subsequent aerobic recovery by Orchomene obtusus in Saanich Inlet, British Columbia, Canada. Marine Ecology Progress Series, 219:221–227, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p221.pdf.

Dittel:2000:CNS

[DES^{+}00] A. I. Dittel, C. E. Epifanio, S. M. Schwalm, M. S. Fantle, and M. L. Fogel. Carbon and nitrogen sources for juvenile blue crabs Callinectes sapidus in coastal wetlands. Marine Ecology Progress Series, 194:103–112, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/194/m194p103.pdf.

Duggins:2001:IRM

[DESK01] D. Duggins, J. E. Eckman, C. E. Siddon, and T. Klinger. Interactive roles of mesograzers and current flow in survival of kelps. Marine Ecology Progress Series, 223:143–155, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p143.pdf.

Duggins:2003:PMB

[DESK03] D. O. Duggins, J. E. Eckman, C. E. Siddon, and T. Klinger. Population, morphometric and biomechanical studies of three understory kelps along a hydrodynamic gradient. Marine Ecology Progress Series, 265:57–76, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p057.pdf.

Dupont:2007:LGF

[DEV07] L. Dupont, C. Ellien, and F. Viard. Limits to gene flow in the slipper limpet Crepidula fornicata as revealed by microsatellite data and a larval dispersal model. Marine Ecology Progress Series, 349:125–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/349/m349p125.pdf.

Duijnstee:2003:EAV

[DEvdZ03] I. A. P. Duijnstee, S. R. Ernst, and G. J. van der Zwaan. Effect of anoxia on the vertical migration of benthic Foraminifera. Marine Ecology Progress Series, 246:85–94, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p085.pdf.

Danovaro:2002:MVZ

[DF02] R. Danovaro and S. Fraschetti. Meiofaunal vertical zonation on hard-bottoms: comparison with soft-bottom meiofauna. Marine Ecology Progress Series, 230:159–169, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p159.pdf.

DeMartini:2004:SPE

[DF04] E. E. DeMartini and A. M. Friedlander. Spatial patterns of endemism in shallow-water reef fish populations of the Northwestern Hawaiian Islands. Marine Ecology Progress Series, 271:281–296, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/271/m271p281.pdf.

Dufour:2006:SAT

[DF06] S. C. Dufour and H. Felbeck. Symbiont abundance in thyasirids (Bivalvia) is related to particulate food and sulphide availability. Marine Ecology Progress Series, 320:185– 194, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p185.pdf.

Deudero:2005:DDD

[DFCH05] S. Deudero, A. Frau, M. Cerda, and H. Hampelm. Distribution and densities of the decapod crab Percnon gibbesi, an invasive Grapsidae, in western Mediterranean waters. Marine Ecology Progress Series, 285:151–156, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/285/m285p151.pdf.

DeMartini:2005:SSC

[DFH05] E. E. DeMartini, A. M. Friedlander, and S. R. Holzwarth. Size at sex change in protogynous labroids, prey body size distributions, and apex predator densities at NW Hawaiian atolls. Marine Ecology Progress Series, 297:259–271, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p259.pdf.

Dempster:2005:VVW

[DFJSJ^{+}05] T. Dempster, D. Fernandez-Jover, P. Sanchez-Jerez, F. Tuya, J. Bayle-Sempere, A. Boyra, and R. J. Haroun. Vertical variability of wild fish assemblages around sea-cage fish farms: implications for management. Marine Ecology Progress Series, 304:15–29, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p015.pdf.

DellAnno:2000:EHP

[DFMD00] A. Dell’Anno, M. Fabiano, M. L. Mei, and R. Danovaro. Enzymatically hydrolysed protein and carbohydrate pools in deep-sea sediments: estimates of the potentially bioavailable fraction and methodological considerations. Marine Ecology Progress Series, 196:15–23, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/196/m196p015.pdf.

Dahl:2003:LSI

[DFPG^{+}03] T. M. Dahl, S. Falk-Petersen, G. W. Gabrielsen, J. R. Sargent, H. Hop, and R. M. Millar. Lipids and stable isotopes in common eider, black-legged kittiwake and northern fulmar: a trophic study from an Arctic fjord. Marine Ecology Progress Series, 256:257–269, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/256/m256p257.pdf.

Dullo:2008:CWC

[DFR08] Wolf-Christian Dullo, Sascha Flögel, and Andres Rüggeberg. Cold-water coral growth in relation to the hydrography of the Celtic and Nordic European continental margin. Marine Ecology Progress Series, 371:165–176, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/371/
m371p165.pdf.

DeMartini:2008:DFA

[DFSS08] E. E. DeMartini, A. M. Friedlander, S. A. Sandin, and E. Sala. Differences in fish-assemblage structure between fished and unfished atolls in the northern Line Islands, central Pacific. Marine Ecology Progress Series, 365:199–215, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p199.pdf.

Dolbeth:2007:BMI

[DFT^{+}07] M. Dolbeth, Ó. Ferreira, H. Teixeira, J. C. Marques, J. A. Dias, and M. A. Pardal. Beach morphodynamic impact on a macrobenthic community along a subtidal depth gradient. Marine Ecology Progress Series, 352:113–124, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/352/
m352p113.pdf.

DeTroch:2001:ABD

[DFV01] M. De Troch, F. Fiers, and M. Vincx. Alpha and beta diversity of harpacticoid copepods in a tropical seagrass bed: the relation between diversity and species’ range size distribution. Marine Ecology Progress Series, 215:225–236, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p225.pdf.

Davis:2003:CWS

[DFW^{+}03] R. W. Davis, L. A. Fuiman, T. M. Williams, M. Horning, and W. Hagey. Classification of Weddell seal dives based on 3-dimensional movements and video-recorded observations. Marine Ecology Progress Series, 264:109–122, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p109.pdf.

Dai:2000:RIG

[DFY00] C. F. Dai, T. Y. Fan, and J. K. Yu. Reproductive isolation and genetic differentiation of a scleractinian coral Mycedium elephantotus. Marine Ecology Progress Series, 201:179–187, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/201/m201p179.pdf.

Defeo:2005:MHS

[DG05] O. Defeo and J. Gómez. Morphodynamics and habitat safety in sandy beaches: life-history adaptations in a supralittoral amphipod. Marine Ecology Progress Series, 293:143– 153, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p143.pdf.

Durand:2002:FHM

[DGB^{+}02] J. P. Durand, F. Goudard, C. Barbot, J. Pieri, S. W. Fowler, and O. Cotret. Ferritin and hemocyanin: 210Po molecular traps in marine fish, oyster and lobster. Marine Ecology Progress Series, 233:199–205, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/233/m233p199.pdf.

Degerholm:2008:SSF

[DGBS08] J. Degerholm, K. Gundersen, B. Bergman, and E. Söderbäck. Seasonal significance of N_{2} fixation in coastal and offshore waters of the northwestern Baltic Sea. Marine Ecology Progress Series, 360:73–84, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p073.pdf.

Dillane:2005:GVL

[DGC^{+}05a] E. Dillane, P. Galvin, J. Coughlan, M. Lipinski, and T. F. Cross. Genetic variation in the lesser flying squid Todaropsis eblanae (Cephalopoda, Ommastrephidae) in east Atlantic and Mediterranean waters. Marine Ecology Progress Series, 292:225–232, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/292/m292p225.pdf.

Dorenbosch:2005:IPS

[DGC^{+}05b] M. Dorenbosch, M. G. G. Grol, M. J. A. Christianen, I. Nagelkerken, and G. van der Velde. Indo–Pacific seagrass beds and mangroves contribute to fish density and diversity on adjacent coral reefs. Marine Ecology Progress Series, 302:63–76, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p063.pdf.

Duci:2009:ISB

[DGCM09] A. Duci, E. Giacomello, N. Chimento, and C. Mazzoldi. Intertidal and subtidal blennies: assessment of their habitat through individual and nest distribution. Marine Ecology Progress Series, 383:273–283, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/383/m383p273.pdf.

Dorenbosch:2009:PAP

[DGdG^{+}09] M. Dorenbosch, M. G. G. Grol, A. de Groene, G. van der Velde, and I. Nagelkerken. Piscivore assemblages and predation pressure affect relative safety of some back-reef habitats for juvenile fish in a Caribbean bay. Marine Ecology Progress Series, 379:181–196, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m379p181.pdf.

Deagle:2008:VFB

[DGH08] B. E. Deagle, N. J. Gales, and M. A. Hindell. Variability in foraging behaviour of chick-rearing macaroni penguins Eudyptes chrysolophus and its relation to diet. Marine Ecology Progress Series, 359:295–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/359/m359p295.pdf.

DiTullio:2003:PAS

[DGJ^{+}03] G. R. DiTullio, M. E. Geesey, D. R. Jones, K. L. Daly, L. Campbell, and W. O. Smith, Jr. Phytoplankton assemblage structure and primary productivity along 170^{o}W in the South Pacific Ocean. Marine Ecology Progress Series, 255:55–80, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p055.pdf.

Defeo:2001:TSE

[DGL01] O. Defeo, J. Gomez, and D. Lercari. Testing the swash exclusion hypothesis in sandy beach populations: the mole crab Emerita brasiliensis in Uruguay. Marine Ecology Progress Series, 212:159–170, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p159.pdf.

Dolan:2000:DEM

[DGM00] J. R. Dolan, C. L. Gallegos, and A. Moigis. Dilution effects on microzooplankton in dilution grazing experiments. Marine Ecology Progress Series, 200:127–139, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p127.pdf.

Danovaro:2002:MPE

[DGM02] R. Danovaro, C. Gambi, and S. Mirto. Meiofaunal production and energy transfer efficiency in a seagrass Posidonia oceanica bed in the western Mediterranean. Marine Ecology Progress Series, 234:95–104, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p095.pdf.

Dorenbosch:2005:DCR

[DGNvdV05] M. Dorenbosch, M. G. G. Grol, I. Nagelkerken, and G. van der Velde. Distribution of coral reef fishes along a coral reefseagrass gradient: edge effects and habitat segregation. Marine Ecology Progress Series, 299:277–288, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p277.pdf.

deGoeij:2008:MBD

[dGvdBvO^{+}08] J. M. de Goeij, H. van den Berg, M. M. van Oostveen, E. H. G. Epping, and F. C. van Duyl. Major bulk dissolved organic carbon (DOC) removal by encrusting coral reef cavity sponges. Marine Ecology Progress Series, 357:139–151, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p139.pdf.

Dahdouh-Guebas:2004:HIM

[DGVK^{+}04] F. Dahdouh-Guebas, I. Van Pottelbergh, J. G. Kairo, S. Cannicci, and N. Koedam. Human-impacted mangroves in Gazi (Kenya): predicting future vegetation based on retrospective remote sensing, social surveys, and tree distribution. Marine Ecology Progress Series, 272:77–92, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p077.pdf.

Dudley:2001:RBV

[DGW01] B. J. Dudley, A. M. E. Gahnström, and D. I. Walker. The role of benthic vegetation as a sink for elevated inputs of ammonium and nitrate in a mesotrophic estuary. Marine Ecology Progress Series, 219:99–107, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p099.pdf.

Duffy:2001:SSI

[DH01] J. E. Duffy and A. M. Harvilicz. Species-specific impacts of grazing amphipods in an eelgrass-bed community. Marine Ecology Progress Series, 223:201–211, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p201.pdf.

Duncan:2006:HUG

[DH06] K. M. Duncan and K. N. Holland. Habitat use, growth rates and dispersal patterns of juvenile scalloped hammerhead sharks Sphyrna lewini in a nursery habitat. Marine Ecology Progress Series, 312:211–221, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/312/m312p211.pdf.

Dieckmann:2007:PMR

[DH07a] U. Dieckmann and M. Heino. Probabilistic maturation reaction norms: their history, strengths, and limitations. Marine Ecology Progress Series, 335:253–269, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m335p253.pdf.

Dirnwober:2007:MSE

[DH07b] M. Dirnwöber and J. Herler. Microhabitat specialisation and ecological consequences for coral gobies of the genus Gobiodon in the Gulf of Aqaba, northern Red Sea. Marine Ecology Progress Series, 342:265–275, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p265.pdf.

Dijkstra:2009:MDA

[DH09a] J. A. Dijkstra and L. G. Harris. Maintenance of diversity altered by a shift in dominant species: implications for species coexistence. Marine Ecology Progress Series, 387: 71–80, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/387/m387p071.pdf.

Dunn:2009:RBR

[DH09b] D. C. Dunn and P. N. Halpin. Rugosity-based regional modeling of hard-bottom habitat. Marine Ecology Progress Series, 377:1–11, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

DeLaRocha:2000:EIZ

[DHBZ00] C. L. De La Rocha, D. A. Hutchins, M. A. Brzezinski, and Y. Zhang. Effects of iron and zinc deficiency on elemental composition and silica production by diatoms. Marine Ecology Progress Series, 195:71–79, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p071.pdf.

Davenport:2004:DDU

[DHCH04] J. Davenport, A. Aine Healy, N. Casey, and J. J. A. Heffron. Diet-dependent UVAR and UVBR resistance in the high shore harpacticoid copepod Tigriopus brevicornis. Marine Ecology Progress Series, 276:299–303, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p299.pdf.

Davis:2004:RTO

[DHG^{+}04] C. S. Davis, Q. Hu, S. M. Gallager, X. Tang, and C. J. Ashjian. Real-time observation of taxa-specific plankton distributions: an optical sampling method. Marine Ecology Progress Series, 284:77–96, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p077.pdf.

Davies:2009:SST

[DHHY09] W. E. Davies, J. M. Hipfner, K. A. Hobson, and R. C. Ydenberg. Seabird seasonal trophodynamics: isotopic patterns in a community of Pacific alcids. Marine Ecology Progress Series, 382:211–219, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/382/m382p211.pdf.

Drolet:2004:URO

[DHR04a] D. Drolet, J. H. Himmelman, and R. Rochette. Use of refuges by the ophiuroid Ophiopholis aculeata: contrasting effects of substratum complexity on predation risk from two predators. Marine Ecology Progress Series, 284:173–183, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p173.pdf.

Dumont:2004:SSM

[DHR04b] C. Dumont, J. H. Himmelman, and M. P. Russell. Size-specific movement of green sea urchins Strongylocentrotus droebachiensis on urchin barrens in eastern Canada. Marine Ecology Progress Series, 276:93–101, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p093.pdf.

Dumont:2006:DMS

[DHR06] C. P. Dumont, J. H. Himmelman, and M. P. Russell. Daily movement of the sea urchin Strongylocentrotus droebachiensis in different subtidal habitats in eastern Canada. Marine Ecology Progress Series, 317:87–99, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p087.pdf.

Descroix:2005:MCP

[DHRG05] A. Descroix, M. Harvey, S. Roy, and P. S. Galbraith. Macrozooplankton community patterns driven by water circulation in the St. Lawrence marine system, Canada. Marine Ecology Progress Series, 302:103–119, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/302/m302p103.pdf.

Dupont:2008:NFL

[DHT^{+}08] S. Dupont, J. Havenhand, W. Thorndyke, L. Peck, and M. Thorndyke. Near-future level of CO_{2}-driven ocean acidification radically affects larval survival and development in the brittlestar Ophiothrix fragilis. Marine Ecology Progress Series, 373:285–294, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m373p285.pdf.

Dippner:2002:RQA

[DHZ02] J. W. Dippner, R. Heerkloss, and J. P. Zbilut. Recurrence quantification analysis as a tool for characterization of nonlinear mesocosm dynamics. Marine Ecology Progress Series, 242:29–37, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p029.pdf.

Distel:2007:MBB

[Dis07] D. L. Distel. Molecular biorepositories and biomaterials management: enhancing the value of high-throughput molecular methodologies for the natural sciences. Marine Ecology Progress Series, 332:307–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m332p307.pdf.

Dinmore:2004:PAB

[DJ04] T. A. Dinmore and S. Jennings. Predicting abundance-body mass relationships in benthic infaunal communities. Marine Ecology Progress Series, 276:289–292, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p289.pdf.

Davis:2007:DBS

[DJG^{+}07] R. W. Davis, N. Jaquet, D. Gendron, U. Markaida, G. Bazzino, and W. Gilly. Diving behavior of sperm whales in relation to behavior of a major prey species, the jumbo squid, in the Gulf of California, Mexico. Marine Ecology Progress Series, 333:291–302, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p291.pdf.

Dick:2005:PNL

[DJM^{+}05] J. T. A. Dick, M. P. Johnson, S. McCambridge, J. J. Johnson, V. E. E. Carson, D. W. Kelly, and C. MacNeil. Predatory nature of littoral amphipod Echinogammarus marinus: gut contents analysis and effects of alternative food and substrate heterogeneity. Marine Ecology Progress Series, 291: 151–158, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p151.pdf.

Davies:2007:LGL

[DJM07] A. J. Davies, M. P. Johnson, and C. A. Maggs. Limpet grazing and loss of Ascophyllum nodosum canopies on decadal time scales. Marine Ecology Progress Series, 339:131–141, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/339/m339p131.pdf.

Davies:2008:SAN

[DJM08] A. J. Davies, M. P. Johnson, and C. A. Maggs. Subsidy by Ascophyllum nodosum increases growth rate and survivorship of Patella vulgata. Marine Ecology Progress Series, 366: 43–48, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/366/m366p043.pdf.

deJuan:2007:FCI

[dJTD07] S. de Juan, S. F. Thrush, and M. Demestre. Functional changes as indicators of trawling disturbance on a benthic community located in a fishing ground (NW Mediterranean Sea). Marine Ecology Progress Series, 334:117–129, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/334/m334p117.pdf.

Dupont:2003:HGD

[DJV03] L. Dupont, D. Jollivet, and F. Viard. High genetic diversity and ephemeral drift effects in a successful introduced mollusc (Crepidula fornicata: Gastropoda). Marine Ecology Progress Series, 253:183–195, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/253/m253p183.pdf.

Dempster:2003:HPF

[DK03] T. Dempster and M. J. Kingsford. Homing of pelagic fish to fish aggregation devices (FADs): the role of sensory cues. Marine Ecology Progress Series, 258:213–222, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p213.pdf.

Dijkman:2006:PDF

[DK06] N. A. Dijkman and J. C. Kromkamp. Phospholipid-derived fatty acids as chemotaxonomic markers for phytoplankton: application for inferring phytoplankton composition. Marine Ecology Progress Series, 324:113–125, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p113.pdf.

Dahlgren:2006:MNE

[DKA^{+}06] C. P. Dahlgren, G. T. Kellison, A. J. Adams, B. M. Gillanders, M. S. Kendall, C. A. Layman, J. A. Ley, I. Nagelkerken, and J. E. Serafy. Marine nurseries and effective juvenile habitats: concepts and applications. Marine Ecology Progress Series, 312:291–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/312/m312p291.pdf.

Dube:2002:SSG

[DKAH02] D. Dube, K. Kim, A. P. Alker, and C. D. Harvell. Size structure and geographic variation in chemical resistance of sea fan corals Gorgonia ventalina to a fungal pathogen. Marine Ecology Progress Series, 231:139–150, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p139.pdf.

Dreyer:2005:DMC

[DKFV05] J. C. Dreyer, K. E. Knick, W. B. Flickinger, and C. L. Van Dover. Development of macrofaunal community structure in mussel beds on the northern East Pacific Rise. Marine Ecology Progress Series, 302:121–134, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/302/m302p121.pdf.

Darling:2009:HMT

[DKR09] J. A. Darling, A. Kuenzi, and A. M. Reitzel. Humanmediated transport determines the non-native distribution of the anemone Nematostella vectensis, a dispersal-limited estuarine invertebrate. Marine Ecology Progress Series, 380: 137–146, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/380/m380p137.pdf.

Dittmar:2001:DMR

[DL01] T. Dittmar and R. J. Lara. Do mangroves rather than rivers provide nutrients to coastal environments south of the Amazon River? Evidence from long-term flux measurements. Marine Ecology Progress Series, 213:67–77, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
213/m213p067.pdf.

Davis:2002:IPR

[DL02] J. L. D. Davis and L. A. Levin. Importance of pre-recruitment life-history stages to population dynamics of the woolly sculpin Clinocottus analis. Marine Ecology Progress Series, 234:229–246, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p229.pdf.

Devescovi:2003:GTR

[DL03a] M. Devescovi and C. Lucu. Growth of tissues related to haemolymph copper throughout the moult cycle of the lobster Homarus gammarus. Marine Ecology Progress Series, 247:165–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p165.pdf.

Duquesne:2003:ISM

[DL03b] S. Duquesne and M. Liess. Increased sensitivity of the macroinvertebrate Paramorea walkeri to heavy-metal contamination in the presence of solar UV radiation in Antarctic shoreline waters. Marine Ecology Progress Series, 255:183– 191, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/255/m255p183.pdf.

Duineveld:2004:PFF

[DLB04] G. C. A. Duineveld, M. S. S. Lavaleye, and E. M. Berghuis. Particle flux and food supply to a seamount cold-water coral community (Galicia Bank, NW Spain). Marine Ecology Progress Series, 277:13–23, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p013.pdf.

Dyrynda:2000:CIP

[DLD^{+}00] E. A. Dyrynda, R. J. Law, P. E. J. Dyrynda, C. A. Kelly, R. K. Pipe, and N. A. Ratcliffe. Changes in immune parameters of natural mussel Mytilus edulis populations following a major oil spill (‘Sea Empress’, Wales, UK). Marine Ecology Progress Series, 206:155–170, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/206/m206p155.pdf.

Das:2003:MMS

[DLLB03] K. Das, G. Lepoint, Y. Leroy, and J. M. Bouquegneau. Marine mammals from the southern North Sea: feeding ecology data from δ ^{13}C and δ ^{15}N measurements. Marine Ecology Progress Series, 263:287–298, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/263/m263p287.pdf.

delaMoriniere:2003:ODC

[dlMPN^{+}03] E. Cocheret de la Morinière, B. J. A. Pollux, I. Nagelkerken, M. A. Hemminga, A. H. L. Huiskes, and G. van der Velde. Ontogenetic dietary changes of coral reef fishes in the mangrove-seagrass-reef continuum: stable isotopes and gut-content analysis. Marine Ecology Progress Series, 246: 279–289, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p279.pdf.

delosSantos:2009:PGM

[dlSPLV09] C. B. de los Santos, J. L. Pérez-Lloréns, and J. J. Vergara. Photosynthesis and growth in macroalgae: linking functionalform and power-scaling approaches. Marine Ecology Progress Series, 377:113–122, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/377/m377p113.pdf.

Daverat:2006:PPH

[DLT^{+}06] F. Daverat, K. E. Limburg, I. Thibault, J. C. Shiao, J. J. Dodson, F. Caron, W. N. Tzeng, Y. Iizuka, and H. Wickström. Phenotypic plasticity of habitat use by three temperate eel species, Anguilla anguilla, A. japonica and A. rostrata. Marine Ecology Progress Series, 308:231–241, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p231.pdf.

deBruyn:2001:DLC

[dM01a] A. M. H. deBruyn and J. J. Meeuwig. Detecting lunar cycles in marine ecology: periodic regression versus categorical ANOVA. Marine Ecology Progress Series, 214:307–310, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p307.pdf.

Dobretsov:2001:LPL

[DM01b] S. V. Dobretsov and G. Miron. Larval and post-larval vertical distribution of the mussel Mytilus edulis in the White Sea. Marine Ecology Progress Series, 218:179–187, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p179.pdf.

Davoren:2003:SSI

[DM03] G. K. Davoren and W. A. Montevecchi. Signals from seabirds indicate changing biology of capelin stocks. Marine Ecology Progress Series, 258:253–261, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/258/m258p253.pdf. See

comment [Car05].

Davoren:2005:DSS

[DM05a] G. K. Davoren and W. A. Montevecchi. Did signals from seabirds indicate changes in capelin biology? Reply to Carscadden (2004). Marine Ecology Progress Series, 285:

299–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p299.pdf. See [Car05].

Defeo:2005:PPR

[DM05b] O. Defeo and A. McLachlan. Patterns, processes and regulatory mechanisms in sandy beach macrofauna: a multi-scale analysis. Marine Ecology Progress Series, 295:1–20, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Dale:2008:HII

[DM08] R. K. Dale and D. C. Miller. Hydrologic interactions of infaunal polychaetes and intertidal groundwater discharge. Marine Ecology Progress Series, 363:205–215, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p205.pdf.

Duhamel:2009:CPI

[DM09] S. Duhamel and T. Moutin. Carbon and phosphate incorporation rates of microbial assemblages in contrasting environments in the Southeast Pacific. Marine Ecology Progress Series, 375:53–64, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m375p053.pdf.

Davoren:2002:SDA

[DMA02] G. K. Davoren, W. A. Montevecchi, and J. T. Anderson. Scaledependent associations of predators and prey: constraints imposed by flightlessness of common murres. Marine Ecology Progress Series, 245:259–272, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p259.pdf.

Davoren:2003:DPM

[DMA03] G. K. Davoren, W. A. Montevecchi, and J. T. Anderson. Distributional patterns of a marine bird and its prey: habitat selection based on prey and conspecific behaviour. Marine Ecology Progress Series, 256:229–242, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p229.pdf.

Duffy:2009:VSR

[DMC09] T. A. Duffy, A. E. McElroy, and D. O. Conover. Variable susceptibility and response to estrogenic chemicals in Menidia menidia. Marine Ecology Progress Series, 380:245– 254, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/380/m380p245.pdf.

Dauwe:2001:EOD

[DMH01] B. Dauwe, J. J. Middelburg, and P. M. J. Herman. Effect of oxygen on the degradability of organic matter in subtidal and intertidal sediments of the North Sea area. Marine Ecology Progress Series, 215:13–22, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/215/m215p013.pdf.

DeAngelis:2008:FCS

[DMK^{+}08] B. M. DeAngelis, C. T. McCandless, N. E. Kohler, C. W. Recksiek, and G. B. Skomal. First characterization of shark nursery habitat in the United States Virgin Islands: evidence of habitat partitioning by two shark species. Marine Ecology Progress Series, 358:257–271, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/358/m358p257.pdf.

Delany:2003:RPS

[DMM^{+}03] J. Delany, A. A. Myers, D. McGrath, R. M. O’Riordan, and A. M. Power. Role of post-settlement mortality and ’supplyside’ ecology in setting patterns of intertidal distribution in the chthamalid barnacles Chthamalus montagui and C. stellatus. Marine Ecology Progress Series, 249:207–214, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p207.pdf.

DeRobertis:2005:CRP

[DMS^{+}05] A. De Robertis, C. A. Morgan, R. A. Schabetsberger, R. W. Zabel, R. D. Brodeur, R. L. Emmett, C. M. Knight, G. K. Krutzikowsky, and E. Casillas. Columbia River plume fronts. II. Distribution, abundance, and feeding ecology of juvenile salmon. Marine Ecology Progress Series, 299:33–44, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/299/m299p033.pdf.

Dickmann:2007:FEC

[DMV07] M. Dickmann, C. Möllmann, and R. Voss. Feeding ecology of Central Baltic sprat Sprattus sprattus larvae in relation to zooplankton dynamics: implications for survival. Marine Ecology Progress Series, 342:277–289, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p277.pdf.

Domeier:2008:MPW

[DNL08a] M. L. Domeier and N. Nasby-Lucas. Migration patterns of white sharks Carcharodon carcharias tagged at Guadalupe Island, Mexico, and identification of an eastern Pacific shared offshore foraging area. Marine Ecology Progress Series, 370: 221–237, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m370p221.pdf.

Dudgeon:2008:ADS

[DNL08b] C. L. Dudgeon, M. J. Noad, and J. M. Lanyon. Abundance and demography of a seasonal aggregation of zebra sharks Stegostoma fasciatum. Marine Ecology Progress Series, 368:269–281, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p269.pdf.

Dubois:2007:SSS

[DOML^{+}07] S. Dubois, F. Orvain, J. C. Marin-Léal, M. Ropert, and S. Lefebvre. Small-scale spatial variability of food partitioning between cultivated oysters and associated suspensionfeeding species, as revealed by stable isotopes. Marine Ecology Progress Series, 336:151–160, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/336/m336p151.pdf.

Donahue:2004:SDC

[Don04] M. J. Donahue. Size-dependent competition in a gregarious porcelain crab Petrolisthes cinctipes (Anomura: Porcellanidae). Marine Ecology Progress Series, 267:219–231, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/267/m267p219.pdf.

Dove:2004:SCP

[Dov04] S. Dove. Scleractinian corals with photoprotective host pigments are hypersensitive to thermal bleaching. Marine Ecology Progress Series, 272:99–116, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p099.pdf.

DeFelice:2001:PPD

[DP01] R. C. DeFelice and J. D. Parrish. Physical processes dominate in shaping invertebrate assemblages in reef-associated sediments of an exposed Hawaiian coast. Marine Ecology Progress Series, 215:121–131, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/215/m215p121.pdf.

Dale:2002:PBC

[DP02] A. W. Dale and R. Prego. Physico-biogeochemical controls on benthic-pelagic coupling of nutrient fluxes and recycling in a coastal upwelling system. Marine Ecology Progress Series, 235:15–28, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p015.pdf.

Desai:2009:PRH

[DP09] D. V. Desai and S. Prakash. Physiological responses to hypoxia and anoxia in Balanus amphitrite (Cirripedia: Thoracica). Marine Ecology Progress Series, 390:157–166, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/390/m390p157.pdf.

Debier:2003:DPT

[DPD^{+}03a] C. Debier, P. P. Pomeroy, C. Dupont, C. Joiris, V. Comblin, E. Le Boulengé, Y. Larondelle, and J. P. Thomé. Dynamics of PCB transfer from mother to pup during lactation in UK grey seals Halichoerus grypus: differences in PCB profile between compartments of transfer and changes during the lactation period. Marine Ecology Progress Series, 247:249–256, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p249.pdf.

Debier:2003:QDP

[DPD^{+}03b] C. Debier, P. P. Pomeroy, C. Dupont, C. Joiris, V. Comblin, E. Le Boulengé, Y. Larondelle, and J. P. Thomé. Quantitative dynamics of PCB transfer from mother to pup during lactation in UK grey seals Halichoerus grypus. Marine Ecology Progress Series, 247:237–248, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/247/m247p237.pdf.

Drake:2008:PND

[DPD^{+}08] D. C. Drake, B. J. Peterson, L. A. Deegan, L. A. Harris, E. E. Miller, and R. S. Warren. Plant nitrogen dynamics in fertilized and natural New England salt marshes: a paired ^{15}N tracer study. Marine Ecology Progress Series, 354:35–46, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p035.pdf.

Drazen:2008:LSF

[DPGN08] J. C. Drazen, C. F. Phleger, M. A. Guest, and P. D. Nichols. Lipid, sterols and fatty acids of abyssal polychaetes, crustaceans, and a cnidarian from the northeast Pacific Ocean: food web implications. Marine Ecology Progress Series, 372: 157–167, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p157.pdf.

Drazen:2009:LCD

[DPGN09] J. C. Drazen, C. F. Phleger, M. A. Guest, and P. D. Nichols. Lipid composition and diet inferences in abyssal macrourids of the eastern North Pacific. Marine Ecology Progress Series, 387:1–14, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Defew:2002:UNM

[DPH02] E. C. Defew, D. M. Paterson, and S. E. Hagerthey. The use of natural microphytobenthic assemblages as laboratory model systems. Marine Ecology Progress Series, 237:15– 25, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p015.pdf.

Diaz-Pulido:2002:FBC

[DPM02] G. D. Diaz-Pulido and L. J. McCook. The fate of bleached corals: patterns and dynamics of algal recruitment. Marine Ecology Progress Series, 232:115–128, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p115.pdf.

Daunt:2003:RRR

[DPS^{+}03] F. Daunt, G. Peters, B. Scott, D. Grémillet, and S. Wanless. Rapid-response recorders reveal interplay between marine physics and seabird behaviour. Marine Ecology Progress Series, 255:283–288, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p283.pdf.

Doubleday:2008:SES

[DPSD08a] Z. A. Doubleday, G. T. Pecl, J. M. Semmens, and L. Danyushevsky. Stylet elemental signatures indicate population structure in a holobenthic octopus species, Octopus pallidus. Marine Ecology Progress Series, 371:1–10, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Doubleday:2008:USE

[DPSD08b] Z. A. Doubleday, G. T. Pecl, J. M. Semmens, and L. Danyushevsky. Using stylet elemental signatures to determine the population structure of Octopus maorum. Marine Ecology Progress Series, 360:125–133, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p125.pdf.

Dahms:2005:EBM

[DQ05] H. U. Dahms and P. Y. Qian. Exposure of biofilms to meiofaunal copepods affects the larval settlement of Hydroides elegans (Polychaeta). Marine Ecology Progress Series, 297: 203–214, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/297/m297p203.pdf.

Dobretsov:2005:ESU

[DQW05] S. V. Dobretsov, P. Y. Qian, and M. Wahl. Effect of solar ultraviolet radiation on the formation of shallow, early successional biofouling communities in Hong Kong. Marine Ecology Progress Series, 290:55–65, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/290/m290p055.pdf.

Diaz:2000:NRD

[DR00] F. Diaz and P. Raimbault. Nitrogen regeneration and dissolved organic nitrogen release during spring in a NW Mediterranean coastal zone (Gulf of Lions): implications for the estimation of new production. Marine Ecology Progress Series, 197:51–65, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p051.pdf.

Duchene:2001:MBF

[DR01] J. C. Duchêne and R. Rosenberg. Marine benthic faunal activity patterns on a sediment surface assessed by video numerical tracking. Marine Ecology Progress Series, 223:113–119, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p113.pdf.

Diaz:2001:ESP

[DRB^{+}01] F. Diaz, P. Raimbault, B. Boudjellal, N. Garcia, and T. Moutin. Early spring phosphorus limitation of primary productivity in a NW Mediterranean coastal zone (Gulf of Lions). Marine Ecology Progress Series, 211:51–62, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p051.pdf.

DeBruyn:2009:SBP

[DRDD09] C. De Bruyn, T. Rigaud, B. David, and C. De Ridder. Symbiosis between the pea crab Dissodactylus primitivus and its echinoid host Meoma ventricosa: potential consequences for the crab mating system. Marine Ecology Progress Series, 375:173–183, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p173.pdf.

Darnell:2009:LRP

[DRDM09] M. Z. Darnell, D. Rittschof, K. M. Darnell, and R. E. Mc-Dowell. Lifetime reproductive potential of female blue crabs Callinectes sapidus in North Carolina, USA. Marine Ecology Progress Series, 394:153–163, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p153.pdf.

Demarcq:2008:GMP

[DRF08] H. Demarcq, A. J. Richardson, and J. G. Field. Generalised model of primary production in the southern Benguela upwelling system. Marine Ecology Progress Series, 354:59– 74, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p059.pdf.

Drake:2002:MEB

[DRG^{+}02] L. A. Drake, G. M. Ruiz, B. S. Galil, T. L. Mullady, D. O. Friedmann, and F. C. Dobbs. Microbial ecology of ballast water during a transoceanic voyage and the effects of open-ocean exchange. Marine Ecology Progress Series, 233: 13–20, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p013.pdf.

Dedieu:2007:CCN

[DRG^{+}07] K. Dedieu, C. Rabouille, F. Gilbert, K. Soetaert, E. Metzger, C. Simonucci, D. Jézéquel, F. Prévot, P. Anschutz, S. Hulth, S. Ogier, and V. Mesnage. Coupling of carbon, nitrogen and oxygen cycles in sediments from a Mediterranean lagoon: a seasonal perspective. Marine Ecology Progress Series, 346:45–59, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/346/m346p045.pdf.

Dudas:2009:LLA

[DRTM09] S. E. Dudas, G. Rilov, J. Tyburczy, and B. A. Menge. Linking larval abundance, onshore supply and settlement using instantaneous versus integrated methods. Marine Ecology Progress Series, 387:81–95, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/387/m387p081.pdf.

Derycke:2005:MDV

[DRV^{+}05] S. Derycke, T. Remerie, A. Vierstraete, T. Backeljau, J. Vanfleteren, M. Vincx, and T. Moens. Mitochondrial DNA variation and cryptic speciation within the free-living marine nematode Pellioditis marina. Marine Ecology Progress Series, 300:91–103, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p091.pdf.

Dethier:2001:GPD

[DS01] M. N. Dethier and R. S. Steneck. Growth and persistence of diverse intertidal crusts: survival of the slow in a fast-paced world. Marine Ecology Progress Series, 223:89–100, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p089.pdf.

Dethier:2006:TSD

[DS06a] M. N. Dethier and G. C. Schoch. Taxonomic sufficiency in distinguishing natural spatial patterns on an estuarine shoreline. Marine Ecology Progress Series, 306:41–49, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p041.pdf.

Duffy:2006:WBI

[DS06b] J. E. Duffy and J. J. Stachowicz. Why biodiversity is important to oceanography: potential roles of genetic, species, and trophic diversity in pelagic ecosystem processes. Marine Ecology Progress Series, 311:179–189, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m311p179.pdf.

Dytham:2007:EMF

[DS07] C. Dytham and S. D. Simpson. Elevated mortality of fish larvae on coral reefs drives the evolution of larval movement patterns. Marine Ecology Progress Series, 346:255– 264, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/346/m346p255.pdf.

Dawson:2009:MSP

[DS09a] C. L. Dawson and R. M. Starr. Movements of subadult prickly sharks Echinorhinus cookei in the Monterey Canyon. Marine Ecology Progress Series, 386:253–262, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m386p253.pdf.

Duenas:2009:CLD

[DS09b] L. F. Dueñas and J. A. Sánchez. Character lability in deep-sea bamboo corals (Octocorallia, Isididae, Keratoisidinae). Marine Ecology Progress Series, 397:11–23, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p011.pdf.

deStephanis:2008:SSD

[dSCV^{+}08] R. de Stephanis, T. Cornulier, P. Verborgh, J. Salazar Sierra, N. Pérez Gimeno, and C. Guinet. Summer spatial distribution of cetaceans in the Strait of Gibraltar in relation to the oceanographic context. Marine Ecology Progress Series, 353:275–288, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/353/m353p275.pdf.

Druon:2004:CAL

[DSDS04] J. N. Druon, W. Schrimpf, S. Dobricic, and A. Stips. Comparative assessment of large-scale marine eutrophication: North Sea area and Adriatic Sea as case studies. Marine Ecology Progress Series, 272:1–23, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p001.pdf.

Das:2004:EPF

[DSF^{+}04] K. Das, U. Siebert, M. Fontaine, T. Jauniaux, L. Holsbeek, and J. M. Bouquegneau. Ecological and pathological factors related to trace metal concentrations in harbour porpoises Phocoena phocoena from the North Sea and adjacent areas. Marine Ecology Progress Series, 281:283–295, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p283.pdf.

Donlan:2003:CMI

[DSH03] M. Donlan, M. Sperduto, and C. Hebert. Compensatory mitigation for injury to a threatened or endangered species: scaling piping plover restoration. Marine Ecology Progress Series, 264:213–219, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/264/m264p213.pdf.

deSegura:2007:CCA

[dSHCR07] A. Gómez de Segura, P. S. Hammond, A. Cañadas, and J. A. Raga. Comparing cetacean abundance estimates derived from spatial models and design-based line transect methods. Marine Ecology Progress Series, 329:289–299, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p289.pdf.

Dempster:2002:AWF

[DSJBS^{+}02] T. Dempster, P. Sanchez-Jerez, J. T. Bayle-Sempere, F. Giménez-Casalduero, and C. Valle. Attraction of wild fish to sea-cage fish farms in the south-western Mediterranean Sea: spatial and short-term temporal variability. Marine Ecology Progress Series, 242:237–252, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/242/m242p237.pdf.

Dempster:2006:CAC

[DSJT^{+}06] T. Dempster, P. Sanchez-Jerez, F. Tuya, D. Fernandez-Jover, J. Bayle-Sempere, A. Boyra, and R. Haroun. Coastal aquaculture and conservation can work together. Marine Ecology Progress Series, 314:309–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/314/m314p309.pdf.

DAlessandro:2007:PPL

[DSL07] E. D’Alessandro, S. Sponaugle, and T. Lee. Patterns and processes of larval fish supply to the coral reefs of the upper Florida Keys. Marine Ecology Progress Series, 331:85– 100, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/331/m331p085.pdf.

DiBacco:2001:VMB

[DSM01] C. DiBacco, D. Sutton, and L. McConnico. Vertical migration behavior and horizontal distribution of brachyuran larvae in a low-inflow estuary: implications for bay-ocean exchange. Marine Ecology Progress Series, 217:191–206, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p191.pdf.

Dionne:2003:DHS

[DSMBG03] M. Dionne, B. Sainte-Marie, E. Bourget, and D. Gilbert. Distribution and habitat selection of early benthic stages of snow crab Chionoecetes opilio. Marine Ecology Progress Series, 259:117–128, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p117.pdf.

Davenport:2000:MME

[DSP00] J. Davenport, R. J. J. W. Smith, and M. Packer. Mussels Mytilus edulis: significant consumers and destroyers of mesozooplankton. Marine Ecology Progress Series, 198:131– 137, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/198/m198p131.pdf.

Devred:2007:DEP

[DSP07] E. Devred, S. Sathyendranath, and T. Platt. Delineation of ecological provinces using ocean colour radiometry. Marine Ecology Progress Series, 346:1–13, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Darnaude:2004:DVT

[DSPHV04] A. M. Darnaude, C. Salen-Picard, and M. L. Harmelin-Vivien. Depth variation in terrestrial particulate organic matter exploitation by marine coastal benthic communities off the Rhone River delta (NW Mediterranean). Marine Ecology Progress Series, 275:47–57, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/275/m275p047.pdf.

Dando:2004:RSS

[DSS04] P. R. Dando, A. J. Southward, and E. C. Southward. Rates of sediment sulphide oxidation by the bivalve mollusc Thyasira sarsi. Marine Ecology Progress Series, 280: 181–187, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/280/m280p181.pdf.

Dang:2009:DDM

[DSS^{+}09] C. Dang, P. G. Sauriau, N. Savoye, N. Caill-Milly, P. Martinez, C. Millaret, J. Haure, and X. de Montaudouin. Determination of diet in Manila clams by spatial analysis of stable isotopes. Marine Ecology Progress Series, 387:167–177, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p167.pdf.

dosSantos:2007:HVD

[dSSC07] A. dos Santos, A. M. P. Santos, and D. V. P. Conway. Horizontal and vertical distribution of cirripede cyprid larvae in an upwelling system off the Portuguese coast. Marine Ecology Progress Series, 329:145–155, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/329/m329p145.pdf.

dosSantos:2008:DVM

[dSSC^{+}08] A. dos Santos, A. M. P. Santos, D. V. P. Conway, C. Bartilotti, P. Lourenço, and H. Queiroga. Diel vertical migration of decapod larvae in the Portuguese coastal upwelling ecosystem: implications for offshore transport. Marine Ecology Progress Series, 359:171–183, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/359/m359p171.pdf.

Daverat:2006:TDA

[DT06] F. Daverat and J. Tomás. Tactics and demographic attributes in the European eel Anguilla anguilla in the Gironde watershed, SW France. Marine Ecology Progress Series, 307: 247–257, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p247.pdf.

Darby:2008:EES

[DT08] F. A. Darby and R. E. Turner. Effects of eutrophication on salt marsh root and rhizome biomass accumulation. Marine Ecology Progress Series, 363:63–70, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p063.pdf.

Duarte:2000:ETO

[DTAF00] C. M. Duarte, J. Terrados, N. S. R. Agawin, and M. D. Fortes. An experimental test of the occurrence of competitive interactions among SE Asian seagrasses. Marine Ecology Progress Series, 197:231–240, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/197/m197p231.pdf.

Diehl:2007:SVR

[DTB07] J. M. Diehl, R. J. Toonen, and L. W. Botsford. Spatial variability of recruitment in the sand crab Emerita analoga throughout California in relation to wind-driven currents. Marine Ecology Progress Series, 350:1–17, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Dearman:2003:IAM

[DTD03] J. R. Dearman, A. H. Taylor, and K. Davidson. Influence of autotroph model complexity on simulations of microbial communities in marine mesocosms. Marine Ecology Progress Series, 250:13–28, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p013.pdf.

Dong:2000:DSR

[DTNU00] L. F. Dong, D. C. O. Thornton, D. B. Nedwell, and G. J. C. Underwood. Denitrification in sediments of the River Colne estuary, England. Marine Ecology Progress Series, 203:109– 122, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/203/m203p109.pdf.

Diaz:2008:BCC

[DTV08] E. Díaz, J. M. Txurruka, and F. Villate. Biochemical composition and condition in anchovy larvae Engraulis encrasicolus during growth. Marine Ecology Progress Series, 361: 227–238, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p227.pdf.

Diaz:2009:BCS

[DTV09] E. Díaz, J. M. Txurruka, and F. Villate. Biochemical composition and somatic growth of pelagic larvae of three fish species from the Bay of Biscay. Marine Ecology Progress Series, 382:173–183, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/382/m382p173.pdf.

Duffy:2006:BFS

[Duf06] J. E. Duffy. Biodiversity and the functioning of seagrass ecosystems. Marine Ecology Progress Series, 311:233–250, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m311p233.pdf.

Dunne:2002:CBT

[Dun02] R. P. Dunne. Comment on Berkelmans (2002) ‘Timeintegrated thermal bleaching thresholds of reefs and their variation on the Great Barrier Reef’. Marine Ecology Progress Series, 237:307–308, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p307.pdf. See [Ber02b]

and reply [Ber02a].

Dempster:2009:CSF

[DUSJ^{+}09] T. Dempster, I. Uglem, P. Sanchez-Jerez, D. Fernandez-Jover, J. Bayle-Sempere, R. Nilsen, and P. A. Bjørn. Coastal salmon farms attract large and persistent aggregations of wild fish: an ecosystem effect. Marine Ecology Progress Series, 385:1–14, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

DeCaralt:2008:GDS

[DUW08] S. De Caralt, M. J. Uriz, and R. H. Wijffels. Grazing, differential size-class dynamics and survival of the Mediterranean sponge Corticium candelabrum. Marine Ecology Progress Series, 360:97–106, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p097.pdf.

deVoogd:2009:SCC

[dVBC09] N. J. de Voogd, L. E. Becking, and D. F. R. Cleary. Sponge community composition in the Derawan Islands, NE Kalimantan, Indonesia. Marine Ecology Progress Series, 396: 169–180, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m396p169.pdf.

Doniol-Valcroze:2007:ITF

[DVBLS07] T. Doniol-Valcroze, D. Berteaux, P. Larouche, and R. Sears. Influence of thermal fronts on habitat selection by four rorqual whale species in the Gulf of St. Lawrence. Marine Ecology Progress Series, 335:207–216, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/335/m335p207.pdf.

deVoogd:2006:SBD

[dVCH^{+}06] N. J. de Voogd, D. F. R. Cleary, B. W. Hoeksema, A. Noor, and R. W. M. van Soest. Sponge beta diversity in the Spermonde Archipelago, SW Sulawesi, Indonesia. Marine Ecology Progress Series, 309:131–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/309/m309p131.pdf.

Diez:2002:HEH

[DvD02] C. E. Diez and R. P. van Dam. Habitat effect on hawksbill turtle growth rates on feeding grounds at Mona and Monito Islands, Puerto Rico. Marine Ecology Progress Series, 234:301–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p301.pdf.

Dupuy:2000:FRO

[DVLH^{+}00] C. Dupuy, A. Vaquer, T. Lam-Höai, C. Rougier, N. Mazouni, J. Lautier, Y. Collos, and S. Le Gall. Feeding rate of the oyster Crassostrea gigas in a natural planktonic community of the Mediterranean Thau Lagoon. Marine Ecology Progress Series, 205:171–184, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/205/m205p171.pdf.

Dorenbosch:2007:IHC

[DVNvdV07] M. Dorenbosch, W. C. E. P. Verberk, I. Nagelkerken, and G. van der Velde. Influence of habitat configuration on connectivity between fish assemblages of Caribbean seagrass beds, mangroves and coral reefs. Marine Ecology Progress Series, 334:103–116, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p103.pdf.

Davis:2009:EPS

[DW09] A. R. Davis and D. W. Ward. Establishment and persistence of species-rich patches in a species-poor landscape: role of a structure-forming subtidal barnacle. Marine Ecology Progress Series, 380:187–198, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/380/m380p187.pdf.

Denny:2004:RRS

[DWB04] C. M. Denny, T. J. Willis, and R. C. Babcock. Rapid recolonisation of snapper Pagrus auratus: Sparidae within an offshore island marine reserve after implementation of no-take status. Marine Ecology Progress Series, 272:183–190, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p183.pdf.

Duckworth:2009:PAS

[DWL09] A. R. Duckworth, C. W. Wolff, and H. Luter. Patterns of abundance and size across varying spatial scales for the coral reef sponge Coscinoderma matthewsi. Marine Ecology Progress Series, 396:27–33, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/396/m396p027.pdf.

Dunne:2004:NSR

[DWM04] J. A. Dunne, R. J. Williams, and N. D. Martinez. Network structure and robustness of marine food webs. Marine Ecology Progress Series, 273:291–302, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/273/m273p291.pdf.

Dupont:2007:MEG

[DWO^{+}07] S. Dupont, K. Wilson, M. Obst, H. Sköld, H. Nakano, and M. C. Thorndyke. Marine ecological genomics: when genomics meets marine ecology. Marine Ecology Progress Series, 332:257–273, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m332p257.pdf.

Dizon:2005:CRS

[DY05] R. M. Dizon and H. T. Yap. Coral responses in single- and mixed-species plots to nutrient disturbance. Marine Ecology Progress Series, 296:165–172, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/296/m296p165.pdf.

Dominik:2006:DCA

[DZ06] C. M. Dominik and R. C. Zimmerman. Dynamics of carbon allocation in a deep-water population of the deciduous kelp Pleurophycus gardneri (Laminariales). Marine Ecology Progress Series, 309:143–157, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m309p143.pdf.

deZwaan:2001:IIC

[dZCVC01] A. de Zwaan, O. Cattani, G. Vitali, and P. Cortesi. Influence of incubation conditions on the anoxic survival of marine bivalves. Static and semi-static incubations. Marine Ecology Progress Series, 211:169–179, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/211/m211p169.pdf.

Erga:2005:UTN

[EAF^{+}05] S. R. Erga, K. Aursland, Ó. Frette, B. Hamre, J. K. Lotsberg, J. J. Stamnes, J. Aure, F. Rey, and K. Stamnes. UV transmission in Norwegian marine waters: controlling factors and possible effects on primary production and vertical distribution of phytoplankton. Marine Ecology Progress Series, 305:79–100, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p079.pdf.

Elmetri:2004:EPG

[EB04a] I. Elmetri and P. R. F. Bell. Effects of phosphorus on the growth and nitrogen fixation rates of Lyngbya majuscula: implications for management in Moreton Bay, Queensland. Marine Ecology Progress Series, 281:27–35, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p027.pdf.

Elwen:2004:FSR

[EB04b] S. H. Elwen and P. B. Best. Female southern right whales Eubalaena australis: Are there reproductive benefits associated with their coastal distribution off South Africa? Marine Ecology Progress Series, 269:289–295, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p289.pdf.

Ellison:2005:ABA

[EB05] C. K. Ellison and R. S. Burton. Application of bead array technology to community dynamics of marine phytoplankton. Marine Ecology Progress Series, 288:75–85, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p075.pdf.

Ellis:2008:TIF

[EB08] W. L. Ellis and S. S. Bell. Tidal influence on a fringing mangrove intertidal fish community as observed by in situ video recording: implications for studies of tidally migrating nekton. Marine Ecology Progress Series, 370:207–219, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/370/
m370p207.pdf.

Emslie:2003:LHI

[EBA^{+}03] S. D. Emslie, P. A. Berkman, D. G. Ainley, L. Coats, and M. Polito. Late-holocene initiation of ice-free ecosystems in the southern Ross Sea, Antarctica. Marine Ecology Progress Series, 262:19–25, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p019.pdf.

Edmunds:2004:EDM

[EBC04] P. J. Edmunds, J. F. Bruno, and D. B. Carlon. Effects of depth and microhabitat on growth and survivorship of juvenile corals in the Florida Keys. Marine Ecology Progress Series, 278:115–124, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p115.pdf.

Evanson:2000:TVB

[EBG^{+}00] M. Evanson, E. A. Bornhold, R. H. Goldblatt, P. J. Harrison, and A. G. Lewis. Temporal variation in body composition and lipid storage of the overwintering, subarctic copepod Neocalanus plumchrus in the Strait of Georgia, British Columbia (Canada). Marine Ecology Progress Series, 192:239–247, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p239.pdf.

Edeline:2007:DMJ

[EBLE07] E. Edeline, L. Beaulaton, R. Le Barh, and P. Elie. Dispersal in metamorphosing juvenile eel Anguilla anguilla. Marine Ecology Progress Series, 344:213–218, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p213.pdf.

Edwards:2002:OCA

[EBR^{+}02] M. Edwards, G. Beaugrand, P. C. Reid, A. A. Rowden, and M. B. Jones. Ocean climate anomalies and the ecology of the North Sea. Marine Ecology Progress Series, 239: 1–10, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/239/m239p001.pdf.

Esler:2002:HDP

[EBT^{+}02] D. Esler, T. D. Bowman, K. A. Trust, B. E. Ballachey, T. A. Dean, S. C. Jewett, and C. E. Charles O’Clair. Harlequin duck population recovery following the ‘Exxon Valdez’ oil spill: progress, process and constraints. Marine Ecology Progress Series, 241:271–286, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/241/m241p271.pdf.

Eberl:2007:ACM

[EC07a] R. Eberl and E. J. Carpenter. Association of the copepod Macrosetella gracilis with the cyanobacterium Trichodesmium spp. in the North Pacific Gyre. Marine Ecology Progress Series, 333:205–212, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p205.pdf.

Erwin:2007:DDV

[EC07b] C. A. Erwin and B. C. Congdon. Day-to-day variation in sea-surface temperature reduces sooty tern Sterna fuscata foraging success on the Great Barrier Reef, Australia. Marine Ecology Progress Series, 331:255–266, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p255.pdf.

Eckert:2002:DJL

[Eck02] S. A. Eckert. Distribution of juvenile leatherback sea turtle Dermochelys coriacea sightings. Marine Ecology Progress Series, 230:289–293, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p289.pdf.

Eldridge:2005:DSI

[ECK05] P. M. Eldridge, L. A. Cifuentes, and J. E. Kaldy. Development of a stable-isotope constraint system for estuarine foodweb models. Marine Ecology Progress Series, 303:73–90, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m303p073.pdf.

Eby:2005:HDI

[ECM^{+}05] L. A. Eby, L. B. Crowder, C. M. McClellan, C. H. Peterson, and M. J. Powers. Habitat degradation from intermittent hypoxia: impacts on demersal fishes. Marine Ecology Progress Series, 291:249–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/291/m291p249.pdf.

Edwards:2002:FEA

[ED02] M. Edwards and M. S. Davies. Functional and ecological aspects of the mucus trails of the intertidal prosobranch gastropod Littorina littorea. Marine Ecology Progress Series, 239:129–137, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p129.pdf.

Edeline:2004:TSR

[EDB^{+}04] E. Edeline, S. Dufour, C. Briand, D. Fatin, and P. Elie. Thyroid status is related to migratory behavior in Anguilla anguilla glass eels. Marine Ecology Progress Series, 282:261– 270, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p261.pdf.

Eker-Develi:2008:PCD

[EDBvdL08] E. Eker-Develi, J. F. Berthon, and D. van der Linde. Phytoplankton class determination by microscopic and HPLC-CHEMTAX analyses in the southern Baltic Sea. Marine Ecology Progress Series, 359:69–87, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m359p069.pdf.

Edeline:2005:RGE

[EDE05] E. Edeline, S. Dufour, and P. Elie. Role of glass eel salinity preference in the control of habitat selection and growth plasticity in Anguilla anguilla. Marine Ecology Progress Series, 304:191–199, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p191.pdf.

Eker-Develi:2006:RSD

[EDKT06] E. Eker-Develi, A. E. Kideys, and S. Tugrul. Role of Saharan dust on phytoplankton dynamics in the northeastern Mediterranean. Marine Ecology Progress Series, 314: 61–75, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p061.pdf.

Edmunds:2000:PDJ

[Edm00] P. J. Edmunds. Patterns in the distribution of juvenile corals and coral reef community structure in St. John, US Virgin Islands. Marine Ecology Progress Series, 202:113–124, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p113.pdf.

Edmunds:2004:JCP

[Edm04] P. J. Edmunds. Juvenile coral population dynamics track rising seawater temperature on a Caribbean reef. Marine Ecology Progress Series, 269:111–119, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p111.pdf.

Edmunds:2007:EDS

[Edm07] P. J. Edmunds. Evidence for a decadal-scale decline in the growth rates of juvenile scleractinian corals. Marine Ecology Progress Series, 341:1–13, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Edeyer:2000:DVS

[EdPP^{+}00] A. Edeyer, H. de Pontual, P. Payan, H. Troadec, A. Sévère, and N. Mayer-Gostan. Daily variations of the saccular endolymph and plasma compositions in the turbot Psetta maxima: relationship with the diurnal rhythm in otolith formation. Marine Ecology Progress Series, 192:287–294, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p287.pdf.

Eckman:2003:CWD

[EDS03] J. E. Eckman, D. O. Duggins, and C. E. Siddon. Current and wave dynamics in the shallow subtidal: implications to the ecology of understory and surface-canopy kelps. Marine Ecology Progress Series, 265:45–56, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p045.pdf.

Etherington:2000:LSB

[EE00] L. L. Etherington and D. B. Eggleston. Large-scale blue crab recruitment: linking postlarval transport, post-settlement planktonic dispersal, and multiple nursery habitats. Marine Ecology Progress Series, 204:179–198, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p179.pdf.

Edwards:2006:CRR

[EE06] M. S. Edwards and J. A. Estes. Catastrophe, recovery and range limitation in NE Pacific kelp forests: a largescale perspective. Marine Ecology Progress Series, 320:79– 87, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p079.pdf.

Elmgren:2001:IFC

[EEA01] R. Elmgren, G. Ejdung, and S. Ankar. Intraspecific food competition in the deposit-feeding benthic amphipod Monoporeia affinis — a laboratory study. Marine Ecology Progress Series, 210:185–193, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p185.pdf.

Ettinger-Epstein:2008:HSC

[EEWBdN08] P. Ettinger-Epstein, S. Whalan, C. N. Battershill, and R. de Nys. A hierarchy of settlement cues influences larval behaviour in a coral reef sponge. Marine Ecology Progress Series, 365:103–113, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p103.pdf.

Eyre:2002:CCP

[EF02] B. D. Eyre and A. J. P. Ferguson. Comparison of carbon production and decomposition, benthic nutrient fluxes and denitrification in seagrass, phytoplankton, benthic microalgae- and macroalgae-dominated warm-temperate Australian lagoons. Marine Ecology Progress Series, 229:43–59, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p043.pdf.

Esteban:2007:ESR

[EF07] G. F. Esteban and B. J. Finlay. Exceptional species richness of ciliated protozoa in pristine intertidal rock pools. Marine Ecology Progress Series, 335:133–141, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p133.pdf.

Eklof:2009:HEM

[EFL^{+}09] J. S. Eklöf, S. Fröcklin, A. Lindvall, N. Stadlinger, A. Kimathi, J. N. Uku, and T. R. McClanahan. How effective are MPAs? Predation control and ‘spill-in effects’ in seagrass-coral reef lagoons under contrasting fishery management. Marine Ecology Progress Series, 384:83–96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m384p083.pdf.

Elsdon:2003:RBW

[EG03] T. S. Elsdon and B. M. Gillanders. Relationship between water and otolith elemental concentrations in juvenile black bream Acanthopagrus butcheri. Marine Ecology Progress Series, 260:263–272, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p263.pdf.

Elsdon:2005:SIC

[EG05] T. S. Elsdon and B. M. Gillanders. Strontium incorporation into calcified structures: separating the effects of ambient water concentration and exposure time. Marine Ecology Progress Series, 285:233–243, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/285/m285p233.pdf.

Edmunds:2008:ATR

[EG08] P. J. Edmunds and R. D. Gates. Acclimatization in tropical reef corals. Marine Ecology Progress Series, 361:307– 310, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p307.pdf.

Enstipp:2007:IFL

[EGJ07] M. R. Enstipp, D. Grémillet, and D. R. Jones. Investigating the functional link between prey abundance and seabird predatory performance. Marine Ecology Progress Series, 331: 267–279, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/331/m331p267.pdf.

Exadactylos:2003:PSD

[EGPT03] A. Exadactylos, A. J. Geffen, P. Panagiotaki, and J. P. Thorpe. Population structure of Dover sole Solea solea: RAPD and allozyme data indicate divergence in European stocks. Marine Ecology Progress Series, 246:253–264, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/246/m246p253.pdf.

Eckrich:2000:TSA

[EH00] C. E. Eckrich and J. G. Holmquist. Trampling in a seagrass assemblage: direct effects, response of associated fauna, and the role of substrate characteristics. Marine Ecology Progress Series, 201:199–209, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/201/m201p199.pdf.

Engelhard:2004:MCN

[EH04] G. H. Engelhard and M. Heino. Maturity changes in Norwegian spring-spawning herring Clupea harengus: compensatory or evolutionary responses? Marine Ecology Progress Series, 272:245–256, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/272/m272p245.pdf.

Eklof:2006:ETO

[EHK06] J. S. Eklöf, R. Henriksson, and N. Kautsky. Effects of tropical open-water seaweed farming on seagrass ecosystem structure and function. Marine Ecology Progress Series, 325: 73–84, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/325/m325p073.pdf.

Escaravage:2009:DPM

[EHM^{+}09] V. Escaravage, P. M. J. Herman, B. Merckx, M. Włodarska-Kowalczuk, J. M. Amouroux, S. Degraer, A. Grémare, C. H. R. Heip, H. Hummel, I. Karakassis, C. Labrune, and W. Willems. Distribution patterns of macrofaunal species diversity in subtidal soft sediments: biodiversity-productivity relationships from the MacroBen database. Marine Ecology Progress Series, 382:253–264, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m382p253.pdf.

Edwards:2007:UDD

[EHWS07] K. P. Edwards, J. A. Hare, F. E. Werner, and H. Seim. Using 2-dimensional dispersal kernels to identify the dominant influences on larval dispersal on continental shelves. Marine Ecology Progress Series, 352:77–87, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/352/
m352p077.pdf.

Eggleston:2003:IRN

[EJKN03] D. B. Eggleston, E. G. Johnson, G. T. Kellison, and D. A. Nadeau. Intense removal and non-saturating functional responses by recreational divers on spiny lobster Panulirus argus. Marine Ecology Progress Series, 257:197–207, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/257/m257p197.pdf.

Eagle:2001:MSS

[EJM01] J. V. Eagle, G. P. Jones, and M. I. McCormick. A multiscale study of the relationships between habitat use and the distribution and abundance patterns of three coral reef angelfishes (Pomacanthidae). Marine Ecology Progress Series, 214:253–265, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/214/m214p253.pdf.

Eden:2003:DRM

[EKA03] N. Eden, T. Katz, and D. L. Angel. Dynamic response of a mud snail Nassarius sinusigerus to changes in sediment biogeochemistry. Marine Ecology Progress Series, 263:139– 147, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/263/m263p139.pdf.

Evrard:2005:NDS

[EKBM05] V. Evrard, W. Kiswara, T. J. Bouma, and J. J. Middelburg. Nutrient dynamics of seagrass ecosystems: ^{15}N evidence for the importance of particulate organic matter and root systems. Marine Ecology Progress Series, 295:49–55, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/295/m295p049.pdf.

Eddie:2008:CGR

[EKN08] B. Eddie, C. Krembs, and S. Neuer. Characterization and growth response to temperature and salinity of psychrophilic, halotolerant Chlamydomonas sp. ARC isolated from Chukchi Sea ice. Marine Ecology Progress Series, 354: 107–117, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p107.pdf.

Eder:2005:PCE

[EL05] E. B. Eder and M. N. Lewis. Proximate composition and energetic value of demersal and pelagic prey species from the SW Atlantic Ocean. Marine Ecology Progress Series, 291: 43–52, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p043.pdf.

Ellingsen:2001:BCS

[Ell01] K. E. Ellingsen. Biodiversity of a continental shelf softsediment macrobenthos community. Marine Ecology Progress Series, 218:1–15, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/218/m218p001.pdf.

Ellingsen:2002:SSB

[Ell02] K. E. Ellingsen. Soft-sediment benthic biodiversity on the continental shelf in relation to environmental variability. Marine Ecology Progress Series, 232:15–27, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p015.pdf.

Erlandsson:2004:SSR

[EM04] J. Erlandsson and C. D. McQuaid. Spatial structure of recruitment in the mussel Perna perna at local scales: effects of adults, algae and recruit size. Marine Ecology Progress Series, 267:173–185, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/267/m267p173.pdf.

Egner:2005:ASS

[EM05] S. A. Egner and D. A. Mann. Auditory sensitivity of sergeant major damselfish Abudefduf saxatilis from post-settlement juvenile to adult. Marine Ecology Progress Series, 285:213– 222, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p213.pdf.

Elkin:2007:DLI

[EM07a] C. Elkin and D. J. Marshall. Desperate larvae: influence of deferred costs and habitat requirements on habitat selection.

Marine Ecology Progress Series, 335:143–153, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p143.pdf.

Etnoyer:2007:PHM

[EM07b] P. Etnoyer and L. E. Morgan. Predictive habitat model for deep gorgonians needs better resolution: Comment on Bryan & Metaxas (2007). Marine Ecology Progress Series, 339:311–312, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p311.pdf. See [BM07c]

and reply [MB07e].

Einbinder:2009:CMD

[EMB^{+}09] S. Einbinder, T. Mass, E. Brokovich, Z. Dubinsky, J. Erez, and D. Tchernov. Changes in morphology and diet of the coral Stylophora pistillata along a depth gradient. Marine Ecology Progress Series, 381:167–174, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/381/m381p167.pdf.

Enriquez:2001:EST

[EMD^{+}01] S. Enríquez, N. Marbà, C. M. Duarte, B. I. van Tussenbroek, and G. Reyes-Zavala. Effects of seagrass Thalassia testudinum on sediment redox. Marine Ecology Progress Series, 219:149–158, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p149.pdf.

Eastwood:2001:MSV

[EMG01] P. D. Eastwood, G. J. Meaden, and A. Grioche. Modelling spatial variations in spawning habitat suitability for the sole Solea solea using regression quantiles and GIS procedures. Marine Ecology Progress Series, 224:251–266, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

224/m224p251.pdf. See erratum [Ano02d].

Epifanio:2007:ASO

[EMPF07] R. A. Epifanio, L. F. Maia, J. R. Pawlik, and W. Fenical. Antipredatory secosterols from the octocoral Pseudopterogorgia americana. Marine Ecology Progress Series, 329:307– 310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p307.pdf.

Endo:2007:FSU

[ENAT07] H. Endo, N. Nakabayashi, Y. Agatsuma, and K. Taniguchi. Food of the sea urchins Strongylocentrotus nudus and Hemicentrotus pulcherrimus associated with vertical distributions in fucoid beds and crustose coralline flats in northern Honshu, Japan. Marine Ecology Progress Series, 352:125–135, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/352/m352p125.pdf.

Enberg:2005:BTS

[Enb05] K. Enberg. Benefits of threshold strategies and age-selective harvesting in a fluctuating fish stock of Norwegian spring spawning herring Clupea harengus. Marine Ecology Progress Series, 298:277–286, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/298/m298p277.pdf.

Ellis:2004:ETS

[ENC^{+}04] J. Ellis, P. Nicholls, R. Craggs, D. Hofstra, and J. Hewitt. Effect of terrigenous sedimentation on mangrove physiology and associated macrobenthic communities. Marine Ecology Progress Series, 270:71–82, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/270/m270p071.pdf.

Enriquez:2005:LAE

[Enr05] S. Enríquez. Light absorption efficiency and the package effect in the leaves of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 289:141–150, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p141.pdf.

Eiane:2004:SSM

[EO04] K. Eiane and M. D. Ohman. Stage-specific mortality of Calanus finmarchicus, Pseudocalanus elongatus and Oithona similis on Fladen Ground, North Sea, during a spring bloom. Marine Ecology Progress Series, 268:183–193, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p183.pdf.

Espeland:2008:NPF

[EOK^{+}08] S. H. Espeland, E. M. Olsen, H. Knutsen, J. Gjøsæter, D. Danielssen, and N. C. Stenseth. New perspectives on fish movement: kernel and GAM smoothers applied to a century of tagging data on coastal Atlantic cod. Marine Ecology Progress Series, 372:231–241, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/372/m372p231.pdf.

Engstrom-Ost:2002:ETC

[EÖKS^{+}02] J. Engström-Öst, M. Koski, K. Schmidt, M. Viitasalo, S. H. Jónasdóttir, M. Kokkonen, S. Repka, and K. Sivonen. Effects of toxic cyanobacteria on a plankton assemblage: community development during decay of Nodularia spumigena. Marine Ecology Progress Series, 232:1–14, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p001.pdf.

Eisert:2005:DFI

[EOL^{+}05] R. Eisert, O. T. Oftedal, M. Lever, S. Ramdohr, B. H. Breier, and G. K. Barrell. Detection of food intake in a marine mammal using marine osmolytes and their analogues as dietary biomarkers. Marine Ecology Progress Series, 300:213– 228, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/300/m300p213.pdf.

Engstrom-Ost:2008:FGH

[EÖM08] J. Engström-Öst and J. Mattila. Foraging, growth and habitat choice in turbid water: an experimental study with fish larvae in the Baltic Sea. Marine Ecology Progress Series, 359:275–281, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p275.pdf.

Engel:2000:AAS

[EP00a] S. Engel and J. R. Pawlik. Allelopathic activities of sponge extracts. Marine Ecology Progress Series, 207:273–281, 2000.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p273.pdf.

Estes:2000:MER

[EP00b] J. A. Estes and C. H. Peterson. Marine ecological research in seashore and seafloor systems: accomplishments and future directions. Marine Ecology Progress Series, 195:281–289, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p281.pdf.

Engel:2001:CNC

[EP01] A. Engel and U. Passow. Carbon and nitrogen content of transparent exopolymer particles (TEP) in relation to their alcian blue adsorption. Marine Ecology Progress Series, 219: 1–10, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/219/m219p001.pdf.

Eggleston:2008:DIS

[EP08] David B. Eggleston and Darren M. Parsons. Disturbanceinduced ‘spill-in’ of Caribbean spiny lobster to marine reserves. Marine Ecology Progress Series, 371:213–220, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/371/m371p213.pdf.

Erlandsson:2006:RCR

[EPM06] J. Erlandsson, P. Pal, and C. D. McQuaid. Re-colonisation rate differs between co-existing indigenous and invasive intertidal mussels following major disturbance. Marine Ecology Progress Series, 320:169–176, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p169.pdf.

Enriquez:2006:EWF

[ERR06] S. Enríquez and A. Rodríguez-Román. Effect of water flow on the photosynthesis of three marine macrophytes from a fringing-reef lagoon. Marine Ecology Progress Series, 323: 119–132, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p119.pdf.

Evans:2004:PGS

[ESB^{+}04] B. S. Evans, N. A. Sweijd, R. C. K. Bowie, P. A. Cook, and N. G. Elliott. Population genetic structure of the perlemoen Haliotis midae in South Africa: evidence of range expansion and founder events. Marine Ecology Progress Series, 270:163–172, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/270/m270p163.pdf.

Eckes:2008:USR

[ESDG08] M. J. Eckes, U. E. Siebeck, S. Dove, and A. S. Grutter. Ultraviolet sunscreens in reef fish mucus. Marine Ecology Progress Series, 353:203–211, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/353/m353p203.pdf.

El-Sabaawi:2009:IVF

[ESDKM09] R. El-Sabaawi, J. F. Dower, M. Kainz, and A. Mazumder. Interannual variability in fatty acid composition of the copepod Neocalanus plumchrus in the Strait of Georgia, British Columbia. Marine Ecology Progress Series, 382:151–161, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/382/m382p151.pdf.

Eissler:2003:EVI

[ESQ03] Y. Eissler, E. Sahlsten, and R. A. Quiñones. Effects of virus infection on respiration rates of marine phytoplankton and microplankton communities. Marine Ecology Progress Series, 262:71–80, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p071.pdf.

Edgar:2009:EEM

[ESS09] G. J. Edgar and R. D. Stuart-Smith. Ecological effects of marine protected areas on rocky reef communities — a continental-scale analysis. Marine Ecology Progress Series, 388:51–62, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/388/m388p051.pdf.

Estrada:2004:MRM

[Est04] Marta Estrada. In memoriam: Ramon Margalef (1919– 2004). Marine Ecology Progress Series, 276:1, 2004. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/misc/
m276p001.pdf.

Erwin:2008:PNS

[ET08] P. M. Erwin and R. W. Thacker. Phototrophic nutrition and symbiont diversity of two Caribbean sponge-cyanobacteria symbioses. Marine Ecology Progress Series, 362:139–147, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/362/m362p139.pdf.

Evans:2008:AUB

[Eva08] R. D. Evans. Assessment of an underwater biopsy probe for collecting teleost fish tissue samples. Marine Ecology Progress Series, 368:305–308, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/368/m368p305.pdf.

Erftemeijer:2009:VTF

[EvBB^{+}09] P. L. A. Erftemeijer, J. K. L. van Beek, L. J. Bolle, M. Dickey-Collas, and H. F. J. Los. Variability in transport of fish eggs and larvae. I. Modelling the effects of coastal reclamation. Marine Ecology Progress Series, 390:167–181, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p167.pdf.

Erftemeijer:2008:ESD

[EvBO^{+}08] P. L. A. Erftemeijer, J. K. L. van Beek, C. A. Ochieng, Z. Jager, and H. J. Los. Eelgrass seed dispersal via floating generative shoots in the Dutch Wadden Sea: a model approach. Marine Ecology Progress Series, 358:115–124, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/358/m358p115.pdf.

Engelhard:2008:FMV

[EvdKB^{+}08] G. H. Engelhard, J. van der Kooij, E. D. Bell, J. K. Pinnegar, J. L. Blanchard, S. Mackinson, and D. A. Righton. Fishing mortality versus natural predation on diurnally migrating sandeels Ammodytes marinus. Marine Ecology Progress Series, 369:213–227, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p213.pdf.

Ehrenhauss:2004:EAP

[EWBH04] S. Ehrenhauss, U. Witte, S. I. Bühring, and M. Huettel. Effect of advective pore water transport on distribution and degradation of diatoms in permeable North Sea sediments. Marine Ecology Progress Series, 271:99–111, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p099.pdf.

Elliott:2008:SFB

[EWG^{+}08] K. H. Elliott, K. Woo, A. J. Gaston, S. Benvenuti, L. Dall’Antonia, and G. K. Davoren. Seabird foraging behaviour indicates prey type. Marine Ecology Progress Series, 354:289–303, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m354p289.pdf.

Ehlers:2008:IGD

[EWR08] A. Ehlers, B. Worm, and T. B. H. Reusch. Importance of genetic diversity in eelgrass Zostera marina for its resilience to global warming. Marine Ecology Progress Series, 355:1–7, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Esser:2008:EHM

[EWSH08] F. Esser, M. Winterberg, Z. Sebesvari, and T. Harder. Effects of halogenated metabolites from infaunal polychaetes on larval settlement of the spionid polychaete Streblospio benedicti. Marine Ecology Progress Series, 355:161–172, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p161.pdf.

Ellis:2006:PMS

[EYH^{+}06] J. Ellis, T. Ysebaert, T. Hume, A. Norkko, T. Bult, P. Herman, S. Thrush, and J. Oldman. Predicting macrofaunal species distributions in estuarine gradients using logistic regression and classification systems. Marine Ecology Progress Series, 316:69–83, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/316/m316p069.pdf.

Eyre:2000:REN

[Eyr00] B. D. Eyre. Regional evaluation of nutrient transformation and phytoplankton growth in nine river-dominated subtropical east Australian estuaries. Marine Ecology Progress Series, 205:61–83, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p061.pdf.

Edvardsen:2002:ZPD

[EZTZ02] A. Edvardsen, M. Zhou, K. S. Tande, and Y. Zhu. Zooplankton population dynamics: measuring in situ growth and mortality rates using an optical plankton counter. Marine Ecology Progress Series, 227:205–219, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p205.pdf.

Findlay:2002:TPS

[FA02] A. M. Findlay and L. G. Allen. Temporal patterns of settlement in the temperate reef fish Paralabrax clathratus. Marine Ecology Progress Series, 238:237–248, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p237.pdf.

Faria:2006:PSD

[FA06] C. Faria and V. C. Almada. Patterns of spatial distribution and behaviour of fish on a rocky intertidal platform at high tide. Marine Ecology Progress Series, 316:155–164, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/316/m316p155.pdf.

Fetzer:2008:RSB

[FA08] I. Fetzer and W. E. Arntz. Reproductive strategies of benthic invertebrates in the Kara Sea (Russian Arctic): adaptation of reproduction modes to cold water. Marine Ecology Progress Series, 356:189–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/356/m356p189.pdf.

Fujikura:2007:LTL

[FAB^{+}07] K. Fujikura, K. Amaki, J. P. Barry, Y. Fujiwara, Y. Furushima, R. Iwase, H. Yamamoto, and T. Maruyama. Longterm in situ monitoring of spawning behavior and fecundity in Calyptogena spp. Marine Ecology Progress Series, 333:185–193, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p185.pdf.

Foggo:2003:EMS

[FAFR03] A. Foggo, M. J. Attrill, M. T. Frost, and A. A. Rowden. Estimating marine species richness: an evaluation of six extrapolative techniques. Marine Ecology Progress Series, 248: 15–26, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p015.pdf.

Feary:2007:CDS

[FAJM07] D. A. Feary, G. R. Almany, G. P. Jones, and M. I. Mc-Cormick. Coral degradation and the structure of tropical reef fish communities. Marine Ecology Progress Series, 333: 243–248, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p243.pdf.

Ford:2007:SNE

[FAK07] R. B. Ford, M. J. Anderson, and S. Kelly. Subtle and negligible effects of rainfall on estuarine infauna: evidence from three years of event-driven sampling. Marine Ecology Progress Series, 340:17–27, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p017.pdf.

Ferron:2009:BNR

[FAPA^{+}09] S. Ferrón, F. Alonso-Pérez, E. Anfuso, F. J. Murillo, T. Ortega, C. G. Castro, and J. M. Forja. Benthic nutrient recycling on the northeastern shelf of the Gulf of Cádiz (SW Iberian Peninsula). Marine Ecology Progress Series, 390: 79–95, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p079.pdf.

Ferron:2009:BRN

[FAPOF09] S. Ferrón, F. Alonso-Pérez, T. Ortega, and J. M. Forja. Benthic respiration on the northeastern shelf of the Gulf of Cádiz (SW Iberian Peninsula). Marine Ecology Progress Series, 392:69–80, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p069.pdf.

Fauchald:2009:SIB

[Fau09] P. Fauchald. Spatial interaction between seabirds and prey: review and synthesis. Marine Ecology Progress Series, 391: 139–151, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m391p139.pdf.

Fernandez:2000:NSU

[FB00] C. Fernandez and C. F. Boudouresque. Nutrition of the sea urchin Paracentrotus lividus (Echinodermata: Echinoidea) fed different artificial food. Marine Ecology Progress Series, 204:131–141, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p131.pdf.

Fulton:2002:OHU

[FB02a] C. J. Fulton and D. R. Bellwood. Ontogenetic habitat use in labrid fishes: an ecomorphological perspective. Marine Ecology Progress Series, 236:255–262, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p255.pdf.

Fulton:2002:PFL

[FB02b] C. J. Fulton and D. R. Bellwood. Patterns of foraging in labrid fishes. Marine Ecology Progress Series, 226:135– 142, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p135.pdf.

Fisher:2003:USB

[FB03a] R. Fisher and D. R. Bellwood. Undisturbed swimming behaviour and nocturnal activity of coral reef fish larvae. Marine Ecology Progress Series, 263:177–188, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p177.pdf.

Fu:2003:ESG

[FB03b] F. X. Fu and P. R. F. Bell. Effect of salinity on growth, pigmentation, N2 fixation and alkaline phosphatase activity of cultured Trichodesmium sp. Marine Ecology Progress Series, 257:69–76, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p069.pdf.

Fox:2007:QHA

[FB07] R. J. Fox and D. R. Bellwood. Quantifying herbivory across a coral reef depth gradient. Marine Ecology Progress Series, 339:49–59, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p049.pdf.

Field:2007:DRA

[FBBH07] I. C. Field, C. J. A. Bradshaw, H. R. Burton, and M. A. Hindell. Differential resource allocation strategies in juvenile elephant seals in the highly seasonal Southern Ocean. Marine Ecology Progress Series, 331:281–290, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p281.pdf.

Figueira:2009:PTF

[FBBV09] W. F. Figueira, P. Biro, D. J. Booth, and V. C. Valenzuela. Performance of tropical fish recruiting to temperate habitats: role of ambient temperature and implications of climate change. Marine Ecology Progress Series, 384:231– 239, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p231.pdf.

Falk:2002:FBT

[FBD^{+}02] K. Falk, S. Benvenuti, L. Dall’Antonia, G. Gilchrist, and K. Kampp. Foraging behaviour of thick-billed murres breeding in different sectors of the North water polynya: an intercolony comparison. Marine Ecology Progress Series, 231: 293–302, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p293.pdf.

Frid:2006:DRD

[FBD06] A. Frid, G. G. Baker, and L. M. Dill. Do resource declines increase predation rates on North Pacific harbor seals? A behavior-based plausibility model. Marine Ecology Progress Series, 312:265–275, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/312/m312p265.pdf.

Franck:2003:IZE

[FBHB03] V. M. Franck, K. W. Bruland, D. A. Hutchins, and M. A. Brzezinski. Iron and zinc effects on silicic acid and nitrate uptake kinetics in three high-nutrient, low-chlorophyll (HNLC) regions. Marine Ecology Progress Series, 252:15– 33, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/252/m252p015.pdf.

Fisher:2000:DSA

[FBJ00a] R. Fisher, D. R. Bellwood, and S. D. Job. Development of swimming abilities in reef fish larvae. Marine Ecology Progress Series, 202:163–173, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/202/m202p163.pdf.

Fowler:2000:DSA

[FBJ00b] A. J. Fowler, K. P. Black, and G. P. Jenkins. Determination of spawning areas and larval advection pathways for King George whiting in southeastern Australia using otolith microstructure and hydrodynamic modelling. II. South Australia. Marine Ecology Progress Series, 199:243–254, 2000.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p243.pdf.

Fong:2003:IIT

[FBKB03] P. Fong, K. E. Boyer, K. Kamer, and K. A. Boyle. Influence of initial tissue nutrient status of tropical marine algae on response to nitrogen and phosphorus additions. Marine Ecology Progress Series, 262:111–123, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p111.pdf.

Friedlander:2007:DRF

[FBM07] A. M. Friedlander, E. Brown, and M. E. Monaco. Defining reef fish habitat utilization patterns in Hawaii: comparisons between marine protected areas and areas open to fishing. Marine Ecology Progress Series, 351:221–233, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p221.pdf.

Foster:2008:CEM

[FBM08] N. L. Foster, S. J. Box, and P. J. Mumby. Competitive effects of macroalgae on the fecundity of the reef-building coral Montastraea annularis. Marine Ecology Progress Series, 367:143–152, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/367/m367p143.pdf.

Fulford:2007:EOP

[FBN^{+}07] R. S. Fulford, D. L. Breitburg, R. I. E. Newell, W. M. Kemp, and M. Luckenbach. Effects of oyster population restoration strategies on phytoplankton biomass in Chesapeake Bay: a flexible modeling approach. Marine Ecology Progress Series, 336:43–61, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p043.pdf.

Fromentin:2002:DNS

[FBSG02] J. M. Fromentin, O. N. Bjørnstad, N. C. Stenseth, and J. Gjøsæter. Dynamics of the Norwegian Skagerrak cod: intuition vs inference. Marine Ecology Progress Series, 241: 231–234, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/241/m241p231.pdf.

Fraschetti:2001:SVH

[FBT^{+}01] S. Fraschetti, C. N. Bianchi, A. Terlizzi, G. Fanelli, C. Morri, and F. Boero. Spatial variability and human disturbance in shallow subtidal hard substrate assemblages: a regional approach. Marine Ecology Progress Series, 212:1–12, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p001.pdf.

Fernandez:2000:RHP

[FC00] M. Fernández and J. C. Castilla. Recruitment of Homalaspis plana in intertidal habitats of central Chile and implications for the current use of Management and Marine Protected Areas. Marine Ecology Progress Series, 208:157–170, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
208/m208p157.pdf.

Fanelli:2004:FHP

[FC04] E. Fanelli and J. E. Cartes. Feeding habits of pandalid shrimps in the Alboran Sea (SW Mediterranean): influence of biological and environmental factors. Marine Ecology Progress Series, 280:227–238, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p227.pdf.

Fanelli:2008:STC

[FC08] E. Fanelli and J. E. Cartes. Spatio-temporal changes in gut contents and stable isotopes in two deep Mediterranean pandalids: influence on the reproductive cycle. Marine Ecology Progress Series, 355:219–233, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/355/m355p219.pdf.

Fonseca:2008:FIE

[FCC^{+}08] P. Fonseca, P. L. Correia, A. Campos, P. Y. Lau, and V. Henriques. Fishery-independent estimation of benthic species density — a novel approach applied to Norway lobster Nephrops norvegicus. Marine Ecology Progress Series, 369:267–271, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p267.pdf.

Fernandez:2006:BBC

[FCCP06] M. Fernández, R. Calderón, M. Cifuentes, and P. Pappalardo. Brooding behaviour and cost of brooding in small body size brachyuran crabs. Marine Ecology Progress Series, 309:213– 220, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/309/m309p213.pdf.

Fry:2008:SIM

[FCH^{+}08] B. Fry, M. Cieri, J. Hughes, C. Tobias, L. A. Deegan, and B. Peterson. Stable isotope monitoring of benthic-planktonic coupling using salt marsh fish. Marine Ecology Progress Series, 369:193–204, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p193.pdf.

Franklin:2009:EDP

[FCH^{+}09] D. J. Franklin, C. J. Choi, C. Hughes, G. Malin, and J. A. Berges. Effect of dead phytoplankton cells on the apparent efficiency of photosystem II. Marine Ecology Progress Series, 382:35–40, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/382/m382p035.pdf.

Forward:2004:SBC

[FCI^{+}04] R. B. Forward, Jr., J. H. Cohen, R. D. Irvine, J. L. Lax, R. Mitchell, A. M. Schick, M. M. Smith, J. M. Thompson, and J. I. Venezia. Settlement of blue crab Callinectes sapidus megalopae in a North Carolina estuary. Marine Ecology Progress Series, 269:237–247, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p237.pdf.

Flores:2002:TSP

[FCP02a] A. A. V. Flores, J. Cruz, and J. Paula. Temporal and spatial patterns of settlement of brachyuran crab megalopae at a rocky coast in Central Portugal. Marine Ecology Progress Series, 229:207–220, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p207.pdf.

Fraser:2002:FFA

[FCP02b] K. P. P. Fraser, A. Clarke, and L. S. Peck. Feast and famine in Antarctica: seasonal physiology in the limpet Nacella concinna. Marine Ecology Progress Series, 242:169– 177, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p169.pdf.

Faure:2007:STD

[FCP^{+}07] B. Faure, P. Chevaldonné, F. Pradillon, E. Thiébaut, and D. Jollivet. Spatial and temporal dynamics of reproduction and settlement in the Pompeii worm Alvinella pompejana (Polychaeta: Alvinellidae). Marine Ecology Progress Series, 348:197–211, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p197.pdf.

Forster:2006:BEF

[FCS^{+}06] R. M. Forster, V. Créach, K. Sabbe, W. Vyverman, and L. J. Stal. Biodiversity-ecosystem function relationship in microphytobenthic diatoms of the Westerschelde estuary. Marine Ecology Progress Series, 311:191–201, 2006. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m311p191.pdf.

Fontes:2009:NSJ

[FCS^{+}09] J. Fontes, J. E. Caselle, M. S. Sheehy, R. S. Santos, and R. R. Warner. Natal signatures of juvenile Coris julis in the Azores: investigating connectivity scenarios in an oceanic archipelago. Marine Ecology Progress Series, 387:51–59, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p051.pdf.

Fabricius:2000:DSP

[FD00] K. E. Fabricius and M. Dommisse. Depletion of suspended particulate matter over coastal reef communities dominated by zooxanthellate soft corals. Marine Ecology Progress Series, 196:157–167, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p157.pdf.

Friedlander:2002:CDS

[FD02] A. M. Friedlander and E. E. DeMartini. Contrasts in density, size, and biomass of reef fishes between the northwestern and the main Hawaiian islands: the effects of fishing down apex predators. Marine Ecology Progress Series, 230: 253–264, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/230/m230p253.pdf.

Forde:2004:MIM

[FD04] S. E. Forde and D. F. Doak. Multitrophic interactions mediate recruitment variability in a rocky intertidal community. Marine Ecology Progress Series, 275:33–45, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p033.pdf.

Ferguson:2008:HGS

[FD08] A. M. Ferguson and A. R. Davis. Heart of glass: spicule armament and physical defense in temperate reef sponges. Marine Ecology Progress Series, 372:77–86, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p077.pdf.

Fehling:2006:SPN

[FDBT06] J. Fehling, K. Davidson, C. Bolch, and P. Tett. Seasonality of Pseudo-nitzschia spp. (Bacillariophyceae) in western Scottish waters. Marine Ecology Progress Series, 323: 91–105, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p091.pdf.

Fiala:2003:MSD

[FDD^{+}03] M. Fiala, B. Delille, C. Dubreuil, E. Kopczynska, K. Leblanc, J. Morvan, B. Quéguiner, S. Blain, C. Cailliau, P. Conan, R. Corvaisier, M. Denis, M. Frankignoulle, L. Oriol, and S. Roy. Mesoscale surface distribution of biogeochemical characteristics in the Crozet Basin frontal zones (South Indian Ocean). Marine Ecology Progress Series, 249:1–14, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/249/m249p001.pdf.

Fanini:2009:THS

[FDDS09] L. Fanini, O. Defeo, C. Do Santos, and F. Scapini. Testing the habitat safety hypothesis with behavioural field experiments: amphipod orientation on sandy beaches with contrasting morphodynamics. Marine Ecology Progress Series, 392:133–141, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p133.pdf.

Fraschetti:2009:DMP

[FDM^{+}09] S. Fraschetti, P. D’Ambrosio, F. Micheli, F. Pizzolante, S. Bussotti, and A. Terlizzi. Design of marine protected areas in a human-dominated seascape. Marine Ecology Progress Series, 375:13–24, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p013.pdf.

Fitzsimons:2005:EET

[FDRR05] M. F. Fitzsimons, M. Dawit, D. M. Revitt, and C. Rocha. Effects of early tidal inundation on the cycling of methylamines in inter-tidal sediments. Marine Ecology Progress Series, 294:51–61, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/294/m294p051.pdf.

Fauchald:2002:SDP

[FE02] P. Fauchald and K. E. Erikstad. Scale-dependent predator– prey interactions: the aggregative response of seabirds to prey under variable prey abundance and patchiness. Marine Ecology Progress Series, 231:279–291, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p279.pdf.

Ford:2006:SFF

[FE06] J. K. B. Ford and G. M. Ellis. Selective foraging by fisheating killer whales Orcinus orca in British Columbia. Marine Ecology Progress Series, 316:185–199, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/316/
m316p185.pdf.

Ferguson:2007:SDD

[FE07] A. J. P. Ferguson and B. D. Eyre. Seasonal discrepancies in denitrification measured by isotope pairing and N_{2}: Ar techniques. Marine Ecology Progress Series, 350:19–27, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m350p019.pdf.

Franco:2008:LSF

[FEFT08] A. Franco, M. Elliott, P. Franzoi, and P. Torricelli. Life strategies of fishes in European estuaries: the functional guild approach. Marine Ecology Progress Series, 354:219– 228, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p219.pdf.

Frederiksen:2007:RAV

[FEMW07] M. Frederiksen, M. Edwards, R. A. Mavor, and S. Wanless. Regional and annual variation in black-legged kittiwake breeding productivity is related to sea surface temperature. Marine Ecology Progress Series, 350:137–143, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/350/
m350p137.pdf.

Fennel:2008:WIC

[Fen08] K. Fennel. Widespread implementation of controlled upwelling in the North Pacific Subtropical Gyre would counteract diazotrophic N_{2} fixation. Marine Ecology Progress Series, 371:301–303, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m371p301.pdf.

Furness:2007:IMP

[FEO07] R. W. Furness, A. E. Edwards, and D. Oro. Influence of management practices and of scavenging seabirds on availability of fisheries discards to benthic scavengers. Marine Ecology Progress Series, 350:235–244, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m350p235.pdf.

Ferrell:2004:GSS

[Fer04] D. L. Ferrell. Gastropod shell size and morphology influence conspecific interactions in an encrusting hydroid. Marine Ecology Progress Series, 275:153–162, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p153.pdf.

Fisher:2002:CFC

[FF02] J. A. D. Fisher and K. T. Frank. Changes in finfish community structure associated with an offshore fishery closed area on the Scotian Shelf. Marine Ecology Progress Series, 240:249–265, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p249.pdf.

Finley:2003:IED

[FF03] R. J. Finley and G. E. Forrester. Impact of ectoparasites on the demography of a small reef fish. Marine Ecology Progress Series, 248:305–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/248/m248p305.pdf.

Fisher:2004:ADR

[FF04] J. A. D. Fisher and K. T. Frank. Abundance-distribution relationships and conservation of exploited marine fishes. Marine Ecology Progress Series, 279:201–213, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p201.pdf.

Foggo:2003:AOP

[FFA03] A. Foggo, M. T. Frost, and M. J. Attrill. Abundanceoccupancy patterns in British estuarine macroinvertebrates. Marine Ecology Progress Series, 265:297–302, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p297.pdf.

Fox:2003:OMI

[FFG03] C. J. Fox, A. Folkvord, and A. J. Geffen. Otolith microincrement formation in herring Clupea harengus larvae in relation to growth rate. Marine Ecology Progress Series, 264:

83–94, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p083.pdf.

Fisher:2007:LRR

[FFHW07] E. M. Fisher, J. E. Fauth, P. Hallock, and C. M. Woodley. Lesion regeneration rates in reef-building corals Montastraea spp. as indicators of colony condition. Marine Ecology Progress Series, 339:61–71, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/339/m339p061.pdf.

Fortier:2002:CBF

[FFML02] M. Fortier, L. Fortier, C. Michel, and L. Legendre. Climatic and biological forcing of the vertical flux of biogenic particles under seasonal Arctic sea ice. Marine Ecology Progress Series, 225:1–16, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p001.pdf.

Fire:2008:PBP

[FFN^{+}08] S. E. Fire, L. J. Flewelling, J. Naar, M. J. Twiner, M. S. Henry, R. H. Pierce, D. P. Gannon, Z. Wang, L. Davidson, and R. S. Wells. Prevalence of brevetoxins in prey fish of bottlenose dolphins in Sarasota Bay, Florida. Marine Ecology Progress Series, 368:283–294, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/368/m368p283.pdf.

Ford:2003:MSI

[FFR^{+}03] A. T. Ford, T. F. Fernandes, S. A. Rider, P. A. Read, C. D. Robinson, and I. M. Davies. Measuring sublethal impacts of pollution on reproductive output of marine crustacea. Marine Ecology Progress Series, 265:303–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p303.pdf.

Fossette:2007:DDP

[FFT^{+}07] S. Fossette, S. Ferraroli, H. Tanaka, Y. Ropert-Coudert, N. Arai, K. Sato, Y. Naito, Y. Le Maho, and J. Y. Georges. Dispersal and dive patterns in gravid leatherback turtles during the nesting season in French Guiana. Marine Ecology Progress Series, 338:233–247, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/338/m338p233.pdf.

Forget:2007:CPP

[FFYS^{+}07] M. H. Forget, C. Fuentes-Yaco, S. Sathyendranath, T. Platt, J. Pommier, and E. Devred. Computation of primary production from remote sensing of ocean colour at the northwestern Atlantic C-SOLAS Lagrangian site. Marine Ecology Progress Series, 352:27–38, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/352/m352p027.pdf.

Forcada:2004:BDA

[FGA^{+}04] J. Forcada, M. Gazo, A. Aguilar, J. Gonzalvo, and M. Fernández-Contreras. Bottlenose dolphin abundance in the NW Mediterranean: addressing heterogeneity in distribution. Marine Ecology Progress Series, 275:275–287, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p275.pdf.

Fabri:2006:CBD

[FGLM06] M. C. Fabri, J. Galéron, M. Larour, and G. Maudire. Combining the biocean database for deep-sea benthic data with the online ocean biogeographic information system. Marine Ecology Progress Series, 316:215–224, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m316p215.pdf.

Frangopulos:2000:STL

[FGM^{+}00] M. Frangópulos, C. Guisande, I. Maneiro, I. Riveiro, and J. Franco. Short-term and long-term effects of the toxic dinoflagellate Alexandrium minutum on the copepod Acartia clausi. Marine Ecology Progress Series, 203:161–169, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p161.pdf.

Fautin:2009:CDD

[FGO09] D. G. Fautin, J. M. Guinotte, and J. C. Orr. Comparative depth distribution of corallimorpharians and scleractinians (Cnidaria: Anthozoa). Marine Ecology Progress Series, 397:63–70, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p063.pdf.

Fabi:2004:EFC

[FGPS04] G. Fabi, F. Grati, M. Puletti, and G. Scarcella. Effects on fish community induced by installation of two gas platforms in the Adriatic Sea. Marine Ecology Progress Series, 273:187–197, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p187.pdf.

Friedrichs:2000:SFI

[FGS00] M. Friedrichs, G. Graf, and B. Springer. Skimming flow induced over a simulated polychaete tube lawn at low population densities. Marine Ecology Progress Series, 192:219–228, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p219.pdf.

Francois:2002:FAS

[FGS^{+}02] F. François, M. Gerino, G. Stora, J. P. Durbec, and J. C. Poggiale. Functional approach to sediment reworking by galleryforming macrobenthic organisms: modeling and application with the polychaete Nereis diversicolor. Marine Ecology Progress Series, 229:127–136, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/229/m229p127.pdf.

Forero:2005:SIR

[FGSH^{+}05] M. G. Forero, J. González-Solís, K. A. Hobson, J. A. Donázar, M. Bertellotti, G. Blanco, and G. R. Bortolotti. Stable isotopes reveal trophic segregation by sex and age in the southern giant petrel in two different food webs. Marine Ecology Progress Series, 296:107–113, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/296/m296p107.pdf.

Fox:2007:BDS

[FGT^{+}07] C. J. Fox, A. J. Geffen, N. Taylor, P. Davison, H. Rossetti, and R. D. M. Nash. Birth-date selection in early life stages of plaice Pleuronectes platessa in the eastern Irish Sea (British Isles). Marine Ecology Progress Series, 345:255–269, 2007.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m345p255.pdf.

Ford:2002:GIM

[FH02] R. B. Ford and C. Honeywill. Grazing on intertidal microphytobenthos by macrofauna: is pheophorbide-a a useful marker? Marine Ecology Progress Series, 229:33–42, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p033.pdf.

Fraser:2003:PPC

[FH03] W. R. Fraser and E. E. Hofmann. A predator’s perspective on causal links between climate change, physical forcing and ecosystem response. Marine Ecology Progress Series, 265:1–15, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p001.pdf.

First:2008:PBB

[FH08a] M. R. First and J. T. Hollibaugh. Protistan bacterivory and benthic microbial biomass in an intertidal creek mudflat. Marine Ecology Progress Series, 361:59–68, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p059.pdf.

Fodrie:2008:TJF

[FH08b] F. J. Fodrie and S. Z. Herzka. Tracking juvenile fish movement and nursery contribution within arid coastal embayments via otolith microchemistry. Marine Ecology Progress Series, 361:253–265, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/361/m361p253.pdf.

Forsterra:2008:USR

[FH08c] G. Försterra and V. Häussermann. Unusual symbiotic relationships between microendolithic phototrophic organisms and azooxanthellate cold-water corals from Chilean fjords. Marine Ecology Progress Series, 370:121–125, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/370/
m370p121.pdf.

Furman:2008:ENE

[FH08d] B. T. Furman and K. L. Heck, Jr. Effects of nutrient enrichment and grazers on coral reefs: an experimental assessment. Marine Ecology Progress Series, 363:89–101, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p089.pdf.

Forero:2002:FRU

[FHB^{+}02] M. G. Forero, K. A. Hobson, G. R. Bortolotti, J. A. Donázar, M. Bertellotti, and G. Blanco. Food resource utilisation by the Magellanic penguin evaluated through stable-isotope analysis: segregation by sex and age and influence on offspring quality. Marine Ecology Progress Series, 234:289–299, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p289.pdf.

Frederiksen:2007:SIS

[FHDA^{+}07] M. S. Frederiksen, M. Holmer, E. Díaz-Almela, N. Marba, and C. M. Duarte. Sulfide invasion in the seagrass Posidonia oceanica at Mediterranean fish farms: assessment using stable sulfur isotopes. Marine Ecology Progress Series, 345: 93–104, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p093.pdf.

Fisk:2003:IHT

[FHG^{+}03] A. T. Fisk, P. F. Hoekstra, J. M. Gagnon, J. Duffe, R. J. Norstrom, K. A. Hobson, M. Kwan, and D. C. G. Muir. Influence of habitat, trophic ecology and lipids on, and spatial trends of, organochlorine contaminants in Arctic marine invertebrates. Marine Ecology Progress Series, 262:201– 214, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p201.pdf.

Franklin:2004:CDD

[FHGJB04] D. J. Franklin, O. Hoegh-Guldberg, R. J. Jones, and J. A. Berges. Cell death and degeneration in the symbiotic dinoflagellates of the coral Stylophora pistillata during bleaching. Marine Ecology Progress Series, 272:117–130, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p117.pdf.

Fujiki:2008:LSO

[FHK^{+}08] T. Fujiki, T. Hosaka, H. Kimoto, T. Ishimaru, and T. Saino. In situ observation of phytoplankton productivity by an underwater profiling buoy system: use of fast repetition rate fluorometry. Marine Ecology Progress Series, 353:81–88, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/353/m353p081.pdf.

Feng:2009:EIP

[FHL^{+}09] Y. Feng, C. E. Hare, K. Leblanc, J. M. Rose, Y. Zhang, G. R. DiTullio, P. A. Lee, S. W. Wilhelm, J. M. Rowe, J. Sun, N. Nemcek, C. Gueguen, U. Passow, I. Benner, C. Brown, and D. A. Hutchins. Effects of increased pCO_{2} and temperature on the North Atlantic spring bloom. I. The phytoplankton community and biogeochemical response. Marine Ecology Progress Series, 388:13–25, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p013.pdf.

Fach:2002:MSA

[FHM02] B. A. Fach, E. E. Hofmann, and E. J. Murphy. Modeling studies of Antarctic krill Euphausia superba survival during transport across the Scotia Sea. Marine Ecology Progress Series, 231:187–203, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p187.pdf.

Friedlaender:2009:DCH

[FHN^{+}09] A. S. Friedlaender, E. L. Hazen, D. P. Nowacek, P. N. Halpin, C. Ware, M. T. Weinrich, T. Hurst, and D. Wiley. Diel changes in humpback whale Megaptera novaeangliae feeding behavior in response to sand lance Ammodytes spp. behavior and distribution. Marine Ecology Progress Series, 395:91–100, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p091.pdf.

Finelli:2007:IVP

[FHPW07] C. M. Finelli, B. S. Helmuth, N. D. Pentcheff, and D. S. Wethey. Intracolony variability in photosynthesis by corals is affected by water flow: role of oxygen flux. Marine Ecology Progress Series, 349:103–110, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p103.pdf.

Friedlaender:2006:WDR

[FHQ^{+}06] A. S. Friedlaender, P. N. Halpin, S. S. Qian, G. L. Lawson, P. H. Wiebe, D. Thiele, and A. J. Read. Whale distribution in relation to prey abundance and oceanographic processes in shelf waters of the Western Antarctic Peninsula. Marine Ecology Progress Series, 317:297–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p297.pdf.

Ferraton:2007:STV

[FHVMDS07] F. Ferraton, M. Harmelin-Vivien, C. Mellon-Duval, and A. Souplet. Spatio-temporal variation in diet may affect condition and abundance of juvenile European hake in the Gulf of Lions (NW Mediterranean). Marine Ecology Progress Series, 337:197–208, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p197.pdf.

Flynn:2009:AII

[FI09] K. J. Flynn and X. Irigoien. Aldehyde-induced insidious effects cannot be considered as a diatom defence mechanism against copepods. Marine Ecology Progress Series, 377: 79–89, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p079.pdf.

Fiedler:2002:ECE

[Fie02] P. C. Fiedler. Environmental change in the eastern tropical Pacific Ocean: review of ENSO and decadal variability. Marine Ecology Progress Series, 244:265–283, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p265.pdf.

Finelli:2000:VCD

[Fin00] C. M. Finelli. Velocity and concentration distributions in turbulent odor plumes in the presence of vegetation mimics: a flume study. Marine Ecology Progress Series, 207: 297–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/207/m207p297.pdf.

Fine:2002:URP

[FIRL02] M. Fine, T. Banin Israely, E. Rosenberg, and Y. Loya. Ultraviolet radiation prevents bleaching in the Mediterranean coral Oculina patagonica. Marine Ecology Progress Series, 226:249–254, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p249.pdf.

Finkel:2004:RLA

[FIS04] Z. V. Finkel, A. J. Irwin, and O. Schofield. Resource limitation alters the 3/4 size scaling of metabolic rates in phytoplankton. Marine Ecology Progress Series, 273:269–279, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/273/
m273p269.pdf.

Fisher:2005:SSL

[Fis05] R. Fisher. Swimming speeds of larval coral reef fishes: impacts on self-recruitment and dispersal. Marine Ecology Progress Series, 285:223–232, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/285/m285p223.pdf.

Fisher:2006:STD

[Fis06] J. L. Fisher. Seasonal timing and duration of brachyuran larvae in a high-latitude fjord. Marine Ecology Progress Series, 323:213–222, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/323/m323p213.pdf.

Franks:2001:MDP

[FJ01] P. J. S. Franks and J. S. Jaffe. Microscale distributions of phytoplankton: initial results from a two-dimensional imaging fluorometer, OSST. Marine Ecology Progress Series, 220:

59–72, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/220/m220p059.pdf.

Fiksen:2007:LBE

[FJK^{+}07] Ó. Fiksen, C. Jørgensen, T. Kristiansen, F. Vikebø, and G. Huse. Linking behavioural ecology and oceanography: larval behaviour determines growth, mortality and dispersal. Marine Ecology Progress Series, 347:195–205, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p195.pdf.

Figueroa:2002:ESU

[FJV^{+}02] F. L. Figueroa, C. Jiménez, B. Viñegla, E. Pérez-Rodríguez, J. Aguilera, A. Flores-Moya, M. Altamirano, M. Lebert, and D. P. Häder. Effects of solar UV radiation on photosynthesis of the marine angiosperm Posidonia oceanica from southern Spain. Marine Ecology Progress Series, 230:59– 70, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/230/m230p059.pdf.

Fukuda:2000:FCP

[FK00] H. Fukuda and I. Koike. Feeding currents of particle-attached nanoflagellates — a novel mechanism for aggregation of submicron particles. Marine Ecology Progress Series, 202:101– 112, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/202/m202p101.pdf.

Forster:2004:MEC

[FK04] R. M. Forster and J. C. Kromkamp. Modelling the effects of chlorophyll fluorescence from subsurface layers on photosynthetic efficiency measurements in microphytobenthic algae. Marine Ecology Progress Series, 284:9–22, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p009.pdf.

Fassler:2008:PMA

[FK08] S. M. M. Fässler and M. J. Kaiser. Phylogenetically mediated anti-predator responses in bivalve molluscs. Marine Ecology Progress Series, 363:217–225, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/363/m363p217.pdf.

Finenko:2006:ICM

[FKA^{+}06] G. A. Finenko, A. E. Kideys, B. E. Anninsky, T. A. Shiganova, A. Roohi, M. R. Tabari, H. Rostami, and S. Bagheri. Invasive ctenophore Mnemiopsis leidyi in the Caspian Sea: feeding, respiration, reproduction and predatory impact on the zooplankton community. Marine Ecology Progress Series, 314:171–185, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p171.pdf.

Fong:2001:NCM

[FKBB01] P. Fong, K. Kamer, K. E. Boyer, and K. A. Boyle. Nutrient content of macroalgae with differing morphologies may indicate sources of nutrients for tropical marine systems. Marine Ecology Progress Series, 220:137–152, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p137.pdf.

Freitas:2009:DIW

[FKI^{+}09] C. Freitas, K. M. Kovacs, R. A. Ims, M. A. Fedak, and C. Lydersen. Deep into the ice: over-wintering and habitat selection in male Atlantic walruses. Marine Ecology Progress Series, 375:247–261, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p247.pdf.

Fujiwara:2001:DSC

[FKM^{+}01] Y. Fujiwara, C. Kato, N. Masui, K. Fujikura, and S. Kojima. Dual symbiosis in the cold-seep thyasirid clam Maorithyas hadalis from the hadal zone in the Japan Trench, western Pacific. Marine Ecology Progress Series, 214:151–159, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p151.pdf.

Findlay:2009:FHC

[FKSW09] H. S. Findlay, M. A. Kendall, J. I. Spicer, and S. Widdicombe. Future high CO_{2} in the intertidal may compromise adult barnacle Semibalanus balanoides survival and embryonic development rate. Marine Ecology Progress Series, 389:193–202, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p193.pdf.

Farstey:2002:EHV

[FLG02] V. Farstey, B. Lazar, and A. Genin. Expansion and homogeneity of the vertical distribution of zooplankton in a very deep mixed layer. Marine Ecology Progress Series, 238:91– 100, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p091.pdf.

Fistarol:2003:AEP

[FLG03] G. O. Fistarol, C. Legrand, and E. Granéli. Allelopathic effect of Prymnesium parvum on a natural plankton community. Marine Ecology Progress Series, 255:115–125, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p115.pdf.

Fan:2006:DPL

[FLK^{+}06] T. Y. Fan, K. H. Lin, F. W. Kuo, K. Soong, L. L. Liu, and L. S. Fang. Diel patterns of larval release by five brooding scleractinian corals. Marine Ecology Progress Series, 321: 133–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p133.pdf.

Fodrie:2009:UPF

[FLL09] F. J. Fodrie, L. A. Levin, and A. J. Lucas. Use of population fitness to evaluate the nursery function of juvenile habitats. Marine Ecology Progress Series, 385:39–49, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p039.pdf.

Fauchot:2005:DNV

[FLR05] J. Fauchot, M. Levasseur, and S. Roy. Daytime and nighttime vertical migrations of Alexandrium tamarense in the St. Lawrence estuary (Canada). Marine Ecology Progress Series, 296:241–250, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p241.pdf.

Fernandez:2004:REM

[FLUF^{+}04] D. Fernández, Á. López-Urrutia, A. Fernández, J. L. Acuña, and R. Harris. Retention efficiency of 0.2 to 6 μ m particles by the appendicularians Oikopleura dioica and Fritillaria borealis. Marine Ecology Progress Series, 266:89– 101, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/266/m266p089.pdf.

Fiksen:2002:PBM

[FM02] Ó. Fiksen and B. R. MacKenzie. Process-based models of feeding and prey selection in larval fish. Marine Ecology Progress Series, 243:151–164, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/243/m243p151.pdf.

Faugeron:2005:LTC

[FMCB05] S. Faugeron, E. A. Martínez, J. A. Correa, and C. Billot. Long-term copper mine waste disposal in northern Chile associated with gene flow disruption of the intertidal kelp Lessonia nigrescens. Marine Ecology Progress Series, 288:129– 140, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/288/m288p129.pdf.

Fernandez:2003:PCU

[FMMS03] E. Fernández, E. Marañón, X. A. G. Morán, and P. Serret. Potential causes for the unequal contribution of picophytoplankton to total biomass and productivity in oligotrophic waters. Marine Ecology Progress Series, 254:101–109, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p101.pdf.

Fredensborg:2005:ITH

[FMP05] B. L. Fredensborg, K. N. Mouritsen, and R. Poulin. Impact of trematodes on host survival and population density in the intertidal gastropod Zeacumantus subcarinatus. Marine Ecology Progress Series, 290:109–117, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/290/m290p109.pdf.

Freeman:2009:HPC

[FMT09] D. J. Freeman, A. B. MacDiarmid, and R. B. Taylor. Habitat patches that cross marine reserve boundaries: consequences for the lobster Jasus edwardsii. Marine Ecology Progress Series, 388:159–167, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p159.pdf.

Frederiksen:2007:SEI

[FMW07] M. Frederiksen, R. A. Mavor, and S. Wanless. Seabirds as environmental indicators: the advantages of combining data sets. Marine Ecology Progress Series, 352:205–211, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m352p205.pdf.

Fach:2008:BBM

[FMWGB08] B. A. Fach, B. Meyer, D. Wolf-Gladrow, and U. Bathmann. Biochemically based modeling study of Antarctic krill Euphausia superba growth and development. Marine Ecology Progress Series, 360:147–161, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p147.pdf.

Fock:2002:DHD

[FMZvW02] H. O. Fock, B. Matthiessen, H. Zidowitz, and H. v. Westernhagen. Diel and habitat-dependent resource utilisation of deep-sea fishes at the Great Meteor seamount (subtropical NE Atlantic): niche overlap and support for the sound-scattering layer-interception hypothesis. Marine Ecology Progress Series, 244:219–233, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p219.pdf.

Finke:2007:TRT

[FNB07] G. R. Finke, S. A. Navarrete, and F. Bozinovic. Tidal regimes of temperate coasts and their influences on aerial exposure for intertidal organisms. Marine Ecology Progress Series, 343:57–62, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p057.pdf.

Fritz:2008:RAP

[FNDP08] J. J. Fritz, P. J. Neale, R. F. Davis, and J. A. Peloquin. Response of Antarctic phytoplankton to solar UVR exposure: inhibition and recovery of photosynthesis in coastal and pelagic assemblages. Marine Ecology Progress Series, 365:1–16, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Faria:2009:OCS

[FOFG09] A. M. Faria, A. F. Ojanguren, L. A. Fuiman, and E. J. Gonçalves. Ontogeny of critical swimming speed of wildcaught and laboratory-reared red drum Sciaenops ocellatus larvae. Marine Ecology Progress Series, 384:221–230, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/384/m384p221.pdf.

Fukumori:2008:SIR

[FOH^{+}08] K. Fukumori, N. Okuda, H. Hamaoka, T. Fukumoto, D. Takahashi, and K. Omori. Stable isotopes reveal life history polymorphism in the coastal fish Apogon notatus. Marine Ecology Progress Series, 362:279–289, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p279.pdf.

Fine:2002:BER

[FOL02] M. Fine, U. Oren, and Y. Loya. Bleaching effect on regeneration and resource translocation in the coral Oculina patagonica. Marine Ecology Progress Series, 234:119–125, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p119.pdf.

Fujikura:2003:SME

[FON03] K. Fujikura, K. Okoshi, and T. Naganuma. Strontium as a marker for estimation of microscopic growth rates in a bivalve. Marine Ecology Progress Series, 257:295–301, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/257/
m257p295.pdf.

Fox:2004:CRB

[Fox04] H. E. Fox. Coral recruitment in blasted and unblasted sites in Indonesia: assessing rehabilitation potential. Marine Ecology Progress Series, 269:131–139, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p131.pdf.

Fossheim:2008:HCM

[FP08] M. Fossheim and R. Primicerio. Habitat choice by marine zooplankton in a high-latitude ecosystem. Marine Ecology Progress Series, 364:47–56, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/364/m364p047.pdf.

Fernandez:2002:POS

[FPB02] M. Fernández, L. M. Pardo, and J. A. Baeza. Patterns of oxygen supply in embryo masses of brachyuran crabs throughout development: the effect of oxygen availability and chemical cues in determining female brooding behavior. Marine Ecology Progress Series, 245:181–190, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p181.pdf.

Fawcett:2007:CWP

[FPB^{+}07] A. Fawcett, G. C. Pitcher, S. Bernard, A. D. Cembella, and R. M. Kudela. Contrasting wind patterns and toxigenic phytoplankton in the southern Benguela upwelling system. Marine Ecology Progress Series, 348:19–31, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p019.pdf.

Ferrier-Pages:2002:KSU

[FPBAT02] C. Ferrier-Pagès, F. Boisson, D. Allemand, and E. Tambutté. Kinetics of strontium uptake in the scleractinian coral Stylophora pistillata. Marine Ecology Progress Series, 245:93– 100, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p093.pdf.

Falcon:2005:GKM

[FPC05] L. I. Falcón, S. Pluvinage, and E. J. Carpenter. Growth kinetics of marine unicellular N_{2}-fixing cyanobacterial isolates in continuous culture in relation to phosphorus and temperature. Marine Ecology Progress Series, 285:3–9, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p003.pdf.

Falk-Petersen:2002:LBT

[FPDS^{+}02] S. Falk-Petersen, T. M. Dahl, C. L. Scott, J. R. Sargent, B. Gulliksen, S. Kwasniewski, H. Hop, and R. M. Millar. Lipid biomarkers and trophic linkages between ctenophores and copepods in Svalbard waters. Marine Ecology Progress Series, 227:187–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/227/m227p187.pdf.

Ferrier-Pages:2001:PNB

[FPF01] C. Ferrier-Pagès and P. Furla. Pico- and nanoplankton biomass and production in the two largest atoll lagoons of French Polynesia. Marine Ecology Progress Series, 211:63– 76, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/211/m211p063.pdf.

Ferraris:2005:AIR

[FPKC05] J. Ferraris, D. Pelletier, M. Kulbicki, and C. Chauvet. Assessing the impact of removing reserve status on the Abore Reef fish assemblage in New Caledonia. Marine Ecology Progress Series, 292:271–286, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/292/m292p271.pdf.

Finenko:2003:VPM

[FPWM03] Z. Z. Finenko, S. A. Piontkovski, R. Williams, and A. V. Mishonov. Variability of phytoplankton and mesozooplankton biomass in the subtropical and tropical Atlantic Ocean. Marine Ecology Progress Series, 250:125–144, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/250/
m250p125.pdf.

Finlay:2006:OGR

[FR06] K. Finlay and J. C. Roff. Ontogenetic growth rate responses of temperate marine copepods to chlorophyll concentration and light. Marine Ecology Progress Series, 313: 145–156, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/313/m313p145.pdf.

FernandezI:2007:NRA

[FR07] Camilla Fernández I. and Patrick Raimbault. Nitrogen regeneration in the NE Atlantic Ocean and its impact on seasonal new, regenerated and export production. Marine Ecology Progress Series, 337:79–92, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m337p079.pdf.

Freestone:2008:OIF

[FR08a] D. Freestone and R. Rayfuse. Ocean iron fertilization and international law. Marine Ecology Progress Series, 364:227– 233, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m364p227.pdf.

Fujii:2008:SLR

[FR08b] T. Fujii and D. Raffaelli. Sea-level rise, expected environmental changes, and responses of intertidal benthic macrofauna in the Humber estuary, UK. Marine Ecology Progress Series, 371:23–35, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/371/m371p023.pdf.

Franks:2005:PPT

[Fra05] P. J. S. Franks. Plankton patchiness, turbulent transport and spatial spectra. Marine Ecology Progress Series, 294: 295–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/294/m294p295.pdf.

Fredriksen:2003:FWS

[Fre03] S. Fredriksen. Food web studies in a Norwegian kelp forest based on stable isotope (δ ^{13}C and δ ^{1}5N) analysis. Marine Ecology Progress Series, 260:71–81, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p071.pdf.

Freeman:2007:SID

[Fre07] A. S. Freeman. Specificity of induced defenses in Mytilus edulis and asymmetrical predator deterrence. Marine Ecology Progress Series, 334:145–153, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/334/m334p145.pdf.

Foden:2009:RRU

[FRJ09] J. Foden, S. I. Rogers, and A. P. Jones. Recovery rates of UK seabed habitats after cessation of aggregate extraction. Marine Ecology Progress Series, 390:15–26, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p015.pdf.

Frouin:2000:EAD

[Fro00] P. Frouin. Effects of anthropogenic disturbances of tropical soft-bottom benthic communities. Marine Ecology Progress Series, 194:39–53, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p039.pdf.

Finn:2002:FMS

[FRvdMF02] R. N. Finn, I. Rønnestad, T. van der Meeren, and H. J. Fyhn. Fuel and metabolic scaling during the early life stages of Atlantic cod Gadus morhua. Marine Ecology Progress Series, 243:217–234, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/243/m243p217.pdf.

Forrester:2002:TEO

[FS02] G. E. Forrester and S. E. Swearer. Trace elements in otoliths indicate the use of open-coast versus bay nursery habitats by juvenile California halibut. Marine Ecology Progress Series, 241:201–213, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p201.pdf.

Faunce:2006:MFH

[FS06] C. H. Faunce and J. E. Serafy. Mangroves as fish habitat: 50 years of field studies. Marine Ecology Progress Series, 318: 1–18, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Fox:2007:NPD

[FS07] C. H. Fox and A. K. Swanson. Nested PCR detection of microscopic life-stages of laminarian macroalgae and comparison with adult forms along intertidal height gradients. Marine Ecology Progress Series, 332:1–10, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Faunce:2008:SUM

[FS08] C. H. Faunce and J. E. Serafy. Selective use of mangrove shorelines by snappers, grunts, and great barracuda. Marine Ecology Progress Series, 356:153–162, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/356/
m356p153.pdf.

Fisher:2007:TOB

[FSB07] R. Fisher, S. M. Sogard, and S. A. Berkeley. Trade-offs between size and energy reserves reflect alternative strategies for optimizing larval survival potential in rockfish. Marine Ecology Progress Series, 344:257–270, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p257.pdf.

Ferner:2005:CCR

[FSC05] M. C. Ferner, D. L. Smee, and Y. P. Chang. Cannibalistic crabs respond to the scent of injured conspecifics: danger or dinner? Marine Ecology Progress Series, 300:193– 200, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/300/m300p193.pdf.

Fox:2009:EEF

[FSHB09] R. J. Fox, T. L. Sunderland, A. S. Hoey, and D. R. Bellwood. Estimating ecosystem function: contrasting roles of closely related herbivorous rabbitfishes (Siganidae) on coral reefs. Marine Ecology Progress Series, 385:261–269, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p261.pdf.

Fulton:2003:ECM

[FSJ03] E. A. Fulton, A. D. M. Smith, and C. R. Johnson. Effect of complexity on marine ecosystem models. Marine Ecology Progress Series, 253:1–16, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/253/m253p001.pdf.

Ferber:2008:GDC

[FSO08] S. Ferber, W. T. Stam, and J. L. Olsen. Genetic diversity and connectivity remain high in eelgrass Zostera marina populations in the Wadden Sea, despite major impacts. Marine Ecology Progress Series, 372:87–96, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p087.pdf.

Feinberg:2006:LDE

[FSP06] L. R. Feinberg, C. T. Shaw, and W. T. Peterson. Larval development of Euphausia pacifica in the laboratory: variability in developmental pathways. Marine Ecology Progress Series, 316:127–137, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/316/m316p127.pdf.

Fisher:2000:TMM

[FSSW^{+}00] N. S. Fisher, I. Stupakoff, S. Sañudo-Wilhelmy, W. X. Wang, J. L. Teyssié, S. W. Fowler, and J. Crusius. Trace metals in marine copepods: a field test of a bioaccumulation model coupled to laboratory uptake kinetics data. Marine Ecology Progress Series, 194:211–218, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/194/m194p211.pdf.

Franco:2008:DVD

[FSV^{+}08] M. A. Franco, K. Soetaert, D. Van Oevelen, D. Van Gansbeke, M. J. Costa, M. Vincx, and J. Vanaverbeke. Density, vertical distribution and trophic responses of metazoan meiobenthos to phytoplankton deposition in contrasting sediment types. Marine Ecology Progress Series, 358:51–62, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p051.pdf.

Fields:2002:MNR

[FSW02] D. M. Fields, D. S. Shaeffer, and M. J. Weissburg. Mechanical and neural responses from the mechanosensory hairs on the antennule of Gaussia princeps. Marine Ecology Progress Series, 227:173–186, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p173.pdf.

Ferner:2009:HCA

[FSW09] M. C. Ferner, D. L. Smee, and M. J. Weissburg. Habitat complexity alters lethal and non-lethal olfactory interactions between predators and prey. Marine Ecology Progress Series, 374:13–22, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/374/m374p013.pdf.

Furness:2000:SFI

[FT00] R. W. Furness and M. L. Tasker. Seabird-fishery interactions: quantifying the sensitivity of seabirds to reductions in sandeel abundance, and identification of key areas for sensitive seabirds in the North Sea. Marine Ecology Progress Series, 202:253–264, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p253.pdf.

Fischer:2009:EET

[FTB09] S. Fischer, S. Thatje, and T. Brey. Early egg traits in Cancer setosus (Decapoda, Brachyura): effects of temperature and female size. Marine Ecology Progress Series, 377:193– 202, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p193.pdf.

Fraschetti:2005:PDM

[FTBC05] S. Fraschetti, A. Terlizzi, and L. Benedetti-Cecchi. Patterns of distribution of marine assemblages from rocky shores: evidence of relevant scales of variation. Marine Ecology Progress Series, 296:13–29, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p013.pdf.

Farnsworth:2007:HES

[FTDK07] K. D. Farnsworth, U. H. Thygesen, S. Ditlevsen, and N. J. King. How to estimate scavenger fish abundance using baited camera data. Marine Ecology Progress Series, 350:223–234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
theme/m350p223.pdf.

Fujiki:2004:PLP

[FTK^{+}04] T. Fujiki, T. Toda, T. Kikuchi, H. Aono, and S. Taguchi. Phosphorus limitation of primary productivity during the spring-summer blooms in Sagami Bay, Japan. Marine Ecology Progress Series, 283:29–38, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p029.pdf.

Fox:2009:RBC

[FTO^{+}09] S. E. Fox, M. Teichberg, Y. S. Olsen, L. Heffner, and I. Valiela. Restructuring of benthic communities in eutrophic estuaries: lower abundance of prey leads to trophic shifts from omnivory to grazing. Marine Ecology Progress Series, 380: 43–57, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/380/m380p043.pdf.

Farina:2009:SMS

[FTP^{+}09] S. Farina, F. Tomas, P. Prado, J. Romero, and T. Alcoverro. Seagrass meadow structure alters interactions between the sea urchin Paracentrotus lividus and its predators. Marine Ecology Progress Series, 377:131–137, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/377/m377p131.pdf.

Fujiwara:2000:PCE

[FTU^{+}00] Y. Fujiwara, K. Takai, K. Uematsu, S. Tsuchida, J. C. Hunt, and J. Hashimoto. Phylogenetic characterization of endosymbionts in three hydrothermal vent mussels: influence on host distributions. Marine Ecology Progress Series, 208:147–155, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p147.pdf.

Fernandez-Vergaz:2000:MFS

[FVAB00] V. Fernández-Vergaz, L. J. López Abellán, and E. Balguerías. Morphometric, functional and sexual maturity of the deepsea red crab Chaceon affinis inhabiting Canary Island waters: chronology of maturation. Marine Ecology Progress Series, 204:169–178, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p169.pdf.

Forcada:2009:EHS

[FVB^{+}09] A. Forcada, C. Valle, P. Bonhomme, G. Criquet, G. Cadiou, P. Lenfant, and J. L. Sánchez-Lizaso. Effects of habitat on spillover from marine protected areas to artisanal fisheries. Marine Ecology Progress Series, 379:197–211, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m379p197.pdf.

Flores:2008:DAE

[FVS^{+}08] H. Flores, A. P. Van de Putte, V. Siegel, E. A. Pakhomov, J. A. van Franeker, E. H. W. G. Meesters, and F. A. M. Volckaert. Distribution, abundance and ecological relevance of pelagic fishes in the Lazarev Sea, Southern Ocean. Marine Ecology Progress Series, 367:271–282, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p271.pdf.

Fowler-Walker:2002:OSS

[FWC02] M. J. Fowler-Walker and S. D. Connell. Opposing states of subtidal habitat across temperate Australia: consistency and predictability in kelp canopy-benthic associations. Marine Ecology Progress Series, 240:49–56, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p049.pdf.

Friedland:2001:OOO

[FWD^{+}01] K. D. Friedland, R. V. Walker, N. D. Davis, K. W. Myers, G. W. Boehlert, S. Urawa, and Y. Ueno. Open-ocean orientation and return migration routes of chum salmon based on temperature data from data storage tags. Marine Ecology Progress Series, 216:235–252, 2001. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/216/m216p235.pdf.

Finn:2002:DPY

[FWF02] R. N. Finn, M. Wamboldt, and H. J. Fyhn. Differential processing of yolk proteins during oocyte hydration in marine fishes (Labridae) that spawn benthic and pelagic eggs. Marine Ecology Progress Series, 237:217–226, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p217.pdf.

Frederiksen:2005:RPK

[FWH^{+}05] M. Frederiksen, P. J. Wright, M. P. Harris, R. A. Mavor, M. Heubeck, and S. Wanless. Regional patterns of kittiwake Rissa tridactyla breeding success are related to variability in sandeel recruitment. Marine Ecology Progress Series, 300:201–211, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p201.pdf.

Fossheim:2005:IBB

[FZT^{+}05] M. Fossheim, M. Zhou, K. S. Tande, O. P. Pedersen, Y. Zhu, and A. Edvardsen. Interactions between biological and environmental structures along the coast of northern Norway. Marine Ecology Progress Series, 300:147–158, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/300/
m300p147.pdf.

Gimenez:2003:LPE

[GA03] L. Giménez and K. Anger. Larval performance in an estuarine crab, Chasmagnathus granulata, is a consequence of both larval and embryonic experience. Marine Ecology Progress Series, 249:251–264, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p251.pdf.

Galst:2008:FHA

[GA08a] C. J. Galst and T. W. Anderson. Fish-habitat associations and the role of disturbance in surfgrass beds. Marine Ecology Progress Series, 365:177–186, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/365/m365p177.pdf.

Gochfeld:2008:ACD

[GA08b] D. J. Gochfeld and G. S. Aeby. Antibacterial chemical defenses in Hawaiian corals provide possible protection from disease. Marine Ecology Progress Series, 362:119–128, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p119.pdf.

Goni:2008:SSW

[GAAB^{+}08] R. Goñi, S. Adlerstein, D. Alvarez-Berastegui, A. Forcada, O. Reñones, G. Criquet, S. Polti, G. Cadiou, C. Valle, P. Lenfant, P. Bonhomme, A. Pérez-Ruzafa, J. L. Sánchez-Lizaso, J. A. García-Charton, G. Bernard, V. Stelzenmüller, and S. Planes. Spillover from six western Mediterranean marine protected areas: evidence from artisanal fisheries. Marine Ecology Progress Series, 366:159–174, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m366p159.pdf.

Gillanders:2003:ECB

[GAB^{+}03] B. M. Gillanders, K. W. Able, J. A. Brown, D. B. Eggleston, and P. F. Sheridan. Evidence of connectivity between juvenile and adult habitats for mobile marine fauna: an important component of nurseries. Marine Ecology Progress Series, 247:281–295, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p281.pdf.

Gonzalez:2002:LSV

[GAM02] N. González, R. Anadón, and E. Marañón. Large-scale variability of planktonic net community metabolism in the Atlantic Ocean: importance of temporal changes in oligotrophic subtropical waters. Marine Ecology Progress Series, 233: 21–30, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p021.pdf.

Garcia-Aguilar:2005:EHP

[GAMB05] M. C. García-Aguilar and E. Morales-Bojórquez. Estimating the haul-out population size of a colony of northern elephant seals Mirounga angustirostris in Mexico, based on mark-recapture data. Marine Ecology Progress Series, 297: 297–302, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/297/m297p297.pdf.

Gardner:2000:WMC

[Gar00] J. P. A. Gardner. Where are the mussels on Cook Strait (New Zealand) shores? Low seston quality as a possible factor limiting multi-species distributions. Marine Ecology Progress Series, 194:123–132, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p123.pdf.

Gago:2003:PPB

[GÁSPR03] J. Gago, X. A. Álvarez-Salgado, F. F. Pérez, and A. F. Ríos. Partitioning of physical and biogeochemical contributions to short-term variability of pCO_{2} in a coastal upwelling system: a quantitative approach. Marine Ecology Progress Series, 255:43–54, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p043.pdf.

Goffredo:2002:SRM

[GAZ02] S. Goffredo, S. Arnone, and F. Zaccanti. Sexual reproduction in the Mediterranean solitary coral Balanophyllia europaea (Scleractinia, Dendrophylliidae). Marine Ecology Progress Series, 229:83–94, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p083.pdf.

Gibbons:2001:STV

[GB01] M. J. Gibbons and E. Buecher. Short-term variability in the assemblage of medusae and ctenophores following upwelling events in the southern Benguela ecosystem. Marine Ecology Progress Series, 220:169–177, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/220/m220p169.pdf.

Goffredi:2002:SSV

[GB02a] S. K. Goffredi and J. P. Barry. Species-specific variation in sulfide physiology between closely related vesicomyid clams. Marine Ecology Progress Series, 225:227–238, 2002. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p227.pdf.

Groeneveld:2002:LDM

[GB02b] J. C. Groeneveld and G. M. Branch. Long-distance migration of South African deep-water rock lobster Palinurus gilchristi. Marine Ecology Progress Series, 232:225–238, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p225.pdf.

Gobler:2003:IAI

[GB03] C. J. Gobler and G. E. Boneillo. Impacts of anthropogenically influenced groundwater seepage on water chemistry and phytoplankton dynamics within a coastal marine system. Marine Ecology Progress Series, 255:101–114, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p101.pdf.

Gregson:2005:ZPA

[GB05a] M. A. Gregson and D. J. Booth. Zooplankton patchiness and the associated shoaling response of the temperate reef fish Trachinops taeniatus. Marine Ecology Progress Series, 299:269–275, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p269.pdf.

Grove:2005:GRG

[GB05b] M. Grove and D. L. Breitburg. Growth and reproduction of gelatinous zooplankton exposed to low dissolved oxygen. Marine Ecology Progress Series, 301:185–198, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/301/
m301p185.pdf.

Guppy:2006:EEB

[GB06] R. Guppy and J. C. Bythell. Environmental effects on bacterial diversity in the surface mucus layer of the reef coral Montastraea faveolata. Marine Ecology Progress Series, 328: 133–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p133.pdf.

Godley:2007:ESS

[GB07] J. Godley and R. Brodie. Effect of summer storms on early life stages of Uca minax, U. pugnax and U. pugilator in North Inlet Estuary, South Carolina, USA. Marine Ecology Progress Series, 342:197–204, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/342/m342p197.pdf.

Gremillet:2009:SEC

[GB09] D. Grémillet and T. Boulinier. Spatial ecology and conservation of seabirds facing global climate change: a review. Marine Ecology Progress Series, 391:121–137, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m391p121.pdf.

Giere:2001:TSW

[GBB^{+}01] Olav Giere, Ferdinando Boero, Erik Bonsdorff, Tom Fenchel, Carlo Heip, Peter Herman, Jack Middelburg, and Gerd Liebezeit. Theme section: Who needs symposium proceedings? Marine Ecology Progress Series, 211:305–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p305.pdf.

Gazeau:2005:NEM

[GBB^{+}05] F. Gazeau, A. V. Borges, C. Barrón, C. M. Duarte, N. Iversen, J. J. Middelburg, B. Delille, M. D. Pizay, M. Frankignoulle, and J. P. Gattuso. Net ecosystem metabolism in a micro-tidal estuary (Randers Fjord, Denmark): evaluation of methods. Marine Ecology Progress Series, 301:23–41, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/301/
m301p023.pdf.

Guidetti:2004:LRS

[GBC^{+}04] P. Guidetti, C. N. Bianchi, M. Chiantore, S. Schiaparelli, C. Morri, and R. Cattaneo-Vietti. Living on the rocks: substrate mineralogy and the structure of subtidal rocky substrate communities in the Mediterranean Sea. Marine Ecology Progress Series, 274:57–68, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p057.pdf.

Guizien:2006:DOF

[GBD^{+}06] K. Guizien, T. Brochier, J. C. Duchêne, B. S. Koh, and P. Marsaleix. Dispersal of Owenia fusiformis larvae by winddriven currents: turbulence, swimming behaviour and mortality in a three-dimensional stochastic model. Marine Ecology Progress Series, 311:47–66, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/311/m311p047.pdf.

Gaertner:2005:LSS

[GBdS^{+}05] J. C. Gaertner, J. A. Bertrand, L. G. de Sola, J. P. Durbec, E. Ferrandis, and A. Souplet. Large spatial scale variation of demersal fish assemblage structure on the continental shelf of the NW Mediterranean Sea. Marine Ecology Progress Series, 297:245–257, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/297/m297p245.pdf.

Godley:2002:RSS

[GBF^{+}02] B. J. Godley, A. C. Broderick, R. Frauenstein, F. Glen, and G. C. Hays. Reproductive seasonality and sexual dimorphism in green turtles. Marine Ecology Progress Series, 226:125–133, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p125.pdf.

Georges:2000:FHD

[GBG00a] J. Y. Georges, F. Bonadonna, and C. Guinet. Foraging habitat and diving activity of lactating subantarctic fur seals in relation to sea-surface temperatures at Amsterdam Island. Marine Ecology Progress Series, 196:291–304, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p291.pdf.

Guarini:2000:DMS

[GBG^{+}00b] J. M. Guarini, G. F. Blanchard, P. Gros, D. Gouleau, and C. Bacher. Dynamic model of the short-term variability of microphytobenthic biomass on temperate intertidal mudflats. Marine Ecology Progress Series, 195:291–303, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p291.pdf.

Godley:2002:TDS

[GBGH02] B. J. Godley, A. C. Broderick, F. Glen, and G. C. Hays. Temperature-dependent sex determination of Ascension Island green turtles. Marine Ecology Progress Series, 226: 115–124, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p115.pdf.

Garrido:2008:DFI

[GBHO^{+}08] S. Garrido, R. Ben-Hamadou, P. B. Oliveira, M. E. Cunha, M. A. Chícharo, and C. D. van der Lingen. Diet and feeding intensity of sardine Sardina pilchardus: correlation with satellite-derived chlorophyll data. Marine Ecology Progress Series, 354:245–256, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/354/m354p245.pdf.

Guinotte:2006:MHD

[GBI^{+}06] J. M. Guinotte, J. D. Bartley, A. Iqbal, D. G. Fautin, and R. W. Buddemeier. Modeling habitat distribution from organism occurrences and environmental data: case study using anemonefishes and their sea anemone hosts. Marine Ecology Progress Series, 316:269–283, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m316p269.pdf.

Godley:2001:EHS

[GBM01] B. J. Godley, A. C. Broderick, and N. Mrosovsky. Estimating hatchling sex ratios of loggerhead turtles in Cyprus from incubation duration. Marine Ecology Progress Series, 210:195–201, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p195.pdf.

Garcias-Bonet:2008:ESS

[GBMHD08] N. Garcias-Bonet, N. Marbà, M. Holmer, and C. M. Duarte. Effects of sediment sulfides on seagrass Posidonia oceanica meristematic activity. Marine Ecology Progress Series, 372: 1–6, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gullstrom:2008:SSC

[GBNÖ08] M. Gullström, M. Bodin, P. G. Nilsson, and M. C. Öhman. Seagrass structural complexity and landscape configuration as determinants of tropical fish assemblage composition. Marine Ecology Progress Series, 363:241–255, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p241.pdf.

Guichard:2001:SIT

[GBR01] F. Guichard, E. Bourget, and J. L. Robert. Scaling the influence of topographic heterogeneity on intertidal benthic communities: alternate trajectories mediated by hydrodynamics and shading. Marine Ecology Progress Series, 217:27–41, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p027.pdf.

Galbraith:2004:ETE

[GBR^{+}04] P. S. Galbraith, Howard I. Browman, R. G. Racca, A. B. Skiftesvik, and J. F. Saint-Pierre. Effect of turbulence on the energetics of foraging in Atlantic cod Gadus morhua larvae. Marine Ecology Progress Series, 281:241–257, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p241.pdf.

Gascon:2009:PSS

[GBSQ09] S. Gascón, D. Boix, J. Sala, and X. D. Quintana. Patterns in size and species diversity of benthic macroinvertebrates in Mediterranean salt marshes. Marine Ecology Progress Series, 391:21–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/391/m391p021.pdf.

Grover:2002:ENE

[GBW02] J. J. Grover, T. W. Buckley, and D. Woodbury. Effects of the 1997–1998 El Niño on early-juvenile Pacific hake Merluccius productus: age, growth, abundance, and diet in coastal nursery habitats. Marine Ecology Progress Series, 240:235– 247, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/240/m240p235.pdf.

Green:2005:BFD

[GBW^{+}05] J. A. Green, I. L. Boyd, A. J. Woakes, N. L. Warren, and P. J. Butler. Behavioural flexibility during year-round foraging in macaroni penguins. Marine Ecology Progress Series, 296:183–196, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p183.pdf.

Green:2007:FFF

[GBW^{+}07] J. A. Green, I. L. Boyd, A. J. Woakes, C. J. Green, and P. J. Butler. Feeding, fasting and foraging success during chick rearing in macaroni penguins. Marine Ecology Progress Series, 346:299–312, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/346/m346p299.pdf.

Glasby:2001:OPS

[GC01] T. M. Glasby and S. D. Connell. Orientation and position of substrata have large effects on epibiotic assemblages. Marine Ecology Progress Series, 214:127–135, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p127.pdf.

Goodsell:2002:CHL

[GC02] P. J. Goodsell and S. D. Connell. Can habitat loss be treated independently of habitat configuration? Implications for rare and common taxa in fragmented landscapes. Marine Ecology Progress Series, 239:37–44, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/239/m239p037.pdf.

Goulet:2003:SOA

[GC03] T. L. Goulet and M. A. Coffroth. Stability of an octocoralalgal symbiosis over time and space. Marine Ecology Progress Series, 250:117–124, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/250/m250p117.pdf.

Goodsell:2005:HCH

[GC05] P. J. Goodsell and S. D. Connell. Historical configuration of habitat influences the effects of disturbance on mobile invertebrates. Marine Ecology Progress Series, 299:79– 87, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p079.pdf.

Glassom:2006:RGM

[GC06a] D. Glassom and N. E. Chadwick. Recruitment, growth and mortality of juvenile corals at Eilat, northern Red Sea. Marine Ecology Progress Series, 318:111–122, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p111.pdf.

Guest:2006:MCA

[GC06b] M. A. Guest and R. M. Connolly. Movement of carbon among estuarine habitats: the influence of saltmarsh patch size. Marine Ecology Progress Series, 310:15–24, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p015.pdf.

Green:2007:MEV

[GC07] B. S. Green and R. C. Chambers. Maternal effects vary between source populations in the Atlantic tomcod Microgadus tomcod. Marine Ecology Progress Series, 344:185– 195, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/344/m344p185.pdf.

Gevaert:2003:LSP

[GCD^{+}03] F. Gévaert, A. Créach, D. Davoult, A. Migné, G. Levavasseur, P. Arzel, A. C. Holl, and Y. Lemoine. Laminaria saccharina photosynthesis measured in situ: photoinhibition and xanthophyll cycle during a tidal cycle. Marine Ecology Progress Series, 247:43–50, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p043.pdf.

Goldthwait:2005:EPF

[GCHA05] S. A. Goldthwait, C. A. Carlson, G. K. Henderson, and A. L. Alldredge. Effects of physical fragmentation on remineralization of marine snow. Marine Ecology Progress Series, 305:59–65, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p059.pdf.

Grenz:2000:DNC

[GCHC00] C. Grenz, J. E. Cloern, S. W. Hager, and B. E. Cole. Dynamics of nutrient cycling and related benthic nutrient and oxygen fluxes during a spring phytoplankton bloom in South San Francisco Bay (USA). Marine Ecology Progress Series, 197:67–80, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p067.pdf.

Guest:2004:CMA

[GCL04] M. A. Guest, R. M. Connolly, and N. R. Loneragan. Carbon movement and assimilation by invertebrates in estuarine habitats at a scale of metres. Marine Ecology Progress Series, 278:27–34, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p027.pdf.

Gauthier-Clerc:2002:SRU

[GCLC^{+}02] M. Gauthier-Clerc, Y. Le Maho, Y. Clerquin, C. A. Bost, and Y. Handrich. Seabird reproduction in an unpredictable environment: how King penguins provide their young chicks with food. Marine Ecology Progress Series, 237:291–300, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/237/m237p291.pdf.

Gust:2001:SVR

[GCM01] N. Gust, J. H. Choat, and M. I. McCormick. Spatial variability in reef fish distribution, abundance, size and biomass: a multi-scale analysis. Marine Ecology Progress Series, 214: 237–251, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p237.pdf.

Grorud-Colvert:2006:ICB

[GCS06] K. Grorud-Colvert and S. Sponaugle. Influence of condition on behavior and survival potential of a newly settled coral reef fish, the bluehead wrasse Thalassoma bifasciatum. Marine Ecology Progress Series, 327:279–288, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p279.pdf.

Gonzalez-Correa:2007:POM

[GCSSJV07] J. M. González-Correa, J. T. Bayle Sempere, P. Sánchez-Jerez, and C. Valle. Posidonia oceanica meadows are not declining globally. Analysis of population dynamics in marine protected areas of the Mediterranean Sea. Marine Ecology Progress Series, 336:111–119, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p111.pdf.

Goodsell:2007:DBB

[GCU07] P. J. Goodsell, M. G. Chapman, and A. J. Underwood. Differences between biota in anthropogenically fragmented habitats and in naturally patchy habitats. Marine Ecology Progress Series, 351:15–23, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p015.pdf.

Guisande:2001:ECR

[GCVR01] C. Guisande, J. M. Cabanas, A. R. Vergara, and I. Riveiro. Effect of climate on recruitment success of Atlantic Iberian sardine Sardina pilchardus. Marine Ecology Progress Series, 223:243–250, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p243.pdf.

Gimenez:2007:SSC

[GD07] L. Giménez and S. Dick. Settlement of shore crab Carcinus maenas on a mesotidal open habitat as a function of transport mechanisms. Marine Ecology Progress Series, 338:

159–168, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p159.pdf.

Grober-Dunsmore:2006:PIR

[GDBF06] R. Grober-Dunsmore, V. Bonito, and T. K. Frazer. Potential inhibitors to recovery of Acropora palmata populations in St. John, US Virgin Islands. Marine Ecology Progress Series, 321:123–132, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p123.pdf.

Grober-Dunsmore:2007:DSR

[GDBF07] R. Grober-Dunsmore, V. Bonito, and T. K. Frazer. Discernment of sexual recruits is not critical for assessing population recovery of Acropora palmata. Marine Ecology Progress Series, 335:233–236, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/335/m335p233.pdf.

Gattuso:2005:PPE

[GDDM05] J. P. Gattuso, N. A. Dawson, C. M. Duarte, and J. J. Middelburg. Patterns of publication effort in coastal biogeochemistry: a bibliometric survey (1971 to 2003). Marine Ecology Progress Series, 294:9–22, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/294/m294p009.pdf.

Guinet:2007:KWP

[GDdS^{+}07] C. Guinet, P. Domenici, R. de Stephanis, L. Barrett-Lennard, J. K. B. Ford, and P. Verborgh. Killer whale predation on bluefin tuna: exploring the hypothesis of the enduranceexhaustion technique. Marine Ecology Progress Series, 347: 111–119, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/347/m347p111.pdf.

Godinez-Dominguez:2003:ITA

[GDF03] E. Godínez-Domínguez and J. Freire. Information-theoretic approach for selection of spatial and temporal models of community organization. Marine Ecology Progress Series, 253: 17–24, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p017.pdf.

Gannier:2002:DRA

[GDG02] A. Gannier, V. Drouot, and J. C. Goold. Distribution and relative abundance of sperm whales in the Mediterranean Sea. Marine Ecology Progress Series, 243:281–293, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p281.pdf.

Gomez-Diaz:2009:PSH

[GDGSP09] E. Gómez-Díaz, J. González-Solís, and M. A. Peinado. Population structure in a highly pelagic seabird, the Cory’s shearwater Calonectris diomedea: an examination of genetics, morphology and ecology. Marine Ecology Progress Series, 382: 197–209, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/382/m382p197.pdf.

Grol:2008:MSB

[GDKN08] M. G. G. Grol, M. Dorenbosch, E. M. G. Kokkelmans, and I. Nagelkerken. Mangroves and seagrass beds do not enhance growth of early juveniles of a coral reef fish. Marine Ecology Progress Series, 366:137–146, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/366/m366p137.pdf.

Guinet:2001:SDF

[GDL^{+}01] C. Guinet, L. Dubroca, M. A. Lea, S. Goldsworthy, Y. Cherel, G. Duhamel, F. Bonadonna, and J. P. Donnay. Spatial distribution of foraging in female Antarctic fur seals Arctocephalus gazella in relation to oceanographic variables: a scale-dependent approach using geographic information systems. Marine Ecology Progress Series, 219:251–264, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p251.pdf.

Griffen:2004:PRR

[GDL04] B. D. Griffen, T. H. DeWitt, and C. Langdon. Particle removal rates by the mud shrimp Upogebia pugettensis, its burrow, and a commensal clam: effects on estuarine phytoplankton abundance. Marine Ecology Progress Series, 269: 223–236, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/269/m269p223.pdf.

Gerpe:2009:SAH

[GdLB^{+}09] M. S. Gerpe, A. Ponce de León, R. Bastida, V. J. Moreno, and D. H. Rodríguez. Sharp accumulation of heavy metals after weaning in the South American fur seal Arctocephalus australis. Marine Ecology Progress Series, 375:239– 245, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/375/m375p239.pdf.

Gremare:2004:FBF

[GDR^{+}04a] A. Grémare, J. C. Duchêne, R. Rosenberg, E. David, and M. Desmalades. Feeding behaviour and functional response of Abra ovata and A. nitida compared by image analysis. Marine Ecology Progress Series, 267:195–208, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p195.pdf.

Gremillet:2004:ODH

[GDR^{+}04b] D. Grémillet, G. Dell’Omo, P. G. Ryan, G. Peters, Y. Ropert-Coudert, and S. J. Weeks. Offshore diplomacy or how seabirds mitigate intra-specific competition: a case study based on GPS tracking of Cape gannets from neighbouring colonies. Marine Ecology Progress Series, 268:265–279, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p265.pdf.

Gordillo:2002:NPU

[GDS02] F. J. L. Gordillo, M. J. Dring, and G. Savidge. Nitrate and phosphate uptake characteristics of three species of brown algae cultured at low salinity. Marine Ecology Progress Series, 234:111–118, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p111.pdf.

Garrigue:2004:OGC

[GDSB04] C. Garrigue, R. Dodemont, D. Steel, and C. S. Baker. Organismal and ’gametic’ capture-recapture using microsatellite genotyping confirm low abundance and reproductive autonomy of humpback whales on the wintering grounds of New Caledonia. Marine Ecology Progress Series, 274:251–262, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p251.pdf.

Griffiths:2006:RHC

[GDW06] S. P. Griffiths, A. R. Davis, and R. J. West. Role of habitat complexity in structuring temperate rockpool ichthyofaunas. Marine Ecology Progress Series, 313:227–239, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p227.pdf.

Green:2008:IRA

[GEC08] D. H. Green, P. J. Edmunds, and R. C. Carpenter. Increasing relative abundance of Porites astreoides on Caribbean reefs mediated by an overall decline in coral cover. Marine Ecology Progress Series, 359:1–10, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gorokhova:2007:NAL

[GEHZ07] E. Gorokhova, A. Edlund, S. Hajdu, and E. N. Zhivotova. Nucleic acid levels in copepods: dynamic response to phytoplankton blooms in the northern Baltic proper. Marine Ecology Progress Series, 349:213–225, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p213.pdf.

Giraldo:2002:SDC

[GEM02] A. Giraldo, R. Escribano, and V. Marin. Spatial distribution of Calanus chilensis off Mejillones Peninsula (northern Chile): ecological consequences upon coastal upwelling. Marine Ecology Progress Series, 230:225–234, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p225.pdf.

Gilly:2006:SJS

[GES^{+}06] W. F. Gilly, C. A. Elliger, C. A. Salinas, S. Camarilla-Coop, G. Bazzino, and M. Beman. Spawning by jumbo squid Dosidicus gigas in San Pedro Mártir Basin, Gulf of California, Mexico. Marine Ecology Progress Series, 313:125– 133, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/313/m313p125.pdf.

Gwyther:2002:CEM

[GF02] J. Gwyther and P. G. Fairweather. Colonisation by epibionts and meiofauna of real and mimic pneumatophores in a cool temperate mangrove habitat. Marine Ecology Progress Series, 229:137–149, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p137.pdf.

Gregoire:2004:NBN

[GF04] M. Grégoire and J. Friedrich. Nitrogen budget of the northwestern Black Sea shelf inferred from modeling studies and in situ benthic measurements. Marine Ecology Progress Series, 270:15–39, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/270/m270p015.pdf.

Gerrodette:2005:NRT

[GF05] T. Gerrodette and J. Forcada. Non-recovery of two spotted and spinner dolphin populations in the eastern tropical Pacific Ocean. Marine Ecology Progress Series, 291:1–21, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Georges:2006:EBM

[GF06] J. Y. Georges and S. Fossette. Estimating body mass in leatherback turtles Dermochelys coriacea. Marine Ecology Progress Series, 318:255–262, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/318/m318p255.pdf.

Georges:2007:MAM

[GFB^{+}07] J. Y. Georges, S. Fossette, A. Billes, S. Ferraroli, J. Fretey, D. Grémillet, Y. Le Maho, A. E. Myers, H. Tanaka, and G. C.

Hays. Meta-analysis of movements in Atlantic leatherback turtles during the nesting season: conservation implications. Marine Ecology Progress Series, 338:225–232, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/338/
m338p225.pdf.

Guisande:2002:FPS

[GFC^{+}02] C. Guisande, M. Frangópulos, Y. Carotenuto, I. Maneiro, I. Riveiro, and A. R. Vergara. Fate of paralytic shellfish poisoning toxins ingested by the copepod Acartia clausi. Marine Ecology Progress Series, 240:105–115, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p105.pdf.

Galvan:2008:SIA

[GFF08] K. Galván, J. W. Fleeger, and B. Fry. Stable isotope addition reveals dietary importance of phytoplankton and microphytobenthos to saltmarsh infauna. Marine Ecology Progress Series, 359:37–49, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p037.pdf.

Griscom:2002:KMA

[GFL02] S. B. Griscom, N. S. Fisher, and S. N. Luoma. Kinetic modeling of Ag, Cd and Co bioaccumulation in the clam Macoma balthica: quantifying dietary and dissolved sources. Marine Ecology Progress Series, 240:127–141, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p127.pdf.

Guisande:2002:EAT

[GFM^{+}02] C. Guisande, M. Frangópulos, I. Maneiro, A. R. Vergara, and I. Riveiro. Ecological advantages of toxin production by the dinoflagellate Alexandrium minutum under phosphorus limitation. Marine Ecology Progress Series, 225:169–176, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p169.pdf.

Guest:2009:TEF

[GFN^{+}09] M. A. Guest, S. D. Frusher, P. D. Nichols, C. R. Johnson, and K. E. Wheatley. Trophic effects of fishing southern rock lobster Jasus edwardsii shown by combined fatty acid and stable isotope analyses. Marine Ecology Progress Series, 388:169–184, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p169.pdf.

Gibbs:2005:BNF

[GFP^{+}05] M. Gibbs, G. Funnell, S. Pickmere, A. Norkko, and J. Hewitt. Benthic nutrient fluxes along an estuarine gradient: influence of the pinnid bivalve Atrina zelandica in summer. Marine Ecology Progress Series, 288:151–164, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p151.pdf.

Gurney:2001:TPT

[GFPM01] L. J. Gurney, P. W. Froneman, E. A. Pakhomov, and C. D. McQuaid. Trophic positions of three euphausiid species from the Prince Edward Islands (Southern Ocean): implications for the pelagic food web structure. Marine Ecology Progress Series, 217:167–174, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/217/m217p167.pdf.

Grossart:2009:HPA

[GG09] H. P. Grossart and G. Gust. Hydrostatic pressure affects physiology and community structure of marine bacteria during settling to 4000 m: an experimental approach. Marine Ecology Progress Series, 390:97–104, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m390p097.pdf.

Garrison:2005:SIM

[GGC^{+}05] D. L. Garrison, A. Gibson, S. L. Coale, M. M. Gowing, Y. B. Okolodkov, C. H. Fritsen, and M. O. Jeffries. Sea-ice microbial communities in the Ross Sea: autumn and summer biota. Marine Ecology Progress Series, 300:39–52, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/300/
m300p039.pdf.

Guinan:2009:QRB

[GGDB09] J. Guinan, A. J. Grehan, M. F. J. Dolan, and C. Brown. Quantifying relationships between video observations of cold-water coral cover and seafloor features in Rockall Trough, west of Ireland. Marine Ecology Progress Series, 375:125–138, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m375p125.pdf.

Gomez-Gutierrez:2000:HEA

[GGDHRA00] J. Gómez-Gutiérrez, E. Domínguez-Hernández, C. J. Robinson, and V. Arenas. Hydroacoustical evidence of autumn inshore residence of the pelagic red crab Pleuroncodes planipes at Punta Eugenia, Baja California, Mexico. Marine Ecology Progress Series, 208:283–291, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/208/m208p283.pdf.

Gomez-Gutierrez:2006:VBS

[GGFSP06] J. Gómez-Gutiérrez, L. R. Feinberg, T. Shaw, and W. T. Peterson. Variability in brood size and female length of Euphausia pacifica among three populations in the North Pacific. Marine Ecology Progress Series, 323:185–194, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p185.pdf.

Gowing:2004:BVA

[GGG^{+}04] M. M. Gowing, D. L. Garrison, A. H. Gibson, J. M. Krupp, M. O. Jeffries, and C. H. Fritsen. Bacterial and viral abundance in Ross Sea summer pack ice communities. Marine Ecology Progress Series, 279:3–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/279/m279p003.pdf.

Guerra-Garcia:2003:AHP

[GGGVGG03] J. M. Guerra-García, F. J. González-Vila, and J. C. García-Gómez. Aliphatic hydrocarbon pollution and macrobenthic assemblages in Ceuta harbour: a multivariate approach. Marine Ecology Progress Series, 263:127–138, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p127.pdf.

Geffard:2002:RBM

[GGHA02] A. Geffard, O. Geffard, E. His, and J. C. Amiard. Relationships between metal bioaccumulation and metallothionein levels in larvae of Mytilus galloprovincialis exposed to contaminated estuarine sediment elutriate. Marine Ecology Progress Series, 233:131–142, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/233/m233p131.pdf.

Garneau:2007:NRP

[GGK^{+}07] M. Garneau, M. Gosselin, B. Klein, J. Tremblay, and E. Fouilland. New and regenerated production during a late summer bloom in an Arctic polynya. Marine Ecology Progress Series, 345:13–26, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/345/m345p013.pdf.

Gutierrez:2000:EDO

[GGM^{+}00] D. Gutiérrez, V. A. Gallardo, S. Mayor, C. Neira, C. Vásquez, J. Sellanes, M. Rivas, A. Soto, F. Carrasco, and M. Baltazar. Effects of dissolved oxygen and fresh organic matter on the bioturbation potential of macrofauna in sublittoral sediments off Central Chile during the 1997/1998 El Niño. Marine Ecology Progress Series, 202:81–99, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/202/m202p081.pdf.

Gomez-Gutierrez:2007:ITH

[GGMGR07] J. Gómez-Gutiérrez, S. Martínez-Gómez, and C. J. Robinson. Influence of thermo-haline fronts forced by tides on near-surface zooplankton aggregation and community structure in Bahía Magdalena, Mexico. Marine Ecology Progress Series, 346:109–125, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/346/m346p109.pdf.

Gomez-Gutierrez:2006:TCT

[GGR06] J. Gómez-Gutiérrez and C. J. Robinson. Tidal current transport of epibenthic swarms of the euphausiid Nyctiphanes simplex in a shallow, subtropical bay on Baja California peninsula, México. Marine Ecology Progress Series, 320:215– 231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p215.pdf.

Guenther:2008:PTS

[GGRC^{+}08] M. Guenther, E. Gonzalez-Rodriguez, W. F. Carvalho, C. E. Rezende, G. Mugrabe, and J. L. Valentin. Plankton trophic structure and particulate organic carbon production during a coastal downwelling-upwelling cycle. Marine Ecology Progress Series, 363:109–119, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/363/m363p109.pdf.

Goldstien:2009:CCI

[GGS09] S. J. Goldstien, N. J. Gemmell, and D. R. Schiel. Colonisation and connectivity by intertidal limpets among New Zealand, Chatham and Sub-Antarctic Islands. I. Genetic connections. Marine Ecology Progress Series, 388:111–119, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/388/
m388p111.pdf.

Guidi-Guilvard:2009:DBC

[GGTKG09] L. D. Guidi-Guilvard, D. Thistle, A. Khripounoff, and S. Gasparini. Dynamics of benthic copepods and other meiofauna in the benthic boundary layer of the deep NW Mediterranean Sea. Marine Ecology Progress Series, 396:181–195, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p181.pdf.

Gaymer:2008:KPS

[GH08] C. F. Gaymer and J. H. Himmelman. A keystone predatory sea star in the intertidal zone is controlled by a higher-order predatory sea star in the subtidal zone. Marine Ecology Progress Series, 370:143–153, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/370/m370p143.pdf.

Gerdts:2002:FFA

[GHD^{+}02] G. Gerdts, C. Hummert, G. Donner, B. Luckas, and C. Schütt. A fast fluorimetric assay (FFA) for the detection of saxitoxin in natural phytoplankton samples. Marine Ecology Progress Series, 230:29–34, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p029.pdf.

Guenther:2007:PCT

[GHdN07] J. Guenther, K. Heimann, and R. de Nys. Pedicellariae of the crown-of-thorns sea star Acanthaster planci are not an effective defence against fouling. Marine Ecology Progress Series, 340:101–108, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p101.pdf.

Gailhard:2002:VPM

[GhGD^{+}02] I. Gailhard, P. h. Gros, J. P. Durbec, B. Beliaeff, C. Belin, E. Nézan, and P. Lassus. Variability patterns of microphytoplankton communities along the French coasts. Marine Ecology Progress Series, 242:39–50, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/242/
m242p039.pdf.

Glibert:2004:EDO

[GHH^{+}04] P. M. Glibert, C. A. Heil, D. Hollander, M. Revilla, A. Hoare, J. Alexander, and S. Murasko. Evidence for dissolved organic nitrogen and phosphorus uptake during a cyanobacterial bloom in Florida Bay. Marine Ecology Progress Series, 280:73–83, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p073.pdf.

Gaymer:2002:EII

[GHJ02] C. F. Gaymer, J. H. Himmelman, and L. E. Johnson. Effect of intra- and interspecific interactions on the feeding behavior of two subtidal sea stars. Marine Ecology Progress Series, 232:149–162, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/232/m232p149.pdf.

Goffart:2002:CDW

[GHL02] A. Goffart, J. H. Hecq, and L. Legendre. Changes in the development of the winter-spring phytoplankton bloom in the Bay of Calvi (NW Mediterranean) over the last two decades: a response to changing climate? Marine Ecology Progress Series, 236:45–60, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p045.pdf.

Gilfillan:2002:CPA

[GHP02] E. S. Gilfillan, E. J. Harner, and D. S. Page. Comment on Peterson et al. (2001) ’Sampling design begets conclusions’. Marine Ecology Progress Series, 231:303–308, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/

m231p303.pdf. See [PMGE01].

Grossart:2003:MDD

[GHP03] H. P. Grossart, S. Hietanen, and H. Ploug. Microbial dynamics on diatom aggregates in Øresund, Denmark. Marine Ecology Progress Series, 249:69–78, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/249/m249p069.pdf.

Grift:2007:TDM

[GHR^{+}07] R. E. Grift, M. Heino, A. D. Rijnsdorp, S. B. M. Kraak, and U. Dieckmann. Three-dimensional maturation reaction norms for North Sea plaice. Marine Ecology Progress Series, 334:213–224, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p213.pdf.

Granberg:2005:RIM

[GHS05] M. E. Granberg, R. Hansen, and H. Selck. Relative importance of macrofaunal burrows for the microbial mineralization of pyrene in marine sediments: impact of macrofaunal species and organic matter quality. Marine Ecology Progress Series, 288:59–74, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/288/m288p059.pdf.

Goldsworthy:2001:TIB

[GHT^{+}01] S. D. Goldsworthy, X. He, G. N. Tuck, M. Lewis, and R. Williams. Trophic interactions between the Patagonian toothfish, its fishery, and seals and seabirds around Macquarie Island. Marine Ecology Progress Series, 218:283–302, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p283.pdf.

Gage:2002:SSI

[GHV02] J. D. Gage, D. J. Hughes, and J. L. Gonzalez Vecino. Sieve size influence in estimating biomass, abundance and diversity in samples of deep-sea macrobenthos. Marine Ecology Progress Series, 225:97–107, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p097.pdf.

Goecker:2005:ENC

[GHV05] M. E. Goecker, K. L. Heck, Jr., and J. F. Valentine. Effects of nitrogen concentrations in turtlegrass Thalassia testudinum on consumption by the bucktooth parrotfish Sparisoma radians. Marine Ecology Progress Series, 286:239– 248, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/286/m286p239.pdf.

Gillanders:2002:CBJ

[Gil02a] B. M. Gillanders. Connectivity between juvenile and adult fish populations: do adults remain near their recruitment estuaries? Marine Ecology Progress Series, 240:215–223, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/240/m240p215.pdf.

Gilmour:2002:SAR

[Gil02b] J. P. Gilmour. Substantial asexual recruitment of mushroom corals contributes little to population genetics of adults in conditions of chronic sedimentation. Marine Ecology Progress Series, 235:81–91, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p081.pdf.

Gimenez:2004:MCE

[Gim04] L. Giménez. Marine community ecology: importance of traitmediated effects propagating through complex life cycles. Marine Ecology Progress Series, 283:303–310, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p303.pdf.

Gimenez:2009:URE

[Gim09] L. Giménez. Uncertainty relationships in the estimation of settlement rates of mobile marine organisms. Marine Ecology Progress Series, 377:103–111, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/377/m377p103.pdf.

Gislason:2005:SSV

[Gis05] A. Gislason. Seasonal and spatial variability in egg production and biomass of Calanus finmarchicus around Iceland. Marine Ecology Progress Series, 286:177–192, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p177.pdf.

Graneli:2003:ETH

[GJ03] E. Granéli and N. Johansson. Effects of the toxic haptophyte Prymnesium parvum on the survival and feeding of a ciliate: the influence of different nutrient conditions. Marine Ecology Progress Series, 254:49–56, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/254/m254p049.pdf.

Gardiner:2005:HSO

[GJ05] N. M. Gardiner and G. P. Jones. Habitat specialisation and overlap in a guild of coral reef cardinalfishes (Apogonidae). Marine Ecology Progress Series, 305:163–175, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/305/
m305p163.pdf.

Green:2004:AGL

[GJB04] J. Green, R. Jones, and S. Brownell. Age and growth of larval cod and haddock on Georges Bank during 1995 and 1996. Marine Ecology Progress Series, 283:255–268, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p255.pdf.

Garrison:2003:DSI

[GJG^{+}03] D. L. Garrison, M. O. Jeffries, A. Gibson, S. L. Coale, D. Neenan, C. Fritsen, Y. B. Okolodkov, and M. M. Gowing. Development of sea ice microbial communities during autumn ice formation in the Ross Sea. Marine Ecology Progress Series, 259:1–15, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p001.pdf.

Garbary:2009:PEM

[GJT09] D. J. Garbary, M. M. Jamieson, and B. R. Taylor. Population ecology of the marine insect Halocladius variabilis (Diptera: Chironomidae) in the rocky intertidal zone of Nova Scotia, Canada. Marine Ecology Progress Series, 376:193– 202, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p193.pdf.

Gillanders:2000:EFO

[GK00a] B. M. Gillanders and M. J. Kingsford. Elemental fingerprints of otoliths of fish may distinguish estuarine ‘nursery’ habitats. Marine Ecology Progress Series, 201:273–286, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p273.pdf.

Gust:2000:LSS

[GK00b] G. Gust and H. P. Kozerski. In situ sinking-particle flux from collection rates of cylindrical traps. Marine Ecology Progress Series, 208:93–106, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p093.pdf.

Gribsholt:2002:EBP

[GK02] B. Gribsholt and E. Kristensen. Effects of bioturbation and plant roots on salt marsh biogeochemistry: a mesocosm study. Marine Ecology Progress Series, 241:71–87, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p071.pdf.

Gray:2003:VTD

[GK03] C. A. Gray and M. J. Kingsford. Variability in thermocline depth and strength, and relationships with vertical distributions of fish larvae and mesozooplankton in dynamic coastal waters. Marine Ecology Progress Series, 247:211– 224, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p211.pdf.

Goebel:2007:TSV

[GK07] N. L. Goebel and J. N. Kremer. Temporal and spatial variability of photosynthetic parameters and community respiration in Long Island Sound. Marine Ecology Progress Series, 329:23–42, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/329/m329p023.pdf.

Goetze:2008:HMS

[GK08] E. Goetze and T. Kiørboe. Heterospecific mating and species recognition in the planktonic marine copepods Temora stylifera and T. longicornis. Marine Ecology Progress Series, 370:185–198, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/370/m370p185.pdf.

Gotz:2008:EFP

[GKAS08] A. Götz, S. E. Kerwath, C. G. Attwood, and W. H. H. Sauer. Effects of fishing on population structure and life history of roman Chrysoblephus laticeps (Sparidae). Marine Ecology Progress Series, 362:245–259, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p245.pdf.

Gradinger:2009:PRS

[GKB09] R. R. Gradinger, M. R. Kaufman, and B. A. Bluhm. Pivotal role of sea ice sediments in the seasonal development of near-shore Arctic fast ice biota. Marine Ecology Progress Series, 394:49–63, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/394/m394p049.pdf.

Gribsholt:2003:IFC

[GKK03] B. Gribsholt, J. E. Kostka, and E. Kristensen. Impact of fiddler crabs and plant roots on sediment biogeochemistry in a Georgia saltmarsh. Marine Ecology Progress Series, 259:237–251, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p237.pdf.

Granum:2002:CEP

[GKM02] E. Granum, S. Kirkvold, and S. M. Myklestad. Cellular and extracellular production of carbohydrates and amino acids by the marine diatom Skeletonema costatum: diel variations and effects of N depletion. Marine Ecology Progress Series, 242:83–94, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p083.pdf.

Gilg:2007:DVR

[GKS^{+}07] M. R. Gilg, S. E. Kirby, R. Sullivan, L. W. Knapp, and T. J. Hilbish. Dispersal vs. retention: correspondence of species-specific reproductive cycles and settlement periods in a blue mussel hybrid zone. Marine Ecology Progress Series, 351:151–161, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p151.pdf.

Graeve:2002:FAC

[GKWK02] M. Graeve, G. Kattner, C. Wiencke, and U. Karsten. Fatty acid composition of Arctic and Antarctic macroalgae: indicator of phylogenetic and trophic relationships. Marine Ecology Progress Series, 231:67–74, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/231/m231p067.pdf.

Glud:2002:BDH

[GKWR02] R. N. Glud, M. Kühl, F. Wenzhöfer, and S. Rysgaard. Benthic diatoms of a high Arctic fjord (Young Sound, NE Greenland): importance for ecosystem primary production. Marine Ecology Progress Series, 238:15–29, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p015.pdf.

Garrison:2000:DGS

[GL00a] L. P. Garrison and J. S. Link. Dietary guild structure of the fish community in the Northeast United States continental shelf ecosystem. Marine Ecology Progress Series, 202: 231–240, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/202/m202p231.pdf.

Garrison:2000:DFH

[GL00b] L. P. Garrison and J. S. Link. Diets of five hake species in the northeast United States continental shelf ecosystem. Marine Ecology Progress Series, 204:243–255, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p243.pdf.

Grutter:2002:CFL

[GL02] A. S. Grutter and R. J. G. Lester. Cleaner fish Labroides dimidiatus reduce ’temporary’ parasitic corallanid isopods on the coral reef fish Hemigymnus melapterus. Marine Ecology Progress Series, 234:247–255, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/234/m234p247.pdf.

Gascoigne:2004:AEM

[GL04a] J. Gascoigne and R. N. Lipcius. Allee effects in marine systems. Marine Ecology Progress Series, 269:49–59, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/269/m269p049.pdf.

Gascoigne:2004:CPL

[GL04b] J. Gascoigne and R. N. Lipcius. Conserving populations at low abundance: delayed functional maturity and Allee effects in reproductive behaviour of the queen conch Strombus gigas. Marine Ecology Progress Series, 284:185–194, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p185.pdf.

Glasby:2001:DSM

[Gla01] T. M. Glasby. Development of sessile marine assemblages on fixed versus moving substrata. Marine Ecology Progress Series, 215:37–47, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p037.pdf.

Godley:2003:MPG

[GLÅ^{+}03] B. J. Godley, E. H. S. M. Lima, S. Åkesson, A. C. Broderick, F. Glen, M. H. Godfrey, P. Luschi, and G. C. Hays. Movement patterns of green turtles in Brazilian coastal waters described by satellite tracking and flipper tagging. Marine Ecology Progress Series, 253:279–288, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/253/m253p279.pdf.

Gazo:2006:DBM

[GLA06] M. Gazo, C. Lydersen, and A. Aguilar. Diving behaviour of Mediterranean monk seal pups during lactation and post weaning. Marine Ecology Progress Series, 308:303–309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/308/m308p303.pdf.

Gladstone:2007:SSA

[Gla07] W. Gladstone. Selection of a spawning aggregation site by Chromis hypsilepis (Pisces: Pomacentridae): habitat structure, transport potential, and food availability. Marine Ecology Progress Series, 351:235–247, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p235.pdf.

Gateno:2003:STF

[GLB^{+}03] D. Gateño, A. León, Y. Barki, J. Cortés, and B. Rinkevich. Skeletal tumor formations in the massive coral Pavona clavus. Marine Ecology Progress Series, 258:97–108, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p097.pdf.

Gage:2004:LSB

[GLB^{+}04] J. D. Gage, P. J. D. Lambshead, J. D. D. Bishop, C. T. Stuart, and N. S. Jones. Large-scale biodiversity pattern of Cumacea (Peracarida: Crustacea) in the deep Atlantic. Marine Ecology Progress Series, 277:181–196, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p181.pdf.

Gutow:2007:DLH

[GLBF07] L. Gutow, S. Leidenberger, K. Boos, and H. Franke. Differential life history responses of two Idotea species (Crustacea: Isopoda) to food limitation. Marine Ecology Progress Series, 344:159–172, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/344/m344p159.pdf.

Greenfield:2004:EBC

[GLCL04] D. I. Greenfield, D. J. Lonsdale, R. M. Cerrato, and G. R. Lopez. Effects of background concentrations of Aureococcus anophagefferens (brown tide) on growth and feeding in the bivalve Mercenaria mercenaria. Marine Ecology Progress Series, 274:171–181, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p171.pdf.

Gomez:2004:PPS

[GLFU^{+}04] I. Gómez, F. López-Figueroa, N. Ulloa, V. Morales, C. Lovengreen, P. Huovinen, and S. Hess. Patterns of photosynthesis in 18 species of intertidal macroalgae from southern Chile. Marine Ecology Progress Series, 270:103–116, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/270/
m270p103.pdf.

Grutter:2000:ERB

[GLG00] A. S. Grutter, R. J. G. Lester, and J. Greenwood. Emergence rates from the benthos of the parasitic juveniles of gnathiid isopods. Marine Ecology Progress Series, 207:123–127, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p123.pdf.

Govenar:2005:ECS

[GLG^{+}05] B. Govenar, N. Le Bris, S. Gollner, J. Glanville, A. B. Aperghis, S. Hourdez, and C. R. Fisher. Epifaunal community structure associated with Riftia pachyptila aggregations in chemically different hydrothermal vent habitats. Marine Ecology Progress Series, 305:67–77, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/305/
m305p067.pdf.

Granhag:2007:ASS

[GLJ07] L. M. Granhag, A. I. Larsson, and P. R. Jonsson. Algal spore settlement and germling removal as a function of flow speed. Marine Ecology Progress Series, 344:63–69, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p063.pdf.

Goodyear:2008:VHU

[GLP^{+}08] C. P. Goodyear, J. Luo, E. D. Prince, J. P. Hoolihan, D. Snodgrass, E. S. Orbesen, and J. E. Serafy. Vertical habitat use of Atlantic blue marlin Makaira nigricans: interaction with pelagic longline gear. Marine Ecology Progress Series, 365:233–245, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p233.pdf.

Gjosaeter:2000:LTP

[GLS^{+}00] J. Gjøsæter, K. Lekve, N. C. Stenseth, H. P. Leinaas, H. Christie, E. Dahl, D. S. Danielssen, B. Edvardsen, F. Olsgard, E. Oug, and E. Paasche. A long-term perspective on the Chrysochromulina bloom on the Norwegian Skagerrak coast 1988: a catastrophe or an innocent incident? Marine Ecology Progress Series, 207:201–218, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/207/m207p201.pdf.

Gartner:2002:UDS

[GLS02] A. Gartner, P. Lavery, and A. J. Smit. Use of d^{15}N signatures of different functional forms of macroalgae and filter-feeders to reveal temporal and spatial patterns in sewage dispersal. Marine Ecology Progress Series, 235:63–73, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p063.pdf.

Gremare:2009:CPT

[GLV^{+}09] A. Grémare, C. Labrune, E. Vanden Berghe, J. M. Amouroux, G. Bachelet, M. L. Zettler, J. Vanaverbeke, D. Fleischer, L. Bigot, O. Maire, B. Deflandre, J. Craeymeersch, S. Degraer, C. Dounas, G. Duineveld, C. Heip, M. Herrmann, H. Hummel, I. Karakassis, M. Kedra, M. Kendall, P. Kingston, J. Laudien, A. Occhipinti-Ambrogi, E. Rachor, R. Sardá, J. Speybroeck, G. Van Hoey, M. Vincx, P. Whomersley, W. Willems, M. Włodarska-Kowalczuk, and A. Zenetos. Comparison of the performances of two biotic indices based on the MacroBen database. Marine Ecology Progress Series, 382:297–311, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m382p297.pdf.

Garces:2001:PPG

[GM01] E. Garcés and M. Masó. Phytoplankton potential growth rate versus increase in cell numbers: estimation of cell lysis. Marine Ecology Progress Series, 212:297–300, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p297.pdf.

Gagliano:2004:FHI

[GM04] M. Gagliano and M. I. McCormick. Feeding history influences otolith shape in tropical fish. Marine Ecology Progress Series, 278:291–296, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p291.pdf.

Green:2005:MPE

[GM05] B. S. Green and M. I. McCormick. Maternal and paternal effects determine size, growth and performance in larvae of a tropical reef fish. Marine Ecology Progress Series, 289:263–272, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p263.pdf.

Gnanadesikan:2008:EEN

[GM08] A. Gnanadesikan and I. Marinov. Export is not enough: nutrient cycling and carbon sequestration. Marine Ecology Progress Series, 364:289–294, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m364p289.pdf.

Gagnaire:2009:OCD

[GMA^{+}09] P. A. Gagnaire, Y. Minegishi, J. Aoyama, E. Réveillac, T. Robinet, P. Bosc, K. Tsukamoto, E. Feunteun, and P. Berrebi. Ocean currents drive secondary contact between Anguilla marmorata populations in the Indian Ocean. Marine Ecology Progress Series, 379:267–278, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p267.pdf.

Gilly:2006:VHM

[GMB^{+}06] W. F. Gilly, U. Markaida, C. H. Baxter, B. A. Block, A. Boustany, L. Zeidberg, K. Reisenbichler, B. Robison, G. Bazzino, and C. Salinas. Vertical and horizontal migrations by the jumbo squid Dosidicus gigas revealed by electronic tagging. Marine Ecology Progress Series, 324:1–17, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gannon:2009:EKB

[GMC^{+}09] D. P. Gannon, E. J. Berens McCabe, S. A. Camilleri, J. G. Gannon, M. K. Brueggen, A. A. Barleycorn, V. I. Palubok, G. J. Kirkpatrick, and R. S. Wells. Effects of Karenia brevis harmful algal blooms on nearshore fish communities in southwest Florida. Marine Ecology Progress Series, 378:171– 186, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/378/m378p171.pdf.

Gremare:2002:RBS

[GMd^{+}02] A. Grémare, L. Medernach, F. deBovée, J. M. Amouroux, G. Vétion, and P. Albert. Relationships between sedimentary organics and benthic meiofauna on the continental shelf and the upper slope of the Gulf of Lions (NW Mediterranean). Marine Ecology Progress Series, 234:85–94, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p085.pdf.

Gutierrez:2008:IEC

[GMD08] N. L. Gutiérrez, A. Martínez, and O. Defeo. Identifying environmental constraints at the edge of a species’ range: scallop Psychrochlamys patagonica in the SW Atlantic Ocean. Marine Ecology Progress Series, 353:147–156, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p147.pdf.

Garthe:2009:EHM

[GMHA09] S. Garthe, N. Markones, O. Hüppop, and S. Adler. Effects of hydrographic and meteorological factors on seasonal seabird abundance in the southern North Sea. Marine Ecology Progress Series, 391:243–255, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m391p243.pdf.

Giannoulaki:2003:ECT

[GMK^{+}03] M. Giannoulaki, A. Machias, C. Koutsikopoulos, J. Haralabous, S. Somarakis, and N. Tsimenides. Effect of coastal topography on the spatial structure of the populations of small pelagic fish. Marine Ecology Progress Series, 265:243– 253, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p243.pdf.

Grutter:2002:OCB

[GML02] A. S. Grutter, H. McCallum, and R. J. G. Lester. Optimising cleaning behaviour: minimising the costs and maximising ectoparasite removal. Marine Ecology Progress Series, 234:257–264, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p257.pdf.

Graham:2003:LSQ

[GMM03] W. M. Graham, D. L. Martin, and J. C. Martin. In situ quantification and analysis of large jellyfish using a novel video profiler. Marine Ecology Progress Series, 254:129– 140, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p129.pdf.

Garvey:2007:AFA

[GMP07] M. Garvey, B. Moriceau, and U. Passow. Applicability of the FDA assay to determine the viability of marine phytoplankton under different environmental conditions. Marine Ecology Progress Series, 352:17–26, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/352/m352p017.pdf.

Guisande:2002:ECT

[GMR^{+}02] C. Guisande, I. Maneiro, I. Riveiro, A. Barreiro, and Y. Pazos. Estimation of copepod trophic niche in the field using amino acids and marker pigments. Marine Ecology Progress Series, 239:147–156, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/239/m239p147.pdf.

Gowland:2002:DQD

[GMS02] F. C. Gowland, N. A. Moltschaniwskyj, and M. A. Steer. Description and quantification of developmental abnormalities in a natural Sepioteuthis australis spawning population (Mollusca: Cephalopoda). Marine Ecology Progress Series, 243:133–141, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/243/m243p133.pdf.

Gowen:2000:PFT

[GMTN00] R. J. Gowen, D. K. Mills, M. Trimmer, and D. B. Nedwell. Production and its fate in two coastal regions of the Irish Sea: the influence of anthropogenic nutrients. Marine Ecology Progress Series, 208:51–64, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/208/m208p051.pdf.

Gallucci:2008:ACD

[GMVF08] F. Gallucci, T. Moens, A. Vanreusel, and G. Fonseca. Active colonisation of disturbed sediments by deep-sea nematodes: evidence for the patch mosaic model. Marine Ecology Progress Series, 367:173–183, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/367/m367p173.pdf.

Garrido:2007:LIE

[GMZvdL07] S. Garrido, A. Marçalo, J. Zwolinski, and C. D. van der Lingen. Laboratory investigations on the effect of prey size and concentration on the feeding behaviour of Sardina pilchardus. Marine Ecology Progress Series, 330:189–199, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/330/
m330p189.pdf.

Gallucci:2004:EPC

[GN04] F. Gallucci and S. A. Netto. Effects of the passage of cold fronts over a coastal site: an ecosystem approach. Marine Ecology Progress Series, 281:79–92, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p079.pdf.

Grahl-Nielsen:2003:FAC

[GNAD^{+}03] O. Grahl-Nielsen, M. Andersen, A. E. Derocher, C. Lydersen, Ó. Wiig, and K. M. Kovacs. Fatty acid composition of the adipose tissue of polar bears and of their prey: ringed seals, bearded seals and harp seals. Marine Ecology Progress Series, 265:275–282, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p275.pdf. See [TBBI04]

and reply [GNAD^{+}04].

Grahl-Nielsen:2004:RCG

[GNAD^{+}04] O. Grahl-Nielsen, M. Andersen, A. E. Derocher, C. Lydersen, Ó. Wiig, and K. M. Kovacs. Reply to Comment on Grahl-Nielsen et al. (2003): sampling, data treatment and predictions in investigations on fatty acids in marine mammals. Marine Ecology Progress Series, 281:303–306, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/

m281p303.pdf. See [GNAD^{+}03] and comment [TBBI04].

Guzner:2007:PDR

[GNC07] B. Guzner, A. Novoplansky, and N. E. Chadwick. Population dynamics of the reef-building coral Acropora hemprichii as an indicator of reef condition. Marine Ecology Progress Series, 333:143–150, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p143.pdf.

Grahl-Nielsen:2005:FAC

[GNHB^{+}05] O. Grahl-Nielsen, A. K. Halvorsen, N. Bodoev, L. Averina, L. Radnaeva, N. Pronin, R. Käkelä, and E. Petrov. Fatty acid composition of blubber of the Baikal seal Phoca sibirica and its marine relative, the ringed seal P. hispida. Marine Ecology Progress Series, 305:261–274, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p261.pdf.

Galuardi:2008:OSS

[GNL08] B. Galuardi, A. Nielsen, and M. Lutcavage. Optimizing smoothed sea surface temperature for improving archival tag geolocation. Marine Ecology Progress Series, 365:35– 44, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/365/m365p035.pdf.

Gallego:2007:ISF

[GNP07] A. Gallego, E. W. North, and P. Petitgas. Introduction: status and future of modelling physical-biological interactions during the early life of fishes. Marine Ecology Progress Series, 347:121–126, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m347p121.pdf.

Gallego:2007:AMP

[GNPB07] Alejandro Gallego, Elizabeth W. North, Pierre Petitgas, and Howard I. Browman. Advances in modelling physical– biological interactions in fish early life history. Marine Ecology Progress Series, 347:121–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m347_TS.pdf.

Gochfeld:2004:PIM

[Goc04] D. J. Gochfeld. Predation-induced morphological and behavioral defenses in a hard coral: implications for foraging behavior of coral-feeding butterflyfishes. Marine Ecology Progress Series, 267:145–158, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/267/m267p145.pdf.

Goodsell:2009:DFA

[Goo09] P. J. Goodsell. Diversity in fragments of artificial and natural marine habitats. Marine Ecology Progress Series, 384: 23–31, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p023.pdf.

Gorokhova:2003:RBN

[Gor03] E. Gorokhova. Relationships between nucleic acid levels and egg production rates in Acartia bifilosa: implications for growth assessment of copepods in situ. Marine Ecology Progress Series, 262:163–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p163.pdf.

Goulet:2006:MCM

[Gou06] T. L. Goulet. Most corals may not change their symbionts. Marine Ecology Progress Series, 321:1–7, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See comment [BR07].

Goulet:2007:MSC

[Gou07] T. L. Goulet. Most scleractinian corals and octocorals host a single symbiotic zooxanthella clade. Marine Ecology Progress Series, 335:243–248, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m335p243.pdf.

Ginsburg:2001:CDS

[GP01] D. W. Ginsburg and V. J. Paul. Chemical defenses in the sea hare Aplysia parvula: importance of diet and sequestration of algal secondary metabolites. Marine Ecology Progress Series, 215:261–274, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p261.pdf.

Gutt:2003:SDI

[GP03] J. Gutt and D. Piepenburg. Scale-dependent impact on diversity of Antarctic benthos caused by grounding of icebergs. Marine Ecology Progress Series, 253:77–83, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p077.pdf.

Giles:2004:EDS

[GP04a] H. Giles and C. A. Pilditch. Effects of diet on sinking rates and erosion thresholds of mussel Perna canaliculus biodeposits. Marine Ecology Progress Series, 282:205–219, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p205.pdf.

Grabowski:2004:HCM

[GP04b] J. H. Grabowski and S. P. Powers. Habitat complexity mitigates trophic transfer on oyster reefs. Marine Ecology Progress Series, 277:291–295, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/277/m277p291.pdf.

Greenstreet:2008:ASE

[GP08] S. P. R. Greenstreet and G. J. Piet. Assessing the sampling effort required to estimate a species diversity in the groundfish assemblages of the North Sea. Marine Ecology Progress Series, 364:181–197, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p181.pdf.

Gooday:2002:ENL

[GPB02] A. J. Gooday, D. W. Pond, and S. S. Bowser. Ecology and nutrition of the large agglutinated Foraminiferan Bathysiphon capillare in the bathyal NE Atlantic: distribution within the sediment profile and lipid biomarker composition. Marine Ecology Progress Series, 245:69–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p069.pdf.

Geraldi:2009:CHR

[GPHC09] N. R. Geraldi, S. P. Powers, K. L. Heck, and J. Cebrian. Can habitat restoration be redundant? Response of mobile fishes and crustaceans to oyster reef restoration in marsh tidal creeks. Marine Ecology Progress Series, 389:171–180, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m389p171.pdf.

Geertjes:2004:ARD

[GPK^{+}04] G. J. Geertjes, J. Postema, A. Kamping, W. van Delden, J. J. Videler, and L. van de Zande. Allozymes and RAPDs detect little genetic population substructuring in the Caribbean stoplight parrotfish Sparisoma viride. Marine Ecology Progress Series, 279:225–235, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p225.pdf.

Garza-Perez:2004:SPC

[GPLAG04] J. R. Garza-Pérez, A. Lehmann, and J. E. Arias-González. Spatial prediction of coral reef habitats: integrating ecology with spatial modeling and remote sensing. Marine Ecology Progress Series, 269:141–152, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p141.pdf.

Gutowska:2008:GCC

[GPM08] M. A. Gutowska, H. O. Pörtner, and F. Melzner. Growth and calcification in the cephalopod Sepia officinalis under elevated seawater pCO_{2}. Marine Ecology Progress Series, 373:303–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m373p303.pdf.

Guadayol:2009:EMN

[GPM^{+}09] Òscar Guadayol, Francesc Peters, Cèlia Marrasé, Josep M. Gasol, Cristina Roldán, Elisas Berdalet, Ramon Massana, and Anna Sabata. Episodic meteorological and nutrientload events as drivers of coastal planktonic ecosystem dynamics: a time-series analysis. Marine Ecology Progress Series, 381:139–155, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p139.pdf.

Goldbogen:2007:BGR

[GPS07] J. A. Goldbogen, N. D. Pyenson, and R. E. Shadwick. Big gulps require high drag for fin whale lunge feeding. Marine Ecology Progress Series, 349:289–301, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p289.pdf.

Gremillet:2006:CIA

[GPSG06] D. Grémillet, L. Pichegru, F. Siorat, and J. Y. Georges. Conservation implications of the apparent mismatch between population dynamics and foraging effort in French northern gannets from the English Channel. Marine Ecology Progress Series, 319:15–25, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p015.pdf.

Garrabou:2001:MME

[GPSH01] J. Garrabou, T. Perez, S. Sartoretto, and J. G. Harmelin. Mass mortality event in red coral Corallium rubrum populations in the Provence region (France, NW Mediterranean). Marine Ecology Progress Series, 217:263–272, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p263.pdf.

Gilpin:2002:PPC

[GPW^{+}02] L. C. Gilpin, J. Priddle, M. J. Whitehouse, G. Savidge, and A. Atkinson. Primary production and carbon uptake dynamics in the vicinity of south georgia—balancing carbon fixation and removal. Marine Ecology Progress Series, 242: 51–62, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p051.pdf.

Gosselin:2000:CBS

[GQ00] L. A. Gosselin and P. Y. Qian. Can bacterivory sustain survival and growth in early juveniles of the bryozoan Bugula neritina, the polychaete Hydroides elegans and the barnacle Balanus amphitrite? Marine Ecology Progress Series, 192:163–172, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p163.pdf.

Goni:2006:SSL

[GQR06] R. Goñi, A. Quetglas, and O. Reñones. Spillover of spiny lobsters Palinurus elephas from a marine reserve to an adjoining fishery. Marine Ecology Progress Series, 308:207– 219, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/308/m308p207.pdf.

Gosselin:2007:IJS

[GR07a] L. A. Gosselin and R. Rehak. Initial juvenile size and environmental severity: influence of predation and wave exposure on hatching size in Nucella ostrina. Marine Ecology Progress Series, 339:143–155, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p143.pdf.

Granek:2007:PCM

[GR07b] E. F. Granek and B. I. Ruttenberg. Protective capacity of mangroves during tropical storms: a case study from ‘Wilma’ and ‘Gamma’ in Belize. Marine Ecology Progress Series, 343:101–105, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p101.pdf.

Gray:2002:SRM

[Gra02] J. S. Gray. Species richness of marine soft sediments. Marine Ecology Progress Series, 244:285–297, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p285.pdf.

Grift:2003:FIT

[GRB^{+}03] R. E. Grift, A. D. Rijnsdorp, S. Barot, M. Heino, and U. Dieckmann. Fisheries-induced trends in reaction norms for maturation in North Sea plaice. Marine Ecology Progress Series, 257:247–257, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p247.pdf.

Gifford:2007:MOU

[GRBB07] S. M. Gifford, G. Rollwagen-Bollens, and S. M. Bollens. Mesozooplankton omnivory in the upper San Francisco Estuary. Marine Ecology Progress Series, 348:33–46, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p033.pdf.

Garcia-Robledo:2008:BEM

[GRCdLvB08] E. García-Robledo, A. Corzo, J. G. de Lomas, and S. A. van Bergeijk. Biogeochemical effects of macroalgal decomposition on intertidal microbenthos: a microcosm experiment. Marine Ecology Progress Series, 356:139–151, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/356/
m356p139.pdf.

Gowing:2002:LVR

[GRG^{+}02] M. M. Gowing, B. E. Riggs, D. L. Garrison, A. H. Gibson, and M. O. Jeffries. Large viruses in Ross Sea late autumn pack ice habitats. Marine Ecology Progress Series, 241:1–11, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/241/m241p001.pdf. See errata [Ano03b].

Griffiths:2002:RVI

[Gri02] S. P. Griffiths. Retention of visible implant tags in small rockpool fishes. Marine Ecology Progress Series, 236:307– 309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/236/m236p307.pdf.

Griffen:2009:ENI

[Gri09] B. D. Griffen. Effects of a newly invasive parasite on the burrowing mud shrimp, a widespread ecosystem engineer. Marine Ecology Progress Series, 391:73–83, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/391/
m391p073.pdf.

Gall:2006:DRC

[GRIR06] A. E. Gall, D. D. Roby, D. B. Irons, and I. C. Rose. Differential response in chick survival to diet in least and crested auklets. Marine Ecology Progress Series, 308:279–291, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p279.pdf.

Gutierrez-Rodriguez:2009:CBP

[GRLML09] A. Gutiérrez-Rodríguez, M. Latasa, B. Mourre, and E. A. Laws. Coupling between phytoplankton growth and microzooplankton grazing in dilution experiments: potential artefacts. Marine Ecology Progress Series, 383:1–9, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Guisande:2000:CAA

[GRM00] C. Guisande, I. Riveiro, and I. Maneiro. Comparisons among the amino acid composition of females, eggs and food to determine the relative importance of food quantity and food quality to copepod reproduction. Marine Ecology Progress Series, 202:135–142, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p135.pdf.

Gili:2006:NTL

[GRP^{+}06] J. M. Gili, S. Rossi, F. Pagès, C. Orejas, N. Teixidó, P. J. López-González, and W. E. Arntz. A new trophic link between the pelagic and benthic systems on the Antarctic shelf. Marine Ecology Progress Series, 322:43–49, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p043.pdf.

Garcia-Rodriguez:2008:LGD

[GRPE08] F. J. García-Rodríguez and R. Perez-Enriquez. Lack of genetic differentiation of blue spiny lobster Panulirus inflatus along the Pacific coast of Mexico inferred from mtDNA sequences. Marine Ecology Progress Series, 361:203–212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p203.pdf.

Glud:2000:BCM

[GRPT^{+}00] R. N. Glud, N. Risgaard-Petersen, B. Thamdrup, H. Fossing, and S. Rysgaard. Benthic carbon mineralization in a high-Arctic sound (Young Sound, NE Greenland). Marine Ecology Progress Series, 206:59–71, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p059.pdf.

Goetz:2007:HUM

[GRRH07] K. T. Goetz, D. J. Rugh, A. J. Read, and R. C. Hobbs. Habitat use in a marine ecosystem: beluga whales Delphinapterus leucas in Cook Inlet, Alaska. Marine Ecology Progress Series, 330:247–256, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p247.pdf.

Garcia:2007:SNF

[GRS07] N. Garcia, P. Raimbault, and V. Sandroni. Seasonal nitrogen fixation and primary production in the southwest Pacific: nanoplankton diazotrophy and transfer of nitrogen to picoplankton organisms. Marine Ecology Progress Series, 343: 25–33, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p025.pdf.

Grutter:2000:OVD

[Gru00] A. S. Grutter. Ontogenetic variation in the diet of the cleaner fish Labroides dimidiatus and its ecological consequences. Marine Ecology Progress Series, 197:241–246, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p241.pdf.

Grutter:2003:FEF

[Gru03] A. Grutter. Feeding ecology of the fish ectoparasite Gnathia sp. (Crustacea: Isopoda) from the Great Barrier Reef, and its implications for fish cleaning behaviour. Marine Ecology Progress Series, 259:295–302, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/259/m259p295.pdf.

Gibson:2002:OCD

[GRWB02] R. N. Gibson, L. Robb, H. Wennhage, and M. T. Burrows. Ontogenetic changes in depth distribution of juvenile flatfishes in relation to predation risk and temperature on a shallow-water nursery ground. Marine Ecology Progress Series, 229:233–244, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p233.pdf.

Grossart:2002:BDG

[GS02] H. P. Grossart and M. Simon. Bacterioplankton dynamics in the Gulf of Aqaba and the northern Red Sea in early spring. Marine Ecology Progress Series, 239:263–276, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p263.pdf.

Gratwicke:2005:EHC

[GS05] B. Gratwicke and M. R. Speight. Effects of habitat complexity on Caribbean marine fish assemblages. Marine Ecology Progress Series, 292:301–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/292/m292p301.pdf.

Guidetti:2007:CWE

[GS07] P. Guidetti and E. Sala. Community-wide effects of marine reserves in the Mediterranean Sea. Marine Ecology Progress Series, 335:43–56, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/335/m335p043.pdf.

Godbold:2009:RIB

[GS09] J. A. Godbold and M. Solan. Relative importance of biodiversity and the abiotic environment in mediating an ecosystem process. Marine Ecology Progress Series, 396:273–282, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p273.pdf.

Girard:2006:HGT

[GSB^{+}06] C. Girard, J. Sudre, S. Benhamou, D. Roos, and P. Luschi. Homing in green turtles Chelonia mydas: oceanic currents act as a constraint rather than as an information source. Marine Ecology Progress Series, 322:281–289, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p281.pdf.

Granata:2001:FPD

[GSC^{+}01] T. C. Granata, T. Serra, J. Colomer, X. Casamitjana, C. M. Duarte, and E. Gacia. Flow and particle distributions in a nearshore seagrass meadow before and after a storm. Marine Ecology Progress Series, 218:95–106, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

218/m218p095.pdf. See erratum [Ano01d].

Gonzalez-Solis:2000:FPB

[GSCW00] J. González-Solís, J. P. Croxall, and A. G. Wood. Foraging partitioning between giant petrels Macronectes spp. and its relationship with breeding population changes at Bird Island, South Georgia. Marine Ecology Progress Series, 204: 279–288, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p279.pdf.

Gonzalez-Solis:2009:ISS

[GSFF^{+}09] J. González-Solís, A. Felicísimo, J. W. Fox, V. Afanasyev, Y. Kolbeinsson, and J. Muñoz. Influence of sea surface winds on shearwater migration detours. Marine Ecology Progress Series, 391:221–230, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m391p221.pdf.

Gonzalez:2000:CBP

[GSFN00] H. E. González, M. Sobarzo, D. Figueroa, and E. M. Nöthig. Composition, biomass and potential grazing impact of the crustacean and pelagic tunicates in the northern Humboldt Current area off Chile: differences between El Niño and non-El Niño years. Marine Ecology Progress Series, 195:201– 220, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/195/m195p201.pdf.

Goulet:2008:WBS

[GSG08] T. L. Goulet, C. Simmons, and D. Goulet. Worldwide biogeography of Symbiodinium in tropical octocorals. Marine Ecology Progress Series, 355:45–58, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/355/m355p045.pdf.

Geller:2008:SIN

[GSK^{+}08] J. Geller, E. E. Sotka, R. Kado, S. R. Palumbi, and E. Schwindt. Sources of invasions of a northeastern Pacific acorn barnacle, Balanus glandula, in Japan and Argentina. Marine Ecology Progress Series, 358:211–218, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p211.pdf.

Gao:2006:SIF

[GSL^{+}06] Q. F. Gao, P. K. S. Shin, G. H. Lin, S. P. Chen, and S. G. Cheung. Stable isotope and fatty acid evidence for uptake of organic waste by green-lipped mussels Perna viridis in a polyculture fish farm system. Marine Ecology Progress Series, 317:273–283, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p273.pdf.

Golet:2002:LTD

[GSM^{+}02] G. H. Golet, P. E. Seiser, A. D. McGuire, D. D. Roby, J. B. Fischer, K. J. Kuletz, D. B. Irons, T. A. Dean, S. C. Jewett, and S. H. Newman. Long-term direct and indirect effects of the ‘Exxon Valdez’ oil spill on pigeon guillemots in Prince William Sound, Alaska. Marine Ecology Progress Series, 241:287–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p287.pdf.

Gallucci:2005:CFD

[GSM05] F. Gallucci, M. Steyaert, and T. Moens. Can field distributions of marine predacious nematodes be explained by sediment constraints on their foraging success? Marine Ecology Progress Series, 304:167–178, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/304/m304p167.pdf.

Gendron:2006:GJL

[GSM06] L. Gendron and B. Sainte-Marie. Growth of juvenile lobster Homarus americanus off the Magdalen Islands (Quebec, Canada) and projection of instar and age at commercial size. Marine Ecology Progress Series, 326:221–233, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/326/
m326p221.pdf.

Gagern:2008:BRI

[GSM^{+}08] A. Gagern, T. Schürg, N. K. Michiels, G. Schulte, D. Sprenger, and N. Anthes. Behavioural response to interference competition in a sessile suspension feeder. Marine Ecology Progress Series, 353:131–135, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/353/m353p131.pdf.

Galeron:2000:VSB

[GSMD00] J. Galéron, M. Sibuet, M. L. Mahaut, and A. Dinet. Variation in structure and biomass of the benthic communities at three contrasting sites in the tropical Northeast Atlantic.

Marine Ecology Progress Series, 197:121–137, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p121.pdf.

Glover:2002:PSD

[GSP^{+}02] A. G. Glover, C. R. Smith, G. L. J. Paterson, G. D. F. Wilson, L. Hawkins, and M. Sheader. Polychaete species diversity in the central Pacific abyss: local and regional patterns, and relationships with productivity. Marine Ecology Progress Series, 240:157–170, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p157.pdf.

Godhe:2002:OSE

[GSRH02] A. Godhe, S. Svensson, and A. S. Rehnstam-Holm. Oceanographic settings explain fluctuations in Dinophysis spp. and concentrations of diarrhetic shellfish toxin in the plankton community within a mussel farm area on the Swedish west coast. Marine Ecology Progress Series, 240:71–83, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/240/m240p071.pdf.

Gilles:2009:SDH

[GSS09a] A. Gilles, M. Scheidat, and U. Siebert. Seasonal distribution of harbour porpoises and possible interference of offshore wind farms in the German North Sea. Marine Ecology Progress Series, 383:295–307, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/383/m383p295.pdf.

Gonzales-Solis:2009:ISS

[GSS09b] J. Gonzáles-Solís and S. A. Shaffer. Introduction and synthesis: spatial ecology of seabirds at sea. Marine Ecology Progress Series, 391:117–120, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m391p117.pdf.

Gonzales-Solis:2009:TSS

[GSS09c] Jacob Gonzáles-Solís and Scott A. Shaffer. Theme section: Spatiotemporal dynamics of seabirds in the marine environment. Marine Ecology Progress Series, 391:116–306, 2009.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m391_ThemeSection.pdf.

Gonzalez-Solis:2002:MSB

[GSSR02] J. González-Solís, C. Sanpera, and X. Ruiz. Metals and selenium as bioindicators of geographic and trophic segregation in giant petrels Macronectes spp. Marine Ecology Progress Series, 244:257–264, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p257.pdf.

Gobler:2001:EOC

[GSW01a] C. J. Gobler and S. A. Sañudo-Wilhelmy. Effects of organic carbon, organic nitrogen, inorganic nutrients, and iron additions on the growth of phytoplankton and bacteria during a brown tide bloom. Marine Ecology Progress Series, 209: 19–34, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/209/m209p019.pdf.

Gobler:2001:TVG

[GSW01b] C. J. Gobler and S. A. Sañudo-Wilhelmy. Temporal variability of groundwater seepage and brown tide blooms in a Long Island embayment. Marine Ecology Progress Series, 217:299–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/217/m217p299.pdf.

Graneli:2002:TRC

[GT02] E. Granéli and J. T. Turner. Top-down regulation in ctenophore-copepod-ciliate-diatom-phytoflagellate communities in coastal waters: a mesocosm study. Marine Ecology Progress Series, 239:57–68, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/239/m239p057.pdf.

Gregr:2008:NPO

[GT08] E. J. Gregr and A. W. Trites. A novel presence-only validation technique for improved Steller sea lion Eumetopias jubatus critical habitat descriptions. Marine Ecology Progress Series, 365:247–261, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p247.pdf.

Gaertner:2008:VCA

[GTD^{+}08] J. C. Gaertner, M. Taquet, L. Dagorn, B. Mérigot, R. Aumeeruddy, G. Sancho, and D. Itano. Visual censuses around drifting fish aggregating devices (FADs): a new approach for assessing the diversity of fish in open-ocean waters. Marine Ecology Progress Series, 366:175–186, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/366/
m366p175.pdf.

Ganesan:2009:CGA

[GTE^{+}09] M. Ganesan, S. Thiruppathi, K. Eswaran, C. R. K. Reddy, and B. Jha. Cultivation of Gelidiella acerosa in the open sea on the southeastern coast of India. Marine Ecology Progress Series, 382:49–57, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/382/m382p049.pdf.

Guidetti:2003:CMR

[GTFB03] P. Guidetti, A. Terlizzi, S. Fraschetti, and F. Boero. Changes in Mediterranean rocky-reef fish assemblages exposed to sewage pollution. Marine Ecology Progress Series, 253:269– 278, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p269.pdf.

Gopal:2006:MDV

[GTGM06] K. Gopal, K. A. Tolley, J. C. Groeneveld, and C. A. Matthee. Mitochondrial DNA variation in spiny lobster Palinurus delagoae suggests genetically structured populations in the southwestern Indian Ocean. Marine Ecology Progress Series, 319:191–198, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p191.pdf.

Grange:2004:LTI

[GTPC04] L. J. Grange, P. A. Tyler, L. S. Peck, and N. Cornelius. Longterm interannual cycles of the gametogenic ecology of the Antarctic brittle star Ophionotus victoriae. Marine Ecology Progress Series, 278:141–155, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p141.pdf.

Guo:2003:NCI

[GTSS03] L. Guo, N. Tanaka, D. M. Schell, and P. H. Santschi. Nitrogen and carbon isotopic composition of high-molecular-weight dissolved organic matter in marine environments. Marine Ecology Progress Series, 252:51–60, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p051.pdf.

Grob:2007:PAB

[GUL^{+}07] C. Grob, O. Ulloa, W. K. W. Li, G. Alarcón, M. Fukasawa, and S. Watanabe. Picoplankton abundance and biomass across the eastern South Pacific Ocean along latitude 32.5^{o}S. Marine Ecology Progress Series, 332:53–62, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/332/
m332p053.pdf.

Guisande:2004:SAG

[GUT04] Cástor Guisande, Ana Ulla, and Peter Thejll. Solar activity governs abundance of Atlantic Iberian sardine Sardina pilchardus. Marine Ecology Progress Series, 269:297– 301, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/269/m269p297.pdf.

Griffin:2001:DIS

[GV01] M. P. A. Griffin and I. Valiela. d^{15}N isotope studies of life history and trophic position of Fundulus heteroclitus and Menidia menidia. Marine Ecology Progress Series, 214:299–305, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p299.pdf.

Garibotti:2003:PSD

[GVF^{+}03] I. A. Garibotti, M. Vernet, M. E. Ferrario, R. C. Smith, R. M. Ross, and L. B. Quetin. Phytoplankton spatial distribution patterns along the western Antarctic Peninsula (Southern Ocean). Marine Ecology Progress Series, 261: 21–39, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/261/m261p021.pdf.

Garibotti:2003:CBP

[GVKF03] I. A. Garibotti, M. Vernet, W. A. Kozlowski, and M. E. Ferrario. Composition and biomass of phytoplankton assemblages in coastal Antarctic waters: a comparison of chemotaxonomic and microscopic analyses. Marine Ecology Progress Series, 247:27–42, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p027.pdf.

Garces:2005:TSA

[GVM^{+}05] E. Garcés, M. Vila, M. Masó, N. Sampedro, M. G. Giacobbe, and A. Penna. Taxon-specific analysis of growth and mortality rates of harmful dinoflagellates during bloom conditions. Marine Ecology Progress Series, 301:67–79, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/301/
m301p067.pdf.

Gribben:2006:ISE

[GW06] P. E. Gribben and J. T. Wright. Invasive seaweed enhances recruitment of a native bivalve: roles of refuge from predation and the habitat choice of recruits. Marine Ecology Progress Series, 318:177–185, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/318/m318p177.pdf.

Guenther:2009:CAD

[GWBdN09] J. Guenther, A. D. Wright, K. Burns, and R. de Nys. Chemical antifouling defences of sea stars: effects of the natural products hexadecanoic acid, cholesterol, lathosterol and sitosterol. Marine Ecology Progress Series, 385:137–149, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p137.pdf.

Grigg:2007:NDM

[GWF07] N. J. Grigg, I. T. Webster, and P. W. Ford. Non-destructive measurement of the time evolution of burrowing shrimp mound topography. Marine Ecology Progress Series, 329: 157–168, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p157.pdf.

Gagnon:2003:UWT

[GWH03] P. Gagnon, G. Wagner, and J. H. Himmelman. Use of a wave tank to study the effects of water motion and algal movement on the displacement of the sea star Asterias vulgaris towards its prey. Marine Ecology Progress Series, 258:125–132, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p125.pdf.

Gray:2002:EHO

[GWO02] J. S. Gray, R. S. Wu, and Y. Y. Or. Effects of hypoxia and organic enrichment on the coastal marine environment. Marine Ecology Progress Series, 238:249–279, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p249.pdf.

Gonzalez-Wanguemert:2007:TGV

[GWPRC^{+}07] M. González-Wangüemert, Á. Pérez-Ruzafa, F. Cánovas, J. A. García-Charton, and C. Marcos. Temporal genetic variation in populations of Diplodus sargus from the SW Mediterranean Sea. Marine Ecology Progress Series, 334:237–244, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/334/m334p237.pdf.

Gremillet:2000:BBA

[GWWC00] D. Grémillet, R. P. Wilson, S. Wanless, and T. Chater. Blackbrowed albatrosses, international fisheries and the Patagonian Shelf. Marine Ecology Progress Series, 195:269–280, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p269.pdf.

Gao:2008:SCC

[GXL^{+}08] Q. F. Gao, W. Z. Xu, X. S. Liu, S. G. Cheung, and P. K. S. Shin. Seasonal changes in C, N and P budgets of greenlipped mussels Perna viridis and removal of nutrients from fish farming in Hong Kong. Marine Ecology Progress Series, 353:137–146, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/353/m353p137.pdf.

Gimenez:2000:LPD

[GY00] L. Giménez and B. Yannicelli. Longshore patterns of distribution of macroinfauna on a Uruguayan sandy beach: an analysis at different spatial scales and of their potential causes. Marine Ecology Progress Series, 199:111–125, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p111.pdf.

Gallager:2004:CFS

[GYD04] S. M. Gallager, H. Yamazaki, and C. S. Davis. Contribution of fine-scale vertical structure and swimming behavior to formation of plankton layers on Georges Bank. Marine Ecology Progress Series, 267:27–43, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/267/m267p027.pdf. See

erratum [Ano06].

Garpe:2006:LTE

[GYLÖ06] K. C. Garpe, S. A. S. Yahya, U. Lindahl, and M. C. Öhman. Long-term effects of the 1998 coral bleaching event on reef fish assemblages. Marine Ecology Progress Series, 315:237– 247, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/315/m315p237.pdf.

Garland:2002:TIB

[GZ02] E. D. Garland and C. A. Zimmer. Techniques for the identification of bivalve larvae. Marine Ecology Progress Series, 225:299–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p299.pdf.

Gollner:2007:TAP

[GZG^{+}07] S. Gollner, J. Zekely, B. Govenar, N. Le Bris, H. L. Nemeschkal, C. R. Fisher, and M. Bright. Tubewormassociated permanent meiobenthic communities from two chemically different hydrothermal vent sites on the East Pacific Rise. Marine Ecology Progress Series, 337:39–49, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/337/m337p039.pdf.

Gonzalez-Zevallos:2006:SUD

[GZY06] D. González-Zevallos and P. Yorio. Seabird use of discards and incidental captures at the Argentine hake trawl fishery in the Golfo San Jorge, Argentina. Marine Ecology Progress Series, 316:175–183, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/316/m316p175.pdf.

Hughes:2000:FTE

[HAA00] D. J. Hughes, R. J. A. Atkinson, and A. D. Ansell. A field test of the effects of megafaunal burrows on benthic chamber measurements of sediment-water solute fluxes. Marine Ecology Progress Series, 195:189–199, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p189.pdf.

Hauser:2006:EHC

[HAC06a] A. Hauser, M. J. Attrill, and P. A. Cotton. Effects of habitat complexity on the diversity and abundance of macrofauna colonising artificial kelp holdfasts. Marine Ecology Progress Series, 325:93–100, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/325/m325p093.pdf.

Hoogenboom:2006:ECP

[HAC06b] M. O. Hoogenboom, K. R. N. Anthony, and S. R. Connolly. Energetic cost of photoinhibition in corals. Marine Ecology Progress Series, 313:1–12, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hubas:2007:RBC

[HAD07] C. Hubas, L. F. Artigas, and D. Davoult. Role of the bacterial community in the annual benthic metabolism of two contrasted temperate intertidal sites (Roscoff Aber Bay, France). Marine Ecology Progress Series, 344:39–48, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p039.pdf.

Halpin:2000:HUI

[Hal00] P. M. Halpin. Habitat use by an intertidal salt-marsh fish: trade-offs between predation and growth. Marine Ecology Progress Series, 198:203–214, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/198/m198p203.pdf.

Halpern:2004:MLR

[Hal04a] B. S. Halpern. Are mangroves a limiting resource for two coral reef fishes? Marine Ecology Progress Series, 272: 93–98, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/272/m272p093.pdf.

Halpern:2004:HBS

[Hal04b] B. S. Halpern. Habitat bottlenecks in stage-structured species: hermit crabs as a model system. Marine Ecology Progress Series, 276:197–207, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/276/m276p197.pdf.

Humston:2005:MSF

[HALL05] R. Humston, J. S. Ault, M. F. Larkin, and J. Luo. Movements and site fidelity of the bonefish Albula vulpes in northern Florida Keys determined by acoustic telemetry. Marine Ecology Progress Series, 291:237–248, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p237.pdf.

Hansson:2009:CCB

[HAMW09] L. Hansson, M. Agis, C. Maier, and M. G. Weinbauer. Community composition of bacteria associated with cold-water coral Madrepora oculata: within and between colony variability. Marine Ecology Progress Series, 397:89–102, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p089.pdf.

Hannah:2007:FDM

[Han07] C. G. Hannah. Future directions in modelling physicalbiological interactions. Marine Ecology Progress Series, 347:

301–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m347p301.pdf.

Harris:2005:EWA

[HANM^{+}05] M. P. Harris, T. Anker-Nilssen, R. H. McCleery, K. E. Erikstad, D. N. Shaw, and V. Grosbois. Effect of wintering area and climate on the survival of adult Atlantic puffins Fratercula arctica in the eastern Atlantic. Marine Ecology Progress Series, 297:283–296, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/297/m297p283.pdf.

Hegde:2008:IES

[HAP^{+}08] S. Hegde, A. C. Anil, J. S. Patil, S. Mitbavkar, V. Krishnamurthy, and V. V. Gopalakrishna. Influence of environmental settings on the prevalence of Trichodesmium spp. in the Bay of Bengal. Marine Ecology Progress Series, 356: 93–101, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/356/m356p093.pdf.

Hartman:2000:ISS

[Har00] K. J. Hartman. The influence of size on striped bass foraging. Marine Ecology Progress Series, 194:263–268, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p263.pdf.

Harley:2006:EPE

[Har06a] C. D. G. Harley. Effects of physical ecosystem engineering and herbivory on intertidal community structure. Marine Ecology Progress Series, 317:29–39, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/317/m317p029.pdf.

Hart:2006:ESS

[Har06b] D. R. Hart. Effects of sea stars and crabs on sea scallop Placopecten magellanicus recruitment in the Mid-Atlantic Bight (USA). Marine Ecology Progress Series, 306:209– 221, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/306/m306p209.pdf.

Harley:2008:TDT

[Har08] C. D. G. Harley. Tidal dynamics, topographic orientation, and temperature-mediated mass mortalities on rocky shores. Marine Ecology Progress Series, 371:37–46, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/371/
m371p037.pdf.

Hall-Aspland:2005:SCN

[HARC05] S. A. Hall-Aspland, T. L. Rogers, and R. B. Canfield. Stable carbon and nitrogen isotope analysis reveals seasonal variation in the diet of leopard seals. Marine Ecology Progress Series, 305:249–259, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p249.pdf.

Haywood:2003:FRN

[HB03] G. J. Haywood and C. W. Burns. Feeding response of Nyctiphanes australis (Euphausiacea) to various nanoplankton sizes and taxa. Marine Ecology Progress Series, 253:209– 216, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p209.pdf.

Hooff:2004:FRP

[HB04] R. C. Hooff and S. M. Bollens. Functional response and potential predatory impact of Tortanus dextrilobatus, a carnivorous copepod recently introduced to the San Francisco Estuary. Marine Ecology Progress Series, 277:167–179, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p167.pdf.

Helaouet:2007:MCF

[HB07a] P. Helaouët and G. Beaugrand. Macroecology of Calanus finmarchicus and C. helgolandicus in the North Atlantic Ocean and adjacent seas. Marine Ecology Progress Series, 345:147– 165, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p147.pdf.

Hosia:2007:SCG

[HB07b] A. Hosia and U. Båmstedt. Seasonal changes in the gelatinous zooplankton community and hydromedusa abundances in korsfjord and fanafjord, western Norway. Marine Ecology Progress Series, 351:113–127, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/351/m351p113.pdf.

Heath:2008:NPP

[HB08a] M. R. Heath and D. J. Beare. New primary production in northwest European shelf seas, 1960–2003. Marine Ecology Progress Series, 363:183–203, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/363/m363p183.pdf.

Huntington:2008:ERM

[HB08b] B. E. Huntington and K. E. Boyer. Effects of red macroalgal (Gracilariopsis sp.) abundance on eelgrass Zostera marina in Tomales Bay, California, USA. Marine Ecology Progress Series, 367:133–142, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/367/m367p133.pdf.

Herreman:2009:EBC

[HBBD09] J. K. Herreman, G. M. Blundell, and M. Ben-David. Evidence of bottom-up control of diet driven by top-down processes in a declining harbor seal Phoca vitulina richardsi population. Marine Ecology Progress Series, 374:287–300, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p287.pdf.

Hochscheid:2007:OBS

[HBBH07] S. Hochscheid, F. Bentivegna, M. N. Bradai, and G. C. Hays. Overwintering behaviour in sea turtles: dormancy is optional. Marine Ecology Progress Series, 340:287–298, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p287.pdf.

Hannes:2009:SSD

[HBC09] A. R. Hannes, M. Barbeitos, and M. A. Coffroth. Stability of symbiotic dinoflagellate type in the octocoral Briareum asbestinum. Marine Ecology Progress Series, 391: 65–72, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/391/m391p065.pdf.

Helbling:2001:LSI

[HBdBV01] E. W. Helbling, A. G. J. Buma, M. K. de Boer, and V. E. Villafañe. In situ impact of solar ultraviolet radiation on photosynthesis and DNA in temperate marine phytoplankton. Marine Ecology Progress Series, 211:43–49, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p043.pdf.

Harcourt:2002:FEG

[HBDD02] R. G. Harcourt, C. J. A. Bradshaw, K. Dickson, and L. S. Davis. Foraging ecology of a generalist predator, the female New Zealand fur seal. Marine Ecology Progress Series, 227:11–24, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p011.pdf.

Houghton:2002:DBD

[HBG^{+}02] J. D. R. Houghton, A. C. Broderick, B. J. Godley, J. D. Metcalfe, and G. C. Hays. Diving behaviour during the internesting interval for loggerhead turtles Caretta caretta nesting in Cyprus. Marine Ecology Progress Series, 227: 63–70, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/227/m227p063.pdf.

Hays:2003:STS

[HBG^{+}03] G. C. Hays, A. C. Broderick, B. J. Godley, P. Luschi, and W. J. Nichols. Satellite telemetry suggests high levels of fishing-induced mortality in marine turtles. Marine Ecology Progress Series, 262:305–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p305.pdf.

Hays:2004:TTT

[HBG^{+}04] G. C. Hays, A. C. Broderick, B. J. Godley, P. Luschi, and W. J. Nichols. Tracking turtles to their death. Marine Ecology Progress Series, 283:299–300, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p299.pdf. See

reply [CPB04b].

Hirst:2007:SDM

[HBH07] A. G. Hirst, D. Bonnet, and R. P. Harris. Seasonal dynamics and mortality rates of Calanus helgolandicus over two years at a station in the English Channel. Marine Ecology Progress Series, 340:189–205, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m340p189.pdf.

Hyland:2005:OCC

[HBK^{+}05] J. Hyland, L. Balthis, I. Karakassis, P. Magni, A. Petrov, J. Shine, O. Vestergaard, and R. Warwick. Organic carbon content of sediments as an indicator of stress in the marine benthos. Marine Ecology Progress Series, 295:91– 103, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/295/m295p091.pdf.

Hirche:2001:HFT

[HBN01] H. J. Hirche, T. Brey, and B. Niehoff. A high-frequency time series at Ocean Weather Ship Station M (Norwegian Sea): population dynamics of Calanus finmarchicus. Marine Ecology Progress Series, 219:205–219, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/219/m219p205.pdf.

Harris:2004:ESP

[HBNA04] L. A. Harris, B. Buckley, S. W. Nixon, and B. T. Allen. Experimental studies of predation by bluefish Pomatomus saltatrix in varying densities of seagrass and macroalgae. Marine Ecology Progress Series, 281:233–239, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p233.pdf.

Herlory:2005:DSM

[HBP^{+}05] O. Herlory, G. F. Blanchard, S. Planche, V. Huet, and P. Richard. Does the size of the microphytobenthic biofilm on intertidal mudflats depend on the available photosynthetic biomass? Marine Ecology Progress Series, 298:95– 100, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p095.pdf.

Hedd:2006:EIC

[HBRJ06] A. Hedd, D. F. Bertram, J. L. Ryder, and I. L. Jones. Effects of interdecadal climate variability on marine trophic interactions: rhinoceros auklets and their fish prey. Marine Ecology Progress Series, 309:263–278, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/309/m309p263.pdf.

Hughes:2004:REG

[HBRW04] A. R. Hughes, K. J. Bando, L. F. Rodriguez, and S. L. Williams. Relative effects of grazers and nutrients on seagrasses: a meta-analysis approach. Marine Ecology Progress Series, 282:87–99, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p087.pdf.

Hedman:2008:FCB

[HBT^{+}08] J. E. Hedman, C. Bradshaw, M. H. Thorsson, M. Gilek, and J. S. Gunnarsson. Fate of contaminants in Baltic Sea sediments: role of bioturbation and settling organic matter. Marine Ecology Progress Series, 356:25–38, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/356/
m356p025.pdf.

Howell:2004:FED

[HBTD04] K. L. Howell, D. S. M. Billett, P. A. Tyler, and R. Davidson. Feeding ecology of deep-sea seastars (Echinodermata: Asteroidea): a pigment biomarker approach. Marine Ecology Progress Series, 266:103–110, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/266/m266p103.pdf.

Herman:2005:FEE

[HBW^{+}05] D. P. Herman, D. G. Burrows, P. R. Wade, J. W. Durban, C. O. Matkin, R. G. LeDuc, L. G. Barrett-Lennard, and M. M. Krahn. Feeding ecology of eastern North Pacific killer whales Orcinus orca from fatty acid, stable isotope, and organochlorine analyses of blubber biopsies. Marine Ecology Progress Series, 302:275–291, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p275.pdf.

Holloway:2002:WDF

[HC02] M. G. Holloway and S. D. Connell. Why do floating structures create novel habitats for subtidal epibiota? Marine Ecology Progress Series, 235:43–52, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/235/m235p043.pdf.

Hennicke:2005:FPR

[HC05] J. C. Hennicke and B. M. Culik. Foraging performance and reproductive success of Humboldt penguins in relation to prey availability. Marine Ecology Progress Series, 296:173– 181, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/296/m296p173.pdf.

Holmer:2006:BIP

[HCA06] M. Holmer, C. Carta, and F. Ó. Andersen. Biogeochemical implications for phosphorus cycling in sandy and muddy rhizosphere sediments of Zostera marina meadows (Denmark). Marine Ecology Progress Series, 320:141–151, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p141.pdf.

Hanfling:2002:MDS

[HCB02] B. Hänfling, G. R. Carvalho, and R. Brandl. mt-DNA sequences and possible invasion pathways of the Chinese mitten crab. Marine Ecology Progress Series, 238:307–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/238/m238p307.pdf.

Harvey:2007:BAA

[HCB^{+}07] E. S. Harvey, M. Cappo, J. J. Butler, N. Hall, and G. A. Kendrick. Bait attraction affects the performance of remote underwater video stations in assessment of demersal fish community structure. Marine Ecology Progress Series, 350:245–254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m350p245.pdf.

Hopcroft:2005:PMS

[HCBP05] R. R. Hopcroft, C. Clarke, A. G. Byrd, and A. I. Pinchuk. The paradox of Metridia spp. egg production rates: a new technique and measurements from the coastal Gulf of Alaska. Marine Ecology Progress Series, 286:193–201, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p193.pdf.

Hsieh:2002:DBO

[HCCY02] H. L. Hsieh, C. P. Chen, Y. G. Chen, and H. H. Yang. Diversity of benthic organic matter flows through polychaetes and crabs in a mangrove estuary: δ ^{13}C and δ ^{34}S signals. Marine Ecology Progress Series, 227:145–155, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p145.pdf.

Hussey:2009:CCA

[HCD^{+}09] N. E. Hussey, D. T. Cocks, S. F. J. Dudley, I. D. McCarthy, and S. P. Wintner. The condition conundrum: application of multiple condition indices to the dusky shark Carcharhinus obscurus. Marine Ecology Progress Series, 380: 199–212, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/380/m380p199.pdf.

Hazen:2009:VDF

[HCGC09] E. L. Hazen, J. K. Craig, C. P. Good, and L. B. Crowder. Vertical distribution of fish biomass in hypoxic waters on the Gulf of Mexico shelf. Marine Ecology Progress Series, 375:195–207, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p195.pdf.

Hamner:2007:EID

[HCH07] W. M. Hamner, P. L. Colin, and P. P. Hamner. Export– import dynamics of zooplankton on a coral reef in Palau. Marine Ecology Progress Series, 334:83–92, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p083.pdf.

Hsu:2009:ESC

[HCH09] C. B. Hsu, C. P. Chen, and H. L. Hsieh. Effects of sediment compaction on macroinfauna in a protected coastal wetland in Taiwan. Marine Ecology Progress Series, 375:73–83, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m375p073.pdf.

Huettel:2007:TDD

[HCJ^{+}07] M. Huettel, P. Cook, F. Janssen, G. Lavik, and J. J. Middelburg. Transport and degradation of a dinoflagellate bloom in permeable sublittoral sediment. Marine Ecology Progress Series, 340:139–153, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p139.pdf.

Hughes:2005:GFA

[HCK^{+}05] A. D. Hughes, A. I. Catarino, M. S. Kelly, D. K. A. Barnes, and K. D. Black. Gonad fatty acids and trophic interactions of the echinoid Psammechinus miliaris. Marine Ecology Progress Series, 305:101–111, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/305/m305p101.pdf.

Heck:2001:PPS

[HCM01] K. L. Heck, Jr., L. D. Coen, and S. G. Morgan. Pre- and post-settlement factors as determinants of juvenile blue crab Callinectes sapidus abundance: results from the north-central Gulf of Mexico. Marine Ecology Progress Series, 222:163– 176, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/222/m222p163.pdf.

Hunter:2009:LSV

[HCMR09] E. Hunter, R. J. Cotton, J. D. Metcalfe, and J. D. Reynolds. Large-scale variation in seasonal swimming patterns of plaice in the North Sea. Marine Ecology Progress Series, 392:167– 178, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p167.pdf.

Herwig:2006:OTB

[HCP^{+}06] R. P. Herwig, J. R. Cordell, J. C. Perrins, P. A. Dinnel, R. W. Gensemer, W. A. Stubblefield, G. M. Ruiz, J. A. Kopp, M. L. House, and W. J. Cooper. Ozone treatment of ballast water on the oil tanker S/T Tonsina: chemistry, biology and toxicity. Marine Ecology Progress Series, 324:37–55, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p037.pdf.

Hutchinson:2001:MGS

[HCR01] W. F. Hutchinson, G. R. Carvalho, and S. I. Rogers. Marked genetic structuring in localised spawning populations of cod Gadus morhua in the North Sea and adjoining waters, as revealed by microsatellites. Marine Ecology Progress Series, 223:251–260, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p251.pdf.

Haslob:2007:IML

[HCS^{+}07a] H. Haslob, C. Clemmesen, M. Schaber, H. Hinrichsen, J. O. Schmidt, R. Voss, G. Kraus, and F. W. Köster. Invading Mnemiopsis leidyi as a potential threat to Baltic fish. Marine Ecology Progress Series, 349:303–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p303.pdf.

Heupel:2007:SNA

[HCS07b] M. R. Heupel, J. K. Carlson, and C. A. Simpfendorfer. Shark nursery areas: concepts, definition, characterization and assumptions. Marine Ecology Progress Series, 337:287– 297, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p287.pdf.

Hauxwell:2003:EZM

[HCV03a] J. Hauxwell, J. Cebrián, and I. Valiela. Eelgrass Zostera marina loss in temperate estuaries: relationship to land-derived nitrogen loads and effect of light limitation imposed by algae. Marine Ecology Progress Series, 247:59–73, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p059.pdf.

Hermsen:2003:MFG

[HCV03b] J. M. Hermsen, J. S. Collie, and P. C. Valentine. Mobile fishing gear reduces benthic megafaunal production on Georges Bank. Marine Ecology Progress Series, 260:97–108, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p097.pdf.

Hu:2005:API

[HD05] Q. Hu and C. Davis. Automatic plankton image recognition with co-occurrence matrices and support vector machine. Marine Ecology Progress Series, 295:21–31, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p021.pdf.

Hill:2006:MDD

[HD06a] R. W. Hill and J. W. H. Dacey. Metabolism of dimethylsulfoniopropionate (DMSP) by juvenile Atlantic menhaden Brevoortia tyrannus. Marine Ecology Progress Series, 322: 239–248, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p239.pdf.

Hu:2006:AAQ

[HD06b] Q. Hu and C. Davis. Accurate automatic quantification of taxa-specific plankton abundance using dual classification with correction. Marine Ecology Progress Series, 306: 51–61, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/306/m306p051.pdf.

Hill:2007:PID

[HD07a] R. W. Hill and J. W. H. Dacey. Processing of ingested dimethylsulfoniopropionate by mussels Mytilus edulis and scallops Argopecten irradians. Marine Ecology Progress Series, 343:131–140, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p131.pdf.

Hoeinghaus:2007:SBT

[HD07b] D. J. Hoeinghaus and S. E. Davis III. Size-based trophic shifts of saltmarsh dwelling blue crabs elucidated by dual stable C and N isotope analyses. Marine Ecology Progress Series, 334:199–204, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p199.pdf.

Hargrave:2008:IMA

[HDC^{+}08] B. T. Hargrave, L. I. Doucette, P. J. Cranford, B. A. Law, and T. G. Milligan. Influence of mussel aquaculture on sediment organic enrichment in a nutrient-rich coastal embayment. Marine Ecology Progress Series, 365:137–149, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p137.pdf.

Hubas:2006:FCB

[HDCA06] C. Hubas, D. Davoult, T. Cariou, and L. F. Artigas. Factors controlling benthic metabolism during low tide along a granulometric gradient in an intertidal bay (Roscoff Aber Bay, France). Marine Ecology Progress Series, 316:53–68, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/316/m316p053.pdf.

Houghton:2006:DSR

[HDDH06] J. D. R. Houghton, T. K. Doyle, J. Davenport, and G. C. Hays. Developing a simple, rapid method for identifying and monitoring jellyfish aggregations from the air. Marine Ecology Progress Series, 314:159–170, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/314/m314p159.pdf.

Heterier:2008:ECF

[HDDR08] V. Hétérier, B. David, C. De Ridder, and T. Rigaud. Ectosymbiosis is a critical factor in the local benthic biodiversity of the Antarctic deep sea. Marine Ecology Progress Series, 364:67–76, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p067.pdf.

Himmelman:2008:SSA

[HDG^{+}08] J. H. Himmelman, C. P. Dumont, C. F. Gaymer, C. Vallières, and D. Drolet. Spawning synchrony and aggregative behaviour of cold-water echinoderms during multi-species mass spawnings. Marine Ecology Progress Series, 361:161–168, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/361/m361p161.pdf.

Hansson:2001:RDF

[HDGM01] S. Hansson, B. T. De Stasio, E. Gorokhova, and M. A. Mohammadian. Ratio-dependent functional responses — tests with the zooplanktivore Mysis mixta. Marine Ecology Progress Series, 216:181–189, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/216/m216p181.pdf.

Heffernan:2004:DSS

[HDM04] O. A. Heffernan, B. S. Danilowicz, and S. P. Milligan. Determination of species-specific spawning distributions of commercial finfish in the Irish Sea using a biochemical proteinbased method. Marine Ecology Progress Series, 284:279– 291, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p279.pdf.

Hagerthey:2002:ICV

[HDP02] S. E. Hagerthey, E. C. Defew, and D. M. Paterson. Influence of Corophium volutator and Hydrobia ulvae on intertidal benthic diatom assemblages under different nutrient and temperature regimes. Marine Ecology Progress Series, 245:47–59, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p047.pdf.

Hereu:2006:TPS

[HDP^{+}06] B. Hereu, D. Diaz, J. Pasqual, M. Zabala, and E. Sala. Temporal patterns of spawning of the dusky grouper Epinephelus marginatus in relation to environmental factors. Marine Ecology Progress Series, 325:187–194, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p187.pdf.

Hu:2008:SAS

[HDP08] Q. Hu, C. S. Davis, and C. M. Petrik. A simplified age-stage model for copepod population dynamics. Marine Ecology Progress Series, 360:179–187, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p179.pdf.

Harder:2004:WPM

[HDQ04] T. Harder, S. Dobretsov, and P. Y. Qian. Waterborne polar macromolecules act as algal antifoulants in the seaweed Ulva reticulata. Marine Ecology Progress Series, 274:133– 141, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/274/m274p133.pdf.

Hoskins:2008:FBH

[HDRC^{+}08] A. J. Hoskins, P. Dann, Y. Ropert-Coudert, A. Kato, A. Chiaradia, D. P. Costa, and J. P. Y. Arnould. Foraging behaviour and habitat selection of the little penguin Eudyptula minor during early chick rearing in Bass Strait, Australia. Marine Ecology Progress Series, 366:293–303, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/366/
m366p293.pdf.

Haring:2002:DFW

[HDW02] R. N. Haring, M. N. Dethier, and S. L. Williams. Desiccation facilitates wave-induced mortality of the intertidal alga Fucus gardneri. Marine Ecology Progress Series, 232: 75–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/232/m232p075.pdf.

Holmes:2004:PDG

[HDW04] S. P. Holmes, R. Dekker, and I. D. Williams. Population dynamics and genetic differentiation in the bivalve mollusc Abra tenuis: aplanic dispersal. Marine Ecology Progress Series, 268:131–140, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/268/m268p131.pdf.

Herter:2008:TDC

[HE08] H. Herter and G. L. Eckert. Transport of Dungeness crab Cancer magister megalopae into Glacier Bay, Alaska. Marine Ecology Progress Series, 372:181–194, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p181.pdf.

Helle:2000:DMS

[Hel00] K. Helle. Does the midnight sun increase the feeding rate and hence the growth rate of early juvenile Arcto–Norwegian cod Gadus morhua in the Barents Sea? Marine Ecology Progress Series, 197:293–297, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/197/m197p293.pdf.

Hense:2007:RFM

[Hen07] I. Hense. Regulative feedback mechanisms in cyanobacteriadriven systems: a model study. Marine Ecology Progress Series, 339:41–47, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p041.pdf.

Holt:2008:SIA

[HEN^{+}08] B. G. Holt, B. C. Emerson, J. Newton, M. J. G. Gage, and I. M. Côté. Stable isotope analysis of the Hypoplectrus species complex reveals no evidence for dietary niche divergence. Marine Ecology Progress Series, 357:283–289, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p283.pdf.

Henkel:2009:COA

[Hen09] L. A. Henkel. Co-occurrence and aggregation of marine birds and mammals in Monterey Bay, California, USA. Marine Ecology Progress Series, 387:295–303, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p295.pdf.

Hereu:2006:DPA

[Her06] B. Hereu. Depletion of palatable algae by sea urchins and fishes in a Mediterranean subtidal community. Marine Ecology Progress Series, 313:95–103, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/313/m313p095.pdf.

Haussermann:2003:FEC

[HF03] V. Häussermann and G. Försterra. First evidence for coloniality in sea anemones. Marine Ecology Progress Series, 257:291–294, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p291.pdf.

Hovel:2005:ISL

[HF05] K. A. Hovel and M. S. Fonseca. Influence of seagrass landscape structure on juvenile blue crab habitat-survival function. Marine Ecology Progress Series, 300:179–191, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/300/
m300p179.pdf.

Hewson:2006:SVB

[HF06] I. Hewson and J. A. Fuhrman. Spatial and vertical biogeography of coral reef sediment bacterial and diazotroph communities. Marine Ecology Progress Series, 306:79–86, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/306/m306p079.pdf.

Halpern:2008:FDR

[HF08] B. S. Halpern and S. R. Floeter. Functional diversity responses to changing species richness in reef fish communities. Marine Ecology Progress Series, 364:147–156, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p147.pdf.

Hatch:2009:NBB

[HF09] L. T. Hatch and K. M. Fristrup. No barrier at the boundaries: implementing regional frameworks for noise management in protected natural areas. Marine Ecology Progress Series, 395:223–244, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p223.pdf.

Hyrenbach:2002:OHT

[HFA02] K. D. Hyrenbach, P. Fernández, and D. J. Anderson. Oceanographic habitats of two sympatric North Pacific albatrosses during the breeding season. Marine Ecology Progress Series, 233:283–301, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p283.pdf.

Hunter:2009:IFF

[HFA09] K. L. Hunter, M. G. Fox, and K. W. Able. Influence of flood frequency, temperature and population density on migration of Fundulus heteroclitus in semi-isolated marsh pond habitats. Marine Ecology Progress Series, 391:85–96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/391/
m391p085.pdf.

Hjorth:2008:PSR

[HFD08] M. Hjorth, V. E. Forbes, and I. Dahllöf. Plankton stress responses from PAH exposure and nutrient enrichment. Marine Ecology Progress Series, 363:121–130, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p121.pdf.

Hughes:2009:SDS

[HFH09] R. G. Hughes, P. W. Fletcher, and M. J. Hardy. Successional development of saltmarsh in two managed realignment areas in SE England, and prospects for saltmarsh restoration. Marine Ecology Progress Series, 384:13–22, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/384/
m384p013.pdf.

Hovel:2002:ESL

[HFM^{+}02] K. A. Hovel, M. S. Fonseca, D. L. Myer, W. J. Kenworthy, and P. E. Whitfield. Effects of seagrass landscape structure, structural complexity and hydrodynamic regime on macrofaunal densities in North Carolina seagrass beds. Marine Ecology Progress Series, 243:11–24, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/243/m243p011.pdf.

Huvet:2004:NHB

[HFM^{+}04] A. Huvet, C. Fabioux, H. McCombie, S. Lapègue, and P. Boudry. Natural hybridization between genetically differentiated populations of Crassostrea gigas and C. angulata highlighted by sequence variation in flanking regions of a microsatellite locus. Marine Ecology Progress Series, 272: 141–152, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/272/m272p141.pdf.

Huggett:2003:MTS

[HFMP03] J. Huggett, P. Fréon, C. Mullon, and P. Penven. Modelling the transport success of anchovy Engraulis encrasicolus eggs and larvae in the southern Benguela: the effect of spatiotemporal spawning patterns. Marine Ecology Progress Series, 250:247–262, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p247.pdf.

Hazen:2009:FSP

[HFT^{+}09] E. L. Hazen, A. S. Friedlaender, M. A. Thompson, C. R. Ware, M. T. Weinrich, P. N. Halpin, and D. N. Wiley. Finescale prey aggregations and foraging ecology of humpback whales Megaptera novaeangliae. Marine Ecology Progress Series, 395:75–89, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p075.pdf.

Heithaus:2005:BST

[HFW^{+}05] M. R. Heithaus, A. Frid, A. J. Wirsing, L. Bejder, and L. M. Dill. Biology of sea turtles under risk from tiger sharks at a foraging ground. Marine Ecology Progress Series, 288:285–294, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/288/m288p285.pdf.

Hinke:2005:PHU

[HFWW05] J. T. Hinke, D. G. Foley, C. Wilson, and G. M. Watters. Persistent habitat use by Chinook salmon Oncorhynchus tshawytscha in the coastal ocean. Marine Ecology Progress Series, 304:207–220, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p207.pdf.

Hallier:2008:DFA

[HG08] J. P. Hallier and D. Gaertner. Drifting fish aggregation devices could act as an ecological trap for tropical tuna species. Marine Ecology Progress Series, 353:255–264, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p255.pdf.

Henriques:2007:RSM

[HGA07] M. Henriques, E. J. Gonçalves, and V. C. Almada. Rapid shifts in a marine fish assemblage follow fluctuations in winter sea conditions. Marine Ecology Progress Series, 340:259– 270, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/340/m340p259.pdf.

Heil:2001:FRF

[HGAS^{+}01] C. A. Heil, P. M. Glibert, M. A. Al-Sarawi, M. Faraj, M. Behbehani, and M. Husain. First record of a fish-killing Gymnodinium sp. bloom in Kuwait Bay, Arabian Sea: chronology and potential causes. Marine Ecology Progress Series, 214: 15–23, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p015.pdf.

Hedd:2001:FSS

[HGB01] A. Hedd, R. Gales, and N. Brothers. Foraging strategies of shy albatross Thalassarche cauta breeding at Albatross Island, Tasmania, Australia. Marine Ecology Progress Series, 224:267–282, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p267.pdf.

Heyl:2007:CVC

[HGC^{+}07] T. P. Heyl, W. P. Gilhooly, R. M. Chambers, G. W. Gilchrist, S. A. Macko, C. D. Ruppel, and C. L. Van Dover. Characteristics of vesicomyid clams and their environment at the Blake Ridge cold seep, South Carolina, USA. Marine Ecology Progress Series, 339:169–184, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m339p169.pdf.

Helbling:2003:USU

[HGG^{+}03] E. W. Helbling, K. Gao, R. J. Gonçalves, H. Wu, and V. E. Villafañe. Utilization of solar UV radiation by coastal phytoplankton assemblages off SE China when exposed to fast mixing. Marine Ecology Progress Series, 259:59–66, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/259/m259p059.pdf.

Holland:2005:ISA

[HGG^{+}05] G. J. Holland, S. P. R. Greenstreet, I. M. Gibb, H. M. Fraser, and M. R. Robertson. Identifying sandeel Ammodytes marinus sediment habitat preferences in the marine environment. Marine Ecology Progress Series, 303:269–282, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p269.pdf.

Hind:2000:OSC

[HGHB00] A. Hind, W. S. C. Gurney, M. Heath, and A. D. Bryant. Overwintering strategies in Calanus finmarchicus. Marine Ecology Progress Series, 193:95–107, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p095.pdf.

Hillebrand:2009:CDI

[HGJM09] H. Hillebrand, L. Gamfeldt, P. R. Jonsson, and B. Matthiessen. Consumer diversity indirectly changes prey nutrient content. Marine Ecology Progress Series, 380:33–41, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/380/
m380p033.pdf.

Holmer:2002:ESL

[HGK02] M. Holmer, B. Gribsholt, and E. Kristensen. Effects of sea level rise on growth of Spartina anglica and oxygen dynamics in rhizosphere and salt marsh sediments. Marine Ecology Progress Series, 225:197–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/225/m225p197.pdf.

Heyman:2001:WSR

[HGKJ01] W. D. Heyman, R. T. Graham, B. Kjerfve, and R. E. Johannes. Whale sharks Rhincodon typus aggregate to feed on fish spawn in Belize. Marine Ecology Progress Series, 215:275–282, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p275.pdf.

Hoppe:2006:CBG

[HGKK06] H. G. Hoppe, K. Gocke, R. Koppe, and G. Kraus. Changing bacterioplankton growth characteristics on a large spatial scale: oligotrophic versus mesotrophic ocean. Marine Ecology Progress Series, 323:21–33, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/323/m323p021.pdf.

Hoegh-Guldberg:2004:NIP

[HGMG^{+}04] O. Hoegh-Guldberg, L. Muscatine, C. Goiran, D. Siggaard, and G. Marion. Nutrient-induced perturbations to δ ^{13}C and δ ^{15}N in symbiotic dinoflagellates and their coral hosts. Marine Ecology Progress Series, 280:105–114, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p105.pdf.

Huxham:2006:MMS

[HGMH06] M. Huxham, L. Gilpin, M. Mocogni, and S. Harper. Microalgae, macrofauna and sediment stability: an experimental test of a reciprocal relationship. Marine Ecology Progress Series, 310:55–63, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/310/m310p055.pdf.

Herlory:2004:MMB

[HGRB04] O. Herlory, J. M. Guarini, P. Richard, and G. F. Blanchard. Microstructure of microphytobenthic biofilm and its spatio-temporal dynamics in an intertidal mudflat (Aiguillon Bay, France). Marine Ecology Progress Series, 282:33– 44, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p033.pdf.

Hatase:2002:UAB

[HGS^{+}02] H. Hatase, K. Goto, K. Sato, T. Bando, Y. Matsuzawa, and W. Sakamoto. Using annual body size fluctuations to explore potential causes for the decline in a nesting population of the loggerhead turtle Caretta caretta at Senri Beach, Japan. Marine Ecology Progress Series, 245:299–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p299.pdf.

Harnstrom:2007:TPC

[HGS^{+}07] K. Härnström, A. Godhe, V. Saravanan, I. Karunasagar, I. Karunasagar, and A. Rehnstam-Holm. Tropical phytoplankton community development in mesocosms inoculated with different life stages. Marine Ecology Progress Series, 346: 75–88, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m346p075.pdf.

Harvey:2008:TFI

[HGSH08] C. J. Harvey, K. Gross, V. H. Simon, and J. Hastie. Trophic and fishery interactions between Pacific hake and rockfish: effect on rockfish population rebuilding times. Marine Ecology Progress Series, 365:165–176, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/365/m365p165.pdf.

Hopkinson:2001:BMN

[HGT01] C. S. Hopkinson, Jr., A. E. Giblin, and J. Tucker. Benthic metabolism and nutrient regeneration on the continental shelf of Eastern Massachusetts, USA. Marine Ecology Progress Series, 224:1–19, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p001.pdf.

Huntley:2000:ZDM

[HGZ^{+}00] M. E. Huntley, A. González, Y. Zhu, M. Zhou, and X. Irigoien. Zooplankton dynamics in a mesoscale eddy-jet system off California. Marine Ecology Progress Series, 201:165–178, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p165.pdf.

Hodum:2000:TRA

[HH00a] P. J. Hodum and K. A. Hobson. Trophic relationships among Antarctic fulmarine petrels: insights into dietary overlap and chick provisioning strategies inferred from stable-isotope (δ 15N and δ ^{13}C) analyses. Marine Ecology Progress Series, 198:273–281, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p273.pdf.

Holt:2000:VDR

[HH00b] G. J. Holt and S. A. Holt. Vertical distribution and the role of physical processes in the feeding dynamics of two larval sciaenids Sciaenops ocellatus and Cynoscion nebulosus. Marine Ecology Progress Series, 193:181–190, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p181.pdf.

Halsband:2001:RCD

[HH01] C. Halsband and H. J. Hirche. Reproductive cycles of dominant calanoid copepods in the North Sea. Marine Ecology Progress Series, 209:219–229, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p219.pdf.

Hamersley:2003:CDN

[HH03a] M. R. Hamersley and B. L. Howes. Contribution of denitrification to nitrogen, carbon, and oxygen cycling in tidal creek sediments of a New England salt marsh. Marine Ecology Progress Series, 262:55–69, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p055.pdf.

Huang:2003:CDB

[HH03b] S. Huang and M. G. Hadfield. Composition and density of bacterial biofilms determine larval settlement of the polychaete Hydroides elegans. Marine Ecology Progress Series, 260:161–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p161.pdf.

Hamersley:2005:CND

[HH05a] M. R. Hamersley and B. L. Howes. Coupled nitrificationdenitrification measured in situ in a Spartina alterniflora marsh with a ^{15}NH_{4}^{+} tracer. Marine Ecology Progress Series, 299:123–135, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p123.pdf.

Hepburn:2005:CMB

[HH05b] C. D. Hepburn and C. L. Hurd. Conditional mutualism between the giant kelp Macrocystis pyrifera and colonial epifauna. Marine Ecology Progress Series, 302:37–48, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/302/m302p037.pdf.

Hollebone:2007:PDN

[HH07a] A. L. Hollebone and M. E. Hay. Population dynamics of the non-native crab Petrolisthes armatus invading the South Atlantic Bight at densities of thousands m-2. Marine Ecology Progress Series, 336:211–223, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/336/m336p211.pdf.

Hollebone:2007:PPI

[HH07b] A. L. Hollebone and M. E. Hay. Propagule pressure of an invasive crab overwhelms native biotic resistance. Marine Ecology Progress Series, 342:191–196, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p191.pdf.

Holm-Hansen:2000:REC

[HHAH00] O. Holm-Hansen, A. F. Amos, and C. D. Hewes. Reliability of estimating chlorophyll-a concentrations in Antarctic waters by measurement of in situ chlorophyll-a fluorescence. Marine Ecology Progress Series, 196:103–110, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p103.pdf.

Hays:2000:DBG

[HHB^{+}00] G. C. Hays, S. Hochscheid, A. C. Broderick, B. J. Godley, and J. D. Metcalfe. Diving behaviour of green turtles: dive depth, dive duration and activity levels. Marine Ecology Progress Series, 208:297–298, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/208/m208p297.pdf.

Head:2000:IEC

[HHC00] E. J. H. Head, L. R. Harris, and R. W. Campbell. Investigations on the ecology of Calanus spp. in the Labrador Sea. I. Relationship between the phytoplankton bloom and reproduction and development of Calanus finmarchicus in spring. Marine Ecology Progress Series, 193:53–73, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p053.pdf.

Hedger:2009:MSD

[HHD^{+}09] R. D. Hedger, D. Hatin, J. J. Dodson, F. Martin, D. Fournier, F. Caron, and F. G. Whoriskey. Migration and swimming depth of Atlantic salmon kelts Salmo salar in coastal zone and marine habitats. Marine Ecology Progress Series, 392: 179–192, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p179.pdf.

Halsey:2007:FSA

[HHF^{+}07] L. G. Halsey, Y. Handrich, A. Fahlman, A. Schmidt, C. A. Bost, R. L. Holder, A. J. Woakes, and P. J. Butler. Fine-scale analyses of diving energetics in king penguins Aptenodytes patagonicus: how behaviour affects costs of a foraging dive. Marine Ecology Progress Series, 344:299–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p299.pdf.

Hamer:2007:AVD

[HHG^{+}07] K. C. Hamer, E. M. Humphreys, S. Garthe, J. Hennicke, G. Peters, D. Grémillet, R. A. Phillips, M. P. Harris, and S. Wanless. Annual variation in diets, feeding locations and foraging behaviour of gannets in the North Sea: flexibility, consistency and constraint. Marine Ecology Progress Series, 338:295–305, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p295.pdf.

Herzka:2002:CSP

[HHH02] S. Z. Herzka, S. A. Holt, and G. J. Holt. Characterization of settlement patterns of red drum Sciaenops ocellatus larvae to estuarine nursery habitat: a stable isotope approach. Marine Ecology Progress Series, 226:143–156, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p143.pdf.

Hashimoto:2006:MBG

[HHIS06] S. Hashimoto, N. Horimoto, T. Ishimaru, and T. Saino. Metabolic balance of gross primary production and community respiration in Sagami Bay, Japan. Marine Ecology Progress Series, 321:31–40, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/321/m321p031.pdf.

Hemmi:2004:IRH

[HHJ04] A. Hemmi, T. Honkanen, and V. Jormalainen. Inducible resistance to herbivory in Fucus vesiculosus — duration, spreading and variation with nutrient availability. Marine Ecology Progress Series, 273:109–120, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/273/m273p109.pdf.

Holm-Hansen:2005:DCM

[HHKH05] O. Holm-Hansen, M. Kahru, and C. D. Hewes. Deep chlorophyll-a maxima (DCMs) in pelagic Antarctic waters. II. Relation to bathymetric features and dissolved iron concentrations. Marine Ecology Progress Series, 297:71–81, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/297/m297p071.pdf.

Hoover:2006:ZRS

[HHM^{+}06] R. S. Hoover, D. Hoover, M. Miller, M. R. Landry, E. H. De-Carlo, and F. T. Mackenzie. Zooplankton response to storm runoff in a tropical estuary: bottom-up and top-down controls. Marine Ecology Progress Series, 318:187–201, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p187.pdf.

Heck:2003:CEN

[HHO03] K. L. Heck, Jr., G. Hays, and R. J. Orth. Critical evaluation of the nursery role hypothesis for seagrass meadows. Marine Ecology Progress Series, 253:123–136, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p123.pdf.

Hofmann:2009:DCC

[HHT^{+}09] T. Hofmann, A. R. M. Hanlon, J. D. Taylor, A. S. Ball, A. M. Osborn, and G. J. C. Underwood. Dynamics and compositional changes in extracellular carbohydrates in estuarine sediments during degradation. Marine Ecology Progress Series, 379:45–58, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/379/m379p045.pdf.

Hall:2006:FSH

[HHV06] L. M. Hall, M. D. Hanisak, and R. W. Virnstein. Fragments of the seagrasses Halodule wrightii and Halophila johnsonii as potential recruits in Indian River Lagoon, Florida. Marine Ecology Progress Series, 310:109–117, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p109.pdf.

Hepburn:2007:EWE

[HHW^{+}07] C. D. Hepburn, J. D. Holborow, S. R. Wing, R. D. Frew, and C. L. Hurd. Exposure to waves enhances the growth rate and nitrogen status of the giant kelp Macrocystis pyrifera. Marine Ecology Progress Series, 339:99–108, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p099.pdf.

Hindell:2002:FST

[HHWT02] M. A. Hindell, R. Harcourt, J. R. Waas, and D. Thompson. Fine-scale three-dimensional spatial use by diving, lactating female Weddell seals Leptonychotes weddellii. Marine Ecology Progress Series, 242:275–284, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/242/m242p275.pdf.

Hiddink:2003:MAV

[Hid03] J. G. Hiddink. Modelling the adaptive value of intertidal migration and nursery use in the bivalve Macoma balthica. Marine Ecology Progress Series, 252:173–185, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p173.pdf.

Hillebrand:2004:SSV

[Hil04] H. Hillebrand. Strength, slope and variability of marine latitudinal gradients. Marine Ecology Progress Series, 273:251– 267, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/273/m273p251.pdf.

Hildebrand:2009:ANS

[Hil09] J. A. Hildebrand. Anthropogenic and natural sources of ambient noise in the ocean. Marine Ecology Progress Series, 395:5–20, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p005.pdf.

Hasegawa:2005:NDS

[HIM05] N. Hasegawa, H. Iizumi, and H. Mukai. Nitrogen dynamics of the surfgrass Phyllospadix iwatensis. Marine Ecology Progress Series, 293:59–68, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p059.pdf.

Hinga:2002:EPC

[Hin02] K. R. Hinga. Effects of pH on coastal marine phytoplankton. Marine Ecology Progress Series, 238:281–300, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p281.pdf.

Hipfner:2008:MMO

[Hip08] J. M. Hipfner. Matches and mismatches: ocean climate, prey phenology and breeding success in a zooplanktivorous seabird. Marine Ecology Progress Series, 368:295–304, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p295.pdf.

Hipfner:2009:EDN

[Hip09] J. M. Hipfner. Euphausiids in the diet of a North Pacific seabird: annual and seasonal variation and the role of ocean climate. Marine Ecology Progress Series, 390:277– 289, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p277.pdf.

Hirst:2006:ITR

[Hir06] A. J. Hirst. Influence of taxonomic resolution on multivariate analyses of arthropod and macroalgal reef assemblages. Marine Ecology Progress Series, 324:83–93, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p083.pdf.

Hansen:2003:AAP

[HJ03] J. L. S. Hansen and A. B. Josefson. Accumulation of algal pigments and live planktonic diatoms in aphotic sediments during the spring bloom in the transition zone of the North and Baltic Seas. Marine Ecology Progress Series, 248:41–54, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/248/m248p041.pdf.

Hindell:2005:APF

[HJ05] J. S. Hindell and G. P. Jenkins. Assessing patterns of fish zonation in temperate mangroves, with emphasis on evaluating sampling artefacts. Marine Ecology Progress Series, 290:193–205, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/290/m290p193.pdf.

Huff:2007:SAA

[HJ07] T. M. Huff and J. K. Jarett. Sand addition alters the invertebrate community of intertidal coralline turf. Marine Ecology Progress Series, 345:75–82, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/345/m345p075.pdf.

Hamer:2003:OCJ

[HJG03] P. A. Hamer, G. P. Jenkins, and B. M. Gillanders. Otolith chemistry of juvenile snapper Pagrus auratus in Victorian waters: natural chemical tags and their temporal variation. Marine Ecology Progress Series, 263:261–273, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p261.pdf.

Humborstad:2006:ECG

[HJG06] O. B. Humborstad, T. Jørgensen, and S. Grotmol. Exposure of cod Gadus morhua to resuspended sediment: an experimental study of the impact of bottom trawling. Marine Ecology Progress Series, 309:247–254, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/309/m309p247.pdf.

Hardege:2002:NBA

[HJH^{+}02] J. D. Hardege, A. Jennings, D. Hayden, C. T. Müller, D. Pascoe, M. G. Bentley, and A. S. Clare. Novel behavioural assay and partial purification of a female-derived sex pheromone in Carcinus maenas. Marine Ecology Progress Series, 244: 179–189, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p179.pdf.

Hindell:2001:STV

[HJK01] J. S. Hindell, G. P. Jenkins, and M. J. Keough. Spatial and temporal variability in the effects of fish predation on macrofauna in relation to habitat complexity and cage effects. Marine Ecology Progress Series, 224:231–250, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p231.pdf.

Hollows:2007:CRF

[HJM07] C. F. Hollows, E. L. Johnston, and D. J. Marshall. Copper reduces fertilisation success and exacerbates Allee effects in the field. Marine Ecology Progress Series, 333: 51–60, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p051.pdf.

Hindell:2008:HUM

[HJW08] J. S. Hindell, G. P. Jenkins, and B. Womersley. Habitat utilisation and movement of black bream Acanthopagrus butcheri (Sparidae) in an Australian estuary. Marine Ecology Progress Series, 366:219–229, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/366/m366p219.pdf.

Hata:2002:WHD

[HK02a] H. Hata and M. Kato. Weeding by the herbivorous damselfish Stegastes nigricans in nearly monocultural algae farms. Marine Ecology Progress Series, 237:227–231, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p227.pdf.

Hirst:2002:MMP

[HK02b] A. G. Hirst and T. Kiørboe. Mortality of marine planktonic copepods: global rates and patterns. Marine Ecology Progress Series, 230:195–209, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/230/m230p195.pdf.

Holloway:2002:EIP

[HK02c] M. G. Holloway and M. J. Keough. Effects of an introduced polychaete, Sabella spallanzanii, on the development of epifaunal assemblages. Marine Ecology Progress Series, 236:137–154, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p137.pdf.

Hata:2003:DMA

[HK03a] H. Hata and M. Kato. Demise of monocultural algal farms by exclusion of territorial damselfish. Marine Ecology Progress Series, 263:159–167, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/263/m263p159.pdf.

Hazzard:2003:EPC

[HK03b] S. E. Hazzard and G. S. Kleppel. Egg production of the copepod Acartia tonsa in Florida Bay: role of fatty acids in the nutritional composition of the food environment. Marine Ecology Progress Series, 252:199–206, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p199.pdf.

Hyun:2003:BAP

[HK03c] J. H. Hyun and K. H. Kim. Bacterial abundance and production during the unique spring phytoplankton bloom in the central Yellow Sea. Marine Ecology Progress Series, 252: 77–88, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/252/m252p077.pdf.

Henry:2004:DBE

[HK04a] L. A. Henry and E. Kenchington. Differences between epilithic and epizoic hydroid assemblages from commercial scallop grounds in the Bay of Fundy, northwest Atlantic. Marine Ecology Progress Series, 266:123–134, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p123.pdf.

Huckstadt:2004:IBS

[HK04b] L. A. Hückstädt and M. C. Krautz. Interaction between southern sea lions Otaria flavescens and jack mackerel Trachurus symmetricus commercial fishery off Central Chile: a geostatistical approach. Marine Ecology Progress Series, 282: 285–294, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p285.pdf.

Hargrave:2004:BEA

[HKH04] B. T. Hargrave, V. E. Kostylev, and C. M. Hawkins. Benthic epifauna assemblages, biomass and respiration in the gully region on the Scotian Shelf, NW Atlantic Ocean. Marine Ecology Progress Series, 270:55–70, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/270/m270p055.pdf.

Henkel:2009:IIV

[HKH09] S. K. Henkel, H. Kawai, and G. E. Hofmann. Interspecific and interhabitat variation in hsp70 gene expression in native and invasive kelp populations. Marine Ecology Progress Series, 386:1–13, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Henry:2006:IOT

[HKK^{+}06] L. A. Henry, E. L. R. Kenchington, T. J. Kenchington, K. G. MacIsaac, C. Bourbonnais-Boyce, and D. C. Gordon, Jr. Impacts of otter trawling on colonial epifaunal assemblages on a cobble bottom ecosystem on Western Bank (northwest Atlantic). Marine Ecology Progress Series, 306: 63–78, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/306/m306p063.pdf.

Hasegawa:2000:RDO

[HKM00] T. Hasegawa, I. Koike, and H. Mukai. Release of dissolved organic nitrogen by size-fractionated natural planktonic assemblages in coastal waters. Marine Ecology Progress Series, 198:43–49, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p043.pdf.

Hasegawa:2001:FFN

[HKM01] T. Hasegawa, I. Koike, and H. Mukai. Fate of food nitrogen in marine copepods. Marine Ecology Progress Series, 210:167–174, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p167.pdf.

Hyde:2008:MPM

[HKR^{+}08] J. R. Hyde, C. Kimbrell, L. Robertson, K. Clifford, E. Lynn, and R. Vetter. Multiple paternity and maintenance of genetic diversity in the live-bearing rockfishes Sebastes spp. Marine Ecology Progress Series, 357:245–253, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p245.pdf.

Hoyer:2001:PSA

[HKSW01] K. Hoyer, U. Karsten, T. Sawall, and C. Wiencke. Photoprotective substances in Antarctic macroalgae and their variation with respect to depth distribution, different tissues and developmental stages. Marine Ecology Progress Series, 211:117–129, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/211/m211p117.pdf.

Hammerstrom:2007:RRD

[HKWM07] K. K. Hammerstrom, W. J. Kenworthy, P. E. Whitfield, and M. F. Merello. Response and recovery dynamics of seagrasses Thalassia testudinum and Syringodium filiforme and macroalgae in experimental motor vessel disturbances. Marine Ecology Progress Series, 345:83–92, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/345/m345p083.pdf.

Hata:2002:OCF

[HKY^{+}02] H. Hata, S. Kudo, H. Yamano, N. Kurano, and H. Kayanne. Organic carbon flux in Shiraho coral reef (Ishigaki Island, Japan). Marine Ecology Progress Series, 232:129–140, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/232/m232p129.pdf.

Hanson:2004:ECO

[HKZ04] P. J. Hanson, C. C. Koenig, and V. S. Zdanowicz. Elemental composition of otoliths used to trace estuarine habitats of juvenile gag Mycteroperca microlepis along the west coast of Florida. Marine Ecology Progress Series, 267:253– 265, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/267/m267p253.pdf.

Hofmann:2000:MGD

[HL00] E. E. Hofmann and C. M. Lascara. Modeling the growth dynamics of Antarctic krill Euphausia superba. Marine Ecology Progress Series, 194:219–231, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/194/m194p219.pdf.

Hentschel:2005:GRI

[HL05] B. T. Hentschel and A. A. Larson. Growth rates of interfacefeeding polychaetes: combined effects of flow speed and suspended food concentration. Marine Ecology Progress Series, 293:119–129, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p119.pdf.

Heagney:2007:PFA

[HLBS07] E. C. Heagney, T. P. Lynch, R. C. Babcock, and I. M. Suthers. Pelagic fish assemblages assessed using mid-water baited video: standardising fish counts using bait plume size. Marine Ecology Progress Series, 350:255–266, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m350p255.pdf.

Hare:2007:CIT

[HLD^{+}07] C. E. Hare, K. Leblanc, G. R. DiTullio, R. M. Kudela, Y. Zhang, P. A. Lee, S. Riseman, and D. A. Hutchins. Consequences of increased temperature and CO_{2} for phytoplankton community structure in the Bering Sea. Marine Ecology Progress Series, 352:9–16, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/352/m352p009.pdf.

Hoareau:2007:OLL

[HLFG^{+}07] T. B. Hoareau, R. Lecomte-Finiger, H. P. Grondin, C. Conand, and P. Berrebi. Oceanic larval life of La Réunion ‘bichiques’, amphidromous gobiid post-larvae. Marine Ecology Progress Series, 333:303–308, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/333/m333p303.pdf.

Hereu:2006:CPG

[HLGCO06] C. M. Hereu, B. E. Lavaniegos, G. Gaxiola-Castro, and M. D. Ohman. Composition and potential grazing impact of salp assemblages off Baja California during the 1997– 1999 El Niño and La Niña. Marine Ecology Progress Series, 318:123–140, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/318/m318p123.pdf.

Hoglander:2004:VDS

[HLH04] H. Höglander, U. Larsson, and S. Hajdu. Vertical distribution and settling of spring phytoplankton in the offshore NW Baltic Sea proper. Marine Ecology Progress Series, 283: 15–27, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/283/m283p015.pdf.

Hyde:2005:SIF

[HLH^{+}05] J. R. Hyde, E. Lynn, R. Humphreys, Jr., M. Musyl, A. P. West, and R. Vetter. Shipboard identification of fish eggs and larvae by multiplex PCR, and description of fertilized eggs of blue marlin, shortbill spearfish, and wahoo. Marine Ecology Progress Series, 286:269–277, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/286/m286p269.pdf.

Hauser:2007:SDP

[HLH^{+}07] D. D. W. Hauser, M. G. Logsdon, E. E. Holmes, G. R. Van-Blaricom, and R. W. Osborne. Summer distribution patterns of southern resident killer whales Orcinus orca: core areas and spatial segregation of social groups. Marine Ecology Progress Series, 351:301–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/351/m351p301.pdf.

Huang:2008:RPP

[HLH^{+}08] C. H. Huang, H. J. Lin, T. C. Huang, H. M. Su, and J. J. Hung. Responses of phytoplankton and periphyton to systemscale removal of oyster-culture racks from a eutrophic tropical lagoon. Marine Ecology Progress Series, 358:1–12, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hernandez-Leon:2001:DFB

[HLPHA^{+}01] S. Hernández-León, A. Portillo-Hahnefeld, C. Almeida, P. Bécognée, and I. Moreno. Diel feeding behaviour of krill in the Gerlache Strait, Antarctica. Marine Ecology Progress Series, 223:235–242, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p235.pdf.

Harder:2002:ILS

[HLQ02] T. Harder, C. Lam, and P. Y. Qian. Induction of larval settlement in the polychaete Hydroides elegans by marine biofilms: an investigation of monospecific diatom films as settlement cues. Marine Ecology Progress Series, 229:105– 112, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/229/m229p105.pdf.

Hernandez-Llamas:2004:GFC

[HLR04] Alfredo Hernandez-Llamas and David A. Ratkowsky. Growth of fishes, crustaceans and molluscs: estimation of the von Bertalanffy, Logistic, Gompertz and Richards curves and a new growth model. Marine Ecology Progress Series, 282: 237–244, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p237.pdf.

Hansen:2007:GLM

[HLR07] P. J. Hansen, N. Lundholm, and B. Rost. Growth limitation in marine red-tide dinoflagellates: effects of pH versus inorganic carbon availability. Marine Ecology Progress Series, 334:63–71, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m334p063.pdf.

Hines:2009:FIS

[HLTT09] A. H. Hines, W. C. Long, J. R. Terwin, and S. F. Thrush. Facilitation, interference, and scale: the spatial distribution of prey patches affects predation rates in an estuarine benthic community. Marine Ecology Progress Series, 385:127– 135, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/385/m385p127.pdf.

Hirst:2006:EEE

[HLU06] A. Hirst and A. López-Urrutia. Effects of evolution on egg development time. Marine Ecology Progress Series, 326: 29–35, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m326p029.pdf.

Hansson:2000:EST

[HLVU00] M. Hansson, M. Lindegarth, D. Valentinsson, and M. Ulmestrand. Effects of shrimp-trawling on abundance of benthic macrofauna in Gullmarsfjorden, Sweden. Marine Ecology Progress Series, 198:191–201, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/198/m198p191.pdf.

Hill:2005:APD

[HLY05] M. S. Hill, N. A. Lopez, and K. A. Young. Anti-predator defenses in western North Atlantic sponges with evidence of enhanced defense through interactions between spicules and chemicals. Marine Ecology Progress Series, 291:93– 102, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p093.pdf.

Hirst:2001:DEP

[HM01] A. G. Hirst and A. D. McKinnon. Does egg production represent adult female copepod growth? A call to account for body weight changes. Marine Ecology Progress Series, 223:179–199, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p179.pdf.

Harvey:2003:LFE

[HM03] M. Harvey and G. Morrier. Laboratory feeding experiments on zoea of northern shrimp Pandalus borealis fed with natural zooplankton. Marine Ecology Progress Series, 265:165– 174, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p165.pdf.

Hobbs:2004:ICC

[HM04] J. P. A. Hobbs and P. L. Munday. Intraspecific competition controls spatial distribution and social organisation of the coral-dwelling goby Gobiodon histrio. Marine Ecology Progress Series, 278:253–259, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p253.pdf.

Hernaman:2005:LHCa

[HM05a] V. Hernaman and P. L. Munday. Life-history characteristics of coral reef gobies. I. Growth and life-span. Marine Ecology Progress Series, 290:207–221, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/290/m290p207.pdf.

Hernaman:2005:LHCb

[HM05b] V. Hernaman and P. L. Munday. Life-history characteristics of coral reef gobies. II. Mortality rate, mating system and timing of maturation. Marine Ecology Progress Series, 290:223–237, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/290/m290p223.pdf.

Huang:2005:DPA

[HM05c] X. Huang and J. T. Morris. Distribution of phosphatase activity in marsh sediments along an estuarine salinity gradient. Marine Ecology Progress Series, 292:75–83, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p075.pdf.

Hedd:2006:DTP

[HM06a] A. Hedd and W. A. Montevecchi. Diet and trophic position of Leach’s storm-petrel Oceanodroma leucorhoa during breeding and moult, inferred from stable isotope analysis of feathers. Marine Ecology Progress Series, 322:291–301, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p291.pdf.

Holmes:2006:AEP

[HM06b] S. P. Holmes and N. Miller. Aspects of the ecology and population genetics of the bivalve Corbula gibba. Marine Ecology Progress Series, 315:129–140, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/315/m315p129.pdf.

Holmes:2006:LIS

[HM06c] T. H. Holmes and M. I. McCormick. Location influences sizeselective predation on newly settled reef fish. Marine Ecology Progress Series, 317:203–209, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/317/m317p203.pdf.

Holt:2009:DST

[HM09] C. A. Holt and N. Mantua. Defining spring transition: regional indices for the California Current System. Marine Ecology Progress Series, 393:285–299, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m393p285.pdf. See erratum

[Ano10a].

Hansen:2004:GNS

[HMA04] P. J. Hansen, L. Miranda, and R. Azanza. Green Noctiluca scintillans: a dinoflagellate with its own greenhouse. Marine Ecology Progress Series, 275:79–87, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p079.pdf.

Holsman:2006:IMH

[HMA06] K. K. Holsman, P. S. McDonald, and D. A. Armstrong. Intertidal migration and habitat use by subadult Dungeness crab Cancer magister in a NE Pacific estuary. Marine Ecology Progress Series, 308:183–195, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/308/m308p183.pdf.

Heithaus:2001:ECS

[HMBD01] M. R. Heithaus, G. J. Marshall, B. M. Buhleier, and L. M. Dill. Employing crittercam to study habitat use and behavior of large sharks. Marine Ecology Progress Series, 209: 307–310, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/209/m209p307.pdf.

Hjermann:2007:FOL

[HMD^{+}07] D. Ó. Hjermann, A. Melsom, G. E. Dingsør, J. M. Durant, A. M. Eikeset, L. P. Røed, G. Ottersen, G. Storvik, and N. C. Stenseth. Fish and oil in the Lofoten–Barents Sea system: synoptic review of the effect of oil spills on fish populations. Marine Ecology Progress Series, 339:283–299, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p283.pdf.

Hunt:2007:STV

[HMFC07] H. L. Hunt, M. Maltais, D. C. Fugate, and R. J. Chant. Spatial and temporal variability in juvenile bivalve dispersal: effects of sediment transport and flow regime. Marine Ecology Progress Series, 352:145–159, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/352/m352p145.pdf.

Halpin:2004:EDP

[HMH04] P. M. Halpin, B. A. Menge, and G. E. Hofmann. Experimental demonstration of plasticity in the heat shock response of the intertidal mussel Mytilus californianus. Marine Ecology Progress Series, 276:137–145, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p137.pdf.

Hedger:2008:AMW

[HMH^{+}08] R. D. Hedger, F. Martin, D. Hatin, F. Caron, F. G. Whoriskey, and J. J. Dodson. Active migration of wild Atlantic salmon Salmo salar smolt through a coastal embayment. Marine Ecology Progress Series, 355:235–246, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p235.pdf.

Hussy:2003:FDV

[HMHB03a] K. Hüssy, H. Mosegaard, H. H. Hinrichsen, and U. Böttcher. Factors determining variations in otolith microincrement width of demersal juvenile Baltic cod Gadus morhua. Marine Ecology Progress Series, 258:243–251, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p243.pdf.

Hussy:2003:UOM

[HMHB03b] K. Hüssy, H. Mosegaard, H. H. Hinrichsen, and U. Böttcher. Using otolith microstructure to analyse growth of juvenile Baltic cod Gadus morhua. Marine Ecology Progress Series, 258:233–241, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p233.pdf.

Horinouchi:2009:SHC

[HMJ^{+}09] M. Horinouchi, N. Mizuno, Y. Jo, M. Fujita, M. Sano, and Y. Suzuki. Seagrass habitat complexity does not always decrease foraging efficiencies of piscivorous fishes. Marine Ecology Progress Series, 377:43–49, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m377p043.pdf.

Hill:2006:BNO

[HMK06] J. M. Hill, C. D. McQuaid, and S. Kaehler. Biogeographic and nearshore-offshore trends in isotope ratios of intertidal mussels and their food sources around the coast of southern Africa. Marine Ecology Progress Series, 318:63–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p063.pdf.

Hansson:2005:CRJ

[HMKR05] L. J. Hansson, O. Moeslund, T. Kiørboe, and H. U. Riisgård. Clearance rates of jellyfish and their potential predation impact on zooplankton and fish larvae in a neritic ecosystem (Limfjorden, Denmark). Marine Ecology Progress Series, 304:117–131, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/304/m304p117.pdf.

Hietanen:2002:NSD

[HMKT02] S. Hietanen, P. H. Moisander, J. Kuparinen, and L. Tuominen. No sign of denitrification in a Baltic Sea cyanobacterial bloom. Marine Ecology Progress Series, 242:73–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/242/
m242p073.pdf.

Hunter:2004:VAP

[HMO^{+}04] E. Hunter, J. D. Metcalfe, C. M. O’Brien, G. P. Arnold, and J. D. Reynolds. Vertical activity patterns of free-swimming adult plaice in the southern North Sea. Marine Ecology Progress Series, 279:261–273, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/279/m279p261.pdf.

Holm:2000:VAB

[HMR00] E. R. Holm, M. McClary, Jr., and D. Rittschof. Variation in attachment of the barnacle Balanus amphitrite: sensation or something else? Marine Ecology Progress Series, 202:153–162, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p153.pdf.

Herman:2000:SIT

[HMW^{+}00] P. M. J. Herman, J. J. Middelburg, J. Widdows, C. H. Lucas, and C. H. R. Heip. Stable isotopes as trophic tracers: combining field sampling and manipulative labelling of food resources for macrobenthos. Marine Ecology Progress Series, 204:79–92, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p079.pdf.

Herman:2008:AAD

[HMY^{+}08] D. P. Herman, C. O. Matkin, G. M. Ylitalo, J. W. Durban, M. B. Hanson, M. E. Dahlheim, J. M. Straley, P. R. Wade, K. L. Tilbury, R. H. Boyer, R. W. Pearce, and M. M. Krahn. Assessing age distributions of killer whale Orcinus orca populations from the composition of endogenous fatty acids in their outer blubber layers. Marine Ecology Progress Series, 372:289–302, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/372/m372p289.pdf.

Hashimoto:2003:ENL

[HN03] T. Hashimoto and S. I. Nakano. Effect of nutrient limitation on abundance and growth of phytoplankton in a Japanese pearl farm. Marine Ecology Progress Series, 258: 43–50, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/258/m258p043.pdf.

Hussy:2009:UDS

[HNMC09] K. Hüssy, B. Nielsen, H. Mosegaard, and L. W. Clausen. Using data storage tags to link otolith macro-structure in Baltic cod Gadus morhua with environmental conditions. Marine Ecology Progress Series, 378:161–170, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/378/
m378p161.pdf.

Harii:2007:TCS

[HNYI07] S. Harii, K. Nadaoka, M. Yamamoto, and K. Iwao. Temporal changes in settlement, lipid content and lipid composition of larvae of the spawning hermatypic coral Acropora tenuis. Marine Ecology Progress Series, 346:89–96, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p089.pdf.

Holmer:2002:RDM

[HO02] M. Holmer and A. Bachmann Olsen. Role of decomposition of mangrove and seagrass detritus in sediment carbon and nitrogen cycling in a tropical mangrove forest. Marine Ecology Progress Series, 230:87–101, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/230/m230p087.pdf.

Hobday:2000:ADD

[Hob00] A. J. Hobday. Abundance and dispersal of drifting kelp Macrocystis pyrifera rafts in the Southern California Bight. Marine Ecology Progress Series, 195:101–116, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p101.pdf.

Hori:2006:ISA

[Hor06] M. Hori. Intertidal surfgrass as an allochthonous resource trap from the subtidal habitat. Marine Ecology Progress Series, 314:89–96, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p089.pdf.

Hofmann:2008:UFG

[HOT08] Gretchen E. Hofmann, Michael J. O’Donnell, and Anne E. Todgham. Using functional genomics to explore the effects of ocean acidification on calcifying marine organisms. Marine Ecology Progress Series, 373:219–225, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p219.pdf.

Hourigan:2009:MFI

[Hou09] T. F. Hourigan. Managing fishery impacts on deep-water coral ecosystems of the USA: emerging best practices. Marine Ecology Progress Series, 397:333–340, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m397p333.pdf.

Holzman:2007:ADN

[HOVG07] R. Holzman, M. Ohavia, R. Vaknin, and A. Genin. Abundance and distribution of nocturnal fishes over a coral reef during the night. Marine Ecology Progress Series, 342:205– 215, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/342/m342p205.pdf.

Hammer:2006:MAP

[HP06] A. C. Hammer and J. W. Pitchford. Mixotrophy, allelopathy and the population dynamics of phagotrophic algae (cryptophytes) in the darss zingst bodden estuary, southern Baltic. Marine Ecology Progress Series, 328:105–115, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p105.pdf.

Hofmann:2007:GER

[HP07] G. E. Hofmann and S. P. Place. Genomics-enabled research in marine ecology: challenges, risks and pay-offs. Marine Ecology Progress Series, 332:249–255, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m332p249.pdf.

Halford:2009:PRC

[HP09] A. R. Halford and J. Perret. Patterns of recovery in catastrophically disturbed reef fish assemblages. Marine Ecology Progress Series, 383:261–272, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/383/m383p261.pdf.

Hudson:2004:TVF

[HPB^{+}04] I. R. Hudson, D. W. Pond, D. S. M. Billett, P. A. Tyler, R. S. Lampitt, and G. A. Wolff. Temporal variations in fatty acid composition of deep-sea holothurians: evidence of bentho-pelagic coupling. Marine Ecology Progress Series, 281:109–120, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p109.pdf.

Howell:2003:FED

[HPBT03] K. L. Howell, D. W. Pond, D. S. M. Billett, and P. A. Tyler. Feeding ecology of deep-sea seastars (Echinodermata: Asteroidea): a fatty-acid biomarker approach. Marine Ecology Progress Series, 255:193–206, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/255/m255p193.pdf.

Huet:2004:INL

[HPC04] M. Huet, Y. M. Paulet, and J. Clavier. Imposex in Nucella lapillus: a ten year survey in NW Brittany. Marine Ecology Progress Series, 270:153–161, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/270/m270p153.pdf.

Hamer:2001:CFS

[HPH^{+}01] K. C. Hamer, R. A. Phillips, J. K. Hill, S. Wanless, and A. G. Wood. Contrasting foraging strategies of gannets Morus bassanus at two North Atlantic colonies: foraging trip duration and foraging area fidelity. Marine Ecology Progress Series, 224:283–290, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p283.pdf.

Hirst:2005:EJC

[HPR05] A. G. Hirst, W. T. Peterson, and P. Rothery. Errors in juvenile copepod growth rate estimates are widespread: problems with the Moult Rate method. Marine Ecology Progress Series, 296:263–279, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p263.pdf.

Harding:2007:SBI

[HPS07] A. M. A. Harding, J. F. Piatt, and J. A. Schmutz. Seabird behavior as an indicator of food supplies: sensitivity across the breeding season. Marine Ecology Progress Series, 352: 269–274, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m352p269.pdf.

Houdan:2005:CPR

[HPVL05] A. Houdan, I. Probert, K. Van Lenning, and S. Lefebvre. Comparison of photosynthetic responses in diploid and haploid lifecycle phases of Emiliania huxleyi (Prymnesiophyceae). Marine Ecology Progress Series, 292:139–146, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p139.pdf.

Hamer:2000:FRD

[HPW^{+}00] K. C. Hamer, R. A. Phillips, S. Wanless, M. P. Harris, and A. G. Wood. Foraging ranges, diets and feeding locations of gannets Morus bassanus in the North Sea: evidence from satellite telemetry. Marine Ecology Progress Series, 200:257–264, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p257.pdf.

Hernandez:2006:ESR

[HQDH06] K. L. Hernández, R. A. Quiñones, G. Daneri, and E. W. Helbling. Effects of solar radiation on bacterioplankton production in the upwelling system off central-southern Chile. Marine Ecology Progress Series, 315:19–31, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p019.pdf.

Hammerli:2002:LAT

[HR02a] A. Hämmerli and T. B. H. Reusch. Local adaptation and transplant dominance in genets of the marine clonal plant Zostera marina. Marine Ecology Progress Series, 242:111– 118, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p111.pdf.

Hempel:2002:RSP

[HR02b] Gotthilf Hempel and Claudio Richter. The Red Sea programme: sailing a nutshell of hope in Red Sea waters. Marine Ecology Progress Series, 239:231–232, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p231.pdf.

Hempel:2002:TSR

[HR02c] Gotthilf Hempel and Claudio Richter. Theme section: The Red Sea programme: sailing a nutshell of hope in Red Sea waters. Marine Ecology Progress Series, 239:231–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m239_TS.pdf.

Hill:2007:PBV

[HR07] R. Hill and P. J. Ralph. Post-bleaching viability of expelled zooxanthellae from the scleractinian coral Pocillopora damicornis. Marine Ecology Progress Series, 352:137–144, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/352/m352p137.pdf.

Hinton:2008:BTP

[HR08] M. Hinton and J. S. Ramsdell. Brevetoxin in two planktivorous fishes after exposure to Karenia brevis: implications for food-web transfer to bottlenose dolphins. Marine Ecology Progress Series, 356:251–258, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/356/m356p251.pdf.

Halpin:2006:OSD

[HRB^{+}06] P. N. Halpin, A. J. Read, B. D. Best, K. D. Hyrenbach, E. Fujioka, M. S. Coyne, L. B. Crowder, S. A. Freeman, and C. Spoerri. OBIS-SEAMAP: developing a biogeographic research data commons for the ecological studies of marine mammals, seabirds, and sea turtles. Marine Ecology Progress Series, 316:239–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m316p239.pdf.

Huret:2007:DMF

[HRC^{+}07] M. Huret, J. A. Runge, C. Chen, G. Cowles, Q. Xu, and J. M. Pringle. Dispersal modeling of fish early life stages: sensitivity with application to Atlantic cod in the western Gulf of Maine. Marine Ecology Progress Series, 347:261–274, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p261.pdf.

Hedd:2002:IAV

[HRCB02] A. Hedd, J. L. Ryder, L. L. Cowen, and D. F. Bertram. Inter-annual variation in the diet, provisioning and growth of Cassin’s auklet at Triangle Island, British Columbia: responses to variation in ocean climate. Marine Ecology Progress Series, 229:221–232, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/229/m229p221.pdf.

Hutson:2005:ASD

[HRDA05] K. S. Hutson, D. J. Ross, R. W. Day, and J. J. Ahern. Australian scallops do not recognise the introduced predatory seastar Asterias amurensis. Marine Ecology Progress Series, 298:305–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p305.pdf.

Hewes:2008:CPB

[HRK^{+}08] C. D. Hewes, C. S. Reiss, M. Kahru, B. G. Mitchell, and O. Holm-Hansen. Control of phytoplankton biomass by dilution and mixed layer depth in the western Weddell– Scotia Confluence. Marine Ecology Progress Series, 366: 15–29, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/366/m366p015.pdf.

Hobson:2007:VMN

[HRMH07] V. J. Hobson, D. Righton, J. D. Metcalfe, and G. C. Hays. Vertical movements of North Sea cod. Marine Ecology Progress Series, 347:101–110, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/347/m347p101.pdf.

Haberman:2002:GAK

[HRQ^{+}02] K. L. Haberman, R. M. Ross, L. B. Quetin, M. Vernet, G. A. Nevitt, and W. Kozlowski. Grazing by Antarctic krill Euphausia superba on Phaeocystis antarctica: an immunochemical approach. Marine Ecology Progress Series, 241:139–149, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p139.pdf.

Hamm:2001:TFP

[HRR^{+}01] C. Hamm, M. Reigstad, C. Wexels Riser, A. Mühlebach, and P. Wassmann. On the trophic fate of Phaeocystis pouchetii. VII. Sterols and fatty acids reveal sedimentation of P. pouchetii-derived organic matter via krill fecal strings. Marine Ecology Progress Series, 209:55–69, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
209/m209p055.pdf.

Hahn:2007:OCF

[HRR^{+}07] S. Hahn, K. Reinhardt, M. S. Ritz, T. Janicke, D. Montalti, and H. U. Peter. Oceanographic and climatic factors differentially affect reproduction performance of Antarctic skuas. Marine Ecology Progress Series, 334:287–297, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m334p287.pdf.

Hunt:2001:PWD

[HS01] H. L. Hunt and R. E. Scheibling. Predicting wave dislodgment of mussels: variation in attachment strength with body size, habitat, and season. Marine Ecology Progress Series, 213:157–164, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p157.pdf.

Huisman:2002:MSS

[HS02] J. Huisman and B. Sommeijer. Maximal sustainable sinking velocity of phytoplankton. Marine Ecology Progress Series, 244:39–48, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p039.pdf.

Hickford:2003:CDL

[HS03] M. J. H. Hickford and D. R. Schiel. Comparative dispersal of larvae from demersal versus pelagic spawning fishes. Marine Ecology Progress Series, 252:255–271, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p255.pdf.

Harris:2005:SFD

[HS05] B. Harris and S. Sulkin. Significance of feeding to the development of postlarval megalopae in the free-living crab Lophopanopeus bellus and commensal crab Fabia subquadrata. Marine Ecology Progress Series, 291:169–175, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p169.pdf.

Howe:2007:RTF

[HS07] E. R. Howe and C. A. Simenstad. Restoration trajectories and food web linkages in San Francisco Bay’s estuarine marshes: a manipulative translocation experiment. Marine Ecology Progress Series, 351:65–76, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/351/m351p065.pdf.

Hale:2008:OMM

[HS08a] R. Hale and S. E. Swearer. Otolith microstructural and microchemical changes associated with settlement in the diadromous fish Galaxias maculatus. Marine Ecology Progress Series, 354:229–234, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p229.pdf.

Heaven:2008:BCC

[HS08b] C. S. Heaven and R. A. Scrosati. Benthic community composition across gradients of intertidal elevation, wave exposure, and ice scour in Atlantic Canada. Marine Ecology Progress Series, 369:13–23, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p013.pdf.

Holbrook:2008:RRC

[HSB08] S. J. Holbrook, R. J. Schmitt, and A. J. Brooks. Resistance and resilience of a coral reef fish community to changes in coral cover. Marine Ecology Progress Series, 371:263– 271, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/371/m371p263.pdf.

Hendriks:2008:EAM

[HSBD08] I. E. Hendriks, T. Sintes, T. J. Bouma, and C. M. Duarte. Experimental assessment and modeling evaluation of the effects of the seagrass Posidonia oceanica on flow and particle trapping. Marine Ecology Progress Series, 356:163– 173, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/356/m356p163.pdf.

Herbert:2009:PBA

[HSC^{+}09] R. J. H. Herbert, A. J. Southward, R. T. Clarke, M. Sheader, and S. J. Hawkins. Persistent border: an analysis of the geographic boundary of an intertidal species. Marine Ecology Progress Series, 379:135–150, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/379/m379p135.pdf.

Holmes:2001:SWC

[HSCD01] S. P. Holmes, C. J. Sturgess, A. Cherrill, and M. S. Davies. Shell wiping in Calliostoma zizyphinum: the use of pedal mucus as a provendering agent and its contribution to daily energetic requirements. Marine Ecology Progress Series, 212: 171–181, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/212/m212p171.pdf.

Hewson:2006:TSS

[HSCF06] I. Hewson, J. A. Steele, D. G. Capone, and J. A. Fuhrman. Temporal and spatial scales of variation in bacterioplankton assemblages of oligotrophic surface waters. Marine Ecology Progress Series, 311:67–77, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/311/m311p067.pdf.

Henriksen:2007:FAS

[HSCH07] C. I. Henriksen, E. Saiz, A. Calbet, and B. W. Hansen. Feeding activity and swimming patterns of Acartia grani and Oithona davisae nauplii in the presence of motile and non-motile prey. Marine Ecology Progress Series, 331:119–129, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p119.pdf.

Hoskins:2003:UAB

[HSD03] D. L. Hoskins, S. E. Stancyk, and A. W. Decho. Utilization of algal and bacterial extracellular polymeric secretions (EPS) by the deposit-feeding brittlestar Amphipholis gracillima (Echinodermata). Marine Ecology Progress Series, 247: 93–101, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p093.pdf.

Hansen:2000:PMD

[HSGS00] P. J. Hansen, A. Skovgaard, R. N. Glud, and D. K. Stoecker. Physiology of the mixotrophic dinoflagellate Fragilidium subglobosum. II. Effects of time scale and prey concentration on photosyntheticperformance. Marine Ecology Progress Series, 201:137–146, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p137.pdf.

Henson:2006:TND

[HSHA06] S. A. Henson, R. Sanders, C. Holeton, and J. T. Allen. Timing of nutrient depletion, diatom dominance and a lowerboundary estimate of export production for Irminger Basin, North Atlantic. Marine Ecology Progress Series, 313:73–84, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/313/m313p073.pdf. See erratum [tAHSHA07].

Harvey:2008:UEL

[HSJ08] H. R. Harvey, D. H. Secor, and S. J. Ju. The use of extractable lipofuscin for age determination of crustaceans: Reply to Sheehy (2008). Marine Ecology Progress Series, 353: 307–311, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/353/m353p307.pdf. See [She08].

Hawkins:2009:CCD

[HSM^{+}09] S. J. Hawkins, H. E. Sugden, N. Mieszkowska, P. J. Moore, E. Poloczanska, R. Leaper, R. J. H. Herbert, M. J. Genner, P. S. Moschella, R. C. Thompson, S. R. Jenkins, A. J. Southward, and M. T. Burrows. Consequences of climate-driven biodiversity changes for ecosystem functioning of North European rocky shores. Marine Ecology Progress Series, 396: 245–259, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m396p245.pdf.

Hjermann:2004:ICF

[HSO04] D. Ó. Hjermann, N. C. Stenseth, and G. Ottersen. Indirect climatic forcing of the Barents Sea capelin: a cohort effect. Marine Ecology Progress Series, 273:229–238, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/273/
m273p229.pdf.

Hennige:2008:PGD

[HSP^{+}08] S. J. Hennige, D. J. Smith, R. Perkins, M. Consalvey, D. M. Paterson, and D. J. Suggett. Photoacclimation, growth and distribution of massive coral species in clear and turbid waters. Marine Ecology Progress Series, 369:77–88, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/369/
m369p077.pdf.

Heinz:2005:LBF

[HSPH05] P. Heinz, S. Sommer, O. Pfannkuche, and C. Hemleben. Living benthic Foraminifera in sediments influenced by gas hydrates at the Cascadia convergent margin, NE Pacific. Marine Ecology Progress Series, 304:77–89, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p077.pdf.

Heyward:2002:ECR

[HSRF02] A. J. Heyward, L. D. Smith, M. Rees, and S. N. Field. Enhancement of coral recruitment by in situ mass culture of coral larvae. Marine Ecology Progress Series, 230:113– 118, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/230/m230p113.pdf.

Havskum:2004:RQP

[HSS^{+}04] H. Havskum, L. Schlüter, R. Scharek, E. Berdalet, and S. Jacquet. Routine quantification of phytoplankton groups — microscopy or pigment analyses? Marine Ecology Progress Series, 273:31–42, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p031.pdf.

Heifetz:2009:DDC

[HSS09] J. Heifetz, R. P. Stone, and S. K. Shotwell. Damage and disturbance to coral and sponge habitat of the Aleutian Archipelago. Marine Ecology Progress Series, 397:295–303, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p295.pdf.

Hurst:2005:CGE

[HSSS05] T. P. Hurst, M. L. Spencer, S. M. Sogard, and A. W. Stoner. Compensatory growth, energy storage and behavior of juvenile Pacific halibut Hippoglossus stenolepis following thermally induced growth reduction. Marine Ecology Progress Series, 293:233–240, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p233.pdf.

Hinz:2001:DIP

[HSST01] S. Hinz, S. Sulkin, S. Strom, and J. Testermann. Discrimination in ingestion of protistan prey by larval crabs. Marine Ecology Progress Series, 222:155–162, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p155.pdf.

Hlista:2009:SIC

[HST^{+}09] B. L. Hlista, H. M. Sosik, L. V. Martin Traykovski, R. D. Kenney, and M. J. Moore. Seasonal and interannual correlations between right-whale distribution and calving success and chlorophyll concentrations in the Gulf of Maine, USA. Marine Ecology Progress Series, 394:289–302, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/394/
m394p289.pdf.

Hall-Spencer:2009:DMP

[HSTS^{+}09] J. M. Hall-Spencer, M. Tasker, M. Soffker, S. Christiansen, S. Rogers, M. Campbell, and K. Hoydal. Design of Marine Protected Areas on high seas and territorial waters of Rockall Bank. Marine Ecology Progress Series, 397:305–308, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p305.pdf.

Hegaret:2008:PTH

[HSW^{+}08] H. Hégaret, S. E. Shumway, G. H. Wikfors, S. Pate, and J. M. Burkholder. Potential transport of harmful algae via relocation of bivalve molluscs. Marine Ecology Progress Series, 361:169–179, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p169.pdf.

Hall-Spencer:2006:IFF

[HSWG^{+}06] J. Hall-Spencer, N. White, E. Gillespie, K. Gillham, and A. Foggo. Impact of fish farms on maerl beds in strongly tidal areas. Marine Ecology Progress Series, 326:1–9, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Han:2007:SDH

[HT07] Y. S. Han and W. N. Tzeng. Sex-dependent habitat use by the Japanese eel Anguilla japonica in Taiwan. Marine Ecology Progress Series, 338:193–198, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/338/m338p193.pdf.

Hunt:2007:OSS

[HTBT07] C. D. Hunt, D. Tanis, E. Bruce, and M. Taylor. Optical signatures of seawater and potential use for verification of midocean ballast water exchange. Marine Ecology Progress Series, 331:35–47, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/331/m331p035.pdf.

Hansen:2000:AIS

[HTJ00] J. W. Hansen, B. Thamdrup, and B. B. Jørgensen. Anoxic incubation of sediment in gas-tight plastic bags: a method for biogeochemical process studies. Marine Ecology Progress Series, 208:273–282, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/208/m208p273.pdf.

Hewitt:2002:IHA

[HTL^{+}02] J. E. Hewitt, S. F. Thrush, P. Legendre, V. J. Cummings, and A. Norkko. Integrating heterogeneity across spatial scales: interactions between Atrina zelandica and benthic macrofauna. Marine Ecology Progress Series, 239:115–128, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p115.pdf.

Hicks:2001:PDN

[HTM01] D. W. Hicks, J. W. Tunnell, Jr., and R. F. McMahon. Population dynamics of the nonindigenous brown mussel Perna perna in the Gulf of Mexico compared to other world-wide populations. Marine Ecology Progress Series, 211:181–192, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p181.pdf.

Hatase:2002:SRD

[HTM^{+}02] H. Hatase, N. Takai, Y. Matsuzawa, W. Sakamoto, K. Omuta, K. Goto, N. Arai, and T. Fujiwara. Size-related differences in feeding habitat use of adult female loggerhead turtles Caretta caretta around Japan determined by stable isotope analyses and satellite telemetry. Marine Ecology Progress Series, 233:273–281, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p273.pdf.

Houlbreque:2004:IMD

[HTRFP04] F. Houlbrèque, E. Tambutté, C. Richard, and C. Ferrier-Pagès. Importance of a micro-diet for scleractinian corals. Marine Ecology Progress Series, 282:151–160, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p151.pdf.

Hultgren:2006:GVD

[HTS06] K. M. Hultgren, P. D. Thanh, and M. Sato. Geographic variation in decoration selectivity of Micippa platipes and Tiarinia cornigera in Japan. Marine Ecology Progress Series, 326:235–244, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p235.pdf.

Hare:2005:BML

[HTW^{+}05] J. A. Hare, S. Thorrold, H. Walsh, C. Reiss, A. Valle-Levinson, and C. Jones. Biophysical mechanisms of larval fish ingress into Chesapeake Bay. Marine Ecology Progress Series, 303: 295–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/303/m303p295.pdf.

Hung:2005:EUR

[HTWQ05] O. S. Hung, V. Thiyagarajan, R. S. S. Wu, and P. Y. Qian. Effect of ultraviolet radiation on biofilms and subsequent larval settlement of Hydroides elegans. Marine Ecology Progress Series, 304:155–166, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p155.pdf.

Hung:2007:ABA

[HTZ^{+}07] O. S. Hung, V. Thiyagarajan, R. Zhang, R. S. S. Wu, and P. Y. Qian. Attachment of Balanus amphitrite larvae to biofilms originating from contrasting environments. Marine Ecology Progress Series, 333:229–242, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/333/m333p229.pdf.

Huebert:2008:BDR

[Hue08a] K. B. Huebert. Barokinesis and depth regulation by pelagic coral reef fish larvae. Marine Ecology Progress Series, 367: 261–269, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p261.pdf.

Huesemann:2008:OFO

[Hue08b] M. H. Huesemann. Ocean fertilization and other climate change mitigation strategies: an overview. Marine Ecology Progress Series, 364:243–250, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m364p243.pdf.

Humphreys:2000:OBA

[Hum00] R. L. Humphreys, Jr. Otolith-based assessment of recruitment variation in a North Pacific seamount population of armorhead Pseudopentaceros wheeleri. Marine Ecology Progress Series, 204:213–223, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p213.pdf.

Hunt:2004:EEP

[Hun04] H. L. Hunt. Effects of epibenthic predators in flow: transport and mortality of juveniles of the soft shell clam Mya arenaria. Marine Ecology Progress Series, 279:151–160, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p151.pdf.

Hunt:2005:ESS

[Hun05] H. L. Hunt. Effects of sediment source and flow regime on clam and sediment transport. Marine Ecology Progress Series, 296:143–153, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p143.pdf.

Hansen:2000:ESZ

[HUP^{+}00] J. W. Hansen, J. W. Udy, C. J. Perry, W. C. Dennison, and B. A. a. Lomstein. Effect of the seagrass Zostera capricorni on sediment microbial processes. Marine Ecology Progress Series, 199:83–96, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p083.pdf.

Hilario:2009:RAT

[HVCT09] A. Hilário, S. Vilar, M. R. Cunha, and P. Tyler. Reproductive aspects of two bythograeid crab species from hydrothermal vents in the Pacific–Antarctic Ridge. Marine Ecology Progress Series, 378:153–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/378/m378p153.pdf.

Hjorth:2007:FSR

[HVH^{+}07] M. Hjorth, J. Vester, P. Henriksen, V. Forbes, and I. Dahllöf. Functional and structural responses of marine plankton food web to pyrene contamination. Marine Ecology Progress Series, 338:21–31, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p021.pdf.

Heck:2006:ENE

[HVP^{+}06] K. L. Heck, Jr., J. F. Valentine, J. R. Pennock, G. Chaplin, and P. M. Spitze. Effects of nutrient enrichment and grazing on shoalgrass Halodule wrightii and its epiphytes: results of a field experiment. Marine Ecology Progress Series, 326:145–156, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p145.pdf.

Hinrichsen:2007:STH

[HVW^{+}07] H. H. Hinrichsen, R. Voss, K. Wieland, F. Köster, K. H. Andersen, and P. Margonski. Spatial and temporal heterogeneity of the cod spawning environment in the Bornholm Basin, Baltic Sea. Marine Ecology Progress Series, 345:245– 254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p245.pdf.

Houk:2008:DSW

[HvW08] P. Houk and R. van Woesik. Dynamics of shallow-water assemblages in the Saipan Lagoon. Marine Ecology Progress Series, 356:39–50, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p039.pdf.

Hyrenbach:2006:SAM

[HVWH06] K. D. Hyrenbach, R. R. Veit, H. Weimerskirch, and G. L. Hunt, Jr. Seabird associations with mesoscale eddies: the subtropical Indian Ocean. Marine Ecology Progress Series, 324:271–279, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/324/m324p271.pdf.

Huppop:2000:EWF

[HW00] O. Hüppop and S. Wurm. Effects of winter fishery activities on resting numbers, food and body condition of large gulls Larus argentatus and L. marinus in the south-eastern North Sea. Marine Ecology Progress Series, 194:241–247, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p241.pdf.

Hutchinson:2001:STV

[HW01] N. Hutchinson and G. A. Williams. Spatio-temporal variation in recruitment on a seasonal tropical rocky shore: the importance of local versus non-local processes. Marine Ecology Progress Series, 215:57–68, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/215/m215p057.pdf.

Hiddink:2002:CDD

[HW02] J. G. Hiddink and W. J. Wolff. Changes in distribution and decrease in numbers during migration of the bivalve Macoma balthica. Marine Ecology Progress Series, 233:117– 130, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p117.pdf.

Hutchinson:2003:DSR

[HW03] N. Hutchinson and G. A. Williams. Disturbance and subsequent recovery of mid-shore assemblages on seasonal, tropical, rocky shores. Marine Ecology Progress Series, 249: 25–38, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/249/m249p025.pdf.

Hirst:2008:SMC

[HW08] A. G. Hirst and P. Ward. Spring mortality of the cyclopoid copepod Oithona similis in polar waters. Marine Ecology Progress Series, 372:169–180, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/372/m372p169.pdf.

Hamilton:2009:OBP

[HW09] S. L. Hamilton and R. R. Warner. Otolith barium profiles verify the timing of settlement in a coral reef fish. Marine Ecology Progress Series, 385:237–244, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/385/m385p237.pdf.

Hanson:2009:IEF

[HWB^{+}09] N. N. Hanson, C. M. Wurster, M. I. Bird, K. Reid, and I. L. Boyd. Intrinsic and extrinsic forcing in life histories: patterns of growth and stable isotopes in male Antarctic fur seal teeth. Marine Ecology Progress Series, 388:263–272, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/388/
m388p263.pdf.

Hinke:2005:OHU

[HWBZ05] J. T. Hinke, G. M. Watters, G. W. Boehlert, and P. Zedonis. Ocean habitat use in autumn by Chinook salmon in coastal waters of Oregon and California. Marine Ecology Progress Series, 285:181–192, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/285/m285p181.pdf.

Hamilton:2006:CLT

[HWC^{+}06] S. L. Hamilton, J. W. White, J. E. Caselle, S. E. Swearer, and R. R. Warner. Consistent long-term spatial gradients in replenishment for an island population of a coral reef fish. Marine Ecology Progress Series, 306:247–256, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p247.pdf.

Huettel:2006:MTC

[HWG06] M. Huettel, C. Wild, and S. Gonelli. Mucus trap in coral reefs: formation and temporal evolution of particle aggregates caused by coral mucus. Marine Ecology Progress Series, 307:69–84, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p069.pdf.

Hooker:2002:FDP

[HWGB02] S. K. Hooker, H. Whitehead, S. Gowans, and R. W. Baird. Fluctuations in distribution and patterns of individual range use of northern bottlenose whales. Marine Ecology Progress Series, 225:287–297, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/225/m225p287.pdf.

Hillebrand:2000:MMC

[HWL00] H. Hillebrand, B. Worm, and H. K. Lotze. Marine microbenthic community structure regulated by nitrogen loading and grazing pressure. Marine Ecology Progress Series, 204: 27–38, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p027.pdf.

Holmes:2003:PGP

[HWvdM03] S. P. Holmes, R. Witbaard, and J. van der Meer. Phenotypic and genotypic population differentiation in the bivalve mollusc Arctica islandica: results from RAPD analysis. Marine Ecology Progress Series, 254:163–176, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p163.pdf.

Han:2003:SPS

[HYLT03] Y. S. Han, J. Y. L. Yu, I. C. Liao, and W. N. Tzeng. Salinity preference of silvering Japanese eel Anguilla japonica: evidence from pituitary prolactin mRNA levels and otolith Sr:Ca ratios. Marine Ecology Progress Series, 259: 253–261, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/259/m259p253.pdf.

Hyrenbach:2001:ARS

[Hyr01] K. D. Hyrenbach. Albatross response to survey vessels: implications for studies of the distribution, abundance, and prey consumption of seabird populations. Marine Ecology Progress Series, 212:283–295, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/212/m212p283.pdf.

Herman:2009:ADH

[HYR^{+}09] D. P. Herman, G. M. Ylitalo, J. Robbins, J. M. Straley, C. M. Gabriele, P. J. Clapham, R. H. Boyer, K. L. Tilbury, R. W. Pearce, and M. M. Krahn. Age determination of humpback whales Megaptera novaeangliae through blubber fatty acid compositions of biopsy samples. Marine Ecology Progress Series, 392:277–293, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p277.pdf.

Huntley:2004:IAT

[HZ04] M. E. Huntley and M. Zhou. Influence of animals on turbulence in the sea. Marine Ecology Progress Series, 273: 65–79, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/273/m273p065.pdf.

Ikeda:2003:SCR

[IAKS03] Y. Ikeda, N. Arai, H. Kidokoro, and W. Sakamoto. Strontium: calcium ratios in statoliths of Japanese common squid Todarodes pacificus (Cephalopoda: Ommastrephidae) as indicators of migratory behavior. Marine Ecology Progress Series, 251:169–179, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p169.pdf.

Itoh:2007:FOM

[IAT^{+}07] H. Itoh, M. N. Aoki, Y. Tsuchiya, T. Sato, H. Shinagawa, T. Komatsu, A. Mikami, and T. Hama. Fate of organic matter in faecal pellets egested by epifaunal mesograzers in a Sargassum forest and implications for biogeochemical cycling. Marine Ecology Progress Series, 352:101–112, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/352/
m352p101.pdf.

Ichii:2007:PSA

[IBB^{+}07] T. Ichii, J. L. Bengtson, P. L. Boveng, Y. Takao, J. K. Jansen, L. M. Hiruki-Raring, M. F. Cameron, H. Okamura, T. Hayashi, and M. Naganobu. Provisioning strategies of Antarctic fur seals and chinstrap penguins produce different responses to distribution of common prey and habitat. Marine Ecology Progress Series, 344:277–297, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m344p277.pdf.

Irving:2002:SLP

[IC02] A. D. Irving and S. D. Connell. Sedimentation and light penetration interact to maintain heterogeneity of subtidal habitats: algal versus invertebrate dominated assemblages. Marine Ecology Progress Series, 245:83–91, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p083.pdf.

Irving:2006:PDK

[IC06] A. D. Irving and S. D. Connell. Physical disturbance by kelp abrades erect algae from the understorey. Marine Ecology Progress Series, 324:127–137, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/324/m324p127.pdf.

Irigoien:2004:FDV

[ICH04] X. Irigoien, D. V. P. Conway, and R. P. Harris. Flexible diel vertical migration behaviour of zooplankton in the Irish Sea. Marine Ecology Progress Series, 267:85–97, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p085.pdf.

Irwin:2002:HBL

[ID02] S. Irwin and J. Davenport. Hyperoxic boundary layers inhabited by the epiphytic meiofauna of Fucus serratus. Marine Ecology Progress Series, 244:73–79, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p073.pdf.

Idjadi:2006:SCF

[IE06] J. A. Idjadi and P. J. Edmunds. Scleractinian corals as facilitators for other invertebrates on a Caribbean reef. Marine Ecology Progress Series, 319:117–127, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/319/
m319p117.pdf.

Irie:2009:ECE

[IF09] T. Irie and K. Fischer. Ectotherms with a calcareous exoskeleton follow the temperature-size rule—evidence from field survey. Marine Ecology Progress Series, 385:33–37, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/385/m385p033.pdf.

Iverson:2002:FCF

[IFL02] S. J. Iverson, K. J. Frost, and S. L. C. Lang. Fat content and fatty acid composition of forage fish and invertebrates in Prince William Sound, Alaska: factors contributing to among and within species variability. Marine Ecology Progress Series, 241:161–181, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p161.pdf.

Iriarte:2004:PSS

[IG04] J. L. Iriarte and H. E. González. Phytoplankton size structure during and after the 1997/98 El Niño in a coastal upwelling area of the northern Humboldt Current System. Marine Ecology Progress Series, 269:83–90, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p083.pdf.

Iyengar:2002:SCI

[IH02] E. V. Iyengar and C. D. Harvell. Specificity of cues inducing defensive spines in the bryozoan Membranipora membranacea. Marine Ecology Progress Series, 225:205–218, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p205.pdf.

Ishimatsu:2008:FHC

[IHK08] A. Ishimatsu, M. Hayashi, and T. Kikkawa. Fishes in high-CO_{2}, acidified oceans. Marine Ecology Progress Series, 373: 295–302, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m373p295.pdf.

Incze:2001:CCD

[IHW^{+}01] L. S. Incze, D. Hebert, N. Wolff, N. Oakey, and D. Dye. Changes in copepod distributions associated with increased turbulence from wind stress. Marine Ecology Progress Series, 213:229–240, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p229.pdf.

Ikeda:2008:MMB

[Ike08] T. Ikeda. Metabolism in mesopelagic and bathypelagic copepods: Reply to Childress et al. (2008). Marine Ecology Progress Series, 373:193–198, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/373/m373p193.pdf. See

[ISYM06, CST08].

Incera:2006:ESC

[ILL06] M. Incera, M. Lastra, and J. López. Effect of swash climate and food availability on sandy beach macrofauna along the NW coast of the Iberian Peninsula. Marine Ecology Progress Series, 314:25–33, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p025.pdf.

Iwata:2005:DPR

[IMS05] Y. Iwata, H. Munehara, and Y. Sakurai. Dependance of paternity rates on alternative reproductive behaviors in the squid Loligo bleekeri. Marine Ecology Progress Series, 298: 219–228, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p219.pdf.

Ichii:2009:LHN

[IMSO09] T. Ichii, K. Mahapatra, M. Sakai, and Y. Okada. Life history of the neon flying squid: effect of the oceanographic regime in the North Pacific Ocean. Marine Ecology Progress Series, 378:1–11, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ito:2009:SIA

[IMTW09] M. Ito, H. Minami, Y. Tanaka, and Y. Watanuki. Seasonal and inter-annual oceanographic changes induce diet switching in a piscivorous seabird. Marine Ecology Progress Series, 393:273–284, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m393p273.pdf.

Ichii:2002:OJF

[IMW^{+}02] T. Ichii, K. Mahapatra, T. Watanabe, A. Yatsu, D. Inagake, and Y. Okada. Occurrence of jumbo flying squid Dosidicus gigas aggregations associated with the countercurrent ridge off the Costa Rica dome during 1997 El Niño and 1999 La Niña. Marine Ecology Progress Series, 231:151–166, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p151.pdf.

Ichinomiya:2009:RHD

[INH^{+}09] M. Ichinomiya, M. Nakamachi, M. Honda, M. Fukuchi, and A. Taniguchi. Role of heterotrophic dinoflagellates in the fate of diatoms released from fast ice in coastal water of Lützow-Holm Bay, East Antarctica. Marine Ecology Progress Series, 383:27–36, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/383/m383p027.pdf.

Irlandi:2004:DAE

[IOB04] E. A. Irlandi, B. A. Orlando, and P. D. Biber. Drift algaeepiphyte-seagrass interactions in a subtropical Thalassia testudinum meadow. Marine Ecology Progress Series, 279:81– 91, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p081.pdf.

Ibarra-Obando:2004:BMS

[IOSPT^{+}04] S. E. Ibarra-Obando, S. V. Smith, M. Poumian-Tapia, V. Camacho-Ibar, J. D. Carriquiry, and M. Montes-Hugo. Benthic metabolism in San Quintin Bay, Baja California, Mexico. Marine Ecology Progress Series, 283:99–112, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/283/m283p099.pdf.

Incera:2009:RIS

[IOTL09] M. Incera, C. Olabarria, J. S. Troncoso, and J. López. Response of the invader Sargassum muticum to variability in nutrient supply. Marine Ecology Progress Series, 377:91– 101, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p091.pdf.

Izsak:2001:MBD

[IP01] C. Izsak and A. R. G. Price. Measuring b-diversity using a taxonomic similarity index, and its relation to spatial scale. Marine Ecology Progress Series, 215:69–77, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p069.pdf.

Iversen:2007:CCC

[IP07] M. H. Iversen and L. K. Poulsen. Coprorhexy, coprophagy, and coprochaly in the copepods Calanus helgolandicus, Pseudocalanus elongatus, and Oithona similis. Marine Ecology Progress Series, 350:79–89, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/350/m350p079.pdf.

Ingram:2002:ICA

[IR02] S. N. Ingram and E. Rogan. Identifying critical areas and habitat preferences of bottlenose dolphins Tursiops truncatus. Marine Ecology Progress Series, 244:247–255, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p247.pdf.

Ivancic:2009:APA

[IRL^{+}09] I. Ivančić, T. Radić, D. M. Lyons, D. Fuks, R. Precali, and R. Kraus. Alkaline phosphatase activity in relation to nutrient status in the northern Adriatic Sea. Marine Ecology Progress Series, 378:27–35, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p027.pdf.

Izagirre:2008:NVS

[IRM08] U. Izagirre, R. R. Ramos, and I. Marigómez. Natural variability in size and membrane stability of lysosomes in mussel digestive cells: seasonal and tidal zonation. Marine Ecology Progress Series, 372:105–117, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/372/m372p105.pdf.

Iwata:2007:TDE

[IS07] Y. Iwata and Y. Sakurai. Threshold dimorphism in ejaculate characteristics in the squid Loligo bleekeri. Marine Ecology Progress Series, 345:141–146, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/345/m345p141.pdf.

Ieno:2006:HBE

[ISBP06] E. N. Ieno, M. Solan, P. Batty, and G. J. Pierce. How biodiversity effects ecosystem functioning: roles of infaunal species richness, identity and density in the marine benthos. Marine Ecology Progress Series, 311:263–271, 2006. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m311p263.pdf.

Invidia:2004:SCA

[ISG04] M. Invidia, S. Sei, and G. Gorbi. Survival of the copepod Acartia tonsa following egg exposure to near anoxia and to sulfide at different pH values. Marine Ecology Progress Series, 276:187–196, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p187.pdf.

Iverson:2007:SIF

[ISK07] S. J. Iverson, A. M. Springer, and A. S. Kitaysky. Seabirds as indicators of food web structure and ecosystem variability: qualitative and quantitative diet analyses using fatty acids. Marine Ecology Progress Series, 352:235–244, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m352p235.pdf.

Imhoff:2003:RBP

[ISSK03] J. F. Imhoff, H. Sahling, J. Süling, and T. Kath. 16S rDNAbased phylogeny of sulphur-oxidising bacterial endosymbionts in marine bivalves from cold-seep habitats. Marine Ecology Progress Series, 249:39–51, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/249/m249p039.pdf.

Ikeda:2007:RMP

[ISY07] T. Ikeda, F. Sano, and A. Yamaguchi. Respiration in marine pelagic copepods: a global-bathymetric model. Marine Ecology Progress Series, 339:215–219, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p215.pdf.

Ikeda:2006:MMB

[ISYM06] Tsutomu Ikeda, Fumikazu Sano, Atsushi Yamaguchi, and Takashi Matsuishi. Metabolism of mesopelagic and bathypelagic copepods in the western North Pacific Ocean. Marine Ecology Progress Series, 322:199–211, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/

m322p199.pdf. See comment [CST08] and reply [Ike08].

Ichinokawa:2006:SDC

[IT06a] M. Ichinokawa and M. M. Takahashi. Size-dependent carbon flow in the epipelagic food web of the Western Equatorial Pacific. Marine Ecology Progress Series, 313:13–26, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/313/m313p013.pdf.

Islam:2006:ODS

[IT06b] M. Shahidul Islam and M. Tanaka. Ontogenetic dietary shift of Japanese sea bass during larva–juvenile transition in Ariake Bay. Marine Ecology Progress Series, 323:305– 310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p305.pdf.

Ianora:2004:CEP

[ITE^{+}04] A. Ianora, J. T. Turner, F. Esposito, Y. Carotenuto, G. d’Ippolito, G. Romano, A. Fontana, C. Guisande, and A. Miralto. Copepod egg production and hatching success is reduced by maternal diets of a non-neurotoxic strain of the dinoflagellate Alexandrium tamarense. Marine Ecology Progress Series, 280:199–210, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/280/m280p199.pdf.

Irigoien:2003:FCF

[ITH^{+}03] X. Irigoien, J. Titelman, R. P. Harris, D. Harbour, and C. Castellani. Feeding of Calanus finmarchicus nauplii in the Irminger Sea. Marine Ecology Progress Series, 262:193– 200, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p193.pdf.

Ide:2006:CFS

[ITN^{+}06] K. Ide, K. Takahashi, T. Nakano, M. Sato, and M. Omori. Chemoreceptive foraging in a shallow-water scavenging lysianassid amphipod: role of amino acids in the location of carrion in Scopelocheirus onagawae. Marine Ecology Progress Series, 317:193–202, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/317/m317p193.pdf.

Imai:2009:GCA

[ITN09] N. Imai, M. Takyu, and Y. Nakamura. Growth, crown architecture and leaf dynamics of saplings of five mangrove tree species in Ranong, Thailand. Marine Ecology Progress Series, 377:139–148, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/377/m377p139.pdf.

Iyengar:2002:SSW

[Iye02] E. V. Iyengar. Sneaky snails and wasted worms: kleptoparasitism by Trichotropis cancellata (Mollusca, Gastropoda) on Serpula columbiana (Annelida, Polychaeta). Marine Ecology Progress Series, 244:153–162, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/244/m244p153.pdf.

Jenkins:2001:ESA

[JAA^{+}01] S. R. Jenkins, F. Arenas, J. Arrontes, J. Bussell, J. Castro, R. A. Coleman, S. J. Hawkins, S. Kay, B. Martínez, J. Oliveros, M. F. Roberts, S. Sousa, R. C. Thompson, and R. G. Hartnoll. European-scale analysis of seasonal variability in limpet grazing activity and microalgal abundance. Marine Ecology Progress Series, 211:193–203, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p193.pdf.

Jenkins:2001:PDI

[JÅC^{+}01] S. R. Jenkins, P. Åberg, G. Cervin, R. A. Coleman, J. Delany, S. J. Hawkins, K. Hyder, A. A. Myers, J. Paula, A. M. Power, P. Range, and R. G. Hartnoll. Population dynamics of the intertidal barnacle Semibalanus balanoides at three European locations: spatial scales of variability. Marine Ecology Progress Series, 217:207–217, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/217/m217p207.pdf.

Jakobsen:2001:ERP

[Jak01] H. H. Jakobsen. Escape response of planktonic protists to fluid mechanical signals. Marine Ecology Progress Series, 214: 67–78, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p067.pdf.

Jarrett:2000:TVE

[Jar00] J. N. Jarrett. Temporal variation in early mortality of an intertidal barnacle. Marine Ecology Progress Series, 204:305– 308, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p305.pdf.

Jansen:2007:AVW

[JB07a] S. Jansen and U. Bathmann. Algae viability within copepod faecal pellets: evidence from microscopic examinations. Marine Ecology Progress Series, 337:145–153, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/337/
m337p145.pdf.

Johnson:2007:IGP

[JB07b] Stewart C. Johnson and Howard I. Browman. Introducing genomics, proteomics and metabolomics in marine ecology. Marine Ecology Progress Series, 332:247–310, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m332_TS.pdf.

Johnson:2007:I

[JB07c] Stewart C. Johnson and Howard I. Browman. Introduction. Marine Ecology Progress Series, 332:247–248, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m332p247.pdf.

Jones:2002:PRU

[JBA02] R. F. Jones, D. M. Baltz, and R. L. Allen. Patterns of resource use by fishes and macroinvertebrates in Barataria Bay, Louisiana. Marine Ecology Progress Series, 237:271– 289, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p271.pdf.

Jaureguizar:2003:DWC

[JBCL03] A. J. Jaureguizar, J. Bava, C. R. Carozza, and C. A. Lasta. Distribution of whitemouth croaker Micropogonias furnieri in relation to environmental factors at the Río de la plata estuary, South America. Marine Ecology Progress Series, 255:271–282, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p271.pdf.

Johansen:2008:CRF

[JBF08] J. L. Johansen, D. R. Bellwood, and C. J. Fulton. Coral reef fishes exploit flow refuges in high-flow habitats. Marine Ecology Progress Series, 360:219–226, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/360/
m360p219.pdf.

Jackson:2000:DSO

[JBG00] G. D. Jackson, N. G. Buxton, and M. J. A. George. Diet of the southern opah Lampris immaculatus on the Patagonian Shelf; the significance of the squid Moroteuthis ingens and anthropogenic plastic. Marine Ecology Progress Series, 206:261–271, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/206/m206p261.pdf.

Jenkins:2000:DSA

[JBH00] G. P. Jenkins, K. P. Black, and P. A. Hamer. Determination of spawning areas and larval advection pathways for King George whiting in southeastern Australia using otolith microstructure and hydrodynamic modelling. I. Victoria. Marine Ecology Progress Series, 199:231–242, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/199/m199p231.pdf.

Jones:2004:DTR

[JBHGB04] R. J. Jones, J. Bowyer, O. Hoegh-Guldberg, and L. L. Blackall. Dynamics of a temperature-related coral disease outbreak. Marine Ecology Progress Series, 281:63–77, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p063.pdf.

Juanes:2001:PBS

[JBS01] F. Juanes, J. A. Buckel, and F. S. Scharf. Predatory behaviour and selectivity of a primary piscivore: comparison of fish and non-fish prey. Marine Ecology Progress Series, 217:157–165, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/217/m217p157.pdf.

Jaschinski:2008:CST

[JBS08] S. Jaschinski, D. C. Brepohl, and U. Sommer. Carbon sources and trophic structure in an eelgrass Zostera marina bed, based on stable isotope and fatty acid analyses. Marine Ecology Progress Series, 358:103–114, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/358/m358p103.pdf.

Jagannathan:2009:OAW

[JBS^{+}09] S. Jagannathan, I. Bertsatos, D. Symonds, T. Chen, H. T. Nia, A. D. Jain, M. Andrews, Z. Gong, R. Nero, L. Ngor, M. Jech, O. R. Godø, S. Lee, P. Ratilal, and N. Makris. Ocean acoustic waveguide remote sensing (OAWRS) of marine ecosystems. Marine Ecology Progress Series, 395:137–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p137.pdf.

Jenkins:2001:ISD

[JBSB01] S. R. Jenkins, B. D. Beukers-Stewart, and A. R. Brand. Impact of scallop dredging on benthic megafauna: a comparison of damage levels in captured and non-captured organisms. Marine Ecology Progress Series, 215:297–301, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p297.pdf.

Jiang:2007:STR

[JBT^{+}07] M. Jiang, M. W. Brown, J. T. Turner, R. D. Kenney, C. A. Mayo, Z. Zhang, and M. Zhou. Springtime transport and retention of Calanus finmarchicus in Massachusetts and Cape Cod Bays, USA, and implications for right whale foraging. Marine Ecology Progress Series, 349:183–197, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p183.pdf.

Jensen:2009:VNE

[JBW^{+}09] F. H. Jensen, L. Bejder, M. Wahlberg, N. Aguilar Soto, M. Johnson, and P. T. Madsen. Vessel noise effects on delphinid communication. Marine Ecology Progress Series, 395: 161–175, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m395p161.pdf.

Jacinto:2008:TSI

[JC08] D. Jacinto and T. Cruz. Tidal settlement of the intertidal barnacles Chthamalus spp. in SW Portugal: interaction between diel and semi-lunar cycles. Marine Ecology Progress Series, 366:129–135, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/366/m366p129.pdf.

Jenkins:2005:RSD

[JCD^{+}05] S. R. Jenkins, R. A. Coleman, P. Della Santina, S. J. Hawkins, M. T. Burrows, and R. G. Hartnoll. Regional scale differences in the determinism of grazing effects in the rocky intertidal. Marine Ecology Progress Series, 287:77–86, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p077.pdf.

Jonzen:2002:RCE

[JCG^{+}02] N. Jonzén, M. Cardinale, A. Gårdmark, F. Arrhenius, and P. Lundberg. Risk of collapse in the eastern Baltic cod fishery. Marine Ecology Progress Series, 240:225–233, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p225.pdf.

Jones:2003:SCI

[JCK03] W. B. Jones, L. A. Cifuentes, and J. E. Kaldy. Stable carbon isotope evidence for coupling between sedimentary bacteria and seagrasses in a sub-tropical lagoon. Marine Ecology Progress Series, 255:15–25, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/255/m255p015.pdf.

Jiang:2009:TCF

[JCL09] Y. Z. Jiang, J. H. Cheng, and S. F. Li. Temporal changes in the fish community resulting from a summer fishing moratorium in the northern East China Sea. Marine Ecology Progress Series, 387:265–273, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/387/m387p265.pdf.

Jaubert:2003:REE

[JCMR^{+}03] J. M. Jaubert, J. R. M. Chisholm, A. Minghelli-Roman, M. Marchioretti, J. H. Morrow, and H. T. Ripley. Reevaluation of the extent of Caulerpa taxifolia development in the northern Mediterranean using airborne spectrographic sensing. Marine Ecology Progress Series, 263:75–82, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/263/m263p075.pdf.

Jahncke:2005:DFS

[JCZ^{+}05] J. Jahncke, K. O. Coyle, S. I. Zeeman, N. B. Kachel, and G. L. Hunt, Jr. Distribution of foraging shearwaters relative to inner front of SE Bering Sea. Marine Ecology Progress Series, 305:219–233, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/305/m305p219.pdf.

Ji:2009:LHT

[JDCB09] R. Ji, C. S. Davis, C. Chen, and R. C. Beardsley. Life history traits and spatiotemporal distributional patterns of copepod populations in the Gulf of Maine–Georges Bank region. Marine Ecology Progress Series, 384:187–205, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/384/
m384p187.pdf.

Johnson:2009:SBS

[JdSM09] M. Johnson, N. Aguilar de Soto, and P. T. Madsen. Studying the behaviour and sensory ecology of marine mammals using acoustic recording tags: a review. Marine Ecology Progress Series, 395:55–73, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m395p055.pdf.

Jacquet:2006:RPC

[JDTB06] S. Jacquet, B. Delesalle, J. P. Torréton, and J. Blanchot. Response of phytoplankton communities to increased anthropogenic influences (southwestern lagoon, New Caledonia). Marine Ecology Progress Series, 320:65–78, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p065.pdf.

Jones:2008:MPM

[JDVG08] O. A. H. Jones, F. Dondero, A. Viarengo, and J. L. Griffin. Metabolic profiling of Mytilus galloprovincialis and its potential applications for pollution assessment. Marine Ecology Progress Series, 369:169–179, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/369/m369p169.pdf.

Jeffery:2003:DAD

[Jef03] C. J. Jeffery. Determination of abundance and distribution of an intertidal barnacle: settlement or post-settlement mortality? Marine Ecology Progress Series, 246:291–305, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/246/m246p291.pdf.

Jenkins:2005:ICF

[Jen05] G. P. Jenkins. Influence of climate on the fishery recruitment of a temperate, seagrass-associated fish, the King George whiting Sillaginodes punctata. Marine Ecology Progress Series, 288:263–271, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/288/m288p263.pdf.

Johns:2003:IBD

[JERS03] D. G. Johns, M. Edwards, A. Richardson, and J. I. Spicer. Increased blooms of a dinoflagellate in the NW Atlantic. Marine Ecology Progress Series, 265:283–287, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p283.pdf.

John:2002:MCP

[JF02] E. H. John and K. J. Flynn. Modelling changes in paralytic shellfish toxin content of dinoflagellates in response to nitrogen and phosphorus supply. Marine Ecology Progress Series, 225:147–160, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p147.pdf.

Ji:2007:VMD

[JF07] R. Ji and P. J. S. Franks. Vertical migration of dinoflagellates: model analysis of strategies, growth, and vertical distribution patterns. Marine Ecology Progress Series, 344: 49–61, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/344/m344p049.pdf.

Jones:2002:EFQ

[JFA02] R. H. Jones, K. J. Flynn, and T. R. Anderson. Effect of food quality on carbon and nitrogen growth efficiency in the copepod Acartia tonsa. Marine Ecology Progress Series, 235:147–156, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p147.pdf.

Johnson:2009:LSM

[JFD09] D. S. Johnson, J. W. Fleeger, and L. A. Deegan. Largescale manipulations reveal that top-down and bottom-up controls interact to alter habitat utilization by saltmarsh fauna. Marine Ecology Progress Series, 377:33–41, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/377/
m377p033.pdf.

Josefson:2002:FPM

[JFR02] A. B. Josefson, T. L. Forbes, and R. Rosenberg. Fate of phytodetritus in marine sediments: functional importance of macrofaunal community. Marine Ecology Progress Series, 230:71–85, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p071.pdf.

Jorgensen:2005:ODB

[JGH05] B. B. Jørgensen, R. N. Glud, and O. Holby. Oxygen distribution and bioirrigation in Arctic fjord sediments (Svalbard, Barents Sea). Marine Ecology Progress Series, 292: 85–95, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p085.pdf.

Jenkins:2001:SVU

[JH01] G. P. Jenkins and P. A. Hamer. Spatial variation in the use of seagrass and unvegetated habitats by post-settlement King George whiting (Percoidei: Sillaginidae) in relation to meiofaunal distribution and macrophyte structure. Marine Ecology Progress Series, 224:219–229, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/224/m224p219.pdf.

Jung:2004:PBA

[JH04] S. Jung and E. D. Houde. Production of bay anchovy Anchoa mitchilli in Chesapeake Bay: application of size-based theory. Marine Ecology Progress Series, 281:217–232, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p217.pdf.

Johnson:2006:EHF

[JH06] M. W. Johnson and K. L. Heck, Jr. Effects of habitat fragmentation per se on decapods and fishes inhabiting seagrass meadows in the northern Gulf of Mexico. Marine Ecology Progress Series, 306:233–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/306/m306p233.pdf.

Jones:2006:EPD

[JHB06] D. O. B. Jones, I. R. Hudson, and B. J. Bett. Effects of physical disturbance on the cold-water megafaunal communities of the Faroe–Shetland Channel. Marine Ecology Progress Series, 319:43–54, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p043.pdf.

Jormalainen:2001:FPP

[JHH01] V. Jormalainen, T. Honkanen, and N. Heikkilä. Feeding preferences and performance of a marine isopod on seaweed hosts: cost of habitat specialization. Marine Ecology Progress Series, 220:219–230, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p219.pdf.

Jakobsen:2003:ITS

[JHJ^{+}03] T. S. Jakobsen, P. B. Hansen, E. Jeppesen, P. Grønkjær, and M. Søndergaard. Impact of three-spined stickleback Gasterosteus aculeatus on zooplankton and chl-a in shallow, eutrophic, brackish lakes. Marine Ecology Progress Series, 262: 277–284, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p277.pdf.

Jakobsen:2004:CET

[JHJS04] T. S. Jakobsen, P. B. Hansen, E. Jeppesen, and M. Søndergaard. Cascading effect of three-spined stickleback Gasterosteus aculeatus on community composition, size, biomass and diversity of phytoplankton in shallow, eutrophic brackish lagoons. Marine Ecology Progress Series, 279:305–309, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p305.pdf.

Jakobsen:2000:GGR

[JHL00] H. H. Jakobsen, P. J. Hansen, and J. Larsen. Growth and grazing responses of two chloroplast-retaining dinoflagellates: effect of irradiance and prey species. Marine Ecology Progress Series, 201:121–128, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/201/m201p121.pdf.

Jessop:2004:BCP

[JHL04] T. S. Jessop, M. Hamann, and C. J. Limpus. Body condition and physiological changes in male green turtles during breeding. Marine Ecology Progress Series, 276:281–288, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/276/m276p281.pdf.

Jewett:2005:EDP

[JHR05] E. B. Jewett, A. H. Hines, and G. M. Ruiz. Epifaunal disturbance by periodic low levels of dissolved oxygen: native vs. invasive species response. Marine Ecology Progress Series, 304:31–44, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p031.pdf.

Jacquet:2002:EIO

[JHTV02] S. Jacquet, H. Havskum, T. F. Thingstad, and D. Vaulot. Effects of inorganic and organic nutrient addition on a coastal microbial community (Isefjord, Denmark). Marine Ecology Progress Series, 228:3–14, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/228/m228p003.pdf.

Johnston:2003:CMR

[JK03a] E. L. Johnston and M. J. Keough. Competition modifies the response of organisms to toxic disturbance. Marine Ecology Progress Series, 251:15–26, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/251/m251p015.pdf.

Jones:2003:PPI

[JK03b] R. J. Jones and A. P. Kerswell. Phytotoxicity of photosystem II (PSII) herbicides to coral. Marine Ecology Progress Series, 261:149–159, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p149.pdf.

Jackson:2004:ZUC

[JK04] G. A. Jackson and T. Kiørboe. Zooplankton use of chemodetection to find and eat particles. Marine Ecology Progress Series, 269:153–162, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/269/m269p153.pdf.

Janowitz:2006:MKB

[JK06] G. S. Janowitz and D. Kamykowski. Modeled Karenia brevis accumulation in the vicinity of a coastal nutrient front. Marine Ecology Progress Series, 314:49–59, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/314/
m314p049.pdf.

Jorde:2007:SSG

[JKES07] P. E. Jorde, H. Knutsen, S. H. Espeland, and N. C. Stenseth. Spatial scale of genetic structuring in coastal cod Gadus morhua and geographic extent of local populations. Marine Ecology Progress Series, 343:229–237, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/343/
m343p229.pdf.

Jensen:2005:OMR

[JKG^{+}05] S. I. Jensen, M. Kühl, R. N. Glud, L. B. Jørgensen, and A. Priemé. Oxic microzones and radial oxygen loss from roots of Zostera marina. Marine Ecology Progress Series, 293: 49–58, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p049.pdf.

Jeong:2005:FGI

[JKK^{+}05] H. J. Jeong, J. S. Kim, J. H. Kim, S. T. Kim, K. A. Seong, T. H. Kim, J. Y. Song, and S. K. Kim. Feeding and grazing impact of the newly described heterotrophic dinoflagellate Stoeckeria algicida on the harmful alga Heterosigma akashiwo. Marine Ecology Progress Series, 295:69–78, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p069.pdf.

Jorgensen:2006:LMB

[JKK^{+}06] S. J. Jorgensen, D. M. Kaplan, A. P. Klimley, S. G. Morgan, M. R. O’Farrell, and L. W. Botsford. Limited movement in blue rockfish Sebastes mystinus: internal structure of home range. Marine Ecology Progress Series, 327:157– 170, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p157.pdf.

Jago:2007:SSV

[JKNJ07] C. F. Jago, G. M. Kennaway, G. Novarino, and S. E. Jones. Size and settling velocity of suspended flocs during a Phaeocystis bloom in the tidally stirred Irish Sea, NW European shelf.

Marine Ecology Progress Series, 345:51–62, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p051.pdf.

Johnston:2002:MSD

[JKQ02] E. L. Johnston, M. J. Keough, and P. Y. Qian. Maintenance of species dominance through pulse disturbances to a sessile marine invertebrate assemblage in Port Shelter, Hong Kong. Marine Ecology Progress Series, 226:103–114, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p103.pdf.

Jeong:2001:IAT

[JKS^{+}01] H. J. Jeong, H. Kang, J. H. Shim, J. K. Park, J. S. Kim, J. Y. Song, and H. J. Choi. Interactions among the toxic dinoflagellate Amphidinium carterae, the heterotrophic dinoflagellate Oxyrrhis marina, and the calanoid copepods Acartia spp. Marine Ecology Progress Series, 218:77–86, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p077.pdf.

Jeong:2007:FPC

[JKS^{+}07] H. J. Jeong, J. S. Kim, J. Y. Song, J. H. Kim, T. H. Kim, S. K. Kim, and N. S. Kang. Feeding by protists and copepods on the heterotrophic dinoflagellates Pfiesteria piscicida, Stoeckeria algicida, and Luciella masanensis. Marine Ecology Progress Series, 349:199–211, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p199.pdf.

Jonzen:2001:VFM

[JLCA01] N. Jonzén, P. Lundberg, M. Cardinale, and F. Arrhenius. Variable fishing mortality and the possible commercial extinction of the eastern Baltic cod. Marine Ecology Progress Series, 210:291–296, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p291.pdf.

Jonsson:2001:SAB

[JLJ01] L. G. Jonsson, T. Lundälv, and K. Johannesson. Symbiotic associations between anthozoans and crustaceans in a temperate coastal area. Marine Ecology Progress Series, 209: 189–195, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/209/m209p189.pdf.

Jaquemet:2004:SCS

[JLW04] S. Jaquemet, M. Le Corre, and H. Weimerskirch. Seabird community structure in a coastal tropical environment: importance of natural factors and fish aggregating devices (FADs). Marine Ecology Progress Series, 268:281–292, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p281.pdf.

Jackson:2001:TVG

[JM01] G. D. Jackson and N. A. Moltschaniwskyj. Temporal variation in growth rates and reproductive parameters in the small near-shore tropical squid Loliolus noctiluca; is cooler better? Marine Ecology Progress Series, 218:167–177, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p167.pdf.

Jompa:2003:CET

[JM03a] J. Jompa and L. J. McCook. Contrasting effects of turf algae on corals: massive Porites spp. are unaffected by mixedspecies turfs, but killed by the red alga Anotrichium tenue. Marine Ecology Progress Series, 258:79–86, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p079.pdf.

Jompa:2003:CAC

[JM03b] J. Jompa and L. J. McCook. Coral-algal competition: macroalgae with different properties have different effects on corals. Marine Ecology Progress Series, 258:87–95, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/258/m258p087.pdf.

Jessopp:2007:WRL

[JM07] M. J. Jessopp and R. J. McAllen. Water retention and limited larval dispersal: implications for short and long distance dispersers in marine reserves. Marine Ecology Progress Series, 333:27–36, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p027.pdf.

Jensen:2002:EMW

[JMA02] G. C. Jensen, P. S. McDonald, and D. A. Armstrong. East meets west: competitive interactions between green crab Carcinus maenas, and native and introduced shore crab Hemigrapsus spp. Marine Ecology Progress Series, 225:251–262, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/225/m225p251.pdf.

Jacob:2005:SIF

[JMB^{+}05] U. Jacob, K. Mintenbeck, T. Brey, R. Knust, and K. Beyer. Stable isotope food web studies: a case for standardized sample treatment. Marine Ecology Progress Series, 287:251– 253, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/287/m287p251.pdf.

Jara:2006:LEI

[JMdG^{+}06] V. C. Jara, J. H. S. Miyamoto, B. A. P. da Gama, M. Molis, M. Wahl, and R. C. Pereira. Limited evidence of interactive disturbance and nutrient effects on the diversity of macrobenthic assemblages. Marine Ecology Progress Series, 308:37–48, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/308/m308p037.pdf.

jMalta:2005:NLI

[jMFVPL05] E. j. Malta, D. G. Ferreira, J. J. Vergara, and J. L. Pérez-Lloréns. Nitrogen load and irradiance affect morphology, photosynthesis and growth of Caulerpa prolifera (Bryopsidales: Chlorophyta). Marine Ecology Progress Series, 298: 101–114, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p101.pdf.

Jones:2003:EHD

[JMHS03] R. J. Jones, J. Muller, D. Haynes, and U. Schreiber. Effects of herbicides diuron and atrazine on corals of the Great Barrier Reef, Australia. Marine Ecology Progress Series, 251: 153–167, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/251/m251p153.pdf.

Jonsen:2007:ILT

[JMJ07] I. D. Jonsen, R. A. Myers, and M. C. James. Identifying leatherback turtle foraging behaviour from satellite telemetry using a switching state-space model. Marine Ecology Progress Series, 337:255–264, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/337/m337p255.pdf.

Jessopp:2007:CCD

[JMM^{+}07] M. Jessopp, O. R. Mulholland, R. McAllen, M. P. Johnson, T. P. Crowe, and A. L. Allcock. Coastline configuration as a determinant of structure in larval assemblages. Marine Ecology Progress Series, 352:67–75, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/352/
m352p067.pdf.

Jasmine:2009:HPC

[JMM^{+}09] P. Jasmine, K. R. Muraleedharan, N. V. Madhu, C. R. Asha Devi, R. Alagarsamy, C. T. Achuthankutty, Z. Jayan, V. N. Sanjeevan, and S. Sahayak. Hydrographic and productivity characteristics along 45 ^{o}E longitude in the southwestern Indian Ocean and Southern Ocean during austral summer 2004. Marine Ecology Progress Series, 389:97–116, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p097.pdf.

Jennings:2008:BSD

[JMSM08] S. Jennings, T. A. D. Maxwell, M. Schratzberger, and S. P. Milligan. Body-size dependent temporal variations in nitrogen stable isotope ratios in food webs. Marine Ecology Progress Series, 370:199–206, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/370/m370p199.pdf.

Jackson:2009:PEG

[JMU09] A. C. Jackson, R. J. Murphy, and A. J. Underwood. Patiriella exigua: grazing by a starfish in an overgrazed intertidal system. Marine Ecology Progress Series, 376:153–163, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m376p153.pdf.

Jacobsen:2009:FSL

[JN09] H. P. Jacobsen and M. F. Norrbin. Fine-scale layer of hydromedusae is revealed by video plankton recorder (VPR) in a semi-enclosed bay in northern Norway. Marine Ecology Progress Series, 380:129–135, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/380/m380p129.pdf.

Jennings:2002:ECT

[JNDL02] S. Jennings, M. D. Nicholson, T. A. Dinmore, and J. E. Lancaster. Effects of chronic trawling disturbance on the production of infaunal communities. Marine Ecology Progress Series, 243:251–260, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/243/m243p251.pdf.

Jonsson:2004:DPM

[JNFL04] L. G. Jonsson, P. G. Nilsson, F. Floruta, and T. Lundälv. Distributional patterns of macro- and megafauna associated with a reef of the cold-water coral Lophelia pertusa on the Swedish west coast. Marine Ecology Progress Series, 284: 163–171, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p163.pdf.

Jonsson:2009:EYC

[JNH^{+}09] A. Jonsson, T. G. Nielsen, I. Hrubenja, M. Maar, and J. K. Petersen. Eating your competitor: functional triangle between turbulence, copepod escape behavior and predation from mussels. Marine Ecology Progress Series, 376:143–151, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m376p143.pdf.

Jephson:2008:TIZ

[JNMB08] T. Jephson, P. Nyström, P. O. Moksnes, and S. P. Baden. Trophic interactions in Zostera marina beds along the Swedish coast. Marine Ecology Progress Series, 369:63–76, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/369/m369p063.pdf.

Janssen:2004:IAV

[JNS04] F. Janssen, T. Neumann, and M. Schmidt. Inter-annual variability in cyanobacteria blooms in the Baltic Sea controlled by wintertime hydrographic conditions. Marine Ecology Progress Series, 275:59–68, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p059.pdf.

Johnson:2001:MDT

[Joh01] M. P. Johnson. Metapopulation dynamics of Tigriopus brevicornis (Harpacticoida) in intertidal rock pools. Marine Ecology Progress Series, 211:215–224, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/211/m211p215.pdf.

Johannessen:2002:TST

[Joh02] T. Johannessen. Is there short-term periodicity in gadoid recruitment along the Norwegian Skagerrak coast? Marine Ecology Progress Series, 241:227–229, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p227.pdf.

Johnson:2003:EOC

[Joh03] C. L. Johnson. Ecdysteroids in the oceanic copepod Calanus pacificus: variation during molt cycle and change associated with diapause. Marine Ecology Progress Series, 257: 159–165, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/257/m257p159.pdf.

Johnson:2004:DAB

[Joh04] Z. I. Johnson. Description and application of the background irradiance gradient-single turnover fluorometer (BIG-STf).

Marine Ecology Progress Series, 283:73–80, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p073.pdf.

Johnson:2007:RDC

[Joh07] C. L. Johnson. Retention of dormant copepods in deep basins of the Southern California Bight. Marine Ecology Progress Series, 336:203–210, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/336/m336p203.pdf.

Jones:2004:TPM

[Jon04] R. J. Jones. Testing the ‘photoinhibition’ model of coral bleaching using chemical inhibitors. Marine Ecology Progress Series, 284:133–145, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/284/m284p133.pdf.

Josefson:2009:APE

[Jos09] A. B. Josefson. Additive partitioning of estuarine benthic macroinvertebrate diversity across multiple spatial scales. Marine Ecology Progress Series, 396:283–292, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p283.pdf.

Jiang:2004:RBC

[JP04] H. Jiang and G. A. Paffenhöfer. Relation of behavior of copepod juveniles to potential predation by omnivorous copepods: an empirical-modeling study. Marine Ecology Progress Series, 278:225–239, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p225.pdf.

Jiang:2008:HSP

[JP08] H. Jiang and G. A. Paffenhöfer. Hydrodynamic signal perception by the copepod Oithona plumifera. Marine Ecology Progress Series, 373:37–52, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/373/m373p037.pdf.

James-Pirri:2000:ISD

[JPC00] M. J. James-Pirri and J. S. Cobb. Influence of size and delayed settlement on the recapture rate of newly settled American lobsters Homarus americanus. Marine Ecology Progress Series, 208:197–203, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p197.pdf.

Jesus:2006:EVM

[JPC^{+}06] B. Jesus, R. G. Perkins, M. Consalvey, V. Brotas, and D. M. Paterson. Effects of vertical migrations by benthic microalgae on fluorescence measurements of photophysiology. Marine Ecology Progress Series, 315:55–66, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p055.pdf.

Johnson:2005:UNM

[JPG05] D. R. Johnson, H. M. Perry, and W. M. Graham. Using nowcast model currents to explore transport of non-indigenous jellyfish into the Gulf of Mexico. Marine Ecology Progress Series, 305:139–146, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p139.pdf.

Jimenez-Perez:2004:CDC

[JPL04] L. C. Jiménez-Pérez and B. E. Lavaniegos. Changes in dominance of copepods off Baja California during the 1997–1999 El Niño and La Niña. Marine Ecology Progress Series, 277: 147–165, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/277/m277p147.pdf.

Juul-Pedersen:2008:SCS

[JPMGS08] T. Juul-Pedersen, C. Michel, M. Gosselin, and L. Seuthe. Seasonal changes in the sinking export of particulate material under first-year sea ice on the Mackenzie Shelf (western Canadian Arctic). Marine Ecology Progress Series, 353: 13–25, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/353/m353p013.pdf.

Juul-Pedersen:2006:SFS

[JPNM^{+}06] T. Juul-Pedersen, T. G. Nielsen, C. Michel, E. Friis Møller, P. Tiselius, P. Thor, M. Olesen, E. Selander, and S. Gooding. Sedimentation following the spring bloom in Disko Bay, West Greenland, with special emphasis on the role of copepods. Marine Ecology Progress Series, 314:239–255, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m314p239.pdf.

Jennings:2001:ITD

[JPPW01] S. Jennings, J. K. Pinnegar, N. V. C. Polunin, and K. J. Warr. Impacts of trawling disturbance on the trophic structure of benthic invertebrate communities. Marine Ecology Progress Series, 213:127–142, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/213/m213p127.pdf.

Jennings:2002:LSB

[JPPW02] S. Jennings, J. K. Pinnegar, N. V. C. Polunin, and K. J. Warr. Linking size-based and trophic analyses of benthic community structure. Marine Ecology Progress Series, 226: 77–85, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p077.pdf.

Jin:2004:EAA

[JQ04] T. Jin and P. Y. Qian. Effect of amino acids on larval metamorphosis of the polychaete Hydroides elegans. Marine Ecology Progress Series, 267:209–218, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p209.pdf.

Jin:2005:AAE

[JQ05] T. Jin and P. Y. Qian. Amino acid exposure modulates the bioactivity of biofilms for larval settlement of Hydroides elegans by altering bacterial community components. Marine Ecology Progress Series, 297:169–179, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p169.pdf.

Johnston:2001:SDR

[JR01] N. M. Johnston and D. A. Ritz. Synchronous development and release of broods by the swarming mysids Anisomysis mixta australis, Paramesopodopsis rufa and Tenagomysis tasmaniae (Mysidacea: Crustacea). Marine Ecology Progress Series, 223:225–233, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p225.pdf.

Jackson:2002:FCP

[JRA02] A. C. Jackson, S. D. Rundle, and M. J. Attrill. Fitness consequences of prey depletion for the common goby Pomatoschistus microps. Marine Ecology Progress Series, 242: 229–235, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p229.pdf.

Jelbart:2006:EEP

[JRC06] J. E. Jelbart, P. M. Ross, and R. M. Connolly. Edge effects and patch size in seagrass landscapes: an experimental test using fish. Marine Ecology Progress Series, 319: 93–102, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p093.pdf.

Jackson:2001:RLD

[JRP01] C. J. Jackson, P. C. Rothlisberg, and R. C. Pendrey. Role of larval distribution and abundance in overall life-history dynamics: a study of the prawn Penaeus semisulcatus in Albatross Bay, Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 213:241–252, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/213/m213p241.pdf.

Jodice:2006:IEE

[JRS^{+}06] P. G. R. Jodice, D. D. Roby, R. M. Suryan, D. B. Irons, K. R. Turco, E. D. Brown, J. F. Thedinga, and G. H. Visser. Increased energy expenditure by a seabird in response to higher food abundance. Marine Ecology Progress Series, 306:283–293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/306/m306p283.pdf.

Jodice:2006:ANS

[JRT^{+}06] P. G. R. Jodice, D. D. Roby, K. R. Turco, R. M. Suryan, D. B. Irons, J. F. Piatt, M. T. Shultz, D. G. Roseneau, A. B. Kettle, and J. A. Anthony. Assessing the nutritional stress hypothesis: relative influence of diet quantity and quality on seabird productivity. Marine Ecology Progress Series, 325:267–279, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/325/m325p267.pdf.

Johnston:2004:DEP

[JRTJ04] D. Johnston, A. Ritar, C. Thomas, and A. Jeffs. Digestive enzyme profiles of spiny lobster Jasus edwardsii phyllosoma larvae. Marine Ecology Progress Series, 275:219– 230, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/275/m275p219.pdf.

Job:2001:SSL

[JS01] S. D. Job and J. Shand. Spectral sensitivity of larval and juvenile coral reef fishes: implications for feeding in a variable light environment. Marine Ecology Progress Series, 214: 267–277, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p267.pdf.

Johansson:2002:MPR

[JS02] G. Johansson and P. Snoeijs. Macroalgal photosynthetic responses to light in relation to thallus morphology and depth zonation. Marine Ecology Progress Series, 244:63– 72, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p063.pdf.

Johnson:2003:LRP

[JS03] K. B. Johnson and A. L. Shanks. Low rates of predation on planktonic marine invertebrate larvae. Marine Ecology Progress Series, 248:125–139, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/248/m248p125.pdf.

Johnsen:2008:AMF

[JS08a] E. Johnsen and G. Skaret. Adaptive mass formations in fish shoals founded in conflicting motivation. Marine Ecology Progress Series, 358:295–300, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/358/m358p295.pdf.

Johnston:2008:CCD

[JS08b] R. Johnston and M. Sheaves. Cross-channel distribution of small fish in tropical and subtropical coastal wetlands is trophic-, taxonomic-, and wetland depth-dependent. Marine Ecology Progress Series, 357:255–270, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p255.pdf.

Ju:2001:GRV

[JSH01] S. J. Ju, D. H. Secor, and H. R. Harvey. Growth rate variability and lipofuscin accumulation rates in the blue crab Callinectes sapidus. Marine Ecology Progress Series, 224: 197–205, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/224/m224p197.pdf.

Jessop:2002:MBH

[JSIT02] B. M. Jessop, J. C. Shiao, Y. Iizuka, and W. N. Tzeng. Migratory behaviour and habitat use by American eels Anguilla rostrata as revealed by otolith microchemistry. Marine Ecology Progress Series, 233:217–229, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/233/m233p217.pdf.

Jessop:2004:VAG

[JSIT04] B. M. Jessop, J. C. Shiao, Y. Iizuka, and W. N. Tzeng. Variation in the annual growth, by sex and migration history, of silver American eels Anguilla rostrata. Marine Ecology Progress Series, 272:231–244, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/272/m272p231.pdf.

Jessop:2006:MJA

[JSIT06] B. M. Jessop, J. C. Shiao, Y. Iizuka, and W. N. Tzeng. Migration of juvenile American eels Anguilla rostrata between freshwater and estuary, as revealed by otolith microchemistry. Marine Ecology Progress Series, 310:219–233, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p219.pdf.

Jessop:2007:EIH

[JSIT07] B. M. Jessop, J. C. Shiao, Y. Iizuka, and W. N. Tzeng. Effects of inter-habitat migration on the evaluation of growth rate and habitat residence of American eels Anguilla rostrata. Marine Ecology Progress Series, 342:255–263, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p255.pdf.

Jeong:2007:FHD

[JSK^{+}07] H. J. Jeong, J. E. Song, N. S. Kang, S. Kim, Y. D. Yoo, and J. Y. Park. Feeding by heterotrophic dinoflagellates on the common marine heterotrophic nanoflagellate Cafeteria sp. Marine Ecology Progress Series, 333:151–160, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m333p151.pdf.

Jensen:2005:WDB

[JSMB05] O. P. Jensen, R. Seppelt, T. J. Miller, and L. J. Bauer. Winter distribution of blue crab Callinectes sapidus in Chesapeake Bay: application and cross-validation of a two-stage generalized additive model. Marine Ecology Progress Series, 299:239–255, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p239.pdf.

Jones:2006:RUI

[JSMM06] G. P. Jones, L. Santana, L. J. McCook, and M. I. McCormick. Resource use and impact of three herbivorous damselfishes on coral reef communities. Marine Ecology Progress Series, 328:215–224, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p215.pdf.

James:2007:PCS

[JSMM07] M. C. James, S. A. Sherrill-Mix, and R. A. Myers. Population characteristics and seasonal migrations of leatherback sea turtles at high latitudes. Marine Ecology Progress Series, 337:245–254, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p245.pdf.

Jackson:2008:PCC

[JSN08] N. L. Jackson, D. R. Smith, and K. F. Nordstrom. Physical and chemical changes in the foreshore of an estuarine beach: implications for viability and development of horseshoe crab Limulus polyphemus eggs. Marine Ecology Progress Series, 355:209–218, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/355/m355p209.pdf.

Jackson:2003:VAG

[JSWH03] G. D. Jackson, B. McGrath Steer, S. Wotherspoon, and A. J. Hobday. Variation in age, growth and maturity in the Australian arrow squid Nototodarus gouldi over time and spacewhat is the pattern? Marine Ecology Progress Series, 264: 57–71, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p057.pdf.

Jellyman:2002:FUA

[JT02] D. Jellyman and K. Tsukamoto. First use of archival transmitters to track migrating freshwater eels Anguilla dieffenbachii at sea. Marine Ecology Progress Series, 233:207– 215, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p207.pdf.

Jellyman:2005:SDO

[JT05] D. Jellyman and K. Tsukamoto. Swimming depths of offshore migrating longfin eels Anguilla dieffenbachii. Marine Ecology Progress Series, 286:261–267, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/286/m286p261.pdf.

Johnson:2008:IIS

[JT08] M. L. Johnson and G. A. Tarling. Influence of individual state on swimming capacity and behaviour of Antarctic krill Euphausia superba. Marine Ecology Progress Series, 366:99–110, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p099.pdf.

Jones:2003:BDS

[JTB^{+}03] E. G. Jones, A. Tselepides, P. M. Bagley, M. A. Collins, and I. G. Priede. Bathymetric distribution of some benthic and benthopelagic species attracted to baited cameras and traps in the deep eastern Mediterranean. Marine Ecology Progress Series, 251:75–86, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p075.pdf.

Jiang:2009:DCR

[JTLG09] X. Jiang, Y. Tang, D. J. Lonsdale, and C. J. Gobler. Deleterious consequences of a red tide dinoflagellate Cochlodinium polykrikoides for the calanoid copepod Acartia tonsa. Marine Ecology Progress Series, 390:105–116, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p105.pdf.

Johnson:2009:ISC

[JTN09] S. W. Johnson, J. F. Thedinga, and A. D. Neff. Invasion by saffron cod Eleginus gracilis into nearshore habitats of Prince William Sound, Alaska, USA. Marine Ecology Progress Series, 389:203–212, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p203.pdf.

Johnston:2005:FWB

[JTR05] D. W. Johnston, L. H. Thorne, and A. J. Read. Fin whales Balaenoptera physalus and minke whales Balaenoptera acutorostrata exploit a tidally driven island wake ecosystem in the Bay of Fundy. Marine Ecology Progress Series, 305: 287–295, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/305/m305p287.pdf.

Jacobson:2008:PAM

[JTVC08] K. C. Jacobson, D. Teel, D. M. Van Doornik, and E. Casillas. Parasite-associated mortality of juvenile Pacific salmon caused by the trematode Nanophyetus salmincola during early marine residence. Marine Ecology Progress Series, 354:235– 244, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p235.pdf.

Jonasdottir:2009:ARF

[JVJ09] S. H. Jónasdóttir, A. W. Visser, and C. Jespersen. Assessing the role of food quality in the production and hatching of Temora longicornis eggs. Marine Ecology Progress Series, 382:139–150, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/382/m382p139.pdf.

Johnston:2001:FCP

[JVT01] I. A. Johnston, V. L. A. Vieira, and G. K. Temple. Functional consequences and population differences in the developmental plasticity of muscle to temperature in Atlantic herring Clupea harengus. Marine Ecology Progress Series, 213:285–300, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
213/m213p285.pdf.

Jansen:2001:CDC

[JWG01] H. Jansen and D. A. Wolf-Gladrow. Carbonate dissolution in copepod guts: a numerical model. Marine Ecology Progress Series, 221:199–207, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/221/m221p199.pdf.

Jennings:2002:USB

[JWM02] S. Jennings, K. J. Warr, and S. Mackinson. Use of sizebased production and stable isotope analyses to predict trophic transfer efficiencies and predator–prey body mass ratios in food webs. Marine Ecology Progress Series, 240: 11–20, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/240/m240p011.pdf.

Jack:2009:PBS

[JWM09] L. Jack, S. R. Wing, and R. J. McLeod. Prey base shifts in red rock lobster Jasus edwardsii in response to habitat conversion in fiordland marine reserves: implications for effective spatial management. Marine Ecology Progress Series, 381:213–222, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p213.pdf.

Johnston:2005:EFS

[JWR05] D. W. Johnston, A. J. Westgate, and A. J. Read. Effects of fine-scale oceanographic features on the distribution and movements of harbour porpoises Phocoena phocoena in the Bay of Fundy. Marine Ecology Progress Series, 295:279– 293, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/295/m295p279.pdf.

John:2006:GDO

[JWTP06] D. E. John, B. Wawrik, F. R. Tabita, and J. H. Paul. Gene diversity and organization in rbcL-containing genome fragments from uncultivated Synechococcus in the Gulf of Mexico. Marine Ecology Progress Series, 316:23–33, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/316/
m316p023.pdf.

Johnson:2008:MPS

[JY08] S. L. Johnson and P. O. Yund. Multiple paternity and subsequent fusion-rejection interactions in a kin-structured population. Marine Ecology Progress Series, 364:129–134, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/364/m364p129.pdf.

Jeong:2006:SPJ

[JYS06] S. J. Jeong, O. H. Yu, and H. L. Suh. Secondary production of Jassa slatteryi (Amphipoda, Ischyroceridae) on a Zostera marina seagrass bed in southern Korea. Marine Ecology Progress Series, 309:205–211, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/309/m309p205.pdf.

Jackson:2009:MEC

[JZL09] G. Jackson, R. G. Zingmark, and A. J. Lewitus. Modeling epiphytic community production. Marine Ecology Progress Series, 387:61–70, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/387/m387p061.pdf.

Kaartokallio:2004:FWC

[Kaa04] H. Kaartokallio. Food web components, and physical and chemical properties of Baltic Sea ice. Marine Ecology Progress Series, 273:49–63, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p049.pdf.

Keiper:2005:MMO

[KAAH05] C. A. Keiper, D. G. Ainley, S. G. Allen, and J. T. Harvey. Marine mammal occurrence and ocean climate off central California, 1986 to 1994 and 1997 to 1999. Marine Ecology Progress Series, 289:285–306, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/289/m289p285.pdf.

Kado:2003:IJS

[Kad03] R. Kado. Invasion of Japanese shores by the NE Pacific barnacle Balanus glandula and its ecological and biogeographical impact. Marine Ecology Progress Series, 249:199– 206, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/249/m249p199.pdf.

Kamiyama:2000:AVS

[Kam00] T. Kamiyama. Application of a vital staining method to measure feeding rates of field ciliate assemblages on a harmful alga. Marine Ecology Progress Series, 197:299–303, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p299.pdf.

Kuroki:2006:CPG

[KAM^{+}06] M. Kuroki, J. Aoyama, M. J. Miller, S. Wouthuyzen, T. Arai, and K. Tsukamoto. Contrasting patterns of growth and migration of tropical anguillid leptocephali in the western Pacific and Indonesian Seas. Marine Ecology Progress Series, 309:233–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/309/m309p233.pdf.

Kappner:2000:MNF

[KAMR00] I. Kappner, S. M. Al-Moghrabi, and C. Richter. Mucusnet feeding by the vermetid gastropod Dendropoma maxima in coral reefs. Marine Ecology Progress Series, 204: 309–313, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p309.pdf.

Kaplan:2006:AAM

[Kap06] D. M. Kaplan. Alongshore advection and marine reserves: consequences for modeling and management. Marine Ecology Progress Series, 309:11–24, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/309/m309p011.pdf.

Kaplan:2009:FLH

[Kap09] D. M. Kaplan. Fish life histories and marine protected areas: an odd couple? Marine Ecology Progress Series, 377: 213–225, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m377p213.pdf.

Karako:2002:ABA

[KAPTK02] S. Karako, Y. Achituv, R. Perl-Treves, and D. Katcoff. Asterina burtoni (Asteroidea; Echinodermata) in the Mediterranean and the Red Sea: does asexual reproduction enhance colonization? Marine Ecology Progress Series, 234: 139–145, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/234/m234p139.pdf.

Kaunda-Arara:2004:HSF

[KAR04] B. Kaunda-Arara and G. A. Rose. Homing and site fidelity in the greasy grouper Epinephelus tauvina (Serranidae) within a marine protected area in coastal Kenya. Marine Ecology Progress Series, 277:245–251, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p245.pdf.

Karlson:2007:DBA

[Kar07] K. Karlson. Diurnal bioturbating activities of Monoporeia affinis: effects on benthic oxygen and nutrient fluxes. Marine Ecology Progress Series, 331:195–205, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p195.pdf.

Kasamatsu:2000:SDW

[Kas00] F. Kasamatsu. Species diversity of the whale community in the Antarctic. Marine Ecology Progress Series, 200:297–301, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p297.pdf.

Kavanagh:2005:BBG

[Kav05] K. D. Kavanagh. Boom-or-bust growth in coral reef lagoons. Marine Ecology Progress Series, 286:307–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p307.pdf.

Kawamata:2001:AMT

[Kaw01] S. Kawamata. Adaptive mechanical tolerance and dislodgement velocity of the kelp Laminaria japonica in wave-induced water motion. Marine Ecology Progress Series, 211:89–104, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/211/m211p089.pdf.

Koski:2003:IDC

[KB03] M. Koski and W. C. M. Klein Breteler. Influence of diet on copepod survival in the laboratory. Marine Ecology Progress Series, 264:73–82, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/264/m264p073.pdf.

Kvitek:2004:HAB

[KB04] R. Kvitek and C. Bretz. Harmful algal bloom toxins protect bivalve populations from sea otter predation. Marine Ecology Progress Series, 271:233–243, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/271/m271p233.pdf.

Kvitek:2005:SFB

[KB05] R. Kvitek and C. Bretz. Shorebird foraging behavior, diet, and abundance vary with harmful algal bloom toxin concentrations in invertebrate prey. Marine Ecology Progress Series, 293:303–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p303.pdf.

Killen:2006:ECR

[KB06] S. S. Killen and J. A. Brown. Energetic cost of reduced foraging under predation threat in newly hatched ocean pout. Marine Ecology Progress Series, 321:255–266, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p255.pdf.

King:2007:EPI

[KB07] A. L. King and K. Barbeau. Evidence for phytoplankton iron limitation in the southern California Current system. Marine Ecology Progress Series, 342:91–103, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p091.pdf.

Kemp:2005:ECB

[KBA^{+}05] W. M. Kemp, W. R. Boynton, J. E. Adolf, D. F. Boesch, W. C. Boicourt, G. Brush, J. C. Cornwell, T. R. Fisher, P. M. Glibert, J. D. Hagy, L. W. Harding, E. D. Houde, D. G. Kimmel, W. D. Miller, R. I. E. Newell, M. R. Roman, E. M. Smith, and J. C. Stevenson. Eutrophication of Chesapeake Bay: historical trends and ecological interactions. Marine Ecology Progress Series, 303:1–29, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kaltenberg:2009:DBS

[KBB09] A. M. Kaltenberg and K. J. Benoit-Bird. Diel behavior of sardine and anchovy schools in the California Current system. Marine Ecology Progress Series, 394:247–262, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/394/
m394p247.pdf.

Kemp:2001:NEH

[KBH01] W. M. Kemp, M. T. Brooks, and R. R. Hood. Nutrient enrichment, habitat variability and trophic transfer efficiency in simple models of pelagic ecosystems. Marine Ecology Progress Series, 223:73–87, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p073.pdf.

Kirby:2007:CEB

[KBL^{+}07] R. R. Kirby, G. Beaugrand, J. A. Lindley, A. J. Richardson, M. Edwards, and P. C. Reid. Climate effects and benthicpelagic coupling in the North Sea. Marine Ecology Progress Series, 330:31–38, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m330p031.pdf.

King:2006:DZL

[KBP06] N. J. King, P. M. Bagley, and I. G. Priede. Depth zonation and latitudinal distribution of deep-sea scavenging demersal fishes of the Mid-Atlantic Ridge, 42 to 53^{o}N. Marine Ecology Progress Series, 319:263–274, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/319/m319p263.pdf.

King:2007:I

[KBP07] N. J. King, D. M. Bailey, and I. G. Priede. Introduction. Marine Ecology Progress Series, 350:175–178, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m350p175.pdf.

King:2007:RSM

[KBPB07] N. J. King, D. M. Bailey, I. G. Priede, and Howard I. Browman. Role of scavengers in marine ecosystems. Marine Ecology Progress Series, 350:175–298, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m350_TS.pdf.

Kragefsky:2009:RSC

[KBSWG09] S. Krägefsky, U. Bathmann, V. Strass, and D. Wolf-Gladrow. Response of small copepods to an iron-induced phytoplankton bloom: a model to address the mechanisms of aggregation.

Marine Ecology Progress Series, 374:181–198, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p181.pdf.

Kiorboe:2005:BDM

[KBT05] T. Kiørboe, E. Bagøien, and U. H. Thygesen. Blind dating — mate finding in planktonic copepods. II. The pheromone cloud of Pseudocalanus elongatus. Marine Ecology Progress Series, 300:117–128, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/300/m300p117.pdf.

Kuguru:2008:MHS

[KCA^{+}08] B. Kuguru, N. E. Chadwick, Y. Achituv, K. Zandbank, and D. Tchernov. Mechanisms of habitat segregation between corallimorpharians: photosynthetic parameters and Symbiodinium types. Marine Ecology Progress Series, 369:115– 129, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/369/m369p115.pdf.

Katz:2006:SEE

[KCB06a] S. Katz, C. M. Cavanaugh, and M. Bright. Symbiosis of epi- and endocuticular bacteria with Helicoradomenia spp. (Mollusca, Aplacophora, Solenogastres) from deep-sea hydrothermal vents. Marine Ecology Progress Series, 320:89– 99, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p089.pdf.

Knights:2006:MLT

[KCB06b] A. M. Knights, T. P. Crowe, and G. Burnell. Mechanisms of larval transport: vertical distribution of bivalve larvae varies with tidal conditions. Marine Ecology Progress Series, 326:167–174, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p167.pdf.

Kubis:2009:GRJ

[KCEB09] S. Kubis, M. Chaloupka, L. Ehrhart, and M. Bresette. Growth rates of juvenile green turtles Chelonia mydas from three ecologically distinct foraging habitats along the east central coast of Florida, USA. Marine Ecology Progress Series, 389:257–269, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p257.pdf.

Krautz:2007:ITK

[KCG07] M. C. Krautz, L. R. Castro, and M. González. Interaction of two key pelagic species in the Humboldt Current: euphausiid predation on anchoveta eggs estimated by immunoassays. Marine Ecology Progress Series, 335:175–185, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p175.pdf.

Koschinski:2003:BRF

[KCH^{+}03] S. Koschinski, B. M. Culik, O. Damsgaard Henriksen, N. Tregenza, G. Ellis, C. Jansen, and G. Kathe. Behavioural reactions of free-ranging porpoises and seals to the noise of a simulated 2 MW windpower generator. Marine Ecology Progress Series, 265:263–273, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/265/m265p263.pdf.

Kaiser:2006:GAR

[KCH^{+}06] M. J. Kaiser, K. R. Clarke, H. Hinz, M. C. V. Austen, P. J. Somerfield, and I. Karakassis. Global analysis of response and recovery of benthic biota to fishing. Marine Ecology Progress Series, 311:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kawaguchi:2006:MGA

[KCK^{+}06] S. S. Kawaguchi, S. G. Candy, R. King, M. Naganobu, and S. Nicol. Modelling growth of Antarctic krill. I. Growth trends with sex, length, season, and region. Marine Ecology Progress Series, 306:1–15, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kang:2007:CPO

[KCP^{+}07] C. K. Kang, E. J. Choy, S. K. Paik, H. J. Park, K. S. Lee, and S. An. Contributions of primary organic matter sources to macroinvertebrate production in an intertidal salt marsh (Scirpus triqueter) ecosystem. Marine Ecology Progress Series, 334:131–143, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p131.pdf.

Kelaher:2004:ITG

[KCS04] B. P. Kelaher, J. C. Castilla, and R. Seed. Intercontinental test of generality for spatial patterns among diverse molluscan assemblages in coralline algal turf. Marine Ecology Progress Series, 271:221–231, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/271/m271p221.pdf.

Kruck:2009:EPS

[KCSPT09] N. C. Krück, C. A. Chargulaf, U. Saint-Paul, and I. R. Tibbetts. Early post-settlement habitat and diet shifts and the nursery function of tidepools during Sillago spp. recruitment in Moreton Bay, Australia. Marine Ecology Progress Series, 384:207–219, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/384/m384p207.pdf.

Koschinski:2006:BRF

[KCTG06] S. Koschinski, B. M. Culik, E. A. Trippel, and L. Ginzkey. Behavioral reactions of free-ranging harbor porpoises Phocoena phocoena encountering standard nylon and BaSO_{4} mesh gillnets and warning sound. Marine Ecology Progress Series, 313:285–294, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/313/m313p285.pdf.

Kakela:2006:FAS

[KCV^{+}06] A. Käkelä, J. Crane, S. C. Votier, R. W. Furness, and R. Käkelä. Fatty acid signatures as indicators of diet in great skuas Stercorarius skua, Shetland. Marine Ecology Progress Series, 319:297–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/319/m319p297.pdf.

Kaldy:2000:ABG

[KD00] J. E. Kaldy and K. H. Dunton. Above- and below-ground production, biomass and reproductive ecology of Thalassia testudinum (turtle grass) in a subtropical coastal lagoon.

Marine Ecology Progress Series, 193:271–283, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p271.pdf.

Kowalski:2001:PEH

[KDAK01] J. L. Kowalski, H. R. DeYoe, T. C. Allison, and J. E. Kaldy. Productivity estimation in Halodule wrightii Aschers: comparison of leaf-clipping and leaf-marking techniques and the importance of clip height. Marine Ecology Progress Series, 220:131–136, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p131.pdf.

Kuparinen:2009:GHP

[KdLCM09] A. Kuparinen, C. Garcia de Leaniz, S. Consuegra, and J. Merilä. Growth-history perspective on the decreasing age and size at maturation of exploited Atlantic salmon. Marine Ecology Progress Series, 376:245–252, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p245.pdf.

Kendrick:2005:CSE

[KDM05] G. A. Kendrick, C. M. Duarte, and N. Marbà. Clonality in seagrasses, emergent properties and seagrass landscapes. Marine Ecology Progress Series, 290:291–296, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p291.pdf.

Kirk:2007:MDM

[KEB07] M. Kirk, D. Esler, and W. S. Boyd. Morphology and density of mussels on natural and aquaculture structure habitats: implications for sea duck predators. Marine Ecology Progress Series, 346:179–187, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/346/m346p179.pdf.

Kaldy:2006:UDS

[KECJ06] J. E. Kaldy, P. M. Eldridge, L. A. Cifuentes, and W. B. Jones. Utilization of DOC from seagrass rhizomes by sediment bacteria: ^{13}C-tracer experiments and modeling. Marine Ecology Progress Series, 317:41–55, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m317p041.pdf.

Kelaher:2002:IPC

[Kel02] B. P. Kelaher. Influence of physical characteristics of coralline turf on associated macrofaunal assemblages. Marine Ecology Progress Series, 232:141–148, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/232/m232p141.pdf.

Kellogg:2004:TAD

[Kel04] C. A. Kellogg. Tropical archaea: diversity associated with the surface microlayer of corals. Marine Ecology Progress Series, 273:81–88, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p081.pdf.

Karez:2000:CCP

[KES00] R. Karez, S. Engelbert, and U. Sommer. ’Co-consumption’ and ‘protective coating’: two new proposed effects of epiphytes on their macroalgal hosts in mesograzer-epiphyte-host interactions. Marine Ecology Progress Series, 205:85–93, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p085.pdf.

Kellison:2003:PES

[KET^{+}03] G. T. Kellison, D. B. Eggleston, J. C. Taylor, J. S. Burke, and J. A. Osborne. Pilot evaluation of summer flounder stock enhancement potential using experimental ecology. Marine Ecology Progress Series, 250:263–278, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/250/
m250p263.pdf.

Kevekordes:2001:TTU

[Kev01] K. Kevekordes. Toxicity tests using developmental stages of Hormosira banksii (Phaeophyta) identify ammonium as a damaging component of secondary treated sewage effluent discharged into Bass Strait, Victoria, Australia. Marine Ecology Progress Series, 219:139–148, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p139.pdf.

Kamer:2001:NEA

[KF01] K. Kamer and P. Fong. Nitrogen enrichment ameliorates the negative effects of reduced salinity on the green macroalga Enteromorpha intestinalis. Marine Ecology Progress Series, 218: 87–93, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/218/m218p087.pdf.

Kaehler:2002:HMI

[KF02] S. Kaehler and P. W. Froneman. Herbivore-mediated increase in the photosynthetic capacity of marine biofilms: indirect effects of changing microalgal assemblage composition. Marine Ecology Progress Series, 234:15–22, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p015.pdf.

Kideys:2004:PCC

[KFA^{+}04] A. E. Kideys, G. A. Finenko, B. E. Anninsky, T. A. Shiganova, A. Roohi, M. R. Tabari, M. Youseffyan, M. T. Rostamian, H. Rostami, and H. Negarestan. Physiological characteristics of the ctenophore Beroe ovata in Caspian Sea water. Marine Ecology Progress Series, 266:111–121, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p111.pdf.

Kjeldstad:2003:UNO

[KFE^{+}03] B. Kjeldstad, Ó. Frette, S. R. Erga, Howard I. Browman, P. Kuhn, R. Davis, W. Miller, and J. J. Stamnes. UV (280 to 400 nm) optical properties in a Norwegian fjord system and an intercomparison of underwater radiometers. Marine Ecology Progress Series, 256:1–11, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p001.pdf.

Kicklighter:2004:CDH

[KFH04] C. E. Kicklighter, C. R. Fisher, and M. E. Hay. Chemical defense of hydrothermal vent and hydrocarbon seep organisms: a preliminary assessment using shallow-water consumers. Marine Ecology Progress Series, 275:11–19, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p011.pdf.

Kawaguchi:2007:MKG

[KFJ^{+}07] S. Kawaguchi, L. A. Finley, S. Jarman, S. G. Candy, R. M. Ross, L. B. Quetin, V. Siegel, W. Trivelpiece, M. Naganobu, and S. Nicol. Male krill grow fast and die young. Marine Ecology Progress Series, 345:199–210, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p199.pdf.

Kakela:2007:FAS

[KFK^{+}07] A. Käkelä, R. W. Furness, A. Kelly, U. Strandberg, S. Waldron, and R. Käkelä. Fatty acid signatures and stable isotopes as dietary indicators in North Sea seabirds. Marine Ecology Progress Series, 342:291–301, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p291.pdf.

Kelly:2009:DRS

[KFL09] J. E. Kelly, K. T. Frank, and W. C. Leggett. Degraded recruitment synchrony in Northwest Atlantic cod stocks. Marine Ecology Progress Series, 393:131–146, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m393p131.pdf.

Kristiansen:2001:SBN

[KFN01] S. Kristiansen, T. Farbrot, and L. J. Naustvoll. Spring bloom nutrient dynamics in the Oslofjord. Marine Ecology Progress Series, 219:41–49, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p041.pdf.

Kolasinski:2009:FEO

[KFS^{+}09] J. Kolasinski, P. Frouin, A. Sallon, K. Rogers, H. J. Bruggemann, and M. Potier. Feeding ecology and ontogenetic dietary shift of yellowstripe goatfish Mulloidichthys flavolineatus (Mullidae) at Reunion Island, SW Indian Ocean. Marine Ecology Progress Series, 386:181–195, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/386/
m386p181.pdf.

Kempf:2008:PPO

[KFT08] A. Kempf, J. Floeter, and A. Temming. Predator–prey overlap induced Holling type III functional response in the North Sea fish assemblage. Marine Ecology Progress Series, 367: 295–308, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p295.pdf.

Kristensen:2008:ECC

[KFU^{+}08] E. Kristensen, M. R. Flindt, S. Ulomi, A. V. Borges, G. Abril, and S. Bouillon. Emission of CO_{2} and CH_{4} to the atmosphere by sediments and open waters in two Tanzanian mangrove forests. Marine Ecology Progress Series, 370:53– 67, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p053.pdf.

Korb:2000:ECL

[KG00a] R. E. Korb and V. A. Gerard. Effects of concurrent low temperature and low nitrogen supply on polar and temperate seaweeds. Marine Ecology Progress Series, 198:73–82, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p073.pdf.

Korb:2000:NAC

[KG00b] R. E. Korb and V. A. Gerard. Nitrogen assimilation characteristics of polar seaweeds from differing nutrient environments. Marine Ecology Progress Series, 198:83–92, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p083.pdf.

Kurle:2007:RDF

[KG07] C. M. Kurle and C. J. Gudmundson. Regional differences in foraging of young-of-the-year Steller sea lions Eumetopias jubatus in Alaska: stable carbon and nitrogen isotope ratios in blood. Marine Ecology Progress Series, 342:303– 310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/342/m342p303.pdf.

Kuhl:2001:PPS

[KGB^{+}01] M. Kühl, R. N. Glud, J. Borum, R. Roberts, and S. Rysgaard. Photosynthetic performance of surface-associated algae below sea ice as measured with a pulse-amplitude-modulated (PAM) fluorometer and O_{2} microsensors. Marine Ecology Progress Series, 223:1–14, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p001.pdf.

Kubetzki:2009:IMS

[KGF^{+}09] U. Kubetzki, S. Garthe, D. Fifield, B. Mendel, and R. W. Furness. Individual migratory schedules and wintering areas of northern gannets. Marine Ecology Progress Series, 391:257–265, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m391p257.pdf.

Koslow:2001:SBM

[KGHL^{+}01] J. A. Koslow, K. Gowlett-Holmes, J. K. Lowry, T. O’Hara, G. C. B. Poore, and A. Williams. Seamount benthic macrofauna off southern Tasmania: community structure and impacts of trawling. Marine Ecology Progress Series, 213:111– 125, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/213/m213p111.pdf.

Kelly-Gerreyn:2005:CDB

[KGHW05] B. A. Kelly-Gerreyn, D. J. Hydes, and J. J. Waniek. Control of the diffusive boundary layer on benthic fluxes: a model study. Marine Ecology Progress Series, 292:61–74, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p061.pdf.

Kroger:2006:RSS

[KGRW06] K. Kröger, J. P. A. Gardner, A. A. Rowden, and R. G. Wear. Recovery of a subtidal soft-sediment macroinvertebrate assemblage following experimentally induced effects of a harmful algal bloom. Marine Ecology Progress Series, 326:

85–98, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/326/m326p085.pdf.

Kvitek:2008:DAC

[KGS^{+}08] R. G. Kvitek, J. D. Goldberg, G. J. Smith, G. J. Doucette, and M. W. Silver. Domoic acid contamination within eight representative species from the benthic food web of Monterey Bay, California, USA. Marine Ecology Progress Series, 367:35–47, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/367/m367p035.pdf.

Kelly-Gerreyn:2001:MDD

[KGTH01] B. A. Kelly-Gerreyn, M. Trimmer, and D. J. Hydes. A model discriminating denitrification and dissimilatory nitrate reduction to ammonium in a temperate estuarine sediment. Marine Ecology Progress Series, 220:33–46, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/220/m220p033.pdf.

Karakassis:2000:BDD

[KH00a] I. Karakassis and E. Hatziyanni. Benthic disturbance due to fish farming analyzed under different levels of taxonomic resolution. Marine Ecology Progress Series, 203:247–253, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p247.pdf.

Kim:2000:SRB

[KH00b] Y. O. Kim and M. S. Han. Seasonal relationships between cyst germination and vegetative population of Scrippsiella trochoidea (Dinophyceae). Marine Ecology Progress Series, 204:111–118, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p111.pdf.

Knowlton:2000:CTS

[KH00c] Ann L. Knowlton and Raymond C. Highsmith. Convergence in the time-space continuum: a predator–prey interaction. Marine Ecology Progress Series, 197:285–291, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p285.pdf.

Kuhlmann:2005:DDP

[KH05] M. L. Kuhlmann and A. H. Hines. Density-dependent predation by blue crabs Callinectes sapidus on natural prey populations of infaunal bivalves. Marine Ecology Progress Series, 295:215–228, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p215.pdf.

Kaiser:2007:FSF

[KH07a] M. J. Kaiser and J. G. Hiddink. Food subsidies from fisheries to continental shelf benthic scavengers. Marine Ecology Progress Series, 350:267–276, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m350p267.pdf.

Kramer:2007:TTS

[KH07b] K. Linday Kramer and Kenneth L. Heck, Jr. Top-down trophic shifts in Florida Keys patch reef marine protected areas. Marine Ecology Progress Series, 349:111–123, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/

m349p111.pdf. See erratum [KH07c].

Kramer:2007:EKA

[KH07c] K. Lindsey Kramer and Kenneth L. Heck, Jr. Erratum: Kramer et al., 349:111–123 (2007). Marine Ecology Progress Series, 352:310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/352/m352p310.pdf. See [KH07b].

Kimmel:2008:URB

[KH08a] D. G. Kimmel and S. Hameed. Update on the relationship between the North Atlantic Oscillation and Calanus finmarchicus. Marine Ecology Progress Series, 366:111–117, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/366/m366p111.pdf.

Kiorboe:2008:ODT

[KH08b] T. Kiørboe and A. G. Hirst. Optimal development time in pelagic copepods. Marine Ecology Progress Series, 367: 15–22, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p015.pdf.

Kooyman:2000:MEP

[KHA^{+}00] G. L. Kooyman, E. C. Hunke, S. F. Ackley, R. P. van Dam, and G. Robertson. Moult of the emperor penguin: travel, location, and habitat selection. Marine Ecology Progress Series, 204:269–277, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p269.pdf.

Keister:2000:EBL

[KHB00a] J. E. Keister, E. D. Houde, and D. L. Breitburg. Effects of bottom-layer hypoxia on abundances and depth distributions of organisms in Patuxent River, Chesapeake Bay. Marine Ecology Progress Series, 205:43–59, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/205/m205p043.pdf.

Klap:2000:RLS

[KHB00b] V. A. Klap, M. A. Hemminga, and J. J. Boon. Retention of lignin in seagrasses: angiosperms that returned to the sea. Marine Ecology Progress Series, 194:1–11, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p001.pdf.

Karle:2007:BRI

[KHD07] I. M. Karle, P. O. J. Hall, and I. Dahllöf. Biogeochemical response of an intact coastal sediment to organic matter input: a multivariate approach. Marine Ecology Progress Series, 342:15–25, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/342/m342p015.pdf.

Kitaysky:2000:RAB

[KHF^{+}00] A. S. Kitaysky, G. L. Hunt, Jr., E. N. Flint, M. A. Rubega, and M. B. Decker. Resource allocation in breeding seabirds: responses to fluctuations in their food supply. Marine Ecology Progress Series, 206:283–296, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p283.pdf.

Kent:2006:PEG

[KHG06] R. Kent, R. Holzman, and A. Genin. Preliminary evidence on group-size dependent feeding success in the damselfish Dascyllus marginatus. Marine Ecology Progress Series, 323: 299–303, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p299.pdf.

Katz:2002:GMA

[KHGA02] T. Katz, B. Herut, A. Genin, and D. L. Angel. Gray mullets ameliorate organically enriched sediments below a fish farm in the oligotrophic Gulf of Aqaba (Red Sea). Marine Ecology Progress Series, 234:205–214, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/234/m234p205.pdf.

Kirkwood:2008:STV

[KHH08] R. Kirkwood, F. Hume, and M. Hindell. Sea temperature variations mediate annual changes in the diet of Australian fur seals in Bass Strait. Marine Ecology Progress Series, 369:297–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p297.pdf.

Kimmerer:2007:EJC

[KHHM07] W. J. Kimmerer, A. G. Hirst, R. R. Hopcroft, and A. D. McKinnon. Estimating juvenile copepod growth rates: corrections, inter-comparisons and recommendations. Marine Ecology Progress Series, 336:187–202, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/336/m336p187.pdf.

Kayanne:2002:RCP

[KHIA02] H. Kayanne, S. Harii, Y. Ide, and F. Akimoto. Recovery of coral populations after the 1998 bleaching on Shiraho Reef, in the southern Ryukyus, NW Pacific. Marine Ecology Progress Series, 239:93–103, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p093.pdf.

Karnovsky:2008:SCD

[KHIH08] N. J. Karnovsky, K. A. Hobson, S. Iverson, and G. L. Hunt, Jr. Seasonal changes in diets of seabirds in the North Water Polynya: a multiple-indicator approach. Marine Ecology Progress Series, 357:291–299, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/357/m357p291.pdf.

Kojima:2004:PID

[KHK^{+}04] S. Kojima, I. Hayashi, D. Kim, A. Iijima, and T. Furota. Phylogeography of an intertidal direct-developing gastropod Batillaria cumingi around the Japanese Islands. Marine Ecology Progress Series, 276:161–172, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p161.pdf.

Kodama:2006:EHE

[KHK^{+}06] K. Kodama, T. Horiguchi, G. Kume, S. Nagayama, T. Shimizu, H. Shiraishi, M. Morita, and M. Shimizu. Effects of hypoxia on early life history of the stomatopod Oratosquilla oratoria in a coastal sea. Marine Ecology Progress Series, 324:197–206, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/324/m324p197.pdf.

Kingsford:2009:OCN

[KHP09] M. J. Kingsford, J. M. Hughes, and H. M. Patterson. Otolith chemistry of the non-dispersing reef fish Acanthochromis polyacanthus: cross-shelf patterns from the central Great Barrier Reef. Marine Ecology Progress Series, 377:279–288, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/377/m377p279.pdf.

Kamermans:2002:GED

[KHT^{+}02] P. Kamermans, M. A. Hemminga, J. F. Tack, M. A. Mateo, N. Marbà, M. Mtolera, J. Stapel, A. Verheyden, and T. Van Daele. Groundwater effects on diversity and abundance of lagoonal seagrasses in Kenya and on Zanzibar Island (East Africa). Marine Ecology Progress Series, 231:

75–83, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p075.pdf.

Kloppmann:2002:WME

[KHvW02] M. H. F. Kloppmann, N. Hillgruber, and H. von Westernhagen. Wind-mixing effects on feeding success and condition of blue whiting larvae in the Porcupine Bank area. Marine Ecology Progress Series, 235:263–277, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p263.pdf.

Kirchman:2003:RGE

[KHWH03] D. L. Kirchman, K. A. Hoffman, R. Weaver, and D. A. Hutchins. Regulation of growth and energetics of a marine bacterium by nitrogen source and iron availability. Marine Ecology Progress Series, 250:291–296, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/250/
m250p291.pdf.

Kobari:2001:LCN

[KI01] T. Kobari and T. Ikeda. Life cycle of Neocalanus flemingeri (Crustacea: Copepoda) in the Oyashio region, western subarctic Pacific, with notes on its regional variations. Marine Ecology Progress Series, 209:243–255, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/209/m209p243.pdf.

Kimmerer:2002:EFF

[Kim02] W. J. Kimmerer. Effects of freshwater flow on abundance of estuarine organisms: physical effects or trophic linkages? Marine Ecology Progress Series, 243:39–55, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p039.pdf.

Kimmerer:2006:RAD

[Kim06] W. J. Kimmerer. Response of anchovies dampens effects of the invasive bivalve Corbula amurensis on the San Francisco Estuary foodweb. Marine Ecology Progress Series, 324:207–218, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/324/m324p207.pdf. See erratum

[Ano07d].

King:2001:ACM

[Kin01] G. M. King. Aspects of carbon monoxide production and oxidation by marine macroalgae. Marine Ecology Progress Series, 224:69–75, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p069.pdf.

Kinne:2002:IOD

[Kin02] O. Kinne. Importance and organisation of direct postpublication critique MEPS 228:1. Marine Ecology Progress Series, ??:??, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/misc/m228p001.pdf.

Kinne:2003:SPN

[Kin03] O. Kinne. The scientific process: new forces attempt to enter the scene MEPS 254:1. Marine Ecology Progress Series, ??:??, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/misc/m254p001.pdf.

Kinne:2004:QAI

[Kin04] O. Kinne. Quality assessment and improvement of manuscripts MEPS 279:1. Marine Ecology Progress Series, ??:??, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/misc/
m279p001.pdf.

Kinne:2005:YIR

[Kin05a] O. Kinne. 25 years inter-research 1979–2004. Marine Ecology Progress Series, 285:1–2, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kinne:2005:VME

[Kin05b] O. Kinne. 300 volumes of marine ecology progress series MEPS 300:1. Marine Ecology Progress Series, ??:??, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/misc/
m300p001.pdf.

Kiorboe:2003:HTR

[Kiø03] T. Kiørboe. High turnover rates of copepod fecal pellets due to Noctiluca scintillans grazing. Marine Ecology Progress Series, 258:181–188, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p181.pdf.

Kirchman:2002:CMG

[Kir02a] D. L. Kirchman. Calculating microbial growth rates from data on production and standing stocks. Marine Ecology Progress Series, 233:303–306, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/233/m233p303.pdf.

Kirchman:2002:RCO

[Kir02b] D. L. Kirchman. Reply comment: Overcoming linearisation errors in calculating bacterial growth rates. Marine Ecology Progress Series, 245:309–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p309.pdf. See

[Pet02a, Kir02c].

Kirchman:2002:RPO

[Kir02c] David L. Kirchman. Reply to Peters (2002) Overcoming linearisation errors in calculating bacterial growth rates. Marine Ecology Progress Series, 245:309–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/

m245p309.pdf. See [Pet02a, Kir02b].

Kerswell:2003:EHO

[KJ03] A. P. Kerswell and R. J. Jones. Effects of hypo-osmosis on the coral Stylophora pistillata: nature and cause of ’lowsalinity bleaching’. Marine Ecology Progress Series, 253:145– 154, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p145.pdf.

Kim:2004:FHD

[KJ04] J. S. Kim and H. J. Jeong. Feeding by the heterotrophic dinoflagellates Gyrodinium dominans and G. spirale on the redtide dinoflagellate Prorocentrum minimum. Marine Ecology Progress Series, 280:85–94, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/280/m280p085.pdf.

Kemp:2006:CLB

[KJB^{+}06] K. M. Kemp, A. J. Jamieson, P. M. Bagley, H. McGrath, D. M. Bailey, M. A. Collins, and I. G. Priede. Consumption of large bathyal food fall, a six month study in the NE Atlantic. Marine Ecology Progress Series, 310:65–76, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p065.pdf.

Karlsson:2003:DLS

[KJL03] Ö. Karlsson, P. R. Jonsson, and A. I. Larsson. Do large seston particles contribute to the diet of the bivalve Cerastoderma edule? Marine Ecology Progress Series, 261:161– 173, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/261/m261p161.pdf.

Keister:2005:BIT

[KJMP05] J. E. Keister, T. B. Johnson, C. A. Morgan, and W. T. Peterson. Biological indicators of the timing and direction of warm-water advection during the 1997/1998 El Niño off the central Oregon coast, USA. Marine Ecology Progress Series, 295:43–48, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p043.pdf.

Kuhn:2009:ATM

[KJRG09] C. E. Kuhn, D. S. Johnson, R. R. Ream, and T. S. Gelatt. Advances in the tracking of marine species: using GPS locations to evaluate satellite track data and a continuous-time movement model. Marine Ecology Progress Series, 393:97– 109, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/393/m393p097.pdf.

Khalaf:2002:CTC

[KK02] M. A. Khalaf and M. Kochzius. Changes in trophic community structure of shore fishes at an industrial site in the Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 239:287–299, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p287.pdf. See erratum

[Ano02e].

Koh:2006:TRM

[KKA^{+}06] C. H. Koh, J. S. Khim, H. Araki, H. Yamanishi, H. Mogi, and K. Koga. Tidal resuspension of microphytobenthic chlorophyll-a in a Nanaura mudflat, Saga, Ariake Sea, Japan: flood-ebb and spring-neap variations. Marine Ecology Progress Series, 312:85–100, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/312/m312p085.pdf.

Kline:2006:REO

[KKB^{+}06] D. I. Kline, N. M. Kuntz, M. Breitbart, N. Knowlton, and F. Rohwer. Role of elevated organic carbon levels and microbial activity in coral mortality. Marine Ecology Progress Series, 314:119–125, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p119.pdf.

Kicklighter:2003:PDS

[KKBH03] C. E. Kicklighter, J. Kubanek, T. Barsby, and M. E. Hay. Palatability and defense of some tropical infaunal worms: alkylpyrrole sulfamates as deterrents to fish feeding. Marine Ecology Progress Series, 263:299–306, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p299.pdf.

Karplus:2006:PRS

[KKG06] I. Karplus, R. Katzenstein, and M. Goren. Predator recognition and social facilitation of predator avoidance in coral reef fish Dascyllus marginatus juveniles. Marine Ecology Progress Series, 319:215–223, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/319/m319p215.pdf.

Kon:2007:RMF

[KKH07] K. Kon, H. Kurokura, and K. Hayashizaki. Role of microhabitats in food webs of benthic communities in a mangrove forest. Marine Ecology Progress Series, 340:55–62, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p055.pdf.

Kuwae:2006:OEF

[KKI^{+}06] T. Kuwae, K. Kamio, T. Inoue, E. Miyoshi, and Y. Uchiyama. Oxygen exchange flux between sediment and water in an intertidal sandflat, measured in situ by the eddy-correlation method. Marine Ecology Progress Series, 307:59–68, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m307p059.pdf.

Kharlamenko:2001:IFS

[KKIV01] V. I. Kharlamenko, S. I. Kiyashko, A. B. Imbs, and D. I. Vyshkvartzev. Identification of food sources of invertebrates from the seagrass Zostera marina community using carbon and sulfur stable isotope ratio and fatty acid analyses. Marine Ecology Progress Series, 220:103–117, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p103.pdf.

Kakela:2005:FAS

[KKK^{+}05] R. Käkelä, A. Käkelä, S. Kahle, P. H. Becker, A. Kelly, and R. W. Furness. Fatty acid signatures in plasma of captive herring gulls as indicators of demersal or pelagic fish diet. Marine Ecology Progress Series, 293:191–200, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/293/
m293p191.pdf.

Kang:2003:PCD

[KKKK03] J. S. Kang, S. H. Kang, D. Kim, and D. Y. Kim. Planktonic centric diatom Minidiscus chilensis dominated sediment trap material in eastern Bransfield Strait, Antarctica. Marine Ecology Progress Series, 255:93–99, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p093.pdf.

Kang:2003:TIB

[KKL^{+}03] C. K. Kang, J. B. Kim, K. S. Lee, J. B. Kim, P. Y. Lee, and J. S. Hong. Trophic importance of benthic microalgae to macrozoobenthos in coastal bay systems in Korea: dual stable C and N isotope analyses. Marine Ecology Progress Series, 259:79–92, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p079.pdf.

Krafft:2007:DSA

[KKL07] B. A. Krafft, K. M. Kovacs, and C. Lydersen. Distribution of sex and age groups of ringed seals Pusa hispida in the fast-ice breeding habitat of Kongsfjorden, Svalbard. Marine Ecology Progress Series, 335:199–206, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/335/m335p199.pdf.

Kang:2002:SVM

[KKLL02] J. S. Kang, S. H. Kang, J. H. Lee, and S. H. Lee. Seasonal variation of microalgal assemblages at a fixed station in King George Island, Antarctica, 1996. Marine Ecology Progress Series, 229:19–32, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p019.pdf.

Kim:2008:NCD

[KKMH08] J. H. Kim, S. J. Kim, G. S. Min, and K. N. Han. Nutritional condition determined using RNA/DNA ratios of the river pufferfish Takifugu obscurus under different salinities. Marine Ecology Progress Series, 372:243–252, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p243.pdf.

Koster:2008:TSM

[KKP08] M. Köster, C. Krause, and G. A. Paffenhöfer. Time-series measurements of oxygen consumption of copepod nauplii. Marine Ecology Progress Series, 353:157–164, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p157.pdf.

Kuntz:2005:PMR

[KKSR05] N. M. Kuntz, D. I. Kline, S. A. Sandin, and F. Rohwer. Pathologies and mortality rates caused by organic carbon and nutrient stressors in three Caribbean coral species. Marine Ecology Progress Series, 294:173–180, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p173.pdf.

Katano:2005:SCA

[KKTN05] T. Katano, A. Kaneda, H. Takeoka, and S. Nakano. Seasonal changes in the abundance and composition of picophytoplankton in relation to the occurrence of ‘Kyucho ’ and bottom intrusion in Uchiumi Bay, Japan. Marine Ecology Progress Series, 298:59–67, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p059.pdf.

Karnovsky:2003:FBL

[KKW^{+}03] N. J. Karnovsky, S. Kwaśniewski, J. M. Wesławski, W. Walkusz, and A. Beszczyńska-Möller. Foraging behavior of little auks in a heterogeneous environment. Marine Ecology Progress Series, 253:289–303, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p289.pdf.

Kelaher:2003:VDE

[KL03] B. P. Kelaher and J. S. Levinton. Variation in detrital enrichment causes spatio-temporal variation in soft-sediment assemblages. Marine Ecology Progress Series, 261:85–97, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/261/m261p085.pdf.

Kang:2006:MSD

[KLC^{+}06] C. K. Kang, Y. W. Lee, E. J. Choy, J. K. Shin, I. S. Seo, and J. S. Hong. Microphytobenthos seasonality determines growth and reproduction in intertidal bivalves. Marine Ecology Progress Series, 315:113–127, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p113.pdf.

Kubanek:2004:ARP

[KLFH04] J. Kubanek, S. E. Lester, W. Fenical, and M. E. Hay. Ambiguous role of phlorotannins as chemical defenses in the brown alga Fucus vesiculosus. Marine Ecology Progress Series, 277:79–93, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/277/m277p079.pdf.

Kaartvedt:2002:OKM

[KLHO02] S. Kaartvedt, T. Larsen, K. Hjelmseth, and M. S. R. Onsrud. Is the omnivorous krill Meganyctiphanes norvegica primarily a selectively feeding carnivore? Marine Ecology Progress Series, 228:193–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p193.pdf.

Kanouse:2006:NUR

[KLN06] S. Kanouse, M. K. La Peyre, and J. A. Nyman. Nekton use of Ruppia maritima and non-vegetated bottom habitat types within brackish marsh ponds. Marine Ecology Progress Series, 327:61–69, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/327/m327p061.pdf.

Keppler:2006:SCE

[KLR^{+}06] C. J. Keppler, A. J. Lewitus, A. H. Ringwood, J. Hoguet, and T. Staton. Sublethal cellular effects of short-term raphidophyte and brevetoxin exposures on the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 312:141– 147, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/312/m312p141.pdf.

Kahru:2000:CBG

[KLRS00] M. Kahru, J. M. Leppänen, O. Rud, and O. P. Savchuk. Cyanobacteria blooms in the Gulf of Finland triggered by saltwater inflow into the Baltic Sea. Marine Ecology Progress Series, 207:13–18, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p013.pdf.

Kristiansen:2009:IBM

[KLW^{+}09] T. Kristiansen, R. G. Lough, F. E. Werner, E. A. Broughton, and L. J. Buckley. Individual-based modeling of feeding ecology and prey selection of larval cod on Georges Bank. Marine Ecology Progress Series, 376:227–243, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p227.pdf.

Koster:2000:ECB

[KM00] F. W. Köster and C. Möllmann. Egg cannibalism in Baltic sprat Sprattus sprattus. Marine Ecology Progress Series, 196: 269–277, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/196/m196p269.pdf.

Koch:2001:PPP

[KM01] M. S. Koch and C. J. Madden. Patterns of primary production and nutrient availability in a Bahamas lagoon with fringing mangroves. Marine Ecology Progress Series, 219: 109–119, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/219/m219p109.pdf.

Kido:2003:VOL

[KM03a] J. S. Kido and S. N. Murray. Variation in owl limpet Lottia gigantea population structures, growth rates, and gonadal production on southern California rocky shores. Marine Ecology Progress Series, 257:111–124, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/257/m257p111.pdf.

Kristensen:2003:IBD

[KM03b] E. Kristensen and O. L. Mikkelsen. Impact of the burrowdwelling polychaete Nereis diversicolor on the degradation of fresh and aged macroalgal detritus in a coastal marine sediment. Marine Ecology Progress Series, 265:141–153, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/265/m265p141.pdf.

Kuropat:2002:ERC

[KMAC^{+}02] C. Kuropat, R. Mercaldo-Allen, E. Caldarone, R. Goldberg, B. Phelan, and F. Thurberg. Evaluation of RNA concentration as an indicator of growth in young-of-the-year winter flounder Pseudopleuronectes americanus and tautog Tautoga onitis. Marine Ecology Progress Series, 230:265–274, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p265.pdf.

Kojadinovic:2008:TEM

[KMB^{+}08] J. Kojadinovic, F. Ménard, P. Bustamante, R. P. Cosson, and M. Le Corre. Trophic ecology of marine birds and pelagic fishes from Reunion Island as determined by stable isotope analysis. Marine Ecology Progress Series, 361:239– 251, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p239.pdf.

Kamio:2002:CPC

[KMF02] M. Kamio, S. Matsunaga, and N. Fusetani. Copulation pheromone in the crab Telmessus cheiragonus (Brachyura: Decapoda). Marine Ecology Progress Series, 234:183–190, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p183.pdf.

Kamenos:2004:NAF

[KMHS04] N. A. Kamenos, P. G. Moore, and J. M. Hall-Spencer. Nursery-area function of maerl grounds for juvenile queen scallops Aequipecten opercularis and other invertebrates. Marine Ecology Progress Series, 274:183–189, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p183.pdf.

Kurata:2001:SIA

[KMK01] K. Kurata, H. Minami, and E. Kikuchi. Stable isotope analysis of food sources for salt marsh snails. Marine Ecology Progress Series, 223:167–177, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/223/m223p167.pdf.

Karakassis:2006:CCC

[KMP^{+}06] I. Karakassis, A. Machias, P. Pitta, K. N. Papadopoulou, C. J. Smith, E. T. Apostolaki, M. Giannoulaki, D. Koutsoubas, and S. Somarakis. Cross-community congruence of patterns in a marine ecosystem: Do the parts reflect the whole? Marine Ecology Progress Series, 310:47–54, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p047.pdf.

Koster:2001:CCM

[KMR01] M. Köster and L. A. Meyer-Reil. Characterization of carbon and microbial biomass pools in shallow water coastal sediments of the southern Baltic Sea (Nordrügensche Bodden). Marine Ecology Progress Series, 214:25–41, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p025.pdf.

Kimball:2004:TTP

[KMWH04] M. E. Kimball, J. M. Miller, P. E. Whitfield, and J. A. Hare. Thermal tolerance and potential distribution of invasive lionfish (Pterois volitans/miles complex) on the east coast of the United States. Marine Ecology Progress Series, 283:269–278, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/283/m283p269.pdf.

Kokita:2000:SVD

[KN00] T. Kokita and A. Nakazono. Seasonal variation in the diel spawning time of the coral reef fish Oxymonacanthus longirostris (Monacanthidae): parental control of progeny development. Marine Ecology Progress Series, 199:263–270, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p263.pdf.

Kneib:2003:BLC

[Kne03] R. T. Kneib. Bioenergetic and landscape considerations for scaling expectations of nekton production from intertidal marshes. Marine Ecology Progress Series, 264:279– 296, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p279.pdf.

Kneib:2009:GVD

[Kne09] R. T. Kneib. Genotypic variation does not explain differences in growth of mummichogs Fundulus heteroclitus from simple and complex tidal marsh landscapes. Marine Ecology Progress Series, 386:207–219, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/386/m386p207.pdf.

Kvamme:2003:MPN

[KNF^{+}03] C. Kvamme, L. Nøttestad, A. Fernö, O. A. Misund, A. Dommasnes, B. E. Axelsen, P. Dalpadado, and W. Melle. Migration patterns in Norwegian spring-spawning herring: why young fish swim away from the wintering area in late summer. Marine Ecology Progress Series, 247:197–210, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/

m247p197.pdf. See erratum [Ano03d].

Kitagawa:2000:EAT

[KNK^{+}00] T. Kitagawa, H. Nakata, S. Kimura, T. Itoh, S. Tsuji, and A. Nitta. Effect of ambient temperature on the vertical distribution and movement of Pacific bluefin tuna Thunnus thynnus orientalis. Marine Ecology Progress Series, 206:251–260, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p251.pdf.

Kitagawa:2001:TMI

[KNKT01] T. Kitagawa, H. Nakata, S. Kimura, and S. Tsuji. Thermoconservation mechanisms inferred from peritoneal cavity temperature in free-swimming Pacific bluefin tuna Thunnus thynnus orientalis. Marine Ecology Progress Series, 220:253–263, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p253.pdf.

Koike:2007:ADF

[KNT^{+}07] K. Koike, A. Nishiyama, K. Takishita, A. Kobiyama, and T. Ogata. Appearance of Dinophysis fortii following blooms of certain cryptophyte species. Marine Ecology Progress Series, 337:303–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p303.pdf.

Kanaya:2005:EDF

[KNTK05] G. Kanaya, E. Nobata, T. Toya, and E. Kikuchi. Effects of different feeding habits of three bivalve species on sediment characteristics and benthic diatom abundance. Marine Ecology Progress Series, 299:67–78, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/299/m299p067.pdf.

Kocum:2002:RPP

[KNU02] E. Kocum, D. B. Nedwell, and G. J. C. Underwood. Regulation of phytoplankton primary production along a hypernutrified estuary. Marine Ecology Progress Series, 231: 13–22, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p013.pdf.

Knutsen:2007:EDB

[KOC^{+}07] H. Knutsen, E. M. Olsen, L. Ciannelli, S. H. Espeland, J. A. Knutsen, J. H. Simonsen, S. Skreslet, and N. C. Stenseth. Egg distribution, bottom topography and small-scale cod population structure in a coastal marine system. Marine Ecology Progress Series, 333:249–255, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/333/m333p249.pdf.

Koeve:2005:MEC

[Koe05] W. Koeve. Magnitude of excess carbon sequestration into the deep ocean and the possible role of TEP. Marine Ecology Progress Series, 291:53–64, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/291/m291p053.pdf.

Konar:2000:LEK

[Kon00] B. Konar. Limited effects of a keystone species: trends of sea otters and kelp forests at the Semichi Islands, Alaska. Marine Ecology Progress Series, 199:271–280, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p271.pdf.

Koranteng:2001:SDD

[Kor01] K. A. Koranteng. Structure and dynamics of demersal assemblages on the continental shelf and upper slope off Ghana, West Africa. Marine Ecology Progress Series, 220:1–12, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p001.pdf.

Kangro:2007:SSR

[KOTL07] K. Kangro, K. Olli, T. Tamminen, and R. Lignell. Speciesspecific responses of a cyanobacteria-dominated phytoplankton community to artificial nutrient limitation in the Baltic Sea. Marine Ecology Progress Series, 336:15–27, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m336p015.pdf.

Kotake:2005:SVM

[KOY^{+}05] A. Kotake, A. Okamura, Y. Yamada, T. Utoh, T. Arai, M. J. Miller, H. P. Oka, and K. Tsukamoto. Seasonal variation in the migratory history of the Japanese eel Anguilla japonica in Mikawa Bay, Japan. Marine Ecology Progress Series, 293:213–221, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p213.pdf.

Kang:2000:RSC

[KP00] H. K. Kang and S. A. Poulet. Reproductive success in Calanus helgolandicus as a function of diet and egg cannibalism. Marine Ecology Progress Series, 201:241–250, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p241.pdf.

Kaehler:2001:ESP

[KP01a] S. Kaehler and E. A. Pakhomov. Effects of storage and preservation on the δ ^{13}C and δ ^{15}N signatures of selected marine organisms. Marine Ecology Progress Series, 219:299–304, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p299.pdf.

Kuffner:2001:ENP

[KP01b] I. B. Kuffner and V. J. Paul. Effects of nitrate, phosphate and iron on the growth of macroalgae and benthic cyanobacteria from Cocos Lagoon, Guam. Marine Ecology Progress Series, 222:63–72, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p063.pdf.

Kalman:2009:AHS

[KPA^{+}09] J. Kalman, F. Palais, J. C. Amiard, C. Mouneyrac, A. Muntz, J. Blasco, I. Riba, and C. Amiard-Triquet. Assessment of the health status of populations of the ragworm Nereis diversicolor using biomarkers at different levels of biological organisation. Marine Ecology Progress Series, 393:55– 67, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/393/m393p055.pdf.

Kubanek:2000:TGD

[KPEF00] J. Kubanek, J. R. Pawlik, T. M. Eve, and W. Fenical. Triterpene glycosides defend the Caribbean reef sponge Erylus formosus from predatory fishes. Marine Ecology Progress Series, 207:69–77, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p069.pdf.

Kaehler:2006:TIK

[KPKD06] S. Kaehler, E. A. Pakhomov, R. M. Kalin, and S. Davis. Trophic importance of kelp-derived suspended particulate matter in a through-flow sub-Antarctic system. Marine Ecology Progress Series, 316:17–22, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/316/
m316p017.pdf.

Kaehler:2000:TSM

[KPM00] S. Kaehler, E. A. Pakhomov, and C. D. McQuaid. Trophic structure of the marine food web at the Prince Edward Islands (Southern Ocean) determined by δ ^{13}C and δ ^{15}N analysis. Marine Ecology Progress Series, 208:13–20, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
208/m208p013.pdf.

Kibirige:2003:GRF

[KPN03] I. Kibirige, R. Perissinotto, and C. Nozais. Grazing rates and feeding preferences of the mysid shrimp Gastrosaccus brevifissura in a temporarily open estuary in South Africa. Marine Ecology Progress Series, 251:201–210, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/251/
m251p201.pdf.

Kiorboe:2001:FMSa

[KPT01] T. Kiørboe, H. Ploug, and U. H. Thygesen. Fluid motion and solute distribution around sinking aggregates. I. Smallscale fluxes and heterogeneity of nutrients in the pelagic environment. Marine Ecology Progress Series, 211:1–13, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p001.pdf.

Keller:2003:ROA

[KPW03] T. A. Keller, I. Powell, and M. J. Weissburg. Role of olfactory appendages in chemically mediated orientation of blue crabs. Marine Ecology Progress Series, 261:217–231, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/261/
m261p217.pdf.

Kitaysky:2007:SHL

[KPW07] A. S. Kitaysky, J. F. Piatt, and J. C. Wingfield. Stress hormones link food availability and population processes in seabirds. Marine Ecology Progress Series, 352:245–258, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m352p245.pdf.

King:2001:EMC

[KR01] C. K. King and M. J. Riddle. Effects of metal contaminants on the development of the common Antarctic sea urchin Sterechinus neumayeri and comparisons of sensitivity with tropical and temperate echinoids. Marine Ecology Progress Series, 215:143–154, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p143.pdf.

Kimmel:2004:LTT

[KR04] D. G. Kimmel and M. R. Roman. Long-term trends in mesozooplankton abundance in Chesapeake Bay, USA: influence of freshwater input. Marine Ecology Progress Series, 267:71–83, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/267/m267p071.pdf.

Kittner:2005:ETF

[KR05] C. Kittner and H. U. Riisgård. Effect of temperature on filtration rate in the mussel Mytilus edulis: no evidence for temperature compensation. Marine Ecology Progress Series, 305:147–152, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p147.pdf.

Kraak:2007:DPM

[Kra07] S. B. M. Kraak. Does the probabilistic maturation reaction norm approach disentangle phenotypic plasticity from genetic change? Marine Ecology Progress Series, 335:295–300, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m335p295.pdf.

Kato:2006:LFS

[KRCGC06] A. Kato, Y. Ropert-Coudert, D. Grémillet, and B. Cannell. Locomotion and foraging strategy in foot-propelled and wingpropelled shallow-diving seabirds. Marine Ecology Progress Series, 308:293–301, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/308/m308p293.pdf.

Kremp:2001:ECR

[Kre01] A. Kremp. Effects of cyst resuspension on germination and seeding of two bloom-forming dinoflagellates in the Baltic Sea. Marine Ecology Progress Series, 216:57–66, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p057.pdf.

Kaartvedt:2005:PFP

[KRF^{+}05] S. Kaartvedt, A. Røstad, Ó. Fiksen, W. Melle, T. Torgersen, M. T. Breien, and T. A. Klevjer. Piscivorous fish patrol krill swarms. Marine Ecology Progress Series, 299:1–5, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kritzer:2002:VPB

[Kri02] J. P. Kritzer. Variation in the population biology of stripey bass Lutjanus carponotatus within and between two island groups on the Great Barrier Reef. Marine Ecology Progress Series, 243:191–207, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/243/m243p191.pdf.

Kaartvedt:2009:SSS

[KRK09] S. Kaartvedt, A. Røstad, and T. A. Klevjer. Sprat Sprattus sprattus can exploit low oxygen waters for overwintering. Marine Ecology Progress Series, 390:237–249, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m390p237.pdf.

Kaartvedt:2009:UBM

[KRKS09] S. Kaartvedt, A. Røstad, T. A. Klevjer, and A. Staby. Use of bottom-mounted echo sounders in exploring behavior of mesopelagic fishes. Marine Ecology Progress Series, 395:109– 118, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m395p109.pdf.

Kroon:2005:BAA

[Kro05] F. J. Kroon. Behavioural avoidance of acidified water by juveniles of four commercial fish and prawn species with migratory life stages. Marine Ecology Progress Series, 285: 193–204, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p193.pdf.

Klimpel:2006:DMP

[KRP^{+}06] S. Klimpel, S. Rückert, U. Piatkowski, H. W. Palm, and R. Hanel. Diet and metazoan parasites of silver scabbard fish Lepidopus caudatus from the Great Meteor Seamount (North Atlantic). Marine Ecology Progress Series, 315:249– 257, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/315/m315p249.pdf.

Krug:2001:BHD

[Kru01] P. J. Krug. Bet-hedging dispersal strategy of a specialist marine herbivore: a settlement dimorphism among sibling larvae of Alderia modesta. Marine Ecology Progress Series, 213:177–192, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p177.pdf.

Klein:2005:SID

[KRVB05] J. Klein, S. Ruitton, M. Verlaque, and C. F. Boudouresque. Species introductions, diversity and disturbances in marine macrophyte assemblages of the northwestern Mediterranean Sea. Marine Ecology Progress Series, 290:79–88, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/290/m290p079.pdf.

Klomp:2000:STS

[KS00a] N. I. Klomp and M. A. Schultz. Short-tailed shearwaters breeding in Australia forage in Antarctic waters. Marine Ecology Progress Series, 194:307–310, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/194/m194p307.pdf.

Kneib:2000:DTM

[KS00b] R. T. Kneib and C. E. H. Scheele. Does tethering of mobile prey measure relative predation potential? An empirical test using mummichogs and grass shrimp. Marine Ecology Progress Series, 198:181–190, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/198/m198p181.pdf.

Kiyofuji:2004:UNV

[KS04a] H. Kiyofuji and S. I. Saitoh. Use of nighttime visible images to detect Japanese common squid Todarodes pacificus fishing areas and potential migration routes in the Sea of Japan. Marine Ecology Progress Series, 276:173–186, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p173.pdf.

Kraus:2004:DWP

[KS04b] R. T. Kraus and D. H. Secor. Dynamics of white perch Morone americana population contingents in the Patuxent River estuary, Maryland, USA. Marine Ecology Progress Series, 279:247–259, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p247.pdf.

Kurihara:2004:EIA

[KS04c] H. Kurihara and Y. Shirayama. Effects of increased atmospheric CO_{2} on sea urchin early development. Marine Ecology Progress Series, 274:161–169, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p161.pdf.

Kraus:2005:ANR

[KS05] R. T. Kraus and D. H. Secor. Application of the nurseryrole hypothesis to an estuarine fish. Marine Ecology Progress Series, 291:301–305, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/291/m291p301.pdf. See erra-

tum [Ano05c].

Kamiyama:2006:LAP

[KS06] T. Kamiyama and T. Suzuki. Lack of accumulation of paralytic shellfish poisoning (PSP) toxins in the tintinnid ciliate Favella taraikaensis feeding on the toxic dinoflagellate Alexandrium tamarense. Marine Ecology Progress Series, 317: 57–65, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/317/m317p057.pdf.

Kai:2008:ECS

[KS08] S. Kai and K. Sakai. Effect of colony size and age on resource allocation between growth and reproduction in the corals Goniastrea aspera and Favites chinensis. Marine Ecology Progress Series, 354:133–139, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/354/m354p133.pdf.

Kuo:2009:GVU

[KS09] E. S. L. Kuo and E. Sanford. Geographic variation in the upper thermal limits of an intertidal snail: implications for climate envelope models. Marine Ecology Progress Series, 388:137–146, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m388p137.pdf.

Kremp:2003:EDF

[KSA03] A. Kremp, D. H. Shull, and D. M. Anderson. Effects of deposit-feeder gut passage and fecal pellet encapsulation on germination of dinoflagellate resting cysts. Marine Ecology Progress Series, 263:65–73, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/263/m263p065.pdf.

Klopper:2003:DST

[KSG03] S. Klöpper, R. Scharek, and G. Gerdts. Diarrhetic shellfish toxicity in relation to the abundance of Dinophysis spp. in the German Bight near Helgoland. Marine Ecology Progress Series, 259:93–102, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p093.pdf.

Kochzius:2009:GPS

[KSH^{+}09] M. Kochzius, C. Seidel, J. Hauschild, S. Kirchhoff, P. Mester, I. Meyer-Wachsmuth, A. Nuryanto, and J. Timm. Genetic population structures of the blue starfish Linckia laevigata and its gastropod ectoparasite Thyca crystallina. Marine Ecology Progress Series, 396:211–219, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p211.pdf.

Kojima:2001:PDS

[KSHO01] S. Kojima, R. Segawa, I. Hayashi, and M. Okiyama. Phylogeography of a deep-sea demersal fish, Bothrocara hollandi, in the Japan Sea. Marine Ecology Progress Series, 217:135– 143, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/217/m217p135.pdf.

Kimura:2000:EDY

[KSHP00] R. Kimura, D. H. Secor, E. D. Houde, and P. M. Piccoli. Upestuary dispersal of young-of-the-year bay anchovy Anchoa mitchilli in the Chesapeake Bay: inferences from microprobe analysis of strontium in otoliths. Marine Ecology Progress Series, 208:217–227, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/208/m208p217.pdf.

Kelly:2009:FAT

[KSI09] J. R. Kelly, R. E. Scheibling, and S. J. Iverson. Fatty acids tracers for native and invasive macroalgae in an experimental food web. Marine Ecology Progress Series, 391: 53–63, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/391/m391p053.pdf.

Kelly:2008:FAP

[KSIG08] J. R. Kelly, R. E. Scheibling, S. J. Iverson, and P. Gagnon. Fatty acid profiles in the gonads of the sea urchin Strongylocentrotus droebachiensis on natural algal diets. Marine Ecology Progress Series, 373:1–9, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kerr:2007:SIS

[KSK07] L. A. Kerr, D. H. Secor, and R. T. Kraus. Stable isotope (δ 13C and δ ^{18}O) and Sr/Ca composition of otoliths as proxies for environmental salinity experienced by an estuarine fish. Marine Ecology Progress Series, 349:245–253, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/

m349p245.pdf. See erratum [Ano10b].

Kozlowsky-Suzuki:2007:DMF

[KSKK^{+}07] B. Kozlowsky-Suzuki, M. Karjalainen, M. Koski, P. Carlsson, W. Stolte, M. Balode, and E. Granéli. Disruption of the microbial food web and inhibition of metazooplankton development in the presence of iron- and DOM-stimulated Baltic Sea cyanobacteria. Marine Ecology Progress Series, 337:15–26, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p015.pdf.

Kozlowsky-Suzuki:2003:FRT

[KSKL^{+}03] B. Kozlowsky-Suzuki, M. Karjalainen, M. Lehtiniemi, J. Engström-Öst, M. Koski, and P. Carlsson. Feeding, reproduction and toxin accumulation by the copepods Acartia bifilosa and Eurytemora affinis in the presence of the toxic cyanobacterium Nodularia spumigena. Marine Ecology Progress Series, 249:237–249, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p237.pdf.

Koehl:2007:IBM

[KSR^{+}07] M. A. R. Koehl, J. A. Strother, M. A. Reidenbach, J. R. Koseff, and M. G. Hadfield. Individual-based model of larval transport to coral reefs in turbulent, wave-driven flow: behavioral responses to dissolved settlement inducer. Marine Ecology Progress Series, 335:1–18, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

King:2005:STA

[KSS^{+}05] C. K. King, S. L. Simpson, S. V. Smith, J. L. Stauber, and G. E. Batley. Short-term accumulation of Cd and Cu from water, sediment and algae by the amphipod Melita plumulosa and the bivalve Tellina deltoidalis. Marine Ecology Progress Series, 287:177–188, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/287/m287p177.pdf.

Kamenos:2008:RCA

[KSS^{+}08] N. A. Kamenos, S. C. Strong, D. M. Shenoy, S. T. Wilson, A. D. Hatton, and P. G. Moore. Red coralline algae as a source of marine biogenic dimethylsulphoniopropionate. Marine Ecology Progress Series, 372:61–66, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p061.pdf.

Kooyman:2004:MHP

[KSSB04] G. L. Kooyman, D. B. Siniff, I. Stirling, and J. L. Bengtson. Moult habitat, pre- and post-moult diet and post-moult travel of Ross Sea emperor penguins. Marine Ecology Progress Series, 267:281–290, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/267/m267p281.pdf.

Katechakis:2002:CPC

[KSSH02] A. Katechakis, H. Stibor, U. Sommer, and T. Hansen. Changes in the phytoplankton community and microbial food web of Blanes Bay (Catalan Sea, NW Mediterranean) under prolonged grazing pressure by doliolids (Tunicata), cladocerans or copepods (Crustacea). Marine Ecology Progress Series, 234:55–69, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p055.pdf.

Kuramochi:2008:LTC

[KST^{+}08] T. Kuramochi, Y. Sudo, E. Tamaki, K. Shiroma, and T. Naganuma. Long-term collection of benthic and benthopelagic organisms from a deep-water inlet offshore from Okinawa, Japan. Marine Ecology Progress Series, 358:301–307, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/358/m358p301.pdf.

Kroncke:2004:RBS

[KSWP04] I. Kröncke, T. Stoeck, G. Wieking, and A. Palojärvi. Relationship between structural and functional aspects of microbial and macrofaunal communities in different areas of the North Sea. Marine Ecology Progress Series, 282:13– 31, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p013.pdf.

Kiorboe:2001:FMSb

[KT01] T. Kiørboe and U. H. Thygesen. Fluid motion and solute distribution around sinking aggregates. II. Implications for remote detection by colonizing zooplankters. Marine Ecology Progress Series, 211:15–25, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/211/m211p015.pdf.

Kroncke:2003:SFA

[KT03] I. Kröncke and M. Türkay. Structural and functional aspects of the benthic communities in the deep Angola Basin. Marine Ecology Progress Series, 260:43–53, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p043.pdf.

Kawai:2004:VMF

[KT04] T. Kawai and M. Tokeshi. Variable modes of facilitation in the upper intertidal: goose barnacles and mussels. Marine Ecology Progress Series, 272:203–213, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p203.pdf.

Kamimura:2006:EOF

[KT06] S. Kamimura and M. Tsuchiya. Effects of opportunistic feeding by the intertidal gastropods Batillaria zonalis and B. flectosiphonata on material flux on a tidal flat. Marine Ecology Progress Series, 318:203–211, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/318/m318p203.pdf.

Kostylev:2001:BHM

[KTF^{+}01] V. E. Kostylev, B. J. Todd, G. B. J. Fader, R. C. Courtney, G. D. M. Cameron, and R. A. Pickrill. Benthic habitat mapping on the Scotian Shelf based on multibeam bathymetry, surficial geology and sea floor photographs. Marine Ecology Progress Series, 219:121–137, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p121.pdf.

Kattner:2004:SIG

[KTH^{+}04] G. Kattner, D. N. Thomas, C. Haas, H. Kennedy, and G. S. Dieckmann. Surface ice and gap layers in Antarctic sea ice: highly productive habitats. Marine Ecology Progress Series, 277:1–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/277/m277p001.pdf.

Kennedy:2002:POM

[KTK^{+}02] H. Kennedy, D. N. Thomas, G. Kattner, C. Haas, and G. S. Dieckmann. Particulate organic matter in Antarctic summer sea ice: concentration and stable isotopic composition. Marine Ecology Progress Series, 238:1–13, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p001.pdf.

Kaartvedt:2008:BIM

[KTK^{+}08a] S. Kaartvedt, T. Torgersen, T. A. Klevjer, A. Røstad, and J. A. Devine. Behavior of individual mesopelagic fish in acoustic scattering layers of Norwegian fjords. Marine Ecology Progress Series, 360:201–209, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p201.pdf.

Kanaya:2008:SDV

[KTK08b] G. Kanaya, S. Takagi, and E. Kikuchi. Spatial dietary variations in Laternula marilina (Bivalva) and Hediste spp. (Polychaeta) along environmental gradients in two brackish lagoons. Marine Ecology Progress Series, 359:133–144, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/359/m359p133.pdf.

Kanaya:2007:SDS

[KTNK07] G. Kanaya, S. Takagi, E. Nobata, and E. Kikuchi. Spatial dietary shift of macrozoobenthos in a brackish lagoon revealed by carbon and nitrogen stable isotope ratios. Marine Ecology Progress Series, 345:117–127, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/345/m345p117.pdf.

Karakassis:2002:FFI

[KTSR02] I. Karakassis, M. Tsapakis, C. J. Smith, and H. Rumohr. Fish farming impacts in the Mediterranean studied through sediment profiling imagery. Marine Ecology Progress Series, 227:125–133, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p125.pdf.

Karaiskou:2004:SGS

[KTT04] N. Karaiskou, A. Triantafyllidis, and C. Triantaphyllidis. Shallow genetic structure of 3 species of the genus Trachurus in European waters. Marine Ecology Progress Series, 281: 193–205, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/281/m281p193.pdf.

Kuffner:2001:EUU

[Kuf01] I. B. Kuffner. Effects of ultraviolet (UV) radiation on larval settlement of the reef coral Pocillopora damicornis. Marine Ecology Progress Series, 217:251–261, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p251.pdf.

Kumagai:2006:DEP

[Kum06] N. H. Kumagai. Distance effects on patterns and processes of dispersal in an octocoral-associated amphipod. Marine Ecology Progress Series, 321:203–214, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p203.pdf.

Kocum:2002:SMP

[KUN02] E. Kocum, G. J. C. Underwood, and D. B. Nedwell. Simultaneous measurement of phytoplanktonic primary production, nutrient and light availability along a turbid, eutrophic UK east coast estuary (the Colne Estuary). Marine Ecology Progress Series, 231:1–12, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/231/m231p001.pdf.

Kawahara:2006:UPE

[KUOI06] M. Kawahara, S. I. Uye, K. Ohtsu, and H. Iizumi. Unusual population explosion of the giant jellyfish Nemopilema nomurai (Scyphozoa: Rhizostomeae) in East Asian waters. Marine Ecology Progress Series, 307:161–173, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/307/
m307p161.pdf.

Kupschus:2003:DES

[Kup03] S. Kupschus. Development and evaluation of statistical habitat suitability models: an example based on juvenile spotted seatrout Cynoscion nebulosus. Marine Ecology Progress Series, 265:197–212, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p197.pdf.

Kurihara:2008:ECD

[Kur08] H. Kurihara. Effects of CO_{2}-driven ocean acidification on the early developmental stages of invertebrates. Marine Ecology Progress Series, 373:275–284, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m373p275.pdf.

Kraan:2009:PMI

[KvdMDP09] C. Kraan, J. van der Meer, A. Dekinga, and T. Piersma. Patchiness of macrobenthic invertebrates in homogenized intertidal habitats: hidden spatial structure at a landscape scale. Marine Ecology Progress Series, 383:211–224, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/383/m383p211.pdf.

Kirsch:2002:PGT

[KVH02] K. D. Kirsch, J. F. Valentine, and K. L. Heck, Jr. Parrotfish grazing on turtlegrass Thalassia testudinum: evidence for the importance of seagrass consumption in food web dynamics of the Florida Keys National Marine Sanctuary. Marine Ecology Progress Series, 227:71–85, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p071.pdf.

Kater:2006:DCC

[KvKB06] B. J. Kater, A. J. M. Geurts van Kessel, and J. J. M. D. Baars. Distribution of cockles Cerastoderma edule in the Eastern Scheldt: habitat mapping with abiotic variables. Marine Ecology Progress Series, 318:221–227, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p221.pdf.

Kuipers:2000:DPB

[KvNVH00] B. Kuipers, G. J. van Noort, J. Vosjan, and G. J. Herndl. Diel periodicity of bacterioplankton in the euphotic zone of the subtropical Atlantic Ocean. Marine Ecology Progress Series, 201:13–25, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p013.pdf.

Kahl:2008:EPP

[KVS08a] L. A. Kahl, A. Vardi, and O. Schofield. Effects of phytoplankton physiology on export flux. Marine Ecology Progress Series, 354:3–19, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/feature/m354p003.pdf.

Klaas:2008:DCG

[KVS08b] C. Klaas, P. G. Verity, and S. Schultes. Determination of copepod grazing on natural plankton communities: correcting for trophic cascade effects. Marine Ecology Progress Series, 357:195–206, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/357/m357p195.pdf.

Kroncke:2000:DBS

[KVV^{+}00] I. Kröncke, A. Vanreusel, M. Vincx, J. Wollenburg, A. Mackensen, G. Liebezeit, and B. Behrends. Different benthic sizecompartments and their relationship to sediment chemistry in the deep Eurasian Arctic Ocean. Marine Ecology Progress Series, 199:31–41, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p031.pdf.

Kihslinger:2000:FPS

[KW00a] R. L. Kihslinger and S. A. Woodin. Food patches and a surface deposit feeding spionid polychaete. Marine Ecology Progress Series, 201:233–239, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/201/m201p233.pdf.

Kuipers:2000:PSP

[KW00b] B. R. Kuipers and H. J. Witte. Prochlorophytes as secondary prey for heterotrophic nanoflagellates in the deep chlorophyll maximum layer of the (sub)tropical North Atlantic. Marine Ecology Progress Series, 204:53–63, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/204/m204p053.pdf.

Koch:2002:EBE

[KW02a] V. Koch and M. Wolff. Energy budget and ecological role of mangrove epibenthos in the Caeté estuary, North Brazil.

Marine Ecology Progress Series, 228:119–130, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p119.pdf.

Kurle:2002:SNC

[KW02b] C. M. Kurle and G. A. J. Worthy. Stable nitrogen and carbon isotope ratios in multiple tissues of the northern fur seal Callorhinus ursinus: implications for dietary and migratory reconstructions. Marine Ecology Progress Series, 236: 289–300, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/236/m236p289.pdf.

Kamp:2005:PLP

[KW05] A. Kamp and U. Witte. Processing of ^{13}C-labelled phytoplankton in a fine-grained sandy-shelf sediment (North Sea): relative importance of different macrofauna species. Marine Ecology Progress Series, 297:61–70, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p061.pdf.

Korb:2008:MMP

[KWAT08] R. E. Korb, M. J. Whitehouse, A. Atkinson, and S. E. Thorpe. Magnitude and maintenance of the phytoplankton bloom at South Georgia: a naturally iron-replete environment. Marine Ecology Progress Series, 368:75–91, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p075.pdf.

Kuffner:2006:ICR

[KWB^{+}06] I. B. Kuffner, L. J. Walters, M. A. Becerro, V. J. Paul, R. Ritson-Williams, and K. S. Beach. Inhibition of coral recruitment by macroalgae and cyanobacteria. Marine Ecology Progress Series, 323:107–117, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m323p107.pdf.

Koch:2005:CPD

[KWD05] V. Koch, M. Wolff, and K. Diele. Comparative population dynamics of four fiddler crabs (ocypodidae, genus Uca) from a North Brazilian mangrove ecosystem. Marine Ecology Progress Series, 291:177–188, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/291/m291p177.pdf.

Kerford:2008:DRD

[KWHD08] M. R. Kerford, A. J. Wirsing, M. R. Heithaus, and L. M. Dill. Danger on the rise: diurnal tidal state mediates an exchange of food for safety by the bar-bellied sea snake Hydrophis elegans. Marine Ecology Progress Series, 358:289–294, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p289.pdf.

Kramarsky-Winter:2006:IPC

[KWHS^{+}06] E. Kramarsky-Winter, M. Harel, N. Siboni, E. Ben Dov, I. Brickner, Y. Loya, and A. Kushmaro. Identification of a protist-coral association and its possible ecological role. Marine Ecology Progress Series, 317:67–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p067.pdf.

Kinoshita:2003:MSB

[KWKF03] K. Kinoshita, M. Wada, K. Kogure, and T. Furota. Mud shrimp burrows as dynamic traps and processors of tidalflat materials. Marine Ecology Progress Series, 247:159– 164, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p159.pdf.

Kulkarni:2005:DSM

[KWL^{+}05] N. R. Kulkarni, D. L. White, A. J. Lewitus, R. G. Tymowski, and D. C. Yoch. Dimethylsulfoniopropionate in a salt marsh estuary: correlation to tidal cycle and phytoplankton assemblage composition. Marine Ecology Progress Series, 289:13–25, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p013.pdf.

Karpouzi:2007:MMR

[KWP07] V. S. Karpouzi, R. Watson, and D. Pauly. Modelling and mapping resource overlap between seabirds and fisheries on a global scale: a preliminary assessment. Marine Ecology Progress Series, 343:87–99, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/343/m343p087.pdf.

Kaschner:2006:MWW

[KWTP06] K. Kaschner, R. Watson, A. W. Trites, and D. Pauly. Mapping world-wide distributions of marine mammal species using a relative environmental suitability (RES) model. Marine Ecology Progress Series, 316:285–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m316p285.pdf.

Kilminster:2008:CGI

[KWTR08] K. L. Kilminster, D. I. Walker, P. A. Thompson, and J. A. Raven. Changes in growth, internode distance and nutrient concentrations of the seagrass Halophila ovalis with exposure to sediment sulphide. Marine Ecology Progress Series, 361:83–91, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p083.pdf.

Kauppila:2005:TPR

[KWV^{+}05] P. Kauppila, K. Weckström, S. Vaalgamaa, A. Korhola, H. Pitkänen, N. Reuss, and S. Drew. Tracing pollution and recovery using sediments in an urban estuary, northern Baltic Sea: are we far from ecological reference conditions? Marine Ecology Progress Series, 290:35–53, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p035.pdf.

Kendall:2002:IMS

[KWWH02] M. S. Kendall, D. L. Wolcott, T. G. Wolcott, and A. H. Hines. Influence of male size and mating history on sperm content of ejaculates of the blue crab Callinectes sapidus. Marine Ecology Progress Series, 230:235–240, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/230/m230p235.pdf.

Kim:2002:EWT

[KYML02] W. S. Kim, S. J. Yoon, H. T. Moon, and T. W. Lee. Effects of water temperature changes on the endogenous and exogenous rhythms of oxygen consumption in glass eels Anguilla japonica. Marine Ecology Progress Series, 243:209– 216, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p209.pdf.

Kudo:2000:ENT

[KYYM00] I. Kudo, T. Yoshimura, M. Yanada, and K. Matsunaga. Exhaustion of nitrate terminates a phytoplankton bloom in Funka Bay, Japan: change in SiO_{4} : NO_{3} consumption rate during the bloom. Marine Ecology Progress Series, 193: 45–51, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/193/m193p045.pdf.

Krug:2000:LSC

[KZ00] P. J. Krug and R. K. Zimmer. Larval settlement: chemical markers for tracing production, transport, and distribution of a waterborne cue. Marine Ecology Progress Series, 207:283–296, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p283.pdf.

Langlet:2006:ENC

[LABR^{+}06] D. Langlet, M. Alunno-Bruscia, M. Rafélis, M. Renard, M. Roux, E. Schein, and D. Buestel. Experimental and natural cathodoluminescence in the shell of Crassostrea gigas from Thau lagoon (France): ecological and environmental implications. Marine Ecology Progress Series, 317:143– 156, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/317/m317p143.pdf.

Lindegarth:2001:EGS

[LÅCN01] M. Lindegarth, P. Åberg, G. Cervin, and P. G. Nilsson. Effects of grazing on the structure of mid-shore, intertidal assemblages on moderately exposed rocky shores of the Swedish west coast. Marine Ecology Progress Series, 212:29–38, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p029.pdf.

Luque:2008:TSD

[LAG08] S. P. Luque, J. P. Y. Arnould, and C. Guinet. Temporal structure of diving behaviour in sympatric Antarctic and subantarctic fur seals. Marine Ecology Progress Series, 372: 277–287, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p277.pdf.

Littnan:2007:EPS

[LAH07] C. L. Littnan, J. P. Y. Arnould, and R. G. Harcourt. Effect of proximity to the shelf edge on the diet of female Australian fur seals. Marine Ecology Progress Series, 338: 257–267, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p257.pdf.

Lindholm:2007:SFM

[LAK07] J. Lindholm, P. J. Auster, and A. Knight. Site fidelity and movement of adult Atlantic cod Gadus morhua at deep boulder reefs in the western Gulf of Maine, USA. Marine Ecology Progress Series, 342:239–247, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p239.pdf.

Lenihan:2008:RSR

[LAK^{+}08] H. S. Lenihan, M. Adjeroud, M. J. Kotchen, J. L. Hench, and T. Nakamura. Reef structure regulates small-scale spatial variation in coral bleaching. Marine Ecology Progress Series, 370:127–141, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/370/m370p127.pdf.

Luo:2008:SMP

[LALB08] J. Luo, J. S. Ault, M. F. Larkin, and L. R. Barbieri. Salinity measurements from pop-up archival transmitting (PAT) tags and their application to geolocation estimation for Atlantic tarpon. Marine Ecology Progress Series, 357:101– 109, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/357/m357p101.pdf.

Lam:2000:EGP

[Lam00a] K. K. Y. Lam. Early growth of a pioneer recruited coral Oulastrea crispata (Scleractinia, Faviidae) on PFA-concrete blocks in a marine park in Hong Kong, China. Marine Ecology Progress Series, 205:113–121, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/205/m205p113.pdf.

Lam:2000:SRL

[Lam00b] K. K. Y. Lam. Sexual reproduction of a low-temperature tolerant coral Oulastrea crispata (Scleractinia, Faviidae) in Hong Kong, China. Marine Ecology Progress Series, 205: 101–111, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p101.pdf.

Lozano-Alvarez:2009:OHS

[LÁMBF09] E. Lozano-Álvarez, C. Meiners, and P. Briones-Fourzán. Ontogenetic habitat shifts affect performance of artificial shelters for Caribbean spiny lobsters. Marine Ecology Progress Series, 396:85–97, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m396p085.pdf.

Lapolla:2001:BAAa

[Lap01a] A. E. Lapolla. Bay anchovy Anchoa mitchilli in Narragansett Bay, Rhode Island. I. Population structure, growth and mortality. Marine Ecology Progress Series, 217:93–102, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p093.pdf.

Lapolla:2001:BAAb

[Lap01b] A. E. Lapolla. Bay anchovy Anchoa mitchilli in Narragansett Bay, Rhode Island. II. Spawning season, hatch-date distribution and young-of-the-year growth. Marine Ecology Progress Series, 217:103–109, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/217/m217p103.pdf.

Larned:2003:EIN

[Lar03] S. T. Larned. Effects of the invasive, nonindigenous seagrass Zostera japonica on nutrient fluxes between the water column and benthos in a NE Pacific estuary. Marine Ecology Progress Series, 254:69–80, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p069.pdf.

Lasiak:2003:ITR

[Las03] T. Lasiak. Influence of taxonomic resolution, biological attributes and data transformations on multivariate comparisons of rocky macrofaunal assemblages. Marine Ecology Progress Series, 250:29–34, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/250/m250p029.pdf.

Latasa:2007:IEP

[Lat07] M. Latasa. Improving estimations of phytoplankton class abundances using CHEMTAX. Marine Ecology Progress Series, 329:13–21, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/329/m329p013.pdf.

Lindholm:2004:RLM

[LAV04] J. Lindholm, P. Auster, and P. Valentine. Role of a large marine protected area for conserving landscape attributes of sand habitats on Georges Bank (NW Atlantic). Marine Ecology Progress Series, 269:61–68, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p061.pdf.

Lindsey:2006:IBH

[LAW06] E. L. Lindsey, A. H. Altieri, and J. D. Witman. Influence of biogenic habitat on the recruitment and distribution of a subtidal xanthid crab. Marine Ecology Progress Series, 306:223–231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p223.pdf.

Law:2007:FIE

[Law07] R. Law. Fisheries-induced evolution: present status and future directions. Marine Ecology Progress Series, 335:271– 277, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m335p271.pdf.

Law:2008:PME

[Law08] C. S. Law. Predicting and monitoring the effects of largescale ocean iron fertilization on marine trace gas emissions. Marine Ecology Progress Series, 364:283–288, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m364p283.pdf.

Lamare:2001:SRN

[LB01] M. D. Lamare and M. F. Barker. Settlement and recruitment of the New Zealand sea urchin Evechinus chloroticus. Marine Ecology Progress Series, 218:153–166, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p153.pdf.

Leon:2002:SFH

[LB02] Y. L. León and K. A. Bjorndal. Selective feeding in the hawksbill turtle, an important predator in coral reef ecosystems. Marine Ecology Progress Series, 245:249–258, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p249.pdf.

Lapolla:2005:HDD

[LB05a] A. Lapolla and L. J. Buckley. Hatch date distributions of young-of-year haddock Melanogrammus aeglefinus in the Gulf of Maine/Georges Bank region: implications for recruitment. Marine Ecology Progress Series, 290:239–249, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p239.pdf.

Lescroel:2005:FUC

[LB05b] A. Lescroël and C. A. Bost. Foraging under contrasting oceanographic conditions: the gentoo penguin at Kerguelen Archipelago. Marine Ecology Progress Series, 302:245– 261, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p245.pdf.

LaPeyre:2008:PVN

[LB08a] M. K. La Peyre and T. Birdsong. Physical variation of nonvegetated marsh edge habitats, and use patterns by nekton in Barataria Bay, Louisiana, USA. Marine Ecology Progress Series, 356:51–61, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p051.pdf.

Lewis:2008:ESP

[LB08b] D. M. Lewis and S. I. Bala. An examination of saltatory predation strategies employed by fish larvae foraging in a variety of different turbulent regimes. Marine Ecology Progress Series, 359:261–274, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p261.pdf.

Lin:2008:PRI

[LBB08] D. T. Lin and J. H. Bailey-Brock. Partial recovery of infaunal communities during a fallow period at an openocean aquaculture. Marine Ecology Progress Series, 371: 65–72, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/371/m371p065.pdf.

Lake:2008:ARR

[LBBH08] S. Lake, H. Burton, R. Barker, and M. Hindell. Annual reproductive rates of Weddell seals in eastern Antarctica from 1973 to 2000. Marine Ecology Progress Series, 366:259– 270, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/366/m366p259.pdf.

Lee:2006:TCP

[LBBM06] H. W. Lee, J. H. Bailey-Brock, and M. M. McGurr. Temporal changes in the polychaete infaunal community surrounding a Hawaiian mariculture operation. Marine Ecology Progress Series, 307:175–185, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/307/m307p175.pdf.

Langston:2007:FMC

[LBC07] W. J. Langston, G. R. Burt, and B. S. Chesman. Feminisation of male clams Scrobicularia plana from estuaries in southwest UK and its induction by endocrine-disrupting chemicals. Marine Ecology Progress Series, 333:173–184, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/333/
m333p173.pdf.

Lundsten:2009:BIC

[LBC^{+}09] L. Lundsten, J. P. Barry, G. M. Cailliet, D. A. Clague, A. P. DeVogelaere, and J. B. Geller. Benthic invertebrate communities on three seamounts off southern and central California, USA. Marine Ecology Progress Series, 374:23– 32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/374/m374p023.pdf.

Lye:2005:EDS

[LBCS05] C. M. Lye, M. G. Bentley, A. S. Clare, and E. M. Sefton. Endocrine disruption in the shore crab Carcinus maenas — a biomarker for benthic marine invertebrates? Marine Ecology Progress Series, 288:221–232, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/288/m288p221.pdf.

Lima:2000:PDY

[LBD00] M. Lima, A. Brazeiro, and O. Defeo. Population dynamics of the yellow clam Mesodesma mactroides: recruitment variability, density-dependence and stochastic processes. Marine Ecology Progress Series, 207:97–108, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p097.pdf.

Lambshead:2002:LDP

[LBF^{+}02] P. J. D. Lambshead, C. J. Brown, T. J. Ferrero, N. J. Mitchell, C. R. Smith, L. E. Hawkins, and J. Tietjen. Latitudinal diversity patterns of deep-sea marine nematodes and organic fluxes: a test from the central equatorial Pacific. Marine Ecology Progress Series, 236:129–135, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/236/m236p129.pdf.

Labenia:2007:BII

[LBF^{+}07] J. S. Labenia, D. H. Baldwin, B. L. French, J. W. Davis, and N. L. Scholz. Behavioral impairment and increased predation mortality in cutthroat trout exposed to carbaryl. Marine Ecology Progress Series, 329:1–11, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lucas:2001:BPE

[LBH01] C. H. Lucas, C. Banham, and P. M. Holligan. Benthic-pelagic exchange of microalgae at a tidal flat. 2. Taxonomic analysis. Marine Ecology Progress Series, 212:39–52, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p039.pdf.

LaJeunesse:2004:CRS

[LBH^{+}04] T. C. LaJeunesse, R. Bhagooli, M. Hidaka, L. deVantier, T. Done, G. W. Schmidt, W. K. Fitt, and O. Hoegh-Guldberg. Closely related Symbiodinium spp. differ in relative dominance in coral reef host communities across environmental, latitudinal and biogeographic gradients. Marine Ecology Progress Series, 284:147–161, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p147.pdf.

Lowerre-Barbieri:2009:DSP

[LBHL^{+}09] S. K. Lowerre-Barbieri, N. Henderson, J. Llopiz, S. Walters, J. Bickford, and R. Muller. Defining a spawning population (spotted seatrout Cynoscion nebulosus) over temporal, spatial, and demographic scales. Marine Ecology Progress Series, 394:231–245, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/394/m394p231.pdf.

Litzow:2006:CRS

[LBPH06] M. A. Litzow, K. M. Bailey, F. G. Prahl, and R. Heintz. Climate regime shifts and reorganization of fish communities: the essential fatty acid limitation hypothesis. Marine Ecology Progress Series, 315:1–11, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ladah:2003:FSR

[LBPS03] L. Ladah, R. Bermudez, G. Pearson, and E. Serrão. Fertilization success and recruitment of dioecious and hermaphroditic fucoid seaweeds with contrasting distributions near their southern limit. Marine Ecology Progress Series, 262:173– 183, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p173.pdf.

Laurion:2004:PML

[LBR04] I. Laurion, F. Blouin, and S. Roy. Packaging of mycosporinelike amino acids in dinoflagellates. Marine Ecology Progress Series, 279:297–303, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/279/m279p297.pdf.

Louzao:2009:CVB

[LBR^{+}09] M. Louzao, J. Bécares, B. Rodríguez, K. D. Hyrenbach, A. Ruiz, and J. M. Arcos. Combining vessel-based surveys and tracking data to identify key marine areas for seabirds. Marine Ecology Progress Series, 391:183–197, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m391p183.pdf.

Lahaye:2005:LTD

[LBS^{+}05] V. Lahaye, P. Bustamante, J. Spitz, W. Dabin, K. Das, G. J. Pierce, and F. Caurant. Long-term dietary segregation of common dolphins Delphinus delphis in the Bay of Biscay, determined using cadmium as an ecological tracer. Marine Ecology Progress Series, 305:275–285, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p275.pdf.

Lake:2003:RTF

[LBvdH03] S. Lake, H. Burton, and J. van den Hoff. Regional, temporal and fine-scale spatial variation in Weddell seal diet at four coastal locations in east Antarctica. Marine Ecology Progress Series, 254:293–305, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/254/m254p293.pdf.

Landaeta:2002:SSE

[LC02] M. F. Landaeta and L. R. Castro. Spring spawning and early nursery zone of the mesopelagic fish Maurolicus parvipinnis at the coastal upwelling zone off talcahuano, central Chile. Marine Ecology Progress Series, 226:179–191, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p179.pdf.

Lefrancois:2003:IAO

[LC03] C. Lefrançois and G. Claireaux. Influence of ambient oxygenation and temperature on metabolic scope and scope for heart rate in the common sole Solea solea. Marine Ecology Progress Series, 259:273–284, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/259/m259p273.pdf.

Lapid:2006:LTE

[LC06] E. D. Lapid and N. E. Chadwick. Long-term effects of competition on coral growth and sweeper tentacle development. Marine Ecology Progress Series, 313:115–123, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p115.pdf.

Lambert:2007:CSE

[LC07a] A. M. Lambert and R. A. Casagrande. Characteristics of a successful estuarine invader: evidence of self-compatibility in native and non-native lineages of Phragmites australis. Marine Ecology Progress Series, 337:299–301, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/337/
m337p299.pdf.

Lee:2007:MMT

[LC07b] J. E. Lee and S. L. Chown. Mytilus on the move: transport of an invasive bivalve to the Antarctic. Marine Ecology Progress Series, 339:307–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/339/m339p307.pdf.

Llopiz:2008:PSS

[LC08] J. K. Llopiz and R. K. Cowen. Precocious, selective and successful feeding of larval billfishes in the oceanic Straits of Florida. Marine Ecology Progress Series, 358:231–244, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/358/m358p231.pdf.

Lee:2009:TDH

[LC09a] J. E. Lee and S. L. Chown. Temporal development of hull-fouling assemblages associated with an Antarctic supply vessel. Marine Ecology Progress Series, 386:97–105, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/386/m386p097.pdf.

Llopiz:2009:VTR

[LC09b] J. K. Llopiz and R. K. Cowen. Variability in the trophic role of coral reef fish larvae in the oceanic plankton. Marine Ecology Progress Series, 381:259–272, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/381/m381p259.pdf.

Lovvorn:2005:OMP

[LCB^{+}05] J. R. Lovvorn, L. W. Cooper, M. L. Brooks, C. C. De Ruyck, J. K. Bump, and J. M. Grebmeier. Organic matter pathways to zooplankton and benthos under pack ice in late winter and open water in late summer in the north-central Bering Sea. Marine Ecology Progress Series, 291:135–150, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p135.pdf.

Lampman:2001:MMP

[LCC01] G. G. Lampman, N. F. Caraco, and J. J. Cole. A method for the measurement of particulate C and P on the same filtered sample. Marine Ecology Progress Series, 217:59–65, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p059.pdf.

LeGoff-Vitry:2007:LSH

[LCC07] M. C. Le Goff-Vitry, A. D. Chipman, and T. Comtet. In situ hybridization on whole larvae: a novel method for monitoring bivalve larvae. Marine Ecology Progress Series, 343: 161–172, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p161.pdf.

Londono-Cruz:2003:IBM

[LCCTFO03] E. Londoño-Cruz, J. R. Cantera, G. Toro-Farmer, and C. Orozco. Internal bioerosion by macroborers in Pocillopora spp. in the tropical eastern Pacific. Marine Ecology Progress Series, 265:289–295, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/265/m265p289.pdf.

Linares:2005:IDE

[LCD^{+}05] C. Linares, R. Coma, D. Diaz, M. Zabala, B. Hereu, and L. Dantart. Immediate and delayed effects of a mass mortality event on gorgonian population dynamics and benthic community structure in the NW Mediterranean Sea. Marine Ecology Progress Series, 305:127–137, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p127.pdf.

Leis:2003:OPL

[LCE03] J. M. Leis and B. M. Carson-Ewart. Orientation of pelagic larvae of coral-reef fishes in the ocean. Marine Ecology Progress Series, 252:239–253, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/252/m252p239.pdf.

Leis:2002:SDL

[LCEC02] J. M. Leis, B. M. Carson-Ewart, and D. H. Cato. Sound detection in situ by the larvae of a coral-reef damselfish (Pomacentridae). Marine Ecology Progress Series, 232:259– 268, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/232/m232p259.pdf.

Lennert-Cody:2002:FPH

[LCF02] C. E. Lennert-Cody and P. J. S. Franks. Fluorescence patches in high-frequency internal waves. Marine Ecology Progress Series, 235:29–42, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p029.pdf.

Lea:2002:ASF

[LCGN02] M. A. Lea, Y. Cherel, C. Guinet, and P. D. Nichols. Antarctic fur seals foraging in the polar frontal zone: inter-annual shifts in diet as shown from fecal and fatty acid analyses. Marine Ecology Progress Series, 245:281–297, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/

m245p281.pdf. See erratum [Ano03e].

Lafrance:2003:CCW

[LCHG03] M. Lafrance, G. Cliche, G. A. Haugum, and H. Guderley. Comparison of cultured and wild sea scallops Placopecten magellanicus, using behavioral responses and morphometric and biochemical indices. Marine Ecology Progress Series, 250:183–195, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p183.pdf.

Landry:2000:BRIb

[LCL^{+}00] M. R. Landry, J. Constantinou, M. Latasa, S. L. Brown, R. R. Bidigare, and M. E. Ondrusek. Biological response to iron fertilization in the eastern equatorial Pacific (IronEx II). III. Dynamics of phytoplankton growth and microzooplankton grazing. Marine Ecology Progress Series, 201:57–72, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p057.pdf.

LeCorre:2003:SIA

[LCL^{+}03] M. Le Corre, Y. Cherel, F. Lagarde, H. Lormée, and P. Jouventin. Seasonal and inter-annual variation in the feeding ecology of a tropical oceanic seabird, the red-tailed tropicbird Phaethon rubricauda. Marine Ecology Progress Series, 255:289–301, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p289.pdf.

Larsen:2001:PDD

[LCS^{+}01] A. Larsen, T. Castberg, R. A. Sandaa, C. P. D. Brussaard, J. Egge, M. Heldal, A. Paulino, R. Thyrhaug, E. J. van Hannen, and G. Bratbak. Population dynamics and diversity of phytoplankton, bacteria and viruses in a seawater enclosure. Marine Ecology Progress Series, 221:47–57, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p047.pdf.

Libralato:2008:NIQ

[LCT^{+}08] S. Libralato, M. Coll, S. Tudela, I. Palomera, and F. Pranovi. Novel index for quantification of ecosystem effects of fishing as removal of secondary production. Marine Ecology Progress Series, 355:107–129, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/355/m355p107.pdf.

Loh:2006:DSD

[LCW06] W. K. W. Loh, M. Cowlishaw, and N. G. Wilson. Diversity of Symbiodinium dinoflagellate symbionts from the Indo– Pacific sea slug Pteraeolidia ianthina (Gastropoda: Mollusca). Marine Ecology Progress Series, 320:177–184, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p177.pdf.

Lee:2000:ENE

[LD00] K. S. Lee and K. H. Dunton. Effects of nitrogen enrichment on biomass allocation, growth, and leaf morphology of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 196:39–48, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p039.pdf.

Liu:2003:IBN

[LD03] H. Liu and M. Dagg. Interactions between nutrients, phytoplankton growth, and micro- and mesozooplankton grazing in the plume of the Mississippi River. Marine Ecology Progress Series, 258:31–42, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p031.pdf.

Lehmann:2004:SWT

[LDHC04] M. K. Lehmann, R. F. Davis, Y. Huot, and J. J. Cullen. Spectrally weighted transparency in models of water-column photosynthesis and photoinhibition by ultraviolet radiation. Marine Ecology Progress Series, 269:101–110, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p101.pdf.

Layman:2006:MNE

[LDK^{+}06] C. A. Layman, C. P. Dahlgren, G. T. Kellison, A. J. Adams, B. M. Gillanders, M. S. Kendall, J. A. Ley, I. Nagelkerken, and J. E. Serafy. Marine nurseries and effective juvenile habitats. Marine Ecology Progress Series, 318:307–308, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p307.pdf.

Lawrence:2004:WEB

[LDL^{+}04] D. Lawrence, M. J. Dagg, H. Liu, S. R. Cummings, P. B. Ortner, and C. Kelble. Wind events and benthic-pelagic coupling in a shallow subtropical bay in Florida. Marine Ecology Progress Series, 266:1–13, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/266/m266p001.pdf.

Lobry:2008:DSE

[LDP^{+}08] J. Lobry, V. David, S. Pasquaud, M. Lepage, B. Sautour, and E. Rochard. Diversity and stability of an estuarine trophic network. Marine Ecology Progress Series, 358:13– 25, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/358/m358p013.pdf.

Luckhurst:2000:AGR

[LDR00] B. E. Luckhurst, J. M. Dean, and M. Reichert. Age, growth and reproduction of the lane snapper Lutjanus synagris (Pisces: Lutjanidae) at Bermuda. Marine Ecology Progress Series, 203:255–261, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/203/m203p255.pdf.

Levy:2004:DHC

[LDS^{+}04] O. Levy, Z. Dubinsky, K. Schneider, Y. Achituv, D. Zakai, and M. Y. Gorbunov. Diurnal hysteresis in coral photosynthesis. Marine Ecology Progress Series, 268:105–117, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p105.pdf.

Lopez-Duarte:2007:CSB

[LDT07] P. C. López-Duarte and R. A. Tankersley. Circatidal swimming behaviors of fiddler crab Uca pugilator larvae from different tidal regimes. Marine Ecology Progress Series, 343: 207–220, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p207.pdf.

Liu:2005:MCM

[LDWC05] H. Liu, M. J. Dagg, C. J. Wu, and K. P. Chiang. Mesozooplankton consumption of microplankton in the Mississippi River plume, with special emphasis on planktonic ciliates. Marine Ecology Progress Series, 286:133–144, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p133.pdf.

Lewis:2007:EPS

[LEB07] T. L. Lewis, D. Esler, and W. S. Boyd. Effects of predation by sea ducks on clam abundance in soft-bottom intertidal habitats. Marine Ecology Progress Series, 329:131– 144, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p131.pdf.

Lecchini:2005:SBP

[Lec05] D. Lecchini. Spatial and behavioural patterns of reef habitat settlement by fish larvae. Marine Ecology Progress Series, 301:247–252, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p247.pdf.

Lee:2000:CDD

[Lee00] S. Y. Lee. Carbon dynamics of Deep Bay, eastern Pearl River estuary, China. II: Trophic relationship based on carbonand nitrogen-stable isotopes. Marine Ecology Progress Series, 205:1–10, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p001.pdf.

Leathwick:2006:VDF

[LEF^{+}06] J. R. Leathwick, J. Elith, M. P. Francis, T. Hastie, and P. Taylor. Variation in demersal fish species richness in the oceans surrounding New Zealand: an analysis using boosted regression trees. Marine Ecology Progress Series, 321:267– 281, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p267.pdf.

Leising:2001:CFP

[Lei01] A. W. Leising. Copepod foraging in patchy habitats and thin layers using a 2-d individual-based model. Marine Ecology Progress Series, 216:167–179, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/216/m216p167.pdf.

Leis:2007:BIM

[Lei07] J. M. Leis. Behaviour as input for modelling dispersal of fish larvae: behaviour, biogeography, hydrodynamics, ontogeny, physiology and phylogeny meet hydrography. Marine Ecology Progress Series, 347:185–193, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m347p185.pdf.

Leinen:2008:BRB

[Lei08] M. Leinen. Building relationships between scientists and business in ocean iron fertilization. Marine Ecology Progress Series, 364:251–256, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m364p251.pdf.

Lovell:2002:RMD

[LELC02] C. R. Lovell, N. T. Eriksen, A. J. Lewitus, and Y. P. Chen. Resistance of the marine diatom Thalassiosira sp. to toxicity of phenolic compounds. Marine Ecology Progress Series, 229:11–18, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/229/m229p011.pdf.

Lehtiniemi:2002:FCT

[LEÖK^{+}02] M. Lehtiniemi, J. Engström-Öst, M. Karjalainen, B. Kozlowsky-Suzuki, and M. Viitasalo. Fate of cyanobacterial toxins in the pelagic food web: transfer to copepods or to faecal pellets? Marine Ecology Progress Series, 241:13–21, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p013.pdf.

Lesser:2000:DDP

[Les00] M. P. Lesser. Depth-dependent photoacclimatization to solar ultraviolet radiation in the Caribbean coral Montastraea faveolata. Marine Ecology Progress Series, 192:137–151, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p137.pdf.

Lesen:2005:RBB

[Les05] A. E. Lesen. Relationship between benthic Foraminifera and food resources in South San Francisco Bay, California, USA. Marine Ecology Progress Series, 297:131–145, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p131.pdf.

Levenbach:2008:BMA

[Lev08] S. Levenbach. Behavioral mechanism for an associational refuge for macroalgae on temperate reefs. Marine Ecology Progress Series, 370:45–52, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/370/m370p045.pdf.

Ley:2005:LFA

[Ley05] J. A. Ley. Linking fish assemblages and attributes of mangrove estuaries in tropical Australia: criteria for regional marine reserves. Marine Ecology Progress Series, 305:41– 57, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/305/m305p041.pdf.

Lardies:2002:EOA

[LF02] M. A. Lardies and M. Fernández. Effect of oxygen availability in determining clutch size in Acanthina monodon. Marine Ecology Progress Series, 239:139–146, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p139.pdf.

Link:2006:WPI

[LF06] J. S. Link and M. D. Ford. Widespread and persistent increase of Ctenophora in the continental shelf ecosystem off NE USA. Marine Ecology Progress Series, 320:153–159, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p153.pdf.

Lapierre:2008:AFP

[LF08] J. F. Lapierre and J. J. Frenette. Advection of freshwater phytoplankton in the St. Lawrence River estuarine turbidity maximum as revealed by sulfur-stable isotopes. Marine Ecology Progress Series, 372:19–29, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/372/m372p019.pdf.

Lauer:2009:BFN

[LFF^{+}09] P. R. Lauer, M. Fernandes, P. G. Fairweather, J. Tanner, and A. Cheshire. Benthic fluxes of nitrogen and phosphorus at southern bluefin tuna Thunnus maccoyii sea-cages. Marine Ecology Progress Series, 390:251–263, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p251.pdf.

LeGac:2004:IAC

[LFP^{+}04] M. Le Gac, J. P. Féral, E. Poulin, M. Veyret, and A. Chenuil. Identification of allopatric clades in the cosmopolitan ophiuroid species complex Amphipholis squamata (Echinodermata). The end of a paradox? Marine Ecology Progress Series, 278:171–178, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p171.pdf.

Labarta:2005:RMR

[LFRG^{+}05] U. Labarta, M. J. Fernández-Reiriz, J. L. Garrido, J. M. F. Babarro, J. M. Bayona, and J. Albaigés. Response of mussel recruits to pollution from the ‘Prestige’ oil spill along the Galicia coast. A biochemical approach. Marine Ecology Progress Series, 302:135–145, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/302/m302p135.pdf.

Lesser:2007:NFS

[LFRR^{+}07] M. P. Lesser, L. I. Falcón, A. Rodríguez-Román, S. Enríquez, O. Hoegh-Guldberg, and R. Iglesias-Prieto. Nitrogen fixation by symbiotic cyanobacteria provides a source of nitrogen for the scleractinian coral Montastraea cavernosa. Marine Ecology Progress Series, 346:143–152, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p143.pdf.

Longhi:2006:VPR

[LFSR06] M. L. Longhi, G. Ferreyra, I. Schloss, and S. Roy. Variable phytoplankton response to enhanced UV-B and nitrate addition in mesocosm experiments at three latitudes (Canada, Brazil and Argentina). Marine Ecology Progress Series, 313: 57–72, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/313/m313p057.pdf.

Larsen:2000:IAV

[LG00a] J. K. Larsen and M. Guillemette. Influence of annual variation in food supply on abundance of wintering common eiders Somateria mollissima. Marine Ecology Progress Series, 201:301–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p301.pdf.

Llewellyn:2000:ICV

[LG00b] C. A. Llewellyn and S. W. Gibb. Intra-class variability in the carbon, pigment and biomineral content of prymnesiophytes and diatoms. Marine Ecology Progress Series, 193: 33–44, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/193/m193p033.pdf.

Lesser:2001:DBC

[LG01] M. P. Lesser and M. Y. Gorbunov. Diurnal and bathymetric changes in chlorophyll fluorescence yields of reef corals measured in situ with a fast repetition rate fluorometer. Marine Ecology Progress Series, 212:69–77, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p069.pdf.

Link:2002:TEA

[LG02] J. S. Link and L. P. Garrison. Trophic ecology of Atlantic cod Gadus morhua on the northeast US continental shelf. Marine Ecology Progress Series, 227:109–123, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p109.pdf.

Leichter:2006:IUS

[LG06] J. J. Leichter and S. J. Genovese. Intermittent upwelling and subsidized growth of the scleractinian coral Madracis mirabilis on the deep fore-reef slope of Discovery Bay, Jamaica. Marine Ecology Progress Series, 316:95–103, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/316/
m316p095.pdf.

Lehtiniemi:2008:PIC

[LG08a] M. Lehtiniemi and E. Gorokhova. Predation of the introduced cladoceran Cercopagis pengoi on the native copepod Eurytemora affinis in the northern Baltic Sea. Marine Ecology Progress Series, 362:193–200, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p193.pdf.

Liuzzi:2008:MAA

[LG08b] M. G. Liuzzi and J. López Gappa. Macrofaunal assemblages associated with coralline turf: species turnover and changes in structure at different spatial scales. Marine Ecology Progress Series, 363:147–156, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/363/m363p147.pdf.

Laurel:2003:PDH

[LGB03a] B. J. Laurel, R. S. Gregory, and J. A. Brown. Predator distribution and habitat patch area determine predation rates on age-0 juvenile cod Gadus spp. Marine Ecology Progress Series, 251:245–254, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p245.pdf.

Laurel:2003:SDA

[LGB03b] B. J. Laurel, R. S. Gregory, and J. A. Brown. Settlement and distribution of age-0 juvenile cod, Gadus morhua and G. ogac, following a large-scale habitat manipulation. Marine Ecology Progress Series, 262:241–252, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/262/m262p241.pdf.

Laurel:2004:BCD

[LGB^{+}04] B. J. Laurel, R. S. Gregory, J. A. Brown, J. K. Hancock, and D. C. Schneider. Behavioural consequences of densitydependent habitat use in juvenile cod Gadus morhua and G. ogac: the role of movement and aggregation. Marine Ecology Progress Series, 272:257–270, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p257.pdf.

Lauzon-Guay:2006:MGS

[LGBWH06] J. S. Lauzon-Guay, M. A. Barbeau, J. Watmough, and D. J. Hamilton. Model for growth and survival of mussels Mytilus edulis reared in Prince Edward Island, Canada. Marine Ecology Progress Series, 323:171–183, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p171.pdf.

Lea:2006:ICA

[LGC^{+}06] M. A. Lea, C. Guinet, Y. Cherel, G. Duhamel, L. Dubroca, P. Pruvost, and M. Hindell. Impacts of climatic anomalies on provisioning strategies of a Southern Ocean predator. Marine Ecology Progress Series, 310:77–94, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p077.pdf.

Lea:2008:CBF

[LGC^{+}08] M. A. Lea, C. Guinet, Y. Cherel, M. Hindell, L. Dubroca, and S. Thalmann. Colony-based foraging segregation by Antarctic fur seals at the Kerguelen Archipelago. Marine Ecology Progress Series, 358:273–287, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/358/m358p273.pdf.

LePape:2004:EIM

[LGD04] O. Le Pape, D. Guérault, and Y. Désaunay. Effect of an invasive mollusc, American slipper limpet Crepidula fornicata, on habitat suitability for juvenile common sole Solea solea in the Bay of Biscay. Marine Ecology Progress Series, 277:107–115, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/277/m277p107.pdf.

Linke-Gamenick:2000:ECF

[LGFM00] I. Linke-Gamenick, V. E. Forbes, and N. Méndez. Effects of chronic fluoranthene exposure on sibling species of Capitella with different development modes. Marine Ecology Progress Series, 203:191–203, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/203/m203p191.pdf.

Lorrain:2009:NCI

[LGM^{+}09] A. Lorrain, B. Graham, F. Ménard, B. Popp, S. Bouillon, P. van Breugel, and Y. Cherel. Nitrogen and carbon isotope values of individual amino acids: a tool to study foraging ecology of penguins in the Southern Ocean. Marine Ecology Progress Series, 391:293–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m391p293.pdf.

Lucifora:2006:FHS

[LGME06] L. O. Lucifora, V. B. García, R. C. Menni, and A. H. Escalante. Food habits, selectivity, and foraging modes of the school shark Galeorhinus galeus. Marine Ecology Progress Series, 315:259–270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p259.pdf.

Lawrence:2000:CGT

[LGQ^{+}00] J. E. Lawrence, J. Grant, M. A. Quilliam, A. G. Bauder, and A. D. Cembella. Colonization and growth of the toxic dinoflagellate Prorocentrum lima and associated fouling macroalgae on mussels in suspended culture. Marine Ecology Progress Series, 201:147–154, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p147.pdf.

Lasker:2008:MRS

[LGRB^{+}08] H. R. Lasker, C. Gutiérrez-Rodríguez, K. Bala, A. Hannes, and J. P. Bilewitch. Male reproductive success during spawning events of the octocoral Pseudopterogorgia elisabethae. Marine Ecology Progress Series, 367:153–161, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/367/
m367p153.pdf.

Lauzon-Guay:2007:BSU

[LGS07] J. S. Lauzon-Guay and R. E. Scheibling. Behaviour of sea urchin Strongylocentrotus droebachiensis grazing fronts: foodmediated aggregation and density-dependent facilitation. Marine Ecology Progress Series, 329:191–204, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p191.pdf.

Lauzon-Guay:2009:MPS

[LGSB09] J. S. Lauzon-Guay, R. E. Scheibling, and M. A. Barbeau. Modelling phase shifts in a rocky subtidal ecosystem. Marine Ecology Progress Series, 375:25–39, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p025.pdf.

Linke-Gamenick:2000:EFA

[LGVF00] I. Linke-Gamenick, B. Vismann, and V. E. Forbes. Effects of fluoranthene and ambient oxygen levels on survival and metabolism in three sibling species of Capitella (Polychaeta). Marine Ecology Progress Series, 194:169–177, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p169.pdf.

Li:2009:RBN

[LGZ^{+}09] J. Li, P. M. Glibert, M. Zhou, S. Lu, and D. Lu. Relationships between nitrogen and phosphorus forms and ratios and the development of dinoflagellate blooms in the East China Sea. Marine Ecology Progress Series, 383:11–26, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m383p011.pdf.

Levin:2002:FSA

[LH02] P. S. Levin and M. E. Hay. Fish-seaweed association on temperate reefs: do small-scale experiments predict largescale patterns? Marine Ecology Progress Series, 232:239– 246, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/232/m232p239.pdf.

Long:2006:FLA

[LH06] J. D. Long and M. E. Hay. Fishes learn aversions to a nudibranch’s chemical defense. Marine Ecology Progress Series, 307:199–208, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p199.pdf.

Lester:2008:BRM

[LH08] S. E. Lester and B. S. Halpern. Biological responses in marine no-take reserves versus partially protected areas. Marine Ecology Progress Series, 367:49–56, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/367/
m367p049.pdf.

Lavaniegos:2009:SVH

[LH09] B. E. Lavaniegos and C. M. Hereu. Seasonal variation in hyperiid amphipod abundance and diversity and influence of mesoscale structures off Baja California. Marine Ecology Progress Series, 394:137–152, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p137.pdf.

Lough:2006:SPV

[LHB^{+}06] R. G. Lough, C. G. Hannah, P. Berrien, D. Brickman, J. W. Loder, and J. A. Quinlan. Spawning pattern variability and its effect on retention, larval growth and recruitment in Georges Bank cod and haddock. Marine Ecology Progress Series, 310:193–212, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/310/m310p193.pdf.

Lefebvre:2009:TTC

[LHB09] S. Lefebvre, C. Harma, and J. L. Blin. Trophic typology of coastal ecosystems based on δ ^{13}C and δ ^{15}N ratios in an opportunistic suspension feeder. Marine Ecology Progress Series, 390:27–37, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/390/m390p027.pdf.

Labry:2001:IWP

[LHD^{+}01] C. Labry, A. Herbland, D. Delmas, P. Laborde, P. Lazure, J. M. Froidefond, A. M. Jegou, and B. Sautour. Initiation of winter phytoplankton blooms within the Gironde plume waters in the Bay of Biscay. Marine Ecology Progress Series, 212:117–130, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p117.pdf.

Lester:2009:BEW

[LHGC^{+}09] S. E. Lester, B. S. Halpern, K. Grorud-Colvert, J. Lubchenco, B. I. Ruttenberg, S. D. Gaines, S. Airamé, and R. R. Warner. Biological effects within no-take marine reserves: a global synthesis. Marine Ecology Progress Series, 384: 33–46, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p033.pdf.

Leis:2009:OLS

[LHH09] J. M. Leis, A. C. Hay, and G. J. Howarth. Ontogeny of in situ behaviours relevant to dispersal and population connectivity in larvae of coral-reef fishes. Marine Ecology Progress Series, 379:163–179, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/379/m379p163.pdf.

Lynam:2005:EIJ

[LHHB05] C. P. Lynam, M. R. Heath, S. J. Hay, and A. S. Brierley. Evidence for impacts by jellyfish on North Sea herring recruitment. Marine Ecology Progress Series, 298:157– 167, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p157.pdf.

Laidre:2003:DDN

[LHJD^{+}03] K. L. Laidre, M. P. Heide-Jørgensen, R. Dietz, R. C. Hobbs, and O. A. Jørgensen. Deep-diving by narwhals Monodon monoceros: differences in foraging behavior between wintering areas? Marine Ecology Progress Series, 261:269– 281, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/261/m261p269.pdf.

Laidre:2007:RBW

[LHJN07] K. L. Laidre, M. P. Heide-Jørgensen, and T. G. Nielsen. Role of the bowhead whale as a predator in West Greenland. Marine Ecology Progress Series, 346:285–297, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p285.pdf.

Lesage:2001:MMC

[LHK01] V. Lesage, M. O. Hammill, and K. M. Kovacs. Marine mammals and the community structure of the Estuary and Gulf of St Lawrence, Canada: evidence from stable isotope analysis. Marine Ecology Progress Series, 210:203–221, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
210/m210p203.pdf.

Lundholm:2004:EPG

[LHK04] N. Lundholm, P. J. Hansen, and Y. Kotaki. Effect of pH on growth and domoic acid production by potentially toxic diatoms of the genera Pseudo-nitzschia and Nitzschia. Marine Ecology Progress Series, 273:1–15, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/273/
m273p001.pdf.

Lundholm:2005:LAE

[LHK05] N. Lundholm, P. J. Hansen, and Y. Kotaki. Lack of allelopathic effects of the domoic acid-producing marine diatom Pseudo-nitzschia multiseries. Marine Ecology Progress Series, 288:21–33, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/288/m288p021.pdf.

Lee:2006:LSM

[LHK06] R. F. Lee, W. Hagen, and G. Kattner. Lipid storage in marine zooplankton. Marine Ecology Progress Series, 307: 273–306, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p273.pdf.

Leis:2007:OSS

[LHL^{+}07] J. M. Leis, A. C. Hay, M. M. Lockett, J. Chen, and L. Fang. Ontogeny of swimming speed in larvae of pelagic-spawning, tropical, marine fishes. Marine Ecology Progress Series, 349: 255–267, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/349/m349p255.pdf.

Lam:2003:ILS

[LHQ03] C. Lam, T. Harder, and P. Y. Qian. Induction of larval settlement in the polychaete Hydroides elegans by surfaceassociated settlement cues of marine benthic diatoms. Marine Ecology Progress Series, 263:83–92, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p083.pdf.

Lam:2005:GCB

[LHQ05a] C. Lam, T. Harder, and P. Y. Qian. Growth conditions of benthic diatoms affect quality and quantity of extracellular polymeric larval settlement cues. Marine Ecology Progress Series, 294:109–116, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/294/m294p109.pdf.

Lam:2005:ILS

[LHQ05b] C. Lam, T. Harder, and P. Y. Qian. Induction of larval settlement in the polychaete Hydroides elegans by extracellular polymers of benthic diatoms. Marine Ecology Progress Series, 286:145–154, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p145.pdf.

Ley:2002:EEF

[LHT^{+}02] J. A. Ley, I. A. Halliday, A. J. Tobin, R. N. Garrett, and N. A. Gribble. Ecosystem effects of fishing closures in mangrove estuaries of tropical Australia. Marine Ecology Progress Series, 245:223–238, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/245/m245p223.pdf.

Lohrer:2006:AFF

[LHT06] A. M. Lohrer, J. E. Hewitt, and S. F. Thrush. Assessing far-field effects of terrigenous sediment loading in the coastal marine environment. Marine Ecology Progress Series, 315:13–18, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p013.pdf.

Lidgard:2008:PMB

[Lid08] S. Lidgard. Predation on marine bryozoan colonies: taxa, traits and trophic groups. Marine Ecology Progress Series, 359:117–131, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p117.pdf.

Louzao:2006:SPF

[LIM^{+}06] M. Louzao, J. M. Igual, M. McMinn, J. S. Aguilar, R. Triay, and D. Oro. Small pelagic fish, trawling discards and breeding performance of the critically endangered Balearic shearwater: improving conservation diagnosis. Marine Ecology Progress Series, 318:247–254, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/318/m318p247.pdf.

Link:2002:DFW

[Lin02] J. Link. Does food web theory work for marine ecosystems? Marine Ecology Progress Series, 230:1–9, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p001.pdf.

Link:2004:UFS

[Lin04] J. S. Link. Using fish stomachs as samplers of the benthos: integrating long-term and broad scales. Marine Ecology Progress Series, 269:265–275, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/269/m269p265.pdf.

Lirman:2000:LRB

[Lir00] D. Lirman. Lesion regeneration in the branching coral Acropora palmata: effects of colonization, colony size, lesion size, and lesion shape. Marine Ecology Progress Series, 197:209– 215, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/197/m197p209.pdf.

Ling:2005:DSC

[LIT05] Y. J. Ling, Y. Iizuka, and W. N. Tzeng. Decreased Sr/Ca ratios in the otoliths of two marine eels, Gymnothorax reticularis and Muraenesox cinereus, during metamorphosis. Marine Ecology Progress Series, 304:201–206, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p201.pdf.

Lee:2006:SPN

[LJ06] R. Y. Lee and S. B. Joye. Seasonal patterns of nitrogen fixation and denitrification in oceanic mangrove habitats. Marine Ecology Progress Series, 307:127–141, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/307/
m307p127.pdf.

Lamote:2008:TSV

[LJ08] M. Lamote and L. E. Johnson. Temporal and spatial variation in the early recruitment of fucoid algae: the role of microhabitats and temporal scales. Marine Ecology Progress Series, 368:93–102, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p093.pdf.

Ling:2009:PDE

[LJ09] S. D. Ling and C. R. Johnson. Population dynamics of an ecologically important range-extender: kelp beds versus sea urchin barrens. Marine Ecology Progress Series, 374:113– 125, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/374/m374p113.pdf.

Luna-Jorquera:2000:MRS

[LJC00] G. Luna-Jorquera and B. M. Culik. Metabolic rates of swimming Humboldt penguins. Marine Ecology Progress Series, 203:301–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/203/m203p301.pdf.

Liu:2001:BMP

[LJK01a] G. Liu, G. S. Janowitz, and D. Kamykowski. A biophysical model of population dynamics of the autotrophic dinoflagellate Gymnodinium breve. Marine Ecology Progress Series, 210:101–124, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p101.pdf.

Liu:2001:IEN

[LJK01b] G. Liu, G. S. Janowitz, and D. Kamykowski. Influence of environmental nutrient conditions on Gymnodinium breve (Dinophyceae) population dynamics: a numerical study. Marine Ecology Progress Series, 213:13–37, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/213/m213p013.pdf.

Liu:2002:ICS

[LJK02] G. Liu, G. S. Janowitz, and D. Kamykowski. Influence of current shear on Gymnodinium breve (Dinophyceae) population dynamics: a numerical study. Marine Ecology Progress Series, 231:47–66, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p047.pdf.

Levin:2000:DMS

[LJM^{+}00] L. A. Levin, D. W. James, C. M. Martin, A. E. Rathburn, L. H. Harris, and R. H. Michener. Do methane seeps support distinct macrofaunal assemblages? Observations on community structure and nutrition from the northern California slope and shelf. Marine Ecology Progress Series, 208:21–39, 2000.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
208/m208p021.pdf.

Levesque:2003:FRP

[LJM03] C. Levesque, S. K. Juniper, and J. Marcus. Food resource partitioning and competition among alvinellid polychaetes of Juan de Fuca Ridge hydrothermal vents. Marine Ecology Progress Series, 246:173–182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/246/m246p173.pdf.

Ludeking:2002:ISM

[LK02] A. Lüdeking and A. Köhler. Identification of six mRNA sequences of genes related to multixenobiotic resistance (MXR) and biotransformation in Mytilus edulis. Marine Ecology Progress Series, 238:115–124, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/238/m238p115.pdf.

Laroche:2007:EPE

[LKDO07] R. K. Laroche, A. A. Kock, L. M. Dill, and W. H. Oosthuizen. Effects of provisioning ecotourism activity on the behaviour of white sharks Carcharodon carcharias. Marine Ecology Progress Series, 338:199–209, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/338/m338p199.pdf.

Lassen:2006:MPA

[LKR^{+}06] J. Lassen, M. Kortegård, H. U. Riisgård, M. Friedrichs, G. Graf, and P. S. Larsen. Down-mixing of phytoplankton above filter-feeding mussels — interplay between water flow and biomixing. Marine Ecology Progress Series, 314: 77–88, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p077.pdf.

Loya:2003:LSE

[LKW03] Y. Loya and E. Kramarsky-Winter. In situ eutrophication caused by fish farms in the northern Gulf of Eilat (Aqaba) is beneficial for its coral reefs: a critique. Marine Ecology Progress Series, 261:299–303, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p299.pdf.

Lambert:2003:PDD

[LL03] C. C. Lambert and G. Lambert. Persistence and differential distribution of nonindigenous ascidians in harbors of the Southern California Bight. Marine Ecology Progress Series, 259:145–161, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p145.pdf.

Loh:2001:GVS

[LLCHG01] W. K. W. Loh, T. Loi, D. Carter, and O. Hoegh-Guldberg. Genetic variability of the symbiotic dinoflagellates from the wide ranging coral species Seriatopora hystrix and Acropora longicyathus in the Indo-West Pacific. Marine Ecology Progress Series, 222:97–107, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p097.pdf.

Longphuirt:2006:DMM

[LLG^{+}06] S. Ni Longphuirt, A. Leynaert, J. M. Guarini, L. Chauvaud, P. Claquin, O. Herlory, E. Amice, P. Huonnic, and O. Ragueneau. Discovery of microphytobenthos migration in the subtidal zone. Marine Ecology Progress Series, 328: 143–154, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p143.pdf.

Lambert:2006:AEB

[LLH06] D. M. Lambert, R. N. Lipcius, and J. M. Hoenig. Assessing effectiveness of the blue crab spawning stock sanctuary in Chesapeake Bay using tag-return methodology. Marine Ecology Progress Series, 321:215–225, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p215.pdf.

Levesque:2005:OCN

[LLJ05] C. Levesque, H. Limén, and S. K. Juniper. Origin, composition and nutritional quality of particulate matter at deepsea hydrothermal vents on Axial Volcano, NE Pacific. Marine Ecology Progress Series, 289:43–52, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p043.pdf.

Longphuirt:2009:DIN

[LLL^{+}09] S. Ní Longphuirt, J. H. Lim, A. Leynaert, P. Claquin, E. J. Choy, C. K. Kang, and S. An. Dissolved inorganic nitrogen uptake by intertidal microphytobenthos: nutrient concentrations, light availability and migration. Marine Ecology Progress Series, 379:33–44, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/379/m379p033.pdf.

Lambardi:2008:IOC

[LLM^{+}08] P. Lambardi, J. R. E. Lutjeharms, R. Mencacci, G. C. Hays, and P. Luschi. Influence of ocean currents on longdistance movement of leatherback sea turtles in the southwest Indian Ocean. Marine Ecology Progress Series, 353: 289–301, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/353/m353p289.pdf.

Lopez-Legentil:2005:LCG

[LLRDT05] S. López-Legentil, M. Ruchty, A. Domenech, and X. Turon. Life cycles and growth rates of two morphotypes of Cystodytes (Ascidiacea) in the western Mediterranean. Marine Ecology Progress Series, 296:219–228, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/296/m296p219.pdf.

Lehtonen:2006:BPE

[LLSL06] K. K. Lehtonen, S. Leiniö, R. Schneider, and M. Leivuori. Biomarkers of pollution effects in the bivalves Mytilus edulis and Macoma balthica collected from the southern coast of Finland (Baltic Sea). Marine Ecology Progress Series, 322: 155–168, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p155.pdf.

Love:2009:STP

[LLW09] J. W. Love, D. F. Luers, and B. D. Williams. Spatio-temporal patterns of larval fish ingress to Chincoteague Bay, Maryland, USA during winter and spring 2004 to 2007. Marine Ecology Progress Series, 377:203–212, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/377/m377p203.pdf.

Lee:2001:IRS

[LLY^{+}01] J. S. Lee, B. G. Lee, H. Yoo, C. H. Koh, and S. N. Luoma. Influence of reactive sulfide (AVS) and supplementary food on Ag, Cd and Zn bioaccumulation in the marine polychaete Neanthes arenaceodentata. Marine Ecology Progress Series, 216:129–140, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/216/m216p129.pdf.

Larsen:2009:ITC

[LM09] M. H. Larsen and K. N. Mouritsen. Increasing temperature counteracts the impact of parasitism on periwinkle consumption. Marine Ecology Progress Series, 383:141–149, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/383/m383p141.pdf.

Lundsten:2009:ITS

[LMB^{+}09] L. Lundsten, C. R. McClain, J. P. Barry, G. M. Cailliet, D. A. Clague, and A. P. DeVogelaere. Ichthyofauna on three seamounts off southern and central California, USA. Marine Ecology Progress Series, 389:223–232, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p223.pdf.

Lardies:2008:IBP

[LMC08] M. A. Lardies, M. H. Medina, and J. A. Correa. Intraspecific biogeographic pattern breakage in the snapping shrimp Betaeus emarginatus caused by coastal copper mine tailings. Marine Ecology Progress Series, 358:203–210, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p203.pdf.

Lau:2002:BBS

[LMCQ02] S. C. K. Lau, K. K. W. Mak, F. Chen, and P. Y. Qian. Bioactivity of bacterial strains isolated from marine biofilms in Hong Kong waters for the induction of larval settlement in the marine polychaete Hydroides elegans. Marine Ecology Progress Series, 226:301–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/226/m226p301.pdf.

Lepoint:2002:ANB

[LMD^{+}02] G. Lepoint, S. Millet, P. Dauby, S. Gobert, and J. M. Bouquegneau. Annual nitrogen budget of the seagrass Posidonia oceanica as determined by in situ uptake experiments. Marine Ecology Progress Series, 237:87–96, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p087.pdf.

Lucifora:2005:RAF

[LME05] L. O. Lucifora, R. C. Menni, and A. H. Escalante. Reproduction, abundance and feeding habits of the broadnose sevengill shark Notorynchus cepedianus in north Patagonia, Argentina. Marine Ecology Progress Series, 289:237–244, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p237.pdf.

Lopez:2001:TDG

[LMF^{+}01] J. L. López, E. Mosquera, J. Fuentes, A. Marina, J. Vázquez, and G. Alvarez. Two-dimensional gel electrophoresis of Mytilus galloprovincialis differences in protein expression between intertidal and cultured mussels. Marine Ecology Progress Series, 224:149–156, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/224/m224p149.pdf.

Lydersen:2001:SAM

[LMKG01] C. Lydersen, A. R. Martin, K. M. Kovacs, and I. Gjertz. Summer and autumn movements of white whales Delphinapterus leucas in Svalbard, Norway. Marine Ecology Progress Series, 219:265–274, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p265.pdf.

Lavery:2009:IET

[LMMT09] P. S. Lavery, K. McMahon, M. Mulligan, and A. Tennyson. Interactive effects of timing, intensity and duration of experimental shading on Amphibolis griffithii. Marine Ecology Progress Series, 394:21–33, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p021.pdf.

Laureillard:2004:ULS

[LMS04] J. Laureillard, L. Méjanelle, and M. Sibuet. Use of lipids to study the trophic ecology of deep-sea xenophyophores. Marine Ecology Progress Series, 270:129–140, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/270/
m270p129.pdf.

Leung:2001:GRF

[LMW^{+}01] K. M. Y. Leung, I. J. Morgan, R. S. S. Wu, T. C. Lau, J. Svavarsson, and R. W. Furness. Growth rate as a factor confounding the use of the dogwhelk Nucella lapillus as biomonitor of heavy metal contamination. Marine Ecology Progress Series, 221:145–159, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/221/m221p145.pdf.

Lenz:2004:TID

[LMW04] M. Lenz, M. Molis, and M. Wahl. Testing the intermediate disturbance hypothesis: response of fouling communities to various levels of emersion intensity. Marine Ecology Progress Series, 278:53–65, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p053.pdf.

Litchman:2005:UEP

[LN05] E. Litchman and P. J. Neale. UV effects on photosynthesis, growth and acclimation of an estuarine diatom and cryptomonad. Marine Ecology Progress Series, 300:53–62, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/300/m300p053.pdf.

Leguerrier:2003:NAF

[LNB^{+}03] D. Leguerrier, N. Niquil, N. Boileau, J. Rzeznik, P. G. Sauriau, O. Le Moine, and C. Bacher. Numerical analysis of the food web of an intertidal mudflat ecosystem on the Atlantic coast of France. Marine Ecology Progress Series, 246:17–37, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/246/m246p017.pdf.

Levinsen:2000:ASM

[LNH00] H. Levinsen, T. G. Nielsen, and B. W. Hansen. Annual succession of marine pelagic protozoans in Disko Bay, West Greenland, with emphasis on winter dynamics. Marine Ecology Progress Series, 206:119–134, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p119.pdf.

Langer:2007:DCL

[LNJ07] G. Langer, G. Nehrke, and S. Jansen. Dissolution of Calcidiscus leptoporus coccoliths in copepod guts? A morphological study. Marine Ecology Progress Series, 331:139– 146, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/331/m331p139.pdf.

Leguerrier:2004:MIO

[LNPB04] D. Leguerrier, N. Niquil, A. Petiau, and A. Bodoy. Modeling the impact of oyster culture on a mudflat food web in Marennes–Oléron Bay (France). Marine Ecology Progress Series, 273:147–162, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p147.pdf.

LaPeyre:2003:ESF

[LNV^{+}03] M. K. La Peyre, A. D. Nickens, A. K. Volety, G. S. Tolley, and J. F. La Peyre. Environmental significance of freshets in reducing Perkinsus marinus infection in eastern oysters Crassostrea virginica: potential management applications. Marine Ecology Progress Series, 248:165–176, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/248/m248p165.pdf.

Lagos:2005:MSS

[LNV^{+}05] N. A. Lagos, S. A. Navarrete, F. Véliz, A. Masuero, and J. C. Castilla. Meso-scale spatial variation in settlement and recruitment of intertidal barnacles along the coast of central Chile. Marine Ecology Progress Series, 290:165– 178, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/290/m290p165.pdf.

Lewis:2003:PEC

[LOB03] C. Lewis, P. J. W. Olive, and M. G. Bentley. Pre-emptive competition as a selective pressure for early reproduction in the polychaete Nereis virens. Marine Ecology Progress Series, 254:199–211, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p199.pdf.

Lopez:2007:APM

[Lóp07] J. L. López. Applications of proteomics in marine ecology. Marine Ecology Progress Series, 332:275–279, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m332p275.pdf.

Landry:2000:BRIa

[LOT^{+}00] M. R. Landry, M. E. Ondrusek, S. J. Tanner, S. L. Brown, J. Constantinou, R. R. Bidigare, K. H. Coale, and S. Fitzwater. Biological response to iron fertilization in the eastern equatorial Pacific (IronEx II). I. Microplankton community abundances and biomass. Marine Ecology Progress Series, 201:27–42, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p027.pdf.

Lough:2008:CCS

[Lou08] J. M. Lough. Coral calcification from skeletal records revisited. Marine Ecology Progress Series, 373:257–264, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p257.pdf.

Lawrence:2001:ICP

[LP01a] A. J. Lawrence and C. Poulter. Impact of copper, pentachlorophenol and benzo[a]pyrene on the swimming efficiency and embryogenesis of the amphipod Chaetogammarus marinus. Marine Ecology Progress Series, 223:213–223, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
223/m223p213.pdf.

Lisbjerg:2001:FAR

[LP01b] D. Lisbjerg and J. K. Petersen. Feeding activity, retention efficiency, and effects of temperature and particle concentration on clearance rate in the marine bryozoan Electra crustulenta. Marine Ecology Progress Series, 215:133–141, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p133.pdf.

Lloret:2003:CFR

[LP03] J. Lloret and S. Planes. Condition, feeding and reproductive potential of white seabream Diplodus sargus as indicators of habitat quality and the effect of reserve protection in the northwestern Mediterranean. Marine Ecology Progress Series, 248:197–208, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/248/m248p197.pdf.

Lombarte:2004:QCO

[LP04] A. Lombarte and A. N. Popper. Quantitative changes in the otolithic organs of the inner ear during the settlement period in European hake Merluccius merluccius. Marine Ecology Progress Series, 267:233–240, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/267/m267p233.pdf.

Li:2007:EIR

[LP07] W. Li and J. A. Pechenik. Effect of inbreeding on reproduction and juvenile performance in two marine gastropods with contrasting reproductive patterns. Marine Ecology Progress Series, 346:219–234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/346/m346p219.pdf.

Loret:2000:PCS

[LPBD00] P. Loret, A. Pastoureaud, C. Bacher, and B. Delesalle. Phytoplankton composition and selective feeding of the pearl oyster Pinctada margaritifera in the Takapoto lagoon (Tuamotu Archipelago, French Polynesia): in situ study using optical microscopy and HPLC pigment analysis. Marine Ecology Progress Series, 199:55–67, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/199/m199p055.pdf.

Leduc:2009:MMA

[LPD09] D. Leduc, P. K. Probert, and A. Duncan. A multi-method approach for identifying meiofaunal trophic connections. Marine Ecology Progress Series, 383:95–111, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p095.pdf.

Limborg:2009:GPS

[LPHH^{+}09] M. T. Limborg, J. S. Pedersen, J. Hemmer-Hansen, J. Tomkiewicz, and D. Bekkevold. Genetic population structure of European sprat Sprattus sprattus: differentiation across a steep environmental gradient in a small pelagic fish. Marine Ecology Progress Series, 379:213–224, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p213.pdf.

Lenihan:2003:VMB

[LPK^{+}03] H. S. Lenihan, C. H. Peterson, S. L. Kim, K. E. Conlan, R. Fairey, C. McDonald, J. H. Grabowski, and J. S. Oliver. Variation in marine benthic community composition allows discrimination of multiple stressors. Marine Ecology Progress Series, 261:63–73, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p063.pdf.

Lee:2007:RZM

[LPK^{+}07] K. S. Lee, J. I. Park, Y. K. Kim, S. R. Park, and J. H. Kim. Recolonization of Zostera marina following destruction caused by a red tide algal bloom: the role of new shoot recruitment from seed banks. Marine Ecology Progress Series, 342:105–115, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/342/m342p105.pdf.

Lopes:2006:IMM

[LPM^{+}06] R. J. Lopes, M. A. Pardal, T. Múrias, J. A. Cabral, and J. C. Marques. Influence of macroalgal mats on abundance and distribution of dunlin Calidris alpina in estuaries: a long-term approach. Marine Ecology Progress Series, 323: 11–20, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p011.pdf.

Lindsay:2008:HSM

[LPS08] M. J. Lindsay, H. M. Patterson, and S. E. Swearer. Habitat as a surrogate measure of reef fish diversity in the zoning of the Lord Howe Island Marine Park, Australia. Marine Ecology Progress Series, 353:265–273, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/353/m353p265.pdf.

Leandro:2006:TDD

[LQRGT06] S. M. Leandro, H. Queiroga, L. Rodríguez-Graña, and P. Tiselius. Temperature-dependent development and somatic growth in two allopatric populations of Acartia clausi (Copepoda: Calanoida). Marine Ecology Progress Series, 322:189–197, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p189.pdf.

Legendre:2002:FCU

[LR02] L. Legendre and R. B. Rivkin. Fluxes of carbon in the upper ocean: regulation by food-web control nodes. Marine Ecology Progress Series, 242:95–109, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/242/m242p095.pdf.

Lemke:2006:RPV

[LR06] J. L. Lemke and C. H. Ryer. Relative predation vulnerability of three juvenile (age-0) North Pacific flatfish species: possible influence of nursery-specific predation pressures. Marine Ecology Progress Series, 328:267–273, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p267.pdf.

Legendre:2008:PFW

[LR08] L. Legendre and R. B. Rivkin. Planktonic food webs: microbial hub approach. Marine Ecology Progress Series, 365: 289–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m365p289.pdf.

Llanos-Rivera:2006:IPD

[LRC06] A. Llanos-Rivera and L. R. Castro. Inter-population differences in temperature effects on Engraulis ringens yolksac larvae. Marine Ecology Progress Series, 312:245–253, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/312/m312p245.pdf.

Lombarte:2000:SST

[LRGdS00] A. Lombarte, L. Recasens, M. González, and L. G. de Sola. Spatial segregation of two species of Mullidae (Mullus surmuletus and M. barbatus) in relation to habitat. Marine Ecology Progress Series, 206:239–249, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/206/m206p239.pdf.

Lavery:2007:ELM

[LRHV07] P. S. Lavery, T. Reid, G. A. Hyndes, and B. R. Van Elven. Effect of leaf movement on epiphytic algal biomass of seagrass leaves. Marine Ecology Progress Series, 338:97– 106, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p097.pdf.

Lane:2009:DLR

[LRK09] J. Q. Lane, P. T. Raimondi, and R. M. Kudela. Development of a logistic regression model for the prediction of toxigenic Pseudo-nitzschia blooms in Monterey Bay, California. Marine Ecology Progress Series, 383:37–51, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m383p037.pdf.

Lourey:2000:RDR

[LRM00] M. J. Lourey, D. A. J. Ryan, and I. R. Miller. Rates of decline and recovery of coral cover on reefs impacted by, recovering from and unaffected by crown-of-thorns starfish Acanthaster planci: a regional perspective of the Great Barrier Reef. Marine Ecology Progress Series, 196:179–186, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p179.pdf.

Lee:2009:EIP

[LRN^{+}09] P. A. Lee, J. R. Rudisill, A. R. Neeley, J. M. Maucher, D. A. Hutchins, Y. Feng, C. E. Hare, K. Leblanc, J. M. Rose, S. W. Wilhelm, J. M. Rowe, and G. R. DiTullio. Effects of increased pCO_{2} a and temperature on the North Atlantic spring bloom. III. Dimethylsulfoniopropionate. Marine Ecology Progress Series, 388:41–49, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p041.pdf.

Laruelle:2009:BPC

[LRR^{+}09] G. G. Laruelle, P. Regnier, O. Ragueneau, M. Kempa, B. Moriceau, S. Ni Longphuirt, A. Leynaert, G. Thouzeau, and L. Chauvaud. Benthic-pelagic coupling and the seasonal silica cycle in the Bay of Brest (France): new insights from a coupled physical-biological model. Marine Ecology Progress Series, 385:15–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m385p015.pdf.

Lipcius:2002:CDS

[LS02] R. N. Lipcius and W. T. Stockhausen. Concurrent decline of the spawning stock, recruitment, larval abundance, and size of the blue crab Callinectes sapidus in Chesapeake Bay. Marine Ecology Progress Series, 226:45–61, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p045.pdf.

Lowry:2004:PSA

[LS04] M. Lowry and I. Suther. Population structure of aggregations, and response to spear fishing, of a large temperate reef fish Cheilodactylus fuscus. Marine Ecology Progress Series, 273:199–210, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p199.pdf.

Lartigue:2005:RES

[LS05] J. Lartigue and T. D. Sherman. Response of Enteromorpha sp. (Chlorophyceae) to a nitrate pulse: nitrate uptake, inorganic nitrogen storage and nitrate reductase activity. Marine Ecology Progress Series, 292:147–157, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p147.pdf.

Loher:2006:SME

[LS06] T. Loher and A. Seitz. Seasonal migration and environmental conditions of Pacific halibut Hippoglossus stenolepis, elucidated from pop-up archival transmitting (PAT) tags. Marine Ecology Progress Series, 317:259–271, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p259.pdf.

Larson:2008:RMR

[LS08a] F. Larson and K. Sundbäck. Role of microphytobenthos in recovery of functions in a shallow-water sediment system after hypoxic events. Marine Ecology Progress Series, 357:1–16, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Long:2008:TIU

[LS08b] W. C. Long and R. D. Seitz. Trophic interactions under stress: hypoxia enhances foraging in an estuarine food web. Marine Ecology Progress Series, 362:59–68, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p059.pdf.

Larson:2009:CDS

[LS09a] A. A. Larson and J. J. Stachowicz. Chemical defense of a softsediment dwelling phoronid against local epibenthic predators. Marine Ecology Progress Series, 374:101–111, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p101.pdf.

Lin:2009:SBT

[LS09b] Y. Lin and K. Soong. Spawning behaviors of two congeneric brittle stars. Marine Ecology Progress Series, 387: 157–165, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/387/m387p157.pdf.

Limen:2008:TES

[LSBJ08] H. Limén, C. J. Stevens, Z. Bourass, and S. K. Juniper. Trophic ecology of siphonostomatoid copepods at deep-sea hydrothermal vents in the northeast Pacific. Marine Ecology Progress Series, 359:161–170, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/359/m359p161.pdf.

Lee:2003:ICG

[LSdM^{+}03] P. A. Lee, P. A. Saunders, S. J. de Mora, D. Deibel, and M. Levasseur. Influence of copepod grazing on concentrations of dissolved dimethylsulfoxide and related sulfur compounds in the North Water, northern Baffin Bay. Marine Ecology Progress Series, 255:235–248, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/255/m255p235.pdf.

Leterme:2006:DCD

[LSE06] S. C. Leterme, L. Seuront, and M. Edwards. Differential contribution of diatoms and dinoflagellates to phytoplankton biomass in the NE Atlantic Ocean and the North Sea. Marine Ecology Progress Series, 312:57–65, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p057.pdf.

Lancelot:2005:MDP

[LSG^{+}05a] C. Lancelot, Y. Spitz, N. Gypens, K. Ruddick, S. Becquevort, V. Rousseau, G. Lacroix, and G. Billen. Modelling diatom and Phaeocystis blooms and nutrient cycles in the Southern Bight of the North Sea: the MIRO model. Marine Ecology Progress Series, 289:63–78, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p063.pdf.

Lombard:2005:IBM

[LSG05b] F. Lombard, A. Sciandra, and G. Gorsky. Influence of body mass, food concentration, temperature and filtering activity on the oxygen uptake of the appendicularian Oikopleura dioica. Marine Ecology Progress Series, 301:149– 158, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/301/m301p149.pdf.

Leichter:2008:DWM

[LSG08] J. J. Leichter, M. D. Stokes, and S. J. Genovese. Deep water macroalgal communities adjacent to the Florida Keys reef tract. Marine Ecology Progress Series, 356:123–138, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/356/m356p123.pdf.

Layman:2000:SVI

[LSH00] C. A. Layman, D. E. Smith, and J. D. Herod. Seasonally varying importance of abiotic and biotic factors in marshpond fish communities. Marine Ecology Progress Series, 207: 155–169, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/207/m207p155.pdf.

Lagus:2007:IEP

[LSH^{+}07a] A. Lagus, J. Suomela, H. Helminen, J. M. Lehtimäki, J. Sipura, K. Sivonen, and L. Suominen. Interaction effects of N: P ratios and frequency of nutrient supply on the plankton community in the northern Baltic Sea. Marine Ecology Progress Series, 332:77–92, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p077.pdf.

Laurel:2007:DDH

[LSH07b] B. J. Laurel, A. W. Stoner, and T. P. Hurst. Densitydependent habitat selection in marine flatfish: the dynamic role of ontogeny and temperature. Marine Ecology Progress Series, 338:183–192, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/338/m338p183.pdf.

Lutz:2003:VPC

[LSHL03] V. A. Lutz, S. Sathyendranath, E. J. H. Head, and W. K. W. Li. Variability in pigment composition and optical characteristics of phytoplankton in the Labrador Sea and the Central North Atlantic. Marine Ecology Progress Series, 260: 1–18, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p001.pdf.

Letelier:2008:PEU

[LSK08] R. M. Letelier, P. G. Strutton, and D. M. Karl. Physical and ecological uncertainties in the widespread implementation of controlled upwelling in the North Pacific subtropical gyre. Marine Ecology Progress Series, 371:305–308, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m371p305.pdf.

Liu:2002:PCS

[LSM^{+}02] H. Liu, K. Suzuki, C. Minami, T. Saino, and M. Watanabe. Picoplankton community structure in the subarctic Pacific Ocean and the Bering Sea during summer 1999. Marine Ecology Progress Series, 237:1–14, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/237/m237p001.pdf.

Lee:2005:RSF

[LSMR05] S. H. Lee, D. M. Schell, T. L. McDonald, and W. J. Richardson. Regional and seasonal feeding by bowhead whales Balaena mysticetus as indicated by stable isotope ratios. Marine Ecology Progress Series, 285:271–287, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p271.pdf.

Leporati:2008:DAG

[LSP08] S. C. Leporati, J. M. Semmens, and G. T. Pecl. Determining the age and growth of wild octopus using stylet increment analysis. Marine Ecology Progress Series, 367:213– 222, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p213.pdf.

Luo:2009:MGS

[LSS^{+}09] J. Luo, J. E. Serafy, S. Sponaugle, P. B. Teare, and D. Kieckbusch. Movement of gray snapper Lutjanus griseus among subtropical seagrass, mangrove, and coral reef habitats. Marine Ecology Progress Series, 380:255–269, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/380/
m380p255.pdf.

Linton:2000:FGF

[LT00] D. L. Linton and G. L. Taghon. Feeding, growth, and fecundity of Capitella sp. I in relation to sediment organic concentration. Marine Ecology Progress Series, 205:229–240, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p229.pdf.

Long:2007:GVS

[LT07] J. D. Long and G. C. Trussell. Geographic variation in seaweed induced responses to herbivory. Marine Ecology Progress Series, 333:75–80, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p075.pdf.

Lowe:2003:MPH

[LTCP03] C. G. Lowe, D. T. Topping, D. P. Cartamil, and Y. P. Papastamatiou. Movement patterns, home range, and habitat utilization of adult kelp bass Paralabrax clathratus in a temperate no-take marine reserve. Marine Ecology Progress Series, 256:205–216, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p205.pdf.

Litaker:2002:SNS

[LTD^{+}02] R. W. Litaker, P. A. Tester, C. S. Duke, B. E. Kenney, J. L. Pinckney, and J. Ramus. Seasonal niche strategy of the bloomforming dinoflagellate Heterocapsa triquetra. Marine Ecology Progress Series, 232:45–62, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/232/m232p045.pdf.

Lindsay:2006:REC

[LTFH06] H. Lindsay, C. Todd, T. Fernandes, and M. Huxham. Recruitment in epifaunal communities: an experimental test of the effects of species composition and age. Marine Ecology Progress Series, 307:49–57, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/307/m307p049.pdf.

Lambshead:2000:LDG

[LTFJ00] P. J. D. Lambshead, J. Tietjen, T. Ferrero, and P. Jensen. Latitudinal diversity gradients in the deep sea with special reference to North Atlantic nematodes. Marine Ecology Progress Series, 194:159–167, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/194/m194p159.pdf.

Lambshead:2001:ILS

[LTG^{+}01] P. J. D. Lambshead, J. Tietjen, A. Glover, T. Ferrero, D. Thistle, and A. J. Gooday. Impact of large-scale natural physical disturbance on the diversity of deep-sea North Atlantic nematodes. Marine Ecology Progress Series, 214:121–126, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p121.pdf.

Lohrer:2004:TDS

[LTH^{+}04] A. M. Lohrer, S. F. Thrush, J. E. Hewitt, K. Berkenbusch, M. Ahrens, and V. J. Cummings. Terrestrially derived sediment: response of marine macrobenthic communities to thin terrigenous deposits. Marine Ecology Progress Series, 273: 121–138, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/273/m273p121.pdf.

Lundquist:2006:SVR

[LTH^{+}06] C. J. Lundquist, S. F. Thrush, J. E. Hewitt, J. Halliday, I. MacDonald, and V. J. Cummings. Spatial variability in recolonisation potential: influence of organism behaviour and hydrodynamics on the distribution of macrofaunal colonists. Marine Ecology Progress Series, 324:67–81, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p067.pdf.

Lleonart:2006:IIM

[LTL06a] J. Lleonart, M. Taconet, and M. Lamboeuf. Integrating information on marine species identification for fishery purposes. Marine Ecology Progress Series, 316:231–238, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m316p231.pdf.

Lohrer:2006:DTS

[LTL^{+}06b] A. M. Lohrer, S. F. Thrush, C. J. Lundquist, K. Vopel, J. E. Hewitt, and P. E. Nicholls. Deposition of terrigenous sediment on subtidal marine macrobenthos: response of two contrasting community types. Marine Ecology Progress Series, 307:115–125, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p115.pdf.

Lopanik:2006:OSP

[LTL06c] N. B. Lopanik, N. M. Targett, and N. Lindquist. Ontogeny of a symbiont-produced chemical defense in Bugula neritina (Bryozoa). Marine Ecology Progress Series, 327: 183–191, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p183.pdf.

Laporte:2002:BIM

[LTMDB02] J. M. Laporte, J. P. Truchot, N. Mesmer-Dudons, and A. Boudou. Bioaccumulation of inorganic and methylated mercury by the gills of the shore crab Carcinus maenas: transepithelial fluxes and histochemical localization. Marine Ecology Progress Series, 231:215–228, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p215.pdf.

Lambshead:2001:NAL

[LTMF01] P. J. D. Lambshead, J. Tietjen, C. B. Moncrieff, and T. J. Ferrero. North Atlantic latitudinal diversity patterns in deep-sea marine nematode data: a reply to Rex et al. Marine Ecology Progress Series, 210:299–301, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/210/m210p299.pdf.

Lagos:2007:SSR

[LTNC07] N. A. Lagos, F. J. Tapia, S. A. Navarrete, and J. C. Castilla. Spatial synchrony in the recruitment of intertidal invertebrates along the coast of central Chile. Marine Ecology Progress Series, 350:29–39, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/350/m350p029.pdf.

Levinsen:2000:TCB

[LTNH00] H. Levinsen, J. T. Turner, T. G. Nielsen, and B. W. Hansen. On the trophic coupling between protists and copepods in Arctic marine ecosystems. Marine Ecology Progress Series, 204: 65–77, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p065.pdf.

Ladah:2005:SHS

[LTPL05] L. B. Ladah, F. J. Tapia, J. Pineda, and M. López. Spatially heterogeneous, synchronous settlement of Chthamalus spp. larvae in northern Baja California. Marine Ecology Progress Series, 302:177–185, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/302/m302p177.pdf.

Lass:2001:FNK

[LTV^{+}01] S. Lass, G. A. Tarling, P. Virtue, J. B. L. Matthews, P. Mayzaud, and F. Buchholz. On the food of northern krill Meganyctiphanes norvegica in relation to its vertical distribution. Marine Ecology Progress Series, 214:177–200, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p177.pdf.

Lopez-Urrutia:1999:GTD

[LUA99] Á. López-Urrutia and J. L. Acuña. Gut throughput dynamics in the appendicularian Oikopleura dioica. Marine Ecology Progress Series, 191:195–205, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:/

/www.int-res.com/articles/meps/191/m191p195.pdf. See

erratum [Ano00c].

Lopez-Urrutia:2003:FLG

[LUAIH03] Á. López-Urrutia, J. L. Acuña, X. Irigoien, and R. Harris. Food limitation and growth in temperate epipelagic appendicularians (Tunicata). Marine Ecology Progress Series, 252:143–157, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p143.pdf.

Lopez-Urrutia:2003:LSF

[LUIAH03] Á. López-Urrutia, X. Irigoien, J. L. Acuña, and R. Harris. In situ feeding physiology and grazing impact of the appendicularian community in temperate waters. Marine Ecology Progress Series, 252:125–141, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p125.pdf.

Lusseau:2003:MFB

[Lus03] D. Lusseau. Male and female bottlenose dolphins Tursiops spp. have different strategies to avoid interactions with tour boats in Doubtful Sound, New Zealand. Marine Ecology Progress Series, 257:267–274, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/257/m257p267.pdf.

Lusseau:2005:RPB

[Lus05] D. Lusseau. Residency pattern of bottlenose dolphins Tursiops spp. in Milford Sound, New Zealand, is related to boat traffic. Marine Ecology Progress Series, 295:265–272, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p265.pdf.

Lavery:2002:CST

[LV02] P. S. Lavery and M. A. Vanderklift. A comparison of spatial and temporal patterns in epiphytic macroalgal assemblages of the seagrasses Amphibolis griffithii and Posidonia coriacea. Marine Ecology Progress Series, 236:99–112, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p099.pdf.

Lever:2005:RMB

[LV05] M. A. Lever and I. Valiela. Response of microphytobenthic biomass to experimental nutrient enrichment and grazer exclusion at different land-derived nitrogen loads. Marine Ecology Progress Series, 294:117–129, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/294/m294p117.pdf.

Lopez:2007:SVA

[LVA07] E. López, L. Viesca, and R. Anadón. Seasonal variation in abundance and feeding rates of the first stages of copepods in a temperate sea. Marine Ecology Progress Series, 352:161–175, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/352/m352p161.pdf.

Luedeking:2005:ICM

[LVK05] A. Luedeking, C. J. F. Van Noorden, and A. Koehler. Identification and characterisation of a multidrug resistance-related protein mRNA in the blue mussel Mytilus edulis. Marine Ecology Progress Series, 286:167–175, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p167.pdf.

Lacap:2002:PDS

[LVRN02] C. D. A. Lacap, J. E. Vermaat, R. N. Rollon, and H. M. Nacorda. Propagule dispersal of the SE Asian seagrasses Enhalus acoroides and Thalassia hemprichii. Marine Ecology Progress Series, 235:75–80, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p075.pdf.

Lett:2007:AEB

[LVvdLH07] C. Lett, J. Veitch, C. D. van der Lingen, and L. Hutchings. Assessment of an environmental barrier to transport of ichthyoplankton from the southern to the northern Benguela ecosystems. Marine Ecology Progress Series, 347:247–259, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p247.pdf.

Lopez-Victoria:2006:CSB

[LVZW06] M. López-Victoria, S. Zea, and E. Weil. Competition for space between encrusting excavating Caribbean sponges and other coral reef organisms. Marine Ecology Progress Series, 312:113–121, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/312/m312p113.pdf.

Lotze:2000:VCE

[LW00] H. K. Lotze and B. Worm. Variable and complementary effects of herbivores on different life stages of bloom-forming macroalgae. Marine Ecology Progress Series, 200:167–175, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p167.pdf.

Lohrer:2002:RIT

[LW02] A. M. Lohrer and R. B. Whitlatch. Relative impacts of two exotic brachyuran species on blue mussel populations in Long Island Sound. Marine Ecology Progress Series, 227: 135–144, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/227/m227p135.pdf.

Luzier:2004:AMH

[LW04] C. W. Luzier and R. R. Wilson, Jr. Analysis of mtDNA haplotypes of kelp bass tests for sibling-dominated recruitment near marine protected areas of the California Channel Islands. Marine Ecology Progress Series, 277:221–230, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p221.pdf.

Long:2005:MIB

[LW05] A. Long and W. X. Wang. Metallothionein induction and bioaccumulation kinetics of Cd and Ag in the marine fish Terapon jarbua challenged with dietary or waterborne Ag and Cu. Marine Ecology Progress Series, 291:215–226, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p215.pdf.

Lusseau:2006:ILP

[LW06] S. M. Lusseau and S. R. Wing. Importance of local production versus pelagic subsidies in the diet of an isolated population of bottlenose dolphins Tursiops sp. Marine Ecology Progress Series, 321:283–293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/321/m321p283.pdf.

Lucas:2000:BPE

[LWB^{+}00] C. H. Lucas, J. Widdows, M. D. Brinsley, P. N. Salkeld, and P. M. J. Herman. Benthic-pelagic exchange of microalgae at a tidal flat. 1. Pigment analysis. Marine Ecology Progress Series, 196:59–73, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p059.pdf.

Littman:2009:BCJ

[LWB09] R. A. Littman, B. L. Willis, and D. G. Bourne. Bacterial communities of juvenile corals infected with different Symbiodinium (dinoflagellate) clades. Marine Ecology Progress Series, 389:45–59, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p045.pdf.

Lapid:2004:STB

[LWCF04] E. D. Lapid, J. Wielgus, and N. E. Chadwick-Furman. Sweeper tentacles of the brain coral Platygyra daedalea: induced development and effects on competitors. Marine Ecology Progress Series, 282:161–171, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p161.pdf.

Leys:2004:PGS

[LWH^{+}04] S. P. Leys, K. Wilson, C. Holeton, H. M. Reiswig, W. C. Austin, and V. Tunnicliffe. Patterns of glass sponge (Porifera, Hexactinellida) distribution in coastal waters of British Columbia, Canada. Marine Ecology Progress Series, 283: 133–149, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/283/m283p133.pdf.

Liu:2006:PDA

[LWH06] S. Liu, W. X. Wang, and L. M. Huang. Phosphorus dietary assimilation and efflux in the marine copepod Acartia erythraea. Marine Ecology Progress Series, 321:193–202, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p193.pdf.

Li:2008:TEP

[LWH08] B. Li, J. E. Ward, and B. A. Holohan. Transparent exopolymer particles (TEP) from marine suspension feeders enhance particle aggregation. Marine Ecology Progress Series, 357:67–77, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/357/m357p067.pdf.

Lee:2004:SME

[LWJC04] J. T. Lee, J. Widdows, M. B. Jones, and R. A. Coleman. Settlement of megalopae and early juveniles of the velvet swimming crab Necora puber (Decapoda: Portunidae) in flow conditions. Marine Ecology Progress Series, 272:191– 202, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/272/m272p191.pdf.

Lin:2007:MAN

[LWK^{+}07] H. J. Lin, C. Y. Wu, S. J. Kao, W. Y. Kao, and P. J. Meng. Mapping anthropogenic nitrogen through point sources in coral reefs using δ ^{15}N in macroalgae. Marine Ecology Progress Series, 335:95–109, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/335/m335p095.pdf.

Lotze:2002:EUR

[LWMW02] H. K. Lotze, B. Worm, M. Molis, and M. Wahl. Effects of UV radiation and consumers on recruitment and succession of a marine macrobenthic community. Marine Ecology Progress Series, 243:57–66, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/243/m243p057.pdf.

Litaker:2002:EDI

[LWR^{+}02] R. W. Litaker, V. E. Warner, C. Rhyne, C. S. Duke, B. E. Kenney, J. Ramus, and P. A. Tester. Effect of diel and interday variations in light on the cell division pattern and in situ growth rates of the bloom-forming dinoflagellate Heterocapsa triquetra. Marine Ecology Progress Series, 232: 63–74, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/232/m232p063.pdf.

Lewis:2001:DBP

[LWW^{+}01] S. Lewis, S. Wanless, P. J. Wright, M. P. Harris, J. Bull, and D. A. Elston. Diet and breeding performance of blacklegged kittiwakes Rissa tridactyla at a North Sea colony. Marine Ecology Progress Series, 221:277–284, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p277.pdf.

Lee:2007:SBC

[LYL^{+}07] O. O. Lee, L. H. Yang, X. Li, J. R. Pawlik, and P. Y. Qian. Surface bacterial community, fatty acid profile, and antifouling activity of two congeneric sponges from Hong Kong and the Bahamas. Marine Ecology Progress Series, 339: 25–40, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/339/m339p025.pdf.

Lo-Yat:2006:LSD

[LYMCG06] A. Lo-Yat, M. G. Meekan, J. H. Carleton, and R. Galzin. Large-scale dispersal of the larvae of nearshore and pelagic fishes in the tropical oceanic waters of French Polynesia. Marine Ecology Progress Series, 325:195–203, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p195.pdf.

Liu:2005:ESB

[LZL05] S. M. Liu, J. Zhang, and R. X. Li. Ecological significance of biogenic silica in the East China Sea. Marine Ecology Progress Series, 290:15–26, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/290/m290p015.pdf.

Levin:2003:SHM

[LZM^{+}03] L. A. Levin, W. Ziebis, G. F. Mendoza, V. A. Growney, M. D. Tryon, K. M. Brown, C. Mahn, J. M. Gieskes, and A. E. Rathburn. Spatial heterogeneity of macrofauna at northern California methane seeps: influence of sulfide concentration and fluid flow. Marine Ecology Progress Series, 265:123–139, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p123.pdf.

Lloyd:2008:EST

[LZW^{+}08] D. C. Lloyd, D. C. Zacherl, S. Walker, G. Paradis, M. Sheehy, and R. R. Warner. Egg source, temperature and culture seawater affect elemental signatures in Kelletia kelletii larval statoliths. Marine Ecology Progress Series, 353:115– 130, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/353/m353p115.pdf.

Miller:2000:SPD

[MA00] A. W. Miller and R. F. Ambrose. Sampling patchy distributions: comparison of sampling designs in rocky intertidal habitats. Marine Ecology Progress Series, 196:1–14, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p001.pdf.

Mokievsky:2002:RES

[MA02] V. Mokievsky and A. Azovsky. Re-evaluation of species diversity patterns of free-living marine nematodes. Marine Ecology Progress Series, 238:101–108, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/238/m238p101.pdf.

Moody:2007:THS

[MA07] R. M. Moody and R. B. Aronson. Trophic heterogeneity in salt marshes of the northern Gulf of Mexico. Marine Ecology Progress Series, 331:49–65, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/331/m331p049.pdf.

Medina:2007:ISG

[MAA^{+}07] A. Medina, F. J. Abascal, L. Aragón, G. Mourente, G. Aranda, T. Galaz, A. Belmonte, J. M. de la Serna, and S. García. Influence of sampling gear in assessment of reproductive parameters for bluefin tuna in the western Mediterranean. Marine Ecology Progress Series, 337:221–230, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m337p221.pdf.

McManus:2003:CDP

[MAB^{+}03] M. A. McManus, A. L. Alldredge, A. H. Barnard, E. Boss, J. F. Case, T. J. Cowles, P. L. Donaghay, L. B. Eisner, D. J. Gifford, C. F. Greenlaw, C. M. Herren, D. V. Holliday, D. Johnson, S. MacIntyre, D. M. McGehee, T. R. Osborn, M. J. Perry, R. E. Pieper, J. E. B. Rines, D. C. Smith, J. M. Sullivan, M. K. Talbot, M. S. Twardowski, A. Weidemann, and J. R. Zaneveld. Characteristics, distribution and persistence of thin layers over a 48 hour period. Marine Ecology Progress Series, 261:1–19, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p001.pdf.

Meyer:2003:FEB

[MABB03] B. Meyer, A. Atkinson, B. Blume, and U. V. Bathmann. Feeding and energy budgets of larval Antarctic krill Euphausia superba in summer. Marine Ecology Progress Series, 257:167–178, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p167.pdf.

Mehta:2007:BWI

[MAC^{+}07] A. V. Mehta, J. M. Allen, R. Constantine, C. Garrigue, B. Jann, C. Jenner, M. K. Marx, C. O. Matkin, D. K. Mattila, G. Minton, S. A. Mizroch, C. Olavarría, J. Robbins, K. G. Russell, R. E. Seton, G. H. Steiger, G. A. Víkingsson, P. R. Wade, B. H. Witteveen, and P. J. Clapham. Baleen whales are not important as prey for killer whales Orcinus orca in high-latitude regions. Marine Ecology Progress Series, 348:297–307, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p297.pdf.

Mapstone:2007:MES

[MACS07] B. D. Mapstone, N. L. Andrew, Y. Chancerelle, and B. Salvat. Mediating effects of sea urchins on interactions among corals, algae and herbivorous fish in the Moorea lagoon, French Polynesia. Marine Ecology Progress Series, 332:143– 153, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/332/m332p143.pdf.

Madon:2002:EJC

[Mad02] S. P. Madon. Ecophysiology of juvenile California halibut Paralichthys californicus in relation to body size, water temperature and salinity. Marine Ecology Progress Series, 243: 235–249, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p235.pdf.

Marta-Almeida:2006:IVM

[MADPQ06] M. S. Marta-Almeida, J. Dubert, Á. Peliz, and H. Queiroga. Influence of vertical migration pattern on retention of crab larvae in a seasonal upwelling system. Marine Ecology Progress Series, 307:1–19, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Magnusson:2003:DTD

[MAE03] K. Magnusson, S. Agrenius, and R. Ekelund. Distribution of a tetrabrominated diphenyl ether and its metabolites in softbottom sediment and macrofauna species. Marine Ecology Progress Series, 255:155–170, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/255/m255p155.pdf.

Magnussen:2002:DFA

[Mag02] E. Magnussen. Demersal fish assemblages of Faroe Bank: species composition, distribution, biomass spectrum and diversity. Marine Ecology Progress Series, 238:211–225, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/238/m238p211.pdf.

McClanahan:2007:WIO

[MAG^{+}07a] T. R. McClanahan, M. Ateweberhan, N. A. J. Graham, S. K. Wilson, C. R. Sebastián, M. M. M. Guillaume, and J. H. Bruggemann. Western Indian Ocean coral communities: bleaching responses and susceptibility to extinction. Marine Ecology Progress Series, 337:1–13, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Myhre:2007:RSU

[MAG07b] S. Myhre and A. Acevedo-Gutiérrez. Recovery of sea urchin Diadema antillarum populations is correlated to increased coral and reduced macroalgal cover. Marine Ecology Progress Series, 329:205–210, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/329/m329p205.pdf.

McClintock:2008:IAA

[MAH^{+}08] J. B. McClintock, R. A. Angus, C. P. Ho, C. D. Amsler, and B. J. Baker. Intraspecific agonistic arm-fencing behavior in the Antarctic keystone sea star Odontaster validus influences prey acquisition. Marine Ecology Progress Series, 371:297–300, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/371/m371p297.pdf.

McMahon:2006:BCR

[MAJ^{+}06] K. W. McMahon, W. G. Ambrose, Jr., B. J. Johnson, M. Y. Sun, G. R. Lopez, L. M. Clough, and M. L. Carroll. Benthic community response to ice algae and phytoplankton in Ny Ålesund, Svalbard. Marine Ecology Progress Series, 310:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moses:2009:RER

[MAKMK09] C. S. Moses, S. Andréfouët, C. J. Kranenburg, and F. E. Muller-Karger. Regional estimates of reef carbonate dynamics and productivity using Landsat 7 ETM+, and potential impacts from ocean acidification. Marine Ecology Progress Series, 380:103–115, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p103.pdf.

Maltagliati:2002:GMB

[Mal02] F. Maltagliati. Genetic monitoring of brackish-water populations: the Mediterranean toothcarp Aphanius fasciatus (Cyprinodontidae) as a model. Marine Ecology Progress Series, 235:257–262, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p257.pdf.

Marui:2001:CEL

[MAM^{+}01] M. Marui, T. Arai, M. J. Miller, D. J. Jellyman, and K. Tsukamoto. Comparison of early life history between New Zealand temperate eels and Pacific tropical eels revealed by otolith microstructure and microchemistry. Marine Ecology Progress Series, 213:273–284, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/213/m213p273.pdf.

McClanahan:2009:RST

[MAOP09] T. R. McClanahan, M. Ateweberhan, J. Omukoto, and L. Pearson. Recent seawater temperature histories, status, and predictions for Madagascar’s coral reefs. Marine Ecology Progress Series, 380:117–128, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m380p117.pdf.

Marshall:2000:SDR

[Mar00] P. A. Marshall. Skeletal damage in reef corals: relating resistance to colony morphology. Marine Ecology Progress Series, 200:177–189, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p177.pdf.

MZC2:2001:TSF

[Mar01] Marine Zooplankton Colloquium 2. Theme section: Future marine zooplankton research — a perspective. Marine Ecology Progress Series, 222:297–308, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/222/m222p297.pdf.

Marshall:2002:LSM

[Mar02] D. J. Marshall. In situ measures of spawning synchrony and fertilization success in an intertidal, free-spawning invertebrate. Marine Ecology Progress Series, 236:113–119, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/236/m236p113.pdf.

Marsden:2004:ERS

[Mar04] I. D. Marsden. Effects of reduced salinity and seston availability on growth of the New Zealand little-neck clam Austrovenus stutchburyi. Marine Ecology Progress Series, 266: 157–171, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/266/m266p157.pdf.

Miner:2006:RFS

[MARM06] C. M. Miner, J. M. Altstatt, P. T. Raimondi, and T. E. Minchinton. Recruitment failure and shifts in community structure following mass mortality limit recovery prospects of black abalone. Marine Ecology Progress Series, 327:107– 117, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p107.pdf.

Moncoiffe:2000:SST

[MASFS00] G. Moncoiffé, X. A. Alvarez-Salgado, F. G. Figueiras, and G. Savidge. Seasonal and short-time-scale dynamics of microplankton community production and respiration in an inshore upwelling system. Marine Ecology Progress Series, 196: 111–126, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/196/m196p111.pdf.

McGlathery:2001:MVB

[MAT01] K. J. McGlathery, I. C. Anderson, and A. C. Tyler. Magnitude and variability of benthic and pelagic metabolism in a temperate coastal lagoon. Marine Ecology Progress Series, 216:1–15, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/216/m216p001.pdf.

Martinez-Abrain:2006:SSM

[MAVO^{+}06] A. Martínez-Abraín, A. Velando, D. Oro, M. Genovart, C. Gerique, M. A. Bartolomé, E. Villuendas, and B. Sarzo. Sex-specific mortality of European shags after the Prestige oil spill: demographic implications for the recovery of colonies.

Marine Ecology Progress Series, 318:271–276, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p271.pdf.

Meekan:2001:DAS

[MAW01] M. G. Meekan, J. L. Ackerman, and G. M. Wellington. Demography and age structures of coral reef damselfishes in the tropical eastern Pacific Ocean. Marine Ecology Progress Series, 212:223–232, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p223.pdf.

Milazzo:2006:RFF

[MAW06] M. Milazzo, I. Anastasi, and T. J. Willis. Recreational fish feeding affects coastal fish behavior and increases frequency of predation on damselfish Chromis chromis nests. Marine Ecology Progress Series, 310:165–172, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p165.pdf.

Minello:2003:SMN

[MAWH03] T. J. Minello, K. W. Able, M. P. Weinstein, and C. G. Hays. Salt marshes as nurseries for nekton: testing hypotheses on density, growth and survival through meta-analysis. Marine Ecology Progress Series, 246:39–59, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p039.pdf.

Mazouni:2004:ISO

[Maz04] N. Mazouni. Influence of suspended oyster cultures on nitrogen regeneration in a coastal lagoon (Thau, France). Marine Ecology Progress Series, 276:103–113, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p103.pdf.

McKindsey:2000:EMV

[MB00a] C. W. McKindsey and E. Bourget. Explaining mesoscale variation in intertidal mussel community structure. Marine Ecology Progress Series, 205:155–170, 2000. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/205/m205p155.pdf.

Mundy:2000:VDP

[MB00b] C. Mundy and R. Babcock. Are vertical distribution patterns of scleractinian corals maintained by pre- or postsettlement processes? A case study of three contrasting species. Marine Ecology Progress Series, 198:109–119, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p109.pdf.

McKindsey:2001:BSS

[MB01] C. W. McKindsey and E. Bourget. Body size and spatial variation of community structure in subarctic intertidal boulder fields. Marine Ecology Progress Series, 216:17–30, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p017.pdf.

Marnane:2002:DNF

[MB02a] M. J. Marnane and D. R. Bellwood. Diet and nocturnal foraging in cardinalfishes (Apogonidae) at One Tree Reef, Great Barrier Reef, Australia. Marine Ecology Progress Series, 231:261–268, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p261.pdf.

Marteinsdottir:2002:ERI

[MB02b] G. Marteinsdottir and G. A. Begg. Essential relationships incorporating the influence of age, size and condition on variables required for estimation of reproductive potential in Atlantic cod Gadus morhua. Marine Ecology Progress Series, 235:235–256, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p235.pdf.

Mori:2004:SFB

[MB04] Y. Mori and I. L. Boyd. Segregation of foraging between two sympatric penguin species: does rate maximisation make the difference? Marine Ecology Progress Series, 275:241– 249, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/275/m275p241.pdf.

Manning:2005:MAC

[MB05a] C. A. Manning and A. Bucklin. Multivariate analysis of the copepod community of near-shore waters in the western Gulf of Maine. Marine Ecology Progress Series, 292: 233–249, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p233.pdf.

Moran:2005:SGT

[MB05b] K. L. Moran and K. A. Bjorndal. Simulated green turtle grazing affects structure and productivity of seagrass pastures. Marine Ecology Progress Series, 305:235–247, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/305/
m305p235.pdf.

Mantyka:2007:MGS

[MB07a] C. S. Mantyka and D. R. Bellwood. Macroalgal grazing selectivity among herbivorous coral reef fishes. Marine Ecology Progress Series, 352:177–185, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/352/m352p177.pdf.

Marko:2007:BSP

[MB07b] P. B. Marko and K. R. Barr. Basin-scale patterns of mtDNA differentiation and gene flow in the bay scallop Argopecten irradians concentricus. Marine Ecology Progress Series, 349: 139–150, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/349/m349p139.pdf.

Marshall:2007:I

[MB07c] C. T. Marshall and Howard I. Browman. Introduction. Marine Ecology Progress Series, 335:249–251, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m335p249.pdf.

Marshall:2007:DCM

[MB07d] C. Tare Marshall and Howard I. Browman. Disentangling the causes of maturation trends in exploited fish populations. Marine Ecology Progress Series, 335:249–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m335_TS.pdf.

Metaxas:2007:PHM

[MB07e] A. Metaxas and T. Bryan. Predictive habitat model for deep gorgonians needs better resolution: Reply to Etnoyer & Morgan. Marine Ecology Progress Series, 339:313–314, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/339/m339p313.pdf. See [BM07c, EM07b].

MacQuarrie:2008:BPR

[MB08] S. P. MacQuarrie and V. M. Bricelj. Behavioral and physiological responses to PSP toxins in Mya arenaria populations in relation to previous exposure to red tides. Marine Ecology Progress Series, 366:59–74, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/366/m366p059.pdf.

McClanahan:2005:RCB

[MBA^{+}05] T. R. McClanahan, A. C. Baker, M. Ateweberhan, J. Maina, and K. R. Moothien-Pillay. Refining coral bleaching experiments and models through reiterative field studies. Marine Ecology Progress Series, 305:301–303, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/305/
m305p301.pdf.

Munk:2004:AFL

[MBB^{+}04] Peter Munk, Peter Koefoed Bjørnsen, Pensri Boonruang, Michael Fryd, Per Juel Hansen, Vudhichai Janekarn, Vitthaya Limtrakulvong, Torkel Gissel Nielsen, Ole Schou Hansen, Suree Satapoomin, Suchat Sawangarreruks, Helge Abilhauge Thomsen, and Jette Buch Østergaard. Assemblages of fish larvae and mesozooplankton across the continental shelf and shelf slope of the Andaman Sea (NE Indian Ocean). Marine Ecology Progress Series, 274:87–97, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p087.pdf.

McCormick:2008:DMP

[MBBP08] T. B. McCormick, L. M. Buckley, J. Brogan, and L. M. Perry. Drift macroalgae as a potential dispersal mechanism for the white abalone Haliotis sorenseni. Marine Ecology Progress Series, 362:225–232, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m362p225.pdf.

McManus:2008:BEP

[MBBW08] M. A. McManus, K. J. Benoit-Bird, and C. B. Woodson. Behavior exceeds physical forcing in the diel horizontal migration of the midwater sound-scattering layer in Hawaiian waters. Marine Ecology Progress Series, 365:91–101, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p091.pdf.

McAndrew:2007:MRO

[MBC^{+}07] P. M. McAndrew, K. M. Björkman, M. J. Church, P. J. Morris, N. Jachowski, P. J. le B. Williams, and D. M. Karl. Metabolic response of oligotrophic plankton communities to deep water nutrient enrichment. Marine Ecology Progress Series, 332:63–75, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m332p063.pdf.

Maranon:2003:HVP

[MBG^{+}03] E. Marañón, M. J. Behrenfeld, N. González, B. Mouriño, and M. V. Zubkov. High variability of primary production in oligotrophic waters of the Atlantic Ocean: uncoupling from phytoplankton biomass and size structure. Marine Ecology Progress Series, 257:1–11, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p001.pdf.

Muhling:2008:SLA

[MBG^{+}08] B. A. Muhling, L. E. Beckley, D. J. Gaughan, C. M. Jones, A. G. Miskiewicz, and S. A. Hesp. Spawning, larval abundance and growth rate of Sardinops sagax off southwestern Australia: influence of an anomalous eastern boundary current.

Marine Ecology Progress Series, 364:157–167, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p157.pdf.

Montevecchi:2009:FFT

[MBG^{+}09] W. A. Montevecchi, S. Benvenuti, S. Garthe, G. K. Davoren, and D. Fifield. Flexible foraging tactics by a large opportunistic seabird preying on forage- and large pelagic fishes. Marine Ecology Progress Series, 385:295–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p295.pdf.

McClanahan:2000:RFA

[MBH^{+}00] T. R. McClanahan, K. Bergman, M. Huitric, M. McField, T. Elfwing, M. Nyström, and I. Nordemar. Response of fishes to algae reduction on Glovers Reef, Belize. Marine Ecology Progress Series, 206:273–282, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p273.pdf.

McMahon:2007:STR

[MBH07] C. R. McMahon, C. J. A. Bradshaw, and G. C. Hays. Satellite tracking reveals unusual diving characteristics for a marine reptile, the olive ridley turtle Lepidochelys olivacea. Marine Ecology Progress Series, 329:239–252, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p239.pdf.

Metian:2009:DHM

[MBH^{+}09] M. Metian, P. Bustamente, L. Hédouin, F. Oberhänsli, and M. Warnau. Delineation of heavy metal uptake pathways (seawater and food) in the variegated scallop Chlamys varia, using radiotracer techniques. Marine Ecology Progress Series, 375:161–171, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p161.pdf.

Markey:2007:IFA

[MBHN07] K. L. Markey, A. H. Baird, C. Humphrey, and A. P. Negri. Insecticides and a fungicide affect multiple coral life stages. Marine Ecology Progress Series, 330:127–137, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m330p127.pdf.

Morley:2007:WES

[MBL07] J. W. Morley, J. A. Buckel, and T. E. Lankford, Jr. Winter energy storage dynamics and cohort structure of youngof-the-year bluefish Pomatomus saltatrix off North Carolina. Marine Ecology Progress Series, 334:273–286, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p273.pdf.

Marubini:2001:DCL

[MBLA01] F. Marubini, H. Barnett, C. Langdon, and M. J. Atkinson. Dependence of calcification on light and carbonate ion concentration for the hermatypic coral Porites compressa. Marine Ecology Progress Series, 220:153–162, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/220/m220p153.pdf.

Moodley:2000:ESB

[MBM^{+}00] L. Moodley, H. T. S. Boschker, J. J. Middelburg, R. Pel, P. M. J. Herman, E. de Deckere, and C. H. R. Heip. Ecological significance of benthic Foraminifera: ^{13}C labelling experiments. Marine Ecology Progress Series, 202:289–295, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p289.pdf.

Murphy:2009:DBW

[MBN^{+}09] K. R. Murphy, J. R. Boehme, M. Noble, G. Smith, and G. M. Ruiz. Deducing ballast water sources in ships arriving in New Zealand from southeastern Australia. Marine Ecology Progress Series, 390:39–53, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/390/m390p039.pdf.

Meekan:2006:PSS

[MBP^{+}06] M. G. Meekan, C. J. A. Bradshaw, M. Press, C. McLean, A. Richards, S. Quasnichka, and J. G. Taylor. Population size and structure of whale sharks Rhincodon typus at Ningaloo Reef, Western Australia. Marine Ecology Progress Series, 319:275–285, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p275.pdf.

Meleder:2005:STC

[MBR^{+}05] V. Méléder, L. Barillé, Y. Rincé, M. Morançais, P. Rosa, and P. Gaudin. Spatio-temporal changes in microphytobenthos structure analysed by pigment composition in a macrotidal flat (Bourgneuf Bay, France). Marine Ecology Progress Series, 297:83–99, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/297/m297p083.pdf.

Mountain:2003:DAM

[MBS03] D. Mountain, P. Berrien, and J. Sibunka. Distribution, abundance and mortality of cod and haddock eggs and larvae on Georges Bank in 1995 and 1996. Marine Ecology Progress Series, 263:247–260, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/263/m263p247.pdf. See erra-

tum [MBS07].

Mountain:2007:EDA

[MBS07] David Mountain, Peter Berrien, and John Sibunka. Erratum: Distribution, abundance and mortality of cod and haddock eggs and larvae on Georges Bank in 1995 and 1996 263:247–260 (2003). Marine Ecology Progress Series, ??: ??, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/263/m341p309.pdf. See [MBS03].

Maggi:2009:ECN

[MBVBC09] E. Maggi, I. Bertocci, S. Vaselli, and L. Benedetti-Cecchi. Effects of changes in number, identity and abundance of habitatforming species on assemblages of rocky seashores. Marine Ecology Progress Series, 381:39–49, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m381p039.pdf.

Magalhaes:2002:TSP

[MBW02] C. M. Magalhães, A. A. Bordalo, and W. J. Wiebe. Temporal and spatial patterns of intertidal sediment-water nutrient and oxygen fluxes in the Douro River estuary, Portugal. Marine Ecology Progress Series, 233:55–71, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p055.pdf.

Magalhaes:2003:IBR

[MBW03] Catarina M. Magalhães, Adriano A. Bordalo, and William J. Wiebe. Intertidal biofilms on rocky substratum can play a major role in estuarine carbon and nutrient dynamics. Marine Ecology Progress Series, 258:275–281, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p275.pdf.

Miller:2007:VDS

[MBW07] M. W. Miller, I. B. Baums, and D. E. Williams. Visual discernment of sexual recruits is not feasible for Acropora palmata. Marine Ecology Progress Series, 335:227–231, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p227.pdf.

Manriquez:2001:SMP

[MC01a] P. H. Manríquez and J. C. Castilla. Significance of marine protected areas in central Chile as seeding grounds for the gastropod Concholepas concholepas. Marine Ecology Progress Series, 215:201–211, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/215/m215p201.pdf.

Milton:2001:COC

[MC01b] D. A. Milton and S. R. Chenery. Can otolith chemistry detect the population structure of the shad hilsa Tenualosa ilisha? Comparison with the results of genetic and morphological studies. Marine Ecology Progress Series, 222:239–251, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p239.pdf.

Manriquez:2005:SFA

[MC05a] P. H. Manríquez and J. C. Castilla. Self-fertilization as an alternative mode of reproduction in the solitary tunicate Pyura chilensis. Marine Ecology Progress Series, 305:113– 125, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/305/m305p113.pdf.

McKiness:2005:UMB

[MC05b] Z. P. McKiness and C. M. Cavanaugh. The ubiquitous mussel: Bathymodiolus aff. brevior symbiosis at the Central Indian Ridge hydrothermal vents. Marine Ecology Progress Series, 295:183–190, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p183.pdf.

Milton:2005:MPB

[MC05c] D. A. Milton and S. R. Chenery. Movement patterns of barramundi Lates calcarifer, inferred from ^{87}Sr/^{86}Sr and Sr/Ca ratios in otoliths, indicate non-participation in spawning. Marine Ecology Progress Series, 301:279–291, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/301/
m301p279.pdf.

McClain:2006:IER

[MC06] C. R. McClain and J. Crouse. Influence of ecological role on bathymetric patterns of deep-sea species: size clines in parasitic gastropods. Marine Ecology Progress Series, 320: 161–167, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p161.pdf.

Manriquez:2007:RLB

[MC07] P. H. Manríquez and J. C. Castilla. Roles of larval behaviour and microhabitat traits in determining spatial aggregations in the ascidian Pyura chilensis. Marine Ecology Progress Series, 332:155–165, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/332/m332p155.pdf.

Mohammed:2008:EDS

[MC08a] S. W. C. Mohammed and S. E. Coppard. Ecology and distribution of soft-sediment benthic communities off Viti Levu (Fiji). Marine Ecology Progress Series, 371:91–107, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/371/
m371p091.pdf.

Morrow:2008:SKC

[MC08b] K. Morrow and R. Carpenter. Shallow kelp canopies mediate macroalgal composition: effects on the distribution and abundance of Corynactis californica (Corallimorpharia). Marine Ecology Progress Series, 361:119–127, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p119.pdf.

Mann:2007:ALF

[MCBT07] D. A. Mann, B. M. Casper, K. S. Boyle, and T. C. Tricas. On the attraction of larval fishes to reef sounds. Marine Ecology Progress Series, 338:307–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/338/m338p307.pdf.

Mathieson:2000:FAE

[MCC^{+}00] S. Mathieson, A. Cattrijsse, M. J. Costa, P. Drake, M. Elliott, J. Gardner, and J. Marchand. Fish assemblages of European tidal marshes: a comparison based on species, families and functional guilds. Marine Ecology Progress Series, 204:225–242, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p225.pdf.

Montory:2009:IIC

[MCCP09] J. A. Montory, O. R. Chaparro, V. M. Cubillos, and J. A. Pechenik. Isolation of incubation chambers during brooding: effect of reduced pH on protoconch development in the estuarine gastropod Crepipatella dilatata (Calyptraeidae). Marine Ecology Progress Series, 374:157–166, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p157.pdf.

Martin:2007:CMTa

[MCCT07a] S. Martin, J. Clavier, L. Chauvaud, and G. Thouzeau. Community metabolism in temperate maerl beds. I. Carbon and carbonate fluxes. Marine Ecology Progress Series, 335:19– 29, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/335/m335p019.pdf.

Martin:2007:CMTb

[MCCT07b] S. Martin, J. Clavier, L. Chauvaud, and G. Thouzeau. Community metabolism in temperate maerl beds. II. Nutrient fluxes. Marine Ecology Progress Series, 335:31–41, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/335/m335p031.pdf.

Mangialajo:2008:LFA

[MCCV08] L. Mangialajo, M. Chiantore, and R. Cattaneo-Vietti. Loss of fucoid algae along a gradient of urbanisation, and structure of benthic assemblages. Marine Ecology Progress Series, 358:63–74, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/358/m358p063.pdf.

McManus:2005:EPP

[MCD^{+}05] M. A. McManus, O. M. Cheriton, P. T. Drake, D. V. Holliday, C. D. Storlazzi, P. L. Donaghay, and C. F. Greenlaw. Effects of physical processes on structure and transport of thin zooplankton layers in the coastal ocean. Marine Ecology Progress Series, 301:199–215, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/301/m301p199.pdf.

Marsh:2001:LEM

[MCE01a] A. G. Marsh, S. Cohen, and C. E. Epifanio. Larval energy metabolism and physiological variability in the Asian shore crab Hemigrapsus sanguineus. Marine Ecology Progress Series, 218:303–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/218/m218p303.pdf.

Mumby:2001:CWM

[MCE^{+}01b] P. J. Mumby, J. R. M. Chisholm, A. L. Edwards, S. Andrefouet, and J. Jaubert. Cloudy weather may have saved Society Island reef corals during the 1998 ENSO event. Marine Ecology Progress Series, 222:209–216, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p209.pdf.

MacAvoy:2002:UCB

[MCFM02] S. E. MacAvoy, R. S. Carney, C. R. Fisher, and S. A. Macko. Use of chemosynthetic biomass by large, mobile, benthic predators in the Gulf of Mexico. Marine Ecology Progress Series, 225:65–78, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p065.pdf.

McGovern:2002:PSA

[McG02] T. M. McGovern. Patterns of sexual and asexual reproduction in the brittle star Ophiactis savignyi in the Florida Keys. Marine Ecology Progress Series, 230:119–126, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p119.pdf.

Marino:2006:ECP

[MCH^{+}06] R. Marino, F. Chan, R. W. Howarth, M. L. Pace, and G. E. Likens. Ecological constraints on planktonic nitrogen fixation in saline estuaries. I. Nutrient and trophic controls. Marine Ecology Progress Series, 309:25–39, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m309p025.pdf.

McIlwain:2002:LBR

[McI02] J. L. McIlwain. Link between reproductive output and larval supply of a common damselfish species, with evidence of replenishment from outside the local population. Marine Ecology Progress Series, 236:219–232, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p219.pdf.

McIlwain:2003:FST

[McI03] J. L. McIlwain. Fine-scale temporal and spatial patterns of larval supply to a fringing reef in Western Australia. Marine Ecology Progress Series, 252:207–222, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p207.pdf.

Mackey:2007:EER

[MCL07] P. Mackey, C. J. Collier, and P. S. Lavery. Effects of experimental reduction of light availability on the seagrass Amphibolis griffithii. Marine Ecology Progress Series, 342:117– 126, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/342/m342p117.pdf.

Meekan:2003:WDG

[MCM^{+}03] M. G. Meekan, J. H. Carleton, A. D. McKinnon, K. Flynn, and M. Furnas. What determines the growth of tropical reef fish larvae in the plankton: food or temperature? Marine Ecology Progress Series, 256:193–204, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p193.pdf.

Miller:2005:SVM

[MCM05] C. B. Miller, J. A. Crain, and N. H. Marcus. Seasonal variation of male-type antennular setation in female Calanus finmarchicus. Marine Ecology Progress Series, 301:217–229, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/301/m301p217.pdf.

Munday:2009:IEE

[MCN09] P. L. Munday, N. E. Crawley, and G. E. Nilsson. Interacting effects of elevated temperature and ocean acidification on the aerobic performance of coral reef fishes. Marine Ecology Progress Series, 388:235–242, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p235.pdf.

Maranon:2005:CPP

[MCP05] E. Marañón, P. Cermeño, and V. Pérez. Continuity in the photosynthetic production of dissolved organic carbon from eutrophic to oligotrophic waters. Marine Ecology Progress Series, 299:7–17, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p007.pdf.

Martins:2006:MCF

[MCP^{+}06] I. Martins, V. Costa, F. M. Porteiro, A. Colaço, and R. S. Santos. Mercury concentrations in fish species caught at Mid-Atlantic Ridge hydrothermal vent fields. Marine Ecology Progress Series, 320:253–258, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p253.pdf.

Meyer:2009:LTM

[MCP^{+}09] C. G. Meyer, T. B. Clark, Y. P. Papastamatiou, N. M. Whitney, and K. N. Holland. Long-term movement patterns of tiger sharks Galeocerdo cuvier in Hawaii. Marine Ecology Progress Series, 381:223–235, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/381/m381p223.pdf.

McQuoid:2005:ISS

[McQ05] M. R. McQuoid. Influence of salinity on seasonal germination of resting stages and composition of microplankton on the Swedish west coast. Marine Ecology Progress Series, 289:151–163, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p151.pdf.

Malzahn:2003:TEG

[MCR03] A. M. Malzahn, C. Clemmesen, and H. Rosenthal. Temperature effects on growth and nucleic acids in laboratoryreared larval coregonid fish. Marine Ecology Progress Series, 259:285–293, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p285.pdf.

Miller:2005:SVZ

[MCS05] C. J. Miller, S. J. Campbell, and S. Scudds. Spatial variation of Zostera tasmanica morphology and structure across an environmental gradient. Marine Ecology Progress Series, 304:45–53, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p045.pdf.

Maloy:2009:UPD

[MCS09] A. P. Maloy, S. C. Culloty, and J. W. Slater. Use of PCR-DGGE to investigate the trophic ecology of marine suspension feeding bivalves. Marine Ecology Progress Series, 381: 109–118, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/381/m381p109.pdf.

Moreira:2006:SDS

[MCU06] J. Moreira, M. G. Chapman, and A. J. Underwood. Seawalls do not sustain viable populations of limpets. Marine Ecology Progress Series, 322:179–188, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/322/m322p179.pdf.

Martinez-Crego:2008:SMS

[MCVAR08] B. Martínez-Crego, A. Vergés, T. Alcoverro, and J. Romero. Selection of multiple seagrass indicators for environmental biomonitoring. Marine Ecology Progress Series, 361:93–109, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/361/m361p093.pdf.

Malzahn:2007:CNC

[MCW^{+}07] A. M. Malzahn, C. Clemmesen, K. H. Wiltshire, S. Laakmann, and M. Boersma. Comparative nutritional condition of larval dab Limanda limanda and lesser sandeel Ammodytes marinus in a highly variable environment. Marine Ecology Progress Series, 334:205–212, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/334/m334p205.pdf.

Marba:2001:GSS

[MD01] N. Marbà and C. M. Duarte. Growth and sediment space occupation by seagrass Cymodocea nodosa roots. Marine Ecology Progress Series, 224:291–298, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/224/m224p291.pdf.

Miller:2007:SIM

[MD07] H. L. Miller III and K. H. Dunton. Stable isotope (^{13}C) and O_{2} micro-optode alternatives for measuring photosythesis in seaweeds. Marine Ecology Progress Series, 329:85– 97, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p085.pdf.

MacKenzie:2008:HHI

[MD08a] R. A. MacKenzie and M. Dionne. Habitat heterogeneity: importance of salt marsh pools and high marsh surfaces to fish production in two Gulf of Maine salt marshes. Marine Ecology Progress Series, 368:217–230, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/368/m368p217.pdf.

Menden-Deuer:2008:STC

[MD08b] S. Menden-Deuer. Spatial and temporal characteristics of plankton-rich layers in a shallow, temperate fjord. Marine Ecology Progress Series, 355:21–30, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p021.pdf.

Marvin-DiPasquale:2003:BSR

[MDBC03] M. C. Marvin-DiPasquale, W. R. Boynton, and D. G. Capone. Benthic sulfate reduction along the Chesapeake Bay central channel. II. Temporal controls. Marine Ecology Progress Series, 260:55–70, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p055.pdf.

Maire:2007:LFA

[MDBG07] O. Maire, J. Duchêne, L. Bigot, and A. Grémare. Linking feeding activity and sediment reworking in the depositfeeding bivalve Abra ovata with image analysis, laser telemetry, and luminophore tracers. Marine Ecology Progress Series, 351:139–150, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p139.pdf.

Melia:2004:DTD

[MDG04] P. Melià, G. A. De Leo, and M. Gatto. Density and temperature-dependence of vital rates in the Manila clam Tapes philippinarum: a stochastic demographic model. Marine Ecology Progress Series, 272:153–164, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p153.pdf.

Metcalf:2008:ITI

[MDHL08] S. J. Metcalf, J. M. Dambacher, A. J. Hobday, and J. M. Lyle. Importance of trophic information, simplification and aggregation error in ecosystem models. Marine Ecology Progress Series, 360:25–36, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p025.pdf.

Meziane:2006:FMO

[MdL06] T. Meziane, F. d’Agata, and S. Y. Lee. Fate of mangrove organic matter along a subtropical estuary: small-scale exportation and contribution to the food of crab communities. Marine Ecology Progress Series, 312:15–27, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p015.pdf.

Menden-Deuer:2001:EPD

[MDLS01] S. Menden-Deuer, E. J. Lessard, and J. Satterberg. Effect of preservation on dinoflagellate and diatom cell volume, and consequences for carbon biomass predictions. Marine Ecology Progress Series, 222:41–50, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/222/m222p041.pdf.

Michaud:2004:ESD

[MdMR04] M. Michaud, T. de Meeûs, and F. Renaud. Environmental sex determination in a parasitic copepod: checking heterogeneity and unpredictability of the environment. Marine Ecology Progress Series, 269:163–171, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/269/m269p163.pdf.

Maire:2006:EFA

[MDR^{+}06] O. Maire, J. C. Duchêne, R. Rosenberg, J. B. de Mendonça, Jr., and A. Grémare. Effects of food availability on sediment reworking in Abra ovata and A. nitida. Marine Ecology Progress Series, 319:135–153, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/319/m319p135.pdf.

Menge:2007:MZN

[MDS^{+}07] B. A. Menge, B. A. Daley, E. Sanford, E. P. Dahlhoff, and J. Lubchenco. Mussel zonation in New Zealand: an integrative eco-physiological approach. Marine Ecology Progress Series, 345:129–140, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/345/m345p129.pdf.

Migliaccio:2005:MSG

[MDSP05] M. Migliaccio, F. De Martino, F. Silvestre, and G. Procaccini. Meadow-scale genetic structure in Posidonia oceanica. Marine Ecology Progress Series, 304:55–65, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p055.pdf.

Meyer:2009:GVR

[MDW^{+}09] E. Meyer, S. Davies, S. Wang, B. L. Willis, D. Abrego, T. E. Juenger, and M. V. Matz. Genetic variation in responses to a settlement cue and elevated temperature in the reefbuilding coral Acropora millepora. Marine Ecology Progress Series, 392:81–92, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p081.pdf.

Morgan:2005:CRP

[MDZ05] C. A. Morgan, A. De Robertis, and R. W. Zabel. Columbia River plume fronts. I. Hydrography, zooplankton distribution, and community composition. Marine Ecology Progress Series, 299:19–31, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p019.pdf.

Moran:2001:STV

[ME01] X. A. G. Morán and M. Estrada. Short-term variability of photosynthetic parameters and particulate and dissolved primary production in the Alboran Sea (SW Mediterranean). Marine Ecology Progress Series, 212:53–67, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p053.pdf.

Mass:2007:PSP

[MEB^{+}07] T. Mass, S. Einbinder, E. Brokovich, N. Shashar, R. Vago, J. Erez, and Z. Dubinsky. Photoacclimation of Stylophora pistillata to light extremes: metabolism and calcification. Marine Ecology Progress Series, 334:93–102, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p093.pdf.

Mattern:2007:CFR

[MEHD07] T. Mattern, U. Ellenberg, D. M. Houston, and L. S. Davis. Consistent foraging routes and benthic foraging behaviour in yellow-eyed penguins. Marine Ecology Progress Series, 343:295–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p295.pdf.

Mills:2003:MPE

[MF03] K. E. Mills and M. S. Fonseca. Mortality and productivity of eelgrass Zostera marina under conditions of experimental burial with two sediment types. Marine Ecology Progress Series, 255:127–134, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p127.pdf.

Morita:2007:WAS

[MF07] K. Morita and M. Fukuwaka. Why age and size at maturity have changed in Pacific salmon. Marine Ecology Progress Series, 335:289–294, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m335p289.pdf.

Mathews:2008:TTS

[MF08] T. Mathews and N. S. Fisher. Trophic transfer of seven trace metals in a four-step marine food chain. Marine Ecology Progress Series, 367:23–33, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p023.pdf.

McCarthy:2003:AAG

[MFA03] I. D. McCarthy, L. A. Fuiman, and M. C. Alvarez. Aroclor 1254 affects growth and survival skills of Atlantic croaker Micropogonias undulatus larvae. Marine Ecology Progress Series, 252:295–301, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p295.pdf.

Mulholland:2002:PDC

[MFCC02] M. R. Mulholland, S. Floge, E. J. Carpenter, and D. G. Capone. Phosphorus dynamics in cultures and natural populations of Trichodesmium spp. Marine Ecology Progress Series, 239:45–55, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p045.pdf.

MacAvoy:2005:NAA

[MFCM05] S. E. MacAvoy, C. R. Fisher, R. S. Carney, and S. A. Macko. Nutritional associations among fauna at hydrocarbon seep communities in the Gulf of Mexico. Marine Ecology Progress Series, 292:51–60, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/292/m292p051.pdf.

Molina:2005:ACU

[MFE^{+}05] V. Molina, L. Farías, Y. Eissler, L. A. Cuevas, C. E. Morales, and R. Escribano. Ammonium cycling under a strong oxygen gradient associated with the Oxygen Minimum Zone off northern Chile (≈23^{o}S). Marine Ecology Progress Series, 288: 35–43, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/288/m288p035.pdf.

Maneiro:2000:ZPV

[MFG^{+}00] I. Maneiro, M. Frangópulos, C. Guisande, M. Fernández, B. Reguera, and I. Riveiro. Zooplankton as a potential vector of diarrhetic shellfish poisoning toxins through the food web. Marine Ecology Progress Series, 201:155–163, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p155.pdf.

Macleod:2004:SDM

[MFG^{+}04] K. Macleod, R. Fairbairns, A. Gill, B. Fairbairns, J. Gordon, C. Blair-Myers, and E. C. M. Parsons. Seasonal distribution of minke whales Balaenoptera acutorostrata in relation to physiography and prey off the Isle of Mull, Scotland. Marine Ecology Progress Series, 277:263–274, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p263.pdf.

Mathews:2008:ARM

[MFJT08] T. Mathews, N. S. Fisher, R. A. Jeffree, and J. L. Teyssié. Assimilation and retention of metals in teleost and elasmobranch fishes following dietary exposure. Marine Ecology Progress Series, 360:1–12, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Magnusson:2007:LME

[MFK07] S. H. Magnusson, M. Fine, and M. Kühl. Light microclimate of endolithic phototrophs in the scleractinian corals Montipora monasteriata and Porites cylindrica. Marine Ecology Progress Series, 332:119–128, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/332/m332p119.pdf.

Morgan:2009:LRR

[MFM09a] S. G. Morgan, J. L. Fisher, and A. J. Mace. Larval recruitment in a region of strong, persistent upwelling and recruitment limitation. Marine Ecology Progress Series, 394:79–99, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m394p079.pdf.

Morgan:2009:CSD

[MFM^{+}09b] S. G. Morgan, J. L. Fisher, A. J. Mace, L. Akins, A. M. Slaughter, and S. M. Bollens. Cross-shelf distributions and recruitment of crab postlarvae in a region of strong upwelling. Marine Ecology Progress Series, 380:173–185, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m380p173.pdf.

Moran:2004:SFP

[MFP04] X. A. G. Morán, E. Fernández, and V. Pérez. Size-fractionated primary production, bacterial production and net community production in subtropical and tropical domains of the oligotrophic NE Atlantic in autumn. Marine Ecology Progress Series, 274:17–29, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p017.pdf.

Mourino:2005:CEM

[MFP^{+}05] B. Mouriño, E. Fernández, R. Pingree, B. Sinha, J. Escánez, and D. de Armas. Constraining effect of mesoscale features on carbon budget of photic layer in the NE subtropical Atlantic. Marine Ecology Progress Series, 287:45–52, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p045.pdf.

Mistri:2001:RML

[MFR01] M. Mistri, E. A. Fano, and R. Rossi. Redundancy of macrobenthos from lagoonal habitats in the Adriatic Sea. Marine Ecology Progress Series, 215:289–296, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p289.pdf.

Martin:2007:SCF

[MFS07] M. A. Martin, J. P. Fram, and M. T. Stacey. Seasonal chlorophyll-a fluxes between the coastal Pacific Ocean and San Francisco Bay. Marine Ecology Progress Series, 337: 51–61, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p051.pdf.

Madden:2008:IFD

[MFWD08] K. M. Madden, L. A. Fuiman, T. M. Williams, and R. W. Davis. Identification of foraging dives in free-ranging Weddell seals Leptonychotes weddellii: confirmation using video records. Marine Ecology Progress Series, 365:263–275, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/365/m365p263.pdf.

McCarthy:2002:OPG

[MFY02] D. A. McCarthy, R. B. Forward, Jr., and C. M. Young. Ontogeny of phototaxis and geotaxis during larval development of the sabellariid polychaete Phragmatopoma lapidosa. Marine Ecology Progress Series, 241:215–220, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p215.pdf.

Mock:2000:CPC

[MG00] T. Mock and R. Gradinger. Changes in photosynthetic carbon allocation in algal assemblages of Arctic sea ice with decreasing nutrient concentrations and irradiance. Marine Ecology Progress Series, 202:1–11, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/202/m202p001.pdf.

McClanahan:2005:RTC

[MG05a] T. R. McClanahan and N. A. J. Graham. Recovery trajectories of coral reef fish assemblages within Kenyan marine protected areas. Marine Ecology Progress Series, 294: 241–248, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/294/m294p241.pdf.

Musco:2005:MSA

[MG05b] L. Musco and A. Giangrande. Mediterranean syllidae (Annelida: Polychaeta) revisited: biogeography, diversity and species fidelity to environmental features. Marine Ecology Progress Series, 304:143–153, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/304/m304p143.pdf.

Mass:2008:EVI

[MG08] T. Mass and A. Genin. Environmental versus intrinsic determination of colony symmetry in the coral Pocillopora verrucosa. Marine Ecology Progress Series, 369:131–137, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/369/m369p131.pdf.

Mann:2009:STU

[MG09a] D. A. Mann and T. M. Grothues. Short-term upwelling events modulate fish sound production at a mid-Atlantic Ocean observatory. Marine Ecology Progress Series, 375: 65–71, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/375/m375p065.pdf.

Minutoli:2009:ZRE

[MG09b] R. Minutoli and L. Guglielmo. Zooplankton respiratory electron transport system (ETS) activity in the Mediterranean Sea: spatial and diel variability. Marine Ecology Progress Series, 381:199–211, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p199.pdf.

Medernach:2001:TCA

[MGA^{+}01] L. Medernach, A. Grémare, J. M. Amouroux, J. C. Colomines, and G. Vétion. Temporal changes in the amino acid contents of particulate organic matter sedimenting in the Bay of Banyuls-sur-Mer (northwestern Mediterranean). Marine Ecology Progress Series, 214:55–65, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/214/m214p055.pdf.

McDougall:2006:CLC

[MGBB06] K. E. McDougall, S. W. Gibb, K. G. Boyd, and B. E. Brown. ‘Chlorophyll-like’ compounds as novel biomarkers of stress in corals. Marine Ecology Progress Series, 325:137– 144, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/325/m325p137.pdf.

McCay:2003:SRA

[MGC03] D. P. French McCay, M. Gibson, and J. S. Cobb. Scaling restoration of American lobsters: combined demographic and discounting model for an exploited species. Marine Ecology Progress Series, 264:177–196, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p177.pdf.

McDonald:2009:EMT

[MGC^{+}09] B. I. McDonald, M. E. Goebel, D. E. Crocker, Y. Tremblay, and D. P. Costa. Effects of maternal traits and individual behavior on the foraging strategies and provisioning rates of an income breeder, the Antarctic fur seal. Marine Ecology Progress Series, 394:277–288, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p277.pdf.

McKinnon:2009:DSS

[MGD^{+}09] J. G. McKinnon, P. E. Gribben, A. R. Davis, D. F. Jolley, and J. T. Wright. Differences in soft-sediment macrobenthic assemblages invaded by Caulerpa taxifolia compared to uninvaded habitats. Marine Ecology Progress Series, 380:59–71, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p059.pdf.

Mazouni:2001:CBC

[MGDP01] N. Mazouni, J. C. Gaertner, and J. M. Deslous-Paoli. Composition of biofouling communities on suspended oyster cultures: an in situ study of their interactions with the water column. Marine Ecology Progress Series, 214:93–102, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
214/m214p093.pdf.

Marubini:2009:HPI

[MGE^{+}09] F. Marubini, A. Gimona, P. G. H. Evans, P. J. Wright, and G. J. Pierce. Habitat preferences and interannual variability in occurrence of the harbour porpoise Phocoena phocoena off northwest Scotland. Marine Ecology Progress Series, 381:297–310, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p297.pdf.

Meiners:2003:VDE

[MGF^{+}03] K. Meiners, R. Gradinger, J. Fehling, G. Civitarese, and M. Spindler. Vertical distribution of exopolymer particles in sea ice of the Fram Strait (Arctic) during autumn. Marine Ecology Progress Series, 248:1–13, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/248/
m248p001.pdf.

Myrand:2000:RSM

[MGH00] B. Myrand, H. Guderley, and J. H. Himmelman. Reproduction and summer mortality of blue mussels Mytilus edulis in the Magdalen Islands, southern Gulf of St. Lawrence. Marine Ecology Progress Series, 197:193–207, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p193.pdf.

Montiel:2005:PAM

[MGHA05] A. Montiel, D. Gerdes, B. Hilbig, and W. E. Arntz. Polychaete assemblages on the Magellan and Weddell Sea shelves: comparative ecological evaluation. Marine Ecology Progress Series, 297:189–202, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/297/m297p189.pdf.

Mousseau:2000:EUB

[MGL^{+}00] L. Mousseau, M. Gosselin, M. Levasseur, S. Demers, J. Fauchot, S. Roy, P. Z. Villegas, and B. Mostajir. Effects of ultraviolet-B radiation on simultaneous carbon and nitrogen transport rates by estuarine phytoplankton during a weeklong mesocosm study. Marine Ecology Progress Series, 199: 69–81, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/199/m199p069.pdf.

Malcolm:2007:STV

[MGL^{+}07] H. A. Malcolm, W. Gladstone, S. Lindfield, J. Wraith, and T. P. Lynch. Spatial and temporal variation in reef fish assemblages of marine parks in New South Wales, Australia — baited video observations. Marine Ecology Progress Series, 350:277–290, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m350p277.pdf.

Meysman:2005:IPP

[MGM05] F. J. R. Meysman, O. S. Galaktionov, and J. J. Middelburg. Irrigation patterns in permeable sediments induced by burrow ventilation: a case study of Arenicola marina. Marine Ecology Progress Series, 303:195–212, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/303/m303p195.pdf.

McClanahan:2007:IIV

[MGM^{+}07] T. R. McClanahan, N. A. J. Graham, J. Maina, P. Chabanet, J. H. Bruggemann, and N. V. C. Polunin. Influence of instantaneous variation on estimates of coral reef fish populations and communities. Marine Ecology Progress Series, 340:221–234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p221.pdf.

Murdoch:2007:LSD

[MGO^{+}07] T. J. T. Murdoch, A. F. Glasspool, M. Outerbridge, J. Ward, S. Manuel, J. Gray, A. Nash, K. A. Coates, J. Pitt, J. W. Fourqurean, P. A. Barnes, M. Vierros, K. Holzer, and S. R. Smith. Large-scale decline in offshore seagrass meadows in Bermuda. Marine Ecology Progress Series, 339:123–130, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/339/m339p123.pdf.

Moran:2001:DPP

[MGPAE01] X. A. G. Morán, J. M. Gasol, C. Pedrós-Alió, and M. Estrada. Dissolved and particulate primary production and bacterial production in offshore Antarctic waters during austral summer: coupled or uncoupled? Marine Ecology Progress Series, 222:25–39, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p025.pdf.

McQuatters-Gollop:2007:SPD

[MGREA07] A. McQuatters-Gollop, D. E. Raitsos, M. Edwards, and M. J. Attrill. Spatial patterns of diatom and dinoflagellate seasonal cycles in the NE Atlantic Ocean. Marine Ecology Progress Series, 339:301–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/339/m339p301.pdf.

Moriceau:2007:ERB

[MGRP07] B. Moriceau, M. Garvey, O. Ragueneau, and U. Passow. Evidence for reduced biogenic silica dissolution rates in diatom aggregates. Marine Ecology Progress Series, 333:129– 142, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p129.pdf.

Mountain:2008:GMA

[MGSJ08] D. Mountain, J. Green, J. Sibunka, and D. Johnson. Growth and mortality of Atlantic cod Gadus morhua and haddock Melanogrammus aeglefinus eggs and larvae on Georges Bank, 1995 to 1999. Marine Ecology Progress Series, 353:225– 242, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/353/m353p225.pdf.

Meng:2000:UWF

[MGTK00] L. Meng, C. Gray, B. Taplin, and E. Kupcha. Using winter flounder growth rates to assess habitat quality in Rhode Island’s coastal lagoons. Marine Ecology Progress Series, 201:287–299, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p287.pdf.

McClanahan:2009:EFC

[MGW^{+}09] T. R. McClanahan, N. A. J. Graham, S. K. Wilson, Y. Letourneur, and R. Fisher. Effects of fisheries closure size, age, and history of compliance on coral reef fish communities in the western Indian Ocean. Marine Ecology Progress Series, 396:99–109, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m396p099.pdf.

Miralto:2003:IPG

[MGZ^{+}03] A. Miralto, L. Guglielmo, G. Zagami, I. Buttino, A. Granata, and A. Ianora. Inhibition of population growth in the copepods Acartia clausi and Calanus helgolandicus during diatom blooms. Marine Ecology Progress Series, 254:253– 268, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p253.pdf.

Maxwell:2000:FRO

[MH00] M. R. Maxwell and R. T. Hanlon. Female reproductive output in the squid Loligo pealeii: multiple egg clutches and implications for a spawning strategy. Marine Ecology Progress Series, 199:159–170, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p159.pdf.

Moksnes:2006:RIH

[MH06a] P. O. Moksnes and K. L. Heck, Jr. Relative importance of habitat selection and predation for the distribution of blue crab megalopae and young juveniles. Marine Ecology Progress Series, 308:165–181, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/308/m308p165.pdf.

Myers:2006:DLT

[MH06b] A. E. Myers and G. C. Hays. Do leatherback turtles Dermochelys coriacea forage during the breeding season? A combination of data-logging devices provide new insights. Marine Ecology Progress Series, 322:259–267, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p259.pdf.

Middelboe:2007:HPS

[MH07] A. L. Middelboe and P. J. Hansen. High pH in shallowwater macroalgal habitats. Marine Ecology Progress Series, 338:107–117, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p107.pdf.

Mallela:2008:RSD

[MH08a] J. Mallela and C. Harrod. δ ^{13}C and δ ^{15}N reveal significant differences in the coastal foodwebs of the seas surrounding Trinidad and Tobago. Marine Ecology Progress Series, 368:41–51, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p041.pdf.

Mangubhai:2008:ACS

[MH08b] Sangeeta Mangubhai and Peter L. Harrison. Asynchronous coral spawning patterns on equatorial reefs in Kenya. Marine Ecology Progress Series, 360:85–96, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/360/

m360p085.pdf. See comment [BG09] and reply [MH09].

McKay:2008:PJC

[MH08c] K. M. McKay and K. L. Heck, Jr. Presence of the Jonah crab Cancer borealis significantly reduces kelp consumption by the green sea urchin Strongylocentrotus droebachiensis. Marine Ecology Progress Series, 356:295–298, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/356/
m356p295.pdf.

Mangubhai:2009:RCE

[MH09] S. Mangubhai and P. L. Harrison. Reply comment: Extended breeding seasons and asynchronous spawning among equatorial reef corals in Kenya. Marine Ecology Progress Series, 374:305–310, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/374/m374p305.pdf. See

[MH08b, BG09].

Montes-Hugo:2004:APP

[MHABGC04] M. A. Montes-Hugo, S. Alvarez-Borrego, and G. Gaxiola-Castro. Annual phytoplankton production in a coastal lagoon of the southern California Current system. Marine Ecology Progress Series, 277:51–60, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p051.pdf.

Manriquez:2001:ADL

[MHB01a] P. H. Manríquez, R. N. Hughes, and J. D. D. Bishop. Agedependent loss of fertility in water-borne sperm of the bryozoan Celleporella hyalina. Marine Ecology Progress Series, 224:87–92, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p087.pdf.

Maranon:2001:PPS

[MHB^{+}01b] E. Marañón, P. M. Holligan, R. Barciela, N. González, B. Mouriño, M. J. Pazó, and M. Varela. Patterns of phytoplankton size structure and productivity in contrasting openocean environments. Marine Ecology Progress Series, 216: 43–56, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/216/m216p043.pdf.

Massom:2009:FID

[MHB^{+}09] R. A. Massom, K. Hill, C. Barbraud, N. Adams, A. Ancel, L. Emmerson, and M. J. Pook. Fast ice distribution in Adélie Land, East Antarctica: interannual variability and implications for emperor penguins Aptenodytes forsteri. Marine Ecology Progress Series, 374:243–257, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p243.pdf.

McMahon:2005:CSE

[MHBB05] C. R. McMahon, M. A. Hindell, H. R. Burton, and M. N. Bester. Comparison of southern elephant seal populations, and observations of a population on a demographic knife-edge. Marine Ecology Progress Series, 288:273–283, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p273.pdf.

Marion:2008:FEP

[MHCB08] A. Marion, M. Harvey, D. Chabot, and J. C. Brêthes. Feeding ecology and predation impact of the recently established amphipod, Themisto libellula, in the St. Lawrence marine system, Canada. Marine Ecology Progress Series, 373:53– 70, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/373/m373p053.pdf.

Montes-Hugo:2009:CDM

[MHDS09] M. A. Montes-Hugo, H. Ducklow, and O. M. Schofield. Contribution by different marine bacterial communities to particulate beam attenuation. Marine Ecology Progress Series, 379: 13–22, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/379/m379p013.pdf.

Mariani:2005:NSH

[MHH^{+}05] S. Mariani, W. F. Hutchinson, E. M. C. Hatfield, D. E. Ruzzante, E. J. Simmonds, T. G. Dahlgren, C. Andre, J. Brigham, E. Torstensen, and G. R. Carvalho. North Sea herring population structure revealed by microsatellite analysis. Marine Ecology Progress Series, 303:245–257, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p245.pdf.

Mowitt:2006:GPB

[MHHS06] W. P. Mowitt, E. D. Houde, D. C. Hinkle, and A. Sanford. Growth of planktivorous bay anchovy Anchoa mitchilli, topdown control, and scale-dependence in estuarine mesocosms. Marine Ecology Progress Series, 308:255–269, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p255.pdf.

Meesters:2001:CSF

[MHK^{+}01] E. H. Meesters, M. Hilterman, E. Kardinaal, M. Keetman, M. de Vries, and R. P. M. Bak. Colony size-frequency distributions of scleractinian coral populations: spatial and interspecific variation. Marine Ecology Progress Series, 209: 43–54, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/209/m209p043.pdf.

Mortimer:2004:ANM

[MHK^{+}04] R. J. G. Mortimer, S. J. Harris, M. D. Krom, T. E. Freitag, J. I. Prosser, J. Barnes, P. Anschutz, P. J. Hayes, and I. M. Davies. Anoxic nitrification in marine sediments. Marine Ecology Progress Series, 276:37–52, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p037.pdf.

Marba:2002:CNT

[MHM^{+}02] N. Marbà, M. A. Hemminga, M. A. Mateo, M. D. Duarte, Y. E. M. Mass, J. Terrados, and E. Gacia. Carbon and nitrogen translocation between seagrass ramets. Marine Ecology Progress Series, 226:287–300, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/226/m226p287.pdf.

Mattila:2008:IHS

[MHM^{+}08] J. Mattila, K. L. Heck, Jr., E. Millstein, E. Miller, C. Gustafsson, S. Williams, and D. Byron. Increased habitat structure does not always provide increased refuge from predation. Marine Ecology Progress Series, 361:15–20, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p015.pdf.

Mangubhai:2007:PCL

[MHO07] S. Mangubhai, P. L. Harrison, and D. O. Obura. Patterns of coral larval settlement on lagoon reefs in the Mombasa Marine National Park and Reserve, Kenya. Marine Ecology Progress Series, 348:149–159, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/348/m348p149.pdf.

MacKenzie:2000:QEH

[MHP^{+}00] B. R. MacKenzie, H. H. Hinrichsen, M. Plikshs, K. Wieland, and A. S. Zezera. Quantifying environmental heterogeneity: habitat size necessary for successful development of cod Gadus morhua eggs in the Baltic Sea. Marine Ecology Progress Series, 193:143–156, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p143.pdf.

Maravelias:2003:PDF

[MHP03] C. D. Maravelias, J. Haralabous, and C. Papaconstantinou. Predicting demersal fish species distributions in the Mediterranean Sea using artificial neural networks. Marine Ecology Progress Series, 255:249–258, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/255/m255p249.pdf.

Meyer:2007:SDM

[MHP07] C. G. Meyer, K. N. Holland, and Y. P. Papastamatiou. Seasonal and diel movements of giant trevally Caranx ignobilis at remote Hawaiian atolls: implications for the design of Marine Protected Areas. Marine Ecology Progress Series, 333: 13–25, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p013.pdf.

Moksnes:2003:SBS

[MHR03] Per-Olav Moksnes, Ola Hedvall, and Torsten Reinwald. Settlement behavior in shore crabs Carcinus maenas: why do postlarvae emigrate from nursery habitats? Marine Ecology Progress Series, 250:215–230, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/250/m250p215.pdf.

Martinson:2008:DDG

[MHSS08] E. C. Martinson, J. H. Helle, D. L. Scarnecchia, and H. H. Stokes. Density-dependent growth of Alaska sockeye salmon in relation to climate-oceanic regimes, population abundance, and body size, 1925 to 1998. Marine Ecology Progress Series, 370:1–18, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moore:2007:RBH

[MHT07] P. Moore, S. J. Hawkins, and R. C. Thompson. Role of biological habitat amelioration in altering the relative responses of congeneric species to climate change. Marine Ecology Progress Series, 334:11–19, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/334/m334p011.pdf.

Martins:2007:CSF

[MHTJ07] G. M. Martins, S. J. Hawkins, R. C. Thompson, and S. R. Jenkins. Community structure and functioning in intertidal rock pools: effects of pool size and shore height at different successional stages. Marine Ecology Progress Series, 329: 43–55, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/329/m329p043.pdf.

Moens:2000:PRP

[MHV^{+}00] T. Moens, P. Herman, L. Verbeeck, M. Steyaert, and M. Vincx. Predation rates and prey selectivity in two predacious estuarine nematode species. Marine Ecology Progress Series, 205:185–193, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/205/m205p185.pdf.

Mann:2006:OIS

[MHW06] R. Mann, J. M. Harding, and E. Westcott. Occurrence of imposex and seasonal patterns of gameteogenesis in the invading veined rapa whelk Rapana venosa from Chesapeake Bay, USA. Marine Ecology Progress Series, 310:129–138, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p129.pdf.

Miles:2007:HES

[MHW07] C. M. Miles, M. G. Hadfield, and M. L. Wayne. Heritability for egg size in the serpulid polychaete Hydroides elegans. Marine Ecology Progress Series, 340:155–162, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p155.pdf.

Miao:2006:MIL

[MHY^{+}06] A. J. Miao, D. A. Hutchins, K. Yin, F. X. Fu, P. J. Harrison, and W. X. Wang. Macronutrient and iron limitation of phytoplankton growth in Hong Kong coastal waters. Marine Ecology Progress Series, 318:141–152, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/318/m318p141.pdf.

Mortazavi:2000:NBA

[MIH^{+}00] B. Mortazavi, R. L. Iverson, W. Huang, F. G. Lewis, and J. M. Caffrey. Nitrogen budget of Apalachicola Bay, a bar-built estuary in the northeastern Gulf of Mexico. Marine Ecology Progress Series, 195:1–14, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p001.pdf.

Mortazavi:2001:DON

[MIH01] B. Mortazavi, R. L. Iverson, and W. Huang. Dissolved organic nitrogen and nitrate in Apalachicola Bay, Florida: spatial distributions and monthly budgets. Marine Ecology Progress Series, 214:79–91, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/214/m214p079.pdf.

Mortazavi:2000:CPP

[MIL^{+}00] B. Mortazavi, R. L. Iverson, W. M. Landing, F. G. Lewis, and W. Huang. Control of phytoplankton production and biomass in a river-dominated estuary: Apalachicola Bay, Florida, USA. Marine Ecology Progress Series, 198:19–31, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p019.pdf.

Miller:2007:FEF

[Mil07a] L. P. Miller. Feeding in extreme flows: behavior compensates for mechanical constraints in barnacle cirri. Marine Ecology Progress Series, 349:227–234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p227.pdf.

Miller:2007:CIB

[Mil07b] T. J. Miller. Contribution of individual-based coupled physical-biological models to understanding recruitment in marine fish populations. Marine Ecology Progress Series, 347:127–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m347p127.pdf.

Mortazavi:2000:PBA

[MILH00] B. Mortazavi, R. L. Iverson, W. M. Landing, and W. Huang. Phosphorus budget of Apalachicola Bay: a river-dominated estuary in the northeastern Gulf of Mexico. Marine Ecology Progress Series, 198:33–42, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/198/m198p033.pdf.

Minchinton:2002:PDN

[Min02] T. E. Minchinton. Precipitation during El Niño correlates with increasing spread of Phragmites australis in New England, USA, coastal marshes. Marine Ecology Progress Series, 242:305–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p305.pdf.

Mistri:2004:EHP

[Mis04] M. Mistri. Effects of hypoxia on predator–prey interactions between juvenile Carcinus aestuarii and Musculista senhousia.

Marine Ecology Progress Series, 275:211–217, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p211.pdf.

Miller:2007:CFS

[MJ07] W. D. Miller and L. W. Harding Jr. Climate forcing of the spring bloom in Chesapeake Bay. Marine Ecology Progress Series, 331:11–22, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m331p011.pdf.

Medernach:2000:PDS

[MJG^{+}00] L. Medernach, E. Jordana, A. Grémare, C. Nozais, F. Charles, and J. M. Amouroux. Population dynamics, secondary production and calcification in a Mediterranean population of Ditrupa arietina (Annelida: Polychaeta). Marine Ecology Progress Series, 199:171–184, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/199/m199p171.pdf.

Meyer:2001:CNU

[MJG01] D. L. Meyer, J. M. Johnson, and J. W. Gill. Comparison of nekton use of Phragmites australis and Spartina alterniflora marshes in the Chesapeake Bay, USA. Marine Ecology Progress Series, 209:71–83, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p071.pdf.

Mills:2008:BLT

[MJG08] D. J. Mills, C. R. Johnson, and C. Gardner. Bias in lobster tethering experiments conducted for selecting low-predation release sites. Marine Ecology Progress Series, 364:1–13, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moran:2003:RPD

[MJKH03] S. M. Moran, G. P. Jenkins, M. J. Keough, and J. S. Hindell. Role of physical disturbance in structuring fish assemblages in seagrass beds in Port Phillip Bay, Australia. Marine Ecology Progress Series, 251:127–139, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/251/m251p127.pdf.

Moran:2004:SSD

[MJKH04] S. M. Moran, G. P. Jenkins, M. J. Keough, and J. S. Hindell. Small-scale dynamics of secondary dispersal in a seagrass associated fish: a caging study. Marine Ecology Progress Series, 272:271–280, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/272/m272p271.pdf.

Marino:2007:ECA

[MJS07] M. C. Marino II, F. Juanes, and K. D. E. Stokesbury. Effect of closed areas on populations of sea star Asterias spp. on Georges Bank. Marine Ecology Progress Series, 347: 39–49, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/347/m347p039.pdf.

Marino:2009:STV

[MJS09] M. C. Marino II, F. Juanes, and K. D. E. Stokesbury. Spatio-temporal variations of sea star Asterias spp. distributions between sea scallop Placopecten magellanicus beds on Georges Bank. Marine Ecology Progress Series, 382: 59–68, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/382/m382p059.pdf.

Morris:2002:OPE

[MK02] L. Morris and M. J. Keough. Organic pollution and its effects: a short-term transplant experiment to assess the ability of biological endpoints to detect change in a soft sediment environment. Marine Ecology Progress Series, 225:109–121, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p109.pdf.

Marshall:2003:ESS

[MK03a] D. J. Marshall and M. J. Keough. Effects of settler size and density on early post-settlement survival of Ciona intestinalis in the field. Marine Ecology Progress Series, 259: 139–144, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/259/m259p139.pdf.

Marshall:2003:VDP

[MK03b] D. J. Marshall and M. J. Keough. Variation in the dispersal potential of non-feeding invertebrate larvae: the desperate larva hypothesis and larval size. Marine Ecology Progress Series, 255:145–153, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p145.pdf.

Morris:2003:VRI

[MK03c] L. Morris and M. J. Keough. Variation in the response of intertidal infaunal invertebrates to nutrient additions: field manipulations at two sites within Port Phillip Bay, Australia. Marine Ecology Progress Series, 250:35–49, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/250/
m250p035.pdf.

Marshall:2004:VEL

[MK04a] D. J. Marshall and M. J. Keough. Variable effects of larval size on post-metamorphic performance in the field. Marine Ecology Progress Series, 279:73–80, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/279/m279p073.pdf.

Marshall:2004:WGG

[MK04b] D. J. Marshall and M. J. Keough. When the going gets rough: effect of maternal size manipulation on larval quality. Marine Ecology Progress Series, 272:301–305, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p301.pdf.

Machias:2005:RDF

[MKG^{+}05] A. Machias, I. Karakassis, M. Giannoulaki, K. N. Papadopoulou, C. J. Smith, and S. Somarakis. Response of demersal fish communities to the presence of fish farms. Marine Ecology Progress Series, 288:241–250, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p241.pdf.

Mortimer:2002:ESN

[MKH^{+}02] R. J. G. Mortimer, M. D. Krom, S. J. Harris, P. J. Hayes, I. M. Davies, W. Davison, and H. Zhang. Evidence for suboxic nitrification in recent marine sediments. Marine Ecology Progress Series, 236:31–35, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/236/m236p031.pdf.

Miyake:2009:SLD

[MKK^{+}09] Y. Miyake, S. Kimura, T. Kawamura, T. Horii, H. Kurogi, and T. Kitagawa. Simulating larval dispersal processes for abalone using a coupled particle-tracking and hydrodynamic model: implications for refugium design. Marine Ecology Progress Series, 387:205–222, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/387/m387p205.pdf.

Molis:2008:SIS

[MKKK08] M. Molis, J. Körner, Y. W. Ko, and J. H. Kim. Specificity of inducible seaweed anti-herbivory defences depends on identity of macroalgae and herbivores. Marine Ecology Progress Series, 354:97–105, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p097.pdf.

May:2003:EST

[MKL^{+}03] C. L. May, J. R. Koseff, L. V. Lucas, J. E. Cloern, and D. H. Schoellhamer. Effects of spatial and temporal variability of turbidity on phytoplankton blooms. Marine Ecology Progress Series, 254:111–128, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/254/m254p111.pdf.

Mollet:2007:FIE

[MKR07] F. M. Mollet, S. B. M. Kraak, and A. D. Rijnsdorp. Fisheriesinduced evolutionary changes in maturation reaction norms in North Sea sole Solea solea. Marine Ecology Progress Series, 351:189–199, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p189.pdf.

Moku:2000:FHT

[MKWO00] M. Moku, K. Kawaguchi, H. Watanabe, and A. Ohno. Feeding habits of three dominant myctophid fishes, Diaphus theta, Stenobrachius leucopsarus and S. nannochir, in the subarctic and transitional waters of the western North Pacific. Marine Ecology Progress Series, 207:129–140, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p129.pdf.

McQuaid:2000:EWE

[ML00] C. D. McQuaid and T. L. Lindsay. Effect of wave exposure on growth and mortality rates of the mussel Perna perna: bottom-up regulation of intertidal populations. Marine Ecology Progress Series, 206:147–154, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/206/m206p147.pdf.

McQuaid:2005:IEW

[ML05] C. D. McQuaid and J. R. Lindsay. Interacting effects of wave exposure, tidal height and substratum on spatial variation in densities of mussel Perna perna plantigrades. Marine Ecology Progress Series, 301:173–184, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/301/m301p173.pdf.

Maricle:2006:EES

[ML06a] Brian R. Maricle and Raymond W. Lee. Effects of environmental salinity on carbon isotope discrimination and stomatal conductance in Spartina grasses. Marine Ecology Progress Series, 313:305–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/313/m313p305.pdf.

Methratta:2006:SVG

[ML06b] E. T. Methratta and J. S. Link. Seasonal variation in groundfish habitat associations in the Gulf of Maine–Georges Bank region. Marine Ecology Progress Series, 326:245–256, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/326/m326p245.pdf.

Methratta:2007:OVH

[ML07] E. T. Methratta and J. S. Link. Ontogenetic variation in habitat association for four groundfish species in the Gulf of Maine–Georges Bank region. Marine Ecology Progress Series, 338:169–181, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p169.pdf.

McGee:2008:LBD

[MLC08] D. McGee, R. A. Laws, and L. B. Cahoon. Live benthic diatoms from the upper continental slope: extending the limits of marine primary production. Marine Ecology Progress Series, 356:103–112, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p103.pdf.

Mei:2003:PPN

[MLG^{+}03a] Z. P. Mei, L. Legendre, Y. Gratton, J. E. Tremblay, B. LeBlanc, B. Klein, and M. Gosselin. Phytoplankton production in the North Water Polynya: size-fractions and carbon fluxes, April to July 1998. Marine Ecology Progress Series, 256:13–27, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p013.pdf.

Mulholland:2003:EEA

[MLG03b] M. R. Mulholland, C. Lee, and P. M. Glibert. Extracellular enzyme activity and uptake of carbon and nitrogen along an estuarine salinity and nutrient gradient. Marine Ecology Progress Series, 258:3–17, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/258/m258p003.pdf.

Maros:2003:SDO

[MLGG03] A. Maros, A. Louveaux, M. H. Godfrey, and M. Girondot. Scapteriscus didactylus (Orthoptera, Gryllotalpidae), predator of leatherback turtle eggs in French Guiana. Marine Ecology Progress Series, 249:289–296, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p289.pdf.

Moens:2002:TOM

[MLM^{+}02] T. Moens, C. Luyten, J. J. Middelburg, P. M. J. Herman, and M. Vincx. Tracing organic matter sources of estuarine tidal flat nematodes with stable carbon isotopes. Marine Ecology Progress Series, 234:127–137, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/234/m234p127.pdf.

Mulsow:2002:BMR

[MLR02] S. Mulsow, P. F. Landrum, and J. A. Robbins. Biological mixing responses to sublethal concentrations of DDT in sediments by Heteromastus filiformis using a 137Cs marker layer technique. Marine Ecology Progress Series, 239:181– 191, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/239/m239p181.pdf.

Menard:2006:OPT

[MLS^{+}06] F. Ménard, C. Labrune, Y. J. Shin, A. S. Asine, and F. X. Bard. Opportunistic predation in tuna: a size-based approach. Marine Ecology Progress Series, 323:223–231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p223.pdf.

Merzouk:2008:BDM

[MLS^{+}08] A. Merzouk, M. Levasseur, M. Scarratt, S. Michaud, M. Lizotte, R. B. Rivkin, and R. P. Kiene. Bacterial DMSP metabolism during the senescence of the spring diatom bloom in the Northwest Atlantic. Marine Ecology Progress Series, 369:1–11, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Marinelli:2002:EIC

[MLW^{+}02] R. L. Marinelli, C. R. Lovell, S. G. Wakeham, D. B. Ringelberg, and D. C. White. Experimental investigation of the control of bacterial community composition in macrofaunal burrows. Marine Ecology Progress Series, 235:1–13, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/235/m235p001.pdf.

Molis:2003:REA

[MLW03] M. Molis, M. Lenz, and M. Wahl. Radiation effects along a UV-b gradient on species composition and diversity of a shallow-water macrobenthic community in the western Baltic. Marine Ecology Progress Series, 263:113–125, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p113.pdf.

McEdward:2003:FTM

[MM03] L. R. McEdward and B. G. Miner. Fecundity-time models of reproductive strategies in marine benthic invertebrates: fitness differences under fluctuating environmental conditions. Marine Ecology Progress Series, 256:111–121, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p111.pdf.

Matozzo:2005:NII

[MM05] V. Matozzo and M. G. Marin. 4-nonylphenol induces immunomodulation and apoptotic events in the clam Tapes philippinarum. Marine Ecology Progress Series, 285:97–106, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/285/m285p097.pdf.

Mace:2006:BPC

[MM06a] A. J. Mace and S. G. Morgan. Biological and physical coupling in the lee of a small headland: contrasting transport mechanisms for crab larvae in an upwelling region. Marine Ecology Progress Series, 324:185–196, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/324/m324p185.pdf.

Mace:2006:LAL

[MM06b] A. J. Mace and S. G. Morgan. Larval accumulation in the lee of a small headland: implications for the design of marine reserves. Marine Ecology Progress Series, 318: 19–29, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/318/m318p019.pdf.

Marshall:2007:IPG

[MM07a] C. T. Marshall and B. J. McAdam. Integrated perspectives on genetic and environmental effects on maturation can reduce potential for errors of inference. Marine Ecology Progress Series, 335:301–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m335p301.pdf.

Miller:2007:SAL

[MM07b] M. J. Miller and J. D. McCleave. Species assemblages of leptocephali in the southwestern Sargasso Sea. Marine Ecology Progress Series, 344:197–212, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/344/m344p197.pdf.

Minchinton:2008:NEA

[MM08] T. E. Minchinton and L. A. McKenzie. Nutrient enrichment affects recruitment of oysters and barnacles in a mangrove forest. Marine Ecology Progress Series, 354:181–189, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/354/m354p181.pdf.

Montani:2003:SIP

[MMA03a] S. Montani, P. Magni, and N. Abe. Seasonal and interannual patterns of intertidal microphytobenthos in combination with laboratory and areal production estimates. Marine Ecology Progress Series, 249:79–91, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/249/m249p079.pdf.

Moore:2003:RBS

[MMA03b] T. S. Moore II, J. Marra, and A. Alkatiri. Response of the Banda Sea to the southeast monsoon. Marine Ecology Progress Series, 261:41–49, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p041.pdf.

Minagawa:2004:CAL

[MMA^{+}04] G. Minagawa, M. J. Miller, J. Aoyama, S. Wouthuyzen, and K. Tsukamoto. Contrasting assemblages of leptocephali in the western Pacific. Marine Ecology Progress Series, 271:

245–259, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/271/m271p245.pdf.

McDonald:2009:EOA

[MMA^{+}09] M. R. McDonald, J. B. McClintock, C. D. Amsler, D. Rittschof, R. A. Angus, B. Orihuela, and K. Lutostanski. Effects of ocean acidification over the life history of the barnacle Amphibalanus amphitrite. Marine Ecology Progress Series, 385:179–187, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m385p179.pdf.

Moodley:2002:BFK

[MMB^{+}02] L. Moodley, J. J. Middelburg, H. T. S. Boschker, G. C. A. Duineveld, R. Pel, P. M. J. Herman, and C. H. R. Heip. Bacteria and Foraminifera: key players in a short-term deep-sea benthic response to phytodetritus. Marine Ecology Progress Series, 236:23–29, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p023.pdf. See erratum

[Ano02f].

MacKenzie:2003:SRR

[MMB03] B. R. MacKenzie, R. A. Myers, and K. G. Bowen. Spawnerrecruit relationships and fish stock carrying capacity in aquatic ecosystems. Marine Ecology Progress Series, 248: 209–220, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p209.pdf.

Musyl:2009:FIM

[MMBF09] M. K. Musyl, C. D. Moyes, R. W. Brill, and N. M. Fragoso. Factors influencing mortality estimates in post-release survival studies. Marine Ecology Progress Series, 396:157–159, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/396/m396p157.pdf.

Mfilinge:2003:FAD

[MMBT03] P. L. Mfilinge, T. Meziane, Z. Bachok, and M. Tsuchiya. Fatty acids in decomposing mangrove leaves: microbial activity, decay and nutritional quality. Marine Ecology Progress Series, 265:97–105, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p097.pdf.

Mayor:2007:CIA

[MMC^{+}07] D. J. Mayor, C. Matthews, K. Cook, A. F. Zuur, and S. Hay. CO_{2}-induced acidification affects hatching success in Calanus finmarchicus. Marine Ecology Progress Series, 350: 91–97, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/350/m350p091.pdf.

Macleod:2008:EFC

[MMC08a] C. K. Macleod, N. A. Moltschaniwskyj, and C. M. Crawford. Ecological and functional changes associated with long-term recovery from organic enrichment. Marine Ecology Progress Series, 365:17–24, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p017.pdf.

Meynier:2008:TSD

[MMC^{+}08b] L. Meynier, P. C. H. Morel, B. L. Chilvers, D. D. S. Mackenzie, A. MacGibbon, and P. J. Duignan. Temporal and sex differences in the blubber fatty acid profiles of the New Zealand sea lion Phocarctos hookeri. Marine Ecology Progress Series, 366:271–279, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p271.pdf.

Macleod:2007:BRO

[MMCF07] C. K. Macleod, N. A. Moltschaniwskyj, C. M. Crawford, and S. E. Forbes. Biological recovery from organic enrichment: some systems cope better than others. Marine Ecology Progress Series, 342:41–53, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p041.pdf.

Minier:2006:MRP

[MMGC06] C. Minier, M. N. Moore, F. Galgani, and D. Claisse. Multixenobiotic resistance protein expression in Mytilus edulis, M. galloprovincialis and Crassostrea gigas from the French coasts. Marine Ecology Progress Series, 322:143–154, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p143.pdf.

Morato:2008:ADS

[MMK^{+}08] T. Morato, M. Machete, A. Kitchingman, F. Tempera, S. Lai, G. Menezes, T. J. Pitcher, and R. S. Santos. Abundance and distribution of seamounts in the Azores. Marine Ecology Progress Series, 357:17–21, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/357/m357p017.pdf.

Molina-Montenegro:2005:PAB

[MMMB^{+}05] M. A. Molina-Montenegro, A. A. Muñoz, E. I. Badano, B. W. Morales, K. M. Fuentes, and L. A. Cavieres. Positive associations between macroalgal species in a rocky intertidal zone and their effects on the physiological performance of Ulva lactuca. Marine Ecology Progress Series, 292:173–180, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p173.pdf.

McClanahan:2005:EGT

[MMMPB05] T. R. McClanahan, J. Maina, R. Moothien-Pillay, and A. C. Baker. Effects of geography, taxa, water flow, and temperature variation on coral bleaching intensity in Mauritius. Marine Ecology Progress Series, 298:131–142, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/

m298p131.pdf. See comment [vWNYS05].

McKee:2009:PMI

[MMPD09] K. L. McKee, B. A. Middleton, C. E. Proffitt, and D. J. Devlin. Perspectives on mosquito impoundments in eastern Florida, USA: Reply to Rey et al. (2009). Marine Ecology Progress Series, 389:301–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m389p301.pdf.

Mullineaux:2005:VLT

[MMS^{+}05] L. S. Mullineaux, S. W. Mills, A. K. Sweetman, A. H. Beaudreau, A. Metaxas, and H. L. Hunt. Vertical, lateral and temporal structure in larval distributions at hydrothermal vents. Marine Ecology Progress Series, 293:1–16, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/293/
m293p001.pdf.

Metaxas:2009:DEL

[MMS09] A. Metaxas, L. S. Mullineaux, and J. Sisson. Distribution of echinoderm larvae relative to the halocline of a salt wedge. Marine Ecology Progress Series, 377:157–168, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/377/
m377p157.pdf.

Magni:2000:TSR

[MMTT00] P. Magni, S. Montani, C. Takada, and H. Tsutsumi. Temporal scaling and relevance of bivalve nutrient excretion on a tidal flat of the Seto Inland Sea, Japan. Marine Ecology Progress Series, 198:139–155, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/198/m198p139.pdf.

Macho:2005:TLR

[MMV05] G. Macho, J. Molares, and E. Vázquez. Timing of larval release by three barnacles from the NW Iberian Peninsula. Marine Ecology Progress Series, 298:251–260, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p251.pdf.

Mariani:2007:IBS

[MMVB07] P. Mariani, B. R. MacKenzie, A. W. Visser, and V. Botte. Individual-based simulations of larval fish feeding in turbulent environments. Marine Ecology Progress Series, 347:155– 169, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m347p155.pdf.

Middelburg:2000:NUH

[MN00a] J. J. Middelburg and J. Nieuwenhuize. Nitrogen uptake by heterotrophic bacteria and phytoplankton in the nitraterich Thames estuary. Marine Ecology Progress Series, 203: 13–21, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/203/m203p013.pdf.

Middelburg:2000:UDI

[MN00b] J. J. Middelburg and J. Nieuwenhuize. Uptake of dissolved inorganic nitrogen in turbid, tidal estuaries. Marine Ecology Progress Series, 192:79–88, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/192/m192p079.pdf.

Miyamoto:2004:EMC

[MN04] Y. Miyamoto and T. Noda. Effects of mussels on competitively inferior species: competitive exclusion to facilitation. Marine Ecology Progress Series, 276:293–298, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p293.pdf.

Maar:2007:GEB

[MNB^{+}07] M. Maar, T. G. Nielsen, K. Bolding, H. Burchard, and A. W. Visser. Grazing effects of blue mussel Mytilus edulis on the pelagic food web under different turbulence conditions. Marine Ecology Progress Series, 339:199–213, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m339p199.pdf.

Morais:2002:LDD

[MNC^{+}02] S. Morais, L. Narciso, R. Calado, M. L. Nunes, and R. Rosa. Lipid dynamics during the embryonic development of Plesionika martia martia (Decapoda; Pandalidae), Palaemon serratus and P. elegans (Decapoda; Palaemonidae): relation to metabolic consumption. Marine Ecology Progress Series, 242:195–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p195.pdf.

Meesters:2002:RDR

[MND^{+}02] E. H. Meesters, G. Nieuwland, G. C. A. Duineveld, A. Kok, and R. P. M. Bak. RNA/DNA ratios of scleractinian corals suggest acclimatisation/adaptation in relation to light gradients and turbidity regimes. Marine Ecology Progress Series, 227:233–239, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p233.pdf.

Mullers:2009:ECF

[MND^{+}09] R. H. E. Mullers, R. A. Navarro, S. Daan, J. M. Tinbergen, and H. A. J. Meijer. Energetic costs of foraging in breeding Cape gannets Morus capensis. Marine Ecology Progress Series, 393:161–171, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/393/m393p161.pdf.

Miller:2002:IDD

[MNP02] D. C. Miller, A. Norkko, and C. A. Pilditch. Influence of diet on dispersal of horse mussel Atrina zelandica biodeposits. Marine Ecology Progress Series, 242:153–167, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/242/
m242p153.pdf.

Maar:2002:STV

[MNR^{+}02] M. Maar, T. G. Nielsen, K. Richardson, U. Christaki, O. S. Hansen, S. Zervoudaki, and E. D. Christou. Spatial and temporal variability of food web structure during the spring bloom in the Skagerrak. Marine Ecology Progress Series, 239: 11–29, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/239/m239p011.pdf.

Miller:2009:RAD

[MNT09] K. Miller, H. Neil, and D. Tracey. Recent advances in deep-sea coral science and emerging links to conservation and management of deep-sea ecosystems. Marine Ecology Progress Series, 397:1–5, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p001.pdf.

Madsen:2008:IFE

[MNTS08] S. J. Madsen, T. G. Nielsen, O. M. Tervo, and J. Söderkvist. Importance of feeding for egg production in Calanus finmarchicus and C. glacialis during the Arctic spring. Marine Ecology Progress Series, 353:177–190, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m353p177.pdf.

Meyer:2005:EST

[MO05] B. Meyer and B. Oettl. Effects of short-term starvation on composition and metabolism of larval Antarctic krill Euphausia superba. Marine Ecology Progress Series, 292:263– 270, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p263.pdf.

Moksnes:2004:ICS

[Mok04] P. O. Moksnes. Interference competition for space in nursery habitats: density-dependent effects on growth and dispersal in juvenile shore crabs Carcinus maenas. Marine Ecology Progress Series, 281:181–191, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/281/m281p181.pdf.

Miller:2002:DLK

[MOM^{+}02] M. J. Miller, T. Otake, G. Minagawa, T. Inagaki, and K. Tsukamoto. Distribution of leptocephali in the Kuroshio Current and East China Sea. Marine Ecology Progress Series, 235:279–288, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p279.pdf.

Montevecchi:2007:BDR

[Mon07] W. A. Montevecchi. Binary dietary responses of northern gannets Sula bassana indicate changing food web and oceanographic conditions. Marine Ecology Progress Series, 352: 213–220, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m352p213.pdf.

Melrose:2006:CFR

[MOOB06] D. C. Melrose, C. A. Oviatt, J. E. O’Reilly, and M. S. Berman. Comparisons of fast repetition rate fluorescence estimated primary production and ^{14}C uptake by phytoplankton. Marine Ecology Progress Series, 311:37–46, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/311/
m311p037.pdf.

Morote:2008:TEB

[MOP^{+}08] E. Morote, M. P. Olivar, P. M. Pankhurst, F. Villate, and I. Uriarte. Trophic ecology of bullet tuna Auxis rochei larvae and ontogeny of feeding-related organs. Marine Ecology Progress Series, 353:243–254, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m353p243.pdf.

Michaelidis:2005:ELT

[MOPP05] B. Michaelidis, C. Ouzounis, A. Paleras, and H. O. Pörtner. Effects of long-term moderate hypercapnia on acid-base balance and growth rate in marine mussels Mytilus galloprovincialis. Marine Ecology Progress Series, 293:109–118, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/293/m293p109.pdf.

Mouritsen:2004:IFI

[Mou04] K. N. Mouritsen. Intertidal facilitation and indirect effects: causes and consequences of crawling in the New Zealand cockle. Marine Ecology Progress Series, 271:207– 220, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/271/m271p207.pdf.

McQuaid:2000:LWD

[MP00] C. D. McQuaid and T. E. Phillips. Limited wind-driven dispersal of intertidal mussel larvae: in situ evidence from the plankton and the spread of the invasive species Mytilus galloprovincialis in South Africa. Marine Ecology Progress Series, 201:211–220, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p211.pdf.

Mengelt:2005:UEC

[MP05] C. Mengelt and B. B. Prézelin. UVA enhancement of carbon fixation and resilience to UV inhibition in the genus Pseudo-nitzschia may provide a competitive advantage in high UV surface waters. Marine Ecology Progress Series, 301: 81–93, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/301/m301p081.pdf.

McQuaid:2006:MVR

[MP06] C. D. McQuaid and T. E. Phillips. Mesoscale variation in reproduction, recruitment and population structure of intertidal mussels with low larval input: a bay/open coast comparison. Marine Ecology Progress Series, 327:193–206, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p193.pdf.

Morris:2009:CPS

[MPB^{+}09] E. P. Morris, G. Peralta, J. Benavente, R. Freitas, A. M. Rodrigues, V. Quintino, O. Alvarez, N. Valcárcel-Pérez, J. J. Vergara, I. Hernández, and J. L. Pérez-Lloréns. Caulerpa prolifera stable isotope ratios reveal anthropogenic nutrients within a tidal lagoon. Marine Ecology Progress Series, 390: 117–128, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p117.pdf.

McCay:2003:RTF

[MPDC03] D. P. French McCay, C. H. Peterson, J. T. DeAlteris, and J. Catena. Restoration that targets function as opposed to structure: replacing lost bivalve production and filtration. Marine Ecology Progress Series, 264:197–212, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p197.pdf.

Moisander:2007:PLD

[MPDS07] P. H. Moisander, H. W. Paerl, J. Dyble, and K. Sivonen. Phosphorus limitation and diel control of nitrogen-fixing cyanobacteria in the Baltic Sea. Marine Ecology Progress Series, 345: 41–50, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p041.pdf.

Morgan:2003:OOV

[MPE03] C. A. Morgan, W. T. Peterson, and R. L. Emmett. Onshoreoffshore variations in copepod community structure off the Oregon coast during the summer upwelling season. Marine Ecology Progress Series, 249:223–236, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p223.pdf.

Marshall:2003:LAL

[MPK03] D. J. Marshall, J. A. Pechenik, and M. J. Keough. Larval activity levels and delayed metamorphosis affect post-larval performance in the colonial ascidian Diplosoma listerianum. Marine Ecology Progress Series, 246:153–162, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p153.pdf.

Manderson:2002:SDH

[MPM^{+}02] J. P. Manderson, B. A. Phelan, C. Meise, L. L. Stehlik, A. J. Bejda, J. Pessutti, L. Arlen, A. Draxler, and A. W. Stoner. Spatial dynamics of habitat suitability for the growth of newly settled winter flounder Pseudopleuronectes americanus in an estuarine nursery. Marine Ecology Progress Series, 228:227–239, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p227.pdf.

Manderson:2003:WFS

[MPM^{+}03] J. P. Manderson, J. Pessutti, C. Meise, D. Johnson, and P. Shaheen. Winter flounder settlement dynamics and the modification of settlement patterns by post-settlement processes in a NW Atlantic estuary. Marine Ecology Progress Series, 253:253–267, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p253.pdf.

Mazzoldi:2002:ESV

[MPR02] C. Mazzoldi, C. Poltronieri, and M. B. Rasotto. Egg size variability and mating system in the marbled goby Pomatoschistus marmoratus (Pisces: Gobiidae). Marine Ecology Progress Series, 233:231–239, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p231.pdf.

Maes:2006:MCB

[MPR^{+}06] G. E. Maes, J. M. Pujolar, J. A. M. Raeymaekers, J. Dannewitz, and F. A. M. Volckaert. Microsatellite conservation and Bayesian individual assignment in four Anguilla species.

Marine Ecology Progress Series, 319:251–261, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/319/
m319p251.pdf.

Meynier:2008:IDV

[MPS^{+}08] L. Meynier, C. Pusineri, J. Spitz, M. B. Santos, G. J. Pierce, and V. Ridoux. Intraspecific dietary variation in the shortbeaked common dolphin Delphinus delphis in the Bay of Biscay: importance of fat fish. Marine Ecology Progress Series, 354:277–287, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p277.pdf.

Manderson:2006:DEJ

[MPSJ06] J. P. Manderson, J. Pessutti, P. Shaheen, and F. Juanes. Dynamics of early juvenile winter flounder predation risk on a North West Atlantic estuarine nursery ground. Marine Ecology Progress Series, 328:249–265, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p249.pdf.

Murphy:2001:MSO

[MR01] E. J. Murphy and K. Reid. Modelling Southern Ocean krill population dynamics: biological processes generating fluctuations in the South Georgia ecosystem. Marine Ecology Progress Series, 217:175–189, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/217/m217p175.pdf.

McCay:2003:HRM

[MR03] D. P. French McCay and J. J. Rowe. Habitat restoration as mitigation for lost production at multiple trophic levels. Marine Ecology Progress Series, 264:233–247, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p233.pdf.

Mari:2004:RTMa

[MR04] X. Mari and F. Rassoulzadegan. Role of TEP in the microbial food web structure. I. Grazing behavior of a bacterivorous pelagic ciliate. Marine Ecology Progress Series, 279:

13–22, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p013.pdf.

Macia:2005:EHH

[MR05] S. Maciá and M. P. Robinson. Effects of habitat heterogeneity in seagrass beds on grazing patterns of parrotfishes. Marine Ecology Progress Series, 303:113–121, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p113.pdf.

Macpherson:2006:RBP

[MR06a] E. Macpherson and N. Raventos. Relationship between pelagic larval duration and geographic distribution of Mediterranean littoral fishes. Marine Ecology Progress Series, 327:257– 265, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p257.pdf.

Mizroch:2006:NPK

[MR06b] S. A. Mizroch and D. W. Rice. Have North Pacific killer whales switched prey species in response to depletion of the great whale populations? Marine Ecology Progress Series, 310:235–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/310/m310p235.pdf.

Moller:2006:FFB

[MR06c] L. F. Møller and H. U. Riisgård. Filter feeding in the burrowing amphipod Corophium volutator. Marine Ecology Progress Series, 322:213–224, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/322/m322p213.pdf.

Moller:2007:FBG

[MR07a] L. F. Møller and H. U. Riisgård. Feeding, bioenergetics and growth in the common jellyfish Aurelia aurita and two hydromedusae, Sarsia tubulosa and Aequorea vitrina. Marine Ecology Progress Series, 346:167–177, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p167.pdf.

Moller:2007:PDG

[MR07b] L. F. Møller and H. U. Riisgård. Population dynamics, growth and predation impact of the common jellyfish Aurelia aurita and two hydromedusae, Sarsia tubulosa and Aequorea vitrina, in Limfjorden (Denmark). Marine Ecology Progress Series, 346:153–165, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/346/m346p153.pdf.

Moller:2007:RSJ

[MR07c] L. F. Møller and H. U. Riisgård. Respiration in the scyphozoan jellyfish Aurelia aurita and two hydromedusae Sarsia tubulosa and Aequorea vitrina: effect of size, temperature and growth. Marine Ecology Progress Series, 330:149–154, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/330/
m330p149.pdf.

Mari:2004:RTMb

[MRB04] X. Mari, F. Rassoulzadegan, and C. P. D. Brussaard. Role of TEP in the microbial food web structure. II. Influence on the ciliate community structure. Marine Ecology Progress Series, 279:23–32, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/279/m279p023.pdf.

Morata:2008:SSV

[MRB^{+}08] N. Morata, P. E. Renaud, S. Brugel, K. A. Hobson, and B. J. Johnson. Spatial and seasonal variations in the pelagicbenthic coupling of the southeastern Beaufort Sea revealed by sedimentary biomarkers. Marine Ecology Progress Series, 371:47–63, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/371/m371p047.pdf.

Miller:2009:CSP

[MRB09] R. J. Miller, D. C. Reed, and M. A. Brzezinski. Community structure and productivity of subtidal turf and foliose algal assemblages. Marine Ecology Progress Series, 388: 1–11, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/feature/m388p001.pdf.

Mercado:2008:PEC

[MRC^{+}08] J. M. Mercado, T. Ramírez, D. Cortés, M. Sebastián, E. Liger, and B. Bautista. Partitioning the effects of changes in nitrate availability and phytoplankton community structure on relative nitrate uptake. Marine Ecology Progress Series, 359: 51–68, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/359/m359p051.pdf.

Mikkelsen:2008:MCP

[MRG08] D. M. Mikkelsen, S. Rysgaard, and R. N. Glud. Microalgal composition and primary production in Arctic sea ice: a seasonal study from Kobbefjord (Kangerluarsunnguaq), West Greenland. Marine Ecology Progress Series, 368:65– 74, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/368/m368p065.pdf.

Molloy:2009:EAF

[MRGC09] P. P. Molloy, J. D. Reynolds, M. J. G. Gage, and I. M. Côté. Effects of an artisanal fishery on non-spawning grouper populations. Marine Ecology Progress Series, 392:253–262, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/392/m392p253.pdf.

Monticelli:2008:JST

[MRH^{+}08] D. Monticelli, J. A. Ramos, J. E. Hines, J. D. Nichols, and J. A. Spendelow. Juvenile survival in a tropical population of roseate terns: interannual variation and effect of tick parasitism. Marine Ecology Progress Series, 365:277– 287, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/365/m365p277.pdf.

Mateo:2001:MSP

[MRHP01] M. A. Mateo, P. Renom, M. A. Hemminga, and J. Peene. Measurement of seagrass production using the ^{13}C stable isotope compared with classical O_{2} and ^{14}C methods. Marine Ecology Progress Series, 223:157–165, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/223/m223p157.pdf.

McClain:2005:DBB

[MRJ05] C. R. McClain, M. A. Rex, and R. Jabbour. Deconstructing bathymetric body size patterns in deep-sea gastropods. Marine Ecology Progress Series, 297:181–187, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p181.pdf.

Macreadie:2006:DMS

[MRLK06] P. I. Macreadie, D. J. Ross, A. R. Longmore, and M. J. Keough. Denitrification measurements of sediments using cores and chambers. Marine Ecology Progress Series, 326: 49–59, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/326/m326p049.pdf.

Miller:2007:DSF

[MRMA07] W. A. Miller, G. M. Ruiz, M. S. Minton, and R. F. Ambrose. Differentiating successful and failed molluscan invaders in estuarine ecosystems. Marine Ecology Progress Series, 332:41–51, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/332/m332p041.pdf.

Macia:2007:EDR

[MRN07] S. Maciá, M. P. Robinson, and A. Nalevanko. Experimental dispersal of recovering Diadema antillarum increases grazing intensity and reduces macroalgal abundance on a coral reef. Marine Ecology Progress Series, 348:173–182, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p173.pdf.

McCall:2009:EGS

[MRP09] D. Drury McCall, C. F. Rakocinski, and J. L. Pinckney. Effects of grass shrimp versus nutrient addition on epiphytic algae associated with the ephemeral widgeongrass Ruppia maritima. Marine Ecology Progress Series, 379:151–162, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p151.pdf.

Masotti:2007:PCT

[MRPL07] I. Masotti, D. Ruiz-Pino, and A. Le Bouteiller. Photosynthetic characteristics of Trichodesmium in the southwest Pacific Ocean: importance and significance. Marine Ecology Progress Series, 338:47–59, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/338/m338p047.pdf.

Monticelli:2007:EAC

[MRQ07] D. Monticelli, J. A. Ramos, and G. D. Quartly. Effects of annual changes in primary productivity and ocean indices on breeding performance of tropical roseate terns in the western Indian Ocean. Marine Ecology Progress Series, 351:273– 286, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/351/m351p273.pdf.

Martin:2006:MAP

[MRQM06] D. L. Martin, R. M. Ross, L. B. Quetin, and A. E. Murray. Molecular approach (PCR-DGGE) to diet analysis in young Antarctic krill Euphausia superba. Marine Ecology Progress Series, 319:155–165, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/319/m319p155.pdf.

Maravelias:2000:SSW

[MRS00] C. D. Maravelias, D. G. Reid, and G. Swartzman. Seabed substrate, water depth and zooplankton as determinants of the prespawning spatial aggregation of North Atlantic herring. Marine Ecology Progress Series, 195:249–259, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p249.pdf.

Maps:2005:EPH

[MRZJ05] F. Maps, J. A. Runge, B. Zakardjian, and P. Joly. Egg production and hatching success of Temora longicornis (Copepoda, Calanoida) in the southern Gulf of St. Lawrence. Marine Ecology Progress Series, 285:117–128, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p117.pdf.

Meissner:2000:ESC

[MS00a] K. Meißner and T. Schaarschmidt. Ecophysiological studies of Corophium volutator (Amphipoda) infested by microphallid trematodes. Marine Ecology Progress Series, 202:143– 151, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/202/m202p143.pdf.

Miles:2000:DVM

[MS00b] A. Miles and K. Sundbäck. Diel variation in microphytobenthic productivity in areas of different tidal amplitude. Marine Ecology Progress Series, 205:11–22, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/205/m205p011.pdf.

Meidel:2001:VES

[MS01a] S. K. Meidel and R. E. Scheibling. Variation in egg spawning among subpopulations of sea urchins Strongylocentrotus droebachiensis: a theoretical approach. Marine Ecology Progress Series, 213:97–110, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p097.pdf.

Moncreiff:2001:TIE

[MS01b] C. A. Moncreiff and M. J. Sullivan. Trophic importance of epiphytic algae in subtropical seagrass beds: evidence from multiple stable isotope analyses. Marine Ecology Progress Series, 215:93–106, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p093.pdf.

Murphy:2003:UAF

[MS03] N. E. Murphy and B. Schaffelke. Use of amplified fragment length polymorphism (AFLP) as a new tool to explore the invasive green alga Caulerpa taxifolia in Australia. Marine Ecology Progress Series, 246:307–310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p307.pdf.

Miller:2004:RRL

[MS04a] A. K. Miller and W. J. Sydeman. Rockfish response to lowfrequency ocean climate change as revealed by the diet of a marine bird over multiple time scales. Marine Ecology Progress Series, 281:207–216, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/281/m281p207.pdf.

Miller:2004:OEC

[MS04b] J. A. Miller and A. L. Shanks. Ocean-estuary coupling in the Oregon upwelling region: abundance and transport of juvenile fish and of crab megalopae. Marine Ecology Progress Series, 271:267–279, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/271/m271p267.pdf.

Miller:2005:ADL

[MS05] J. A. Miller and A. L. Shanks. Abundance and distribution of larval and juvenile fish in Coos Bay, Oregon: time-series analysis based on light-trap collections. Marine Ecology Progress Series, 305:177–191, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/305/m305p177.pdf.

McDermott:2006:OSM

[MS06] C. J. McDermott and J. S. Shima. Ontogenetic shifts in microhabitat preference of the temperate reef fish Forsterygion lapillum: implications for population limitation. Marine Ecology Progress Series, 320:259–266, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p259.pdf.

McKeown:2008:SPW

[MS08] N. J. McKeown and P. W. Shaw. Single paternity within broods of the brown crab Cancer pagurus: a highly fecund species with long-term sperm storage. Marine Ecology Progress Series, 368:209–215, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/368/m368p209.pdf.

Malecha:2009:RSW

[MS09] P. W. Malecha and R. P. Stone. Response of the sea whip Halipteris willemoesi to simulated trawl disturbance and its vulnerability to subsequent predation. Marine Ecology Progress Series, 388:197–206, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p197.pdf.

McGlathery:2007:ESC

[MSA07] K. J. McGlathery, K. Sundbäck, and I. C. Anderson. Eutrophication in shallow coastal bays and lagoons: the role of plants in the coastal filter. Marine Ecology Progress Series, 348:1–18, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Marshall:2004:CSL

[MSC04] D. J. Marshall, D. Semmens, and C. Cook. Consequences of spawning at low tide: limited gamete dispersal for a rockpool anemone. Marine Ecology Progress Series, 266:135– 142, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/266/m266p135.pdf.

Mincks:2005:PLO

[MSD05] S. L. Mincks, C. R. Smith, and D. J. DeMaster. Persistence of labile organic matter and microbial biomass in Antarctic shelf sediments: evidence of a sediment ‘food bank’. Marine Ecology Progress Series, 300:3–19, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/300/m300p003.pdf.

Mutchler:2004:PIE

[MSF04] T. Mutchler, M. J. Sullivan, and B. Fry. Potential of ^{14}N isotope enrichment to resolve ambiguities in coastal trophic relationships. Marine Ecology Progress Series, 266:27–33, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/266/m266p027.pdf.

MacNeil:2005:SIM

[MSF05] M. A. MacNeil, G. B. Skomal, and A. T. Fisk. Stable isotopes from multiple tissues reveal diet switching in sharks. Marine Ecology Progress Series, 302:199–206, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p199.pdf.

Malet:2007:ESV

[MSF^{+}07] N. Malet, P. G. Sauriau, N. Faury, P. Soletchnik, and G. Guillou. Effect of seasonal variation in trophic conditions and the gametogenic cycle on δ ^{13}C and δ ^{15}N levels of diploid and triploid Pacific oysters Crassostrea gigas. Marine Ecology Progress Series, 346:203–217, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/346/m346p203.pdf.

Meager:2009:VDR

[MSF^{+}09] J. J. Meager, J. E. Skjæraasen, A. Fernö, Ó. Karlsen, S. Løkkeborg, K. Michalsen, and S. O. Utskot. Vertical dynamics and reproductive behaviour of farmed and wild Atlantic cod Gadus morhua. Marine Ecology Progress Series, 389:233–243, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p233.pdf.

Mangubhai:2007:PVC

[MSG07] S. Mangubhai, P. Souter, and M. Grahn. Phenotypic variation in the coral Platygyra daedalea in Kenya: morphometry and genetics. Marine Ecology Progress Series, 345:105– 115, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p105.pdf.

Moisander:2003:VNP

[MSH^{+}03a] P. H. Moisander, T. F. Steppe, N. S. Hall, J. Kuparinen, and H. W. Paerl. Variability in nitrogen and phosphorus limitation for Baltic Sea phytoplankton during nitrogen-fixing cyanobacterial blooms. Marine Ecology Progress Series, 262: 81–95, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p081.pdf.

Moore:2003:PCP

[MSH^{+}03b] C. M. Moore, D. Suggett, P. M. Holligan, J. Sharples, E. R. Abraham, M. I. Lucas, T. P. Rippeth, N. R. Fisher, J. H. Simpson, and D. J. Hydes. Physical controls on phytoplankton physiology and production at a shelf sea front: a fast repetition-rate fluorometer based field study. Marine Ecology Progress Series, 259:29–45, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/259/m259p029.pdf.

Magalhaes:2008:DFC

[MSH08] M. C. Magalhães, R. S. Santos, and K. C. Hamer. Dualforaging of Cory’s shearwaters in the Azores: feeding locations, behaviour at sea and implications for food provisioning of chicks. Marine Ecology Progress Series, 359:283– 293, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/359/m359p283.pdf.

Moe:2009:CCP

[MSJ^{+}09] B. Moe, L. Stempniewicz, D. Jakubas, F. Angelier, O. Chastel, F. Dinessen, G. W. Gabrielsen, F. Hanssen, N. J. Karnovsky, B. Rønning, J. Welcker, K. Wojczulanis-Jakubas, and C. Bech. Climate change and phenological responses of two seabird species breeding in the high-Arctic. Marine Ecology Progress Series, 393:235–246, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m393p235.pdf.

Marshall:2000:ICV

[MSK00a] D. J. Marshall, C. A. Styan, and M. J. Keough. Intraspecific co-variation between egg and body size affects fertilisation kinetics of free-spawning marine invertebrates. Marine Ecology Progress Series, 195:305–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p305.pdf.

Miyazaki:2000:DCL

[MSK^{+}00b] T. Miyazaki, S. Shiozawa, T. Kogane, R. Masuda, K. Maruyama, and K. Tsukamoto. Developmental changes of the light intensity threshold for school formation in the striped jack Pseudocaranx dentex. Marine Ecology Progress Series, 192:267–275, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/192/m192p267.pdf.

MacLeod:2006:RPS

[MSLP06] C. D. MacLeod, M. B. Santos, A. López, and G. J. Pierce. Relative prey size consumption in toothed whales: implications for prey selection and level of specialisation. Marine Ecology Progress Series, 326:295–307, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/326/m326p295.pdf.

Morgan:2006:STM

[MSP^{+}06] S. G. Morgan, S. A. Spilseth, H. M. Page, A. J. Brooks, and E. D. Grosholz. Spatial and temporal movement of the lined shore crab Pachygrapsus crassipes in salt marshes and its utility as an indicator of habitat condition. Marine Ecology Progress Series, 314:271–281, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/314/m314p271.pdf.

McClanahan:2003:IBN

[MSS^{+}03] T. R. McClanahan, E. Sala, P. A. Stickels, B. A. Cokos, A. C. Baker, C. J. Starger, and S. H. Jones IV. Interaction between nutrients and herbivory in controlling algal communities and coral condition on Glover’s Reef, Belize. Marine Ecology Progress Series, 261:135–147, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p135.pdf.

Menezes:2006:SZD

[MSSP06] G. M. Menezes, M. F. Sigler, H. M. Silva, and M. R. Pinho. Structure and zonation of demersal fish assemblages off the Azores Archipelago (mid-Atlantic). Marine Ecology Progress Series, 324:241–260, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/324/m324p241.pdf.

McPherson:2003:GDA

[MST03] A. A. McPherson, R. L. Stephenson, and C. T. Taggart. Genetically different Atlantic herring Clupea harengus spawning waves. Marine Ecology Progress Series, 247:303–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/247/m247p303.pdf.

Muniz-Salazar:2006:GSE

[MSTS^{+}06] R. Muñiz-Salazar, S. L. Talbot, G. K. Sage, D. H. Ward, and A. Cabello-Pasini. Genetic structure of eelgrass Zostera marina meadows in an embayment with restricted water flow. Marine Ecology Progress Series, 309:107–116, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p107.pdf.

Miyajima:2001:MNT

[MSUK01] T. Miyajima, M. Suzumura, Y. Umezawa, and I. Koike. Microbiological nitrogen transformation in carbonate sediments of a coral-reef lagoon and associated seagrass beds. Marine Ecology Progress Series, 217:273–286, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/217/m217p273.pdf.

Metaxas:2002:EFS

[MSY02] A. Metaxas, R. E. Scheibling, and C. M. Young. Estimating fertilization success in marine benthic invertebrates: a case study with the tropical sea star Oreaster reticulatus. Marine Ecology Progress Series, 226:87–101, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p087.pdf.

Meziane:2000:FAT

[MT00] T. Meziane and M. Tsuchiya. Fatty acids as tracers of organic matter in the sediment and food web of a mangrove/intertidal flat ecosystem, Okinawa, Japan. Marine Ecology Progress Series, 200:49–57, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p049.pdf.

Martin:2005:TSE

[MT05] G. Bath Martin and S. R. Thorrold. Temperature and salinity effects on magnesium, manganese, and barium incorporation in otoliths of larval and early juvenile spot Leiostomus xanthurus. Marine Ecology Progress Series, 293:223– 232, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p223.pdf.

McKone:2009:RSS

[MT09] K. L. McKone and C. E. Tanner. Role of salinity in the susceptibility of eelgrass Zostera marina to the wasting disease pathogen Labyrinthula zosterae. Marine Ecology Progress Series, 377:123–130, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/377/m377p123.pdf.

Murphy:2008:SVC

[MTCU08] R. J. Murphy, T. J. Tolhurst, M. G. Chapman, and A. J. Underwood. Spatial variation of chlorophyll on estuarine mudflats determined by field-based remote sensing. Marine Ecology Progress Series, 365:45–55, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p045.pdf.

MacNeil:2008:ADR

[MTF^{+}08] M. A. MacNeil, E. H. M. Tyler, C. J. Fonnesbeck, S. P. Rushton, N. V. C. Polunin, and M. J. Conroy. Accounting for detectability in reef-fish biodiversity estimates. Marine Ecology Progress Series, 367:249–260, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p249.pdf.

Marchetti:2004:ECP

[MTH04] A. Marchetti, V. L. Trainer, and P. J. Harrison. Environmental conditions and phytoplankton dynamics associated with Pseudo-nitzschia abundance and domoic acid in the Juan de Fuca eddy. Marine Ecology Progress Series, 281: 1–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/281/m281p001.pdf.

Martins:2008:RIC

[MTH^{+}08] G. M. Martins, R. C. Thompson, S. J. Hawkins, A. I. Neto, and S. R. Jenkins. Rocky intertidal community structure in oceanic islands: scales of spatial variability. Marine Ecology Progress Series, 356:15–24, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/356/m356p015.pdf.

Marchand:2003:REF

[MTL^{+}03] J. Marchand, A. Tanguy, J. Laroche, L. Quiniou, and D. Moraga. Responses of European flounder Platichthys flesus populations to contamination in different estuaries along the Atlantic coast of France. Marine Ecology Progress Series, 260: 273–284, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p273.pdf.

Musco:2009:TSE

[MTLG09] L. Musco, A. Terlizzi, M. Licciano, and A. Giangrande. Taxonomic structure and the effectiveness of surrogates in environmental monitoring: a lesson from polychaetes. Marine Ecology Progress Series, 383:199–210, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/383/m383p199.pdf.

Mendes:2002:ITC

[MTLT02] S. Mendes, W. Turrell, T. Lütkebohle, and P. Thompson. Influence of the tidal cycle and a tidal intrusion front on the spatio-temporal distribution of coastal bottlenose dolphins. Marine Ecology Progress Series, 239:221–229, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p221.pdf.

Moller:2003:PDC

[MTN03] E. F. Møller, P. Thor, and T. G. Nielsen. Production of DOC by Calanus finmarchicus, C. glacialis and C. hyperboreus through sloppy feeding and leakage from fecal pellets. Marine Ecology Progress Series, 262:185–191, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/262/
m262p185.pdf.

Martin:2007:BCR

[MTR^{+}07] S. Martin, G. Thouzeau, M. Richard, L. Chauvaud, F. Jean, and J. Clavier. Benthic community respiration in areas impacted by the invasive mollusk Crepidula fornicata. Marine Ecology Progress Series, 347:51–60, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/347/
m347p051.pdf.

Miyaji:2007:ECD

[MTS07] T. Miyaji, K. Tanabe, and B. R. Schöne. Environmental controls on daily shell growth of Phacosoma japonicum (Bivalvia: Veneridae) from Japan. Marine Ecology Progress Series, 336:141–150, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p141.pdf.

Martinetto:2006:CEB

[MTV06] P. Martinetto, M. Teichberg, and I. Valiela. Coupling of estuarine benthic and pelagic food webs to land-derived nitrogen sources in Waquoit Bay, Massachusetts, USA. Marine Ecology Progress Series, 307:37–48, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/307/m307p037.pdf.

Munk:2007:CFV

[Mun07] P. Munk. Cross-frontal variation in growth rate and prey availability of larval North Sea cod Gadus morhua. Marine Ecology Progress Series, 334:225–235, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/334/m334p225.pdf.

Murphy:2006:QIM

[MUP06] R. J. Murphy, A. J. Underwood, and M. H. Pinkerton. Quantitative imaging to measure photosynthetic biomass on an intertidal rock-platform. Marine Ecology Progress Series, 312: 45–55, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/312/m312p045.pdf.

Mariani:2003:MBE

[MUT03] S. Mariani, M. J. Uriz, and X. Turon. Methodological bias in the estimations of important meroplanktonic components from near-shore bottoms. Marine Ecology Progress Series, 253:67–75, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/253/m253p067.pdf.

Marti:2004:SSV

[MUT04] R. Martí, M. J. Uriz, and X. Turon. Seasonal and spatial variation of species toxicity in Mediterranean seaweed communities: correlation to biotic and abiotic factors. Marine Ecology Progress Series, 282:73–85, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/282/m282p073.pdf.

Moutin:2005:PAC

[MVB^{+}05] T. Moutin, N. Van Den Broeck, B. Beker, C. Dupouy, P. Rimmelin, and A. Le Bouteiller. Phosphate availability controls Trichodesmium spp. biomass in the SW Pacific Ocean. Marine Ecology Progress Series, 297:15–21, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p015.pdf.

Martel:2004:IMG

[MVBGM04] C. Martel, F. Viard, D. Bourguet, and P. Garcia-Meunier. Invasion by the marine gastropod Ocinebrellus inornatus in France. II. Expansion along the Atlantic coast. Marine Ecology Progress Series, 273:163–172, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/273/
m273p163.pdf.

Montserrat:2008:BCM

[MVD^{+}08a] F. Montserrat, C. Van Colen, S. Degraer, T. Ysebaert, and P. M. J. Herman. Benthic community-mediated sediment dynamics. Marine Ecology Progress Series, 372:43–59, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m372p043.pdf.

Morato:2008:ESE

[MVD^{+}08b] T. Morato, D. A. Varkey, C. Damaso, M. Machete, M. Santos, R. Prieto, R. S. Santos, and T. J. Pitcher. Evidence of a seamount effect on aggregating visitors. Marine Ecology Progress Series, 357:23–32, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/357/m357p023.pdf.

Modig:2000:UPT

[MvdBÓ00] H. Modig, W. J. van de Bund, and E. Ólafsson. Uptake of phytodetritus by three ostracod species from the Baltic Sea: effects of amphipod disturbance and ostracod density. Marine Ecology Progress Series, 202:125–134, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p125.pdf.

Minte-Vera:2008:NSI

[MVdMFF08] C. V. Minte-Vera, R. L. de Moura, and R. B. Francini-Filho. Nested sampling: an improved visual-census technique for studying reef fish assemblages. Marine Ecology Progress Series, 367:283–293, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/367/m367p283.pdf.

Meekan:2006:BBS

[MVH^{+}06] M. G. Meekan, L. Vigliola, A. Hansen, P. J. Doherty, A. Halford, and J. H. Carleton. Bigger is better: size-selective mortality throughout the life history of a fast-growing clupeid, Spratelloides gracilis. Marine Ecology Progress Series, 317:237–244, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p237.pdf.

Maes:2006:SDE

[MVHV06] J. Maes, A. Van de Putte, J. H. Hecq, and F. A. M. Volckaert. State-dependent energy allocation in the pelagic Antarctic silverfish Pleuragramma antarcticum: trade-off between winter reserves and buoyancy. Marine Ecology Progress Series, 326:269–282, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p269.pdf.

Miller:2009:ABA

[MVK^{+}09] M. W. Miller, A. Valdivia, K. L. Kramer, B. Mason, D. E. Williams, and L. Johnston. Alternate benthic assemblages on reef restoration structures and cascading effects on coral settlement. Marine Ecology Progress Series, 387:147–156, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/387/m387p147.pdf.

Molinet:2006:ESP

[MVLM^{+}06] C. Molinet, A. Valle-Levinson, C. A. Moreno, M. Cáceres, M. Bello, and M. Castillo. Effects of sill processes on the distribution of epineustonic competent larvae in a stratified system of Southern Chile. Marine Ecology Progress Series, 324: 95–104, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/324/m324p095.pdf.

Maar:2006:TFB

[MVN^{+}06] M. Maar, A. W. Visser, T. G. Nielsen, A. Stips, and H. Saito. Turbulence and feeding behaviour affect the vertical distributions of Oithona similis and Microsetella norwegica. Marine Ecology Progress Series, 313:157–172, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p157.pdf.

Moksnes:2001:MED

[MW01] P. O. Moksnes and H. Wennhage. Methods for estimating decapod larval supply and settlement: importance of larval behavior and development stage. Marine Ecology Progress Series, 209:257–273, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/209/m209p257.pdf.

Mendes:2002:EBE

[MW02a] J. M. Mendes and J. D. Woodley. Effect of the 1995–1996 bleaching event on polyp tissue depth, growth, reproduction and skeletal band formation in Montastraea annularis. Marine Ecology Progress Series, 235:93–102, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p093.pdf.

Mendes:2002:TRM

[MW02b] J. M. Mendes and J. D. Woodley. Timing of reproduction in Montastraea annularis: relationship to environmental variables. Marine Ecology Progress Series, 227:241–251, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/227/m227p241.pdf.

Marinelli:2004:DRT

[MW04a] R. L. Marinelli and S. A. Woodin. Disturbance and recruitment: a test of solute and substrate specificity using Mercenaria mercenaria and Capitella sp. 1. Marine Ecology Progress Series, 269:209–221, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/269/m269p209.pdf.

Miao:2004:RBC

[MW04b] A. J. Miao and W. X. Wang. Relationships between cellspecific growth rate, and uptake rate of cadmium and zinc by a coastal diatom. Marine Ecology Progress Series, 275: 103–113, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/275/m275p103.pdf.

Martin:2006:ETS

[MW06] G. B. Martin and M. J. Wuenschel. Effect of temperature and salinity on otolith element incorporation in juvenile gray snapper Lutjanus griseus. Marine Ecology Progress Series, 324:229–239, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/324/m324p229.pdf.

McKenzie:2008:STV

[MW08] J. McKenzie and K. M. Wynne. Spatial and temporal variation in the diet of Steller sea lions in the Kodiak Archipelago, 1999 to 2005. Marine Ecology Progress Series, 360:265– 283, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/360/m360p265.pdf.

Mosher:2009:PLB

[MW09] C. V. Mosher and L. Watling. Partners for life: a brittle star and its octocoral host. Marine Ecology Progress Series, 397:81–88, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p081.pdf.

Murphy:2009:IBP

[MWD^{+}09] S. Murphy, A. Winship, W. Dabin, P. D. Jepson, R. Deaville, R. J. Reid, C. Spurrier, E. Rogan, A. López, A. F. González, F. L. Read, M. Addink, M. Silva, V. Ridoux, J. A. Learmonth, G. J. Pierce, and S. P. Northridge. Importance of biological parameters in assessing the status of Delphinus delphis. Marine Ecology Progress Series, 388:273–291, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/388/
m388p273.pdf.

Marston:2002:PAD

[MWG02] B. H. Marston, M. F. Willson, and S. M. Gende. Predator aggregations during eulachon Thaleichthys pacificus spawning runs. Marine Ecology Progress Series, 231:229–236, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/231/m231p229.pdf.

Mori:2005:CPR

[MWM^{+}05] Y. Mori, Y. Watanabe, Y. Mitani, K. Sato, M. F. Cameron, and Y. Naito. A comparison of prey richness estimates for Weddell seals using diving profiles and image data. Marine Ecology Progress Series, 295:257–263, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p257.pdf.

McKee:2005:PTE

[MWMH05] M. P. McKee, J. E. Ward, B. A. MacDonald, and B. A. Holohan. Production of transparent exopolymer particles (TEP) by the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 288:141–149, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/288/m288p141.pdf.

Marcoux:2007:SWF

[MWR07] M. Marcoux, H. Whitehead, and L. Rendell. Sperm whale feeding variation by location, year, social group and clan: evidence from stable isotopes. Marine Ecology Progress Series, 333:309–314, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p309.pdf.

Mitani:2004:DBW

[MWS^{+}04] Y. Mitani, Y. Watanabe, K. Sato, M. F. Cameron, and Y. Naito. 3D diving behavior of Weddell seals with respect to prey accessibility and abundance. Marine Ecology Progress Series, 281:275–281, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p275.pdf.

Madsen:2006:WTU

[MWT^{+}06] P. T. Madsen, M. Wahlberg, J. Tougaard, K. Lucke, and P. Tyack. Wind turbine underwater noise and marine mammals: implications of current knowledge and data needs. Marine Ecology Progress Series, 309:279–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p279.pdf.

Miron:2000:PCB

[MWTB00] G. Miron, L. J. Walters, R. Tremblay, and E. Bourget. Physiological condition and barnacle larval behavior: a preliminary look at the relationship between TAG/DNA ratio and larval substratum exploration in Balanus amphitrite. Marine Ecology Progress Series, 198:303–310, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/198/m198p303.pdf.

McCarthy:2002:GRB

[MY02] D. A. McCarthy and C. M. Young. Gametogenesis and reproductive behavior in the echinoid Lytechinus variegatus. Marine Ecology Progress Series, 233:157–168, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p157.pdf.

McCarthy:2004:EWB

[MY04] D. A. McCarthy and C. M. Young. Effects of water-borne gametes on the aggregation behavior of Lytechinus variegatus. Marine Ecology Progress Series, 283:191–198, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p191.pdf.

McCarthy:2003:IWI

[MYE03] D. A. McCarthy, C. M. Young, and R. H. Emson. Influence of wave-induced disturbance on seasonal spawning patterns in the sabellariid polychaete Phragmatopoma lapidosa. Marine Ecology Progress Series, 256:123–133, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p123.pdf.

Mercier:2007:LTS

[MYH07] Annie Mercier, Roberto H. Ycaza, and Jean-François Hamel. Long-term study of gamete release in a broadcast-spawning holothurian: predictable lunar and diel periodicities. Marine Ecology Progress Series, 329:179–189, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p179.pdf.

Miki:2008:IPL

[MYNY08] T. Miki, T. Yokokawa, T. Nagata, and N. Yamamura. Immigration of prokaryotes to local environments enhances remineralization efficiency of sinking particles: a metacommunity model. Marine Ecology Progress Series, 366:1–14, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moreno:2005:DHC

[MZD^{+}05] I. B. Moreno, A. N. Zerbini, D. Danilewicz, M. C. de Oliveira Santos, P. C. Simões-Lopes, J. Lailson-Brito, Jr., and A. F. Azevedo. Distribution and habitat characteristics of dolphins of the genus Stenella (Cetacea: Delphinidae) in the southwest Atlantic Ocean. Marine Ecology Progress Series, 300:229–240, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p229.pdf.

Mews:2006:SSD

[MZJ06] M. Mews, M. Zimmer, and D. E. Jelinski. Species-specific decomposition rates of beach-cast wrack in Barkley Sound, British Columbia, Canada. Marine Ecology Progress Series, 328:155–160, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p155.pdf.

Nemerson:2004:SPD

[NA04a] D. M. Nemerson and K. W. Able. Spatial patterns in diet and distribution of juveniles of four fish species in Delaware Bay marsh creeks: factors influencing fish abundance. Marine Ecology Progress Series, 276:249–262, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p249.pdf.

Nishizaki:2004:JAA

[NA04b] M. T. Nishizaki and J. D. Ackerman. Juvenile-adult associations in sea urchins Strongylocentrotus franciscanus and S. droebachiensis: Is nutrition involved? Marine Ecology Progress Series, 268:93–103, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/268/m268p093.pdf.

Nieto:2006:TVL

[NAB^{+}06] Ó. Nieto, J. Aboigor, R. Buján, M. N’Diaye, J. Graña, L. Saco-Álvarez, Á. Franco, J. A. Soriano, and R. Beiras. Temporal variation in the levels of polycyclic aromatic hydrocarbons (PAHs) off the Galician Coast after the ‘Prestige’ oil spill. Marine Ecology Progress Series, 328:41–49, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p041.pdf.

Nordstrom:2009:TVB

[NAB09] M. Nordström, K. Aarnio, and E. Bonsdorff. Temporal variability of a benthic food web: patterns and processes in a low-diversity system. Marine Ecology Progress Series, 378: 13–26, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/378/m378p013.pdf.

Naeem:2006:ESB

[Nae06] S. Naeem. Expanding scales in biodiversity-based research: challenges and solutions for marine systems. Marine Ecology Progress Series, 311:273–283, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m311p273.pdf.

Ng:2005:PCF

[NATR^{+}05] T. Y. T. Ng, C. Amiard-Triquet, P. S. Rainbow, J. C. Amiard, and W. X. Wang. Physico-chemical form of trace metals accumulated by phytoplankton and their assimilation by filter-feeding invertebrates. Marine Ecology Progress Series, 299:179–191, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p179.pdf.

Neuenfeldt:2006:EDP

[NB06a] S. Neuenfeldt and J. E. Beyer. Environmentally driven predator–prey overlaps determine the aggregate diet of the cod Gadus morhua in the Baltic Sea. Marine Ecology Progress Series, 310:151–163, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/310/m310p151.pdf.

Nugues:2006:DCA

[NB06b] M. M. Nugues and R. P. M. Bak. Differential competitive abilities between Caribbean coral species and a brown alga: a year of experiments and a long-term perspective. Marine Ecology Progress Series, 315:75–86, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/315/m315p075.pdf.

Nielsen:2004:HBP

[NBB^{+}04] T. G. Nielsen, P. K. Bjørnsen, P. Boonruang, M. Fryd, P. J. Hansen, V. Janekarn, V. Limtrakulvong, P. Munk, O. S. Hansen, S. Satapoomin, S. Sawangarreruks, H. A. Thomsen, and J. B. Østergaard. Hydrography, bacteria and protist communities across the continental shelf and shelf slope of the Andaman Sea (NE Indian Ocean). Marine Ecology Progress Series, 274:69–86, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p069.pdf.

Nevitt:2006:PRA

[NBB06] G. A. Nevitt, D. M. Bergstrom, and F. Bonadonna. The potential role of ammonia as a signal molecule for procellariiform seabirds. Marine Ecology Progress Series, 315:271– 277, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/315/m315p271.pdf.

Nyssen:2005:TPA

[NBDG05] F. Nyssen, T. Brey, P. Dauby, and M. Graeve. Trophic position of Antarctic amphipods — enhanced analysis by a 2dimensional biomarker assay. Marine Ecology Progress Series, 300:135–145, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p135.pdf.

Nagelkerken:2008:IBC

[NBN^{+}08] I. Nagelkerken, J. Bothwell, R. S. Nemeth, J. M. Pitt, and G. van der Velde. Interlinkage between Caribbean coral reefs and seagrass beds through feeding migrations by grunts (Haemulidae) depends on habitat accessibility. Marine Ecology Progress Series, 368:155–164, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p155.pdf.

Nieto-Cid:2004:CDP

[NCÁSBP04] M. Nieto-Cid, X. A. Álvarez-Salgado, S. Brea, and F. F. Pérez. Cycling of dissolved and particulate carbohydrates in a coastal upwelling system (NW Iberian Peninsula). Marine Ecology Progress Series, 283:39–54, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p039.pdf.

Nieto-Cid:2005:DFT

[NCÁSGP05] M. Nieto-Cid, X. A. Álvarez-Salgado, J. Gago, and F. F. Pérez. DOM fluorescence, a tracer for biogeochemical processes in a coastal upwelling system (NW Iberian Peninsula). Marine Ecology Progress Series, 297:33–50, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/297/
m297p033.pdf.

Newman:2007:ABD

[NCC^{+}07] S. H. Newman, A. Chmura, K. Converse, A. M. Kilpatrick, N. Patel, E. Lammers, and P. Daszak. Aquatic bird disease and mortality as an indicator of changing ecosystem health. Marine Ecology Progress Series, 352:299–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m352p299.pdf.

Nylund:2005:CIB

[NCHP05] G. M. Nylund, G. Cervin, M. Hermansson, and H. Pavia. Chemical inhibition of bacterial colonization by the red alga Bonnemaisonia hamifera. Marine Ecology Progress Series, 302:27–36, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p027.pdf.

Nelson:2009:PPZ

[NCMC09] R. J. Nelson, E. C. Carmack, F. A. McLaughlin, and G. A. Cooper. Penetration of Pacific zooplankton into the western Arctic Ocean tracked with molecular population genetics. Marine Ecology Progress Series, 381:129–138, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m381p129.pdf.

Newton:2009:SMS

[NCN^{+}09] K. M. Newton, D. A. Croll, H. M. Nevins, S. R. Benson, J. T. Harvey, and B. R. Tershy. At-sea mortality of seabirds based on beachcast and offshore surveys. Marine Ecology Progress Series, 392:295–305, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/392/m392p295.pdf.

Nylund:2008:SDA

[NCP^{+}08] G. M. Nylund, G. Cervin, F. Persson, M. Hermansson, P. D. Steinberg, and H. Pavia. Seaweed defence against bacteria: a poly-brominated 2-heptanone from the red alga Bonnemaisonia hamifera inhibits bacterial colonisation. Marine Ecology Progress Series, 369:39–50, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/369/m369p039.pdf.

Norkko:2001:LDJ

[NCT^{+}01] A. Norkko, V. J. Cummings, S. F. Thrush, J. E. Hewitt, and T. Hume. Local dispersal of juvenile bivalves: implications for sandflat ecology. Marine Ecology Progress Series, 212:131–144, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p131.pdf.

Nugues:2004:CDA

[NDB04] M. M. Nugues, L. Delvoye, and R. P. M. Bak. Coral defence against macroalgae: differential effects of mesenterial filaments on the green alga Halimeda opuntia. Marine Ecology Progress Series, 278:103–114, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p103.pdf.

Ninio:2003:ECB

[NDOS03] R. Ninio, S. Delean, K. Osborne, and H. Sweatman. Estimating cover of benthic organisms from underwater video images: variability associated with multiple observers. Marine Ecology Progress Series, 265:107–116, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/265/m265p107.pdf.

Nagelkerken:2000:DNS

[NDV^{+}00a] I. Nagelkerken, M. Dorenbosch, W. C. E. P. Verberk, E. Cocheret de la Morinière, and G. van der Velde. Daynight shifts of fishes between shallow-water biotopes of a Caribbean bay, with emphasis on the nocturnal feeding of Haemulidae and Lutjanidae. Marine Ecology Progress Series, 194:55–64, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p055.pdf.

Nagelkerken:2000:ISW

[NDV^{+}00b] I. Nagelkerken, M. Dorenbosch, W. C. E. P. Verberk, E. Cocheret de la Morinière, and G. van der Velde. Importance of shallow-water biotopes of a Caribbean bay for juvenile coral reef fishes: patterns in biotope association, community structure and spatial distribution. Marine Ecology Progress Series, 202:175–192, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/202/m202p175.pdf.

Natunewicz:2001:STS

[NE01] C. C. Natunewicz and C. E. Epifanio. Spatial and temporal scales of patches of crab larvae in coastal waters. Marine Ecology Progress Series, 212:217–222, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p217.pdf.

Nedved:2004:OPC

[Ned04] O. Nedved. Occurrence of the phylum Cycliophora in the Mediterranean. Marine Ecology Progress Series, 277:297– 299, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/277/m277p297.pdf.

Natunewicz:2001:TCL

[NEG01] C. C. Natunewicz, C. E. Epifanio, and R. W. Garvine. Transport of crab larval patches in the coastal ocean. Marine Ecology Progress Series, 222:143–154, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/222/m222p143.pdf.

Newsome:2007:HDP

[NEK^{+}07] S. D. Newsome, M. A. Etnier, C. M. Kurle, J. R. Waldbauer, C. P. Chamberlain, and P. L. Koch. Historic decline in primary productivity in western Gulf of Alaska and eastern Bering Sea: isotopic analysis of northern fur seal teeth. Marine Ecology Progress Series, 332:211–224, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p211.pdf.

Nemeth:2005:PCR

[Nem05] R. S. Nemeth. Population characteristics of a recovering US Virgin Islands red hind spawning aggregation following protection. Marine Ecology Progress Series, 286:81–97, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/286/m286p081.pdf.

Newsome:2009:RCO

[NEMF09] S. D. Newsome, M. A. Etnier, D. H. Monson, and M. L. Fogel. Retrospective characterization of ontogenetic shifts in killer whale diets via δ ^{13}C and δ ^{15}N analysis of teeth. Marine Ecology Progress Series, 374:229–242, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p229.pdf.

Norton:2006:MT

[NF06] D. W. Norton and H. M. Feder. Mytilus thermophily? Marine Ecology Progress Series, 309:301–303, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p301.pdf.

Norderhaug:2003:TIL

[NFN03] K. M. Norderhaug, S. Fredriksen, and K. Nygaard. Trophic importance of Laminaria hyperborea to kelp forest consumers and the importance of bacterial degradation to food quality. Marine Ecology Progress Series, 255:135–144, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p135.pdf.

Newland:2009:BFA

[NFN^{+}09] C. Newland, I. C. Field, P. D. Nichols, C. J. A. Bradshaw, and M. A. Hindell. Blubber fatty acid profiles indicate dietary resource partitioning between adult and juvenile southern elephant seals. Marine Ecology Progress Series, 384:303–312, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/384/m384p303.pdf.

Nejstgaard:2006:PDT

[NFV^{+}06] J. C. Nejstgaard, M. E. Frischer, P. G. Verity, J. T. Anderson, A. Jacobsen, M. J. Zirbel, A. Larsen, J. Martínez-Martínez, A. F. Sazhin, T. Walters, D. A. Bronk, S. J. Whipple, S. R. Borrett, B. C. Patten, and J. D. Long. Plankton development and trophic transfer in seawater enclosures with nutrients and Phaeocystis pouchetii added. Marine Ecology Progress Series, 321:99–121, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p099.pdf.

Nash:2007:DSW

[NGBG07] R. D. M. Nash, A. J. Geffen, M. T. Burrows, and R. N. Gibson. Dynamics of shallow-water juvenile flatfish nursery grounds: application of the self-thinning rule. Marine Ecology Progress Series, 344:231–244, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/344/m344p231.pdf.

Nylund:2007:SCV

[NGdN^{+}07] G. M. Nylund, P. E. Gribben, R. de Nys, P. D. Steinberg, and H. Pavia. Surface chemistry versus whole-cell extracts: antifouling tests with seaweed metabolites. Marine Ecology Progress Series, 329:73–84, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/329/m329p073.pdf.

Naviaux:2005:SDG

[NGM^{+}05] R. K. Naviaux, B. Good, J. D. McPherson, D. L. Steffen, D. Markusic, B. Ransom, and J. Corbeil. Sand DNA — a genetic library of life at the water’s edge. Marine Ecology Progress Series, 301:9–22, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/301/m301p009.pdf.

Nozais:2001:ABC

[NGMT01] C. Nozais, M. Gosselin, C. Michel, and G. Tita. Abundance, biomass, composition and grazing impact of the sea-ice meiofauna in the North Water, northern Baffin Bay. Marine Ecology Progress Series, 217:235–250, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/217/m217p235.pdf.

Navarro:2009:EDF

[NGS09] J. Navarro and J. González-Solís. Environmental determinants of foraging strategies in Cory’s shearwaters Calonectris diomedea. Marine Ecology Progress Series, 378:259– 267, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/378/m378p259.pdf.

Navarro:2007:NFE

[NGSV07] J. Navarro, J. González-Solís, and G. Viscor. Nutritional and feeding ecology in Cory’s shearwater Calonectris diomedea during breeding. Marine Ecology Progress Series, 351:261– 271, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/351/m351p261.pdf.

North:2003:ELE

[NH03a] E. W. North and E. D. Houde. Erratum: Linking ETM physics, zooplankton prey, and fish early-life histories tostriped bass Morone saxatilis and white perch M. americana recruitment, E. W. North, E. D. Houde, 260:219–236 (2003). Marine Ecology Progress Series, 264:309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p309.pdf. See [NH03b].

North:2003:LEP

[NH03b] E. W. North and E. D. Houde. Linking ETM physics, zooplankton prey, and fish early-life histories to striped bass Morone saxatilis and white perch M. americana recruitment. Marine Ecology Progress Series, 260:219–236, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/

m260p219.pdf. See erratum [NH03a].

Niehoff:2005:RCG

[NH05] B. Niehoff and H. J. Hirche. Reproduction of Calanus glacialis in the Lurefjord (western Norway): indication for temperature-induced female dormancy. Marine Ecology Progress Series, 285:107–115, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/285/m285p107.pdf.

Nomaki:2006:DIP

[NHN^{+}06] H. Nomaki, P. Heinz, T. Nakatsuka, M. Shimanaga, N. Ohkouchi, N. O. Ogawa, K. Kogure, E. Ikemoto, and H. Kitazato. Different ingestion patterns of ^{13}C-labeled bacteria and algae by deep-sea benthic Foraminifera. Marine Ecology Progress Series, 310:95–108, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m310p095.pdf.

Nejstgaard:2001:ZGD

[NHNB01] J. C. Nejstgaard, B. H. Hygum, L. J. Naustvoll, and U. Båmstedt. Zooplankton growth, diet and reproductive success compared in simultaneous diatom- and flagellatemicrozooplankton-dominated plankton blooms. Marine Ecology Progress Series, 221:77–91, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/221/m221p077.pdf.

Nakamura:2008:EOM

[NHS^{+}08] Y. Nakamura, M. Horinouchi, T. Shibuno, Y. Tanaka, T. Miyajima, I. Koike, H. Kurokura, and M. Sano. Evidence of ontogenetic migration from mangroves to coral reefs by black-tail snapper Lutjanus fulvus: stable isotope approach.

Marine Ecology Progress Series, 355:257–266, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/355/
m355p257.pdf.

Nogueira:2000:EMH

[NIF00] E. Nogueira, F. Ibanez, and F. G. Figueiras. Effect of meteorological and hydrographic disturbances on the microplankton community structure in the Ría de Vigo (NW Spain). Marine Ecology Progress Series, 203:23–45, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/203/m203p023.pdf.

Nizinski:2007:PSS

[Niz07] M. S. Nizinski. Predation in subtropical soft-bottom systems: spiny lobster and molluscs in Florida Bay. Marine Ecology Progress Series, 345:185–197, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/345/m345p185.pdf.

Nahas:2003:HMS

[NJPI03] E. L. Nahas, G. Jackson, C. B. Pattiaratchi, and G. N. Ivey. Hydrodynamic modelling of snapper Pagrus auratus egg and larval dispersal in Shark Bay, Western Australia: reproductive isolation at a fine spatial scale. Marine Ecology Progress Series, 265:213–226, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/265/m265p213.pdf.

Ng:2003:DEL

[NK03] T. Y. T. Ng and M. J. Keough. Delayed effects of larval exposure to Cu in the bryozoan Watersipora subtorquata. Marine Ecology Progress Series, 257:77–85, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/257/
m257p077.pdf.

Norling:2007:SFE

[NK07] P. Norling and N. Kautsky. Structural and functional effects of Mytilus edulis on diversity of associated species and ecosystem functioning. Marine Ecology Progress Series, 351: 163–175, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/351/m351p163.pdf.

Nagelkerken:2001:DCR

[NKK^{+}01] I. Nagelkerken, S. Kleijnen, T. Klop, R. A. C. J. van den Brand, E. C. de la Morinière, and G. van der Velde. Dependence of Caribbean reef fishes on mangroves and seagrass beds as nursery habitats: a comparison of fish faunas between bays with and without mangroves/seagrass beds. Marine Ecology Progress Series, 214:225–235, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p225.pdf.

Nishikawa:2003:LSR

[NKS03] A. Nishikawa, M. Katoh, and K. Sakai. Larval settlement rates and gene flow of broadcast-spawning (Acropora tenuis) and planula-brooding (Stylophora pistillata) corals. Marine Ecology Progress Series, 256:87–97, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p087.pdf.

Nozawa:2005:GRM

[NL05] Y. Nozawa and Y. Loya. Genetic relationship and maturity state of the allorecognition system affect contact reactions in juvenile Seriatopora corals. Marine Ecology Progress Series, 286:115–123, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p115.pdf.

Navarro:2008:ESC

[NLB08] N. Navarro, R. J. G. Leakey, and K. D. Black. Effect of salmon cage aquaculture on the pelagic environment of temperate coastal waters: seasonal changes in nutrients and microbial community. Marine Ecology Progress Series, 361: 47–58, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p047.pdf.

Ning:2005:CAB

[NLCS05] X. Ning, W. K. W. Li, Y. Cai, and J. Shi. Comparative analysis of bacterioplankton and phytoplankton in three ecological provinces of the northern South China Sea. Marine Ecology Progress Series, 293:17–28, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/293/m293p017.pdf.

Neira:2005:BMC

[NLG05] Carlos Neira, Lisa A. Levin, and Edwin Grosholz. Benthic macrofaunal communities of three sites in San Francisco Bay invaded by hybrid Spartina, with comparison to uninvaded habitats. Marine Ecology Progress Series, 292:111– 126, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p111.pdf.

Nye:2009:CSD

[NLHO09] J. A. Nye, J. S. Link, J. A. Hare, and W. J. Overholtz. Changing spatial distribution of fish stocks in relation to climate and population size on the Northeast United States continental shelf. Marine Ecology Progress Series, 393:111– 129, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/393/m393p111.pdf.

Nilsson:2000:MMO

[NLK00] P. G. Nilsson, J. S. Levinton, and J. P. Kurdziel. Migration of a marine oligochaete: induction of dispersal and microhabitat choice. Marine Ecology Progress Series, 207:89–96, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p089.pdf.

Nel:2001:EMO

[NLP^{+}01] D. C. Nel, J. R. E. Lutjeharms, E. A. Pakhomov, I. J. Ansorge, P. G. Ryan, and N. T. W. Klages. Exploitation of mesoscale oceanographic features by grey-headed albatross Thalassarche chrysostoma in the southern Indian Ocean. Marine Ecology Progress Series, 217:15–26, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/217/m217p015.pdf.

Nielsen:2004:GPG

[NM04] R. Nielsen and P. Munk. Growth pattern and growth dependent mortality of larval and pelagic juvenile North Sea cod Gadus morhua. Marine Ecology Progress Series, 278:

261–270, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/278/m278p261.pdf.

Nielsen:2007:EBM

[NM07] Torkel Gissel Nielsen and Marie Maar. Effects of a blue mussel Mytilus edulis bed on vertical distribution and composition of the pelagic food web. Marine Ecology Progress Series, 339:185–198, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p185.pdf.

Nonaka:2009:RHR

[NM09] M. Nonaka and K. Muzik. Recent harvest records of commercially valuable precious corals in the Ryukyu Archipelago. Marine Ecology Progress Series, 397:269–278, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p269.pdf.

Neubauer:2000:CCT

[NMA00] S. C. Neubauer, W. D. Miller, and I. Cofman Anderson. Carbon cycling in a tidal freshwater marsh ecosystem: a carbon gas flux study. Marine Ecology Progress Series, 199: 13–30, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/199/m199p013.pdf.

Nisbet:2002:USI

[NMBH02] I. C. T. Nisbet, J. P. Montoya, J. Burger, and J. J. Hatch. Use of stable isotopes to investigate individual differences in diets and mercury exposures among common terns Sterna hirundo in breeding and wintering grounds. Marine Ecology Progress Series, 242:267–274, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/242/m242p267.pdf.

Nice:2003:LTT

[NMCT03] H. E. Nice, D. Morritt, M. Crane, and M. Thorndyke. Longterm and transgenerational effects of nonylphenol exposure at a key stage in the development of Crassostrea gigas. Possible endocrine disruption? Marine Ecology Progress Series, 256:293–300, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p293.pdf.

Ninio:2000:TPC

[NMDS00] R. Ninio, M. Meekan, T. Done, and H. Sweatman. Temporal patterns in coral assemblages on the Great Barrier Reef from local to large spatial scales. Marine Ecology Progress Series, 194:65–74, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p065.pdf.

Nitschke:2002:EDD

[NMJ02] P. Nitschke, M. Mather, and F. Juanes. Evidence for densitydependent mortality in recruitment of a temperate reef fish, cunner Tautogolabrus adspersus, among similar reefs in the vicinity of an anthropogenic disturbance. Marine Ecology Progress Series, 226:165–178, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/226/m226p165.pdf.

Nakano:2003:ATB

[NMOO03] H. Nakano, H. Matsunaga, H. Okamoto, and M. Okazaki. Acoustic tracking of bigeye thresher shark Alopias superciliosus in the eastern Pacific Ocean. Marine Ecology Progress Series, 265:255–261, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/265/m265p255.pdf.

Nielsen:2002:EHR

[NMS^{+}02] T. G. Nielsen, E. F. Møller, S. Satapoomin, M. Ringuette, and R. R. Hopcroft. Egg hatching rate of the cyclopoid copepod Oithona similis in Arctic and temperate waters. Marine Ecology Progress Series, 236:301–306, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p301.pdf.

Nielsen:2006:PTM

[NMVH06] P. Nielsen, J. Mortensen, B. Vismann, and B. W. Hansen. Physiological tolerance of marine calanoid copepod eggs to sulphide. Marine Ecology Progress Series, 328:171–182, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/328/m328p171.pdf.

Nausch:2006:BDO

[NN06] M. Nausch and G. Nausch. Bioavailability of dissolved organic phosphorus in the Baltic Sea. Marine Ecology Progress Series, 321:9–17, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p009.pdf.

Narayanaswamy:2003:ALT

[NNG03] B. E. Narayanaswamy, T. D. Nickell, and J. D. Gage. Appropriate levels of taxonomic discrimination in deep-sea studies: species vs family. Marine Ecology Progress Series, 257: 59–68, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/257/m257p059.pdf.

Norstrom:2009:ASC

[NNLF09] A. V. Norström, M. Nyström, J. Lokrantz, and C. Folke. Alternative states on coral reefs: beyond coral-macroalgal phase shifts. Marine Ecology Progress Series, 376:295–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m376p295.pdf.

Narvaez:2006:OAW

[NNLV06] D. A. Narváez, S. A. Navarrete, J. Largier, and C. A. Vargas. Onshore advection of warm water, larval invertebrate settlement, and relaxation of upwelling off central Chile. Marine Ecology Progress Series, 309:159–173, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p159.pdf.

Nejstgaard:2001:CUM

[NNS01] J. C. Nejstgaard, L. J. Naustvoll, and A. Sazhin. Correcting for underestimation of microzooplankton grazing in bottle incubation experiments with mesozooplankton. Marine Ecology Progress Series, 221:59–75, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/221/m221p059.pdf.

Nausch:2004:PDD

[NNW04] M. Nausch, G. Nausch, and N. Wasmund. Phosphorus dynamics during the transition from nitrogen to phosphate limitation in the central Baltic Sea. Marine Ecology Progress Series, 266:15–25, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p015.pdf.

Noda:2004:LSV

[Nod04] T. Noda. Large-scale variability in recruitment of the barnacle Semibalanus cariosus: its cause and effects on the population density and predator. Marine Ecology Progress Series, 278:241–252, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p241.pdf.

Nakagawa:2004:ICP

[NOE^{+}04] Y. Nakagawa, T. Ota, Y. Endo, K. Taki, and H. Sugisaki. Importance of ciliates as prey of the euphausiid Euphausia pacifica in the NW North Pacific. Marine Ecology Progress Series, 271:261–266, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/271/m271p261.pdf.

Nomaki:2008:BFT

[NOO^{+}08] H. Nomaki, N. O. Ogawa, N. Ohkouchi, H. Suga, T. Toyofuku, M. Shimanaga, T. Nakatsuka, and H. Kitazato. Benthic Foraminifera as trophic links between phytodetritus and benthic metazoans: carbon and nitrogen isotopic evidence. Marine Ecology Progress Series, 357:153–164, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m357p153.pdf.

Nishida:2002:FMF

[NOP02] S. Nishida, S. Ohtsuka, and A. R. Parker. Functional morphology and food habits of deep-sea copepods of the genus Cephalophanes (Calanoida: Phaennidae): perception of bioluminescence as a strategy for food detection. Marine Ecology Progress Series, 227:157–171, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p157.pdf.

Nylund:2005:CVM

[NP05] G. M. Nylund and H. Pavia. Chemical versus mechanical inhibition of fouling in the red alga Dilsea carnosa. Marine Ecology Progress Series, 299:111–121, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p111.pdf.

Nevitt:2000:SSL

[NPLZ00] G. Nevitt, N. D. Pentcheff, K. J. Lohmann, and R. K. Zimmer. Den selection by the spiny lobster Panulirus argus: testing attraction to conspecific odors in the field. Marine Ecology Progress Series, 203:225–231, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/203/m203p225.pdf.

Nozais:2001:ACM

[NPM01] C. Nozais, R. Perissinotto, and S. Mundree. Annual cycle of microalgal biomass in a South African temporarily-open estuary: nutrient versus light limitation. Marine Ecology Progress Series, 223:39–48, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p039.pdf.

Nilsen:2006:MBP

[NPN06] M. Nilsen, T. Pedersen, and E. M. Nilssen. Macrobenthic biomass, productivity (P/B) and production in a highlatitude ecosystem, North Norway. Marine Ecology Progress Series, 321:67–77, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p067.pdf.

Norkko:2005:EFA

[NPTW05] J. Norkko, C. A. Pilditch, S. F. Thrush, and R. M. G. Wells. Effects of food availability and hypoxia on bivalves: the value of using multiple parameters to measure bivalve condition in environmental studies. Marine Ecology Progress Series, 298:205–218, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p205.pdf.

Newman:2006:ITL

[NPVW06] B. K. Newman, I. Papadopoulos, J. Vorsatz, and T. H. Wooldridge. Influence of temperature on the larval development of Upogebia africana and U. capensis (Decapoda: Thalassinidae: Upogebiidae) in the laboratory. Marine Ecology Progress Series, 325:165–180, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p165.pdf.

Nilsson:2000:SMB

[NR00] H. C. Nilsson and R. Rosenberg. Succession in marine benthic habitats and fauna in response to oxygen deficiency: analysed by sediment profile-imaging and by grab samples. Marine Ecology Progress Series, 197:139–149, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p139.pdf.

Neer:2007:SET

[NRC07] J. A. Neer, K. A. Rose, and E. Cortés. Simulating the effects of temperature on individual and population growth of Rhinoptera bonasus: a coupled bioenergetics and matrix modeling approach. Marine Ecology Progress Series, 329:211–223, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/329/m329p211.pdf. See erratum

[Ano07e].

Naumann:2009:CME

[NReZW09] M. S. Naumann, C. Richter, M. el Zibdah, and C. Wild. Coral mucus as an efficient trap for picoplanktonic cyanobacteria: implications for pelagic-benthic coupling in the reef ecosystem. Marine Ecology Progress Series, 385:65–76, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p065.pdf.

Norling:2007:IFB

[NRH^{+}07] K. Norling, R. Rosenberg, S. Hulth, A. Grémare, and E. Bonsdorff. Importance of functional biodiversity and speciesspecific traits of benthic fauna for ecosystem functions in marine sediment. Marine Ecology Progress Series, 332:11– 23, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m332p011.pdf.

Niemann:2004:RSG

[NRJB04] H. Niemann, C. Richter, H. M. Jonkers, and M. I. Badran. Red Sea gravity currents cascade near-reef phytoplankton to the twilight zone. Marine Ecology Progress Series, 269: 91–99, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/269/m269p091.pdf.

Necaise:2005:EHE

[NRM05] A. M. D. Necaise, S. W. Ross, and J. M. Miller. Estuarine habitat evaluation measured by growth of juvenile summer flounder Paralichthys dentatus in a North Carolina estuary. Marine Ecology Progress Series, 285:157–168, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p157.pdf.

Nicholls:2002:FNT

[NRP^{+}02] D. G. Nicholls, C. J. R. Robertson, P. A. Prince, M. D. Murray, K. J. Walker, and G. P. Elliott. Foraging niches of three Diomedea albatrosses. Marine Ecology Progress Series, 231:269–277, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p269.pdf.

Nielsen:2001:NPR

[NRPS^{+}01] K. Nielsen, N. Risgaard-Petersen, B. Sømod, S. Rysgaard, and T. Bergø. Nitrogen and phosphorus retention estimated independently by flux measurements and dynamic modelling in the estuary, Randers Fjord, Denmark. Marine Ecology Progress Series, 219:25–40, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p025.pdf.

Nagelkerken:2002:HIM

[NRvdV^{+}02] I. Nagelkerken, C. M. Roberts, G. van der Velde, M. Dorenbosch, M. C. van Riel, E. Cocheret de la Morinière, and P. H. Nienhuis. How important are mangroves and seagrass beds for coral-reef fish? The nursery hypothesis tested on an island scale. Marine Ecology Progress Series, 244:299– 305, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p299.pdf.

Nichol:2002:DVM

[NS02] D. G. Nichol and D. A. Somerton. Diurnal vertical migration of the Atka mackerel Pleurogrammus monopterygius as shown by archival tags. Marine Ecology Progress Series, 239:193–207, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p193.pdf.

Nydam:2007:PEF

[NS07] M. Nydam and J. J. Stachowicz. Predator effects on fouling community development. Marine Ecology Progress Series, 337:93–101, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/337/m337p093.pdf.

Neckles:2005:DEZ

[NSBK05] H. A. Neckles, F. T. Short, S. Barker, and B. S. Kopp. Disturbance of eelgrass Zostera marina by commercial mussel Mytilus edulis harvesting in Maine: dragging impacts and habitat recovery. Marine Ecology Progress Series, 285: 57–73, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p057.pdf.

Newman:2000:SSG

[NSEL00] S. J. Newman, R. A. Steckis, J. S. Edmonds, and J. Lloyd. Stock structure of the goldband snapper Pristipomoides multidens (Pisces: Lutjanidae) from the waters of northern and western Australia by stable isotope ratio analysis of sagittal otolith carbonate. Marine Ecology Progress Series, 198: 239–247, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/198/m198p239.pdf.

Negri:2009:BIA

[NSFW09] A. P. Negri, R. M. Soo, F. Flores, and N. S. Webster. Bacillus insecticides are not acutely harmful to corals and sponges.

Marine Ecology Progress Series, 381:157–165, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m381p157.pdf.

North:2008:VSB

[NSH^{+}08] E. W. North, Z. Schlag, R. R. Hood, M. Li, L. Zhong, T. Gross, and V. S. Kennedy. Vertical swimming behavior influences the dispersal of simulated oyster larvae in a coupled particle-tracking and hydrodynamic model of Chesapeake Bay. Marine Ecology Progress Series, 359:99–115, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/359/m359p099.pdf.

Naqvi:2002:CCN

[NSJ02] S. W. A. Naqvi, V. V. S. S. Sarma, and D. A. Jayakumar. Carbon cycling in the northern Arabian Sea during the northeast monsoon: significance of salps. Marine Ecology Progress Series, 226:35–44, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p035.pdf.

Neulinger:2006:PTP

[NSSI06] S. C. Neulinger, H. Sahling, J. Süling, and J. F. Imhoff. Presence of two phylogenetically distinct groups in the deepsea mussel Acharax (Mollusca: Bivalvia: Solemyidae). Marine Ecology Progress Series, 312:161–168, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p161.pdf.

Norkko:2006:EEI

[NT06] J. Norkko and S. F. Thrush. Ecophysiology in environmental impact assessment: implications of spatial differences in seasonal variability of bivalve condition. Marine Ecology Progress Series, 326:175–186, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/326/m326p175.pdf.

Nunn:2007:MP

[NT07] B. L. Nunn and A. T. Timperman. Marine proteomics. Marine Ecology Progress Series, 332:281–289, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m332p281.pdf.

Newell:2009:RBR

[NTGK09] R. I. E. Newell, S. T. Tettelbach, C. J. Gobler, and D. G. Kimmel. Relationships between reproduction in suspensionfeeding hard clams Mercenaria mercenaria and phytoplankton community structure. Marine Ecology Progress Series, 387:179–196, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/387/m387p179.pdf.

Norkko:2002:SES

[NTH^{+}02] A. Norkko, S. F. Thrush, J. E. Hewitt, V. J. Cummings, J. Norkko, J. I. Ellis, G. A. Funnell, D. Schultz, and I. Mac-Donald. Smothering of estuarine sandflats by terrigenous clay: the role of wind-wave disturbance and bioturbation in sitedependent macrofaunal recovery. Marine Ecology Progress Series, 234:23–42, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p023.pdf.

Nozawa:2008:SDH

[NTN08] Y. Nozawa, M. Tokeshi, and S. Nojima. Structure and dynamics of a high-latitude scleractinian coral community in amakusa, southwestern Japan. Marine Ecology Progress Series, 358:151–160, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/358/m358p151.pdf.

Nugues:2002:ICD

[Nug02] M. M. Nugues. Impact of a coral disease outbreak on coral communities in St. Lucia: What and how much has been lost? Marine Ecology Progress Series, 229:61–71, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p061.pdf.

Nagelkerken:2002:DNE

[NvdV02] I. Nagelkerken and G. van der Velde. Do non-estuarine mangroves harbour higher densities of juvenile fish than adjacent shallow-water and coral reef habitats in Curaçao (Netherlands Antilles)? Marine Ecology Progress Series, 245:191– 204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p191.pdf.

Nagelkerken:2004:CMI

[NvdV04a] I. Nagelkerken and G. van der Velde. Are Caribbean mangroves important feeding grounds for juvenile reef fish from adjacent seagrass beds? Marine Ecology Progress Series, 274:143–151, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p143.pdf.

Nagelkerken:2004:RII

[NvdV04b] I. Nagelkerken and G. van der Velde. Relative importance of interlinked mangroves and seagrass beds as feeding habitats for juvenile reef fish on a Caribbean island. Marine Ecology Progress Series, 274:153–159, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p153.pdf.

Nagelkerken:2006:SAM

[NvdVVD06] I. Nagelkerken, G. van der Velde, W. C. E. P. Verberk, and M. Dorenbosch. Segregation along multiple resource axes in a tropical seagrass fish community. Marine Ecology Progress Series, 308:79–89, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/308/m308p079.pdf.

Nakamura:2001:WFR

[NvW01] T. Nakamura and R. van Woesik. Water-flow rates and passive diffusion partially explain differential survival of corals during the 1998 bleaching event. Marine Ecology Progress Series, 212:301–304, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p301.pdf.

Nakamura:2005:PPR

[NvWY05] T. Nakamura, R. van Woesik, and H. Yamasaki. Photoinhibition of photosynthesis is reduced by water flow in the reefbuilding coral Acropora digitifera. Marine Ecology Progress Series, 301:109–118, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/301/m301p109.pdf.

Nishimoto:2002:PCE

[NW02] M. M. Nishimoto and L. Washburn. Patterns of coastal eddy circulation and abundance of pelagic juvenile fish in the Santa Barbara Channel, California, USA. Marine Ecology Progress Series, 241:183–199, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/241/m241p183.pdf.

Ng:2004:DEA

[NW04] T. Y. T. Ng and W. X. Wang. Detoxification and effects of Ag, Cd, and Zn pre-exposure on metal uptake kinetics in the clam Ruditapes philippinarum. Marine Ecology Progress Series, 268:161–172, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/268/m268p161.pdf.

Neilson:2005:MSE

[NW05a] M. E. Neilson and R. R. Wilson, Jr. mtDNA singletons as evidence of a post-invasion genetic bottleneck in yellowfin goby Acanthogobius flavimanus from San Francisco Bay, California. Marine Ecology Progress Series, 296:197–208, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/296/
m296p197.pdf.

Ng:2005:SCS

[NW05b] T. Y. T. Ng and W. X. Wang. Subcellular controls of silver biokinetics in the green mussel Perna viridis from two hydrographic zones. Marine Ecology Progress Series, 299: 193–204, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p193.pdf.

Ng:2006:IVF

[NW06] J. S. S. Ng and G. A. Williams. Intraspecific variation in foraging behaviour: influence of shore height on temporal organization of activity in the chiton Acanthopleura japonica. Marine Ecology Progress Series, 321:183–192, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p183.pdf.

Nizzoli:2006:ICM

[NWFV06] D. Nizzoli, D. T. Welsh, E. A. Fano, and P. Viaroli. Impact of clam and mussel farming on benthic metabolism and nitrogen cycling, with emphasis on nitrate reduction pathways. Marine Ecology Progress Series, 315:151–165, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/315/
m315p151.pdf.

Needoba:2003:NIF

[NWHC03] J. A. Needoba, N. A. Waser, P. J. Harrison, and S. E. Calvert. Nitrogen isotope fractionation in 12 species of marine phytoplankton during growth on nitrate. Marine Ecology Progress Series, 255:81–91, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p081.pdf.

Negri:2001:MBS

[NWHH01] A. P. Negri, N. S. Webster, R. T. Hill, and A. J. Heyward. Metamorphosis of broadcast spawning corals in response to bacteria isolated from crustose algae. Marine Ecology Progress Series, 223:121–131, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/223/m223p121.pdf.

Nielsen:2009:MPS

[NWHH^{+}09] E. E. Nielsen, P. J. Wright, J. Hemmer-Hansen, N. A. Poulsen, I. M. Gibb, and D. Meldrup. Microgeographical population structure of cod Gadus morhua in the North Sea and west of Scotland: the role of sampling loci and individuals. Marine Ecology Progress Series, 376:213–225, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p213.pdf.

Nordstrom:2008:EQF

[NWIT08] C. A. Nordstrom, L. J. Wilson, S. J. Iverson, and D. J. Tollit. Evaluating quantitative fatty acid signature analysis (QFASA) using harbour seals Phoca vitulina richardsi in captive feeding studies. Marine Ecology Progress Series, 360:245–263, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p245.pdf.

Natanson:2008:OVG

[NWJ^{+}08] L. J. Natanson, S. P. Wintner, F. Johansson, A. Piercy, P. Campbell, A. De Maddalena, S. J. B. Gulak, B. Human, F. C. Fulgosi, D. A. Ebert, F. Hemida, F. H. Mollen, S. Vanni, G. H. Burgess, L. J. V. Compagno, and A. Wedderburn-Maxwell. Ontogenetic vertebral growth patterns in the basking shark Cetorhinus maximus. Marine Ecology Progress Series, 361:267–278, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p267.pdf.

Neideman:2003:CBI

[NWÓ03] R. Neideman, J. Wenngren, and E. Ólafsson. Competition between the introduced polychaete Marenzelleria sp. and the native amphipod Monoporeia affinis in Baltic soft bottoms. Marine Ecology Progress Series, 264:49–55, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p049.pdf.

Ni:2000:TCC

[NWT00] I. H. Ni, W. X. Wang, and Y. K. Tam. Transfer of Cd, Cr and Zn from zooplankton prey to mudskipper Periophthalmus cantonensis and glassy Ambassis urotaenia fishes. Marine Ecology Progress Series, 194:203–210, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/194/m194p203.pdf.

Nagai:2003:LPM

[NYK03] T. Nagai, H. Yamazaki, and D. Kamykowski. A Lagrangian photoresponse model coupled with 2nd-order turbulence closure. Marine Ecology Progress Series, 265:17– 30, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p017.pdf.

Nakamura:2003:WFF

[NYvW03] T. Nakamura, H. Yamasaki, and R. van Woesik. Water flow facilitates recovery from bleaching in the coral Stylophora pistillata. Marine Ecology Progress Series, 256:287– 291, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/256/m256p287.pdf.

Neely:2008:AMP

[NZ08] A. E. Neely and R. N. Zajac. Applying marine protected area design models in large estuarine systems. Marine Ecology Progress Series, 373:11–23, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/373/m373p011.pdf.

Nicastro:2007:BRI

[NZM07] K. R. Nicastro, G. I. Zardi, and C. D. McQuaid. Behavioural response of invasive Mytilus galloprovincialis and indigenous Perna perna mussels exposed to risk of predation. Marine Ecology Progress Series, 336:169–175, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p169.pdf.

Nicastro:2008:MBM

[NZM08a] K. R. Nicastro, G. I. Zardi, and C. D. McQuaid. Movement behaviour and mortality in invasive and indigenous mussels: resilience and resistance strategies at different spatial scales. Marine Ecology Progress Series, 372:119–126, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p119.pdf.

Nicastro:2008:CTD

[NZM^{+}08b] K. R. Nicastro, G. I. Zardi, C. D. McQuaid, P. R. Teske, and N. P. Barker. Coastal topography drives genetic structure in marine mussels. Marine Ecology Progress Series, 368:189–195, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p189.pdf.

Ouellet:2002:SIV

[OA02] P. Ouellet and J. P. Allard. Seasonal and interannual variability in larval lobster Homarus americanus size, growth and condition in the Magdalen Islands, southern Gulf of St. Lawrence. Marine Ecology Progress Series, 230:241–251, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p241.pdf.

Olsen:2007:PMZ

[OAGV07] Yngvar Olsen, Tom Andersen, Ingrid Gismervik, and Olav Vadstein. Protozoan and metazoan zooplankton-mediated carbon flows in nutrient-enriched coastal planktonic communities. Marine Ecology Progress Series, 331:67–83, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m331p067.pdf.

Ohji:2002:ETE

[OAM02] M. Ohji, T. Arai, and N. Miyazaki. Effects of tributyltin exposure in the embryonic stage on sex ratio and survival rate in the caprellid amphipod Caprella danilevskii. Marine Ecology Progress Series, 235:171–176, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/235/m235p171.pdf.

Ohji:2006:TTP

[OAM06] M. Ohji, T. Arai, and N. Miyazaki. Transfer of tributyltin from parental female to offspring in the viviparous surfperch Ditrema temmincki. Marine Ecology Progress Series, 307: 307–310, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p307.pdf.

Oedekoven:2001:VRS

[OAS01] C. S. Oedekoven, D. G. Ainley, and L. B. Spear. Variable responses of seabirds to change in marine climate: California Current, 1985–1994. Marine Ecology Progress Series, 212:265–281, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p265.pdf.

Otero:2008:BCC

[OÁSG^{+}08] J. Otero, X. A. Álvarez-Salgado, Á. F. González, A. Miranda, S. B. Groom, J. M. Cabanas, G. Casas, B. Wheatley, and Á. Guerra. Bottom-up control of common octopus Octopus vulgaris in the Galician upwelling system, northeast Atlantic Ocean. Marine Ecology Progress Series, 362:

181–192, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/362/m362p181.pdf.

Otero:2009:HFC

[OÁSG^{+}09] J. Otero, X. A. Álvarez-Salgado, Á. F. González, M. Gilcoto, and Á. Guerra. High-frequency coastal upwelling events influence Octopus vulgaris larval dynamics on the NW Iberian shelf. Marine Ecology Progress Series, 386:123–132, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/386/m386p123.pdf.

Oosterhuis:2000:REC

[OBB00] S. S. Oosterhuis, M. A. Baars, and W. C. M. Klein Breteler. Release of the enzyme chitobiase by the copepod Temora longicornis: characteristics and potential tool for estimating crustacean biomass production in the sea. Marine Ecology Progress Series, 196:195–206, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/196/m196p195.pdf.

Osterblom:2001:DFB

[ÖBFO01] H. Österblom, A. Bignert, T. Fransson, and O. Olsson. A decrease in fledging body mass in common guillemot Uria aalge chicks in the Baltic Sea. Marine Ecology Progress Series, 224:305–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/224/m224p305.pdf.

Olavarria:2007:PSS

[OBG^{+}07] C. Olavarría, C. S. Baker, C. Garrigue, M. Poole, N. Hauser, S. Caballero, L. Flórez-González, M. Brasseur, J. Bannister, J. Capella, P. Clapham, R. Dodemont, M. Donoghue, C. Jenner, M. N. Jenner, D. Moro, M. Oremus, D. Paton, H. Rosenbaum, and K. Russell. Population structure of South Pacific humpback whales and the origin of the eastern Polynesian breeding grounds. Marine Ecology Progress Series, 330:257–268, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p257.pdf.

Olabarria:2001:CPS

[OC01] C. Olabarria and M. G. Chapman. Comparison of patterns of spatial variation of microgastropods between 2 contrasting intertidal habitats. Marine Ecology Progress Series, 220: 201–211, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/220/m220p201.pdf.

OConnor:2007:SMM

[O’C07] N. J. O’Connor. Stimulation of molting in megalopae of the Asian shore crab Hemigrapsus sanguineus: physical and chemical cues. Marine Ecology Progress Series, 352:1–8, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Oleson:2007:BCC

[OCB^{+}07] E. M. Oleson, J. Calambokidis, W. C. Burgess, M. A. Mc-Donald, C. A. LeDuc, and J. A. Hildebrand. Behavioral context of call production by eastern North Pacific blue whales. Marine Ecology Progress Series, 330:269–284, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m330p269.pdf.

Osterblom:2006:FST

[ÖCOB06] H. Österblom, M. Casini, O. Olsson, and A. Bignert. Fish, seabirds and trophic cascades in the Baltic Sea. Marine Ecology Progress Series, 323:233–238, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m323p233.pdf.

ODonnell:2008:RWF

[O’D08] M. J. O’Donnell. Reduction of wave forces within bare patches in mussel beds. Marine Ecology Progress Series, 362:157– 167, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m362p157.pdf.

Olive:2009:NGL

[ODA^{+}09] P. J. W. Olive, T. Duangchinda, E. Ashforth, S. Craig, A. C. Ward, and S. J. Davies. Net gain of long-chain polyunsaturated fatty acids (PUFA) in a lugworm Arenicola marina bioturbated mesocosm. Marine Ecology Progress Series, 387:223–239, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/387/m387p223.pdf.

Ogburn:2009:MRE

[ODF09] M. B. Ogburn, H. Diaz, and R. B. Forward, Jr. Mechanisms regulating estuarine ingress of blue crab Callinectes sapidus megalopae. Marine Ecology Progress Series, 389: 181–192, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/389/m389p181.pdf.

Ouellette:2004:ETL

[ODG^{+}04] D. Ouellette, G. Desrosiers, J. P. Gagne, F. Gilbert, J. C. Poggiale, P. U. Blier, and G. Stora. Effects of temperature on in vitro sediment reworking processes by a gallery biodiffusor, the polychaete Neanthes virens. Marine Ecology Progress Series, 266:185–193, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p185.pdf.

Olesen:2002:DAP

[OEDSJ02] B. Olesen, S. Enríquez, C. M. Duarte, and K. Sand-Jensen. Depth-acclimation of photosynthesis, morphology and demography of Posidonia oceanica and Cymodocea nodosa in the Spanish Mediterranean Sea. Marine Ecology Progress Series, 236:89–97, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p089.pdf.

Orth:2003:SDE

[OFHM03] R. J. Orth, J. R. Fishman, M. C. Harwell, and S. R. Marion. Seed-density effects on germination and initial seedling establishment in eelgrass Zostera marina in the Chesapeake Bay region. Marine Ecology Progress Series, 250:71–79, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/250/m250p071.pdf.

Orejas:2003:RSP

[OGA03] C. Orejas, J. M. Gili, and W. Arntz. The role of the small planktonic communities in the diet of two Antarctic octocorals (Primnoisis antarctica and Primnoella sp.). Marine Ecology Progress Series, 250:105–116, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/250/m250p105.pdf.

Orejas:2009:CWC

[OGL^{+}09a] C. Orejas, A. Gori, C. Lo Iacono, P. Puig, J. M. Gili, and M. R. T. Dale. Cold-water corals in the cap de creus canyon, northwestern Mediterranean: spatial distribution, density and anthropogenic impact. Marine Ecology Progress Series, 397: 37–51, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m397p037.pdf.

Ostrovsky:2009:IEM

[OGL09b] A. N. Ostrovsky, D. P. Gordon, and S. Lidgard. Independent evolution of matrotrophy in the major classes of Bryozoa: transitions among reproductive patterns and their ecological background. Marine Ecology Progress Series, 378: 113–124, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/378/m378p113.pdf.

Orbea:2006:EVP

[OGMC06] A. Orbea, L. Garmendia, I. Marigómez, and M. P. Cajaraville. Effects of the ‘Prestige’ oil spill on cellular biomarkers in intertidal mussels: results of the first year of studies. Marine Ecology Progress Series, 306:177–189, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p177.pdf.

Oguz:2007:NRB

[Ogu07] T. Oguz. Nonlinear response of Black Sea pelagic fish stocks to over-exploitation. Marine Ecology Progress Series, 345: 211–228, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p211.pdf.

Orth:2000:RIS

[OHB^{+}00] R. J. Orth, M. C. Harwell, E. M. Bailey, A. Bartholomew, J. T. Jawad, A. V. Lombana, K. A. Moore, J. M. Rhode, and H. E. Woods. A review of issues in seagrass seed dormancy and germination: implications for conservation and restoration. Marine Ecology Progress Series, 200:277–288, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p277.pdf.

ODea:2004:LHV

[OHCFJ04] A. O’Dea, A. Herrera-Cubilla, H. Fortunato, and J. B. C. Jackson. Life history variation in cupuladriid bryozoans from either side of the Isthmus of Panama. Marine Ecology Progress Series, 280:145–161, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/280/m280p145.pdf.

Ou:2006:PSP

[OHL^{+}06] L. Ou, B. Huang, L. Lin, H. Hong, F. Zhang, and Z. Chen. Phosphorus stress of phytoplankton in the Taiwan Strait determined by bulk and single-cell alkaline phosphatase activity assays. Marine Ecology Progress Series, 327:95–106, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/327/m327p095.pdf.

Oh:2001:FEC

[OHN01] C. W. Oh, R. G. Hartnoll, and R. D. M. Nash. Feeding ecology of the common shrimp Crangon crangon in Port Erin Bay, Isle of Man, Irish Sea. Marine Ecology Progress Series, 214:211–223, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/214/m214p211.pdf.

Ohtsuka:2004:MHS

[OHS^{+}04] S. Ohtsuka, M. Hora, T. Suzaki, M. Arikawa, G. Omura, and K. Yamada. Morphology and host-specificity of the apostome ciliate Vampyrophrya pelagica infecting pelagic copepods in the Seto Inland Sea, Japan. Marine Ecology Progress Series, 282:129–142, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p129.pdf.

Ohtsuka:2007:TPD

[OHS^{+}07] S. Ohtsuka, S. Harada, M. Shimomura, G. A. Boxshall, R. Yoshizaki, D. Ueno, Y. Nitta, S. Iwasaki, H. Okawachi, and T. Sakakihara. Temporal partitioning: dynamics of alternating occupancy of a host microhabitat by two different crustacean parasites. Marine Ecology Progress Series, 348:261–272, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p261.pdf.

Orth:2002:PSS

[OHT02] R. J. Orth, K. L. Heck, Jr., and D. J. Tunbridge. Predation on seeds of the seagrass Posidonia australis in Western Australia. Marine Ecology Progress Series, 244:81–88, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p081.pdf.

Oldham:2004:ECF

[OIP04] C. E. Oldham, G. N. Ivey, and C. Pullin. Estimation of a characteristic friction velocity in stirred benthic chambers. Marine Ecology Progress Series, 279:291–295, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p291.pdf.

Osovitz:2002:BIB

[OJ02] C. J. Osovitz and D. Julian. Burrow irrigation behavior of Urechis caupo, a filter-feeding marine invertebrate, in its natural habitat. Marine Ecology Progress Series, 245:149– 155, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p149.pdf.

OConnor:2004:MFC

[OJ04] N. J. O’Connor and M. L. Judge. Molting of fiddler crab Uca minax megalopae: stimulatory cues are specific to salt marshes. Marine Ecology Progress Series, 282:229– 236, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p229.pdf.

OFoighil:2001:PRM

[OJPM01] D. O’Foighil, R. Jennings, J. K. Park, and D. A. Merriwether. Phylogenetic relationships of mid-oceanic ridge and continental lineages of Lasaea spp. (Mollusca: Bivalvia) in the northeastern Atlantic. Marine Ecology Progress Series, 213:165–175, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p165.pdf.

Ohizumi:2000:PSD

[OKAM00] H. Ohizumi, T. Kuramochi, M. Amano, and N. Miyazaki. Prey switching of Dall’s porpoise Phocoenoides dalli with population decline of Japanese pilchard Sardinops melanostictus around Hokkaido, Japan. Marine Ecology Progress Series, 200:265–275, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p265.pdf.

Onitsuka:2007:SSA

[OKH^{+}07] T. Onitsuka, T. Kawamura, T. Horii, N. Takiguchi, H. Takami, and Y. Watanabe. Synchronized spawning of abalone Haliotis diversicolor triggered by typhoon events in Sagami Bay, Japan. Marine Ecology Progress Series, 351: 129–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/351/m351p129.pdf.

Orth:2006:PPA

[OKM06] R. J. Orth, G. A. Kendrick, and S. R. Marion. Predation on Posidonia australis seeds in seagrass habitats of Rottnest Island, Western Australia: patterns and predators. Marine Ecology Progress Series, 313:105–114, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p105.pdf.

OReilly:2006:DCB

[OKP06] R. O’Reilly, R. Kennedy, and A. Patterson. Destruction of conspecific bioturbation structures by Amphiura filiformis (Ophiuroida): evidence from luminophore tracers and in situ time-lapse sediment-profile imagery. Marine Ecology Progress Series, 315:99–111, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p099.pdf.

Ochwada:2009:CAH

[OLG^{+}09] F. Ochwada, N. R. Loneragan, C. A. Gray, I. M. Suthers, and M. D. Taylor. Complexity affects habitat preference and predation mortality in postlarval Penaeus plebejus: implications for stock enhancement. Marine Ecology Progress Series, 380:161–171, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p161.pdf.

Orejas:2002:DRE

[OLGG^{+}02] C. Orejas, P. J. López-González, J. M. Gili, N. Teixidó, J. Gutt, and W. E. Arntz. Distribution and reproductive ecology of the Antarctic octocoral Ainigmaptilon antarcticum in the Weddell Sea. Marine Ecology Progress Series, 231: 101–114, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p101.pdf.

Olson:2006:CFS

[OLWB06] M. B. Olson, E. J. Lessard, C. H. J. Wong, and M. J. Bernhardt. Copepod feeding selectivity on microplankton, including the toxigenic diatoms Pseudo-nitzschia spp., in the coastal Pacific northwest. Marine Ecology Progress Series, 326:207–220, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p207.pdf.

Overholtzer-McLeod:2004:VRS

[OM04] K. L. Overholtzer-McLeod. Variance in reef spatial structure masks density dependence in coral-reef fish populations on natural versus artificial reefs. Marine Ecology Progress Series, 276:269–280, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p269.pdf.

Ordines:2005:VSR

[OMP^{+}05] F. Ordines, J. Moranta, M. Palmer, A. Lerycke, A. Suau, B. Morales-Nin, and A. M. Grau. Variations in a shallow rocky reef fish community at different spatial scales in the western Mediterranean Sea. Marine Ecology Progress Series, 304:221–233, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p221.pdf.

Ohi:2006:LAI

[OMT06] N. Ohi, A. Mizobuchi, and S. Taguchi. Light absorption of Isochrysis galbana (Prymnesiophyceae) under a day-night cycle with ultraviolet radiation. Marine Ecology Progress Series, 316:85–93, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/316/m316p085.pdf.

Olli:2004:EPE

[ONA04] K. Olli, M. G. Neubert, and D. M. Anderson. Encystment probability and encystment rate: new terms to quantitatively describe formation of resting cysts in planktonic microbial populations. Marine Ecology Progress Series, 273: 43–48, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/273/m273p043.pdf.

Ostlund-Nilsson:2005:PIA

[ÖNCNG05] S. Östlund-Nilsson, L. Curtis, G. E. Nilsson, and A. S. Grutter. Parasitic isopod Anilocra apogonae, a drag for the cardinal fish Cheilodipterus quinquelineatus. Marine Ecology Progress Series, 287:209–216, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/287/m287p209.pdf.

Okazaki:2003:OEA

[ONKK03] Y. Okazaki, H. Nakata, S. Kimura, and A. Kasai. Offshore entrainment of anchovy larvae and its implication for their survival in a frontal region of the Kuroshio. Marine Ecology Progress Series, 248:237–244, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/248/m248p237.pdf.

Olseng:2002:GHD

[ONP02] C. D. Olseng, L. J. Naustvoll, and E. Paasche. Grazing by the heterotrophic dinoflagellate Protoperidinium steinii on a Ceratium bloom. Marine Ecology Progress Series, 225:161– 167, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/225/m225p161.pdf.

ONeal:2002:RCP

[OP02] W. O’Neal and J. R. Pawlik. A reappraisal of the chemical and physical defenses of Caribbean gorgonian corals against predatory fishes. Marine Ecology Progress Series, 240:117– 126, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/240/m240p117.pdf.

Oppel:2008:AKE

[OP08] S. Oppel and A. N. Powell. Assigning king eiders to wintering regions in the Bering Sea using stable isotopes of feathers and claws. Marine Ecology Progress Series, 373:149–156, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m373p149.pdf.

Oliver:2009:DSR

[OP09] T. A. Oliver and S. R. Palumbi. Distributions of stressresistant coral symbionts match environmental patterns at local but not regional scales. Marine Ecology Progress Series, 378:93–103, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p093.pdf.

Oremus:2007:IIA

[OPSB07] M. Oremus, M. M. Poole, D. Steel, and C. S. Baker. Isolation and interchange among insular spinner dolphin communities in the South Pacific revealed by individual identification and genetic diversity. Marine Ecology Progress Series, 336:275–289, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/336/m336p275.pdf.

Orsini:2004:MRT

[OPSM04] L. Orsini, G. Procaccini, D. Sarno, and M. Montresor. Multiple rDNA ITS-types within the diatom Pseudo-nitzschia delicatissima (Bacillariophyceae) and their relative abundances across a spring bloom in the Gulf of Naples. Marine Ecology Progress Series, 271:87–98, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/271/m271p087.pdf.

Orbach:2008:CCO

[Orb08] M. K. Orbach. Cultural context of ocean fertilization. Marine Ecology Progress Series, 364:235–242, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m364p235.pdf.

Olaso:2000:TEF

[ORD00] I. Olaso, M. Rauschert, and C. De Broyer. Trophic ecology of the family artedidraconidae (Pisces: Osteichthyes) and its impact on the eastern Weddell Sea benthic system. Marine Ecology Progress Series, 194:143–158, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/194/m194p143.pdf.

Oresland:2000:DFC

[Øre00] V. Øresland. Diel feeding of the chaetognath Sagitta enflata in the Zanzibar Channel, western Indian Ocean. Marine Ecology Progress Series, 193:117–123, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p117.pdf.

Ottosen:2001:DEI

[ORPND01] L. D. M. Ottosen, N. Risgaard-Petersen, L. P. Nielsen, and T. Dalsgaard. Denitrification in exposed intertidal mud-flats, measured with a new^{15}N-ammonium spray technique. Marine Ecology Progress Series, 209:35–42, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/209/m209p035.pdf.

Olsen:2002:PDE

[ORS02] R. B. Olsen, K. Richardson, and V. Simonsen. Population differentiation of eelpout Zoarces viviparus in a Danish fjord. Marine Ecology Progress Series, 227:97–107, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p097.pdf.

Orvain:2005:MST

[Orv05] F. Orvain. A model of sediment transport under the influence of surface bioturbation: generalisation to the facultative suspension-feeder Scrobicularia plana. Marine Ecology Progress Series, 286:43–56, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/286/m286p043.pdf.

Olli:2001:VNS

[OS01] K. Olli and J. Seppälä. Vertical niche separation of phytoplankton: large-scale mesocosm experiments. Marine Ecology Progress Series, 217:219–233, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/217/m217p219.pdf.

Olesen:2005:ENP

[OSA05] M. Olesen, S. Strake, and A. Andrushaitis. Egestion of nonpellet-bound fecal material from the copepod Acartia tonsa: implication for vertical flux and degradation. Marine Ecology Progress Series, 293:131–142, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/293/m293p131.pdf.

Osborn:2007:ATM

[Osb07] T. Osborn. Applicability of turbulence measurement technology to small-scale plankton studies. Marine Ecology Progress Series, 347:139–143, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m347p139.pdf.

Oguz:2008:CPA

[OSF08] T. Oguz, B. Salihoglu, and B. Fach. A coupled planktonanchovy population dynamics model assessing nonlinear controls of anchovy and gelatinous biomass in the Black Sea. Marine Ecology Progress Series, 369:229–256, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/369/
m369p229.pdf.

Ouellet:2007:SOC

[OSL07] P. Ouellet, L. Savard, and P. Larouche. Spring oceanographic conditions and northern shrimp Pandalus borealis recruitment success in the north-western Gulf of St. Lawrence. Marine Ecology Progress Series, 339:229–241, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p229.pdf.

Olivar:2001:CSS

[OSP01] M. P. Olivar, J. Salat, and I. Palomera. Comparative study of spatial distribution patterns of the early stages of anchovy and pilchard in the NW Mediterranean Sea. Marine Ecology Progress Series, 217:111–120, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/217/m217p111.pdf.

Orvain:2004:IEH

[OSS^{+}04] F. Orvain, P. G. Sauriau, A. Sygut, L. Joassard, and P. Le Hir. Interacting effects of Hydrobia ulvae bioturbation and microphytobenthos on the erodibility of mudflat sediments. Marine Ecology Progress Series, 278:205–223, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/278/
m278p205.pdf.

Olsson:2007:FBL

[OST07] M. Olsson, L. Svärdh, and G. B. Toth. Feeding behaviour in Littorina littorea: the red seaweed Osmundea ramosissima may not prevent trematode infection. Marine Ecology Progress Series, 348:221–228, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/348/m348p221.pdf.

Ortiz:2008:OPH

[OT08] D. M. Ortiz and B. N. Tissot. Ontogenetic patterns of habitat use by reef-fish in a marine protected area network: a multiscaled remote sensing and in situ approach. Marine Ecology Progress Series, 365:217–232, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/365/m365p217.pdf.

Oltra:2000:LHF

[OTB^{+}00] R. Oltra, R. Todolí, T. Bosque, L. M. Lubián, and J. C. Navarro. Life history and fatty acid composition of the marine rotifer Synchaeta cecilia valentina fed different algae. Marine Ecology Progress Series, 193:125–133, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p125.pdf.

Olivier:2000:BSF

[OTBR00] F. Olivier, R. Tremblay, E. Bourget, and D. Rittschof. Barnacle settlement: field experiments on the influence of larval supply, tidal level, biofilm quality and age on Balanus amphitrite cyprids. Marine Ecology Progress Series, 199:185–204, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p185.pdf.

Olafsson:2005:CMB

[ÓUA05] E. Ólafsson, J. Ullberg, and N. L. Arroyo. The clam Macoma balthica prevents in situ growth of microalgal mats: implications for meiofaunal assemblages. Marine Ecology Progress Series, 298:179–188, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p179.pdf.

Orth:2002:HQP

[OvM02] R. J. Orth and J. van Montfrans. Habitat quality and prey size as determinants of survival in post-larval and early juvenile instars of the blue crab Callinectes sapidus. Marine Ecology Progress Series, 231:205–213, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/231/m231p205.pdf.

Ortiz:2002:ALA

[OW02] M. Ortiz and M. Wolff. Application of loop analysis to benthic systems in northern Chile for the elaboration of sustainable management strategies. Marine Ecology Progress Series, 242:15–27, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p015.pdf.

ODonnell:2007:SRT

[OWBD07] K. P. O’Donnell, R. A. Wahle, M. Bell, and M. Dunnington. Spatially referenced trap arrays detect sediment disposal impacts on lobsters and crabs in a New England estuary. Marine Ecology Progress Series, 348:249–260, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p249.pdf.

Okubo:2008:OCS

[OYNO08] N. Okubo, H. H. Yamamoto, F. Nakaya, and K. Okaji. Oxygen consumption of a single embryo/planula in the reef-building coral Acropora intermedia. Marine Ecology Progress Series, 366:305–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p305.pdf.

Okamura:2002:ADF

[OYT^{+}02] A. Okamura, A. Yamada, S. Tanaka, N. Horie, T. Utoh, N. Mikawa, A. Akazawa, and H. P. Oka. Atmospheric depression as the final trigger for the seaward migration of the Japanese eel Anguilla japonica. Marine Ecology Progress Series, 234:281–288, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p281.pdf.

Patil:2005:IDE

[PA05] J. S. Patil and A. C. Anil. Influence of diatom exopolymers and biofilms on metamorphosis in the barnacle Balanus amphitrite. Marine Ecology Progress Series, 301:231– 245, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/301/m301p231.pdf.

Polte:2006:ISB

[PA06] P. Polte and H. Asmus. Intertidal seagrass beds (Zostera noltii) as spawning grounds for transient fishes in the Wadden Sea. Marine Ecology Progress Series, 312:235–243, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/312/m312p235.pdf.

Power:2007:TDT

[PA07] M. Power and M. J. Attrill. Temperature-dependent temporal variation in the size and growth of Thames estuary smelt Osmerus eperlanus. Marine Ecology Progress Series, 330:213–222, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p213.pdf.

Pond:2002:OLC

[PAB^{+}02] D. W. Pond, C. E. Allen, M. V. Bell, C. L. Van Dover, A. E. Fallick, D. R. Dixon, and J. R. Sargent. Origins of long-chain polyunsaturated fatty acids in the hydrothermal vent worms Ridgea piscesae and Protis hydrothermica. Marine Ecology Progress Series, 225:219–226, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/225/m225p219.pdf.

Persson:2008:TRO

[PABM08] M. Persson, S. Andersson, S. Baden, and P. O. Moksnes. Trophic role of the omnivorous grass shrimp Palaemon elegans in a Swedish eelgrass system. Marine Ecology Progress Series, 371:203–212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/371/m371p203.pdf.

Pawlik:2000:TSM

[PAC^{+}00] Joseph R. Pawlik, M. Assmann, G. L. Cetrulo, R. de Nys, S. Engel, C. M. Finelli, M. E. Hay, M. Köck, P. J. Krug, E. Lichte, V. J. Paul, J. R. Pawlik, M. P. Puglisi, M. Slattery, P. D. Steinberg, J. T. Wright, and R. K. Zimmer. Theme section: Marine chemical ecology. Marine Ecology Progress Series, 207:225–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m207_TS.pdf.

Paffenhofer:2006:OCR

[Paf06] G. A. Paffenhöfer. Oxygen consumption in relation to motion of marine planktonic copepods. Marine Ecology Progress Series, 317:187–192, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p187.pdf.

Pahlow:2005:LCN

[Pah05] M. Pahlow. Linking chlorophyll-nutrient dynamics to the RedTeld N: C ratio with a model of optimal phytoplankton growth. Marine Ecology Progress Series, 287:33–43, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/287/m287p033.pdf.

Paul:2000:FIG

[PAK^{+}00] J. H. Paul, A. Alfreider, J. B. Kang, R. A. Stokes, D. Griffin, L. Campbell, and E. Ornolfsdottir. Form IA rbcL transcripts associated with a low salinity/high chlorophyll plume (’Green River’) in the eastern Gulf of Mexico. Marine Ecology Progress Series, 198:1–8, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/198/m198p001.pdf.

Peters:2009:PCD

[PAM^{+}09] K. J. Peters, C. D. Amsler, J. B. McClintock, R. W. M. van Soest, and B. J. Baker. Palatability and chemical defenses of sponges from the western Antarctic Peninsula. Marine Ecology Progress Series, 385:77–85, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m385p077.pdf.

Posey:2003:BCC

[PAMJ03] M. H. Posey, T. D. Alphin, D. L. Meyer, and J. M. Johnson. Benthic communities of common reed Phragmites australis and marsh cordgrass Spartina alterniflora marshes in Chesapeake Bay. Marine Ecology Progress Series, 261:51– 61, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/261/m261p051.pdf.

Parry:2006:FBT

[Par06] M. Parry. Feeding behavior of two ommastrephid squids Ommastrephes bartramii and Sthenoteuthis oualaniensis off Hawaii. Marine Ecology Progress Series, 318:229–235, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/318/m318p229.pdf.

Prado:2008:SRP

[PAR08] P. Prado, T. Alcoverro, and J. Romero. Seasonal response of Posidonia oceanica epiphyte assemblages to nutrient increase. Marine Ecology Progress Series, 359:89–98, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/359/
m359p089.pdf.

Passow:2000:FTE

[Pas00] U. Passow. Formation of transparent exopolymer particles, TEP, from dissolved precursor material. Marine Ecology Progress Series, 192:1–11, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/192/m192p001.pdf.

Passow:2002:PTE

[Pas02] U. Passow. Production of transparent exopolymer particles (TEP) by phyto- and bacterioplankton. Marine Ecology Progress Series, 236:1–12, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/236/m236p001.pdf.

Paul:2000:MMR

[PAW00a] J. H. Paul, A. Alfreider, and B. Wawrik. Micro- and macrodiversity in rbcL sequences in ambient phytoplankton populations from the southeastern Gulf of Mexico. Marine Ecology Progress Series, 198:9–18, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/198/m198p009.pdf.

Pawlik:2000:MCE

[Paw00b] J. R. Pawlik. Marine chemical ecology. Marine Ecology Progress Series, 207:225–226, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/207/m207p225.pdf.

Pavia:2000:EFI

[PB00a] H. Pavia and E. Brock. Extrinsic factors influencing phlorotannin production in the brown alga Ascophyllum nodosum. Marine Ecology Progress Series, 193:285–294, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p285.pdf.

Plaganyi:2000:DLP

[PB00b] E. E. Plagányi and G. M. Branch. Does the limpet Patella cochlear fertilize its own algal garden? Marine Ecology Progress Series, 194:113–122, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/194/m194p113.pdf.

Pane:2007:EAB

[PB07] E. F. Pane and J. P. Barry. Extracellular acid-base regulation during short-term hypercapnia is effective in a shallowwater crab, but ineffective in a deep-sea crab. Marine Ecology Progress Series, 334:1–9, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pandolfi:2008:MEZ

[PB08a] J. M. Pandolfi and A. F. Budd. Morphology and ecological zonation of Caribbean reef corals: the Montastraea annularis species complex. Marine Ecology Progress Series, 369: 89–102, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/369/m369p089.pdf.

Planque:2008:QRM

[PB08b] B. Planque and L. Buffaz. Quantile regression models for fish recruitment-environment relationships: four case studies. Marine Ecology Progress Series, 357:213–223, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p213.pdf.

Pasternak:2004:ISB

[PBAA04] Z. Pasternak, A. Bachar, A. Abelson, and Y. Achituv. Initiation of symbiosis between the soft coral Heteroxenia fuscescens and its zooxanthellae. Marine Ecology Progress Series, 279: 113–116, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p113.pdf.

Palsboll:2005:HLS

[PBAD05] P. J. Palsbøll, M. Bérubé, E. C. Anderson, and K. K. Dunham. High levels of statistical uncertainty in ‘gametic’ recapture estimates of male abundance in humpback whales. Marine Ecology Progress Series, 295:305–307, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p305.pdf.

Peterson:2002:NVP

[PBB^{+}02] Charles H. Peterson, B. E. Ballachey, J. L. Bodkin, T. D. Bowman, T. A. Dean, D. Esler, J. B. Fischer, A. K. Fukuyama, G. H. Golet, L. Holland-Bartels, D. B. Irons, S. C. Jewett, K. J. Kuletz, L. McDonald, D. A. McGuire, D. H. Monson, S. H. Newman, C. E. O’Clair, C. H. Peterson, D. D. Roby, P. E. Seiser, K. A. Trust, and G. R. VanBlaricom. Nearshore vertebrate predators: constraints to recovery from oil pollution. Marine Ecology Progress Series, 241:235–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m241_TS.pdf.

Phlips:2006:FAD

[PBBC06] J. E. Phlips, S. Badylak, E. Bledsoe, and M. Cichra. Factors affecting the distribution of Pyrodinium bahamense var. bahamense in coastal waters of Florida. Marine Ecology Progress Series, 322:99–115, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p099.pdf.

Peck:2003:EFC

[PBCB03] M. A. Peck, L. J. Buckley, E. M. Caldarone, and D. A. Bengtson. Effects of food consumption and temperature on growth rate and biochemical-based indicators of growth in early juvenile Atlantic cod Gadus morhua and haddock Melanogrammus aeglefinus. Marine Ecology Progress Series, 251:233–243, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p233.pdf.

Pillay:2007:ECK

[PBF07] D. Pillay, G. M. Branch, and A. T. Forbes. Effects of Callianassa kraussi on microbial biofilms and recruitment of macrofauna: a novel hypothesis for adult-juvenile interactions. Marine Ecology Progress Series, 347:1–14, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pillay:2008:HCE

[PBF08] D. Pillay, G. M. Branch, and A. T. Forbes. Habitat change in an estuarine embayment: anthropogenic influences and a regime shift in biotic interactions. Marine Ecology Progress Series, 370:19–31, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/370/m370p019.pdf.

Pederson:2008:EPP

[PBFB08] H. G. Pederson, N. S. Barrett, S. D. Frusher, and C. D. Buxton. Effect of predator–prey and competitive interactions on size at emergence in the black-lip abalone Haliotis rubra in a Tasmanian MPA. Marine Ecology Progress Series, 366:91–98, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p091.pdf.

Parsons:2003:SPA

[PBH^{+}03] D. M. Parsons, R. C. Babcock, R. K. S. Hankin, T. J. Willis, J. P. Aitken, R. K. O’Dor, and G. D. Jackson. Snapper Pagrus auratus (Sparidae) home range dynamics: acoustic tagging studies in a marine reserve. Marine Ecology Progress Series, 262:253–265, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p253.pdf.

Philipp:2006:PAT

[PBH^{+}06] E. Philipp, T. Brey, O. Heilmayer, D. Abele, and H. O. Pörtner. Physiological ageing in a temperate and a polar swimming scallop. Marine Ecology Progress Series, 307:187– 198, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p187.pdf.

Ponton:2005:EPM

[PBJ^{+}05] F. Ponton, D. G. Biron, C. Joly, S. Helluy, D. Duneau, and F. Thomas. Ecology of parasitically modified populations: a case study from a gammarid-trematode system. Marine Ecology Progress Series, 299:205–215, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p205.pdf.

Purser:2009:UML

[PBL^{+}09] A. Purser, M. Bergmann, T. Lundälv, J. Ontrup, and T. W. Nattkemper. Use of machine-learning algorithms for the automated detection of cold-water coral habitats: a pilot study. Marine Ecology Progress Series, 397:241–251, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p241.pdf.

Parrish:2007:BBP

[PBN^{+}07] J. K. Parrish, N. Bond, H. Nevins, N. Mantua, R. Loeffel, W. T. Peterson, and J. T. Harvey. Beached birds and physical forcing in the California Current system. Marine Ecology Progress Series, 352:275–288, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m352p275.pdf.

Peralta:2006:DEC

[PBPLB06] G. Peralta, F. G. Brun, J. L. Pérez-Lloréns, and T. J. Bouma. Direct effects of current velocity on the growth, morphometry and architecture of seagrasses: a case study on Zostera noltii. Marine Ecology Progress Series, 327:135–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p135.pdf.

Purcell:2000:AJA

[PBS^{+}00] J. E. Purcell, E. D. Brown, K. D. E. Stokesbury, L. H. Haldorson, and T. C. Shirley. Aggregations of the jellyfish Aurelia labiata: abundance, distribution, association with age-0 walleye pollock, and behaviors promoting aggregation in Prince William Sound, Alaska, USA. Marine Ecology Progress Series, 195:145–158, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/195/m195p145.pdf.

Petersen:2004:IMM

[PBS^{+}04] J. K. Petersen, S. Bougrier, A. C. Smaal, P. Garen, S. Robert, J. E. N. Larsen, and E. Brummelhuis. Intercalibration of mussel Mytilus edulis clearance rate measurements. Marine Ecology Progress Series, 267:187–194, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p187.pdf.

Powell:2005:ENS

[PBS^{+}05] L. M. Powell, J. P. Bowman, J. H. Skerratt, P. D. Franzmann, and H. R. Burton. Ecology of a novel Synechococcus clade occurring in dense populations in saline Antarctic lakes.

Marine Ecology Progress Series, 291:65–80, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/291/
m291p065.pdf.

Pelc:2009:QLE

[PBT^{+}09] R. A. Pelc, M. L. Baskett, T. Tanci, S. D. Gaines, and R. R. Warner. Quantifying larval export from South African marine reserves. Marine Ecology Progress Series, 394:65– 78, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/394/m394p065.pdf.

Piggott:2008:GED

[PBTB08] M. P. Piggott, S. C. Banks, P. Tung, and L. B. Beheregaray. Genetic evidence for different scales of connectivity in a marine mollusc. Marine Ecology Progress Series, 365:127–136, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/365/m365p127.pdf.

Paul:2009:DEI

[PBVP09] C. Paul, A. Barofsky, C. Vidoudez, and G. Pohnert. Diatom exudates influence metabolism and cell growth of cocultured diatom species. Marine Ecology Progress Series, 389: 61–70, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/389/m389p061.pdf.

Prast:2007:ECN

[PBW^{+}07] M. Prast, A. A. Bischoff, U. Waller, R. Amann, and U. G. Berninger. Effect of ciliates on nitrification and nitrifying bacteria in Baltic Sea sediments. Marine Ecology Progress Series, 350:55–61, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/350/m350p055.pdf.

Park:2002:AHN

[PC02a] J. S. Park and B. C. Cho. Active heterotrophic nanoflagellates in the hypoxic water-column of the eutrophic Masan Bay, Korea. Marine Ecology Progress Series, 230:35–45, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/230/m230p035.pdf.

Piazzi:2002:ECB

[PC02b] L. Piazzi and G. Ceccherelli. Effects of competition between two introduced Caulerpa. Marine Ecology Progress Series, 225:189–195, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p189.pdf.

Peck:2005:CSF

[PC05] D. R. Peck and B. C. Congdon. Colony-specific foraging behaviour and co-ordinated divergence of chick development in the wedge-tailed shearwater Puffinus pacificus. Marine Ecology Progress Series, 299:289–296, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p289.pdf.

Prescott:2008:MRN

[PCB08] R. Prescott and R. Cudney-Bueno. Mobile ‘reefs’ in the northeastern Gulf of California: aggregations of black murex snails Hexaplex nigritus as habitat for invertebrates. Marine Ecology Progress Series, 367:185–192, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p185.pdf.

Pospelova:2005:SDM

[PCBL05] V. Pospelova, G. L. Chmura, W. S. Boothman, and J. S. Latimer. Spatial distribution of modern dinoflagellate cysts in polluted estuarine sediments from Buzzards Bay (Massachusetts, USA) embayments. Marine Ecology Progress Series, 292:23–40, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/292/m292p023.pdf.

Pikitch:2005:HUD

[PCBS05] E. K. Pikitch, D. D. Chapman, E. A. Babcock, and M. S. Shivji. Habitat use and demographic population structure of elasmobranchs at a Caribbean atoll (glover’s Reef, Belize). Marine Ecology Progress Series, 302:187–197, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p187.pdf.

Piazzi:2001:TMD

[PCC01] L. Piazzi, G. Ceccherelli, and F. Cinelli. Threat to macroalgal diversity: effects of the introduced green alga Caulerpa racemosa in the Mediterranean. Marine Ecology Progress Series, 210:149–159, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p149.pdf.

Peck:2006:MDP

[PCC06] L. S. Peck, A. Clarke, and A. L. Chapman. Metabolism and development of pelagic larvae of Antarctic gastropods with mixed reproductive strategies. Marine Ecology Progress Series, 318:213–220, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/318/m318p213.pdf.

Paris:2007:SSD

[PCC07] C. B. Paris, L. M. Chérubin, and R. K. Cowen. Surfing, spinning, or diving from reef to reef: effects on population connectivity. Marine Ecology Progress Series, 347:285–300, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p285.pdf.

Paris:2005:LTP

[PCCL05] C. B. Paris, R. K. Cowen, R. Claro, and K. C. Lindeman. Larval transport pathways from Cuban snapper (Lutjanidae) spawning aggregations based on biophysical modeling. Marine Ecology Progress Series, 296:93–106, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/296/
m296p093.pdf.

Perez-Castaneda:2004:SDS

[PCD04] R. Pérez-Castañeda and O. Defeo. Spatial distribution and structure along ecological gradients: penaeid shrimps in a tropical estuarine habitat of Mexico. Marine Ecology Progress Series, 273:173–185, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/273/m273p173.pdf.

Porter:2005:EFI

[PCH^{+}05] S. M. Porter, L. Ciannelli, N. Hillgruber, K. M. Bailey, K. S. Chan, M. F. Canino, and L. J. Haldorson. Environmental factors influencing larval walleye pollock Theragra chalcogramma feeding in Alaskan waters. Marine Ecology Progress Series, 302:207–217, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p207.pdf.

Pittman:2007:UST

[PCHM07] S. J. Pittman, C. Caldow, S. D. Hile, and M. E. Monaco. Using seascape types to explain the spatial patterns of fish in the mangroves of SW Puerto Rico. Marine Ecology Progress Series, 348:273–284, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/348/m348p273.pdf.

Pemberton:2006:QUA

[PCJ06] K. L. Pemberton, K. R. Clarke, and I. Joint. Quantifying uncertainties associated with the measurement of primary production. Marine Ecology Progress Series, 322:51– 59, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p051.pdf.

Peterson:2006:PRS

[PCJF06] B. J. Peterson, C. M. Chester, F. J. Jochem, and J. W. Fourqurean. Potential role of sponge communities in controlling phytoplankton blooms in Florida Bay. Marine Ecology Progress Series, 328:93–103, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/328/m328p093.pdf.

Prado:2008:TWA

[PCL08] P. Prado, C. J. Collier, and P. S. Lavery. ^{13}C and ^{15}N translocation within and among shoots in two Posidonia species from Western Australia. Marine Ecology Progress Series, 361: 69–82, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p069.pdf.

Pansch:2009:CLR

[PCL^{+}09] C. Pansch, O. Cerda, M. Lenz, M. Wahl, and M. Thiel. Consequences of light reduction for anti-herbivore defense and bioactivity against mussels in four seaweed species from northern-central Chile. Marine Ecology Progress Series, 381: 83–97, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/381/m381p083.pdf.

Pears:2006:DLG

[PCMB06] R. J. Pears, J. H. Choat, B. D. Mapstone, and G. A. Begg. Demography of a large grouper, Epinephelus fuscoguttatus, from Australia’s Great Barrier Reef: implications for fishery management. Marine Ecology Progress Series, 307:259– 272, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p259.pdf.

Pusineri:2008:FNS

[PCRR08] C. Pusineri, O. Chancollon, J. Ringelstein, and V. Ridoux. Feeding niche segregation among the Northeast Atlantic community of oceanic top predators. Marine Ecology Progress Series, 361:21–34, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p021.pdf.

Pardal:2004:AEQ

[PCS^{+}04a] M. A. Pardal, P. G. Cardoso, J. P. Sousa, J. C. Marques, and D. Raffaelli. Assessing environmental quality: a novel approach. Marine Ecology Progress Series, 267:1–8, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/267/m267p001.pdf.

Porter:2004:EOC

[PCS04b] E. T. Porter, J. C. Cornwell, and L. P. Sanford. Effect of oysters Crassostrea virginica and bottom shear velocity on benthic-pelagic coupling and estuarine water quality: Erratum. Marine Ecology Progress Series, 271:61–75, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/271/m271p061.pdf. See erratum [Ano04e].

Phillips:2007:MWD

[PCS^{+}07] R. A. Phillips, P. Catry, J. R. D. Silk, S. Bearhop, R. McGill, V. Afanasyev, and I. J. Strange. Movements, winter distribution and activity patterns of Falkland and brown skuas: insights from loggers and isotopes. Marine Ecology Progress Series, 345:281–291, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/345/m345p281.pdf.

Pires:2009:RBD

[PCS09] D. O. Pires, C. B. Castro, and J. C. Silva. Reproductive biology of the deep-sea pennatulacean Anthoptilum murrayi (Cnidaria, Octocorallia). Marine Ecology Progress Series, 397:103–112, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p103.pdf.

Pagola-Carte:2001:CSF

[PCSS01] S. Pagola-Carte and J. I. Saiz-Salinas. Changes in the sublittoral faunal biomass induced by the discharge of a polluted river along the adjacent rocky coast (N. Spain). Marine Ecology Progress Series, 212:13–27, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p013.pdf.

Pereira:2000:ESM

[PCT00] R. C. Pereira, D. N. Cavalcanti, and V. L. Teixeira. Effects of secondary metabolites from the tropical Brazilian brown alga Dictyota menstrualis on the amphipod Parhyale hawaiensis. Marine Ecology Progress Series, 205:95–100, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p095.pdf.

Pinaud:2005:EEV

[PCW05] D. Pinaud, Y. Cherel, and H. Weimerskirch. Effect of environmental variability on habitat selection, diet, provisioning behaviour and chick growth in yellow-nosed albatrosses. Marine Ecology Progress Series, 298:295–304, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p295.pdf.

Park:2002:ETS

[PCYC02] M. G. Park, S. K. Cooney, W. Yih, and D. W. Coats. Effects of two strains of the parasitic dinoflagellate Amoebophrya on growth, photosynthesis, light absorption, and quantum yield of bloom-forming dinoflagellates. Marine Ecology Progress Series, 227:281–292, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p281.pdf.

Poggiale:2001:LTD

[PD01] J. C. Poggiale and J. C. Dauvin. Long-term dynamics of three benthic Ampelisca (Crustacea–Amphipoda) populations from the Bay of Morlaix (western English Channel) related to their disappearance after the ‘Amoco Cadiz’ oil spill. Marine Ecology Progress Series, 214:201–209, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/214/m214p201.pdf.

Proffitt:2005:GIG

[PD05] C. E. Proffitt and D. J. Devlin. Grazing by the intertidal gastropod Melampus coffeus greatly increases mangrove leaf litter degradation rates. Marine Ecology Progress Series, 296:209–218, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p209.pdf.

Peck:2007:PBL

[PD07] M. A. Peck and U. Daewel. Physiologically based limits to food consumption, and individual-based modeling of foraging and growth of larval fishes. Marine Ecology Progress Series, 347:171–183, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m347p171.pdf.

Przeslawski:2004:EUR

[PDB04] R. Przeslawski, A. R. Davis, and K. Benkendorff. Effects of ultraviolet radiation and visible light on the development of encapsulated molluscan embryos. Marine Ecology Progress Series, 268:151–160, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/268/m268p151.pdf.

Page:2006:EIS

[PDCH06] H. M. Page, J. E. Dugan, C. S. Culver, and J. C. Hoesterey. Exotic invertebrate species on offshore oil platforms. Marine Ecology Progress Series, 325:101–107, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p101.pdf.

Pusceddu:2009:QBS

[PDFD09] A. Pusceddu, A. Dell’Anno, M. Fabiano, and R. Danovaro. Quantity and bioavailability of sediment organic matter as signatures of benthic trophic status. Marine Ecology Progress Series, 375:41–52, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p041.pdf.

Post:2002:STD

[PDG^{+}02] A. F. Post, Z. Dedej, R. Gottlieb, H. Li, D. N. Thomas, M. El-Absawi, A. El-Naggar, M. El-Gharabawi, and U. Sommer. Spatial and temporal distribution of Trichodesmium spp. in the stratified Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 239:241–250, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/239/m239p241.pdf.

Perez-Dominguez:2006:EVS

[PDHH06] R. Perez-Dominguez, S. A. Holt, and G. J. Holt. Environmental variability in seagrass meadows: effects of nursery environment cycles on growth and survival in larval red drum Sciaenops ocellatus. Marine Ecology Progress Series, 321: 41–53, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p041.pdf.

Paz:2003:GSB

[PDM^{+}03] G. Paz, J. Douek, C. Mo, M. Goren, and B. Rinkevich. Genetic structure of Botryllus schlosseri (Tunicata) populations from the Mediterranean coast of Israel. Marine Ecology Progress Series, 250:153–162, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/250/m250p153.pdf.

Panfili:2004:ISL

[PDM^{+}04] J. Panfili, J. D. Durand, A. Mbow, B. Guinand, K. Diop, J. Kantoussan, D. Thior, O. T. Thiaw, J. J. Albaret, and R. Laë. Influence of salinity on life history traits of the bonga shad Ethmalosa fimbriata (Pisces, Clupeidae): comparison between the Gambia and Saloum estuaries. Marine Ecology Progress Series, 270:241–257, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/270/m270p241.pdf.

Paul:2006:CDA

[PdNS06a] N. A. Paul, R. de Nys, and P. D. Steinberg. Chemical defence against bacteria in the red alga Asparagopsis armata: linking structure with function. Marine Ecology Progress Series, 306:87–101, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p087.pdf.

Paul:2006:SHI

[PdNS06b] N. A. Paul, R. de Nys, and P. D. Steinberg. Seaweed-herbivore interactions at a small scale: direct tests of feeding deterrence by filamentous algae. Marine Ecology Progress Series, 323: 1–9, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Parker:2001:PSD

[PDO01] J. D. Parker, J. E. Duffy, and R. J. Orth. Plant species diversity and composition: experimental effects on marine epifaunal assemblages. Marine Ecology Progress Series, 224: 55–67, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/224/m224p055.pdf.

Paula:2001:IOI

[PDQ01] J. Paula, T. Dray, and H. Queiroga. Interaction of offshore and inshore processes controlling settlement of brachyuran megalopae in Saco mangrove creek, Inhaca Island (South Mozambique). Marine Ecology Progress Series, 215:251–260, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p251.pdf.

Plourde:2002:STV

[PDRT02] S. Plourde, J. J. Dodson, J. A. Runge, and J. C. Therriault. Spatial and temporal variations in copepod community structure in the lower St. Lawrence Estuary, Canada. Marine Ecology Progress Series, 230:211–224, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/230/
m230p211.pdf.

Page:2007:TLC

[PDS^{+}07] H. M. Page, J. E. Dugan, D. M. Schroeder, M. M. Nishimoto, M. S. Love, and J. C. Hoesterey. Trophic links and condition of a temperate reef fish: comparisons among offshore oil platform and natural reef habitats. Marine Ecology Progress Series, 344:245–256, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/344/m344p245.pdf.

Parrish:2009:ESS

[PDT09] C. C. Parrish, D. Deibel, and R. J. Thompson. Effect of sinking spring phytoplankton blooms on lipid content and composition in suprabenthic and benthic invertebrates in a cold ocean coastal environment. Marine Ecology Progress Series, 391:33–51, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/391/m391p033.pdf.

Parsons:2005:IER

[PE05] D. M. Parsons and D. B. Eggleston. Indirect effects of recreational fishing on behavior of the spiny lobster Panulirus argus. Marine Ecology Progress Series, 303:235–244, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/303/m303p235.pdf.

Polikarpov:2009:RBS

[PE09] G. G. Polikarpov and V. N. Egorov. Response of the Black Sea ecosystem to the Chernobyl nuclear accident. Marine Ecology Progress Series, 376:307–308, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/376/m376p307.pdf.

Pearre:2000:LTC

[Pea00] S. Pearre, Jr. Long-term changes in diel vertical migration behavior: more ups and downs. Marine Ecology Progress Series, 197:305–307, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p305.pdf.

Pearre:2003:WLL

[Pea03] S. Pearre, Jr. When lights are low: Comment on ’Midnight sinking behaviour in Calanus finmarchicus: a response to satiation or krill predation?’ by Tarling et al. (2002). Marine Ecology Progress Series, 252:303–305, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/

m252p303.pdf. See [TJEM02].

Porte:2002:ACP

[PEdlP^{+}02] C. Porte, E. Escartín, L. M. García de la Parra, X. Biosca, and J. Albaigés. Assessment of coastal pollution by combined determination of chemical and biochemical markers in Mullus barbatus. Marine Ecology Progress Series, 235:205– 216, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/235/m235p205.pdf.

Perovich:2003:STP

[PEDT03] G. M. Perovich, C. E. Epifanio, A. I. Dittel, and P. A. Tyler. Spatial and temporal patterns in development of eggs in the vent crab Bythograea thermydron. Marine Ecology Progress Series, 251:211–220, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/251/m251p211.pdf.

Peebles:2002:TRB

[Pee02] E. B. Peebles. Temporal resolution of biological and physical influences on bay anchovy Anchoa mitchilli egg abundance near a river-plume frontal zone. Marine Ecology Progress Series, 237:257–269, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p257.pdf.

Porte:2000:XME

[PEG^{+}00] C. Porte, E. Escartín, L. M. García, M. Solé, and J. Albaigés. Xenobiotic metabolising enzymes and antioxidant defences in deep-sea fish: relationship with contaminant body burden. Marine Ecology Progress Series, 192:259–266, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p259.pdf.

Pizarro:2008:GBC

[PEGG^{+}08] G. Pizarro, L. Escalera, S. González-Gil, J. M. Franco, and B. Reguera. Growth, behaviour and cell toxin quota of Dinophysis acuta during a daily cycle. Marine Ecology Progress Series, 353:89–105, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/353/m353p089.pdf.

Penry:2000:DKS

[Pen00] D. L. Penry. Digestive kinematics of suspension-feeding bivalves: modeling and measuring particle-processing in the gut of Potamocorbula amurensis. Marine Ecology Progress Series, 197:181–192, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p181.pdf.

Peperzak:2006:MVG

[Pep06] L. Peperzak. Modelling vegetative growth, gamete production and encystment of dinoflagellates in batch culture. Marine Ecology Progress Series, 306:143–152, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p143.pdf.

Peters:2002:OLE

[Pet02a] F. Peters. Overcoming linearisation errors in calculating bacterial growth rates. Marine Ecology Progress Series, 245: 305–308, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p305.pdf. See reply [Kir02b,

Kir02c].

Petraitis:2002:EIC

[Pet02b] P. S. Petraitis. Effects of intraspecific competition and scavenging on growth of the periwinkle Littorina littorea. Marine Ecology Progress Series, 236:179–187, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p179.pdf.

Petersen:2004:MMB

[Pet04] J. K. Petersen. Methods for measurement of bivalve clearance rate-hope for common understanding. Marine Ecology Progress Series, 276:309–310, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p309.pdf.

Puyana:2003:TAC

[PFP03] M. Puyana, W. Fenical, and J. R. Pawlik. Are there activated chemical defenses in sponges of the genus Aplysina from the Caribbean? Marine Ecology Progress Series, 246: 127–135, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p127.pdf.

Prado:2008:MPM

[PFT^{+}08] P. Prado, S. Farina, F. Tomas, J. Romero, and T. Alcoverro. Marine protection and meadow size alter fish herbivory in seagrass ecosystems. Marine Ecology Progress Series, 371:11–21, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/371/m371p011.pdf.

Pakhomov:2002:CAS

[PFW^{+}02] E. A. Pakhomov, P. W. Froneman, P. Wassmann, T. Ratkova, and E. Arashkevich. Contribution of algal sinking and zooplankton grazing to downward flux in the Lazarev Sea (Southern Ocean) during the onset of phytoplankton bloom: a Lagrangian study. Marine Ecology Progress Series, 233:73– 88, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p073.pdf.

Perkol-Finkel:2006:FFA

[PFZS^{+}06] S. Perkol-Finkel, G. Zilman, I. Sella, T. Miloh, and Y. Benayahu. Floating and fixed artificial habitats: effects of substratum motion on benthic communities in a coral reef environment. Marine Ecology Progress Series, 317:9–20, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/317/m317p009.pdf.

Pitta:2000:PCO

[PG00] P. Pitta and A. Giannakourou. Planktonic ciliates in the oligotrophic Eastern Mediterranean: vertical, spatial distribution and mixotrophy. Marine Ecology Progress Series, 194:269–282, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p269.pdf.

Patti:2001:PIP

[PG01] F. P. Patti and M. C. Gambi. Phylogeography of the invasive polychaete Sabella spallanzanii (Sabellidae) based on the nucleotide sequence of internal transcribed spacer 2 (ITS2) of nuclear rDNA. Marine Ecology Progress Series, 215:169– 177, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/215/m215p169.pdf.

Paschke:2004:SVS

[PGBA04] K. A. Paschke, P. Gebauer, F. Buchholz, and K. Anger. Seasonal variation in starvation resistance of early larval North Sea shrimp Crangon crangon (Decapoda: Crangonidae). Marine Ecology Progress Series, 279:183–191, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p183.pdf.

Pallas:2006:DHU

[PGCC^{+}06] A. Pallas, B. Garcia-Calvo, A. Corgos, C. Bernardez, and J. Freire. Distribution and habitat use patterns of benthic decapod crustaceans in shallow waters: a comparative approach. Marine Ecology Progress Series, 324:173–184, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p173.pdf.

Penfold:2008:IBJ

[PGK^{+}08] R. Penfold, A. S. Grutter, A. M. Kuris, M. I. McCormick, and C. M. Jones. Interactions between juvenile marine fish and gnathiid isopods: predation versus micropredation. Marine Ecology Progress Series, 357:111–119, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p111.pdf.

Palardy:2005:EUD

[PGM05] J. E. Palardy, A. G. Grottoli, and K. A. Matthews. Effects of upwelling, depth, morphology and polyp size on feeding in three species of Panamanian corals. Marine Ecology Progress Series, 300:79–89, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/300/m300p079.pdf.

Peterson:2003:EEF

[PGP03] C. H. Peterson, J. H. Grabowski, and S. P. Powers. Estimated enhancement of fish production resulting from restoring oyster reef habitat: quantitative valuation. Marine Ecology Progress Series, 264:249–264, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/264/m264p249.pdf.

Powers:2003:EEF

[PGPL03] S. P. Powers, J. H. Grabowski, C. H. Peterson, and W. J. Lindberg. Estimating enhancement of fish production by offshore artificial reefs: uncertainty exhibited by divergent scenarios. Marine Ecology Progress Series, 264:265–277, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p265.pdf.

Pages:2001:GZA

[PGR^{+}01] F. Pagès, H. E. González, M. Ramón, M. Sobarzo, and J. M. Gili. Gelatinous zooplankton assemblages associated with water masses in the Humboldt Current System, and potential predatory impact by Bassia bassensis (Siphonophora: Calycophorae). Marine Ecology Progress Series, 210:13–24, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
210/m210p013.pdf.

Pond:2000:UFA

[PGS^{+}00] D. W. Pond, A. Gebruk, E. C. Southward, A. J. Southward, A. E. Fallick, M. V. Bell, and J. R. Sargent. Unusual fatty acid composition of storage lipids in the bresilioid shrimp Rimicaris exoculata couples the photic zone with MAR hydrothermal vent sites. Marine Ecology Progress Series, 198:171–179, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p171.pdf.

Peterson:2001:ETM

[PH01a] B. J. Peterson and K. L. Heck, Jr. An experimental test of the mechanism by which suspension feeding bivalves elevate seagrass productivity. Marine Ecology Progress Series, 218:115–125, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/218/m218p115.pdf.

Peterson:2001:PIB

[PH01b] B. J. Peterson and K. L. Heck, Jr. Positive interactions between suspension-feeding bivalves and seagrass — a facultative mutualism. Marine Ecology Progress Series, 213: 143–155, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/213/m213p143.pdf.

Pedersen:2003:EHPa

[PH03a] F. M. Pedersen and P. J. Hansen. Effects of high pH on a natural marine planktonic community. Marine Ecology Progress Series, 260:19–31, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p019.pdf.

Pedersen:2003:EHPb

[PH03b] F. M. Pedersen and P. J. Hansen. Effects of high pH on the growth and survival of six marine heterotrophic protists. Marine Ecology Progress Series, 260:33–41, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p033.pdf.

Phillips:2003:NEI

[PH03c] J. C. Phillips and C. L. Hurd. Nitrogen ecophysiology of intertidal seaweeds from New Zealand: N uptake, storage and utilisation in relation to shore position and season. Marine Ecology Progress Series, 264:31–48, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p031.pdf.

Paramor:2005:EII

[PH05] O. A. L. Paramor and R. G. Hughes. Effects of the invertebrate infauna on early saltmarsh plant colonisation of managed realignment areas in south-east England. Marine Ecology Progress Series, 303:61–71, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/303/m303p061.pdf.

Peterson:2002:CIO

[PHB02] C. H. Peterson and L. Holland-Bartels. Chronic impacts of oil pollution in the sea: risks to vertebrate predators. Marine Ecology Progress Series, 241:235–236, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p235.pdf.

Pomeroy:2005:MNA

[PHH^{+}05] P. P. Pomeroy, J. A. Hammond, A. J. Hall, M. Lonergan, C. D. Duck, V. J. Smith, and H. Thompson. Morbillivirus neutralising antibodies in Scottish grey seals Halichoerus grypus: assessing the effects of the 1988 and 2002 PDV epizootics. Marine Ecology Progress Series, 287:241–250, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p241.pdf.

Phillips:2005:GFF

[Phi05] N. E. Phillips. Growth of filter-feeding benthic invertebrates from a region with variable upwelling intensity. Marine Ecology Progress Series, 295:79–89, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p079.pdf.

Pawlik:2008:PSR

[PHM^{+}08] J. R. Pawlik, T. P. Henkel, S. E. McMurray, S. López-Legentil, T. L. Loh, and S. Rohde. Patterns of sponge recruitment and growth on a shipwreck corroborate chemical defense resource trade-off. Marine Ecology Progress Series, 368:137–143, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p137.pdf.

Post:2007:HBC

[PHO^{+}07] A. L. Post, M. A. Hemer, P. E. O’Brien, D. Roberts, and M. Craven. History of benthic colonisation beneath the Amery Ice Shelf, East Antarctica. Marine Ecology Progress Series, 344:29–37, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/344/m344p029.pdf.

Petes:2003:PCR

[PHP^{+}03a] L. E. Petes, C. D. Harvell, E. C. Peters, M. A. H. Webb, and K. M. Mullen. Pathogens compromise reproduction and induce melanization in Caribbean sea fans. Marine Ecology Progress Series, 264:167–171, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p167.pdf.

Petes:2003:RSM

[PHP^{+}03b] L. E. Petes, C. D. Harvell, E. C. Peters, M. A. H. Webb, and K. M. Mullen. Restoration scaling in the marine environment. Marine Ecology Progress Series, 264:173–307, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m264_TS.pdf.

Piatt:2007:SIMb

[PHS^{+}07] J. F. Piatt, A. M. A. Harding, M. Shultz, S. G. Speckman, T. I. van Pelt, G. S. Drew, and A. B. Kettle. Seabirds as indicators of marine food supplies: Cairns revisited. Marine Ecology Progress Series, 352:221–234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m352p221.pdf.

Purcell:2009:IVS

[PHS09] J. E. Purcell, R. A. Hoover, and N. T. Schwarck. Interannual variation of strobilation by the scyphozoan Aurelia labiata in relation to polyp density, temperature, salinity, and light conditions in situ. Marine Ecology Progress Series, 375:139–149, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m375p139.pdf.

Putland:2007:EAT

[PI07a] J. N. Putland and R. L. Iverson. Ecology of Acartia tonsa in Apalachicola Bay, Florida, and implications of river water diversion. Marine Ecology Progress Series, 340:173–187, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/340/m340p173.pdf.

Putland:2007:MMH

[PI07b] J. N. Putland and R. L. Iverson. Microzooplankton: major herbivores in an estuarine planktonic food web. Marine Ecology Progress Series, 345:63–73, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p063.pdf.

Poulsen:2008:DCF

[PI08] L. K. Poulsen and M. H. Iversen. Degradation of copepod fecal pellets: key role of protozooplankton. Marine Ecology Progress Series, 367:1–13, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Paffenhofer:2005:CDC

[PIM^{+}05] G. A. Paffenhöfer, A. Ianora, A. Miralto, J. T. Turner, G. S. Kleppel, M. R. d’Alcalà, R. Casotti, G. S. Caldwell, G. Pohnert, A. Fontana, D. Müller-Navarra, S. Jónasdóttir, V. Armbrust, U. Båmstedt, S. Ban, M. G. Bentley, M. Boersma, M. Bundy, I. Buttino, A. Calbet, F. Carlotti, Y. Carotenuto, G. d’Ippolito, B. Frost, C. Guisande, W. Lampert, R. F. Lee, S. Mazza, M. G. Mazzocchi, J. C. Nejstgaard, S. A. Poulet, G. Romano, V. Smetacek, S. Uye, S. Wakeham, S. Watson, and T. Wichard. Colloquium on diatom-copepod interactions. Marine Ecology Progress Series, 286:293–305, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p293.pdf.

Putz:2002:WDR

[PISL02] K. Pütz, R. J. Ingham, J. G. Smith, and B. H. Lüthi. Winter dispersal of rockhopper penguins Eudyptes chrysocome from the Falkland Islands and its implications for conservation. Marine Ecology Progress Series, 240:273–284, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p273.pdf.

Padmavati:2004:LCP

[PIY04] G. Padmavati, T. Ikeda, and A. Yamaguchi. Life cycle, population structure and vertical distribution of Metridia spp. (Copepoda: Calanoida) in the Oyashio region (NW Pacific Ocean). Marine Ecology Progress Series, 270:181– 198, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/270/m270p181.pdf.

Pardo:2006:IWM

[PJ06a] L. M. Pardo and L. E. Johnson. Influence of water motion and reproductive attributes on movement and shelter use in the marine snail Littorina saxatilis. Marine Ecology Progress Series, 315:177–186, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/315/m315p177.pdf.

Piola:2006:DRE

[PJ06b] R. F. Piola and E. L. Johnston. Differential resistance to extended copper exposure in four introduced bryozoans. Marine Ecology Progress Series, 311:103–114, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/311/
m311p103.pdf.

Plourde:2008:CLS

[PJ08] S. Plourde and P. Joly. Comparison of in situ egg production rate in Calanus finmarchicus and Metridia longa: discriminating between methodological and species-specific effects. Marine Ecology Progress Series, 353:165–175, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p165.pdf.

Pitchford:2003:OFP

[PJB03] J. W. Pitchford, A. James, and J. Brindley. Optimal foraging in patchy turbulent environments. Marine Ecology Progress Series, 256:99–110, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p099.pdf.

Petrone:2005:DLP

[PJJ^{+}05] C. Petrone, L. B. Jancaitis, M. B. Jones, C. C. Natunewicz, C. E. Tilburg, and C. E. Epifanio. Dynamics of larval patches: spatial distribution of fiddler crab larvae in Delaware Bay and adjacent waters. Marine Ecology Progress Series, 293: 177–190, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p177.pdf.

Phillips:2001:PMS

[PJN01] K. L. Phillips, G. D. Jackson, and P. D. Nichols. Predation on myctophids by the squid Moroteuthis ingens around Macquarie and Heard Islands: stomach contents and fatty acid analyses. Marine Ecology Progress Series, 215:179–189, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p179.pdf.

Phillips:2003:TVD

[PJN03] K. L. Phillips, G. D. Jackson, and P. D. Nichols. Temporal variations in the diet of the squid Moroteuthis ingens at Macquarie Island: stomach contents and fatty acid analyses. Marine Ecology Progress Series, 256:135–149, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/256/
m256p135.pdf.

Perrett:2006:IAD

[PJP06] L. A. Perrett, E. L. Johnston, and A. G. B. Poore. Impact by association: direct and indirect effects of copper exposure on mobile invertebrate fauna. Marine Ecology Progress Series, 326:195–205, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p195.pdf.

Plourde:2003:LCC

[PJR^{+}03] S. Plourde, P. Joly, J. A. Runge, J. Dodson, and B. Zakardjian. Life cycle of Calanus hyperboreus in the lower St. Lawrence Estuary and its relationship to local environmental conditions. Marine Ecology Progress Series, 255:219–233, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p219.pdf.

Pitt:2000:GSS

[PK00] K. A. Pitt and M. J. Kingsford. Geographic separation of stocks of the edible jellyfish Catostylus mosaicus (Rhizostomeae) in New South Wales, Australia. Marine Ecology Progress Series, 196:143–155, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/196/m196p143.pdf.

Pueschel:2001:SNF

[PK01] C. M. Pueschel and R. E. Korb. Storage of nitrogen in the form of protein bodies in the kelp Laminaria solidungula. Marine Ecology Progress Series, 218:107–114, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p107.pdf.

Paffenhofer:2005:DDP

[PK05a] G. A. Paffenhöfer and M. Köster. Digestion of diatoms by planktonic copepods and doliolids. Marine Ecology Progress Series, 297:303–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/297/m297p303.pdf.

Poulsen:2005:CCC

[PK05b] L. K. Poulsen and T. Kiørboe. Coprophagy and coprorhexy in the copepods Acartia tonsa and Temora longicornis: clearance rates and feeding behaviour. Marine Ecology Progress Series, 299:217–227, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/299/m299p217.pdf.

Poulsen:2006:VFD

[PK06] L. K. Poulsen and T. Kiørboe. Vertical flux and degradation rates of copepod fecal pellets in a zooplankton community dominated by small copepods. Marine Ecology Progress Series, 323:195–204, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/323/m323p195.pdf.

Premke:2006:AAD

[PKA06] K. Premke, M. Klages, and W. E. Arntz. Aggregations of Arctic deep-sea scavengers at large food falls: temporal distribution, consumption rates and population structure. Marine Ecology Progress Series, 325:121–135, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p121.pdf.

Permata:2000:HSO

[PKH00] W. D. Permata, R. A. Kinzie III, and M. Hidaka. Histological studies on the origin of planulae of the coral Pocillopora damicornis. Marine Ecology Progress Series, 200:191–200, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p191.pdf.

Petrik:2009:PSL

[PKLD09] C. M. Petrik, T. Kristiansen, R. G. Lough, and C. S. Davis. Prey selection by larval haddock and cod on copepods with species-specific behavior: an individual-based model analysis. Marine Ecology Progress Series, 396:123–143, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/396/
m396p123.pdf.

Peterson:2003:SRA

[PKM03] C. H. Peterson, R. T. Kneib, and C. A. Manen. Scaling restoration actions in the marine environment to meet quantitative targets of enhanced ecosystem services. Marine Ecology Progress Series, 264:173–175, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/264/m264p173.pdf.

Patterson:2004:ESP

[PKM04] H. M. Patterson, M. J. Kingsford, and M. T. McCulloch. Elemental signatures of Pomacentrus coelestis otoliths at multiple spatial scales on the Great Barrier Reef, Australia. Marine Ecology Progress Series, 270:229–239, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/270/
m270p229.pdf.

Perkins:2007:ILO

[PKR07] R. G. Perkins, J. C. Kromkamp, and R. P. Reid. Importance of light and oxygen for photochemical reactivation in photosynthetic stromatolite communities after natural sand burial. Marine Ecology Progress Series, 349:23–32, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p023.pdf.

Pitt:2007:JMR

[PKRK07] K. A. Pitt, M. J. Kingsford, D. Rissik, and K. Koop. Jellyfish modify the response of planktonic assemblages to nutrient pulses. Marine Ecology Progress Series, 351:1–13, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pearson:2000:REB

[PKS00] G. Pearson, L. Kautsky, and E. Serrão. Recent evolution in Baltic Fucus vesiculosus: reduced tolerance to emersion stresses compared to intertidal (North Sea) populations. Marine Ecology Progress Series, 202:67–79, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p067.pdf.

Price:2007:RMI

[PKS07] A. R. G. Price, M. J. Keeling, and I. N. Stewart. A robustness metric integrating spatial and temporal information: application to coral reefs exposed to local and regional disturbances. Marine Ecology Progress Series, 331:101–108, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p101.pdf.

Paffenhofer:2000:UCA

[PL00] G. A. Paffenhöfer and P. A. Loyd. Ultrastructure of cephalic appendage setae of marine planktonic copepods. Marine Ecology Progress Series, 203:171–180, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/203/m203p171.pdf.

Peterson:2003:CPT

[PL03] C. H. Peterson and R. N. Lipcius. Conceptual progress towards predicting quantitative ecosystem benefits of ecological restorations. Marine Ecology Progress Series, 264:297– 307, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p297.pdf.

Pechenik:2007:EPL

[PL07] J. A. Pechenik and S. H. Levine. Estimates of planktonic larval mortality using the marine gastropods Crepidula fornicata and C. plana. Marine Ecology Progress Series, 344: 107–118, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/344/m344p107.pdf.

Pohnert:2002:VUA

[PLC^{+}02] G. Pohnert, O. Lumineau, A. Cueff, S. Adolph, C. Cordevant, M. Lange, and S. Poulet. Are volatile unsaturated aldehydes from diatoms the main line of chemical defence against copepods? Marine Ecology Progress Series, 245: 33–45, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p033.pdf.

Parnell:2005:ESM

[PLCG^{+}05] P. E. Parnell, C. E. Lennert-Cody, L. Geelen, L. D. Stanley, and P. K. Dayton. Effectiveness of a small marine reserve in southern California. Marine Ecology Progress Series, 296:39–52, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p039.pdf.

Pringle:2009:GAF

[PLG09] J. M. Pringle, F. Lutscher, and E. Glick. Going against the flow: effects of non-Gaussian dispersal kernels and reproduction over multiple generations. Marine Ecology Progress Series, 377:13–17, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/377/m377p013.pdf.

Piniak:2003:APP

[PLM03] G. A. Piniak, F. Lipschultz, and J. McClelland. Assimilation and partitioning of prey nitrogen within two anthozoans and their endosymbiotic zooxanthellae. Marine Ecology Progress Series, 262:125–136, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/262/m262p125.pdf.

Peterson:2004:AWF

[PLMS04] C. H. Peterson, R. A. Luettich, Jr., F. Micheli, and G. A. Skilleter. Attenuation of water flow inside seagrass canopies of differing structure. Marine Ecology Progress Series, 268: 81–92, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/268/m268p081.pdf.

Peters:2000:ETP

[PM00] F. Peters and C. Marrasé. Effects of turbulence on plankton: an overview of experimental evidence and some theoretical considerations. Marine Ecology Progress Series, 205: 291–306, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p291.pdf.

Poulin:2003:LSD

[PM03] R. Poulin and K. N. Mouritsen. Large-scale determinants of trematode infections in intertidal gastropods. Marine Ecology Progress Series, 254:187–198, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/254/m254p187.pdf.

Page:2005:DRP

[PMG05a] B. Page, J. McKenzie, and S. D. Goldsworthy. Dietary resource partitioning among sympatric New Zealand and Australian fur seals. Marine Ecology Progress Series, 293:283– 302, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p283.pdf.

Page:2005:ISD

[PMG05b] B. Page, J. McKenzie, and S. D. Goldsworthy. Inter-sexual differences in New Zealand fur seal diving behaviour. Marine Ecology Progress Series, 304:249–264, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p249.pdf.

Pigeot:2006:CPE

[PMG^{+}06] J. Pigeot, P. Miramand, T. Guyot, P. G. Sauriau, D. Fichet, O. Le Moine, and V. Huet. Cadmium pathways in an exploited intertidal ecosystem with chronic cadmium inputs (Marennes– Oléron, Atlantic coast, France). Marine Ecology Progress Series, 307:101–114, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p101.pdf.

Peterson:2001:SDB

[PMGE01] C. H. Peterson, L. L. McDonald, R. H. Green, and W. P. Erickson. Sampling design begets conclusions: the statistical basis for detection of injury to and recovery of shoreline communities after the ‘Exxon Valdez’ oil spill. Marine Ecology Progress Series, 210:255–283, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:/

/www.int-res.com/articles/meps/210/m210p255.pdf. See

comment [GHP02].

Peterson:2002:JCM

[PMGE02] C. H. Peterson, L. L. McDonald, R. H. Green, and W. P. Erickson. The joint consequences of multiple components of statistical sampling designs. Marine Ecology Progress Series, 231:309–314, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p309.pdf.

Pawlik:2007:AFC

[PMH07] J. R. Pawlik, S. E. McMurray, and T. P. Henkel. Abiotic factors control sponge ecology in Florida mangroves. Marine Ecology Progress Series, 339:93–98, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p093.pdf.

Parker:2008:SDL

[PMH08] J. D. Parker, J. P. Montoya, and M. E. Hay. A specialist detritivore links Spartina alterniflora to salt marsh food webs. Marine Ecology Progress Series, 364:87–95, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p087.pdf.

Pardal:2000:IEL

[PMM^{+}00] M. A. Pardal, J. C. Marques, I. Metelo, A. I. Lillebø, and M. R. Flindt. Impact of eutrophication on the life cycle, population dynamics and production of Ampithoe valida (Amphipoda) along an estuarine spatial gradient (Mondego estuary, Portugal). Marine Ecology Progress Series, 196:207–219, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p207.pdf.

Perissinotto:2007:GPA

[PMNL07] R. Perissinotto, P. Mayzaud, P. D. Nichols, and J. P. Labat. Grazing by Pyrosoma atlanticum (Tunicata, Thaliacea) in the south Indian Ocean. Marine Ecology Progress Series, 330:1– 11, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Polunin:2001:FRM

[PMNP^{+}01] A. N. V. C. Polunin, B. Morales-Nin, W. E. Pawsey, J. E. Cartes, J. K. Pinnegar, and J. Moranta. Feeding relationships in Mediterranean bathyal assemblages elucidated by carbon and nitrogen stable-isotope data. Marine Ecology Progress Series, 220:13–23, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p013.pdf.

Pittman:2004:LFP

[PMP04] S. J. Pittman, C. A. McAlpine, and K. M. Pittman. Linking fish and prawns to their environment: a hierarchical landscape approach. Marine Ecology Progress Series, 283:233– 254, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/283/m283p233.pdf.

Palomo:2003:API

[PMPI03] G. Palomo, P. Martinetto, C. Perez, and O. Iribarne. Ant predation on intertidal polychaetes in a SW Atlantic estuary. Marine Ecology Progress Series, 253:165–173, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p165.pdf.

Porri:2006:STV

[PMR06] F. Porri, C. D. McQuaid, and S. Radloff. Spatio-temporal variability of larval abundance and settlement of Perna perna: differential delivery of mussels. Marine Ecology Progress Series, 315:141–150, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p141.pdf.

Page:2006:SSF

[PMS^{+}06] B. Page, J. McKenzie, M. D. Sumner, M. Coyne, and S. D. Goldsworthy. Spatial separation of foraging habitats among New Zealand fur seals. Marine Ecology Progress Series, 323: 263–279, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/323/m323p263.pdf.

Pande:2008:MRI

[PMS^{+}08] A. Pande, A. B. MacDiarmid, P. J. Smith, R. J. Davidson, R. G. Cole, D. Freeman, S. Kelly, and J. P. A. Gardner. Marine reserves increase the abundance and size of blue cod and rock lobster. Marine Ecology Progress Series, 366:147–158, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p147.pdf.

Pichavant:2002:EHS

[PMT^{+}02] K. Pichavant, V. Maxime, M. T. Thébault, H. Ollivier, J. P. Garnier, B. Bousquet, M. Diouris, G. Boeuf, and G. Nonnotte. Effects of hypoxia and subsequent recovery on turbot Scophthalmus maximus: hormonal changes and anaerobic metabolism. Marine Ecology Progress Series, 225:275– 285, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/225/m225p275.pdf.

Pujolar:2006:GPA

[PMV06] J. M. Pujolar, G. E. Maes, and F. A. M. Volckaert. Genetic patchiness among recruits in the European eel Anguilla anguilla. Marine Ecology Progress Series, 307:209– 217, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/307/m307p209.pdf.

Piedracoba:2008:PBCa

[PNCS^{+}08] S. Piedracoba, M. Nieto-Cid, C. Souto, M. Gilcoto, G. Rosón, X. A. Álvarez-Salgado, R. Varela, and F. G. Figueiras. Physical-biological coupling in the coastal upwelling system of the Ría de Vigo (NW Spain). I: In situ approach. Marine Ecology Progress Series, 353:27–40, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p027.pdf.

Piedracoba:2008:PBCb

[PNCT^{+}08] S. Piedracoba, M. Nieto-Cid, I. G. Teixeira, J. L. Garrido, X. A. Álvarez-Salgado, G. Rosón, C. G. Castro, and F. F. Pérez. Physical-biological coupling in the coastal upwelling system of the Ría de Vigo (NW Spain). II: In vitro approach. Marine Ecology Progress Series, 353:41–53, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p041.pdf.

Parry:2002:CAS

[PNK^{+}02] D. M. Parry, L. A. Nickell, M. A. Kendall, M. T. Burrows, D. A. Pilgrim, and M. B. Jones. Comparison of abundance and spatial distribution of burrowing megafauna from diver and remotely operated vehicle observations. Marine Ecology Progress Series, 244:89–93, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/244/m244p089.pdf.

Pahlow:2009:CMP

[PO09] M. Pahlow and A. Oschlies. Chain model of phytoplankton P, N and light colimitation. Marine Ecology Progress Series, 376:69–83, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/376/m376p069.pdf.

Perkins:2002:CCF

[POH^{+}02] R. G. Perkins, K. Oxborough, A. R. M. Hanlon, G. J. C. Underwood, and N. R. Baker. Can chlorophyll fluorescence be used to estimate the rate of photosynthetic electron transport within microphytobenthic biofilms? Marine Ecology Progress Series, 228:47–56, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p047.pdf.

Place:2008:GEI

[POH08] S. P. Place, M. J. O’Donnell, and G. E. Hofmann. Gene expression in the intertidal mussel Mytilus californianus: physiological response to environmental factors on a biogeographic scale. Marine Ecology Progress Series, 356:1–14, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Poltermann:2000:GPP

[Pol00] M. Poltermann. Growth, production and productivity of the Arctic sympagic amphipod Gammarus wilkitzkii. Marine Ecology Progress Series, 193:109–116, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p109.pdf.

Portner:2008:EEO

[Pör08] H. O. Pörtner. Ecosystem effects of ocean acidification in times of ocean warming: a physiologist’s view. Marine Ecology Progress Series, 373:203–217, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m373p203.pdf.

Pepin:2000:FLF

[PP00] P. Pepin and R. Penney. Feeding by a larval fish community: impact on zooplankton. Marine Ecology Progress Series, 204:199–212, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/204/m204p199.pdf.

Phillips:2002:HDM

[PP02] J. A. Phillips and I. R. Price. How different is Mediterranean Caulerpa taxifolia (Caulerpales: Chlorophyta) to other populations of the species? Marine Ecology Progress Series, 238:61–71, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/238/m238p061.pdf.

Pan:2001:PNS

[PPB^{+}01] Y. Pan, M. L. Parsons, M. Busman, P. D. R. Moeller, Q. Dortch, C. L. Powell, and G. J. Doucette. Pseudo-nitzschia sp. cf. pseudodelicatissima — a confirmed producer of domoic acid from the northern Gulf of Mexico. Marine Ecology Progress Series, 220:83–92, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/220/m220p083.pdf.

Pendleton:2009:RSM

[PPB^{+}09a] D. E. Pendleton, A. J. Pershing, M. W. Brown, C. A. Mayo, R. D. Kenney, N. R. Record, and T. V. N. Cole. Regional-scale mean copepod concentration indicates relative abundance of North Atlantic right whales. Marine Ecology Progress Series, 378:211–225, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m378p211.pdf.

Penney:2009:PMI

[PPB09b] A. J. Penney, S. J. Parker, and J. H. Brown. Protection measures implemented by New Zealand for vulnerable marine ecosystems in the South Pacific Ocean. Marine Ecology Progress Series, 397:341–354, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m397p341.pdf.

Puglisi:2002:COC

[PPBS02] M. P. Puglisi, V. J. Paul, J. Biggs, and M. Slattery. Cooccurrence of chemical and structural defenses in the gorgonian corals of Guam. Marine Ecology Progress Series, 239:105–114, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/239/m239p105.pdf.

Powers:2005:EEB

[PPC^{+}05] S. P. Powers, C. H. Peterson, R. R. Christian, E. Sullivan, M. J. Powers, M. J. Bishop, and C. P. Buzzelli. Effects of eutrophication on bottom habitat and prey resources of demersal fishes. Marine Ecology Progress Series, 302:233– 243, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p233.pdf.

Parker:2009:DCM

[PPC09] S. J. Parker, A. J. Penney, and M. R. Clark. Detection criteria for managing trawl impacts on vulnerable marine ecosystems in high seas fisheries of the South Pacific Ocean. Marine Ecology Progress Series, 397:309–317, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m397p309.pdf.

Pujo-Pay:2006:IPB

[PPCJ^{+}06] M. Pujo-Pay, P. Conan, F. Joux, L. Oriol, J. J. Naudin, and G. Cauwet. Impact of phytoplankton and bacterial production on nutrient and DOM uptake in the Rhône River plume (NW Mediterranean). Marine Ecology Progress Series, 315:43–54, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p043.pdf.

Palomo:2007:SEP

[PPCU07] M. G. Palomo, J. People, M. G. Chapman, and A. J. Underwood. Separating the effects of physical and biological aspects of mussel beds on their associated assemblages. Marine Ecology Progress Series, 344:131–142, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p131.pdf.

Powers:2009:SCO

[PPGL09] S. P. Powers, C. H. Peterson, J. H. Grabowski, and H. S. Lenihan. Success of constructed oyster reefs in no-harvest sanctuaries: implications for restoration. Marine Ecology Progress Series, 389:159–170, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m389p159.pdf.

Pinckney:2000:RPP

[PPHT00] J. L. Pinckney, H. W. Paerl, E. Haugen, and P. A. Tester. Responses of phytoplankton and Pfiesteria-like dinoflagellate zoospores to nutrient enrichment in the Neuse River Estuary, North Carolina, USA. Marine Ecology Progress Series, 192:65–78, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p065.pdf.

Podgorsek:2004:CGD

[PPI04] Lilijana Podgorsek, Ralf Petri, and Johannes F. Imhoff. Cultured and genetic diversity, and activities of sulfur-oxidizing bacteria in low-temperature hydrothermal fluids of the North Fiji Basin. Marine Ecology Progress Series, 266:65–76, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/266/m266p065.pdf.

Pillans:2005:EMR

[PPJ^{+}05] S. Pillans, R. D. Pillans, R. W. Johnstone, P. G. Kraft, M. D. E. Haywood, and H. P. Possingham. Effects of marine reserve protection on the mud crab Scylla serrata in a sex-biased fishery in subtropical Australia. Marine Ecology Progress Series, 295:201–213, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/295/m295p201.pdf.

Poulin:2002:TSV

[PPL^{+}02] E. Poulin, A. T. Palma, G. Leiva, E. Hernández, P. Martínez, S. A. Navarrete, and J. C. Castilla. Temporal and spatial variation in the distribution of epineustonic competent larvae of Concholepas concholepas along the central coast of Chile. Marine Ecology Progress Series, 229:95–104, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/229/
m229p095.pdf.

Puglisi:2000:BCC

[PPS00] M. P. Puglisi, V. J. Paul, and M. Slattery. Biogeographic comparisons of chemical and structural defenses of the Pacific gorgonians Annella mollis and A. reticulata. Marine Ecology Progress Series, 207:263–272, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/207/m207p263.pdf.

Phalan:2007:FBF

[PPS^{+}07] B. Phalan, R. A. Phillips, J. R. D. Silk, V. Afanasyev, A. Fukuda, J. Fox, P. Catry, H. Higuchi, and J. P. Croxall. Foraging behaviour of four albatross species by night and day. Marine Ecology Progress Series, 340:271–286, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p271.pdf.

Powers:2007:MGB

[PPSP07] M. J. Powers, C. H. Peterson, H. C. Summerson, and S. P. Powers. Macroalgal growth on bivalve aquaculture netting enhances nursery habitat for mobile invertebrates and juvenile fishes. Marine Ecology Progress Series, 339:109–122, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/339/m339p109.pdf.

Palma:2006:CBD

[PPV^{+}06] A. T. Palma, L. M. Pardo, R. Veas, C. Cartes, M. Silva, K. Manriquez, A. Diaz, C. Muñoz, and F. P. Ojeda. Coastal brachyuran decapods: settlement and recruitment under contrasting coastal geometry conditions. Marine Ecology Progress Series, 316:139–153, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/316/m316p139.pdf.

Pineira:2008:GDA

[PQRAC08] J. Piñeira, H. Quesada, E. Rolán-Alvarez, and A. Caballero. Genetic discontinuity associated with an environmentally induced barrier to gene exchange in the marine snail Littorina saxatilis. Marine Ecology Progress Series, 357:175– 184, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/357/m357p175.pdf.

Poulin:2001:SSS

[PR01] R. Poulin and S. R. Rate. Small-scale spatial heterogeneity in infection levels by symbionts of the amphipod Talorchestia quoyana (Talitridae). Marine Ecology Progress Series, 212:211–216, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p211.pdf.

Purkis:2005:STD

[PR05] S. J. Purkis and B. Riegl. Spatial and temporal dynamics of Arabian Gulf coral assemblages quantified from remotesensing and in situ monitoring data. Marine Ecology Progress Series, 287:99–113, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/287/m287p099.pdf.

Parrish:2009:GVG

[PR09] F. A. Parrish and E. B. Roark. Growth validation of gold coral Gerardia sp. in the Hawaiian Archipelago. Marine Ecology Progress Series, 397:163–172, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m397p163.pdf.

Pratchett:2001:ICS

[Pra01] M. S. Pratchett. Influence of coral symbionts on feeding preferences of crown-of-thorns starfish Acanthaster planci in the western Pacific. Marine Ecology Progress Series, 214:111– 119, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p111.pdf.

Pratt:2005:CCI

[Pra05] M. C. Pratt. Consequences of coloniality: influence of colony form and size on feeding success in the bryozoan Membranipora membranacea. Marine Ecology Progress Series, 303: 153–165, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/303/m303p153.pdf.

Page:2008:AIL

[PRB^{+}08] H. M. Page, D. C. Reed, M. A. Brzezinski, J. M. Melack, and J. E. Dugan. Assessing the importance of land and marine sources of organic matter to kelp forest food webs. Marine Ecology Progress Series, 360:47–62, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/360/
m360p047.pdf.

Price:2002:SHC

[Pri02] A. R. G. Price. Simultaneous ’hotspots’ and ’coldspots’ of marine biodiversity and implications for global conservation.

Marine Ecology Progress Series, 241:23–27, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p023.pdf.

Pichegru:2009:OBV

[PRL^{+}09] L. Pichegru, P. G. Ryan, C. Le Bohec, C. D. van der Lingen, R. Navarro, S. Petersen, S. Lewis, J. van der Westhuizen, and D. Grémillet. Overlap between vulnerable top predators and fisheries in the Benguela upwelling system: implications for marine protected areas. Marine Ecology Progress Series, 391:199–208, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m391p199.pdf.

Pershing:2009:MBEb

[PRM^{+}09a] A. J. Pershing, N. R. Record, B. C. Monger, C. A. Mayo, M. W. Brown, T. V. N. Cole, R. D. Kenney, D. E. Pendleton, and L. A. Woodard. Model-based estimates of right whale habitat use in the Gulf of Maine. Marine Ecology Progress Series, 378:245–257, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m378p245.pdf.

Pershing:2009:MBEa

[PRM^{+}09b] A. J. Pershing, N. R. Record, B. C. Monger, D. E. Pendleton, and L. A. Woodard. Model-based estimates of Calanus finmarchicus abundance in the Gulf of Maine. Marine Ecology Progress Series, 378:227–243, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m378p227.pdf.

Pichegru:2007:FBE

[PRvdL^{+}07] L. Pichegru, P. G. Ryan, C. D. van der Lingen, J. Coetzee, Y. Ropert-Coudert, and D. Grémillet. Foraging behaviour and energetics of Cape gannets Morus capensis feeding on live prey and fishery discards in the Benguela upwelling system. Marine Ecology Progress Series, 350:127–136, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/350/
m350p127.pdf.

Przeslawski:2005:CES

[Prz05] R. Przeslawski. Combined effects of solar radiation and desiccation on the mortality and development of encapsulated embryos of rocky shore gastropods. Marine Ecology Progress Series, 298:169–177, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/298/m298p169.pdf.

Purcell:2001:PSD

[PS01] J. E. Purcell and M. V. Sturdevant. Prey selection and dietary overlap among zooplanktivorous jellyfish and juvenile fishes in Prince William Sound, Alaska. Marine Ecology Progress Series, 210:67–83, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p067.pdf.

Phillips:2006:DES

[PS06] N. E. Phillips and J. S. Shima. Differential effects of suspended sediments on larval survival and settlement of New Zealand urchins Evechinus chloroticus and abalone Haliotis iris. Marine Ecology Progress Series, 314:149–158, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/314/
m314p149.pdf.

Panagiotopoulos:2007:SDL

[PS07] C. Panagiotopoulos and R. Sempéré. Sugar dynamics in large particles during in vitro incubation experiments. Marine Ecology Progress Series, 330:67–74, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/330/
m330p067.pdf.

Paddack:2008:RHS

[PS08] M. J. Paddack and S. Sponaugle. Recruitment and habitat selection of newly settled Sparisoma viride to reefs with low coral cover. Marine Ecology Progress Series, 369:205– 212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/369/m369p205.pdf.

Phillips:2004:AGE

[PSC^{+}04] R. A. Phillips, J. R. D. Silk, J. P. Croxall, V. Afanasyev, and D. R. Briggs. Accuracy of geolocation estimates for flying seabirds. Marine Ecology Progress Series, 266:265– 272, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/266/m266p265.pdf.

Phillips:2005:FPS

[PSC05] R. A. Phillips, J. R. D. Silk, and J. P. Croxall. Foraging and provisioning strategies of the light-mantled sooty albatross at South Georgia: competition and co-existence with sympatric pelagic predators. Marine Ecology Progress Series, 285:259–270, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/285/m285p259.pdf.

Platt:2003:NSD

[PSE^{+}03] T. Platt, S. Sathyendranath, A. M. Edwards, D. S. Broomhead, and O. Ulloa. Nitrate supply and demand in the mixed layer of the ocean. Marine Ecology Progress Series, 254:3–9, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p003.pdf.

Porter:2004:CWC

[PSGP04] E. T. Porter, L. P. Sanford, G. Gust, and F. C. Porter. Combined water-column mixing and benthic boundary-layer flow in mesocosms: key for realistic benthic-pelagic coupling studies. Marine Ecology Progress Series, 271:43–60, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p043.pdf.

Peck:2004:SST

[PSKC04] D. R. Peck, B. V. Smithers, A. K. Krockenberger, and B. C. Congdon. Sea surface temperature constrains wedge-tailed shearwater foraging success within breeding seasons. Marine Ecology Progress Series, 281:259–266, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p259.pdf.

Pradillon:2007:SIM

[PSPD07] F. Pradillon, A. Schmidt, J. Peplies, and N. Dubilier. Species identification of marine invertebrate early stages by wholelarvae in situ hybridisation of 18S ribosomal RNA. Marine Ecology Progress Series, 333:103–116, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/333/
m333p103.pdf.

Papetti:2009:STB

[PSPZ09] C. Papetti, E. Susana, T. Patarnello, and L. Zane. Spatial and temporal boundaries to gene flow between Chaenocephalus aceratus populations at South Orkney and South Shetlands. Marine Ecology Progress Series, 376:269–281, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p269.pdf.

Pinedo:2000:ESP

[PSRB00] S. Pinedo, R. Sardá, C. Rey, and M. Bhaud. Effect of sediment particle size on recruitment of Owenia fusiformis in the Bay of Blanes (NW Mediterranean Sea): an experimental approach to explain field distribution. Marine Ecology Progress Series, 203:205–213, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/203/m203p205.pdf.

Pineda:2006:TSS

[PSS06] J. Pineda, V. Starczak, and T. A. Stueckle. Timing of successful settlement: demonstration of a recruitment window in the barnacle Semibalanus balanoides. Marine Ecology Progress Series, 320:233–237, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/320/m320p233.pdf.

Piatt:2007:SIMa

[PSS^{+}07] John F. Piatt, William J. Sydeman, William J. Sydeman, John F. Piatt, and Howard I. Browman. Seabirds as indicators of marine ecosystems. Marine Ecology Progress Series, 352:199–309, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m352_TS.pdf.

Piatt:2007:IMR

[PSW07] I. J. F. Piatt, W. J. Sydeman, and F. Wiese. Introduction: a modern role for seabirds as indicators. Marine Ecology Progress Series, 352:199–204, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m352p199.pdf.

Painter:2008:UDU

[PSW^{+}08] S. C. Painter, R. Sanders, H. N. Waldron, M. I. Lucas, and S. Torres-Valdes. Urea distribution and uptake in the Atlantic Ocean between 50^{o}N and 50^{o}S. Marine Ecology Progress Series, 368:53–63, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p053.pdf.

Plourde:2004:CVV

[PT04] J. Plourde and J. C. Therriault. Climate variability and vertical advection of nitrates in the Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 279:33–43, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p033.pdf.

Pitta:2009:GNF

[PTA^{+}09] P. Pitta, M. Tsapakis, E. T. Apostolaki, T. Tsagaraki, M. Holmer, and I. Karakassis. ’Ghost nutrients’ from fish farms are transferred up the food web by phytoplankton grazers. Marine Ecology Progress Series, 374:1–6, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Prado:2007:EDM

[PTAR07] P. Prado, F. Tomas, T. Alcoverro, and J. Romero. Extensive direct measurements of Posidonia oceanica defoliation confirm the importance of herbivory in temperate seagrass meadows. Marine Ecology Progress Series, 340:63–71, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p063.pdf.

Pernet:2003:SSS

[PTB03] F. Pernet, R. Tremblay, and E. Bourget. Settlement success, spatial pattern and behavior of mussel larvae Mytilus spp. in experimental ’downwelling’ systems of varying velocity and turbulence. Marine Ecology Progress Series, 260: 125–140, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p125.pdf.

Peterson:2001:CPP

[PTCH01] B. J. Peterson, K. R. Thompson, J. H. Cowan, Jr., and K. L. Heck, Jr. Comparison of predation pressure in temperate and subtropical seagrass habitats based on chronographic tethering. Marine Ecology Progress Series, 224:77–85, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p077.pdf.

Parrish:2005:LCF

[PTD05] C. C. Parrish, R. J. Thompson, and D. Deibel. Lipid classes and fatty acids in plankton and settling matter during the spring bloom in a cold ocean coastal environment. Marine Ecology Progress Series, 286:57–68, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p057.pdf.

Pinn:2008:PMB

[PTH08] E. H. Pinn, R. C. Thompson, and S. J. Hawkins. Piddocks (Mollusca: Bivalvia: Pholadidae) increase topographical complexity and species diversity in the intertidal. Marine Ecology Progress Series, 355:173–182, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/355/m355p173.pdf.

Papadimitriou:2009:ICR

[PTK^{+}09] S. Papadimitriou, D. N. Thomas, H. Kennedy, H. Kuosa, and G. S. Dieckmann. Inorganic carbon removal and isotopic enrichment in Antarctic sea ice gap layers during early austral summer. Marine Ecology Progress Series, 386:15– 27, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/386/m386p015.pdf.

Papaspyrou:2004:IPT

[PTLK04] S. Papaspyrou, M. Thessalou-Legaki, and E. Kristensen. Impact of Pestarella tyrrhena on benthic metabolism in sediment microcosms enriched with seagrass and macroalgal detritus. Marine Ecology Progress Series, 281:165–179, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p165.pdf.

Pardoe:2008:STT

[PTM08] H. Pardoe, G. Thórdarson, and G. Marteinsdóttir. Spatial and temporal trends in condition of Atlantic cod Gadus morhua on the Icelandic shelf. Marine Ecology Progress Series, 362:261–277, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/362/m362p261.pdf.

Powell:2001:ESC

[PTP01] D. K. Powell, P. A. Tyler, and L. S. Peck. Effect of sperm concentration and sperm ageing on fertilisation success in the Antarctic soft-shelled clam Laternula elliptica and the Antarctic limpet Nacella concinna. Marine Ecology Progress Series, 215:191–200, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p191.pdf.

Pedersen:2000:SSM

[PTS00] O. P. Pedersen, K. S. Tande, and D. Slagstad. A synoptic sampling method applied to Calanus finmarchicus population on the Norwegian mid-shelf in 1997. Marine Ecology Progress Series, 204:143–157, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/204/m204p143.pdf.

Pecl:2006:UAT

[PTSJ06] G. T. Pecl, S. R. Tracey, J. M. Semmens, and G. D. Jackson. Use of acoustic telemetry for spatial management of southern calamary Sepioteuthis australis, a highly mobile inshore squid species. Marine Ecology Progress Series, 328:1–15, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Perkins:2001:RML

[PUB^{+}01] R. G. Perkins, G. J. C. Underwood, V. Brotas, G. C. Snow, B. Jesus, and L. Ribeiro. Responses of microphytobenthos to light: primary production and carbohydrate allocation over an emersion period. Marine Ecology Progress Series, 223:101–112, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/223/m223p101.pdf.

Pudovkin:2005:MVK

[Pud05] Alexander I. Pudovkin. In memoriam: Vladimir Kasyanov (1940–2005). Marine Ecology Progress Series, ??:??, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/misc/
m304memoriam.pdf.

Purcell:2007:ACJ

[PUL07] J. E. Purcell, S. Uye, and W. T. Lo. Anthropogenic causes of jellyfish blooms and their direct consequences for humans: a review. Marine Ecology Progress Series, 350:153– 174, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/350/m350p153.pdf.

Purcell:2003:PZL

[Pur03] J. E. Purcell. Predation on zooplankton by large jellyfish, Aurelia, Cyanea, and Aequorea, in Prince William Sound, Alaska. Marine Ecology Progress Series, 246:137– 152, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p137.pdf.

Purcell:2007:EEA

[Pur07] J. E. Purcell. Environmental effects on asexual reproduction rates of the scyphozoan Aurelia labiata. Marine Ecology Progress Series, 348:183–196, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/348/m348p183.pdf.

Petersen:2005:PPE

[PV05] D. Petersen and G. W. N. M. Van Moorsel. Pre-planular external development in the brooding coral Agaricia humilis.

Marine Ecology Progress Series, 289:307–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p307.pdf.

Peralta:2008:CSD

[PvDMB08] G. Peralta, L. A. van Duren, E. P. Morris, and T. J. Bouma. Consequences of shoot density and stiffness for ecosystem engineering by benthic macrophytes in flow dominated areas: a hydrodynamic flume study. Marine Ecology Progress Series, 368:103–115, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m368p103.pdf.

Pohn:2001:MBA

[PVGO01] M. Pöhn, K. Vopel, E. Grünberger, and J. Ott. Microclimate of the brown alga feldmannia caespitula interstitium under zero-flow conditions. Marine Ecology Progress Series, 210:285–290, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p285.pdf.

Price:2006:CBD

[PVV^{+}06] A. R. G. Price, L. P. A. Vincent, A. J. Venkatachalam, J. J. Bolton, and P. W. Basson. Concordance between different measures of biodiversity in Indian Ocean macroalgae. Marine Ecology Progress Series, 319:85–91, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/319/
m319p085.pdf.

Pan:2004:UHI

[PW04] J. F. Pan and W. X. Wang. Uptake of Hg(II) and methylmercury by the green mussel Perna viridis under different organic carbon conditions. Marine Ecology Progress Series, 276:125–136, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p125.pdf.

Pringle:2007:GAF

[PW07] J. M. Pringle and J. P. Wares. Going against the flow: maintenance of alongshore variation in allele frequency in a coastal ocean. Marine Ecology Progress Series, 335:69– 84, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/335/m335p069.pdf.

Pan:2008:ACZ

[PW08] K. Pan and W. X. Wang. Allometry of cadmium and zinc concentrations and bioaccumulation in the scallop Chlamys nobilis. Marine Ecology Progress Series, 365:115–126, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/365/m365p115.pdf.

Pan:2009:IIV

[PW09] K. Pan and W. X. Wang. Inter-individual variations in cadmium and zinc biodynamics in the scallop Chlamys nobilis. Marine Ecology Progress Series, 383:151–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p151.pdf.

Parslow-Williams:2001:NAI

[PWAT01] P. J. Parslow-Williams, R. J. A. Atkinson, and A. C. Taylor. Nucleic acids as indicators of nutritional condition in the Norway lobster Nephrops norvegicus. Marine Ecology Progress Series, 211:235–243, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/211/m211p235.pdf.

Pinhassi:2003:SVB

[PWB^{+}03] J. Pinhassi, A. Winding, S. J. Binnerup, U. L. Zweifel, B. Riemann, and Å. Hagström. Spatial variability in bacterioplankton community composition at the Skagerrak–Kattegat Front. Marine Ecology Progress Series, 255:1–13, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p001.pdf.

Phillips:2009:AFB

[PWC^{+}09] R. A. Phillips, E. D. Wakefield, J. P. Croxall, A. Fukuda, and H. Higuchi. Albatross foraging behaviour: no evidence for dual foraging, and limited support for anticipatory regulation of provisioning at South Georgia. Marine Ecology Progress Series, 391:279–292, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m391p279.pdf.

Poore:2000:PHU

[PWdN^{+}00] A. G. B. Poore, M. J. Watson, R. de Nys, J. K. Lowry, and P. D. Steinberg. Patterns of host use among alga- and sponge-associated amphipods. Marine Ecology Progress Series, 208:183–196, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p183.pdf.

Praebel:2008:CGP

[PWFC08] K. Præbel, J. I. Westgaard, S. E. Fevolden, and J. S. Christiansen. Circumpolar genetic population structure of capelin Mallotus villosus. Marine Ecology Progress Series, 360:189– 199, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/360/m360p189.pdf.

Pinckney:2009:PCS

[PWHW09] J. L. Pinckney, J. L. Wee, A. Hou, and N. D. Walker. Phytoplankton community structure responses to urban effluent inputs following Hurricanes Katrina and Rita. Marine Ecology Progress Series, 387:137–146, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/387/m387p137.pdf.

Poulet:2006:IDC

[PWL^{+}06] S. A. Poulet, T. Wichard, J. B. Ledoux, B. Lebreton, J. Marchetti, C. Dancie, D. Bonnet, A. Cueff, P. Morin, and G. Pohnert. Influence of diatoms on copepod reproduction. I. Field and laboratory observations related to Calanus helgolandicus egg production. Marine Ecology Progress Series, 308:129–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/308/m308p129.pdf.

Papastamatiou:2006:DDF

[PWLC06] Y. P. Papastamatiou, B. M. Wetherbee, C. G. Lowe, and G. L. Crow. Distribution and diet of four species of carcharhinid shark in the Hawaiian Islands: evidence for resource partitioning and competitive exclusion. Marine Ecology Progress Series, 320:239–251, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/320/m320p239.pdf.

Peck:2008:TLA

[PWM^{+}08] L. S. Peck, K. E. Webb, A. Miller, M. S. Clark, and T. Hill. Temperature limits to activity, feeding and metabolism in the Antarctic starfish Odontaster validus. Marine Ecology Progress Series, 358:181–189, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/358/m358p181.pdf.

Palacios:2007:REZ

[PZ07] S. L. Palacios and R. C. Zimmerman. Response of eelgrass Zostera marina to CO_{2} enrichment: possible impacts of climate change and potential for remediation of coastal habitats. Marine Ecology Progress Series, 344:1–13, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Przeslawski:2009:EAS

[PZA09] R. Przeslawski, Q. Zhu, and R. Aller. Effects of abiotic stressors on infaunal burrowing and associated sediment characteristics. Marine Ecology Progress Series, 392: 33–42, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p033.pdf.

Papetti:2005:GDL

[PZB^{+}05] C. Papetti, L. Zane, E. Bortolotto, A. Bucklin, and T. Patarnello. Genetic differentiation and local temporal stability of population structure in the euphausiid Meganyctiphanes norvegica. Marine Ecology Progress Series, 289:225–235, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/289/m289p225.pdf.

Pradillon:2005:CNO

[PZMG05] F. Pradillon, M. Zbinden, L. S. Mullineaux, and F. Gaill. Colonisation of newly-opened habitat by a pioneer species, Alvinella pompejana (Polychaeta: Alvinellidae), at East Pacific Rise vent sites. Marine Ecology Progress Series, 302: 147–157, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/302/m302p147.pdf.

Quevedo:2001:PCP

[QA01] M. Quevedo and R. Anadón. Protist control of phytoplankton growth in the subtropical north-east Atlantic. Marine Ecology Progress Series, 221:29–38, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/221/m221p029.pdf.

Queiroga:2006:TWC

[QAA^{+}06] H. Queiroga, M. J. Almeida, T. Alpuim, A. A. V. Flores, S. Francisco, I. Gonzàlez-Gordillo, A. I. Miranda, I. Silva, and J. Paula. Tide and wind control of megalopal supply to estuarine crab populations on the Portuguese west coast. Marine Ecology Progress Series, 307:21–36, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/feature/m307p021.pdf.

Queric:2008:PCD

[QASA08] N. V. Quéric, J. M. Arrieta, T. Soltwedel, and W. E. Arntz. Prokaryotic community dynamics in the sedimentary microenvironment of the demosponge Tentorium semisuberites from deep Arctic waters. Marine Ecology Progress Series, 370: 87–95, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p087.pdf.

Qian:2006:AAM

[QDDP06] P. Y. Qian, S. Dobretsov, H. U. Dahms, and J. Pawlik. Antifouling activity and microbial diversity of two congeneric sponges Callyspongia spp. from Hong Kong and the Bahamas. Marine Ecology Progress Series, 324:151–165, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/324/
m324p151.pdf.

Quan:2007:TME

[QFJ^{+}07] W. Quan, C. Fu, B. Jin, Y. Luo, B. Li, J. Chen, and J. Wu. Tidal marshes as energy sources for commercially important nektonic organisms: stable isotope analysis. Marine Ecology Progress Series, 352:89–99, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/352/m352p089.pdf.

Quinnell:2004:PMP

[QHD04] S. Quinnell, K. Hulsman, and P. J. F. Davie. Protein model for pollutant uptake and elimination by living organisms and its implications for ecotoxicology. Marine Ecology Progress Series, 274:1–16, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p001.pdf.

Quillfeldt:2005:DFA

[QMF05] P. Quillfeldt, R. A. R. McGill, and R. W. Furness. Diet and foraging areas of Southern Ocean seabirds and their prey inferred from stable isotopes: review and case study of Wilson’s storm-petrel. Marine Ecology Progress Series, 295:295–304, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p295.pdf.

Queiroga:2002:VMB

[QMM02] H. Queiroga, P. O. Moksnes, and S. Meireles. Vertical migration behaviour in the larvae of the shore crab Carcinus maenas from a microtidal system (Gullmarsfjord, Sweden). Marine Ecology Progress Series, 237:195–207, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p195.pdf.

Quillfeldt:2008:SIA

[QMM^{+}08] P. Quillfeldt, R. A. R. McGill, J. F. Masello, F. Weiss, I. J. Strange, P. Brickle, and R. W. Furness. Stable isotope analysis reveals sexual and environmental variability and individual consistency in foraging of thin-billed prions. Marine Ecology Progress Series, 373:137–148, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/373/m373p137.pdf.

Quetin:2003:ERA

[QR03] L. B. Quetin and R. M. Ross. Episodic recruitment in Antarctic krill Euphausia superba in the Palmer LTER study region. Marine Ecology Progress Series, 259:185–200, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/259/
m259p185.pdf.

Qian:2000:MUF

[QRS00] P. Y. Qian, D. Rittschof, and B. Sreedhar. Macrofouling in unidirectional flow: miniature pipes as experimental models for studying the interaction of flow and surface characteristics on the attachment of barnacle, bryozoan and polychaete larvae. Marine Ecology Progress Series, 207:109–121, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p109.pdf.

Quinones-Rivera:2007:POS

[QRWJF07] Zoraida J. Quiñones-Rivera, Björn Wissel, Dubravko Justić, and Brian Fry. Partitioning oxygen sources and sinks in a stratified, eutrophic coastal ecosystem using stable oxygen isotopes. Marine Ecology Progress Series, 342:69–83, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/342/m342p069.pdf.

Quijon:2005:DRR

[QS05] P. A. Quijón and P. V. R. Snelgrove. Differential regulatory roles of crustacean predators in a sub-arctic, softsediment system. Marine Ecology Progress Series, 285:137– 149, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/285/m285p137.pdf.

Quijon:2008:TCM

[QS08] P. A. Quijon and P. V. R. Snelgrove. Trophic complexity in marine sediments: new evidence from the Gulf of St. Lawrence. Marine Ecology Progress Series, 371:85–89, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/371/m371p085.pdf.

Qiu:2002:OCH

[QTB02] J. W. Qiu, R. Tremblay, and E. Bourget. Ontogenetic changes in hyposaline tolerance in the mussels Mytilus edulis and M. trossulus: implications for distribution. Marine Ecology Progress Series, 228:143–152, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/228/m228p143.pdf.

Quintana:2007:DDP

[QWY07] F. Quintana, R. P. Wilson, and P. Yorio. Dive depth and plumage air in wettable birds: the extraordinary case of the imperial cormorant. Marine Ecology Progress Series, 334: 299–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/334/m334p299.pdf.

Robles:2002:OCC

[RAA02] R. Robles, F. Alvarez, and G. Alcaraz. Oxygen consumption of the crab Callinectes rathbunae parasitized by the rhizocephalan barnacle Loxothylacus texanus as a function of salinity. Marine Ecology Progress Series, 235:189–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p189.pdf.

Rainbow:2007:TTT

[RAAT^{+}07] P. S. Rainbow, J. C. Amiard, C. Amiard-Triquet, M. S. Cheung, L. Zhang, H. Zhong, and W. X. Wang. Trophic transfer of trace metals: subcellular compartmentalization in bivalve prey, assimilation by a gastropod predator and in vitro digestion simulations. Marine Ecology Progress Series, 348:125–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p125.pdf.

Robertson:2005:OSAa

[RAC^{+}05] D. R. Robertson, J. L. Ackerman, J. H. Choat, J. M. Posada, and J. Pitt. Ocean surgeonfish Acanthurus bahianus. I. The geography of demography. Marine Ecology Progress Series, 295:229–244, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p229.pdf.

Rolff:2007:NFA

[RAE07] C. Rolff, L. Almesjö, and R. Elmgren. Nitrogen fixation and abundance of the diazotrophic cyanobacterium Aphanizomenon sp. in the Baltic proper. Marine Ecology Progress Series, 332:107–118, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m332p107.pdf.

Raffaelli:2006:BEF

[Raf06] D. Raffaelli. Biodiversity and ecosystem functioning: issues of scale and trophic complexity. Marine Ecology Progress Series, 311:285–294, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m311p285.pdf.

Rosenberg:2002:RMB

[RAH^{+}02] R. Rosenberg, S. Agrenius, B. Hellman, H. C. Nilsson, and K. Norling. Recovery of marine benthic habitats and fauna in a Swedish fjord following improved oxygen conditions. Marine Ecology Progress Series, 234:43–53, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/234/
m234p043.pdf.

Russ:2003:SMR

[RAM03] G. R. Russ, A. C. Alcala, and A. P. Maypa. Spillover from marine reserves: the case of Naso vlamingii at Apo Island, the Philippines. Marine Ecology Progress Series, 264: 15–20, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p015.pdf.

Rand:2002:MFG

[Ran02] P. S. Rand. Modeling feeding and growth in Gulf of Alaska sockeye salmon: implications for high-seas distribution and migration. Marine Ecology Progress Series, 234:265–280, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p265.pdf.

Roegner:2007:WTI

[RAS07] G. C. Roegner, D. A. Armstrong, and A. L. Shanks. Wind and tidal influences on larval crab recruitment to an Oregon estuary. Marine Ecology Progress Series, 351:177–188, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/351/m351p177.pdf.

Rinkevich:2003:FFF

[RASB03] B. Rinkevich, D. Angel, S. Shafir, and L. Bongiorni. ’Fair is foul and foul is fair’: response to a critique. Marine Ecology Progress Series, 261:305–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p305.pdf.

Richard:2007:SIO

[RATD07] M. Richard, P. Archambault, G. Thouzeau, and G. Desrosiers. Summer influence of 1 and 2 yr old mussel cultures on benthic fluxes in Grande–Entrée lagoon, Îles-de-la-Madeleine (Québec, Canada). Marine Ecology Progress Series, 338:131– 143, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p131.pdf.

Rainbow:2002:ECS

[RB02] P. S. Rainbow and W. H. Black. Effects of changes in salinity and osmolality on the rate of uptake of zinc by three crabs of different ecologies. Marine Ecology Progress Series, 244:205–217, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p205.pdf.

Robbins:2004:RBH

[RB04] B. D. Robbins and S. S. Bell. Relationships between a hermit crab and its shell resource: spatial patterns within a seagrass-dominated landscape. Marine Ecology Progress Series, 282:221–227, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p221.pdf.

Rainbow:2005:PPC

[RB05a] P. S. Rainbow and W. H. Black. Physicochemistry or physiology: cadmium uptake and effects of salinity and osmolality in three crabs of different ecologies. Marine Ecology Progress Series, 286:217–229, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p217.pdf.

Ruxton:2005:CMB

[RB05b] G. D. Ruxton and D. M. Bailey. Combining motility and bioluminescent signalling aids mate finding in deep-sea fish: a simulation study. Marine Ecology Progress Series, 293: 253–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/293/m293p253.pdf.

Ronconi:2008:LFF

[RB08] R. A. Ronconi and A. E. Burger. Limited foraging flexibility: increased foraging effort by a marine predator does not buffer against scarce prey. Marine Ecology Progress Series, 366:245–258, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p245.pdf.

Redfern:2008:ASD

[RBB^{+}08] J. V. Redfern, J. Barlow, L. T. Ballance, T. Gerrodette, and E. A. Becker. Absence of scale dependence in dolphin-habitat models for the eastern tropical Pacific Ocean. Marine Ecology Progress Series, 363:1–14, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Riebesell:2000:CIF

[RBDK00] U. Riebesell, S. Burkhardt, A. Dauelsberg, and B. Kroon. Carbon isotope fractionation by a marine diatom: dependence on the growth-rate-limiting resource. Marine Ecology Progress Series, 193:295–303, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/193/m193p295.pdf.

Rilov:2001:LAS

[RBG01] G. Rilov, Y. Benayahu, and A. Gasith. Low abundance and skewed population structure of the whelk Stramonita haemastoma along the Israeli Mediterranean coast. Marine Ecology Progress Series, 218:189–202, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/218/m218p189.pdf.

Rilov:2004:LED

[RBG04] G. Rilov, Y. Benayahu, and A. Gasith. Life on the edge: do biomechanical and behavioral adaptations to wave-exposure correlate with habitat partitioning in predatory whelks? Marine Ecology Progress Series, 282:193–204, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p193.pdf.

Ruiz:2005:BOP

[RBG05] J. M. Ruiz, R. Barreiro, and J. J. González. Biomonitoring organotin pollution with gastropods and mussels. Marine Ecology Progress Series, 287:169–176, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p169.pdf.

Robinson:2007:CSA

[RBG^{+}07] T. B. Robinson, G. M. Branch, C. L. Griffiths, A. Govender, and P. A. R. Hockey. Changes in South African rocky intertidal invertebrate community structure associated with the invasion of the mussel Mytilus galloprovincialis. Marine Ecology Progress Series, 340:163–171, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/340/m340p163.pdf.

Ruchonnet:2006:FAC

[RBGM06] D. Ruchonnet, M. Boutoute, C. Guinet, and P. Mayzaud. Fatty acid composition of Mediterranean fin whale Balaenoptera physalus blubber with respect to body heterogeneity and trophic interaction. Marine Ecology Progress Series, 311:165–174, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/311/m311p165.pdf.

Ramp:2006:SAB

[RBHS06] C. Ramp, M. Bérubé, W. Hagen, and R. Sears. Survival of adult blue whales Balaenoptera musculus in the Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 319: 287–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p287.pdf.

Raaberg:2005:EPL

[RBJB^{+}05] S. Råberg, R. Berger-Jönsson, A. Björn, E. Granéli, and L. Kautsky. Effects of Pilayella littoralis on Fucus vesiculosus recruitment: implications for community composition. Marine Ecology Progress Series, 289:131–139, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p131.pdf.

Riisgaard:2004:PRJ

[RBL^{+}04] Hans Ulrik Riisgård, Hartmut Barth, Poul Scheel Larsen, Peter Roepstorff, Ferdinando Boero, John Dolan, Ulrich Sommer, Valerio Zupo, Jan Vermaat, Thomas R. Anderson, Ron T. Kneib, Bente Aa. Lomstein, Ian Jenkinson, and Sandra E. Shumway. Peer review: journal articles versus research proposals. Marine Ecology Progress Series, 277:301–309, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m277p301.pdf.

Rodary:2000:BDM

[RBMB00] D. Rodary, W. Bonneau, Y. L. Maho, and C. A. Bost. Benthic diving in male emperor penguins Aptenodytes forsteri foraging in winter. Marine Ecology Progress Series, 207:171–181, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p171.pdf.

Rasheed:2002:ERF

[RBRH02] M. Rasheed, M. I. Badran, C. Richter, and M. Huettel. Effect of reef framework and bottom sediment on nutrient enrichment in a coral reef of the Gulf of Aqaba, Red Sea. Marine Ecology Progress Series, 239:277–285, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p277.pdf.

Rodriguez:2000:ISU

[RBRSP00] C. A. Rodriguez, Howard I. Browman, J. A. Runge, and J. F. St-Pierre. Impact of solar ultraviolet radiation on hatching of a marine copepod, Calanus finmarchicus. Marine Ecology Progress Series, 193:85–93, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p085.pdf.

Rodriguez:2000:HSN

[RBSP00] C. A. Rodriguez, Howard I. Browman, and J. F. St-Pierre. High survival of neustonic zoea I larvae of American lobster Homarus americanus following short-term exposure to ultraviolet radiation (280 to 400 nm). Marine Ecology Progress Series, 193:305–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p305.pdf.

Rainbow:2003:EPF

[RBW03] P. S. Rainbow, G. Blackmore, and W. X. Wang. Effects of previous field-exposure history on the uptake of trace metals from water and food by the barnacle Balanus amphitrite. Marine Ecology Progress Series, 259:201–213, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/259/
m259p201.pdf.

Rossi:2000:MIH

[RC00] L. Rossi and M. L. Costantini. Mapping the intra-habitat variation of leaf mass loss rate in a brackish Mediterranean lake. Marine Ecology Progress Series, 203:145–159, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p145.pdf.

Richardson:2004:DLL

[RC04a] D. E. Richardson and R. K. Cowen. Diversity of leptocephalus larvae around the island of Barbados (West Indies): relevance to regional distributions. Marine Ecology Progress Series, 282:271–284, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p271.pdf.

Roberson:2004:VBM

[RC04b] L. M. Roberson and J. A. Coyer. Variation in blade morphology of the kelp Eisenia arborea: incipient speciation due to local water motion? Marine Ecology Progress Series, 282:115–128, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p115.pdf.

Russell:2005:NIB

[RC05] B. D. Russell and S. D. Connell. A novel interaction between nutrients and grazers alters relative dominance of marine habitats. Marine Ecology Progress Series, 289:5–11, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p005.pdf.

Russell:2007:RGS

[RC07] B. D. Russell and S. D. Connell. Response of grazers to sudden nutrient pulses in oligotrophic versus eutrophic conditions. Marine Ecology Progress Series, 349:73–80, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p073.pdf.

Roberts:2008:DBR

[RC08] B. K. Roberts and S. D. Connell. Detecting benthic responses to human-induced change: effectiveness of alternate taxonomic classification and indices. Marine Ecology Progress Series, 358:75–84, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/358/m358p075.pdf.

Robertson:2009:SFB

[RC09] D. R. Robertson and K. L. Cramer. Shore fishes and biogeographic subdivisions of the Tropical Eastern Pacific. Marine Ecology Progress Series, 380:1–17, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ribes:2003:PRC

[RCAK03] M. Ribes, R. Coma, M. J. Atkinson, and R. A. Kinzie III. Particle removal by coral reef communities: picoplankton is a major source of nitrogen. Marine Ecology Progress Series, 257:13–23, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p013.pdf.

Ryan:2005:PBC

[RCB05] J. P. Ryan, F. P. Chavez, and J. G. Bellingham. Physicalbiological coupling in Monterey Bay, California: topographic influences on phytoplankton ecology. Marine Ecology Progress Series, 287:23–32, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/287/m287p023.pdf.

Richardson:2001:AHT

[RCC01] C. A. Richardson, S. R. N. Chenery, and J. M. Cook. Assessing the history of trace metal (Cu, Zn, Pb) contamination in the North Sea through laser ablation-ICP-MS of horse mussel Modiolus modiolus shells. Marine Ecology Progress Series, 211:157–167, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/211/m211p157.pdf.

Regoli:2004:SVP

[RCC^{+}04] F. Regoli, C. Cerrano, E. Chierici, M. C. Chiantore, and G. Bavestrello. Seasonal variability of prooxidant pressure and antioxidant adaptation to symbiosis in the Mediterranean demosponge Petrosia ficiformis. Marine Ecology Progress Series, 275:129–137, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p129.pdf.

Ravaux:2009:HVS

[RCC^{+}09] J. Ravaux, D. Cottin, T. Chertemps, G. Hamel, and B. Shillito. Hydrothermal vent shrimps display low expression of the heat-inducible hsp70 gene in nature. Marine Ecology Progress Series, 396:153–156, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/396/m396p153.pdf.

Regula:2009:PSM

[RCCK09] C. Regula, S. P. Colin, J. H. Costello, and H. Kordula. Prey selection mechanism of ambush-foraging hydromedusae. Marine Ecology Progress Series, 374:135–144, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p135.pdf.

Roberts:2006:SDS

[RCDC06] D. E. Roberts, S. P. Cummins, A. R. Davis, and M. G. Chapman. Structure and dynamics of sponge-dominated assemblages on exposed and sheltered temperate reefs. Marine Ecology Progress Series, 321:19–30, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p019.pdf.

Robert:2007:EGR

[RCF07] D. Robert, M. Castonguay, and L. Fortier. Early growth and recruitment in Atlantic mackerel Scomber scombrus: discriminating the effects of fast growth and selection for fast growth. Marine Ecology Progress Series, 337:209–219, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m337p209.pdf.

Robert:2009:EPP

[RCF09] D. Robert, M. Castonguay, and L. Fortier. Effects of preferred prey density and temperature on feeding success and recent growth in larval mackerel of the southern Gulf of St. Lawrence. Marine Ecology Progress Series, 377:227–237, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/377/
m377p227.pdf.

Ruiz-Cooley:2004:TRB

[RCGA^{+}04] R. I. Ruiz-Cooley, D. Gendron, S. Aguíñiga, S. Mesnick, and J. D. Carriquiry. Trophic relationships between sperm whales and jumbo squid using stable isotopes of C and N. Marine Ecology Progress Series, 277:275–283, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p275.pdf.

Riascos:2009:TDC

[RCL^{+}09] J. M. Riascos, D. Carstensen, J. Laudien, W. E. Arntz, M. E. Oliva, A. Güntner, and O. Heilmayer. Thriving and declining: climate variability shaping life-history and population persistence of Mesodesma donacium in the Humboldt Upwelling System. Marine Ecology Progress Series, 385:151– 163, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/385/m385p151.pdf.

Robertson:2005:OSAb

[RCP^{+}05] D. R. Robertson, J. H. Choat, J. M. Posada, J. Pitt, and J. L. Ackerman. Ocean surgeonfish Acanthurus bahianus. II. Fishing effects on longevity, size and abundance? Marine Ecology Progress Series, 295:245–256, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/295/m295p245.pdf.

Ribes:2003:NFT

[RCR03] M. Ribes, R. Coma, and S. Rossi. Natural feeding of the temperate asymbiotic octocoral-gorgonian Leptogorgia sarmentosa (Cnidaria: Octocorallia). Marine Ecology Progress Series, 254:141–150, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p141.pdf.

Ropert-Coudert:2006:ERF

[RCWG^{+}06] Y. Ropert-Coudert, R. P. Wilson, D. Grémillet, A. Kato, S. Lewis, and P. G. Ryan. Electrocardiogram recordings in free-ranging gannets reveal minimum difference in heart rate during flapping versus gliding flight. Marine Ecology Progress Series, 328:275–284, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/328/m328p275.pdf. See erra-

tum [Ano07f].

Ropert-Coudert:2007:APC

[RCWYK07] Y. Ropert-Coudert, R. P. Wilson, K. Yoda, and A. Kato. Assessing performance constraints in penguins with externallyattached devices. Marine Ecology Progress Series, 333:281– 289, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p281.pdf.

Roberts:2006:EMS

[RDC06] D. E. Roberts, A. R. Davis, and S. P. Cummins. Experimental manipulation of shade, silt, nutrients and salinity on the temperate reef sponge Cymbastela concentrica. Marine Ecology Progress Series, 307:143–154, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/307/m307p143.pdf.

Rines:2002:TLC

[RDD^{+}02] J. E. B. Rines, P. L. Donaghay, M. M. Dekshenieks, J. M. Sullivan, and M. S. Twardowski. Thin layers and camouflage: hidden Pseudo-nitzschia spp. (Bacillariophyceae) populations in a fjord in the San Juan Islands, Washington, USA.

Marine Ecology Progress Series, 225:123–137, 2002. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p123.pdf.

Rosenberg:2007:ACA

[RDG^{+}07] R. Rosenberg, E. Davey, J. Gunnarsson, K. Norling, and M. Frank. Application of computer-aided tomography to visualize and quantify biogenic structures in marine sediments. Marine Ecology Progress Series, 331:23–34, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/331/
m331p023.pdf.

Roberts:2009:MRC

[RDH^{+}09] J. M. Roberts, A. J. Davies, L. A. Henry, L. A. Dodds, G. C. A. Duineveld, M. S. S. Lavaleye, C. Maier, R. W. M. van Soest, M. J. N. Bergman, V. Hühnerbach, V. A. I. Huvenne, D. J. Sinclair, T. Watmough, D. Long, S. L. Green, and H. van Haren. Mingulay reef complex: an interdisciplinary study of cold-water coral habitat, hydrography and biodiversity. Marine Ecology Progress Series, 397:139–151, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p139.pdf.

Rogers:2002:EAD

[RdNS02] C. N. Rogers, R. de Nys, and P. D. Steinberg. Effects of algal diet on the performance and susceptibility to predation of the sea hare Aplysia parvula. Marine Ecology Progress Series, 236:241–254, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p241.pdf.

Riding:2009:TFU

[RDS^{+}09] T. A. C. Riding, T. E. Dennis, C. L. Stewart, M. M. Walker, and J. C. Montgomery. Tracking fish using ‘buoy-based’ GPS telemetry. Marine Ecology Progress Series, 377:255– 262, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p255.pdf.

Rodriguez:2000:MCI

[RE00a] R. A. Rodriguez and C. E. Epifanio. Multiple cues for induction of metamorphosis in larvae of the common mud crab Panopeus herbstii. Marine Ecology Progress Series, 195:221– 229, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/195/m195p221.pdf.

Rolff:2000:URO

[RE00b] C. Rolff and R. Elmgren. Use of riverine organic matter in plankton food webs of the Baltic Sea. Marine Ecology Progress Series, 197:81–101, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/197/m197p081.pdf.

Rebstock:2001:LTS

[Reb01] G. A. Rebstock. Long-term stability of species composition in calanoid copepods off southern California. Marine Ecology Progress Series, 215:213–224, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/215/m215p213.pdf.

Rees:2003:SFN

[Ree03] T. A. V. Rees. Safety factors and nutrient uptake by seaweeds. Marine Ecology Progress Series, 263:29–42, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/263/
m263p029.pdf.

Rex:2006:GBP

[REM^{+}06] M. A. Rex, R. J. Etter, J. S. Morris, J. Crouse, C. R. McClain, N. A. Johnson, C. T. Stuart, J. W. Deming, R. Thies, and R. Avery. Global bathymetric patterns of standing stock and body size in the deep-sea benthos. Marine Ecology Progress Series, 317:1–8, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Reusch:2001:NMO

[Reu01] T. B. H. Reusch. New markers-old questions: population genetics of seagrasses. Marine Ecology Progress Series, 211: 261–274, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/211/m211p261.pdf.

Royer:2006:RDD

[RF06] F. Royer and J. M. Fromentin. Recurrent and densitydependent patterns in long-term fluctuations of Atlantic bluefin tuna trap catches. Marine Ecology Progress Series, 319:237–249, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p237.pdf.

Ramskov:2008:LHP

[RF08a] T. Ramskov and V. E. Forbes. Life history and population dynamics of the opportunistic polychaete Capitella sp. I in relation to sediment organic matter. Marine Ecology Progress Series, 369:181–192, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/369/m369p181.pdf.

Richoux:2008:TED

[RF08b] N. B. Richoux and P. W. Froneman. Trophic ecology of dominant zooplankton and macrofauna in a temperate, oligotrophic South African estuary: a fatty acid approach. Marine Ecology Progress Series, 357:121–137, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p121.pdf.

Redfern:2006:TCH

[RFB^{+}06] J. V. Redfern, M. C. Ferguson, E. A. Becker, K. D. Hyrenbach, C. Good, J. Barlow, K. Kaschner, M. F. Baumgartner, K. A. Forney, L. T. Ballance, P. Fauchald, P. Halpin, T. Hamazaki, A. J. Pershing, S. S. Qian, A. Read, S. B. Reilly, L. Torres, and F. Werne. Techniques for cetacean-habitat modeling. Marine Ecology Progress Series, 310:271–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p271.pdf.

Robinson:2007:TLC

[RFB07] H. E. Robinson, C. M. Finelli, and E. J. Buskey. The turbulent life of copepods: effects of water flow over a coral reef on their ability to detect and evade predators. Marine Ecology Progress Series, 349:171–181, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p171.pdf.

Roed:2003:ICD

[RFCM03] K. H. Røed, S. E. Fevolden, J. S. Christiansen, and L. Midthjell. Isolation and characterization of DNA microsatellites for capelin Mallotus villosus population genetic studies. Marine Ecology Progress Series, 262:299–303, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/262/
m262p299.pdf.

Royer:2004:ABB

[RFG04] F. Royer, J. M. Fromentin, and P. Gaspar. Association between bluefin tuna schools and oceanic features in the western Mediterranean. Marine Ecology Progress Series, 269: 249–263, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/269/m269p249.pdf.

Rose:2009:EIP

[RFG^{+}09] J. M. Rose, Y. Feng, C. J. Gobler, R. Gutierrez, C. E. Hare, K. Leblanc, and D. A. Hutchins. Effects of increased pCO_{2} and temperature on the North Atlantic spring bloom. II. Microzooplankton abundance and grazing. Marine Ecology Progress Series, 388:27–40, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p027.pdf.

Ruesink:2006:CPA

[RFH^{+}06] J. L. Ruesink, B. E. Feist, C. J. Harvey, J. S. Hong, A. C. Trimble, and L. M. Wisehart. Changes in productivity associated with four introduced species: ecosystem transformation of a ‘pristine’ estuary. Marine Ecology Progress Series, 311:203–215, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m311p203.pdf.

Rilov:2007:CHM

[RFLC07] G. Rilov, W. F. Figueira, S. J. Lyman, and L. B. Crowder. Complex habitats may not always benefit prey: linking visual field with reef fish behavior and distribution. Marine Ecology Progress Series, 329:225–238, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/329/m329p225.pdf.

Rosel:2003:SAS

[RFLK03] P. E. Rosel, A. Frantzis, C. Lockyer, and A. Komnenou. Source of Aegean Sea harbour porpoises. Marine Ecology Progress Series, 247:257–261, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/247/m247p257.pdf.

Rivadeneira:2002:LTS

[RFN02] M. M. Rivadeneira, M. Fernández, and S. A. Navarrete. Latitudinal trends of species diversity in rocky intertidal herbivore assemblages: spatial scale and the relationship between local and regional species richness. Marine Ecology Progress Series, 245:123–131, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p123.pdf.

Reynaud:2004:ELT

[RFPB^{+}04] S. Reynaud, C. Ferrier-Pagès, F. Boisson, D. Allemand, and R. G. Fairbanks. Effect of light and temperature on calcification and strontium uptake in the scleractinian coral Acropora verweyi. Marine Ecology Progress Series, 279:105– 112, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p105.pdf.

Reynaud:2002:EFC

[RFPS^{+}02] S. Reynaud, C. Ferrier-Pagès, R. Sambrotto, A. Juillet-Leclerc, J. Jaubert, and J. P. Gattuso. Effect of feeding on the carbon and oxygen isotopic composition in the tissues and skeleton of the zooxanthellate coral Stylophora pistillata. Marine Ecology Progress Series, 238:81–89, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p081.pdf.

Rouyer:2008:AMT

[RFSC08] T. Rouyer, J. M. Fromentin, N. C. Stenseth, and B. Cazelles. Analysing multiple time series and extending significance testing in wavelet analysis. Marine Ecology Progress Series, 359:

11–23, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/359/m359p011.pdf.

Ross:2009:NCB

[RG09] O. N. Ross and R. J. Geider. New cell-based model of photosynthesis and photo-acclimation: accumulation and mobilisation of energy reserves in phytoplankton. Marine Ecology Progress Series, 383:53–71, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m383p053.pdf.

Riisgaard:2002:CSW

[RGA^{+}02] H. U. Riisgård, A. Grémare, J. M. Amouroux, F. Charles, G. Vétion, R. Rosenberg, and C. Nielsen. Comparative study of water-processing in two ciliary filter-feeding polychaetes (Ditrupa arietina and Euchone papillosa) from two different habitats. Marine Ecology Progress Series, 229:113– 126, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/229/m229p113.pdf.

Rocha:2002:RTS

[RGB02] C. Rocha, H. Galvão, and A. Barbosa. Role of transient silicon limitation in the development of cyanobacteria blooms in the Guadiana estuary, south-western Iberia. Marine Ecology Progress Series, 228:35–45, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/228/m228p035.pdf.

Russell:2005:PSN

[RGC05] B. D. Russell, B. M. Gillanders, and S. D. Connell. Proximity and size of neighbouring habitat affects invertebrate diversity. Marine Ecology Progress Series, 296:31–38, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/296/
m296p031.pdf.

Rodriguez:2006:SFP

[RGC^{+}06] F. Rodríguez, J. L. Garrido, B. G. Crespo, B. Arbones, and F. G. Figueiras. Size-fractionated phytoplankton pigment groups in the NW Iberian upwelling system: impact of the Iberian Poleward Current. Marine Ecology Progress Series, 323:59–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/323/m323p059.pdf.

Rodriguez-Grana:2008:FWS

[RGCC^{+}08] L. Rodríguez-Graña, D. Calliari, D. Conde, J. Sellanes, and R. Urrutia. Food web of a SW Atlantic shallow coastal lagoon: spatial environmental variability does not impose substantial changes in the trophic structure. Marine Ecology Progress Series, 362:69–83, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/362/m362p069.pdf.

Rodriguez-Grana:2005:FED

[RGCL^{+}05] L. Rodríguez-Graña, L. Castro, M. Loureiro, H. E. González, and D. Calliari. Feeding ecology of dominant larval myctophids in an upwelling area of the Humboldt Current. Marine Ecology Progress Series, 290:119–134, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p119.pdf.

Redondo-Gomez:2007:FND

[RGCL^{+}07] S. Redondo-Gómez, J. M. Castillo, C. J. Luque, T. Luque, M. E. Figueroa, and A. J. Davy. Fundamental niche differentiation in subspecies of Sarcocornia perennis on a salt marsh elevational gradient. Marine Ecology Progress Series, 347:15–20, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/347/m347p015.pdf.

Rosenberg:2008:VQM

[RGD^{+}08] R. Rosenberg, A. Grémare, J. C. Duchêne, E. Davey, and M. Frank. 3D visualization and quantification of marine benthic biogenic structures and particle transport utilizing computer-aided tomography. Marine Ecology Progress Series, 363:171–182, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/363/m363p171.pdf.

Roark:2006:RBA

[RGDI06] E. B. Roark, T. P. Guilderson, R. B. Dunbar, and B. L. Ingram. Radiocarbon-based ages and growth rates of Hawaiian deep-sea corals. Marine Ecology Progress Series, 327:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Roux:2002:ERU

[RGDN02] R. Roux, M. Gosselin, G. Desrosiers, and C. Nozais. Effects of reduced UV radiation on a microbenthic community during a microcosm experiment. Marine Ecology Progress Series, 225:29–43, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p029.pdf.

Rilov:2000:UUT

[RGEB00] G. Rilov, A. Gasith, S. M. Evans, and Y. Benayahu. Unregulated use of TBT-based antifouling paints in Israel (eastern Mediterranean): high contamination and imposex levels in two species of marine gastropods. Marine Ecology Progress Series, 192:229–238, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p229.pdf.

Riveiro:2003:ELA

[RGF^{+}03] I. Riveiro, C. Guisande, C. Franco, A. Lago de Lanzós, I. Maneiro, and A. R. Vergara. Egg and larval amino acid composition as indicators of niche resource partitioning in pelagic fish species. Marine Ecology Progress Series, 260: 255–262, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p255.pdf.

Rainbow:2004:EFC

[RGJ^{+}04] P. S. Rainbow, A. Geffard, A. Y. Jeantet, B. D. Smith, J. C. Amiard, and C. Amiard-Triquet. Enhanced food-chain transfer of copper from a diet of copper-tolerant estuarine worms. Marine Ecology Progress Series, 271:183–191, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p183.pdf.

Riveiro:2000:IPE

[RGL^{+}00] I. Riveiro, C. Guisande, M. Lloves, L. Maneiro, and J. M. Cabanas. Importance of parental effects on larval survival in Sardina pilchardus. Marine Ecology Progress Series, 205: 249–258, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/205/m205p249.pdf.

Ralph:2001:ZEB

[RGL01] P. J. Ralph, R. Gademann, and A. W. D. Larkum. Zooxanthellae expelled from bleached corals at 33^{o}C are photosynthetically competent. Marine Ecology Progress Series, 220:163–168, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p163.pdf.

Rindorf:2006:PSC

[RGL06] A. Rindorf, H. Gislason, and P. Lewy. Prey switching of cod and whiting in the North Sea. Marine Ecology Progress Series, 325:243–253, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/325/m325p243.pdf.

Rossi:2008:CDE

[RGMH08] F. Rossi, B. Gribsholt, J. J. Middelburg, and C. Heip. Context-dependent effects of suspension feeding on intertidal ecosystem functioning. Marine Ecology Progress Series, 354: 47–57, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p047.pdf.

Ravago-Gotanco:2007:INE

[RGMJM07] R. G. Ravago-Gotanco, R. M. Magsino, and M. A. Juinio-Meñez. Influence of the North Equatorial Current on the population genetic structure of Tridacna crocea (Mollusca: Tridacnidae) along the eastern Philippine seaboard. Marine Ecology Progress Series, 336:161–168, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p161.pdf.

Riveiro:2004:PEE

[RGMV04] I. Riveiro, C. Guisande, I. Maneiro, and A. R. Vergara. Parental effects in the European sardine Sardina pilchardus. Marine Ecology Progress Series, 274:225–234, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p225.pdf.

Reguera:2003:CCP

[RGP^{+}03] B. Reguera, E. Garcés, Y. Pazos, I. Bravo, I. Ramilo, and S. González-Gil. Cell cycle patterns and estimates of in situ division rates of dinoflagellates of the genus Dinophysis by a postmitotic index. Marine Ecology Progress Series, 249:117–131, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p117.pdf.

Rozanska:2009:IEF

[RGP^{+}09] M. Rozanska, M. Gosselin, M. Poulin, J. M. Wiktor, and C. Michel. Influence of environmental factors on the development of bottom ice protist communities during the winterspring transition. Marine Ecology Progress Series, 386:43– 59, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m386p043.pdf.

Reinsel:2001:EFA

[RGR^{+}01] K. A. Reinsel, P. S. Glas, J. R. Rayburn, M. K. Pritchard, and W. S. Fisher. Effects of food availability on survival, growth, and reproduction of the grass shrimp Palaemonetes pugio: a laboratory study. Marine Ecology Progress Series, 220:231–239, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p231.pdf.

Reiss:2009:EFD

[RGS^{+}09] H. Reiss, S. P. R. Greenstreet, K. Sieben, S. Ehrich, G. J. Piet, F. Quirijns, L. Robinson, W. J. Wolff, and I. Kröncke. Effects of fishing disturbance on benthic communities and secondary production within an intensively fished area. Marine Ecology Progress Series, 394:201–213, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p201.pdf.

Rouleau:2001:CAS

[RGT01] C. Rouleau, C. Gobeil, and H. Tjälve. Cadmium accumulation in the snow crab Chionoecetes opilio. Marine Ecology Progress Series, 224:207–217, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/224/m224p207.pdf.

Rouleau:2006:CAC

[RGT06] C. Rouleau, C. Gobeil, and H. Tjälve. Cadmium accumulation in coastal demersal fish. Marine Ecology Progress Series, 311:131–143, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/311/m311p131.pdf.

Rorandelli:2007:FMS

[RGVC07] Rocco Rorandelli, Marina Gomei, Marco Vannini, and Stefano Cannicci. Feeding and masking selection in Inachus phalangium (Decapoda, Majidae): dressing up has never been so complicated. Marine Ecology Progress Series, 336:225– 233, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/336/m336p225.pdf.

Riera:2003:TEN

[RH03] P. Riera and C. Hubas. Trophic ecology of nematodes from various microhabitats of the Roscoff Aber Bay (France): importance of stranded macroalgae evidenced through δ ^{13}C and δ ^{15}N. Marine Ecology Progress Series, 260:151–159, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/260/m260p151.pdf.

Ruxton:2004:EFO

[RH04] G. D. Ruxton and D. C. Houston. Energetic feasibility of an obligate marine scavenger. Marine Ecology Progress Series, 266:59–63, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p059.pdf.

Reed:2006:SHD

[RH06] B. J. Reed and K. A. Hovel. Seagrass habitat disturbance: how loss and fragmentation of eelgrass Zostera marina influences epifaunal abundance and diversity. Marine Ecology Progress Series, 326:133–143, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/326/m326p133.pdf.

Ryer:2008:IPE

[RH08] C. H. Ryer and T. P. Hurst. Indirect predator effects on age-0 northern rock sole Lepidopsetta polyxystra: growth suppression and temporal reallocation of feeding. Marine Ecology Progress Series, 357:207–212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/357/m357p207.pdf.

Riisgaard:2009:RFU

[RH09] K. Riisgaard and P. J. Hansen. Role of food uptake for photosynthesis, growth and survival of the mixotrophic dinoflagellate Dinophysis acuminata. Marine Ecology Progress Series, 381:51–62, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/381/m381p051.pdf.

Reed:2009:PSV

[RHBW09] D. C. Reed, S. J. Holbrook, C. A. Blanchette, and S. Worcester. Patterns and sources of variation in flowering, seed supply and seedling recruitment in surfgrass Phyllospadix torreyi. Marine Ecology Progress Series, 384:97–106, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/384/
m384p097.pdf.

Rayner:2008:FEC

[RHC^{+}08] M. J. Rayner, M. E. Hauber, M. N. Clout, D. S. Seldon, S. Van Dijken, S. Bury, and R. A. Phillips. Foraging ecology of the Cook’s petrel Pterodroma cookii during the austral breeding season: a comparison of its two populations. Marine Ecology Progress Series, 370:271–284, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/370/m370p271.pdf.

Reid:2005:MIC

[RHDC05] K. Reid, S. L. Hill, T. C. D. Diniz, and M. A. Collins. Mackerel icefish Champsocephalus gunnari in the diet of upper trophic level predators at South Georgia: implications for fisheries management. Marine Ecology Progress Series, 305:153–161, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p153.pdf.

Raymond:2004:SDB

[RHG04] J. F. Raymond, J. H. Himmelman, and H. E. Guderley. Sex differences in biochemical composition, energy content and allocation to reproductive effort in the brooding sea star Leptasterias polaris. Marine Ecology Progress Series, 283:179– 190, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/283/m283p179.pdf.

Ribak:2005:MNF

[RHG05] G. Ribak, J. Heller, and A. Genin. Mucus-net feeding on organic particles by the vermetid gastropod Dendropoma maximum in and below the surf zone. Marine Ecology Progress Series, 293:77–87, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p077.pdf.

Rydin:2002:NLC

[RHGB02] E. Rydin, P. Hyenstrand, M. Gunnerhed, and P. Blomqvist. Nutrient limitation of cyanobacterial blooms: an enclosure experiment from the coastal zone of the NW Baltic proper. Marine Ecology Progress Series, 239:31–36, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p031.pdf.

Ruttenberg:2005:ELT

[RHH^{+}05] B. I. Ruttenberg, S. L. Hamilton, M. J. H. Hickford, G. L. Paradis, M. S. Sheehy, J. D. Standish, O. Ben-Tzvi, and R. R. Warner. Elevated levels of trace elements in cores of otoliths and their potential for use as natural tags. Marine Ecology Progress Series, 297:273–281, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/297/m297p273.pdf.

Reiss:2009:IAT

[RHHH09] C. S. Reiss, C. D. Hewes, and O. Holm-Hansen. Influence of atmospheric teleconnections and Upper Circumpolar Deep Water on phytoplankton biomass around Elephant Island, Antarctica. Marine Ecology Progress Series, 377: 51–62, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p051.pdf.

Rey:2001:IAD

[RHI^{+}01] C. Rey, R. Harris, X. Irigoien, R. Head, and F. Carlotti. Influence of algal diet on growth and ingestion of Calanus helgolandicus nauplii. Marine Ecology Progress Series, 216: 151–165, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/216/m216p151.pdf.

Renner:2008:SDP

[RHPB08] M. Renner, G. L. Hunt, Jr., J. F. Piatt, and G. V. Byrd. Seasonal and distributional patterns of seabirds along the Aleutian Archipelago. Marine Ecology Progress Series, 357: 301–311, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/357/m357p301.pdf.

Reid:2001:PEM

[RHS01a] P. C. Reid, N. P. Holliday, and T. J. Smyth. Pulses in the eastern margin current and warmer water off the north west European shelf linked to North Sea ecosystem changes. Marine Ecology Progress Series, 215:283–287, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p283.pdf.

Roman:2001:TSP

[RHS01b] M. R. Roman, D. V. Holliday, and L. P. Sanford. Temporal and spatial patterns of zooplankton in the Chesapeake Bay turbidity maximum. Marine Ecology Progress Series, 213:215–227, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p215.pdf.

Rowe:2008:MBC

[RHSB08] S. Rowe, J. A. Hutchings, J. E. Skjæraasen, and L. Bezanson. Morphological and behavioural correlates of reproductive success in Atlantic cod Gadus morhua. Marine Ecology Progress Series, 354:257–265, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/354/m354p257.pdf.

Robinette:2007:SPR

[RHSN07] D. P. Robinette, J. Howar, W. J. Sydeman, and N. Nur. Spatial patterns of recruitment in a demersal fish as revealed by seabird diet. Marine Ecology Progress Series, 352:259– 268, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m352p259.pdf.

Riisgaard:2001:IBC

[Rii01a] H. U. Riisgård. Inaccurate bivalve clearance rate measurements: a reply. Marine Ecology Progress Series, 221:307– 309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/221/m221p307.pdf. See [Wid01].

Riisgaard:2001:MFR

[Rii01b] H. U. Riisgård. On measurement of filtration rate in bivalvesthe stony road to reliable data: review and interpretation. Marine Ecology Progress Series, 211:275–291, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p275.pdf.

Riisgaard:2001:SRR

[Rii01c] H. U. Riisgård. The stony road to reliable filtration rate measurements in bivalves: a reply. Marine Ecology Progress Series, 215:307–310, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p307.pdf. See [Cra01].

Anonymous:2003:RHM

[Rii03] Hans Ulrich Riisgård. Theme section: Misuse of the peerreview system: time for countermeasures? Marine Ecology Progress Series, 258:297–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m258p297.pdf.

Riisgaard:2004:IMM

[Rii04a] H. U. Riisgård. Intercalibration of methods for measurement of bivalve filtration rates — a turning point. Marine Ecology Progress Series, 276:307–308, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p307.pdf.

Riisgaard:2004:TSP

[Rii04b] Hans Ulrik Riisgård. Theme section: Peer review: journal articles versus research proposals. Marine Ecology Progress Series, 277:301–309, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m277p301.pdf.

Rijstenbil:2003:EUR

[Rij03] J. W. Rijstenbil. Effects of UVB radiation and salt stress on growth, pigments and antioxidative defence of the marine diatom Cylindrotheca closterium. Marine Ecology Progress Series, 254:37–48, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p037.pdf.

Rindorf:2003:DFP

[Rin03] A. Rindorf. Diel feeding pattern of whiting in the North Sea. Marine Ecology Progress Series, 249:265–276, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p265.pdf.

Ritz:2002:SSG

[Rit02] D. A. Ritz. Survival strategies and growth of krill: avoiding predators in space and time. Marine Ecology Progress Series, 244:307–308, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p307.pdf.

Ritchie:2006:RMP

[Rit06] K. B. Ritchie. Regulation of microbial populations by coral surface mucus and mucus-associated bacteria. Marine Ecology Progress Series, 322:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Roast:2003:MGF

[RJ03] S. D. Roast and M. B. Jones. Morphometrics and gill function in Gammarus duebeni (Peracarida: Amphipoda) inhabiting a sewage treatment works. Marine Ecology Progress Series, 260:189–194, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p189.pdf.

Riser:2003:GCH

[RJBW03] C. Wexels Riser, S. Jansen, U. Bathmann, and P. Wassmann. Grazing of Calanus helgolandicus on Dinophysis norvegica during bloom conditions in the North Sea: evidence from investigations of faecal pellets. Marine Ecology Progress Series, 256:301–304, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/256/m256p301.pdf.

Ross:2002:IIS

[RJH02] D. J. Ross, C. R. Johnson, and C. L. Hewitt. Impact of introduced seastars Asterias amurensis on survivorship of juvenile commercial bivalves Fulvia tenuicostata. Marine Ecology Progress Series, 241:99–112, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/241/m241p099.pdf.

Reeves:2004:PHO

[RJS04] R. R. Reeves, E. Josephson, and T. D. Smith. Putative historical occurrence of North Atlantic right whales in mid-latitude offshore waters: ‘Maury’s Smear’ is likely apocryphal. Marine Ecology Progress Series, 282:295–305, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p295.pdf.

Radford:2008:RSU

[RJTM08] C. Radford, A. Jeffs, C. Tindle, and J. C. Montgomery. Resonating sea urchin skeletons create coastal choruses. Marine Ecology Progress Series, 362:37–43, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p037.pdf.

Raaberg:2008:GIC

[RK08] S. Råberg and L. Kautsky. Grazer identity is crucial for facilitating growth of the perennial brown alga Fucus vesiculosus. Marine Ecology Progress Series, 361:111–118, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p111.pdf.

Roth:2009:DAM

[RK09] M. S. Roth and N. Knowlton. Distribution, abundance, and microhabitat characterization of small juvenile corals at Palmyra Atoll. Marine Ecology Progress Series, 376:133– 142, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m376p133.pdf.

Rysgaard:2001:BPH

[RKGH01] S. Rysgaard, M. Kühl, R. N. Glud, and J. Würgler Hansen. Biomass, production and horizontal patchiness of sea ice algae in a high-Arctic fjord (Young Sound, NE Greenland). Marine Ecology Progress Series, 223:15–26, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/223/m223p015.pdf.

Rumohr:2001:ESR

[RKJ01] H. Rumohr, I. Karakassis, and J. N. Jensen. Estimating species richness, abundance and diversity with 70 macrobenthic replicates in the Western Baltic Sea. Marine Ecology Progress Series, 214:103–110, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p103.pdf.

Ross:2007:ITI

[RKLK07] D. J. Ross, M. J. Keough, A. R. Longmore, and N. A. Knott. Impacts of two introduced suspension feeders in Port Phillip Bay, Australia. Marine Ecology Progress Series, 340:41–53, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p041.pdf.

Richman:2003:ECS

[RL03] S. E. Richman and J. R. Lovvorn. Effects of clam species dominance on nutrient and energy acquisition by spectacled eiders in the Bering Sea. Marine Ecology Progress Series, 261:283–297, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p283.pdf.

Riisgaard:2005:WFA

[RL05a] H. U. Riisgård and P. S. Larsen. Water flow analysis and particle capture in ciliary suspension-feeding scallops (Pectinidae). Marine Ecology Progress Series, 303:177–193, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p177.pdf.

Rotjan:2005:SPP

[RL05b] R. D. Rotjan and S. M. Lewis. Selective predation by parrotfishes on the reef coral Porites astreoides. Marine Ecology Progress Series, 305:193–201, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p193.pdf.

Riisgaard:2007:VSC

[RL07] H. U. Riisgård and P. S. Larsen. Viscosity of seawater controls beat frequency of water-pumping cilia and filtration rate of mussels Mytilus edulis. Marine Ecology Progress Series, 343:141–150, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m343p141.pdf.

Rotjan:2008:ICP

[RL08] R. D. Rotjan and S. M. Lewis. Impact of coral predators on tropical reefs. Marine Ecology Progress Series, 367: 73–91, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p073.pdf.

Ragnarsson:2009:THA

[RL09a] S. Á. Ragnarsson and M. Lindegarth. Testing hypotheses about temporary and persistent effects of otter trawling on infauna: changes in diversity rather than abundance.

Marine Ecology Progress Series, 385:51–64, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p051.pdf.

Rinke:2009:PAT

[RL09b] C. Rinke and R. W. Lee. Pathways, activities and thermal stability of anaerobic and aerobic enzymes in thermophilic vent paralvinellid worms. Marine Ecology Progress Series, 382: 99–112, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/382/m382p099.pdf.

Raniello:2004:PPI

[RLBB04] R. Raniello, M. Lorenti, C. Brunet, and M. C. Buia. Photosynthetic plasticity of an invasive variety of Caulerpa racemosa in a coastal Mediterranean area: light harvesting capacity and seasonal acclimation. Marine Ecology Progress Series, 271:113–120, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/271/m271p113.pdf.

Rousseau:2002:DSS

[RLDL02] V. Rousseau, A. Leynaert, N. Daoud, and C. Lancelot. Diatom succession, silicification and silicic acid availability in Belgian coastal waters (Southern North Sea). Marine Ecology Progress Series, 236:61–73, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p061.pdf.

Robinet:2003:TEA

[RLFEF03] T. Robinet, R. Lecomte-Finiger, K. Escoubeyrou, and E. Feunteun. Tropical eels Anguilla spp. recruiting to Réunion Island in the Indian Ocean: taxonomy, patterns of recruitment and early life histories. Marine Ecology Progress Series, 259:263–272, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p263.pdf.

Rodriguez-Lanetty:2004:DTS

[RLKW04] M. Rodriguez-Lanetty, D. A. Krupp, and V. M. Weis. Distinct ITS types of Symbiodinium in Clade C correlate with cnidarian/dinoflagellate specificity during onset of symbiosis.

Marine Ecology Progress Series, 275:97–102, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p097.pdf.

Richner:2000:FSM

[RM00] Heinz Richner and Manfred Milinski. On the functional significance of masking behaviour in sea urchins — an experiment with Paracentrotus lividus. Marine Ecology Progress Series, 205:307–308, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/205/m205p307.pdf.

Raventos:2005:EPL

[RM05] N. Raventos and E. Macpherson. Effect of pelagic larval growth and size-at-hatching on post-settlement survivorship in two temperate labrid fish of the genus Symphodus. Marine Ecology Progress Series, 285:205–211, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p205.pdf.

Rogers:2006:PPS

[RM06] C. S. Rogers and J. Miller. Permanent ‘phase shifts’ or reversible declines in coral cover? Lack of recovery of two coral reefs in St. John, US Virgin Islands. Marine Ecology Progress Series, 306:103–114, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/306/m306p103.pdf.

Rooper:2009:PPA

[RM09a] C. N. Rooper and M. H. Martin. Predicting presence and abundance of demersal fishes: a model application to shortspine thornyhead Sebastolobus alascanus. Marine Ecology Progress Series, 379:253–266, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/379/m379p253.pdf.

Rozas:2009:UNG

[RM09b] L. P. Rozas and T. J. Minello. Using nekton growth as a metric for assessing habitat restoration by marsh terracing. Marine Ecology Progress Series, 394:179–193, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m394p179.pdf.

Ramos:2002:ILL

[RMA^{+}02] J. A. Ramos, A. M. Maul, V. Ayrton, I. Bullock, J. Hunter, J. Bowler, G. Castle, R. Mileto, and C. Pacheco. Influence of local and large-scale weather events and timing of breeding on tropical roseate tern reproductive parameters. Marine Ecology Progress Series, 243:271–279, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/243/m243p271.pdf.

Rufino:2004:SSN

[RMAY04] M. M. Rufino, F. Maynou, P. Abelló, and A. B. Yule. Smallscale non-linear geostatistical analysis of Liocarcinus depurator (Crustacea: Brachyura) abundance and size structure in a western Mediterranean population. Marine Ecology Progress Series, 276:223–235, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/276/m276p223.pdf.

Ramos:2005:BMD

[RMC^{+}05] A. G. Ramos, A. Martel, G. A. Codd, E. Soler, J. Coca, A. Redondo, L. F. Morrison, J. S. Metcalf, A. Ojeda, S. Suárez, and M. Petit. Bloom of the marine diazotrophic cyanobacterium Trichodesmium erythraeum in the Northwest African upwelling. Marine Ecology Progress Series, 301:303–305, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/301/m301p303.pdf.

Ruiz:2007:DFA

[RMC^{+}07] O. Ruiz, G. R. Medina, R. G. Cohen, F. Amat, and J. C. Navarro. Diversity of the fatty acid composition of Artemia spp. cysts from Argentinean populations. Marine Ecology Progress Series, 335:155–165, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/335/m335p155.pdf.

Riedel:2007:ENE

[RMGL07] A. Riedel, C. Michel, M. Gosselin, and B. LeBlanc. Enrichment of nutrients, exopolymeric substances and microorganisms in newly formed sea ice on the Mackenzie shelf.

Marine Ecology Progress Series, 342:55–67, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p055.pdf.

Reglero:2007:DOF

[RMH07] P. Reglero, H. Mosegaard, and H. H. Hinrichsen. Date of otolith first increment formation in Baltic sprat Sprattus sprattus and its relation to onset of maturity. Marine Ecology Progress Series, 330:223–233, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/330/m330p223.pdf.

Reynaud:2009:ELF

[RMH^{+}09] S. Reynaud, P. Martinez, F. Houlbrèque, I. Billy, D. Allemand, and C. Ferrier-Pagès. Effect of light and feeding on the nitrogen isotopic composition of a zooxanthellate coral: role of nitrogen recycling. Marine Ecology Progress Series, 392:103–110, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p103.pdf.

Riera:2000:UEO

[RMKR00] P. Riera, P. A. Montagna, R. D. Kalke, and P. Richard. Utilization of estuarine organic matter during growth and migration by juvenile brown shrimp Penaeus aztecus in a South Texas estuary. Marine Ecology Progress Series, 199: 205–216, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/199/m199p205.pdf.

Reitzel:2004:RBS

[RMM04] A. M. Reitzel, B. G. Miner, and L. R. McEdward. Relationships between spawning date and larval development time for benthic marine invertebrates: a modeling approach. Marine Ecology Progress Series, 280:13–23, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p013.pdf.

Ryan:2008:PTL

[RMPC08] J. P. Ryan, M. A. McManus, J. D. Paduan, and F. P. Chavez. Phytoplankton thin layers caused by shear in frontal zones of a coastal upwelling system. Marine Ecology Progress Series, 354:21–34, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p021.pdf.

Rodolfo-Metalpa:2008:TDR

[RMRAFP08] R. Rodolfo-Metalpa, S. Reynaud, D. Allemand, and C. Ferrier-Pagès. Temporal and depth responses of two temperate corals, Cladocora caespitosa and Oculina patagonica, from the North Mediterranean Sea. Marine Ecology Progress Series, 369:103–114, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/369/m369p103.pdf.

Riffaut:2005:PGC

[RMT^{+}05] L. Riffaut, K. D. McCoy, C. Tirard, V. L. Friesen, and T. Boulinier. Population genetics of the common guillemot, Uria aalge, in the North Atlantic: geographic impact of oil spills. Marine Ecology Progress Series, 291:263– 273, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p263.pdf.

Raateoja:2009:EWC

[RMWO09] M. Raateoja, B. G. Mitchell, H. Wang, and E. Olivo. Effect of water column light gradient on phytoplankton fluorescence transients. Marine Ecology Progress Series, 376: 85–101, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p085.pdf.

Rozas:2007:NPL

[RMZC07] L. P. Rozas, T. J. Minello, R. J. Zimmerman, and P. Caldwell. Nekton populations, long-term wetland loss, and the effect of recent habitat restoration in Galveston Bay, Texas, USA. Marine Ecology Progress Series, 344:119–130, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/344/
m344p119.pdf.

Riisgaard:2006:FMP

[RN06] H. U. Riisgård and C. Nielsen. Feeding mechanism of the polychaete Sabellaria alveolata revisited: comment on Dubois et al. (2005). Marine Ecology Progress Series, 328:295– 305, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p295.pdf. See [DBCB05].

Rochelle-Newall:2004:CDO

[RNDF^{+}04] E. Rochelle-Newall, B. Delille, M. Frankignoulle, J. P. Gattuso, S. Jacquet, U. Riebesell, A. Terbruggen, and I. Zondervan. Chromophoric dissolved organic matter in experimental mesocosms maintained under different pCO_{2} levels. Marine Ecology Progress Series, 272:25–31, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/272/
m272p025.pdf.

Riisgaard:2000:DCC

[RNL00] H. U. Riisgård, C. Nielsen, and P. S. Larsen. Downstream collecting in ciliary suspension feeders: the catch-up principle. Marine Ecology Progress Series, 207:33–51, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p033.pdf.

Robbins:2006:ETU

[Rob06] W. D. Robbins. Evaluation of two underwater biopsy probes for in situ collection of shark tissue samples. Marine Ecology Progress Series, 310:213–217, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/310/m310p213.pdf.

Rzeznik-Orignac:2008:SIA

[ROBFR08] J. Rzeznik-Orignac, G. Boucher, D. Fichet, and P. Richard. Stable isotope analysis of food source and trophic position of intertidal nematodes and copepods. Marine Ecology Progress Series, 359:145–150, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/359/m359p145.pdf.

Rey:2009:CMS

[ROCB09] J. R. Rey, S. M. O’Connell, D. B. Carlson, and R. E. Brockmeyer, Jr. Characteristics of mangrove swamps managed for mosquito control in eastern Florida, USA: a re-examination.

Marine Ecology Progress Series, 389:295–300, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m389p295.pdf.

Rodriguez:2003:CDK

[Rod03] S. R. Rodríguez. Consumption of drift kelp by intertidal populations of the sea urchin Tetrapygus niger on the central Chilean coast: possible consequences at different ecological levels. Marine Ecology Progress Series, 251:141–151, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/251/m251p141.pdf.

Rolff:2000:SVS

[Rol00] C. Rolff. Seasonal variation in δ ^{13}C and δ ^{15}N of sizefractionated plankton at a coastal station in the northern Baltic proper. Marine Ecology Progress Series, 203:47–65, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/203/m203p047.pdf.

Ross:2001:LSS

[Ros01] P. M. Ross. Larval supply, settlement and survival of barnacles in a temperate mangrove forest. Marine Ecology Progress Series, 215:237–249, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p237.pdf.

Roy:2002:BBS

[Roy02] K. Roy. Bathymetry and body size in marine gastropods: a shallow water perspective. Marine Ecology Progress Series, 237:143–149, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p143.pdf.

Raniello:2002:ADS

[RP02] R. Raniello and G. Procaccini. Ancient DNA in the seagrass Posidonia oceanica. Marine Ecology Progress Series, 227:269–273, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/227/m227p269.pdf.

Ritter:2006:SVS

[RP06] A. F. Ritter and R. K. Preisler. Spatial variation in structure of an intertidal fish assemblage reflects daily settlement patterns. Marine Ecology Progress Series, 317:211– 223, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/317/m317p211.pdf.

Rivalan:2006:ECF

[RPC^{+}06] P. Rivalan, R. Pradel, R. Choquet, M. Girondot, and A. C. Prévot-Julliard. Estimating clutch frequency in the sea turtle Dermochelys coriacea using stopover duration. Marine Ecology Progress Series, 317:285–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/317/
m317p285.pdf.

Reay:2001:RLT

[RPN^{+}01] D. S. Reay, J. Priddle, D. B. Nedwell, M. J. Whitehouse, J. C. Ellis-Evans, C. Deubert, and D. P. Connelly. Regulation by low temperature of phytoplankton growth and nutrient uptake in the Southern Ocean. Marine Ecology Progress Series, 219:51–64, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p051.pdf.

Risgaard-Petersen:2000:NCT

[RPO00] N. Risgaard-Petersen and L. D. M. Ottosen. Nitrogen cycling in two temperate Zostera marina beds: seasonal variation. Marine Ecology Progress Series, 198:93–107, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p093.pdf.

Robbart:2004:PRD

[RPS^{+}04] M. L. Robbart, P. Peckol, S. P. Scordilis, H. A. Curran, and J. Brown-Saracino. Population recovery and differential heat shock protein expression for the corals Agaricia agaricites and A. tenuifolia in Belize. Marine Ecology Progress Series, 283:151–160, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/283/m283p151.pdf.

Rainbow:2006:TTTb

[RPS^{+}06a] P. S. Rainbow, L. Poirier, B. D. Smith, K. V. Brix, and S. N. Luoma. Trophic transfer of trace metals from the polychaete worm Nereis diversicolor to the polychaete N. virens and the decapod crustacean Palaemonetes varians. Marine Ecology Progress Series, 321:167–181, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/321/m321p167.pdf.

Rainbow:2006:TTTa

[RPS^{+}06b] P. S. Rainbow, L. Poirier, B. D. Smith, K. V. Brix, and S. N. Luoma. Trophic transfer of trace metals: subcellular compartmentalization in a polychaete and assimilation by a decapod crustacean. Marine Ecology Progress Series, 308: 91–100, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/308/m308p091.pdf.

Riisgaard:2000:TSP

[RQF^{+}00] Hans Ulrik Riisgård, Gerry Quinn, Everett Fee, Poul Scheel Larsen, Sandra E. Shumway, Josep-Maria Gili, Thomas Kiørboe, Lars Hagerman, Peter Beninger, Alan Tessier, Carlos Duarte, John Raven, Jack J. Middelburg, Michael Lesser, Antoine Grémare, Jon Cole, Ole Naesbye Larsen, Jan J. Beukema, Karsten Reise, Don Canfield, and Otto Kinne. Theme section: The peer-review system: time for re-assessment? Marine Ecology Progress Series, 192:305– 313, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/riisgard.pdf.

Ruiz:2001:ELS

[RR01] J. M. Ruiz and J. Romero. Effects of in situ experimental shading on the Mediterranean seagrass Posidonia oceanica. Marine Ecology Progress Series, 215:107–120, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
215/m215p107.pdf.

Renegar:2005:ENE

[RR05] D. A. Renegar and B. M. Riegl. Effect of nutrient enrichment and elevated CO_{2} partial pressure on growth rate of Atlantic scleractinian coral Acropora cervicornis. Marine Ecology Progress Series, 293:69–76, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/293/m293p069.pdf.

Rideout:2006:SRA

[RR06] R. M. Rideout and G. A. Rose. Suppression of reproduction in Atlantic cod Gadus morhua. Marine Ecology Progress Series, 320:267–277, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/320/m320p267.pdf.

Ridgway:2008:GCP

[RRDHG08] T. Ridgway, C. Riginos, J. Davis, and O. Hoegh-Guldberg. Genetic connectivity patterns of Pocillopora verrucosa in southern African Marine Protected Areas. Marine Ecology Progress Series, 354:161–168, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/354/m354p161.pdf.

Romero:2009:EHC

[RRH09] O. E. Romero, T. Rixen, and B. Herunadi. Effects of hydrographic and climatic forcing on diatom production and export in the tropical southeastern Indian Ocean. Marine Ecology Progress Series, 384:69–82, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/384/m384p069.pdf.

Rey-Rassat:2002:EPR

[RRIH^{+}02a] C. Rey-Rassat, X. Irigoien, R. Harris, R. Head, and F. Carlotti. Egg production rates of Calanus helgolandicus females reared in the laboratory: variability due to present and past feeding conditions. Marine Ecology Progress Series, 238:139–151, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/238/m238p139.pdf.

Rey-Rassat:2002:GDC

[RRIH^{+}02b] C. Rey-Rassat, X. Irigoien, R. Harris, R. Head, and F. Carlotti. Growth and development of Calanus helgolandicus reared in the laboratory. Marine Ecology Progress Series, 238:125–138, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/238/m238p125.pdf.

Rey-Rassat:2002:ECG

[RRIHC02] C. Rey-Rassat, X. Irigoien, R. Harris, and F. Carlotti. Energetic cost of gonad development in Calanus finmarchicus and C. helgolandicus. Marine Ecology Progress Series, 238: 301–306, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p301.pdf.

Roth:2008:RIH

[RRRM08] B. M. Roth, K. A. Rose, L. P. Rozas, and T. J. Minello. Relative influence of habitat fragmentation and inundation on brown shrimp Farfantepenaeus aztecus production in northern Gulf of Mexico salt marshes. Marine Ecology Progress Series, 359:185–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p185.pdf.

Ren:2000:FDF

[RRS00] J. S. Ren, A. H. Ross, and D. R. Schiel. Functional descriptions of feeding and energetics of the Pacific oyster Crassostrea gigas in New Zealand. Marine Ecology Progress Series, 208:119–130, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p119.pdf.

Reigstad:2005:FCF

[RRS05] M. Reigstad, C. W. Riser, and C. Svensen. Fate of copepod faecal pellets and the role of Oithona spp. Marine Ecology Progress Series, 304:265–270, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/304/m304p265.pdf.

Ramos:2009:FEY

[RRS^{+}09] R. Ramos, F. Ramírez, C. Sanpera, L. Jover, and X. Ruiz. Feeding ecology of yellow-legged gulls Larus michahellis in the western Mediterranean: a comparative assessment using conventional and isotopic methods. Marine Ecology Progress Series, 377:289–297, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/377/m377p289.pdf.

Rissik:2000:EFP

[RS00] D. Rissik and I. M. Suthers. Enhanced feeding by pelagic juvenile myctophid fishes within a region of island-induced flow disturbance in the Coral Sea. Marine Ecology Progress Series, 203:263–273, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/203/m203p263.pdf.

Ralph:2002:IWD

[RS02] P. J. Ralph and F. T. Short. Impact of the wasting disease pathogen, Labyrinthula zosterae, on the photobiology of eelgrass Zostera marina. Marine Ecology Progress Series, 226:265–271, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/226/m226p265.pdf.

Ramey:2003:SPS

[RS03] P. A. Ramey and P. V. R. Snelgrove. Spatial patterns in sedimentary macrofaunal communities on the south coast of Newfoundland in relation to surface oceanography and sediment characteristics. Marine Ecology Progress Series, 262: 215–227, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p215.pdf.

Rule:2005:SVR

[RS05] M. J. Rule and S. D. A. Smith. Spatial variation in the recruitment of benthic assemblages to artificial substrata. Marine Ecology Progress Series, 290:67–78, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p067.pdf.

Rilov:2006:TLA

[RS06] G. Rilov and D. R. Schiel. Trophic linkages across seascapes: subtidal predators limit effective mussel recruitment in rocky intertidal communities. Marine Ecology Progress Series, 327: 83–93, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p083.pdf.

Rivers:2007:EGC

[RS07a] D. O. Rivers and F. T. Short. Effect of grazing by Canada geese Branta canadensis on an intertidal eelgrass Zostera marina meadow. Marine Ecology Progress Series, 333:271– 279, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p271.pdf.

Ross:2007:PMC

[RS07b] O. N. Ross and J. Sharples. Phytoplankton motility and the competition for nutrients in the thermocline. Marine Ecology Progress Series, 347:21–38, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/347/m347p021.pdf.

Russ:2005:IVM

[RSA05] G. R. Russ, B. Stockwell, and A. C. Alcala. Inferring versus measuring rates of recovery in no-take marine reserves. Marine Ecology Progress Series, 292:1–12, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rohwer:2002:DDC

[RSAK02] F. Rohwer, V. Seguritan, F. Azam, and N. Knowlton. Diversity and distribution of coral-associated bacteria. Marine Ecology Progress Series, 243:1–10, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/243/m243p001.pdf.

Reyier:2008:REF

[RSC08] E. A. Reyier, J. M. Shenker, and D. Christian. Role of an estuarine fisheries reserve in the production and export of ichthyoplankton. Marine Ecology Progress Series, 359:249– 260, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/359/m359p249.pdf.

Rooker:2008:ETA

[RSD^{+}08] J. R. Rooker, D. H. Secor, G. DeMetrio, A. J. Kaufman, A. Belmonte Ríos, and V. Ticina. Evidence of trans-Atlantic movement and natal homing of bluefin tuna from stable isotopes in otoliths. Marine Ecology Progress Series, 368:231– 239, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m368p231.pdf.

Rex:2001:DDS

[RSE01] M. A. Rex, C. T. Stuart, and R. J. Etter. Do deep-sea nematodes show a positive latitudinal gradient of species diversity? The potential role of depth. Marine Ecology Progress Series, 210:297–298, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/210/m210p297.pdf.

Ryer:2009:ESU

[RSIS09] C. H. Ryer, A. W. Stoner, P. J. Iseri, and M. L. Spencer. Effects of simulated underwater vehicle lighting on fish behavior. Marine Ecology Progress Series, 391:97–106, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/391/
m391p097.pdf.

Raateoja:2004:BOM

[RSK04] M. Raateoja, J. Seppälä, and H. Kuosa. Bio-optical modelling of primary production in the SW Finnish coastal zone, Baltic Sea: fast repetition rate fluorometry in case 2 waters. Marine Ecology Progress Series, 267:9–26, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p009.pdf.

Rainbow:2009:BMP

[RSL09a] P. S. Rainbow, B. D. Smith, and S. N. Luoma. Biodynamic modelling and the prediction of Ag, Cd and Zn accumulation from solution and sediment by the polychaete Nereis diversicolor. Marine Ecology Progress Series, 390: 145–155, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/390/m390p145.pdf.

Rainbow:2009:DTM

[RSL09b] P. S. Rainbow, B. D. Smith, and S. N. Luoma. Differences in trace metal bioaccumulation kinetics among populations of the polychaete Nereis diversicolor from metalcontaminated estuaries. Marine Ecology Progress Series, 376: 173–184, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p173.pdf.

Riera:2002:VCO

[RSN02] P. Riera, L. J. Stal, and J. Nieuwenhuize. δ ^{13}C versus δ ^{15}N of co-occurring molluscs within a community dominated by Crassostrea gigas and Crepidula fornicata (Oosterschelde, The Netherlands). Marine Ecology Progress Series, 240:291–295, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p291.pdf.

Risk:2009:TRS

[RSNG09] M. J. Risk, O. A. Sherwood, R. Nairn, and C. Gibbons. Tracking the record of sewage discharge off Jeddah, Saudi Arabia, since 1950, using stable isotope records from antipatharians. Marine Ecology Progress Series, 397:219–226, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p219.pdf.

Redden:2002:EAD

[RSR02] A. M. Redden, B. G. Sanderson, and D. Rissik. Extending the analysis of the dilution method to obtain the phytoplankton concentration at which microzooplankton grazing becomes saturated. Marine Ecology Progress Series, 226: 27–33, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p027.pdf.

Ryer:2007:PLF

[RSSA07] C. H. Ryer, A. W. Stoner, M. L. Spencer, and A. A. Abookire. Presence of larger flatfish modifies habitat preference by age-0 northern rock sole Lepidopsetta polyxystra. Marine Ecology Progress Series, 342:227–238, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p227.pdf.

Ryer:2004:BMU

[RST04] C. H. Ryer, A. W. Stoner, and R. H. Titgen. Behavioral mechanisms underlying the refuge value of benthic habitat structure for two flatfishes with differing anti-predator strategies. Marine Ecology Progress Series, 268:231–243, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p231.pdf.

Rosenfeld:2006:IBE

[RSY^{+}06] M. Rosenfeld, A. Shemesh, R. Yam, K. Sakai, and Y. Loya. Impact of the 1998 bleaching event on δ ^{18}O records of Okinawa corals. Marine Ecology Progress Series, 314:127– 133, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p127.pdf.

Rooker:2001:DNB

[RSZI01] J. R. Rooker, D. H. Secor, V. S. Zdanowicz, and T. Itoh. Discrimination of northern bluefin tuna from nursery areas in the Pacific Ocean using otolith chemistry. Marine Ecology Progress Series, 218:275–282, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/218/m218p275.pdf.

Roa:2000:CSG

[RT00a] R. Roa and F. Tapia. Cohorts in space: geostatistical mapping of the age structure of the squat lobster Pleuroncodes monodon population off central Chile. Marine Ecology Progress Series, 196:239–251, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/196/m196p239.pdf.

Robinson:2000:SVC

[RT00b] M. Robinson and O. Tully. Seasonal variation in community structure and recruitment of benthic decapods in a subtidal cobble habitat. Marine Ecology Progress Series, 206: 181–191, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/206/m206p181.pdf.

Robinson:2000:SVD

[RT00c] M. Robinson and O. Tully. Spatial variability in decapod community structure and recruitment in sub-tidal habitats. Marine Ecology Progress Series, 194:133–141, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p133.pdf.

Rutherford:2005:MPS

[RT05] L. D. Rutherford, Jr. and E. V. Thuesen. Metabolic performance and survival of medusae in estuarine hypoxia. Marine Ecology Progress Series, 294:189–200, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p189.pdf.

Riemann:2006:LBJ

[RTB06] L. Riemann, J. Titelman, and U. Båmstedt. Links between jellyfish and microbes in a jellyfish dominated fjord. Marine Ecology Progress Series, 325:29–42, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p029.pdf.

Rooker:2006:TES

[RTH06] J. R. Rooker, J. P. Turner, and S. A. Holt. Trophic ecology of sargassum-associated fishes in the Gulf of Mexico determined from stable isotopes and fatty acids. Marine Ecology Progress Series, 313:249–259, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/313/m313p249.pdf.

Rideout:2005:EES

[RTL05] R. M. Rideout, E. A. Trippel, and M. K. Litvak. Effects of egg size, food supply and spawning time on early life history success of haddock Melanogrammus aeglefinus. Marine Ecology Progress Series, 285:169–180, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/285/m285p169.pdf.

Rey:2007:EOC

[RTPS07] A. Raya Rey, P. Trathan, K. Pütz, and A. Schiavini. Effect of oceanographic conditions on the winter movements of rockhopper penguins Eudyptes chrysocome chrysocome from Staten Island, Argentina. Marine Ecology Progress Series, 330:285–295, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/330/m330p285.pdf.

Rossi:2002:SSD

[RU02] F. Rossi and A. J. Underwood. Small-scale disturbance and increased nutrients as influences on intertidal macrobenthic assemblages: experimental burial of wrack in different intertidal environments. Marine Ecology Progress Series, 241: 29–39, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/241/m241p029.pdf.

Rueda:2001:SDF

[Rue01] M. Rueda. Spatial distribution of fish species in a tropical estuarine lagoon: a geostatistical appraisal. Marine Ecology Progress Series, 222:217–226, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/222/m222p217.pdf.

Ruesink:2007:BRF

[Rue07] J. L. Ruesink. Biotic resistance and facilitation of a non-native oyster on rocky shores. Marine Ecology Progress Series, 331: 1–9, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ruiz:2004:OSV

[Rui04] J. M. Ruiz. Oil spills versus shifting baselines. Marine Ecology Progress Series, 282:307–309, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/282/m282p307.pdf.

Richardson:2005:CCP

[RV05] L. L. Richardson and J. D. Voss. Changes in a coral population on reefs of the northern Florida Keys following a coral disease epizootic. Marine Ecology Progress Series, 297: 147–156, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/297/m297p147.pdf.

Roleda:2004:PUE

[RvdPHW04] M. Y. Roleda, W. H. van de Poll, D. Hanelt, and C. Wiencke. PAR and UVBR effects on photosynthesis, viability, growth and DNA in different life stages of coexisting gigartinales: implications for recruitment and zonation pattern. Marine Ecology Progress Series, 281:37–50, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p037.pdf.

Richardson:2003:SES

[RVMI^{+}03] A. J. Richardson, H. M. Verheye, B. A. Mitchell-Innes, J. L. Fowler, and J. G. Field. Seasonal and event-scale variation in growth of Calanus agulhensis (Copepoda) in the Benguela upwelling system and implications for spawning of sardine Sardinops sagax. Marine Ecology Progress Series, 254:239– 251, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p239.pdf.

Rainbow:2001:CAC

[RW01] P. S. Rainbow and W. X. Wang. Comparative assimilation of Cd, Cr, Se, and Zn by the barnacle Elminius modestus from phytoplankton and zooplankton diets. Marine Ecology Progress Series, 218:239–248, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/218/m218p239.pdf.

Raymond:2003:PSS

[RW03] B. Raymond and E. J. Woehler. Predicting seabirds at sea in the Southern Indian Ocean. Marine Ecology Progress Series, 263:275–285, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p275.pdf.

Reich:2006:IAF

[RW06a] K. J. Reich and G. A. J. Worthy. An isotopic assessment of the feeding habits of free-ranging manatees. Marine Ecology Progress Series, 322:303–309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/322/m322p303.pdf.

Rutger:2006:EFI

[RW06b] S. M. Rutger and S. R. Wing. Effects of freshwater input on shallow-water infaunal communities in Doubtful Sound, New Zealand. Marine Ecology Progress Series, 314:35–47, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/314/m314p035.pdf.

Ruttenberg:2006:SVC

[RW06c] B. I. Ruttenberg and R. R. Warner. Spatial variation in the chemical composition of natal otoliths from a reef fish in the Galápagos Island. Marine Ecology Progress Series, 328:225–236, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p225.pdf.

Raymond:2007:BRT

[RW07] E. H. Raymond and E. A. Widder. Behavioral responses of two deep-sea fish species to red, far-red, and white light. Marine Ecology Progress Series, 350:291–298, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m350p291.pdf.

Rodgers:2008:SSM

[RW08] K. L. Rodgers and S. R. Wing. Spatial structure and movement of blue cod Parapercis colias in Doubtful Sound, New Zealand, inferred from δ ^{13}C and δ ^{15}N. Marine Ecology Progress Series, 359:239–248, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/359/m359p239.pdf.

Raimondi:2002:CDB

[RWA^{+}02] P. T. Raimondi, C. M. Wilson, R. F. Ambrose, J. M. Engle, and T. E. Minchinton. Continued declines of black abalone along the coast of California: are mass mortalities related to El Niño events? Marine Ecology Progress Series, 242:143–152, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p143.pdf.

Renaud:2009:CSP

[RWB^{+}09] P. E. Renaud, T. J. Webb, A. Bjørgesæter, I. Karakassis, M. Kedra, M. A. Kendall, C. Labrune, N. Lampadariou, P. J. Somerfield, M. Włodarska-Kowalczuk, E. Vanden Berghe, S. Claus, I. F. Aleffi, J. M. Amouroux, K. H. Bryne, S. J. Cochrane, S. Dahle, S. Degraer, S. G. Denisenko, T. Deprez, C. Dounas, D. Fleischer, J. Gil, A. Grémare, U. Janas, A. S. Y. Mackie, R. Palerud, H. Rumohr, R. Sardá, J. Speybroeck, S. Taboada, G. Van Hoey, J. M. Wesawski, P. Whomersley, and M. L. Zettler. Continental-scale patterns in benthic invertebrate diversity: insights from the MacroBen database. Marine Ecology Progress Series, 382:239–252, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m382p239.pdf.

Romano:2003:IEI

[RWBS03] C. Romano, J. Widdows, M. D. Brinsley, and F. J. Staff. Impact of Enteromorpha intestinalis mats on near-bed currents and sediment dynamics: flume studies. Marine Ecology Progress Series, 256:63–74, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p063.pdf.

Rendell:2004:SWH

[RWE04] L. Rendell, H. Whitehead, and R. Escribano. Sperm whale habitat use and foraging success off northern Chile: Evidence of ecological links between coastal and pelagic systems. Marine Ecology Progress Series, 275:289–295, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p289.pdf.

Rindorf:2000:ECS

[RWH00] A. Rindorf, S. Wanless, and M. P. Harris. Effects of changes in sandeel availability on the reproductive output of seabirds. Marine Ecology Progress Series, 202:241–252, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p241.pdf.

Roast:2002:DSB

[RWJ02] S. D. Roast, J. Widdows, and M. B. Jones. Distribution and swimming behaviour of Neomysis integer (Peracarida: Mysidacea) in response to gradients of dissolved oxygen following exposure to cadmium at environmental concentrations. Marine Ecology Progress Series, 237:185–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p185.pdf.

Ritson-Williams:2007:MBI

[RWP07] R. Ritson-Williams and V. J. Paul. Marine benthic invertebrates use multimodal cues for defense against reef fish. Marine Ecology Progress Series, 340:29–39, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p029.pdf.

Roast:2004:SBI

[RWPJ04] S. D. Roast, J. Widdows, N. Pope, and M. B. Jones. Sedimentbiota interactions: mysid feeding activity enhances water turbidity and sediment erodability. Marine Ecology Progress Series, 281:145–154, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/281/m281p145.pdf.

Reigstad:2000:CSV

[RWR^{+}00] M. Reigstad, P. Wassmann, T. Ratkova, E. Arashkevich, A. Pasternak, and S. Øygarden. Comparison of the springtime vertical export of biogenic matter in three northern Norwegian fjords. Marine Ecology Progress Series, 201:73–89, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p073.pdf.

Ritson-Williams:2003:HSF

[RWSP03] R. Ritson-Williams, S. Shjegstad, and V. Paul. Host specificity of four corallivorous Phestilla nudibranchs (Gastropoda: Opisthobranchia). Marine Ecology Progress Series, 255:207– 218, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/255/m255p207.pdf.

Ryer:2002:TSE

[Rye02] C. H. Ryer. Trawl stress and escapee vulnerability to predation in juvenile walleye pollock: Is there an unobserved bycatch of behaviorally impaired escapees? Marine Ecology Progress Series, 232:269–279, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/232/m232p269.pdf.

Ramaiah:2001:TVT

[RYF01] N. Ramaiah, T. Yoshikawa, and K. Furuya. Temporal variations in transparent exopolymer particles (TEP) associated with a diatom spring bloom in a subarctic ria in Japan. Marine Ecology Progress Series, 212:79–88, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p079.pdf.

Rypien:2008:ADU

[Ryp08] K. L. Rypien. African dust is an unlikely source of Aspergillus sydowii, the causative agent of sea fan disease. Marine Ecology Progress Series, 367:125–131, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/367/m367p125.pdf.

Rozas:2000:SSP

[RZ00] L. P. Rozas and R. J. Zimmerman. Small-scale patterns of nekton use among marsh and adjacent shallow nonvegetated areas of the Galveston Bay Estuary, Texas (USA). Marine Ecology Progress Series, 193:217–239, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p217.pdf.

Ruiz-Zarate:2000:RBM

[RZEACGF00] M. A. Ruiz-Zárate, J. Espinoza-Avalos, J. P. Carricart-Ganivet, and D. Fragoso. Relationships between Manicina areolata (Cnidaria: Scleractinia), Thalassia testudinum (Anthophyta) and Neogoniolithon sp. (Rhodophyta). Marine Ecology Progress Series, 206:135–146, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p135.pdf.

Rhyne:2009:TWD

[RZLS09] A. L. Rhyne, D. Zhang, J. Lin, and N. V. Schizas. Not any two will do: DNA divergence and interpopulation reproductive compatibility in the simultaneous hermaphroditic shrimp Lysmata wurdemanni. Marine Ecology Progress Series, 388:185–195, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p185.pdf.

Rooper:2005:UEB

[RZS05] C. N. Rooper, M. Zimmermann, and P. D. Spencer. Using ecologically based relationships to predict distribution of flathead sole Hippoglossoides elassodon in the eastern Bering Sea. Marine Ecology Progress Series, 290:251–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/290/
m290p251.pdf.

Rost:2008:SPF

[RZWG08] B. Rost, I. Zondervan, and D. Wolf-Gladrow. Sensitivity of phytoplankton to future changes in ocean carbonate chemistry: current knowledge, contradictions and research directions. Marine Ecology Progress Series, 373:227–237, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p227.pdf.

Schanz:2003:IHD

[SA03a] A. Schanz and H. Asmus. Impact of hydrodynamics on development and morphology of intertidal seagrasses in the Wadden Sea. Marine Ecology Progress Series, 261:123–134, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2003/261/m261p123.pdf.

Smith:2003:DOD

[SA03b] K. J. Smith and K. W. Able. Dissolved oxygen dynamics in salt marsh pools and its potential impacts on fish assemblages. Marine Ecology Progress Series, 258:223–232, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p223.pdf.

Sornes:2006:CTP

[SA06] T. A. Sørnes and D. L. Aksnes. Concurrent temporal patterns in light absorbance and fish abundance. Marine Ecology Progress Series, 325:181–186, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m325p181.pdf.

Simeone:2002:OCF

[SAB^{+}02] A. Simeone, B. Araya, M. Bernal, E. N. Diebold, K. Grzybowski, M. Michaels, J. A. Teare, R. S. Wallace, and M. J. Willis. Oceanographic and climatic factors influencing breeding and colony attendance patterns of Humboldt penguins Spheniscus humboldti in central Chile. Marine Ecology Progress Series, 227:43–50, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p043.pdf.

Sammarco:2004:ECC

[SAB04] P. W. Sammarco, A. D. Atchison, and G. S. Boland. Expansion of coral communities within the Northern Gulf of Mexico via offshore oil and gas platforms. Marine Ecology Progress Series, 280:129–143, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/280/m280p129.pdf.

Sala:2005:FDB

[SAB^{+}05] M. Montserrat Sala, L. Arin, V. Balagué, J. Felipe, Ò. Guadayol, and D. Vaqué. Functional diversity of bacterioplankton assemblages in western Antarctic seawaters during late spring. Marine Ecology Progress Series, 292:13–21, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p013.pdf.

Shimeta:2003:RBP

[SABK03] J. Shimeta, C. L. Amos, S. E. Beaulieu, and S. L. Katz. Resuspension of benthic protists at subtidal coastal sites with differing sediment composition. Marine Ecology Progress Series, 259:103–115, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/259/m259p103.pdf.

Schories:2000:FGH

[SAC^{+}00] D. Schories, J. Anibal, A. S. Chapman, E. Herre, I. Isaksson, A. I. Lillebø, L. Pihl, K. Reise, M. Sprung, and M. Thiel. Flagging greens: hydrobiid snails as substrata for the development of green algal mats (Enteromorpha spp.) on tidal flats of North Atlantic coasts. Marine Ecology Progress Series, 199:127–136, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p127.pdf.

Stevick:2003:NAH

[SAC^{+}03] P. T. Stevick, J. Allen, P. J. Clapham, N. Friday, S. K. Katona, F. Larsen, J. Lien, D. K. Mattila, P. J. Palsbøll, J. Sigurjónsson, T. D. Smith, N. Øien, and P. S. Hammond. North Atlantic humpback whale abundance and rate of increase four decades after protection from whaling. Marine Ecology Progress Series, 258:263–273, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/258/m258p263.pdf.

Safi:2003:MPT

[Saf03] K. A. Safi. Microalgal populations of three New Zealand coastal locations: forcing functions and benthic-pelagic links. Marine Ecology Progress Series, 259:67–78, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/259/
m259p067.pdf.

Somerfield:2009:AER

[SAF^{+}09] P. J. Somerfield, C. Arvanitidis, S. Faulwetter, G. Chatzigeorgiou, A. Vasileiadou, J. M. Amouroux, N. Anisimova, S. J. Cochrane, J. Craeymeersch, S. Dahle, S. Denisenko, K. Dounas, G. Duineveld, A. Grémare, C. H. R. Heip, M. Herrmann, I. Karakassis, M. Kedra, M. A. Kendall, P. Kingston, L. Kotwicki, C. Labrune, J. Laudien, H. Nevrova, A. Nicolaidou, A. Occhipinti-Ambrogi, R. Palerud, A. Petrov, E. Rachor, N. Revkov, H. Rumohr, R. Sardá, U. Janas, E. Vanden Berghe, and M. Włodarska-Kowalczuk. Assessing evidence for random assembly of marine benthic communities from regional species pools. Marine Ecology Progress Series, 382:279–286, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m382p279.pdf.

Star:2003:GSA

[SAG03] B. Star, S. Apte, and J. P. A. Gardner. Genetic structuring among populations of the greenshell mussel Perna canaliculus revealed by analysis of randomly amplified polymorphic DNA. Marine Ecology Progress Series, 249:171–182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p171.pdf.

Sackett:2008:HDS

[SAG08] D. K. Sackett, K. W. Able, and T. M. Grothues. Habitat dynamics of summer flounder Paralichthys dentatus within a shallow USA estuary, based on multiple approaches using acoustic telemetry. Marine Ecology Progress Series, 364: 199–212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/364/m364p199.pdf.

Swimmer:2005:FCM

[SAH^{+}05] Y. Swimmer, R. Arauz, B. Higgins, L. McNaughton, M. Mc-Cracken, J. Ballestero, and R. Brill. Food color and marine turtle feeding behavior: Can blue bait reduce turtle bycatch in commercial fisheries? Marine Ecology Progress Series, 295:273–278, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/295/m295p273.pdf.

Sharples:2009:SSS

[SAH09] R. J. Sharples, B. Arrizabalaga, and P. S. Hammond. Seals, sandeels and salmon: diet of harbour seals in St. Andrews Bay and the Tay Estuary, southeast Scotland. Marine Ecology Progress Series, 390:265–276, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/390/m390p265.pdf.

Santelices:2003:FTI

[SAHF03] B. Santelices, D. Aedo, M. Hormazábal, and V. Flores. Field testing of inter- and intraspecific coalescence among midintertidal red algae. Marine Ecology Progress Series, 250: 91–103, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p091.pdf.

Sommer:2006:ACC

[SAHK06] F. Sommer, C. Agurto, P. Henriksen, and T. Kiørboe. Astaxanthin in the calanoid copepod Calanus helgolandicus: dynamics of esterification and vertical distribution in the German Bight, North Sea. Marine Ecology Progress Series, 319: 167–173, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p167.pdf.

Sato:2006:VSA

[SAJG06] T. Sato, M. Ashidate, T. Jinbo, and S. Goshima. Variation of sperm allocation with male size and recovery rate of sperm numbers in spiny king crab Paralithodes brevipes. Marine Ecology Progress Series, 312:189–199, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p189.pdf.

Sandberg:2004:PFW

[SAJW04] J. Sandberg, A. Andersson, S. Johansson, and J. Wikner. Pelagic food web structure and carbon budget in the northern Baltic Sea: potential importance of terrigenous carbon. Marine Ecology Progress Series, 268:13–29, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p013.pdf.

Sugeha:2001:IMT

[SAM^{+}01] H. Y. Sugeha, T. Arai, M. J. Miller, D. Limbong, and K. Tsukamoto. Inshore migration of the tropical eels Anguilla spp. recruiting to the Poigar River estuary on north Sulawesi Island. Marine Ecology Progress Series, 221:233–243, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p233.pdf.

Swimmer:2006:DBD

[SAM^{+}06] Y. Swimmer, R. Arauz, M. McCracken, L. McNaughton, J. Ballestero, M. Musyl, K. Bigelow, and R. Brill. Diving behavior and delayed mortality of olive ridley sea turtles Lepidochelys olivacea after their release from longline fishing gear. Marine Ecology Progress Series, 323:253–261, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p253.pdf.

Sebastian:2004:KAP

[SAMN04] M. Sebastián, J. Arístegui, M. F. Montero, and F. X. Niell. Kinetics of alkaline phosphatase activity, and effect of phosphate enrichment: a case study in the NW African upwelling region. Marine Ecology Progress Series, 270:1–13, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/270/
m270p001.pdf.

Sano:2000:SRF

[San00] M. Sano. Stability of reef fish assemblages: responses to coral recovery after catastrophic predation by Acanthaster planci. Marine Ecology Progress Series, 198:121–130, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p121.pdf.

Sukhotin:2002:GML

[SAP02] A. A. Sukhotin, D. Abele, and H. O. Pörtner. Growth, metabolism and lipid peroxidation in Mytilus edulis: age and size effects. Marine Ecology Progress Series, 226:223– 234, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/226/m226p223.pdf.

Sara:2006:HEO

[Sar06] G. Sarà. Hydrodynamic effects on the origin and quality of organic matter for bivalves: an integrated isotopic, biochemical and transplant study. Marine Ecology Progress Series, 328:65–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p065.pdf.

Shepard:2006:DTR

[SAS^{+}06] E. L. C. Shepard, M. Z. Ahmed, E. J. Southall, M. J. Witt, J. D. Metcalfe, and D. W. Sims. Diel and tidal rhythms in diving behaviour of pelagic sharks identified by signal processing of archival tagging data. Marine Ecology Progress Series, 328:205–213, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p205.pdf.

Saunders:2003:SVE

[SASR03] J. E. Saunders, M. J. Attrill, S. M. Shaw, and A. A. Rowden. Spatial variability in the epiphytic algal assemblages of Zostera marina seagrass beds. Marine Ecology Progress Series, 249:107–115, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p107.pdf.

Savoye:2003:DPO

[SAT^{+}03] N. Savoye, A. Aminot, P. Tréguer, M. Fontugne, N. Naulet, and R. Kérouel. Dynamics of particulate organic matter δ ^{15}N and δ ^{13}C during spring phytoplankton blooms in a macrotidal ecosystem (Bay of Seine, France). Marine Ecology Progress Series, 255:27–41, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/255/m255p027.pdf.

Somerfield:2009:MDL

[SAV^{+}09a] P. J. Somerfield, C. Arvanitidis, E. Vanden Berghe, P. Van Avesaath, H. Hummel, and C. H. R. Heip. MarBEF, databases, and the legacy of John Gray. Marine Ecology Progress Series, 382:221–224, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m382p221.pdf.

Somerfield:2009:LSS

[SAV09b] Paul J. Somerfield, Christos Arvanitidis, and Edward Vanden Bergh. Large-scale studies of the European benthos: the MacroBen database. Marine Ecology Progress Series, 382:221–311, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m382_ThemeSection.pdf.

Somerfield:2009:TSL

[SAV^{+}09c] Paul J. Somerfield, Christos Arvanitidis, Edward Vanden Berghe, I. F. Aleffi, J. M. Amouroux, N. Anisimova, W. Appeltans, C. Arvanitidis, G. Bachelet, L. Bigot, A. Bjørgesæter, K. H. Bryne, M. Büntzow, G. Chatzigeorgiou, S. Claus, S. J. Cochrane, M. J. Costello, J. A. Craeymeersch, S. Dahle, S. Degraer, S. G. Denisenko, T. Deprez, C. Dounas, G. Duineveld, C. Emblow, V. Escaravage, M. C. Fabri, S. Faulwetter, D. Fleischer, J. Gil, J. S. Gray, A. Grémare, C. H. R. Heip, P. M. J. Herman, M. Herrmann, H. Hummel, U. Janas, I. Karakassis, M. Kedra, M. A. Kendall, P. Kingston, L. Kotwicki, C. Labrune, N. Lampadariou, J. Laudien, A. S. Y. Mackie, O. Maire, B. Merckx, E. L. Nevrova, A. Nicolaidou, A. Occhipinti-Ambrogi, G. P. Oliver, F. Olsgard, R. Palerud, A. Petrov, E. Rachor, P. E. Renaud, N. K. Revkov, A. Rose, H. Rumohr, R. Sardá, N. Simboura, W. C. H. Sistermans, P. J. Somerfield, J. Speybroeck, S. Taboada, V. Valavanis, J. Vanaverbeke, E. Van den Berghe, G. Van Hoey, K. Vasileiadou, M. Vincx, T. J. Webb, P. Whomersley, W. Willems, M. Włodarska-Kowalczuk, J. M. Weslawski, A. Zenetos, and M. L. Zettler. Theme section: Large-scale studies of the European benthos: the MacroBen database. Marine Ecology Progress Series, 382:221–311, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m382_ThemeSection.pdf.

Sasai:2001:OMS

[SAW^{+}01] S. Sasai, J. Aoyama, S. Watanabe, T. Kaneko, M. J. Miller, and K. Tsukamoto. Occurrence of migrating silver eels Anguilla japonica in the East China Sea. Marine Ecology Progress Series, 212:305–310, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/212/m212p305.pdf.

Sato:2005:EMM

[SAWG05] T. Sato, M. Ashidate, S. Wada, and S. Goshima. Effects of male mating frequency and male size on ejaculate size and reproductive success of female spiny king crab Paralithodes brevipes. Marine Ecology Progress Series, 296:251– 262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/296/m296p251.pdf.

Stewart:2000:BTI

[SB00] B. D. Stewart and J. S. Beukers. Baited technique improves censuses of cryptic fish in complex habitats. Marine Ecology Progress Series, 197:259–272, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/197/m197p259.pdf.

Shears:2003:CTC

[SB03a] N. T. Shears and R. C. Babcock. Continuing trophic cascade effects after 25 years of no-take marine reserve protection. Marine Ecology Progress Series, 246:1–16, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p001.pdf.

Steffani:2003:GRC

[SB03b] C. N. Steffani and G. M. Branch. Growth rate, condition, and shell shape of Mytilus galloprovincialis: responses to wave exposure. Marine Ecology Progress Series, 246:197– 209, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/246/m246p197.pdf.

Shanks:2005:UDC

[SB05a] A. L. Shanks and L. Brink. Upwelling, downwelling, and crossshelf transport of bivalve larvae: test of a hypothesis. Marine Ecology Progress Series, 302:1–12, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Stergiou:2005:TSB

[SB05b] Konstantinos I. Stergiou and Howard I. Browman. Theme section: Bridging the gap between aquatic and terrestrial ecology. Marine Ecology Progress Series, 304:271–307, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m304p271.pdf.

Stachowicz:2006:SDI

[SB06] J. J. Stachowicz and J. E. Byrnes. Species diversity, invasion success, and ecosystem functioning: disentangling the influence of resource competition, facilitation, and extrinsic factors. Marine Ecology Progress Series, 311:251–262, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m311p251.pdf.

Stensland:2007:BCF

[SB07] E. Stensland and P. Berggren. Behavioural changes in female Indo–Pacific bottlenose dolphins in response to boatbased tourism. Marine Ecology Progress Series, 332:225– 234, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/332/m332p225.pdf.

Suntsov:2008:TET

[SB08] A. V. Suntsov and R. D. Brodeur. Trophic ecology of three dominant myctophid species in the northern California Current region. Marine Ecology Progress Series, 373: 81–96, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/373/m373p081.pdf.

Sydeman:2009:MEC

[SB09] W. J. Sydeman and S. J. Bograd. Marine ecosystems, climate and phenology: introduction. Marine Ecology Progress Series, 393:185–188, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m393p185.pdf.

Spear:2001:RST

[SBA01] L. B. Spear, L. T. Ballance, and D. G. Ainley. Response of seabirds to thermal boundaries in the tropical Pacific: the thermocline versus the equatorial front. Marine Ecology Progress Series, 219:275–289, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p275.pdf.

Semesi:2009:SPC

[SBB09] I. S. Semesi, S. Beer, and M. Björk. Seagrass photosynthesis controls rates of calcification and photosynthesis of calcareous macroalgae in a tropical seagrass meadow. Marine Ecology Progress Series, 382:41–47, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/382/m382p041.pdf.

Sommer:2002:GDE

[SBBS^{+}02] U. Sommer, U. G. Berninger, R. Böttger-Schnack, A. Cornils, W. Hagen, T. Hansen, T. Al-Najjar, A. F. Post, S. B. Schnack-Schiel, H. Stibor, D. Stübing, and S. Wickham. Grazing during early spring in the Gulf of Aqaba and the northern Red Sea. Marine Ecology Progress Series, 239:251–261, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/239/
m239p251.pdf.

Santini:2004:BFR

[SBC04] G. Santini, M. T. Burrows, and G. Chelazzi. Bioeconomics of foraging route selection by limpets. Marine Ecology Progress Series, 280:189–198, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/280/m280p189.pdf.

Stolte:2006:SNF

[SBC^{+}06] W. Stolte, M. Balode, P. Carlsson, D. Grzebyk, S. Janson, I. Lips, R. Panosso, C. J. Ward, and E. Granéli. Stimulation of nitrogen-fixing cyanobacteria in a Baltic Sea plankton community by land-derived organic matter or iron addition. Marine Ecology Progress Series, 327:71–82, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p071.pdf.

Sogard:2008:MER

[SBF08] S. M. Sogard, S. A. Berkeley, and R. Fisher. Maternal effects in rockfishes Sebastes spp.: a comparison among species. Marine Ecology Progress Series, 360:227–236, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/360/
m360p227.pdf.

Smale:2008:BCR

[SBFP08] D. A. Smale, D. K. A. Barnes, K. P. P. Fraser, and L. S. Peck. Benthic community response to iceberg scouring at an intensely disturbed shallow water site at Adelaide Island, Antarctica. Marine Ecology Progress Series, 355:85– 94, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/355/m355p085.pdf.

Skerratt:2002:ABA

[SBH^{+}02] J. H. Skerratt, J. P. Bowman, G. Hallegraeff, S. James, and P. D. Nichols. Algicidal bacteria associated with blooms of a toxic dinoflagellate in a temperate Australian estuary. Marine Ecology Progress Series, 244:1–15, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/244/
m244p001.pdf.

Stien:2005:PDS

[SBHE05] A. Stien, P. A. Bjørn, P. A. Heuch, and D. A. Elston. Population dynamics of salmon lice Lepeophtheirus salmonis on Atlantic salmon and sea trout. Marine Ecology Progress Series, 290:263–275, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/290/m290p263.pdf.

Springer:2007:HOP

[SBI07] A. M. Springer, G. V. Byrd, and S. J. Iverson. Hot oceanography: planktivorous seabirds reveal ecosystem responses to warming of the Bering Sea. Marine Ecology Progress Series, 352:289–297, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m352p289.pdf.

Scharf:2002:SDV

[SBJ02] F. S. Scharf, J. A. Buckel, and F. Juanes. Size-dependent vulnerability of juvenile bay anchovy Anchoa mitchilli to bluefish predation: Does large body size always provide a refuge? Marine Ecology Progress Series, 233:241–252, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p241.pdf.

Sheaves:2006:MNE

[SBJ06] M. Sheaves, R. Baker, and R. Johnston. Marine nurseries and effective juvenile habitats: an alternative view. Marine Ecology Progress Series, 318:303–306, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p303.pdf.

Sanders:2000:HMN

[SBL^{+}00] R. W. Sanders, U. G. Berninger, E. L. Lim, P. F. Kemp, and D. A. Caron. Heterotrophic and mixotrophic nanoplankton predation on picoplankton in the Sargasso Sea and on Georges Bank. Marine Ecology Progress Series, 192:103–118, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p103.pdf.

Shenkar:2008:PDC

[SBL08] N. Shenkar, O. Bronstein, and Y. Loya. Population dynamics of a coral reef ascidian in a deteriorating environment. Marine Ecology Progress Series, 367:163–171, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/367/
m367p163.pdf.

Stentiford:2009:SSD

[SBLF09] G. D. Stentiford, J. P. Bignell, B. P. Lyons, and S. W. Feist. Site-specific disease profiles in fish and their use in environmental monitoring. Marine Ecology Progress Series, 381:1–15, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Skold:2002:SVS

[SBM02] M. Sköld, M. F. Barker, and P. V. Mladenov. Spatial variability in sexual and asexual reproduction of the fissiparous seastar Coscinasterias muricata: the role of food and fluctuating temperature. Marine Ecology Progress Series, 233: 143–155, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/233/m233p143.pdf.

Sasso:2006:ETE

[SBMAE06] C. R. Sasso, J. Braun-McNeill, L. Avens, and S. P. Epperly. Effects of transients on estimating survival and population growth in juvenile loggerhead turtles. Marine Ecology Progress Series, 324:287–292, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/324/m324p287.pdf.

Sejr:2009:SIC

[SBR09] M. K. Sejr, M. E. Blicher, and S. Rysgaard. Sea ice cover affects inter-annual and geographic variation in growth of the Arctic cockle Clinocardium ciliatum (Bivalvia) in Greenland. Marine Ecology Progress Series, 389:149–158, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p149.pdf.

Semmens:2004:HNN

[SBSPS04] B. X. Semmens, E. R. Buhle, A. K. Salomon, and C. V. Pattengill-Semmens. A hotspot of non-native marine fishes: evidence for the aquarium trade as an invasion pathway. Marine Ecology Progress Series, 266:239–244, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p239.pdf.

Stoner:2002:GVB

[SBSRW02] D. S. Stoner, R. Ben-Shlomo, B. Rinkevich, and I. L. Weissman. Genetic variability of Botryllus schlosseri invasions to the east and west coasts of the USA. Marine Ecology Progress Series, 243:93–100, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/243/m243p093.pdf.

Scheffers:2005:WBG

[SBvD05] S. R. Scheffers, R. P. M. Bak, and F. C. van Duyl. Why is bacterioplankton growth in coral reef framework cavities enhanced? Marine Ecology Progress Series, 299:89– 99, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p089.pdf.

Serodio:2000:MPP

[SC00a] J. Serôdio and F. Catarino. Modelling the primary productivity of intertidal microphytobenthos: time scales of variability and effects of migratory rhythms. Marine Ecology Progress Series, 192:13–30, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p013.pdf.

Steves:2000:SGM

[SC00b] B. P. Steves and R. K. Cowen. Settlement, growth, and movement of silver hake Merluccius bilinearis in nursery habitat on the New York Bight continental shelf. Marine Ecology Progress Series, 196:279–290, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/196/m196p279.pdf.

Seddon:2001:PRA

[SC01] S. Seddon and A. C. Cheshire. Photosynthetic response of Amphibolis antarctica and Posidonia australis to temperature and desiccation using chlorophyll fluorescence. Marine Ecology Progress Series, 220:119–130, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/220/m220p119.pdf.

Schils:2003:SVS

[SC03] T. Schils and E. Coppejans. Spatial variation in subtidal plant communities around the Socotra Archipelago and their biogeographic affinities within the Indian Ocean. Marine Ecology Progress Series, 251:103–114, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/251/m251p103.pdf.

Shearer:2006:GIC

[SC06a] T. L. Shearer and M. A. Coffroth. Genetic identification of Caribbean scleractinian coral recruits at the Flower Garden Banks and the Florida Keys. Marine Ecology Progress Series, 306:133–142, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p133.pdf.

Sohm:2006:PDT

[SC06b] J. A. Sohm and D. G. Capone. Phosphorus dynamics of the tropical and subtropical north Atlantic: Trichodesmium spp. versus bulk plankton. Marine Ecology Progress Series, 317:21–28, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p021.pdf.

Smoothey:2007:SSV

[SC07] A. F. Smoothey and M. G. Chapman. Small-scale variability in the dispersion of the sea urchin Heliocidaris erythrogramma among boulders. Marine Ecology Progress Series, 340:89–99, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p089.pdf.

Sullivan:2000:SSR

[SCAF00] M. C. Sullivan, R. K. Cowen, K. W. Able, and M. P. Fahay. Spatial scaling of recruitment in four continental shelf fishes. Marine Ecology Progress Series, 207:141–154, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p141.pdf.

Sullivan:2003:EAN

[SCAF03] M. C. Sullivan, R. K. Cowen, K. W. Able, and M. P. Fahay. Effects of anthropogenic and natural disturbance on a recently settled continental shelf flatfish. Marine Ecology Progress Series, 260:237–253, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p237.pdf.

Svensson:2000:NCS

[SCB00] J. M. Svensson, G. M. Carrer, and M. Bocci. Nitrogen cycling in sediments of the Lagoon of Venice, Italy. Marine Ecology Progress Series, 199:1–11, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/199/m199p001.pdf.

Steele:2005:SPI

[SCBA05] L. T. Steele, M. Caldwell, A. Boettcher, and T. Arnold. Seagrass-pathogen interactions: ‘pseudo-induction’ of turtlegrass phenolics near wasting disease lesions. Marine Ecology Progress Series, 303:123–131, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/303/m303p123.pdf.

Soong:2005:PSI

[SCC05] K. Soong, D. Chang, and S. M. Chao. Presence of spawninducing pheromones in two brittle stars (Echinodermata: Ophiuroidea). Marine Ecology Progress Series, 292:195–201, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/292/m292p195.pdf.

Stutes:2006:ENE

[SCC06] A. L. Stutes, J. Cebrian, and A. A. Corcoran. Effects of nutrient enrichment and shading on sediment primary production and metabolism in eutrophic estuaries. Marine Ecology Progress Series, 312:29–43, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/312/m312p029.pdf.

Son:2005:PPY

[SCD^{+}05] S. H. Son, J. Campbell, M. Dowell, S. Yoo, and J. Noh. Primary production in the Yellow Sea determined by ocean color remote sensing. Marine Ecology Progress Series, 303: 91–103, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/303/m303p091.pdf.

Spivak:2009:ECC

[SCD^{+}09] A. C. Spivak, E. A. Canuel, J. E. Duffy, J. G. Douglass, and J. P. Richardson. Epifaunal community composition and nutrient addition alter sediment organic matter composition in a natural eelgrass Zostera marina bed: a field experiment. Marine Ecology Progress Series, 376:55–67, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p055.pdf.

Shipe:2007:LPP

[SCGC07] R. F. Shipe, E. J. Carpenter, S. R. Govil, and D. G. Capone. Limitation of phytoplankton production by Si and N in the western Atlantic Ocean. Marine Ecology Progress Series, 338: 33–45, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p033.pdf.

Svendsen:2007:PHC

[SCH^{+}07] T. C. Svendsen, L. Camus, B. Hargrave, A. Fisk, D. C. G. Muir, and K. Borgå. Polyaromatic hydrocarbons, chlorinated and brominated organic contaminants as tracers of feeding ecology in polar benthic amphipods. Marine Ecology Progress Series, 337:155–164, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/337/m337p155.pdf.

Shen:2009:FHS

[SCIT09] K. N. Shen, C. W. Chang, Y. Iizuka, and W. N. Tzeng. Facultative habitat selection in Pacific tarpon Megalops cyprinoides as revealed by otolith Sr:Ca ratios. Marine Ecology Progress Series, 387:255–263, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/387/m387p255.pdf.

Simmons:2007:LFB

[SCKC07] S. E. Simmons, D. E. Crocker, R. M. Kudela, and D. P. Costa. Linking foraging behaviour of the northern elephant seal with oceanography and bathymetry at mesoscales. Marine Ecology Progress Series, 346:265–275, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/346/m346p265.pdf.

Stafford:2009:ECB

[SCM^{+}09a] K. M. Stafford, J. J. Citta, S. E. Moore, M. A. Daher, and J. E. George. Environmental correlates of blue and fin whale call detections in the North Pacific Ocean from 1997 to 2002. Marine Ecology Progress Series, 395:37–53, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m395p037.pdf.

Steller:2009:CAR

[SCM09b] D. L. Steller and C. Cáceres-Martínez. Coralline algal rhodoliths enhance larval settlement and early growth of the Pacific calico scallop Argopecten ventricosus. Marine Ecology Progress Series, 396:49–60, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/396/m396p049.pdf.

Santangelo:2003:RPS

[SCMB03] G. Santangelo, E. Carletti, E. Maggi, and L. Bramanti. Reproduction and population sexual structure of the overexploited Mediterranean red coral Corallium rubrum. Marine Ecology Progress Series, 248:99–108, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/248/m248p099.pdf.

Smalley:2003:FMD

[SCS03] G. W. Smalley, D. W. Coats, and D. K. Stoecker. Feeding in the mixotrophic dinoflagellate Ceratium furca is influenced by intracellular nutrient concentrations. Marine Ecology Progress Series, 262:137–151, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/262/m262p137.pdf.

Stutes:2007:BMA

[SCS^{+}07] J. Stutes, J. Cebrian, A. L. Stutes, A. Hunter, and A. A. Corcoran. Benthic metabolism across a gradient of anthropogenic impact in three shallow coastal lagoons in NW Florida. Marine Ecology Progress Series, 348:55–70, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p055.pdf.

Shreeve:2009:FEM

[SCT^{+}09] R. S. Shreeve, M. A. Collins, G. A. Tarling, C. E. Main, P. Ward, and N. M. Johnston. Feeding ecology of myctophid fishes in the northern Scotia Sea. Marine Ecology Progress Series, 386:221–236, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/386/m386p221.pdf.

Skilleter:2006:EPD

[SCZ^{+}06] G. A. Skilleter, B. Cameron, Y. Zharikov, D. Boland, and D. P. McPhee. Effects of physical disturbance on infaunal and epifaunal assemblages in subtropical, intertidal seagrass beds. Marine Ecology Progress Series, 308:61–78, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p061.pdf.

Sirois:2000:CPG

[SD00a] P. Sirois and J. J. Dodson. Critical periods and growthdependent survival of larvae of an estuarine fish, the rainbow smelt Osmerus mordax. Marine Ecology Progress Series, 203:233–245, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/203/m203p233.pdf.

Sirois:2000:ITF

[SD00b] P. Sirois and J. J. Dodson. Influence of turbidity, food density and parasites on the ingestion and growth of larval rainbow smelt Osmerus mordax in an estuarine turbidity maximum. Marine Ecology Progress Series, 193:167–179, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p167.pdf.

Shaughnessy:2001:SSP

[SD01a] F. J. Shaughnessy and R. E. DeWreede. Size, survival and the potential for reproduction in transplants of Mazzaella splendens and M. linearis (Rhodophyta). Marine Ecology Progress Series, 222:109–118, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/222/m222p109.pdf.

Skov:2001:SAF

[SD01b] H. Skov and J. Durinck. Seabird attraction to fishing vessels is a local process. Marine Ecology Progress Series, 214: 289–298, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/214/m214p289.pdf.

Sastri:2009:IVC

[SD09] A. R. Sastri and J. F. Dower. Interannual variability in chitobiase-based production rates of the crustacean zooplankton community in the Strait of Georgia. Marine Ecology Progress Series, 388:147–157, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/388/m388p147.pdf.

Sara:2007:EBN

[SDD^{+}07] G. Sarà, J. M. Dean, D. D’Amato, G. Buscaino, A. Oliveri, S. Genovese, S. Ferro, G. Buffa, M. L. Martire, and S. Mazzola. Effect of boat noise on the behaviour of bluefin tuna Thunnus thynnus in the Mediterranean Sea. Marine Ecology Progress Series, 331:243–253, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/331/m331p243.pdf.

Sasal:2000:EAI

[SDFM00] P. Sasal, P. Durand, E. Faliex, and S. Morand. Experimental approach to the importance of parasitism in biological conservation. Marine Ecology Progress Series, 198:293–302, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p293.pdf.

Solow:2001:ETC

[SDH01] A. Solow, C. Davis, and Q. Hu. Estimating the taxonomic composition of a sample when individuals are classified with error. Marine Ecology Progress Series, 216:309–311, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p309.pdf.

Saxby:2003:PRC

[SDHG03] T. Saxby, W. C. Dennison, and O. Hoegh-Guldberg. Photosynthetic responses of the coral Montipora digitata to cold temperature stress. Marine Ecology Progress Series, 248: 85–97, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p085.pdf.

Short:2002:SSM

[SDK^{+}02] F. T. Short, R. C. Davis, B. S. Kopp, C. A. Short, and D. M. Burdick. Site-selection model for optimal transplantation of eelgrass Zostera marina in the northeastern US. Marine Ecology Progress Series, 227:253–267, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p253.pdf.

Suchman:2008:FPP

[SDK^{+}08] C. L. Suchman, E. A. Daly, J. E. Keister, W. T. Peterson, and R. D. Brodeur. Feeding patterns and predation potential of scyphomedusae in a highly productive upwelling region. Marine Ecology Progress Series, 358:161–172, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p161.pdf.

Strong:2006:CMS

[SDM06] J. A. Strong, M. J. Dring, and C. A. Maggs. Colonisation and modification of soft substratum habitats by the invasive macroalga Sargassum muticum. Marine Ecology Progress Series, 321:87–97, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/321/m321p087.pdf.

Swanson:2006:SQN

[SdNH^{+}06] R. L. Swanson, R. de Nys, M. J. Huggett, J. K. Green, and P. D. Steinberg. In situ quantification of a natural settlement cue and recruitment of the Australian sea urchin Holopneustes purpurascens. Marine Ecology Progress Series, 314:

1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Scolardi:2006:FEM

[SDPT06] K. M. Scolardi, K. L. Daly, E. A. Pakhomov, and J. J. Torres. Feeding ecology and metabolism of the Antarctic cydippid ctenophore Callianira antarctica. Marine Ecology Progress Series, 317:111–126, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p111.pdf.

Sponer:2001:LGD

[SDR01] R. Sponer, D. Deheyn, and M. S. Roy. Large genetic distances within a population of Amphipholis squamata (Echinodermata; Ophiuroidea) do not support colour varieties as sibling species. Marine Ecology Progress Series, 219:169–175, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
219/m219p169.pdf.

Soria-Dengg:2001:SMC

[SDRH01] S. Soria-Dengg, R. Reissbrodt, and U. Horstmann. Siderophores in marine coastal waters and their relevance for iron uptake by phytoplankton: experiments with the diatom Phaeodactylum tricornutum. Marine Ecology Progress Series, 220:73–82, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/220/m220p073.pdf.

Saunders:2003:CHR

[SDS^{+}03] P. A. Saunders, D. Deibel, C. J. Stevens, R. B. Rivkin, S. H. Lee, and B. Klein. Copepod herbivory rate in a large Arctic polynya and its relationship to seasonal and spatial variation in copepod and phytoplankton biomass. Marine Ecology Progress Series, 261:183–199, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/261/m261p183.pdf.

Serodio:2001:ULV

[SdSC01] J. Serôdio, J. Marques da Silva, and F. Catarino. Use of in vivo chlorophyll-a fluorescence to quantify short-term variations in the productive biomass of intertidal microphytobenthos. Marine Ecology Progress Series, 218:45–61, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p045.pdf.

Shaked:2006:PPR

[SdV06] Y. Shaked and C. de Vargas. Pelagic photosymbiosis: rDNA assessment of diversity and evolution of dinoflagellate symbionts and planktonic Foraminiferal hosts. Marine Ecology Progress Series, 325:59–71, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p059.pdf.

Schmaljohann:2001:OMZ

[SDW^{+}01] R. Schmaljohann, M. Drews, S. Walter, P. Linke, U. von Rad, and J. F. Imhoff. Oxygen-minimum zone sediments in the northeastern Arabian Sea off Pakistan: a habitat for the bacterium Thioploca. Marine Ecology Progress Series, 211: 27–42, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/211/m211p027.pdf. See erratum [Ano01e].

Somerfield:2007:TDS

[SDW07a] P. J. Somerfield, S. L. Dashfield, and R. M. Warwick. Threedimensional spatial structure: nematodes in a sandy tidal flat. Marine Ecology Progress Series, 336:177–186, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p177.pdf.

Stafford:2007:CSH

[SDW07b] R. Stafford, M. S. Davies, and G. A. Williams. Computer simulations of high shore littorinids predict small-scale spatial and temporal distribution patterns on rocky shores. Marine Ecology Progress Series, 342:151–161, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/342/
m342p151.pdf.

Shin:2004:SPS

[SE04] P. K. S. Shin and K. E. Ellingsen. Spatial patterns of softsediment benthic diversity in subtropical Hong Kong waters. Marine Ecology Progress Series, 276:25–35, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p025.pdf.

Searcy:2007:EIR

[SEH07a] S. P. Searcy, D. B. Eggleston, and J. A. Hare. Environmental influences on the relationship between juvenile and larval growth of Atlantic croaker Micropogonias undulatus. Marine Ecology Progress Series, 349:81–88, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/349/
m349p081.pdf.

Sjotun:2007:GCR

[SEH07b] K. Sjøtun, S. F. Eggereide, and T. Høisæter. Grazercontrolled recruitment of the introduced Sargassum muticum (Phaeophyceae, Fucales) in northern Europe. Marine Ecology Progress Series, 342:127–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p127.pdf.

Savage:2002:ESD

[SEL02] C. Savage, R. Elmgren, and U. Larsson. Effects of sewagederived nutrients on an estuarine macrobenthic community. Marine Ecology Progress Series, 243:67–82, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p067.pdf.

Sgrosso:2001:TDR

[SEM01] S. Sgrosso, F. Esposito, and M. Montresor. Temperature and daylength regulate encystment in calcareous cyst-forming dinoflagellates. Marine Ecology Progress Series, 211:77–87, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/211/m211p077.pdf.

Svensen:2001:CST

[SES01] C. Svensen, J. K. Egge, and J. E. Stiansen. Can silicate and turbulence regulate the vertical flux of biogenic matter? A mesocosm study. Marine Ecology Progress Series, 217: 67–80, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/217/m217p067.pdf.

Salita:2003:FEI

[SESP03] J. T. Salita, W. Ekau, and U. Saint-Paul. Field evidence on the influence of seagrass landscapes on fish abundance in Bolinao, northern Philippines. Marine Ecology Progress Series, 247:183–195, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p183.pdf.

Seuront:2005:HTC

[Seu05] L. Seuront. Hydrodynamic and tidal controls of small-scale phytoplankton patchiness. Marine Ecology Progress Series, 302:93–101, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p093.pdf.

Schwamborn:2002:HIM

[SEVSP02] R. Schwamborn, W. Ekau, M. Voss, and U. Saint-Paul. How important are mangroves as a carbon source for decapod crustacean larvae in a tropical estuary? Marine Ecology Progress Series, 229:195–205, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/229/m229p195.pdf.

Sewell:2005:ULD

[Sew05] M. A. Sewell. Utilization of lipids during early development of the sea urchin Evechinus chloroticus. Marine Ecology Progress Series, 304:133–142, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/304/m304p133.pdf.

Shima:2002:PLG

[SF02a] J. S. Shima and A. M. Findlay. Pelagic larval growth rate impacts benthic settlement and survival of a temperate reef fish. Marine Ecology Progress Series, 235:303–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p303.pdf.

Steele:2002:VRI

[SF02b] M. A. Steele and G. E. Forrester. Variation in the relative importance of sublethal effects of predators and competitors on growth of a temperate reef fish. Marine Ecology Progress Series, 237:233–245, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/237/m237p233.pdf.

Stentiford:2005:FRC

[SF05] G. D. Stentiford and S. W. Feist. First reported cases of intersex (ovotestis) in the flatfish species dab Limanda limanda: Dogger Bank, North Sea. Marine Ecology Progress Series, 301:307–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/301/m301p307.pdf.

Sanchez:2006:RAI

[SF06] Í. Sánchez and C. Fernández. Resource availability and invasibility in an intertidal macroalgal assemblage. Marine Ecology Progress Series, 313:85–94, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/313/m313p085.pdf.

Santos:2007:IIV

[SF07a] R. Santos and P. Flammang. Intra- and interspecific variation of attachment strength in sea urchins. Marine Ecology Progress Series, 332:129–142, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p129.pdf.

Shackell:2007:CEM

[SF07b] N. L. Shackell and K. T. Frank. Compensation in exploited marine fish communities on the Scotian Shelf, Canada. Marine Ecology Progress Series, 336:235–247, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p235.pdf.

Schloss:2007:RPC

[SFF^{+}07] I. R. Schloss, G. A. Ferreyra, M. E. Ferrario, G. O. Almandoz, R. Codina, A. A. Bianchi, C. F. Balestrini, H. A. Ochoa, D. Ruiz Pino, and A. Poisson. Role of plankton communities in sea-air variations in pCO_{2} in the SW Atlantic Ocean. Marine Ecology Progress Series, 332:93–106, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/332/
m332p093.pdf.

Suikkanen:2005:ECA

[SFG05] S. Suikkanen, G. O. Fistarol, and E. Granéli. Effects of cyanobacterial allelochemicals on a natural plankton community. Marine Ecology Progress Series, 287:1–9, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p001.pdf.

Sandnes:2000:BIN

[SFH^{+}00a] J. Sandnes, T. Forbes, R. Hansen, B. Sandnes, and B. Rygg. Bioturbation and irrigation in natural sediments, described by animal-community parameters. Marine Ecology Progress Series, 197:169–179, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/197/m197p169.pdf.

Sikkel:2000:HSD

[SFH00b] P. C. Sikkel, C. A. Fuller, and W. Hunte. Habitat/sex differences in time at cleaning stations and ectoparasite loads in a Caribbean reef fish. Marine Ecology Progress Series, 193:191–199, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p191.pdf.

Sandnes:2000:IPT

[SFHS00] J. Sandnes, T. Forbes, R. Hansen, and B. Sandnes. Influence of particle type and faunal activity on mixing of di(2-ethylhexyl)phthalate (DEHP) in natural sediments. Marine Ecology Progress Series, 197:151–167, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p151.pdf.

Shiozaki:2009:CFN

[SFKT09] T. Shiozaki, K. Furuya, T. Kodama, and S. Takeda. Contribution of N_{2} fixation to new production in the western North Pacific Ocean along 155^{o}E. Marine Ecology Progress Series, 377:19–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m377p019.pdf.

Shenkar:2005:SMB

[SFL05] N. Shenkar, M. Fine, and Y. Loya. Size matters: bleaching dynamics of the coral Oculina patagonica. Marine Ecology Progress Series, 294:181–188, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/294/m294p181.pdf.

Soetaert:2009:FAN

[SFL^{+}09] K. Soetaert, M. Franco, N. Lampadariou, A. Muthumbi, M. Steyaert, L. Vandepitte, E. vanden Berghe, and J. Vanaverbeke. Factors affecting nematode biomass, length and width from the shelf to the deep sea. Marine Ecology Progress Series, 392:123–132, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p123.pdf.

Shick:2005:ESA

[SFPGA05] J. M. Shick, C. Ferrier-Pagès, R. Grover, and D. Allemand. Effects of starvation, ammonium concentration, and photosynthesis on the UV-dependent accumulation of mycosporinelike amino acids (MAAs) in the coral Stylophora pistillata. Marine Ecology Progress Series, 295:135–156, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/295/
m295p135.pdf.

Stewart:2009:DGM

[SFR^{+}09] H. L. Stewart, J. P. Fram, D. C. Reed, S. L. Williams, M. A. Brzezinski, S. MacIntyre, and B. Gaylord. Differences in growth, morphology and tissue carbon and nitrogen of Macrocystis pyrifera within and at the outer edge of a giant kelp forest in California, USA. Marine Ecology Progress Series, 375:101–112, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p101.pdf.

Stewart:2005:CTP

[SFT^{+}05] G. M. Stewart, S. W. Fowler, J. L. Teyssié, O. Cotret, J. K. Cochran, and N. S. Fisher. Contrasting transfer of polonium-210 and lead-210 across three trophic levels in marine plankton. Marine Ecology Progress Series, 290:27–33, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/290/m290p027.pdf.

Savage:2001:CMA

[SFW01] C. Savage, J. G. Field, and R. M. Warwick. Comparative meta-analysis of the impact of offshore marine mining on macrobenthic communities versus organic pollution studies. Marine Ecology Progress Series, 221:265–275, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p265.pdf.

Stocks:2001:EMF

[SG01] K. I. Stocks and J. F. Grassle. Effects of microalgae and food limitation on the recolonization of benthic macrofauna into in situ saltmarsh-pond mesocosms. Marine Ecology Progress Series, 221:93–104, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/221/m221p093.pdf.

Starmans:2002:MED

[SG02a] A. Starmans and J. Gutt. Mega-epibenthic diversity: a polar comparison. Marine Ecology Progress Series, 225:45–52, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

225/m225p045.pdf. See erratum [Ano02g].

Stoecker:2002:PGM

[SG02b] D. K. Stoecker and D. E. Gustafson. Predicting grazing mortality of an estuarine dinoflagellate, Pfiesteria piscicida. Marine Ecology Progress Series, 233:31–38, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p031.pdf.

Schwemmer:2005:SDB

[SG05] P. Schwemmer and S. Garthe. At-sea distribution and behaviour of a surface-feeding seabird, the lesser black-backed gull Larus fuscus, and its association with different prey. Marine Ecology Progress Series, 285:245–258, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p245.pdf.

Sato:2006:IMO

[SG06a] T. Sato and S. Goshima. Impacts of male-only fishing and sperm limitation in manipulated populations of an unfished crab, Hapalogaster dentata. Marine Ecology Progress Series, 313:193–204, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/313/m313p193.pdf.

Scheibling:2006:CIB

[SG06b] R. E. Scheibling and P. Gagnon. Competitive interactions between the invasive green alga Codium fragile ssp. tomentosoides and native canopy-forming seaweeds in Nova Scotia (Canada). Marine Ecology Progress Series, 325:1–14, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Scheibling:2009:TMO

[SG09] R. E. Scheibling and P. Gagnon. Temperature-mediated outbreak dynamics of the invasive bryozoan Membranipora membranacea in Nova Scotian kelp beds. Marine Ecology Progress Series, 390:1–13, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/feature/m390p001.pdf.

Salo:2009:EPD

[SGB09] T. Salo, C. Gustafsson, and C. Boström. Effects of plant diversity on primary production and species interactions in brackish water angiosperm communities. Marine Ecology Progress Series, 396:261–272, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m396p261.pdf.

Stribling:2006:MVP

[SGCC06] J. M. Stribling, O. A. Glahn, X. M. Chen, and J. C. Cornwell. Microtopographic variability in plant distribution and biogeochemistry in a brackish-marsh system. Marine Ecology Progress Series, 320:121–129, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/320/m320p121.pdf.

Sirabella:2001:BCE

[SGCD01] P. Sirabella, A. Giuliani, A. Colosimo, and J. W. Dippner. Breaking down the climate effects on cod recruitment by principal component analysis and canonical correlation. Marine Ecology Progress Series, 216:213–222, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p213.pdf.

Sponaugle:2006:TMV

[SGCP06] S. Sponaugle, K. Grorud-Colvert, and D. Pinkard. Temperaturemediated variation in early life history traits and recruitment success of the coral reef fish Thalassoma bifasciatum in the Florida Keys. Marine Ecology Progress Series, 308:1–15, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Speirs:2006:OSM

[SGH^{+}06] D. C. Speirs, W. S. C. Gurney, M. R. Heath, W. Horbelt, S. N. Wood, and B. A. de Cuevas. Ocean-scale modelling of the distribution, abundance, and seasonal dynamics of the copepod Calanus finmarchicus. Marine Ecology Progress Series, 313:173–192, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/313/m313p173.pdf.

Sorensen:2007:SSM

[SGHG07] K. B. Sørensen, B. Glazer, A. Hannides, and E. Gaidos. Spatial structure of the microbial community in sandy carbonate sediment. Marine Ecology Progress Series, 346:61– 74, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/346/m346p061.pdf.

Schluter:2004:DTC

[SGK04] L. Schlüter, K. Garde, and H. Kaas. Detection of the toxic cyanobacteria Nodularia spumigena by means of a 4keto-myxoxanthophyll-like pigment in the Baltic Sea. Marine Ecology Progress Series, 275:69–78, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p069.pdf.

Seuront:2002:SSN

[SGL02a] L. Seuront, V. Gentilhomme, and Y. Lagadeuc. Smallscale nutrient patches in tidally mixed coastal waters. Marine Ecology Progress Series, 232:29–44, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/232/
m232p029.pdf.

Sorgo:2002:GSR

[SGL^{+}02b] A. Sorgo, F. Gaill, J. P. Lechaire, C. Arndt, and M. Bright. Glycogen storage in the Riftia pachyptila trophosome: contribution of host and symbionts. Marine Ecology Progress Series, 231:115–120, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/231/m231p115.pdf.

Sommer:2007:SHG

[SGP07] S. Sommer, E. Gutzmann, and O. Pfannkuche. Sediments hosting gas hydrates: oases for metazoan meiofauna. Marine Ecology Progress Series, 337:27–37, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/337/
m337p027.pdf.

Savage:2002:MDS

[SGV^{+}02] A. M. Savage, M. S. Goodson, S. Visram, H. Trapido-Rosenthal, J. Wiedenmann, and A. E. Douglas. Molecular diversity of symbiotic algae at the latitudinal margins of their distribution: dinoflagellates of the genus Symbiodinium in corals and sea anemones. Marine Ecology Progress Series, 244:17–26, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p017.pdf.

Swartzman:2000:SAB

[SH00] G. Swartzman and G. Hunt. Spatial association between murres (Uria spp.), puffins (Fratercula spp.) and fish shoals near Pribilof Islands, Alaska. Marine Ecology Progress Series, 206:297–309, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/206/m206p297.pdf.

Schmidt:2001:APC

[SH01] L. E. Schmidt and P. J. Hansen. Allelopathy in the prymnesiophyte Chrysochromulina polylepis: effect of cell concentration, growth phase and pH. Marine Ecology Progress Series, 216:67–81, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/216/m216p067.pdf.

Slezak:2003:EUV

[SH03] D. Slezak and G. J. Herndl. Effects of ultraviolet and visible radiation on the cellular concentrations of dimethylsulfoniopropionate (DMSP) in Emiliania huxleyi (strain L). Marine Ecology Progress Series, 246:61–71, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p061.pdf.

Shiao:2004:THN

[SH04] J. C. Shiao and P. P. Hwang. Thyroid hormones are necessary for teleostean otolith growth. Marine Ecology Progress Series, 278:271–278, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p271.pdf.

Smith:2005:AED

[SH05] T. M. Smith and J. S. Hindell. Assessing effects of diel period, gear selectivity and predation on patterns of microhabitat use by fish in a mangrove dominated system in SE Australia. Marine Ecology Progress Series, 294:257–270, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p257.pdf.

Scrosati:2006:FTQ

[SH06a] R. Scrosati and C. Heaven. Field technique to quantify intensity of scouring by sea ice in rocky intertidal habitats. Marine Ecology Progress Series, 320:293–295, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p293.pdf.

Sirota:2006:SEZ

[SH06b] L. Sirota and K. A. Hovel. Simulated eelgrass Zostera marina structural complexity: effects of shoot length, shoot density, and surface area on the epifaunal community of San Diego Bay, California, USA. Marine Ecology Progress Series, 326:115–131, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/326/m326p115.pdf.

Stokesbury:2006:ILS

[SH06c] K. D. E. Stokesbury and B. P. Harris. Impact of limited short-term sea scallop fishery on epibenthic community of Georges Bank closed areas. Marine Ecology Progress Series, 307:85–100, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/307/m307p085.pdf.

Schneider:2007:SVH

[SH07a] K. R. Schneider and B. Helmuth. Spatial variability in habitat temperature may drive patterns of selection between an invasive and native mussel species. Marine Ecology Progress Series, 339:157–167, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p157.pdf.

Scrosati:2007:STC

[SH07b] R. Scrosati and C. Heaven. Spatial trends in community richness, diversity, and evenness across rocky intertidal environmental stress gradients in eastern Canada. Marine Ecology Progress Series, 342:1–14, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Smith:2007:IBM

[SH07c] M. Smith and P. J. Hansen. Interaction between Mesodinium rubrum and its prey: importance of prey concentration, irradiance and pH. Marine Ecology Progress Series, 338:61–70, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p061.pdf.

Soderberg:2007:GLD

[SH07d] L. M. Søderberg and P. J. Hansen. Growth limitation due to high pH and low inorganic carbon concentrations in temperate species of the dinoflagellate genus Ceratium. Marine Ecology Progress Series, 351:103–112, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/351/m351p103.pdf.

Shafer:2000:EPA

[Sha00] D. J. Shafer. Evaluation of periodic and aperiodic otolith structure and somatic-otolith scaling for use in retrospective life history analysis of a tropical marine goby, Bathygobius coalitus. Marine Ecology Progress Series, 199:217–229, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p217.pdf.

Schwarz:2005:PPP

[SHA^{+}05] A. M. Schwarz, I. Hawes, N. Andrew, S. Mercer, V. Cummings, and S. Thrush. Primary production potential of non-geniculate coralline algae at Cape Evans, Ross Sea, Antarctica. Marine Ecology Progress Series, 294:131–140, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/294/m294p131.pdf.

Shanks:2009:BSVa

[Sha09a] A. L. Shanks. Barnacle settlement versus recruitment as indicators of larval delivery. I. Effects of post-settlement mortality and recruit density. Marine Ecology Progress Series, 385:205–216, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/385/m385p205.pdf.

Shanks:2009:BSVb

[Sha09b] A. L. Shanks. Barnacle settlement versus recruitment as indicators of larval delivery. II. Time-series analysis and hypothesized delivery mechanisms. Marine Ecology Progress Series, 385:217–226, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/385/m385p217.pdf.

Schallenberg:2003:CCI

[SHB03] M. Schallenberg, C. J. Hall, and C. W. Burns. Consequences of climate-induced salinity increases on zooplankton abundance and diversity in coastal lakes. Marine Ecology Progress Series, 251:181–189, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p181.pdf.

Sheaves:2001:TRF

[She01] M. Sheaves. Are there really few piscivorous fishes in shallow estuarine habitats? Marine Ecology Progress Series, 222:279–290, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p279.pdf.

Sheaves:2005:NCB

[She05] M. Sheaves. Nature and consequences of biological connectivity in mangrove systems. Marine Ecology Progress Series, 302:293–305, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p293.pdf.

Sheaves:2006:SDV

[She06] M. Sheaves. Scale-dependent variation in composition of fish fauna among sandy tropical estuarine embayments. Marine Ecology Progress Series, 310:173–184, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p173.pdf.

Sheehy:2008:QUB

[She08] M. R. J. Sheehy. Questioning the use of biochemical extraction to measure lipofuscin for age determination of crabs: Comment on Ju et al. (1999, 2001). Marine Ecology Progress Series, 353:303–306, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/353/m353p303.pdf. See reply

[HSJ08].

Sheaves:2009:CEC

[She09] M. Sheaves. Consequences of ecological connectivity: the coastal ecosystem mosaic. Marine Ecology Progress Series, 391:107–115, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/391/m391p107.pdf.

Soreide:2003:MCE

[SHFP^{+}03] J. E. Søreide, H. Hop, S. Falk-Petersen, B. Gulliksen, and E. Hansen. Macrozooplankton communities and environmental variables in the Barents Sea marginal ice zone in late winter and spring. Marine Ecology Progress Series, 263: 43–64, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/263/m263p043.pdf.

Suwa:2008:SFZ

[SHH08] R. Suwa, M. Hirose, and M. Hidaka. Seasonal fluctuation in zooxanthellar genotype composition and photophysiology in the corals Pavona divaricata and P. decussata. Marine Ecology Progress Series, 361:129–137, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m361p129.pdf.

Samuelsen:2009:SRC

[SHH09] A. Samuelsen, G. Huse, and C. Hansen. Shelf recruitment of Calanus finmarchicus off the west coast of Norway: role of physical processes and timing of diapause termination. Marine Ecology Progress Series, 386:163–180, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/386/
m386p163.pdf.

Smith:2008:EEF

[SHJC08] T. M. Smith, J. S. Hindell, G. P. Jenkins, and R. M. Connolly. Edge effects on fish associated with seagrass and sand patches. Marine Ecology Progress Series, 359:203–213, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/359/
m359p203.pdf.

Sikkel:2005:CCS

[SHK05] P. C. Sikkel, S. E. Herzlieb, and D. L. Kramer. Compensatory cleaner-seeking behavior following spawning in female yellowtail damselfish. Marine Ecology Progress Series, 296:1– 11, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sciandra:2003:RCE

[SHL^{+}03] A. Sciandra, J. Harlay, D. Lefèvre, R. Lemée, P. Rimmelin, M. Denis, and J. P. Gattuso. Response of coccolithophorid Emiliania huxleyi to elevated partial pressure of CO_{2} under nitrogen limitation. Marine Ecology Progress Series, 261: 111–122, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/261/m261p111.pdf.

Steyaert:2001:TMN

[SHM^{+}01] M. Steyaert, P. M. J. Herman, T. Moens, J. Widdows, and M. Vincx. Tidal migration of nematodes on an estuarine tidal flat (the Molenplaat, Schelde Estuary, SW Netherlands). Marine Ecology Progress Series, 224:299–304, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p299.pdf.

Staehr:2002:PFM

[SHM02] P. A. Staehr, P. Henriksen, and S. Markager. Photoacclimation of four marine phytoplankton species to irradiance and nutrient availability. Marine Ecology Progress Series, 238: 47–59, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p047.pdf.

Stokesbury:2007:SSM

[SHMN07] K. D. E. Stokesbury, B. P. Harris, M. C. Marino II, and J. I. Nogueira. Sea scallop mass mortality in a marine protected area. Marine Ecology Progress Series, 349:151– 158, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/349/m349p151.pdf.

Schaber:2009:HRS

[SHNV09] M. Schaber, H. H. Hinrichsen, S. Neuenfeldt, and R. Voss. Hydroacoustic resolution of small-scale vertical distribution in Baltic cod Gadus morhua —habitat choice and limits during spawning. Marine Ecology Progress Series, 377:239– 253, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p239.pdf.

Skovgaard:2000:PMD

[SHS00] A. Skovgaard, P. J. Hansen, and D. K. Stoecker. Physiology of the mixotrophic dinoflagellate Fragilidium subglobosum. I. Effects of phagotrophy and irradiance on photosynthesis and carbon content. Marine Ecology Progress Series, 201:129–136, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/201/m201p129.pdf.

Sherwood:2005:SIC

[SHS^{+}05] O. A. Sherwood, J. M. Heikoop, D. B. Scott, M. J. Risk, T. P. Guilderson, and R. A. McKinney. Stable isotopic composition of deep-sea gorgonian corals Primnoa spp.: a new archive of surface processes. Marine Ecology Progress Series, 301:135–148, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p135.pdf.

Sommer:2006:TDN

[SHS06] F. Sommer, T. Hansen, and U. Sommer. Transfer of diazotrophic nitrogen to mesozooplankton in Kiel Fjord, Western Baltic Sea: a mesocosm study. Marine Ecology Progress Series, 324:105–112, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/324/m324p105.pdf.

Sheppard:2008:AWC

[SHS08] C. R. C. Sheppard, A. Harris, and A. L. S. Sheppard. Archipelago-wide coral recovery patterns since 1998 in the Chagos Archipelago, central Indian Ocean. Marine Ecology Progress Series, 362:109–117, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/362/m362p109.pdf.

Suzuki:2008:ESS

[SHSF08] G. Suzuki, T. Hayashibara, Y. Shirayama, and H. Fukami. Evidence of species-specific habitat selectivity of Acropora corals based on identification of new recruits by two molecular markers. Marine Ecology Progress Series, 355:149–159, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m355p149.pdf.

Sommer:2004:PPR

[SHSV04] U. Sommer, T. Hansen, H. Stibor, and O. Vadstein. Persistence of phytoplankton responses to different Si: N ratios under mesozooplankton grazing pressure: a mesocosm study with NE Atlantic plankton. Marine Ecology Progress Series, 278:67–75, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p067.pdf.

Seuront:2004:IVS

[SHT^{+}04a] L. Seuront, J. S. Hwang, L. C. Tseng, F. G. Schmitt, S. Souissi, and C. K. Wong. Individual variability in the swimming behavior of the sub-tropical copepod Oncaea venusta (Copepoda: Poecilostomatoida). Marine Ecology Progress Series, 283:199–217, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/283/m283p199.pdf.

Staahl:2004:RSO

[SHT^{+}04b] H. Ståhl, P. O. J. Hall, A. Tengberg, A. B. Josefson, N. Streftaris, A. Zenetos, and A. P. Karageorgis. Respiration and sequestering of organic carbon in shelf sediments of the oligotrophic northern Aegean Sea. Marine Ecology Progress Series, 269:33–48, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/269/m269p033.pdf.

Szathmary:2009:CCR

[SHW09] P. L. Szathmary, B. Helmuth, and D. S. Wethey. Climate change in the rocky intertidal zone: predicting and measuring the body temperature of a keystone predator. Marine Ecology Progress Series, 374:43–56, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/374/m374p043.pdf.

Shi:2005:GSI

[SHZ^{+}05] H. H. Shi, C. J. Huang, S. X. Zhu, X. J. Yu, and W. Y. Xie. Generalized system of imposex and reproductive failure in female gastropods of coastal waters of mainland China. Marine Ecology Progress Series, 304:179–189, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/304/
m304p179.pdf.

Shiao:2003:DHU

[SICT03] J. C. Shiao, Y. Iizuka, C. W. Chang, and W. N. Tzeng. Disparities in habitat use and migratory behavior between tropical eel Anguilla marmorata and temperate eel A. japonica in four Taiwanese rivers. Marine Ecology Progress Series, 261:233–242, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p233.pdf.

Suryan:2002:STF

[SIK^{+}02] R. M. Suryan, D. B. Irons, M. Kaufman, J. Benson, P. G. R. Jodice, D. D. Roby, and E. D. Brown. Short-term fluctuations in forage fish availability and the effect on prey selection and brood-rearing in the black-legged kittiwake Rissa tridactyla. Marine Ecology Progress Series, 236:273–287, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p273.pdf.

Stevick:2008:TRO

[SIK^{+}08] P. T. Stevick, L. S. Incze, S. D. Kraus, S. Rosen, N. Wolff, and A. Baukus. Trophic relationships and oceanography on and around a small offshore bank. Marine Ecology Progress Series, 363:15–28, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/363/m363p015.pdf.

Sims:2000:CTF

[Sim00] D. W. Sims. Can threshold foraging responses of basking sharks be used to estimate their metabolic rate? Marine Ecology Progress Series, 200:289–296, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/200/m200p289.pdf.

Strand:2002:ITS

[SJ02] J. Strand and J. A. Jacobsen. Imposex in two sublittoral neogastropods from the Kattegat and Skagerrak: the common whelk Buccinum undatum and the red whelk Neptunea antiqua. Marine Ecology Progress Series, 244:171– 177, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p171.pdf.

Syms:2004:HSD

[SJ04] C. Syms and G. P. Jones. Habitat structure, disturbance and the composition of sand-dwelling goby assemblages in a coral reef lagoon. Marine Ecology Progress Series, 268: 221–230, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/268/m268p221.pdf.

Semmens:2005:EBI

[SJ05] J. M. Semmens and G. D. Jackson. Evaluation of biochemical indices for assessing growth and condition of the deepwater squid Moroteuthis ingens. Marine Ecology Progress Series, 289:215–223, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p215.pdf.

Smith:2006:SOC

[SJ06] Nathan G. Smith and Cynthia M. Jones. Substituting otoliths for chemical analyses: Does sagitta = lapillus? Marine Ecology Progress Series, 313:241–247, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/313/
m313p241.pdf.

Sheaves:2008:IMF

[SJ08] M. Sheaves and R. Johnston. Influence of marine and freshwater connectivity on the dynamics of subtropical estuarine wetland fish metapopulations. Marine Ecology Progress Series, 357:225–243, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/357/m357p225.pdf.

Sheaves:2009:EDS

[SJ09a] M. Sheaves and R. Johnston. Ecological drivers of spatial variability among fish fauna of 21 tropical Australian estuaries. Marine Ecology Progress Series, 385:245–260, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p245.pdf.

Strain:2009:CBI

[SJ09b] E. M. A. Strain and C. R. Johnson. Competition between an invasive urchin and commercially fished abalone: effect on body condition, reproduction and survivorship. Marine Ecology Progress Series, 377:169–182, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/377/
m377p169.pdf.

Stockwell:2009:TBR

[SJA^{+}09] B. Stockwell, C. R. L. Jadloc, R. A. Abesamis, A. C. Alcala, and G. R. Russ. Trophic and benthic responses to no-take marine reserve protection in the Philippines. Marine Ecology Progress Series, 389:1–15, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Smith:2001:MWE

[SJCM01] I. P. Smith, A. C. Jensen, K. J. Collins, and E. L. Mattey. Movement of wild European lobsters Homarus gammarus in natural habitat. Marine Ecology Progress Series, 222:177–186, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p177.pdf.

Seminoff:2006:SID

[SJE^{+}06] J. A. Seminoff, T. T. Jones, T. Eguchi, D. R. Jones, and P. H. Dutton. Stable isotope discrimination (δ ^{13}C and δ 15N) between soft tissues of the green sea turtle Chelonia mydas and its diet. Marine Ecology Progress Series, 308: 271–278, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/308/m308p271.pdf.

Svensson:2004:MOP

[SJH^{+}04] C. J. Svensson, S. R. Jenkins, S. J. Hawkins, A. A. Myers, P. Range, J. Paula, R. M. O’Riordan, and P. Åberg. Models of open populations with space-limited recruitment in stochastic environments: relative importance of recruitment and survival in populations of Semibalanus balanoides. Marine Ecology Progress Series, 275:185–197, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/275/m275p185.pdf.

Seong:2006:BCO

[SJK^{+}06] K. A. Seong, H. J. Jeong, S. Kim, G. H. Kim, and J. H. Kang. Bacterivory by co-occurring red-tide algae, heterotrophic nanoflagellates, and ciliates. Marine Ecology Progress Series, 322:85–97, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p085.pdf.

Stark:2005:TAS

[SJL05] K. E. Stark, G. D. Jackson, and J. M. Lyle. Tracking arrow squid movements with an automated acoustic telemetry system. Marine Ecology Progress Series, 299:167–177, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/299/m299p167.pdf.

Seminoff:2006:UBG

[SJM06] J. A. Seminoff, T. T. Jones, and G. J. Marshall. Underwater behaviour of green turtles monitored with video-time-depth recorders: what’s missing from dive profiles? Marine Ecology Progress Series, 322:269–280, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/322/m322p269.pdf.

St-Jean:2002:HFB

[SJPC02a] S. D. St-Jean, É. Pelletier, and S. C. Courtenay. Hemocyte functions and bacterial clearance affected in vivo by TBT and DBT in the blue mussel Mytilus edulis. Marine Ecology Progress Series, 236:163–178, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/236/m236p163.pdf.

St-Jean:2002:VLL

[SJPC02b] S. D. St-Jean, É. Pelletier, and S. C. Courtenay. Very low levels of waterborne butyltins modulate hemocyte function in the blue mussel Mytilus edulis. Marine Ecology Progress Series, 236:155–161, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p155.pdf.

Scharf:2000:PSP

[SJR00] F. S. Scharf, F. Juanes, and R. A. Rountree. Predator size-prey size relationships of marine fish predators: interspecific variation and effects of ontogeny and body size on trophic-niche breadth. Marine Ecology Progress Series, 208: 229–248, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/208/m208p229.pdf.

Shillito:2001:TRH

[SJS^{+}01] B. Shillito, D. Jollivet, P. M. Sarradin, P. Rodier, F. Lallier, D. Desbruyeres, and F. Gaill. Temperature resistance of Hesiolyra bergi, a polychaetous annelid living on deepsea vent smoker walls. Marine Ecology Progress Series, 216: 141–149, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/216/m216p141.pdf.

Solan:2002:OQL

[SK02] M. Solan and R. Kennedy. Observation and quantification of in situ animal-sediment relations using time-lapse sediment profile imagery (t-SPI). Marine Ecology Progress Series, 228:179–191, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p179.pdf.

Skarra:2003:VDF

[SK03] H. Skarra and S. Kaartvedt. Vertical distribution and feeding of the carnivorous copepod Paraeuchaeta norvegica. Marine Ecology Progress Series, 249:215–222, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p215.pdf.

Sassa:2004:LFH

[SK04] C. Sassa and K. Kawaguchi. Larval feeding habits of Diaphus garmani and Myctophum asperum (Pisces: Myctophidae) in the transition region of the western North Pacific. Marine Ecology Progress Series, 278:279–290, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/278/
m278p279.pdf.

Sassa:2005:LFH

[SK05] C. Sassa and K. Kawaguchi. Larval feeding habits of Diaphus theta, Protomyctophum thompsoni, and Tarletonbeania taylori (Pisces: Myctophidae) in the transition region of the western North Pacific. Marine Ecology Progress Series, 298:261–276, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/298/m298p261.pdf.

Sentchev:2007:MDF

[SK07] A. Sentchev and K. Korotenko. Modelling distribution of flounder larvae in the eastern English Channel: sensitivity to physical forcing and biological behaviour. Marine Ecology Progress Series, 347:233–245, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m347p233.pdf.

Schultz:2009:ISM

[SKBP09] S. T. Schultz, C. Kruschel, and T. Bakran-Petricioli. Influence of seagrass meadows on predator–prey habitat segregation in an Adriatic lagoon. Marine Ecology Progress Series, 374:85–99, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/374/m374p085.pdf.

Stoeck:2002:PFA

[SKDP02] T. Stoeck, I. Kröncke, G. C. A. Duineveld, and A. Palojärvi. Phospholipid fatty acid profiles at depositional and nondepositional sites in the North Sea. Marine Ecology Progress Series, 241:57–70, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p057.pdf.

Southwell:2005:PDC

[SKE05] C. J. Southwell, K. R. Kerry, and P. H. Ensor. Predicting the distribution of crabeater seals Lobodon carcinophaga off east Antarctica during the breeding season. Marine Ecology Progress Series, 299:297–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/299/m299p297.pdf.

Scott:2002:SDO

[SKFPS02] C. L. Scott, S. Kwasniewski, S. Falk-Petersen, and J. R. Sargent. Species differences, origins and functions of fatty alcohols and fatty acids in the wax esters and phospholipids of Calanus hyperboreus, C. glacialis and C. finmarchicus from Arctic waters. Marine Ecology Progress Series, 235:127–134, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p127.pdf.

Siegel:2003:LDM

[SKGG03] D. A. Siegel, B. P. Kinlan, B. Gaylord, and S. D. Gaines. Lagrangian descriptions of marine larval dispersion. Marine Ecology Progress Series, 260:83–96, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p083.pdf.

Suwa:2006:CPC

[SKH06] R. Suwa, M. N. I. Khan, and A. Hagihara. Canopy photosynthesis, canopy respiration and surplus production in a subtropical mangrove Kandelia candel forest, Okinawa Island, Japan. Marine Ecology Progress Series, 320:131–139, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p131.pdf.

Sopanen:2006:THP

[SKK^{+}06] S. Sopanen, M. Koski, P. Kuuppo, P. Uronen, C. Legrand, and T. Tamminen. Toxic haptophyte Prymnesium parvum affects grazing, survival, egestion and egg production of the calanoid copepods Eurytemora affinis and Acartia bifilosa. Marine Ecology Progress Series, 327:223–232, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p223.pdf.

Sinclair:2006:NUKa

[SKMS06a] G. A. Sinclair, D. Kamykowski, E. Milligan, and B. Schaeffer. Nitrate uptake by Karenia brevis. I. Influences of prior environmental exposure and biochemical state on diel uptake of nitrate. Marine Ecology Progress Series, 328:117– 124, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p117.pdf.

Sinclair:2006:NUKb

[SKMS06b] G. A. Sinclair, D. Kamykowski, E. Milligan, and B. Schaeffer. Nitrate uptake by Karenia brevis. II. Behavior and uptake physiology in a nitrate-depleted mesocosm with a bottom nutrient source. Marine Ecology Progress Series, 328: 125–131, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p125.pdf.

Suzuki:2008:MJT

[SKO^{+}08] K. W. Suzuki, A. Kasai, T. Ohta, K. Nakayama, and M. Tanaka. Migration of Japanese temperate bass Lateolabrax japonicus juveniles within the Chikugo River estuary revealed by δ ^{13}C analysis. Marine Ecology Progress Series, 358:245–256, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/358/m358p245.pdf.

Schofield:2007:FFA

[SKP^{+}07] G. Schofield, K. A. Katselidis, J. D. Pantis, P. Dimopoulos, and G. C. Hays. Female-female aggression: structure of interaction and outcome in loggerhead sea turtles. Marine Ecology Progress Series, 336:267–274, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/336/m336p267.pdf.

Sekar:2008:MCC

[SKR08] R. Sekar, L. T. Kaczmarsky, and L. L. Richardson. Microbial community composition of black band disease on the coral host Siderastrea siderea from three regions of the wider Caribbean. Marine Ecology Progress Series, 362:85–98, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p085.pdf.

Scheibling:2009:HCO

[SKR09] R. E. Scheibling, N. E. Kelly, and B. G. Raymond. Herbivory and community organization on a subtidal cobble bed. Marine Ecology Progress Series, 382:113–128, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/382/
m382p113.pdf.

Sopanen:2008:PPE

[SKU^{+}08] S. Sopanen, M. Koski, P. Uronen, P. Kuuppo, S. Lehtinen, C. Legrand, and T. Tamminen. Prymnesium parvum exotoxins affect the grazing and viability of the calanoid copepod Eurytemora affinis. Marine Ecology Progress Series, 361:191–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p191.pdf.

Suzuki:2006:SHS

[SKW^{+}06] Y. Suzuki, S. Kojima, H. Watanabe, M. Suzuki, S. Tsuchida, T. Nunoura, H. Hirayama, K. Takai, K. H. Nealson, and K. Horikoshi. Single host and symbiont lineages of hydrothermal-vent gastropods Ifremeria nautilei (Provannidae): biogeography and evolution. Marine Ecology Progress Series, 315:167–175, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/315/m315p167.pdf.

Schonberg:2005:MIU

[SL05a] C. H. L. Schönberg and W. K. W. Loh. Molecular identity of the unique symbiotic dinoflagellates found in the bioeroding demosponge Cliona orientalis. Marine Ecology Progress Series, 299:157–166, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/299/m299p157.pdf.

Sherrard:2005:FFJa

[SL05b] K. M. Sherrard and M. LaBarbera. Form and function in juvenile ascidians. I. Implications of early juvenile morphologies for performance. Marine Ecology Progress Series, 287: 127–138, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/287/m287p127.pdf.

Sherrard:2005:FFJb

[SL05c] K. M. Sherrard and M. LaBarbera. Form and function in juvenile ascidians. II. Ontogenetic scaling of volumetric flow rates. Marine Ecology Progress Series, 287:139–148, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/287/
m287p139.pdf.

Soong:2005:AMB

[SL05d] K. Soong and Y. Leu. Adaptive mechanism of the bimodal emergence dates in the intertidal midge Pontomyia oceana. Marine Ecology Progress Series, 286:107–114, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/286/
m286p107.pdf.

Schiel:2007:GDA

[SL07] D. R. Schiel and S. A. Lilley. Gradients of disturbance to an algal canopy and the modification of an intertidal community. Marine Ecology Progress Series, 339:1–11, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Slater:2007:OAY

[SLB07] J. J. Slater, T. E. Lankford, Jr., and J. A. Buckel. Overwintering ability of young-of-the-year bluefish Pomatomus saltatrix: effect of ration and cohort of origin on survival. Marine Ecology Progress Series, 339:259–269, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p259.pdf.

Stockin:2008:TAB

[SLB^{+}08] K. A. Stockin, D. Lusseau, V. Binedell, N. Wiseman, and M. B. Orams. Tourism affects the behavioural budget of the common dolphin Delphinus sp. in the Hauraki Gulf, New Zealand. Marine Ecology Progress Series, 355:287– 295, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/355/m355p287.pdf.

Scheinberg:2005:GTC

[SLC05] R. D. Scheinberg, M. R. Landry, and A. Calbet. Grazing of two common appendicularians on the natural prey assemblage of a tropical coastal ecosystem. Marine Ecology Progress Series, 294:201–212, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/294/m294p201.pdf.

Sciandra:2000:RGR

[SLCB00] A. Sciandra, L. Lazzara, H. Claustre, and M. Babin. Responses of growth rate, pigment composition and optical properties of Cryptomonas sp. to light and nitrogen stresses. Marine Ecology Progress Series, 201:107–120, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p107.pdf.

Soong:2009:STT

[SLCC09] K. Soong, Y. J. Lin, S. M. Chao, and D. Chang. Spawning time of two shallow-water brittle stars. Marine Ecology Progress Series, 376:165–171, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/376/m376p165.pdf.

Sutherland:2001:EBW

[SLEH01] T. F. Sutherland, C. D. Levings, C. C. Elliott, and W. W. Hesse. Effect of ballast water treatment system on survivorship of natural populations of marine plankton. Marine Ecology Progress Series, 210:139–148, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/210/m210p139.pdf.

Schultz:2003:MPU

[SLFM03] E. T. Schultz, K. M. M. Lwiza, M. C. Fencil, and J. M. Martin. Mechanisms promoting upriver transport of larvae of two fish species in the Hudson River estuary. Marine Ecology Progress Series, 251:263–277, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/251/m251p263.pdf.

Sakka:2000:SOP

[SLGD00] A. Sakka, L. Legendre, M. Gosselin, and B. Delesalle. Structure of the oligotrophic planktonic food web under low grazing of heterotrophic bacteria: Takapoto Atoll, French Polynesia. Marine Ecology Progress Series, 197:1–17, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p001.pdf.

Sponaugle:2009:SVL

[SLHR09] S. Sponaugle, J. K. Llopiz, L. N. Havel, and T. L. Rankin. Spatial variation in larval growth and gut fullness in a coral reef fish. Marine Ecology Progress Series, 383:239– 249, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/383/m383p239.pdf.

Sunda:2005:DDR

[SLHT05] W. G. Sunda, R. W. Litaker, D. R. Hardison, and P. A. Tester. Dimethylsulfoniopropionate (DMSP) and its relation to algal pigments in diverse waters of the Belize coastal lagoon and barrier reef system. Marine Ecology Progress Series, 287:11–22, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/287/m287p011.pdf.

Solomon:2003:CEP

[SLKF03] C. M. Solomon, E. J. Lessard, R. G. Keil, and M. S. Foy. Characterization of extracellular polymers of Phaeocystis globosa and P. antarctica. Marine Ecology Progress Series, 250: 81–89, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p081.pdf.

Schlekat:2002:ASP

[SLL02] C. E. Schlekat, B. G. Lee, and S. N. Luoma. Assimilation of selenium from phytoplankton by three benthic invertebrates: effect of phytoplankton species. Marine Ecology Progress Series, 237:79–85, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p079.pdf.

Sonnenholzner:2007:CEF

[SLL07] J. I. Sonnenholzner, L. B. Ladah, and K. D. Lafferty. Cascading effects of fishing on Galapagos rocky reef communities [Article Retracted]. Marine Ecology Progress Series, 343:77–85, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See retraction [Ano07h] and reanalysis [SLL09].

Sonnenholzner:2009:CEF

[SLL09] J. I. Sonnenholzner, L. B. Ladah, and K. D. Lafferty. Cascading effects of fishing on Galapagos rocky reef communities: reanalysis using corrected data. Marine Ecology Progress Series, 375:209–218, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m375p209.pdf. See [SLL07].

Schultes:2000:DDP

[SLM^{+}00] S. Schultes, M. Levasseur, S. Michaud, G. Cantin, G. Wolfe, M. Gosselin, and S. de Mora. Dynamics of dimethylsulfide production from dissolved dimethylsulfoniopropionate in the Labrador Sea. Marine Ecology Progress Series, 202: 27–40, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/202/m202p027.pdf.

Scarratt:2002:IPT

[SLM^{+}02] M. G. Scarratt, M. Levasseur, S. Michaud, G. Cantin, M. Gosselin, and S. J. de Mora. Influence of phytoplankton taxonomic profile on the distribution of dimethylsulfide and dimethylsulfoniopropionate in the northwest Atlantic. Marine Ecology Progress Series, 244:49–61, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/244/m244p049.pdf.

Sosa-Lopez:2007:SSR

[SLM07] A. Sosa-Lopez and D. Mouillot. Spatiotemporal segregation regulates species richness along biomass gradients in estuarine fish assemblages. Marine Ecology Progress Series, 342: 217–226, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/342/m342p217.pdf.

Sommer:2009:SME

[SLP^{+}09] S. Sommer, P. Linke, O. Pfannkuche, T. Schleicher, J. Schneider v. Deimling, A. Reitz, M. Haeckel, S. Flögel, and C. Hensen. Seabed methane emissions and the habitat of frenulate tubeworms on the Captain Arutyunov mud volcano (Gulf of Cadiz). Marine Ecology Progress Series, 382: 69–86, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m382p069.pdf.

Scarratt:2000:PCD

[SLS^{+}00] M. G. Scarratt, M. Levasseur, S. Schultes, S. Michaud, G. Cantin, A. Vézina, M. Gosselin, and S. J. de Mora. Production and consumption of dimethylsulfide (DMS) in North Atlantic waters. Marine Ecology Progress Series, 204: 13–26, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/204/m204p013.pdf.

Sengco:2001:RRB

[SLT^{+}01] M. R. Sengco, A. Li, K. Tugend, D. Kulis, and D. M. Anderson. Removal of red- and brown-tide cells using clay flocculation. I. Laboratory culture experiments with Gymnodinium breve and Aureococcus anophagefferens. Marine Ecology Progress Series, 210:41–53, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/210/m210p041.pdf.

Semmens:2000:EVG

[SM00a] J. M. Semmens and N. A. Moltschaniwskyj. An examination of variable growth in the loliginid squid Sepioteuthis lessoniana: a whole animal and reductionist approach. Marine Ecology Progress Series, 193:135–141, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/193/m193p135.pdf.

Sheaves:2000:SCM

[SM00b] M. Sheaves and B. Molony. Short-circuit in the mangrove food chain. Marine Ecology Progress Series, 199:97–109, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
199/m199p097.pdf.

Sanford:2001:STV

[SM01a] E. Sanford and B. A. Menge. Spatial and temporal variation in barnacle growth in a coastal upwelling system. Marine Ecology Progress Series, 209:143–157, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/209/m209p143.pdf.

Swain:2001:SST

[SM01b] D. P. Swain and M. J. Morgan. Sex-specific temperature distribution in four populations of American plaice Hippoglossoides platessoides. Marine Ecology Progress Series, 212: 233–246, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/212/m212p233.pdf.

Szela:2005:MPO

[SM05] T. L. Szela and A. G. Marsh. Microtiter plate, optode respirometry, and inter-individual variance in metabolic rates among nauplii of Artemia sp. Marine Ecology Progress Series, 296:281–289, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p281.pdf.

Sanford:2007:ROC

[SM07a] E. Sanford and B. A. Menge. Reproductive output and consistency of source populations in the sea star Pisaster ochraceus. Marine Ecology Progress Series, 349:1–12, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Saunders:2007:TES

[SM07b] M. Saunders and A. Metaxas. Temperature explains settlement patterns of the introduced bryozoan Membranipora membranacea in Nova Scotia, Canada. Marine Ecology Progress Series, 344:95–106, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/344/m344p095.pdf.

Saunders:2008:HRI

[SM08a] M. Saunders and A. Metaxas. High recruitment of the introduced bryozoan Membranipora membranacea is associated with kelp bed defoliation in Nova Scotia, Canada. Marine Ecology Progress Series, 369:139–151, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/369/
m369p139.pdf.

Saunders:2008:PBV

[SM08b] T. Saunders and S. Mayfield. Predicting biological variation using a simple morphometric marker in the sedentary marine invertebrate Haliotis rubra. Marine Ecology Progress Series, 366:75–89, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/366/m366p075.pdf.

Shinen:2009:MIR

[SM09] J. S. Shinen and S. G. Morgan. Mechanisms of invasion resistance: competition among intertidal mussels promotes establishment of invasive species and displacement of native species. Marine Ecology Progress Series, 383:187–197, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m383p187.pdf.

Smith:2003:UTF

[SMA^{+}03] C. J. Smith, S. J. Marrs, R. J. A. Atkinson, K. N. Papadopoulou, and J. M. Hills. Underwater television for fisheries-independent stock assessment of Nephrops norvegicus from the Aegean (eastern Mediterranean) Sea. Marine Ecology Progress Series, 256:161–170, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p161.pdf.

Smale:2007:IDI

[Sma07] D. A. Smale. Ice disturbance intensity structures benthic communities in nearshore Antarctic waters. Marine Ecology Progress Series, 349:89–102, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p089.pdf.

Smoot:2003:SCS

[SMB^{+}03] J. C. Smoot, L. M. Mayer, M. J. Bock, P. C. Wood, and R. H. Findlay. Structures and concentrations of surfactants in gut fluid of the marine polychaete Arenicola marina. Marine Ecology Progress Series, 258:161–169, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/258/
m258p161.pdf.

Sargian:2005:NSE

[SMC^{+}05] P. Sargian, B. Mostajir, K. Chatila, G. A. Ferreyra, E. Pelletier, and S. Demers. Non-synergistic effects of water-soluble crude oil and enhanced ultraviolet-B radiation on a natural plankton assemblage. Marine Ecology Progress Series, 294: 63–77, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/294/m294p063.pdf.

Shinen:2009:IRR

[SMC09] J. S. Shinen, S. G. Morgan, and A. L. Chan. Invasion resistance on rocky shores: direct and indirect effects of three native predators on an exotic and a native prey species. Marine Ecology Progress Series, 378:47–54, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m378p047.pdf.

Stergiou:2007:TSS

[SMCE07] Konstantinos I. Stergiou, D. K. Moutopoulos, H. J. A. Casal, and K. Erzini. Trophic signatures of small-scale fishing gears: implications for conservation and management. Marine Ecology Progress Series, 333:117–128, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m333p117.pdf.

Strom:2000:WSL

[SMF00] S. L. Strom, C. B. Miller, and B. W. Frost. What sets lower limits to phytoplankton stocks in high-nitrate, low-chlorophyll regions of the open ocean? Marine Ecology Progress Series, 193:19–31, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p019.pdf.

Steele:2002:SME

[SMF^{+}02] M. A. Steele, J. C. Malone, A. M. Findlay, M. H. Carr, and G. E. Forrester. A simple method for estimating larval supply in reef fishes and a preliminary test of population limitation by larval delivery in the kelp bass Paralabrax clathratus. Marine Ecology Progress Series, 235:195–203, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p195.pdf.

Sotelo:2008:GVS

[SMFP08] G. Sotelo, P. Morán, L. Fernández, and D. Posada. Genetic variation of the spiny spider crab Maja brachydactyla in the northeastern Atlantic. Marine Ecology Progress Series, 362:211–223, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/362/m362p211.pdf.

Sundback:2000:NFD

[SMG00] K. Sundbäck, A. Miles, and E. Göransson. Nitrogen fluxes, denitrification and the role of microphytobenthos in microtidal shallow-water sediments: an annual study. Marine Ecology Progress Series, 200:59–76, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/200/m200p059.pdf.

Steer:2002:TVE

[SMG02] M. A. Steer, N. A. Moltschaniwskyj, and F. C. Gowland. Temporal variability in embryonic development and mortality in the southern calamary Sepioteuthis australis: a field assessment. Marine Ecology Progress Series, 243:143–150, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/243/m243p143.pdf.

Stillman:2000:PMP

[SMGCW00] R. A. Stillman, S. McGrorty, J. D. Goss-Custard, and A. D. West. Predicting mussel population density and age structure: the relationship between model complexity and predictive power. Marine Ecology Progress Series, 208:131–145, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
208/m208p131.pdf.

Soetaert:2002:SSO

[SMH02] K. Soetaert, A. Muthumbi, and C. Heip. Size and shape of ocean margin nematodes: morphological diversity and depthrelated patterns. Marine Ecology Progress Series, 242:179– 193, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p179.pdf.

Sundback:2003:IBN

[SMH^{+}03] K. Sundbäck, A. Miles, S. Hulth, L. Pihl, P. Engström, E. Selander, and A. Svenson. Importance of benthic nutrient regeneration during initiation of macroalgal blooms in shallow bays. Marine Ecology Progress Series, 246:115–126, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p115.pdf.

Sharples:2009:ESA

[SMH09] R. J. Sharples, M. L. Mackenzie, and P. S. Hammond. Estimating seasonal abundance of a central place forager using counts and telemetry data. Marine Ecology Progress Series, 378:289–298, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/378/m378p289.pdf.

Seitz:2003:EHP

[SMHC03] R. D. Seitz, L. S. Marshall, Jr., A. H. Hines, and K. L. Clark. Effects of hypoxia on predator–prey dynamics of the blue crab Callinectes sapidus and the Baltic clam Macoma balthica in Chesapeake Bay. Marine Ecology Progress Series, 257:179–188, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p179.pdf.

Suggett:2009:IFR

[SMHG09] D. J. Suggett, C. M. Moore, A. E. Hickman, and R. J. Geider. Interpretation of fast repetition rate (FRR) fluorescence: signatures of phytoplankton community structure versus physiological state. Marine Ecology Progress Series, 376:1–19, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Schluter:2000:UPP

[SMHL00] L. Schlüter, F. Møhlenberg, H. Havskum, and S. Larsen. The use of phytoplankton pigments for identifying and quantifying phytoplankton groups in coastal areas: testing the influence of light and nutrients on pigment/chlorophyll-a ratios. Marine Ecology Progress Series, 192:49–63, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/192/m192p049.pdf.

Smith:2004:BDC

[Smi04] L. D. Smith. Biogeographic differences in claw size and performance in an introduced crab predator Carcinus maenas. Marine Ecology Progress Series, 276:209–222, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/276/
m276p209.pdf.

Sainte-Marie:2002:GSR

[SML02a] B. Sainte-Marie and M. Lafrance. Growth and survival of recently settled snow crab Chionoecetes opilio in relation to intra- and intercohort competition and cannibalism: a laboratory study. Marine Ecology Progress Series, 244:191– 203, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p191.pdf.

Stunz:2002:GNS

[SML02b] G. W. Stunz, T. J. Minello, and P. S. Levin. Growth of newly settled red drum Sciaenops ocellatus in different estuarine habitat types. Marine Ecology Progress Series, 238:

227–236, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p227.pdf.

Sobrino:2005:EUU

[SML05] C. Sobrino, O. Montero, and L. M. Lubián. Effect of UV-a and UV-b on diel patterns of growth and metabolic activity in Nannochloris atomus cultures assessed by flow cytometry. Marine Ecology Progress Series, 293:29–35, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/293/
m293p029.pdf.

Sabarros:2009:MEI

[SML^{+}09] P. S. Sabarros, F. Ménard, J. J. Lévénez, E. Tew-Kai, and J. F. Ternon. Mesoscale eddies influence distribution and aggregation patterns of micronekton in the Mozambique Channel. Marine Ecology Progress Series, 395:101–107, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p101.pdf.

Schmidt:2004:TLI

[SMM^{+}04] K. Schmidt, J. W. McClelland, E. Mente, J. P. Montoya, A. Atkinson, and M. Voss. Trophic-level interpretation based on d^{15}N values: implications of tissue-specific fractionation and amino acid composition. Marine Ecology Progress Series, 266:43–58, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p043.pdf.

Skjaeraasen:2009:SCC

[SMM^{+}09] J. E. Skjæraasen, I. Mayer, J. J. Meager, G. Rudolfsen, Ó. Karlsen, T. Haugland, and O. Kleven. Sperm characteristics and competitive ability in farmed and wild cod. Marine Ecology Progress Series, 375:219–228, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p219.pdf.

Sampey:2007:GGO

[SMMM07] A. Sampey, A. D. McKinnon, M. G. Meekan, and M. I. McCormick. Glimpse into guts: overview of the feeding of larvae of tropical shorefishes. Marine Ecology Progress Series, 339:243–257, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p243.pdf.

Stoner:2001:SEA

[SMP01] A. W. Stoner, J. P. Manderson, and J. P. Pessutti. Spatially explicit analysis of estuarine habitat for juvenile winter flounder: combining generalized additive models and geographic information systems. Marine Ecology Progress Series, 213:253–271, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/213/m213p253.pdf.

Silva:2009:ESA

[SMP^{+}09] M. A. Silva, S. Magalhães, R. Prieto, R. S. Santos, and P. S. Hammond. Estimating survival and abundance in a bottlenose dolphin population taking into account transience and temporary emigration. Marine Ecology Progress Series, 392:263–276, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p263.pdf.

Stevens:2006:FFP

[SMS06] P. W. Stevens, C. L. Montague, and K. J. Sulak. Fate of fish production in a seasonally flooded saltmarsh. Marine Ecology Progress Series, 327:267–277, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/327/m327p267.pdf.

Spilmont:2007:STV

[SMSD07] N. Spilmont, A. Migné, L. Seuront, and D. Davoult. Shortterm variability of intertidal benthic community production during emersion and the implication in annual budget calculation. Marine Ecology Progress Series, 333:95– 101, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/333/m333p095.pdf.

Staehr:2004:PSL

[SMSJ04] P. A. Stæhr, S. Markager, and K. Sand-Jensen. Pigment specific in vivo light absorption of phytoplankton from estuarine, coastal and oceanic waters. Marine Ecology Progress Series, 275:115–128, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p115.pdf.

Strasser:2008:TSE

[SMT08] C. A. Strasser, L. S. Mullineaux, and S. R. Thorrold. Temperature and salinity effects on elemental uptake in the shells of larval and juvenile softshell clams Mya arenaria. Marine Ecology Progress Series, 370:155–169, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/370/m370p155.pdf.

Stratoudakis:2003:CSN

[SMVF03] Y. Stratoudakis, A. Marçalo, C. Vale, and M. Falcão. Changes in seawater nutrient concentrations during purse seine fishing for sardine Sardina pilchardus off northern Portugal. Marine Ecology Progress Series, 265:235–242, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/265/
m265p235.pdf.

Shin:2002:URB

[SN02] H. C. Shin and S. Nicol. Using the relationship between eye diameter and body length to detect the effects of long-term starvation on Antarctic krill Euphausia superba. Marine Ecology Progress Series, 239:157–167, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/239/m239p157.pdf.

Sebastian:2004:APA

[SN04] M. Sebastián and F. X. Niell. Alkaline phosphatase activity in marine oligotrophic environments: implications of singlesubstrate addition assays for potential activity estimations. Marine Ecology Progress Series, 277:285–290, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/277/
m277p285.pdf.

Sakamaki:2006:DES

[SN06] T. Sakamaki and O. Nishimura. Dynamic equilibrium of sediment carbon content in an estuarine tidal flat: characterization and mechanisms. Marine Ecology Progress Series, 328:29–40, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p029.pdf.

Southall:2009:AME

[SN09] B. L. Southall and D. P. Nowacek. Acoustics in marine ecology: innovation in technology expands the use of sound in ocean science. Marine Ecology Progress Series, 395:1–3, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p001.pdf.

Steinberg:2004:PCD

[SNCP04] D. K. Steinberg, N. B. Nelson, C. A. Carlson, and A. Prusak. Production of chromophoric dissolved organic matter (CDOM) in the open ocean by zooplankton and the colonial cyanobacterium Trichodesmium spp. Marine Ecology Progress Series, 267:45–56, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/267/m267p045.pdf.

Swadling:2000:RLL

[SNGR00] K. M. Swadling, P. D. Nichols, J. A. E. Gibson, and D. A. Ritz. Role of lipid in the life cycles of ice-dependent and ice-independent populations of the copepod Paralabidocera antarctica. Marine Ecology Progress Series, 208:171–182, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/208/m208p171.pdf.

Satapoomin:2004:ASC

[SNH04] S. Satapoomin, T. G. Nielsen, and P. J. Hansen. Andaman Sea copepods: spatio-temporal variations in biomass and production, and role in the pelagic food web. Marine Ecology Progress Series, 274:99–122, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/274/m274p099.pdf.

Stewart:2003:OCA

[SNP03] R. R. Stewart, T. Noyce, and H. P. Possingham. Opportunity cost of ad hoc marine reserve design decisions: an example from South Australia. Marine Ecology Progress Series, 253:25–38, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p025.pdf.

Shulman:2005:FCB

[SNY^{+}05] G. E. Shulman, V. N. Nikolsky, T. V. Yuneva, G. S. Minyuk, V. Y. Shchepkin, A. M. Shchepkina, E. V. Ivleva, O. A. Yunev, I. S. Dobrovolov, F. Bingel, and A. E. Kideys. Fat content in Black Sea sprat as an indicator of fish food supply and ecosystem condition. Marine Ecology Progress Series, 293:201–212, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/293/m293p201.pdf.

Sogard:2001:GBR

[SO01a] S. M. Sogard and B. L. Olla. Growth and behavioral responses to elevated temperatures by juvenile sablefish Anoplopoma fimbria and the interactive role of food availability. Marine Ecology Progress Series, 217:121–134, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p121.pdf.

Stone:2001:SMD

[SO01b] R. P. Stone and C. E. O’Clair. Seasonal movements and distribution of Dungeness crabs Cancer magister in a glacial southeastern Alaska estuary. Marine Ecology Progress Series, 214:167–176, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/214/m214p167.pdf.

Sogard:2002:CCU

[SO02] S. M. Sogard and B. L. Olla. Contrasts in the capacity and underlying mechanisms for compensatory growth in two pelagic marine fishes. Marine Ecology Progress Series, 243: 165–177, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p165.pdf.

Seppala:2008:MAP

[SO08] J. Seppälä and K. Olli. Multivariate analysis of phytoplankton spectral in vivo fluorescence: estimation of phytoplankton biomass during a mesocosm study in the Baltic Sea. Marine Ecology Progress Series, 370:69–85, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/370/
m370p069.pdf.

Stige:2006:CCE

[SOB^{+}06] L. C. Stige, G. Ottersen, K. Brander, K. S. Chan, and C. h. r. Stenseth N. Cod and climate: effect of the North Atlantic Oscillation on recruitment in the North Atlantic. Marine Ecology Progress Series, 325:227–241, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p227.pdf.

Solan:2009:I

[Sol09] M. Solan. Introduction. Marine Ecology Progress Series, 396:166–168, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m396p166.pdf.

Sommer:2000:SMS

[Som00] U. Sommer. Scarcity of medium-sized phytoplankton in the northern Red Sea explained by strong bottom-up and weak top-down control. Marine Ecology Progress Series, 197:19– 25, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/197/m197p019.pdf.

Somerfield:2008:IBC

[Som08] P. J. Somerfield. Identification of the Bray–Curtis similarity index: Comment on Yoshioka (2008). Marine Ecology Progress Series, 372:303–306, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m372p303.pdf. See [Yos08b].

Strom:2006:CSG

[SOMM06] S. L. Strom, M. B. Olson, E. L. Macri, and C. W. Mord. Crossshelf gradients in phytoplankton community structure, nutrient utilization, and growth rate in the coastal Gulf of Alaska. Marine Ecology Progress Series, 328:75–92, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p075.pdf.

Sotka:2003:GCF

[Sot03] E. E. Sotka. Genetic control of feeding preference in the herbivorous amphipod Ampithoe longimana. Marine Ecology Progress Series, 256:305–310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p305.pdf.

Strzepek:2000:IIT

[SP00] R. F. Strzepek and N. M. Price. Influence of irradiance and temperature on the iron content of the marine diatom Thalassiosira weissflogii (Bacillariophyceae). Marine Ecology Progress Series, 206:107–117, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/206/m206p107.pdf.

Shaw:2001:PPI

[SP01a] P. J. Shaw and D. A. Purdie. Phytoplankton photosynthesisirradiance parameters in the near-shore UK coastal waters of the North Sea: temporal variation and environmental control. Marine Ecology Progress Series, 216:83–94, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p083.pdf.

Skov:2001:IEF

[SP01b] H. Skov and E. Prins. Impact of estuarine fronts on the dispersal of piscivorous birds in the German Bight. Marine Ecology Progress Series, 214:279–287, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/214/m214p279.pdf.

Sokolova:2001:PAH

[SP01c] I. M. Sokolova and H. O. Pörtner. Physiological adaptations to high intertidal life involve improved water conservation abilities and metabolic rate depression in Littorina saxatilis. Marine Ecology Progress Series, 224:171–186, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
224/m224p171.pdf.

Sommer:2002:MCA

[SP02] A. M. Sommer and H. O. Pörtner. Metabolic cold adaptation in the lugworm Arenicola marina: comparison of a North Sea and a White Sea population. Marine Ecology Progress Series, 240:171–182, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p171.pdf.

Saburova:2003:DAW

[SP03a] M. A. Saburova and I. G. Polikarpov. Diatom activity within soft sediments: behavioural and physiological processes. Marine Ecology Progress Series, 251:115–126, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/251/
m251p115.pdf.

Simonini:2003:LHD

[SP03b] R. Simonini and D. Prevedelli. Life history and demography of three populations of Ophryotrocha japonica (Polychaeta: Dorvilleidae). Marine Ecology Progress Series, 258: 171–180, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/258/m258p171.pdf.

Sponaugle:2004:LCP

[SP04] S. Sponaugle and D. Pinkard. Lunar cyclic population replenishment of a coral reef fish: shifting patterns following oceanic events. Marine Ecology Progress Series, 267:267– 280, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/267/m267p267.pdf.

Sheehy:2008:POQ

[SP08a] M. R. J. Sheehy and A. E. Prior. Progress on an old question for stock assessment of the edible crab Cancer pagurus. Marine Ecology Progress Series, 353:191–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/353/
m353p191.pdf.

Slattery:2008:IEB

[SP08b] M. Slattery and V. J. Paul. Indirect effects of bleaching on predator deterrence in the tropical Pacific soft coral Sinularia maxima. Marine Ecology Progress Series, 354:169– 179, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p169.pdf.

Simo:2000:BTD

[SPAMG00] R. Simó, C. Pedrós-Alió, G. Malin, and J. O. Grimalt. Biological turnover of DMS, DMSP and DMSO in contrasting opensea waters. Marine Ecology Progress Series, 203:1–11, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p001.pdf.

Soars:2009:OPP

[SPB09] N. A. Soars, T. A. A. Prowse, and M. Byrne. Overview of phenotypic plasticity in echinoid larvae, ‘Echinopluteus transversus’ type vs. typical echinoplutei. Marine Ecology Progress Series, 383:113–125, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/383/m383p113.pdf.

Seymour:2005:SDV

[SPBM05] J. R. Seymour, N. Patten, D. G. Bourne, and J. G. Mitchell. Spatial dynamics of virus-like particles and heterotrophic bacteria within a shallow coral reef system. Marine Ecology Progress Series, 288:1–8, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/288/m288p001.pdf.

Solidoro:2000:MGT

[SPC^{+}00] C. Solidoro, R. Pastres, D. Melaku Canu, M. Pellizzato, and R. Rossi. Modelling the growth of Tapes philippinarum in Northern Adriatic lagoons. Marine Ecology Progress Series, 199:137–148, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/199/m199p137.pdf.

Stat:2009:SCS

[SPCG09] M. Stat, X. Pochon, R. O. M. Cowie, and R. D. Gates. Specificity in communities of Symbiodinium in corals from Johnston Atoll. Marine Ecology Progress Series, 386:83– 96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/386/m386p083.pdf.

Sperduto:2003:SRA

[SPD03] M. B. Sperduto, S. P. Powers, and M. Donlan. Scaling restoration to achieve quantitative enhancement of loon, seaduck, and other seabird populations. Marine Ecology Progress Series, 264:221–232, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/264/m264p221.pdf.

Suhr:2003:SFB

[SPGS03] S. B. Suhr, D. W. Pond, A. J. Gooday, and C. R. Smith. Selective feeding by benthic Foraminifera on phytodetritus on the western Antarctic Peninsula shelf: evidence from fatty acid biomarker analysis. Marine Ecology Progress Series, 262:153–162, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p153.pdf.

Shultz:2009:TBR

[SPH^{+}09] M. T. Shultz, J. F. Piatt, A. M. A. Harding, A. B. Kettle, and T. I. Van Pelt. Timing of breeding and reproductive performance in murres and kittiwakes reflect mismatched seasonal prey dynamics. Marine Ecology Progress Series, 393:247–258, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m393p247.pdf.

Serra:2007:CIS

[SPI^{+}07] I. A. Serra, G. Procaccini, M. C. Intrieri, M. Migliaccio, S. Mazzuca, and A. M. Innocenti. Comparison of ISSR and SSR markers for analysis of genetic diversity in the seagrass Posidonia oceanica. Marine Ecology Progress Series, 338: 71–79, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p071.pdf.

Sancho:2000:PPR

[SPL00] G. Sancho, C. W. Petersen, and P. S. Lobel. Predator– prey relations at a spawning aggregation site of coral reef fishes. Marine Ecology Progress Series, 203:275–288, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
203/m203p275.pdf.

Steer:2003:BCS

[SPM03] M. A. Steer, G. T. Pecl, and N. A. Moltschaniwskyj. Are bigger calamary Sepioteuthis australis hatchlings more likely to survive? A study based on statolith dimensions. Marine Ecology Progress Series, 261:175–182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/261/
m261p175.pdf.

Seeyave:2009:NNA

[SPP^{+}09] S. Seeyave, T. A. Probyn, G. C. Pitcher, M. I. Lucas, and D. A. Purdie. Nitrogen nutrition in assemblages dominated by Pseudo-nitzschia spp., Alexandrium catenella and Dinophysis acuminata off the west coast of South Africa. Marine Ecology Progress Series, 379:91–107, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/379/m379p091.pdf.

Sommer:2002:EIS

[SPRK02] S. Sommer, O. Pfannkuche, D. Rickert, and A. Kähler. Ecological implications of surficial marine gas hydrates for the associated small-sized benthic biota at the Hydrate Ridge (Cascadia Convergent Margin, NE Pacific). Marine Ecology Progress Series, 243:25–38, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/243/m243p025.pdf.

Staehr:2000:ISM

[SPT^{+}00] P. A. Stæhr, M. F. Pedersen, M. S. Thomsen, T. Wernberg, and D. Krause-Jensen. Invasion of Sargassum muticum in Limfjorden (Denmark) and its possible impact on the indigenous macroalgal community. Marine Ecology Progress Series, 207:79–88, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p079.pdf.

Sutherland:2004:DIC

[SPT04] K. P. Sutherland, J. W. Porter, and C. Torres. Disease and immunity in Caribbean and Indo–Pacific zooxanthellate corals. Marine Ecology Progress Series, 266:273–302, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p273.pdf.

Stieglitz:2001:TMP

[SR01] T. Stieglitz and P. V. Ridd. Trapping of mangrove propagules due to density-driven secondary circulation in the Normanby River estuary, NE Australia. Marine Ecology Progress Series, 211:131–142, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/211/m211p131.pdf.

Sherwood:2003:ISF

[SR03] G. D. Sherwood and G. A. Rose. Influence of swimming form on otolith δ ^{13}C in marine fish. Marine Ecology Progress Series, 258:283–289, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p283.pdf.

Schiavini:2004:LDL

[SR04a] A. Schiavini and A. R. Rey. Long days, long trips: foraging ecology of female rockhopper penguins Eudyptes chrysocome chrysocome at Tierra del Fuego. Marine Ecology Progress Series, 275:251–262, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p251.pdf.

Smith:2004:PPL

[SR04b] N. F. Smith and G. M. Ruiz. Phenotypic plasticity in the life history of the mangrove snail Cerithidea scalariformis. Marine Ecology Progress Series, 284:195–209, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p195.pdf.

Sakamaki:2008:ESR

[SR08a] T. Sakamaki and J. S. Richardson. Effects of small rivers on chemical properties of sediment and diets for primary consumers in estuarine tidal flats. Marine Ecology Progress Series, 360:13–24, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p013.pdf.

Semeniuk:2008:CGL

[SR08b] C. A. D. Semeniuk and K. D. Rothley. Costs of group-living for a normally solitary forager: effects of provisioning tourism on southern stingrays Dasyatis americana. Marine Ecology Progress Series, 357:271–282, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/357/m357p271.pdf.

Sakamaki:2009:DRT

[SR09] T. Sakamaki and J. S. Richardson. Dietary responses of tidal flat macrobenthos to reduction of benthic microalgae: a test for potential use of allochthonous organic matter. Marine Ecology Progress Series, 386:107–113, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/386/
m386p107.pdf.

Staniland:2004:CIS

[SRB04] I. J. Staniland, K. Reid, and I. L. Boyd. Comparing individual and spatial influences on foraging behaviour in Antarctic fur seals Arctocephalus gazella. Marine Ecology Progress Series, 275:263–274, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p263.pdf.

Stoner:2000:EAE

[SRC00] A. W. Stoner and M. Ray-Culp. Evidence for Allee effects in an over-harvested marine gastropod: density-dependent mating and egg production. Marine Ecology Progress Series, 202:297–302, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p297.pdf.

Skilleter:2005:LPT

[SRDG05] G. A. Skilleter, B. D. Russell, B. M. Degnan, and M. J. Garson. Living in a potentially toxic environment: comparisons of endofauna in two congeneric sponges from the Great Barrier Reef. Marine Ecology Progress Series, 304: 67–75, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/304/m304p067.pdf.

Saborido-Rey:2007:GJT

[SRDPT^{+}07] F. Saborido-Rey, R. Domínguez-Petit, J. Tomás, B. Morales-Nin, and A. Alonso-Fernandez. Growth of juvenile turbot in response to food pellets contaminated by fuel oil from the tanker ‘Prestige’. Marine Ecology Progress Series, 345: 271–279, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/345/m345p271.pdf.

Schumann:2001:SPA

[SRGS01] R. Schumann, D. Rentsch, S. Görs, and U. Schiewer. Seston particles along a eutrophication gradient in coastal waters of the Southern Baltic Sea: Significance of detritus and transparent mucoid material. Marine Ecology Progress Series, 218:17–31, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/218/m218p017.pdf.

Svedang:2007:DNH

[SRJ07a] H. Svedäng, D. Righton, and P. Jonsson. Defining ‘natal homing’ in marine fish populations; need for inference in fishery science: reply to Bradbury & Laurel (2007). Marine Ecology Progress Series, 349:309–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/349/m349p309.pdf.

Svedang:2007:MBA

[SRJ07b] H. Svedäng, D. Righton, and P. Jonsson. Migratory behaviour of Atlantic cod Gadus morhua: natal homing is the prime stock-separating mechanism. Marine Ecology Progress Series, 345:1–12, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See comment [BL07].

Schaffler:2009:PLA

[SRJ09a] J. J. Schaffler, C. S. Reiss, and C. M. Jones. Patterns of larval Atlantic croaker ingress into Chesapeake Bay, USA. Marine Ecology Progress Series, 378:187–197, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/378/
m378p187.pdf.

Schaffler:2009:SVO

[SRJ09b] J. J. Schaffler, C. S. Reiss, and C. M. Jones. Spatial variation in otolith chemistry of Atlantic croaker larvae in the Mid-Atlantic Bight. Marine Ecology Progress Series, 382: 185–195, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m382p185.pdf.

Sahling:2002:MCS

[SRL^{+}02] H. Sahling, D. Rickert, R. W. Lee, P. Linke, and E. Suess. Macrofaunal community structure and sulfide flux at gas hydrate deposits from the Cascadia convergent margin, NE Pacific. Marine Ecology Progress Series, 231:121–138, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/231/m231p121.pdf.

Styrishave:2004:VEL

[SRLA04] B. Styrishave, K. Rewitz, T. Lund, and O. Andersen. Variations in ecdysteroid levels and cytochrome p 450 expression during moult and reproduction in male shore crabs Carcinus maenas. Marine Ecology Progress Series, 274:215– 224, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/274/m274p215.pdf.

Seminoff:2002:HRG

[SRN02] J. A. Seminoff, A. Resendiz, and W. J. Nichols. Home range of green turtles Chelonia mydas at a coastal foraging area in the Gulf of California, Mexico. Marine Ecology Progress Series, 242:253–265, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/242/m242p253.pdf.

Solan:2006:MBE

[SRP^{+}06] M. Solan, D. G. Raffaelli, D. M. Paterson, P. C. L. White, and G. J. Pierce. Marine biodiversity and ecosystem function: empirical approaches and future research needs. Marine Ecology Progress Series, 311:175–309, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m311_TS.pdf.

Stevens:2006:CPR

[SRR06] D. Stevens, A. J. Richardson, and P. C. Reid. Continuous plankton recorder database: evolution, current uses and future directions. Marine Ecology Progress Series, 316:247– 255, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m316p247.pdf.

Stark:2004:IAW

[SRS04] J. S. Stark, M. J. Riddle, and S. D. A. Smith. Influence of an Antarctic waste dump on recruitment to nearshore marine soft-sediment assemblages. Marine Ecology Progress Series, 276:53–70, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p053.pdf.

Smith:2005:CLS

[SRS05] J. E. Smith, J. W. Runcie, and C. M. Smith. Characterization of a large-scale ephemeral bloom of the green alga Cladophora sericea on the coral reefs of West Maui, Hawai’i. Marine Ecology Progress Series, 302:77–91, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p077.pdf.

Secor:2001:IRE

[SRZZ01] D. H. Secor, J. R. Rooker, E. Zlokovitz, and V. S. Zdanowicz. Identification of riverine, estuarine, and coastal contingents of Hudson River striped bass based upon otolith elemental fingerprints. Marine Ecology Progress Series, 211:245–253, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p245.pdf.

Sabates:2000:IIV

[SS00a] A. Sabatés and E. Saiz. Intra- and interspecific variability in prey size and niche breadth of myctophiform fish larvae. Marine Ecology Progress Series, 201:261–271, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p261.pdf.

Searcy:2000:VLG

[SS00b] S. P. Searcy and S. Sponaugle. Variable larval growth in a coral reef fish. Marine Ecology Progress Series, 206:213–226, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p213.pdf.

Schnetzer:2002:ATP

[SS02a] A. Schnetzer and D. K. Steinberg. Active transport of particulate organic carbon and nitrogen by vertically migrating zooplankton in the Sargasso Sea. Marine Ecology Progress Series, 234:71–84, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p071.pdf.

Spicer:2002:DVM

[SS02b] J. I. Spicer and J. O. Strömberg. Diel vertical migration and the haemocyanin of krill Meganyctiphanes norvegica. Marine Ecology Progress Series, 238:153–162, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p153.pdf.

Silva:2003:DVP

[SS03] J. Silva and R. Santos. Daily variation patterns in seagrass photosynthesis along a vertical gradient. Marine Ecology Progress Series, 257:37–44, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/257/m257p037.pdf.

Shaw:2004:MPC

[SS04a] P. W. Shaw and W. H. H. Sauer. Multiple paternity and complex fertilisation dynamics in the squid Loligo vulgaris reynaudii. Marine Ecology Progress Series, 270:173–179, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/270/m270p173.pdf.

Smith:2004:ECM

[SS04b] J. Smith and S. E. Shackley. Effects of a commercial mussel Mytilus edulis lay on a sublittoral, soft sediment benthic community. Marine Ecology Progress Series, 282:185– 191, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/282/m282p185.pdf.

Sumi:2005:RGS

[SS05] C. B. T. Sumi and R. E. Scheibling. Role of grazing by sea urchins Strongylocentrotus droebachiensis in regulating the invasive alga Codium fragile ssp. tomentosoides in Nova Scotia.

Marine Ecology Progress Series, 292:203–212, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p203.pdf.

Skovgaard:2006:SOR

[SS06] A. Skovgaard and E. Saiz. Seasonal occurrence and role of protistan parasites in coastal marine zooplankton. Marine Ecology Progress Series, 327:37–49, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p037.pdf.

Sherr:2007:HDS

[SS07] E. B. Sherr and B. F. Sherr. Heterotrophic dinoflagellates: a significant component of microzooplankton biomass and major grazers of diatoms in the sea. Marine Ecology Progress Series, 352:187–197, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/352/m352p187.pdf.

Sauchyn:2009:FPS

[SS09a] L. K. Sauchyn and R. E. Scheibling. Fecal production by sea urchins in native and invaded algal beds. Marine Ecology Progress Series, 396:35–48, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/396/m396p035.pdf.

Shanks:2009:PLC

[SS09b] A. L. Shanks and R. K. Shearman. Paradigm lost? Cross-shelf distributions of intertidal invertebrate larvae are unaffected by upwelling or downwelling. Marine Ecology Progress Series, 385:189–204, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/385/m385p189.pdf.

Shima:2009:SVL

[SS09c] J. S. Shima and S. E. Swearer. Spatially variable larval histories may shape recruitment rates of a temperate reef fish. Marine Ecology Progress Series, 394:223–229, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/394/
m394p223.pdf.

Sebastian:2002:HMP

[SSB02] C. Ruiz Sebastián, C. N. Steffani, and G. M. Branch. Homing and movement patterns of a South African limpet Scutellastra argenvillei in an area invaded by an alien mussel Mytilus galloprovincialis. Marine Ecology Progress Series, 243:111– 122, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p111.pdf.

Saba:2009:EDR

[SSB09] G. K. Saba, D. K. Steinberg, and D. A. Bronk. Effects of diet on release of dissolved organic and inorganic nutrients by the copepod Acartia tonsa. Marine Ecology Progress Series, 386:147–161, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m386p147.pdf.

Springer:2002:IBT

[SSBG02] J. J. Springer, S. E. Shumway, J. A. M. Burkholder, and H. B. Glasgow. Interactions between the toxic estuarine dinoflagellate Pfiesteria piscicida and two species of bivalve molluscs. Marine Ecology Progress Series, 245:1–10, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p001.pdf.

Starr:2007:SDN

[SSBZ07] R. M. Starr, E. Sala, E. Ballesteros, and M. Zabala. Spatial dynamics of the Nassau grouper Epinephelus striatus in a Caribbean atoll. Marine Ecology Progress Series, 343:239– 249, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p239.pdf.

Staats:2000:OPD

[SSdWM00] N. Staats, L. J. Stal, B. de Winder, and L. R. Mur. Oxygenic photosynthesis as driving process in exopolysaccharide production of benthic diatoms. Marine Ecology Progress Series, 193:261–269, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p261.pdf.

Schroder:2000:PCC

[SSG^{+}00] H. C. Schröder, K. Shostak, V. Gamulin, M. Lacorn, A. Skorokhod, V. Kavsan, and W. E. G. Müller. Purification, cDNA cloning and expression of a cadmium-inducible cysteine-rich metallothionein-like protein from the marine sponge Suberites domuncula. Marine Ecology Progress Series, 200:149–157, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/200/m200p149.pdf.

Schoor:2008:PDP

[SSGB^{+}08] A. Schoor, U. Selig, U. Geiß-Brunschweiger, R. Schaible, R. Schumann, B. Schubert, P. Feuerpfeil, M. Hagemann, and H. Schubert. Phytoplankton diversity and photosynthetic acclimation along a longitudinal transect through a shallow estuary in summer. Marine Ecology Progress Series, 364:31–46, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p031.pdf.

Schulz:2000:PSF

[SSGJ00] H. N. Schulz, B. Strotmann, V. A. Gallardo, and B. B. Jørgensen. Population study of the filamentous sulfur bacteria Thioploca spp. off the Bay of Concepción, Chile. Marine Ecology Progress Series, 200:117–126, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/200/m200p117.pdf.

Stuart:2000:BOC

[SSH^{+}00] V. Stuart, S. Sathyendranath, E. J. H. Head, T. Platt, B. Irwin, and H. Maass. Bio-optical characteristics of diatom and prymnesiophyte populations in the Labrador Sea. Marine Ecology Progress Series, 201:91–106, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
201/m201p091.pdf.

Sampei:2004:FSP

[SSH^{+}04] M. Sampei, H. Sasaki, H. Hattori, M. Fukuchi, and B. T. Hargrave. Fate of sinking particles, especially fecal pellets, within the epipelagic zone in the North Water (NOW) polynya of northern Baffin Bay. Marine Ecology Progress Series, 278:17–25, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p017.pdf.

Sunda:2009:AUG

[SSH09] W. G. Sunda, K. W. Shertzer, and D. R. Hardison. Ammonium uptake and growth models in marine diatoms: Monod and Droop revisited. Marine Ecology Progress Series, 386: 29–41, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/386/m386p029.pdf.

Schlacher:2007:RDS

[SSHW^{+}07] T. A. Schlacher, M. A. Schlacher-Hoenlinger, A. Williams, F. Althaus, J. N. A. Hooper, and R. Kloser. Richness and distribution of sponge megabenthos in continental margin canyons off southeastern Australia. Marine Ecology Progress Series, 340:73–88, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/340/m340p073.pdf.

Sejr:2002:GPH

[SSJ^{+}02] M. K. Sejr, M. K. Sand, K. T. Jensen, J. K. Petersen, P. B. Christensen, and S. Rysgaard. Growth and production of Hiatella arctica (Bivalvia) in a high-Arctic fjord (Young Sound, Northeast Greenland). Marine Ecology Progress Series, 244:163–169, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/244/m244p163.pdf.

Sancho:2000:EID

[SSL00a] G. Sancho, A. R. Solow, and P. S. Lobel. Environmental influences on the diel timing of spawning in coral reef fishes. Marine Ecology Progress Series, 206:193–212, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p193.pdf.

Sclafani:2000:OCB

[SSL00b] M. Sclafani, G. Stirling, and W. C. Leggett. Osmotic condition, buoyancy change and mortality in larval cod Gadus morhua. A bioassay for assessing near-term mortality. Marine Ecology Progress Series, 193:157–166, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p157.pdf.

Sonin:2007:NDF

[SSL^{+}07] O. Sonin, E. Spanier, D. Levi, B. Patti, P. Rizzo, and M. G. Andreoli. Nanism (dwarfism) in fish: a comparison between red mullet Mullus barbatus from the southeastern and the central Mediterranean. Marine Ecology Progress Series, 343: 221–228, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p221.pdf.

Sellner:2001:ENS

[SSLM01] K. G. Sellner, S. G. Sellner, R. V. Lacouture, and R. E. Magnien. Excessive nutrients select for dinoflagellates in the stratified Patapsco River estuary: Margalef reigns. Marine Ecology Progress Series, 220:93–102, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/220/m220p093.pdf.

Sugeha:2001:VOD

[SSM^{+}01] H. Y. Sugeha, A. Shinoda, M. Marui, T. Arai, and K. Tsukamoto. Validation of otolith daily increments in the tropical eel Anguilla marmorata. Marine Ecology Progress Series, 220:291–294, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p291.pdf.

Seymour:2005:MSS

[SSM05] J. R. Seymour, L. Seuront, and J. G. Mitchell. Microscale and small-scale temporal dynamics of a coastal planktonic microbial community. Marine Ecology Progress Series, 300: 21–37, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/300/m300p021.pdf.

Sara:2006:USI

[SSMM06] G. Sarà, D. Scilipoti, M. Milazzo, and A. Modica. Use of stable isotopes to investigate dispersal of waste from fish farms as a function of hydrodynamics. Marine Ecology Progress Series, 313:261–270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/313/m313p261.pdf.

Salvanes:2004:SPC

[SSN04] A. G. V. Salvanes, J. E. Skjæraasen, and T. Nilsen. Subpopulations of coastal cod with different behaviour and lifehistory strategies. Marine Ecology Progress Series, 267:241– 251, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/267/m267p241.pdf.

Sathyendranath:2009:CCR

[SSN^{+}09] S. Sathyendranath, V. Stuart, A. Nair, K. Oka, T. Nakane, H. Bouman, M. H. Forget, H. Maass, and T. Platt. Carbonto-chlorophyll ratio and growth rate of phytoplankton in the sea. Marine Ecology Progress Series, 383:73–84, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m383p073.pdf.

Sekiguchi:2001:ADK

[SSO^{+}01] K. Sekiguchi, S. Sato, T. Ogata, S. Kaga, and M. Kodama. Accumulation and depuration kinetics of paralytic shellfish toxins in the scallop Patinopecten yessoensis fed Alexandrium tamarense. Marine Ecology Progress Series, 220:213– 218, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/220/m220p213.pdf.

Sims:2003:SMB

[SSR^{+}03] D. W. Sims, E. J. Southall, A. J. Richardson, P. C. Reid, and J. D. Metcalfe. Seasonal movements and behaviour of basking sharks from archival tagging: no evidence of winter hibernation. Marine Ecology Progress Series, 248:187– 196, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p187.pdf.

Sabates:2009:SEC

[SSR^{+}09] A. Sabatés, J. Salat, V. Raya, M. Emelianov, and M. Segura-Noguera. Spawning environmental conditions of Sardinella aurita at the northern limit of its distribution range, the western Mediterranean. Marine Ecology Progress Series, 385: 227–236, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/385/m385p227.pdf.

Sherwood:2005:REA

[SSRG05] O. A. Sherwood, D. B. Scott, M. J. Risk, and T. P. Guilderson. Radiocarbon evidence for annual growth rings in the deep-sea octocoral Primnoa resedaeformis. Marine Ecology Progress Series, 301:129–134, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/301/m301p129.pdf.

Sommer:2005:LFS

[SSS^{+}05] F. Sommer, A. Saage, B. Santer, T. Hansen, and U. Sommer. Linking foraging strategies of marine calanoid copepods to patterns of nitrogen stable isotope signatures in a mesocosm study. Marine Ecology Progress Series, 286:99– 106, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/286/m286p099.pdf.

Sourisseau:2006:KAS

[SSS06] M. Sourisseau, Y. Simard, and F. J. Saucier. Krill aggregation in the St. Lawrence system, and supply of krill to the whale feeding grounds in the estuary from the gulf. Marine Ecology Progress Series, 314:257–270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/314/m314p257.pdf.

Schroeder:2009:WPC

[SSS^{+}09] I. D. Schroeder, W. J. Sydeman, N. Sarkar, S. A. Thompson, S. J. Bograd, and F. B. Schwing. Winter pre-conditioning of seabird phenology in the California Current. Marine Ecology Progress Series, 393:211–223, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m393p211.pdf.

Sommer:2000:GMK

[SSSV00] F. Sommer, H. Stibor, U. Sommer, and B. Velimirov. Grazing by mesozooplankton from Kiel Bight, Baltic Sea, on different sized algae and natural seston size fractions. Marine Ecology Progress Series, 199:43–53, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/199/m199p043.pdf.

Standish:2008:UON

[SSW08] J. D. Standish, M. Sheehy, and R. R. Warner. Use of otolith natal elemental signatures as natural tags to evaluate connectivity among open-coast fish populations. Marine Ecology Progress Series, 356:259–268, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/356/m356p259.pdf.

Shen:2002:FMC

[ST02] K. N. Shen and W. N. Tzeng. Formation of a metamorphosis check in otoliths of the amphidromous goby Sicyopterus japonicus. Marine Ecology Progress Series, 228:205– 211, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/228/m228p205.pdf.

Sedlacek:2006:ECS

[ST06a] L. Sedlacek and D. Thistle. Emergence on the continental shelf: differences among species and between microhabitats. Marine Ecology Progress Series, 311:29–36, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/311/
m311p029.pdf.

Shoji:2006:GSS

[ST06b] J. Shoji and M. Tanaka. Growth-selective survival in piscivorous larvae of Japanese Spanish mackerel Scomberomorus niphonius: early selection and significance of ichthyoplankton prey supply. Marine Ecology Progress Series, 321:245– 254, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p245.pdf.

Shoji:2007:DDP

[ST07] J. Shoji and M. Tanaka. Density-dependence in post-recruit Japanese seaperch Lateolabrax japonicus in the Chikugo River, Japan. Marine Ecology Progress Series, 334:255–262, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/334/m334p255.pdf.

Skov:2008:RFS

[ST08] H. Skov and F. Thomsen. Resolving fine-scale spatio-temporal dynamics in the harbour porpoise Phocoena phocoena. Marine Ecology Progress Series, 373:173–186, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/373/

m373p173.pdf. See comment [TW10].

Shreeve:2005:RPC

[STA^{+}05] R. S. Shreeve, G. A. Tarling, A. Atkinson, P. Ward, C. Goss, and J. Watkins. Relative production of Calanoides acutus (Copepoda: Calanoida) and Euphausia superba (Antarctic krill) at South Georgia, and its implications at wider scales. Marine Ecology Progress Series, 298:229–239, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p229.pdf.

Stark:2008:PHT

[Sta08] J. S. Stark. Patterns of higher taxon colonisation and development in sessile marine benthic assemblages at Casey Station, Antarctica, and their use in environmental monitoring. Marine Ecology Progress Series, 365:77–89, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p077.pdf.

Stallings:2009:PIR

[Sta09] C. D. Stallings. Predator identity and recruitment of coralreef fishes: indirect effects of fishing. Marine Ecology Progress Series, 383:251–259, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/383/m383p251.pdf.

Soto:2006:CDM

[STAS06] K. H. Soto, A. W. Trites, and M. Arias-Schreiber. Changes in diet and maternal attendance of South American sea lions indicate changes in the marine environment and prey abundance. Marine Ecology Progress Series, 312:277–290, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/312/m312p277.pdf.

Shiao:2001:DPG

[STCJ01] J. C. Shiao, W. N. Tzeng, A. Collins, and D. J. Jellyman. Dispersal pattern of glass eel stage of Anguilla australis revealed by otolith growth increments. Marine Ecology Progress Series, 219:241–250, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p241.pdf.

Sahade:2004:REA

[STE04a] R. Sahade, M. Tatián, and G. B. Esnal. Reproductive ecology of the ascidian Cnemidocarpa verrucosa at Potter Cove, South Shetland Islands, Antarctica. Marine Ecology Progress Series, 272:131–140, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/272/m272p131.pdf.

Steen:2004:ICB

[Ste04b] H. Steen. Interspecific competition between Enteromorpha (Ulvales: Chlorophyceae) and Fucus (Fucales: Phaeophyceae) germlings: effects of nutrient concentration, temperature, and settlement density. Marine Ecology Progress Series, 278:89– 101, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/278/m278p089.pdf.

Sherwood:2009:MCT

[STF^{+}09] O. A. Sherwood, R. E. Thresher, S. J. Fallon, D. M. Davies, and T. W. Trull. Multi-century time-series of ^{15}N and ^{14}C in bamboo corals from deep Tasmanian seamounts: evidence for stable oceanographic conditions. Marine Ecology Progress Series, 397:209–218, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m397p209.pdf.

Soreide:2006:SPE

[STH^{+}06] J. E. Søreide, T. Tamelander, H. Hop, K. A. Hobson, and I. Johansen. Sample preparation effects on stable C and N isotope values: a comparison of methods in Arctic marine food web studies. Marine Ecology Progress Series, 328: 17–28, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p017.pdf.

Sato:2003:SSH

[STI03] R. Sato, Y. Tanaka, and T. Ishimaru. Species-specific house productivity of appendicularians. Marine Ecology Progress Series, 259:163–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/259/m259p163.pdf.

Stead:2003:AEI

[STJ03] R. A. Stead, R. J. Thompson, and J. R. Jaramillo. Absorption efficiency, ingestion rate, gut passage time and scope for growth in suspension- and deposit-feeding Yoldia hyperborea. Marine Ecology Progress Series, 252:159–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/252/
m252p159.pdf.

Stierhoff:2006:ERJ

[STM06] K. L. Stierhoff, T. E. Targett, and K. L. Miller. Ecophysiological responses of juvenile summer and winter flounder to hypoxia: experimental and modeling analyses of effects on estuarine nursery quality. Marine Ecology Progress Series, 325:255–266, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/325/m325p255.pdf.

Stoner:2003:WCE

[Sto03] A. W. Stoner. What constitutes essential nursery habitat for a marine species? A case study of habitat form and function for queen conch. Marine Ecology Progress Series, 257:275–289, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/257/m257p275.pdf.

Schmidt:2004:MCR

[STPA04] K. Schmidt, G. A. Tarling, N. Plathner, and A. Atkinson. Moult cycle-related changes in feeding rates of larval krill Meganyctiphanes norvegica and Thysanoessa spp. Marine Ecology Progress Series, 281:131–143, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p131.pdf.

Strasser:2002:REP

[Str02] M. Strasser. Reduced epibenthic predation on intertidal bivalves after a severe winter in the European Wadden Sea. Marine Ecology Progress Series, 241:113–123, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/241/
m241p113.pdf.

Suthers:2006:DNI

[STRB06] I. M. Suthers, C. T. Taggart, D. Rissik, and M. E. Baird. Day and night ichthyoplankton assemblages and zooplankton biomass size spectrum in a deep ocean island wake. Marine Ecology Progress Series, 322:225–238, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/322/
m322p225.pdf.

Savage:2002:FSM

[STRD02] A. M. Savage, H. Trapido-Rosenthal, and A. E. Douglas. On the functional significance of molecular variation in Symbiodinium, the symbiotic algae of Cnidaria: photosynthetic response to irradiance. Marine Ecology Progress Series, 244: 27–37, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/244/m244p027.pdf.

Sigler:2009:SSL

[STV^{+}09] M. F. Sigler, D. J. Tollit, J. J. Vollenweider, J. F. Thedinga, D. J. Csepp, J. N. Womble, M. A. Wong, M. J. Rehberg, and A. W. Trites. Steller sea lion foraging response to seasonal changes in prey availability. Marine Ecology Progress Series, 388:243–261, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/388/m388p243.pdf.

Soltwedel:2001:BAB

[SV01] T. Soltwedel and K. Vopel. Bacterial abundance and biomass in response to organism-generated habitat heterogeneity in deep-sea sediments. Marine Ecology Progress Series, 219: 291–298, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/219/m219p291.pdf.

Sheader:2007:TSV

[SV07] M. Sheader and C. L. Van Dover. Temporal and spatial variation in the reproductive ecology of the vent-endemic amphipod Ventiella sulfuris in the eastern Pacific. Marine Ecology Progress Series, 331:181–194, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/331/m331p181.pdf.

Seuront:2008:ISV

[SV08] L. Seuront and D. Vincent. Increased seawater viscosity, Phaeocystis globosa spring bloom and Temora longicornis feeding and swimming behaviours. Marine Ecology Progress Series, 363:131–145, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/363/m363p131.pdf.

Solan:2009:TSMb

[SvAHH09a] M. Solan, P. van Avesaath, H. Hummel, and C. H. R. Heip. Theme section: Marine biodiversity: current understanding and future research. Marine Ecology Progress Series, 396: 166–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m396_ThemeSection.pdf.

Solan:2009:TSMa

[SvAHH09b] Martin Solan, Pim van Avesaath, Herman Hummel, and Carlo Heip. Theme section: Marine biodiversity: current understanding and future research. Marine Ecology Progress Series, 396:166–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m396_ThemeSection.pdf.

Souchu:2001:ISF

[SVC^{+}01] P. Souchu, A. Vaquer, Y. Collos, S. Landrein, J. M. Deslous-Paoli, and B. Bibent. Influence of shellfish farming activities on the biogeochemical composition of the water column in Thau lagoon. Marine Ecology Progress Series, 218:141– 152, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/218/m218p141.pdf.

Svensen:2002:EVF

[Sve02] C. Svensen. Eutrophication and vertical flux: a critical evaluation of silicate addition. Marine Ecology Progress Series, 240:21–26, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/240/m240p021.pdf.

Struyf:2005:BST

[SVG^{+}05] E. Struyf, S. Van Damme, B. Gribsholt, J. J. Middelburg, and P. Meire. Biogenic silica in tidal freshwater marsh sediments and vegetation (Schelde estuary, Belgium). Marine Ecology Progress Series, 303:51–60, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/303/m303p051.pdf.

Selkoe:2007:EEC

[SVG07] K. A. Selkoe, A. Vogel, and S. D. Gaines. Effects of ephemeral circulation on recruitment and connectivity of nearshore fish populations spanning Southern and Baja California. Marine Ecology Progress Series, 351:209–220, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p209.pdf.

Stibor:2004:CCN

[SVL^{+}04] H. Stibor, O. Vadstein, B. Lippert, W. Roederer, and Y. Olsen. Calanoid copepods and nutrient enrichment determine population dynamics of the appendicularian Oikopleura dioica: a mesocosm experiment. Marine Ecology Progress Series, 270:209–215, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/270/m270p209.pdf.

Silverman:2004:NNF

[SVN04] E. D. Silverman, R. R. Veit, and G. A. Nevitt. Nearest neighbors as foraging cues: information transfer in a patchy environment. Marine Ecology Progress Series, 277: 25–36, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/277/m277p025.pdf.

Smallegange:2009:SDS

[SvNvdMvdV09] I. M. Smallegange, C. G. E. van Noordwijk, J. van der Meer, and H. W. van der Veer. Spatial distribution of shore crabs Carcinus maenas in an intertidal environment in relation to their morphology, prey availability and competition. Marine Ecology Progress Series, 392:143–155, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/392/
m392p143.pdf.

Sin:2002:EMAb

[SW02a] Y. Sin and R. L. Wetzel. Ecosystem modeling analysis of size-structured phytoplankton dynamics in the York River estuary, Virginia (USA). II. Use of a plankton ecosystem model for investigating controlling factors on phytoplankton and nutrient dynamics. Marine Ecology Progress Series, 228:91–101, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p091.pdf.

Sin:2002:EMAa

[SW02b] Y. Sin and R. L. Wetzel. Ecosystem modeling analysis of size-structured phytoplankton dynamics in the York River estuary, Virginia (USA). I. Development of a plankton ecosystem model with explicit feedback controls and hydrodynamics. Marine Ecology Progress Series, 228:75–90, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p075.pdf.

Siddon:2003:ICL

[SW03] C. E. Siddon and J. D. Witman. Influence of chronic, lowlevel hydrodynamic forces on subtidal community structure. Marine Ecology Progress Series, 261:99–110, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/261/
m261p099.pdf.

Seebaugh:2004:IMB

[SW04] D. R. Seebaugh and W. G. Wallace. Importance of metalbinding proteins in the partitioning of Cd and Zn as trophically available metal (TAM) in the brine shrimp Artemia franciscana. Marine Ecology Progress Series, 272:215–230, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/272/m272p215.pdf.

Sassa:2008:TOC

[SW08a] S. Sassa and Y. Watabe. Threshold, optimum and critical geoenvironmental conditions for burrowing activity of sand bubbler crab, Scopimera globosa. Marine Ecology Progress Series, 354:191–199, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m354p191.pdf.

Smee:2008:HPR

[SW08b] D. L. Smee and M. J. Weissburg. Heightened prey responses in risky habitats: predation pressure increases prey sensitivity to predation risk. Marine Ecology Progress Series, 363: 39–50, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/363/m363p039.pdf.

Sweetman:2008:RAM

[SW08c] A. K. Sweetman and U. Witte. Response of an abyssal macrofaunal community to a phytodetrital pulse. Marine Ecology Progress Series, 355:73–84, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/355/m355p073.pdf.

Smale:2009:SDS

[SW09] D. A. Smale and T. Wernberg. Satellite-derived SST data as a proxy for water temperature in nearshore benthic ecology. Marine Ecology Progress Series, 387:27–37, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p027.pdf.

Swanson:2004:ESA

[Swa04] C. A. Swanson. Effect of substrate availability and conspecific cues on communal oviposition in the apple murex snail Phyllonotus pomum. Marine Ecology Progress Series, 275:175–184, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/275/m275p175.pdf.

Spagnol:2002:EFM

[SWD^{+}02] S. Spagnol, E. Wolanski, E. Deleersnijder, R. Brinkman, F. McAllister, B. Cushman-Roisin, and E. Hanert. An error frequently made in the evaluation of advective transport in two-dimensional Lagrangian models of advection-diffusion in coral reef waters. Marine Ecology Progress Series, 235: 299–302, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/235/m235p299.pdf.

Sathyendranath:2004:DDO

[SWD^{+}04] S. Sathyendranath, L. Watts, E. Devred, T. Platt, C. Caverhill, and H. Maass. Discrimination of diatoms from other phytoplankton using ocean-colour data. Marine Ecology Progress Series, 272:59–68, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/272/m272p059.pdf.

Stelfox-Widdicombe:2000:MGA

[SWEBS00] C. E. Stelfox-Widdicombe, E. S. Edwards, P. H. Burkill, and M. A. Sleigh. Microzooplankton grazing activity in the temperate and sub-tropical NE Atlantic: summer 1996. Marine Ecology Progress Series, 208:1–12, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/208/m208p001.pdf.

Stobart:2009:LTS

[SWG^{+}09] B. Stobart, R. Warwick, C. González, S. Mallol, D. Díaz, O. Reñones, and R. Goñi. Long-term and spillover effects of a marine protected area on an exploited fish community. Marine Ecology Progress Series, 384:47–60, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m384p047.pdf.

Stillman:2005:PSQ

[SWGC^{+}05] R. A. Stillman, A. D. West, J. D. Goss-Custard, S. McGrorty, N. J. Frost, D. J. Morrisey, A. J. Kenny, and A. L. Drewitt. Predicting site quality for shorebird communities: a case study on the Humber estuary, UK. Marine Ecology Progress Series, 305:203–217, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/305/m305p203.pdf.

Shimeta:2004:INS

[SWH04a] J. Shimeta, P. F. Witucki, and K. R. Hippe. Influences of nutritional state and temperature on suspension-feeding rates and mechanics in the spionid polychaete Polydora cornuta. Marine Ecology Progress Series, 280:173–180, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p173.pdf.

Solan:2004:LSQ

[SWH^{+}04b] M. Solan, B. D. Wigham, I. R. Hudson, R. Kennedy, C. H. Coulon, K. Norling, H. C. Nilsson, and R. Rosenberg. In situ quantification of bioturbation using time-lapse fluorescent sediment profile imaging (f-SPI), luminophore tracers and model simulation. Marine Ecology Progress Series, 271: 1–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/271/m271p001.pdf.

Storch:2005:CBD

[SWHSA05] S. Storch, R. P. Wilson, Z. M. Hillis-Starr, and D. Adelung. Cold-blooded divers: temperature-dependent dive performance in the wild hawksbill turtle Eretmochelys imbricata. Marine Ecology Progress Series, 293:263–271, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/293/
m293p263.pdf.

Shiah:2006:TEH

[SWL^{+}06] F. K. Shiah, T. H. Wu, K. Y. Li, S. J. Kao, Y. F. Tseng, J. L. Chung, and S. Jan. Thermal effects on heterotrophic processes in a coastal ecosystem adjacent to a nuclear power plant. Marine Ecology Progress Series, 309:55–65, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p055.pdf.

Skilleter:2007:EGM

[SWL07] G. A. Skilleter, C. Wegscheidl, and J. M. Lanyon. Effects of grazing by a marine mega-herbivore on benthic assemblages in a subtropical seagrass bed. Marine Ecology Progress Series, 351:287–300, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p287.pdf.

Skold:2003:GSS

[SWM03] M. Sköld, S. R. Wing, and P. V. Mladenov. Genetic subdivision of a sea star with high dispersal capability in relation to physical barriers in a fjordic seascape. Marine Ecology Progress Series, 250:163–174, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/250/m250p163.pdf.

Shaffer:2009:SHU

[SWS^{+}09] S. A. Shaffer, H. Weimerskirch, D. Scott, D. Pinaud, D. R. Thompson, P. M. Sagar, H. Moller, G. A. Taylor, D. G. Foley, Y. Tremblay, and D. P. Costa. Spatiotemporal habitat use by breeding sooty shearwaters Puffinus griseus. Marine Ecology Progress Series, 391:209–220, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m391p209.pdf.

Shreeve:2002:CGD

[SWW02] R. S. Shreeve, P. Ward, and M. J. Whitehouse. Copepod growth and development around South Georgia: relationships with temperature, food and krill. Marine Ecology Progress Series, 233:169–183, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/233/m233p169.pdf.

Somerfield:2002:RBS

[SYA02] P. J. Somerfield, S. Yodnarasri, and C. Aryuthaka. Relationships between seagrass biodiversity and infaunal communities: implications for studies of biodiversity effects. Marine Ecology Progress Series, 237:97–109, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/237/m237p097.pdf.

Sumida:2005:BRU

[SYCG05] P. Y. G. Sumida, M. Y. Yoshinaga, A. M. Ciotti, and S. A. Gaeta. Benthic response to upwelling events off the SE Brazilian coast. Marine Ecology Progress Series, 291: 35–42, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p035.pdf.

Sydeman:2009:TSM

[Syd09] William J. Sydeman. Theme section: Marine ecosystems, climate and phenology: impacts on top predators. Marine Ecology Progress Series, 393:184–299, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m393_ThemeSection.pdf.

Svetlichny:2009:DCE

[SYH^{+}09] L. S. Svetlichny, T. V. Yuneva, E. S. Hubareva, A. M. Schepkina, S. Besiktepe, A. E. Kideys, L. Bat, and F. Sahin. Development of Calanus euxinus during spring cold homothermy in the Black Sea. Marine Ecology Progress Series, 374:199–213, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/374/m374p199.pdf.

Smith:2009:OUK

[SYPO09] S. L. Smith, Y. Yamanaka, M. Pahlow, and A. Oschlies. Optimal uptake kinetics: physiological acclimation explains the pattern of nitrate uptake by phytoplankton in the ocean. Marine Ecology Progress Series, 384:1–12, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sempere:2000:MDL

[SYVC00] R. Sempéré, S. C. Yoro, F. Van Wambeke, and B. Charrière. Microbial decomposition of large organic particles in the northwestern Mediterranean Sea: an experimental approach. Marine Ecology Progress Series, 198:61–72, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p061.pdf.

Simpson:2005:REC

[SYWM05] S. D. Simpson, H. Y. Yan, M. L. Wittenrich, and M. G. Meekan. Response of embryonic coral reef fishes (Pomacentridae: Amphiprion spp.) to noise. Marine Ecology Progress Series, 287:201–208, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/287/m287p201.pdf.

Schwarzschild:2008:ADI

[SZ08a] A. C. Schwarzschild and J. C. Zieman. Apical dominance and the importance of clonal integration to apical growth in the seagrass Syringodium filiforme. Marine Ecology Progress Series, 360:37–46, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p037.pdf.

Schwarzschild:2008:EPI

[SZ08b] A. C. Schwarzschild and J. C. Zieman. Effects of physiological integration on the survival and growth of ramets and clonal fragments in the seagrass Syringodium filiforme. Marine Ecology Progress Series, 372:97–104, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p097.pdf.

Silina:2009:TTR

[SZ09] A. V. Silina and N. V. Zhukova. Topical and trophic relationships in a boring polychaete-scallop association: fatty acid biomarker approach. Marine Ecology Progress Series, 394:125–136, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/394/m394p125.pdf.

Stabenau:2004:RST

[SZBZ04] E. R. Stabenau, R. G. Zepp, E. Bartels, and R. G. Zika. Role of the seagrass Thalassia testudinum as a source of chromophoric dissolved organic matter in coastal south Florida. Marine Ecology Progress Series, 282:59–72, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p059.pdf.

Sabates:2008:VDE

[SZGS08] A. Sabatés, N. Zaragoza, C. Grau, and J. Salat. Vertical distribution of early developmental stages in two coexisting clupeoid species, Sardinella aurita and Engraulis encrasicolus. Marine Ecology Progress Series, 364:169–180, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p169.pdf.

Sun:2005:HSR

[SZL^{+}05] L. Sun, R. Zhu, X. Liu, Z. Xie, X. Yin, S. Zhao, and Y. Wang. HCl-soluble ^{87}Sr/^{86}Sr ratio in sediments impacted by penguin or seal excreta as a proxy for historical population size in the maritime Antarctic. Marine Ecology Progress Series, 303:43–50, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/303/m303p043.pdf.

Sumaila:2007:PCB

[SZW^{+}07] U. R. Sumaila, D. Zeller, R. Watson, J. Alder, and D. Pauly. Potential costs and benefits of marine reserves in the high seas. Marine Ecology Progress Series, 345:305–310, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/345/
m345p305.pdf.

Shang:2008:RIV

[SZZ08] X. Shang, G. S. Zhang, and J. Zhang. Relative importance of vascular plants and algal production in the food web of a Spartina-invaded salt marsh in the Yangtze River estuary. Marine Ecology Progress Series, 367:93–107, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/367/
m367p093.pdf.

Triantafillos:2001:AAR

[TA01a] L. Triantafillos and M. Adams. Allozyme analysis reveals a complex population structure in the southern calamary Sepioteuthis australis from Australia and New Zealand. Marine Ecology Progress Series, 212:193–209, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p193.pdf.

Tsukamoto:2001:FCE

[TA01b] K. Tsukamoto and T. Arai. Facultative catadromy of the eel Anguilla japonica between freshwater and seawater habitats. Marine Ecology Progress Series, 220:265–276, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
220/m220p265.pdf.

Timmermann:2003:BPD

[TA03] K. Timmermann and O. Andersen. Bioavailability of pyrene to the deposit-feeding polychaete Arenicola marina: importance of sediment versus water uptake routes. Marine Ecology Progress Series, 246:163–172, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/246/m246p163.pdf.

Tamminen:2007:SPN

[TA07a] T. Tamminen and T. Andersen. Seasonal phytoplankton nutrient limitation patterns as revealed by bioassays over Baltic Sea gradients of salinity and eutrophication. Marine Ecology Progress Series, 340:121–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m340p121.pdf.

Thygesen:2007:SVT

[TÅ07b] U. H. Thygesen and B. Ådlandsvik. Simulating vertical turbulent dispersal with finite volumes and binned random walks. Marine Ecology Progress Series, 347:145–153, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m347p145.pdf.

Tarrant:2004:ESE

[TAA04] A. M. Tarrant, M. J. Atkinson, and S. Atkinson. Effects of steroidal estrogens on coral growth and reproduction. Marine Ecology Progress Series, 269:121–129, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p121.pdf.

Teira:2003:PCB

[TAÁO^{+}03] E. Teira, J. Abalde, M. T. Álvarez-Ossorio, A. Bode, C. Cariño, Á. Cid, E. Fernández, N. González, J. Lorenzo, J. Valencia, and M. Varela. Plankton carbon budget in a coastal wind-driven upwelling station off A Coruña (NW Iberian Peninsula). Marine Ecology Progress Series, 265: 31–43, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p031.pdf.

Tobias:2001:NCT

[TACM01] C. R. Tobias, I. C. Anderson, E. A. Canuel, and S. A. Macko. Nitrogen cycling through a fringing marsh-aquifer ecotone. Marine Ecology Progress Series, 210:25–39, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
210/m210p025.pdf.

Tomas:2006:DEA

[TÁCTR06] F. Tomas, D. Álvarez-Cascos, X. Turon, and J. Romero. Differential element assimilation by sea urchins Paracentrotus lividus in seagrass beds: implications for trophic interactions. Marine Ecology Progress Series, 306:125–131, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/306/
m306p125.pdf.

Terlizzi:2007:SSV

[TAFBC07] A. Terlizzi, M. J. Anderson, S. Fraschetti, and L. Benedetti-Cecchi. Scales of spatial variation in Mediterranean subtidal sessile assemblages at different depths. Marine Ecology Progress Series, 332:25–39, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p025.pdf.

Henson:2007:ETN

[tAHSHA07] tephanie A. Henson, Richard Sanders, Claire Holeton, and John T. Allen. Erratum: Timing of nutrient depletion, diatom dominance and a lower-boundary estimate of export production for Irminger Basin, North Atlantic, Mar Ecol Prog Ser 313:73–84, 2006. Marine Ecology Progress Series, 343: 307, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p307.pdf. See [HSHA06].

Takagi:2002:OFE

[Tak02] Y. Takagi. Otolith formation and endolymph chemistry: a strong correlation between the aragonite saturation state and pH in the endolymph of the trout otolith organ. Marine Ecology Progress Series, 231:237–245, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/231/
m231p237.pdf.

Tallqvist:2001:BBB

[Tal01] M. Tallqvist. Burrowing behaviour of the Baltic clam Macoma balthica: effects of sediment type, hypoxia and predator presence. Marine Ecology Progress Series, 212:183–191, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
212/m212p183.pdf.

Tallis:2009:KRS

[Tal09] H. Tallis. Kelp and rivers subsidize rocky intertidal communities in the Pacific Northwest (USA). Marine Ecology Progress Series, 389:85–96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p085.pdf.

Takasuka:2003:EGS

[TAM03] A. Takasuka, I. Aoki, and I. Mitani. Evidence of growthselective predation on larval Japanese anchovy Engraulis japonicus in Sagami Bay. Marine Ecology Progress Series, 252:223–238, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p223.pdf.

Takasuka:2004:TSG

[TAM04] A. Takasuka, I. Aoki, and I. Mitani. Three synergistic growth-related mechanisms in the short-term survival of larval Japanese anchovy Engraulis japonicus in Sagami Bay. Marine Ecology Progress Series, 270:217–228, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/270/
m270p217.pdf.

Tanner:2002:CDD

[Tan02] J. E. Tanner. Consequences of density-dependent heterotrophic feeding for a partial autotroph. Marine Ecology Progress Series, 227:293–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/227/m227p293.pdf.

Takasuka:2007:PSG

[TAO07] A. Takasuka, I. Aoki, and Y. Oozeki. Predator-specific growthselective predation on larval Japanese anchovy Engraulis japonicus. Marine Ecology Progress Series, 350:99–107, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m350p099.pdf.

Tarling:2003:SDD

[Tar03] G. A. Tarling. Sex-dependent diel vertical migration in northern krill Meganyctiphanes norvegica and its consequences for population dynamics. Marine Ecology Progress Series, 260: 173–188, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p173.pdf.

Tentori:2004:CGC

[TAS04] E. Tentori, D. Allemand, and R. Shepherd. Cell growth and calcification result from uncoupled physiological processes in the soft coral Litophyton arboreum. Marine Ecology Progress Series, 276:85–92, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p085.pdf.

Takada:2007:CPM

[TAS07] Y. Takada, O. Abe, and T. Shibuno. Colonization patterns of mobile cryptic animals into interstices of coral rubble. Marine Ecology Progress Series, 343:35–44, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/343/
m343p035.pdf.

Tolimieri:2009:HRS

[TAW^{+}09] N. Tolimieri, K. Andrews, G. Williams, S. Katz, and P. S. Levin. Home range size and patterns of space use by lingcod, copper rockfish and quillback rockfish in relation to diel and tidal cycles. Marine Ecology Progress Series, 380:229– 243, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m380p229.pdf.

Taylor:2003:SDP

[Tay03] D. L. Taylor. Size-dependent predation on post-settlement winter flounder Pseudopleuronectes americanus by sand shrimp Crangon septemspinosa. Marine Ecology Progress Series, 263:197–215, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p197.pdf.

Taylor:2005:PPS

[Tay05] D. L. Taylor. Predation on post-settlement winter flounder Pseudopleuronectes americanus by sand shrimp Crangon septemspinosa in NW Atlantic estuaries. Marine Ecology Progress Series, 289:245–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/289/m289p245.pdf.

Taylor:2008:IOR

[TB08] J. Taylor and D. Bushek. Intertidal oyster reefs can persist and function in a temperate North American Atlantic estuary. Marine Ecology Progress Series, 361:301–306, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/361/
m361p301.pdf.

Troedsson:2002:RAB

[TBAT02] C. Troedsson, J. M. Bouquet, D. L. Aksnes, and E. M. Thompson. Resource allocation between somatic growth and reproductive output in the pelagic chordate Oikopleura dioica allows opportunistic response to nutritional variation. Marine Ecology Progress Series, 243:83–91, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p083.pdf.

Teo:2004:VGE

[TBB^{+}04] S. L. H. Teo, A. Boustany, S. Blackwell, A. Walli, K. C. Weng, and B. A. Block. Validation of geolocation estimates based on light level and sea surface temperature from electronic tags. Marine Ecology Progress Series, 283:81–98, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p081.pdf.

Tiwari:2006:EDD

[TBBB06] M. Tiwari, K. A. Bjorndal, A. B. Bolten, and B. M. Bolker. Evaluation of density-dependent processes and green turtle Chelonia mydas hatchling production at Tortuguero, Costa Rica. Marine Ecology Progress Series, 326:283–293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/326/m326p283.pdf.

Thiemann:2004:CGN

[TBBI04] G. W. Thiemann, S. M. Budge, W. D. Bowen, and S. J. Iverson. Comment on Grahl-Nielsen et al. (2003) ’Fatty acid composition of the adipose tissue of polar bears and of their prey: ringed seals, bearded seals and harp seals’. Marine Ecology Progress Series, 281:297–301, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/281/m281p297.pdf. See

[GNAD^{+}03] and reply [GNAD^{+}04].

Turley:2000:RBP

[TBC^{+}00] C. M. Turley, M. Bianchi, U. Christaki, P. Conan, J. R. W. Harris, S. Psarra, G. Ruddy, E. D. Stutt, A. Tselepides, and F. Van Wambeke. Relationship between primary producers and bacteria in an oligotrophic sea — the Mediterranean and biogeochemical implications. Marine Ecology Progress Series, 193:11–18, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/193/m193p011.pdf.

Terlizzi:2005:MUA

[TBCB^{+}05] A. Terlizzi, L. Benedetti-Cecchi, S. Bevilacqua, S. Fraschetti, P. Guidetti, and M. J. Anderson. Multivariate and univariate asymmetrical analyses in environmental impact assessment: a case study of Mediterranean subtidal sessile assemblages. Marine Ecology Progress Series, 289:27–42, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p027.pdf.

Tornero:2004:TDR

[TBFA04] V. Tornero, A. Borrell, J. Forcada, and Á. Aguilar. Tissue distribution of retinoids in common dolphins Delphinus delphis.

Marine Ecology Progress Series, 280:275–283, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/280/
m280p275.pdf.

Turner:2006:ZMB

[TBH06] J. T. Turner, D. G. Borkman, and C. D. Hunt. Zooplankton of Massachusetts Bay, USA, 1992–2003: relationships between the copepod Calanus finmarchicus and the North Atlantic Oscillation. Marine Ecology Progress Series, 311: 115–124, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/311/m311p115.pdf.

Thums:2008:TCR

[TBH08] M. Thums, C. J. A. Bradshaw, and M. A. Hindell. Tracking changes in relative body composition of southern elephant seals using swim speed data. Marine Ecology Progress Series, 370:249–261, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/370/m370p249.pdf.

Tremblay:2009:AAI

[TBH^{+}09] Y. Tremblay, S. Bertrand, R. W. Henry, M. A. Kappes, D. P. Costa, and S. A. Shaffer. Analytical approaches to investigating seabird-environment interactions: a review. Marine Ecology Progress Series, 391:153–163, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/theme/m391p153.pdf.

Tucker:2007:DDS

[TBI07] S. Tucker, W. D. Bowen, and S. J. Iverson. Dimensions of diet segregation in grey seals Halichoerus grypus revealed through stable isotopes of carbon (d ^{13}C) and nitrogen (d ^{1}5N). Marine Ecology Progress Series, 339:271–282, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/339/
m339p271.pdf.

Tucker:2008:CDE

[TBI08] S. Tucker, W. D. Bowen, and S. J. Iverson. Convergence of diet estimates derived from fatty acids and stable isotopes within individual grey seals. Marine Ecology Progress Series, 354:267–276, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/354/m354p267.pdf.

Tucker:2009:SVD

[TBI^{+}09] S. Tucker, W. D. Bowen, S. J. Iverson, W. Blanchard, and G. B. Stenson. Sources of variation in diets of harp and hooded seals estimated from quantitative fatty acid signature analysis (QFASA). Marine Ecology Progress Series, 384:287–302, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/384/m384p287.pdf.

Thompson:2009:LTC

[TBID09] P. A. Thompson, M. E. Baird, T. Ingleton, and M. A. Doblin. Long-term changes in temperate Australian coastal waters: implications for phytoplankton. Marine Ecology Progress Series, 394:1–19, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Treude:2003:AOM

[TBK^{+}03] T. Treude, A. Boetius, K. Knittel, K. Wallmann, and B. Barker Jørgensen. Anaerobic oxidation of methane above gas hydrates at Hydrate Ridge, NE Pacific Ocean. Marine Ecology Progress Series, 264:1–14, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/264/
m264p001.pdf.

Turner:2009:PSB

[TBL^{+}09] J. T. Turner, D. G. Borkman, J. A. Lincoln, D. A. Gauthier, and C. M. Petitpas. Plankton studies in Buzzards Bay, Massachusetts, USA. VI. Phytoplankton and water quality, 1987 to 1998. Marine Ecology Progress Series, 376:103– 122, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p103.pdf.

Trathan:2008:LTR

[TBM^{+}08] P. N. Trathan, C. Bishop, G. Maclean, P. Brown, A. Fleming, and M. A. Collins. Linear tracks and restricted temperature ranges characterise penguin foraging pathways. Marine Ecology Progress Series, 370:285–294, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m370p285.pdf.

Tanguy:2003:MGE

[TBR^{+}03] A. Tanguy, I. Boutet, R. Riso, P. Boudry, M. Auffret, and D. Moraga. Metallothionein genes in the European flat oyster Ostrea edulis: a potential ecological tool for environmental monitoring? Marine Ecology Progress Series, 257: 87–97, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/257/m257p087.pdf.

Tuya:2004:RBR

[TBSJ^{+}04] F. Tuya, A. Boyra, P. Sanchez-Jerez, C. Barbera, and R. J. Haroun. Relationships between rocky-reef fish assemblages, the sea urchin Diadema antillarum and macroalgae throughout the Canarian Archipelago. Marine Ecology Progress Series, 278:157–169, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p157.pdf.

Tarrant:2008:DGE

[TBVJ08] A. M. Tarrant, M. F. Baumgartner, T. Verslycke, and C. L. Johnson. Differential gene expression in diapausing and active Calanus finmarchicus (Copepoda). Marine Ecology Progress Series, 355:193–207, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/355/m355p193.pdf.

Tremblay:2000:BPD

[TC00] Y. Tremblay and Y. Cherel. Benthic and pelagic dives: a new foraging behaviour in rockhopper penguins. Marine Ecology Progress Series, 204:257–267, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/204/m204p257.pdf.

Thomas:2001:FUS

[TC01a] B. E. Thomas and R. M. Connolly. Fish use of subtropical saltmarshes in Queensland, Australia: relationships with vegetation, water depth and distance onto the marsh. Marine Ecology Progress Series, 209:275–288, 2001. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/209/m209p275.pdf.

Thomas:2001:CNC

[TC01b] C. R. Thomas and R. R. Christian. Comparison of nitrogen cycling in salt marsh zones related to sea-level rise. Marine Ecology Progress Series, 221:1–16, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/221/m221p001.pdf.

Taylor:2003:ETF

[TC03a] D. L. Taylor and J. S. Collie. Effect of temperature on the functional response and foraging behavior of the sand shrimp Crangon septemspinosa preying on juvenile winter flounder Pseudopleuronectes americanus. Marine Ecology Progress Series, 263:217–234, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p217.pdf.

Thomas:2003:AUS

[TC03b] F. I. M. Thomas and C. D. Cornelisen. Ammonium uptake by seagrass communities: effects of oscillatory versus unidirectional flow. Marine Ecology Progress Series, 247: 51–57, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p051.pdf.

Tremblay:2003:GVF

[TC03c] Y. Tremblay and Y. Cherel. Geographic variation in the foraging behaviour, diet and chick growth of rockhopper penguins. Marine Ecology Progress Series, 251:279–297, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/251/
m251p279.pdf.

Tester:2004:RVM

[TCC04] P. A. Tester, J. H. Cohen, and G. Cervetto. Reverse vertical migration and hydrographic distribution of Anomalocera ornata (Copepoda: Pontellidae) in the US South Atlantic Bight. Marine Ecology Progress Series, 268:195–204, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p195.pdf.

Teegarden:2001:ZFB

[TCD01] G. J. Teegarden, R. G. Campbell, and E. G. Durbin. Zooplankton feeding behavior and particle selection in natural plankton assemblages containing toxic Alexandrium spp. Marine Ecology Progress Series, 218:213–226, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
218/m218p213.pdf.

Touratier:2003:IGM

[TCG03] F. Touratier, F. Carlotti, and G. Gorsky. Individual growth model for the appendicularian Oikopleura dioica. Marine Ecology Progress Series, 248:141–163, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/248/
m248p141.pdf.

Tarling:2007:RAK

[TCRT^{+}07] G. A. Tarling, J. Cuzin-Roudy, S. E. Thorpe, R. S. Shreeve, P. Ward, and E. J. Murphy. Recruitment of Antarctic krill Euphausia superba in the South Georgia region: adult fecundity and the fate of larvae. Marine Ecology Progress Series, 331:161–179, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/331/m331p161.pdf.

Thistle:2005:DOS

[TCS^{+}05] D. Thistle, K. R. Carman, L. Sedlacek, P. G. Brewer, J. W. Fleeger, and J. P. Barry. Deep-ocean, sediment-dwelling animals are sensitive to sequestered carbon dioxide. Marine Ecology Progress Series, 289:1–4, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tian:2009:MCB

[TCS^{+}09] R. C. Tian, C. Chen, K. D. E. Stokesbury, B. Rothschild, G. W. Cowles, Q. Xu, S. Hu, B. P. Harris, and M. C. Marino II. Modeling the connectivity between sea scallop populations in the Middle Atlantic Bight and over Georges Bank. Marine Ecology Progress Series, 380:147–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/380/
m380p147.pdf.

Turon:2000:MRO

[TCT^{+}00] X. Turon, M. Codina, I. Tarjuelo, M. J. Uriz, and M. A. Becerro. Mass recruitment of Ophiothrix fragilis (Ophiuroidea) on sponges: settlement patterns and post-settlement dynamics. Marine Ecology Progress Series, 200:201–212, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p201.pdf.

Tzeng:2007:MMH

[TCW^{+}07] W. Tzeng, C. Chang, C. Wang, J. Shiao, Y. Iizuka, Y. Yang, C. You, and L. Lozys. Misidentification of the migratory history of anguillid eels by Sr/Ca ratios of vaterite otoliths. Marine Ecology Progress Series, 348:285–295, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/348/
m348p285.pdf.

Tsang:2008:PDB

[TCW^{+}08] L. M. Tsang, B. K. K. Chan, T. H. Wu, W. C. Ng, T. Chatterjee, G. A. Williams, and K. H. Chu. Population differentiation in the barnacle Chthamalus malayensis: postglacial colonization and recent connectivity across the Pacific and Indian Oceans. Marine Ecology Progress Series, 364:107–118, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p107.pdf.

Trip:2008:IOA

[TCWR08] E. L. Trip, J. H. Choat, D. T. Wilson, and D. R. Robertson. Inter-oceanic analysis of demographic variation in a widely distributed Indo–Pacific coral reef fish. Marine Ecology Progress Series, 373:97–109, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/373/m373p097.pdf.

Tang:2001:PIC

[TD01a] K. W. Tang and H. G. Dam. Phytoplankton inhibition of copepod egg hatching: test of an exudate hypothesis. Marine Ecology Progress Series, 209:197–202, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/209/m209p197.pdf.

Tsai:2001:LHP

[TD01b] M. L. Tsai and C. F. Dai. Life history plasticity and the reproductive strategy enabling the invasion of Ligia exotica (Crustacea: Isopoda) from the littoral zone to an inland creek. Marine Ecology Progress Series, 210:175–184, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
210/m210p175.pdf.

Troncoso:2003:BCF

[TDC^{+}03] V. A. Troncoso, G. Daneri, L. A. Cuevas, B. Jacob, and P. Montero. Bacterial carbon flow in the Humboldt Current System off Chile. Marine Ecology Progress Series, 250: 1–12, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p001.pdf.

Thibault:2007:FCA

[TDC^{+}07] I. Thibault, J. J. Dodson, F. Caron, W. Tzeng, Y. Iizuka, and J. Shiao. Facultative catadromy in American eels: testing the conditional strategy hypothesis. Marine Ecology Progress Series, 344:219–229, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/344/m344p219.pdf.

Turra:2005:PGS

[TDL05] A. Turra, M. R. Denadai, and F. P. P. Leite. Predation on gastropods by shell-breaking crabs: effects on shell availability to hermit crabs. Marine Ecology Progress Series, 286:279–291, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p279.pdf.

Teal:2008:ECC

[TdLvdVR08] L. R. Teal, J. J. de Leeuw, H. W. van der Veer, and A. D. Rijnsdorp. Effects of climate change on growth of 0-group sole and plaice. Marine Ecology Progress Series, 358:219– 230, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/358/m358p219.pdf.

Thieltges:2008:PMT

[TdMF^{+}08] D. W. Thieltges, X. de Montaudouin, B. Fredensborg, K. T. Jensen, J. Koprivnikar, and R. Poulin. Production of marine trematode cercariae: a potentially overlooked path of energy flow in benthic systems. Marine Ecology Progress Series, 372:147–155, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/372/m372p147.pdf.

Turner:2000:ART

[TDP^{+}00] J. T. Turner, G. J. Doucette, C. L. Powell, D. M. Kulis, B. A. Keafer, and D. M. Anderson. Accumulation of red tide toxins in larger size fractions of zooplankton assemblages from Massachusetts Bay, USA. Marine Ecology Progress Series, 203:95–107, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/203/m203p095.pdf.

Thompson:2008:VFF

[TDRM08] R. J. Thompson, D. Deibel, A. M. Redden, and C. H. McKenzie. Vertical flux and fate of particulate matter in a Newfoundland fjord at sub-zero water temperatures during spring. Marine Ecology Progress Series, 357:33–49, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/357/
m357p033.pdf.

Tortell:2002:CET

[TDSM02] P. D. Tortell, G. R. DiTullio, D. M. Sigman, and F. M. M. Morel. CO_{2} effects on taxonomic composition and nutrient utilization in an Equatorial Pacific phytoplankton assemblage. Marine Ecology Progress Series, 236:37–43, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p037.pdf.

Takahashi:2003:FSC

[TDT^{+}03] A. Takahashi, M. J. Dunn, P. N. Trathan, K. Sato, Y. Naito, and J. P. Croxall. Foraging strategies of chinstrap penguins at Signy Island, Antarctica: importance of benthic feeding on Antarctic krill. Marine Ecology Progress Series, 250:279– 289, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/250/m250p279.pdf.

Tian:2003:MCF

[TDTR03] R. C. Tian, D. Deibel, R. J. Thompson, and R. B. Rivkin. Modeling of climate forcing on a cold-ocean ecosystem, Conception Bay, Newfoundland. Marine Ecology Progress Series, 262:1–17, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p001.pdf.

Thornton:2007:SWI

[TDUN07] D. C. O. Thornton, L. F. Dong, G. J. C. Underwood, and D. B. Nedwell. Sediment-water inorganic nutrient exchange and nitrogen budgets in the Colne Estuary, UK. Marine Ecology Progress Series, 337:63–77, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/337/m337p063.pdf.

Timmermans:2001:CLI

[TDvdW^{+}01] K. R. Timmermans, M. S. Davey, B. van der Wagt, J. Snoek, R. J. Geider, M. J. W. Veldhuis, L. J. A. Gerringa, and H. J. W. de Baar. Co-limitation by iron and light of Chaetoceros brevis, C. dichaeta and C. calcitrans (Bacillariophyceae). Marine Ecology Progress Series, 217:287–297, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
217/m217p287.pdf.

Taylor:2000:EHE

[TE00] D. L. Taylor and D. B. Eggleston. Effects of hypoxia on an estuarine predator–prey interaction: foraging behavior and mutual interference in the blue crab Callinectes sapidus and the infaunal clam prey Mya arenaria. Marine Ecology Progress Series, 196:221–237, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p221.pdf.

Thomas:2007:IGP

[TEB^{+}07] T. Thomas, S. Egan, D. Burg, C. Ng, L. Ting, and R. Cavicchioli. Integration of genomics and proteomics into marine microbial ecology. Marine Ecology Progress Series, 332: 291–299, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m332p291.pdf.

Titelman:2004:OVD

[TF04] J. Titelman and Ó. Fiksen. Ontogenetic vertical distribution patterns in small copepods: field observations and model predictions. Marine Ecology Progress Series, 284: 49–63, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p049.pdf.

Teichberg:2008:MRE

[TFA^{+}08] M. Teichberg, S. E. Fox, C. Aguila, Y. S. Olsen, and I. Valiela. Macroalgal responses to experimental nutrient enrichment in shallow coastal waters: growth, internal nutrient pools, and isotopic signatures. Marine Ecology Progress Series, 368:117–126, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/368/m368p117.pdf.

Trenkel:2004:ADW

[TFL^{+}04] V. M. Trenkel, R. I. C. C. Francis, P. Lorance, S. Mahévas, M. J. Rochet, and D. M. Tracey. Availability of deep-water fish to trawling and visual observation from a remotely operated vehicle (ROV). Marine Ecology Progress Series, 284: 293–303, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p293.pdf.

Tilstone:2003:PCP

[TFLA03] G. H. Tilstone, F. G. Figueiras, L. M. Lorenzo, and B. Arbones. Phytoplankton composition, photosynthesis and primary production during different hydrographic conditions at the northwest Iberian upwelling system. Marine Ecology Progress Series, 252:89–104, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p089.pdf. See

erratum [Ano03f].

Temple:2000:VMG

[TFSJ00] G. K. Temple, C. J. Fox, R. Stewart, and I. A. Johnston. Variability in muscle growth characteristics during the spawning season in a natural population of Atlantic herring Clupea harengus. Marine Ecology Progress Series, 205:271–281, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p271.pdf.

Thieltges:2009:LSP

[TFSP09] D. W. Thieltges, B. L. Fredensborg, A. Studer, and R. Poulin. Large-scale patterns in trematode richness and infection levels in marine crustacean hosts. Marine Ecology Progress Series, 389:139–147, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/389/m389p139.pdf.

Thornber:2003:STV

[TG03] C. S. Thornber and S. D. Gaines. Spatial and temporal variation of haploids and diploids in populations of four congeners of the marine alga Mazzaella. Marine Ecology Progress Series, 258:65–77, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p065.pdf.

Teixido:2002:SPQ

[TGA02] N. Teixidó, J. Garrabou, and W. E. Arntz. Spatial pattern quantification of Antarctic benthic communities using landscape indices. Marine Ecology Progress Series, 242:1– 14, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/242/m242p001.pdf.

Teixido:2004:RAB

[TGGA04] N. Teixidó, J. Garrabou, J. Gutt, and W. E. Arntz. Recovery in Antarctic benthos after iceberg disturbance: trends in benthic composition, abundance and growth forms. Marine Ecology Progress Series, 278:1–16, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/278/m278p001.pdf.

Tolley:2005:MDP

[TGGM05] K. A. Tolley, J. C. Groeneveld, K. Gopal, and C. A. Matthee. Mitochondrial DNA panmixia in spiny lobster Palinurus gilchristi suggests a population expansion. Marine Ecology Progress Series, 297:225–231, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/297/m297p225.pdf.

Tobias:2003:SDF

[TGM^{+}03] C. Tobias, A. Giblin, J. McClelland, J. Tucker, and B. Peterson. Sediment DIN fluxes and preferential recycling of benthic microalgal nitrogen in a shallow macrotidal estuary. Marine Ecology Progress Series, 257:25–36, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/257/
m257p025.pdf.

Tewfik:2007:IAC

[TGM07] A. Tewfik, F. Guichard, and K. S. McCann. Influence of acute and chronic disturbance on macrophyte landscape zonation. Marine Ecology Progress Series, 335:111–121, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/335/
m335p111.pdf.

Troedsson:2005:VFA

[TGNT05] C. Troedsson, O. Grahl-Nielsen, and E. M. Thompson. Variable fatty acid composition of the pelagic appendicularian Oikopleura dioica in response to dietary quality and quantity. Marine Ecology Progress Series, 289:165–176, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/289/
m289p165.pdf.

Taylor:2006:SCD

[TGSW06] G. T. Taylor, C. J. Gobler, and S. A. Sañudo-Wilhelmy. Speciation and concentrations of dissolved nitrogen as determinants of brown tide Aureococcus anophagefferens bloom initiation. Marine Ecology Progress Series, 312:67–83, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p067.pdf.

Trathan:2006:FDM

[TGT^{+}06] P. N. Trathan, C. Green, J. Tanton, H. Peat, J. Poncet, and A. Morton. Foraging dynamics of macaroni penguins Eudyptes chrysolophus at South Georgia during brood-guard. Marine Ecology Progress Series, 323:239–251, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p239.pdf.

Tang:2006:MSP

[TGY^{+}06] K. W. Tang, H. P. Grossart, E. M. Yam, G. A. Jackson, H. W. Ducklow, and T. Kiørboe. Mesocosm study of particle dynamics and control of particle-associated bacteria by flagellate grazing. Marine Ecology Progress Series, 325:15– 27, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/325/m325p015.pdf.

Tolley:2002:IES

[TH02] K. A. Tolley and H. E. Heldal. Inferring ecological separation from regional differences in radioactive caesium in harbour porpoises Phocoena phocoena. Marine Ecology Progress Series, 228:301–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p301.pdf.

Tuya:2006:SPR

[TH06] F. Tuya and R. J. Haroun. Spatial patterns and response to wave exposure of shallow water algal assemblages across the Canarian Archipelago: a multi-scaled approach. Marine Ecology Progress Series, 311:15–28, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/311/
m311p015.pdf.

Tuya:2005:SUD

[THBSJ05] F. Tuya, R. J. Haroun, A. Boyra, and P. Sanchez-Jerez. Sea urchin Diadema antillarum: different functions in the structure and dynamics of reefs on both sides of the Atlantic. Marine Ecology Progress Series, 302:307–310, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p307.pdf.

Thiyagarajan:2005:GSJ

[THC^{+}05] V. Thiyagarajan, O. S. Hung, J. M. Y. Chiu, R. S. S. Wu, and P. Y. Qian. Growth and survival of juvenile barnacle Balanus amphitrite: interactive effects of cyprid energy reserve and habitat. Marine Ecology Progress Series, 299: 229–237, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p229.pdf.

Thrush:2001:FDMa

[THF^{+}01a] S. F. Thrush, J. E. Hewitt, G. A. Funnell, V. J. Cummings, J. Ellis, D. Schultz, D. Talley, and A. Norkko. Fishing disturbance and marine biodiversity: role of habitat structure in simple soft-sediment systems. Marine Ecology Progress Series, 221:255–264, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/221/m221p255.pdf. See erratum

[Ano02h].

Thrush:2001:FDMb

[THF^{+}01b] S. F. Thrush, J. E. Hewitt, G. A. Funnell, V. J. Cummings, J. Ellis, D. Schultz, D. Talley, and A. Norkko. Fishing disturbance and marine biodiversity: the role of habitat structure in simple soft-sediment systems. Marine Ecology Progress Series, 223:277–286, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/223/m223p277.pdf. See erratum

[Ano02h].

Thrush:2005:MSA

[THHY05] S. F. Thrush, J. E. Hewitt, P. M. J. Herman, and T. Ysebaert. Multi-scale analysis of species-environment relationships. Marine Ecology Progress Series, 302:13–26, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/302/
m302p013.pdf.

Thistle:2003:HCE

[Thi03] D. Thistle. Harpacticoid copepod emergence at a shelf site in summer and winter: implications for hydrodynamic and mating hypotheses. Marine Ecology Progress Series, 248: 177–185, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p177.pdf.

Thieltges:2005:IIE

[Thi05] D. W. Thieltges. Impact of an invader: epizootic American slipper limpet Crepidula fornicata reduces survival and growth in European mussels. Marine Ecology Progress Series, 286:13–19, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p013.pdf.

Thieltges:2006:EIM

[Thi06] D. W. Thieltges. Effect of infection by the metacercarial trematode Renicola roscovita on growth in intertidal blue mussel Mytilus edulis. Marine Ecology Progress Series, 319: 129–134, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p129.pdf.

Tillin:2006:CBT

[THJK06] H. M. Tillin, J. G. Hiddink, S. Jennings, and M. J. Kaiser. Chronic bottom trawling alters the functional composition of benthic invertebrate communities on a sea-basin scale. Marine Ecology Progress Series, 318:31–45, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/318/
m318p031.pdf.

Taggart:2003:LED

[THM^{+}03] S. J. Taggart, P. N. Hooge, J. Mondragon, E. R. Hooge, and A. G. Andrews. Living on the edge: distribution of Dungeness crab Cancer magister in a recently deglaciated fjord. Marine Ecology Progress Series, 246:241–252, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/246/
m246p241.pdf.

Thrush:2003:HCE

[THN^{+}03] S. F. Thrush, J. E. Hewitt, A. Norkko, P. E. Nicholls, G. A. Funnell, and J. I. Ellis. Habitat change in estuaries: predicting broad-scale responses of intertidal macrofauna to sediment mud content. Marine Ecology Progress Series, 263: 101–112, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/263/m263p101.pdf.

Thornton:2004:FTE

[Tho04] D. C. O. Thornton. Formation of transparent exopolymeric particles (TEP) from macroalgal detritus. Marine Ecology Progress Series, 282:1–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/282/m282p001.pdf.

Thorpe:2007:MRS

[Tho07] J. E. Thorpe. Maturation responses of salmonids to changing developmental opportunities. Marine Ecology Progress Series, 335:285–288, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m335p285.pdf.

Teske:2005:MEL

[THP^{+}05] P. R. Teske, H. Hamilton, P. J. Palsbøll, C. K. Choo, H. Gabr, S. A. Lourie, M. Santos, A. Sreepada, M. I. Cherry, and C. A. Matthee. Molecular evidence for long-distance colonization in an Indo–Pacific seahorse lineage. Marine Ecology Progress Series, 286:249–260, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/286/m286p249.pdf.

Thresher:2009:ECC

[Thr09] R. E. Thresher. Environmental and compositional correlates of growth rate in deep-water bamboo corals (Gorgonacea; Isididae). Marine Ecology Progress Series, 397:187–196, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p187.pdf.

Thiebaut:2002:STV

[THS^{+}02] E. Thiébaut, X. Huther, B. Shillito, D. Jollivet, and F. Gaill. Spatial and temporal variations of recruitment in the tube worm Riftia pachyptila on the East Pacific Rise (9^{o}50’ N and 13^{o}N). Marine Ecology Progress Series, 234:147–157, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/234/m234p147.pdf.

Thurber:2007:DAS

[Thu07] A. R. Thurber. Diets of Antarctic sponges: links between the pelagic microbial loop and benthic metazoan food web.

Marine Ecology Progress Series, 351:77–89, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/351/
m351p077.pdf.

Toledo-Hernandez:2009:IDD

[THYBS09] C. Toledo-Hernández, P. Yoshioka, P. Bayman, and A. M. Sabat. Impact of disease and detachment on growth and survivorship of sea fans (Gorgonia ventalina). Marine Ecology Progress Series, 393:47–54, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/393/m393p047.pdf.

Turner:2001:DCG

[TIM^{+}01] J. T. Turner, A. Ianora, A. Miralto, M. Laabir, and F. Esposito. Decoupling of copepod grazing rates, fecundity and egg-hatching success on mixed and alternating diatom and dinoflagellate diets. Marine Ecology Progress Series, 220: 187–199, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/220/m220p187.pdf.

Titelman:2001:SEB

[Tit01] J. Titelman. Swimming and escape behavior of copepod nauplii: implications for predator–prey interactions among copepods. Marine Ecology Progress Series, 213:203–213, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
213/m213p203.pdf.

Tillmann:2002:TEA

[TJ02] U. Tillmann and U. John. Toxic effects of Alexandrium spp. on heterotrophic dinoflagellates: an allelochemical defence mechanism independent of PSP-toxin content. Marine Ecology Progress Series, 230:47–58, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/230/m230p047.pdf.

Tweedley:2008:ZMS

[TJA08] J. R. Tweedley, E. L. Jackson, and M. J. Attrill. Zostera marina seagrass beds enhance the attachment of the invasive alga Sargassum muticum in soft sediments. Marine Ecology Progress Series, 354:305–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/354/m354p305.pdf.

Tarling:2002:MSB

[TJEM02] G. A. Tarling, T. Jarvis, S. M. Emsley, and J. B. L. Matthews. Midnight sinking behaviour in Calanus finmarchicus: a response to satiation or krill predation? Marine Ecology Progress Series, 240:183–194, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/240/m240p183.pdf. See

response [Pea03].

Tolimieri:2000:ASC

[TJM00] N. Tolimieri, A. Jeffs, and J. C. Montgomery. Ambient sound as a cue for navigation by the pelagic larvae of reef fishes. Marine Ecology Progress Series, 207:219–224, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p219.pdf.

Tarling:2003:CFD

[TJM03] G. A. Tarling, T. Jarvis, and J. B. L. Matthews. Calanus finmarchicus descends in response to the arrival of krill-better unfed than dead. Marine Ecology Progress Series, 252:307– 310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/252/m252p307.pdf.

Thornber:2008:DHF

[TJS08] C. S. Thornber, E. Jones, and J. J. Stachowicz. Differences in herbivore feeding preferences across a vertical rocky intertidal gradient. Marine Ecology Progress Series, 363: 51–62, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/363/m363p051.pdf.

Thomasson:2003:SCP

[TJSG03] M. A. Thomasson, M. L. Johnson, J. O. Strömberg, and E. Gaten. Swimming capacity and pleopod beat rate as a function of sex, size and moult stage in northern krill Meganyctiphanes norvegica. Marine Ecology Progress Series, 250:205–213, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p205.pdf.

Talman:2001:IEC

[TK01] S. G. Talman and M. J. Keough. Impact of an exotic clam, Corbula gibba, on the commercial scallop Pecten fumatus in Port Phillip Bay, south-east Australia: evidence of resourcerestricted growth in a subtidal environment. Marine Ecology Progress Series, 221:135–143, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/221/m221p135.pdf.

Titelman:2003:MCN

[TK03a] J. Titelman and T. Kiørboe. Motility of copepod nauplii and implications for food encounter. Marine Ecology Progress Series, 247:123–135, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/247/m247p123.pdf.

Titelman:2003:PAN

[TK03b] J. Titelman and T. Kiørboe. Predator avoidance by nauplii. Marine Ecology Progress Series, 247:137–149, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p137.pdf.

Takahashi:2004:RBI

[TK04] K. Takahashi and K. Kawaguchi. Reproductive biology of the intertidal and infralittoral mysids Archaeomysis kokuboi and A. japonica on a sandy beach in NE Japan. Marine Ecology Progress Series, 283:219–231, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/283/m283p219.pdf.

Testa:2008:VBP

[TK08] J. M. Testa and W. M. Kemp. Variability of biogeochemical processes and physical transport in a partially stratified estuary: a box-modeling analysis. Marine Ecology Progress Series, 356:63–79, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p063.pdf.

Thornber:2004:PEI

[TKGS04] C. S. Thornber, B. P. Kinlan, M. H. Graham, and J. J. Stachowicz. Population ecology of the invasive kelp Undaria pinnatifida in California: environmental and biological controls on demography. Marine Ecology Progress Series, 268:69–80, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/268/m268p069.pdf.

Trippel:2005:MPI

[TKK05] E. A. Trippel, G. Kraus, and F. W. Köster. Maternal and paternal influences on early life history traits and processes of Baltic cod Gadus morhua. Marine Ecology Progress Series, 303:259–267, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/303/m303p259.pdf.

Tomaru:2000:CPB

[TKN00] Y. Tomaru, Z. Kawabata, and S. Nakano. Consumption of picoplankton by the bivalve larvae of Japanese pearl oyster Pinctada fucata martensii. Marine Ecology Progress Series, 192:195–202, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p195.pdf.

Takasuka:2008:SOA

[TKO08] A. Takasuka, H. Kubota, and Y. Oozeki. Spawning overlap of anchovy and sardine in the western North Pacific. Marine Ecology Progress Series, 366:231–244, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/366/
m366p231.pdf.

Teuten:2006:NSB

[TKR06] E. L. Teuten, G. M. King, and C. M. Reddy. Natural ^{14}C in Saccoglossus bromophenolosus compared to ^{14}C in surrounding sediments. Marine Ecology Progress Series, 324: 167–172, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/324/m324p167.pdf.

Thor:2007:SED

[TKTJ07] P. Thor, M. Koski, K. W. Tang, and S. H. Jónasdóttir. Supplemental effects of diet mixing on absorption of ingested organic carbon in the marine copepod Acartia tonsa. Marine Ecology Progress Series, 331:131–138, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/331/m331p131.pdf.

Tang:2004:OPB

[TKVL04] D. L. Tang, H. Kawamura, T. Van Dien, and M. A. Lee. Offshore phytoplankton biomass increase and its oceanographic causes in the South China Sea. Marine Ecology Progress Series, 268:31–41, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/268/m268p031.pdf.

Turra:2003:MHL

[TL03] A. Turra and F. P. P. Leite. The molding hypothesis: linking shell use with hermit crab growth, morphology, and shellspecies selection. Marine Ecology Progress Series, 265:155– 163, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/265/m265p155.pdf.

Tew:2005:DTS

[TL05] K. S. Tew and W. T. Lo. Distribution of Thaliacea in SW Taiwan coastal water in 1997, with special reference to Doliolum denticulatum, Thalia democratica and T. orientalis. Marine Ecology Progress Series, 292:181–193, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p181.pdf.

Toth:2008:WSC

[TL08] G. B. Toth and M. Lindeborg. Water-soluble compounds from the breadcrumb sponge Halichondria panicea deter attachment of the barnacle Balanus improvisus. Marine Ecology Progress Series, 354:125–132, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/354/m354p125.pdf.

Topping:2006:SFS

[TLC06] D. T. Topping, C. G. Lowe, and J. E. Caselle. Site fidelity and seasonal movement patterns of adult California sheephead Semicossyphus pulcher (Labridae): an acoustic monitoring study. Marine Ecology Progress Series, 326:257– 267, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/326/m326p257.pdf.

Taylor:2006:KHH

[TLFS06] M. W. Taylor, S. D. Laffan, S. Fielder, and I. M. Suthers. Key habitat and home range of mulloway Argyrosomus japonicus in a south-east Australian estuary: finding the estuarine niche to optimise stocking. Marine Ecology Progress Series, 328:237–247, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/328/m328p237.pdf.

Tunin-Ley:2009:PBN

[TLIL^{+}09] A. Tunin-Ley, F. Ibañez, J. P. Labat, A. Zingone, and R. Lemée. Phytoplankton biodiversity and NW Mediterranean Sea warming: changes in the dinoflagellate genus Ceratium in the 20th century. Marine Ecology Progress Series, 375:85–99, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/375/m375p085.pdf.

Tarjuelo:2002:DMA

[TLLCT02] I. Tarjuelo, S. López-Legentil, M. Codina, and X. Turon. Defence mechanisms of adults and larvae of colonial ascidians: patterns of palatability and toxicity. Marine Ecology Progress Series, 235:103–115, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/235/m235p103.pdf.

Todd:2004:GEI

[TLLKC04] P. A. Todd, R. J. Ladle, N. J. I. Lewin-Koh, and L. M. Chou. Genotype × environment interactions in transplanted clones of the massive corals Favia speciosa and Diploastrea heliopora. Marine Ecology Progress Series, 271:167–182, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p167.pdf.

Turner:2001:ZFE

[TLNH01] J. T. Turner, H. Levinsen, T. G. Nielsen, and B. W. Hansen. Zooplankton feeding ecology: grazing on phytoplankton and predation on protozoans by copepod and barnacle nauplii in Disko Bay, West Greenland. Marine Ecology Progress Series, 221:209–219, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/221/m221p209.pdf.

Turner:2003:AFG

[TLP^{+}03] V. L. G. Turner, S. M. Lynch, L. Paterson, J. L. León-Cortés, and J. P. Thorpe. Aggression as a function of genetic relatedness in the sea anemone Actinia equina (Anthozoa: Actiniaria). Marine Ecology Progress Series, 247: 85–92, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/247/m247p085.pdf.

Trussell:2006:DSD

[TLPG06] G. C. Trussell, M. P. Lesser, M. R. Patterson, and S. J. Genovese. Depth-specific differences in growth of the reef sponge Callyspongia vaginalis: role of bottom-up effects. Marine Ecology Progress Series, 323:149–158, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p149.pdf.

Thyrhaug:2003:SCM

[TLTB03] R. Thyrhaug, A. Larsen, T. F. Thingstad, and G. Bratbak. Stable coexistence in marine algal host-virus systems. Marine Ecology Progress Series, 254:27–35, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p027.pdf.

Tanaka:2002:EES

[TM02a] M. O. Tanaka and C. A. Magalhães. Edge effects and succession dynamics in Brachidontes mussel beds. Marine Ecology Progress Series, 237:151–158, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/237/m237p151.pdf.

Thompson:2002:IVI

[TM02b] A. A. Thompson and B. D. Mapstone. Intra- versus interannual variation in counts of reef fishes and interpretations of long-term monitoring studies. Marine Ecology Progress Series, 232:247–257, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/232/m232p247.pdf.

Taggart:2008:SPM

[TMAN08] S. J. Taggart, J. Mondragon, A. G. Andrews, and J. K. Nielsen. Spatial patterns and movements of red king and Tanner crabs: implications for the design of marine protected areas. Marine Ecology Progress Series, 365:151–163, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/365/m365p151.pdf.

Taylor:2007:TPS

[TMD^{+}07] M. D. Taylor, L. M. MacKenzie, T. J. Dodgshun, G. A. Hopkins, E. J. de Zwart, and C. D. Hunt. Trans-Pacific shipboard trials on planktonic communities as indicators of open ocean ballast water exchange. Marine Ecology Progress Series, 350:41–54, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/350/m350p041.pdf.

Teske:2006:IMB

[TMFB06] P. R. Teske, C. D. McQuaid, P. W. Froneman, and N. P. Barker. Impacts of marine biogeographic boundaries on phylogeographic patterns of three South African estuarine crustaceans. Marine Ecology Progress Series, 314:283–293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/314/m314p283.pdf.

Tilstone:2000:DDC

[TMFF00] G. H. Tilstone, B. M. Míguez, F. G. Figueiras, and E. G. Fermín. Diatom dynamics in a coastal ecosystem affected by upwelling: coupling between species succession, circulation and biogeochemical processes. Marine Ecology Progress Series, 205:23–41, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p023.pdf.

Thompson:2005:DPM

[TMJ^{+}05] R. C. Thompson, P. S. Moschella, S. R. Jenkins, T. A. Norton, and S. J. Hawkins. Differences in photosynthetic marine biofilms between sheltered and moderately exposed rocky shores. Marine Ecology Progress Series, 296:53–63, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/296/m296p053.pdf.

Thompson:2003:FBS

[TML03] D. Thompson, S. E. W. Moss, and P. Lovell. Foraging behaviour of South American fur seals Arctocephalus australis: extracting fine scale foraging behaviour from satellite tracks. Marine Ecology Progress Series, 260:285–296, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p285.pdf.

Tyler:2005:UUA

[TMM05] A. C. Tyler, K. J. McGlathery, and S. A. Macko. Uptake of urea and amino acids by the macroalgae Ulva lactuca (Chlorophyta) and Gracilaria vermiculophylla (Rhodophyta). Marine Ecology Progress Series, 294:161–172, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p161.pdf.

Tanaka:2004:ESM

[TN04] Y. Tanaka and M. Nakaoka. Emergence stress and morphological constraints affect the species distribution and growth of subtropical intertidal seagrasses. Marine Ecology Progress Series, 284:117–131, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p117.pdf.

Tanaka:2007:IVP

[TN07] Y. Tanaka and M. Nakaoka. Interspecific variation in photosynthesis and respiration balance of three seagrasses in relation to light availability. Marine Ecology Progress Series, 350:63–70, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/350/m350p063.pdf.

Tanaka:2005:FRS

[TND^{+}05] H. Tanaka, Y. Naito, N. D. Davis, S. Urawa, H. Ueda, and M. Fukuwaka. First record of the at-sea swimming speed of a Pacific salmon during its oceanic migration. Marine Ecology Progress Series, 291:307–312, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/291/m291p307.pdf.

Thingstad:2002:CBP

[TNHL02] T. F. Thingstad, T. G. Nielsen, A. Skjoldborg Hansen, and H. Levinsen. Control of bacterial production in cold waters. A theoretical analysis of mechanisms relating bacterial production and zooplankton biomass in Disko Bay, Western Greenland. Marine Ecology Progress Series, 228:15–24, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/228/m228p015.pdf.

Trimmer:2000:SBO

[TNSM00a] M. Trimmer, D. B. Nedwell, D. B. Sivyer, and S. J. Malcolm. Seasonal benthic organic matter mineralisation measured by oxygen uptake and denitrification along a transect of the inner and outer River Thames estuary, UK. Marine Ecology Progress Series, 197:103–119, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/197/m197p103.pdf.

Trimmer:2000:SOM

[TNSM00b] M. Trimmer, D. B. Nedwell, D. B. Sivyer, and S. J. Malcolm. Seasonal organic mineralisation and denitrification in intertidal sediments and their relationship to the abundance of Enteromorpha sp. and Ulva sp. Marine Ecology Progress Series, 203:67–80, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/203/m203p067.pdf.

Tonnesson:2006:FPC

[TNT06] K. Tönnesson, T. G. Nielsen, and P. Tiselius. Feeding and production of the carnivorous copepod Pareuchaeta norvegica in the Skagerrak. Marine Ecology Progress Series, 314:213– 225, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/314/m314p213.pdf.

Thor:2008:MRE

[TNT08] P. Thor, T. G. Nielsen, and P. Tiselius. Mortality rates of epipelagic copepods in the post-spring bloom period in Disko Bay, western Greenland. Marine Ecology Progress Series, 359:151–160, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/359/m359p151.pdf.

Talman:2004:HSS

[TNTH04] S. G. Talman, A. Norkko, S. F. Thrush, and J. E. Hewitt. Habitat structure and the survival of juvenile scallops Pecten novaezelandiae: comparing predation in habitats with varying complexity. Marine Ecology Progress Series, 269:197–207, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/269/m269p197.pdf.

Tremblay:2007:PCB

[TOBR07] R. Tremblay, F. Olivier, E. Bourget, and D. Rittschof. Physiological condition of Balanus amphitrite cyprid larvae determines habitat selection success. Marine Ecology Progress Series, 340:1–8, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Takasuka:2004:GSP

[TOK^{+}04] A. Takasuka, Y. Oozeki, R. Kimura, H. Kubota, and I. Aoki. Growth-selective predation hypothesis revisited for larval anchovy in offshore waters: cannibalism by juveniles versus predation by skipjack tunas. Marine Ecology Progress Series, 278:297–302, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/278/m278p297.pdf.

Takasuka:2008:MSR

[TOK08] A. Takasuka, Y. Oozeki, and H. Kubota. Multi-species regime shifts reflected in spawning temperature optima of small pelagic fish in the western North Pacific. Marine Ecology Progress Series, 360:211–217, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p211.pdf.

Tomasovych:2008:SER

[Tom08] A. Tomasovych. Substrate exploitation and resistance to biotic disturbance in the brachiopod Terebratalia transversa and the bivalve Pododesmus macrochisma. Marine Ecology Progress Series, 363:157–170, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/363/m363p157.pdf.

Toonen:2004:GEM

[Too04] R. J. Toonen. Genetic evidence of multiple paternity of broods in the intertidal crab Petrolisthes cinctipes. Marine Ecology Progress Series, 270:259–263, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/270/m270p259.pdf.

Torgersen:2001:VPE

[Tor01] T. Torgersen. Visual predation by the euphausiid Meganyctiphanes norvegica. Marine Ecology Progress Series, 209: 295–299, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/209/m209p295.pdf.

Torres:2009:KMB

[Tor09] L. G. Torres. A kaleidoscope of mammal, bird and fish: habitat use patterns of top predators and their prey in Florida Bay. Marine Ecology Progress Series, 375:289–304, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p289.pdf.

Toth:2000:LPI

[TP00a] G. Toth and H. Pavia. Lack of phlorotannin induction in the brown seaweed Ascophyllum nodosum in response to increased copper concentrations. Marine Ecology Progress Series, 192:119–126, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p119.pdf.

Tuck:2000:MPA

[TP00b] G. N. Tuck and H. P. Possingham. Marine protected areas for spatially structured exploited stocks. Marine Ecology Progress Series, 192:89–101, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p089.pdf.

Toonen:2001:SGTa

[TP01a] R. J. Toonen and J. R. Pawlik. Settlement of the gregarious tube worm Hydroides dianthus (Polychaeta: Serpulidae). I. Gregarious and nongregarious settlement. Marine Ecology Progress Series, 224:103–114, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/224/m224p103.pdf.

Toonen:2001:SGTb

[TP01b] R. J. Toonen and J. R. Pawlik. Settlement of the gregarious tube worm Hydroides dianthus (Polychaeta: Serpulidae). II. Testing the desperate larva hypothesis. Marine Ecology Progress Series, 224:115–131, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/224/m224p115.pdf.

Tapia:2007:SSD

[TP07] F. J. Tapia and J. Pineda. Stage-specific distribution of barnacle larvae in nearshore waters: potential for limited dispersal and high mortality rates. Marine Ecology Progress Series, 342:177–190, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/342/m342p177.pdf.

Thoma:2009:DSO

[TPBF09] J. N. Thoma, E. Pante, M. R. Brugler, and S. C. France. Deep-sea octocorals and antipatharians show no evidence of seamount-scale endemism in the NW Atlantic. Marine Ecology Progress Series, 397:25–35, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m397p025.pdf.

Trenkel:2005:STS

[TPD^{+}05] V. M. Trenkel, J. K. Pinnegar, W. A. Dawson, M. H. du Buit, and A. N. Tidd. Spatial and temporal structure of predator– prey relationships in the Celtic Sea fish community. Marine Ecology Progress Series, 299:257–268, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p257.pdf.

Teixido:2009:DDT

[TPG09] N. Teixidó, M. C. Pineda, and J. Garrabou. Decadal demographic trends of a long-lived temperate encrusting sponge. Marine Ecology Progress Series, 375:113–124, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p113.pdf.

Thiyagarajan:2007:JGB

[TPGQ07] V. Thiyagarajan, J. A. Pechenik, L. A. Gosselin, and P. Y. Qian. Juvenile growth in barnacles: combined effect of delayed metamorphosis and sub-lethal exposure of cyprids to low-salinity stress. Marine Ecology Progress Series, 344:173– 184, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/344/m344p173.pdf.

Teira:2003:RDO

[TPQ^{+}03] E. Teira, M. J. Pazó, M. Quevedo, M. V. Fuentes, F. X. Niell, and E. Fernández. Rates of dissolved organic carbon production and bacterial activity in the eastern North Atlantic Subtropical Gyre during summer. Marine Ecology Progress Series, 249:53–67, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p053.pdf.

Trimmer:2005:ILT

[TPS^{+}05] M. Trimmer, J. Petersen, D. B. Sivyer, C. Mills, E. Young, and E. R. Parker. Impact of long-term benthic trawl disturbance on sediment sorting and biogeochemistry in the southern North Sea. Marine Ecology Progress Series, 298: 79–94, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/298/m298p079.pdf.

Thompson:2000:LSP

[TPSW00] D. R. Thompson, R. A. Phillips, F. M. Stewart, and S. Waldron. Low δ ^{13}C signatures in pelagic seabirds: lipid ingestion as a potential source of ^{13}C-depleted carbon in the Procellariiformes. Marine Ecology Progress Series, 208:265–271, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
208/m208p265.pdf.

Thompson:2005:BBS

[TR05] B. A. W. Thompson and M. J. Riddle. Bioturbation behaviour of the spatangoid urchin Abatus ingens in Antarctic marine sediments. Marine Ecology Progress Series, 290: 135–143, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/290/m290p135.pdf.

Tolley:2006:PSH

[TR06] K. A. Tolley and P. E. Rosel. Population structure and historical demography of eastern North Atlantic harbour porpoises inferred through mtDNA sequences. Marine Ecology Progress Series, 327:297–308, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2007/327/m327p297.pdf.

Tang:2000:SDD

[TRDV00] K. W. Tang, D. R. Rogers, H. G. Dam, and P. T. Visscher. Seasonal distribution of DMSP among seston, dissolved matter and zooplankton along a transect in the Long Island Sound estuary. Marine Ecology Progress Series, 206:1–11, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
206/m206p001.pdf.

Tamelander:2006:TRP

[TRH^{+}06] T. Tamelander, P. E. Renaud, H. Hop, M. L. Carroll, W. G. Ambrose, Jr., and K. A. Hobson. Trophic relationships and pelagic-benthic coupling during summer in the Barents Sea marginal ice zone, revealed by stable carbon and nitrogen isotope measurements. Marine Ecology Progress Series, 310: 33–46, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/310/m310p033.pdf.

Trnski:2002:BSS

[Trn02] T. Trnski. Behaviour of settlement-stage larvae of fishes with an estuarine juvenile phase: in situ observations in a warm-temperate estuary. Marine Ecology Progress Series, 242:205–214, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/242/m242p205.pdf.

Trimmer:2006:DMA

[TRPNE06] M. Trimmer, N. Risgaard-Petersen, J. C. Nicholls, and P. Engström. Direct measurement of anaerobic ammonium oxidation (anammox) and denitrification in intact sediment cores. Marine Ecology Progress Series, 326:37–47, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/326/
m326p037.pdf.

Titelman:2006:TDJ

[TRS^{+}06] J. Titelman, L. Riemann, T. A. Sørnes, T. Nilsen, P. Griekspoor, and U. Båmstedt. Turnover of dead jellyfish: stimulation and retardation of microbial activity. Marine Ecology Progress Series, 325:43–58, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/325/m325p043.pdf.

Tomas:2004:SRS

[TRT04] F. Tomas, J. Romero, and X. Turon. Settlement and recruitment of the sea urchin Paracentrotus lividus in two contrasting habitats in the Mediterranean. Marine Ecology Progress Series, 282:173–184, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/282/m282p173.pdf.

Trussell:2002:ECV

[Tru02] G. C. Trussell. Evidence of countergradient variation in the growth of an intertidal snail in response to water velocity. Marine Ecology Progress Series, 243:123–131, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/243/
m243p123.pdf.

Thistle:2004:ENE

[TS04] D. Thistle and L. Sedlacek. Emergent and non-emergent species of harpacticoid copepods can be recognized morphologically. Marine Ecology Progress Series, 266:195–200, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/266/m266p195.pdf.

Taylor:2005:SRC

[TS05] D. I. Taylor and D. R. Schiel. Self-replacement and community modification by the southern bull kelp Durvillaea antarctica. Marine Ecology Progress Series, 288:87–102, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p087.pdf.

Travis:2006:GSN

[TS06] S. E. Travis and P. Sheridan. Genetic structure of natural and restored shoalgrass Halodule wrightii populations in the NW Gulf of Mexico. Marine Ecology Progress Series, 322:117–127, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p117.pdf.

Thayer:2007:STV

[TS07] J. A. Thayer and W. J. Sydeman. Spatio-temporal variability in prey harvest and reproductive ecology of a piscivorous seabird, Cerorhinca monocerata, in an upwelling system. Marine Ecology Progress Series, 329:253–265, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p253.pdf.

Thistle:2007:ECD

[TSC^{+}07] D. Thistle, L. Sedlacek, K. R. Carman, J. W. Fleeger, P. G. Brewer, and J. P. Barry. Exposure to carbon dioxide-rich seawater is stressful for some deep-sea species: an in situ, behavioral study. Marine Ecology Progress Series, 340:9– 16, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/340/m340p009.pdf.

Takami:2008:OMY

[TSE^{+}08] H. Takami, T. Saido, T. Endo, T. Noro, T. Musashi, and T. Kawamura. Overwinter mortality of young-of-the-year Ezo abalone in relation to seawater temperature on the North Pacific coast of Japan. Marine Ecology Progress Series, 367: 203–212, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/367/m367p203.pdf.

Tierney:2008:EUS

[TSEH08] M. Tierney, C. Southwell, L. M. Emmerson, and M. A. Hindell. Evaluating and using stable-isotope analysis to infer diet composition and foraging ecology of Adélie penguins Pygoscelis adeliae. Marine Ecology Progress Series, 355:297– 307, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/355/m355p297.pdf.

Teira:2001:PSS

[TSF01] E. Teira, P. Serret, and E. Fernández. Phytoplankton sizestructure, particulate and dissolved organic carbon production and oxygen fluxes through microbial communities in the NW Iberian coastal transition zone. Marine Ecology Progress Series, 219:65–83, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p065.pdf.

Tozzi:2004:HCC

[TSF04] S. Tozzi, O. Schofield, and P. Falkowski. Historical climate change and ocean turbulence as selective agents for two key phytoplankton functional groups. Marine Ecology Progress Series, 274:123–132, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/274/m274p123.pdf.

Thrush:2002:HSS

[TSHT02] S. F. Thrush, D. Schultz, J. E. Hewitt, and D. Talley. Habitat structure in soft-sediment environments and abundance of juvenile snapper Pagrus auratus. Marine Ecology Progress Series, 245:273–280, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p273.pdf.

Tzeng:2002:UOS

[TSI02] W. N. Tzeng, J. C. Shiao, and Y. Iizuka. Use of otolith Sr:Ca ratios to study the riverine migratory behaviors of Japanese eel Anguilla japonica. Marine Ecology Progress Series, 245:213–221, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/245/m245p213.pdf.

Tanaka:2006:GDR

[TSIY06] Y. Tanaka, K. Satoh, M. Iwahashi, and H. Yamada. Growthdependent recruitment of Pacific bluefin tuna Thunnus orientalis in the northwestern Pacific Ocean. Marine Ecology Progress Series, 319:225–235, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/319/m319p225.pdf.

Tweddle:2005:PCF

[TSJ05] J. F. Tweddle, J. H. Simpson, and C. D. Janzen. Physical controls of food supply to benthic filter feeders in the Menai Strait, UK. Marine Ecology Progress Series, 289: 79–88, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/289/m289p079.pdf.

Tuya:2005:IFF

[TSJH05] F. Tuya, P. Sanchez-Jerez, and R. J. Haroun. Influence of fishing and functional group of algae on sea urchin control of algal communities in the eastern Atlantic. Marine Ecology Progress Series, 287:255–260, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/287/m287p255.pdf.

Tovar:2009:MVC

[TSMC09] E. Tovar, E. Suárez-Morales, and L. Carrillo. Multiscale variability of the Chaetognatha along a Caribbean reef lagoon system. Marine Ecology Progress Series, 375:151–160, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/375/m375p151.pdf.

Tod:2005:PEC

[TSRR05] C. D. Tod, R. J. Stevenson, H. Reinardy, and M. G. Ritchie. Polyandry in the ectoparasitic copepod Lepeophtheirus salmonis despite complex precopulatory and postcopulatory mate-guarding. Marine Ecology Progress Series, 303:225–234, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/303/m303p225.pdf.

Thatje:2003:DTO

[TSSA03] S. Thatje, S. Schnack-Schiel, and W. E. Arntz. Developmental trade-offs in subantarctic meroplankton communities and the enigma of low decapod diversity in high southern latitudes. Marine Ecology Progress Series, 260:195–207, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p195.pdf.

Thiyagarajan:2006:STV

[TSSQ06] V. Thiyagarajan, L. Soo, P. K. S. Shin, and P. Y. Qian. Spatio-temporal variation in sediment biochemistry alters larval habitat selection and juvenile performance in the polychaete Capitella sp. I. Marine Ecology Progress Series, 327: 207–222, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/327/m327p207.pdf.

Tarling:2004:LCP

[TSW^{+}04] G. A. Tarling, R. S. Shreeve, P. Ward, A. Atkinson, and A. G. Hirst. Life-cycle phenotypic composition and mortality of Calanoides acutus (Copepoda: Calanoida) in the Scotia Sea: a modelling approach. Marine Ecology Progress Series, 272:165–181, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/272/m272p165.pdf.

Tarling:2007:LCP

[TSW07] G. A. Tarling, R. S. Shreeve, and P. Ward. Life-cycle and population dynamics of Rhincalanus gigas (Copepoda: Calanoida) in the Scotia Sea. Marine Ecology Progress Series, 338:145–158, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p145.pdf.

Treude:2009:BDS

[TSW^{+}09] T. Treude, C. R. Smith, F. Wenzhöfer, E. Carney, A. F. Bernardino, A. K. Hannides, M. Krüger, and A. Boetius. Biogeochemistry of a deep-sea whale fall: sulfate reduction, sulfide efflux and methanogenesis. Marine Ecology Progress Series, 382:1–21, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Townsend:2002:SNP

[TT02] D. W. Townsend and M. Thomas. Springtime nutrient and phytoplankton dynamics on Georges Bank. Marine Ecology Progress Series, 228:57–74, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/228/m228p057.pdf.

Takahashi:2005:OCF

[TT05a] K. Takahashi and P. Tiselius. Ontogenetic change of foraging behaviour during copepodite development of Acartia clausi. Marine Ecology Progress Series, 303:213–223, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/303/
m303p213.pdf.

Tonnesson:2005:DCS

[TT05b] K. Tönnesson and P. Tiselius. Diet of the chaetognaths Sagitta setosa and S. elegans in relation to prey abundance and vertical distribution. Marine Ecology Progress Series, 289:177–190, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/289/m289p177.pdf.

Towanda:2006:EBC

[TT06] T. Towanda and E. V. Thuesen. Ectosymbiotic behavior of Cancer gracilis and its trophic relationships with its host Phacellophora camtschatica and the parasitoid Hyperia medusarum. Marine Ecology Progress Series, 315:221– 236, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/315/m315p221.pdf.

Toonen:2007:ILW

[TT07a] R. J. Toonen and A. J. Tyre. If larvae were smart: a simple model for optimal settlement behavior of competent larvae. Marine Ecology Progress Series, 349:43–61, 2007. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m349p043.pdf.

Tyler:2007:JWC

[TT07b] R. M. Tyler and T. E. Targett. Juvenile weakfish Cynoscion regalis distribution in relation to diel-cycling dissolved oxygen in an estuarine tributary. Marine Ecology Progress Series, 333:257–269, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p257.pdf.

Taquet:2006:FGS

[TTD^{+}06] C. Taquet, M. Taquet, T. Dempster, M. Soria, S. Ciccione, D. Roos, and L. Dagorn. Foraging of the green sea turtle Chelonia mydas on seagrass beds at Mayotte Island (Indian Ocean), determined by acoustic transmitters. Marine Ecology Progress Series, 306:295–302, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/306/m306p295.pdf.

Tseng:2006:PGS

[TTL06] M. C. Tseng, W. N. Tzeng, and S. C. Lee. Population genetic structure of the Japanese eel Anguilla japonica in the northwest Pacific Ocean: evidence of non-panmictic populations. Marine Ecology Progress Series, 308:221–230, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/308/
m308p221.pdf.

Takagi:2005:DCE

[TTM^{+}05] Y. Takagi, H. Tohse, E. Murayama, T. Ohira, and H. Nagasawa. Diel changes in endolymph aragonite saturation rate and mRNA expression of otolith matrix proteins in the trout otolith organ. Marine Ecology Progress Series, 294: 249–256, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/294/m294p249.pdf.

Tioho:2001:EAR

[TTN01] H. Tioho, M. Tokeshi, and S. Nojima. Experimental analysis of recruitment in a scleractinian coral at high latitude. Marine Ecology Progress Series, 213:79–86, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/213/m213p079.pdf.

Tomas:2005:EHP

[TTR05a] F. Tomas, X. Turon, and J. Romero. Effects of herbivores on a Posidonia oceanica seagrass meadow: importance of epiphytes. Marine Ecology Progress Series, 287:

115–125, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/287/m287p115.pdf.

Tomas:2005:SSS

[TTR05b] F. Tomas, X. Turon, and J. Romero. Seasonal and smallscale spatial variability of herbivory pressure on the temperate seagrass Posidonia oceanica. Marine Ecology Progress Series, 301:95–107, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p095.pdf.

Taira:2009:SEV

[TUK09] Y. Taira, M. Uchimiya, and I. Kudo. Simultaneous estimation of viral lysis and protozoan grazing on bacterial mortality using a modified virus-dilution method. Marine Ecology Progress Series, 379:23–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/379/m379p023.pdf.

Tupper:2007:INH

[Tup07] M. Tupper. Identification of nursery habitats for commercially valuable humphead wrasse Cheilinus undulatus and large groupers (Pisces: Serranidae) in Palau. Marine Ecology Progress Series, 332:189–199, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p189.pdf.

Thampanya:2002:CSC

[TVD02] U. Thampanya, J. E. Vermaat, and C. M. Duarte. Colonization success of common Thai mangrove species as a function of shelter from water movement. Marine Ecology Progress Series, 237:111–120, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/237/m237p111.pdf.

Thiermann:2000:STM

[TVG00] F. Thiermann, B. Vismann, and O. Giere. Sulphide tolerance of the marine nematode Oncholaimus campylocercoides — a result of internal sulphur formation? Marine Ecology Progress Series, 193:251–259, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/193/m193p251.pdf.

Tankersley:2000:PRP

[TW00] R. A. Tankersley and M. G. Wieber. Physiological responses of postlarval and juvenile blue crabs Callinectes sapidus to hypoxia and anoxia. Marine Ecology Progress Series, 194: 179–191, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/194/m194p179.pdf.

Takahashi:2004:DGR

[TW04a] M. Takahashi and Y. Watanabe. Developmental and growth rates of Japanese anchovy Engraulis japonicus during metamorphosis in the kuroshio-oyashio transitional waters. Marine Ecology Progress Series, 282:253–260, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/282/
m282p253.pdf.

Takahashi:2004:GRD

[TW04b] M. Takahashi and Y. Watanabe. Growth rate-dependent recruitment of Japanese anchovy Engraulis japonicus in the kuroshio-oyashio transitional waters. Marine Ecology Progress Series, 266:227–238, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/266/m266p227.pdf.

Tougaard:2010:GMS

[TW10] J. Tougaard and M. S. Wisz. General models of the spatial distribution of porpoises require representative data and parsimony: Comment on Skov & Thomsen (2008). Marine Ecology Progress Series, 399:295–297, 2010. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m399p295.pdf. See

[ST08].

Tsutsumi:2001:ECF

[TWM^{+}01] H. Tsutsumi, S. Wainright, S. Montani, M. Saga, S. Ichihara, and K. Kogure. Exploitation of a chemosynthetic food resource by the polychaete Capitella sp. Marine Ecology Progress Series, 216:119–127, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/216/m216p119.pdf.

Todd:2006:ESL

[TWMW06] C. D. Todd, B. D. M. Whyte, J. C. MacLean, and A. M. Walker. Ectoparasitic sea lice (Lepeophtheirus salmonis and Caligus elongatus) infestations of wild, adult, one sea-winter Atlantic salmon Salmo salar returning to Scotland. Marine Ecology Progress Series, 328:183–193, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/328/
m328p183.pdf.

Tuya:2008:TAC

[TWT08] F. Tuya, T. Wernberg, and M. S. Thomsen. Testing the ‘abundant centre’ hypothesis on endemic reef fishes in southwestern Australia. Marine Ecology Progress Series, 372:225– 230, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p225.pdf.

Taylor:2003:EAS

[TWYS03] G. T. Taylor, J. Way, Y. Yu, and M. I. Scranton. Ectohydrolase activity in surface waters of the Hudson River and western Long Island Sound estuaries. Marine Ecology Progress Series, 263:1–15, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p001.pdf.

Tyack:2009:APE

[Tya09] P. Tyack. Acoustic playback experiments to study behavioral responses of free-ranging marine animals to anthropogenic sound. Marine Ecology Progress Series, 395:187–200, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p187.pdf.

Tyler:2000:TPT

[TYC00] P. A. Tyler, C. M. Young, and A. Clarke. Temperature and pressure tolerances of embryos and larvae of the Antarctic sea urchin Sterechinus neumayeri (Echinodermata: Echinoidea): potential for deep-sea invasion from high latitudes. Marine Ecology Progress Series, 192:173–180, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
192/m192p173.pdf.

Tyler:2007:SGR

[TYD07] P. A. Tyler, C. M. Young, and F. Dove. Settlement, growth and reproduction in the deep-sea wood-boring bivalve mollusc Xylophaga depalmai. Marine Ecology Progress Series, 343:151–159, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p151.pdf.

Tynan:2004:CPS

[Tyn04] C. T. Tynan. Cetacean populations on the SE Bering Sea shelf during the late 1990s: implications for decadal changes in ecosystem structure and carbon flow. Marine Ecology Progress Series, 272:281–300, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/272/m272p281.pdf.

Takai:2004:TPM

[TYT^{+}04] N. Takai, A. Yorozu, T. Tanimoto, A. Hoshika, and K. Yoshihara. Transport pathways of microphytobenthos-originating organic carbon in the food web of an exposed hard bottom shore in the Seto Inland Sea, Japan. Marine Ecology Progress Series, 284:97–108, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p097.pdf.

Uthicke:2001:RGF

[UB01] S. Uthicke and J. A. H. Benzie. Restricted gene flow between Holothuria scabra (Echinodermata: Holothuroidea) populations along the north-east coast of Australia and the Solomon Islands. Marine Ecology Progress Series, 216:109–117, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p109.pdf.

Uthicke:2002:GFR

[UB02] S. Uthicke and J. A. H. Benzie. A genetic fingerprint recapture technique for measuring growth in ’unmarkable’ invertebrates: negative growth in commercially fished holothurians (Holothuria nobilis). Marine Ecology Progress Series, 241:221–226, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p221.pdf.

Umani:2003:SVD

[UB03] S. Fonda Umani and A. Beran. Seasonal variations in the dynamics of microbial plankton communities: first estimates from experiments in the Gulf of Trieste, Northern Adriatic Sea. Marine Ecology Progress Series, 247:1–16, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p001.pdf.

Urcuyo:2007:GLT

[UBM^{+}07] I. A. Urcuyo, D. C. Bergquist, I. R. MacDonald, M. Van-Horn, and C. R. Fisher. Growth and longevity of the tubeworm Ridgeia piscesae in the variable diffuse flow habitats of the Juan de Fuca Ridge. Marine Ecology Progress Series, 344:143–157, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/344/m344p143.pdf.

Ulstrup:2006:VBS

[UBRvO06] K. E. Ulstrup, R. Berkelmans, P. J. Ralph, and M. J. H. van Oppen. Variation in bleaching sensitivity of two coral species across a latitudinal gradient on the Great Barrier Reef: the role of zooxanthellae. Marine Ecology Progress Series, 314:135–148, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/314/m314p135.pdf.

Uglem:2009:HCS

[UDB^{+}09] I. Uglem, T. Dempster, P. A. Bjørn, P. Sanchez-Jerez, and F. Økland. High connectivity of salmon farms revealed by aggregation, residence and repeated movements of wild fish among farms. Marine Ecology Progress Series, 384:251– 260, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/384/m384p251.pdf.

Ugland:2004:ESR

[UG04] K. I. Ugland and J. S. Gray. Estimation of species richness: analysis of the methods developed by Chao and Karakassis. Marine Ecology Progress Series, 284:1–8, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/284/
m284p001.pdf.

Uhrin:2003:EPS

[UH03] A. V. Uhrin and J. G. Holmquist. Effects of propeller scarring on macrofaunal use of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 250:61–70, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/250/
m250p061.pdf.

Ulstrup:2005:PIH

[UHR05] K. E. Ulstrup, R. Hill, and P. J. Ralph. Photosynthetic impact of hypoxia on in hospite zooxanthellae in the scleractinian coral Pocillopora damicornis. Marine Ecology Progress Series, 286:125–132, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/286/m286p125.pdf.

Ulstrup:2008:SVP

[UHvO^{+}08] K. E. Ulstrup, R. Hill, M. J. H. van Oppen, A. W. D. Larkum, and P. J. Ralph. Seasonal variation in the photo-physiology of homogeneous and heterogeneous Symbiodinium consortia in two scleractinian corals. Marine Ecology Progress Series, 361:139–150, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/361/m361p139.pdf.

Uiblein:2003:BHU

[ULL03] F. Uiblein, P. Lorance, and D. Latrouite. Behaviour and habitat utilization of seven demersal fish species on the Bay of Biscay continental slope, NE Atlantic. Marine Ecology Progress Series, 257:223–232, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/257/m257p223.pdf.

Uronen:2005:HAA

[ULL^{+}05] P. Uronen, S. Lehtinen, C. Legrand, P. Kuuppo, and T. Tamminen. Haemolytic activity and allelopathy of the haptophyte Prymnesium parvum in nutrient-limited and balanced growth conditions. Marine Ecology Progress Series, 299:137– 148, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/299/m299p137.pdf.

Underwood:2008:UPF

[UM08] A. J. Underwood and R. J. Murphy. Unexpected patterns of facilitatory grazing revealed by quantitative imaging. Marine Ecology Progress Series, 358:85–94, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p085.pdf.

Underwood:2004:LOF

[Und04] A. J. Underwood. Landing on one’s foot: small-scale topographic features of habitat and the dispersion of juvenile intertidal gastropods. Marine Ecology Progress Series, 268: 173–182, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/268/m268p173.pdf.

Urrutia:2001:PSP

[UNII01] M. B. Urrutia, E. Navarro, I. Ibarrola, and J. I. P. Iglesias. Preingestive selection processes in the cockle Cerastoderma edule: mucus production related to rejection of pseudofaeces. Marine Ecology Progress Series, 209:177–187, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
209/m209p177.pdf.

Ullberg:2003:FLM

[UÓ03] J. Ullberg and E. Ólafsson. Free-living marine nematodes actively choose habitat when descending from the water column. Marine Ecology Progress Series, 260:141–149, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/260/
m260p141.pdf.

Utne-Palm:2001:RNV

[UP01] A. C. Utne-Palm. Response of naïve two-spotted gobies Gobiusculus flavescens to visual and chemical stimuli of their natural predator, cod Gadus morhua. Marine Ecology Progress Series, 218:267–274, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/218/m218p267.pdf.

Untersee:2007:LAM

[UP07] S. Untersee and J. A. Pechenik. Local adaptation and maternal effects in two species of marine gastropod (genus Crepidula) that differ in dispersal potential. Marine Ecology Progress Series, 347:79–85, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/347/m347p079.pdf.

Urban:2002:MGD

[Urb02] H. J. Urban. Modeling growth of different developmental stages in bivalves. Marine Ecology Progress Series, 238:109– 114, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p109.pdf.

Urban-Rich:2006:GIC

[URFA06] J. Urban-Rich, D. Fernández, and J. L. Acuña. Grazing impact on chromophoric dissolved organic matter (CDOM) by the larvacean Oikopleura dioica. Marine Ecology Progress Series, 317:101–110, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p101.pdf.

Unsworth:2008:HCI

[USG^{+}08] R. K. F. Unsworth, P. Salinas De León, S. L. Garrard, J. Jompa, D. J. Smith, and J. J. Bell. High connectivity of Indo–Pacific seagrass fish assemblages with mangrove and coral reef habitats. Marine Ecology Progress Series, 353:213–224, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/353/m353p213.pdf.

Usher:2004:BPC

[UST^{+}04] K. M. Usher, D. C. Sutton, S. Toze, J. Kuo, and J. Fromont. Biogeography and phylogeny of Chondrilla species (Demospongiae) in Australia. Marine Ecology Progress Series, 270: 117–127, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/270/m270p117.pdf.

Uthicke:2001:IBS

[Uth01] S. Uthicke. Interactions between sediment-feeders and microalgae on coral reefs: grazing losses versus production enhancement. Marine Ecology Progress Series, 210:125–138, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/210/m210p125.pdf.

Uthicke:2006:PER

[Uth06] S. Uthicke. Photosynthetic efficiency and rapid light curves of sediment-biofilms along a water quality gradient in the Great Barrier Reef, Australia. Marine Ecology Progress Series, 322:61–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/322/m322p061.pdf.

Urakawa:2001:CMC

[UYNO01] H. Urakawa, T. Yoshida, M. Nishimura, and K. Ohwada. Characterization of microbial communities in marine surface sediments by terminal-restriction fragment length polymorphism (T-RFLP) analysis and quinone profiling. Marine Ecology Progress Series, 220:47–57, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/220/m220p047.pdf.

Viejo:2001:EDV

[VÅ01] R. M. Viejo and P. Åberg. Effects of density on the vital rates of a modular seaweed. Marine Ecology Progress Series, 221:105–115, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/221/m221p105.pdf.

Verges:2009:RFH

[VAB09] A. Vergés, T. Alcoverro, and E. Ballesteros. Role of fish herbivory in structuring the vertical distribution of canopy algae Cystoseira spp. in the Mediterranean Sea. Marine Ecology Progress Series, 375:1–11, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Villegas-Amtmann:2008:MFS

[VACT^{+}08] S. Villegas-Amtmann, D. P. Costa, Y. Tremblay, S. Salazar, and D. Aurioles-Gamboa. Multiple foraging strategies in a marine apex predator, the Galapagos sea lion Zalophus wollebaeki. Marine Ecology Progress Series, 363:299–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/363/m363p299.pdf.

Villanueva:2003:ELL

[VAJ^{+}03] R. Villanueva, A. Arkhipkin, P. Jereb, E. Lefkaditou, M. R. Lipinski, C. Perales-Raya, J. Riba, and F. Rocha. Embryonic life of the loliginid squid Loligo vulgaris: comparison between statoliths of Atlantic and Mediterranean populations. Marine Ecology Progress Series, 253:197–208, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p197.pdf.

VanDover:2002:CSM

[Van02] C. L. Van Dover. Community structure of mussel beds at deep-sea hydrothermal vents. Marine Ecology Progress Series, 230:137–158, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p137.pdf.

VanDover:2003:VCS

[Van03] C. L. Van Dover. Variation in community structure within hydrothermal vent mussel beds of the East Pacific Rise. Marine Ecology Progress Series, 253:55–66, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/253/
m253p055.pdf.

Vazquez:2007:DMC

[VAP^{+}07] E. Vázquez, J. Ameneiro, S. Putzeys, C. Gordo, and P. Sangrà. Distribution of meroplankton communities in the Bransfield Strait, Antarctica. Marine Ecology Progress Series, 338:119–129, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p119.pdf.

Vermeij:2002:HCP

[VB02] M. J. A. Vermeij and R. P. M. Bak. How are coral populations structured by light? Marine light regimes and the distribution of Madracis. Marine Ecology Progress Series, 233:105–116, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/233/m233p105.pdf.

Virgilio:2006:MDA

[VBA06] M. Virgilio, T. Backeljau, and M. Abbiati. Mitochondrial DNA and allozyme patterns of Hediste diversicolor (Polychaeta: Nereididae): the importance of small scale genetic structuring. Marine Ecology Progress Series, 326:157–165, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/326/m326p157.pdf.

Verges:2007:EEC

[VBAR07] A. Vergés, M. A. Becerro, T. Alcoverro, and J. Romero. Experimental evidence of chemical deterrence against multiple herbivores in the seagrass Posidonia oceanica. Marine Ecology Progress Series, 343:107–114, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/343/m343p107.pdf.

Veliz:2006:FEA

[VBB06] D. Véliz, E. Bourget, and L. Bernatchez. Field evidence for an association between growth and protein polymorphism in the acorn barnacle Semibalanus balanoides. Marine Ecology Progress Series, 308:197–206, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/308/m308p197.pdf.

Vilchis:2006:PHS

[VBF06] L. I. Vilchis, L. T. Ballance, and P. C. Fiedler. Pelagic habitat of seabirds in the eastern tropical Pacific: effects of foraging ecology on habitat selection. Marine Ecology Progress Series, 315:279–292, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/315/m315p279.pdf.

Votier:2008:TSV

[VBFF08] S. C. Votier, S. Bearhop, R. Fyfe, and R. W. Furness. Temporal and spatial variation in the diet of a marine top predator — links with commercial fisheries. Marine Ecology Progress Series, 367:223–232, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/367/m367p223.pdf.

Vaselli:2008:ACS

[VBMBC08a] S. Vaselli, I. Bertocci, E. Maggi, and L. Benedetti-Cecchi. Assessing the consequences of sea level rise: effects of changes in the slope of the substratum on sessile assemblages of rocky seashores. Marine Ecology Progress Series, 368:9–22, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/368/m368p009.pdf.

Vaselli:2008:EMI

[VBMBC08b] S. Vaselli, I. Bertocci, E. Maggi, and L. Benedetti-Cecchi. Effects of mean intensity and temporal variance of sediment scouring events on assemblages of rocky shores. Marine Ecology Progress Series, 364:57–66, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p057.pdf.

Visintainer:2006:CCD

[VBS06] T. A. Visintainer, S. M. Bollens, and C. Simenstad. Community composition and diet of fishes as a function of tidal channel geomorphology. Marine Ecology Progress Series, 321: 227–243, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/321/m321p227.pdf.

vanBergeijk:2002:PCD

[vBSSH02] S. A. van Bergeijk, K. Schönefeldt, L. J. Stal, and J. Huisman. Production and consumption of dimethylsulfide (DMS) and dimethylsulfoniopropionate (DMSP) in a diatom-dominated intertidal sediment. Marine Ecology Progress Series, 231: 37–46, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/231/m231p037.pdf.

Vanhove:2000:SVB

[VBV^{+}00] S. Vanhove, M. Beghyn, D. Van Gansbeke, L. W. Bullough, and M. Vincx. A seasonally varying biotope at Signy Island, Antarctic: implications for meiofaunal structure. Marine Ecology Progress Series, 202:13–25, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/202/m202p013.pdf.

vanBergeijk:2006:SVD

[vBWSS06] S. A. van Bergeijk, U. Wollenzien, K. Schönefeldt, and L. J. Stal. Seasonal variation in dimethylsulfoniopropionate related to microphytobenthos composition in intertidal estuarine sediments. Marine Ecology Progress Series, 320:55– 63, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/320/m320p055.pdf.

VandenBerghe:2009:MID

[VCA^{+}09] E. Vanden Berghe, S. Claus, W. Appeltans, S. Faulwetter, C. Arvanitidis, P. J. Somerfield, I. F. Aleffi, J. M. Amouroux, N. Anisimova, G. Bachelet, S. J. Cochrane, M. J. Costello, J. Craeymeersch, S. Dahle, S. Degraer, S. Denisenko, C. Dounas, G. Duineveld, C. Emblow, V. Escaravage, M. C. Fabri, D. Fleischer, A. Grémare, M. Herrmann, H. Hummel, I. Karakassis, M. Kedra, M. A. Kendall, P. Kingston, L. Kotwicki, C. Labrune, J. Laudien, E. L. Nevrova, A. Occhipinti-Ambrogi, F. Olsgard, R. Palerud, A. Petrov, E. Rachor, N. Revkov, H. Rumohr, R. Sardá, W. C. H. Sistermans, J. Speybroeck, U. Janas, G. Van Hoey, M. Vincx, P. Whomersley, W. Willems, M. Włodarska-Kowalczuk, A. Zenetos, M. L. Zettler, and C. H. R. Heip. MacroBen integrated database on benthic invertebrates of European continental shelves: a tool for large-scale analysis across Europe. Marine Ecology Progress Series, 382:225–238, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m382p225.pdf.

Voss:2006:BSL

[VCBH06] R. Voss, C. Clemmesen, H. Baumann, and H. H. Hinrichsen. Baltic sprat larvae: coupling food availability, larval condition and survival. Marine Ecology Progress Series, 308: 243–254, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/308/m308p243.pdf.

VanParijs:2009:MRA

[VCSL^{+}09] S. M. Van Parijs, C. W. Clark, R. S. Sousa-Lima, S. E. Parks, S. Rankin, D. Risch, and I. C. Van Opzeeland. Management and research applications of real-time and archival passive acoustic sensors over varying temporal and spatial scales. Marine Ecology Progress Series, 395:21–36, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m395p021.pdf.

Vidal:2000:NAD

[VD00] M. Vidal and C. M. Duarte. Nutrient accumulation at different supply rates in experimental Mediterranean planktonic communities. Marine Ecology Progress Series, 207:1–11, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
207/m207p001.pdf.

Vidal:2003:APA

[VDA^{+}03] M. Vidal, C. M. Duarte, S. Agustí, J. M. Gasol, and D. Vaqué. Alkaline phosphatase activities in the central Atlantic Ocean indicate large areas with phosphorus deficiency. Marine Ecology Progress Series, 262:43–53, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/262/
m262p043.pdf.

vandeBund:2001:ECB

[vdBÓME01] W. J. van de Bund, E. Ólafsson, H. Modig, and R. Elmgren. Effects of the coexisting Baltic amphipods Monoporeia affinis and Pontoporeia femorata on the fate of a simulated spring diatom bloom. Marine Ecology Progress Series, 212:107–115, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/212/m212p107.pdf.

vanDuyl:2008:DCF

[vDHHM08] F. C. van Duyl, J. Hegeman, A. Hoogstraten, and C. Maier. Dissolved carbon fixation by sponge-microbe consortia of deep water coral mounds in the northeastern Atlantic Ocean. Marine Ecology Progress Series, 358:137–150, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p137.pdf.

vanderMeeren:2003:GMP

[vdMM03] T. van der Meeren and E. Moksness. Growth and mortality patterns evaluated from otolith microstructure in Atlantic cod Gadus morhua larvae reared on different feeding regimes in mesocosms. Marine Ecology Progress Series, 264: 95–107, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/264/m264p095.pdf.

vonderMeden:2008:CTA

[vdMPEM08] C. E. O. von der Meden, F. Porri, J. Erlandsson, and C. D. McQuaid. Coastline topography affects the distribution of indigenous and invasive mussels. Marine Ecology Progress Series, 372:135–145, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/372/m372p135.pdf.

vonDassow:2005:FEN

[vDS05] Y. J. von Dassow and R. R. Strathmann. Full of eggs and no place to lay them: hidden cost of benthic development. Marine Ecology Progress Series, 294:23–34, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p023.pdf.

vanDongen:2002:CIV

[vDSD02] B. E. van Dongen, S. Schouten, and J. S. Sinninghe Damsté. Carbon isotope variability in monosaccharides and lipids of aquatic algae and terrestrial plants. Marine Ecology Progress Series, 232:83–92, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/232/m232p083.pdf.

vanderVeer:2009:VTF

[vdVBGW09] H. W. van der Veer, L. J. Bolle, A. J. Geffen, and J. I. J. Witte. Variability in transport of fish eggs and larvae. IV. Interannual variability in larval stage duration of immigrating plaice in the Dutch Wadden Sea. Marine Ecology Progress Series, 390:213–223, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/390/m390p213.pdf.

vanderVeer:2000:SGM

[vdVBW00] H. W. van der Veer, B. Bies, and J. I. J. Witte. Selective growth and mortality of juvenile 0-group plaice Pleuronectes platessa in the Dutch Wadden Sea: a consequence of irreversible non-genetic adaptation during early pelagic life. Marine Ecology Progress Series, 197:273–283, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p273.pdf.

vanDeurs:2009:RLS

[vDvHT^{+}09] M. van Deurs, R. van Hal, M. T. Tomczak, S. H. Jónasdóttir, and P. Dolmer. Recruitment of lesser sandeel Ammodytes marinus in relation to density dependence and zooplankton composition. Marine Ecology Progress Series, 381:249–258, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m381p249.pdf.

vanDijk:2009:HLG

[vDvTJD^{+}09] J. K. van Dijk, B. I. van Tussenbroek, K. Jiménez-Durán, G. J. Márquez-Guzmán, and J. Ouborg. High levels of gene flow and low population genetic structure related to high dispersal potential of a tropical marine angiosperm. Marine Ecology Progress Series, 390:67–77, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/390/m390p067.pdf.

vanderWal:2008:DDI

[vdWHF^{+}08] D. van der Wal, P. M. J. Herman, R. M. Forster, T. Ysebaert, F. Rossi, E. Knaeps, Y. M. G. Plancke, and S. J. Ides. Distribution and dynamics of intertidal macrobenthos predicted from remote sensing: response to microphytobenthos and environment. Marine Ecology Progress Series, 367:57–72, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m367p057.pdf.

Ventura:2006:LBE

[Ven06] M. Ventura. Linking biochemical and elemental composition in freshwater and marine crustacean zooplankton. Marine Ecology Progress Series, 327:233–246, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/327/
m327p233.pdf.

Vermeij:2005:SCA

[Ver05] M. J. A. Vermeij. Substrate composition and adult distribution determine recruitment patterns in a Caribbean brooding coral. Marine Ecology Progress Series, 295:123–133, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/295/m295p123.pdf.

Vogel:2009:TTF

[VF09] C. Vogel and N. S. Fisher. Trophic transfer of Fe, Zn and Am from marine bacteria to a planktonic ciliate. Marine Ecology Progress Series, 384:61–68, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/384/m384p061.pdf.

Verdy:2009:OPC

[VFF09] A. Verdy, M. Follows, and G. Flierl. Optimal phytoplankton cell size in an allometric model. Marine Ecology Progress Series, 379:1–12, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Vandieken:2006:PCO

[VFJ06] V. Vandieken, N. Finke, and B. B. Jørgensen. Pathways of carbon oxidation in an Arctic fjord sediment (Svalbard) and isolation of psychrophilic and psychrotolerant Fe(III)- reducing bacteria. Marine Ecology Progress Series, 322:29– 41, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p029.pdf.

Vermeij:2007:ERM

[VFJ^{+}07] M. J. A. Vermeij, P. R. Frade, R. I. R. Jacinto, A. O. Debrot, and R. P. M. Bak. Effects of reproductive mode on habitat-related differences in the population structure of eight Caribbean coral species. Marine Ecology Progress Series, 351:91–102, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/351/m351p091.pdf.

Vermeij:2006:PCA

[VFM06] M. J. A. Vermeij, N. D. Fogarty, and M. W. Miller. Pelagic conditions affect larval behavior, survival, and settlement patterns in the Caribbean coral Montastraea faveolata. Marine Ecology Progress Series, 310:119–128, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p119.pdf.

Vitaliano:2007:RBM

[VFP^{+}07] J. J. Vitaliano, S. A. Fromm, D. B. Packer, R. N. Reid, and R. A. Pikanowski. Recovery of benthic macrofauna from sewage sludge disposal in the New York Bight. Marine Ecology Progress Series, 342:27–40, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/342/m342p027.pdf.

Viitasalo:2001:ZBS

[VFV01] M. Viitasalo, J. Flinkman, and M. Viherluoto. Zooplanktivory in the Baltic Sea: a comparison of prey selectivity by Clupea harengus and Mysis mixta, with reference to prey escape reactions. Marine Ecology Progress Series, 216:191–200, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p191.pdf.

Vasquez:2005:STU

[VGB^{+}05] E. A. Vasquez, E. P. Glenn, J. J. Brown, G. R. Guntenspergen, and S. G. Nelson. Salt tolerance underlies the cryptic invasion of North American salt marshes by an introduced haplotype of the common reed Phragmites australis (Poaceae). Marine Ecology Progress Series, 298:1–8, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Vila:2001:DTD

[VGMC01] M. Vila, E. Garcés, M. Masó, and J. Camp. Is the distribution of the toxic dinoflagellate Alexandrium catenella expanding along the NW Mediterranean coast? Marine Ecology Progress Series, 222:73–83, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/222/m222p073.pdf.

VanWambeke:2001:BDD

[VGS^{+}01] F. Van Wambeke, M. Goutx, L. Striby, R. Sempéré, and F. Vidussi. Bacterial dynamics during the transition from spring bloom to oligotrophy in the northwestern Mediterranean Sea: relationships with particulate detritus and dissolved organic matter. Marine Ecology Progress Series, 212: 89–105, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/212/m212p089.pdf.

VanderVeer:2000:ESY

[VGW00] H. W. Van der Veer, A. J. Geffen, and J. I. J. Witte. Exceptionally strong year classes in plaice Pleuronectes platessa: are they generated during the pelagic stage only, or also in the juvenile stage? Marine Ecology Progress Series, 199: 255–262, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/199/m199p255.pdf.

VanAlstyne:2003:DRD

[VH03] K. L. Van Alstyne and L. T. Houser. Dimethylsulfide release during macroinvertebrate grazing and its role as an activated chemical defense. Marine Ecology Progress Series, 250:175–181, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/250/m250p175.pdf.

Valentine:2007:FWI

[VHB^{+}07] J. F. Valentine, K. L. Heck, Jr., D. Blackmon, M. E. Goecker, J. Christian, R. M. Kroutil, K. D. Kirsch, B. J. Peterson, M. Beck, and M. A. Vanderklift. Food web interactions along seagrass-coral reef boundaries: effects of piscivore reductions on cross-habitat energy exchange. Marine Ecology Progress Series, 333:37–50, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p037.pdf.

Vezina:2008:I

[VHG08] A. F. Vézina and O. Hoegh-Guldberg. Introduction. Marine Ecology Progress Series, 373:199–201, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m373p199.pdf.

vonHerbing:2001:MCP

[vHGH01] I. H. von Herbing, S. M. Gallager, and W. Halteman. Metabolic costs of pursuit and attack in early larval Atlantic cod. Marine Ecology Progress Series, 216:201–212, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
216/m216p201.pdf.

Vance:2002:DJP

[VHH^{+}02] D. J. Vance, M. D. E. Haywood, D. S. Heales, R. A. Kenyon, N. R. Loneragan, and R. C. Pendreyne. Distribution of juvenile penaeid prawns in mangrove forests in a tropical Australian estuary, with particular reference to Penaeus merguiensis. Marine Ecology Progress Series, 228:165– 177, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/228/m228p165.pdf.

Valles:2009:VTR

[VHK09] H. Vallès, W. Hunte, and D. L. Kramer. Variable temporal relationships between environment and recruitment in coral reef fishes. Marine Ecology Progress Series, 379:225– 240, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/379/m379p225.pdf.

Valentine:2000:RSU

[VHKW00] J. F. Valentine, K. L. Heck, Jr., K. D. Kirsch, and D. Webb. Role of sea urchin Lytechinus variegatus grazing in regulating subtropical turtlegrass Thalassia testudinum meadows in the Florida Keys (USA). Marine Ecology Progress Series, 200:213–228, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/200/m200p213.pdf.

Virnstein:2009:PPM

[VHM09] R. W. Virnstein, L. C. Hayek, and L. J. Morris. Pulsating patches: a model for the spatial and temporal dynamics of the threatened seagrass Halophila johnsonii. Marine Ecology Progress Series, 385:97–109, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/385/m385p097.pdf.

Votier:2009:CTE

[VHMB09] S. C. Votier, B. J. Hatchwell, M. Mears, and T. R. Birkhead. Changes in the timing of egg-laying of a colonial seabird in relation to population size and environmental conditions. Marine Ecology Progress Series, 393:225–233, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m393p225.pdf.

Valdivia:2005:EDD

[VHT^{+}05] N. Valdivia, A. Heidemann, M. Thiel, M. Molis, and M. Wahl. Effects of disturbance on the diversity of hard-bottom macrobenthic communities on the coast of Chile. Marine Ecology Progress Series, 299:45–54, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/299/m299p045.pdf.

Villinski:2004:CIS

[VHV^{+}04] J. C. Villinski, J. M. Hayes, J. T. Villinski, S. C. Brassell, and R. A. Raff. Carbon-isotopic shifts associated with heterotrophy and biosynthetic pathways in direct- and indirectdeveloping sea urchins. Marine Ecology Progress Series, 275: 139–151, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/275/m275p139.pdf.

Vanderklift:2007:PRI

[VHW^{+}07] M. A. Vanderklift, J. How, T. Wernberg, L. D. MacArthur, K. L. Heck, Jr., and J. F. Valentine. Proximity to reef influences density of small predatory fishes, while type of seagrass influences intensity of their predation on crabs. Marine Ecology Progress Series, 340:235–243, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/340/
m340p235.pdf.

Viant:2007:MAO

[Via07] M. R. Viant. Metabolomics of aquatic organisms: the new ‘omics’ on the block. Marine Ecology Progress Series, 332: 301–306, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m332p301.pdf.

Viejo:2009:RIR

[Vie09] R. M. Viejo. Resilience in intertidal rocky shore assemblages across the stress gradient created by emersion times. Marine Ecology Progress Series, 390:55–65, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/390/
m390p055.pdf.

Viitasalo:2007:EBT

[Vii07] S. Viitasalo. Effects of bioturbation by three macrozoobenthic species and predation by necto-benthic mysids on cladoceran benthic eggs. Marine Ecology Progress Series, 336: 131–140, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/336/m336p131.pdf.

Visser:2001:HSP

[Vis01] A. W. Visser. Hydromechanical signals in the plankton. Marine Ecology Progress Series, 222:1–24, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p001.pdf.

Vanderklift:2003:PFA

[VJ03] M. A. Vanderklift and C. A. Jacoby. Patterns in fish assemblages 25 years after major seagrass loss. Marine Ecology Progress Series, 247:225–235, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/247/m247p225.pdf.

Visser:2004:CCP

[VJ04] A. W. Visser and G. A. Jackson. Characteristics of the chemical plume behind a sinking particle in a turbulent water column. Marine Ecology Progress Series, 283:55–71, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/283/m283p055.pdf.

Valentine:2005:PEK

[VJ05] J. P. Valentine and C. R. Johnson. Persistence of the exotic kelp Undaria pinnatifida does not depend on sea urchin grazing. Marine Ecology Progress Series, 285:43–55, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2004/285/m285p043.pdf.

Vikebo:2007:DGS

[VJKF07] F. Vikebø, C. Jørgensen, T. Kristiansen, and Ó. Fiksen. Drift, growth, and survival of larval Northeast Arctic cod with simple rules of behaviour. Marine Ecology Progress Series, 347:207–219, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m347p207.pdf.

VanMooy:2002:SVS

[VK02] B. A. S. Van Mooy and R. G. Keil. Seasonal variation in sedimentary amino acids and the association of organic matter with mineral surfaces in a sandy eelgrass meadow. Marine Ecology Progress Series, 227:275–280, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p275.pdf.

Vanderklift:2005:CIS

[VK05] M. A. Vanderklift and G. A. Kendrick. Contrasting influence of sea urchins on attached and drift macroalgae. Marine Ecology Progress Series, 299:101–110, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/299/
m299p101.pdf.

vanKatwijk:2000:EWD

[vKH00] M. M. van Katwijk and D. C. R. Hermus. Effects of water dynamics on Zostera marina: transplantation experiments in the intertidal Dutch Wadden Sea. Marine Ecology Progress Series, 208:107–118, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/208/m208p107.pdf.

Valles:2008:TSP

[VKH08] H. Vallès, D. L. Kramer, and W. Hunte. Temporal and spatial patterns in the recruitment of coral-reef fishes in Barbados. Marine Ecology Progress Series, 363:257–272, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/363/
m363p257.pdf.

Vanderklift:2000:PAE

[VL00] M. A. Vanderklift and P. S. Lavery. Patchiness in assemblages of epiphytic macroalgae on Posidonia coriacea at a hierarchy of spatial scales. Marine Ecology Progress Series, 192:127–135, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p127.pdf.

Vlietstra:2005:SAB

[Vli05] L. S. Vlietstra. Spatial associations between seabirds and prey: effects of large-scale prey abundance on small-scale seabird distribution. Marine Ecology Progress Series, 291: 275–287, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/291/m291p275.pdf.

VanElven:2004:RCD

[VLK04] B. R. Van Elven, P. S. Lavery, and G. A. Kendrick. Reefs as contributors to diversity of epiphytic macroalgae assemblages in seagrass meadows. Marine Ecology Progress Series, 276:71–83, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/276/m276p071.pdf.

VanWambeke:2004:DMB

[VLP^{+}04] F. Van Wambeke, D. Lefèvre, L. Prieur, R. Sempéré, M. Bianchi, K. Oubelkheir, and F. Bruyant. Distribution of microbial biomass, production,respiration, dissolved organic carbon and factors controlling bacterial production across a geostrophic front (Almeria-Oran, SW Mediterranean Sea). Marine Ecology Progress Series, 269:1–15, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p001.pdf.

Vasas:2007:EOT

[VLRJ07] V. Vasas, C. Lancelot, V. Rousseau, and F. Jordán. Eutrophication and overfishing in temperate nearshore pelagic food webs: a network perspective. Marine Ecology Progress Series, 336:1–14, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

vanLeeuwe:2005:TSD

[vLvSGS05] M. A. van Leeuwe, B. van Sikkelerus, W. W. C. Gieskes, and J. Stefels. Taxon-specific differences in photoacclimation to fluctuating irradiance in an Antarctic diatom and a green flagellate. Marine Ecology Progress Series, 288: 9–19, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/288/m288p009.pdf.

Vanderklift:2009:IHK

[VLW09] M. A. Vanderklift, P. S. Lavery, and K. I. Waddington. Intensity of herbivory on kelp by fish and sea urchins differs between inshore and offshore reefs. Marine Ecology Progress Series, 376:203–211, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/376/m376p203.pdf.

Vainikka:2009:SVG

[VMCG09] A. Vainikka, F. Mollet, M. Casini, and A. Gårdmark. Spatial variation in growth, condition and maturation reaction norms of the Baltic herring Clupea harengus membras. Marine Ecology Progress Series, 383:285–294, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/383/
m383p285.pdf.

VanGaever:2006:MAH

[VMdBV06] S. Van Gaever, L. Moodley, D. de Beer, and A. Vanreusel. Meiobenthos at the Arctic Håkon Mosby Mud Volcano, with a parental-caring nematode thriving in sulphide-rich sediments. Marine Ecology Progress Series, 321:143–155, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/321/
m321p143.pdf.

Villafane:2004:PVI

[VMH04] V. E. Villafañe, M. A. Marcoval, and E. W. Helbling. Photosynthesis versus irradiance characteristics in phytoplankton assemblages off Patagonia (Argentina): temporal variability and solar UVR effects. Marine Ecology Progress Series, 284:23–34, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p023.pdf.

Villeger:2008:STS

[VMH^{+}08] S. Villéger, J. Ramos Miranda, D. Flores Hernandez, A. Sosa Lopez, and D. Mouillot. Stable trophic structure across coastal nekton assemblages despite high species turnover. Marine Ecology Progress Series, 364:135–146, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/364/
m364p135.pdf.

Velando:2005:STI

[VML05] A. Velando, I. Munilla, and P. M. Leyenda. Short-term indirect effects of the ‘Prestige’ oil spill on European shags: changes in availability of prey. Marine Ecology Progress Series, 302:263–274, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/302/m302p263.pdf.

Vita:2004:EWF

[VMM^{+}04] R. Vita, A. Marín, J. A. Madrid, B. Jiménez-Brinquis, A. Cesar, and L. Marín-Guirao. Effects of wild fishes on waste exportation from a Mediterranean fish farm. Marine Ecology Progress Series, 277:253–261, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/277/m277p253.pdf.

Vasseur:2003:EBO

[VMN^{+}03] C. Vasseur, B. Mostajir, C. Nozais, M. Denis, E. Fouilland, B. Klein, and S. Demers. Effects of bio-optical factors on the attenuation of ultraviolet and photosynthetically available radiation in the North Water Polynya, northern Baffin Bay: ecological implications. Marine Ecology Progress Series, 252:1–13, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/252/m252p001.pdf.

VanDenBroeck:2004:SVP

[VMRL04] N. Van Den Broeck, T. Moutin, M. Rodier, and A. Le Bouteiller. Seasonal variations of phosphate availability in the SW Pacific Ocean near New Caledonia. Marine Ecology Progress Series, 268:1–12, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/268/m268p001.pdf.

vanMontfrans:2003:SSB

[vMRO03] J. van Montfrans, C. H. Ryer, and R. J. Orth. Substrate selection by blue crab Callinectes sapidus megalopae and first juvenile instars. Marine Ecology Progress Series, 260:209– 217, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/260/m260p209.pdf.

VanColen:2008:MRH

[VMV^{+}08] C. Van Colen, F. Montserrat, M. Vincx, P. M. J. Herman, T. Ysebaert, and S. Degraer. Macrobenthic recovery from hypoxia in an estuarine tidal mudflat. Marine Ecology Progress Series, 372:31–42, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m372p031.pdf.

VanColen:2009:TFN

[VMV^{+}09] C. Van Colen, F. Montserrat, K. Verbist, M. Vincx, M. Steyaert, J. Vanaverbeke, P. M. J. Herman, S. Degraer, and T. Ysebaert. Tidal flat nematode responses to hypoxia and subsequent macrofauna-mediated alterations of sediment properties. Marine Ecology Progress Series, 381:189–197, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m381p189.pdf.

Viaud-Martinez:2007:MGD

[VMVG^{+}07] K. A. Viaud-Martínez, M. Martínez Vergara, P. E. Gol’din, V. Ridoux, A. A. Öztürk, B. Öztürk, P. E. Rosel, A. Frantzis, A. Komnenou, and A. J. Bohonak. Morphological and genetic differentiation of the Black Sea harbour porpoise Phocoena phocoena. Marine Ecology Progress Series, 338:281– 294, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/338/m338p281.pdf.

Velasco:2002:FPI

[VN02] L. A. Velasco and J. M. Navarro. Feeding physiology of infaunal (Mulinia edulis) and epifaunal (Mytilus chilensis) bivalves under a wide range of concentrations and qualities of seston. Marine Ecology Progress Series, 240:143–155, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/240/
m240p143.pdf.

Velasco:2005:FPT

[VN05] L. A. Velasco and J. M. Navarro. Feeding physiology of two bivalves under laboratory and field conditions in response to variable food concentrations. Marine Ecology Progress Series, 291:115–124, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/291/m291p115.pdf.

Verweij:2006:SFS

[VNdG^{+}06] M. C. Verweij, I. Nagelkerken, D. de Graaff, M. Peeters, E. J. Bakker, and G. van der Velde. Structure, food and shade attract juvenile coral reef fish to mangrove and seagrass habitats: a field experiment. Marine Ecology Progress Series, 306:257–268, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p257.pdf.

Vandieken:2006:CMA

[VNJ06] V. Vandieken, M. Nickel, and B. B. Jørgensen. Carbon mineralization in Arctic sediments northeast of Svalbard: Mn(IV) and Fe(III) reduction as principal anaerobic respiratory pathways. Marine Ecology Progress Series, 322:15– 27, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/322/m322p015.pdf.

vanOevelen:2009:OMI

[vOSF^{+}09] D. van Oevelen, K. Soetaert, M. A. Franco, L. Moodley, L. van IJzerloo, M. Vincx, and J. Vanaverbeke. Organic matter input and processing in two contrasting North Sea sediments: insights from stable isotope and biomass data. Marine Ecology Progress Series, 380:19–32, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/380/m380p019.pdf.

VanAlstyne:2000:ENE

[VP00] K. L. Van Alstyne and K. N. Pelletreau. Effects of nutrient enrichment on growth and phlorotannin production in Fucus gardneri embryos. Marine Ecology Progress Series, 206:33–43, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/206/m206p033.pdf.

Viscido:2004:IBE

[VPG04] S. V. Viscido, J. K. Parrish, and D. Grünbaum. Individual behavior and emergent properties of fish schools: a comparison of observation and theory. Marine Ecology Progress Series, 273:239–249, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p239.pdf.

Varela:2004:VBP

[VPP04] M. Varela, R. Prego, and Y. Pazos. Vertical biogenic particle flux in a western Galician ria (NW Iberian Peninsula). Marine Ecology Progress Series, 269:17–32, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/269/
m269p017.pdf.

Visser:2002:DVD

[VPS^{+}02] P. M. Visser, J. J. Poos, B. B. Scheper, P. Boelen, and F. C. van Duyl. Diurnal variations in depth profiles of UVinduced DNA damage and inhibition of bacterioplankton production in tropical coastal waters. Marine Ecology Progress Series, 228:25–33, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p025.pdf.

Vonk:2008:LSQ

[VPS08] J. A. Vonk, M. H. J. Pijnappels, and J. Stapel. In situ quantification of Tripneustes gratilla grazing and its effects on three co-occurring tropical seagrass species. Marine Ecology Progress Series, 360:107–114, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/360/m360p107.pdf.

Vopel:2001:CGA

[VPSO01] K. Vopel, M. Pöhn, A. Sorgo, and J. Ott. Ciliate-generated advective seawater transport supplies chemoautotrophic ectosymbionts. Marine Ecology Progress Series, 210:93–99, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/210/m210p093.pdf.

Vopel:2009:OSS

[VPWE09] K. Vopel, C. A. Pilditch, P. Wilson, and M. J. Ellwood. Oxidation of surface sediment: effects of disturbance depth and seawater flow speed. Marine Ecology Progress Series, 392: 43–55, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/392/m392p043.pdf.

Valentine:2002:CNC

[VRJ02] J. W. Valentine, K. Roy, and D. Jablonski. Carnivore/noncarnivore ratios in northeastern Pacific marine gastropods. Marine Ecology Progress Series, 228:153–163, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/228/
m228p153.pdf.

Vasconcelos:2007:DEN

[VRST^{+}07] R. P. Vasconcelos, P. Reis-Santos, S. Tanner, V. Fonseca, C. Latkoczy, D. Günther, M. J. Costa, and H. Cabral. Discriminating estuarine nurseries for five fish species through otolith elemental fingerprints. Marine Ecology Progress Series, 350:117–126, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/350/m350p117.pdf.

Vonk:2008:RNS

[VS08] J. A. Vonk and J. Stapel. Regeneration of nitrogen (^{1}5N) from seagrass litter in tropical Indo–Pacific meadows. Marine Ecology Progress Series, 368:165–175, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/368/
m368p165.pdf.

Vergani:2004:ENN

[VSB04] D. F. Vergani, Z. B. Stanganelli, and D. Bilenca. Effects of El Niño and La Niña events on the sex ratio of southern elephant seals at King George Island. Marine Ecology Progress Series, 268:293–300, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2004/268/m268p293.pdf.

Vermeij:2003:VPR

[VSBB03] M. J. A. Vermeij, E. Sampayo, K. Bröker, and R. P. M. Bak. Variation in planulae release of closely related coral species. Marine Ecology Progress Series, 247:75–84, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/247/
m247p075.pdf.

Vadstein:2004:MIB

[VSL^{+}04] O. Vadstein, H. Stibor, B. Lippert, K. Løseth, W. Roederer, L. Sundt-Hansen, and Y. Olsen. Moderate increase in the biomass of omnivorous copepods may ease grazing control of planktonic algae. Marine Ecology Progress Series, 270:199–207, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/270/m270p199.pdf.

Velo-Suarez:2009:VDD

[VSRGW09] L. Velo-Suárez, B. Reguera, E. Garcés, and T. Wyatt. Vertical distribution of division rates in coastal dinoflagellate Dinophysis spp. populations: implications for modelling. Marine Ecology Progress Series, 385:87–96, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/385/
m385p087.pdf.

Vanaverbeke:2004:CMC

[VSV04] J. Vanaverbeke, K. Soetaert, and M. Vincx. Changes in morphometric characteristics of nematode communities during a spring phytoplankton bloom deposition. Marine Ecology Progress Series, 273:139–146, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/273/m273p139.pdf.

Vanaverbeke:2003:NBS

[VSVV03] J. Vanaverbeke, M. Steyaert, A. Vanreusel, and M. Vincx. Nematode biomass spectra as descriptors of functional changes due to human and natural impact. Marine Ecology Progress Series, 249:157–170, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/249/m249p157.pdf.

VanDover:2000:DDS

[VT00] C. L. Van Dover and J. L. Trask. Diversity at deep-sea hydrothermal vent and intertidal mussel beds. Marine Ecology Progress Series, 195:169–178, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/195/m195p169.pdf.

VanWambeke:2009:DVH

[VTDS09] F. Van Wambeke, M. Tedetti, S. Duhamel, and R. Sempéré. Diel variability of heterotrophic bacterial production and underwater UV doses in the eastern South Pacific. Marine Ecology Progress Series, 387:97–108, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/387/
m387p097.pdf.

Vargas:2002:ICV

[VTS^{+}02] C. A. Vargas, K. Tönnesson, A. Sell, M. Maar, E. F. Møller, T. Zervoudaki, A. Giannakourou, E. Christou, S. Satapoomin, J. K. Petersen, T. G. Nielsen, and P. Tiselius. Importance of copepods versus appendicularians in vertical carbon fluxes in a Swedish fjord. Marine Ecology Progress Series, 241:125–138, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/241/m241p125.pdf.

vanTussenbroek:2008:SAP

[vTWMG08] Brigitta I. van Tussenbroek, J. G. Ricardo Wong, and Judith Márquez-Guzman. Synchronized anthesis and predation on pollen in the marine angiosperm Thalassia testudinum (Hydrocharitaceae). Marine Ecology Progress Series, 354:119– 124, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/354/m354p119.pdf.

Villalobos:2006:TPT

[VTY06] F. Benitez Villalobos, P. A. Tyler, and C. M. Young. Temperature and pressure tolerance of embryos and larvae of the Atlantic seastars Asterias rubens and Marthasterias glacialis (Echinodermata: Asteroidea): potential for deep-sea invasion. Marine Ecology Progress Series, 314:109–117, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/314/
m314p109.pdf.

Viitasalo:2004:PMS

[VV04] S. Viitasalo and M. Viitasalo. Predation by the mysid shrimps Mysis mixta and M. relicta on benthic eggs of Bosmina longispina maritima (Cladocera) in the northern Baltic Sea. Marine Ecology Progress Series, 281:155–163, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/281/
m281p155.pdf.

Vopel:2007:ESH

[VVTH07] K. Vopel, A. Vopel, D. Thistle, and N. Hancock. Effects of spatangoid heart urchins on O_{2} supply into coastal sediment. Marine Ecology Progress Series, 333:161–171, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/333/
m333p161.pdf.

Victor:2000:EPL

[VW00] B. C. Victor and G. M. Wellington. Endemism and the pelagic larval duration of reef fishes in the eastern Pacific Ocean. Marine Ecology Progress Series, 205:241–248, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
205/m205p241.pdf.

vanWesenbeeck:2007:DHT

[vWCAB07] B. K. van Wesenbeeck, C. M. Crain, A. H. Altieri, and M. D. Bertness. Distinct habitat types arise along a continuous hydrodynamic stress gradient due to interplay of competition and facilitation. Marine Ecology Progress Series, 349: 63–71, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/349/m349p063.pdf.

VanAlstyne:2001:ADS

[VWF^{+}01] K. L. Van Alstyne, G. V. Wolfe, T. L. Freidenburg, A. Neill, and C. Hicken. Activated defense systems in marine macroalgae: evidence for an ecological role for DMSP cleavage. Marine Ecology Progress Series, 213:53–65, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/213/m213p053.pdf.

Verity:2000:FDM

[VWH00] P. G. Verity, S. C. Williams, and Y. Hong. Formation, degradation, and mass: volume ratios of detritus derived from decaying phytoplankton. Marine Ecology Progress Series, 207:53–68, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/207/m207p053.pdf.

vanWoesik:2005:CEG

[vWNYS05] R. van Woesik, T. Nakamura, H. Yamasaki, and C. Sheppard. Comment on ‘Effects of geography, taxa, water flow, and temperature variation on coral bleaching intensity in Mauritius’ (McClanahan et al. 2005). Marine Ecology Progress Series, 305:297–299, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/305/m305p297.pdf. See

[MMMPB05].

Villanueva:2006:IFF

[VYM06] R. D. Villanueva, H. T. Yap, and M. N. E. Montaño. Intensive fish farming in the Philippines is detrimental to the reef-building coral Pocillopora damicornis. Marine Ecology Progress Series, 316:165–174, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/316/m316p165.pdf.

Villanueva:2008:TPP

[VYM08] R. D. Villanueva, H. T. Yap, and M. N. E. Montaño. Timing of planulation by pocilloporid corals in the northwestern Philippines. Marine Ecology Progress Series, 370:111– 119, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/370/m370p111.pdf.

Waddy:2000:PTP

[WA00] S. L. Waddy and D. E. Aiken. Problems with the terminology proposed by Lavalli & Lawton for certain life history phases of the American lobster Homarus americanus. Marine Ecology Progress Series, 197:309–310, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/197/m197p309.pdf.

Willis:2003:SCR

[WA03] T. J. Willis and M. J. Anderson. Structure of cryptic reef fish assemblages: relationships with habitat characteristics and predator density. Marine Ecology Progress Series, 257: 209–221, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/257/m257p209.pdf.

Werner:2005:SVA

[WA05] I. Werner and H. Auel. Seasonal variability in abundance, respiration and lipid composition of Arctic under-ice amphipods. Marine Ecology Progress Series, 292:251–262, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/292/
m292p251.pdf.

Webb:2009:MES

[WAA^{+}09] T. J. Webb, I. F. Aleffi, J. M. Amouroux, G. Bachelet, S. Degraer, C. Dounas, D. Fleischer, A. Grémare, M. Herrmann, H. Hummel, I. Karakassis, M. Kedra, M. A. Kendall, L. Kotwicki, C. Labrune, E. L. Nevrova, A. Occhipinti-Ambrogi, A. Petrov, N. K. Revkov, R. Sardá, N. Simboura, J. Speybroeck, G. Van Hoey, M. Vincx, P. Whomersley, W. Willems, and M. Włodarska-Kowalczuk. Macroecology of the European soft sediment benthos: insights from the MacroBen database. Marine Ecology Progress Series, 382: 287–296, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/theme/m382p287.pdf.

Warwick:2002:IAC

[WAB^{+}02] R. M. Warwick, C. M. Ashman, A. R. Brown, K. R. Clarke, B. Dowell, B. Hart, R. e. Lewis, N. Shillabeer, P. J. Somerfield, and J. F. Tapp. Inter-annual changes in the biodiversity and community structure of the macrobenthos in Tees Bay and the Tees estuary, UK, associated with local and regional environmental events. Marine Ecology Progress Series, 234:1–13, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/234/m234p001.pdf.

Widdicombe:2004:IBB

[WAK^{+}04] Stephen Widdicombe, Melanie C. Austen, Michael A. Kendall, Frode Olsgard, Morten T. Schaanning, Sarah L. Dashfield, and Hazel R. Needham. Importance of bioturbators for biodiversity maintenance: indirect effects of fishing disturbance. Marine Ecology Progress Series, 275:1–10, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p001.pdf.

Watson:2009:EPF

[WAK^{+}09] D. L. Watson, M. J. Anderson, G. A. Kendrick, K. Nardi, and E. S. Harvey. Effects of protection from fishing on the lengths of targeted and non-targeted fish species at the Houtman Abrolhos Islands, Western Australia. Marine Ecology Progress Series, 384:241–249, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/384/m384p241.pdf.

Walker:2007:EFS

[Wal07] J. W. Walker. Effects of fine sediments on settlement and survival of the sea urchin Evechinus chloroticus in northeastern New Zealand. Marine Ecology Progress Series, 331: 109–118, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/331/m331p109.pdf.

Wada:2008:BMD

[WAM^{+}08] S. Wada, M. N. Aoki, A. Mikami, T. Komatsu, Y. Tsuchiya, T. Sato, H. Shinagawa, and T. Hama. Bioavailability of macroalgal dissolved organic matter in seawater. Marine Ecology Progress Series, 370:33–44, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/370/
m370p033.pdf.

Wildish:2008:PIB

[WAMP08] D. J. Wildish, H. M. Akagi, D. L. McKeown, and G. W. Pohle. Pockmarks influence benthic communities in Passamaquoddy Bay, Bay of Fundy, Canada. Marine Ecology Progress Series, 357:51–66, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/357/m357p051.pdf.

Wilson:2001:APP

[WAN^{+}01] P. R. Wilson, D. G. Ainley, N. Nur, S. S. Jacobs, K. J. Barton, G. Ballard, and J. C. Comiso. Adélie penguin population change in the Pacific sector of Antarctica: relation to sea-ice extent and the Antarctic Circumpolar Current. Marine Ecology Progress Series, 213:301–309, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/213/m213p301.pdf.

Wang:2002:CSA

[Wan02a] W. X. Wang. Cd and Se aqueous uptake and exposure of green mussels Perna viridis: influences of seston quantity. Marine Ecology Progress Series, 226:211–221, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/226/
m226p211.pdf.

Wang:2002:ITM

[Wan02b] W. X. Wang. Interactions of trace metals and different marine food chains. Marine Ecology Progress Series, 243:295– 309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/243/m243p295.pdf.

Ward:2005:TVN

[War05] B. B. Ward. Temporal variability in nitrification rates and related biogeochemical factors in Monterey Bay, California, USA. Marine Ecology Progress Series, 292:97–109, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/292/m292p097.pdf.

Witte:2003:RRD

[WASW03] U. Witte, N. Aberle, M. Sand, and F. Wenzhöfer. Rapid response of a deep-sea benthic community to POM enrichment: an in situ experimental study. Marine Ecology Progress Series, 251:27–36, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p027.pdf.

Whitehouse:2009:RKV

[WAW^{+}09] M. J. Whitehouse, A. Atkinson, P. Ward, R. E. Korb, P. Rothery, and S. Fielding. Role of krill versus bottom-up factors in controlling phytoplankton biomass in the northern Antarctic waters of South Georgia. Marine Ecology Progress Series, 393:69–82, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/393/m393p069.pdf.

Winemiller:2007:PSF

[WAZ07] K. O. Winemiller, S. Akin, and S. C. Zeug. Production sources and food web structure of a temperate tidal estuary: integration of dietary and stable isotope data. Marine Ecology Progress Series, 343:63–76, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/343/m343p063.pdf.

Ward:2001:NNU

[WB01] B. B. Ward and D. A. Bronk. Net nitrogen uptake and DON release in surface waters: importance of trophic interactions implied from size fractionation experiments. Marine Ecology Progress Series, 219:11–24, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/219/m219p011.pdf.

Weissburg:2005:SBL

[WB05a] M. J. Weissburg and Howard I. Browman. Sensory biology: linking the internal and external ecologies of marine organisms. Marine Ecology Progress Series, ??:??, 2005. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m287p263.pdf.

Weissburg:2005:TSS

[WB05b] Marc J. Weissburg and Howard I. Browman. Theme section: Sensory biology: linking the internal and external ecologies of marine organisms. Marine Ecology Progress Series, 287:263–307, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m287p263.pdf.

Waggett:2006:CSF

[WB06] R. J. Waggett and E. J. Buskey. Copepod sensitivity to flow fields: detection by copepods of predatory ctenophores. Marine Ecology Progress Series, 323:205–211, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/323/
m323p205.pdf.

Waggett:2007:CEB

[WB07] R. J. Waggett and E. J. Buskey. Copepod escape behavior in non-turbulent and turbulent hydrodynamic regimes. Marine Ecology Progress Series, 334:193–198, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p193.pdf.

Wilson:2001:SDL

[WBC01] S. K. Wilson, K. Burns, and S. Codi. Sources of dietary lipids in the coral reef blenny Salarias patzneri. Marine Ecology Progress Series, 222:291–296, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/222/m222p291.pdf.

Wellenreuther:2007:EDH

[WBC07] M. Wellenreuther, P. T. Barrett, and K. D. Clements. Ecological diversification in habitat use by subtidal triplefin fishes (Tripterygiidae). Marine Ecology Progress Series, 330:235– 246, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/330/m330p235.pdf.

Williams:2008:CHI

[WBC08] D. A. Williams, S. D. Brown, and D. L. Crawford. Contemporary and historical influences on the genetic structure of the estuarine-dependent Gulf killifish Fundulus grandis. Marine Ecology Progress Series, 373:111–121, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/373/
m373p111.pdf.

Wong:2006:BMU

[WBDR06] M. C. Wong, M. A. Barbeau, M. Dowd, and K. R. Richard. Behavioural mechanisms underlying functional response of sea stars Asterias vulgaris preying on juvenile sea scallops Placopecten magellanicus. Marine Ecology Progress Series, 317: 75–86, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/317/m317p075.pdf.

Webb:2009:BEP

[WBG09] K. E. Webb, D. K. A. Barnes, and J. S. Gray. Benthic ecology of pockmarks in the Inner Oslofjord, Norway. Marine Ecology Progress Series, 387:15–25, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/387/m387p015.pdf.

Widdows:2004:RPB

[WBH^{+}04] J. Widdows, A. Blauw, C. H. R. Heip, P. M. J. Herman, C. H. Lucas, J. J. Middelburg, S. Schmidt, M. D. Brinsley, F. Twisk, and H. Verbeek. Role of physical and biological processes in sediment dynamics of a tidal flat in Westerschelde Estuary, SW Netherlands. Marine Ecology Progress Series, 274:41–56, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/274/m274p041.pdf.

Witt:2007:PLH

[WBJ^{+}07] M. J. Witt, A. C. Broderick, D. J. Johns, C. Martin, R. Penrose, M. S. Hoogmoed, and B. J. Godley. Prey landscapes help identify potential foraging habitats for leatherback turtles in the NE Atlantic. Marine Ecology Progress Series, 337:231–243, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m337p231.pdf.

Watanabe:2006:SMC

[WBL^{+}06] Y. Watanabe, H. Bornemann, N. Liebsch, J. Plötz, K. Sato, Y. Naito, and N. Miyazaki. Seal-mounted cameras detect invertebrate fauna on the underside of an Antarctic ice shelf. Marine Ecology Progress Series, 309:297–300, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p297.pdf.

Welsh:2000:DNF

[WBN^{+}00] D. T. Welsh, M. Bartoli, D. Nizzoli, G. Castaldelli, S. A. Riou, and P. Viaroli. Denitrification, nitrogen fixation, community primary productivity and inorganic-n and oxygen fluxes in an intertidal Zostera noltii meadow. Marine Ecology Progress Series, 208:65–77, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/208/m208p065.pdf.

Widdows:2006:CBS

[WBP^{+}06] J. Widdows, M. D. Brinsley, N. D. Pope, F. J. Staff, S. G. Bolam, and P. J. Somerfield. Changes in biota and sediment erodability following the placement of fine dredged material on upper intertidal shores of estuaries. Marine Ecology Progress Series, 319:27–41, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/319/m319p027.pdf.

Widdows:2009:END

[WBP09] J. Widdows, M. D. Brinsley, and N. D. Pope. Effect of Nereis diversicolor density on the erodability of estuarine sediment. Marine Ecology Progress Series, 378:135–143, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/378/
m378p135.pdf.

Whaley:2007:UEL

[WBR07] S. D. Whaley, J. J. Burd, Jr., and B. A. Robertson. Using estuarine landscape structure to model distribution patterns in nekton communities and in juveniles of fishery species. Marine Ecology Progress Series, 330:83–99, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m330p083.pdf.

Watanabe:2004:FTB

[WBS^{+}04] Y. Watanabe, E. A. Baranov, K. Sato, Y. Naito, and N. Miyazaki. Foraging tactics of Baikal seals differ between day and night. Marine Ecology Progress Series, 279:283– 289, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/279/m279p283.pdf.

Wall:2007:CSD

[WBS^{+}07] S. M. Wall, C. J. A. Bradshaw, C. J. Southwell, N. J. Gales, and M. A. Hindell. Crabeater seal diving behaviour in eastern Antarctica. Marine Ecology Progress Series, 337:265– 277, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/337/m337p265.pdf.

Widdows:2000:IBS

[WBSL00] J. Widdows, M. D. Brinsley, P. N. Salkeld, and C. H. Lucas. Influence of biota on spatial and temporal variation in sediment erodability and material flux on a tidal flat (Westerschelde, The Netherlands). Marine Ecology Progress Series, 194:23–37, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/194/m194p023.pdf.

Watson:2008:DNG

[WBT^{+}08] A. J. Watson, P. W. Boyd, S. M. Turner, T. D. Jickells, and P. S. Liss. Designing the next generation of ocean iron fertilization experiments. Marine Ecology Progress Series, 364:303–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m364p303.pdf.

Werner:2006:SPA

[WBW^{+}06] U. Werner, P. Bird, C. Wild, T. Ferdelman, L. Polerecky, G. Eickert, R. Jonstone, O. Hoegh-Guldberg, and D. de Beer. Spatial patterns of aerobic and anaerobic mineralization rates and oxygen penetration dynamics in coral reef sediments. Marine Ecology Progress Series, 309:93–105, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p093.pdf.

Winters:2009:STP

[WBZ^{+}09] G. Winters, S. Beer, B. B. Zvi, I. Brickner, and Y. Loya. Spatial and temporal photoacclimation of Stylophora pistillata: zooxanthella size, pigmentation, location and clade. Marine Ecology Progress Series, 384:107–119, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/384/
m384p107.pdf.

Wong:2003:SRD

[WC03] W. H. Wong and S. G. Cheung. Site-related differences in the feeding physiology of the green mussel Perna viridis: a reciprocal transplantation experiment. Marine Ecology Progress Series, 258:147–159, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/258/m258p147.pdf.

Wagner:2001:NAG

[WCBD01] M. M. Wagner, R. G. Campbell, C. A. Boudreau, and E. G. Durbin. Nucleic acids and growth of Calanus finmarchicus in the laboratory under different food and temperature conditions. Marine Ecology Progress Series, 221:185–197, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p185.pdf.

Weston:2009:CFF

[WCCA09] J. Weston, S. P. Colin, J. H. Costello, and E. Abbott. Changing form and function during development in rowing hydromedusae. Marine Ecology Progress Series, 374:127– 134, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/374/m374p127.pdf.

Wells:2008:FER

[WCF08] R. J. D. Wells, J. H. Cowan, Jr., and B. Fry. Feeding ecology of red snapper Lutjanus campechanus in the northern Gulf of Mexico. Marine Ecology Progress Series, 361:213– 225, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/361/m361p213.pdf.

Woodby:2009:PMC

[WCH09] D. Woodby, D. Carlile, and L. Hulbert. Predictive modeling of coral distribution in the Central Aleutian Islands, USA. Marine Ecology Progress Series, 397:227–240, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m397p227.pdf.

Whitehead:2008:MSW

[WCJL08] H. Whitehead, A. Coakes, N. Jaquet, and S. Lusseau. Movements of sperm whales in the tropical Pacific. Marine Ecology Progress Series, 361:291–300, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/361/m361p291.pdf.

Wiedenmann:2008:TDG

[WCM08] J. Wiedenmann, K. Cresswell, and M. Mangel. Temperaturedependent growth of Antarctic krill: predictions for a changing climate from a cohort model. Marine Ecology Progress Series, 358:191–202, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/358/m358p191.pdf.

Wilber:2003:RSZ

[WCRB03] D. H. Wilber, D. G. Clarke, G. L. Ray, and M. Burlas. Response of surf zone fish to beach nourishment operations on the northern coast of New Jersey, USA. Marine Ecology Progress Series, 250:231–246, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/250/m250p231.pdf.

Wisshak:2009:SRA

[WCZ^{+}09] M. Wisshak, M. López Correa, H. Zibrowius, J. Jakobsen, and A. Freiwald. Skeletal reorganisation affects geochemical signals, exemplified in the stylasterid hydrocoral Errina dabneyi (Azores Archipelago). Marine Ecology Progress Series, 397:197–208, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m397p197.pdf.

Winkler:2003:TCA

[WDB^{+}03] G. Winkler, J. J. Dodson, N. Bertrand, D. Thivierge, and W. F. Vincent. Trophic coupling across the St. Lawrence River estuarine transition zone. Marine Ecology Progress Series, 251:59–73, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p059.pdf.

Wooldridge:2005:PRC

[WDB^{+}05] S. Wooldridge, T. Done, R. Berkelmans, R. Jones, and P. Marshall. Precursors for resilience in coral communities in a warming climate: a belief network approach. Marine Ecology Progress Series, 295:157–169, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/295/m295p157.pdf.

Wolfstein:2002:BPP

[WdBS02] K. Wolfstein, J. F. C. de Brouwer, and L. J. Stal. Biochemical partitioning of photosynthetically fixed carbon by benthic diatoms during short-term incubations at different irradiances. Marine Ecology Progress Series, 245:21–31, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/245/
m245p021.pdf.

Williams:2005:PRH

[WDLM05] G. A. Williams, M. De Pirro, K. M. Y. Leung, and D. Morritt. Physiological responses to heat stress on a tropical shore: the benefits of mushrooming behaviour in the limpet Cellana grata. Marine Ecology Progress Series, 292:213– 224, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/292/m292p213.pdf.

Widdicombe:2009:ECI

[WDM^{+}09] S. Widdicombe, S. L. Dashfield, C. L. McNeill, H. R. Needham, A. Beesley, A. McEvoy, S. Øxnevad, K. R. Clarke, and J. A. Berge. Effects of CO_{2} induced seawater acidification on infaunal diversity and sediment nutrient fluxes. Marine Ecology Progress Series, 379:59–75, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p059.pdf.

Williams:2000:PSS

[WDN00] G. A. Williams, M. S. Davies, and S. Nagarkar. Primary succession on a seasonal tropical rocky shore: the relative roles of spatial heterogeneity and herbivory. Marine Ecology Progress Series, 203:81–94, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/203/m203p081.pdf. See erra-

tum [Ano01f].

Wright:2000:GVH

[WdNS00] J. T. Wright, R. de Nys, and P. D. Steinberg. Geographic variation in halogenated furanones from the red alga Delisea pulchra and associated herbivores and epiphytes. Marine Ecology Progress Series, 207:227–241, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/207/m207p227.pdf.

Wright:2005:DDS

[WDR^{+}05] J. T. Wright, S. A. Dworjanyn, C. N. Rogers, P. D. Steinberg, J. E. Williamson, and A. G. B. Poore. Density-dependent sea urchin grazing: differential removal of species, changes in community composition and alternative community states. Marine Ecology Progress Series, 298:143–156, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p143.pdf.

Wisehart:2007:IEE

[WDRH07] L. M. Wisehart, B. R. Dumbauld, J. L. Ruesink, and S. D. Hacker. Importance of eelgrass early life history stages in response to oyster aquaculture disturbance. Marine Ecology Progress Series, 344:71–80, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/344/m344p071.pdf.

Wiklund:2009:EWS

[WDSA09] A. K. Eriksson Wiklund, K. Dahlgren, B. Sundelin, and A. Andersson. Effects of warming and shifts of pelagic food web structure on benthic productivity in a coastal marine system. Marine Ecology Progress Series, 396:13–25, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/396/m396p013.pdf.

Watanabe:2008:DML

[WDT^{+}08a] Y. Watanabe, G. E. Dingsør, Y. Tian, I. Tanaka, and N. C. Stenseth. Determinants of mean length at age of spring spawning herring off the coast of Hokkaido, Japan. Marine Ecology Progress Series, 366:209–217, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2008/366/m366p209.pdf.

Watanuki:2008:MUP

[WDT^{+}08b] Y. Watanuki, F. Daunt, A. Takahashi, M. Newell, S. Wanless, K. Sato, and N. Miyazaki. Microhabitat use and prey capture of a bottom-feeding top predator, the European shag, shown by camera loggers. Marine Ecology Progress Series, 356:283–293, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/356/m356p283.pdf.

Wang:2001:CUT

[WDX01] W. X. Wang, R. C. H. Dei, and Y. Xu. Cadmium uptake and trophic transfer in coastal plankton under contrasting nitrogen regimes. Marine Ecology Progress Series, 211:293–298, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
211/m211p293.pdf.

Webb:2004:ENP

[WE04] A. P. Webb and B. D. Eyre. Effect of natural populations of burrowing thalassinidean shrimp on sediment irrigation, benthic metabolism, nutrient fluxes and denitrification. Marine Ecology Progress Series, 268:205–220, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p205.pdf.

Whalan:2008:LVM

[WEEBdN08] S. Whalan, P. Ettinger-Epstein, C. Battershill, and R. de Nys. Larval vertical migration and hierarchical selectivity of settlement in a brooding marine sponge. Marine Ecology Progress Series, 368:145–154, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/368/m368p145.pdf.

Wennhage:2000:VSP

[Wen00] H. Wennhage. Vulnerability of settling plaice Pleuronectes platessa to predation: effects of developmental stage and alternative prey. Marine Ecology Progress Series, 203:289– 299, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/203/m203p289.pdf.

Westin:2003:MFS

[Wes03] L. Westin. Migration failure in stocked eels Anguilla anguilla. Marine Ecology Progress Series, 254:307–311, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/254/
m254p307.pdf.

Wells:2003:IPO

[WFC03] B. K. Wells, K. D. Friedland, and L. M. Clarke. Increment patterns in otoliths and scales from mature Atlantic salmon Salmo salar. Marine Ecology Progress Series, 262: 293–298, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/262/m262p293.pdf.

Weng:2008:MUT

[WFG^{+}08] K. C. Weng, D. G. Foley, J. E. Ganong, C. Perle, G. L. Shillinger, and B. A. Block. Migration of an upper trophic level predator, the salmon shark Lamna ditropis, between distant ecoregions. Marine Ecology Progress Series, 372:253– 264, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p253.pdf.

Warwick:2008:MJS

[WFHS08] Richard Warwick, John Field, Carlo Heip, and Tony Stebbing. In memoriam: John Stuart Gray (1941–2007). Marine Ecology Progress Series, 354:1, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2008/354/m354p001.pdf.

Wells:2008:URA

[WFT^{+}08] B. K. Wells, J. C. Field, J. A. Thayer, C. B. Grimes, S. J. Bograd, W. J. Sydeman, F. B. Schwing, and R. Hewitt. Untangling the relationships among climate, prey and top predators in an ocean ecosystem. Marine Ecology Progress Series, 364:15–29, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/364/m364p015.pdf.

Wang:2000:BCB

[WG00] W. X. Wang and L. Guo. Bioavailability of colloid-bound Cd, Cr, and Zn to marine plankton. Marine Ecology Progress Series, 202:41–49, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/202/m202p041.pdf.

Wear:2001:BET

[WG01] R. G. Wear and J. P. A. Gardner. Biological effects of the toxic algal bloom of February and March 1998 on the benthos of Wellington Harbour, New Zealand. Marine Ecology Progress Series, 218:63–76, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/218/m218p063.pdf.

Williams:2002:PRC

[WG02a] S. L. Williams and E. D. Grosholz. Preliminary reports from the Caulerpa taxifolia invasion in southern California. Marine Ecology Progress Series, 233:307–310, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/233/
m233p307.pdf.

Winkler:2002:LSE

[WG02b] G. Winkler and W. Greve. Laboratory studies of the effect of temperature on growth, moulting and reproduction in the co-occurring mysids Neomysis integer and Praunus flexuosus. Marine Ecology Progress Series, 235:177–188, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/235/
m235p177.pdf.

Wolcott:2002:FIE

[WG02c] B. D. Wolcott and B. Gaylord. Flow-induced energetic bounds to growth in an intertidal sea anemone. Marine Ecology Progress Series, 245:101–109, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p101.pdf.

Wood:2007:SSS

[WG07] A. R. Wood and J. P. A. Gardner. Small spatial scale population genetic structure in two limpet species endemic to the Kermadec Islands, New Zealand. Marine Ecology Progress Series, 349:159–170, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/349/m349p159.pdf.

Wang:2009:ESS

[WG09] Z. Wang and L. Goodman. Evolution of the spatial structure of a thin phytoplankton layer into a turbulent field. Marine Ecology Progress Series, 374:57–74, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/374/
m374p057.pdf.

West:2003:BSO

[WGCM^{+}03] Andrew D. West, John D. Goss-Custard, Selwyn McGrorty, Richard A. Stillman, Sarah E. A. le V.dit Durell, Barry Stewart, Peter Walker, David W. Palmer, and Philip J. Coates. The Burry shellfishery and oystercatchers: using a behaviourbased model to advise on shellfishery management policy. Marine Ecology Progress Series, 248:279–292, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/248/
m248p279.pdf.

Waite:2005:BSS

[WGD05] L. Waite, J. Grant, and J. Davidson. Bay-scale spatial growth variation of mussels Mytilus edulis in suspended culture, Prince Edward Island, Canada. Marine Ecology Progress Series, 297:157–167, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/297/m297p157.pdf. See erra-

tum [Ano05d].

Wahle:2009:DDI

[WGF09] R. A. Wahle, M. Gibson, and M. Fogarty. Distinguishing disease impacts from larval supply effects in a lobster fishery collapse. Marine Ecology Progress Series, 376:185– 192, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/376/m376p185.pdf.

Wilson:2000:COE

[WGH^{+}00] B. Wilson, K. Grellier, P. S. Hammond, G. Brown, and P. M. Thompson. Changing occurrence of epidermal lesions in wild bottlenose dolphins. Marine Ecology Progress Series, 205:283–290, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/205/m205p283.pdf.

Wing:2003:RSS

[WGL03] S. R. Wing, M. T. Gibbs, and M. D. Lamare. Reproductive sources and sinks within a sea urchin, Evechinus chloroticus, population of a New Zealand fjord. Marine Ecology Progress Series, 248:109–123, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/248/m248p109.pdf.

Wilding:2002:GMR

[WGM02] C. S. Wilding, J. Grahame, and P. J. Mill. A GTT microsatellite repeat motif and differentiation between morphological forms of Littorina saxatilis: speciation in progress? Marine Ecology Progress Series, 227:195–204, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/227/
m227p195.pdf.

Wiencke:2000:IUR

[WGP^{+}00] C. Wiencke, I. Gómez, H. Pakker, A. Flores-Moya, M. Altamirano, D. Hanelt, K. Bischof, and F. L. Figueroa. Impact of UV-radiation on viability, photosynthetic characteristics and DNA of brown algal zoospores: implications for depth zonation. Marine Ecology Progress Series, 197:217–229, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
197/m197p217.pdf.

Wetherbee:2007:MPJ

[WGR07] B. M. Wetherbee, S. H. Gruber, and R. S. Rosa. Movement patterns of juvenile lemon sharks Negaprion brevirostris within Atol das Rocas, Brazil: a nursery characterized by tidal extremes. Marine Ecology Progress Series, 343:283– 293, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/343/m343p283.pdf.

Wilson:2002:RSS

[WGS^{+}02] R. P. Wilson, D. Grémillet, J. Syder, M. A. M. Kierspel, S. Garthe, H. Weimerskirch, C. Schäfer-Neth, J. A. Scolaro, C. A. Bost, J. Plötz, and D. Nel. Remote-sensing systems and seabirds: their use, abuse and potential for measuring marine environmental variables. Marine Ecology Progress Series, 228:241–261, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/228/m228p241.pdf.

Whitfield:2002:BII

[WGV^{+}02] P. E. Whitfield, T. Gardner, S. P. Vives, M. R. Gilligan, W. R. Courtenay, Jr., G. C. Ray, and J. A. Hare. Biological invasion of the Indo–Pacific lionfish Pterois volitans along the Atlantic coast of North America. Marine Ecology Progress Series, 235:289–297, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/235/m235p289.pdf.

Webster:2000:MIC

[WH00] M. S. Webster and M. A. Hixon. Mechanisms and individual consequences of intraspecific competition in a coral-reef fish. Marine Ecology Progress Series, 196:187–194, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
196/m196p187.pdf.

Wahl:2002:IBS

[WH02] W. Wahl and K. Hoppe. Interactions between substratum rugosity, colonization density and periwinkle grazing efficiency. Marine Ecology Progress Series, 225:239–249, 2002. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
225/m225p239.pdf.

Wilson:2003:SPS

[WH03] J. R. Wilson and P. L. Harrison. Spawning patterns of scleractinian corals at the Solitary Islands — a high latitude coral community in eastern Australia. Marine Ecology Progress Series, 260:115–123, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/260/m260p115.pdf.

Weise:2008:TVO

[WH08] M. J. Weise and J. T. Harvey. Temporal variability in ocean climate and California sea lion diet and biomass consumption: implications for fisheries management. Marine Ecology Progress Series, 373:157–172, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m373p157.pdf.

Wirsing:2009:OHS

[WH09] A. J. Wirsing and M. R. Heithaus. Olive-headed sea snakes Disteria major shift seagrass microhabitats to avoid shark predation. Marine Ecology Progress Series, 387:287–293, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2009/387/m387p287.pdf.

Weetman:2006:GPS

[WHB^{+}06] D. Weetman, L. Hauser, M. K. Bayes, J. R. Ellis, and P. W. Shaw. Genetic population structure across a range of geographic scales in the commercially exploited marine gastropod Buccinum undatum. Marine Ecology Progress Series, 317:157–169, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/317/m317p157.pdf.

Wismer:2009:CSB

[WHB09] S. Wismer, A. S. Hoey, and D. R. Bellwood. Cross-shelf benthic community structure on the Great Barrier Reef: relationships between macroalgal cover and herbivore biomass. Marine Ecology Progress Series, 376:45–54, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/376/
m376p045.pdf.

Wildish:2004:ADO

[WHCP^{+}04] D. J. Wildish, J. E. Hughes-Clarke, G. W. Pohle, B. T. Hargrave, and L. M. Mayer. Acoustic detection of organic enrichment in sediments at a salmon farm is confirmed by independent groundtruthing methods. Marine Ecology Progress Series, 267:99–105, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/267/m267p099.pdf.

Waniek:2005:FCO

[WHD^{+}05] J. J. Waniek, N. P. Holliday, R. Davidson, L. Brown, and S. A. Henson. Freshwater control of onset and species composition of Greenland shelf spring bloom. Marine Ecology Progress Series, 288:45–57, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/288/m288p045.pdf.

Whitehead:2002:ECG

[Whi02] H. Whitehead. Estimates of the current global population size and historical trajectory for sperm whales. Marine Ecology Progress Series, 242:295–304, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/242/m242p295.pdf.

Whitaker:2006:GEM

[Whi06] K. Whitaker. Genetic evidence for mixed modes of reproduction in the coral Pocillopora damicornis and its effect on population structure. Marine Ecology Progress Series, 306:115–124, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p115.pdf.

Wathne:2000:PPN

[WHL00] J. A. Wathne, T. Haug, and C. Lydersen. Prey preference and niche overlap of ringed seals Phoca hispida and harp seals P. groenlandica in the Barents Sea. Marine Ecology Progress Series, 194:233–239, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/194/m194p233.pdf.

Wetz:2006:VPP

[WHL^{+}06] M. S. Wetz, K. C. Hayes, A. J. Lewitus, J. L. Wolny, and D. L. White. Variability in phytoplankton pigment biomass and taxonomic composition over tidal cycles in a salt marsh estuary. Marine Ecology Progress Series, 320:109–120, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/320/
m320p109.pdf.

Wanless:2008:LBN

[WHL^{+}08] S. Wanless, M. P. Harris, S. Lewis, M. Frederiksen, and S. Murray. Later breeding in northern gannets in the eastern Atlantic. Marine Ecology Progress Series, 370:263–269, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2008/370/m370p263.pdf.

Ward:2001:RMM

[WHM^{+}01] T. M. Ward, F. Hoedt, L. McLeay, W. F. Dimmlich, G. Jackson, P. J. Rogers, and K. Jones. Have recent mass mortalities of the sardine Sardinops sagax facilitated an expansion in the distribution and abundance of the anchovy Engraulis australis in South Australia? Marine Ecology Progress Series, 220:241–251, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/220/m220p241.pdf.

Walton:2000:UBF

[WHP00] M. J. Walton, R. J. Henderson, and P. P. Pomeroy. Use of blubber fatty acid profiles to distinguish dietary differences between grey seals Halichoerus grypus from two UK breeding colonies. Marine Ecology Progress Series, 193:201–208, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
193/m193p201.pdf.

Wanless:2005:LEV

[WHRS05] S. Wanless, M. P. Harris, P. Redman, and J. R. Speakman. Low energy values of fish as a probable cause of a major seabird breeding failure in the North Sea. Marine Ecology Progress Series, 294:1–8, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Warner:2009:GVN

[WHS^{+}09] R. R. Warner, S. L. Hamilton, M. S. Sheehy, L. D. Zeidberg, B. C. Brady, and J. E. Caselle. Geographic variation in natal and early larval trace-elemental signatures in the statoliths of the market squid Doryteuthis (formerly Loligo) opalescens. Marine Ecology Progress Series, 379:109–121, 2009. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m379p109.pdf.

Williams:2008:SIF

[WIB08] C. T. Williams, S. J. Iverson, and C. L. Buck. Stable isotopes and fatty acid signatures reveal age- and stage-dependent foraging niches in tufted puffins. Marine Ecology Progress Series, 363:287–298, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/363/m363p287.pdf.

Widdows:2001:BCR

[Wid01] J. Widdows. Bivalve clearance rates: inaccurate measurements or inaccurate reviews and misrepresentation? Marine Ecology Progress Series, 221:303–305, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

221/m221p303.pdf. See reply [Rii01a].

Watanuki:2009:CFS

[WIDM09] Y. Watanuki, M. Ito, T. Deguchi, and S. Minobe. Climateforced seasonal mismatch between the hatching of rhinoceros auklets and the availability of anchovy. Marine Ecology Progress Series, 393:259–271, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m393p259.pdf.

Wielgus:2003:CRE

[Wie03] J. Wielgus. The coral reef of Eilat (northern Red Sea) requires immediate protection MEPS 263:307. Marine Ecology Progress Series, ??:??, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/misc/m263p307.pdf.

Wieters:2005:UCP

[Wie05] E. A. Wieters. Upwelling control of positive interactions over mesoscales: a new link between bottom-up and topdown processes on rocky shores. Marine Ecology Progress Series, 301:43–54, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/301/m301p043.pdf.

Wilson:2003:PPT

[Wil03] R. P. Wilson. Penguins predict their performance. Marine Ecology Progress Series, 249:305–310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/249/
m249p305.pdf.

Wilson:2004:GMT

[Wil04] S. K. Wilson. Growth, mortality and turnover rates of a small detritivorous fish. Marine Ecology Progress Series, 284:253–259, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/284/m284p253.pdf.

Wilson:2009:ESNa

[Wil09] M. T. Wilson. Ecology of small neritic fishes in the western Gulf of Alaska. I. Geographic distribution in relation to prey density and the physical environment. Marine Ecology Progress Series, 392:223–237, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/392/m392p223.pdf.

Wing:2009:DSD

[Win09] S. R. Wing. Decadal-scale dynamics of sea urchin population networks in Fiordland, New Zealand are driven by juxtaposition of larval transport against benthic productivity gradients. Marine Ecology Progress Series, 378:125–134, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/378/
m378p125.pdf.

Wang:2009:STD

[WISH09] S. W. Wang, S. J. Iverson, A. M. Springer, and S. A. Hatch. Spatial and temporal diet segregation in northern fulmars Fulmarus glacialis breeding in Alaska: insights from fatty acid signatures. Marine Ecology Progress Series, 377:299– 307, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/377/m377p299.pdf.

Westerbom:2006:EWE

[WJ06] M. Westerbom and S. Jattu. Effects of wave exposure on the sublittoral distribution of blue mussels Mytilus edulis in a heterogeneous archipelago. Marine Ecology Progress Series, 306:191–200, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/306/m306p191.pdf.

Wilson:2009:ESNb

[WJB09] M. T. Wilson, C. M. Jump, and A. Buchheister. Ecology of small neritic fishes in the western Gulf of Alaska. II. Consumption of krill in relation to krill standing stock and the physical environment. Marine Ecology Progress Series, 392:239–251, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/392/m392p239.pdf.

Wilson:2006:CAF

[WJDA06] M. T. Wilson, C. M. Jump, and J. T. Duffy-Anderson. Comparative analysis of the feeding ecology of two pelagic forage fishes: capelin Mallotus villosus and walleye pollock Theragra chalcogramma. Marine Ecology Progress Series, 317: 245–258, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/317/m317p245.pdf.

Wang:2005:DDS

[WJWL05] G. Wang, X. Jiang, L. Wu, and S. Li. Differences in the density, sinking rate and biochemical composition of Centropages tenuiremis (Copepoda: Calanoida) subitaneous and diapause eggs. Marine Ecology Progress Series, 288:165– 171, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/288/m288p165.pdf.

Wieland:2000:STT

[WK00] A. Wieland and M. Kühl. Short-term temperature effects on oxygen and sulfide cycling in a hypersaline cyanobacterial mat (Solar Lake, Egypt). Marine Ecology Progress Series, 196:87–102, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/196/m196p087.pdf.

Webb:2002:ADJ

[WK02] S. R. Webb and R. T. Kneib. Abundance and distribution of juvenile white shrimp Litopenaeus setiferus within a tidal marsh landscape. Marine Ecology Progress Series, 232:213–223, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/232/m232p213.pdf.

Whitfield:2004:RTT

[WKD^{+}04] P. E. Whitfield, W. J. Kenworthy, M. J. Durako, K. K. Hammerstrom, and M. F. Merello. Recruitment of Thalassia testudinum seedlings into physically disturbed seagrass beds. Marine Ecology Progress Series, 267:121–131, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/267/
m267p121.pdf.

Watanabe:2002:FHJ

[WKH02] H. Watanabe, K. Kawaguchi, and A. Hayashi. Feeding habits of juvenile surface migratory myctophid fishes (family Myctophidae) in the Kuroshio region of the western North Pacific. Marine Ecology Progress Series, 236:263–272, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/236/
m236p263.pdf.

Ward:2007:TAC

[WKH07] J. R. Ward, K. Kim, and C. D. Harvell. Temperature affects coral disease resistance and pathogen growth. Marine Ecology Progress Series, 329:115–121, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/329/m329p115.pdf.

Watanabe:2004:FHN

[WKIK04] H. Watanabe, T. Kubodera, T. Ichii, and S. Kawahara. Feeding habits of neon flying squid Ommastrephes bartramii in the transitional region of the central North Pacific. Marine Ecology Progress Series, 266:173–184, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/266/
m266p173.pdf.

Wlodarska-Kowalczuk:2007:SNP

[WKK07] Maria Włodarska-Kowalczuk and Monika Kedra. Surrogacy in natural patterns of benthic distribution and diversity: selected taxa versus lower taxonomic resolution. Marine Ecology Progress Series, 351:53–63, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/351/m351p053.pdf.

Williams:2001:DDC

[WKL01] A. Williams, J. A. Koslow, and P. R. Last. Diversity, density and community structure of the demersal fish fauna of the continental slope off western Australia (20 to 35^{o}S). Marine Ecology Progress Series, 212:247–263, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/212/m212p247.pdf.

Waite:2003:RNP

[WKL^{+}03] T. D. Waite, J. Kazumi, P. V. Z. Lane, L. L. Farmer, S. G. Smith, S. L. Smith, G. Hitchcock, and T. R. Capo. Removal of natural populations of marine plankton by a large-scale ballast water treatment system. Marine Ecology Progress Series, 258:51–63, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/258/m258p051.pdf.

Watanabe:2006:DVM

[WKMK06] H. Watanabe, T. Kubodera, M. Moku, and K. Kawaguchi. Diel vertical migration of squid in the warm core ring and cold water masses in the transition region of the western North Pacific. Marine Ecology Progress Series, 315:187– 197, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/315/m315p187.pdf.

Wieters:2003:ATV

[WKN^{+}03] E. A. Wieters, D. M. Kaplan, S. A. Navarrete, A. Sotomayor, J. Largier, K. J. Nielsen, and F. Véliz. Alongshore and temporal variability in chlorophyll-a concentration in Chilean nearshore waters. Marine Ecology Progress Series, 249:93– 105, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/249/m249p093.pdf.

Wicks:2009:ESO

[WKOE09] E. C. Wicks, E. W. Koch, J. M. O’Neil, and K. Elliston. Effects of sediment organic content and hydrodynamic conditions on the growth and distribution of Zostera marina.

Marine Ecology Progress Series, 378:71–80, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/378/
m378p071.pdf.

Wlodarska-Kowalczuk:2005:BRC

[WKPK05] M. Włodarska-Kowalczuk, T. H. Pearson, and M. A. Kendall. Benthic response to chronic natural physical disturbance by glacial sedimentation in an Arctic fjord. Marine Ecology Progress Series, 303:31–41, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/303/m303p031.pdf.

Wallace:2006:EBC

[WKPS06] B. P. Wallace, S. S. Kilham, F. V. Paladino, and J. R. Spotila. Energy budget calculations indicate resource limitation in Eastern Pacific leatherback turtles. Marine Ecology Progress Series, 318:263–270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/318/m318p263.pdf.

Watanabe:2009:FES

[WKY09] H. Watanabe, T. Kubodera, and K. Yokawa. Feeding ecology of the swordfish Xiphias gladius in the subtropical region and transition zone of the western North Pacific. Marine Ecology Progress Series, 396:111–122, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/396/m396p111.pdf.

Wang:2000:MRB

[WKYL00] W. X. Wang, C. Ke, K. N. Yu, and P. K. S. Lam. Modeling radiocesium bioaccumulation in a marine food chain. Marine Ecology Progress Series, 208:41–50, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/208/m208p041.pdf.

Wallace:2003:SCCb

[WL03] W. G. Wallace and S. N. Luoma. Subcellular compartmentalization of Cd and Zn in two bivalves. II. Significance of trophically available metal (TAM). Marine Ecology Progress Series, 257:125–137, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2003/257/m257p125.pdf.

Werry:2005:GCM

[WL05] J. Werry and S. Y. Lee. Grapsid crabs mediate link between mangrove litter production and estuarine planktonic food chains. Marine Ecology Progress Series, 293:165–176, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/293/m293p165.pdf.

Waite:2006:BDT

[WL06] A. M. Waite and O. Lindahl. Bloom and decline of the toxic flagellate Chattonella marina in a Swedish fjord. Marine Ecology Progress Series, 326:77–83, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/326/
m326p077.pdf.

Weimerskirch:2008:FSM

[WLB08] H. Weimerskirch, M. Le Corre, and C. A. Bost. Foraging strategy of masked boobies from the largest colony in the world: relationship to environmental conditions and fisheries. Marine Ecology Progress Series, 362:291–302, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p291.pdf.

Weidner:2004:EMN

[WLdG^{+}04] K. Weidner, B. G. Lages, B. A. P. da Gama, M. Molis, M. Wahl, and R. C. Pereira. Effect of mesograzers and nutrient levels on induction of defenses in several Brazilian macroalgae. Marine Ecology Progress Series, 283:113–125, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/283/
m283p113.pdf.

Weimerskirch:2004:FST

[WLJ^{+}04] H. Weimerskirch, M. Le Corre, S. Jaquemet, M. Potier, and F. Marsac. Foraging strategy of a top predator in tropical waters: great frigatebirds in the Mozambique Channel. Marine Ecology Progress Series, 275:297–308, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/275/
m275p297.pdf.

Williams:2008:BTC

[WLJ08] C. J. Williams, P. J. Lavrentyev, and F. J. Jochem. Bottomup and top-down control of heterotrophic bacterioplankton growth in a phosphorus-depleted subtropical estuary, Florida Bay, USA. Marine Ecology Progress Series, 372: 7–18, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p007.pdf.

Weimerskirch:2005:FST

[WLJM05] H. Weimerskirch, M. Le Corre, S. Jaquemet, and F. Marsac. Foraging strategy of a tropical seabird, the red-footed booby, in a dynamic marine environment. Marine Ecology Progress Series, 288:251–261, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2005/288/m288p251.pdf.

Wallace:2003:SCCa

[WLL03] W. G. Wallace, B. G. Lee, and S. N. Luoma. Subcellular compartmentalization of Cd and Zn in two bivalves. I. Significance of metal-sensitive fractions (MSF) and biologically detoxified metal (BDM). Marine Ecology Progress Series, 249:183–197, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/249/m249p183.pdf.

Westerbom:2006:ISS

[WLM06] M. Westerbom, A. Lappalainen, and O. Mustonen. Invariant size selection of blue mussels by roach despite variable prey size distributions. Marine Ecology Progress Series, 328: 161–170, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/328/m328p161.pdf.

Wang:2009:SAR

[WLS^{+}09a] S. Wang, C. Li, S. Sun, X. Ning, and W. Zhang. Spring and autumn reproduction of Calanus sinicus in the Yellow Sea. Marine Ecology Progress Series, 379:123–133, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p123.pdf.

Watanabe:2009:DBS

[WLS^{+}09b] Y. Watanabe, C. Lydersen, K. Sato, Y. Naito, N. Miyazaki, and K. M. Kovacs. Diving behavior and swimming style of nursing bearded seal pups. Marine Ecology Progress Series, 380:287–294, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/380/m380p287.pdf.

Wong:2003:ACR

[WLT^{+}03] W. H. Wong, J. S. Levinton, B. S. Twining, N. S. Fisher, B. P. Kelaher, and A. K. Alt. Assimilation of carbon from a rotifer by the mussels Mytilus edulis and Perna viridis: a potential food-web link. Marine Ecology Progress Series, 253:175–182, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p175.pdf.

Weijerman:2005:RSM

[WLZ05] M. Weijerman, H. Lindeboom, and A. F. Zuur. Regime shifts in marine ecosystems of the North Sea and Wadden Sea. Marine Ecology Progress Series, 298:21–39, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/298/
m298p021.pdf.

Wilson:2001:EIP

[WM01] D. T. Wilson and M. G. Meekan. Environmental influences on patterns of larval replenishment of coral reef fishes. Marine Ecology Progress Series, 222:197–207, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p197.pdf.

Waters:2002:CSS

[WM02a] R. L. Waters and J. G. Mitchell. Centimetre-scale spatial structure of estuarine in vivo fluorescence profiles. Marine Ecology Progress Series, 237:51–63, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/237/
m237p051.pdf.

Wilson:2002:GRA

[WM02b] D. T. Wilson and M. G. Meekan. Growth-related advantages for survival to the point of replenishment in the coral reef fish Stegastes partitus (Pomacentridae). Marine Ecology Progress Series, 231:247–260, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/231/m231p247.pdf.

Walker:2004:OCF

[WM04] S. P. W. Walker and M. I. McCormick. Otolith-check formation and accelerated growth associated with sex change in an annual protogynous tropical fish. Marine Ecology Progress Series, 266:201–212, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/266/m266p201.pdf.

Williams:2005:CDO

[WM05] D. E. Williams and M. W. Miller. Coral disease outbreak: pattern, prevalence and transmission in Acropora cervicornis. Marine Ecology Progress Series, 301:119–128, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/301/
m301p119.pdf.

Waldbusser:2006:MMP

[WM06] G. G. Waldbusser and R. L. Marinelli. Macrofaunal modification of porewater advection: role of species function, species interaction, and kinetics. Marine Ecology Progress Series, 311:217–231, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m311p217.pdf.

Willis:2000:DSV

[WMB00] T. J. Willis, R. B. Millar, and R. C. Babcock. Detection of spatial variability in relative density of fishes: comparison of visual census, angling, and baited underwater video. Marine Ecology Progress Series, 198:249–260, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
198/m198p249.pdf.

Wing:2008:PDT

[WMCF08] S. R. Wing, R. J. McLeod, K. L. Clark, and R. D. Frew. Plasticity in the diet of two echinoderm species across an ecotone: microbial recycling of forest litter and bottom-up forcing of population structure. Marine Ecology Progress Series, 360:115–123, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/360/m360p115.pdf.

Winkler:2007:TDT

[WMD^{+}07] G. Winkler, C. Martineau, J. J. Dodson, W. F. Vincent, and L. E. Johnson. Trophic dynamics of two sympatric mysid species in an estuarine transition zone. Marine Ecology Progress Series, 332:171–187, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/332/m332p171.pdf.

Watson:2002:SME

[WMG02] M. Watson, J. L. Munro, and F. R. Gell. Settlement, movement and early juvenile mortality of the yellowtail snapper Ocyurus chrysurus. Marine Ecology Progress Series, 237: 247–256, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p247.pdf.

Watanabe:2003:DDW

[WMS^{+}03a] Y. Watanabe, Y. Mitani, K. Sato, M. F. Cameron, and Y. Naito. Dive depths of Weddell seals in relation to vertical prey distribution as estimated by image data. Marine Ecology Progress Series, 252:283–288, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/252/m252p283.pdf.

Waters:2003:GCC

[WMS03b] R. L. Waters, J. G. Mitchell, and J. Seymour. Geostatistical characterisation of centimetre-scale spatial structure of in vivo fluorescence. Marine Ecology Progress Series, 251: 49–58, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/251/m251p049.pdf.

Wild:2008:OMR

[WMW^{+}08] C. Wild, C. Mayr, L. Wehrmann, S. Schöttner, M. Naumann, F. Hoffmann, and H. T. Rapp. Organic matter release by cold water corals and its implication for fauna-microbe interaction. Marine Ecology Progress Series, 372:67–75, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/372/
m372p067.pdf.

Wasmund:2005:CNF

[WNS^{+}05] N. Wasmund, G. Nausch, B. Schneider, K. Nagel, and M. Voss. Comparison of nitrogen fixation rates determined with different methods: a study in the Baltic proper. Marine Ecology Progress Series, 297:23–31, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/297/m297p023.pdf.

Webster:2002:SBP

[WNWH02] N. S. Webster, A. P. Negri, R. I. Webb, and R. T. Hill. A spongin-boring a-proteobacterium is the etiological agent of disease in the Great Barrier Reef sponge Rhopaloeides odorabile. Marine Ecology Progress Series, 232:305–309, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/232/m232p305.pdf.

Wenngren:2002:ICF

[WÓ02] J. Wenngren and E. Ólafsson. Intraspecific competition for food within and between year classes in the deposit-feeding amphipod Monoporeia affinis — the cause of population fluctuations? Marine Ecology Progress Series, 240:205– 213, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/240/m240p205.pdf.

Wu:2005:BGR

[WO05] R. S. S. Wu and Y. Y. Or. Bioenergetics, growth and reproduction of amphipods are affected by moderately low oxygen regimes. Marine Ecology Progress Series, 297:215– 223, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/297/m297p215.pdf.

Wonham:2005:PED

[WOH05] M. J. Wonham, M. O’Connor, and C. D. G. Harley. Positive effects of a dominant invader on introduced and native mudflat species. Marine Ecology Progress Series, 289: 109–116, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2005/289/m289p109.pdf.

Weng:2007:MBH

[WOL^{+}07a] K. C. Weng, J. B. O’Sullivan, C. G. Lowe, C. E. Winkler, H. Dewar, and B. A. Block. Movements, behavior and habitat preferences of juvenile white sharks Carcharodon carcharias in the eastern Pacific. Marine Ecology Progress Series, 338:211–224, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/338/m338p211.pdf.

Wolcott:2007:MSL

[Wol07b] B. D. Wolcott. Mechanical size limitation and life-history strategy of an intertidal seaweed. Marine Ecology Progress Series, 338:1–10, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wooldridge:2009:NCM

[Woo09] S. A. Wooldridge. A new conceptual model for the enhanced release of mucus in symbiotic reef corals during ‘bleaching’ conditions. Marine Ecology Progress Series, 396:145– 152, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/396/m396p145.pdf.

Waddell:2000:DCSa

[WP00a] B. Waddell and J. R. Pawlik. Defenses of Caribbean sponges against invertebrate predators. I. Assays with hermit crabs. Marine Ecology Progress Series, 195:125–132, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p125.pdf.

Waddell:2000:DCSb

[WP00b] B. Waddell and J. R. Pawlik. Defenses of Caribbean sponges against invertebrate predators. II. Assays with sea stars.

Marine Ecology Progress Series, 195:133–144, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p133.pdf.

Walton:2003:UBF

[WP03] M. Walton and P. Pomeroy. Use of blubber fatty acid profiles to detect inter-annual variations in the diet of grey seals Halichoerus grypus. Marine Ecology Progress Series, 248: 257–266, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/248/m248p257.pdf.

Wennhage:2007:FSA

[WP07] H. Wennhage and L. Pihl. From flatfish to sticklebacks: assemblage structure of epibenthic fauna in relation to macroalgal blooms. Marine Ecology Progress Series, 335:187–198, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/335/m335p187.pdf.

Widdows:2008:ESA

[WPB08a] J. Widdows, N. D. Pope, and M. D. Brinsley. Effect of Spartina anglica stems on near-bed hydrodynamics, sediment erodability and morphological changes on an intertidal mudflat. Marine Ecology Progress Series, 362:45–57, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/362/
m362p045.pdf.

Widdows:2008:ESB

[WPB^{+}08b] J. Widdows, N. D. Pope, M. D. Brinsley, H. Asmus, and R. M. Asmus. Effects of seagrass beds (Zostera noltii and Z. marina) on near-bed hydrodynamics and sediment resuspension. Marine Ecology Progress Series, 358:125–136, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p125.pdf.

Wawrik:2003:VSP

[WPC^{+}03] B. Wawrik, J. H. Paul, L. Campbell, D. Griffin, L. Houchin, A. Fuentes-Ortega, and F. Muller-Karger. Vertical structure of the phytoplankton community associated with a coastal plume in the Gulf of Mexico. Marine Ecology Progress Series, 251:87–101, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/251/m251p087.pdf.

Wong:2002:TII

[WPD02] G. T. F. Wong, A. U. Piumsomboon, and W. M. Dunstan. The transformation of iodate to iodide in marine phytoplankton cultures. Marine Ecology Progress Series, 237: 27–39, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/237/m237p027.pdf.

Wall:2008:FSZ

[WPG08] C. C. Wall, B. J. Peterson, and C. J. Gobler. Facilitation of seagrass Zostera marina productivity by suspensionfeeding bivalves. Marine Ecology Progress Series, 357:165– 174, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/357/m357p165.pdf.

Williams:2001:LGH

[WPH01] I. D. Williams, N. V. C. Polunin, and V. J. Hendrick. Limits to grazing by herbivorous fishes and the impact of low coral cover on macroalgal abundance on a coral reef in Belize. Marine Ecology Progress Series, 222:187–196, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p187.pdf.

Wijekoon:2009:PCV

[WPIB09] M. P. Wijekoon, V. Puvanendran, D. W. Ings, and J. A. Brown. Possible countergradient variation in growth of juvenile cod Gadus morhua from the northwest Atlantic. Marine Ecology Progress Series, 375:229–238, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/375/
m375p229.pdf.

Wakefield:2009:QHU

[WPM09] E. D. Wakefield, R. A. Phillips, and J. Matthiopoulos. Quantifying habitat use and preferences of pelagic seabirds using individual movement data: a review. Marine Ecology Progress Series, 391:165–182, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/theme/m391p165.pdf.

Williams:2004:DGY

[WPPS04] K. Williams, N. Papanikos, R. P. Phelps, and J. D. Shardo. Development, growth, and yolk utilization of hatchery-reared red snapper Lutjanus campechanus larvae. Marine Ecology Progress Series, 275:231–239, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/275/m275p231.pdf.

Ward-Paige:2005:RNS

[WPRS05] C. A. Ward-Paige, M. J. Risk, and O. A. Sherwood. Reconstruction of nitrogen sources on coral reefs: δ ^{15}N and δ 13C in gorgonians from Florida Reef Tract. Marine Ecology Progress Series, 296:155–163, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/296/m296p155.pdf.

Wu:2007:STC

[WPTS07] Y. Wu, T. Platt, C. C. L. Tang, and S. Sathyendranath. Shortterm changes in chlorophyll distribution in response to a moving storm: a modelling study. Marine Ecology Progress Series, 335:57–68, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/335/m335p057.pdf.

Wu:2008:RDT

[WPTS08] Y. Wu, T. Platt, C. C. L. Tang, and S. Sathyendranath. Regional differences in the timing of the spring bloom in the Labrador Sea. Marine Ecology Progress Series, 355: 9–20, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/355/m355p009.pdf.

Wang:2000:DUC

[WR00a] W. X. Wang and P. S. Rainbow. Dietary uptake of Cd, Cr, and Zn by the barnacle Balanus trigonus: influence of diet composition. Marine Ecology Progress Series, 204:159–168, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p159.pdf.

Worm:2000:DNA

[WR00b] B. Worm and T. B. H. Reusch. Do nutrient availability and plant density limit seagrass colonization in the Baltic Sea? Marine Ecology Progress Series, 200:159–166, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p159.pdf.

Waluda:2006:RSM

[WR06] C. M. Waluda and P. G. Rodhouse. Remotely sensed mesoscale oceanography of the Central Eastern Pacific and recruitment variability in Dosidicus gigas. Marine Ecology Progress Series, 310:25–32, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/310/m310p025.pdf.

White:2007:DFA

[WR07] J. W. White and B. I. Ruttenberg. Discriminant function analysis in marine ecology: some oversights and their solutions. Marine Ecology Progress Series, 329:301–305, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/329/
m329p301.pdf.

Wild:2005:RPM

[WRH05] C. Wild, H. Røy, and M. Huettel. Role of pelletization in mineralization of fine-grained coastal sediments. Marine Ecology Progress Series, 291:23–33, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/291/m291p023.pdf.

Wright:2007:UMP

[Wri07] P. J. Wright. Understanding the maturation process for field investigations of fisheries-induced evolution. Marine Ecology Progress Series, 335:279–283, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/theme/m335p279.pdf.

Wild:2005:BMD

[WRJ^{+}05] C. Wild, M. Rasheed, C. Jantzen, P. Cook, U. Struck, M. Huettel, and A. Boetius. Benthic metabolism and degradation of natural particulate organic matter in carbonate and silicate reef sands of the northern Red Sea. Marine Ecology Progress Series, 298:69–78, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/298/m298p069.pdf.

Woehler:2003:DSS

[WRW03] E. J. Woehler, B. Raymond, and D. J. Watts. Decadal-scale seabird assemblages in Prydz Bay, East Antarctica. Marine Ecology Progress Series, 251:299–310, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/251/
m251p299.pdf.

Wild:2004:DMC

[WRW^{+}04] C. Wild, M. Rasheed, U. Werner, U. Franke, R. Johnstone, and M. Huettel. Degradation and mineralization of coral mucus in reef environments. Marine Ecology Progress Series, 267:159–171, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/267/m267p159.pdf.

Woehler:2006:CDW

[WRW06] E. J. Woehler, B. Raymond, and D. J. Watts. Convergence or divergence: where do short-tailed shearwaters forage in the Southern Ocean? Marine Ecology Progress Series, 324:261–270, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/324/m324p261.pdf.

Worm:2000:RDI

[WS00] B. Worm and U. Sommer. Rapid direct and indirect effects of a single nutrient pulse in a seaweed-epiphyte-grazer system. Marine Ecology Progress Series, 202:283–288, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
202/m202p283.pdf.

Wiklund:2001:IRA

[WS01] A. K. Eriksson Wiklund and B. Sundelin. Impaired reproduction in the amphipods Monoporeia affinis and Pontoporeia femorata as a result of moderate hypoxia and increased temperature. Marine Ecology Progress Series, 222:131–141, 2001.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p131.pdf.

Wolfstein:2002:PEP

[WS02] K. Wolfstein and L. J. Stal. Production of extracellular polymeric substances (EPS) by benthic diatoms: effect of irradiance and temperature. Marine Ecology Progress Series, 236:13–22, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/236/m236p013.pdf.

Wiklund:2004:BST

[WS04a] A. K. Eriksson Wiklund and B. Sundelin. Biomarker sensitivity to temperature and hypoxia — a seven year field study. Marine Ecology Progress Series, 274:209–214, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/274/
m274p209.pdf.

Williams:2004:EIS

[WS04b] S. L. Williams and S. L. Schroeder. Eradication of the invasive seaweed Caulerpa taxifolia by chlorine bleach. Marine Ecology Progress Series, 272:69–76, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p069.pdf.

Womble:2006:SAA

[WS06] J. N. Womble and M. F. Sigler. Seasonal availability of abundant, energy-rich prey influences the abundance and diet of a marine predator, the Steller sea lion Eumetopias jubatus. Marine Ecology Progress Series, 325:281–293, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/325/
m325p281.pdf.

Winterbottom:2008:SLH

[WS08] Richard Winterbottom and Laura Southcott. Short lifespan and high mortality in the western Pacific coral reef goby Trimma nasa. Marine Ecology Progress Series, 366:203– 208, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/366/m366p203.pdf.

Wendel:2009:AMC

[WS09] C. M. Wendel and D. L. Smee. Ambient malathion concentrations modify behavior and increase mortality in blue crabs. Marine Ecology Progress Series, 392:157–165, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/392/
m392p157.pdf.

Ward:2006:PCS

[WSA^{+}06] P. Ward, R. Shreeve, A. Atkinson, B. Korb, M. Whitehouse, S. Thorpe, D. Pond, and N. Cunningham. Plankton community structure and variability in the Scotia Sea: austral summer 2003. Marine Ecology Progress Series, 309: 75–91, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/309/m309p075.pdf.

White:2006:MCB

[WSL06] A. E. White, Y. H. Spitz, and R. M. Letelier. Modeling carbohydrate ballasting by Trichodesmium spp. Marine Ecology Progress Series, 323:35–45, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m323p035.pdf.

Weimerskirch:2009:SSS

[WST^{+}09] H. Weimerskirch, S. A. Shaffer, Y. Tremblay, D. P. Costa, H. Gadenne, A. Kato, Y. Ropert-Coudert, K. Sato, and D. Aurioles. Species- and sex-specific differences in foraging behaviour and foraging zones in blue-footed and brown boobies in the Gulf of California. Marine Ecology Progress Series, 391:267–278, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/theme/m391p267.pdf.

Walther:2006:WFC

[WT06] B. D. Walther and S. R. Thorrold. Water, not food, contributes the majority of strontium and barium deposited in the otoliths of a marine fish. Marine Ecology Progress Series, 311:125–130, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/311/m311p125.pdf.

Weersing:2009:PGL

[WT09] K. Weersing and R. J. Toonen. Population genetics, larval dispersal, and connectivity in marine systems. Marine Ecology Progress Series, 393:1–12, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Watanabe:2005:PHA

[WTF^{+}05] H. Watanabe, S. Tsuchida, K. Fujikura, H. Yamamoto, F. Inagaki, M. Kyo, and S. Kojima. Population history associated with hydrothermal vent activity inferred from genetic structure of neoverrucid barnacles around Japan. Marine Ecology Progress Series, 288:233–240, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/288/m288p233.pdf.

Wild:2004:RRC

[WTH04] C. Wild, R. Tollrian, and M. Huettel. Rapid recycling of coral mass-spawning products in permeable reef sediments. Marine Ecology Progress Series, 271:159–166, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p159.pdf.

Waggett:2008:AGP

[WTP08] R. J. Waggett, P. A. Tester, and A. R. Place. Anti-grazing properties of the toxic dinoflagellate Karlodinium veneficum during predator–prey interactions with the copepod Acartia tonsa. Marine Ecology Progress Series, 366:31–42, 2008. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m366p031.pdf.

Winship:2002:BME

[WTR02] A. J. Winship, A. W. Trites, and D. A. S. Rosen. A bioenergetic model for estimating the food requirements of Steller sea lions Eumetopias jubatus in Alaska, USA. Marine Ecology Progress Series, 229:291–312, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/229/m229p291.pdf.

Webb:2009:LHM

[WTS09] T. J. Webb, E. H. M. Tyler, and P. J. Somerfield. Life history mediates large-scale population ecology in marine benthic taxa. Marine Ecology Progress Series, 396:293–306, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/theme/
m396p293.pdf.

Wasmund:2001:ENF

[WVL01] N. Wasmund, M. Voss, and K. Lochte. Evidence of nitrogen fixation by non-heterocystous cyanobacteria in the Baltic Sea and re-calculation of a budget of nitrogen fixation. Marine Ecology Progress Series, 214:1–14, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/214/m214p001.pdf.

Wang:2003:BKE

[WW03] W. X. Wang and R. S. K. Wong. Bioaccumulation kinetics and exposure pathways of inorganic mercury and methylmercury in a marine fish, the sweetlips Plectorhinchus gibbosus. Marine Ecology Progress Series, 261:257–268, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/261/
m261p257.pdf.

Wetz:2004:RBS

[WW04] M. S. Wetz and P. A. Wheeler. Response of bacteria to simulated upwelling phytoplankton blooms. Marine Ecology Progress Series, 272:49–57, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/272/m272p049.pdf.

Wahlberg:2005:HFT

[WW05] M. Wahlberg and H. Westerberg. Hearing in fish and their reactions to sounds from offshore wind farms. Marine Ecology Progress Series, 288:295–309, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2005/288/m288p295.pdf.

Wai:2006:EGC

[WW06a] T. C. Wai and G. A. Williams. Effect of grazing on coralline algae in seasonal, tropical, low-shore rock pools: spatio-temporal variation in settlement and persistence. Marine Ecology Progress Series, 326:99–113, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/326/
m326p099.pdf.

Wang:2006:BTB

[WW06b] X. Wang and W. X. Wang. Bioaccumulation and transfer of benzo(a)pyrene in a simplified marine food chain. Marine Ecology Progress Series, 312:101–111, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/312/
m312p101.pdf.

Ward:2007:PGC

[WW07] B. A. Ward and J. J. Waniek. Phytoplankton growth conditions during autumn and winter in the Irminger Sea, North Atlantic. Marine Ecology Progress Series, 334:47– 61, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2007/334/m334p047.pdf.

Wainright:2000:RIB

[WWAC00] S. C. Wainright, M. P. Weinstein, K. W. Able, and C. A. Currin. Relative importance of benthic microalgae, phytoplankton and the detritus of smooth cordgrass Spartina alterniflora and the common reed Phragmites australis to brackish-marsh food webs. Marine Ecology Progress Series, 200:77–91, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
200/m200p077.pdf.

Westerman:2009:VBP

[WWDH09] E. L. Westerman, R. Whitlatch, J. A. Dijkstra, and L. G. Harris. Variation in brooding period masks similarities in response to changing temperatures. Marine Ecology Progress Series, 391:13–19, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/391/m391p013.pdf.

Wade:2007:DSS

[WWGR07] P. R. Wade, G. M. Watters, T. Gerrodette, and S. B. Reilly. Depletion of spotted and spinner dolphins in the eastern tropical Pacific: modeling hypotheses for their lack of recovery. Marine Ecology Progress Series, 343:1–14, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wild:2005:ICM

[WWH05] C. Wild, H. Woyt, and M. Huettel. Influence of coral mucus on nutrient fluxes in carbonate sands. Marine Ecology Progress Series, 287:87–98, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/287/m287p087.pdf.

Wanless:2004:EDS

[WWHE04] S. Wanless, P. J. Wright, M. P. Harris, and D. A. Elston. Evidence for decrease in size of lesser sandeels Ammodytes marinus in a North Sea aggregation over a 30-yr period. Marine Ecology Progress Series, 279:237–246, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/279/
m279p237.pdf.

Wetz:2004:LIG

[WWL04] M. S. Wetz, P. A. Wheeler, and R. M. Letelier. Light-induced growth of phytoplankton collected during the winter from the benthic boundary layer off Oregon, USA. Marine Ecology Progress Series, 280:95–104, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/280/m280p095.pdf.

Williams:2006:ERC

[WWMF06] I. D. Williams, W. J. Walsh, A. Miyasaka, and A. M. Friedlander. Effects of rotational closure on coral reef fishes in Waikiki– Diamond Head Fishery Management Area, Oahu, Hawaii. Marine Ecology Progress Series, 310:139–149, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/310/
m310p139.pdf.

Wonham:2001:GSR

[WWR^{+}01] M. J. Wonham, W. C. Walton, G. M. Ruiz, A. M. Frese, and B. S. Galil. Going to the source: role of the invasion pathway in determining potential invaders. Marine Ecology Progress Series, 215:1–12, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/215/m215p001.pdf.

Witteveen:2009:TMM

[WWR09] B. H. Witteveen, G. A. J. Worthy, and J. D. Roth. Tracing migratory movements of breeding North Pacific humpback whales using stable isotope analysis. Marine Ecology Progress Series, 393:173–183, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2009/393/m393p173.pdf.

Womble:2005:DSS

[WWS^{+}05] J. N. Womble, M. F. Willson, M. F. Sigler, B. P. Kelly, and G. R. VanBlaricom. Distribution of Steller sea lions Eumetopias jubatus in relation to spring-spawning fish in SE Alaska. Marine Ecology Progress Series, 294:271–282, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p271.pdf.

Williams:2006:IOE

[WWTH06] I. D. Williams, W. J. Walsh, B. N. Tissot, and L. E. Hallacher. Impact of observers’ experience level on counts of fishes in underwater visual surveys. Marine Ecology Progress Series, 310:185–191, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/310/m310p185.pdf.

Wernberg:2008:PST

[WWV08] T. Wernberg, M. White, and M. A. Vanderklift. Population structure of turbinid gastropods on wave-exposed subtidal reefs: effects of density, body size and algae on grazing behaviour. Marine Ecology Progress Series, 362:169– 179, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/362/m362p169.pdf.

Woodd-Walker:2002:LSP

[WWWC02] R. S. Woodd-Walker, P. Ward, and A. Clarke. Large-scale patterns in diversity and community structure of surface water copepods from the Atlantic Ocean. Marine Ecology Progress Series, 236:189–203, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2002/236/m236p189.pdf.

Witteveen:2009:PSN

[WWWR09] B. H. Witteveen, G. A. J. Worthy, K. M. Wynne, and J. D. Roth. Population structure of North Pacific humpback whales on their feeding grounds revealed by stable carbon and nitrogen isotope ratios. Marine Ecology Progress Series, 379: 299–310, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/379/m379p299.pdf.

Woodson:2007:CHF

[WWWY07] C. B. Woodson, D. R. Webster, M. J. Weissburg, and J. Yen. Cue hierarchy and foraging of calanoid copepods: ecological implications of oceanographic structure. Marine Ecology Progress Series, 330:163–177, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2007/330/m330p163.pdf.

Wegley:2004:CAA

[WYB^{+}04] L. Wegley, Y. Yu, M. Breitbart, V. Casas, D. I. Kline, and F. Rohwer. Coral-associated archaea. Marine Ecology Progress Series, 273:89–96, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/273/m273p089.pdf.

Wang:2002:BSM

[WYFX02] W. X. Wang, Q. L. Yan, W. Fan, and Y. Xu. Bioavailability of sedimentary metals from a contaminated bay. Marine Ecology Progress Series, 240:27–38, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/240/m240p027.pdf.

Wilson:2004:IMP

[WZ04] R. P. Wilson and I. Zimmer. Inspiration by Magellanic penguins: reduced swimming effort when under pressure. Marine Ecology Progress Series, 278:303–307, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/278/
m278p303.pdf.

Xavier:2007:AGR

[XBW07] B. M. Xavier, G. M. Branch, and E. Wieters. Abundance, growth and recruitment of Mytilus galloprovincialis on the west coast of South Africa in relation to upwelling. Marine Ecology Progress Series, 346:189–201, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/346/
m346p189.pdf.

Xu:2001:IRT

[XW01] Y. Xu and W. X. Wang. Individual responses of trace-element assimilation and physiological turnover by the marine copepod Calanus sinicus to changes in food quantity. Marine Ecology Progress Series, 218:227–238, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/218/m218p227.pdf.

Xu:2002:EPF

[XW02] Y. Xu and W. X. Wang. Exposure and potential food chain transfer factor of Cd, Se and Zn in marine fish Lutjanus argentimaculatus. Marine Ecology Progress Series, 238:173– 186, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/238/m238p173.pdf.

Xu:2003:FDC

[XW03] Y. Xu and W. X. Wang. Fates of diatom carbon and trace elements by the grazing of a marine copepod. Marine Ecology Progress Series, 254:225–238, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/254/m254p225.pdf.

Xie:2005:RPH

[XWQQ05] Z. C. Xie, N. C. Wong, P. Y. Qian, and J. W. Qiu. Responses of polychaete Hydroides elegans life stages to copper stress. Marine Ecology Progress Series, 285:89–96, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/285/
m285p089.pdf.

Xu:2009:NLH

[XYH^{+}09] J. Xu, K. Yin, A. Y. T. Ho, J. H. W. Lee, D. M. Anderson, and P. J. Harrison. Nutrient limitation in Hong Kong waters inferred from comparison of nutrient ratios, bioassays and ^{33}P turnover times. Marine Ecology Progress Series, 388:81–97, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p081.pdf.

Xie:2009:MCM

[XZLP09] H. Xie, Y. Zhang, K. Lemarchand, and P. Poulin. Microbial carbon monoxide uptake in the St. Lawrence estuarine system. Marine Ecology Progress Series, 389:17–29, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/389/
m389p017.pdf.

Xie:2005:BCC

[XZUK05] H. Xie, O. C. Zafiriou, T. P. Umile, and D. J. Kieber. Biological consumption of carbon monoxide in Delaware Bay, NW Atlantic and Beaufort Sea. Marine Ecology Progress Series, 290:1–14, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/290/m290p001.pdf.

Yasuda:2009:CFB

[YA09] T. Yasuda and N. Arai. Changes in flipper beat frequency, body angle and swimming speed of female green turtles Chelonia mydas. Marine Ecology Progress Series, 386:275– 286, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2009/386/m386p275.pdf.

Yakovis:2007:EHA

[YAF^{+}07] E. L. Yakovis, A. V. Artemieva, M. V. Fokin, M. A. Varfolomeeva, and N. N. Shunatova. Effect of habitat architecture on mobile benthic macrofauna associated with patches of barnacles and ascidians. Marine Ecology Progress Series, 348:117–124, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/348/m348p117.pdf.

Yap:2009:LCC

[Yap09] H. T. Yap. Local changes in community diversity after coral transplantation. Marine Ecology Progress Series, 374: 33–41, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/374/m374p033.pdf.

York:2006:FAH

[YBGP06] P. H. York, D. J. Booth, T. M. Glasby, and B. C. Pease. Fish assemblages in habitats dominated by Caulerpa taxifolia and native seagrasses in south-eastern Australia. Marine Ecology Progress Series, 312:223–234, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/312/m312p223.pdf.

Yakovleva:2009:ASI

[YBY^{+}09] I. M. Yakovleva, A. H. Baird, H. H. Yamamoto, R. Bhagooli, M. Nonaka, and M. Hidaka. Algal symbionts increase oxidative damage and death in coral larvae at high temperatures. Marine Ecology Progress Series, 378:105–112, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m378p105.pdf.

Yannicelli:2006:CLD

[YCSS06] B. Yannicelli, L. R. Castro, W. Schneider, and M. Sobarzo. Crustacean larvae distribution in the coastal upwelling zone off Central Chile. Marine Ecology Progress Series, 319:175– 189, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2006/319/m319p175.pdf.

Yoshikawa:2004:LTM

[YF04] T. Yoshikawa and K. Furuya. Long-term monitoring of primary production in coastal waters by an improved natural fluorescence method. Marine Ecology Progress Series, 273: 17–30, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/273/m273p017.pdf.

Yamamoto:2002:PCB

[YFH^{+}02] H. Yamamoto, K. Fujikura, A. Hiraishi, K. Kato, and Y. Maki. Phylogenetic characterization and biomass estimation of bacterial endosymbionts associated with invertebrates dwelling in chemosynthetic communities of hydrothermal vent and cold seep fields. Marine Ecology Progress Series, 245:61– 67, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2002/245/m245p061.pdf.

Yamane:2009:ORI

[YG09] L. Yamane and S. E. Gilman. Opposite responses by an intertidal predator to increasing aquatic and aerial temperatures. Marine Ecology Progress Series, 393:27–36, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/393/
m393p027.pdf.

Yin:2000:IFE

[YH00] K. Yin and P. J. Harrison. Influences of flood and ebb tides on nutrient fluxes and chlorophyll on an intertidal flat. Marine Ecology Progress Series, 196:75–85, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/196/m196p075.pdf.

Ysebaert:2002:STV

[YH02] T. Ysebaert and P. M. J. Herman. Spatial and temporal variation in benthic macrofauna and relationships with environmental variables in an estuarine, intertidal soft-sediment environment. Marine Ecology Progress Series, 244:105–124, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2002/244/m244p105.pdf.

Yakovleva:2004:DRP

[YH04] I. Yakovleva and M. Hidaka. Differential recovery of PSII function and electron transport rate in symbiotic dinoflagellates as a possible determinant of bleaching susceptibility of corals. Marine Ecology Progress Series, 268:43–53, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/268/
m268p043.pdf.

Yamamura:2002:DWP

[YHSH02] O. Yamamura, S. Honda, O. Shida, and T. Hamatsu. Diets of walleye pollock Theragra chalcogramma in the Doto area, northern Japan: ontogenetic and seasonal variations. Marine Ecology Progress Series, 238:187–198, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2002/238/
m238p187.pdf.

Yamada:2003:MCC

[YI03a] Y. Yamada and T. Ikeda. Metabolism and chemical composition of four pelagic amphipods in the Oyashio region, western subarctic Pacific Ocean. Marine Ecology Progress Series, 253:233–241, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/253/m253p233.pdf.

Yokoyama:2003:FBT

[YI03b] H. Yokoyama and Y. Ishihi. Feeding of the bivalve Theora lubrica on benthic microalgae: isotopic evidence. Marine Ecology Progress Series, 255:303–309, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2003/255/
m255p303.pdf.

Yamada:2006:PMT

[YI06] Y. Yamada and T. Ikeda. Production, metabolism and trophic importance of four pelagic amphipods in the Oyashio region, western subarctic Pacific. Marine Ecology Progress Series, 308:155–163, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/308/m308p155.pdf.

Yokoyama:2007:VFS

[YI07] H. Yokoyama and Y. Ishihi. Variation in food sources of the macrobenthos along a land–sea transect: a stable isotope study. Marine Ecology Progress Series, 346:127–141, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/346/m346p127.pdf.

Yin:2002:MIS

[Yin02] K. Yin. Monsoonal influence on seasonal variations in nutrients and phytoplankton biomass in coastal waters of Hong Kong in the vicinity of the Pearl River estuary. Marine Ecology Progress Series, 245:111–122, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/245/m245p111.pdf.

Yin:2003:IMO

[Yin03] K. Yin. Influence of monsoons and oceanographic processes on red tides in Hong Kong waters. Marine Ecology Progress Series, 262:27–41, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/262/m262p027.pdf.

Yokoyama:2008:DTI

[YIY08] H. Yokoyama, Y. Ishihi, and S. Yamamoto. Diet-tissue isotopic fractionation of the Pacific oyster Crassostrea gigas. Marine Ecology Progress Series, 358:173–179, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/358/
m358p173.pdf.

Yukihira:2000:CET

[YLK00] H. Yukihira, J. S. Lucas, and D. W. Klumpp. Comparative effects of temperature on suspension feeding and energy budgets of the pearl oysters Pinctada margaritifera and P. maxima. Marine Ecology Progress Series, 195:179–188, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
195/m195p179.pdf.

Yoo:2004:UPA

[YLL^{+}04] H. Yoo, J. S. Lee, B. G. Lee, I. T. Lee, C. E. Schlekat, C. H. Koh, and S. N. Luoma. Uptake pathway for Ag bioaccumulation in three benthic invertebrates exposed to contaminated sediments. Marine Ecology Progress Series, 270: 141–152, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/270/m270p141.pdf.

Young:2005:NRA

[YLv^{+}05] E. B. Young, P. S. Lavery, B. van Elven, M. J. Dring, and J. A. Berges. Nitrate reductase activity in macroalgae and its vertical distribution in macroalgal epiphytes of seagrasses. Marine Ecology Progress Series, 288:103–114, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/288/
m288p103.pdf.

Yeung:2006:CSE

[YM06] C. Yeung and R. A. McConnaughey. Community structure of eastern Bering Sea epibenthic invertebrates from summer bottom-trawl surveys 1982 to 2002. Marine Ecology Progress Series, 318:47–62, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/318/m318p047.pdf.

Yost:2009:DDL

[YM09] D. M. Yost and C. L. Mitchelmore. Dimethylsulfoniopropionate (DMSP) lyase activity in different strains of the symbiotic alga Symbiodinium microadriaticum. Marine Ecology Progress Series, 386:61–70, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/386/m386p061.pdf.

Yunev:2005:LTV

[YMC05] O. A. Yunev, S. Moncheva, and J. Carstensen. Long-term variability of vertical chlorophyll-a and nitrate profiles in the open Black Sea: eutrophication and climate change. Marine Ecology Progress Series, 294:95–107, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/294/
m294p095.pdf.

Ysebaert:2002:MSR

[YMHV02] T. Ysebaert, P. Meire, P. M. J. Herman, and H. Verbeek. Macrobenthic species response surfaces along estuarine gradients: prediction by logistic regression. Marine Ecology Progress Series, 225:79–95, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/225/m225p079.pdf.

Yamasaki:2007:AIB

[YNM^{+}07] Y. Yamasaki, S. Nagasoe, T. Matsubara, T. Shikata, Y. Shimasaki, Y. Oshima, and T. Honjo. Allelopathic interactions between the bacillariophyte Skeletonema costatum and the raphidophyte Heterosigma akashiwo. Marine Ecology Progress Series, 339:83–92, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/339/m339p083.pdf.

Yamada:2009:OCP

[YOM^{+}09] Y. Yamada, A. Okamura, N. Mikawa, T. Utoh, N. Horie, S. Tanaka, M. J. Miller, and K. Tsukamoto. Ontogenetic changes in phototactic behavior during metamorphosis of artificially reared Japanese eel Anguilla japonica larvae. Marine Ecology Progress Series, 379:241–251, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/379/
m379p241.pdf.

Yoshioka:2008:BCS

[Yos08a] P. M. Yoshioka. The Bray–Curtis similarity index remains misidentified: Reply to Somerfield (2008). Marine Ecology Progress Series, 372:307–309, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m372p307.pdf.

Yoshioka:2008:MBC

[Yos08b] P. M. Yoshioka. Misidentification of the Bray–Curtis similarity index. Marine Ecology Progress Series, 368:309–310, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m368p309.pdf. See comment [Som08].

Yin:2000:DNP

[YQC^{+}00] K. Yin, P. Y. Qian, J. C. Chen, D. P. H. Hsieh, and P. J. Harrison. Dynamics of nutrients and phytoplankton biomass in the Pearl River estuary and adjacent waters of Hong Kong during summer: preliminary evidence for phosphorus and silicon limitation. Marine Ecology Progress Series, 194:295–305, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
194/m194p295.pdf.

Yin:2001:SPL

[YQW^{+}01] K. Yin, P. Y. Qian, M. C. S. Wu, J. C. Chen, L. Huang, X. Song, and W. Jian. Shift from P to N limitation of phytoplankton growth across the Pearl River estuarine plume during summer. Marine Ecology Progress Series, 221:17–28, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
221/m221p017.pdf.

Yu:2003:FET

[YSS03] O. H. Yu, H. L. Suh, and Y. Shirayama. Feeding ecology of three amphipod species Synchelidium lenorostralum, S. trioostegitum and Gitanopsis japonica in the surf zone of a sandy shore. Marine Ecology Progress Series, 258:189– 199, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/258/m258p189.pdf.

Yokoyama:2005:VDT

[YTH^{+}05] H. Yokoyama, A. Tamaki, K. Harada, K. Shimoda, K. Koyama, and Y. Ishihi. Variability of diet-tissue isotopic fractionation in estuarine macrobenthos. Marine Ecology Progress Series, 296:115–128, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/296/m296p115.pdf.

Yokoyama:2005:IEP

[YTK^{+}05] H. Yokoyama, A. Tamaki, K. Koyama, Y. Ishihi, K. Shimoda, and K. Harada. Isotopic evidence for phytoplankton as a major food source for macrobenthos on an intertidal sandflat in Ariake Sound, Japan. Marine Ecology Progress Series, 304:101–116, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2005/304/m304p101.pdf.

Yunev:2002:LTV

[YVB^{+}02] O. A. Yunev, V. I. Vedernikov, O. Basturk, A. Yilmaz, A. E. Kideys, S. Moncheva, and S. K. Konovalov. Longterm variations of surface chlorophyll-a and primary production in the open Black Sea. Marine Ecology Progress Series, 230:11–28, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p011.pdf.

Yoneda:2004:TSV

[YW04] M. Yoneda and P. J. Wright. Temporal and spatial variation in reproductive investment of Atlantic cod Gadus morhua in the northern North Sea and Scottish west coast. Marine Ecology Progress Series, 276:237–248, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2004/276/m276p237.pdf.

Yahel:2008:FAM

[YYK^{+}08] G. Yahel, R. Yahel, T. Katz, B. Lazar, B. Herut, and V. Tunnicliffe. Fish activity: a major mechanism for sediment resuspension and organic matter remineralization in coastal marine sediments. Marine Ecology Progress Series, 372:195– 209, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/372/m372p195.pdf.

Yanada:2000:SVT

[YYL^{+}00] M. Yanada, T. Yokokawa, C. W. Lee, H. Tanaka, I. Kudo, and Y. Maita. Seasonal variation of two different heterotrophic bacterial assemblages in subarctic coastal seawater. Marine Ecology Progress Series, 204:289–292, 2000. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
204/m204p289.pdf.

Zabin:2007:HLF

[ZA07] C. J. Zabin and A. Altieri. A Hawaiian limpet facilitates recruitment of a competitively dominant invasive barnacle. Marine Ecology Progress Series, 337:175–185, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/337/
m337p175.pdf.

Zabin:2009:BBN

[Zab09] C. J. Zabin. Battle of the barnacle newcomers: niche compression in invading species in Kaneohe Bay, Oahu, Hawaii. Marine Ecology Progress Series, 381:175–182, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2009/381/
m381p175.pdf.

Zacherl:2005:STVa

[Zac05a] D. C. Zacherl. Spatial and temporal variation in statolith and protoconch trace elements as natural tags to track larval dispersal. Marine Ecology Progress Series, 290:145–163, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Zacherl:2005:STVb

[Zac05b] D. C. Zacherl. Spatial and temporal variation in statolith and protoconch trace elements as natural tags to track larval dispersal. Marine Ecology Progress Series, 290:145–163, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
feature/m290p145.pdf.

Zhang:2003:MPP

[ZAH03] X. Zhang, J. T. Anderson, and R. R. Hood. Modeling Pfiesteria piscicida population dynamics: a new approach for tracking size and mass in mixotrophic species. Marine Ecology Progress Series, 256:29–44, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2003/256/m256p029.pdf.

Zerbini:2006:SMM

[ZAHJ^{+}06] A. N. Zerbini, A. Andriolo, M. P. Heide-Jørgensen, J. L. Pizzorno, Y. G. Maia, G. R. VanBlaricom, D. P. DeMaster, P. C. Simões-Lopes, S. Moreira, and C. Bethlem. Satellitemonitored movements of humpback whales Megaptera novaeangliae in the Southwest Atlantic Ocean. Marine Ecology Progress Series, 313:295–304, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2006/313/m313p295.pdf.

Zamon:2002:TCC

[Zam02] J. E. Zamon. Tidal changes in copepod abundance and maintenance of a summer Coscinodiscus bloom in the southern San Juan Channel, San Juan Islands, USA. Marine Ecology Progress Series, 226:193–210, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2002/226/m226p193.pdf.

Zamon:2003:MSA

[Zam03a] J. E. Zamon. Mixed species aggregations feeding upon herring and sandlance schools in a nearshore archipelago depend on flooding tidal currents. Marine Ecology Progress Series, 261:243–255, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/261/m261p243.pdf.

Zamzow:2003:UAC

[Zam03b] J. P. Zamzow. Ultraviolet-absorbing compounds in the mucus of temperate Pacific tidepool sculpins: variation over local and geographic scales. Marine Ecology Progress Series, 263:169–175, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/263/m263p169.pdf.

Zamzow:2007:UAC

[Zam07] J. P. Zamzow. Ultraviolet-absorbing compounds in the mucus of shallow-dwelling tropical reef fishes correlate with environmental water clarity. Marine Ecology Progress Series, 343:263–271, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/343/m343p263.pdf.

Zvuloni:2008:SDB

[ZAZL08] A. Zvuloni, R. Armoza-Zvuloni, and Y. Loya. Structural deformation of branching corals associated with the vermetid gastropod Dendropoma maxima. Marine Ecology Progress Series, 363:103–108, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/363/m363p103.pdf.

Zavalaga:2007:DBB

[ZBDE07] C. B. Zavalaga, S. Benvenuti, L. Dall’Antonia, and S. D. Emslie. Diving behavior of blue-footed boobies Sula nebouxii in northern Peru in relation to sex, body size and prey type. Marine Ecology Progress Series, 336:291–303, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/336/
m336p291.pdf.

Zakas:2009:RGF

[ZBNW09] C. Zakas, J. Binford, S. A. Navarrete, and J. P. Wares. Restricted gene flow in Chilean barnacles reflects an oceanographic and biogeographic transition zone. Marine Ecology Progress Series, 394:165–177, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/394/m394p165.pdf.

Zipperle:2009:EPS

[ZCR^{+}09] A. M. Zipperle, J. A. Coyer, K. Reise, W. T. Stam, and J. L. Olsen. Evidence for persistent seed banks in dwarf eelgrass Zostera noltii in the German Wadden Sea. Marine Ecology Progress Series, 380:73–80, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2009/380/m380p073.pdf.

Zakai:2006:LPP

[ZDA^{+}06] D. Zakai, Z. Dubinsky, A. Avishai, T. Caaras, and N. E. Chadwick. Lunar periodicity of planula release in the reefbuilding coral Stylophora pistillata. Marine Ecology Progress Series, 311:93–102, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2006/311/m311p093.pdf.

Zilberberg:2001:CAS

[ZE01] C. Zilberberg and P. J. Edmunds. Competition among small colonies of Agaricia: the importance of size asymmetry in determining competitive outcome. Marine Ecology Progress Series, 221:125–133, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/221/m221p125.pdf.

Zapata:2001:CCM

[ZER^{+}01] M. Zapata, B. Edvardsen, F. Rodríguez, M. A. Maestro, and J. L. Garrido. Chlorophyll-c_{2} monogalactosyldiacylglyceride ester (chl c_{2}-MGDG). A novel marker pigment for Chrysochromulina species (Haptophyta). Marine Ecology Progress Series, 219:85–98, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p085.pdf.

Zakardjian:2000:LSP

[ZGV00] B. A. Zakardjian, Y. Gratton, and A. F. Vézina. Late spring phytoplankton bloom in the Lower St. Lawrence Estuary: the flushing hypothesis revisited. Marine Ecology Progress Series, 192:31–48, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/192/m192p031.pdf.

Zapata:2002:PLD

[ZH02] F. A. Zapata and P. A. Herrón. Pelagic larval duration and geographic distribution of tropical eastern Pacific snappers (Pisces: Lutjanidae). Marine Ecology Progress Series, 230:295–300, 2002. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2002/230/m230p295.pdf.

Zemmelink:2005:GDD

[ZHS^{+}05] H. J. Zemmelink, L. Houghton, S. M. Sievert, N. M. Frew, and J. W. H. Dacey. Gradients in dimethylsulfide, dimethylsulfoniopropionate, dimethylsulfoxide, and bacteria near the sea surface. Marine Ecology Progress Series, 295:33–42, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2005/295/m295p033.pdf.

Zhuang:2006:SRB

[Zhu06] S. Zhuang. Species richness, biomass and diversity of macro-algal assemblages in tidepools of different sizes. Marine Ecology Progress Series, 309:67–73, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2006/309/
m309p067.pdf.

Zeldis:2008:ERC

[ZHWCS08] J. R. Zeldis, C. Howard-Williams, C. M. Carter, and D. R. Schiel. ENSO and riverine control of nutrient loading, phytoplankton biomass and mussel aquaculture yield in Pelorus Sound, New Zealand. Marine Ecology Progress Series, 371: 131–142, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2008/371/m371p131.pdf.

Zarauz:2007:MPD

[ZIUG07] L. Zarauz, X. Irigoien, A. Urtizberea, and M. Gonzalez. Mapping plankton distribution in the Bay of Biscay during three consecutive spring surveys. Marine Ecology Progress Series, 345:27–39, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/345/m345p027.pdf.

Zapata:2004:PPS

[ZJW^{+}04] M. Zapata, S. W. Jeffrey, S. W. Wright, F. Rodríguez, J. L. Garrido, and L. Clementson. Photosynthetic pigments in 37 species (65 strains) of Haptophyta: implications for oceanography and chemotaxonomy. Marine Ecology Progress Series, 270:83–102, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2004/270/m270p083.pdf.

Zimmermann:2001:BRS

[ZK01] C. Zimmermann and A. Kunzmann. Baseline respiration and spontaneous activity of sluggish marine tropical fish of the family Scorpaenidae. Marine Ecology Progress Series, 219:229–239, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/219/m219p229.pdf.

Zaitsev:2009:GPV

[ZK09] Yuvenaly P. Zaitsev and Otto Kinne. Gennady Polikarpov, Vice President of the Eco-Ethics International Union, celebrates his 80th birthday. Marine Ecology Progress Series, 388:i–ii, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2009/388/m388p000.pdf.

Zumholz:2007:SMT

[ZKHP07] K. Zumholz, A. Klügel, T. Hansteen, and U. Piatkowski. Statolith microchemistry traces the environmental history of the boreoatlantic armhook squid Gonatus fabricii. Marine Ecology Progress Series, 333:195–204, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/333/
m333p195.pdf.

Zubkov:2004:TPB

[ZLAK04] M. Zubkov, L. J. Linn, R. Amann, and R. P. Kiene. Temporal patterns of biological dimethylsulfide (DMS) consumption during laboratory-induced phytoplankton bloom cycles. Marine Ecology Progress Series, 271:77–86, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2004/271/
m271p077.pdf.

Zbinden:2004:DBA

[ZLGC04] M. Zbinden, N. Le Bris, F. Gaill, and P. Compère. Distribution of bacteria and associated minerals in the gill chamber of the vent shrimp Rimicaris exoculata and related biogeochemical processes. Marine Ecology Progress Series, 284: 237–251, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/284/m284p237.pdf.

Zaleski:2006:TDG

[ZM06] S. Frisch Zaleski and S. N. Murray. Taxonomic diversity and geographic distributions of aquarium-traded species of Caulerpa (Chlorophyta: Caulerpaceae) in southern California, USA. Marine Ecology Progress Series, 314:97–108, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2006/314/m314p097.pdf.

Zardi:2007:BSR

[ZMN07] G. I. Zardi, C. D. McQuaid, and K. R. Nicastro. Balancing survival and reproduction: seasonality of wave action, attachment strength and reproductive output in indigenous Perna perna and invasive Mytilus galloprovincialis mussels. Marine Ecology Progress Series, 334:155–163, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2007/334/
m334p155.pdf.

Zacherl:2003:TEF

[ZMP^{+}03] D. C. Zacherl, P. H. Manríquez, G. Paradis, R. W. Day, J. C. Castilla, R. R. Warner, D. W. Lea, and S. D. Gaines. Trace elemental fingerprinting of gastropod statoliths to study larval dispersal trajectories. Marine Ecology Progress Series, 248:297–303, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/248/m248p297.pdf.

Zardi:2007:UGS

[ZMTB07] G. I. Zardi, C. D. McQuaid, P. R. Teske, and N. P. Barker. Unexpected genetic structure of mussel populations in South Africa: indigenous Perna perna and invasive Mytilus galloprovincialis. Marine Ecology Progress Series, 337:135–144, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2007/337/m337p135.pdf.

Zador:2009:FIS

[ZPP09] S. G. Zador, J. K. Parrish, and A. E. Punt. Factors influencing subcolony colonization and persistence in a colonial seabird, the common murre Uria aalge. Marine Ecology Progress Series, 376:283–293, 2009. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2008/376/m376p283.pdf.

Zapata:2000:SCC

[ZRG00] M. Zapata, F. Rodríguez, and J. L. Garrido. Separation of chlorophylls and carotenoids from marine phytoplankton: a new HPLC method using a reversed phase C8 column and pyridine-containing mobile phases. Marine Ecology Progress Series, 195:29–45, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/195/m195p029.pdf.

Zharikov:2003:DBI

[ZS03] Y. Zharikov and G. A. Skilleter. Depletion of benthic invertebrates by bar-tailed godwits Limosa lapponica in a subtropical estuary. Marine Ecology Progress Series, 254:151– 162, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2003/254/m254p151.pdf.

Zhao:2004:SIR

[ZS04] L. Zhao and D. M. Schell. Stable isotope ratios in harbor seal Phoca vitulina vibrissae: effects of growth patterns on ecological records. Marine Ecology Progress Series, 281: 267–273, 2004. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2004/281/m281p267.pdf.

Zimmerman:2001:TIT

[ZSKA01] R. C. Zimmerman, D. L. Steller, D. G. Kohrs, and R. S. Alberte. Top-down impact through a bottom-up mechanism. In situ effects of limpet grazing on growth, light requirements and survival of the eelgrass Zostera marina. Marine Ecology Progress Series, 218:127–140, 2001. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/218/m218p127.pdf.

Zeller:2003:MRF

[ZSR03] D. Zeller, S. L. Stoute, and G. R. Russ. Movements of reef fishes across marine reserve boundaries: effects of manipulating a density gradient. Marine Ecology Progress Series, 254:269–280, 2003. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2003/254/m254p269.pdf.

Zheng:2007:ONC

[ZT07] G. M. Zheng and D. L. Tang. Offshore and nearshore chlorophyll increases induced by typhoon winds and subsequent terrestrial rainwater runoff. Marine Ecology Progress Series, 333:61–74, 2007. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2007/333/m333p061.pdf.

Zhao:2008:CPB

[ZTW08] H. Zhao, D. Tang, and Y. Wang. Comparison of phytoplankton blooms triggered by two typhoons with different intensities and translation speeds in the South China Sea. Marine Ecology Progress Series, 365:57–65, 2008. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2008/365/
m365p057.pdf.

Zupo:2000:EMF

[Zup00] V. Zupo. Effect of microalgal food on the sex reversal of Hippolyte inermis (Crustacea: Decapoda). Marine Ecology Progress Series, 201:251–259, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/201/m201p251.pdf.

Zhao:2001:BRM

[ZWYL01] X. Zhao, W. X. Wang, K. N. Yu, and P. K. S. Lam. Biomagnification of radiocesium in a marine piscivorous fish. Marine Ecology Progress Series, 222:227–237, 2001. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
222/m222p227.pdf.

Zhou:2006:IFB

[ZYZ^{+}06] Y. Zhou, H. Yang, T. Zhang, S. Liu, S. Zhang, Q. Liu, J. Xiang, and F. Zhang. Influence of filtering and biodeposition by the cultured scallop Chlamys farreri on benthic-pelagic coupling in a eutrophic bay in China. Marine Ecology Progress Series, 317:127–141, 2006. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2006/317/m317p127.pdf.

Zhou:2005:CBE

[ZZT05] M. Zhou, Y. Zhu, and K. S. Tande. Circulation and behavior of euphausiids in two Norwegian sub-Arctic fjords. Marine Ecology Progress Series, 300:159–178, 2005. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2005/300/
m300p159.pdf.